A Complete Bibliography of Publications in Applied Mathematics and Computation: 2010–2014

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 December 2023 Version 1.80

Title word cross-reference

(0, +∞) [Occ11]. (1 + 1) [Waz10c, Waz11a]. (2 + 1) [Bai13, DGDA10, Ibr14, LzJZ12, LDLM13, ROK12, SS10f, THZF12, Waz13, WX13d, Zha13a, Asl10b, Asl11a, BZ10a, lChH14, DL12a, GMST10, Kud10c, KRS10, Li10d, LWD10, LLWW10, LLL12, LLL10b, LDLL10, MGYQ12, Pan10, QW10a, SF11a, TW10b, WCL14, Waz10b, Waz10c, WM10, Waz11a, Waz12a, Wen10a, Wu12d, XZT10, qYqLtW10, YM10c, Yu12a, Zha10b, Zha10a, ZX11a, yZqLwW12, ZGW14a]. (2, 2) [WTFC13]. (2, 2, 0) [PH10b]. (3 + 1) [AM12c, AM12d, HCD14, KA10a, MAA11, MZ12, Pan12a, TMXG11, TTM12, TW10b, ZgJlK12, Zha10d, Zha12b, ZZR14]. (A)(C, α) [ÇE11]. (a, c, −a) [Rim13a, Rim13c].

(AB)^{(1,2,3)} [WC13, YL11c]. (AB)^{(1,2,4)} [YL11c]. (AB)^{(1,3)} [WC13]. (AB)^{(1,3,4)} [WC13]. (Aη) [HZ10]. (αγ) [AKH13]. (C, 1) [Bel14, TÇ13]. (C^{1} · N_{p}) [MMKD14]. (δ′)^{k} [LL14c]. (dρηθ) [TD13]. (E.A) [AN– D12]. (η, θ) [MN11]. (F, α, β, ρ, σ, d) [Tri14]. (\frac{G′}{G}) [Asl10a, Asl10c, GZ10d, GZZ10, Hay10, KET10, LLN10, ML10b, MCM10, ÖA10, She10, Zuo10]. (g, f) [RRR11]. (H, ϕ) [pLjH10]. (h_{0}, h) [LLX10]. (K, F) [Kas12a]. (M, N) [MPC14, TI10b]. (N) [Waz12a]. (N + 1) [GW11b, ARG14]. (n, k) [LX14b]. (n − k + 1) [GC11d]. (nk) [Zho12a]. (p(x), q(x)) [MTT14]. (P, Q) [ZW11a, ZJ14a, KSD13, Kra14a, MD14]. (p, r) [MN11]. (ϕ, ψ, p) [LAR13]. (Φ, ρ) [AZ14c, Ant11b]. (r ≥ 1) [MMKD14]. (ρηA) [Ver12]. (RS) [CWS10, YH13]. (ST) [ARR12]. (u, v) [SV12]. 0 [GHLF14, HLTU12, RFK12, ZHH^{+}13]. 0 < p ≤ 1 [CGJ12]. 1 [AMB10, BI10, BK10d, BK10c, BK10b, BZGK10, BT10b, BR10b, BT11b, BTHA11, CG12b, CG14c, DZ14, GHLF14, GFMP13, GVT^{+}13, IB10a, IPB10, IB10b, KUTY14, RFK12, SL14b, SB10, SB11, Tad12, TD13, ZHH^{+}13]. 1 + 1 [ELS13]. {1, 2, 3} [CIN14]. {1, 3, 4} [CIP10, LY10b, LHCI12]. 1.5 [ZZ13c]. 14 [YNBV11]. 150 [dRS13]. 1: −q [CGRS12]. 2 [AdFY13, Bab12, BS12a, BB14a, EL13, BWHQ14, Che10d, DG10c, GVT^{+}13, HWPY14, Hua10, HLS^{+}14, JM14b, JL10, KR13c, Koç13, LZ12c, LW10b, LMT11, Mit13, NTPVNHLX14, RV13a, RPW14, Sav10, SEL14, SK11b, SZ14c, TS10a, TVZ10, WFZL12, WZ13b, WW13i, Zha11a, ZLG^{+}14, bZpZJ10]. 2 + 1 [ELS13]. {2, 3} [CISM11, SCIMM11]. {2, 4} [CISM11, SCIMM11]. 2 − (ν, k, 1) [Han14, LHL14]. 2: −3 [DGOR13]. 2 cos(π/q) [BC12c]. 2m [FNG11]. 2n [LL11c, Wan10f]. 2s + 1 [BS14b]. 2 × 2 [CL11b, GD10b, LO14, SGZC13, ZCZP14]. 3 [AMA14a, ADD12, AP10, Bel11a, Bel12, BTRZ14, CVP^{+}10, CKK12, CEXG12, DTN14, GZW12, GG10b, HLS^{+}14, IN10, LP12a, LKY10, LWY14, LS11b, MG10a, Mam13, Med12, MMS14, MSS14a, Ou10, PMP11, RV13a, RVS10, SSK13, SR10, SHSC13, TÜ12, WHL11, YNBV11, YLTB12, YLJ14, ZCL11]. 3/8 [PVV11]. 3n [HN14b]. 3 × 3 [ÁGVZ12, GMA13, ZW10b]. 4 [ARL13, CL14d, DWZ12, ESNE^{+}14, GGM^{+}13, LLS^{+}13a, SIR14, TZS14, WWS^{+}14, YNBV11]. 4L [BX12]. 5 [LZM^{+}12, Rec14, SSA14b]. 6 [HWL11]. 6R [DESW10]. 7 [pSbL14, ZZ10d]. 8 [Jat12]. [0, 1] [OÖ13]. [x] = [A][x] + [b] [BI14]. ^{1} [ZCC13]. ^{2} [HH14, Hua10, XDZ14]. ^{2}E_{6}(q) [Han14]. _{a} [AH10]. _{n} [Rim13a, Rim13c]. A [BR10a, Kir11, yL14c, MKSN13, SMNN14]. A − BX [Wan14c]. A − BXC [GH10c]. A_{T,S}^{(2)} [SY14d]. A_{T,S}^{(2,3)} [SG14]. A^{β} [GGM^{+}13]. A_{M,N}^{†} [Ji14a]. a^{†}a [BCI13]. A_{1}XB_{1} + C_{1}X^{T} D_{1} = M_{1} [LL11b]. A_{2}XB_{2} + C_{2}X^{T} D_{2} = M_{2} [LL11b]. A_{T,S}^{(2)} [LLCX10, LQY11, SC11, Yu11b]. aa^{†} [BCI13]. AA^{†} − A^{†}A [BR10a]. a ∈ R[??]{0} [Rim13a, Rim13c]. α [CFS11, DRS11b, Fer12a, MÇ12a, SJ12, Ste10g, UOK10]. αβγ [ZH13]. AX + C [Liu14d]. AX = B [QW11b, Yua10, YWD13, Zho12b]. AX = C [XQ11a]. AXB + CX^{T} D = F [XLY10]. AXB + CY D = D [LGZ14a]. AXB = C [TW13]. AXB = E [CPZ10]. AX ± X^{⋆}B^{⋆} = C [CCL12]. B [Kud10c, Li14b, Zha10a, ĆARA12, RPA14]. B(2, 2) [JLM^{+}12]. B_{2} [KGH12]. BE(m + n + 1) [Waz13]. C [HM14b, CN14b]. C(R_{+}) [Ols12]. C^{∗} [MD11b, MD12a]. C^{0} [XSL10]. C^{2} [AMA12a, BAA13]. C^{n+1} [DHM14]. C_{1} [DM10b]. C_{g} [Yue14]. H_{∞} [XPLZ14]. P [PH10a]. c ∈ R [Rim13a, Rim13c]. coth [GS10b, Par10b]. CXD = F [CPZ10]. D [Mao11a, XSX12]. D(m, n) [BT11b, ZTW10]. D^{m} [CHS14a]. ẍ + (1 + ẋ^{2})x = 0 [Fer10b]. Δ [ÇÇD10, CWL^{+}12, sCWW^{+}14]. δ^{k} [LL14c]. Δ^{m} [Bra12, Sav10]. Δ_{ν}^{r} [DB12b]. δ_{±} [AKR14]. Δ_{uv} [SK12b]. Ḃ_{∞,∞}^{−1} [BTRZ14]. ẋ = Ax [Koh13]. ẏ(t) = ∑_{i=1}^{n} β_{i}(t)[y(t − δ_{i}) − y(t − τ_{i})] [DCR13]. E [GT11b, LW14e, MEBYG11]. E(m) [GMMIDM13]. E − v [MMY11]. E_{2}^{4} [KGH12]. ℓ^{p} [AKZ14]. ℓ_{1} [jL12, ZL13d, DB12b]. ϵ [BLK11, BC12d, MF13, Ver12, Ver13]. η [pLjH10, Ver10, ZW13d]. exp [DGDA10]. exp(−Ax^{2n}) [Boy14]. exp(−Ax^{4}) [Boy14]. F [LTZ12, Ery14, WN12a, Alo13]. F (p, q, s) [YM10a, Yan11e]. F (pqs) [Ste12d, Yan12d]. f(x + y) = f(x)f(y) [Jum12]. f′(x_{n}) = 0 [KBK13]. F_{4}(q) [LHL14]. F_{p} + uF_{p} + u^{2}F_{p} [Cen11]. F † F [BT11a]. F F † [BT11a]. \frac{G′}{G} [FLW11, Kud10a, LTW10, Par10c, Zhe11, ZL10g, ZZ11c]. G [ANR12, GS12a, GG11b, GS14d, HRX14, KGM14, MAK13, PLTH12, ZZX14, ZJ14c]. G′/G [HBDC14, QMM13]. G^{1} [KZ14c, LWZ11]. G^{2} [BAA13, Lu12d, RM11]. γ [BÖ11, DWZ12, JLZB14]. Γ_{3} [CS11c]. GM(1, 1) [CH13, LXGM14]. H [Alo13, HP12, QHYW14, ZZGL14a, ZDC12, ZCC13, BDR12b, BRGP14, DHMU14, Mat14d, Mat14c]. H(·, ·) [KKS11, LjH11]. H(div) [JMY13]. H^{1} [LL12e]. H^{1}(R^{3}) [WWQ14]. H^{∞} [Ste10a, Ste10n]. H_{x,ν}^{s} [ML10a]. H_{∞} [AS14b, BWHQ14, FGFRV11, HZ12a, JPK12, LJKP11, LLP14a, LZLF10, LSSZ14, LCL^{+}13a, NZ12, RZ12, WY11b, WZH13, XX14b, YMZS10, ZLM12, ZLS12b, ZZ13b, ZW14b, ZZY^{+}14a, ZCZ^{+}13a, Zim14]. Ê [Yu12a]. antitridiag_{n}(a, c, −a) [Rim13b]. I [Gu10, Tri14, pYX10]. ∫_{a}^{b} f(t) du(t) [Alo13]. J [SSN^{+}12, WLLD13]. K [CHC12, Hou14, Öke11, SN12, AG14f, BGV10, Che11f, EMS10a, Ery14, Jun10b, Jun11a, SEM11, SY14b, Ste13f, SXH11, WL11, WFZL12, YT13a, YL11f, Yua12]. K(2, 2) [DPC10, ZCT11, ZTF10]. K(3, 2) [HM10a]. K(m, n) [Yan11a, TW10b]. K(mn) [BGGH12]. kp [CZL14, GZL14]. L [PVV11, HH14, SÖCS12, XDZ14, Yun11b, RNZH13]. L^{1} [LZ12a]. L^{2} [JMY13, RM12a, Yan12b]. L^{∞} [BH11]. L^{p} [BMA11, KJ10, LLW10a, Lim12, CDCS13, GH11]. L^{q} [YLZ13]. L_{1} [LXK13, Ant12, SK12b, WY10c]. L_{2} [BRP13, FF11, FLZ14, Lee11, LDL11]. L_{2}^{(m)}(0, 1) [CHS14a]. L_{∞} [BRP13, HX13]. l_{p} [CGJ12]. L_{p}(a, b) [HHT13]. L_{q} [SMDZG13]. Λ [MD13a, BS10c, SW14c, VMRA14]. λ − ω [AA14a]. LM [TSP10b]. M [LCL^{+}13b, LCF^{+}14, MN14b, Waz10c, WLXL10, YDTW11, YZ12b, CL14d, Kar14, LLW10b, WCZ11, XYL10]. M(a, b, z) [LS14c]. M/M/1 [Amm14]. M^{[x]}/G/1 [KHP11]. M^{x}/G(a, b)/1 [JS12]. m_{1} = m_{2} > m_{3} = m_{4} = m > 0 [CL14d]. m_{1} = m_{2} ≠ m_{3} = m_{4} [ARL13]. C^{n} [Ste10m]. C^{n×m} [SG10]. C^{⋆} [DESW10]. F_{p} [ÖTC11, dRS13]. R^{3} [DAHAK10, Khe13, Kim14b]. R^{4} [YYL14]. R^{m} [Sch11a]. R^{N} [DH14, CLW14, OGJLCCB14]. Z [ÖTC11]. Z_{6} [pSbL14]. Z_{p} [KK13a]. A_{∈\+∞}(§) [AA14b]. H_{∞} [KPP^{+}13, LLKP14]. Ni(x) [HK11]. μ [LS10a]. N [Waz11a, WG12, Yu12b, Asl12b, Don14, GX12, Ji14a, KCLC13, LMS10, Liu14c, MM13a, MGWG11, SF11a, SR14c, WWZ13, Waz10e, YL11e, YZZ10, CALR11, DX14b, Ery14, GC11d, Kha14c, KZ12, LJR14, LWY14, LYB12, NTPVNHLX14, ÖKU12, PPK11, PC13b, PP12b, PP15, QHZ^{+}13, SPT12, Shi14b, Ste10a, Ste10n, Ste12w, SMPA13, UMMT14, Wan11d, WW13j, XZ12, Yan11e, Yua12]. n + 1 [YL14a]. n ≥ 4 [SMPA13]. O(1/t) [Hou13, SX14]. ω [DN12b, LS10a]. P [Pat13a, SDN14, YZC14, ABG11, BD11, BDR12b, BR14b, CL11c, CLL14, DX14a, DHMU14, GLZ10a, HLMW13, HR12, HRX14, Irm11, JŞ11, KK13a, Kon14, KLW14b, KP11b, LW12a, xLhYlZ13, LL12g, MTER14, Ozd11, SI11a, SS10d, TKS11, XC10b, XC10a, ZT12, ZLWW14, ZF13b]. p(t) [MAO10]. p(x) [Yos13, DWC11]. P_{ϕ}^{g} [Ste10k]. P_{0} [Fan10b, LWW11, HM12]. P_{1} [KCBH14]. Φ [Fer12a, CS12a, sCLCY11, sCCL11, CLCZ12, sCWT^{+}12, GS11d, ICKV14, QCK10a, RRG11, SK11a, YZK12, Yan12a]. ϕ_{0} [WFW10]. P m, H_{j1}// ∑ w_{i}C_{i} [MKSC13]. ψ [KCZK11, VR12]. P T [Vir13]. P X = XP [QW11b]. Q [BS10a, YGZW14, AAC11, AGK13, AKS14, AR13, ÁNA12, AAS14a, AAGR13, BB14b, CS13b, Dha13, DJ14b, EDGC14, ESG11, Fan14d, Fan14c, Fin14, GK12a, GA10, GA11b, Kam14a, KKS11, KK13a, KK14c, LWN13, LGMZ12, Mah14a, Mah14b, MP14d, MKSN13, ÖD11a, OÖ13, OO10, SS10b, Ste13k, Ste14b, Yan14b, YD14, ZLL10]. q < 1 [Mah14a]. q > 1 [OÖ13, OO10]. Q_{k}(p, q) [Ren14a]. q ∈ (0, 1) [ZZZ14]. QR [SPKS12, Su13, XBC^{+}13, AR11]. R [HYC^{+}10, Sha13b, Alo13, DB12b, SC10a]. r − d [Wu14a]. R^{1+1} [XX11]. R^{2} [KSM11]. R^{k} [KP12]. R^{n} [Xia13]. ρ [MN11]. S [AAMADH11, CGJ12, CKD12, DN12b, ZPXX14, AKÖ11, Fra10, HXQ14, ISÖ14]. S_{3} [KP13]. S_{w} [qGyFzX13]. SB^{3}A [DMR10]. σ [Sav11]. SL(n, R) [Sag11]. so(3, R) [Ma13a]. ST [SSRA12]. ⋆ [CCL12]. T [BLK^{+}13, Cao10b, Che14j, KCZK11, ZPXX14, TW11b]. t/k [YLQ13]. tanh [GS10b, Par10b]. τ [CMGR13, GJGSNRÁ12]. θ [BCÁV14]. Ẽ [Yu12a]. U [Kek11, Alo13]. u′(t) = Lu(t) + Mu([t]) [LLY14b]. u′′′ + qu′′ + pu = ϵF (t, u, u′, u′′, u′′′) [LM12]. u = a_{1}v^{m} + a_{0} [SV12]. V [AZ14c]. φ [sCCYW13, ZS15]. ∨ [Tri14]. W [LZ10d, SY12, AKR10a, LAR13, LLK13]. W (L^{r}ξ(t)) [MMKD14]. w^{∗} [NKR12]. x(t_{0}) = x_{0} [Koh13]. X + A^{∗}X^{q}A = Q(q > 0) [fZwXyZ11]. X + A^{H} X̄^{−1}A = I [ZCL13b]. X + ∑_{i=1}^{m} A_{i}^{∗}X^{−1}X_{a} = I [DLL11b]. X + ∑_{i=1}^{m} A_{i}^{⋆}X^{−1}A_{i} = I [mHhL10]. x = 0 [BX12]. X = Af(X)B + C [Chi14]. x = L [BX12]. X^{s} + A ∗ X^{−t}A = Q [CC10a]. X^{s} + A^{∗}X^{−q}A = I [WWH13]. X^{s} + A^{∗}X^{−t}A = Q [ZCZ13b]. X^{s} + ∑_{i=1}^{m} A^{i}^{∗}X^{t_{i}}A_{i} = Q [LC14b]. X^{s} − A^{⋆}X^{−t}A = Q [yYyL10]. x^{x−1} [Mor14a]. x_{n+1} = A_{n} + x_{n−1}^{p}/x_{n}^{q} [PSS11]. x_{n+1} = \frac{1}{B_{n}x_{n}+x_{n−1}} [HSX10]. x_{n+1} = p_{n} + x_{n−3k+1}/x_{n−k+1} [SXH11]. x_{n+1} = p_{n} + x_{n−k}/x_{n} [Öca14]. x_{n+1} = y_{n}x_{n−k}/(y_{n−k+1}(a_{n} + b_{n}y_{n}x_{n−k})) [Ste13m]. x_{n} = c_{n}y_{n−1}/(a_{n} + b_{n}y_{n−1}x_{n−2}y_{n−3}) [Ste13n]. x_{n} = \frac{x_{n−k}}{b+cx_{n−1}···x_{n−k}} [Ste12q]. x_{n} = max\{\frac{1}{x_{n−m}} , \frac{A_{n}}{x_{n−r}} \} [SHWX14]. x_{n} = x_{n−2}/(b_{n} + c_{n}x_{n−1}x_{n−2}) [Ste11f]. XB = D [Liu14d, XQ11a]. XC = D [Yua10]. XH = sX [QW11b]. X ± A^{⋆}X^{−1}A = Q [Has10]. y′′ = f(x, y) [SMIS14]. y′′ = f(x, y, y′) [Jat12]. y_{n+1} = x_{n}y_{n−k}/(x_{n−k+1}(c_{n} + d_{n}x_{n}y_{n−k})) [Ste13m]. y_{n} = γ_{n}x_{n−3}/(α_{n} + β_{n}x_{n−2}x_{n−3}) [Ste13n]. Z(x, y, n) = Z(a_{11}x + a_{12}y, a_{21}x + a_{22}y, n + 1) [RV14]. Z_{2} [LZM^{+}12]. Z_{4} [XH10, ZZ10d]. Z_{n} [DLC10]. ||A||_{k} [Sul10].

* [ZCZP14]. *-regular [ZCZP14].

-3 [DFG14]. -accretive [KKS11, LjH11]. -actions [DESW10]. -adic [KK13a, SS10d, Ozd11]. -algebras [MD11b, MD12a]. -algorithm [GT11b]. -analogue [AGK13]. -Analogues [EDGC14]. -Appell [Mah14b]. -Approximate [BC12d]. -ary [LWY14, WFZL12]. -asymptotic [BH11]. -asymptotically [sCLCY11, CLCZ12, sCWT^{+}12, sCCYW13, SK11a, sCCL11, DN12b, Gu10, pYX10]. -band [SHSC13, WLXL10]. -based [qGyFzX13]. -Baskakov [GA11b]. -Bernstein [AAC11, Fin14, OÖ13, OO10, ZH13]. -Bézier [QHYW14]. -Bloch [CFS11, LS10a, Ste10g]. -body [ARL13, ARG14, CL14d]. -boundedness [CDCS13]. -braids [IN10]. -Brownian [HRX14]. -circulant [SC10a, ZJ14c]. -class [MMKD14, DHM14]. -Coherent [MPC14]. -compatible [AKR10a, NKR12]. -component [Waz10c]. -condition [LSLL14]. -conformal [MN14b]. -conjugate [WLLD13, CWS10]. -constraint [MF13]. -contour [Fan14c]. -contraction [RRG11]. -contractive [SSRA12, ARR12]. -control [Zim14]. -convergence [Sav11, CWL^{+}12, sCWW^{+}14, Pat13a].

-convex [AKÖ11, DRS11b, HXQ14, ISÖ14, Mat14d, GG11b, MEBYG11]. -coupled [Don14, WWZ13]. -cube [LWY14]. -cubes [WFZL12]. -cyclic [TS10a]. -D [Bel11a, Bel12, BWHQ14, CVP^{+}10, CEXG12, LZ12c, SL14b, SZ14c, Tad12, WZ13b, YLTB12, Zha11a]. -demicontractive [BS10c]. -dependent [TW11b]. -derivatives [GK12a]. -designs [BS14b]. -diagnosability [YLQ13]. -Difference [Fan14d, AAGR13, DJ14b, SS10b, Ste13k, Yan14b]. -differential [ESG11]. -Dimensional [Waz13, AM12c, AM12d, Asl10b, Asl11a, BZ10a, Bai13, BS12a, lChH14, CKK12, CS11d, DL12a, DWZ12, GMST10, HCD14, Ibr14, KA10a, Kud10c, KRS10, LJR14, Li10d, LWD10, LLWW10, LLL12, LLL10b, LDLL10, LDLM13, MAA11, MZ12, MGYQ12, Pan10, Pan12a, PMP11, QW10a, SF11a, TMXG11, TTM12, TZS14, TW10b, Wan11d, WCL14, Waz10b, WM10, Wen10a, WX13d, Wu10c, Wu12d, XZT10, XGG^{+}11, qYqLtW10, YM10c, Yu12a, Zha10b, Zha10a, ZX11a, ZgJlK12, yZqLwW12, Zha13a, ZGW14a, ZDW10, Zha10d, Zha12b, ZZR14, ZZ13c, PPK11, Waz12a, Liu14c]. -dimensions [Waz10c]. -distance [LAR13, LLK13]. -distribution [WN12a]. -Durrmeyer [Mah14a, ZLL10]. -efficiency [Ver13]. -eigenvalue [YGZW14]. -equivariant [DLC10, LZM^{+}12, pSbL14, XH10, ZZ10d]. -Euler [KK13a, KK14c]. -expansion [LTZ12, Par10c, Asl10a, Asl10c, FLW11, GZ10d, GZZ10, Hay10, HBDC14, Kud10a, KET10, LTW10, LLN10, ML10b, MCM10, ÖA10, QMM13, She10, Zhe11, ZL10g, Zuo10, ZZ11c]. -factorization [XBC^{+}13]. -Fibonacci [EMS10a]. -fold [KZ12, SPT12, YL11f, MM13a, MGWG11]. -Fréchet [AG14f]. -free [Ji14a]. -function [AH10, DGDA10, BS10a, Fan10b, HM14b]. -functional-difference [Ste14b].

-functions [SÖCS12]. -gain [LXK13, LDL11]. -Galerkin [LL12e, ZCC13]. -gamma [Kra14a, KSD13]. -Gaussian [Kam14a]. -generalized [CISM11, CIN14, LHCI12]. -generating [CS13b]. -Genocchi [AAS14a]. -gon [KCLC13]. -good-neighbor [PLTH12]. -grids [SY12]. -groups [ABG11]. -Hankel [AdFY13]. -harmonic [LW12a]. -Hausdorff [AR13]. -Henon [xLhYlZ13]. -Hermicity [ZW13d]. -Hölder [Alo13]. -Horadam [YT13a]. -infinity [ZDC12]. -integers [AKS14, MP14d]. -integral [KK13a, AR11]. -inverse [LY10b, ZZX14, TSP10b]. -inverses [CIP10, MD14, SCIMM11]. -invexity [Ant11b, MN11, AZ14c, Ver12]. -Iterated [KR13c]. -Jacobi [LWN13, SSN^{+}12]. -junction [BLK^{+}13]. -Kirchhoff [MTT14]. -Lagrange [MKSN13]. -Laplace [KLW14b]. -Laplacian [BD11, BR14b, CS12a, CLL14, DWC11, DX14a, GLZ10a, HLMW13, HR12, JŞ11, Kon14, LL12g, MAO10, ZT12, ZLWW14, ZF13b]. -Laplacians [Yos13]. -largest [CGJ12]. -like [Fer12a]. -linear [AAGR13, CMGR13]. -logistic [KP11b]. -Lucas [BB14b]. -machine [XYL10]. -manifold [LMT11]. -matrices [YZ12b, Yun11b, ZZGL14a]. -matrix [YDTW11]. -means [CHC12]. -method [BCÁV14]. -metric [RPA14, ANR12, GS12a, MAK13]. -metrics [ĆARA12]. -Moment [HRX14]. -monotone [pLjH10, HZ10, Hou14, yL14c, SMNN14]. -multi [Lu12d, RM11]. -NCP [HM12]. -Nekrasov [CKD12]. -noded [Kha14c]. -nonexpansive [QCK10a, YZK12]. -norm [WY10c, FF11]. -normed [Sav10]. -numbers [Kek11]. -numerical [CN14b]. -operator [HP12, PH10a]. -optimality [Ver12]. -orthogonal [AAGR13]. -out-of- [Ery14, GC11d, Yua12]. -periodic [DN12b]. -ply [CALR11]. -point [Bab12, HWL11, LLW10b, LLS^{+}13a, SR10, SSA14b, SIR14, WCZ11, SR14c]. -polynomials [Fan14d, TKS11]. -power [GJGSNRÁ12, XC10b, XC10a]. -preinvex [Mat14c]. -projected [XDZ14]. -projection [JMY13]. -projections [HH14]. -Prolate [Dha13]. -property [LWW11, LW14e]. -proximal [Ver10]. -pseudoconvexity [Kas12a]. -quasi-contractions [VR12]. -quasi-slowly [ÇÇD10]. -quasinorm [CGJ12]. -Racah [ÁNA12]. -reflexive [ZJ14a]. -regular [Che11f]. -restricted [BGV10, Ou10]. -Ritz [CZL14, GZL14]. -sampling [Dha13]. -section [PP12b, PP15, SMPA13]. -set [GFMP13, LCL^{+}13b, LCF^{+}14]. -singular [Kar14]. -skew [HYC^{+}10]. -Soliton [AMB10, BI10, BK10d, BK10c, BK10b, BZGK10, BT10b, BT11b, BTHA11, IB10a, IPB10, IB10b, BR10b, SB10, SB11, Waz11a, WG12, Yu12b, Asl12b, GX12, LMS10, SF11a, Waz10e, YZZ10]. -solutions [GH11, BMA11]. -space [TÜ12, YLJ14]. -spaces [LKY10, LZ12a, Lim12]. -species [ADD12, DX14b, GGM^{+}13, YL11e]. -splines [CHS14a]. -stability [WFW10, Cao10b, Mao11a, XSX12]. -stabilization [LLX10, Mao11a]. -stable [PVV11, YZC14]. -stage [YNBV11]. -star [Zho12a, LX14b]. -starlike [BÖ11]. -state [RNZH13]. -statistical [MKSN13]. -Steps [UMMT14]. -stress [Che14j]. -strict [SW14c]. -strictly [Jun10b, Jun11a]. -strong [MD13a]. -strongly [GS11d, Yan12a, ZS15, Sav10]. -summable [BRGP14]. -Sylvester [CCL12]. -symmetric [Vir13, YH13]. -symmetries [VMRA14]. -systems [SDN14]. -Szász [ÖD11a]. -Szász-beta [YD14]. -th [Fra10, Yan11e]. -theory [AKH13]. -tridiagonal [SEM11, SY14b]. -type [TD13, GS14d, Kud10c, KGM14, Tri14, XSL10, Zha10a]. -uniform [BLK11]. -uniformly [KKS11]. -valent [SI11a]. -valently [Irm11]. -vex [MMY11].

-Volkenborn [AAC11]. -wave

[AAMADH11]. -weak [ICKV14]. -weakly [LAR13]. -weighted [LZ10d]. -Wiener [LGMZ12]. -Zygmund [LS10a].

//doi.org/10.1016/j.amc.2014.03.101 [WHS14a]. /complex [AQ11].

1 [BM14a, BPP12b, CGL12, CCJ14b, LT11, wMwWsT11, TLW11, TX14, WSY14]. 10.1016/j.amc.2009.07.009 [Asl10d]. 113-120 [BDG13]. 160 [Pan12b]. 1684 [EM11]. 174 [Mar14]. 196 [WZ13c]. 199 [SQ11].

2 [Boy13]. 202 [BDG13]. 212 [LL12e]. 215 [Tua13]. 216 [Gür12, Web12]. 217 [Aba12, JDJ12, Xie12a]. 218 [AFSMRL14, Ban14, BM12b]. 219 [AKB13a, Pan15, PP15, Rim13a, TJ13, Tom15]. 232 [AZ19, BCNP14a, EH15]. 233 [Rho19]. 236 [CX14b, ZZ15]. 237 [WHS14a]. 242 [YZZP14a]. 246 [PK18].

3-propanediol [YZZ^{+}14]. 3D [GXC^{+}11]. 3RPR [CR11a].

500 [Web12]. 5604 [BM12b].

7001-7008 [Mos14].

= [ZLGZ11].

A-posteriori [CYZ14b]. A.K [Ban14]. A/H1N1 [RG14]. A2 [Rec13]. Aarts’ [AGZ12]. ABC [FGIV12]. Abel [Bou10, BRM11b, DS13a, GÖS11, PSSS14, SHMCM12, Yan14a]. abelian [ABG11, WZL14b]. Ablowitz [Asl10a]. abrupt [SSM13, dLDP13]. absence [BM14c, CRC13]. Absolute [WL13a, Bor11, EG11, Et13, GSSS13, HD14b, KSM11, Kra14b, MT10, NP10a, WW11d]. absolutely [CF13, HD14b]. Absorbing [Dea13, zSWZW12]. absorption [DAA14, LJCI12, SBS10, VV11a]. Abstract [SS13b, DN14b, HPB11, HOB12, POR13]. abundant [SHY11]. ac [GARK12, LS13b]. ac-driven [GARK12, LS13b]. academic [CSSSV13]. accelerated [LWYY14, LZ13b, MKP13, SGC10, SSU13, Yan13d, ZM14d]. Accelerating [lJCWpW12, Bum12, NL12a]. acceleration [Koç11, PP14c, Sie14, SJ13]. acceptability [BPDZ12]. acceptance [XZZK12]. accidents [SFL^{+}11]. according [SFL^{+}11]. accounting [NDZ14]. accounting-based [NDZ14]. accretive [GS11d, fHcXH11, KKS11, LjH11, SKC10, ST14, Yan12a, ZS15]. Accuracy [GKS12, MH10a, CCJT11, CC14c, Chu11, Dub13, EJ12, GC14, GS14b, HZ13, JW12, Lin12, MG11b, MS12b, MG14c, MSJ14, Nie13, OSRFF14, OSRPC12, SZ11, TLDZ13, Zhu10d]. Accurate [ABF14, DP13a, Fer12b, JBL^{+}11, SKB14, BB14a, Bra13, CZC11, DWLL10, DNSM11, DP14, DKK^{+}13, Jha13, KV14, MVGS11, NN11a, SYK10, SU12, SGA14, TK14b]. acid [HRH14, LKY14, LY14c, ZLC10]. acid-mediated [HRH14, LKY14]. acoustic [Apa10, GO13, Kim14b, OPD12, OR14]. acquisition [TDSA14]. across [LCD12]. action [AS11c, DG14b, FW12, IN10, RH14, dOL13]. action-angle [DG14b]. actions [DESW10]. activation [ASMA12, HZW12, HZW14]. activations [GLY14]. Active [KSY12, RA10a, YDZ11, Ber10, ES12a, PWZZ13, Yu14]. active-set [ES12a, Yu14]. activities [HLS10]. activity [KMC12, WW11d]. actuator [DKK^{+}13]. acute [aBS14]. acyclic [KI10a]. Adams [KMR14]. adaptation [Gal11d, HJVC14, Zhe13]. adapted [GR11, Kum14, MQ14, NBYHV11, XL11, YCF13]. Adaptive [CYZ14a, Che10e, FK12, GY13, JJPW13, JJP^{+}12, LLJF12, NRWF12, NP14, SKD12, THX14, WXWY14, WT11d, XWMW12, Yan10a, ZSL11, ZWD12, ZLZ14b, Zhe12, Ach13, AA12a, ARS14, BS12a, Bie11, BCGV11, CBO12, CZZ^{+}11, Che11b, Che13c, dSCAM14, ESE14, For11, GPP^{+}10, GTPM14, GXF13, GL13c, HVD^{+}14, HM10b, HJZC13, Jia12, JFHZ14, KP14a, KA13a, Ko13, KPBC14, LDWY14, LS11b, MZXZ10, MHA12, Mit13, NTdCJL12, NA14b, PSTL10, SCC12, SCK13, SWLP13, TZZX11, Tsa13a, Tsa14, WHW10, Wan11b, WH13, WW13a, WWW13b, WLY^{+}12, Xu13a, XZZX14, YC14, YSW11, ZZQ10, ZZW^{+}12, ESE13, FB10]. adaptivity [DM13b]. Addendum [Web12]. additional [DMGK13, SP14b]. additive [CI11, DRGR10, Gal11e, Ge14, LJR14, PP14a, RMMM11, YLZ13, Zha14e]. ADI [AB12b, Yan10b, ZWT12]. adic [KK13a, SS10d, Ozd11]. Adini [YWL13]. adjacent [WC11]. Adjoint [RP14b, AS11a, AY11c, Car13a, Car13b, Fre11, GB12, PT11a, PB14a, WT11b, SQJ11]. Adjoint-based [RP14b]. adjointness [IKV11]. adjustment [SVV13]. Adleman [WZZL12]. ADM [CZL11, GC11a]. ADM-Padé [CZL11]. Admissibility [SSM13, WPSC12, DLZ14, LZL12]. admission [Dim13a]. admitting [Van10]. Adomian [BDG13, DCR12, ARH^{+}12, ASRI12, AZAK10, BCEAV12, BRM11b, BMR13, BR13, Dua10a, Dua10b, Dua11a, DR11b, DR11c, EK10, Els12, HBB11, MP10, MP12a, QS11, QS14, RD11, SMH14, SK10c, WZZ11, Waz10a, WRD13]. adults [LD13a]. advance [CT11a]. advanced [AZL13, Fig12, Jan10, Jan14c]. advantage [TDSA14]. Advantages [NN13]. advection [AP14, CLAT11, JZG13, LGN10, SGDK11, Sar12c, SXC^{+}12, ZLZ^{+}14a]. advection-diffusion [JZG13, SXC^{+}12]. advection-diffusion-reaction [Sar12c]. advection-dispersion [SGDK11, ZLZ^{+}14a]. advection-reaction-diffusion [AP14]. advertising [DLP10, JMHZ10]. aerodynamical [XBC^{+}13, XBC^{+}14]. aeroelastic [LCC12, Svá11]. aerofoil [RJ12]. Aerospace [BHJ13, DHDB12]. aesthetic [GTPM14]. affairs [BBF12b]. Affine [Yam14, Abd13, BDH^{+}14, CPOdCN11, HH13b, MVS13, PQW10, Sta14c, WZ14b, ZL12b, ZH10b]. affine-scaling [WZ14b]. affinity [Sza13]. after [yWZhL14]. again [Asl10c]. Agarwal [CS13b]. AGDC [AGHdD13]. Age [VDLEK14, DYL14, Fuk12, KLY12, LZG10, LZ11c, LH14b, LYB12, LH14c, LF14, MZW14, MCL14a, McC10, RWkPk10, SGE13, SKC14, ZR13, ZW13b, ZLL14, ZMC10]. Age-dependency [VDLEK14]. age-dependent [LYB12, LH14c, LF14, MZW14, ZR13, ZW13b, ZLL14]. age-of-infection [McC10]. age-period-cohort [Fuk12]. age-size-structure [LH14b]. age-structured [KLY12, LZ11c, RWkPk10, SGE13]. agent [BMI^{+}11, CWCW11, HL14c, LWM10, LZWZ14, Liu13b, LSL^{+}14, Pal13, WLLY14, XLLL14, YCC^{+}12, ZL14d, ZPZ14]. agent-based [BMI^{+}11, HL14c, Liu13b, Pal13]. agents [LCH13, LYZG13]. agglomerated [HA12]. aggregate [XY10]. aggregated [Lin11d]. aggregation [AGM13, Cel14, EB12, GGM^{+}13, SGK14, Sun14]. agility [LDR10]. aging [Rud12, Yan10e]. ahead [UMMT14]. AHP [BDGIPG12, BCF13, GRKP10, SS11d]. aided [BP10b, ESAH10, LW06]. AIDS [BMH12, CGW14, NTS11a, NTS11b]. Air [MSGN^{+}14, BGP^{+}13, LPY11, NCdCDM13, NA14b, WY10a]. aircraft [AHS14]. Airfoil [Men12a, LHL^{+}13, Svá11, ZTX10]. Airy [CCCRR13, HK11, XX14c, ZD11b]. Aitken [Bum12, Ger12, PC11]. AKNS [lChH14, LzJZ12, LLL10b, QW10a]. al [MK08, MK10a]. al. [Pak11b, PS11]. alcohol [RLW14]. Aleph [SP11]. algal [Mis11]. algebra [CCSM13, DS11, DWZ12, HT11b, MM14a, MD13c, ZDW11]. Algebraic [BZ10a, Ben14a, DESW10, NW12, TKL14, AMPT14, BDH^{+}14, BJ11, CB12a, DSPZ13, Gem11, GMA13, GL10b, HS13, HW13a, HKH10, JJL14, KX10, Kek11, Kim10, LC13, Liu12c, dLfX14, LS14d, MX11, PEMS13, QHL10, RBMA12, SSV11, SCK14b, SSG13, UT13, WFwC12, XKH10, YWCH14, ZZC10, ZCZS12, Zha10e]. Algebraically [PRV13]. algebras [CDS13, MD13a, MD11b, MD12a, MD12c, NSV13, NRAV14, TZS14]. Algorithm [EH15, UJH14, Ach10a, AG14a, Ach13, AOTV12, AOTV14, Ahn13, AGR14b, AGZ12, AAA13a, ÁGVZ12, AC14, ANR14, And12b, And13b, AAS14b, ARS14, BMD13, BSR10, BDDK14, BLR13b, aBS14, qClLyW10, CGTCF13, sCCL11, CLCZ12, CY11a, CkPqL12, CSD12, Che14i, CXD14, CL10b, Cho11b, CTS14, dSCAM14, CM10, CAC14, DWY14, DHI13, DM13a, DAV12a, DGS11, Din11, DQL^{+}12, Din12, DHS10, DSZ14, Dua10a, DSPZ13, DHDB12, EA13, ES12a, EH14a, ESE14, EYAD13, EH14b, FYL13, FZ13, FOZX14, FH11a, FC14, FHK13, FS12, FWZW10, GWX12, GL11a, GW11a, fGyLJ11, GxFX^{+}12, GjLxF^{+}13, GVW^{+}14, Gar11, GS10a, Gee12, GM14a, GT11b, GZ12a, Gu13, GZ14b, GC11d, HLeH12, HAA^{+}14, HM10b, HCW11, HjH12, HjH14, Hsi14, Hu13b, HLZ14, HZW13, HLW13, HD14a, IN10, ICS^{+}13, IO12, JM11, JLL12, JK14b, JLHW10, JFHZ14, JPM12, KMW12]. algorithm [KRJ11, KA13a, Kay11, KAS12b, Kel12, KKSS14, KudKL14, KCK08, KGB12, Ko13, KZS12, La 13, yL14c, LNdCJV14, LG10b, LZ10a, LPX12, LZ13a, LTHG14, LPP^{+}12, LL11d, LK11, LWH14, LS11a, LZGZ11, cLYrL11, LMS^{+}12, LRM12, LLS^{+}13a, LYG14, LWS14, LP13a, LJ11, LWX13, LS14d, Luo14, MZ12, MLM^{+}14, MM10a, MB12, MIM^{+}14, MHS12, MT12b, MK14b, NR14, NBP10, Nem10a, NW10, NSYA14, Niu11, Oss10, ÖBT10, PSTL10, PZ14b, PJFL14, PAC13, PRK13, PZ11, PT11b, PBP14, PP14c, QTYS14, dejW14, Qj10, QC12, Rad11, RY14a, RV10a, RBMA12, Ral12, RWBA13, RCSH10, RAZES13, SLM14, SSC11, SAA14, DAV14b, SN12, Sev13, SW14a, SYl14, SKC11, SM10, SY14b, SW14b, SCIMM11, SLZ11, Su13, SNH10, SXGR11, TNB11, TZYH14, Tan14, TSN13, TL14, TMJ14, TPA10, LM13a, Ver10, VA12, VC12, WzH10, WW11b, WZ11f, WMH12]. algorithm [WDL12a, WY13b, WM13a, WTH^{+}14, Wan14e, WH11, Wu12c, XMA14, XBYF14, XZF11, XM11, Xu10c, Xu13a, XZZX14, YKI10, YCS11, YWYY12, pYmK13, bOYqGL^{+}14, YY14a, YNL10, YWCH11, YDZ11, YHL08, YYW12, YY13b, YJCC14, YDZH14, Yus10, ZS14a, ZDZ10, ZJD12, ZHH^{+}13, ZAY^{+}14, ZH10b, ZBM10, ZHZ12, ZHW11, Zhe13, ZZ12a, ZSQ10, ZWT12, ZF13b, ZGY14, ZYG14, ZK10b, ZYM11, ZL13d, ZP13, ZDH14, ZK14, yZxGyM10, BB11b, Civ13, KL13a, SKB14, STM10, Tsa14]. algorithmic [AQ11, DZ12, Kot14]. Algorithms [BCDI14, ASSPA10, AHS14, Ala10, AGM13, AK10b, And13a, BDH^{+}14, BA13, CÁGGLP14, Cha14a, CKP^{+}13, CAP14, CPOdCN11, DAV12b, DP14, DC14, qDG10, Dua11a, ESE13, GT13, HDZW11, HJZC13, Jia12, JW10b, KZ14a, LFLS14, LZLW11, LL11b, LKH12, Lin10, LPY11, Liu10d, LWS11b, LJKU11, Liu11b, jLLW12, MVQDE12, Mat12, Mit13, Örk13, PS14b, PEMS13, RFK12, SSN^{+}12, TSP10a, Wan13c, WY12c, Woź12, WLH11, WLD11, XLY10, YZ13, ZTY^{+}10, bZlD11, ZZYX12, ZM14d, Zil12, ZANA12]. Allee [MG12b, PP14a, SW13, Xu13b, ZM10c]. allelopathic [MAB14]. allelopathy [LCH12]. Allen [SPK14]. alliances [ALP11]. Allocating [Lin11c, Lin11b, SLJ^{+}13]. allocation [FYL13, HY11b, RTW13, SSTM14, TQWZ10, YH12, YH14e]. allowable [JGM13]. Allowing [AS13b]. Almost [CSS11, HV12a, LY14b, LHC14, ZS13b, ZjHO13, ZXF14, ARR12, CYHL11, ĆASH11, Dia13, DLX13, GWZ14, HPB11, LDL13, LW14h, Mil14a, MB13a, Mop11, PZ14a, PR10, RSRA14, SSRA12, SS13c, Sta14c, Wan13e, Wan14a, ZLJZ14, ZWZ12, ZZW11b, ZZQ13]. almost-periodic [GWZ14]. along [IBS11, LWWZ11, MP12b, YA14]. alpha [THNTNXR11]. Alsaedi [PK18]. Alternating [Yua13, ANS14, BBK13, BA14a, CC14a, GL13b, LL11d, MT12b, PL10, SX14, Sze10, Sze11, Sze12, YLTB12, ZY14b, LIC12]. alternating-direction [PL10]. Alternative [Lin10, YZ13, BB14b, Dos11, MB12, SP14c, THNTNXR11]. alternatives [LTSV14]. Altithermal [WDBM14]. AM [HSX^{+}11]. AM-SG [HSX^{+}11]. ambient [EASK10]. ambiguity [TP12]. Ambystoma [WDBM14]. AMC [EH15, Rho14b]. American [Ben14b, CZ13, GO10, WZ11g, ZY13c, ZZ11b, ZC11]. American-style [GO10]. ammonia [CBA^{+}14]. amoeboid [ZHH^{+}13]. among [ÇM13b, Koj13, LD13a, NR12, WT11c]. amphibians [EWS10]. amplitude [EK11, Fer10b, LLLZ11, WDZ13]. analog [GGTCdlF13, Łen11, Len14b, TP12]. Analogs [Kyr12]. analogue [AGK13, GA10, Suc14]. Analogues [EDGC14]. analyse [OSRPC12]. Analysis [Amm14, BC12a, BLZZ14, BSZ13, BDAD10, Cai10b, CGW14, CCG10, CLF13, DdGL10, DC13a, FP13c, FdBCC14, GL13a, Guo10, GTX11, HWL11, HLMW13, HM11, JK14c, Kim14b, KS12b, KPP^{+}13, LR11, LY10a, LG10a, LWW12, LCL14, LC14e, Liu14b, Mad14a, ML11b, NAL12, OD11b, QGD14, RMN14, RG14, RLV10, Sal11a, Sam11, SS14a, SMR14, Van11a, VCĆV12, WCB14, WSY14, WFZ13, XGZW13, XM12, Zak10, Ach10a, Ach13, AM14a, ABKG^{+}13, AAAH12, All11, AS14c, AHDD13, AH13b, AM14e, AHM14, AMPT14, AJPZ14, AM12e, BPDZ12, BB13a, BPP12b, BGV10, BY14, BM13b, BAE13, BVTC13, BHG10, BCÁV14, BRVR12, BRAS12, BA14b, Boy13, Bra13, dCBdFM14, BIT12b, BGGH12, BGP^{+}13, CL10a, qClLyW10, ÇM13a, CHK10, Cha14c, CL14b, CCCRR13, CDLZ11, CCX11, CZZ12, CCJ14b, CVP^{+}10, CN14c, CLS12, CYZ14b, CCSM13, CST14, DRJ11]. analysis [DWQG12, Dan14, Das13, DGS14a, DMGK13, DB11a, DMW13, DWP14, DRGR10, DISS14a, dCDNCFRH10, dCDNVS11, dCDNRH12, dCDNV^{+}13, DQL^{+}12, qDG10, DLL11b, DW13b, DHQ14, EJ12, EK11, EDYA14, EEZ14, EMH10, EEM12, EMAM13, EHSM14, EKSO14, EMM^{+}12, EGMPMRBC12, EÖ10, ET12b, Fal14, FXD12, FPC14, FHR13, FA12, FKH10a, FKH10b, FGO12, GGFM14, Ge14, GGM^{+}13, GGTCdlF13, Gün10, GYG11, GSDW14, GZL14, GSC14, GYL14, GB13, GH11, HB14, HN14a, HHM10, HSX^{+}11, HL11a, HET12, HYT13, HSZ14, HMMK11, HOMA13, HYZH14, HG13a, HLC13, HYSZ14, JE11, JPK12, Jia12, JK10, KR13a, gKW12, KRJ11, KJ12, KfM11b, KM14b, KB14, KDS13, KA10a, Kim11, Kos11a, KK13d, LLLZ11, yL14c, Lee11, LJ12b, LL14b, LYS10, LWL10, LT11, Li12b, LYS12, LY13a, LL13b, LIX14, LFR14, LZLB14, LLY14b, Lin11c, Lin11b, LDL11, sL10a, LWTC10, LZ10e, LLZY10, LCC12]. analysis [LML12, LSJZ13, LJ13, LZW13b, Liu13a, LYXP13, LWG13, LH14b, LZLY14, LYL11, Ma12, MH14a, MR13a, MTSN14, MS13a, MK14a, MP10, MP12a, MSGN^{+}14, MP11a, MH11a, MNOB10, MG12b, MO12a, MMCD12, MM10c, MK13b, Mor14b, Nar12, Neu12b, NBYHV11, NTPVNHLX14, NZW11, NL10, NW12, NN19, Odi10, OS11, OSAG13, ÖB11, Ou10, PB13a, PWZZ13, PJFL14, QLCF10, RS11a, RSK14, RPW14, RRLA14, RGL12, Rud13, RV13b, Sad11, SGE13, SAMA11, SSM13, SMI12, SA12a, SVGC10, SSP11, SLBS11, SL13a, SSS14, SW14a, SZ11, SZ14b, SSST14, Sie11, Sie14, SF11b, SGK11, SJPS14, Son10, STM10, Sun11b, SC13, SSM11, TZZX11, TA13, TFMB12, THNTNXR11, TW11a, TXX14, TJZ14, TT13, TRE13, Tsa11, VDL12b, VDLEK14, VG14a, VRZ11, Wan10e, WzH10, WY10a, WZ11c, WQ11a, WH12, WMH12, WW13f, WSD13, WW13e, WDZ13]. analysis [WWG13, WZ14a, WWZ14a, WD14b, Wei10a, WY10c, WPSC11, WW11g, WPSC12, Wu12b, XWWF10, XZF11, XXZH10, XTL10, XHXH12, XZ13a, XH13b, YS11a, YLS13, YJ14a, YWW^{+}14, Yua12, Zen12, ZJS10, pZJ11, ZZZ11, ZF11, ZS12, ZCL13a, ZC13a, ZS14b, ZT14, ZLZ^{+}14a, ZZWF13, ZZ10c, ZLD11, ZLY11b, ZLLL12, ZLZ14d, ZW14c, ZYC14, ZF10, ZW10d, ZCC13, ZZL14, ZSD10, ZZS14]. Analytic [AB11, AAT^{+}13, Asl11a, AMA10, BRGP14, FK13, GP12, KCC12, TZJZ12, Aja11, CFZ14b, CBS12, DCL11, Dua11a, EK10, Fer13, GGYW11, GO10, IJ11, Irm11, JVK10, Jun11b, KU11, KL10, LWL10, LLW13, LLL13, Liu11d, LOS13, MCS11, Mil12, NM08, OP14a, PP12a, PP14b, PS12a, Shi12, SKM11, SB12b, WXA10, WZ13c, XXS12, YL10a, YL11b]. Analytical [AE11, ASVV10, AKMM12, BVTC13, DT10, DG10b, DZ14, FL11, Fuk14a, GADK14, HB11a, HBB11, ITLA10, KKH14, KCLC13, LLLZ11, Mam11, MV12, Meh13, SM14a, SV10a, VSV10, Xia10a, ANHA14, AM14b, AGHdD13, AA12d, CLA^{+}12, CS12b, DRJ11, DR11a, DVGY11, DK11a, EZ14a, FKH10a, GDH10, HN14a, HLM14, HRAA12, HLZ13, JZG13, KAB11, MV14, NN19, QS11, QS14, UW10, XG13, Yun10a, Afz10]. analyticity [OO10]. Analyzing [HT14, BDH^{+}14, MSR14]. anchor [GH10a]. Anderson [Rad11]. aneurysm [ML11b]. angiogenesis [FP13c]. angle [DG14b, LL11a, RDFL13]. angled [GVT^{+}13]. angles [BR12, dOL13]. angular [LLW13]. Anharmonic [Ped13, LL13f, Sin12, ZK11]. animal [SF11b]. Anisotropic [CZM11, GS11b, Sha11a, AAMADH11, Aou11a, Den14, HJVC14, MATA14, MPS14a, MGC12, OR14, SCK13, YS10a, ZA14b, ZCZ14]. annealing [AGZ12, AK10b, CEM11, CSD12, Ko13, MT13, Örk13, SKC11, RRRRTR12]. annuity [AC13a]. annular [GAC12, HA12]. Annulus [DG14a, Mah13, TRE13]. anomalies [Mag14a]. ANOVA [GH10a]. Ant [EdCVB^{+}14, SEED13, ASS13, CMP13, KGB12, LKH12, MMM12, RAZES13, ZYM11]. Antev [WDBM14]. Anti [RKA^{+}13, WY10b, WQ13, AdFY13, AIAD12, BAP^{+}14, Çak13, Che12a, FZLZ13, GG11c, GG14g, HSC14, LY12c, MVS14b, NB11, Rim13a, Rim13c, Rim13b, SD10, Wu10a, WLY^{+}12]. anti-bidiagonal [NB11]. Anti-control [WY10b]. anti-diagonals [GG11c, GG14g, Rim13a, Rim13c]. anti-ideal [Che12a, FZLZ13]. anti-malware [MVS14b]. Anti-periodic [WQ13, AIAD12, Çak13, SD10]. anti-symmetric [BAP^{+}14]. anti-synchronization [WLY^{+}12]. anti-triangular [HSC14, LY12c]. anti-tridiagonal [AdFY13, GG11c, Rim13a, Rim13c, Rim13b, Wu10a]. Anti-windup-based [RKA^{+}13]. Anticipated [Xu13c]. anticipation [SW10b]. anticommutator [Den13]. antidiagonals [CDGR13]. antimaximum [HT11a]. Antiperiodic [NV12]. antitridiag [Rim13a, Rim13c]. antivirus [SSST14]. Ants [ZYM11]. Any [RNZH13, CDMS14, CZM11, LTW12, LL10h, Wan10c, ZL10g]. any-order [ZL10g]. AOR [TS10a, Yun11b]. aperiodic [ZT12]. Apostol [EDG14, LS11c, OS14]. Apostol-type [OS14]. appearance [Sta12b]. Appell [Çet13, KR13c, Mah14b]. Appl [Aba12, AFSMRL14, Asl10d, Ban14, BM12b, ESAH10, HLSS10, LL12e, MK10a, Pak11a, Rim13a, Tom15, Tua13, WZ13c, Web12, Wil10a, Xie12a]. Appl. [Mar14]. appliances [SW14a]. applicability [AHVR14, AH11d, AKH13, AG13, AH13a, AM14f, AG14f, MA14b]. Application [Ach13, AGHdD13, AY14, Bek10, BCW13, BBFK12, CX14a, CX14b, CK12b, EZRMR12, HOMA13, IJ11, mJ10, JLHW13, LTW10, Liu11c, LIC12, Man12b, MSJ14, NZW11, ÖA10, RSK11, SEL14, SF11b, Ste12b, SD13b, VDK13, WT11b, WW12, WaZW13, WSD12, WL10b, XWWC10, XG13, ZLC^{+}13, ZY14b, Zuo10, ZZ11c, AMIR14, Aha14, AAAH12, Aou11b, Asl12a, Asl12b, Bar10, BM13a, BMMS13, BCK14, BC10b, BGP^{+}13, CEM11, Cha11, CIVS14, CS10a, Che12b, CCJ14b, ĆSAV11, DSGB12, DK14a, ESG11, Ger12, GR11, GAV14, GML13, GC11d, Haj11, HK11, HZ12a, HjH14, HWPY14, Hu13c, Hwa11, Hwa14, ICKV14, Jum12, KRJ11, Kuj14, La 10, LKP12, LZ10a, LjH11, LTZ12, LSZ12, Li13, rL11, LLZ14b, LZX14, LD11, pLjH10, ML10b, Mar12, MGCFMM13, MO12a, MNP13, MJ13a, MKEG12, NA14a, OM11, PMC11, ROK10, Ray12, RM12b, Sag11]. application [SS13a, SS13b, SHBN13, SLZ12, Shi12, TPA10, Tom14, Van10, VDK10, WY13b, WW13b, Wei12, WY14, YH10, YWK11, YL11d, YCK13, YXZ10, YDZH14, DAV14a, FFOC14, HB11b, NA10a]. Applications [Dzi13, ET11, FY12, MJPB11, O’R11b, dOL13, ABK14a, AR14c, Alo13, ABGGS13, Ara14, AH11c, AH12a, AKÖ11, AS11c, BD14a, BZ10a, BAA13, BBI14, BCNP14a, BCNP14b, CGJ12, sCLCY11, sCLCaL11, CLCZ12, Cha14a, CK12a, CI11, DI13b, Den10, DLX13, EMS10b, EMA13b, EV14, EWS10, Fan14b, FMS11, FZ12, GGPT14, GT11b, GG14f, GZ10d, GZZ10, GG11b, Guo13a, GMSH13, Hao10, HjH12, HW12, HOB12, Hu13a, HXQ14, Hus10b, HP12, HD14b, HYT14b, IO12, Jar10, JHN11, KWS10, KT14, KS11d, Kyr14, yL14c, LZ10b, LL10b, LAF13, LDWY14, LH14a, Lin10, LZZ10, LG14, LC14d, LYLS14, LO12, LLN10, MMX10, MGC12, MMS– D13, NM14, ÖZ10, OS14, PMB12, hPrLgZ10, PH10a, RDA14, RPA14, SS10a, SI11a, Shk14, Son12, SY14d, Sri11b, SÇK14a, SC14b, Sta14b, SSV13, SS13d, TQ11, TFM13, TQWZ10, Vej13, VVK12]. applications [WZ10, WZG12, WHW13b, WT11d, XT14, XXS14, YP14, Yan12f, YG11, YWD13, YWCH14, ZTD10, ZM14a, Zha12c, ZZZM10, ZZH11, FMSGS12, GLSJCBS10, JZ13, KP12, MSCK11]. Applied [AKB13a, AZ19, BDG13, BCNP14a, CX14b, EM11, Gür12, JDJ12, Mos14, Pan12b, Pan15, PK18, PP15, Rho19, SQ11, TJ13, WHS14a, YZZP14a, ZZ15, AK10a, BB14c, Cha14b, DJ14b, FLMR14, FNVLPD^{+}13, GLSV10, GS10b, KR11, LNdCJV14, MVRVSCMV14, Rem10, RCSH10, SGY^{+}14, XK14, Zay11, dFGAN12]. Apply [SNH10]. Applying [FNV10, PEMS13, CCCRR13, HCLP11]. apportionment [Tsi11b]. appraisal [BPDZ12]. approach [AS12a, ABLZ13, AAT^{+}13, ACG13, AM14c, Aou11a, AA12d, Bar10, Bas13, BB14a, BB14b, Ben14a, Ber10, BSKL13, BS12c, Car12, CC12a, CN14a, CMP13, CGVR14, CZZ^{+}11, CLW12, CGLM^{+}12, CJRV12, CSSSV13, DE11, Ema13d, EYAD13, EH14b, FP13a, FBGH12, GWM10, GLX13, GXSH14, GGM^{+}13, GMR11, GD11, GTU13, GGÖS11, GÖ14, GK12b, GS14f, GJNN14, HN14a, Has14, HZ14b, HL12e, HY11b, HMH13, HL14c, Iç14, JM13, JYCY10, JPK12, Kan12a, Kat11, KHK^{+}11, KK11c, KSFG14, KUKU14, KS13b, KL14b, KS10b, KS10a, KS12a, KH14a, KBG11, Kos11b, KRC10, KSM14, KPLC11, LP10, Lee11, LD13a, LP14, LTL11, LLXZ12, LY13a, LSSZ14, LCL^{+}13a, LPP^{+}12, LLC11, LHS12, LTU13, Liu13b, LLT^{+}14, LPPG14, LdS10, LWX13, MR13a, MTSN14, MB12, Mak14, Mal12a, Man12a, Mao11a, MLGdSC11, MCR11, MIM^{+}14, MPSA14, MEBYG11]. approach [MT13, MT12b, NMM14, NZ12, NN19, OO12, Ort14b, OGO14, Oul11, OD11b, PM14a, Pal13, hPnZrL12, PKPL11, PS14b, RRSS14, Ryo10, SR14a, SS14b, SZG12, SKSK12, SA10a, SA10b, SGY^{+}14, Sid10, Sin10b, SWLP13, TP12, TPTW14, Tur14a, VTM14, Ver10, VL12, Wag10, Wan11b, XL11, XKH10, XD11b, YYL^{+}13, YH12, YH14e, YJCC14, YAAS10, Yun10a, YK11b, Yüz14, ZB11, bZjZS11, ZLSW13, ZLG^{+}14, ZL11c, ZM14c, Zhe14, ZW10d]. approaches [AAM13, CGGMG12, CGGMG13, GVT^{+}13, IC11, KS11b, KLPC12, LCCA11, ÖB11, TDT12, WHI12]. approximant [GT11b]. Approximants [BE13, KZ14c]. Approximate [AG10, Aja11, DK11a, DT14, EALS10, HMIZ13, HJW11, IS10a, IS12, Ji14b, JZW12, KAB11, MF14, MPB14, QS11, QS14, Ram10b, SMR11a, SGA13, SK10c, Wan10e, WCN11, ZH14, BAP^{+}14, BRM11a, CC10b, CZL11, DR11a, DVGY11, DMS12, EZ14a, GL12, GS13c, GZZ10, HH12, HMH13, KJ10, KMR14, Kim14b, KAY12, Kov10, LL10c, LLL13, MDR14, MO12b, PSSS14, Reu14a, SMH14, TS13, WZZ11, Yun10a, ZYC12, ZC13b, BC12d]. approximated [LWS14, OO10]. Approximating [DLL14, GSSS13, yLbLjH10, LG14, ÁD10, BCR10, CP12b, DHM14, Ji14b, KR12, LQ13, NN13, TKC12]. Approximation [CMGR14, sCLCY11, DSGB12, EMK13, GG10a, Gal10a, GG14a, GS14e, GRY14, JHP11, Jun11b, LGCX13, Mah14a, Mal13b, MRSS11, NA10b, Ola11, PK12, SMNN14, SI11b, Wal10, Yad14, YWW13, YTS12b, AE11, AQ11, AG14b, Agr14a, AHJ14, AG14d, AM12a, BPPS14, Bas12a, Bhr14, Bir13, BD14d, BH11, BLR10, BSM14, CSC10, CYZ14a, CCJT11, CRX12, CS13e, Cop14, CY12, CG14e, DMS12, DAV12a, DAV12b, DJ14a, DB11b, DMR10, EGOM10, EDD11, EJ11, FF11, FK12, Fer10b, Fer11, FG13a, GL13a, GL11a, GCM10, AA14a, GA11b, GY12, HH14, fHcXH11, Hu13b, Hu14, HYC^{+}10, HP12, ISG11b, IRS14, Izg11, JS14b, JM14b, JDD14, KCBH14, Kre13, Kuj14, KPW11, LCD12, Łen11, Len14b, LL12c, LR14a, jL12, LKS14, LLW11, LqTyCZ12, Man12b, Men11, Mil14a, MMKD14, MG11b, Mor10e, MV11, MKEG12]. approximation [MKSN13, MKK14, OM11, OO14, ÖD11a, PC14b, RPW14, ROA11, RRSPTR12, Sas12, DAV14b, SZW14, SLW12, TW14b, TK14b, VGA12, WCZ13, WYZ14, WFwC12, WL10c, XRBZ13, Xia14, XFW^{+}12, Xu14a, Yam14, YY13a, YH13, YCJ^{+}14, YD14, Zah11, ZLL10, ZLA10, ZCL11, Zha14b, ZM10a, ZM14b, ZCU14, ZWT12, ZCC13, ZR14]. Approximations [Gue14a, AH11a, AK13, Bog11, CC10b, CD12, Dar11, DR10b, DU13, DBBEE13, GG13a, Gat14, GKS12, JVK10, JMY13, LS12c, MM14b, MP11a, MP14a, PP12a, PP14b, PC11, Pre11a, RD11, RRM12, SFF11, SBDB13, ZCZ14, ZL12c]. aquaculture [BLR13b]. aqueous [FdOdSM14]. aquifer [CJM14]. arbitrary [BS10c, CZ14a, CRP13, CP14a, DWZ12, GM12, Gue14a, HMIZ13, Kri14, MT14, MRCD12, NSNEV14, RNZH13, TKL14, WQX^{+}10, Wu10a, ZCŠ10, Zha10b]. arbitrary-state [HMIZ13]. Archimedean [GGM^{+}14a]. architecture [JHN11]. architectures [ABD^{+}12, FTAB14]. Arclength [CRC14]. arcs [KZ14c, LqTyCZ12, NN13]. Arctic [NATQ14]. area [AK13, MSGN^{+}14, MRCD12, NCdCDM13, NA14b, Zak10, Zhu10e]. areas [BC14a]. argument [BCC11a, DS14b, Kum11, LSL10, Ste12c, Ste12l, Ste12f, Ste12g, Ste12m, Ste12r, Ste12n, Ste13o, Yus10]. arguments [DKS14, Fig12, HLX13, Jan14c, JMB12, LW14g, LYL11, Ma14a, Mor10b, NRL12, SY10b, SL10b, Sta14d, Ste13a]. arise [FAM11]. arising [Ara14, BLR13b, CCJ10, Cha14b, Che11g, Cor10, ELS13, HN14a, IN10, JSF14, KLBR13, KP13, KR11, Mez13, MO12b, MK10b, NN19, OP14b, PD12, VSV10, VV10, WKLW12, WXLW13]. arithmetic [CK13b, MO14, MPW12, Pet14, SY13a, Sin10b, MS11]. Arnold [ER13]. Arnoldi [HLW13, LJ13]. Arnoldi-type [HLW13]. array [BSS14, Son10]. arrays [CCJ14b, CCSM13]. arrest [BH12a]. arrhythmias [ZDH14]. arrival [Dim13a, JHZC13, WCB14]. arrowhead [LGZ14a]. arsenous [LY14c]. artery [ML11b, NAHH11]. articular [Gra13]. Artificial [aBS14, AAA13a, BDAD10, Boy10, Che10e, CSD12, fGyLJ11, GxFX^{+}12, GjLxF^{+}13, GHLF14, GRV13, LG10b, LWX13, WZ11g, XMA14, XZF11, YWYY12, ZK10b, BDDK14, MMCD12]. Artificial-free [aBS14]. ary [LWY14, WFZL12]. Asad [PK18]. ascent [VO12, VO14]. Asian [CLX13, GO10, HL11c, MPW12, SY14a, WWWJ14, XLY14]. aspects [CGM13, DT10, GS12c, GJGSNRÁ12, MH10b]. assembling [NW12]. assembly [RV13a, YH12]. assess [Iç14, SKIH12]. Assessing [GVMNS13]. assessment [CHK10, Che12a, DZ14, MMM12, MMM13, NMM14, NNG10]. asset [KLW14a, LL12b, MA11d, TQWZ10]. asset-liability [LL12b]. assignable [YCC^{+}12]. Assignment [WYL11, BH13, GM13a, KHD12, LY11a, LMS^{+}12, MB12, NH12, ÖBT10, WTH^{+}14, WYBT12, ZHC^{+}14]. assisted [DJ12]. Associated [CK12a, AAV10, AJR12, Aou10, AAS14a, BSS14, CS13b, CLR12, CB12b, CM13c, CKS10, CY11b, CBS12, Cvi12, DS11, KM12b, Kud12, Ma13a, MM14a, MCL14b, RY14c, Sch11a, SS10d, SB12b, Sri14, UOK10, WY12b, Wen12, XXS12, YL11f, dACSS12]. Associative [LLS11, HZW12, SAL12, TZS14]. assumed [CC11b]. assumptions [Lin11a]. astrophysical [LIC12]. asymmetric [MZX10, Son10, WHW13a, Waz12b]. asymptote [LCF^{+}14]. Asymptotic [AC10a, BW10, Che14a, CS11c, CSMT12, DRS12a, DRS12b, DV14, DK14c, Dum10, EV14, EZ13a, FD12a, FXL14, GGP^{+}13, GGXH12, GH10e, HV12b, JS13b, KM12b, KM14c, Kro11, KTS11, KH14b, LS13a, MZW14, MMMMB13, MH14b, MM13b, MO12b, PEP14, PP13b, SZ14b, Ste14b, Sun11a, SCK14b, WZ13a, WFZL14, Wei11a, Wu10c, Wu13, WZW14, YJ14a, ZZ11a, ZCY14, dlS10, Abe10, ABT^{+}12, BH11, CDMS14, CC14b, Che14f, Che14b, CO13, DC14, DZZ14, FO11, GCM10, HKI12, HLTT14, HLW12, HHX12, HX14, Ign13, IISA10, IJ11, Izg11, Kar13b, KKV12, Kot14, Kre13, LLXZ12, LSL10, LDL13, LML12, LB14c, LW10b, MPSD12, MX11, NAL12, Oni11, PRS12, PS13, RD11, RS13, RTV12, Sid10, Ste13k, TM11, TK14b, Vod13, VO12, WFL14b, Wu12d, YY10, ZW10d, ZYZ14c, PCT^{+}10]. asymptotic-numerical [KKV12]. Asymptotical [ABK^{+}14b, Men14, Yan13c, YL14a]. Asymptotically [Bac14, CYHL11, CHS14b, DI13a, HPB11, Ste13a, Bra12, sCLCY11, sCCL11, sCLCaL11, CWL^{+}12, CLCZ12, sCWT^{+}12, sCCYW13, CLC^{+}13, CWWC14, CM14, Che14d, CQ14, CMW10, DTP14, DS12b, DN12b, Gu10, KudKL14, LWZD14, PK12, QCK10a, SK11a, pYX10, YZK12, YK12, ZT12]. Asymptotics [HMPVQ14, Boy14, Plo14]. Asynchronous [Che10b, HX13, XC10c, XWC11, XX14b, ZWZH13]. atmosphere [MV13]. atmospheric [SS11a, SS10g]. atom [Dim13b]. atom-bond [Dim13b]. attachments [LL13g]. attacked [Wu12b]. attenuation [HYJ14, JLLD13]. attitude [XWWF10]. Attitudes [WHI12]. Attiya [CKS10, SKW11, YL11f]. attracting [Dia13]. attraction [Ard11, BBSC13, CD10b, GG12c, JHP11, MÇ12a, NSC12b, NCS14, ZLC^{+}13, ZL10e]. Attractivity [Bal14, BBJ10, CW14b, FP12, GC11c, JH10, LS12b, LZTZ13, ZX13]. Attractor [SS10e, EKZ11, GZW13, HLeH12, JZY12, SFP^{+}11, Yan11b, Yan12e]. Attractors [Dan10, DL13, GX13, Kan13, LLW10a, LY13b, MWC11, NSC12a, OGJLCCB14, QZS13, SL11a, SNC11, SW14d, WWQ14, Xia13, YWZ10, YWLX11, YLZ13, Zha14e]. attribute [AEESMA10, FZLZ13, GK11, HS14b, SS12a, SM14c]. attributes [HS14a, Koy12, MS13b]. AUGMECON2 [MF13]. augmentation [Che14c]. Augmented [Kal12, Lee11, CHX13, CSF12, HZ11, Hou13, JYL14, KPLC11, KPP^{+}13, MF13, PKPL11, PW10, Siu10]. Auto [IKU10, CGGZ11, PMB12, WT11c, BD12b, GGYW11]. Auto-Bäcklund [IKU10, GGYW11]. auto-correlations [PMB12]. auto-oscillators [CGGZ11]. autocatalytic [PT13]. autoignition [BGP^{+}13]. automata [D’A12, dRS11a, dRS13]. Automated [SBM12]. Automatic [LWS11b, MGC12, SL11a, Asa10a, NR14, ZHZ12, ZDH14]. automorphic [Dia13, LY14b, MB13a, Mop11, RSRA14]. autonomous [AIAD12, Bel11b, CDR13, CGGZ11, Das12, GZW13, JZ13, JV13, JK14c, Kan13, KLW14b, LM12, LD14, LO14, MAB14, MS14a, PB14a, QZS13, RGL10, RGL11, RHSC14, RSRA14, THP14, WF11]. autonomously [PBP14]. autopilot [HZ12a]. autoregressive [CW11c, HLZ14, ZWP13]. Auxiliary [Din11, GZ10c, AEESMA10, BCK14, CMM12, GK11, GKV12, HS14a, HS14b, Koy12, MS13b, SS12a, SM14c, VG14b, YK14a, gZrLfM10, ZJD12]. availability [ZL13a]. Avalanche [ZYZ14a]. Average [HHX11, AY14, Bog11, HLZ14, JMS13, Sha14c, WWWJ14, WP14, ZCZ^{+}13a]. averaged [HLeH12, Liu11g, Tom13]. Averaging [Med12, MTD10]. aversion [BO13]. avian [TGL14]. avian-human [TGL14]. aviation [SKIH12]. avoidance [HSS10]. avoiding [AS13b]. aware [DXC^{+}14]. awareness [Goy14, MG14b, SRS^{+}13, SC14a]. AXB [ZLGZ11]. Axial [Cha13b, KG14, Lac14]. axially [AS11b, AS14c, GYBR10]. axiom [KL12, Koj13]. axis [EAAN^{+}14, KZ11, KZ12]. Axisymmetric [RRP14, VPS12, AY11b]. Ayala [JV13, VJ11, WZW14].

B [AMIR14, ABJ10, AAM13, AS13a, AM12d, BJS14, GK12b, HWPY14, KS11e, Lee12, MMSZ14, MJ11a, MJ12, MB13b, MB14, Moh14, RV10b, SR14a, ZHY13, ZK10a]. B-spline [AMIR14, ABJ10, AAM13, AS13a, GK12b, KS11e, MMSZ14, MB13b, MB14, Moh14, SR14a, ZHY13, ZK10a]. B-spline-FD [HWPY14]. B-splines [BJS14, MJ11a, MJ12, RV10b]. B-type [AM12d]. background [UB13]. backlogging [AAKB13, BJSS14, DGS14c, San10, SGC12]. Bäcklund [BD12b, CC11a, GGYW11, GMST10, IKU10, LTJ^{+}11, MG12a, MH10b, QTL^{+}12, SHY11, Wan10b, WTL^{+}11, XGG^{+}11, ZlDxX11, ZHT13, ZZR14]. backordering [GKC11, HLL14, WHWC14]. backorders [WHWC14]. backstepping [THX14]. Backtracking [Civ13]. Backward [Lin11a, RSK14, HMD14, HR11, Hu12b, HD14a, HHDM13, JDJ11, JDJ12, dLfX14, MFZ12, Mop14, NTT10, PJFL14, QW11a, TSS^{+}14, TTQ11, Tua13, WaZW13, Xu13c, ZS12, Zha14a, ZLL14, ZXF14]. backward-facing [TSS^{+}14]. bacteria [JZY10, KHA12]. bacteria-fish [KHA12]. bacteria-immunity [JZY10]. bacterial [BDAD10, Hsi14, Sie14]. baffle [HMD14]. baffles [SK14a]. Baghel [Ban14]. Bagley [Ray12]. Baier [NSNEV14]. balance [CC12c, CZC11, EZRMR12, GXC^{+}11, GL10c, GM14c, KK13d, LJ12a, LCC12, ROK10, SGK14]. Balanced [Mat14b, CHF14, HZFZ14, HG13a, SHBN13, Shi14a, TK12, WHZ14, Yan12c]. balancing [BPT12, Ste14a]. balking [LWZ14, SLL14, WZ11c]. ball [Ste10d, Ste10j, Ste10g, SU10a, Ste10k, SU10b, Ste10l, Ste11b, Ste12v, Ste12w, ZTD10]. BAM [BC12b, BBI14, DZZ^{+}13, DZZ14, LPL^{+}13, LZZL10, SAMA11, WQ13, Yan12e, ZW10d, ZLY11c]. Banach [AG14f, BS10c, CDS13, sCCL11, CLCL11, sCCYW13, CWWC14, CCA11, CS13e, CS10c, CM12c, Din11, Din12, DHS10, Fer13, fHcXH11, HMSW14, HKK12, JW10a, KKS11, KAS12b, KD13, KPW11, LjH11, pLjH10, Mal12a, MD13a, Man11, MD12c, MD13c, NV12, NS11, PG11, Sad11, She14a, SK13a, SKC10, WFZL14, WY12c, XZ12, YZK12, bZlD11, ZX11b]. band [Che14g, SHSC13, WZY14, WLXL10]. band-limited [Che14g]. banded [Che14i, Fuy14a]. bandwidth [HLWH12, RMMM11, Rem10]. bang [qWH11]. bank [KUKU14]. bankrupt [LL12b]. bans [Goy14]. bar [B– D11, Kha14c]. Bare [LDW14]. Bare-bones [LDW14]. bargaining [KI10a]. Bargmann [YZZ10]. Barnes [FK13]. barrier [BP11a, Cho11b, KMH^{+}13, NN11a]. barycentric [Zha10c]. base [KYH10, PPMCG14]. based [Abb14, Abd13, AQ11, AR11, AKS14, AOTV12, AOTV14, AST12, AC14, Alz14, ASB10, AY14, AP10, Bel11a, BCDI14, BOR14a, BDDK14, aBS14, BMI^{+}11, Bur12, CC11b, ÇM13b, CHK10, Cha13b, Cha13a, CGVR14, CZZ^{+}11, Che11c, CLW12, CLD12, CSD12, CZX^{+}14a, CZX^{+}14b, CFX14, CWCW11, CRX12, CVP^{+}10, Cho11b, CBRTA11, CHYT12, CHT14, CM12c, CCSM13, CHMT11, DW14, DWY14, DM10a, DS13b, DK14b, DR11b, DSPZ13, Dub13, EL14, EYAD13, EVDG12, EH14b, FZLZ13, Fau14, FFOC14, FPP11b, FGD13, GS13b, GWX12, qGyFzX13, GXSH14, GWW14, Gen11, GRKP10, GCLL10, GS13c, GG14c, GGTCdlF13, GWML12, HLeH12, HS14b, HZ14b, HL10, HH13a, HB11b, HR13, HL13c, HL12e, HY11b, Hua11b, HW11, HM12, HL14c, HSC14, HM14b, IN10, Iç14, ISG11b, JM13, JSWJ13, JKW^{+}11, JLHW13, JFHZ14, JWW14, JMS13, JS14c, KBAvBW14]. based [Kal12, KB13, KUTY14, KM14b, KHK^{+}11, Kha14c, KASY10, KUKU14, KL14b, KAK11, KBYL14, KGB12, KH14a, LMZ13, LPAE12, LY10a, LG10b, LWZY11, LZLW11, LZ12b, LYPC12, LW14e, LL10d, LZWZ14, LPP^{+}12, LW11c, LS11a, LZGZ11, LZW13b, LLS^{+}13a, Liu13b, LYZG13, LP13a, LXZ12, LWX13, LLS13b, LD13c, LZ14d, LXLY14, Luo14, MMX10, MR13a, MC14b, MLGdSC11, MIM^{+}14, MKSC13, MO14, MP14d, MMM11, MK14b, MYL^{+}10b, MJ13b, NMM14, Nar12, NR14, NKM^{+}10, NCdCDM13, NDZ14, Pal13, Pan12a, PTK14, PR12a, PAL14, PRK13, dFPF14, PB14b, QTYS14, QHYW14, dejW14, QC12, RBMA12, Ram10b, RKA^{+}13, RP14b, RRSPTR12, RFK12, Rud12, RX14, RV10b, SGDK11, SE12, SIDR11, SKB14, SCC12, Sie11, SDN14, SF13, SHZ10, SW14b, SPKS12, SNZL13, SXGR11, SBDB13, Tad10, TZZX11, THDF11, TKLH10, TIT13, Tom12, Tom14, Tom15, TDT12, TN14, VC14a]. based [VMRA14, WLJ11, WzH10, WQX^{+}10, WY11a, WDL12a, WY13b, dWcFwWjZ13, WTH^{+}14, WC14b, WZH13, WZ11h, WS13, WWZ14b, XMA14, XYY12, XBC^{+}13, XBYF14, XLY10, YKI10, YD13, YH14a, YLD14, YCZ14, YTY^{+}14, Yin10, YG13, YWL13, Yua11, Zah11, Zei11, ZZQ10, ZX10, ZZYX12, ZLMG12, ZJD12, ZHH^{+}13, ZC13a, ZL14b, ZCZ14, ZBM10, ZHZ12, ZLZ14d, ZYLS12, ZS14c, ZZC13, ZL10f, ZnXwY14, ZZZL14, Zil12, dlSAQ11, AGM14]. bases [AJR11, DP13a, GP13, SS10e, ZTD10]. Bashforth [KMR14]. Basic [AGLR13, LWZ13b, GLSJCBS10, Par10c, PW14, SJL10, ZZ14c, ZZ15]. Basin [NSC12a, SNC11, GG12c, MG14a, ZLC^{+}13, ZL10e]. Basins [NSC12b, NCS14, Ard11, BBSC13, CD10b]. basis [AB12a, BW10, Boy13, Duk12, GA11a, HXY12, JM11, Ker13, KAY12, Ko13, KGO10, LY10a, LLSY11, jLLW12, MJ11a, OPD12, PR12a, Ped13, QHZ^{+}13, Sar12c, WQW^{+}10, WWC12, ZH13]. Baskakov [AGKK14, Ere11, GA11b, VGA12]. Bass [Juk13]. batch [Dim13a, KBG11, WCB14, YZZ^{+}14]. batch-arrival [WCB14]. batching [Liu11b]. bathtub [SSS14]. battery [SBM12]. Baxter [TZS14]. Bayesian [BTT13, KEH14, LCCA11, MH11a, PAL14, SSS13a]. Bazilevic [Den11b]. BBH [KPT11]. BBH-Kantorovich [KPT11]. BBM [ROK10, SGK11, JLM^{+}12, LL10h, Mam13, MK13a, SSF10, SL14c, THH10, TI10b, YM10c]. BBM-like [JLM^{+}12]. BBS [GY12]. BC [YLQ13]. BCs [ZZW13]. BDD [CHK10]. be [KNS14, Ste13i, Wan14d]. Beam [Moh14, Ahn12, ECG14, EAAN^{+}14, Fuy14b, Gal10b, JY12, KDS13, KSY12, Li10b, MBR10, RDFL13, TJZ14, WD14b]. beams [AGM14, AS11b, Cha13c, DW13b, GG13a, GYBR10, HCL^{+}13, MP10, MP12a, Ngu14, RA10a, WW13e, ZL10a]. bearings [CZX^{+}14b, WY10a]. Beavers [FHWZ12]. become [NNPH14]. bed [QS11]. Beddington [GC11c, Kha14b, LB14c, QLWW12, Zuo13]. Bedient [LLLS11]. Bee [MMCD12, AAA13a, BDDK14, CSD12, fGyLJ11, GxFX^{+}12, GjLxF^{+}13, Hor10, LWX13, XMA14, ZK10b]. Bees [ÖBT10, Tsa14]. Beesack [MAO10]. behaved [JSWJ13, SFP^{+}11]. behavior [ABT^{+}12, ABLZ13, Bas12b, BH11, CDMS14, CGVR14, CC12b, CTLD12, CSMT12, DV14, DGM^{+}14, DCL11, FD12a, FW12, HT10, HHX12, JZZ13, JS13b, JL10, Kha14b, Koh13, LZ11a, LX14c, LJX14, MMMMB13, Pan10, PRS12, PEP14, PPMCG14, Rao14, SKC11, Ste14b, SAAS14b, SCK14b, SHWX14, TM11, WFL14b, Wei11a, Wu10c, Wu12d, Wu13, XX10, YY10, ZL10a, ZJLO14, Zhi11]. Behavioral [LH14b]. behaviors [Bai14, DL12a, JSH10, KM12b, MCWC13, WZ13a, Zha12b, ZS14e]. Behaviour [Ste13b, dCDNCFRH10, GH10e, JS12, KG11, Kar13b, KM14c, Mal14, Ngu14, Svá11]. behaviours [HLW12]. beliefs [QGD14]. Bell [SGY^{+}14]. Bell-polynomial [SGY^{+}14]. Bellman [AY11a, WZ10, FZ12]. Bellman-type [FZ12]. belong [HHT13]. belonging [MMKD14, ROA11]. belt [ARH^{+}12]. BEM [KS13a, ZD10, ZhGQ14]. benchmark [ASSPA10, MATA14, SCK13]. benchmarks [AGNAL14]. bending [Li10b, Naj14, Nar12, VCNV13, YXZ10, Zen12]. benefit [KR13a]. benefits [SSCA13]. Benjamin [MA13c, SYZ10, SY10c, Esf11b, LTW12, ROK10, RLV10, S.14, SGK11]. Benjiamin [hXqXlC10]. Bergman [GS11c, JS10, SS14d, Ste10c, SU10b, Ste10m, Ste10l, SSB11a, SSB11b, Ste12v]. Bergman-Orlicz [SS14d]. Bernoulli [Moh14, AJR12, BTS12, CK14a, CK14b, EDG14, FNVLPD^{+}13, KP13, KS11f, Ngu14, Ozd11, SS10d, TLW11, TT13, Zho11c]. Bernstein [BCC11b, AAC11, AGK13, AKR10b, Ata11, Bez13, Bir13, CMGR14, CMG14, CLLS14, CG14d, Fin14, GG10a, GI14, Içö12, ISG11a, ISG11b, JDD14, LW11a, Man12b, MP14d, OÖ13, OO10, Ren14b, Sim11, SPP12, WYZ14, Woź12, Woź13, Yüz13, ZH13]. Bernstein-type [BCC11b]. Bers [LS10b]. Bertrand [CKK12, TÜ12]. Berwald [AMO^{+}10]. Besicovitch [GG12c]. Besov [GZW12, Ste10g]. Bessel [BDGP14b, BEdAdSK14, CPKT13, CA14, Che14f, Meh11, Meh13, MO12a, Mur13, OP13, Tom13, UY11, XH13a, YS13a, Zhu10c, Zhu11b]. Best [Bas12a, Mor10a, ABK14a, CSC10, CEP11, DI13b, EA13, PSTL10, SK14b, TIT13, WY13b, JS14b, SA10a, SA10b]. beta [FMS11, LMA14, Mat14a, WN12b, YD14, AAR14, GA10, GRY14, SSS13a]. Beta-binomial [SSS13a]. better [Bir13]. between [AJF13, Ard11, BBT10, BX12, BCF13, CFS11, CK12c, DL12a, DW13a, Elo14, Fer13, GGM14b, GSN12, JK13, KM14a, Koj14b, KBB12, LTG^{+}12, Liu13b, LXGM14, MD13a, Man11, MPPM12, MVS14b, MSC^{+}13, NN14, SAO12, Sha10b, Sin11a, Ste10d, Ste10g, SU10a, Ste11a, SS11f, Ste11e, SS11h, Ste12v, Sza13, TW13, TDT12, THX14, VSV10, Vis14, WC11, WLY^{+}12, WE13, YHL14, ZWL14a, Zho11c]. betweenness [AOTV14]. Bézier [BAA13, BB13b, BS11c, Bez13, CMD13, Den11a, DLL14, HHM10, Hu13b, HJW11, JK14a, JDLL14, KPT11, LW11a, LW11b, LWZ11, LDJZ13, Lu12d, QHZ^{+}13, QHYW14, RM11, SR14b, Woź12, Wu12a, YL11a, YHL14, ZTD10, ZH13]. Bezoutian [YC12]. BFGS [ZC12, ASB10, PM11b, VL12]. BFGS-trained [PM11b]. BGM [STV12]. Bi [DS14a, KHD12, MM14a, SS11d, WLYC13, Xu12a, AMA14a, Ema13d, JWZ^{+}12, LCH13, dejW14, Tsi11a, WY13a, XXS12, YHWZ11, AS12a]. Bi-CG [AS12a]. Bi-criteria [SS11d]. bi-criterion [LCH13]. bi-cubic [AMA14a]. bi-directional [JWZ^{+}12]. bi-Hamiltonian [YHWZ11]. Bi-integrable [MM14a]. Bi-level [WLYC13, Ema13d]. Bi-objective [KHD12, Xu12a, dejW14]. bi-proportional [Tsi11a]. Bi-seasonal [DS14a]. bi-univalent [XXS12]. bias [EGMPMRBC12]. BIBO [AGH12]. biclustering [Hua11a]. bicriteria [COPR10, EG06, XY11]. bicriterion [WL14a]. bidiagonal [NB11]. bidiagonalization [NY12]. bidirectional [HZW12, LD11, MG12a, SAL12, Waz10e, LLS11]. BIEM [CTLL11]. bifurcated [Par12, Zha12a]. bifurcating [GLM14]. Bifurcation [BY14, BZZ14, CLL10, CC12b, EDYA14, EEZ14, ES12d, GX11, JLM^{+}12, LSJ10, LZM^{+}12, xLhYlZ13, Liu11e, LCC12, LY14d, MH14a, Pan11, PWZZ13, Sha14b, Vil12, WZ12, XM10, ZJS10, ZYP13, ZLD11, ZZL14, ZHW10, AA12b, BPP12b, BH12b, BC14b, ÇM13a, CCK11, CJK12, DH13a, DWP14, DJY13, DLC10, DCL11, FP13c, FL11, GLM14, HL11a, JZY10, KHA12, KJ12, KB14, KP11b, LL13b, Li14b, LX11b, Ma12, MLY12, Mao12, Muk10, NW12, QL12, SZ14a, SHZ14, SHM13, TX14, WDZ13, WZ13b, WY10b, Wei10a, WC14c, XTL10, XH11, Yan13b, YJ14b, ZZC10, ZLJ11, ZZZ11, ZLY11a, ZCZS12, ZZ13a, ZT14]. Bifurcations [GZ10b, HHR13, JLZB14, LS13b, NP14, pSbL14, THH10, WCL14, XZT10, ZTW10, ZZ10d, ZS10b, AKMM12, ESNE^{+}14, Gin12a, Gin12b, LR14b, TRE13, WLL14, Wei14, WW13i, XH14, YTC13, ZZW11a, ZL10d, ZC14b, Zuo13]. big [HZZ14, LWN13]. biharmonic [EGMEM14, PQW10, Zit11]. Bilateral [LLSW12, PC11, LWL14]. bilevel [AC14, Eto11, JLHW13, JYF13, Mak14, WY12c]. Bilinear [bZjZS11, AM14a, AAB^{+}13, CvW10, Guo14, KWJ13, MZTT12, dYyR13, YWCH11, ZHT13, ZMG13, jZzFbL10]. bilingualism [BM12d]. billets [KMD13]. bimetric [BCK14]. bimodal [LK12]. bin [KMW12]. binarization [JM13]. Binary [BD12c, CMP13, CLW12, FYL13, HWL11, KUTY14, KI11a, MIM^{+}14, MZL07, Pak11a, SR10, TYCZ14, Yor14]. Binomial [GG13b, BS13, CM11, GTU13, Imo14, NOS10, SSS13a, Xia10c]. bio [ADD12, ZAY^{+}14]. bio-control [ADD12]. bio-inspired [ZAY^{+}14]. biochemical [DDWW12, Xu12a]. bioeconomic [PM14a]. bioethanol [VD14a]. Biogeography [KBYL14, GCLL10, IS11, LWZY11]. Biogeography-based [KBYL14, GCLL10]. bioheat [BB14a]. biologic [LX11a]. Biological [GLL13, NH12, AS13b, CC12b, CZL14, DISS14a, DK14b, MVRVSCMV14, SS13d, WTH^{+}14, ZZC10, ZCZS12]. biology [OP14b]. biomathematics [DS14e]. bioreactor [AVF10]. biosensor [RLW14]. biosketch [Udw10]. Biot [CLF13]. bipanconnected [STH10]. Bipanconnectivity [Sun12]. biparametric [HCL^{+}13, Kim10]. bipartite [LTT^{+}13]. bipolar [SBM12, ZL13b]. Birikh [AS14d]. Birkhoff [BNBV10, YNBV11, AB12a, MCS11]. Birkhoffian [KWM13]. Birnbaum [CL14e]. birth [HLM14, LTY11, MYJ10, NNG10, XLS13a, YJ14b]. Biswas [Kha11]. bisymmetric [WLY11]. bits [LMT11]. Bitsadze [AO12]. Bivariate [Cao10a, CG14e, LM11, TKS11, WQ11b, AAGR13, GKS10, HL11d, HW11, JK14a, LWN13, NOS10, PMB12, QFCZ13, SPA12, SGK14, WZ11a, YK11b]. BKP [KZJ13, MZ12, pZLC10]. Black [Cha11, LW13a, TM13, TKL14]. blade [LSF^{+}13, Neu12b]. blade-fluid [LSF^{+}13]. BLAS [ACQOR11, DFTQOB14]. Blasius [MH14b, XG13, Yun10a]. Blending [LR10, HW11]. Blind [RLW14, MO14, SNZL13]. Bloch [CFS11, CCLF13, Fer10c, LS10a, LS10b, Ste10a, Ste10j, Ste10g, SU10a, Ste10k, Ste10n, Ste11a, SS11e, SS11g, Ste11b, Ste11e, Ste12d, Ste12i, YM10a, Yan12d, YLTB12, Zhu10e]. Bloch-type [Ste10j, SU10a, Ste10k, Ste11a, SS11e, SS11g, Ste11e, Ste12d, Ste12i, YM10a, Yan12d]. Block [GK10a, Han14, LHL14, YHW13, ÁGVZ12, BZT11, BFB12, CL11b, CJZ12, sCCL11, CD14, DGM^{+}14, Fuy14a, GXSH14, GWLZ10, GD10b, HN10a, Jat12, Jia13a, JCY10, KZ14a, tLpOjCW10, Li11a, LYS12, LZ13b, LLBW11, LZFR10, LY12c, Lu14c, MC12b, MN14a, MHB14, MD12c, MD13c, RY10, SGZC13, SCC12, SEM11, SY14d, Wan14d, WL13c, XT14, wX13c, YH14a, Yan13d]. block-based [SCC12]. block-diagonally [LZ13b]. block-matrix [MC12b]. block-pulse [MHB14]. Block-transitive [Han14, LHL14]. blockability [KI11a]. blocks [LZFR10, YH14a]. blood [BSP11, KMAG12, LJ12b, MNP13, ML11b, NAHH11, RLW14]. bloom [CPC11, Mis11]. Blow [CYZ13, Lia11, LZ14c, ZH10a, ZYZ14c, Gal11a, LL10e, MN12, XP12, ZCF13]. Blow-up [CYZ13, Lia11, ZH10a, ZYZ14c, Gal11a, LL10e, MN12, ZCF13]. blowflies [WD13, Wan14a, ZZQ13]. blowing [Afz10]. Blowup [WZ14a, HW13b, SL14d]. BMOA [SS11g]. BO [BW14]. Board [Ano11, Ano12]. Bochner [ABBR12, KM14c]. body [ARL13, ARG14, CRP13, Cop14, CL14d, Jun12, PRT11]. Boehmians [AM12b, RN10b]. Bogdanov [XM12]. Bogoyavlenskii [MESAN13, Waz10d]. Bogoyavlensky [pXqLlZ10]. Boiti [WM10, qYqLtW10]. Boltzmann [CBO12, Che10c, FZP12, GLS13, ML10a, NKM^{+}10]. Bona [LTW12, MA13c, ROK10, RLV10, S.14, SGK11, SYZ10, SY10c]. Bond [Siu10, Dim13b, Tch11]. bonds [Tch11]. bones [LDW14]. Bonus [LWM10]. Boolean [CS13c, LS12a, LYC12]. Boosting [Chu11]. boosts [Koç11]. bordered [XQ11b]. Bose [HL14b]. both [BX12, CT11a, Mat12, Yan10d, Yan10e]. Bottom [CPC11, HLT14, YTCY12]. bottoms [NATQ14]. Boubaker [AYS13, MJ13a]. bound [AÇ13b, CLTL13, DMP13, AA14a, IS10a, IS12, JZ10, JDLL14, KPBC14, LXGM14, MZL07, Pak11a, Tan14]. Boundaries [BO13, CPS14, CTLL11, NN13, Oss10, SJR10]. Boundary [AHMS14, Ali12, AM14c, BRAS12, CZ11a, CW12, CXG10, Dol14, FZZ12, HAMA14, HMMK11, Jan11, NMIA11, PK18, AE11, AB11, AAV10, AS10, AV14, Ami14, AL10a, And12b, And13b, AY11b, jAzBS14, jAS14, AAS14b, ASKA13, AY11c, AO12, AMA10, Bai10, BZ11, BD11, BC13a, BU12, BP11a, BB14a, BKW11, BCK14, Bes10a, BCC11a, BRGP14, BD10, BAJK10, BRW13, BR13, BCNP14a, BCNP14b, CCMV12, CS12a, CT13a, CH14, Çak13, CLZmF11, CF12, Cha14b, CSLL10, CZZ^{+}11, CT12c, CLL14, Che10f, Che14k, CD11a, CRP13, CEXG12, Cor10, CJRV12, CRC14, CYZ13, DN14a, Dea13, DNI^{+}10, DV14, DU13, DX14a, Dri13, DW11, DR11c, DAA14, EJ12, EGOM10, ECG14, FHLW11, FZY10, FKS13, FNG11, FFB13, Fer12a, FFOC14, Fit12, FBGH12, Fra10, GS12b, GL13a, GX11, Gao14, GWS13, GWA14, GC11a]. boundary [Gen11, GQ14, GSH12, GDH10, GS12c, GYD14, Goo11, Goo13, GS11a, GG12c, GK12a, GKW14, Haj11, Han11, HLZ12, HZ14a, HL12a, HL12b, HL12c, HL14a, Her11, HHM14, HZ12b, HLX13, Hus10a, Ish10, ISG11b, IKL11, JYCY10, Jan13c, Jha13, JY12, JL14, JW10a, Jia13b, JLW13, JPM12, Kan12b, Kar13a, KK11c, KS13b, Khe13, KAK11, Kle12, Kos11b, KAB11, KSS12, LW11b, LSJ11, LLJ11, LZ12b, Lia11, LK14, rL11, LNC12, LLL13, Lin13c, LGN10, LNG12, LLW10b, Liu11a, LLC11, LL13e, LS12c, LP13d, MXH12, Maa12, Maa14, MNP13, MO12b, MT12b, Mol14, MAHAS14, MM12, NKM^{+}10, NSYA14, NRL12, NSZ^{+}14, Pal10, Pan15, Par12, PT14a, PC13b, Plo14, PP11, RB11a, RDKU14, Reu14b, RLV10, SK12a, SMI12, SJR10, Sha12b, She12b, SY12, SGA14, Sta12a, SSK12, SD13b, SS10g]. boundary [zSWZW12, TA11, TJZ13, Tun13, VMS10, Van10, VS13, WLH10, Wan10f, WLCW10, WCZ11, WCL11, WL11, WLW14, Waz11b, WRD13, Web10, Web12, WTFC13, WZ11g, WL10b, WZ11i, XL11, XWWC10, XLGW12, XCH14, XM10, XD11b, XZ12, XL14, YW11a, YZ12a, Yan14b, Yao10, Yig11, YCJ^{+}14, Zah11, ZH10a, ZL10b, Zha11a, ZLZ12, ZG12, ZLWW14, ZWS14, ZLG^{+}14, ZGZG10, ZSHZ11, ZM14c, ZhGQ14, Zhi11, ZH14, ZMZ11, ZLW10, Zit11, ZS14f, AGM12, MAN10a]. boundary-layer [AV14, FFOC14]. boundary-value [AMA10, DN14a, Haj11, LLC11, Plo14, Tun13]. Bounded [CFS11, Fer13, LZZZ10, Ste12c, Alo13, BM11, Bes10a, Cel14, DLGJ14, DM10b, HL13a, JT13, LZ12a, LKP12, LDL11, LN10, LZKK11, LG14, MZ13a, MDR14, NM08, PMP11, SMN11, San14, Sht11, SIV10, Ste10h, Ste10i, Ste12f, Ste12g, Ste12u, Ste13j, Ste13d, Ste14c, Ste14d, THYC10, Tse12, WZ13c, XZH13]. Boundedness [AOAA10, MZX10, Ste12d, ZZ14b, CZD^{+}14, CK13a, CDCS13, DI13a, Iri10a, LDL11, MD13a, SL11b, SXH11, WW11e, WL13d, ZLSW13, Zha13b, Zha13c]. Bounding [HH11]. Bounds [BS10a, CHM^{+}13, DM11, Hla13, HS11b, Neu13, SL14d, XFS12, And12b, ÁB14b, Aza14, BB13b, ÇDO11, CZM11, CDESS14, Che14b, Chi10, CKD12, CDDL13, Das11, DMGK12, Den11a, DHM14, GD10a, GÇ11b, GQ11, HW14b, HMZ14, LW14a, LH10a, Len14a, LDJZ13, LQY11, LJT13, MDÇ13, MP14b, MKSC13, Mor14a, PB13b, Ram11a, RTPH14, SATdV13, Seg14, SC10a, SB12a, TH14b, UT13, Vej13, WC11, wX13c, dFGAN12]. Boussinesq [AK10a, BCW13, CP14b, EASK10, Fer14, HV12b, HSZ14, LGS^{+}11, LDLL10, LDLM13, MM13a, MGWG11, ROK12, SS10f, WTL^{+}11, WW14c, Waz10f, Wen10b, XZT10, YZC13, ZnZ14]. Boussinesq-like [Fer14]. box [BGRS11, ESV12, PT11b, WZ14b, YLWH14, YHL08, ZJD12]. box-constrained [BGRS11, ESV12, PT11b]. Box-Jenkins-like [ZJD12]. Boyd [AKR11]. brachistochrone [FNVLPD^{+}13]. brachistochrone-like [FNVLPD^{+}13]. bracket [IB10c]. bracketing [Suh13]. braids [IN10]. brake [dCDNCFRH10]. branch [DPC10, MZL07, Pak11a]. branching [WL13b]. Bratu [Boy11]. breakage [KK13d]. breakdown [JS12, mL12b, SY14b]. breakdown-free [SY14b]. breakdowns [LWZ14, LK11]. Breaking [BRAB11, Pin10, Waz10d]. Breakout [BH13]. breakup [HBB11]. Breather [WDLM10a, WD14a, WqXlC11]. Breather-type [WDLM10a, WD14a]. breather-wave [WqXlC11]. breeding [PFG^{+}14]. Bregman [CWWC14, YPSW11, ZS14a]. Brézis [Bah13a]. bricks [dCDNRH12]. bridge [Cha14c, Kan13, LL13g, MWC11]. bridge-vehicle [Cha14c, LL13g]. brief [KG10, Lol12, Shi14a]. Bright [Lat14, TW10a, WYTL14, WD12, YY14b]. Brinkman [LL14e, LDL10, MPPM12]. Broer [GZ10d, Pin10, BZ10a, DGDA10, GZZ10, LLWW10, Wen10a]. Brownian [BW14, HRX14, ZHY11]. Broyden [ACK14]. brucellosis [LSW^{+}14a, SZ14d]. Bruggeman [ZCL11]. Bruno [DGS11]. Brusselator [Ma12, MH14a, MJ11b]. BS [MN14a]. bubble [SHZ10, YAAS10, ZYLS12]. bubbly [KS10c]. Buckling [DB11a, KAF13, EKSO14, JE11, WD14b]. budget [LY11a, LC12]. buffer [JU11]. Building [BM12d]. buildings [dCDNV^{+}13, SW14a]. buoyancy [AS14d]. buoyant [EASK10]. Burg [Mat12]. Burger [Bes10a, IKU10, KCC12, Liu10b, ZYZ12]. Burgers [Asl10d, FL10, Ibr14, Mos14, AS13a, GK14, GLS13, GC14, HHU12, JHA14, JMS13, MJ12, PVV11, RSN11, RB11b, SZS13, WLH12, ZW12, AHJ14, AHRV11, Asa10a, BTHA11, Boy10, CS11d, GZ10a, GS14a, GZXJD10, GM11b, Hay10, HHX11, IU10, IRS14, JPB10c, Kud12, LZLB14, LWD10, LH11c, MM13a, PZ12, PPV14, RAIM14, SSF10, S.14, Sal10c, SKD12, TW10a, Wan10c, WTL^{+}11, Waz10c, Waz13, XWZF11, XZ13c, Yad13, YH14c, YM10b, YCJ^{+}14, Zha13a, ZC14b, ZK10a]. Burgers’-type [HHU12, SKD12]. Burmeister [HM12, TC14b]. Burr [WHL14, ZANA12]. bursting [BZZ14, SHWT14b]. Burton [MAD14b]. bus [LXH13]. business [CFGRV13, YS11a]. buyer [SM13, TNB11, TCBL11, WLYC13]. BVMs [GM10]. BVP [GS13a, PV10, WZ10, WW13i]. BVPS [LLL11a, Goć14, OO12, Van11a]. by-catch [ZT11]. bypass [DGS14c].

C [LRM12, BB11b, YL14d, ZLGZ11, GCG14]. C-Algorithm [BB11b]. C/D [LRM12]. CA [CNTG12]. cable [CLLS14]. CAEV [PFG^{+}14]. CAGM [GXF13]. Cahn [SPK14]. calcium [NRWF12]. calculating [BR14a, LW14f]. Calculation [MÇ12a, ABF14, DHI13, dCDNVS11, Fer12b, HMH13, IB10c, KX10, OR14, SSV13, Zhe14]. calculations [HZ14a]. calculus [AA14b, Ara14, MT10, MT12a, UK11, XLY14]. Calibration [MKP13, FGRLSV14]. call [RA12a]. called [AGZ12]. Calogero [MESAN13]. Camassa [CQJ10, Den11d, IKV11, Lai10, LC14a, LZH10, LLR10, TW14a, VDK13, XZH13]. cameras [Cor11a]. campaign [KK14b]. Can [KNS14, ZWL14b, BBF12b, Ste13i]. cancer [KP10]. cannibalism [SJL10]. Canonical [BR12, ZGY14, HMR12, KM14c]. cantilever [MBR10, Ngu14]. cap [STV12, ZZHZ14]. capacitated [KZS12, SSV13]. Capacity [WYL11, DS14d, LHT14a, YWK11]. capillary [BB11a, Bar13, HLT14, KMH^{+}13, TH14a]. capital [ZR13, ZLL14]. Capturing [MVS14b]. Caputo [FT10, MT12a, YLAT14]. Carathéodory [RHSC14]. carbon [CTLD12, Dan14, Kia14, Liu13b, MV13, SRA10, SK10c]. cardinal [AB12a, BJS14, WLXL10]. cardinalities [CHM^{+}13]. cardiopulmonary [Jun12]. care [Liu14a]. Carlitz [Cao12, CS13b, LLLS13]. Carlo [ÁB14b, DJT10, Sie14, XLY14, ZZ14a, ZA14a]. Carreau [PK18, HAMA14, VSS13]. carrier [MMP^{+}12, Rem10]. cartesian [DTN14, Oss10, YRV10, ZZR11]. cartilage [Gra13]. cascade [HSW11, LSF^{+}13, NZ12]. cascaded [CL13a]. cascadic [lL13c]. case [ÁD10, BIT12a, DFG14, FP13c, GL10b, HLC13, IG11, JWG^{+}14, KZS12, Mah14a, MH12, Mat14a, NCdCDM13, OÖ13, WT11b, ZL14a, NKM^{+}10]. cases [MH13b, MOZ11, Set14]. CAT [CWL^{+}12]. Catalan [Sta11]. catalog [KVP14]. catalytic [QS11]. catastrophes [GNdC12, GNS14]. catch [ZT11]. Categorical [Kau12]. catfish [CTY11]. Cattaneo [FMS14]. Cauchy [AMA12b, BA11, BMR13, DL12b, Dri13, Kel12, LWW10, LF11c, LZW13b, LS14e, MG11a, Men12b, NAA13a, NAA13b, OÖ13, PKS10, Sch11a, Set14, SM10, SS11e, TTQ10, XX11]. Cauchy-type [PKS10]. causal [Jan11, Jan13b, WT11c]. cause [dOS13]. caused [Shi14a]. cavitating [GP11a]. cavities [LZ13a]. cavity [Abb14, HN10a, SK14a]. CDO [CBLC12]. Celerities [Aub10]. Cell [AY14, SDN14, BC12a, CCG10, JM10a, KMAG12, LKP12, PPC13, TK14a, WPKE14, Xu14c]. Cell-average [AY14]. Cell-centered [Xu14c]. Cell-like [SDN14]. cell-mediated [WPKE14]. cells [KMAG12, VDL12b, ZK14]. Cellular [D’A12, BC13b, GWZ14, HYJ14, RSPK13, YZW^{+}10, ZS13b, ZCY14, dRS11a, dRS13]. censored [SSS13a]. Center [DCL11, LW14c, CGRS12, DGOR13, DFG14, Gin12b, GMR13, LLW13, MÇ12a, RA12c, Tsa13b, WHF14, WZ12, XH11, ZLJ11, ZC14a, ZS14d]. center-Lipschitz [GMR13, RA12c]. centered [Xu14c]. Centers [GLV14, Gin12a, Lip13, WHF14]. Central [CL14d, SA10a, ARL13, DQL^{+}12, EDGC14, Fer12b, LP12a, LC14a, TK12, WG11, YGL11, For11]. central-field [Fer12b]. central-upwind [WG11]. centrality [AOTV14]. centralized [GLHCLP12]. centrally [Gue14b]. centrifugal [PHPD11]. centroaffine [YYL14]. centrosymmetric [LZDS12]. Certain [Cho11c, GGS13, NRAV14, WXA10, YL11b, Abe10, AOAA10, AC10a, Bes10a, Çak10, CGRS12, CS13e, Cor10, Cur14, FLZ14, HI14, Hus10b, IHM14, Irm11, KÖ11, Kov10, LZW13b, LOS13, MT10, MSC^{+}13, ÖA13, PAC13, RN10a, SI11a, Sek11, SYW13, SC12, SB12b, SÖCS12, SÇK14a, Sri14, Wan13c, WYK11, XXS12, XLS13b, XCS13, ZZ11a, ZCZP14]. CES [VV11b]. Cesáro [AR12a, Bor11, ÇT10b, DM12, Et13, LLLS11]. CG [AS12a, Mat12, yZxGyM10]. CH [FYYT11, JLZB14, yOZ14, XW10]. CH- [JLZB14]. chain [DJT10, ESAH10, FL11, HQ14, HLL14, LW06, LX14c, LB14d, MHTH12, MLY12, PSC12, SCN14, SAuAH14, SM13, SY10a, SYB13, TNB11, TC14a, VG14b, ZLD11]. chained [JZ13]. chains [DO11, MP11a, MP14a, TL10, TLL13, ZZ12a]. Challenges [URR11]. Chan [EDD11]. chance [Din13b, LQY10]. chance-constrained [LQY10]. change [DR11a, HT14, Naj14, WZ11i, YT13b]. changes [MMM12, SSM13, ZLC^{+}13]. changing [DH13a, FHLW11, KSM14, LLW10b, Sha14c, YP14, ZLZ12, ZTZ14, ZL10e]. channel [FHAL14, MSCK11, SD13a, TPD10, VSS13, YWK11, YHH11, YCZ14]. channels [BLK^{+}13]. Chaos [GPCCA14, LL10a, WJ12, ZWL14a, Ben14b, Cha13a, CCCRR13, Che11b, DHSZ10, EDYA14, ESNE^{+}14, EEZ14, EAE13, FdBCC14, GVW^{+}14, LSJ10, NP14, LM13a, Waz14, WT11d, YWYY12, YCZ14, YYW12, ZLD11, ZC14b, dOS13]. chaos-based [YCZ14]. Chaotic [Ala10, CTY11, DRSG14, Jan13a, Abd13, AA12a, AA12b, AFSMRL14, ADBT13, Bel11a, Bel12, CGTCF13, Cha13c, Che11a, CC12b, CC11c, CHC12, Che12b, dSCAM14, EDYSA12, EDYA14, fGyLJ11, GxFX^{+}12, GM12, HLeH12, HCC^{+}12, JJPW13, LZ11a, LLYH10, MTD10, MGCFMM13, MPB13, Pai14, PPMCG14, SL11a, TIT13, TGTCCGRG13, THX14, WJ12, WY10b, WS13, Yan10a, YWLX11, YS11b, YTCY12, Yan13a, YCZ14, ZX10, ZSL11, ZSWS12, ZLZ14b, ZZ14b, ZZ10b, ZLL11b, ZW14b, Zhe12, dCCL14, DAV14a]. Chaplygin [HW13b]. Chapman [ACG14]. character [Iri10a]. characteristic [AS14a, BU12, CB10, CZX^{+}14a, FR11, GBA14, HN14a, KI10b, KI12b, Koj13, Koj14b, Koj14a, NN19, WRZ12, XQ11b]. characteristic-mixed [FR11]. characteristics [AMM11, KML14, LCF^{+}14, SZS13, WZW12, WW13i, YY13a]. characteristics-mixed [YY13a]. Characterization [Fan14b, JS14c, PS13, YLMZ13, Zha14c, BM11, Das11, Jos13, LZ11b, MATA14, Pat13a, SGN13, YLJ14]. Characterizations [Den13, Ste11a, GMP14, HJMY14, MD12b, RT14]. characterizing [WWW13d]. charge [AAAH12, GGYW11]. charged [PM12]. charts [WW11h]. chatter [ZLZ14b]. chattering [AA12a]. chattering-free [AA12a]. cheap [KH14b]. Chebyshev [Pan12b, AR14a, ADÇY11, AMS14, AKR10b, BCDI14, BS12c, Boy12, CLR12, CTV13, EEEE05, Elo14, ELIGMB11, GKW14, GÖS11, JBL^{+}11, LHZ^{+}11, LLK14, MC14b, MCS11, PSSS14, PS12a, PC14b, RAIM14, fRbLJyS13, SSO11, WF12, YC12, Zha10c, ZX14]. Chebyshev-like [PC14b]. Chebyshev-tau [fRbLJyS13]. Chebyshev-type [LHZ^{+}11]. checking [CFN12, HS11b, Kot14]. chemical [BD14b, EM11]. Chemically [MU12, SS10g]. Chemically-reacting [MU12]. Chemo [DM13a]. Chemo-inspired [DM13a]. chemostat [Cai10b, DK14b, GZ11, QL12]. chemotaxis [MYT10]. chemotherapy [KP10]. Chen [YZZP14a, YZZP14b]. Chevalley [LHL14]. chi [Seg14, YK11b]. chi-square [Seg14, YK11b]. China [SW14b]. Chinese [YS11a]. Chip [LG10b]. chirp [LWL12]. CHKS [JLHW13]. Chlamydia} [SS14a]. Chlodowsky [SI11b]. Choice [Mur11, Mur13, BBC14, BBF12a, KL12, PD14]. choices [SSTM14]. cholera [NN14, MMP^{+}12]. Choleski [RSK14]. Cholesky [DKMM11, LYS10, LYS12, WM13a]. choose [GSN12]. Choosing [CNKS14, GH10a]. Choquet [JSWJ13, SSO11]. Choquet-like [SSO11]. Chree [Zuo10]. Christensen [ZYZ14a]. Christov [CP14b]. chromatograms [ANR14]. chromium [LNdCJV14]. Chrystal [Jor10]. Chua [MGCFMM13, DAV14a]. Chyan [EDD11]. Cigler [BB14b]. Cimmino [BRAS12]. CIP [Mat10]. CIP-method [Mat10]. circadian [FW10]. circle [AGLR13, BCM12, CGM13, HZ14c, MR14a, Sol12]. Circuit [ESNE^{+}14, RBMA12]. circuits [GGTCdlF13, TP12, TPTW14]. Circulant [GHZ^{+}14, BT12, Che14i, Fen13, Fuy11, Fuy14a, GMM^{+}14, HW14a, IB11, İpe11, KB11b, LG11, SC10a, SCH11b, YT13a, ZJ14c, fuy13]. circulant-like [fuy13]. circular [Che10c, CXG10, KM12b, KZ14c, LqTyCZ12, MM10b, MWY13, SMN11, SM14a, XWWC10, Yan11d, YS13b]. Ćirić [GS12a, VR12, GIR10]. Ćirić-type [VR12]. cities [Sun14]. claims [HY11a]. clamped [AS11b]. class [ALNGMGMM14, Ala14, AB12a, ACCT14, Bah13a, BZ10b, BMA11, BLR13b, BBK13, BRM11a, BWHQ14, Cha11, CYZ14a, Cha14b, CJ11a, CLW12, CTA13, CLW14, Che14e, CT11c, CPOdCN11, DP13b, DISS14b, fDxZxChT12, DHM14, DCL11, EDAM10, EMM^{+}12, FV14, GLY14, GB12, GLX13, GWS13, Gen11, GLV14, GK12a, GW12, HW14a, HOB12, HRAA12, Hou13, ISG11a, JDJ11, JDJ12, JSMA11, JY12, KÖ11, KX11, KBK13, Kau14, Kir11, KV14, KAB11, KL13b, LJKP11, LZ11b, LL12c, LZW13a, LW14c, LFR14, Li14b, LZL14b, LLL11a, LYAH13, LZ14b, LYD10, LTOX11, LL11c, LHL11, LXC12, Liu14c, LR14b, LP13b, LS13c, LXZ12, LZL12, LHTN13, MA10a, MS12a, MIHS14, Mil14a, MMKD14, Mis10, MRSS11, Mor13, MEL13, NCL11, NR13, NOS10, NWZ13, NZ12, Pan11, Par13, PPD10, QHL10, RSG14, RS11b, RZ12, ROA11, RD14, RMJM12, RX14, SHT14, SMNN14, SMN12]. class [SDAN14, SKSK12, SY13a, SYl14, SGZC13, SP13a, Sin10b, SPP12, SC12, SIV10, Ste11g, Ste12l, Ste12m, Ste12u, Ste13l, Ste13o, Ste14c, SHM12, SW12, TK14a, TFMB12, Tom12, Tom15, VV10, VSM14, VRMT14, Vil10, VG14b, VOL14, Wal10, WLCW10, WHW10, WY11b, WMH12, WY12a, WWW13c, WDT^{+}13, WZL14a, WZSW14, WLD11, Wu14a, XM11, XH14, XLS13b, XC14, XX14c, YK13a, YL10a, YL10b, Yan10b, YS11b, YCXL12, YKL12a, YH14b, YKL14, YLD14, Yin10, YHL08, ZOA10, yZ14a, ZH10a, ZS10a, ZZYX12, ZT12, ZZW^{+}12, ZR13, ZLZ14b, ZLWW14, ZWS14, ZQ14, Zha14d, ZH10b, ZZWY14, ZYZ14b, ZZY^{+}14a, ZM14d, ZCZ^{+}13a, ZZW11b, ZH13, ZWZH13]. class- [Li14b]. Classes [Fau14, AJPZ14, Çak13, CBS12, DN12b, DLLW13, EK10, Fre11, Gem11, GGS13, Hus10b, HA11, Iri10b, JS14c, Lu14a, Lu14b, Mal14, O’R10, RGL11, SB12b, Sri14, Ste14e, XZY14a, wX13c, YWW^{+}14, ZZ11a]. Classical [MM10c, Sho14, ASSPA10, AHDD13, BT10c, CNTG12, İpe11, KM12b, KS12a, LKY14, LCCA11, MDP12, MVQDE12, MM13a, NLLC10, RS12, Tra12, WW10, WDT^{+}13, ZJ14c]. Classification [CDMS14, Hu12a, SQJ11, ZYC12, BD14e, CC11a, yDpY14, Dub13, lF12, FAM11, GXSH14, HZ14c, KUTY14, MIM^{+}14, MK13a, NDZ14, PPV14, Qiu10, SSC11, VDL14, YWY^{+}13]. classifications [CV13]. classifier [CYW^{+}14, QTYS14, ZZF12]. classifiers [MIM^{+}14]. Clausen [AAS14a]. Clifford [Sch11a]. climate [GR11, HT14]. Clinical [GS12c]. cloaking [Kim14b]. cloaks [WQX^{+}10]. close [KOP11, Sek11, SYW13, UOK10]. close-to-convex [KOP11, Sek11, SYW13, UOK10]. Closed [CCJ14a, S.14, Tch11, Ban14, BRGP14, Bur12, DI13b, DBS12, JWW14, MRCD12, Ste13h, Ste13i, TJZ14, WWWJ14, WZL12]. Closed-form [Tch11, WZL12]. closedown [JS12]. closeness [AK10d, AK10c]. closer [AS12a]. closest [AP13]. Closure [BAR11, CDK13]. cloud [DXC^{+}14, KBYL14, Kos11a, LC12]. clouds [dALORPGC14]. cluster [Lip13, WXWY14, ZZW^{+}12]. Clustering [WZZ14, Hua11b, LWS11b, RBMA12, Sáf14, Yan11d, ZDH14]. clusters [CM10, SA10b]. CNT [HA12]. CNT-reinforced [HA12]. CNTs [HA12]. Co [BDDK14, ZDC12, BR12, BMH12, Car12, CW11a, GC12, lHyLfG10, Hua11a, WY13b]. co-coercive [lHyLfG10]. Co-design [ZDC12]. co-EP [BR12]. co-evolution [CW11a, WY13b]. co-evolutionary [Car12]. Co-evolving [BDDK14]. co-infection [BMH12, GC12]. co-movement [Hua11a]. coagulation [SP14d]. coalgebraic [CCG^{+}12a]. Coalition [KI12b, LIX14, KI10b, Koj13, Koj14b, Koj14a]. coalitions [KI11a, KI10a]. Coarse [Sie14]. Coarse-grained [Sie14]. coarsening [MGC12]. coating [CO13]. Coatmèlec [GMGV^{+}11]. coded [DHDB12, GCLL10, SAA14, TMJ14, YK13b]. codes [CHM^{+}13, Cen11]. Codimension [WW13i]. Codimension- [WW13i]. Coefficient [BB14c, ÇDO11, XLS13b, ZWP13, AS11a, BO13, CCJ14a, DE12, DYYL13, FZP12, GGYW11, GLZ10a, GMZL11, Guo14, HL12d, HHDM13, JZG13, KZ11, LWL10, LGS^{+}11, LLWW10, Liu10b, Liu11a, MTZ10, MGWG11, MMGM13, MK13a, Öca14, Ort14a, Ort14b, PPV14, Tad10, WLCW10, WS14, WTZ11, XXS12, XCS13, ZX11a, ZGY^{+}10, fuy13]. coefficients [AMB10, BS13, BA10, Bhr13, BCWM10, BC12d, BW10, Cha14a, CS13d, CSL14, Che14k, CM11, CP10, CTLD12, Cui14, DL12b, Den11b, DN12a, Ech12, EM13, GBA14, GTU13, Gün10, GZ10c, GB13, GG13b, HT11a, HB11a, Ibr14, JLL12, JS13a, JKB10, JK10, Kek11, KH14b, Lin11a, Liu11f, LLL10b, LYL11, MS13d, NSZ^{+}14, Nov14, PR12b, PKS10, RMJM12, Sal10b, Sha11a, SGK11, Ste11d, SXH11, Sza13, TW10a, TW11b, WL13a, Waz10g, Wei10b, Woź12, Woź13, Yan11a, YL10a, YH14d, YM10b, Yu13b, YY14b, ZA11, Zha13a, ZZW14, ZYC14, ZH14, ZWH14, Zou12]. coercive [lHyLfG10, NWZ13]. Coexistence [BKA12, BM12d]. Cohen [TJ13, DZZ14, HLTT14, JHP11, LZZL10, LX12, LX11b, MPSA14, SAMA11, SZ14b, TJ12, TJ14, WZ13a, ZLSW13]. Coherent [MPC14, MWJW10, Pan12a]. cohort [Fuk12]. Coiflet [SNMD13]. coil [KHK^{+}11]. Coincidence [EVR12, EVR13, O’R12b, BB12, Bor12, CIVS14, CMD13, – Dor13b, Rad14, RPA14, SR14b]. coinfection [PC14a]. cold [WZ11a, WX12, WW11g, Wu12b, YLS13]. Cole [SPV10]. collaborative [YWY^{+}13]. collection [ADGP10, DGCFM11, KZS12, Mit13]. collective [Sie14]. collinear [LZ10c, MWY13]. collision [ARG14, HSS10, VLMHK^{+}13, WL13b, WSL^{+}13]. collision-free [VLMHK^{+}13]. collisions [GRV13]. Collocation [ZLS12a, ZL14e, AMS14, Bhr13, Boy11, CC13, DP13a, FL13, FLZ14, GGP^{+}13, GK14, GHPR12, GÖS10, GK12b, JK14a, Kha14a, KS10b, KS10a, KS12a, LDWY14, LZW13b, MH13a, MHA12, MH14c, MJ12, MB13b, Moh14, PR12a, PT14a, QFT^{+}14, Rah13, SR14a, WZ11e, WW13c, YS13b, YS13a, ZHY13, ZC11]. collocation-path [AMS14]. Collusion [Koj14a]. colonies [BDDK14]. Colony [EdCVB^{+}14, ASS13, AAA13a, CMP13, CSD12, fGyLJ11, GxFX^{+}12, GjLxF^{+}13, KGB12, LKH12, LWX13, MMM12, RAZES13, SEED13, XMA14, ZK10b, ZYM11, BDDK14, MMCD12]. color [SNZL13, YWD13]. column [EÖ10, MKSC13]. columns [Tom14]. comb [GG13c]. combination [CX14a, CX14b, DV14, LC11, TCI12, GC11a]. Combinations [Gup13, WCL12, BM12a, CM11, ÖKU12, Tos14]. combinatorial [Sri11b]. Combined [Cui14, Dan14, SBS10, VSS13, BC11a, BAE13, DWQG12, Els12, ESE14, GVW^{+}14, GS10b, HZW13, Lu12b, MT12a, MEBYG11, MT12b, RCSH10, TW11b, WDLM10b, Waz10a, YM10b, YXZ10]. Combining [NLLC10, NDZ14, Men11, PZ14b, GYD14]. coming [TFMB12]. commensalism [Zha12a]. Comment [Aba12, AFSMRL14, EM11, Gün10, LLS14, Pan12b, Pan15, PK18, ZLMG12, Che12a, CIP10, Ma11a, NH12]. Comments [Ban14, BAFP12, CISM11, Ge14, Ibr14, Liu14d, Mos14, NR14, Rho19, SQ11, WDLM10b, WZ13c, Woo20, XYL10, XY11, Rho14b, Rho14a]. commerce [Sev11]. commodity [SS13c]. Common [AKR10a, ARN10, AN– D12, ANR12, ANR13, ARR12, ĆASH11, ĆSAV11, DV12b, DV13, HA11, HP12, PH10a, RRR11, SSRA12, AKN11, AKR11, Bas12a, BS10c, CIVS14, CPZ10, EVR13, GS11d, Hao10, JS14b, KR12, LWZ13a, Lin11b, LW11c, yLbLjH10, Liu14d, LZW12b, NKR12, Ola11, PK12, Rad14, TLDZ13, WLY11, WLH11, XQ11a, Yan12a, YK12, ZS14a, ZS15]. communicability [RV10a]. Communication [Ber14, Abd13, Che12b, Has14, HYZH14, HCC^{+}12, WZH13, YCZ14]. communications [DAV14a]. communities [NN14]. community [RV10a, pZJ12]. commutative [– Dor13b, ÖKU12, YLD14]. commutators [Den13]. commuting [BHP13, Tos14]. Compact [ACZ12, JS10, LL13e, Xia13, Zha13d, AHDD13, AS14e, BM11, BS12a, BBFK12, CC10b, Cui14, DV13, Gal10a, GG11a, GG14a, Gür10, Gür12, LZW12a, LNPS10, LHTN13, LWN14b, MWC11, Mah14a, MMS14, PN11b, SYK10, WZJ13, WW14b, XLGW12, ZFH14, ZW12, ZWT12]. compactification [PBG14]. Compactness [HH13b, Fan14b, Ste12d]. Compacton [XW10, GRV13]. compactons [CHX14, RV10b]. companies [CFGRV13]. companion [Alo13, BS12c, LG14]. company [Iç14]. comparative [GVT^{+}13, MAB14, PKK10, SGK14]. compare [KI11a]. Comparing [HD14b, Ste12a]. Comparison [BDH^{+}14, CK12c, Hou14, MPPM12, MLG13, RDA14, YZ12b, Yun11b, Ard11, Che13d, DC13a, dCDNCFRH10, HB14, KI10a, KI10b, KI12b, MAD11, Sal11b, SDK10, TDT12, Yos13]. comparisons [JY12]. compartmental [uiH14]. compatible [AKR10a, NKR12, Ola11, WY11a, Zei11]. compensator [LZWZ14]. competing [SVV13]. competition [ADD12, ALP11, CHC11, DRSG14, DX14b, GW13, GZ11, JV13, LDL13, LYB12, VJ11]. competition-predator [CHC11, LDL13]. competitive [Bal14, EYAD13, LTX11, LW14h, LF14, Sam10, TDSA14, WLQ^{+}10, YZ14c, Zha13b, ZZJ14]. complement [CN12, Li11a, LHZ10, LJCI13, SY13a, Sin10b, XQ12]. complementarity [Ach10a, qDG10, Fan10b, HM14a, JHX14, KM14b, LTL12, LW14e, LTM10, LLL10a, LHL11, LZ10f, LZ10g, Ném10b, NW10, Soo11a, TDZF14, XM11, YDZ11, ZZGL14b, ZxHW14, ZS14c]. complementarity-constrained [Soo11a]. Complementary [Vej13, CF12, DT10]. complements [tLpOjCW10]. Complete [GMP14, Jos13, LOKB12, RS12, YSZ12, Abe10, CLS12, GCMP14, KSD13, Kra14a, PDBRFER14, VDL14, WYL10b]. Completely [MCL14b, CZ14a, GQ12, Guo13a, Guo13b, Mor10b]. Completeness [Tun14]. completion [AA12c, DMP13, EG11, YW11b, ZB14]. Complex [Bai14, CS13d, CCGT13, EAE13, FXD12, Li13, LXC14, MSC^{+}13, SE12, ZLL11a, AQ11, AMA12b, AMW12, AY11b, AMB10, Bhr14, BZR10, ÇDO11, CL11c, CBRTA11, CC12c, CFG13, DWZW14, DW14, Dra14, EVE12, FWZW13, FS14, Gal10a, GG11a, GG14a, GWLZ10, GAK12, GARK12, GÇ11b, GL13c, GS14e, GG11c, GMV11, HRL14, Hla13, HYZH14, HZW14, HG13b, JJP^{+}12, JT13, JJL14, KNS14, KJW10, KC13b, Kud10c, LJPJ12, LPJ^{+}12, LWP12, LYJP13, LS13b, LDS14, LW14b, LW14f, LTJ^{+}11, LLJF12, LYL11, MG14a, MLGdSC11, MP14b, Mat14a, NAA13a, NWNL10, ONBN10, ÖA13, PKP^{+}12b, RV10a, RCH12, SWL12, SK10a, SPWZ12, SCLC10, SL13b, TP12, TPTW14, TZS14, TDSA14, TV13, Vir13, WW13d, WZWW13, WXWY14, yWZhL14, WJ11, WLH11, WLD11, WLY^{+}12, wX13b, XCS13, YYL^{+}13, YC14, YS10b, YS13b, Zha10a, ZXCL10, pZJ12, ZZ14b, MAN10a, SB13]. complex-network [YYL^{+}13]. Complex-valued [ZLL11a, FS14, HZW14]. complexities [Che11c]. complexiton [CYM13, HLZ13, LLS14, ZM14a]. complexitons [ZCG10, ZGY^{+}10]. Complexity [Ach10a, BB11b, GK13, CGAK10, CG14a, CHT14, DZ12, Goć14, IS14, IO12, LS11b, Rud13]. complexly [Sar12c]. compliant [Gad14, JHA14]. complicated [FWZW10, Ste12f, Ste12m, Ste12u]. Component [JYYJ10, BU12, BSR10, EGMPMRBC12, FB10, GZW13, GX13, JZY12, Sáf14, SSTM14, Waz10c, XZH13, ZLZ14d]. components [ABG14, BDH^{+}14, LL14a, LY11a, SLM14, WX12, ZAY^{+}14]. componentwise [Liu12c]. composed [Sab13, Sab14]. composite [AM14f, BDGP14a, CCS13, CP12b, Dub10, ITLA10, KDS13, Kuj14, LY12b, LCLW10, Moo12, MS11, NSYA14, RMN14, SXDJ11, SXCD11, THNTNXR11, WHC10, ZL10a]. Composition [CCLF13, Fer10c, Ste10a, Ste10b, Ste10c, SS11e, Yan11e, AS13b, JS10, LS10b, Man11, Ste10d, Ste10h, SU10b, Ste10m, Ste10l, Ste10n, Ste11a, SS11f, SSB11a, SS11g, SSB11b, SS11h, SS12c, YM10a, Yan12d, Zhu10e]. compound [BHJ13, HY11a, LGMZ12, Wan10c, YK11b, ZC14b]. comprehensive [CZX^{+}14a, GXF13, OYCCB13, YZ14b]. compressible [CFH^{+}13, WYL10a, WZ14a, YWW13, Zei11, ZCF13]. compression [AMM11, CVP^{+}10, LMT11, MG14c, SCC12]. compressions [Ada11]. compromise [MCR11]. Comput [Aba12, AFSMRL14, Asl10d, BM12b, ESAH10, HLSS10, LL12e, MK10a, Pak11a, Rim13a, Tom15, Tua13, WZ13c, Web12, Wil10a, Xie12a]. Comput. [Ban14, Mar14]. Computability [GZB12]. Computable [MR14b, PRV13]. Computation [AMFM12, Aou11b, AZ19, BDG13, BCNP14a, CCMV12, CCG12b, Che14g, CX14b, Din13a, EM11, Fuk14b, Gür12, HI14, JDJ12, KP12, LPAE12, LH10a, Mos14, Pan12b, Pan15, PK18, PKK10, PP15, Rho19, SQ11, TJ13, YZZP14a, ZZ15, BBB12, Bez13, Bot12, Che11c, Che14e, Dim13b, Fuk13, Fuk14a, GMST10, GP13, GC11d, KPP11, KK13b, La 10, LU12a, LWL10, LLL13, LG11, MSJ14, NR14, NAA13a, NH12, ÖA10, QQXY12, QFCZ13, QTL^{+}12, Sal10c, SIRD13, SSZ11, SBR^{+}12, SU12, SGA14, SXDJ11, SXC^{+}12, TES14, Tur14b, WSL^{+}13, WTH^{+}14, ZHZ12, AKB13a]. Computational [AMPT14, CCG12b, IO12, SSK10, SGC10, WHL14, AGHdD13, BLR13b, CGM13, CDLZ11, CHT14, GSF12, Gra13, GMR11, Hen11, IC11, KMAG12, MAHAS14, Pal13, Rud13, Sal11b, SS14b, Sie11, Sie14, Wah13, GJGSNRÁ12, LK11]. computationally [BJ11, CC11b, YLTB12]. Computations [WHS14a, ABK^{+}14b, CRW11, D’A12, DP13a, GT13, ISH12, KM12b, LWZD14, OGO14, WWZ13, dV12a]. compute [GH10d]. Computer [BP10b, DJ12, ESAH10, FGIV12, LW06, Ser13, Ama14b, AS12b, BP13, BX12, GYL^{+}13, HT10, LC12, MP11b, MVS14b, SSST14, WC14b, YY12, YYL^{+}13, ZYYZ12]. Computing [BS12c, GGM^{+}14a, LLW13, MCR11, MMSZ14, Moo12, PTW11, PW14, Sal10a, SI14, Sta12b, Van11b, Zhi12, BD14c, BAE13, BS11c, CHK10, CGVR14, CW11b, CY11a, Che14i, CDMT12, DGS11, DXC^{+}14, DWZ12, ELIGMB11, Fen13, GM14a, GM10, HL14b, JC14, Kel12, KSFG14, KBYL14, LL10c, xLhYlZ13, LWZZ14, LC12, Liu11a, MMS11, NY12, ONBN10, PTS11, SIDR11, SS10c, SS11c, SY14b, SSU13, SP13b, TS13, TN14, USAF14, WLZ12a, Woź12, Wu10a, YZ13, jZzFbL10, yZxGyM10, LM11]. comrade [JK14b]. concave [Ask14, FNV10, ZL10e]. concavity [SMBI14]. concentrated [TJZ14]. Concentration [CGJ12, RLW14, SWD11, VSS13, ZYZ14c]. concentrations [LPY11]. concept [AOTV12, ZZZ14, Zhi12]. concepts [FZLZ13, LWZD10b]. Concerning [Rho19, AR14a, CIP10, CISM11, Cvi10, DD13, GI14, Rho14b, Rho14a]. concrete [BSZ13, dCDNRH12]. Concurrent [Jea13, MKP13]. condensates [HL14b].

Condensed [SW11]. Condition

[dLfX14, ARR12, BT10c, ÇT10b, CIVS14, CH13, CT13b, CYZ13, DX14a, FKS13, FHWZ12, GO12, HR11, Hu12b, HMGC18, Ish10, JYCY10, KS10a, LS14a, Lia11, LLCX10, Liu12c, dLL13, Mol14, NS13b, PP10, PRK13, Rag12, RA12c, RRP14, SK12a, SSRA12, Ste13l, Tun13, VS13, WTFC13, XLLL14, Yig11, ZL10b, ZLWW14, ZM14c, ZS14d, LSLL14]. conditional [DD11, EGMPMRBC12, GVMNS13, PSWW14, PLTH12, Yan10c, Zho12a]. conditionally [MDR14, Zhi12]. conditioned [GTM13, PMC11, Sal11b]. conditioning [Che14c]. Conditions [ABR^{+}11, CMD13, AAB^{+}14, jASzZ11, jAS14, AH11e, AH12a, AY11c, BB11a, BB14a, BJS14, Bes10a, BAJK10, CH14, Çak13, CLZmF11, CF12, CD11a, CRP13, Dea13, DT14, DN14b, DU13, Dol14, EJ12, EÖ10, FNV10, FT11, GG14a, GL13a, GWA14, GSH12, GKC11, GS12c, Goo13, GS11a, GLB12, GMR13, HZ14a, HL14a, HHM14, HZ12b, Hus10a, ICKV14, IKL11, Jan13c, Jan14e, JLW11, JLW13, JPM12, Kar13b, Khe13, KS12a, Kos11b, LLW13, LAF13, LW14c, LLL13, Lin13c, LGN10, LNG12, LLC11, LL13e, MDD12, Mao11b, MM14b, MVQDE12, MVS14a, MM12, NKM^{+}10, NRL12, Par12, RB11a, RDKU14, RS11b, RHSC14, SR14b, SFL^{+}11, SMI12, SR12, SGA14, zSWZW12, Tag14, TTM12, TJZ13, VSS13, Ver12, Ver13, WLH10, WCL11, Wan13c, WPSW13, WWW13a, WCZ13, WDT^{+}13, Wan14d, WRD13, WYBT12, WTQU14, XL11, XH11]. conditions [XD11b, XL14, Yan14b, ZH10a, ZL10b, ZLJ11, ZG12, ZW13b, ZWS14, Zha14b, Zha10d, Zha13e, ZYZ14b, ZZJ14, ZL13e, ZS14f]. conducting [CL13a, JKP14b, VV11a]. conduction [AY11b, AMA10, DZLW13, HDK^{+}10, HWPY14, LY12b, LW13c, MFZ12, MSR14, QW11a, SZW14, WaZW13]. conductive [JYCY10, KCBH14, LCD12, SWL12, SdAT14, WZ14a]. conductive-radiative [KCBH14, SdAT14]. conductivity [HWPY14, LX11a, MWAM14, PC13a, TH14a]. cone [AKR10a, ARN10, ABR^{+}11, Alz14, AKR11, GIR10, HM14a, KM11a, LW14e, LLL10a, NKR12, RRR11, RRG11, THDF11, TDZF14, WYT13, ZF13a]. coneigen [JJL14]. coneigen-problem [JJL14]. confidence [PTW11]. Configuration [ESE13]. configurations [ARL13, CL14d, GMGV^{+}11, KI11b]. confined [CJM14]. conflict [XKH10]. confluent [LS14c, Res11, Res13]. conformal [MN14b, SMN11]. conforming [BY11b, DLS13]. congestion [ZX12]. conic [AR12b, JDLL14, ZZQ10, ZZC13]. conjecture [GI14, SvB10]. conjectures [Cvi10]. conjugate [And13a, CALR11, CWS10, DW12, Don12, FLMR14, HZW13, HL13d, HL14d, Iid11, LF11a, LWN14a, LL14f, LZX14, LP12b, LP13a, LWM11, LLZC11, MMKD14, Nar13, QLL12, RMJM12, SWX10a, VL12, Wag10, WZDW10, WLD11, WZL12, WLLD13, ZDZ10, ZL11b, ZJ13, WFDL11]. Connected [Mat14b, BGV10, Boz13, CLW11, DRS11b, NMIA11, SMN11, San14, Sht11]. Connectedness [GJ14]. connection [Bel12, BS11a, LWZ11, MVS13, MSC^{+}13, Sha10b, Sta14c, Sza13, Woź13, ZGH10]. Connections [BX12, CTLL11, HLS^{+}14, Koj14b]. connectivity [Che11f, Dim13b, LMY12, Ou10, SSG13, Yan12c, ZZR11, Zhe14]. conquer [Wu12c]. consecutive [Ery14, WE13]. consecutive- [Ery14]. consensus [DdGL10, Sha14c, XLLL14, ZL14d, ZPZ14]. consequence [CPC11]. consequences [Irm11]. Conservation [GB12, Ibr14, Shi12, WX13a, Zha13a, ZYC14, AK12, AMW12, AP10, AZAK10, BK10d, BK10c, BK10b, CL14a, CG12a, CRW11, Dub10, Guo14, JA14, KS10a, LP12a, LCH10, MK10b, PTY12, PTY13, RYÖ11, SLZ12, TF10, WG11, Wei12, Wen12, WHG13, XLZ10, qYqLtW10, YL14d, ZB13, Zha11a]. conservatism [AKC12]. Conservative [YL14b, Zhu10d, DRGR10, Dub10, FBGH12, Ge14, GL10c, HRAA12, Lia14, Nie13, SXL13, WGZ10, WRZ12, ZCZ14]. conserved [lQZWC14]. conserving [Ito13]. consideration [ABKG^{+}13, DWQG12, KS10c, MG14a, Nar12, Yan10e]. considering [EAAN^{+}14, HCCT14, KS13a, WHWC14]. consistency [BDGIPG12, BCF13]. consistent [Sun10, WX13a, WLHZ10]. consolidation [CLF13, MNOB10]. constancy [ZWP13]. Constant [LP13d, AANA10, BE13, CI13, CLTY13, CD12, CM13c, Che10f, CLS12, CP10, DL12b, DWP10, DWZ12, DL11, EM13, ECS13, GI13, GG11c, GG13b, JS13a, JS13b, KSM11, LLW13, LSL10, LY14c, Lu14a, Lu14b, LYL11, Mor10c, NRL12, Ort14a, Ort14b, PKS10, Rim13a, Rim13c, SGC10, THX14, Vít14, Wan13d, Wan14f, Yan13c, ZCT11]. constants [ABK14a, CC14b, Che14b, DWZ12, GGM^{+}13, LLXZ12, NN11b, WW13j]. Constitution [IG11]. Constitutive [BA11, MZWZ12]. Constrained [Hu14, BGRS11, CBTGWW14, CKP^{+}13, CSF12, DS13c, ESV12, ES12a, ESE13, ESE14, GPP^{+}10, Gu13, GS14b, Has14, HZ11, HCW11, Hsi14, JHX14, bJLyZlL14, JW10b, JYL14, KB13, KK12a, KcZ14, LW11a, LQY10, LL13a, jL12, LZZP10, LY12a, LWH14, LLCX10, MMCD12, PZ14b, PZ11, PT11b, QW14, Soo11a, SA12b, Tan14, TMJ14, WZZH11, WMH12, WY13b, WZ14b, Woź12, XZY14b, YKI10, YCTD11, YLWH14, ZHW11, ZP13, XBC^{+}13]. constraint [aBS14, Din13b, LXWW14, LLZ14c, LR13c, MF13, MSS14b, PX13, TNB11, VFGP13, XX14b, YLMZ13, YZZ10, ZZ10a, bZjZS11, ZZ15, ZCD10]. Constraints [GP11a, Ach13, AWS10, BPT12, BDK^{+}14, DS14d, DI13b, GLV10, HHA14, KFM11a, LW11b, LQY10, LTL12, LHT14a, MAN10a, NL12b, QW11b, XSX12, XBC^{+}14, YCXL12, YDZ11, YCJ^{+}14, YHL08, ZCŠ10, ZZ14c, ZL13e]. construct [GA11a, WLZ11, dYyR13, Zha10e]. constructed [EDD11, GG14d, MKSN13, YHWZ11, ZLX12]. Constructing [Hay10, ZL10g, FWZW10, ML12a, MV14, RV14]. Construction [AG14b, BSS12, BS10c, CHS14a, FG13b, KBAvBW14, NP10b, NL12a, Pan12a, QHYW14, SF11a, Sim11, TS14, WT11a, Wen10a, CGAK10, DB12a, ES12c, HL13c, ML11a, OCN11, PRT11, ZLX12, DEP12]. constructive [DW14]. consumption [CC12a, RZHH13, RZHH14, SW14a]. consumption/investment [CC12a]. contact [CLW11, Cop11, Cop14, DU13, NTS11b, RSBK14, Sam11]. contain [VD14b]. contained [YS11b]. container [DJY13]. containing [Bry11, CT11a, CTLL11, DG13, DG14a, KG14, MVS13, Mur13]. contaminant [KMH^{+}13]. content [DM10a, Hen11, RS10]. context [Fal14, SMR11b]. contingent [LSZ12]. continuant [Sah11a]. continuation [AB11, CLW12, CFZ14b, CRC14, FK13, She12c, TDZF14, VLMHK^{+}13]. Continued [Che13b, IT12, DBG11, WQ11b]. continuity [BIT12a, ÇÇD10, LL14d, TW14a, ZHGP14]. Continuous [BFM13, EWS10, LL14e, Sha14c, SVV13, XT14, YL13, AGZ12, BC11a, BCM12, CEM11, CC12a, CD11b, DRGR10, EGO13, FKS13, Ge14, HAA^{+}14, HLC13, HMGC18, HD14b, ICS^{+}13, Jat12, KJ10, KA13a, KcZ14, LFLS14, LR11, LWZD10b, LTL11, LLC12, LDS14, LR14a, LKH12, LL11c, LDL10, Ma11c, MS14b, MA13d, MSS14b, Nak13, ÖP12b, PSTL10, PG11, RY14a, Rec14, RA10b, SEED13, SL10b, Ste12l, Ste12h, Ste12n, Ste12u, Ste13o, Ste14d, Tsa13a, UT13, Üna13, VTM14, VC12, XDZ14, YSB11, ZS14d, ZYM11, ZXZ11, ZWH14, Zuf12]. continuous-discontinuous [BC11a]. continuous-discrete [CC12a]. continuous-review [MSS14b]. Continuous-time [EWS10, Sha14c, SVV13, HLC13, LWZD10b, LLC12, Nak13, Rec14, Tsa13a, Zuf12]. continuously [Li10a, LFC^{+}12, Ste13j]. continuum [GAC12, JWZ^{+}12, MCR14, ZMD^{+}14, ZM14c]. continuum-kinetic [ZM14c]. contour [DAHAK10, Fan14c, MRCD12, WH13]. contract [HLL14]. contracting [SZZ^{+}11]. contraction [ANR13, CR14, DWY14, GxFX^{+}12, MZWZ12, RRG11, SYl14, ZC13a]. Contractions [Rho19, ĆSAV11, DHS10, EVR12, GIR10, LLK13, MAK13, DFTQOB14, Rho14b, Rho14a, VR12]. contractive [ANR12, ARR12, BB12, Bor12, sCCLY10, CLC^{+}13, CIVS14, ĆASH11, CMM12, DI13b,

– Dor13b, GIR10, GS12a, GFMP13, ICKV14, Jun10b, Jun11a, LAR13, SSRA12, pYmK13]. contrast [PP13b]. Contrasts [BBSC13, LPPG14]. Contributions [Ram10a]. Control [B– D11, DNSM11, MB10, RV13b, ZCŠ10, Zhu10a, AS14b, AVF10, ADD12, AABV12, ALP11, ADBT13, AGH12, BYM14, BBI13b, BTV12, BLR13b, Bot12, BH11, BWHQ14, CCK11, Che10e, Che11b, CkPqL12, Che13c, CS14a, CL14c, DdGL10, DT14, DISS14a, Dim13a, DK14b, DKK^{+}13, DS13c, DHSZ10, EDYA14, EAE13, Fan14e, FNV10, FGFRV11, FB10, FR11, Gal10c, GS13b, GWW14, GHG14, GLL13, GC10b, Gra13, GS14b, GC10c, GFZ13, HFS12, HL10, HX13, Ilt11, JZ13, JMB12, JDN14, JKW^{+}11, JQ14, JXK14, Kam14b, KK14b, KC12, KWJ13, Kes13, KLKA14, KH14a, KP10, KSY12, KH14b, KLY12, LJPJ12, LPJ^{+}12, LKP12, LWP12, LYJP13, LL14b, LZLF10, LS12a, LL12d, LL12b, LXC14, LXWW14, LZL14a, LSSZ14, LLYH10, LZG10, LLZ14c, LS14b, LD11, LR13c, LTR11, LYB12, LZ14d, LH14c, LF14, MYJ10, MLH11, MVRVSCMV14, MS10, MG14b, MM13c]. control [NP10a, NLLC10, NSNEV14, NZ12, Pai14, PZ14a, PLF13, PWZZ13, PM14b, RA10a, RZ12, RM12b, SMN12, SSM13, SZH13, SZ14c, SIV10, SSK12, SD14, SWLP13, SL13b, TPT12, Tsa13a, WF10, WHW10, WY11b, WZ11d, WiN11, WJ12, WY12a, WiN12, WL13a, Wan13c, WY13b, WDZ13, Wan13e, WLLY14, WLXX14, WY10b, WTZ11, WC14c, WZH13, Wil11, WT11d, WW11h, Wu12a, XC10c, XWMW12, XSX12, XX14b, Xu14a, YTS12a, Yan10a, Yan13a, YS10b, YCJ^{+}14, ZZ14a, ZLMG12, ZX12, ZLM12, ZLS12b, ZZ13a, ZW13b, ZLZ14b, ZL14d, ZLD11, ZW14b, ZCD10, ZDC12, ZYG14, ZYYZ12, Zim14, dlSAQ11]. Controllability [Ala14, JLW11, LS12a, Mop14, SMN12, VSM14, VRMT14, ZZWY14, APJ14, CFCLM13, DT14, Ji14b, LSL^{+}14, LO13, LR13c, MF14, SMR11a, SGA13, Shk14, YSZ12]. controllable [HH13b, TL14]. controlled [MD10, PPMCG14, ZZWY14]. controller [AA12a, DGH11, FP13b, HZ14b, HSD11, JLLD13, KEGH13, KPBC14, LP10, LXK13, MC13, RKA^{+}13, WS12, XPLZ14, ZX10]. controllers [Che13c, DGS14b, MZXZ10, PQRR10, SHWT14b, Zhu10b]. Controlling [MB11, Zhu10b, BYM14, MMP^{+}12, Wan13b]. controls [DN14b, Mur14, Shk14, YL11e]. Convection [MA13b, ACZ12, ABE^{+}12, AC10b, AS14d, ANS14, Bac14, CLZmF11, CB10, Che11d, CFX14, Che10f, Cui14, DM13b, DNI^{+}10, EM13, FFOC14, FR11, Gal11d, GLR12, HJVC14, Har14, Hay11, HMD14, HCCT14, KM11a, KRMM14, KKV12, MG10a, Mah13, MMGM13, MMS14, Pal10, Par12, SK14a, SBS10, SCC11, SGC10, SHI13b, SW10a, SHZ10, SBDB13, VZ11, WT11b, YS13a, Zha11c, ZYC12, ZX14, ZXZ11]. convection-diffusion [Bac14, CLZmF11, CFX14, Cui14, EM13, FR11, MMS14, SBDB13, WT11b, YS13a, Zha11c]. convection-diffusion-reaction [DM13b, Hay11]. convection-dominated [CFX14, HJVC14, SHZ10]. convection/diffusion [ACZ12]. Convective [VV11a, BB14a, BT10c, EM11, Ish10, KS13d, PC13a, VV10]. convectively [Das13, HAMA14, PK18]. Convenient [Dua11a, BRM11b]. conventional [CAC14]. converge [BCM12, LWX13]. converge-onlookers [LWX13]. Convergence [Afu12, AH11c, AJPZ14, BCÁV14, CCA11, DQL^{+}12, DHZ11, qDG10, GGFM14, GA10, GS11d, HSW12, HQ14, HMGC18, KAS12b, KudKL14, yL14c, LTL11, Li12b, LSL10, LDL10, LZ11c, LJ13, MM11a, OS11, QHL10, RA10b, Ren14b, RWkPk10, Sad11, SK11a, Sau11, She14a, SS13d, WYL10a, WWLL11, WH12, WMH12, Wan13c, WZJ13, WHZ14, WD14c, XZF11, Yan12a, YK12, ZOS10, Zha11a, ZR13, ZZGL14a, ZS15, ZCC13, ZS14d, ZWH14, dV12a, AT12, AK11, AH11e, AH12a, AH12b, AH13b, ACK14, AM14e, AHM14, AZAK10, BM13a, BCDI14, BTV12, BD14d, Bog11, Buo10, CL11c, sCYW10, sCLCaL11, CWL^{+}12, sCWT^{+}12, sCCYW13, sCWW^{+}14, CWWC14, CC14a, CD10a, CQ14, CMW10, CG10, CKP^{+}13, CT11c, Deo12, DTP14, DRS12a, DRS12b, DL10, DHS10, Don14, EMK13, GGP^{+}13, Gho14, Gu10, GD11, GMR13, Hou13, HMH13, lJCWpW12, Jun10b, gKW12, KPT11]. convergence [Kar14, KM14b, KW10, Kim14a, Koç13, KK13d, LjH11, Li14a, LZ14b, LZZ10, LZGZ11, LW12b, LM13b, LWM11, LLZC11, Lu14a, Mag14b, MD13a, Mor10c, MKEG12, NS11, Odi10, Pat13a, PRK13, PMC11, PPD10, PC14b, QLL12, RWB10, RA10c, RA12b, RA12c, RMJM12, Sav10, Sav11, She12a, She14b, SX14, SWX10a, SW14c, SA12b, Su13, TT13, Van11b, WHC10, WL11, WZZH11, WWG13, WYW14, WZ13d, Xia10c, pYX10, bOYqGL^{+}14, YLD14, Zha13b, ZZGL14b, ZLL14, ZY13c]. Convergences [Wan10a]. convergent [CW11b, CG14c, JHX14, bJLyZlL14, LP12b, Lu14b, ONBN10, RK12, Ste13a, ZC12, Zhi12, ZYM11, ZXZ11]. converging [EBR13, Ste12r, SDS14]. converse [BPP12a, GCM10]. Conversion [YHL14, LWL12]. converters [BSKL13]. Convex [PP14b, ZTY^{+}10, AG14a, AG14b, AG14d, Ant12, AM14f, AKÖ11, BAR11, BDK^{+}14, CHX13, sCYW10, CKP^{+}13, Dra11, DRS11b, Eto11, FGS14, GWM10, Gue14a, GG11b, HXQ14, Hwa11, HD14b, Hwa14, Iid11, IW14, ISÖ14, JKW^{+}11, JFHZ14, KOP11, KI12a, KudKL14, KZ11, KZ12, KKVCS12, Kwo11, LP14, LL14d, yLbLjH10, Mat14d, MEBYG11, OP12a, Ota14, PS10, RDA14, Sek11, SYW13, UOK10, WYT13, YK12, ZC12, ZL10e]. Convexity [AR12b, TYCZ14, ADT13, GG14f, Gue14b, ZX11b]. conveying [DWQG12, NZW11]. convolution [AR14c, CRX12, CT11b, Dar13, DW13b, GAC12, LOS13, LM11, SPT12, SB12b, WD14b, XXS14, ZLS12a]. convolutional [Bho10]. Convolutions [CALR11, BDN^{+}14]. convolved [Ram14]. Conway [Imo14]. cooperation [RB11a, XMWD14, YL11e]. Cooperative [RK14b, Far14b, VDL12a]. coordinate [TLDZ13]. coordinated [Liu11b]. coordinates [DTN14, DG14b, LNG12, YTY^{+}14]. Coordination [HSS10]. coprime [TL10, TLL13]. copula [CBLC12, ZL11c]. copulas [dACSS12]. Core [RDD14, BT14, Mal13a, QS14, UJH14, Zen12]. corneal [OP13]. Cornish [WN12a]. corona [AAAH12]. correct [SDL10b]. correcting [KHK^{+}11]. correction [Chu11, FZ13, JFHZ14, MMS– D13, Wan10d, Jun10a, dV12a]. corrector [CTV12, DGSB14, Gal11e, HZW13, LLL10a, LW13d, YLD14, ZW12, ZZ11b]. correlate [OSAG13]. correlated [CÁGGLP14, HCLP11, HLW12, WLP^{+}14]. correlation [CCX11, CBLC12, Gro13, YZ13, ZL11c]. correlations [PMB12]. corresponding [DWZ12, Xu10b]. corrige [RZHH14]. Corrigendum [AKB13a, BM12b, BDG13, BCNP14a, CX14b, JDJ12, Kra14a, Len14b, LL12e, NC14, PP15, Rim13a, Sab14, TJ13, Tom15, Tua13, Wil10a, WHS14a, Xie12a, YZZP14a, ZZ15]. cosine [AG14e, BTVC10b, mJ10, LL11a, SIRD13, ZA11]. Cost [KR13a, AK10b, Ask14, DA11a, DG10a, HZ14d, KCGH14, KWJ13, KMC12, LJPJ12, LPJ^{+}12, LZ14d, SM13, SLJ^{+}13, TPT12, XSX12]. cost-side [DG10a]. costs [COPR10, HT14, KKR11, LW13a, Lin11c, Lin11b, WHWC14]. cotangent [Ras14, RY14b]. Cotes [Bur12]. Coulomb [Woo20, Ahn12, IS10a, IS10b, IHF13]. Coulomb-like [Woo20, IS10a, IS10b, IHF13]. countable [BS10c, sCLCaL11, sCWT^{+}12]. countably [ZOS10]. counterpart [AM12d]. counterparts [SMM14]. counterparty [LWL14]. country [ZL11c]. Couple [ZPZ14, Sin11a]. Couple-group [ZPZ14]. Coupled [AKN11, NKR12, VFGP13, Aba12, AKR10a, eMA13a, AK12, AZ14b, jAS14, AZ19, BSZ13, BK14, BCW13, BI10, Cai10a, ÇM13a, CSZ12, CYY14, DKU13, Don14,

– Dor13b, Elb12, ES12d, FPC14, Fan14e, GH10b, GZ10d, HL14a, Hon10a, HLZ13, IKU10, Jan13b, JZZ13, JLW13, JHN11, Kan13, KS13a, KW13, KS13c, Lac14, LYZ11, LZ11a, LL10e, Liu10b, LH11c, LL13g, MWC11, Mao12, MT12b, Mos14, MAD11, NTPVNHLX14, PKP^{+}12a, PX13, PB14a, QW10b, QTL^{+}12, RK12, RAIM14, RnZ10, SWL12, SdAT14, Sha10a, SZ14a, SCLC10, Sun10, SSS13b, TI10a, VSV10, VV10, Wan10b, WGZ10, WqXlC11, WKLW12, WWZ13, WSL^{+}13, WYTL14, WT10, WG12, Wu13, WW13i, XMWD14, XD11a, XL14, YZC13, YH14c, Yan14b, YLGW14, Yu11a, YY14b, ZGH10, ZZW11a, Zha11b, ZZW13, Zha13b, ZWL14a, ZMZ11, ZGY^{+}10, ZZ11c]. Coupling [EMAM13, FWZW13, JJP^{+}12, LWP12, LR14a, LLJF12, SHWT14a, Son10, WLMJ10, WZWW13, WZZ13, Wen12, WP14, WLY^{+}12, XDG12, Xu10b, YL10c, YL13, ZD10, ZW10b, ZhGQ14]. couplings [BSS12, Ma11c, MM14a, YL14d]. Cournot [Ask14, CG14a, CN14a, EM12, SVV13]. covariance [EGMPMRBC12, GYL14, Nak13]. cover [Man12a]. cover-up [Man12a]. coverage [KL14a, PPL12, WSY14]. Covering [FM14, GM11a, KZ12, GM14a, Ni13]. Crab [ÁME^{+}12]. crack [BH12a, GYS13]. cracks [AA13, BH12a, Che14j, GL11b, JKP14b, LZ10c, LWZ13b, MWY13]. CRAGE [MT12b]. Cramer [Kyr10, Kyr12, SW11, LSS11]. crane [DJY13]. Crank [KMR14, LKS14, LLS13b, SW10a]. crashable [LHT14b]. creation [WJ11]. Credit [JGM13, CCO12, CLS12, LWL14, LHT14a, LOD12, NDZ14, OYCCB13, Tch11, ZZ12a]. credit-risky [Tch11]. credits [Tsa13b]. crime [Mis14]. Criteria [Kir11, AZL13, ASMA12, BBI14, CW11a, CZZL13, DZS^{+}14, DE11, DLZ14, DF11, FO11, GHG14, GZW12, HR10, Has11, HYYG12, KZ14b, KLPC12, LPL^{+}13, LLP14a, LLC12, LL13d, MI10, MH11b, PKP^{+}12a, PKP^{+}12b, SDAN14, SGJ10, Shi14b, Sit14, SL11b, SS11d, SDO10, Sta14d, TLZZ12, WS11, WLX11, WWW12, WT11c, YH14b, YG13, ZPXX14, ZGW14b]. criterion [BTRZ14, Che11e, Che14h, DO11, Gal11a, Ign13, JQ14, LCH13, SY13a, SZH13, Sin11b, TZ11, TK13, WLJ11, ZCF13]. Critical [FG14a, Vít14, Aja11, Aou11a, yCfW12, DH13b, FG13a, GL10b, Kan11, Li11b, LWZD14, Vil12, WR10, XH11, YYY10, ZAY^{+}14]. Cross [BWaY14, Liu14a, WWZ10b, Fan14a, FLMR14, Gal11d, Gui14, Ibr14, LSW^{+}14a, ÖB11, SSCA13, Zha13a, GL13a]. Cross-diffusion [BWaY14, Liu14a, Gal11d, Gui14]. cross-entropy [SSCA13]. cross-validation [FLMR14]. crossed [CKH11, CLTL13]. crossing [WL10d]. crossing-monostability [WL10d]. crossover [Gwi11]. Crouzeix [BY11b]. crowding [lLW14, YLC12]. Crowley [MHXyY14]. crystal [BRVR12, PPC13, ZCF13]. crystals [DG10c]. CT [ZL13d]. CTL [TX14, WES12]. CTLs [MEL13]. cubature [DMS12, SSV11]. cube [CKH11, CLTL13, LWY14]. cubeness [MOZ11]. cubes [WFZL12]. Cubic [CF12, CF14, AMA12a, AMIR14, AMA14a, AS13a, CGRS12, DGOR13, EZ14b, GGP^{+}13, GK14, GK12b, HHM10, HXY12, HJW11, JZ11, LL13f, MMX10, MJ12, MB13b, MG11b, MYL^{+}10b, NN13, ÖP12b, Pan11, RV10b, Sha14b, She10, Xu11, YZ10, ZLJ11, ZM10a, ZK10a]. cubic-quintic [EZ14b, LL13f]. cubically [ONBN10]. culture [YZZ^{+}14]. cure [PT13, VDL12b]. currencies [Hua11a]. current [DKK^{+}13, KCZK11, SM10, TPA10]. currents [Gad11, Gad14, LZZZ10, YA14]. Curvature [JZ11, BCGK14, BDK^{+}14, CL11a, CCLW10, LWZ13a, LP13d, Sol12, VC14a, Wan14f]. curve [BAA13, CK12a, LY10a, ZHGP14]. curved [AHKS12, BCC11b, dCDNV^{+}13, NN13]. curves [BS11a, BS11c, Bez13, BRGP14, CMD13, CK12a, CKK12, CDMT12, Den11a, Den11c, GTPM14, HHM10, HJW11, JDLL14, LDJZ13, LR10, Lu12d, PDBRFER14, QHZ^{+}13, QHYW14, RM11, Sag11, SR14b, SIR14]. curvilinear [ASB10, AS14e, SSV11, WW13a]. customers [SLL14]. cutdown [DLL14]. CVaR [TQWZ10]. Cycle [DD11, Yan10c, CKH11, DMGK13, EK11, Gin12a, Gin12b, HZFZ14, Jar10, LR14b, STM10, WZSW14, WLL14, XH14, YS11a]. cycles [CLTL13, DLC10, LZM^{+}12, LW14c, LL11c, Liu12b, Pan11, SGC12, pSbL14, WHL11, WFZL12, WHF14, WZ12, XH10, Yan10d, ZLJ11, ZZ10d]. Cyclic [Rho19, SY14c, DI13b, DK14a, Rho14b, Rho14a, TS10a, YCS11]. cyclicity [ZC14a]. cyclohexane [BGP^{+}13]. cyclohexane/air [BGP^{+}13]. cylinder [AAMAS11, Che10f, DMS12, VPS12, XDG12, Zen12]. cylinders [CL11a, DRJ11, LL11c]. cylindrical [BMMS13, fD12b, KMD13, KS11b, NPA11]. cytokine [Gra13]. cytokine-induced [Gra13].

D [Hua10, JL10, LRM12, AMA14a, AP10, BB14a, Bel11a, Bel12, BTRZ14, EL13, BWHQ14, Che10d, CVP^{+}10, CEXG12, CG12b, CG14c, DTN14, DG10c, ESNE^{+}14, GVT^{+}13, GZW12, GG10b, HWPY14, JM14b, KUTY14, LP12a, LZ12c, LS11b, LW10b, MG10a, Mam13, Med12, Mit13, MMS14, MSS14a, NTPVNHLX14, RV13a, Rec14, RVS10, RPW14, SSK13, SEL14, SMM14, SK11b, SL14b, SZ14c, Tad12, TVZ10, WHL11, WZ13b, WWS^{+}14, YLTB12, Zha11a, ZCL11, ZLG^{+}14, bZpZJ10]. D-local [KUTY14]. D-measurement [RVS10]. DAE [NBYHV11]. DAGSVM [MIM^{+}14]. d’Alembert [Bah13b]. Damage [MMM12, HTX^{+}12, MMM13, NMM14, ZCL13a, ZCL14a, ZCL14b]. Damped [CG14b, AH12b, CTA13, Che14d, Dia14, EZ14b, GSH12, HL11b, LPAE12, LY13b, LZLB14, Ma14a, Oni11, PT11a, SDAN14, WW14c, Zha13c]. damping [CLA^{+}12, Cve14, HM08, KTT12, LZW13a, MH11b, NA10b, PPV14, SQ11, SY13b, Wu13, XHXH12, YlTA10, ZGW14b]. Danchick [ACCT14]. Darboux [YZZP14a, GH10b, HC14, MG12a, MM13a, MGWG11, TZJZ12, Val11, WGW11, Wen12, XTAJ12, YZZP14b, ZHT13, ZLL13]. Darcian [PMD14]. Darcy [ABBC10, FHWZ12, KM11a, LR14a, MPPM12]. Dark [TW11b, Lat14, WD12, YY14b]. dark-bright [YY14b]. Data [RMMM11, AMA12a, AMA14a, AAM13, AWS10, ABCD11, ACQOR11, Ben14b, Ber14, BLR10, CHYT12, Esm12, FY12, Fuk12, GYL14, HZZ14, IG11, JHN11, KG13, KH14a, KN13, LWP12, LYGZ13, Lin11c, Lin11b, LGCX13, Liu11a, LW13c, MTSN14, MSGN^{+}14, OSAG13, ÖB11, PMB12, PJFL14, RSPK13, RAW13, RLW14, RGA14, Sáf14, SHN10, SHA12a, Sev11, SSS14, SBS11, WY10c, Yua11, ZL14a, ZL14d, ZF10]. Data-driven [RMMM11]. date [ZHC^{+}14]. dates [LWW14, YCC^{+}12]. Dattoli [Cvi10]. Daubechies [PM11a]. Davidson [BHP13]. DC [MAD11]. DCT [LPH11]. DDE [YNBV11]. de-noising [SCC12]. De-quantisation [Abb12]. DEA [AP13, AM12e, Che12a, Esm12, SLJ^{+}13, SW14b, WYL10b]. deal [CST14]. dealing [AP13, DE11, DRT11]. DeAngelis [GC11c, Kha14b, LB14c, QLWW12, Zuo13]. death [DL11, HLM14]. Debye [YJ14a]. decagonal [LF11b, Li13]. Decay [WW11c, WW14c, CP12a, fD12b, Kan12b, MP11a, MP14a, MM12, Mus14, eT12a, TJ12, TJ13, WW10]. decaying [RZ11]. decentralised [NLLC10]. Decentralized [LJPJ12, LL12d, WHW10]. Decision [Tsa11, ZBM10, Bas13, BDGIPG12, Che12a, DZS^{+}14, DRT11, FZLZ13, For11, GPP^{+}10, LIX14, LLZ14b, NP14, PAL14, SO10, WHI12]. decision-making [BDGIPG12, DRT11]. decisions [HLL14, OYCCB13, SYB13, THCW11]. decomposable [JFHZ14]. decomposing [Dua11b]. Decomposition [ABG14, ARH^{+}12, Alz14, AP11, ASRI12, AZAK10, BDG13, BCEAV12, BRM11b, BMR13, BR13, BS14c, BCGV11, CKSV14, CT14a, CS13d, CBRTA11, CM12c, DKMM11, DR11b, DR11c, DCR12, FL13, FHWZ12, Fuk12, GXSH14, GSRR13, GG13c, GG14g, HZ13, HBB11, HB11b, HMZ14, Jun10a, KCG10, LWZZ14, MTSN14, MP10, MP12a, NL10, ÖKU12, QS11, QS14, RD11, SCC12, SMH14, SW14b, SPKS12, SNZL13, SK10c, WZZ11, WM13a, Waz10a, WRD13, XRBZ13, XBC^{+}14, YKI10, YLWH14, ZY13c, Zhu10d, DAA13, XBC^{+}13]. Decomposition-based [Alz14]. decompositions [FDHK14, GMX10, MTER14, MMRS13, XZ13b]. deconvolution [ANR14, KMR14, RLW14, Wil12]. decoupled [Nie13]. decoupling [LR13a]. decreasing [LTHG14, LWW14]. decreasingness [eT12a]. dedicated [Kos13]. deduce [ZFH14]. deep [eMBHA13]. deep-water [eMBHA13]. default [LWL14]. Defaults [And12a]. Defect [RS11b, Chu11, EKZ11, Wan10d, dV12a]. Defect-Correction [dV12a]. defective [GL13a, JGM13, LOD12, SGCG14, XXZH10]. deferential [ZZW^{+}12]. defined [AP13, EKK14, HLT13, LOS13, MPS14b, Sav10, SKM11, WXA10, ZH12]. defining [LL14c]. Definite [BDD^{+}12a, AG14a, Bry11, CC10a, CMX14, GW12, GZF12, GFZ13, mHhL10, Kat14, KM13c, LW14d, LZL14b, LC14b, RSK14, WLY11, WWY10, WYW14, YAW10, yYyL10, fZwXyZ11, ZGZ14, ZCL13b, ZZW13]. definiteness [RS11b]. deflated [LLBW11]. deflation [PR11]. deflection [HCL^{+}13]. deflections [MBR10]. defocusing [DPC10]. deformation [CCLW10, CMS11, Nar12, Yu11a]. deformations [VCĆV12]. Deformed [BR14c, HMIZ13, SMR11b]. Degasperis [Den11d, LZH10]. degeneracy [PP12b, PP13a, PP15, PFS10]. degenerate [AST12, ASST11, Che11g, EJ11, Gin12a, Gin12b, GG14e, KCLC13, LKY10, SS11a, Vod13, WLMJ10, WZ12, YLZ13, Yu14, Zha14e]. degradating [MCPO12]. Degree [GG14c, BOR14a, ÇBS14, CSMT12, DLL14, FGD13, GQ12, LW11b, LZM^{+}12, LHZ10, Lu12d, ML11a, Moo12, QHZ^{+}13, RM11, ROA11, Sha13b, pSbL14, Sun12, XLW11, ZWD12, ZZ10d]. Degree-based [GG14c, FGD13]. degree-elevation [DLL14]. Dekkers [AGZ12]. Delay [BBF12b, BNP11, Che10d, DLZ14, LFR14, MPSA14, McC10, TLZZ12, WPSW13, ABLZ13, AYS13, AM14b, Ala14, AS14b, AC10b, APJ14, BDR12a, Bar10, BP11b, BBDS14, BLK11, BBI14, BJSS14, BCEAV12, Cao10b, CJ14, CLTY13, CS11b, CW14b, CW14a, CZ11c, CZZL13, CY12, DGS14b, DT14, DMW13, DL10, DZZZ12, DWC13, DLX13, DP10, DDWW12, DL11, Fan10a, FWZW13, FP13a, FP13b, FS14, Far14a, FZ12, FL14, FGS14, FH13, GHG14, GSH12, GC10a, GGÖS11, GGXH12, GSDW14, GM14c, GH10e, HL11a, HM11, HX13, Jan13c, JPK12, JQ14, JXK14, JS13b, JK14c, KJ12, KfM11b, KB14, KGC11, Kro11, KSY12, Kum14, LZQ11, LTY11, Li12c, LL12d, LCH12, LYWL12, LAF13, LHC14, LSSZ14, LLY14a, LCL^{+}13a, LDL11, LZG10, LLZY10, LJKU11, LY11b, LHS12, LXC12, LCL14, LC14e, LC14c, LL10i, LYP14, MLH11, MLY12, MYW13, MA10a, MHA12, Mal12b, Mao11a, Mao11b]. delay [MA10b, MG12b, MAD14b, Mil14b, MMP^{+}12, MF14, MN10, NP10a, NCL11, NTS11a, Ngo12, NNG10, NSZ^{+}14, PKP^{+}12a, PWZZ13, QaC13, QHL10, Qui10, RS11a, RCS14, RHSC14, RRLA14, RRM12, SP14b, SHL10, Sar12a, SGCG14, SDL10a, SSA14a, SY10a, SZ14a, SHWT14b, SS13c, Sta14d, SW14d, Sun11a, Sun11b, SHM12, SC13, SCK14b, TGL14, TZ11, TXX14, VDL12b, VJ11, VSM14, WCB14, WL10a, WY11b, WZ11b, WS11, WXLW13, WZWW13, Wan13d, WWS^{+}14, Wei11a, WZH13, WW11f, WC11, WLZ12b, WE13, WHZ14, WD14c, XPLZ14, XC10c, XX14a, XM12, Xu14b, Xu14a, YMZS10, YG11, YYJ14, Yue14, ZB11, ZPXX14, yZ14a, ZHX11, ZZZ11, ZL11a, ZZW11a, ZMZS12, ZCZS12, ZX12, ZLS12b, ZS14b, ZSL10, Zha13c, ZW14b, ZL14f, ZW10d, ZHC11, ZGW14b, ZMC10, ZZS14, ZWZH13, ZFCW13, dS10]. Delay-dependent [BNP11, Che10d, LFR14, TLZZ12, WPSW13, AS14b, CZZL13, LYWL12, TZ11, WS11, ZPXX14]. Delay-independent [DLZ14]. delay-margin [DGS14b]. delay-partitioning [TXX14]. delayed [AC10a, BPP12b, BCÁV14, Cai10b, CGL12, ÇM13a, ÇBS14, CXL12, DGS14b, DJY13, DLZ14, DX14b, EK11, EDYA14, EMM^{+}12, FW14, FA12, Fig12, GLY14, GL13c, HY11a, HL10, HLTT14, HZW12, JMB12, JJP^{+}12, JSH10, JZY10, KG11, LL13b, LWTC10, LZ10e, LB14c, LW14g, LZLY14, LB14a, MI10, MEL13, Mur14, NHCLP10, NL12b, PMP11, QLCF10, RB11a, SS14a, THP14, TX14, WZ11c, WF11, WES12, WPSW13, WD13, WW13g, WW13b, WLP^{+}14, WFL14a, WL10d, WLY^{+}12, XSC11, XLS13a, XTL10, XZ13a, YLWU13, YMG11, Yan13b, ZJS10, ZGH10, ZX10, ZXCL10, ZLMG12, ZZ13a, ZW14c, ZYYZ12]. delayed-differential [ÇM13a]. delaying [CK14b]. delays [AS14b, AGH12, BMR10, BRP13, BC12b, BH12b, BN10, BNP11, BK11b, BK12, CHC11, CZZ12, CZZL13, CZD^{+}14, DGS14a, DGS14b, DWP14, DRGR10, DRS12a, DRS12b, DZZ^{+}13, DZZ14, EGO13, FO11, Far14b, FP13c, GxFX^{+}12, GWZ14, GWW14, GLHCLP12, GRV13, Ge14, GS14c, HYT13, HYJ14, HLC13, HYYG12, JDN14, JJPW13, JV13, KS11a, KH14a, KJW10, KTS11, KPLC11, KLPC12, KPP^{+}13, KPP^{+}14, LPJB12, LPL^{+}13, LWP12, LLP14a, LP14, Li10a, LZZL10, LX12, LLC12, LYC12, LXK13, LY14b, LZL14a, LX14a, LS10c, LTX11, LDL13, Liu10c, LX11b, LLJF12, LLP14b, LSL^{+}14, LLKP14, LZL12, LD14, LDZ10, MS13a, MK14a, Mao12, MM13b, NCÁHCLP13, NP10a, PZ14a, PKPL11, PKP^{+}12a, PKP^{+}12b, RSPK13, RS13, SAMA11, SAL12, Sha14c, Sha11b, SZZH13, SZ14b, Son10, SHZ14, SDB14, SIV10, TPT12, TLZZ12, TXX14, TC14a, TJ12, TJ13, TJ14, WZ13a, WHW10, Wan11d, WW11a]. delays [WY12a, WLLY14, WPKE14, WQ13, WPSC11, WPSC12, XX10, XCT10, YZW^{+}10, YL11e, YLC11, Yan12e, YC14, YH14b, YZ14c, ZaFT11, ZLS12a, Zha12a, ZS13b, ZLJZ14, ZZWY14, ZZW11b, ZZQ13, ZCY14, ZLY11c, Zuo13]. Delimiting [Rec14]. deliveries [LYL13a]. delivery [CBTGWW14, LOD12]. delta [CD12, LM11, OGO14]. delta-evolution [LM11]. demand [AAKB13, Ask14, COPR10, CLS12, DG10a, KGC11, KZS12, LXZX13, Mal14, QGD14, San11a, San11b, Sar12a, SGC12, SMS14, TNB11]. demand-side [DG10a]. demands [Dim13a, Lin13b]. DEMATEL [WT11c]. Demiclosed [CWL^{+}12]. demicontractive [BS10c, MM11a]. demonstration [YS11a]. dendrimers [CDESS14]. dengue [lLW14, MG14b]. denoising [GV14, YPSW11]. denominators [Man12a]. dense [ABD^{+}12, GWLZ10, PM12]. densities [Aub10, MIT12, RRSPTR12, SPT12]. Density [HLMW13, EASK10, HVC12, KRMM14, LTY11, LP13c, Man12b, MCM13, PPK14, SDC10, Üna13, YY14a]. density-dependent [HVC12]. dependable [CHK10]. Dependence [BKW11, DF14b, Gal11b, GWS13, BFM13, Fig11, GS13a, GKS10, HXZ12, LL14e, LDL10, RTV12]. dependency [AMV12, IBS11, OH12, VDLEK14, Yua12]. dependent [AS14b, AKS11, jAS14, APJ14, AMB10, AGH12, ABBC10, BKA12, BLZZ14, BK10a, BRP13, BP11a, Bhr13, BCWM10, BN10, BNP11, CL14a, CFH^{+}13, Che10d, CZX^{+}14b, Che10f, CW11c, CZZL13, CLS12, CFN12, CG10, DLL11a, DWP14, DE11, DMR10, DWC13, EM11, EEM12, ET12b, FW14, FR11, FH13, GWW14, GL10a, Gün10, GB13, HZ12a, HVC12, HMMK11, HHDM13, IS10b, IKL11, JCW12, JKB10, JK10, KXW14, KMC12, KGC11, Kim11, KY14, LL14a, LTY11, LYWL12, LFR14, LYL13a, Liu10c, LLX10, LY11b, LZTZ13, LZL12, LWW14, LDZ10, LYB12, LH14c, LF14, MZW14, Mil14b, MMP^{+}12, NA10a, NP10a, NSZ^{+}14, OP14b, RSBK14, RSPK13, Rud13, Sar12a, SGC12, SMS14, SY10a, TKL14, TZ11, TZJZ12, TLZZ12, TW10a, TW11b, VO12, VO14, WS11, WQ11a, WW11a, WH12, WPSW13, WFL14a, Waz10g, XZY14b, YlTA10, Yan10e, YGL11, Yan13d, YLS13]. dependent [YJ14a, ZPXX14, ZR13, ZW13b, ZLL14, ZLD11, ZD11b, ZS14f, dCCL14, dS10]. depending [DIAQ13, Hil12, Jun12, LfXfC13, OYCCB13, YSI11]. depends [KL12]. depletion [Mis11]. depolymerization [CCG10]. deposition [HCCT14]. depth [Dem11, LXZ14, RTV12]. Derivation [EMA13b, ÖK11, RTV12, BIT12b, GP11b, HMH13, KCLC13, TCBL11]. derivational [MVS14a]. Derivative [Bur12, JDLL14, PID10, Abh14, AGZ12, ACCT14, CCGT13, CLND12, Den11a, ESV12, HD14b, Jan13d, KA11, KM11c, Koç11, LDJZ13, LLK14, LL12f, MT12a, Mer14, MG14c, MSS14a, PG13, PG11, PHGV13, PQW10, RA10b, RA12c, SGG12, SAP14, SK13b, WDZ12, WDS11, WD12, XZ13b, Yun11a, ZL12b]. Derivative-based [Bur12]. derivative-free [AGZ12, ACCT14, CCGT13, ESV12]. Derivatives [MS12a, Akt14b, AJF13, ACG14, ABGGS13, Ans13, Asa10b, Ata11, BDGP14a, BB13b, CKP^{+}13, DWLL10, DHI13, DGMR10, FGRLSV14, FPTS14, Fuk13, GGPT14, GK12a, HHT13, Irm11, Kim10, Kim11, LSZ12, LL14c, LN10, LfXfC13, MXH10, Pre11a, SA10a, Ste13j, Ste13d, TCBL11, VD14b, WZ11g, ZLW12a, ZLW12b]. derived [DC13a, GSDB11, KI12b, TCBL11]. dermal [AAP10]. Descartes’ [HT11b]. descent [BA14a, Boy14, CCA11, DW12,

– Dor12a, DSPZ13, Hou13, HD14a, Jia13a, KcZ14, LF11a, LZGZ11, LWN14a, ZF13b]. descent-like [Hou13]. describe [MPB13]. described [HH13b, KEGH13]. describing [KHA12, KS10c, MM10c, QHZ^{+}13]. description [CGK12, HLT14, KS13c]. descriptor [DLZ14, Hsi10, RZ12, SSM13, WXLW13]. Design [Che13c, DGH11, HCC^{+}12, LPJB12, LWZ11, NHCLP10, Nak13, SA12a, WLP^{+}14, BHJ13, CTV12, DGCFM11, DHDB12, EL14, HZ12a, Iç14, JLLD13, JWWY13, KV12, KAY12, LY10a, LXK13, LX14a, LCL^{+}13a, MZXZ10, MLH11, MS13a, MC13, PQRR10, RDFL13, TGTCCGRG13, Tsa13a, Tsa13b, VFGP13, XPLZ14, ZCŠ10, ZSWS12, ZDC12]. Designing [ZHGP14]. designs [BS14b, Han14, LHL14, TPTW14]. desirability [KI11a]. desires [WHI12]. despite [RGA14]. detailed [GS12c, Zhe14]. detectability [LWZD10b]. Detecting [ZLC^{+}13, RV10a]. Detection [FHK13, CTH14, CBRTA11, CHT14, HTX^{+}12, Jan14b, KRJ11, MSGN^{+}14, SMC12, STM10, Zuf12]. deteriorated [XG10]. deteriorating [BJSS14, CWCW11, KGC11, LL12a, LHT14a, SGC12, WW11d, WL14a, Yan10e, YK11a, YX11, YX12a, YGX14, ZW11b, ZL13a, ZHC^{+}14]. deterioration [AAKB13, Ery14, GL10a, Jea13, LYZG13, San10, WCW^{+}13, YW11b]. determinant [AM12c, CY11a, Chu10, Elo11, Elo13, Fen11, TMXG11, YB12, ZCZ13b]. Determinantal [Kyr14, Son12, Kyr13]. Determinants [BT12, TLL13, Che14i, Elo14, IA14, SCH11b, SY14b, TL10]. Determination [GT12, IKL11, KM11c, AGHdD13, GL12, GKC11, Jea13, XWWF10]. determine [Ahn13, Zou12]. Determining [XwK11, AGR14b, DE12, GGM14b, LPB10, RGL10, RLW14, ZZ12a]. deterministic [CLS12, GK13, MAB14, ZHXM14]. deterrence [Mis14]. developable [LWZ11]. developed [CSQ11, HS10b]. development [KTW10, NCS13, Ram10a]. deviating [BCC11a, DKS14, HLX13, Kum11, Ma14a, Ste12c, Ste12l, Ste12f, Ste12g, Ste12r, Ste13a]. deviation [EG11]. deviations [GVMNS13, GSSS13, Ste12h, Ste12m, Ste12u]. device [LY12b, YY13a]. devices [YY12]. DFSA [AGZ12]. DGMRES [ZW10c]. Dhar [Ban14]. diabetes [FGFRV11]. diagnosability [PLTH12, YLQ13, Zho12a]. diagnosis [CZX^{+}14a, CZX^{+}14b, EdCVB^{+}14, TP12, TL14, ZDH14]. Diagonal [KM13b, CK12b, DSZ14, GWLZ10, KL13b, tLpOjCW10, Nem10a]. diagonal-Schur [tLpOjCW10]. diagonalization [HW14a, SEM11]. Diagonally [Kal13, GK10a, tLpOjCW10, LZ13b, MH10a, WWLL11]. diagonals [GG11c, GG13b, GG14g, Rim13a, Rim13c]. diagrams [ES12d]. diameter [HLS^{+}14]. diameters [SMPA13]. Diamond [BC13b]. Dichotomy [DL14, DGH11]. Dickson [BC12c]. dielectric [LZ10c]. Difference [Fan14d, Mah14b, Ste12e, SQJ11, YL14b, Ali12, And11, AHDD13, AAGR13, ADP11, AY11c, AO12, Asl12a, AS14e, AB12c, Bal13, BP11b, BJS14, Bek10, BKP14, BS11a, BS11b, BBFK12, BS10b, BK11b, CST12, CI13, CCJ10, CLX13, CJ14, CDMS14, CT11a, Chu11, CG10, CG12b, CK13a, CDCS13, Cui14, CK12c, DAHAK10, Das12, DBK14, DL14, Dia13, DS12b, DISS14b, fDxZxChT12, DJ14b, DK14c, DB12b, EM13, EEEE05, EA14, Et13, EKK14, FCJ14, FMF13, FP12, FP13d, Gal11c, Gao10, GGM14b, GG14b, GG12b, GÖS10, GGÖS11, GL13b, GSRR13, Gür10, Gür12, GH10e, GH11, HV12a, Her11, HSX10, HHX11, HZ12c, HHX12, HX14, Ign10, Iri10a, Iri10b, Jan10, Jan11, Jan13b, JŞ11, Jha13, JH10, mJ10, JL10, KPC14, KLBR13, KAK11, Kon14, KS11d, LW13a, LLY11a, LZW12a, LLXZ12, LY13a, LTOX11]. difference [lLZ14a, LZKK11, LY12b, LS12b, LL13e, LYL13b, LT14, LC14c, LZSD14, LO14, MG10a, MDP12, MA13c, MA13d, MM13b, MG11b, Nie13, Öca14, PZ12, Pan12b, PSS11, PRS12, PEP14, Par11, Par13, QZ10, Ram10a, Rao14, RnZ10, SS10b, SMI12, SA10a, SYK10, SXL13, uiH14, SK12b, Sta14d, Ste10e, Ste10f, Ste11f, Ste11g, Ste11c, Ste11d, Ste12b, Ste12l, Ste12q, Ste12n, Ste12s, Ste12o, Ste12j, Ste12p, Ste12k, Ste13k, Ste13c, SAAS13, Ste13n, Ste13h, Ste13i, Ste13e, Ste13f, Ste13g, Ste13o, Ste14b, SAAS14a, Ste14d, SAAS14b, SAAS14c, SDS14, Ste14e, SWX10b, SXH11, zSWZW12, SHWX14, TM13, TYT14, Wan10e, WL10a, WGZ10, WZZ11, WLH12, WL12a, WZJ13, Wan13a, Wei11a, WW14e, WP14, WL13d, XLGW12, Xu14c, XZ13c, YGL11, Yan14b, ZFH14, ZW12, Zha10e, ZL12c, Zha13d, dlS10]. difference-decomposition [GSRR13]. difference/finite [ZL12c]. differences [JS10, LXY10, Nov14, Ste10d, WW11d, XC13]. differencing [BK14, LKS14]. Different [Liu10d, Mag14a, AA12a, AFSMRL14, CFGRV13, Che10b, CCO12, EDYSA12, GM13b, Gwi12, HAA^{+}14, lJLpWxZ14, LL10e, MZXZ10, NOS10, OADL14, SGK14, Ste12v, WX12, WL14b, XLS13a, YLC14, ZH10a, ZX10, Zhe12]. differentiability [AH11c, BM13b, FPTS14, Mal12a]. differentiable [AG14f, BNP11, Gue14a, Hwa11, Hwa14, LFC^{+}12, Mal12b, Ste13j]. Differential [BBS10, CK13b, GL14a, JYCY10, JYL14, Nar12, Shi13a, WW13f, YZZ12, AB11, Ach11, Ach13, AOAA10, AS11a, AC10a, Afu12, ABLZ13, AS10, AISS13, AYS13, ADÇY11, AIAD12, AJF13, AZ14b, AGR14b, APA13, Ali12, AAT^{+}13, AANA10, AM14d, Ans13, AGKR10, AS12b, AS13a, AASS13, AAS14b, ARS14, ADP11, AO12, Asl12a, AJPZ14, AZ19, BRS13, BD12a, BDR12a, BDD12b, BMA11, Bar10, BMD13, BP11b, BLK11, Bek10, Bel13, Bel11b, Bel11a, Bel12, BBI14, Ber12, Bho10, BES10b, BTS12, BCEAV12, BRM11a, BR13, BX12, BK11b, CH14, Çak10, CCJ11, Cao10b, CYHL11, CG12a, CGTCF13, CC13, Car13a, Car13b, ÇM13a, CKN13, Cha11, CCJT11, CCCRR13, CS10a, CZ11a, CT12b, CM12b, CYM13, Che13c, CZZ14, CSL14, CK14a, CD11a, CS10d, CSSN11, CP10, dSCAM14]. differential [CDST14, CTLD12, CJRV12, CSS11, CHS14b, CK12c, DGSB14, DW14, DSGB12, DGSB12, DB11a, DK11b, DX10, Dia14, DISS14b, DL10, DP10, DAEY12, DBBEE13, Du11, DW11, Dua11b, DR11c, DCR12, Ech12, EK10, EN11, ESEEKH12, ESG11, EB14, Els12, ESE14, EÖ10, EZ13a, EH14b, FCJ14, Fan10a, Fan13, FEGEA14, FcDDqZ10, FFOC14, Fer10b, FLP11, Fig11, Fig12, FH13, GLP11, Gal11e, GGP^{+}13, GMX10, GjLxF^{+}13, GP12, GLV14, GG14b, GPCCA14, Goo13, GWH12, GADK14, GÖS10, GÖ14, GZF12, GMSH13, GM14c, GS14f, GMSE13, GG10b, HT11a, Han11, HLZ12, HL11b, Har10, HR10, HET12, HXZ12, HL14a, HPB11, HHM14, HR13, HR11, Hu12b, HLX13, HRX14, Ign13, ISG11a, JKN11, JVK10, JDJ11, JDJ12, Jan13c, Jan13d, Jan14d, Jan14e, JM11, JLW11, Ji14b, JS13a, JL14, mJ10, JS13b, Jia13b, Jia14, JPM12]. differential [JMS13, JMHZ10, KK14a, KEGH13, KC12, Kar13a, KLBR13, Kau14, KL14b, KCG10, Kos13, Kro11, KS11d, KTS11, KS12d, KK12c, Kud13a, Kum14, KM12c, Kyr13, LZ10b, LYZ11, LLJ11, LCH12, LZW13a, LAF13, LDWY14, LSLL14, LH14a, LSL10, LLL11a, Lin11a, rL11, LLL13, Lin13c, LL11e, LJKU11, LH11c, LJ13, LM13b, LZ14c, LLS14, LYG14, LFLH14, LC14c, LW14g, LP13b, LM12, LXZ12, LXLY14, LYL11, MH13a, MYW13, MCH14, Mad14a, Mal12a, Mal12b, Mao11b, MLGdSC11, MGMM14, MN14a, MIHS14, MPS14b, MH11b, Men12a, MAD14b, Mil14a, Mil14b, MB13a, Mis10, MJ11a, MB14, MJPB11, MSC^{+}13, MS12b, MSS14a, MEG10, MF14, Mol14, MAHAS14, Mop11, Mop14, Mor13, Mos14, NCL11, NR14, Ngo12, NZW11, NRL12, OH12, Ort14a, Ort14b, ÖZ10, PSP10, hPrLgZ10, PT11a, PS14a, PJ12, POR13, PB14a, PR12b, QHL10, Rad11]. differential [RMN14, RMY14, RS11a, Ram10b, Ram10a, RGL10, RGL11, RK14a, RS13, RCS14, RSRA14, RN10a, RnZ10, SHT14, SuR13, SAO12, SGA13, SS13a, Sal11c, SDAN14, SB13, SI11a, SSZ11, SMIS14, Ser13, SRP13, She12c, Shi14b, Sho14, SK13b, SGA14, SL10b, Sta14a, Ste12c, Ste12f, Ste12g, Ste12m, Ste12r, Ste12t, Ste13a, Ste13j, Ste13k, Ste13d, Ste14c, SDS14, Ste14f, SW10b, Sun11a, SH12b, SHM12, SCK14b, TM11, TMM13, TMM14, TS14, TSP10a, TA13, TJZ13, TFMB12, TPA10, Tom12, Tom15, TT13, Tur14a, Tur14b, VV10, VMRA14, VSM14, VRMT14, VD14b, Wan10e, WLH10, WzH10, WQX^{+}10, WZZ11, Wan11c, Wan11a, WW12, WF12, WZG12, Wan13a, WW13e, WCZ13, Wan13d, WDT^{+}13, WFZL14, WW14b, WMM14, WZL14a, WYK11, Waz10a, WW14e, WW11f, WC11, Wu12a, WE13, WL13d, WHZ14, WD14c, XL11, XMA14, Xie11b, XD11b]. differential [XM12, XZ12, Xu13c, XL14, Xu14a, XXS14, YL10b, YW11a, YMG11, YKL12a, YKL14, YHWZ11, YG11, YLAT14, YYJ14, YHW13, YH14d, YCF13, YWCH14, Yue14, YS13b, Yüz13, Yüz14, ZLA10, ZZC10, ZL10b, ZZZ11, ZLW12a, ZLW12b, ZCZS12, ZSZ12, ZLWL13, ZHY13, ZZZ14, ZL14e, ZLWW14, ZWS14, ZQ14, Zha14a, ZSL10, Zha10e, ZSHZ11, ZHZ12, Zha13c, Zha13d, ZL14f, ZYSQ11, ZH14, Zho11a, Zho11c, ZZ12b, ZLZY12, Zho12c, ZCC13, ZXF14, ZGW14b, ZSD10, ZR14, ZWH14, dlS10, dS10, uRE13, AGNAL14, Shi13a]. differential-algebraic [SCK14b, ZZC10, ZCZS12]. differential-difference [Asl12a, Bek10, DISS14b, GG14b, RnZ10, SDS14, Wan10e, WZZ11, Wan13a, WW14e]. differential-integral [WZG12, XL14]. Differentiation [TS10b, Asa10a, DSGB12, DFM10, Ger12, LS10b, MNP13, PKBD12, Ste10a, Ste10n, Ste11a, SSB11a, SS11g, SSB11b, Zha10c]. differentiator [Wu14c]. diffuse [XRBZ13]. Diffusion [Yan13c, ACZ12, ABJ10, Alf12, Ali12, AKS11, AC10b, AP14, ANS14, Bac14, Ban14, BC10a, BDG13, BWaY14, BCÁV14, BA13, CLZmF11, CL14a, CB10, CLAT11, CX11, CLA^{+}12, Che13a, CFX14, CC14c, CC10c, CG10, CG12b, CG14c, CS14b, Cui14, DWLL10, DN14a, DJ14a, DYYL13, DM13b, DBS12, DWP10, EM13, EAA11, Fan14a, FGO12, FR11, FIK14, Gal11d, AA14a, GS12c, Gui14, GMZL11, GML13, GK12b, Gür10, Gür12, GJGSNRÁ12, Hay11, HVC12, HZ12c, ITLA10, IS10c, JM13, JXA11, JZG13, Juk13, JBDC12, KKV12, Kos11b, KR10, LS13a, LLW10a, LZQ11, LDWY14, Lia11, LWCS14, lLW14, LZ10e, LTU13, Liu14a, LB14b, LYP14, MG10a, MLY12, MC14a, MVQDE12, MR14b, MPS14a, MO12b, MJ11b, MMGM13, MMS14, MM10c, MYL^{+}10b, NRWF12, NSL14, NKM^{+}10, OP14b, RB11a, RB14, fRbLJyS13, RRRRTR12, RRSPTR12, RRSPTR14]. diffusion [RZHH13, RZHH14, SSL10, SSL12, Sar12c, Sha14a, SL13a, SY10a, SLC^{+}12, SZ14b, SCC11, SF13, SS14e, SW10a, SHZ10, SXC^{+}12, SW14d, SBDB13, Tad10, TGL14, VD14a, VT13, WWZ10b, WT11b, WQ11a, WW11a, WiN12, WZ13b, WW13b, WZZ13, WWQ14, WLQ^{+}10, WL10d, YWZ10, YS13a, Zha11c, ZYC12, ZL12a, ZW13e, ZX13, ZCZ14, ZW11c, ZX14, ZLY11c, Zhu11a, ZXZ11, Zuo13]. diffusion-based [JM13]. diffusion-convection [ZYC12]. diffusion-induced [WWZ10b]. diffusion-reaction [GS12c, MVQDE12]. diffusion-type [SS14e]. diffusion-wave [BDG13, CLA^{+}12, HZ12c]. diffusions [HYX14, RWkPk10, YZ14c]. diffusive [ABE^{+}12, AAA13b, Che11d, SW13, SSM11, TW11a, VS13, WWS10, WD13, WW13g, YLWU13, Gaz12]. digamma [Mor10b]. digestion [CZ11c]. Digital [BSØ11, HCC^{+}12, MO14, SY13a, Tsa13a]. digraphs [GD10a, LW14a, XFS12]. dilemma [XMWD14, ES12b]. dilution [DMGK12]. dimension [BB13a, BF12, CST14, FMSGS12, GGM14b, GM11a, GM14a, GG12c, GMM^{+}14, GL14b, JLTZ13, KKVCS12, LL11c, SI14, Wan11b, YRV10]. dimension-adaptive [Wan11b]. Dimensional [Waz13, AMA12b, ABJ10, AAP10, AHRV11, ANR14, AM12c, AM12d, Asa10b, AWHW11, Asl10b, Asl11a, AC12, Bac14, BHAV10, BAV11, BZ10a, Bai13, Bai14, BPT12, BS12a, Bel12, Boy11, CL11a, CRG11, CTLL11, CLTA13, Che13a, Che14g, lChH14, CKK12, CS11d, DL12a, DWQG12, DR11a, DGDA10, Den14, DWZ12, DHMU14, ESG11, ES12c, Esf11b, ELS13, EKZ11, FAM11, Fer12a, GV14, GT13, GMST10, GL11b, GW11b, GYS13, GM11b, HHHS11, Har14, HB11b, HYT14a, HCD14, Ibr14, ITLA10, JZG13, JW12, JPM12, KMW12, KA10a, KA10b, KX14, Kos11b, Kud10c, KRS10, KCLC13, LzJZ12, LJR14, Li10d, LF11b, LL10d, LWD10, LL13d, LL10f, LLWW10, LY12b, LLL12, LWZ13b, LYL13b, LTU13, Liu14c, LS12c, LLL10b, LDLL10, LDLM13, MAA11, MZ12, MAO10, MGYQ12, MGC12, MHB14, MJ11b, MB13b, MB14, MG11b, MS12b, MG14c, MK14b, MZWZ12]. dimensional [MPB13, NTT10, NAA13b, OGYE13, OPD12, Pan10, Pan12a, PTY12, PTY13, Par12, Pat13a, Pat13b, PMP11, PB14b, PBG14, QZ10, QW10a, ROK12, Reu11, RNEA14, RH14, SWL12, SF11a, SN11, SATdV13, SdAT14, SJR10, SHAO14, Sht11, SY14c, Sol12, SS10f, TMXG11, TTM12, THZF12, TZS14, TW14b, TSN13, TD14a, TD14b, TZJZ12, TK12, TW10b, Tua13, VT13, WLJ11, WDL10, WQX^{+}10, Wan11d, WaZZW11, WL12b, WZ14a, WCL14, WPKE14, WLXX14, Waz10b, WM10, Wen10a, WX13d, Wu10c, Wu12c, Wu12d, XYC11, XZT10, XGG^{+}11, XZZX14, YTS12a, Yan13a, YL14a, Yig11, qYqLtW10, YM10c, Yu12a, Yu13b, YS13a, ZD10, ZFH14, Zha10b, Zha10a, ZHY11, ZX11a, ZMZS12, ZgJlK12, yZqLwW12, Zha13a, ZL13b, ZL14c, ZGW14a, ZDW10, Zha10d, Zha12b, ZZR14, ZZJ14, ZD11a, ZWT12, Zhu11c, ZZ13c, PPK11, Waz12a]. Dimensionless [MATA14]. dimensions [CZM11, ELS13, HSI13, KMR14, Kle12, Kyc11, MRE^{+}12, NSNEV14, PDBRFER14, RNZH13, Sin12, Üna13, WZ11e, Waz10c, Waz11a, Yan12b]. Diophantine [SS14b, ASS13, ÖTC11, PAC13]. dioxide [MV13]. dipolar [HL14b]. Dirac [BDQR11, BK10a, DG14b, Fuk14a, Fuk14b, HMIZ13, IS10a, IS10b, IBS11, IS12, LCH10, WSD12]. Dirac-Hulthén [IBS11]. Direct [CMM10, DTN14, GY12, Kos13, MIHS14, NTT13, Tri12, AT10b, AM14c, FcDDqZ10, Fuy14a, GMR11, LPR10, ML12a, MT10, Sam11, hTlCqL10, WL13a]. directed [AMV12, SXGR11, Xu10a, KL13a]. directing [fGyLJ11]. direction [ANS14, BBK13, BA14a, bJLyZlL14, KZ11, KZ12, Kwo12, LLL11b, PL10, SX14, YLTB12, Yua13, ZY14b]. Directional [KUKU14, AH11c, Esf11a, JWZ^{+}12, TPA10]. directions [AG14b, LWWZ11, VL12, YLD14]. directly [SMIS14]. Dirichlet [Ant13, BV12b, CV12, CT12b, DX14a, Gal10b, HHM14, HR12, JPM12, MZ13a, MQ14, RB11a, SK12a, TJZ13]. disaggregation [SW14a]. disc [LHZ10, Yan11e]. discharge [QS14, ZZ13c]. discipline [Dim13a]. discontinuity [Sar13]. Discontinuous [Fig12, Mea13, lQZWC14, ZYSQ11, BC11a, EL13, CLF13, CYZ14b, DX10, DC13b, DLS13, EVDG12, FGO12, GLY14, GZXJD10, HZW12, JYCY10, KH14b, LC14a, LL11c, TD14a, TD14b, Tem12, TA13, WZ10, WG11, XH13a, ZF11, ZYZ12, ZY13b, ZXZ11, ZZ13c]. discount [TC14a]. discounted [JWG^{+}14, RFK12, SMWR10, ZY12]. discounting [DLP10, RT14]. Discovery [Hua11a, MS10]. discrepancy [LP11]. Discrete [CM12a, Fan10a, LNPS10, LHTN13, Maa14, MZ13b, MS14b, WD14b, WYL11, Yan13d, Zit11, AR14a, ABK14a, ASMA12, BK10a, BC12b, Bel11a, Bel12, Ber10, BR14b, CBTGWW14, CC12a, CGM13, CC12b, CZZ12, CM14, CYG14, CGO14, CHT14, CZ10, Cvi12, DS14a, Den10, DSX11, DRS12a, DRS12b, Dos14, DZZ14, DW13b, DSPZ13, EEZ14, EKZ11, FKS13, Fau14, FLMR14, Gal11b, GS12b, GS13a, GX11, Gao14, Goo11, GAC12, HL11a, HL11c, HLC13, Jan13a, JDN14, JŞ11, JY12, KAY12, KW13, KY14, KV14, LPJB12, Lee11, LJKP11, LWWZ11, Li12c, LYPC12, LHC14, LX14a, LKH12, LCL^{+}13a, LK11, LWFG13, LLZY10, LX11b, LWZD14, LYLS14, LSL^{+}14, LLZ14c, LS14d, MA10a, MK14a, MMMMB13, Mao11a, MG12b, MB11, MCR14, MPB14, NHCLP10, NCÁHCLP13, Pai14, PZ14a, PAC13, RY13b, RY13a, RY14b, RS11b, SE12, SS14c, SKSK12, SBR^{+}12]. discrete [Sin10b, Sin11b, SS10e, SD13b, TSP10a, THNTNXR11, TXX14, WCB14, Wan13e, WLXX14, WF14, WZH13, WR10, WPSC12, WLZ12b, XX14a, XX14b, XX10, XCT10, XGM10, YWT14, YZD10, YL11e, YLC11, YZC13, YT13b, ZMD^{+}14, ZB11, mZT11, ZLS12b, Zha14d, ZW10b, ZZ13b, ZTX10, ZCZ^{+}13a, ZL10f, ZSD10, DMS11]. discrete-continuous [LKH12]. Discrete-time [Yan13d, ASMA12, HL11a, HLC13, Lee11, LJKP11, LYPC12, LX14a, LCL^{+}13a, LK11, LLZY10, LSL^{+}14, LLZ14c, MA10a, MG12b, NHCLP10, NCÁHCLP13, Pai14, Sin10b, Sin11b, WCB14, WF14, WPSC12, WLZ12b, XX14a, XX14b, XCT10, ZB11, ZCZ^{+}13a]. discretised [AGKR10]. Discretization [GL11a, Wu14c, ABBC10, CYZ14b, Don14, DAA13, GMR11, GLB12, MT12b, MS12b, MK14b, MSS14a, MD10, SH12a, SL14b, VT13, WD13, Wil11, ZXF14]. discretizations [CKN13, GPZ12, HS11a, RGL12, SY13b, SGK14, TS12a, Wan13d, ZgJhD12]. Discretized [CPS14, VJBDV13, dlS10]. discriminance [HW14a]. discriminant [LZW12b, PJFL14]. discs [Sab13, Sab14]. disease [LL14b, LG10a, LWB14, LWW12, LSJZ13, MMP^{+}12, SG11, SC14a, SM14b, WQ11a, XGZW13, XZ13a, ZX13, ZJLO14, ÁME^{+}12]. Diseased [SP14a]. diseases [Sam11, WC14a]. disequilibrium [AKMM12]. disjoint [CHF14, JM14b, ÖKU12]. disjointed [Yan11d]. disk [LMS11, Mat14d, SMN11, Ste10a, Ste10b, Ste10n, SS11e]. disks [Gal10a, GG11a, GG14a, Mah14a, VSV10, WS10]. dislocation [Shi14a]. dislocations [ZSMB14]. dispatch [YJCC14]. Dispersal [HL11e, DMGK13, LLY14a, XTL10]. dispersion [EdCVB^{+}14, Hu12a, JB10, Kha14c, LS10c, LGKM13, LDZ10, SS11a, SMR11a, SGDK11, SWX10b, VDOG13, WLZ11, WD14a, YlTA10, ZS13a, ZTW10, ZLZ^{+}14a, ZTF10]. dispersionless [YLGW14]. dispersive [Bai13, LNG12, LLR10, NZ14, SDL10b, WX13d, Xu11, Yan11a, YDT^{+}14, ZS10b]. displacement [LLL12, Ngu14]. disruption [LDW14]. dissimilar [CW12, Sin11a, TSJ13]. dissipation [BD14b, Boy10, Dem14, Esf11a, GRV13, KM11a, Kau14, Kha14a, Pan15, SBS10, Sha12b]. Dissipative [SL13b, AISS13, BU12, HdCOL14, Kud13b, QW10b, QW10a, RZ11, SZ14a, Tun13, Tun14, XHXH12]. Dissipativity [MS13a, MK14a, WPSC11, FB10, WPSW13, WPSC12]. dissipator [dCDNCFRH10]. dissolved [Mis11]. Distance [EVE12, BR14a, ÇM13b, DSPZ13, JM14a, LAR13, Lin11d, LLK13, LS11b, MR14b, TK14a, WE13, ZZW^{+}12]. distance-based [ÇM13b]. Distance-sum [EVE12]. distances [DC13a, WC11]. Distortion [CHYT12, KOP11, MN14b]. Distributed [GLHCLP12, ZL14d, AS14b, AGNAL14, BC12b, BAE13, CHK10, CZZ12, DS13c, DDWW12, DZZ14, EGO13, JDN14, KHD12, LPJB12, Li10a, LZZL10, LTY11, LZWZ14, RLW14, Shk14, SHZ14, SFP^{+}11, TXX14, WL14b, WPKE14, ZMZS12, ZLY11c, yZxGyM10]. Distribution [TRA13, Tsa13b, AŞK10, BM14c, dSCAM14, CL14e, Fuk14b, Gar11, GC11d, GT12, HJZC13, Imo14, JHZC13, KM11c, KVP14, Ma11a, MS12a, Man12b, MLG13, MSS14b, Nad07, Ozd11, SPA12, SL14a, SRA10, SSS13a, SSTM14, TLDZ13, WN12a, WN14a, WHL14, YY14a, YX12a, YK11b, ZZY14b, ZANA12]. distributional [KM12a, TQ11]. distributions [COPP14, GGFM14, GVMNS13, GKS10, HL11d, JS14c, LL14c, LW11c, MC12b, MS13g, NOS10, RAW13, Seg14, Sha13b, SLW12, Tom13, ZZ11a]. disturbance [HYJ14, JJPW13, JLLD13, LDL11, LZ10f, WS12, YCZ14, ZW14b]. disturbances [AA12a, HSS10]. dither [Cha13a]. divergence [Duk12, NMR14, WL11]. divergent [Zhi12]. diverging [dOL13]. Diversified [YLC14, Aha14]. divide [Wu12c]. Divided [XC13, Nov14]. dividends [RA12a]. dividing [GVT^{+}13]. Divisibility [TL10, Ste12b, TLL13]. Division [Lu12b, Asl11b, BC12a, SYl14]. divisor [TL10, TLL13]. Dixit [TES14]. Dixon [FWZW10, YZZ12]. DMFC [SR12]. DMUs [WYL10b]. DNA [CGVR14, IN10, SS14b, WTH^{+}14, ZS13a, ZCL13a, ZCL14a, ZCL14b]. docking [LPP^{+}12]. DOF [DZ14]. doi [Asl10d]. doi.org [WHS14a]. Domain [HZ13, AC10a, AP11, AL10a, Bah13b, BCGV11, CNRY10, CRG11, CKSV14, CS13d, CM12c, DR11a, DGSB12, DLS13, EHRV12, FHWZ12, FMSGS12, FBGH12, GS11a, GMR11, JHP11, Jun10a, KPC14, Khe13, MZ13a, NLLC10, Zhu10d, ZH13, ZnXwY14]. domain-decomposition [BCGV11]. Domains [Ste13c, Cel14, CFH^{+}13, DM10b, DC14, DW11, DHMU14, ESEEKH12, JT13, JM14b, KS11b, KCLC13, Lim12, LGN10, LXLY14, Ros11a, Sab13, Sab14, Sar12c, Ste10h, Ste10i, YS13b, ZB12, Zha10f, Zit11]. dominance [CK12b]. dominant [HB11b, tLpOjCW10, LHZ10, MH10a, WWLL11]. dominated [ABE^{+}12, CFX14, EPMU10, HJVC14, SW10a, SHZ10]. dominating [CHM^{+}13]. domination [NPA11, YGZW14]. Doppler [AWS10]. Dormand [JSSB14]. Double [CYM13, DAV12a, DAV12b, DJY13, LLS14, MK13b, DAV14b, SZ14a, Alo14, BP11a, Che11d, FH11a, HRL14, Hou14, Kia14, Koç14, Li14b, LGS^{+}11, LR14b, Mak14, zM10b, MM11b, PS13, SRA10, SSM11, VG14b, Wu12c, YTCY12, ZYP13, ZW13c, Zhu10c]. double-bottom [YTCY12]. double-diffusive [Che11d, SSM11]. double-walled [Kia14, SRA10]. double-Wronskian-type [LGS^{+}11]. Doubly [SK11b, AR14b, CRR^{+}13, FK13, Hu12b, JDJ11, JDJ12, KKVCS12, Lin11a, LDLM13, WWLL11, WqXlC11, Xu13c]. doubly-periodic [FK13]. down [CPC11, SSS14]. downdating [LYS12]. DP [FYYT11, XW10]. DQ [Tom14, Tom15, Tom12]. DQ-based [Tom14, Tom15, Tom12]. DQMM [Gu11]. Dragomir [PS11]. drainage [ARH^{+}12]. drawdown [YLMZ13]. Drazin [BZT11, BFB12, CI11, GD10b, HSC14, Kyr13, Kyr14, Li11a, LZ10d, LXY10, LY12c, MD12c, MD13c]. drift [GO13]. drifted [CTLD12]. drilling [SHBN13]. Drinfel’d [Hu12a, MG10b, SV10a, SV10b, SV12, ZTW10, BD12b, MK13b]. drive [XWMW12]. drive-response [XWMW12]. driven [Amm14, BLZZ14, BW14, GARK12, HR11, Hu12b, HRX14, KLW14b, LS13b, LCL14, wMwWsT11, RMMM11, RCS14, Sun14, Zha14e]. drop [NNPH14]. droplet [HBB11, MVGS11, Sam11]. dropout [WY11b]. dropouts [WY12a]. Drosophila} [DS14c]. drug [CCJ10, LSJZ13]. drug-resistant [LSJZ13]. DSC [SWLP13]. DSSH [SV10a]. DTM [EÖ10]. Dual [Ker13, KMH^{+}13, UJH14, Ach10a, AG14a, Bal13, CS10b, GS14c, Jar10, LW11a, LC14d, MSR14, SZW14, SK10a, SNZL13, Woź12, ZTD10, ZGY14]. dual-phase-lagging [MSR14]. duality [AZ14c, GV13, GK12c, GS14d, JSMA11, Kas12a, MN11, PN11a, SGN13, TD13, Tri14]. Dubrovsky [SF11a, XGG^{+}11, Zha10b]. duct [RNEA14, XK14]. Due [GM13a, ZHC^{+}14, ASM11, AHMS14, AM14c, AWHW11, Cor11b, dCDNV^{+}13, JCB11, Kha14a, Mis11, PAA14, XWWC10, YCC^{+}12]. Due-window [GM13a]. Duffing [PS14a, DL13, EZ12, EZ14b, GjLxF^{+}13, GL10c, LR13b, LO13, MZX10, NA10b]. Duffing-van [PS14a]. Dufour [HCCT14, PC13a]. Dugdale [LF11b]. Dullin [MHLL11]. Dunkl [Suc14]. duopoly [CN14a, FXD12]. duration [ZYZ14a]. durations [KMC12]. during [GCG14, GSRR13, NATQ14, WY11c]. Durrmeyer [VGA12, CMGR14, CMG14, DB11b, Ere11, Gal10a, GG11a, GGS13, GA11b, Mah14a, ZLL10]. dusty [GGYW11]. dwarf [EN11]. dwell [XX14b, ZCZ^{+}13a]. dwindling [Gu13, GZ14b]. dyadic [MO14]. Dym [MV14]. Dynamic [AS11b, AS14c, AA13, CW11a, DG10a, Hu10, KG11, Kha14b, KKSS14, LXZX13, MCWC13, MT13, MP11b, RFK12, SVGC10, TPT12, XDG12, YT11, YWW^{+}14, AKR14, ABT^{+}12, AZL13, ABLZ14, AM13b, AABV12, BEP13, BAE13, BFM13, BK12, Cha14c, CkPqL12, DK11a, DGS14b, DT14, DI13a, DV14, EM12, FZ12, FA10, GSH12, GM14b, Has11, HMH13, ICKV14, Ito13, JYF13, Kar13b, KA13a, KDS13, KBG11, LFLS14, LJPJ12, LPJ^{+}12, LYPC12, LH14a, MG14a, MS13f, NTPVNHLX14, NWNL10, PKD14, PD12, QGD14, RKA^{+}13, SSA14a, Sit14, SS14e, SVV13, SDB14, Ste14a, SH13, Vít14, WMH12, Wan14e, YWLX11, YY13b, ZW10a, Zhe13, ZDC12, ZnXwY14]. Dynamical [BTVC10a, CST14, HT10, LX14c, LTOX11, SL13a, WWZ14a, Zen12, ZS14b, ZS14e, AMV12, AFSMRL14, BVTC13, BZZ14, EDYSA12, EM12, FWZW13, GC12, GC10a, GL13c, HYZH14, IT12, JZ13, JJP^{+}12, KK12b, Koh13, KJW10, LJPJ12, LPJ^{+}12, LWP12, LYJP13, LLY11a, LZQ11, LFR14, LLJF12, PKP^{+}12b, Rec14, SJ12, Sam11, San11a, San12, SS14c, SS10e, VS13, WMJW11, WZWW13, WXWY14, WLY^{+}12, XWMW12, YS11a, YS11b, YZZ^{+}14, ZLC^{+}13, ZA14a, ZL10a, Zuf12]. dynamically [Sha14c]. Dynamics [BZR10, CTV13, Gaz12, GW13, JBW14, JSH10, JV13, KLW14b, Li12c, LW13b, LTX11, LYL13b, Öca14, Qiu11, QLWW12, SZ14c, WD13, WW14d, WS10, YZC13, ZGH10, ZHXM14, ZLC10, ARG14, Amd10, AAA13b, AKMM12, BBB12, Ban14, BC11b, BKP14, Bel11a, Bel12, BBF12b, CG14a, CN14a, CJK12, CRC13, CM12b, Che13e, CGO14, CCGT13, CLND12, CGGZ11, DRSG14, DMGK12, DBS12, DLGJ14, EMH10, EAE13, ES12b, FXD12, FHW13, Far14a, FGS14, HYT13, HSX10, HYZH14, IKH11, IC11, Jan13a, LXC14, LSW^{+}14a, Lis14, LZG10, LSJZ13, LZLS14, LS13c, Mag14b, MAB14, MS14b, MB11, MG14b, MV13, MPB13, MPB14, NRWF12, NP14, NAL12, NNG10, NN14, PP14a, PRT11, PC14a, Qiu10, RRLA14, RM12b, RTVV11, SC14a, SSP11, Sha13b, SDL10a, Sie11, Sie14, SS12b, STV12, UW10, VDL14, Wah13]. dynamics [WY11c, WC14a, WPKE14, Xu12b, Xu14b, YLWU13, Zak10, ZS13a, ZSMB14, ZL14a, ZMZS12, ZX12, ZMC10]. Dziok [Aou10, Kir11, XXS14, YL11b]. Dzurina [SM06].

E-commerce [Sev11]. E.A [ANR13]. E.F [Pan12b]. Earth [BR14a]. earthquake [KVP14]. eccentric [Mah13]. Eckart [HMIZ13]. eco [Liu11e, Muk10]. eco-epidemic [Muk10]. eco-epidemiological [Liu11e]. ecoepidemic [SP14b]. ecological [GLSV10, MHTH12]. ecology [DRSG14]. Economic [YWK11, AKMM12, CC12b, DHSZ10, JGM13, TCBL11, WWK^{+}14, YJCC14, ZZC10, ZCZS12]. economy [LZLF10]. ecosystem [FHW13]. ecosystems [MB10]. EDAs [PHB14]. Eddy [PHPD11, KCZK11, SM10]. eddy-current [SM10]. Edge [BGV10, Ni13, ZZR11, AAMB14, BS14a, CLTL13, CEXG12, KK13c, LMY12, Ou10, PP12b, PP13a, PP15, PFS10, Shi14a, SMPA13, SSG13, Sze10, Sze12, Yan10c, Yan10d]. edged [WS14]. edges [DD11, WW13f]. Editor [Pol11]. Editorial [Ano11, Udw10, Ano12]. EEG [KUTY14]. Effect [AAMAS11, Gad11, Goy14, KM11a, Lac14, LLHY11, LYP14, MM13d, Pan15, SRS^{+}13, SC14a, Sha12b, SD13a, WSD13, AMM11, BMHM14, BD14b, BRVR12, CPC11, DMGK12, dCDNV^{+}13, GVMNS13, JSH10, JCW12, Lin12, Mam13, MMP^{+}12, Mis14, NKM^{+}10, OADL14, PP14a, Rud12, SMS14, SBS10, SW13, SGC10, WWW13d, WDS10, Xia10a, XLS13a, Xie11b, XG10, XYL10, Xu13b, ZW14c, ZWLW10, ZM10c]. Effective [PTS11, BIT12b, FYL13, HY11b, LPX12, LYG14, Tur14a, Tur14b, WY10c, YCK13, YLTB12, ZCL11, ZL10c]. effectors [TFMB12]. Effects [BJB11, Gad14, GLR12, Gee12, MWAM14, SP14b, SP14c, SJL10, AAV10, AHDD13, BD11, BSP11, CS13c, CWCW11, DIAQ13, EMAM13, GSF12, GP11a, HCCT14, Kay11, lLW14, MG12b, MG14b, PC13a, Sha11b, SSST14, WYTL14, Yan10e, YK11a, YLC12, ZLJZ14, ZWL14b]. Efficiency [LY14a, Wu14a, ABGGS13, AP13, Che12a, DGCFM11, dCDNRH12, LW10a, MTSN14, ÖB11, PRK13, SW14b, Ver13, Wu14b]. Efficient [AQ11, ÁB14b, Bez13, CMT11, CHMT11, Dim13b, DAEY12, HZ14a, HL11c, KX11, Koy12, LV14, OGO14, PMM11, SIDR11, Sas12, SIRD13, Sha13a, TZYH14, XLY14, Ach13, ABCD11, AAS12, Bhr14, CC11b, CM10, DWLL10, Dua10a, GSN12, JE11, JW10b, KKV12, LZZ11, LWL12, LSW14b, LHL^{+}13, Maa12, MLM^{+}14, MCR11, Mol14, DFTQOB14, RY14a, RTW13, RA12b, fRbLJyS13, RWBA13, Res11, Res13, RRSPTR12, SN12, SGG12, SA13, SAP14, TZZX11, USAF14, Wan11b, WDZ12, WZJ13, Yan14a, YLMZ13, YWCH11, Yus10]. effort [LZLY14]. EFG [HB14]. EFGM [HB14]. egress [DV14]. EHTA [ZDH10]. eigen [LNPS10, LHTN13, LWN14b, LfXfC13]. eigen-problems [LNPS10, LHTN13, LWN14b]. eigen-triplet [LfXfC13]. eigenfrequencies [OR14]. eigenfunctions [BEGM12, KM12b]. eigenpair [MXH10]. Eigenparameter [BK10a, jAS14, KY14]. eigenparameter-dependent [jAS14]. eigenproblems [ABD^{+}12]. eigenprojection [HLTU12]. eigensolvers [GPZ12]. Eigenvalue [Car13a, CN12, FFB13, GG14g, Hil12, JY12, Jia13b, OO11, Peñ13, Peñ14, Plo14, ASST11, BDR12b, BY11b, DWC11, HL13b, HDT11, HLW13, Jia13a, JSF14, Kol14b, Len14a, LZFR10, LLW11, LfXfC13, MP14b, NB11, OO12, PN11b, PP10, Wil10a, Wil10b, Wil13, Wu12c, wX13c, YDZS11, YH13, YGZW14, ZLW12a, ZS12, ZYP13, ZLWW14, ZY14b, ZJ14a, ZCZ13b]. Eigenvalues [AP10, BEGM12, Kat14, SW12, AdFY13, Ant13, Chi10, Cra12, EGMPMRBC12, GWS13, Gem11, Hla13, IS11, KM12b, LPAE12, LH10a, LW14b, XQ11b, Zho11b]. Eigenvector [QW11b]. Eigenvector-free [QW11b]. eight [GAC12]. eight-node [GAC12]. Eighth [LW10a, ACCT14, CL13b, CN14c, GK10b, NCS14, Thu10]. Eighth-order [LW10a, ACCT14, CL13b, CN14c, GK10b, Thu10]. eikonal [ACG13]. Einstein [HL14b, PM12]. EIV [ZZ15, ZZ14c]. Elastic [WY14, AGM14, Bas12b, CLW11, CW12, Che14j, CMS11, DB11a, ECG14, Gal10b, HMMK11, Li10b, LWZ13b, Lot14, MNOB10, OADL14, Shi14a, Sin11a, VS10, WG12, ZZHZ14]. elastica [Ito13]. elasticity [Cha13b, Che11g, Dan14, ECS13, LSX10, Li13, MR13b, Tra12, Vod13, Wei12, XYC11, ZL14c, ZLG^{+}14]. elastico [SGC10]. elastico-viscous [SGC10]. Elasto [BHG10]. Elasto-plastic [BHG10]. elastodynamic [DWZW14]. elastoplastic [CKSV14, HB14]. Elbarbary [Pan12b]. Eldabe [Pan12b]. elections [Tsi11a]. Electric [WMJW11, BAO14, LY14c, Ver11, YWW^{+}14]. electrical [KR11]. electrically [YS10a]. electricity [FD12a]. Electro [MSCK11, LWZ13b]. Electro-osmotic [MSCK11]. electrocardiogram [ZDH14]. electrode [QS14]. electrohydrodynamic [DG10b]. electromagnetic [HN10a, LZ13a, PPC13, SSK13]. electromagnetics [DLS13, GWLZ10]. electron [Awa12, RVS10, ZL13b]. electron-positron-ion [Awa12]. electroseismic [SZG12]. electrostatic [AAAH12, KG14, RS12]. element [AAP10, AY11b, Apa10, ASKA13, BC11a, BP11a, BDR12b, BY11b, BH11, BA14b, CYZ14a, Cha14c, CBO12, CB10, CZZ^{+}11, CF13, CS14a, CZX^{+}14b, CRX12, CZL14, Cho11a, Cop14, CRW11, DB11a, DS12a, DX10, DU13, DLS13, DHMU14, EJ12, EAA11, EVDG12, FBGH12, FR11, GT13, GDH10, AA14a, GPZ12, Guo11, GZF12, GFZ13, GZL14, HDK^{+}10, HJVC14, HH14, HMMK11, HOMA13, HL14b, JKN11, JMY13, KCZK11, KP14b, KAK11, Kim12, KSS12, KSM14, KP11a, Lee12, LJ12b, LZ12b, LC14a, LR14a, LL12e, LFLH14, LLS13b, Maa14, MGMM14, Pan15, lQZWC14, RDKU14, SZG12, SA12a, SBS10, Sha12b, SZ11, SZS13, SP13a, SY13b, SZW14, SCC11, SCK13, TS12a, Tem12, THNTNXR11, WRZ12, WW13e, WD14b, WFZ13, XSL10, XHXH12, XDZ14, Yan12b, YY13a, YLS13, YJ14a, YXZ10, YWL13, Yua13, ZLA10, ZF11, Zha11c, ZZL13, ZL14b, ZLG^{+}14, ZCZ14]. element [ZL12c, ZTX10, ZYSQ11, ZCC13, Zhu11c]. element-based [EVDG12]. element-free [CZL14, GZL14, ZL14b]. element-local [XDZ14]. element-Ritz [EJ12]. elementary [GH10d, JK14b, WiN11]. elements [BCI13, Kha14c, LGN10, LNG12, MD12a, MD12b, NN13, Ném10b, RV13a, XYC11, YS12, ZCD10]. elevated [BSZ13]. elevation [DLL14]. Elgazery [Pan12b]. elimination [ADGP10, ACQOR11, JC14, Peñ13, Peñ14, SP13b, YZZ12]. ellipsoidal [JS14a]. Elliptic [KS12c, SK10a, AT10a, AM14d, AO12, AS14e, AZAK10, BDR12b, BA11, Bat14, BEdAdSK14, BDK^{+}14, BA14b, Boy12, CCJ14a, yCfW12, CLW14, CL14c, Chi10, Con13, Cov11, DK11b, DAEY12, DHMU14, EZRMR12, EZ14b, FKS13, FV14, FZP12, FKH14, Fuk13, GMMIDM13, GS14b, GW11b, GFZ13, HH14, Hon10a, Kim12, KSS12, Li11b, MCL14a, Mit13, Neu12a, NSYA14, RDKU14, Ros11a, SK12a, Sha11a, Vej13, WL11, WS14, Wen10b, WFZ13, YL11d, Yos13, YY10, ZX11a, ZYSQ11, ZhGQ14]. elliptical [AAP10, Che10f, GL11b, LK14]. ellipticity [RZ11, ZZ14d]. Elsayed [Pan12b]. Elshehawey [Pan12b]. embankments [NCANCF12]. Embedded [SMIS14, CHC12, HCCT14, Kha14a, Pan15, PMD14, Sha12b]. Embedding [CKH11, WFZL12, Yan10d, DD11, HZFZ14, WLJ11, Yan10c]. Embeddings [ABG11, Lin13a]. Emden [Ben14a, BF12, Gal11b, GS14f, IJ11, KM12c, LML12, MC14b, PKBD12, TM11, Waz11b, WRD13]. emergence [Car12, EN11]. emergency [PSJGLS14]. EMG [MPB14]. emission [AMM11, Mam11]. emotion [YWYY12]. Empirical [AGNAL14, MMCD12, KRJ11, STM10]. Employing [Rem10, MMM13, SL11b, WL13d]. empty [KG14]. EMQ [SMS14]. enclosure [AANA10, KRMM14]. end [BJS14, DE11, MBR10, PHPD11]. end-of-life [DE11]. endogenous [BLR13a]. endoscope [NA10a]. endowed [ĆARA12]. enemies [JDN14]. Energy [Kan12b, MM12, CRW11, dCDNCFRH10, DFGL14, EMAM13, GR11, HTX^{+}12, IB11, Ito13, KBAvBW14, Kam14a, KR13b, Liu13b, MM10c, SHAO14, SW14a, XC14]. energy-stable [KBAvBW14]. Enforcing [CRW11]. engine [AMM11]. engineering [KLBR13]. Enhanced [SSV13, YZ14b, ZLZ14d, MT13, YG13, Esm12]. Enhancing [KPBC14, Lin12, TS10c]. Enriched [HDK^{+}10]. ensemble [TSP10a, XZZX14]. ensure [BM12d]. Entire [LWB14, WW13g, Liu12b, Pat13b, RD14]. entries [İpe11, Neu13, ZLZ14c, ZJ14c]. entropic [CGAK10]. entropies [SMDZG13]. Entropy [Che11d, CDESS14, KRMM14, YK14b, BFH12, DM10a, Duk12, HLWH12, JZY12, LDR10, Mal14, MIT12, SSCA13, YLC14, ZL11c, FT11, Tag14]. entropy-based [DM10a]. enumeration [RBT14]. envelope [PT14b]. envelopment [Lin11c, Lin11b, MTSN14, ÖB11, WY10c, ZF10]. environment [DS14d, DWC13, KHA12, LC12, LZTZ13, LC14e, LH14c, LF14, QK13, ZW14c]. environmental [AAV10, SW14b]. environments [CCO12, Ni13, OSRPC12, PW14, TDSA14, URR11, XTL10, YY13b, ZWL14b]. enzymatic [YZZ^{+}14]. EOQ [DGS14c, GLSJCBS10, KGC11, San11a, San11b, San11c, San12, Sar12a, WT11a]. EP [BR12, HLTU12, MD12a, MD12b]. Epidemic [pZJ12, BR10c, CL10a, CGW14, CCS14, CS11b, CC10c, DMS11, DIAQ13, EMM^{+}12, Fan14a, GYL^{+}13, GHC13, HLM14, HLW12, INE10, KN12, LOKB12, LS13a, LSJ10, LTY11, LWW12, Liu14b, zM10b, Muk10, NM14, NSL14, NTS11a, Qiu11, RG14, SGE13, SRS^{+}13, SS14a, SSL10, SSL12, SSL13, Sha13b, Sie11, SS11i, WW13g, WWZ14a, WL14b, WC14a, Xu12b, YJ14b, YLC12, ZJS10, pZJ11, ZHXM14, ZJ14b, ZZL14, ZZY14b, dlSAQ11]. Epidemics [YYL^{+}13, Goy14, KK14b, LWG13, WC14b]. epidemiological [CSSSV13, Liu11e, PFG^{+}14, PW14, Xu14b]. epileptic [KUTY14]. epistemological [Car12]. EPQ [CBTGWW14, GLSJCBS10, SM11, WHWC14]. EPR [GG14d]. epsilon [Tom13]. Epstein [Lis14, WZ13b]. equal [CFZ14a, DS13b, lF12, FPP11a, HLM14, Ros11b, Udd13]. equalities [GZ12a, Hu10]. equality [BBT10, DJ12, GLV10, LL13a, SA12b, Wri11, ZZ14c, ZZ15, ZP13]. Equation [PB14b, Sie11, Aba12, ABEK11, AMB^{+}14, AK10a, AOAA10, ABP13, Afz10, ARE14, AHJ14, AYS13, ASVV10, Ali12, AAS12, ÁD10, AANA10, ACG13, Aou11a, Aou12, ANS14, AS13a, AM12d, Asa10a, Asa10b, Asl10b, Asl10d, Asl11a, Asl12b, AMB10, Awa12, BP10a, Bah13b, BP10b, BBJ10, BP11b, BB14a, BKP14, Ben14a, Bes10a, BTVC10a, BVTC13, Bho10, Bhr13, BW14, BC10b, BCW13, BZR10, BI10, BK10d, BK10b, BZGK10, BT10b, BPM10, BR10b, BT11b, BTHA11, BCKL11, BD10, BC12d, Bou10, Boy10, Boy11, BD14e, BT10c, BGGH12, BPX14, CST12, CI13, CT13a, CQJ10, CC10a, CL11a, Cai11, CL14a, CNRY10, CRG11, CP12a, CT12a, Cel14, ÇBS14, Cha11, CBO12, CJM14, CCCRR13, CS10a, CLL10, CJ11a, CLAT11, CX11, CLW11, CLA^{+}12, CW12, CSZ12, CYM13, CLTA13, Che13a, Che14j, CC14c, CLLS14]. equation [CLL14, lChH14, CL10b, CCL12, Chi14, CS11d, CC12c, CP10, CK13a, CYZ13, Cui14, DAHAK10, DLFZ10, yDpY14, Dar11, Dar12, Dar13, DS13a, DS14b, DKS14, DVGY11, DGDA10, Dea13, DJ14a, DE12, Dem14, DPC10, DX10, fD12b, DL14, DS13b, DRS12a, DCR13, fDxZxChT12, DP10, DLL11b, DAA14, Ech12, EN11, ESH10, Elb12, EBR13, Esf11b, Esf11a, EALS10, EKZ11, FYYT11, lF12, FYM13, Fan14d, FLW11, Fer10b, FLP11, FP12, FCZ^{+}10, FL10, Gal10b, Gal11b, Gal11c, GGP^{+}13, GK14, Gao10, GZ10a, GLZ10a, GLS13, GC14, GLR12, GS14a, GZ10b, GH10b, GSH12, GDH10, GS10b, GP11a, GAV14, GPP11, GARK12, GWH12, GT11c, GADK14, GG14e, GLB12, GÖS11, Gün10, GZXJD10, GZ10c, GL10b, GMZL11, GW11b, GLW12, GML13, GL13b, Guo14, GZL14, GB13, GMSE13, HK11, HB11a, HXY12, HBTU10, Har10, HA10, HMIZ13]. equation [HBB11, HLMW13, Hay11, HML10, mHhL10, HM10a, HRL14, HSZ14, HSX10, HL12d, HM10c, HHX11, HHX12, HLX13, HX14, HYSZ14, HCD14, HKH10, HH13b, Hus10a, HS12, IKV11, Ibr14, IS10a, IS10b, IS12, ITLA10, IKU10, IU10, IJ11, IB10a, IPB10, IB10b, IRS14, Jan14a, JPB10a, JPB10c, JY12, JH10, JHX14, mJ10, JB10, JZ10, JLM^{+}12, JW12, JS13b, JT13, JLZB14, JZW12, JPM12, JMS13, JKB10, JK10, Jor10, JL10, Jum12, JBDC12, KK14a, KZJ13, KLBR13, Kau14, KA10a, Kha11, KS11b, KL14b, KS10b, KS10a, KE11, KM11c, Kle12, Koç14, KV13, Kon14, KCG10, Kos11b, KLW14b, KCC12, KBB12, Kud10c, KRS10, Kud13b, KSM14, KP11b, Lai10, Lat14, LZ12a, Laz11, LJR14, LW13a, LMS10, Li10c, Li10d, LWL10, LF11c, LZW12a, LLXZ12, LS13b, xLhYlZ13, LY13b, LY13a, LHC14, LC14a, LSLL14, LGZ14a, LZLB14]. equation [LZaY14, LSL10, LWCS14, LYD10, LK14, LWD10, rL11, LC13, Lin14a, LGN10, Liu10a, LLWW10, LJW10, Liu11d, LL11e, Liu11f, LX11a, LTJ^{+}11, Liu12c, LML12, LZW13b, LL13f, dLL13, LY14d, LZ14c, LLS14, dLfX14, LC14b, LFLH14, LC14c, LP13a, LM12, LO13, LLR10, LXZ12, LGKM13, LLL10b, LCLW10, LL10h, Lu12b, LDLL10, LW10b, LDLM13, LD13c, LL12f, MG10a, ML10b, ML10a, MAA11, MG12a, ML12a, Maa14, MKeM13, MDP12, MH12, MH13b, MH14b, Mar11, MCdSS11, MRE^{+}12, MR14b, MC12c, Mel14, MTZ10, MHLL11, MGYQ12, Mer14, MJ13a, MB13a, Mis10, MG10b, MJ12, MB13b, MB14, MESAN13, MMGM13, MK13a, MHA10, MYL^{+}10b, Mop14, MK13b, NMIA11, NC13a, NWZ13, NA10b, NK12, OSRFF14, Öca14, OM11, OPD12, Oul11, yOZ14, Öza14, ÖK11, ÖTC11, PSP10, PP14a, PZ12, PVV11, Pan12a, PZ14a, PSS11, Par10a, PHGV13, PS14a, PV11, PKS10]. equation [QW10b, QW10a, QMM13, QZS13, Qui10, Rad11, ROK10, ROK12, RV14, RB14, RNZH13, Rao14, RK14a, Ray12, RSN11, RM12a, RYÖ11, RLV10, RRM12, RB11b, Ros11b, RJ12, RCYL10, Rya10, S.14, SJ12, SvB10, SMR11a, SK12a, Sal10a, Sal10b, Sal10c, SHMCM12, SB13, SGDK11, SDC10, SATdV13, SSK10, Sch11a, Set14, Sha14a, Sha10a, SHY11, SXL13, She10, SLC^{+}12, SGY^{+}14, SDL10b, SZS13, SPK14, SMC12, Sie10, SGK11, Sin12, SGK14, SL14b, SMH14, SGA14, SYZ10, SS10f, SY10c, SC10b, SHZ10, SHC11, SL14c, Ste10e, Ste11f, Ste12q, Ste12e, SAAS14c, SB10, SXC^{+}12, SB11, SD14, Sun10, SW10b, SXH11, qSyGH11, SHM12, SHWX14, SV10a, SV10b, TM13, TM11, THH10, TMXG11, TTM12, THZF12, TSG10, hTlCqL10, TZJZ12, TW13, TS12b, TK12, TW10a, TI10b, THAB11, TTQ10, TTQ11, Udd13, UT13, Üna13, VDOG13]. equation [VDK10, VDK13, VS13, WDLM10a, WDLM10b, WWS10, WDL10, WLH10, WW10, WQX^{+}10, WZZ11, qWH11, WLH12, WL12b, WH13, WZJ13, WD13, Wan13a, WWH13, WY13a, WWG13, jWtW13, WDT^{+}13, WD14a, WCL14, WW14c, WYTL14, Waz10b, WT10, Waz10f, Waz10e, WM10, Wei11b, WSD12, WTFC13, WGW11, WW14e, WFwC12, WP14, Wu10c, WL10c, WFDL11, WW11f, WZL12, WD12, Wu12d, XC10b, XC10a, Xia10a, Xia10b, XZT10, XW10, XD11a, XHG11, Xie12b, XZH13, pXqLlZ10, XLZ10, hXqXlC10, XWZF11, Xu11, XZ13b, Xu14c, XZ13c, Yad13, YTS12a, YZD10, YlTA10, YDZS11, YL11d, Yan11c, Yan13b, YZZP14a, Yan14a, YZZP14b, YH14c, Yaş10, YZ10, YWCH11, YZ11, YYJ14, YS10b, YHW13, YL14b, yYyL10, YDT^{+}14, YZZ10, YM10b, qYqLtW10, YM10c, YL10c, YDTW11, YLTB12, Yu12b, YWL13, Yu13a, Yu13b, YCJ^{+}14, YY14b, YWD13, ZA11, Zay11, ZFH14, Zha10b, pZLC10, Zha10a, yZhLjMzC10, gZrLX10, gZrLfM10, ZCG10, bZjZS11, ZlDxX11, ZZZ11]. equation [ZLGZ11, Zha11b, fZwXyZ11, ZCT11, ZLW12a, ZgJlK12, ZW12, yZqLwW12, ZYP13, Zha13a, ZHT13, ZW13e, ZS14b, ZTZ14, ZLZ^{+}14a, Zha14a, ZDW10, Zha10d, ZZ10c, ZL12c, ZLLL12, ZZR14, Zhi11, ZYC14, ZS10b, jZzFbL10, ZTF10, ZLYL10, Zho11c, ZWT12, ZLZY12, Zho12b, ZCL13b, ZCZ13b, ZXF14, ZC14b, bZpZJ10, ZK10a, Zhu11a, ZL13f, ZYZ14c, ZR14, ZL10g, Zou12, JCB11, SS14b]. Equation-based [PB14b]. Equation-Free [Sie11]. equations [ABRT14, AB14a, AB11, eMA13a, AG10, ASM11, ASS13, Ach11, AC10a, Afu12, ABT^{+}12, ABLZ13, AZL13, ABLZ14, AS10, AISS13, AH11a, ADÇY11, AIAD12, AJF13, AZ14b, AH10, AT10a, AHRV11, AAT^{+}13, AHVR14, Ami14, ABF14, AL10a, AMW12, AM13b, AM14d, AS14d, Ans13, Aou11b, AGKR10, AAH14, Ard13, AAGR13, ACH10b, AS12b, AASS13, AAS14b, AM12c, ADP11, AO12, Asl11b, Asl12a, AS14e, AJPZ14, ABBC10, AZ19, BRS13, BD12a, BDR12a, Bah13a, BZ10b, BZ10a, BK10a, BP13, BM14b, BMA11, BPPS14, BEP13, Bar10, BLK11, Bec13, Bek10, BA10, BB11c, Bel13, Bel11b, Bel11a, Bel12, BTRZ14, BS11a, BS11b, BTVC10b, BDG13, BK14, BES10b, BTS12, Bie11, BR14b, BFM13, EL13, BRM11a, BRM11b, BMR13, BR13, BX12, Bra13, BPR13, BK11b, BK12, BC12e]. equations [CH14, Çak10, CCJ11, CMM10, CGGMG12, CGGMG13, qClLyW10, Cao10b, CYHL11, CLZmF11, CG12a, CC13, Car13a, Car13b, CPKT13, CK10, ÇM13a, CJ14, Cha11, CWS10, CL14b, CDMS14, CR11b, CPZ10, CB10, CZ11a, CLZ11, Che11e, CZ11b, CM12a, CT12b, CYM13, CLW14, CM14, CT14b, CFZ14a, CZZ14, CFX14, Che14h, CSL14, CL14c, CYY14, Chi10, CCGT13, CS13e, CD13, CK14a, CD11a, CGK12, CG14b, CP14b, CL13b, CYZ14b, CS10d, CSSN11, CDST14, CTV12, CTLD12, CJRV12, CSS11, CDCS13, CHS14b, CY12, CK12c, DGSB14, DW14, DOV^{+}11, DK11a, Das12, DBK14, DL12b, DSGB12, DGSB12, DN14b, DK11b, Den11d, DN12a, Dia13, Dia14, DS12b, DV14, DISS14b, DN12b, DL10, DLX13, DS14e, DAEY12, DBBEE13, DC14, Don14, Dos11, DK14c, Dou10, DJT10, DS13c, Dua11b, DR11c, DCR12, DSPZ13, DLLW13, DC13b, DAA13]. equations [DPP11, Džu11, DPP12, ESV12, EM13, EK10, ESEEKH12, EGJB13, EB14, EMP12, EZ13a, EJ11, EVDG12, EH14b, ES12d, EGH12, EZ13b, Fan10a, FZ13, Fan13, FDHK14, FEGEA14, FH11b, FK12, FcDDqZ10, FZP12, FLZ14, FFOC14, Fer11, Fer14, Fig11, Fig12, FP13d, Fra10, Fre11, FH13, FdBCC14, Fuy14a, GLP11, Gal11a, Gal11e, GB12, GMX10, GP12, GLZ10b, GX12, GPK13, GK10c, GB10, GKS12, GK13, GG14b, GPCCA14, GG12b, GSGN11, GT11b, Gu11, GS14b, GÖS10, GGÖS11, GÖ14, Guo10, GZ10d, GZZ10, GMZL11, Guo11, GWML12, GZW12, GZF12, Gür10, Gür12, GS14f, GMSE13, GS11d, GMV11, GH11, Han11, HLZ12, HL11b, HHU12, HR10, Has10, HKI12, Has11, HET12, HV12b, Hay10, HVC12, HSI13, HL14a, HH13a, HPB11, HOB12, HSW12, HHM14, HR13, HM10d, HR11, Hu12b, HG13a, HRX14]. equations [HW13b, HYW14, IN10, Ign10, Ign13, ICKV14, IISA10, Iri10a, Iri10b, ISG11a, IS10c, JVK10, JDJ11, JDJ12, Jan10, Jan11, Jan13c, Jan13b, Jan13d, Jan14d, Jan14e, JPB10a, JPB10c, JPB10b, JŞ11, JM11, JZZ13, JX10, JS13a, JZG13, JL14, Jia13b, JC13, Jia14, KEGH13, KC12, Kan13, KM11b, KM12a, Kar13a, KJ10, Kar13b, KG14, KSFG14, Khe13, Kho14, KA10b, Kim10, KAK11, KK13b, Koh12, KW13, KY14, Kos13, KL10, KT14, Kro11, KS11d, KD13, KAB11, KS10c, KS12c, KS12d, KK12c, Kud13a, Kum11, KKS12, KK13d, KM12c, KET10, KL13b, Kyc11, Kyr10, Kyr12, Kyr13, Kyr14, La 10, La 13, LZ10b, LWZD10a, Li10b, LLW10a, LMMW10, Li11b, LYZ11, LZLW11, LLJ11, LZ11a, LLY11b, LTZ12, LCH12, LL12c, LZW13a, LAF13, LDWY14, LL14e, Li14a, LH14a, LZL14b, LL10d, Lia11, LLL11a, LL11b]. equations [LYAH13, LKS14, Lia14, LTOX11, Lin11a, LY12a, LL13d, LLL13, Lin13c, LWH14, LTW10, LZZZ10, LDL10, LW10a, LZH10, LL10e, LL10f, Liu10b, Liu11c, LZKK11, LJKU11, LH11c, LL12e, LW12b, LTW12, LS12b, LJ13, LL13e, LM13b, LLS14, Liu14c, Liu14d, LSW14b, LT14, LZSD14, LW14g, LP13b, LR13b, LMB13, LLN10, LZX10, Lu12b, LO14, LHL^{+}13, LLS13b, LS14d, LXLY14, Luo14, LM11, LYL11, MWC11, MZTT12, MZ12, MH13a, MYW13, MZW14, MCH14, MC14a, Mah14b, MZ13b, MCM10, Mal12a, Mal12b, Mal13b, MC14b, MN12, Mam13, Mao11b, MAO10, MMO13, MN14a, MIHS14, Med12, MPS14b, MA13c, MH11b, Men11, MG11a, Men12a, Men12b, MAD14b, Mey14, MZX10, MM13b, MPSD12, Mil14a, Mil14b, MVS14a, MH14c, MHB14, MJ11a, MJPB11, MSC^{+}13, MG11b, MS12b, MG14c, MK14b, MSS14a, MEG10, MM10c, Mol14, MAHAS14, MN10, Mop11, Mos14]. equations [MV11, MS11, NV12, NCL11, Nar13, NMIA11, NR14, NSC12b, NS13a, NC13b, NCS14, NC14, Ngo12, Nie13, NRL12, NSZ^{+}14, NMDWAS10, O’R11b, OR11a, OM11, OO14, ONBN10, Ols12, OH12, Ort14a, Ort14b, OP14b, ÖZ10, ÖK11, ÖA10, hPrLgZ10, hPnZrL12, PRT11, PKBD12, PM12, PRS12, PEP14, PT13, PR12a, Par11, Par13, Par10c, Par10b, PT11a, PD12, PJ12, PW10, PX13, PAC13, PID10, PDN11, Pet13, PNPD14, POR13, Pin10, PPV14, PEMS13, QMZ14, QHL10, QTL^{+}12, RMY14, Rah13, RDKU14, Ram10b, RGL10, RGL11, RW14a, RAIM14, RS13, RWB10, RA12b, fRbLJyS13, RWBA13, RCS14, RHSC14, Rep11, RSRA14, RN10a, RWkPk10, RPA14, RZ11, RnZ10, RSK11, RC12, SSF10, S10, SS10a, SHT14, SNMD13, SuR13, SR14a, SAO12, SF11a, SS13a, Sal11c, SDAN14, SMN11, Sar12c, Sau11, SSZ11, SPV10]. equations [SMIS14, Ser13, SYK10, Sha11a, SH12a, SGG12, SA13, SRP13, SAP14, She14b, She12c, SZ11, SP13a, SY13b, Shi14b, SZW14, Shk14, Sho14, SY10b, SAK10, SF13, SS14e, SZ14c, SK13b, SKD12, SL10b, SW10a, SW11, SL14d, Sta14a, Sta14d, Ste10f, Ste11g, Ste11c, Ste11d, Ste12b, Ste12c, Ste12l, Ste12e, Ste12f, Ste12g, Ste12h, Ste12m, Ste12r, Ste12n, Ste12s, Ste12o, Ste12j, Ste12t, Ste12p, Ste12k, Ste12u, Ste13a, Ste13j, Ste13b, Ste13k, Ste13c, Ste13d, SAAS13, Ste13l, Ste13n, Ste13h, Ste13i, Ste13e, Ste13f, Ste13g, Ste13o, Ste14b, Ste14c, SAAS14a, Ste14d, SAAS14b, SDS14, Ste14e, Ste14f, SWX10b, Suh13, SW14d, SD13b, Sun11a, SH12b, zSWZW12, SCK14b, TMM13, TMM14, TKL14, TF10, TS14, TW14b, TD14a, TD14b, TA13, TFMB12, TH14a, TV13, THP14, Thu10, TYT14, Tom12, Tom15, TT13]. equations [Tri12, TW10b, TI10a, TW11b, TB14, Tur14a, Tur14b, VV10, VA12, VMRA14, VSM14, VS10, VZ11, VD14b, Vít14, Wan10e, WYL10a, WL10a, Wan10d, WGS10, WLY11, WW11c, WqXlC11, Wan11c, WTL^{+}11, Wan11a, WZ11e, WW12, WH12, WDZ12, WDL12a, WF12, WZG12, WW13c, WWZ13, WCZ13, Wan13d, WFL14b, WFZL14, WZ14a, WW14b, WMM14, WZL14a, WS14, WW14a, WWQ14, WYK11, Waz10a, Waz10c, Waz10d, Waz10g, Waz11a, Waz12a, Waz12b, WRD13, Waz13, Wei11a, WWW12, Wen10b, WG12, WFZ13, Wil12, WN11, WN14b, WZDW10, Wu10b, WLHZ10, WL10d, WLH11, WLD11, WC11, Wu12a, WE13, WL13d, WHZ14, WD14c, XX11, XL11, XH13a, XW13, Xia14, XP12, Xia13, XLY10, Xie11b, Xin10, XQ11a, XGG^{+}11, XD11b, XM12, XHXH12, XZ12, Xu13c, XL14, YWZ10, YP14, YDXC10, YL10b, YH10, YCTD11, YMG11, YL11d, Yan12b, Yan12a, Yan12f, YWW13, YZC13, Yan13d, YLLL13, dYyR13]. equations [Yan13e, YKL14, YLS13, YJ14a, YG11, YLAT14, YCK13, YH14d, YLZ13, YLGW14, Yos13, YCF13, YYY10, YuRI14, Yu11a, YY10, Yua10, YWCH14, Yue14, Yun10b, Yun11a, Yus10, YS13b, Yüz13, Yüz14, ZZ14a, ZH12, ZOA10, ZH10a, Zha10b, ZLA10, gZrLfM10, ZW10a, ZW11a, ZX11a, ZYC12, ZLS12a, ZLW12b, ZgJhD12, ZYZ12, ZSZ12, ZL12a, ZW13d, ZLWL13, ZWZ13, ZHY13, Zha14c, ZZZ14, ZL14e, ZLWW14, ZL14b, ZWS14, ZQ14, ZLZ14c, ZGW14a, ZS15, ZDH10, ZSL10, Zha10e, ZSHZ11, ZHZ12, Zha12c, Zha13c, Zha13d, ZF13a, ZYZ14b, ZL14f, Zha14e, ZYSQ11, ZLH11, Zhe11, ZMG13, Zho13, ZH14, jZzFbL10, ZMZ11, Zho12c, ZCS13, ZL13c, ZS14e, ZGW14b, ZX14, ZSD10, Zhu11c, ZWH14, Zuo10, ZZ11c, dFGAN12, dlS10, dS10, fuy13, dlRS14, uRE13, GL14a, Shi13a, SQJ11]. equidistant [CZSV12]. equidistribution [DN14a]. equilibria [TES14]. Equilibrium [FA12, SLL14, WZ11c, sCCLY10, sCCL11, COPR10, CQ14, CS10c, CV11, Din11, Din12, DHZ11, HYZH14, Inc12, KR12, KPK14, LWZ14, LMS^{+}12, Mel14, OS11, PK12, QCK10b, RK14b, RS12, SK11a, She12a, WHC10, WC14b, WY12c, YZK12, ZOS10, Soo11a]. equimaterial [Kos11b]. equitorsion [Sta14c]. equivalence [AMM11, AR12a, Sag11, hTlCqL10]. equivalent [Bra12, DS12b, PS13, Wag10, Wan14d]. equivariant [DLC10, LZM^{+}12, pSbL14, XH10, ZZ10d]. era [SA10a, SA10b]. Erdélyi [DS14b]. Erkus [LLSW12]. Erlang [SSTM14]. Errata [RZHH14]. Erratum [ESAH10, Gür12, MK10a, NN19, Pak11a]. Error [ÁD10, CS14a, DHM14, FGO12, Kau14, Kim11, LU12a, LXGM14, MP11a, RPW14, SATdV13, SH12a, TA13, WSCH14, AQ11, ABE^{+}12, And12b, Bac14, BA14b, Boy13, CZM11, CZZ^{+}11, CL14c, CYZ14b, CCSM13, DN14a, DT10, FLZ14, AA14a, HSW12, HB11b, HHDM13, Jia12, Kim12, LP10, LJT13, LC14c, MJ13b, RSK14, RV13b, SJPS14, TK13, TTQ10, Vej13, WRZ12, WW14a, Yan12b, ZF11, ZS12, ZZL13, ZCZ14, dFGAN12]. errors [CC10b, Gu10, JM10b, yL14c, dLfX14, Maa12, Pre11a]. errors-in-variables [JM10b]. essence [sL10a]. Essential [O’R10, Ste10d, SS11f, SSB11a]. essentially [FHW14]. established [Irm11]. Establishing [BTV12]. Estermann [Ras14]. Estimate [KS11d, CTH14, DN14a, Gar11, MJ13b, SH12a, WRZ12, WW14c, XXS12, dlHV12]. estimates [ABE^{+}12, Bac14, Bar13, Che13b, CS14a, CL14c, FLZ14, GG11a, GY12, Has10, HHDM13, JKP14a, Kau14, LW14b, LC14c, MR14b, Mea13, Mor10a, SL14d, TTQ10, WSCH14, XLS13b, XCS13, ZZL13]. Estimating [GG12c, Maa12, MCM13, Mor10b, RRRRTR12, ZANA12, AEESMA10, HR13, PE10, SS11d, ZF10]. Estimation [AWS10, CWW14, GRKP10, LZZP10, PB13b, ALNGMGMM14, AŞK10, BTT13, CÁGGLP14, CLTY13, CYZ14a, CZZ^{+}11, CXD14, CD14, DNSM11, ELS11, FC10, FL14, GLHCLP12, GC10a, HSX^{+}11, HLWH12, HL13a, HCLP11, HSW12, HB11b, IS14, JSWJ13, JM10b, Juk13, KUKU14, KR11, KG13, KGO10, LD13b, LXGM14, LLKP14, LP13c, MG14a, MLG13, MMM11, MSJ14, MGGS12, RSPK13, RTW13, RRLA14, SSL13, SW14a, SS12a, SSS13a, SM14c, SMPA13, WW14a, YDZS11, Yan12b, YY14a, ZJD12, ZCZ14, ZLY11b, ZFCW13]. estimations [DFGL14, JDLL14, MP14b]. estimator [GK11, HS14a, Kim12, LPJB12, LY14a, MS13b, RGA14, TK13, Wu14a, Wu14b, YK14a]. estimators [DP13b, GKV12, HS14b, Koy12, LTSV14, Mea13, NHCLP10, NCÁHCLP13, SAuAH14, VG14b, YK13a]. Estrada [RV10a]. Eucalyptus [NTdCJL12]. Euclidean [JM14a, KGH12, TK14a, TÜ12, YLJ14]. Euler [KK13a, KK14c, BE13, Cao10b, CM13c, DL10, EDG14, FPP11b, GI13, HW13b, JS13b, KK11a, LPH11, LM13b, Lu14a, Lu14b, Mil14a, Moh14, Mor10c, MD10, Ngu14, Ozd11, PRT11, SE12, SPV10, SCK14b, TH14a, UB13, Vít14, WYL10a, WFL14b, WD14c, YWW13, ZP12, ZL13b, ZLL14, ZSL10, ZXF14, ZWH14]. Euler-type [Vít14]. Eulerian [YL14b]. European [LW13a, RA12a, Xia10c, XLY14]. evader [SGT10]. evaluate [NTdCJL12, Occ11, WT11c]. evaluating [ASSPA10, BSKL13]. Evaluation [Che14j, LWW10, Mat14a, Tad10, AMA12b, AK10b, AGNAL14, Asa10b, CW11a, CA14, DZ12, EGMPMRBC12, GSF12, HWQG11, JBL^{+}11, Lin11d, LC12, Mam11, Meh13, MIT12, NAA13b, ÖB11, SFF11, TSN13, XX14c]. evaluations [CT13c]. evaporating [MVGS11]. evaporation [CO13]. evasion [BM14c]. even [ABLZ13, AKB13a, AKB13b, Rim13b, VD14b, Wu10a, Xia10c, YL10b, YKL14, ZH12]. even-order [ABLZ13, YL10b, YKL14]. event [SSCA13, SKIH12]. Eventual [SAAS14a, DIAQ13]. Evidence [Ben14b]. evidential [DZS^{+}14]. Evolution [HM10c, XMWD14, ZD11b, AH10, APA13, APJ14, AM12c, ARS14, Asl11b, BA10, BTVC10b, BT11b, BTHA11, BGGH12, CGTCF13, CW11a, Che13c, CEP11, dSCAM14, CY12, DW14, DN14b, DN12a, DN12b, DLX13, DG10c, ESE14, FDHK14, Fre11, Gal11d, GjLxF^{+}13, HYW14, IB10a, IPB10, JPB10a, JX10, KS10c, KS12c, Kum11, KET10, LL12c, LTW10, LYG14, LLN10, LM11, ML10b, MLGdSC11, MGYQ12, MN10, NMDWAS10, Par10c, Par10b, RK14a, RJ14, RHSC14, RZ11, RSK11, RC12, SMN12, Shk14, SS11d, SB10, TSP10a, TPA10, TI10b, WzH10, WMJW11, WZ11b, WH13, WY13b, Wu10b, XMA14, YL11d, YuRI14, ZZW^{+}12, ZGW14a, ZDH10, ZZR14, Zhe11, jZzFbL10, AGNAL14]. Evolutionary [ES12b, JYYJ10, LNdCJV14, Ach13, ASSPA10, AGM13, Car12, DWY14, ESE13, EL14, GPP^{+}10, HDZW11, Hua11a, MR14b, SNH10, XMA14, ZDH14]. Evolving [FB10, BDDK14]. Exact [AMB^{+}14, AV14, Asl11b, BA10, BAP^{+}14, BMMS13, CQJ10, CC12c, DLFZ10, DPC10, EZ12, GLZ10b, GL11b, HLM14, Hay11, HSLG10, IS10c, JKB10, JBDC12, KA10a, KW13, KCG10, KL10, KR14, LF11b, LK14, LL11e, Liu11f, LTW12, MCM10, MG10b, MESAN13, QW10b, QW10a, RRM12, Rya10, Sal10b, SPV10, Sha10a, Shk14, Sin12, SYZ10, SS10f, SY10c, SB11, TJZ14, WWS10, WDL10, XCH14, Yan11a, YM10b, qYqLtW10, ZA11, bZpZJ10, Ant11a, Ant12, Asl10d, Awa12, Bac14, Bes10a, BTVC10a, BPM10, Bou10, CG12a, CJ11a, Den11d, Fit12, FM14, FCZ^{+}10, GZ10a, Hay10, HM10a, HRL14, HVC12, HW14b, Hon10a, HL12d, HYW14, IU10, Kud10c, KS12d, KK12c, Kud13b, Li10c, LYZ11, LTZ12, LS13b, LZaY14, jL12, LWD10, LZH10, Liu10a, LLWW10, LJW10, LLL12, LY14d, MC12b, MR14b, MF13, MHLL11, Pin10, dFPF14, PPV14, RK14a]. exact [RnZ10, RSK11, Sal10a, SHY11, Sie10, Van10, Vir13, VDK13, WDLM10b, WM10, Wei10b, Wen10a, XC10b, XC10a, Xin10, XLZ10, Xu11, YH14c, YZ11, Yu13b, YHS13, ZOA10, Zha10b, pZLC10, Zha10a, yZhLjMzC10, gZrLX10, gZrLfM10, ZCT11, yZqLwW12, ZW13d, Zha10e, ZZ10c, ZTF10, ZL10g]. exactly [Ram10a]. exaggerated [Koç13]. example [GJGSNRÁ12]. examples [ADGP10, DK14a]. Exceptional [HM14a, Ném10b]. excess [AH11b, THCW11, YJ10]. excess-of-loss [YJ10]. exchange [And12a, AM12b, BMI^{+}11, DW13a, LSF^{+}13]. exchangeable [DF14b]. exchangers [Che14k]. excitable [Alf10, Alf12, ZLZY12]. excitation [Jan14b]. excitations [Bai13, Bai14, PPMCG14, Zha12b, ZC14b]. excitatory [ZGH10]. excitatory-to-excitatory [ZGH10]. excited [BZZ14]. Existence [ABRT14, AKR14, AIAD12, AFSMRL14, Aou12, AAH14, APJ14, BBJ10, BR14b, Bot12, CI13, CB12a, CJM14, CT11a, CT12b, CLW14, Cov11, DH13b, Den14, DN12a, Dia13, Dia14, DP10, DX14a, Du11, DH14, EDYSA12, FNG11, FP12, FH13, GLY14, Gao10, GWZ14, GSH12, Goo11, Gui14, GC11c, HL12a, HL12b, HLTT14, Jan13c, Jan13b, JXA11, KJ10, Khe13, KPW11, LZ12a, Li11b, LX12, Liu11d, LZKK11, LGZ14b, LDZ10, LL10i, LY14e, Mel14, Mil14a, MN10, NWZ13, PD12, PC13b, POR13, PR10, PV11, RSRA14, SdAT14, SD10, Ste12f, Ste13d, Ste14c, TES14, TJZ13, WLH10, WZ11b, WCZ11, Wan11d, WW11e, Wan14a, WFL14a, WY12c, XL11, XLW11, Xu13b, Yan12e, YDT^{+}14, YY10, ZT12, ZLW12b, ZT14, ZTZ14, ZWS14, ZWZ12, ZL13f, uRE13, AANA10, Bel11b, Bel12, DB12a, DS12b, DISS14b, FT11, GLW12, Hen11, Kam14b, KA10b]. existence [KD13, Lai10, LKY14, LJKU11, Neu12b, QFT^{+}14, Ryo10, Soo11a, Ste12h, Ste12u, Ste13o, Tag14, WW10, WF10, YW11a, Zha12a]. exit [CDLZ11]. exotic [YDT^{+}14]. Exp [Asl10d, IU10, NR14, NMDWAS10, RC12, Zha10b, BA10, MZ12, ZA11, AH10, Asl12a, Asl12b, Bek10, MG10b, NBR10, YM10c, Yu12b, ZHZ12]. Exp-function [Asl10d, IU10, NR14, NMDWAS10, RC12, Zha10b, BA10, MZ12, ZA11, Asl12a, Asl12b, Bek10, MG10b, NBR10, YM10c, Yu12b, ZHZ12]. expand [Rem10]. Expanding [AKH13, AG13, SZZ^{+}11]. Expansion [ZHX11, Abe10, Asl10a, Asl10c, FHR13, FLW11, GZ10d, GZZ10, Hay10, HBDC14, Kud10a, KET10, LYZ11, LTZ12, Lin13a, LTW10, LLN10, ML10b, MCM10, MÇ12a, Meh11, MNP13, ÖA10, Par10c, Par10b, Ped13, PP13b, QMM13, Sha14b, She10, Tom13, TN14, Ver11, WW12, Wan13a, Wen10b, Xin10, XD11b, Xu14a, YJ14a, ZX11a, Zhe11, ZL10g, Zuo10, ZZ11c]. Expansions [WN12b, AAM13, Che14a, CC14b, Che14b, DC14, Dum10, EV14, Fuk14a, GH10a, HSC14, LLXZ12, LS14c, LM11, Sid10, WN12a, dlS10]. expansive [DI13a]. expectation [HJZC13, MLM^{+}14, SO10]. Expectations [BBCR13, LZLF10]. Expected [LXH13, qMWjHpX13, SSS13a, PAL14, ZY12, OGO14]. experiential [KG13]. experiment [Iç14, LWM10]. Experimental [AK10b, BS14b, dCDNCFRH10, TPAB11]. experiments [DGCFM11, KCLC13, dMbJmY11, PCT^{+}10]. Explicit [BJ12, BTVC10b, FF11, GG14b, HML10, Kyr13, LLL10b, LDLL10, LDLM13, MC12c, SS14e, VS10, pXqLlZ10, YM10c, ZnZ14, AT10a, Asl11b, AS14e, BGRS11, CK12c, DGDA10, dCDNCFRH10, Elo11, Elo13, FG11, FG13b, GH10b, GLZ10b, HM10a, Kes13, Liu10b, LL11e, LSW14b, MDP12, MTR13, MTER14, MT12b, OO14, SPV10, SMIS14, SHY11, ST14, WGW11, Wu10b, XZY14a, XC13, YZ10, YCF13]. exploiting [TIT13]. exploration [Bas12a, JS14b]. Exploring [VLMHK^{+}13, lQZWC14]. Explosion [Kos11a, Aja11, NBG10]. exponent [CGTCF13, yCfW12, LLP10, LCL^{+}13b, YYY10]. Exponential [BMR10, BC12b, DZZZ12, DZZ^{+}13, LLY11a, LY13b, LKS14, LLYH10, MMS14, PB14a, RN10a, Sha11b, THP14, WY12a, WLZ12b, XX14a, YZW^{+}10, Yan11b, ZSL10, ZLY11c, AQ11, AC10a, Amm14, Ans13, Ant11a, BBI14, BK14, BN10, BNP11, BK12, CMT11, ÇBS14, CZZ14, CW14a, CM12d, DLZ14, EGO13, FO11, FP12, GG14a, GLX13, GSH12, GK11, GKV12, GGXH12, GT12, KN13, KPLC11, LX12, LTL12, LSSZ14, LCCA11, MI10, wMwWsT11, MPSA14, MPS14b, MA13d, MS13g, Ngo12, PRS12, PEP14, PJ12, PAC13, RH14, SAuAH14, SE12, SIDR11, SS14c, SD10, SZZH13, SCC11, SMR11b, VG14b, WL10a, WZWW13, Wan14a, WW13h, Xin10, XCT10, YMZS10, Yan12e, Yan13a, Zah11, ZWZ12, ZXF14, ZZS14]. exponential-fitting [SCC11]. Exponentially [Abh14, DEP12, KC13a, Pal10, Ram11b]. exponentials [Din13a]. exponentiated [CL14e, GC11d]. exponentiation [ÁGVZ12]. exponents [Aou12, DAA14, Kan11, VVK12]. exposed [YWW^{+}14]. expression [GH10c, SHZ14]. expressions [AKMM12, KWS10, MD13c]. Extend [ZDW10]. Extended [AM12c, CO13, GMSE13, KC13a, LS14d, SXL12, Wan13a, Wen10b, ZM14a, AKR10b, Asl10a, BJ12, BCDI14, BC10b, BCW13, CGW14, Che13d, DQL^{+}12, GD11, GZ10d, GMZL11, Hay10, HLZ14, KS11e, LMS10, LYZ11, Lin13b, LL11e, LMA14, MCR11, NNK14, Öza11, Ram11b, Sah11b, SSZ11, SAK10, SKIH12, Sri14, SC14b, VDK13, WZDW10, XZF11, XGG^{+}11, YZC14, ZA11, ZW11b, ZB14, ZxHW14, ZDW11, Zuo10, Che10c]. Extending [AH11d, AH13a, AM14f, AG14f]. extensibility [BAFP12, Feb11]. Extension [Zhu11b, AR11, AŠA11, BDQR11, Hu11, Hu13c, Jat12, LL13a, O’R10, Öke11, OYCCB13, QHZ^{+}13, SSCT14, SHSC13, Waz10d, YL11a, YXZ10, ZH12, ZF10]. extensions [Boy14, HYT14b, Pic12, Waz10b, ZK11]. exterior [AL10a, Kle12, Maa14]. external [AA12a, DKU13, GC10a, LLY14a, Med12, MGSMRÁ14, WS12, ZW14b, ZC14b]. extinction [CXL12, KN12, LS14a, LZ10e, LD14, SY10a, ZY14a, ZZJ14, dlHV12]. extinction-time [dlHV12]. extra [BLR10]. extra-large [BLR10]. Extracting [RY12]. extraction [AdC14a, DLGJ14, KUTY14, LZW12b, MS10, WLJ11, WH13]. extragradient [BFXZ10, CAY11, FC14, KR12]. extragradient-type [BFXZ10]. Extrapolation [ZWT12, GGTCdlF13, Luo14, MG10a, ZZWF13]. extrapolation-based [GGTCdlF13]. extrapolations [WW14b]. Extrema [KA12, Caz14, MT10]. Extremal [BOR14a, Das11, GH10c, KSM11, OCN11, CHM^{+}13, ET11, WWH13, ZW11a]. extreme [GÇ11b, Has10, WLY11]. extremely [YWW^{+}14]. eye [CE13].

F. [MA14b]. Faà [DGS11]. face

[qGyFzX13, Hua10, KB11a, QZ10]. facility [JS14a, Zhe14]. facing [HMD14, TSS^{+}14]. fact [Koç13]. Factor [Yor14, CBLC12, LW12b, SS11b]. Factor-set [Yor14]. factored [Sar10]. Factorization [Car13b, DJ14b, LZH10, MD12a, ANR14, CT14a, JBDC12, LPB10, LYS10, LYS12, LL11d, XBC^{+}13]. factorizations [Lin10, SR13]. factorized [Car13a]. factors [Man12a, Sul10]. facultative [LWTC10]. fails [Boy10]. failure [Dim13a, DF14b, Gra11, LL14a, RAW13, SSTM14, ZW11b]. failures [WS13]. Falkner [ANHA14, Afz10, PP11]. False [NATQ14]. False-bottoms [NATQ14]. Families [ZCS13, AS11c, CMG14, sCLCaL11, CLS12, FPP11b, Gu10, GKS10, Irm11, KÖ11, KKS12, LLY11b, NRAV14, Öza11, PQW10, Sri11b, SÖCS12, Ste14d]. family [AdFY13, AAS12, ABGGS13, BS10c, sCYW10, sCWT^{+}12, CW11b, CBS12, CS10c, CL13b, CN14c, CMW10, CTV13, DHS10, DHZ11, DPP11, EDG14, EMS10a, FZY10, FP13c, FG11, GK10b, GK10c, GS11d, HS14b, HM14a, Inc12, KA11, Kim10, KS12c, KPK14, LL10c, LWZ13a, LS12b, LYL13b, Mag14a, Mar12, Ném10b, NS13a, NC13b, NC14, OS11, OGJLCCB14, QCK10b, Sah11a, SSTZ14, SAP14, She12a, SS10d, SDO10, Sri14, Ste14f, SL13b, TF10, WDZ12, pYX10, YZK12, Yan12a, YZC14, ZS15, ZZH11, ZLH11]. Fan [YH10]. far [HZ14a, RD11, ZTX10]. far-field [HZ14a, RD11, ZTX10]. farm [LSW^{+}14a]. Fast [And13b, CTA13, Che14e, DP14, GP13, HLeH12, RV13a, SU12, SEM11, XXZH10, BCGV11, CLZ11, Che14i, CM10, DHI13, EBR13, GLM14, JHX14, KCG10, LZ13a, LWZZ14, LL10d, LLS^{+}13a, LWN14b, MK14b, Mur11, Nem10a, NSYA14, QQXY12, RY14a, SC14a, SY14b, WWW13c, Xu10c, YLLL13, ZW13c]. Faster [Deo12]. Fatou [AFFV11, Kar14]. Fault [CZX^{+}14a, CZX^{+}14b, LG10b, LX14b, Sze10, Sze11, Sze12, BGV10, EdCVB^{+}14, HZFZ14, HYZH14, KRJ11, LU12a, TP12, Zho12a, Zuf12]. Fault-tolerance [LX14b]. Fault-tolerant [Sze10]. faults [PKD14, Yan10c, Yan10d]. faulty [DD11, LTT^{+}13, Sun12, WFZL12]. fBm [RCS14]. FD [HWPY14]. FDM [HB14]. FDTD [AAAH12, NZ14]. feasibility [CLC^{+}13, WYY11, ZZY12]. feasible [AG14a, BK11a, bJLyZlL14, JW10b, LHL11, dMbJmY11, RMY14, WYT13, ZZG10]. Feature [WLJ11, CZX^{+}14b, qGyFzX13, KUTY14, LZW12b]. Feder [ZYZ14a]. feedback [AM14b, AGH12, BBI13a, BBI13b, Che11b, DGS14b, DJY13, EDYA14, Fan14e, GHG14, GC10b, Kan12b, KWJ13, LPJ^{+}12, LYPC12, LXC14, LZL14a, Mur14, PZ14a, PWZZ13, SHL10, SWLP13, SL13b, TPT12, WF10, WZ11d, Wan13e, WC14c, XPLZ14, YL11e, yZ14a, ZLM12, ZZ13a, Zhu10b, dlSAQ11, BYM14, NTT13]. feedforward [MHS12]. Fekete [DMS12, Kan12a]. feline [NM14]. Feller [Zhu11a]. FEM [CL14c, dCDNCFRH10, dCDNVS11, dCDNRH12, dCDNV^{+}13, HB14, KS13a, MNOB10, NTPVNHLX14, NCANCF12, RSBK14, RV13a, RGL12, SPK14, ZD10]. femtosecond [LWL12]. FENE [MZWZ12]. FENE-P [MZWZ12]. fermentation [ZLC10]. Fermi [ABF14, Fer11, Fuk14a, Fuk14b, Oul11, ZLLL12]. fermionic [KK13a, ZK11]. ferrofluid [SSM11]. ferrofluids [SHI13b]. fertility [DMGK12]. few [JPB10b]. FGPM [XDG12]. fiber [ZA14b]. fiber-reinforced [ZA14b]. fibers [GQ13, Yu13a]. Fibonacci [Chu10, DRS11b, EMS10a, Fen11, GCMP14, GTU13, İpe11, IA14, mJ10, KA13b, Li12a, LO14, Ma11b, Pet12, Ram14, SC10a, SCH11b, TKS11, Wło13, Yay11, Yor14, ZJ14c, dF14a]. Fibonacci-like [IA14]. Fibonacci-type [Ma11b]. field [AAMADH11, ABG14, AAAH12, dCDNV^{+}13, DFGL14, DKU13, DKK^{+}13, Fer12b, HZ14a, KG14, Kov10, KS13d, Kyr14, LG10b, lLZ14a, LL11c, LX11a, LY14c, Pal10, RD11, SBS10, pSbL14, Shi14a, Son12, SD13a, TV13, TSJ13, Üna13, XK14, YWW^{+}14, Yu11b, YLK14, ZA14b, ZLX12, ZTX10, ZZ10d, WMJW11]. Field-induced [WMJW11]. fields [AS14a, CL11b, DLC10, LZM^{+}12, LLW13, LS11b, Pat13a, XDG12]. fifteenth [GK10c]. fifteenth-order [GK10c]. fifth [Asl10d, BC10b, DAEY12, IU10, Jha13, KS12c, QMM13, YZ11, YZZ10, Yu11a]. fifth-order [BC10b, DAEY12, QMM13, YZ11, YZZ10, Yu11a]. filaments [ZL13c]. Filippov [ZT14]. filled [HCW11, LWF11, LY12a, LWFG13, MYL10a, Mah13, SM14a, SD13a, WR10]. filling [MYT10]. film [ARH^{+}12, ASRI12, ASR^{+}14, KBB12, LWCS14, PS12b, Van11a]. Filon [Che14f, MX11, Sha13a]. Filon-type [Che14f, MX11]. filter [BD14d, Gro13, Gu11, GZ12a, Gu13, GZ14b, GG14d, GG13c, HZ12a, KCGH14, KUKU14, LCL^{+}13a, LZ10g, Nak13, PZ14b, PRK13, PD13, QW11a, SA12b, UMMT14, WZ11f, WZZH11, WLP^{+}14, WZ14b, ZP13, LZ10f]. Filter-Levenberg [LZ10f]. Filtered [MNP13]. Filtering [BRP13, PS14a, AKC12, ELS11, EJ11, HDT11, KCGH14, LJKP11, MMM11, NL10, PD13, WS13, YMZS10, ZZY^{+}14a, ZCZ^{+}13a]. filters [FHR13, cLYrL11, SY13a, WSD13, WZ11h]. filtration [WGS10]. fin [RH14]. final [DYYL13]. finance [GR11, KP12, ZLL11b, ZW14b]. financial [Ben14b, CCO12, DW13a, Hua11b, MCR11, NR12, Val11]. financing [JGM13]. find [BY11a, NSC12b, NS13a, NC13b, NCS14, NC14, hTlCqL10, YS13b]. finder [GSDB11]. Finding [ASS13, BK11a, DL13, DGDA10, GMA13, Lip13, LC11, MRCD12, BBSC13, BDE^{+}14, Che13d, EH14b, KGB12, KSM14, LTZ12, LWFG13, Mag14a, MMRS13, MF13, MMS– D13, Par10c, Par10b, PC14b, RSK11, SSTZ14, Yun10b, ZS14a, ZOA10]. findings [PC14a]. Fine [SK12b, Alt11, DB12b]. Finitary [CCG^{+}12a]. Finite [AZ14a, And11, BJS14, CLX13, DS12a, DU13, EAA11, EKZ11, KPC14, LDL11, SWX10b, WaZZW11, WLH11, YL14b, ZH10b, ZCZ^{+}13a, ZWZH13, AAP10, AHDD13, jASzZ11, jAzBS14, Apa10, AP14, BC11a, BP11b, BAFP12, BH12a, BBFK12, BH11, BA14b, Bra13, CYZ14a, Cha14c, CBO12, CB10, CLD12, CF13, CS14a, CRX12, CZD^{+}14, Cho11a, CS10c, Chu11, CMW10, CG10, CG12b, Cop14, CRW11, Cvi11, DAHAK10, DR11a, DB11a, DX10, DHZ11, EJ12, EM13, ESEEKH12, ESAH10, EEEE05, FCJ14, FY12, Feb11, FF11, FK12, FMF13, FR11, GT13, GM11a, GM14a, AA14a, GKC11, GG12b, GK11, Gu10, GZF12, GFZ13, GMKV12, GSRR13, Gür10, Gür12, GS11d, Gus12, HDK^{+}10, HJVC14, HS14a, HS14b, HH14, HL10, HL13a, Her11, HOMA13, HZ12c, HL14b, HG13b, Inc12, JU11, JKN11, JMY13, Jha13, KCZK11, KP14b, Kha14c]. finite [KAK11, Kim12, KSS12, KK13d, KSM14, KP11a, Lee12, LW13a, LW06, LZW12a, LY13a, LC14a, LL12e, LY12b, LG14, LXZ14, LFLH14, LLS13b, MG10a, Maa14, MDP12, MS12a, MA13c, MG11b, MF14, MJ13b, OS11, PZ12, Pan12b, PR14, lQZWC14, Ram10a, RGA14, SZG12, SPT12, SYK10, SXL13, SZ11, SZS13, SP13a, SY13b, SCC11, Sht11, uiH14, SCK13, TM13, TS12a, Tem12, THNTNXR11, WRZ12, WZJ13, WFZ13, WP14, XYC11, XP12, XLGW12, XHXH12, Xu14c, XZ13c, XDZ14, Yan10b, pYX10, YWK11, Yan12b, Yan12a, YS12, YY13a, YLS13, YJ14a, YXZ10, YWL13, Yua13, ZLA10, ZF11, Zha11c, ZLMG12, ZW12, ZLM12, ZLS12b, ZZL13, ZCZ14, ZS15, ZBM10, ZL12c, Zha13d, ZCD10, ZCC13]. finite-buffer [JU11]. Finite-difference [KPC14, BBFK12, Her11, MDP12, Ram10a]. finite-dimensional [Sht11]. finite-horizon [ZBM10]. Finite-time [AZ14a, LDL11, ZCZ^{+}13a, ZWZH13, CZD^{+}14, HL10, HL13a, ZLMG12, ZLM12, ZLS12b]. finitely [TL10, TLL13]. fins [RH14]. FIR [Nak13]. fireflies [CGVR14]. firefly [RAZES13]. firm [Fuk12]. Firms [CG14a, Liu13b]. First [Jan10, Ami14, BKP14, BMR13, BK12, BPX14, CMH12, DAV12a, DAV12b, DSGB12, Dol14, HLZ12, HMA10, Jan14d, Kim11, LTG^{+}12, LLL11a, LJKU11, LLLS13, LZX10, MYW13, MG11a, MG14c, PSP10, PG11, PAA14, Pre11a, RRSPTR12, RRSPTR14, RTPH14, Ste13j, Ste13d, Wan11a, WW11f, WE13, Xia14, YLLL13, ZF11, dLDP13, AYS13]. First-order [Jan10, BK12, CMH12, Jan14d, LTG^{+}12, MYW13]. first-passage [RRSPTR14]. first-passage-time [RRSPTR12]. Fischer [HM12, TC14b]. Fischer-Burmeister [TC14b]. fish [KHA12]. Fisher [ARE14, CST12, GIR10, HB11a, MDP12, PJFL14, SLW12, WN12a, Yan13b, YCJ^{+}14, ZYZ12, ZK10a]. fishery [BLR13a, CJK12, CDK13, PM14a]. fit [PTW11]. fitness [HWQG11]. fitted [Abh14, AISS13, BP11b, DEP12, FYM13, FG14b, Kal13, KC13a, Tsi14]. fitting [HJZC13, JHZC13, RAW13, SCC11]. Fitzhugh [Bhr13]. Five [CT11b, TYCZ14, WPKE14]. five-dimensional [WPKE14]. five-point [TYCZ14]. Fixed [DI13b, GFMP13, ICKV14, LAR13, LPB10, LLK13, Ols12, AKR10a, ARN10, AKN11, AN– D12, ANR12, ANR13, ARR12, AKR11, Bar10, BOR14a, BCK14, BS10c, Buo10, sCCLY10, CIVS14, CKH11, ĆASH11, ĆSAV11, ĆARA12, DV12b, DV13, DLSS11, – Dor13b, EVR13, Gar11, GCG14, GAV14, Hao10, fHcXH11, HA11, HP12, Iid11, Inc12, KM12a, KP14b, KR12, KI12a, KZ11, KPK14, LPAE12, LXTC13, Lin11c, Lin11b, yLbLjH10, MM11a, MAD14b, NKR12, NA14a, Ola11, PH10a, PK12, RRR11, Rad14, RBT14, She14a, SY13a, SSRA12, SLJ^{+}13, SK13a, Tad12, WHC10, Wei14, XG13, YZK12, YK12, ZS14a, ZL13a, Zha13c, ZZ13b, ZPZ14, ZX11b]. fixed-grid [Tad12]. fixed-lag [ZZ13b]. Fixed-point [LPB10, Bar10, Gar11, LPAE12, MAD14b, SY13a]. flames [VD14a]. flapwise [Nar12]. Flat [YYL14, DNI^{+}10, HL13c, LWM10, RRP14]. flat-top [HL13c]. FLDA [KB11a]. Fletcher [LL13a]. flexibility [RNEA14, VRZ11]. flexible [CC11b, HSD11, LPX12, ZZ13a]. flexural [Kia14]. floating [GO10]. flock [LFLS14]. Floquet [Bel13]. Flow [HFS12, KS12b, PMV14, ZM14c, ANHA14, ASM11, Afz10, AAP10, AHMS14, ARH^{+}12, ASRI12, ASR^{+}14, All11, ABK^{+}14b, AM14c, AHKS12, AM14g, AA12d, AM13c, BAO14, BY14, BD14b, BLK^{+}13, BSP11, CSQ11, CMT11, CFH^{+}13, Che10c, CCLW10, Che11e, Che10f, CRX12, CRP13, CP14a, Cor11b, CRC14, DA11a, DOV^{+}11, Das13, DNI^{+}10, EEEE05, FZZ12, Fer10a, FdOdSM14, FMV14, Fit12, FBGH12, FKH10a, FMSC13, Haj11, HZ14a, HAMA14, HH11, HDZW11, HCCT14, IC11, Ish10, JHA14, JZG13, JWZ^{+}12, JHZC13, JFHZ14, Jos13, KP14b, KK11b, Kha14a, KS13d, KMH^{+}13, Lee12, LFC^{+}12, LYG14, LSF^{+}13, MWAM14, MSCK11, MAD11, MZWZ12, ML11b, NAHH11, Neu12b, NSNEV14, ÖK11, Pan12b, Pan15, PK18, PHPD11, PP11, PMD14, lQZWC14, dejW14, RY10, Rec13, RNEA14, RRP14, SBR^{+}12, SMBI14, SM14a, SBS10, Sha12b, SHAO14, SMC12]. flow [SZZ^{+}11, SFF11, SGC10, SWD11, SD13a, Svá11, TSS^{+}14, TPD10, VPS12, VSV10, Van10, Van11a, Wan14f, XZZK12, XK14, XWWC10, XG13, YW11b, YTY^{+}14, Zei11]. flowing [GW13]. flows [AV14, BMMS13, BBFK12, CGLM^{+}12, Cor10, CWW14, Cur14, FHAL14, FHK13, Gad11, GL13a, GVT^{+}13, GP11a, LJ12b, LR14a, MV12, NR13, RP14b, RW14b, SdAT14, SJR10, SHBN13, Siu10, TKC12, WHG13, XK14, Yan10a, YHH11, ZCF13]. flowshop [EPMU10, EG06, LH10b, Rud13, XY11]. flowshops [AA12c]. fluctuation [FP13b]. Fluctuations [INE10, AKMM12, GGYW11, GS14c, SS13c, ZTX10]. Fluid [KS12b, wMwWsT11, ANHA14, ARH^{+}12, ASRI12, ASR^{+}14, Amm14, AM14c, AA12d, AM13c, Bar13, BBFK12, Che10f, Cor11b, CE13, DWQG12, EEEE05, FMV14, FKH10a, Gad11, GG12a, GZW12, Haj11, HMA10, HAMA14, HYT14a, JHA14, KM11a, Lac14, LFC^{+}12, LSF^{+}13, Mah13, MNOB10, MWAM14, MSCK11, NA10a, NAHH11, NR13, NTPVNHLX14, NZW11, PC13a, Pan12b, PK18, PP11, RNEA14, SM14a, SBS10, SGC10, SWD11, SD13a, TPD10, VSS13, Van10, Van11a, VV11a, VO12, VO14, Wah13, XGG^{+}11, ZGY^{+}10]. fluid-fluid [GG12a]. fluid-solid [NTPVNHLX14]. fluids [AM14g, CS12b, FHAL14, FMSC13, HFS12, HCCT14, KK11b, MV12, MU12, MZWZ12, SFF11]. Flux [Dub13, AE11, BC10a, Che10f, DMGK12, FH11b, FKH10b, JA14, Kos11b, LR14a, RW14b, SGC10, WG11, ZCZ14]. flux-temperature [FKH10b]. fluxes [ZMZ11]. fly [YDZH14]. Focal [Dos14]. foci [WY10b]. Fock [ABP13]. focus [WHW13b]. Foerster [BRW13]. Fokker [WL10c]. fold [BY14, KZ12, MM13a, MGWG11, SPT12, YL11f]. follow [SSS13a]. followed [Ste10a, Ste11a, SS11g]. following [AMS14, WDBM14]. follows [ÇT10b]. Food [DS14c, MHTH12, DMGK13, FL11, Li12c, LX14c, LB14d, MLY12, PBG14, SP14b, SP14c, SY10a, SMR14, WW13b, Wei10a, ZLD11, ZY14a]. food-limited [Li12c, ZY14a]. Food-Search [DS14c]. food-web [PBG14]. Föppl [EGJB13]. forager [BDDK14]. foraging [BDAD10]. Force [For11, ZnXwY14, AS14d, DWQG12, DQL^{+}12, dOS13]. forced [Jan14b, KWM13, MH11b, NA10b, Sal10a, Sal10c, SH12b, Yad13, YL14b]. forces [FHK13, Med12]. Forchheimer [LL14e]. forcing [HL12d, MB13a, SS13d, TW11b, Wan14f, Waz10g]. Forecasting [LPY11, CSSSV13, HDZW11, Hua11b, LXGM14]. foreign [DW13a, Hua11a]. forest [GSC14, OSRPC12]. Forests [PDPGOR13]. Foreword [DSZ12, Sri11a]. form [CCJ14a, DI13b, DWZ12, FP12, JBL^{+}11, JS13b, KZJ13, KM13a, KI10b, KI12b, Koj13, Koj14b, Koj14a, LWL14, MMSZ14, NMR14, PRS12, PPMCG14, Rim13b, S.14, SE12, SH12a, SR10, Ste13h, Ste13i, Tch11, TJZ14, WDLM10b, WL11, WW13d, WWWJ14, Waz10d, WRD13, Wen10a, WZL12, dYyR13, YYL14, YWCH11, Yu12a, Zhu10c]. formalism [Jan14a, WSD12]. formalizing [AS12b]. Formation [WLLY14, Ban14, BWaY14, CPC11, DBS12, Fan14a, Gaz12, KR14, NATQ14, RZHH13, RZHH14, SZS^{+}12]. forms [BES10b, BC10b, CKK12, DAEY12, HC14, HS11b, yL14c, LKY10, QMM13, YWCH14]. formula [BM12a, Che14a, DGS11, DWZ12, Elo11, Elo13, Fan14c, GT11a, GO10, Hwa11, Hwa14, Lin14b, MCS11, Mor10a, Mor10d, Mor10e, Mur11, Mur13, PM11a, SE12, SSV11, UB13, XC13, ZHY11]. formulae [CT11b, Cvi11, GCM10, jLLW12, Mil12, SP11, SPP12]. formulas [BD12b, CS11c, Cho11c, ELIGMB11, FPP11b, Kim11, KC13a, Kyr13, LDS14, MÇ12a, MTR13, MTER14, RSG14, WHW13b, dlRS14]. formulation [AS12a, ASKA13, BC11a, CGGZ11, EJ12, LMB13, LLS13b, RGL10]. formulations [TC14a]. Fornberg [yDpY14, HML10, JB10]. forward [CR11a, HD14a, Tad10]. forward-backward [HD14a]. foundation [DB11a, HA12, KAF13]. foundations [AGM14, UW10]. Four [Pat13a, WW13d, Waz10b, ARN10, ARR12, AS10, AAS12, B– D11, FMF13, GL11b, LTW12, LWZ13b, Meh13, Pat13b, PBG14, WFL14a, Wei10a, Yan13a, ZL10b, ZGZG10, ZS10b, ZANA12]. four-bar [B– D11]. four-dimensional [PBG14, Yan13a]. four-point [AS10, ZL10b, ZGZG10]. four-step [AAS12]. Fourier [ZW13c, Abb12, ABJ10, AM13a, BHAV10, Bez13, BS12c, Boy13, Boy14, Che14g, Cvi12, FY12, LWL12, MMKD14, NL10, ROA11, RGL12, SE12, Sar10, Sid10, TQ11, Ver11, Zou12]. fourteen [TYT14]. Fourth [jAzBS14, Her11, Afu12, AT10a, AM13c, Bai10, BD10, CT13a, CLZmF11, Car13a, CF14, CT11a, CS14a, CLW14, CLND12, DH13a, DK14c, FHLW11, FKH14, Fre11, GWS13, GWA14, GL13b, HSZ14, HZ12c, Iri10a, KK11c, KKBO12, KA10b, KS12d, KM12c, LP12a, LSLL14, LZZ10, LFLH14, LZSD14, LM12, LD13c, MJ11a, Nie13, Sha14a, SS11c, SAP14, She12b, Ste13e, SW10b, SW12, TMM14, Tun13, WCL11, Xia13, XGM10, Xu11, YZ12a, Zah11, ZHY13, ZZH11, ZCS13, ZGY14, ZLW10, Zhu11c, ZR14]. Fourth-order [Her11, Bai10, BD10, CT13a, CLZmF11, Car13a, CF14, CT11a, CLW14, CLND12, DH13a, DK14c, FHLW11, FKH14, Fre11, GWS13, GL13b, HSZ14, HZ12c, Iri10a, KK11c, KKBO12, LSLL14, LZZ10, LFLH14, LZSD14, LM12, LD13c, Nie13, Sha14a, She12b, SW10b, SW12, WCL11, XGM10, Xu11, YZ12a, ZZH11, ZGY14, ZLW10, Zhu11c, ZR14]. Fowler [Gal11b, KM12c, LML12, TM11, Waz11b]. fptdApprox [RRSPTR14]. FPTL [RRSPTR12]. FR [LLZC11]. Fractal [FMSGS12, LCL^{+}13b, PT11a, SSA14b, BB13a, BBCR13, LLT^{+}14, SIR14, VC14b, VOL14, dACSS12]. fraction [Che13b, PSJGLS14, WQ11b]. Fractional [AB14a, Ans13, Els12, FT10, GML13, Jan14c, MRE^{+}12, PHGV13, Ram11b, RTVV11, TS12b, ZL12c, ABRT14, AG10, AS10, AZ14b, AT10b, AAT^{+}13, AANA10, AM12a, AA14b, AZ14c, ADBT13, ADP11, AJPZ14, AZ19, BZ11, BMA11, Bal13, BDG13, BW14, BCKL11, CS12a, CH14, CDS13, CKN13, CLA^{+}12, CT12c, CLTA13, CSL14, CC14c, CLL14, CFSYPP10, CS10d, CDST14, CSS11, CHS14b, Cui14, DGSB14, DVGY11, DK11a, DBK14, Dea13, DT14, DN14b, DJ14a, DBBEE13, DCR12, EN11, ESEEKH12, Ema13d, FIK14, GL11a, GP12, GS14a, GPCCA14, Goo11, GK12a, GKW14, GT10a, GT10b, GT12, HL14a, HHM14, HZ12c, HLX13, HYSZ14, HYT14b, ITLA10, Irm11, IW14, Jan13c, Jan13d, Jan14e, JMB12, Ji14b, JL14, Jia13b, Jia14, Jum12, Kam14b, Kat11, KM11c, KRC10, Kro11, LZ10b, LLJ11, LL14c, LCL^{+}13a, LH11c]. fractional [LFLH14, LW14g, LP13b, MH13a, MZ13a, MT12a, MMMMMG^{+}11, Mat14a, MPS14b, Mer14, MN10, Mop11, Mop14, Mos14, Ort14a, PT14a, Plo14, QLCF10, RB14, RMY14, RSG14, Ray12, RP13, fRbLJyS13, SuR13, SMN12, SGA13, Sal11c, SSK10, SP11, SLC^{+}12, SEL14, Sta14a, Sta12a, SWX10b, Tag11, TJZ13, TPD10, TD13, Tri14, Ver12, Ver13, VSM14, VRMT14, VDOG13, WZ11b, WF12, WDT^{+}13, WXWY14, WMM14, WZL14a, WDS11, WL10c, Wu14c, XL14, YWLX11, Yan14b, YG11, YLAT14, YHW13, YH14d, YLTB12, Yüz13, ZL14a, ZLA10, ZLW12a, ZLW12b, ZL12b, ZLWL13, ZW13e, ZZZ14, ZLWW14, ZWS14, ZLZ^{+}14a, Zha14a, ZSHZ11, ZW11c, Zhu11a, uRE13]. fractional-in-space [ITLA10]. fractional-order [CSL14, KRC10, SMN12, WXWY14, ZL14a]. fractionally [RSG14]. fractioning [MPSA14]. fractions [DBG11, IT12, Man12a]. fracture [MPPM12]. fragile [ADBT13, FP13b, GWW14, LZ14d, RZ12]. fragmentation [BPPS14, SP14d]. frame [LZ11b]. framed [MR13a]. frames [AS14a, SHSC13]. framework [BCR10, BMI^{+}11, BIT12b, CCSM13, DZS^{+}14, FH11a, KDTSN12, KM12c, RDFL13, TMM13, TMM14, TS12a, Zuf12]. frameworks [yL14c]. Frank [Har10]. Fréchet [AG14f, Ols12, RA12c]. Fredholm [PR12a, Tur14b, AASS13, BTS12, CGGMG13, CK10, EZ13b, FEGEA14, FMF13, Fra10, GKS12, GM14b, HKI12, HSW12, KSFG14, Koh12, KD13, LL10d, MMO13, MS13f, MHB14, MAHAS14, MV11, MS11, OR11a, Rah13, SR14a, TW14b, Tur14a, WWG13, WW13c, WW14a, Zho13]. Fredholm-type [Koh12]. Free [DW13b, EEM12, MP10, MP12a, Sie11, WW13e, AAV10, AA12a, AGM14, AGZ12, AAB^{+}13, ACCT14, BHG10, aBS14, CZL14, CCGT13, CLND12, CRC14, Dan14, DOV^{+}11, DB11a, DAA14, ESV12, EJ12, Fer12a, GS10a, GZL14, HCCT14, JE11, Ji14a, KK12a, KA11, Kim10, KS13d, LLK14, LSJZ13, LV14, MYL10a, Mah13, MO12b, MSS14b, Naz10, PG13, Pan15, PID10, QW11b, QC12, SJR10, SBS10, Sha12b, SGG12, SAP14, SGC10, SY14b, VLMHK^{+}13, WiN11, WDZ12, WW13f, WD14b, WCN11, XYC11, Yun11a, ZLZ14b, ZL14b, ZP13]. free-boundary [CRC14]. freedom [ÇBS14, XLW11, ZWD12]. Frenet [CK12a]. frequencies [MMM12]. frequency [FC10, GG12b, Kim11, NLLC10, NL10, SRA10, YWW^{+}14, AC10a]. frequency-dependent [Kim11]. frequency-domain [NLLC10]. fresh [CJM14]. fresh/salt [CJM14]. FRF [MMM13]. friction [Ahn12, Pfe10]. frictional [TS12b]. frictions [GS14c]. Friedmann [MPSD12]. Friedrichs [ZH12]. Frobenius [CD12, DR10b, FDHK14, GMX10, GS13c, LQY11, RBT14]. frog [Aha14, LPX12]. frog-leaping [LPX12]. front [CC12c, Den14]. frontier [YLMZ13]. fronts [FL10, MYT10, ZL12a]. frozen [ABGGS13]. fruit [YDZH14]. fuel [LKP12, NBG10, NAL12]. Fukuyama [HCD14]. Full [LLXZ12, SCIMM11, SPKS12, pZLC10, Abd13, AG14a, LWWZ11, Lin13c, LS11a, WYT13, WSK12, WHG13, ZL10f]. full-NT-step [LS11a]. full-order [Abd13]. Full-rank [SCIMM11, SPKS12]. fullerene [GBA14]. Fully [MC14a, CSQ11, Jos13, MAD11]. Function [CXG10, GT11b, AQ11, AH10, AA14b, Ant11a, Ask14, Asl10d, Asl12a, Asl12b, BS10a, Bek10, BA10, BW10, Che10a, CD12, Che13b, Che14a, CZ14a, Che14j, CX14a, CX14b, Cho11b, CBS12, CS11c, CEXG12, Cvi12, DHI13, DM13a, DGDA10, DAV12a, DT10, DGSB12, DSM^{+}12, ECG14, Fan10b, FMS11, FK13, GW11b, GQ12, GMP14, HK11, HCW11, HW12, HXQ14, HW11, HM12, HM14b, IU10, mJ10, JLHW13, Kam14a, KB13, KAY12, Ko13, KI10b, KI12b, Koj13, Koj14b, Koj14a, KSD13, Kra14a, Kud13a, KGO10, KSM14, KL13b, LSS11, LY10a, LP10, Lee11, jL12, LWF11, LY12a, LWFG13, LZ10e, LG14, LC14d, LS14c, LCLW10, MYL10a, MZ12, MÇ12a, Man12b, Mat14d, MHXyY14, MM11b, MG10b, MCL14b, Mor14a, MGSMRÁ14, NBR10, NR14, Neu11, NMDWAS10, OP13, PM11a, PZ14b, dFPF14, Rag12, RK14b, RY13a, Ras14, RY14b, RY14c]. function [RRSPTR12, RTPH14, RnZ10, RX14, RC12, SPA12, Sah11a, SDC10, SHN10, Sar12c, SATdV13, Sav10, DAV14b, SSZ11, Sha14b, SH12a, Shi12, Sim11, SMWR10, SK10b, SKM11, THDF11, TQ11, TIT13, LM13a, TD13, TC14b, UY11, UJH14, Vil10, VC12, WG11, WYK11, Wen10b, WN12b, WR10, WLXL10, Xin10, YL11d, YK14b, YM10c, Yu12b, ZA11, Zha10b, ZX11a, ZM14a, ZY12, ZHZ12, Zhe12, ZMG13, ZS14c, Zho11a, ZZ12b, ZZL14, ZK10b, ZS14f]. Function-valued [GT11b]. Functional [CS12a, OSAG13, AAM13, BD14a, BP10a, BKA12, Bah13b, BBJ10, BFM13, BMI^{+}11, BC12e, BPX14, CHC11, CYHL11, CZZ14, CK12c, Dar12, Dar13, DKS14, ESH10, FW14, Fig11, Fig12, GZ13, GC11c, GS14f, HLZ12, HXZ12, HL11e, ICKV14, JVK10, Kar13a, Kau14, KR13b, KS11d, KD13, KPLC11, KPP^{+}13, LJR14, LLJ11, Liu11d, LY11b, LB14c, MSGN^{+}14, MAD14b, MN10, Ols12, PSP10, PKPL11, PD12, PJ12, PLF13, QLWW12, RV14, RS13, RCS14, SHMCM12, SDL10a, SL13a, Sho14, SY10b, Ste12c, Ste12l, Ste12f, Ste12g, Ste12h, Ste12m, Ste12r, Ste12t, Ste12u, Ste13a, Ste13j, Ste13b, Ste13k, Ste13d, Ste13l, Ste14b, Ste14c, Ste14d, Ste14f, SH12b, TKLH10, VC14a, WFZL14, Wan14b, Xie11b, YSZ12, Yue14, Yus10, ZOA10, ZjHO13, ZQ14, ZXF14, ZM10c, dS10]. functional-difference [Ste14d]. functional-differential [Ste14f]. functional-integral [Dar13, DKS14]. functional-integro-differential [ZQ14]. functionality [KTW10]. Functionally [ABP13, ASKA13, DRJ11, EEM12, EKSO14, Ngu14, SM12, Zen12, ZZHZ14]. functionals [AG14b, BA11, GZ14a, LLP14a, LLP10, LZL14a, MPC14, RDA14]. functions [AC13a, AAA12, AG14b, AAC11, ASSPA10, AR14c, AG14c, AG14d, AB12a, AMFM12, ASMA12, Ata11, AKÖ11, AS11c, Bab12, BDGP14a, BDGP14b, BAR11, BCM12, BEdAdSK14, BCK14, Bir13, BCEAV12, BÖ11, Boy13, Boy14, BSM14, Bry11, CCMV12, ÇDO11, CEM11, Cao10a, Cao12, CHX13, CS13b, CL11c, Çet13, CALR11, Che10a, CLW12, CB12b, CV13, CA14, CSL14, CKS10, CY11b, CS13f, CFSYPP10, CNKS14, CP12b, Cvi10, DGMR10, DAV12b, DSGB12, DSM^{+}12, Den11b, DC13a, DO11, Dha13, DLX13, Dra11, Dra14, DRS11b, Dzi13, DPP11, EMK13, EMA13b, EV14, EZ14b, FC10, FcDDqZ10, Fer12a, Fer13, FPTS14, Fuk13, Ger12, GA11a, GBD^{+}12, Gue14a, Guo13a, GKS10, GS14d, HVC12, Hon10a, HL13c, HM11, HYT14a, Hsi10, HZW12, HZW14, Hus10b, HD14b, Ign10, Irm11, IW14, ISÖ14, JPI^{+}13, Jun11b, KÖ11, Kal12]. functions [KOP11, KU11, KM12b, KW10, Kel12, Ker13, KYH10, KC13a, KZ11, KZ12, Kwo11, LLSY11, LN10, LS11a, LLSW12, Liu12b, LOS13, LO12, Lu12c, LZ14d, LMA14, MDD12, Mam11, Man11, Mat14c, Meh11, Meh13, MMSZ14, MCS11, Mil12, MHB14, MMY11, MJ11a, Mor10b, MCL14b, MSJ14, MGGS12, Mur13, NS12, Neu12a, Neu14, NZ12, NM08, OP14a, OO10, Ota14, Öza11, PR12a, PP12a, PP14b, PS12a, Pré11b, QHZ^{+}13, RT14, ROA11, RDA14, SI11a, SMDZG13, Sas12, SP11, SEED13, Sek11, SA10a, SYW13, Sid10, SS10d, SHZ10, SL11b, SPP12, SKM11, SC12, SB12b, SÖCS12, SÇK14a, SAJ14, Sri14, Sta11, SMR11b, SKW11, Tom13, Tri14, UOK10, VDL12a, VV11b, WXA10, WZ10, WQW^{+}10, WFW10, WWZ10a, WLX11, WZ13c, WWC12, WL13d, XXS12, XLS13b, XCS13, XX14c, YL10a, YL11b, YL11f, YTS12b, Zha10b]. functions [ZSZ12, ZL12b, ZZ10b, Zhu10c, Zhu11b, Zhu12a, ZH13, Zil12]. Fundamental [Jar10, Tra12, GG12a, HT11b, LKY10, Reu14a, WCL12, YYL14]. fundamentalists [NR12]. funds [BPDZ12]. Further [HY14, LY12c, LZB14, MA13d, MD11a, MD13b, ZJ13, FXD12, MH13b, Pic12]. Fusion [Hua10, CÁGGLP14, GLHCLP12, WX13d]. future [Fig11, NSV13]. fuzziness [CMP13]. Fuzzy [MVRVSCMV14, AY10, CS13a, CZZ^{+}11, Che13c, DGS14c, DS14d, EZ13b, JJPW13, KZS12, LZ14d, Mal13b, MPSA14, OH12, PKP^{+}12b, RSPK13, SAMA11, TR13, ZPXX14]. FVM [DGCFM11].

G [wMwWsT11, WSY14, Yua12, ZA14b, ZHY11, ZZR11]. G-Brownian [ZHY11]. G. [Gal10c, Pak11b]. G2 [ZCL13a, ZCL14a, ZCL14b]. G2/M [ZCL14b]. GA [AKP13, PKD14]. gain [FP13b, HZ14b, Lee11, LXK13, LDL11]. gains [YJ10]. Galerkin [LL12e, Pan15, ZCC13, ABJ10, AB12b, BES10b, EL13, CLF13, CF13, CP14b, CYZ14b, DZLW13, DWZW14, DAEY12, DFM10, DC13b, DLS13, EGMEM14, EVDG12, FEGEA14, FGO12, GZXJD10, HXY12, HH14, JKN11, LZ12b, LC14a, LSL10, LWN14b, MYL^{+}10b, PN11b, lQZWC14, RB11b, Ros11b, RV10b, SNMD13, SJ12, SBS10, Sha12b, SS10e, TD14a, TD14b, Tem12, TA13, VZ11, XH13a, YLLL13, ZLA10, ZYZ12, ZY13b, ZXZ11, ZZ13c]. Galerkin-finite [LC14a]. Galilei [CLAT11]. gambler [KCK08]. gambling [LD13a]. game [Ask14, CN14a, DZS^{+}14, DLGJ14, DLP10, EM12, EAE13, FXD12, GG10b, JMHZ10, MG14a, MTSN14, PD14, PMP11, QGD14, RK14b, LM13a, XMWD14]. games [BBS10, KI10b, KI12b, Koj13, Koj14b, Koj14a, LZLS14, SVV13]. Gamma [Kre13, CB12b, Che14a, CZ14a, dSCAM14, Cor11a, EV14, GQ12, GMP14, GJGSNRÁ12, KSD13, Kra14a, Mam11, Mat14a, MM11b, MCL14b, Mor14a, Neu11, OGO14, RY14c, SPA12, Seg14, TQ11, XwK11, CS11c, CS13f, Izg11, KV13, MS12a]. gamma-spline [XwK11]. gamma-type [SPA12]. GANM [LPP^{+}12]. GAOR [WWLL11, WWT13]. gap [BA11, MMM11, THNTNXR11]. Gaps [SAO12]. garbage [KZS12]. GARCH [MKP13]. Gardan [AA12b]. Gardner [BVTC13, HL12d, LY14d, LL10h, OSRFF14, Waz10g]. garment [GXC^{+}11]. Garnier [ZK11]. gas [DK11a, EASK10, MV13]. gases [HW13b, SJ13]. gasket [BV12b]. gated [KP14b]. gathering [Ber14]. Gaudi [VCĆV12, VC14a]. Gauge [LTG^{+}12]. gauges [lJLpWxZ14]. Gause [DSH12, YTC13]. Gause-type [DSH12].

Gauss [BM12a, Bhr13, BCNP14a, GMV11, ACCT14, AG13, AM14f, AM14e, BCNP14b, DLL11a, DB12a, DWC13, JC14, LY14e, Neu12a, PS12a, SKM11, SAJ14, SP13b, YZ12b, ZYLS12, ZZ11a]. Gauss-type [DLL11a, DB12a, DWC13, LY14e]. Gaussian [ADGP10, BW10, Boy14, CC10b, CDLZ11, CM12d, Kam14a, LW11c, Peñ13, SN11, SPP12, SFP^{+}11, Tom13, WLP^{+}14]. Gaussianity [GVMNS13]. Gbest [ZK10b]. Gbest-guided [ZK10b]. GCD [AT11, TL10, TLL13]. GDHS [KKSS14]. Gear [BI10]. gene [ALNGMGMM14, Hsi14, SHZ14]. General [AY10, AH10, FO11, Kha14c, PSWW14, AAKB13, AZ14b, AM12a, AZ19, BD14d, Bou10, Bra13, BR10c, CK14b, DGDA10, DESW10, DAEY12, EJ12, FYYT11, Fuk14b, Ger12, HYT13, HZ12b, HZW13, ISÖ14, Jha13, JS14c, Jun11a, KudKL14, KK13d, KP11a, yL14c, LW11b, LWZD10a, LL10b, LZW13a, LTM10, LLJF12, LD14, MYT10, MS14a, MG12b, Ném10b, NNK14, O’R10, ÖP12b, PRT11, PM12, PX13, PRK13, Ray12, RRSPTR14, SP14a, SHMCM12, SSTZ14, SS11b, Sha13b, SDO10, Ste11e, TFMB12, TJ12, TJ13, VV10, Wal10, Wan10c, WCL12, WSL^{+}13, WZSW14, WLY^{+}12, XXS12, XLS13b, XC13, Yan12e, pYmK13, YC14, YX12a, Zha11b, Zha14c, ZZZ14, ZZ10b, Zhu10c, dlSAQ11, BZ10a, CBS12]. generalisations [PPS13]. generalist [LYP14]. Generalization [KU11, LSYS14, NBR10, AKS14, Bah13b, BBF12a, BD14e, CZ14b, DT10, FNVLPD^{+}13, GS12a, HT11b, KK11a, KS11f, LMS10, LWZ13a, LJT13, Lu14c, SSK10, Sie10, SG10, TFM13, wX13b, YH10, ZW13a, Zhu12b]. Generalizations [WWY10, BDGP14a, BEdAdSK14, Bil14, JM14a, LS11c, NR13, WV14]. Generalized [AGKK14, Asl10b, BM13a, BCK14, DAA13, Et13, FLMR14, FZ12, Fin14, Fit12, Gho14, GMP13, KRJ11, Lu12c, MGMM14, Mil12, MAK13, Pet12, RRG11, SAuAH14, SN11, SK10b, SS12c, WHS14a, WHS14b, XYY12, XSX12, YM10a, Yan12d, YZZP14b, ZDZ10, ZgJlK12, ZZ12b, Aba12, ANR10, AKN11, AN– D12, ANR12, ANR13, Abb14, eMA13a, ARR12, AR12a, ABD^{+}12, AK13, AL10b, AM12d, BS14a, BC13a, BT14, BY14, BS10a, BAE13, BTVC10a, Bhr13, Bhr14, BZGK10, BT11b, BTHA11, Bor11, Bot12, BSM14, BGGH12, CJZ12, CS13b, CLCL11, CL14b, CR11b, CCCRR13, CSL14, Cho11a, Cho11c, ĆASH11, ĆSAV11, CZSV12, ĆARA12, CL10c, Cop11, CL14e, CG14e, CISM11, CIN14, CK12b, yDpY14, DL13, DGMR12, DGSB12, DV12b, Din11, Din12, DAEY12, DCK11, DB12b, EDGC14, EDG14, Ere11, Esf11b, EVR12, FTAB14]. generalized [lF12, FMF13, Fuk14a, FIK14, GIR10, GYL^{+}13, GZ10a, GLV14, GS10b, GPP11, GTU13, GZL14, GMSE13, HBTU10, HZ10, HRL14, Hon10a, HL12d, HMZ14, HHX11, HLZ14, HM12, IKV11, Imo14, Inc12, IB10a, IPB10, JSWJ13, JCY10, lJLpWxZ14, KK14a, Kat11, KKS11, KAS12b, KK11b, KA10a, KS10b, KS10a, KS12a, KK14c, KCC12, KPW11, LSS11, LOKB12, Len14a, LR13a, LYS10, LF11b, Li11a, LYS12, LLLS11, LLCX10, LLWW10, LJW10, LCH11, LLW11, LH11c, LQY11, LJCI12, LLSW12, LMS^{+}12, LTW12, LHCI12, LML12, LCL^{+}13b, LJCI13, LCF^{+}14, LO12, Lot14, LLN10, LL10h, Lu14c, LW10b, LHL^{+}13, LZ14d, Ma11b, MAA11, ML12a, MZ12, MKeM13, MDR14, MT14, MS13c, MS13d, Mao11b, MTZ10, MGWG11, MC14c, MCS11, MVS14a, MP14d, MMKD14, MHA10, Mor10a, Mor10c, Mor10e, Mos14, MD12b, MD12c, MD13c, NMIA11, Neu14, NM08, OADL14, yOZ14, ÖBT10, Par11]. generalized [PTS11, PS11, PPV14, QCK10b, Ram14, RCYL10, Ryo10, S10, S.14, Sad11, Sal10a, Sar13, SPV10, She14c, SC11, SSRA12, SSS13a, SS10f, SC10b, SHC11, Son12, SY14d, SK12b, SÇK14a, SAJ14, Sri14, SC14b, SB10, Sun14, SV10a, SV10b, SV12, THH10, TSP10a, TSP10b, TS10b, TQ11, hTlCqL10, TZJZ12, LM13a, TN14, TW10a, TI10b, TD13, Tri14, TC14b, VV11b, Vil10, Vil12, VDK13, WZ10, WF10, WZ13c, WY13a, dWcFwWjZ13, WHW13b, WWW13a, WFL14a, WY12c, Waz10d, Waz12b, WTFC13, WLL14, Wen10a, Wło13, WZL12, WS13, WL13c, XZT10, XW10, XD11a, XQ13, XZ13c, YAW10, YH10, YlTA10, YCXL12, YZK12, pYmK13, Yay11, YT13a, Yu11a, Yu11b, Yus10, ZA11, ZZ10a, ZTD10, bZlD11, ZSWS12, ZDW11, ZWP13, ZS14c, Zhu10c, ZL13d, SSF10, YZZP14a]. generalized-Zakharov [BTVC10a]. generalizedDullin [BK10d]. generally [KXW14]. generate [CGVR14]. generated [AKMM12, BSKL13, CMM12, HMR12, LWWZ11, ZL10f]. generates [B– D11]. Generating [AG14c, SC12, SAJ14, AMFM12, Cao10a, Cao12, CS13b, Cvi10, EMA13b, EBR13, KÖ11, LLSW12, MDD12, Man10b, MCR14, Oss10, Öza11, SPA12, Sah11a, SK14a, Sim11, Sri14, ZZW14]. Generation [FM14, GTPM14, SIR14, Che11d, DW14, KRMM14, MKSC13, PMD14, SSA14b, SBM12, TQWZ10, VV11a, Zhe14]. generation-based [MKSC13]. generation/absorption [VV11a]. Generations [YWLX11]. generator [CL12, ÇT10b, COPP14]. generators [Liu10d, PJ10]. Generic [SKC14, GGFM14, LZLF10, SL13a]. Genetic [KL13a, UJH14, AHS14, BSR10, CW14b, CSF12, DM13a, DHDB12, EH14b, Gwi11, Gwi12, HJZC13, HLC13, KMW12, Kay11, KZS12, LPP^{+}12, Liu10d, LWS11b, Liu11b, LMS^{+}12, MB12, Örk13, PTK14, PEMS13, dejW14, RWBA13, SSC11, SAA14, SVGC10, TMJ14, YY14a, YK13b, YHL08, YY13b, ZaFT11, SKB14]. Genocchi [Ara14, AAS14a, DS11, EDG14, LS11c, Ozd11]. genotypes [WY11a]. genuine [Gal10a, GS14e]. genus [ZC14a]. geochemical [OSAG13]. geodesic [CZSV12, DSPZ13, Sta14c]. geodesics [LWZ11]. Geometric [CV13, qSyGH11, TC14b, AŞK10, CGAK10, CZX^{+}14a, CFSYPP10, CK13b, Cur14, EKK14, Jha13, Ral12, VV11b, WWWJ14]. geometrical [ACG13, BRVR12, GJNN14, TLDZ13, Zak10]. Geometricity [YK13b]. geometries [AS14e, SWL12]. Geometry [PBG14, AMPT14, LZ12a, dALORPGC14, PT14b]. geostrophic [Med12, RM12a]. Gerber [LSS11, LC14d, SMWR10]. Ghodousian [MK10a, MK08]. GHS [CEP11]. GI [LT11, TLW11]. GI/M/1 [LT11, TLW11]. Giesekus [MZWZ12]. Gijón [NA14b]. Gilpin [JV13, VJ11, WZW14]. Gin [IN10]. Ginzburg [Bot12, CC12c, GARK12, JT13, KR13b, Li10c, LS13b, LTJ^{+}11, She10, ZL13c]. GIP [GL14a]. given [BCGK14, B– D11, DS12b, GG14c]. GKdV [ZL10g]. GKS [JL10]. glance [DZ12]. Glass [BBI13a, BBI13b]. glassy [MO12b]. glioma [GS12c]. Global [BKP14, CL10a, CGL12, CJK12, CC10b, CS11b, CWLC13, CS13c, CW14b, CFSYPP10, CKP^{+}13, DH13a, DISS14a, DZZ14, DYL14, EGO13, FHW13, Far14a, GSDW14, HHX12, HX14, Iri10b, JH10, JZY12, KcZ14, KKSS14, Laz11, Li10a, LCH12, LKY14, LDL13, LJKU11, LS12b, LB14c, LLP14b, MXH12, MEL13, Mur14, NM14, OR13a, Pai14, PRK13, Rao14, RRLA14, RZ11, Sal11c, She12b, SM14b, SHI13b, SPWZ12, Sta14a, Ste10e, Ste10f, SA12b, SS11i, SHWX14, SZ14d, TPTW14, TX14, WW10, WY11c, WES12, WY13b, WC14a, WPKE14, WYBT12, XP12, XX10, XCT10, Xu12b, Xu14b, XYL14, YTC13, YL14a, yZ14a, ZMZS12, Zha12a, ZLJZ14, ZW10d, ZW14d, Zuo13, AMPT14, BBJ10, BMD13, BC10a, BBI14, BDDK14, BO13, CSD12, CTY11, dSCAM14, CSF12, DMGK12, DS14d, DI13a, EA13, EEA11, EKZ11, Gal10c, GWM10, Gee12, GZ12b, GC11c, GLW12, HCW11, HB13, HLTT14, JLL12]. global [Kam14a, KA10b, KGB12, Kot14, KL13a, LOKB12, LWQ10, LLY11a, LWZY11, Li14a, Lia11, LWF11, LWFG13, Lis14, LZTZ13, LJ11, LWM11, LLZC11, LWX13, MYL10a, MCM13, Pak11b, PSTL10, RY14a, RK12, RCSH10, RMJM12, SYl14, TIT13, TS10c, VTM14, VDLEK14, VDL14, WL10a, WZZH11, WC14b, WLQ^{+}10, XWWF10, Yan13c, YHL08, ZH10a, ZDZ10, ZXCL10, ZX13, Zha13b, ZZS14, Zil12, BGP^{+}13, CEP11, McC10, OSRPC12]. global-best [EA13, CEP11]. globalization [ZL11b, ZSQ10]. Globally [LP12b, SZZH13, Ste12g, Dia13, WFL14b, ZC12]. globulus [NTdCJL12]. glucose [MNP13]. GM [Tie12]. GMRES [LLBW11, ZW14d]. GN [Abb14]. go [AABV12, Koç14]. Goal [ÖB11, AL10b]. goat [PFG^{+}14]. golden [CHMT11]. Gompertz [JK14c, MGGS12, MGSMRÁ14, RRRRTR12]. Gompertz-type [RRRRTR12]. gon [KCLC13]. good [EH14b, PLTH12]. goodness [PTW11]. goodness-of-fit [PTW11]. Gordon [BT10b, IB10b, KRS10, Lia14, SXL12, TB14, WZJ13, ABP13, CYY14, Don14, HRL14, IISA10, Jan14a, JW12, JCB11, KS10b, Li10d, LTZ12, LGN10, Mer14, NC13a, RV13b, Sha10a, SP13a, YZ10]. Gottwald [BK10d, MHLL11]. Gould [KAG12]. Goursat [Dri13, GS11a]. governed [CL14c, GS14b]. governing [LW13a, PP14a]. Gower [LL13b, MM13d, TW11a, YL14a]. GP [DRT11]. GPBi [yZxGyM10]. GPBi-CG [yZxGyM10]. gPC [CCCRR13]. GPIU [MC14c]. GPS [XWWF10]. GPSS [LW14d]. GPUs [CFGRV13]. grade [AA12d, AM13c, FHAL14]. graded [ASKA13, DRJ11, EEM12, EKSO14, HA12, Ngu14, SM12, Zen12, ZZHZ14]. grades [VD14b]. Gradient [XLY10, And13a, CN14a, CC14a, CLD12, CXD14, Con13, Cov11, DW12, Dan14, Don12,

– Dor12a, DSPZ13, EES14, EL14, EHSM14, FLMR14, HCL^{+}13, HLZ14, HZW13, HL13d, HL14d, Iid11, KR13b, LF11a, Lin10, LL11d, LWN14a, LWS14, LL14f, LZX14, LP12b, LP13a, LWM11, LLZC11, Nar13, QLL12, RSN11, RMJM12, SWX10a, WDL12a, YZ13, ZDZ10, ZL11b, ZJ13, ZLZ14c, ZSQ10]. gradient-based [EL14]. gradients [PTS11]. gradually [WT11a]. graduation [EAAN^{+}14]. Graetz [Che10c]. grained [Sie14]. Gram [AM12c]. Gram-type [AM12c]. Grammian [MAA11]. granular [ZM14c]. Graph [LjH11, Aza14, ÇM13b, Das11, DMP13, GGM14b, JM14a, Lin13a, MDÇ13, RJ14, SXGR11, TRA13, WLJ11, ZP12, ZLSW13, ZZR11, Zhe14, NSV13]. graph-theoretic [ZLSW13]. graphene [DFGL14]. graphicable [NSV13, NRAV14]. graphics [BX12]. Graphs [CDIS10, AdC14a, AAMB14, ADÇ14b, AMV12, Das11, DBG11, DD11, EVE12, GBA14, Gol11, GMM^{+}14, HY14, HG14, HLH12, HLS^{+}14, IB11, Jar10, LMY12, LLJF12, MCM13, NPA11, RLC11, Sha13b, SI14, SB12a, Ste12a, Ste14a, Sze10, Sze11, Sze12, TRA13, TH14b, Yan10c, Yan10d, YRV10, Zho12a]. Gravitational [GVW^{+}14, OADL14, TTWL13, MHS12, YJCC14]. gravity [AAMADH11, DGCFM11, HLT14, LGS^{+}11, VS10, Vis14]. gravity-capillary [HLT14]. Gray [GZW13, GX13, JZY12]. grazing [XLW11]. Greatest [Tsi11a]. greedy [WY11a, ZL13d]. Greek [Tsi11a, Tsi11b]. Greeks [XLY14]. Green [CCMV12, Che14j, HYT14a, KL13b, SATdV13]. Gregory [WHS14a, WHS14b]. grey [CH13, GXF13, LXGM14, Tie12]. GRG [EES14]. grid [BY11b, Boy13, CB10, KP14b, Mit13, NPA11, Tad12, Wah13, WFZ13, dFGAN12, PV11]. grids [BA14b, KP11a, LY12b, Ros11a, SY12, ZhGQ14]. Griffith [GYS13]. Grimshaw [BI10]. GRLW [GLR12]. Gröbner [Duk12].

Gronwall [AY11a, FZ12, WZ10, YG11]. grooved [KML14]. Gross [JZ10, Yu13b]. Grossberg [LZZL10, TJ13, DZZ14, HLTT14, JHP11, LX12, LX11b, MPSA14, SAMA11, SZ14b, TJ12, TJ14, WZ13a, ZLSW13]. Grossberg-type [LZZL10]. grossone [DD12, Zhi12, Mar12]. Ground [JŞ11, HL14b, LGZ14b, YY10]. Group [BD14e, EM11, Gün10, JK10, LL12a, MKeM13, Naz10, PPV14, RYÖ11, SGZC13, ZB12, eMA13a, eMBHA13, AT10a, BSKL13, BX12, CL11b, CS11b, Das13, DRT11, GS14a, HZZ14, Jar10, LIX14, LSW^{+}14a, LY12c, LZB14, LLZ14b, LWW14, MT12b, MD12d, MD13b, PH10b, Sah11b, Sht11, SS11i, Sze10, Sze11, Sze12, Tra12, WL14a, WLH11, XLLL14, XG10, XYL14, ZPZ14]. groups [ABG11, AH13a, sCCLY10, sCCYW13, Han14, LHL14, LWZD14, MCL14a, MM13c, OGYE13, pZLC10]. growing [CC11c]. growth [BZ11, BB11a, GG14a, GS12c, HRH14, Li11b, Mad14a, MGSMRÁ14, PKD14, PZ14a, PV11, SS11h, ZK14]. Grüss [AR14a, Alo14, HN14b, Nie11, WV14]. Grüss-type [AR14a]. Guaranteed [KWJ13, Kim12, LPJ^{+}12, LJPJ12, LWN14a, LZ14d, TPT12, XSX12]. guaranteeing [DGH11]. guidance [GG10b]. guided [ZK10b]. guideway [ZZ13a]. Guiding [LO12]. Gulf [AWHW11]. Gupta [Deo12, Yad14]. Guseinov [MÇ12a]. gyroscope [JSF14, MSJ14]. gyroscopic [Yua11].

H [Can11, CZZ^{+}11]. H-R [CZZ^{+}11]. H1N1 [RG14]. Haar [GL11a, LZ10b, Ray12, RP13, SuR13, WMM14]. hABCDE [XMA14]. habit [RZHH13, RZHH14]. habitat [GW13]. Hadamard [AKÖ11, Dra11, HXQ14, HYT14b, IW14, LPH11, LSY14, Mat14d, XZW11]. Hadamard/DCT [LPH11]. haematopoiesis [XX10]. Hager [LWN14a].

Hahn [BPR13, BR14c, CS13b]. hairpin [XK14]. Half [ABK14a, MS11, AJR12, And11, AA12d, CS10a, CLW11, Che14j, CZ10, DF11, FKH10b, GMKV12, GMV12, HL13d, Jia14, KGM14, MK14b, Rah13, RY13b, RY13a, RY14b, Sab13, Sab14, SAO12, SEL14, Sin10a, Sin11a, Ste10b, Ste10c, SS11h, Ste14d, Vis14, ZA14b, ZZHZ14]. Half-discrete [ABK14a, RY13b, RY13a, RY14b]. half-intervals [Ste14d]. half-line [CS10a, Rah13]. half-linear [DF11, SAO12]. half-plane [CLW11, Che14j, Ste10b, Ste10c, SS11h]. half-quadratic [HL13d]. half-space [FKH10b, GMKV12, GMV12, KGM14, SEL14, Sin10a, ZA14b, ZZHZ14]. half-spaces [Sab13, Sab14, Sin11a]. half-step [MK14b]. Half-Sweep [MS11]. Hall [Gad11, Gad14]. Halley [CTV13, LLK14, NCS12, NS13a, RA12c]. Halley-like [NS13a]. halpern [CLCZ12, BL11, NS11]. halpern-type [CLCZ12]. Hamburger [Tag14]. Hamel [UW10]. Hamilton [CK12c, Tsi11b]. Hamiltonian [BDL13, CKH11, CH12, CLTL13, DSX11, DMP13, DMR10, DLSS11, FNV10, HZFZ14, HC14, HZ12b, HHR13, KZ14b, LSW14b, LS14d, Ma11c, MDD12, RS10, RS11b, Sun11b, Vir13, WDL12b, WHG13, Xu10b, XTAJ12, YWT14, YDXC10, YHWZ11, YT13b, mZT11, ZT13, Zha14d, ZGW14a, dOL13]. Hamiltonians [FG14a, WZL14b]. Hammerstein [CLZ11, CXD14, CS13e, CD13, EGH12, KJ10, O’R11b, OM11, OO14, PR12a, She14b, Tur14b, Yan12f, Zha12c]. Hammerstein-type [OM11]. Han [PS14b]. Han-type [PS14b]. hand [EA14, HH12, ZW14d]. handling [Waz10a]. Hankel [AdFY13, AM13a, AM12b, CDGR13, GG14g, PPK11, Rim13b, XWWC10].

Haplotyping [WY11a]. Harary [GMM^{+}14]. hard [DKU13, LNdCJV14, WW13i]. hard-coupled [DKU13]. hardness [Rud12]. Hardy [AY10, Hon10b, Kan11, SS14d, Ste10b, Ste10h, SU10b, SS11f, SS11h, SS12c]. Hardy-Orlicz [SS14d]. Hardy-type [Hon10b]. Harmonic [CK10, PQR14, BÖ11, EL13, BDN^{+}14, BE13, Che11e, CS13d, Cho11c, CC12c, CYZ14b, CL10c, Cvi10, EDGC14, GL10c, GQ11, GM14c, LW12a, LCC12, Maa14, NY12, PPMCG14, dlRS14]. harmonically [IW14]. harmonious [RWBA13]. Harmony [CkPqL12, CAC14, KKSS14, ABDKA12, ABKG^{+}13, Ala10, EA13, GS10a, Gee12, KA13a, PSTL10, TNB11, VTM14, WY13b, bOYqGL^{+}14, CEP11, HAA^{+}14]. harsh [URR11]. harvest [CCK11, LZLY14]. harvesting [BLR13a, CJK12, CDK13, FW14, KG11, LR11, LWS11a, LB14d, LYP14, PM14a, SP14c, SVGC10, WWS10, WF11, Wan14b, WH14, ZW13b, ZT11]. Hassell [Wan14b]. Hatano [RV10a]. Hausdorff [AR13, DM10b, FT11, GG12c, GT10a, ROA11, Tag11]. having [BP10a, EM13, HMD14, Pan11, YX11, Yua12]. Hayat [PK18]. hazard [SSS14, SKIH12]. HBFGen [dYyR13]. HBV [VDL12b]. HDMR [Yam14]. headways [LXH13]. healthy [FdOdSM14]. Heat [Cor11b, FHAL14, GSF12, KS12b, MWAM14, ASM11, AAP10, AHMS14, AY11b, AM14g, BT10c, Che10c, Che11e, Che10f, Che14k, CL10b, DZLW13, EEEE05, FY12, FMV14, Fit12, FKH10b, GSRR13, HDK^{+}10, HMMK11, HMD14, HWPY14, HHDM13, Ish10, IKL11, JHA14, JZZ13, KM11a, Kha14a, KS11b, KM11c, KC13b, KCBH14, KS13c, KS10c, KML14, KL13b, LJ12a, LL10e, LY12b, LW13c, LSF^{+}13, MFZ12, Mel14, MSR14, NTT10, Pal10, PC13a, PMV14, Pan12b, Pan15, PMD14, QW11a, RRP14, SWL12, SK14a, SSP11, SBS10, Sha12b, SM12, SWW14, SZW14, SGC10, SSS13b, TSS^{+}14, TTQ11, Tua13, VPS12, VV10, VV11a, VZ11, WaZW13, WZ14a, WS10, YGL11, ZMZ11]. heat-conductive [WZ14a]. heated [BD14b, Das13, FFOC14, HAMA14, PK18]. Heath [BD14e]. heating [DKU13, FD12a, KMD13, KV12, MVGS11, ZAA^{+}14]. Hecke [AS11c]. Heisenberg [LTG^{+}12]. Hele [Cur14]. helices [KGH12, OGYE13]. helicity [CRW11]. helicoidal [JK13]. helicopter [GG12c]. Helmholtz [CNRY10, EZ12, GG12b, GL10c, Kle12, PB14b, Sch11a]. help [BP13]. hematopoiesis [BBI13a, BBI13b]. hemi [Naz10, SLZ11]. hemi-relatively [SLZ11]. hemi-spherical [Naz10]. Henon [xLhYlZ13, WLXX14]. Henry [YG11]. heptadiagonal [Elo13]. herd [LTHG14]. Hermicity [ZW13d]. Hermite [AG14c, AB12a, AKÖ11, BCM12, BCDI14, BNBV10, CCJ11, Dra11, FBGH12, GGP^{+}13, GK14, HXQ14, HYT14b, IW14, JZ11, KASY10, KYH10, KKA12, LZ12b, LSY14, Lu12c, LD13c, PC11, XZW11, YNBV11, Zha10f]. Hermite-base [KYH10]. Hermite-based [KASY10]. Hermite-Hadamard [HYT14b, LSY14]. Hermitian [CC10a, CMX14, DGM^{+}14, FG14a, GH10c, GW12, mHhL10, KM14a, LW14d, LC14b, MD12b, SSN^{+}12, SBDB13, Wan14c, WYW14, YAW10, ZGZ14, Zhu12b, ZZZL14]. Heronian [LLZ14b]. Hessenberg [CY11a, IA14, Li12a]. Hessian [Gu13, WZ11f]. Heteroclinic [CST12, AFSMRL14, BY11a, EDYSA12]. heterodyne [Rem10]. heterogeneity [EVE12, NR12]. heterogeneous [Alf10, Alf12, ASKA13, qClLyW10, CN14a, CS13d, DE11, EAE13, FXD12, GAK12, GL13c, GMKV12, KHD12, KS13c, KGM14, LZWZ14, Pal13, ZYZ14a]. Heteroscedastic [LCCA11]. Heun [MSJ14]. Heuristic [Oss10, PD13, YH12]. Heuristics [AA12c, WY11a]. hexagonal [BOR14a, Caz14, GL11b, GYS13, KA12, YS12]. hexahedral [WJ11]. hexanomial [KLW14a]. HF [DSGB12, DSM^{+}12]. Hidden [RK14a, ZA14a]. hiding [CHYT12]. Hierarchical [CXD14, ZJD12, RG14, TRA13]. hierarchies [ELS13, GP11b, GMP13, LTG^{+}12, Yu12a, ZGW14a, ZLL13]. hierarchy [DG14b, GG14b, Kud10b, Kud12, LCH11, Ma13a, MM14a, MGYQ12, SMM14, WX13a, Wen12, WHG13, Xu10b, XTAJ12, YDXC10, YL10c, Yu11a, YL13, YL14d, ZHT13, ZDW11]. High [BBB12, EL13, CZC11, DLS13, EJ12, EVDG12, KR10, LC14a, LNG12, MG11b, MG14c, NN11a, NZ14, OSRFF14, SZ11, SL14b, SLW12, Tur14b, Van10, XLGW12, ADÇY11, BTS12, CL11a, CFZ14a, CC14c, CK14a, CG12b, CHMT11, CSSSV13, DNSM11, DFM10, EASK10, EDD11, GT13, GC14, GG12b, GS14b, HZ13, HB11b, ISG11b, JW12, LW06, LY14b, LGN10, LW10a, Liu13b, LSW14b, LPPG14, MG10a, MS12b, MSS14a, MAD11, NSYA14, PP13b, Res11, Res13, RSK11, SDL10b, SDK10, SU12, SMR14, TD14a, Tur14a, WDBM14, Wan10d, WY10a, jWtW13, XBC^{+}14, XZZX14, ZFH14, ZMZS12, ZZWF13, Zha13d, ZHGP14]. high-carbon [Liu13b]. high-density [EASK10]. high-dimensional [GT13, XZZX14, ZMZS12]. High-order [EVDG12, LNG12, NN11a, SLW12, Tur14b, Van10, ADÇY11, BTS12, CHMT11, HZ13, LW06, LY14b, LGN10, LSW14b, SDL10b, Tur14a, ZFH14]. High-precision [BBB12]. high-pressure [SMR14]. high-resolution [MAD11]. Higher [BDD12b, BC10a, BK14, Gin12a, HXZ12, JMB12, OP14b, PN11a, PD13, Tri14, Wei12, AS11a, AS14e, BD12a, Bai14, Çak10, CW11b, Chi10, CGK12, CP10, DHI13, DN12a, Dia13, DR11b, DR11c, ESAH10, EHSM14, FPTS14, GK12a, GMR11, Gro13, GV13, HL12b, Hwa11, Hwa14, Iri10b, JH10, KBK13, Kar13b, KA13b, KK13a, Kyc11, LS12a, LTOX11, LW11c, LLW10b, LW12b, LT14, LL12f, MMS14, MK10b, NN13, OO11, OO12, hPnZrL12, lQZWC14, RSBK14, Rao14, S.14, SH12b, TA13, WDL10, WH12, WW14b, WYTL14, WYK11, Yan13e, Yu13a, ZLW12a, ZLW12b]. higher-dimensional [Bai14]. Higher-order [OP14b, PN11a, CW11b, CGK12, DN12a, Dia13, DR11b, ESAH10, GV13, HL12b, Iri10b, JH10, KBK13, Kar13b, LTOX11, LLW10b, LW12b, LL12f, MMS14, MK10b, lQZWC14, RSBK14, S.14, TA13, WW14b, WYTL14, Yan13e, Yu13a]. highlighting [MGSMRÁ14]. highly [BB14a, Che14e, KX10, KX11, LCD12, MS13e, Mil14a, Mil14b, MX11, TK14b, XH13a, Xia14, XX14c, ZXF14]. Hilbert [ABK14a, AT12, ACK14, AASS13, Buo10, CZ11b, FLP11, GPK13, Hao10, Jun10b, Kim14a, KV14, La 13, LZ10d, LHCI12, LG14, MMS– D13, Occ11, PDBRFER14, RDD14, RY13b, RY13a, RY14b, RY14c, She14b, dlRS14]. Hilbert-type [ABK14a, KV14, RY13a, RY14b, RY14c]. Hilfer [GGPT14]. Hill [Zou12]. Hindmarsh [WS12]. Hirota [Aba12, eMA13a, CP12a, GH10b, GZ10d, Hon10a, MZTT12, MKeM13, XD11a, dYyR13, YWCH11, ZMG13, jZzFbL10, ZZ11c]. histogram [CHYT12]. history [BCEAV12, FMS14, Lol12]. hit [Cor11a]. HIV [BPP12b, BMH12, CGL12, CGW14, CFL14, GC12, HB13, KLY12, NTS11b, PC14a, RM12b, TX14, WES12]. HIV-1 [BPP12b, CGL12, TX14]. HIV-TB [GC12]. HIV/AIDS [BMH12, CGW14, NTS11b]. Hochschild [Sht11]. Hodgkin [MWJW10, WMJW11, YWW^{+}14]. HODIE [CG12b]. Hohenberg [ASVV10, VDOG13]. hold [ZL14a]. Hölder [HMGC18, AOM^{+}11, AR12a, Alo13, LL14d, PG11, RA10b, TW14a, ZS14d]. hole [CC12c, GL11b, MWY13]. holed [dCDNRH12]. holes [Vod13]. Holling [BWaY14, CHC11, HL11e, LZ10e, LY11b, Mis11, PLF13, SP14a, SDL10a, TW11a, ZjHO13, ZM10c, Liu13a, ZL11a]. Holling-type [SDL10a, TW11a]. hollow [DRJ11, XDG12]. Holm [BK10d, CQJ10, Den11d, IKV11, Lai10, LC14a, LZH10, LLR10, MHLL11, TW14a, VDK13, XZH13]. holomorphic [Man11]. Homoclinic [CM14, Kon14, LL12g, WqXlC11, WC14c, AFSMRL14, BY11a, Bel11b, CI13, CT11a, CTA13, EDYSA12, Li10c, LSLL14, LR14b, WLL14, ZT12]. homogeneity [Zil12]. Homogeneous [Ign10, AAMAS11, AAMADH11, BP13, Cha13b, CV13, CT13b, DBK14, FMV14, GNS14, GMKV12, GMV12, HK11, MR11, PDBRFER14, ROK10, SS10b, VS10, ZLY11a, ZF13a, ZZHZ14]. Homogenization [GL14b, SSS13b]. homotopies [BDE^{+}14, HSW11]. Homotopy [BS12b, MH14b, ZC13b, CJ11b, Dum10, EK11, ESEEKH12, Fer14, GL10c, GML13, HKI12, HET12, HSW12, HZW13, IISA10, IJ11, JZW12, KSFG14, sL10a, Liu11c, Mar11, MO12a, NK12, Odi10, RV13b, VLMHK^{+}13, Wan10e, ZLLL12, ZSD10]. honey [Hor10]. honeybee [KHD12]. Hooke [RCSH10]. Hopf [AKMM12, BPP12b, BH12b, ÇM13a, DWP14, DJY13, FP13c, FL11, GLM14, JZY10, KHA12, LL13b, Li14b, Ma12, MLY12, Muk10, QL12, SPV10, SZ14a, SHZ14, SHM13, TX14, WDZ13, WZ13b, WY10b, Yan13b, YJ14b, ZZC10, ZZZ11, ZLY11a, ZZW11a, ZCZS12, ZZ13a, ZL10d, Zuo13]. Hopfield [FOL12, JKW^{+}11, LY14b, SD10, WiN12]. Hopper [KAG12]. Hopscotch [Har10]. Horadam [YT13a]. horizon [KM10, ZBM10]. horizontal [Che10f, DNI^{+}10, FFOC14, LL14b]. host [CL10a, Qiu11, VDLEK14]. host-vector [VDLEK14]. hosts [TFMB12]. hot [GGYW11, VO14]. Hough [CHT14]. Householder [SSCT14]. hp [CL14c]. hp-FEM [CL14c]. HSS [GW12, LW14b, YAW10, ZZZL14]. HSS-based [ZZZL14]. HTLV [MEL13, VDL14]. HTLV-I [MEL13, VDL14]. https [WHS14a]. Hu [ZZ14d]. Hua [Ibr14]. Huber [Bie11]. Hukuhara [Mal12a, Mal12b]. hull [ZTY^{+}10, WZ11g]. Hulthén [IS10a, IBS11]. human [AAP10, CE13, TGL14, ZK14]. Humbert [AŠA11]. humor [FdOdSM14]. humoral [WPKE14]. Hurwicz [SZH13]. Hurwitz [AA12b]. Huxley [GZ10a, KCC12, MWJW10, WMJW11, ZYZ12]. Huygens [Neu14]. Huygens-type [Neu14]. Hybrid [AS12a, BL11, CIVS14, EH15, HDZW11, HJZC13, ICS^{+}13, KZS12, LKH12, NSNEV14, OM11, Tsi11b, WLH12, YWYY12, YZK12, YC14, AISS13, ARS14, AS13b, BLK11, BAE13, CDS13, CNTG12, CR11a, CCA11, CDMT12, CSF12, DSGB12, DGSB12, DSM^{+}12, Dim13a, EEA11, EH14a, GW11a, GS14e, HYD10, HH11, Hsi10, HA11, Jat10, KMW12, KS11a, KR12, KDS13, KI12a, KGB12, LXZX13, LZGZ11, LL14f, LH14c, MH13a, MM10a, MLGdSC11, MHS12, Ni13, Örk13, PS10, QCK10a, dejW14, SPK14, SGT10, SLZ11, SCLC10, VMS10, WFW10, WY10a, WJ12, Wan13c, WZWW13, WWS^{+}14, WZW14, XMA14, YNL10, YXZ10, YHL08, YY13b, ZLD11, ZANA12]. hybridization [PM14a, TPAB11]. Hybridizing [HAA^{+}14, RAZES13]. Hybrids [LJ12a]. hydraulic [KP14b, KML14]. hydrocephalus [WDS10, WDS11]. hydrodynamic [VPS12]. hydrodynamical [EN11]. hydrodynamics [YHH11]. hydrogen [AMM11]. hydrological [JLHW10]. hydromagnetic [ASM11]. hydrothermal [YJCC14]. Hyers [AM13b, AM14d, BC12e, PR12b]. Hylleraas [HMIZ13]. Hyper [TV13, Che11b, DA14]. hyper-chaos [Che11b]. Hyper-Riccati [TV13]. hyperasymptotics [Boy14]. hyperbolic [AMIR14, ANS14, ADP11, AY11c, AZAK10, BR13, sCWW^{+}14, CS14a, CZC11, DS12a, Dub10, GZB12, GS11b, Guo10, HS11a, JPM12, Jun10a, KudKL14, KX14, KCC12, LP12a, Li12b, LL12e, Mar12, MB13b, MB14, MG14c, MK14b, MSS14a, NS12, Neu14, PTY12, PTY13, PP13a, RY14b, SY10b, SL14b, SSK12, SV10b, TV13, WKLW12, Wan14f, WSD12, WLL14, Yan10b, YTY^{+}14, ZB13, ZF11, ZY13b, ZMG13, Zhu12a]. hyperboloidal [DTN14]. hyperchaos [ESNE^{+}14, Zhu10b]. Hyperchaotic [CR12, GS13b, Has14, LL10a, MZXZ10, NSNEV14, OGJLCCB14, Rec14, yWZhL14, WLXX14, WWS^{+}14, ZDB10, Zhu10b, Zhu10a]. hypercomplex [Fau14, SE12]. hypercube [Lin13a, LTT^{+}13, Ou10, PLTH12, STH10]. hypercube-like [LTT^{+}13]. hypercubes [CHF14, HZFZ14, Sun12, Yan12c]. Hyperelliptic [FcDDqZ10]. Hypergeometric [Boy14, HM10d, AGLR13, Çet13, LS14c, LMA14, SMDZG13, SKM11, SC12, SÇK14a, SAJ14, Sri14, SC14b, Sta11, ZZ11a]. hyperplane [YKI10]. hypersingular [GÖ14, ZLYL10]. hypersonic [AHS14]. hyperspectral [CEM11, VOL14]. hypersphere [YD13]. hypersurface [KP14a]. hyperthermia [GSRR13]. hypoglycemic [FGFRV11]. hypothesis [zM10b]. hysteresis [CS10d, LCC12, MB10, MB11].

Iang [WZ10]. ICA [ICS^{+}13]. Ice [NATQ14]. ICVMLS [ZL14b]. ICVMLS-Ritz [ZL14b]. IDA [LP10]. IDA-PBC [LP10]. ideal [Che12a, DU13, FZLZ13, KAF13, Sav10, TR13]. ideas [Dua11b]. idempotent [BBT10, ÖKU12, Tos14, Wri11]. idempotents [Den13]. identical [BBSC13, Gee12, LYJP13, WS12, Yan10a]. Identifiability [VJBDV13]. Identification [SSP11, WL12b, Aha14, BB14c, CLD12, DSM^{+}12, DHQ14, EA14, GxFX^{+}12, GjLxF^{+}13, JYF13, JWW14, LL13g, VJBDV13, WCL12, YZZ^{+}14, ZAY^{+}14, ZYG14, ZnXwY14]. identify [CM10, SBS11]. Identifying [DYYL13, ZW13e, JWWY13, OD11b, Yan11c]. Identities [DD13, KP13, Ma11b, UY11, Ara14, EMA13b, Fen11, FPP11b, KK13a, KK14c, LDS14, Meh11, Sri11b, Sta11, Wło13, YB12, dF14a]. identity [DWZ12, GMMIDM13, Yos13]. IDR [AS12a]. IDRstab [AAI14]. ignition [AMM11]. II [MSC^{+}13, BWaY14, Cra12, FP13c, GAK12, LZ10e, SDL10a, SSS13a, Tse12, Wil13, ZL11a, ZM10c]. III [CHC11, LY11b, Mis11, Mus14, TW11a, ZjHO13, ZANA12]. ill [AG13, FLMR14, GPK13, GHS11, GTM13, gKW12, Liu12a, LW13d, PMC11, Sal11b, SA10a, VG14a, YLLL13, ZM12]. ill-conditioned [GTM13, PMC11, Sal11b]. ill-posed [AG13, FLMR14, GPK13, GHS11, gKW12, Liu12a, LW13d, SA10a, VG14a, YLLL13, ZM12]. illustrated [AM14a]. illustrations [TPAB11]. Image [CSLL10, CL13a, YWY^{+}13, AY14, BB14c, CL12, GP13, GG14d, Hor10, HL13d, LP11, LLP10, LK12, MO14, MD14, PKK10, PTK14, PS10, SCC12, YPSW11, YWD13, ZL13d]. image-kernel [MD14]. imagery [SKB14]. images [GXSH14, HS13, KHK^{+}11, SNZL13, VOL14]. imaginary [HZW14]. Imaging [LXZ14, CvW10, JKP14b, TSN13]. imbeddings [KM13b, KM14a]. immigration [AABV12, SM14b, SZ14d, WL13b]. immiscible [TSJ13]. immobile [LLT^{+}14]. immune [FW10, GW11a, GHLF14, HY11b, KB14, LPY11, LZGZ11, MEL13, RRLA14, TFMB12, TX14, WES12, WPKE14, ZDH14]. immunity [JZY10, ZX13]. Impact [PPK14, Ask14, HT14, SVGC10, SY12, WW11e, XLW11, ZM10c]. impacts [LZ10c, Pfe10]. imperfect [BJB11, CFL14, KL14a, SM11, Sar12a, Sar12b, WSY14]. imperfections [BRVR12]. imperialist [EYAD13]. Implementation [KM13a, AQ11, Ach10a, CFGRV13, CS10b, Che14f, HCC^{+}12, MLM^{+}14, MN14a, SA12a, WLP^{+}14, XH13a, XDZ14]. implementations [And12b]. Implementing [AS13b, MAD11]. implications [GBD^{+}12]. Implicit [Bel12, Mil14b, OO14, WL10c, ANS14, Cai11, Cao10b, CR11b, CLTA13, Che13e, CZC11, CK12c, DWZ12, DC13b, EA14, Gu10, HHX11, HG13a, Kal13, gKW12, KAS12b, Kim14a, Lia14, LR10, MC14a, MS12b, NA14a, NN11a, RP14a, SMNN14, SCK14b, TD14a, TD14b, WP14, XZY14a]. implicit-explicit [XZY14a]. implicity [Din12]. Importance [Gat14]. imposed [GS12c]. improper [TÇ13]. improve [LT12, VL12]. Improved [Den11a, GD10a, JPK12, JLHW10, Jia12, KCZK11, KK13b, LLP14a, LYWL12, MI10, MPM14, Mor10c, QLCF10, Rad11, RMY14, SS12a, SM14c, TZ11, TXX14, UT13, WL10a, WS11, YK13a, YH14b, YYW12, ZB13, ZPXX14, ZZGL14b, ZYZ14b, AH13b, BMD13, BTV12, Che11e, CG14c, CAC14, DW12, DZLW13, EA13, EDAM10, FHW14, FL14, fGyLJ11, Ge14, AA14a, GS10b, GK11, GZZ10, HS14a, HS14b, HA10, HSD11, HJZC13, KPP11, KPP^{+}14, LYGZ13, LTHG14, LWW12, Lu14a, MS13b, MF13, dMbJmY11, SL10b, WWW12, YTCY12, ZL14c, ZLLL12, Zha13e, Zhe13, yZxGyM10]. Improvement [AEESMA10, NP10a, DGS14b, LP10, LLK14, LW12b, NS13b, SK14b, ZZGL14a]. Improvements [THYC10, Tse12, Bab12, Mor10d]. Improving [AHVR14, BDGIPG12, EHRV12, GS13c, Xia10c]. impulse [BH11, CLD12, DSM^{+}12, FY12, Xie11b, ZW10d]. impulses [BDL13, CH12, Hu13a, LLS11, MYW13, POR13, SD10, WL10a, WLZ12b, Yan12e]. Impulsive [CSSN11, LXWW14, MYJ10, RCS14, RM12b, SIV10, ZT11, AZ14a, APJ14, BD11, BDD12b, BFM13, Cai10b, CS10a, CT12b, CS13c, CZZ14, CZ11c, DLL11a, DIAQ13, DL10, DSH12, DX14b, FS14, GWH12, GSC14, HLZ12, HHR13, Hus10a, JLW11, JSH10, JS13b, KS11a, KZ14b, LZZL10, LX12, LZW13a, LAF13, LXC14, LX14a, LS10c, LLL11a, LZTZ13, LW14h, LDZ10, ÖZ10, PZ14a, PLF13, PR10, SDL10a, Sha11b, SL13a, SL11b, SS13c, Sta14a, SHM12, TJ14, WFW10, Wan14a, WLLY14, WZL14a, WH14, WC14c, WD14c, XLS13a, XCT10, XZ12, YWZ10, YZW^{+}10, YSZ12, YC14, YH14b, YXO14, ZSZ12, ZZZ14, ZL14e, ZLJZ14, ZSL10, ZZWY14, ZW14c, ZZQ13, ZGW14b, ZMC10, uRE13]. impulsively [JDN14]. inaccurate [QGD14]. Ince [PHGV13]. incidence [BR10c, CGL12, CS11b, DMS11, DIAQ13, EMM^{+}12, GYL^{+}13, GZ13, HYT13, LOKB12, LSJ10, Liu11e, Liu14b, MS14a, MEL13, PPC13, QaC13, SS11i, TX14, WL14b, Xu12b, ZHXM14, ZZL14]. incidences [CL10a]. inclined [BD14b, MBR10, PC13a]. includes [Imo14]. including [AHDD13, IS10a, Mat12, Öza14, SSS13b]. Inclusion [SKM11, Aou10, DH14, Hao10, yL14c, MPM14, NSL14, PMM11, SMNN14, Ver10, ZL13e]. inclusions [CW12, DT14, Goo13, HZ10, KKS11, KAS12b, pLjH10, MM10a, SGA13, TJZ13, VRMT14]. income [BM12b, BM12c, BM14c, HZ14d, Kar10, ZY12]. incomes [HY11a]. Incomplete [ZMZ11, Boy12, Çet13, DRT11, EM12, Fuk13, Mam11, SC12]. incompressible [AAMADH11, CRW11, DAA13, HZ14a, KK11b, Liu14c, LMB13, MNOB10, MAD11, Neu12b, RY10, RW14a, SYK10, Svá11, TD14b, WYL10a, Wil12]. inconsecutive [Hua11a]. inconsistent [PS14b]. incorporating [CDK13, LL13b, MCWC13, Yan13c]. Increasing [ACQOR11, Bor14b, PSJGLS14]. Incremental [LZW12b, BS14c, LCC12, PB14b, SW10a]. indefinite [Bah13a, Bat14, Cha14a, yCfW12, CT14b, DSX11, GP10, HT11a, Kat14, KW10, Wan10a, WHW13a, wX13c, wX13b]. independence [Ahn13, SY14d]. independent [AB14a, Asa10b, DLZ14, JS14c, Sáf14, STH10, TFM13]. index [AÇ13b, Aza14, Ben14b, Dim13b, FGIV12, HLH12, LW10a, LGMZ12, MT14, MCM13, NOS10, QHL10, Rem10, WHL14, WY14, ZP12]. indicative [Kau12]. indicators [CGAK10, Waz14]. indices [AMFM12, BOR14a, BCF13, FGD13, GG14c, GZW12, HI14, HG14, IHM14, RLC11, RS11b, Ste12a, VKV14]. indirect [CTLL11]. individual [Koj14b, Sie11]. individual-based [Sie11]. individuals [RWBA13]. induced [BWaY14, FGFRV11, Gad11, GRV13, Gra13, Har14, KB14, KS13d, Liu14a, SLYJ13, TC14b, WWZ10b, WLMJ10, WMJW11, YWW^{+}14, ZAA^{+}14, ZX12, ZCL13a]. Induction [DKU13, AH11d, KMD13]. industrial [dCDNV^{+}13, Liu13b]. industry [SLBS11, WT11c]. inefficiency [ZF10]. inelastic [WW14e]. Inequalities [AJF13, CC14b, DM10a, HXQ14, LMS11, Mat14c, NS12, Neu12a, AY11a, AR14a, ABK14a, AOM^{+}11, Akt14a, Alo14, AKÖ11, BD14a, BFXZ10, BBK13, BH11, Buo10, Çak10, Çak13, CAY11, CLCL11, CL14b, CB12b, CM13c, CCA11, CS13f, CAP14, Den10, Dra11, DCK11, Dra14, FH11a, FZ12, FMF13, FPTS14, GZ12a, GM14b, GMSH13, HM10b, Hou13, Hu10, lHyLfG10, Hwa11, HHT13, Hwa14, IW14, ISÖ14, KR12, KV14, LH14a, LN10, LS14e, MAO10, MS13f, MCL14b, Neu11, Neu14, Nie11, NNK14, Ota14, PS11, PS14b, Sad11, Ste12n, SH13, TFM13, WZG12, WYK11, WW13j, XZW11, Xia12, YZL10, YNL10, YG11, YSW11, ZHY11, mZT11, bZlD11, ZL10c, ZCZ13b, Zhu11b, Zhu12a]. inequality [AMO^{+}10, AY10, Alo13, AG14e, BPP12a, BA14a, CS11a, CS13a, ÇT10a, DHZ11, DZ13, DCK11, FC14, Hon10b, Hu11, Hu13c, HN10b, HN14b, HYT14b, JKP14a, JHX14, bJLyZlL14, JW10b, KR12, KPW11, KPK14, LSYS14, LSY14, jL12, LG14, LT14, qMWjHpX13, Mat14d, Mor14b, PPS13, PK12, PS11, RY13b, RY13a, RY14b, RY14c, SK11a, SSO11, THYC10, Tse12, WZ10, WZZH11, WXLW13, WV14, YL10b, YL11b, YKL12b, YKL12a, YKL14, ZOS10, ZS14a, ZH10b, Zhu10c]. inertia [Dan14, Nar12]. inertial [MM10a]. inertial-type [MM10a]. inertias [HW12, WLY11]. Inexact [LH11b, AHM14, BBK13, Bog12, CZ14b, Gu13, LWYY14, LZ13b, LWM11, PW10, QLL12, RX14, WYY11, ZW13a, AH11e, AKH13]. infant [WDS11]. infeasible [LS11a, YLD14]. infeasible-IPM [YLD14]. Infected [HLM14, KMAG12, SSST14, VDL12b]. infection [BPP12b, BMH12, CGL12, DISS14a, FW10, GC12, GZ13, KLY12, LSW^{+}14a, McC10, MEL13, SP14b, Sam11, TX14, VDL12b, VDL14, WY11c, WES12]. infections [lLW14]. infectious [LWG13, MMP^{+}12, NM14, SM14b]. infective [DIAQ13]. inference [GJGSNRÁ12]. infiltration [NCANCF12]. Infinite [Bah13a, DW14, AAMAS11, AR14b, And12a, AZ14c, AAH14, BH12a, Boy13, BIT12a, Bry11, CNTG12, sCYW10, CW12, CA14, Che14f, CY12, D’A12, DHS10, Far14a, FY12, GCMP14, Guo14, GSDW14, JZY12, JZW12, KA13b, KM10, KPK14, KS13d, LAF13, LL14e, LLP14b, Meh11, Meh13, Mez13, MN10, Par12, RCS14, RHSC14, Sav11, SS11b, SBS10, She12a, Sul10, VJ11, WLW14, WT12, XZ12, YZK12, YL14d, Yue14, ZOS10, Zen12, ZZW11b, ZHC11]. infinite-horizon [KM10]. Infinitely [Che14d, CT14b, LSLL14, YZD10, AA13, CT11a]. infinites [Lol12]. Infinitesimal [Naj14, BIT12a, VCNV13]. Infinitesimals [Lol12, Ser13, VDFV12]. Infinity [CDDL13, Ser13, ISH12, KL12, LZ14c, Ste13b, ZLJ11, ZDC12, DSZ12]. inflation [SM11, SMS14]. Influence [HA12, HCCT14, NA10a, TLDZ13, AAMADH11, EGMPMRBC12, KI10b, KI12b, Naj14, RNEA14, Sam10, VSS13, VS10, YY12, Yan13c]. influential [PPK14]. influenza [MM13c, SSL10, SSL12, SSL13, TGL14]. Information [CÁGGLP14, AEESMA10, CGAK10, DM10a, DRT11, EM12, KK14b, LDR10, LMT11, MIT12, Nak13, NNPH14, SK14b, SS12a, SM14c, SS10e, TDSA14, WLLY14, YK14a, ZCŠ10]. information-theoretically [LMT11]. informative [SSC11]. ingress [DV14]. inherent [DW13a]. inheritance [FB10]. inhibition [ES12b, ZLC10]. inhomogeneous [LGKM13, MV12, Xia10a, Xu13b, YS10a, ZLY11a, ZL13c, ZYZ14c]. Initial [Jan13d, Abh14, Ach11, AANA10, Ami14, AMA10, BRW13, BR13, CDS13, CJRV12, DCR12, For11, Gee12, GO12, ISG11b, Jat10, JWWY13, Kos11a, LLY11a, LWL12, Liu10d, LW13c, MV14, MVM10, NA10b, OADL14, RK12, RV13b, SK13b, WW10, Waz11b, WRD13, WL13d, XZY14a, YS10a, Zha11a]. initial-boundary [Ami14, BRW13, BR13, Zha11a]. initial-value [XZY14a]. initialization [APA13]. initially [AAMADH11, CMS11, GMKV12, GMV12, KGM14, Sin10a]. initiatives [San11a]. injection [AA12d]. injury [Gra13]. inmultidimensional [Oss10]. Inner [MD14, WKLW12, CGM13, DCK11, GP13, Ker13, LL10c, LS14e, PS11, SMNN14]. inner-product [GP13]. innermost [DZ12]. innovation [HLZ14]. inpainting [BB14c]. inplane [LZ10c]. input [AZ14a, Cai10b, CLTY13, DGS14b, GWW14, JWWY13, KM10, LZL14a, LSSZ14, LLYH10, LHS12, RKA^{+}13, RM12b, RLW14, WLLY14, Yan10a, Yan11b, Yan13a, ZSL11, ZW14b, ZZY^{+}14a]. input-output [DGS14b]. input-to-state [AZ14a, ZZY^{+}14a]. inputs [BGRS11, LZZP10, LT12, Wei14, ZL14d]. insect [PM14b]. insight [NLLC10]. inspection [dALORPGC14, WW11h]. inspired [DM13a, GW11a, GHLF14, LJ11, ZAY^{+}14]. Instabilities [Sha14a]. instability [AKS11, BWaY14, DW13a, Gaz12, GQ13, PCM14, WZ13b, ZLY11a]. instance [QTYS14]. instant [Ban14, DBS12]. instantaneous [AAV10, POR13, SS10g, WL13b]. insulator [Mam11]. Insurance [Zim14]. Integer [Rim13a, Rim13c, Rim13b, AKB13a, AKB13b, Boz13, CGO14, CSMT12, Ema13d, Fen13, KB11b, LWQ10, LWFG13, Mat14b, MF13, ÖA13, Pak11b, Sol14, Ste14a]. integers [AKS14, CI13, MP14d]. Integrability [GPP11, MVS14a, MD10, Wan10b, WYTL14, AC12, HVC12, MH12, MH13b, PJ12, Val11, pZLC10]. Integrable [ELS13, SMM14, TSG10, XZ13b, ZgJhD12, ZK11, BSS12, CLL10, FDHK14, GMX10, HC14, LLY11b, Ma11c, MM14a, SS13d, TF10, TV13, WDL12b, Waz10b, Waz10f, Waz10e, WGW11, Wen12, Xu10b, YHWZ11, YLGW14, YL10c, Yu12a, YL13, YL14d, ZHT13, ZZW14, ZW10b, ZDW11, ZLL13]. Integral [AS14c, CXG10, HLZ12, LYJP13, LYAH13, LLLS11, LZ10d, Ste10g, ABRT14, AB14a, AY11a, ABEK11, ABBR12, AMO^{+}10, AFFV11, AOM^{+}11, AR11, Agr14a, AH11a, Alo13, AG14e, AL10a, Aou11b, AAH14, BM14b, BBJ10, Bec13, BB11c, Bel13, Bie11, BRM11b, BMR13, Boy12, CH14, Çak10, CMM10, CGGMG12, CGGMG13, CT12a, CK10, CL14b, CZ11a, CJ11a, CZ11b, CLW11, CW12, Che14j, CD13, CBS12, CS11c, CEXG12, DAHAK10, Dar11, Dar12, Dar13, DS14b, DKS14, DL12b, DI13b, DS14e, Dos11, Dou10, DJT10, DHM14, EK10, ESH10, EV14, ECG14, EALS10, EGH12, EZ13b, FP13a, Fan14c, FLZ14, Fra10, FKH10b, Fuk14b, Gal10c, GL11a, GLX13, GMMIDM13, GAV14, GT11b, GM14b, GG14e, GLB12, GWML12, GMSH13, Gup13, GRY14, HK11, HKI12, HL14a, HOB12, HSW12, HS11b, HH13b, HHT13, HD14b, IB11]. integral [ISÖ14, IKL11, Jan14a, Jan14e, JSWJ13, JL14, JC13, KJ10, Kar14, Kat11, KSFG14, KK13a, Kle12, La 10, LJ12a, LLP14a, LS10a, LAF13, LL10d, LX11a, LZW13b, LLLS13, LNPS10, LHTN13, LWN14b, LZX10, Maa12, Mal13b, MN12, MMO13, Mat14a, Meh13, MA10b, MS13f, Men11, MG11a, Men12b, MJ13a, MH14c, MHB14, MP14d, MEG10, Mol14, MV11, MS11, NMIA11, NAL12, O’R11b, OR11a, OO14, Ols12, OPD12, Öza14, PN11b, PR12a, PQW10, PKS10, RB14, Rah13, Ram11b, RGL10, RY14c, Ren14a, RPA14, SNMD13, SR14a, SMN11, Set14, She14b, SSO11, SDO10, SKM11, Ste14a, Ste10j, SU10a, Ste11e, Ste12d, Ste12e, SH13, TW14b, TFM13, THYC10, Tse12, TTQ11, WLH10, WZW12, WZG12, WLZ12a, WWG13, WW13c, WV14, WW14a, WRD13, WN11, WN14b, XH13a, XW13, Xia14, XL14, Yan12f, YLLL13]. integral [Yan14a, Yan14b, YG11, YTS12b, ZZ14a, ZX10, ZLS12a, ZG12, ZLWW14, ZWS14, Zha12c, Zho13, ZCD10, ZL13f, dV12a, AS11b]. Integral-type [Ste10g, LS10a, Ren14a, Ste10j, SU10a, Ste11e, Ste12d]. integrality [GLSJCBS10]. integrals [AMA12b, AY10, Alo14, BDD^{+}12a, BEdAdSK14, Bry11, CS11a, CS13a, Cha14a, CA14, Che14e, Che14f, Fuk13, Fuk14a, Gor10, HLPW14, Hon10b, HYT14b, IW14, KX10, KX11, Kel12, KP13, LWW10, LSYS14, LSY14, MÇ12a, Mat14a, Meh11, MX11, Mur11, Mur13, NAA13a, NAA13b, Sch11a, Sha13a, Shi12, Sid10, SS14e, Tom13, TÇ13, Wei12, WZL14b, XX14c, YSI11]. integrands [DHM14]. Integrated [SM13, AKP13, BES10b, DAEY12, GAK12, KDTSN12, LOD12, SGCG14, TCBL11, TGTCCGRG13, XKH10, YH14e]. Integrating [Hua11b, JWW14]. integration [Abh14, AAC11, AAS12, AM14c, CMT11, DSGB12, EZ13a, GDH10, HR13, KX11, KW10, LZ10b, LHTN13, ML11a, PT13, SP11, qSyGH11, Wan14e, WWZ14b]. integration-based [HR13]. integrations [ZYLS12]. integrator [Don14]. integrators [DLSS11, DHM14, GL14a, KWM13, PB14a]. Integro [Shi13a, WZ13d, ZM14b, AS10, AJF13, AASS13, AAS14b, BRS13, Beh10, Bel13, Bho10, BTS12, BRM11a, CC13, CZ11a, CSS11, Dia14, FCJ14, FEGEA14, FFOC14, FH13, GADK14, GÖS10, GÖ14, GZF12, HLZ12, HPB11, ISG11a, JKN11, LDWY14, LH14a, LLL11a, rL11, LJ13, LZ14c, LXZ12, LXLY14, MH13a, Mad14a, MN14a, MPS14b, Men12a, MB13a, MEG10, MF14, Mol14, MAHAS14, RK14a, RCS14, SK13b, Tom12, Tom15, Tur14a, Tur14b, VSM14, VRMT14, Waz10a, WHZ14, XZ12, Yüz14, ZZZ11, ZHY13, ZQ14, ZM10a, Zha13d, ZCC13, ZR14, dS10]. integro-differential [AS10, AJF13, AASS13, AAS14b, BRS13, Bel13, Bho10, BTS12, BRM11a, CC13, CZ11a, CSS11, Dia14, FCJ14, FEGEA14, FFOC14, FH13, GADK14, GÖ14, GZF12, HLZ12, HPB11, ISG11a, JKN11, LDWY14, LH14a, LLL11a, rL11, LJ13, LZ14c, LXZ12, LXLY14, MH13a, Mad14a, MN14a, MPS14b, Men12a, MB13a, MEG10, MF14, Mol14, MAHAS14, RK14a, RCS14, SK13b, Tom12, Tom15, Tur14a, Tur14b, VSM14, VRMT14, WHZ14, XZ12, Yüz14, ZZZ11, ZHY13, Zha13d, ZCC13, ZR14]. integro-differential-difference [GÖS10]. integrodifferential [GLB12, KD13, Kum11, NV12]. intelligence [Kam14a]. Intelligent [BYM14, BAE13, NNPH14, VTM14]. inter [BC14a, LGKM13, ZZW^{+}12]. inter-cluster [ZZW^{+}12]. inter-population [BC14a]. interacting [LzJZ12]. Interaction [DL12a, KM14a, SJ13, Bai14, BC14a, Gad11, IKH11, JT13, KHA12, KB14, LMS10, Mis11, NTPVNHLX14, SSA14a, Wan13b, WG12, WX13d, WW14e, Zha12b]. Interactions [MWY13, Zha10d, BC14b, DW13a, IHF13, LKY14, LGS^{+}11, Lot14, MGWG11, Pal13, RRLA14, SP14a, SYB13, Woo20]. Interactive [Ema13d, MEBYG11]. intercept [IG11]. interclaim [SMWR10]. interconnected [KLKA14, LL12d, Shk14]. interconnectedness [PPK14]. interconnection [Che11f, CDIS10, IHM14]. interest [CW11c, FGRLSV14, STV12, WZ11g, ZW13c]. interface [BM13b, CJM14, DS12a, DJ14a, FHWZ12, MJ13b, Sin11a]. interfaces [Mam11, MR13b, NZ14, SSS13b, WS14]. interfacial [All11]. interference [BSKL13, DMGK13, GC11c, Liu13a, MCWC13, ZjHO13, ZMC10]. interindustry [GS12d]. interior [AG14a, Alz14, Cho11b, HZW13, LS11a, NY12, SRP13, TDZF14, WYT13, WZ14b, MAN10a]. Interior-Boundary [MAN10a]. interior-point [Cho11b, LS11a, WYT13, WZ14b]. interlacing [ET11]. Intermediate [TMM13, TMM14, Gor10, PK12]. Intermittent [GS13b, BWHQ14, WS13]. Internal [AAA13b, GW13, LZWZ14, MWY13, MGSMRÁ14, SHL10, VV11a]. international [DW13a]. Internet [HT10, ZX12]. interpersonal [BC11b]. interplay [MVS14b]. interpolants [BCM12, BW10, JZ11, LMZ13, XFW^{+}12]. interpolating [BCR10, CMGR13, DGS14c, KFM11a, RCH12, WSCH14]. Interpolation [Beh10, AMA14a, AM14a, ABCD11, AB12a, BJS14, BCDI14, BV12a, CSQ11, CZM11, CFZ14a, FBGH12, GC14, GMGV^{+}11, Ger12, Guo13b, HW11, JW12, Kim11, KC13a, LY10a, LLSY11, LL10d, LD13c, MMX10, MV11, Mur11, Mur13, NP10b, PB14b, SHA12a, SS10d, SJPS14, TZZX11, VC14b, WQ11b, WZ13d, WWZ14b, XYY12, XwK11, ZLX12, ZK10a]. interpolations [BCC11a]. Interpolatory [PDN11, Zha10f, AR12b, ADT13, LZ12b, LLS^{+}13a, LQ13, SSA14b, SIR14, UB13]. Interpretation [CCG12b, ISH12]. interrelationship [KI10a]. interrelationships [KI11a, KI10b, Koj13]. interruption [JS12, TLW11]. intersecting [HW13a]. intersection [CTH14, CDMT12, DA14, ZWZ13]. interspecific [DMGK13]. intertwined [CD10b]. Interval [LZW13a, MH11b, SSTM14, ZGW14b, AKC12, Ahn13, BMR10, BCM12, BSR10, BI14, BNP11, Che13d, CM12d, DA11a, Esm12, GCG14, GGXH12, HDT11, Hla13, JS14a, JSMA11, JPK12, Jia13b, JHZC13, JDD14, KB13, Kol14a, KLPC12, KPP^{+}13, KPP^{+}14, LLP14a, LP14, LH10a, Len14a, Li10a, LCL^{+}13a, LR10, MP14b, MO14, Ols12, PKP^{+}12a, PKP^{+}12b, Pet14, SC13, TPT12, Tsa11, VMS10, WLZ12a, XM10, XZY14b, ZL13a]. interval-valued [JSMA11, KB13]. intervals [BTS12, CC10b, OO11, Ste14d, WLW14]. intracellular [NRWF12, VDL12b]. intraday [Ben14b]. intraventricular [WDS10]. intriguing [SCLC10]. intrinsic [GL13b]. introducing [ACQOR11]. intrusion [MATA14]. intuitionistic [DGS14c, DS14d, LLZ14b]. Intuitive [Yun10a]. invalidates [PJ10]. invariance [BP10b, BP13, WSD12]. Invariant [RS10, ABP13, BS11a, CLAT11, CTLL11, Cur14, Den11c, FV14, GS14a, GB13, KG14, KP13, LR13a, LL11c, MK13a, NAL12, Naz10, PQW10, WT11d, ZLYL10]. Invariants [Kov10, ÇM13b, DMR10, TRA13, Vir13, ZZ14d]. invasion [LKY14]. invasive [BMD13]. Inventory [AAKB13, BM14a, BJB11, BJSS14, CBTGWW14, CLS12, DS14d, GLSJCBS10, LOD12, Lin13b, LHT14b, Mal14, PSC12, PSJGLS14, QK13, Sar12b, SGC12, SGCG14, SYB13, TCBL11, THCW11]. inverse [AGT11, And11, AKB13a, AKB13b, BT14, BD14c, BMMS13, BEGM12, BZT11, CL11b, CW11b, CY11a, CM11, CG14e, CI11, CKD12, CDDL13, CD14, DHI13, DNSM11, DAV12a, DM13b, DESW10, DD13, EdCVB^{+}14, FL13, FIK14, Fuy11, GH10d, GYS13, Gus12, HLT13, HYC^{+}10, HH12, HSC14, HG13b, JC14, Ji14a, JKP14b, KPP11, Kyr13, Kyr14, LL10c, Li11a, LW11c, LZFR10, LZ10d, LLCX10, LXY10, LY10b, LCH11, LLW11, Liu11a, LQY11, LW13d, LW13c, LTU13, LZB14, MS12a, MT14, MP12b, MD11a, MD12c, MD12d, MD13b, MD13c, NMM14, NB11, NP10b, Neu13, OD11b, PH10b, RDD14, Res11, San14, DAV14b, Sev13, SC11, SWW14, SSCT14, SSU13, SY14d, SG10, TSP10b, TSN13, TS13, USAF14, VFGP13, WaZZW11, WCL12, WCN11, Wu12c, XLL11, XQ12, XQ13, YGL11, YT13a, YH13, Yu11b, ZY14b, ZJ14a, ZW11c, ZZX14, ZCZP14]. inverse-free [WCN11]. inverse-of- [Ji14a]. inverse-positive [AGT11]. inverses [BT12, BFB12, CIP10, CISM11, CIN14, GS13c, GD10b, LJCI12, LY12c, LHCI12, LJCI13, MMS11, MD14, PTS11, SCH11b, SGZC13, Son12, SCIMM11, SPKS12, SP13b, TS10b, XLS13b, ZB12]. inversion [Boy12, Che13e, GTM13, KRMM14, LDS14, PSSS14, Res13, TSN13, ZLL11a]. inversions [CL10b]. invertase [IN10]. inverted [PPMCG14, ZHX11]. Invertibility [TCI12]. invertible [dRS13]. invest [BO13]. Investigation [BC14a, CEM11, Soo11a, Asl11a, GO13, HdCOL14, Mat10, MHA10]. investigations [BCGK14]. investment [BM12b, BM12c, CC12a, DG10a, HZ14d, YJ10, YX12b]. investments [AAV10]. invexity [Ant11b, AZ14c, MN11, PN11a, Ver12]. inviscid [EVDG12]. invisible [WQX^{+}10]. involution [MD11a]. involutive [Tos14]. involving [AG14c, AJF13, Aou12, Ara14, BP13, BBC14, BS13, BCK14, BFM13, yCfW12, CZ14a, CA14, CM11, Cho11c, CK13b, Cvi10, Den13, DX14a, EMA13b, Et13, FMF13, GMMIDM13, GQ12, HXQ14, HN10b, HN14b, Jan13d, Jan14e, Kan11, KKS11, KYH10, KM11c, LJCI13, Ma11b, Meh11, MG12b, PTY13, Par11, PPK11, Pré11b, QCK10b, Sad11, SR14a, Soo11a, SKM11, SC14b, WW13j, XX14c, YL11b, YYY10, ZLW12a, ZW13d, ZLWW14]. iodate [LY14c]. iodate-arsenous-acid [LY14c]. ion [Awa12, SBM12]. ionic [YA14]. IP [WYL11]. IPM [YLD14]. iron [QS14]. iron-phosphate [QS14]. IRR [Bas13]. irreducible [Man12a]. irregular [DW11, HVD^{+}14, TVZ10]. irregularity [AAMB14, AÇ13b, BS14a]. irreversible [AAV10]. irrotational [ABG14]. IS-LM [Kal12]. isentropic [WFL14b, YWW13]. Ishikawa [CMW10]. isochronous [WZ12]. Isoelastic [CMS13]. isogeny [Moo12]. isolated [BDE^{+}14]. isolation [SG11]. isometric [LDS14]. Isometries [JK13]. isospectral [ELS13, LTG^{+}12, Sun10]. isothermal [VPS12]. Isotone [GJNN14, SK13a]. isotropic [LWZ13b, Sin10a, YLS13, ZZHZ14]. Issue [CCG12b, DSZ12, KS12b, Udw10]. issues [LLSY11]. Itô [MH14c, LZLW11]. item [AKP13, KGC11, SGC12]. items [BJSS14, GJ14, JGM13, LHT14a, LOD12, PSC12, San11b, SCN14]. iterate [GIR10, GS12a]. Iterated [HLPW14, KR13c, LWN14b, SS11g, FMF13, LW13d, Man12b, Ste12c, Ste12g, Ste12h, Ste12r, Ste12v, Ste13a, LM13a, YLLL13, ZANA12]. iterates [Bog11]. iteration [BEGM12, BL11, CL11c, CLCZ12, Che10b, CFZ14b, DTP14, GS10b, Gu10, GW12, GML13, Inc12, KM14b, Kim14a, LPAE12, LWZD10a, MC13, Ola11, RY10, SSK13, SR13, ST14, SW14c, USAF14, wX13b, Zho12b, Zhu12b]. iterations [HYD10, LPB10, MM11a, MG11a, NS11, SKC10, SBDB13]. Iterative [CS13e, Din12, GLP11, GJNN14, fHcXH11, KLKA14, LL11b, MLH11, QMZ14, RM11, WWG13, dWcFwWjZ13, WXLW13, WZDW10, WLD11, WFDL11, XL14, Ard11, Ard13, Bab12, BBSC13, Bum12, Buo10, CMM10, CAY11, sCCL11, CPZ10, CW11b, CD13, CQ14, CS10c, Chu11, CLND12, CL13b, CNKS14, CMW10, CST14, DK14a, Din11, DHS10, EZ13b, FHWZ12, GK10b, GHZ^{+}14, GT11c, GL10c, GMV11, HZ10, Hu13b, HLX13, HMH13, Jan14d, JM11, JWW14, Jun10b, Jun11a, gKW12, KAS12b, KKH14, KA11, Kim10, Koç11, KC13b, KPW11, KKS12, LL10c, Li10b, LZ10a, LSJ11, LZLW11, Liu11d, LL11e, LJKU11, Lu12d, MXH10, Mag14a, MVRVSCMV14, MP14c, NCS13, NK12, ONBN10, OP12a, PL10, PZ11, RW14a, Sad11, SWW14, SM10, SSU13, SW10b, TA11, Thu10, TS13, Wan10a, Wan10f, WHC10, WMH12, Wan13c, WJ11, WWY10, WYW14, WLH11, XLY10, pYmK13, bOYqGL^{+}14, YNL10, Yao10]. iterative [YDTW11, Yua11, Yun11a, Zha11b, ZJD12, ZLWL13, ZS14e]. iteratively [LWS14]. Ito [CSZ12, S10, SGY^{+}14, ZDW10]. Ito-type [CSZ12]. IV [HL11e]. IVP’s [RP14a, CD11b, FG14b, KP14a]. Iyengar [HN10b, LN10]. Iyengar-type [HN10b, LN10].

J [Ban14, PTY13]. j.amc.2009.07.009 [Asl10d]. j.amc.2014.03.101 [WHS14a]. Jackknife [Lin11d]. Jackson [Len14b, Łen11]. Jacobi [LWN13, Akt14b, BES10b, Bhr13, Bhr14, BHP13, CK12c, DP14, DAEY12, DBBEE13, EZ14b, Gus12, Hon10a, MPC14, SSN^{+}12, Wen10b, Woź13, Wu12c, ZX11a]. Jacobian [Neu12a]. Jacobsthal [BT12]. James [KSM11]. Jarratt [Mag14a, SS10c]. Jaulent [BK10c, MGYQ12, RK14a, XTAJ12, yZqLwW12]. Jeeves [RCSH10]. Jeffrey [SHAO14]. Jenkins [ZJD12]. Jensen [BPP12a, JKP14a, RDA14]. jet [EASK10, IKH11, Naz10, VD14a]. jet-vortex [IKH11]. Jimbo [TMXG11, TTM12, ZgJlK12]. job [EG11, HL12e, JCW12, KBYL14, mL12b, LPX12, Rud13, WL14a]. job-dependent [JCW12]. jobs [CWCW11, GL10a, GM13a, LL12a, Sun14, YK11a, ZL13a, ZHC^{+}14]. Johnson [ARH^{+}12, ASRI12]. joint [CN14b, EGMPMRBC12, GS12d, IS14, LOD12, NMM14]. Jordan [JC14, SP13b, Zhu10c]. Jordan-type [Zhu10c]. Joseph [FHWZ12]. Jost [AB12c]. judgments [GRKP10, Lin11d]. Julia [AK10c, ARC14, SL13b, Yan10f]. jump [DWLL10, GWW14, HL10, HYX14, KXW14, LDWY14, LSSZ14, LZL12, MS13a, RZHH13, RZHH14, Siu10, VSS13, ZLMG12, ZLS12b, ZR13, ZCZ^{+}13a]. jump-augmented [Siu10]. jump-diffusion [DWLL10, LDWY14]. jumping [BMR10, CZD^{+}14, DZZ14, HL13a, RSPK13, TLZZ12, ZaFT11]. jumps [CY12, HG13a, JS13b, LZLW11, LC14e, LB14a, MM14b, SY14a, WCZ13, WZW14, Yue14, ZW13c, ZW14a, ZL14f]. junction [BLK^{+}13]. junctions [GVT^{+}13]. Jungck [DK14a, KKH14, Ola11, OP12a]. Jungck-type [KKH14, OP12a].

K.I [WZ13c]. Kac [YL13]. Kaczmarz [PP14c]. Kadomstev [GB13]. Kadomtsev [XZT10, XLZ10, AM12d, BR10b, Elb12, KA10a, Kud10c, LWL10, Liu10a, MA13c, MTZ10, Sun10, WD14a, Waz10b, Wei11b, Wu10c, XC10b, XC10a, Zha10a, Zha10d]. Kadomtsov [SYZ10]. Kalman [AKC12, Gro13, KCGH14, UMMT14, WFDL11]. Kamenetskii [Har10]. Kamenev [AZL13, Has11]. Kamenev-type [AZL13, Has11]. Kansa [EH14b]. Kantorovich [AAB^{+}14, Dar11, EALS10, EGH12, GMR13, Içö12, KPT11, ÖD11a]. Kantorovich-type [AAB^{+}14]. Kármán [EGJB13, Kan12b]. Karmarkar [NBP10]. Kassem [EM11]. Katayama [Aza14]. Kauffman [IB10c]. Kaup [BZ10a, DGDA10, GZ10d, SMM14, GZZ10, LLWW10, Pin10, S10, Wen10a, XZ13b]. Kawahara [CP14b, ÖA10]. KCL [AP10]. KD [SF11a]. KdV [Aba12, eMA13a, Asl10d, Gün10, WDLM10b, SHC11, AK12, ÁD10, AMB10, BW14, BC10b, BZR10, BI10, BPM10, CSZ12, FL10, GZ10d, GW11b, HXY12, HA10, Hay10, Hon10a, HLZ13, HYSZ14, IU10, IB10a, IRS14, JKB10, JK10, LZ11a, LCH11, LCLW10, QMM13, RCYL10, Sal10a, SMH14, SB10, hTlCqL10, TW10a, TW11b, WDL10, Wan10b, Wan10c, WqXlC11, jWtW13, Waz10g, WGW11, WP14, Xia10a, Xia10b, XD11a, YlTA10, YH14c, YWCH11, YL14b, YZZ10, YM10b, Yu11a, YL14d, bZjZS11, ZgJhD12, ZC14b, bZpZJ10, ZZ11c]. KdV-Burgers [IRS14, YH14c]. KdV-like [SMH14]. KdV-type [HYSZ14, WDL10, YWCH11]. KdV6 [Sie10, GX12, GPP11, SV10a, ZlDxX11]. KdV7 [Sal10a]. Keeler [MAK13]. Kepler [dOL13]. Kernel [LYGZ13, YY14a, AST12, ASST11, AASS13, CJ11a, DS13a, DS13b, DKMM11, Fuy14b, GXSH14, Gen11, GQ14, GKS12, GG14e, Hua11b, ISG11a, JC13, LNC12, LS11a, Men12a, MD14, NMIA11, OÖ13, Sht11, TS12b, WDT^{+}13, WL10b, YP14, YWY^{+}13, ZHY13]. kernel-based [DS13b, GXSH14]. Kernel-improved [LYGZ13]. kernels [Bec13, CC10b, Che14f, Dos11, SPP12, eT12a, UY11, XH13a, XW13].

Kerr [Lat14, yZhLjMzC10]. key

[CXD14, FB10, HZZ14]. keyword [AdC14a]. Khairnar [WZW12]. Khasminskii [Mao11b]. Khasminskii-type [Mao11b]. Khorram [MK08, Shi10]. kind [BM14b, Bie11, Bou10, BMR13, Boy12, CK10, CJ11b, FPP11a, Fra10, GÖS11, GYG11, HKI12, LWW10, LL10d, LR13b, LS11c, MMO13, MG11a, Men12b, MV11, PS12a, PAA14, SR14a, SHMCM12, TW14b, VRZ11, WF12, WWG13, WW14a, WZL14b, XW13, Xia14, YZ12a, YLLL13, ZGZG10, Zho13, ZX14]. kinds [HC14, HYW14, LLLS13, dLL13, ZH10a]. kinematical [AP10]. kinematics [CR11a, DESW10]. kinetic [AGHdD13, BMHM14, Ber10, FOZX14, VD14a, ZM14c]. Kinetics [Sun14, FW12, MYT10]. King [Ma11a, HKH10, Nad07]. Kink [ZLZY12, HCD14, KRS10, Li10d, SC10b, SHC11, WDLM10b, Waz10c, XW10]. kink-shaped [KRS10, Li10d]. KINMODEL [AGHdD13]. Kirchhoff [MTT14, Aou12, DH14, FKH14, MCdSS11, ZTZ14]. Kirchhoff-type [DH14]. Kirsch [LPB10]. KL [Duk12]. KL-divergence [Duk12]. Klein [ABP13, BT10b, CYY14, Don14, IISA10, IB10b, Jan14a, KS10b, Lia14, LGN10, Mer14, NC13a, RV13b, Sha10a, SXL12, TB14, WZJ13, YZ10]. KN [WX13a]. Knapsack [BD12c, BPT12, CMP13, GHLF14, HW14b, RFK12, Sit14, ZHH^{+}13]. knots [HN10b, HN14b, WW13j]. known [DKMM11, MK14a, TLZZ12]. Kober [DS14b, Mat14a]. Kocak [Koç11, Koç13]. Kocak’s [Koç14]. Koch [LCD12]. Koch-type [LCD12]. Kogbetliantz [MH10a]. Kolmogorov [MDP12, ACG14, Bal14, CST12, FLW11, Hu13a, JZY12, LS13c, ZHC11]. Kolmogorov-type [ZHC11]. Konhauser [PAA14]. Konopelchenko [SF11a, XGG^{+}11, Zha10b]. Konoplechenko [pXqLlZ10]. Korean [GSC14]. Korovkin [BSM14, MKEG12]. Korteweg [Asl11a, SB13, eMA13a, Dem14, GGYW11, GMST10, GT11c, Kud10b, LJW10, ÖK11, QW10b, Sie10, THAB11, VDK13, WDLM10a, ZGY^{+}10]. Kotera [SGY^{+}14, XLZ10, GS10b, MG12a, QMM13, Waz10e]. KP [Elb12, Ibr14, Asl12b, DLFZ10, fD12b, IPB10, LY14d, MAA11, MZ12, THH10, WTFC13, YM10c, Yu12b, Zha13a]. KP-BBM [YM10c]. KP-MEW [WTFC13]. KPP [Yan13b]. Krall [ÁNA12]. Krall-type [ÁNA12]. Krasnoselski [MM11a]. Krasnoselskii [BCK14, MAD14b]. Krasovskii [KPLC11, KPP^{+}13, LZL14a, PKPL11]. Kreiss [FdBCC14]. Kriging [EL14, ZD11a]. krill [LTHG14]. Kronecker [AZ19, AZ14b]. Krugman [TES14]. Krull [MM11b]. Krylov [SDK10]. Kudryashov [HBDC14, Rya10, RSK11]. Kullback [Lin11d, SEKR14]. Kummer [GH11]. Kunita [ZHY11]. Kuperberg [BS14b]. Kupershmidt [BZ10a, DGDA10, S10, Yu11a]. Kuramoto [FL10, HBTU10, Kud13b, SJ12]. kurtosis [GVMNS13, RY12]. Kutta [FG11, FG13b, BDD12b, DEP12, FYM13, GTPM14, GSDB11, JSSB14, Kal13, LLY14b, LJ13, MDD12, Mon12, SMIS14, Sho14, TS14, Tsi14, ZQ14]. Kuznetsov [LZLB14, SY10c, Asl10b, Awa12, SC10b, gZrLX10].

labor [BM12b, BM12c]. laceability [LTT^{+}13]. lack [BIT12a]. lactic [ZLC10]. lacunary [Bra12, Et13, Kek11, Sav11]. Ladik [Asl10a]. Lag [WS12, Ach11, JJP^{+}12, Mon12, ZZ13b]. lagging [MSR14, SZW14]. Lagrange [MKSN13, CZM11, jLLW12, MKK14, PZ14b, SR14c, WW13c]. Lagrange-power [jLLW12]. Lagrangian [CHX13, Che14c, CGK12, CGGZ11, CSF12, HZ11, HSS10, Hou13, HLT14, JYL14, MKSC13, PW10, YL14b]. Lagrangians [FAM11]. Laguerre [NC14, AG14c, BPPS14, BPR13, BR14c, GGÖS11, GS14f, HMPVQ14, Jun11b, Kha14a, KAG12, MMMMB13, NC13b, Öza14, PAA14, SS13a, XYY12, Yüz14]. Laguerre-Hahn [BR14c]. Laguerre-Sobolev [MMMMB13]. Lah [JKP14a]. lake [Mis11]. Lamé [DU13, Dos11, Xia10b]. laminar [KK11b, RP14b, SZZ^{+}11, TSS^{+}14]. laminated [CCS13, GYBR10, KAF13, THNTNXR11]. Lanczos [CP12b, NY12]. Landau [AAA12, Bot12, CC14b, Che14b, CC12c, GARK12, JT13, KR13b, Li10c, LLXZ12, LW12a, LS13b, LTJ^{+}11, ML10a, NN11b, She10, SZ14a, ZL13c]. Lane [Ben14a, BF12, GS14f, IJ11, MC14b, PKBD12, WRD13]. Langreo [MSGN^{+}14]. language [BM12d]. Laplace [CRG11, CTLL11, GTM13, Jum12, KLW14b, KCLC13, LF11c, NMIA11, SSCA13, TM13, TTQ10, Waz10a, Xia14]. Laplacian [AAA12, Ant13, BD11, BC13a, BEGM12, BR14b, CS12a, CLL14, CXG10, CEXG12, DWC11, DX14a, FL13, GLZ10a, HLMW13, HR12, JŞ11, Kon14, LW14a, LL12g, MDÇ13, MAO10, ZT12, ZLWW14, ZF13b]. Laplacians [Yos13]. Large [CCX11, CTLD12, MBR10, Ngu14, RD14, BBF12a, BLR10, CR11b, DR10a, dCDNRH12, FOL12, Gu13, GG13b, HCL^{+}13, JLLD13, LLLZ11, LH11a, LH11b, LZ13a, Li14a, MP11a, NY12, PW10, PZ11, RM12a, WN12b, YLWH14, ZCŠ10, ZHW11, PHPD11]. large-amplitude [LLLZ11]. large-scale [BBF12a, CR11b, DR10a, FOL12, LH11a, Li14a, YLWH14, ZCŠ10]. large-time [MP11a]. largest [CGJ12, MRCD12]. laser [Li14b]. Lasota [ZZW11b]. late [LL12a]. latencies [WL14b]. lateness [JCW12, Rud12, XYL10]. latent [XGZW13, Xu12b]. lateral [Cha13c]. Lattice [Che10c, ACG14, Asl10a, DWLL10, FZP12, GLS13, HMSW14, HG13b, KLW14a, KP12, LLY11b, LWB14, LW10b, NKM^{+}10, OCN11, SW14d, TF10, Wen12, WG12, Wu12d, Xin10, YHH11, Yu12a, ZLL13]. lattice-subspaces [KP12]. lattices [AAGR13, JZY12, OCN11, XMWD14]. Lauricella [LLSW12]. Lavrentiev [GPK13, VG14a]. law [Ahn12, AMB10, BA11, BZR10, BCWM10, BK10c, BT10b, BPM10, BR10b, CYY14, CIP10, DD13, FMS14, FKH10a, Hay10, Hay11, IB10b, Lat14, LY10b, LZB14, MP14a, MD11a, MD11b, MD12d, MD13b, NAHH11, PC13a, PP11, SSS13a, SB11, Tom13, TB14, XQ13, Yu11b, yZhLjMzC10]. laws [AK12, AMW12, AP10, AZAK10, BK10d, BK10c, BK10b, CL14a, CG12a, CZC11, CIN14, Dub10, GB12, Guo14, Ibr14, JA14, LP12a, LJCI12, LHCI12, MPPM12, MK10b, PTY12, PTY13, RYÖ11, TF10, WG11, WX13a, WC13, Wen12, WHG13, XLZ10, YL11c, qYqLtW10, YL14d, ZB13, Zha11a, Zha13a, ZYC14]. Lax [MKeM13, Sal10a, Wan10b, WTL^{+}11, WDL12b]. Layer [Kum14, AHMS14, AV14, AHKS12, BC13b, Che10f, DNI^{+}10, FZZ12, FHAL14, FFOC14, Fit12, GO12, GMKV12, GMV12, HAMA14, KAB11, KGM14, LNdCJV14, PSC12, Pan15, PK18, PP13b, PP11, SCN14, Sha12b, SS10g, Vis14, Zha11a, ZGY^{+}10, ZZHZ14]. Layer-adapted [Kum14]. layers [AM14c, Bas12b, DW11, FZY10, Haj11, LCD12, Pal10, SRP13]. layout [Zhe14]. LCM [TL10, TLL13]. LDG [Bac14, LZQ11, ZD10]. lead [LHT14b, MSS14b, SGCG14, SYB13, TNB11]. lead-time [TNB11]. leaders [LSL^{+}14]. leadership [BBS10]. leading [Liu11a]. leakage [GWZ14, LPJB12, LPL^{+}13, LY14b, LFR14, PKP^{+}12a, ZS13b]. leaks [EASK10]. leaky [HLW13]. leaping [Aha14, LPX12]. learnable [CEP11]. learning [BTV12, CWCW11, JCW12, KLKA14, Ko13, LZW12b, MLH11, MVRVSCMV14, QTYS14, WWZ10a, XG10, XYL10, XZZX14, Yan10e, YK11a, YZ14b, ZZYX12]. Least [BDR12b, JJL14, JWW14, LWZD10a, LGZ14a, LIC12, Wu12a, Yua10, ZLGZ11, AM14e, AE13, CBO12, CF13, Cor11a, DR10b, FTAB14, GSSS13, Guo10, Guo11, Hua11b, Juk13, Kyr12, Lee12, LL11b, Mea13, MMS– D13, PX13, PZ11, QW14, RMN14, RA10c, RCH12, RPW14, Sal11b, SZS13, TW13, qWH11, WSCH14, XLY10, Yam14, YWY^{+}13, YGZW14, ZJD12, ZL14c, ZJ14a, ZZX14, Zit11]. least-rank [TW13]. Least-squares [Yua10, CBO12, CF13, FTAB14, Guo10, Guo11, Lee12, LL11b, MMS– D13, PZ11, RCH12, TW13, WSCH14, ZL14c, ZJ14a, Zit11]. Lebesgue [CC14b, Che14b]. Lee [YZZP14a, YZZP14b]. Leffler [TS12b]. Left [YH13, AJR12, ZZW13]. Left-Definite [ZZW13]. left-half [AJR12]. Legendre [Boy10, CSL14, FEGEA14, HHM14, KSY12, MNP13, PKBD12, RAIM14, WZ11h, ZYSQ11, Zhu11c]. Legendre-based [WZ11h]. Leibler [Lin11d, SEKR14]. Leibniz [BDGP14a]. leishmaniasis [EMH10]. Leitmann [AT10b, DLP10, FNV10, Gal10c, JMHZ10, MT10, Wag10]. Leja [DMS12]. LEM [CEP11]. lemma [AFFV11, Hus10b]. lemniscate [AJR12, Neu12a]. length [LLHY11, MHTH12]. lengths [DZ13, GCG14, Yan10d]. Lengyel [Lis14, WZ13b]. Lengyel-Epstein [Lis14]. Leon [WM10, qYqLtW10]. Leray [Wil12]. Leray-deconvolution [Wil12]. Lerch [Cvi12]. Leslie [LL13b, MM13d, TW11a, YL14a]. less [DRGR10, Ge14]. Lethargic [ÁME^{+}12]. Letter [Pol11]. level [AKP13, Car12, DHQ14, Ema13d, MS12b, MSS14b, SH12a, SSN^{+}12, WH13, WLYC13, WP14, XKH10, Yan10b, ZL11c, ACQOR11]. Level-3 [ACQOR11]. levels [DFGL14, FGFRV11, LHT14a, OYCCB13, PTW11].

Levenberg [qDG10, FZ13, HM10b, KR11, LWH14, LZ10f, XM11, Yan13e]. Levin [ZYZ14b]. levitated [Cha13a]. Lévy [BLZZ14, CDLZ11, Gat14, HR11, Hu12b, LCL14, LB14a, WW13h]. Li [Ibr14, Pak11b, SBM12]. Li-Hua [Ibr14]. Li-ion [SBM12]. liability [LL12b]. LIBOR [FGRLSV14]. library [DFTQOB14]. Lidstone [GX11]. Lie [eMA13a, eMBHA13, AH13a, CL14a, Das13, DG14b, GB13, HYSZ14, KA10a, MM14a, MM10c, OGYE13, Sht11, Tra12, YH14c, ZDW11, ZYC14]. Lie-group [eMBHA13]. Liénard [XH11]. life [DE11, GC11d, SEKR14]. lifetime [WHL14]. lifetimes [DF14b]. lifting [ARH^{+}12]. ligand [LPP^{+}12]. light [CC12a, dCDNRH12]. Like [ZM14e, AH12a, Bai13, CBA^{+}14, DW14, Dan10, DGMR12, DA14, DRS11b, Elb12, ER13, EHRV12, Fer12a, Fer14, FNVLPD^{+}13, GCMP14, Hou13, IS10a, IS10b, IHF13, IA14, JLM^{+}12, KAS12b, Koh12, KPW11, LTT^{+}13, Lu12b, MA14b, MB13a, NS13a, PG11, PC14b, RWB10, RCYL10, SSO11, SDK10, SDN14, SMH14, TKS11, Van11b, WSD12, Woo20, WW13j, Xia12, Yan11b, ZJD12, ZY13b, ZGY^{+}10, fuy13, GZ10d, ABGGS13, ZGZG10]. likelihood [LW11c, MC12b]. limbs [AAP10]. Limit [Gin12b, HZ14c, LL11c, LR14b, Sta12a, WHL11, WLL14, XH14, DMGK13, DLC10, EK11, Gin12a, LZM^{+}12, LW14c, MOZ11, Pan11, SA10a, pSbL14, WHF14, WZSW14, WZ12, XH10, ZLJ11, ZL13b, ZZ10d, dOL13]. limitation [JZ10]. limitations [ASSPA10]. limited [Ber14, Che14g, Dub13, GC10c, JKP14b, Li12c, LWZ13b, VL12, ZY14a]. limited-memory [VL12]. limited-permeable [LWZ13b]. limited-view [JKP14b]. limiter [JA14]. limiting [EK11]. Limits [CM13c, BRAB11, BR12, LLHY11, SL14c, YL13]. Linda [AWHW11]. Lindley [MLG13, SSS13a]. line [BPR13, BR14c, CS10a, CD10a, CHT14, CZ10, DMP13, FAM11, GTM13, HBDC14, Jia14, LWZ13a, LWM11, NW10, PZ14b, PKS10, QLL12, Rah13, RLC11, SWX10a, Ste12u, SS10g, WZ11f, WZ14b, WLL14, YA14, ZZYX12, ZC12, ZLZ14d, ZF13b]. Linear [AAGR13, LMT11, SMN11, WHW13b, eMBHA13, Ach10a, AG14b, AC10a, ARE14, Ahn13, AZ14b, AAS12, AC14, And12b, And13b, Aou11b, AJPZ14, AZ19, Bah13a, BM14b, BA11, BM12a, BK11a, BDG13, BTS12, Bie11, BN10, BNP11, BAJK10, BRM11b, BR13, BA14b, BS12c, BPX14, CMX14, CMGR13, CCJ14a, CKN13, CJ14, CWS10, CJ11b, CM14, Che14d, CW14a, CLLS14, Chi14, Cho11b, Cho11a, CP10, CGGZ11, CJRV12, DS14b, DGS14a, Das12, DNSM11, DGSB12, DS12a, dCDNVS11, dCDNRH12, DS12b, DR10b, DLSS11, DF11, DSZ14, EZ14a, FEGEA14, FOL12, Fuy14a, GL12, Gao14, GP10, GQ14, GKV12, GÖS10, GGÖS11, GW12, Gür10, Gür12, GH10e, HVD^{+}14, HN14a, HS13, HZ12b, HM10d, Hsi10, Hu10, HHX11, IISA10, ISG11a, IG11, JK14b, JHX14, JB10, JC13, JPM12, KEGH13, KZ14b, KM14b, KSFG14, KS13b]. linear [KM11c, Kol14a, KM14c, KL10, Kov10, KSS12, KH14b, KPP^{+}13, LOKB12, LLP14a, LR13a, LZLF10, LL13a, LY14a, LTHG14, LZL14a, LW14d, LL11b, LZWZ14, LCL^{+}13a, LLCX10, Liu10c, Liu11c, LHL11, LHS12, Liu12a, LW13d, LWZD14, LLZ14c, MZTT12, MC14a, MAN10a, Man12a, Mao11a, MPC14, MP14c, Mat14b, MA13d, MNOB10, Mey14, MWAM14, MSC^{+}13, MG11b, MS12b, MG14c, MSS14a, MEG10, MF14, MKK14, MS11, NHCLP10, Nak13, NCÁHCLP13, NNG10, NRL12, NL12b, NN19, Oni11, Örk13, ÖKU12, PD14, PKPL11, PT11a, PW10, PJFL14, PAC13, POR13, PPV14, PRV13, PQW10, PR12b, PS14b, QLCF10, RY10, RSK14, RK14b, RS11b, RTPH14, SHT14, SNMD13, SAO12, SGDK11, SDC10, SN11, Sha10b, SHWT14b, SL14b, SL10b, SW11, Ste12n, Ste13b, SD13b, TPT12, TCI12, Tos14, TT13, TN14, VVK12, Wan10a, WW11e, WLZ11, WSD13, WY13a]. linear [WXLW13, WWY10, WYW14, WFZ13, WN14b, WT11d, WCW^{+}13, Wu14a, Wu14b, XX14a, XX14b, XBC^{+}14, XM11, Xu12a, wX13b, YAW10, YW11b, YCXL12, YKL12a, YKL14, Yin10, YDZ11, YLZ13, YZC14, YY10, Yus10, YS13b, mZT11, ZLGZ11, ZZYX12, ZT12, ZM12, ZZ14c, ZZGL14a, ZHXM14, ZZW14, ZZGL14b, ZxHW14, ZZ15, ZZ13b, Zha13e, ZF13a, ZMG13, ZD11b, dRS11a, fuy13]. linear-quadratic [KH14b]. Linear-time [LMT11]. linearisation [AP14, BGP^{+}13]. linearised [Yam14]. linearizability [AC12]. Linearization [LLLS13, Koç14, SMDZG13]. Linearized [JSSB14, AL10a, zSWZW12]. Linearly [Lia14, lLZ14a, Cai11, Cor11b, FPC14, Fan14e, QW14, QTL^{+}12]. lines [BAO14, Bot12, Che11g, DS13b, FFOC14, HBTU10, HHU12]. linguistic [LLZ14b]. link [Sun14]. linkage [B– D11]. linked [NN14]. linking [O’R11b]. Liouville [ABRT14, SQJ11, ZZW13, AGM12, AMS14, And11, jASzZ11, jAS14, BD10, Dou10, Dum10, EDAM10, EMP12, FHLW11, GWS13, GM10, GKW14, Jan13d, KKBO12, KM13c, KY14, KT14, LLY11b, Liu11a, Mer14, OC10, PHGV13, SW12, TF10, Tun14, WSD12, YW11a, ZH12, Zha10c, ZLZ12, ZWZ13, Zha14b, ZGZG10, ZDW11]. Liouville-like [ZGZG10]. Liouvillian [MGCFMM13]. Lip [MRSS11]. Lipschitz [GMR13, Ant11b, Cvi12, Dra14, HR11, Hu12b, MM14b, MMKD14, RA12c, ROA11, Ste12h, Ste13j, Ste13l, THX14, WCZ13, ZB11, ZWH14]. Lipschitz-continuous [Ste12h]. Lipschitzian [sCYW10, yLbLjH10]. Lipton [WZZL12]. liquid [ASM11, CP14a, CGLM^{+}12, PPC13, PS12b, ZCF13]. liquids [KS10c]. lithium [QS14]. little [SS11e]. Littlewood [WW11e]. Liu [YZZP14a, DW12, HL14d, LWM11, Wu14b, YZZP14b, ZJ13]. Liu-type [Wu14b]. living [URR11]. LLT [ZY13a]. LM [Kal12, SDB14]. LMI [LCL^{+}13a, Mao11a, SKSK12, Sin10b, ZB11, ZW10d]. LMMs [IG11]. LN [HJW11]. load [MBR10, SMBI14, Ste14a]. loading [EÖ10]. loadings [Cha14c, CBLC12]. loads [Ste14a]. loans [WW13h]. Lobatto [Bhr13]. Local [AM14e, AC12, CSQ11, CFZ14a, FG13a, FBGH12, FKH10b, GKS10, KAK11, KK13c, LMZ13, MN14b, RA10c, RA12b, Sol12, Ste12h, WZWW13, AS14a, AY14, AHM14, BH13, CN14a, CFCLM13, DZLW13, DWZW14, DMGK12, GLY14, Gal11d, GA11a, GZXJD10, HWPY14, HKK12, KUTY14, KS10a, Lai10, LWQ10, NCdCDM13, NA14b, Pak11b, PTK14, RMMM11, RA12c, RCSH10, SMI12, Sar10, Sar12c, SAA14, SLZ12, SFP^{+}11, Tad12, TL14, Tsa13a, Vil12, WLJ11, WCZ13, WL10d, XH11, XDZ14, ZYZ12, ZM10a, ZD11a, ZYLS12, Zhu12b, ZZ13c, ZANA12]. locality [ACQOR11]. Localization [CG14d, CPS14, CN12, DM13b, Peñ13, Peñ14]. localizations [AKZ14]. localized [lLZ14a, LL10e, Pan12a, Wu13, Zha12b]. Locally [YL14b, Ant11b, HLeH12, LLSY11, Sht11, JSSB14]. Locating [CHM^{+}13, GK10b]. Locating-dominating [CHM^{+}13]. location [FHK13, JS14a, PT11b]. Locking [XYC11]. Locking-free [XYC11]. locomotion [Sie14]. LOD [WW14b]. Loewner [DO11]. Log [GTPM14, BCWM10, LTL12, LW12a]. log- [LW12a]. Log-aesthetic [GTPM14]. log-exponential [LTL12]. logarithm [LPY11, MTR13]. logarithmic [CCLF13, Den11b, LL10i, Men12a, MGGS12, Ste10j, Ste10k, Ste10m, Ste12i, YWCH11, WZ11d, ZWZ12]. Logarithmically [Che11e, Guo13a, WWW12]. logarithms [ZMZ11]. logharmonic [AAA12]. logic [CZZ^{+}11]. Logistic [LB14b, Kud13a, KSM14, KP11b, LHC14, LD14, Rep11, WWS10, ZS14b, ZT11, ZY14a, LWS11a]. logistic-type [Rep11]. Lomax [COPP14]. Long [JL10, LJX14, Nie13, SAAS14b, AA13, Bai13, Cai11, CCJ10, Cha13c, DL12a, GZZ10, GZL14, Lia14, LGS^{+}11, Lin14a, LdS10, Mam13, ÖK11, SXL13, TES14, WFZL12, ZJLO14]. long-run [TES14]. Long-term [SAAS14b, CCJ10, Lia14]. Long-time [JL10, LJX14]. long-wave [Cai11, ÖK11]. longest [PP12b, PP13a, PP15, PFS10, SMPA13]. longest-edge [PP12b, PP13a, PP15, PFS10]. longitudinal [GYL14, PQRR10, Sin11a]. look [LS14e, UB13]. loop [BBS10, DWZ12, FYYT11, JWW14, Koç14, LR14b, MC13, RCYL10, WLL14]. Lord [Sar13]. Lorentz [WWW12]. Lorentzian [GARK12]. Lorentzian-type [GARK12]. Lorenz [EEZ14, WDZ13, WHF14, yWZhL14, WT11d, Yan11b, ZSWS12, ZZ14b, Zhu10b]. Lorenz-like [Yan11b]. loss [AH11b, DAV12a, DAV12b, GT12, RTPH14, DAV14b, YJ10]. losses [BAO14, WZH13]. lost [BM14a, GKC11, KEH14]. lot [BJB11, CBTGWW14, GLSJCBS10, GKC11, SYB13, TCBL11, YPSW11]. Lotka [AC12, BRW13, CDR13, CGRS12, Dan10, DGOR13, DX14b, Far14a, HL11a, KG13, LTX11, LCL14, MCWC13, Mur14, Vil10, Wan11d, WHL11, WLQ^{+}10, XLS13a, Yan13c, YZ14c, ZZJ14, dlHV12]. lotsize [San10]. Loud [Vil12]. loudspeakers [dOS13]. Love [BBF12b, ZZHZ14, MJ13a]. Low [KCGH14, WFwC12, CHT14, EASK10, Liu13b, RW14b, SZS13, SWD11, YWW^{+}14, Zho11b, IS14]. low-carbon [Liu13b]. Low-cost [KCGH14]. low-density [EASK10]. low-frequency [YWW^{+}14]. Low-rank [WFwC12, Zho11b]. Löwdin [MÇ12a]. Löwdin- [MÇ12a]. lower [AG14d, Aza14, CLTL13, Chi10, DMP13, GD10a, JL14, KX14, LSJ11, LLHY11, MP12b, MKSC13, PV10, Ram11a, RTPH14, YSB11]. LPV [HZ12a]. LQP [BA14a]. LS [GWML12, LF11a]. LS-SVR-based [GWML12]. LSQR [CPZ10, KZ14a, QW14]. lubrication [KBB12, MHA10]. Lucas [BB14b, BT12, Bil14, GCMP14, GTU13, İpe11, Özk14, PG13, Ram14, SC10a, SCH11b, Wło13, ZJ14c]. Lundberg [LSS11]. Lupas [GRY14]. Lur’e [DGH11, GHG14, ZCŠ10]. Lurie [NP10a, WL13a, ZC13a]. Lyapunov [Çak10, Çak13, YL10b, YKL12b, Akt14a, BR10c, ÇT10a, CGTCF13, DSPZ13, DLLW13, GZ13, GZ14a, Ign10, KPLC11, KPP^{+}13, Lee11, LZLW11, LSYS14, LZL14a, LT14, LS14d, LZ14d, MAO10, NZ12, PKPL11, PCM14, Sch11a, SL11b, VDL12a, VVK12, WFW10, WYK11, WL13d, YDZS11, YKL12a, YKL14, mZT11, ZSZ12]. Lyapunov-type [Çak10, Çak13, YL10b, YKL12b, ÇT10a, LT14, WYK11, YKL12a, YKL14]. lying [KGM14]. Lyness [DL14].

M [BM14a, Can11, CCJ14b, EM11, GCG14, LT11, PK18, TLW11, BM14a, CCJ14b, GCG14, wMwWsT11, WSY14, ZCL14b]. M. [EM11]. M.E [Pan12b]. M/G/1 [wMwWsT11, WSY14]. M/M/1 [BM14a, CCJ14b]. M/M/c [GCG14]. Ma [BAR11, Nad07]. Mach [RW14b]. machine [CYW^{+}14, EPMU10, HLS10, HL14c, JCW12, LCH13, Liu11b, LYZG13, LWW14, MQ14, Neu12b, WWZ10a, WW11d, WSY14, WL14a, WCW^{+}13, XG10, XYL10, XZY14b, Yan10e, YW11b, YCC^{+}12, ZZW^{+}12, ZL13a, ZHC^{+}14, LYGZ13]. machines [HL14c, NDZ14, SFL^{+}11, Niu11]. machining [KEGH13]. Mackey [BBI13a, BBI13b]. Maclaurin [UB13]. macro [BM14c]. Macroscopic [Aub10, Dol14, PV11]. made [Bas12b, Ngu14]. MAFS [Reu14a]. maglev [ZZ13a]. magma [GZ10b]. magnet [Ver11]. magnetic [BMMS13, CvW10, DFGL14, DKU13, Gwi11, Gwi12, KHK^{+}11, KMAG12, KS13d, lLZ14a, LX11a, MA13b, NZ14, Pal10, SBS10, SD13a, Üna13, XK14, ZA14b]. magnetically [Cha13a, YS10a]. magnetization [CGGZ11]. magnetized [GGYW11, SSM11]. magneto [GZW12, LWZ13b, Lot14, OADL14, VPS12]. magneto-electro-elastic [LWZ13b]. magneto-hydrodynamic [VPS12]. magneto-micropolar [GZW12]. magneto-thermo-microstretch [OADL14]. magneto-thermoelastic [Lot14]. Magnetohydrodynamic [XK14, ASRI12, ASR^{+}14, BD14b, CSQ11]. magnetohydrodynamics [Wil12]. magnetostatic [ZD10]. magnetothermoelastic [DRJ11]. magnitude [Den11a]. magnitudes [KVP14, PE10, ZR13]. Mahler [Kek11]. Mahoney [ROK10]. Mahony [S.14, LTW12, MA13c, RLV10, SGK11, SYZ10, SY10c]. maintenance [HLS10, LC12, WW11d, XH13b, XZY14b, Yan10e]. maker [GPP^{+}10, WHI12]. Makespan [EPMU10, AA12c, HLS10, mL12b]. making [BDGIPG12, Che12a, DZS^{+}14, DRT11, FZLZ13, LLZ14b]. malaria [GLL13, KMAG12, NNG10, TFMB12]. malaria-infected [KMAG12]. Malinowska [HJMY14]. Malliavin [XLY14]. Malthus [ZL10d]. malware [MVS14b]. Man [Can11]. management [GHC13, LL12b, MB10]. Manakov [WHG13]. Mandelbrot [AK10d, AK13, ARC14, ER13]. Mandelbrot-like [ER13]. Manhattan [Zhe14]. manifold [yCfW12, LMT11, Xie12b]. manifolds [AAB^{+}13, AAB^{+}14, AM14f, CDST14, NAL12, WHF14]. manipulator [CR11a, DESW10]. manipulators [AMPT14, KPBC14]. Mann [BL11, HYD10, LJKU11, MM11a, NS11, SW14c]. manpower [GD11]. Mansour [XC13]. manufactured [NTdCJL12]. manufacturers [DS14d, SCN14]. manufacturing [YH14e]. many [CT11a, Che14d, CT14b, DWY14, KTS11, LSLL14, MXH10, TL10, TLL13, Van11b, YZD10]. many-objective [DWY14]. map [MPB13, San14, WLXX14, dOL13]. MAPLE [CZL11, VMRA14, dYyR13, jZzFbL10]. Mapped [Zha10c, FHW14]. mapping [CLC^{+}13, GMZL11, Hwa11, Hwa14, Iid11, KKS11, LLWW10, LYXP13, SMN11, pYmK13, ZS14a]. mappings [AKR10a, AN– D12, ANR13, ARR12, AT12, BS10c, BB12, Bor12, BDN^{+}14, BL11, sCYW10, sCCLY10, sCLCY11, sCCL11, sCLCaL11, CWL^{+}12, CLCZ12, sCWT^{+}12, CLC^{+}13, sCWW^{+}14, CWWC14, CS10c, CMW10, ĆASH11, CZSV12, ĆARA12, CMM12, Den11c, DV13, DM12, DHZ11, – Dor13b, EVR13, GS12a, GFMP13, Gu10, HZ10, HHT13, ICKV14, Inc12, Jun10b, Jun11a, KI12a, KPK14, LAR13, LW12a, LSZ12, yLbLjH10, MM11a, MN14b, NKR12, NS11, OS11, PQW10, PQR14, QCK10a, QCK10b, RRG11, SK11a, She12a, SSRA12, SK13a, Sta14c, SLZ11, THYC10, Tse12, pYX10, YZK12, pYmK13, YK12, ZOS10, dOL13]. maps [ARN10, ANR12, ARG14, Bel12, DI13a, DI13b, GIR10, KudKL14, Mae10, NA14a, O’R10, RRR11, YTCY12]. Marchaud [ZL12b]. marching [BCGV11]. Marcum [BS10a]. margin [DGS14b, WLJ11, ZDH14]. marginals [NOS10]. marine [Gaz12, KHA12]. market [FGRLSV14, JAG11, NDZ14, RZHH13, RZHH14]. marketing [Goy14]. markets [DW13a, NR12, SS13c]. Markov [BCNP14a, CS13a, ESAH10, BCNP14b, CYG14, DJT10, Gra11, HQ14, LW06, LSSZ14, LWL14, SO10, ZA14a, ZBM10]. Markov-type [CS13a]. Markovian [BMR10, BRP13, CZD^{+}14, DZZZ12, DZZ14, GWW14, GLHCLP12, GSDW14, HL13a, KXW14, LZLW11, LC14d, LZL12, MZW14, MS13a, MK14a, MM14b, RSPK13, RWkPk10, Siu10, SLL14, TLZZ12, YMZS10, YCS11, ZaFT11, ZLSW13, ZLL14, ZCZ^{+}13a]. Marquardt [qDG10, FZ13, HM10b, KR11, LWH14, LZ10f, XM11, Yan13e]. Martin [MHXyY14]. Martinelli [ABBR12]. Maruyama [LM13b, Mil14a, ZSL10, ZXF14]. Mascheroni [BE13, CM13c, GI13, Mor10c]. masks [SHSC13]. masonry [SSS13b]. mass [All11, CRW11, FHAL14, FW12, FBGH12, IS10b, KM11a, LL13g, PC13a, WRZ12, ZL13b, ZYZ14c]. mass-action [FW12]. mass-conservative [WRZ12]. masses [ARL13, CL14d, TJZ14]. Master [BYM14, GHG14, GBD^{+}12]. Master-slave [GHG14]. matching [DXC^{+}14, GG11b, QFCZ13, SKB14, TZYH14]. material [AAMAS11, Cha14c, HYT14a, KTT12, LWZ13b, Ngu14, WM13b, XRBZ13]. materials [Caz14, KA12, KS13c, OR14, Sin11a, SXDJ11, SXCD11, TSJ13, Yig11, ZM14c]. Math [Aba12, AFSMRL14, Asl10d, Ban14, BM12b, ESAH10, HLSS10, LL12e, Mar14, MK10a, Pak11a, Rim13a, Tom15, Tua13, WZ13c, Web12, Wil10a, Xie12a]. Mathematica [CCMV12, UK11]. Mathematical [AMM11, AHJ14, BBB12, BAO14, BH12a, EMH10, GAC12, JDN14, JCB11, NNG10, NSV13, TFMB12, ZCL14a, Ant11a, BCÁV14, CCG10, CLS12, DD12, FKH10a, GLL13, KHA12, LD13a, LTL12, MA11d, MV13, PRT11, PLF13, QS11, SRS^{+}13, SMR14, SS10g, URR11, WDBM14, YDZ11, ZCL13a, ZCL14b, Zuf12, CG12a]. Mathematician [Can11]. Mathematics [AKB13a, AZ19, BDG13, BCNP14a, CX14b, EM11, Gür12, JDJ12, Mos14, Pan12b, Pan15, PK18, PP15, Rho19, SQ11, TJ13, WHS14a, YZZP14a, ZZ15, Lol12, Rem10]. Mathieu [CG14b, GHPR12, KK14a]. mating [Hor10, KHD12]. MATLAB [RV13a]. Matrices [BR10a, Cra12, AGT11, Ada11, AdFY13, AR12a, AR13, AR14b, ÁGVZ12, AKB13a, AKB13b, BR12, BHP13, BT12, Boz13, BFB12, CRR^{+}13, CL11b, CDGR13, CT14a, Che14c, Che14i, CKD12, CDDL13, CD14, DGSB12, DP13a, DP14, Din13a, DGM^{+}14, EMA13c, Elo13, Fen13, FWZW10, Gem11, GWLZ10, GRKP10, GD10b, Gus12, GG11c, GG13b, GG13c, GG14g, HW14a, HLTU12, Hla13, HY14, Hou14, HYC^{+}10, HSC14, İpe11, JM10a, JM14a, KZ14a, KB11b, LLSY11, tLpOjCW10, Li11a, Li12a, LWZZ14, LGZ14a, LZFR10, LXY10, LY10b, LLW11, LY12c, LG11, LO14, MT14, MPC14, MP14b, MH10a, MPS14b, MMS11, MD12c, NB11, NY12, ÖA13, ÖKU12, RV13a, Res11, Res13, Rim13a, Rim13c, Rim13b, Sav11, SC10a, SCH11b, SGZC13, SSN^{+}12, SEM11, SY14b, SY14d, SS11d, Su13, TL10, TLL13, TK14a, TS10b, TCI12]. matrices [Tos14, WWLL11, WHW13a, Wan14d, WCN11, WSK12, Wu10a, Wu12c, WLLD13, XQ11b, wX13c, YH14a, YH13, Yor14, YZ12b, Yun11b, ZZ10a, ZB12, ZZGL14a, ZJ14a, ZZX14, Zho11b, ZJ14c, ZCZP14, ZZS14, fuy13]. Matrix [jAS14, GWA14, HGN11, HYC^{+}10, KM11b, AZ14b, AKZ14, AM13a, ANR14, AB12c, AZ19, BBT10, BD14c, BS11c, BTS12, BS12c, BZT11, CC10a, CCJ10, ÇBS14, CWS10, CPZ10, CY11a, Che13e, Chi14, Chu10, CG14e, DWQG12, Das12, DR10a, Din13a, DM10b, DM12, DGM^{+}14, DKMM11, DLL11b, DA11b, ESAH10, Elo11, EGMPMRBC12, Fen11, Fuy11, GXSH14, GM12, GÇ11b, GH10c, mHhL10, HW12, HMSW14, Hu10, IS11, JK14a, KPP11, KM13b, KM14b, KWS10, KCG10, Kra14b, Kyr10, Kyr12, Kyr13, Kyr14, LPH11, LR13a, LW06, LL10c, LZ10b, LWZD10a, LWWZ11, LZDS12, LGZ14a, LW14b, LZL14b, LL11b, LYD10, Lin10, LL11d, LHZ10, LJCI13, dLL13, Liu14d, LC14b, LV14, MDÇ13, MK14a, MD13a, MC12b, MTR13, MTER14, MP12b, MM10b, Mat14a, MMKD14, MD13c, Neu13, OC10, PD14, PKBD12, Pat13a, PX13, PR11, dFPF14, PT14b, QW14]. matrix [Ray12, SE12, SIDR11, Sas12, SIRD13, Sol14, SR13, SCIMM11, Sta11, SG10, SHSC13, TW13, TS13, TT13, USAF14, UT13, WLY11, WWH13, WY13a, WXLW13, Wan14c, WS14, WH11, Wri11, WZDW10, WLH11, WLD11, WFDL11, WZL12, XLY10, XQ11a, XQ13, XZ13b, YC12, YB12, YT13a, YHW13, YH14d, Yin10, yYyL10, YZ13, YDTW11, Yua10, YWD13, Zha10c, ZLL11a, ZLGZ11, Zha11b, fZwXyZ11, ZW11a, ZW13d, Zha14c, ZW10b, ZF13a, Zho12b, ZCL13b, ZCZ13b, ZL10f, GL14a]. matrix-free [LV14]. matrix-valued [AB12c, DGM^{+}14, WS14]. matrix-variate [Mat14a]. Matter [LL13f]. maturation [NNG10]. MAUT [PKD14]. Max [CKD12, CD14, CG14d, CK13a, LS12b, MSS14b, Ste10e, Ste11g, Ste12s, Ste12o, Ste13g, SAAS14a, SAAS14b, Ste14e, SHWX14]. Max-norm [CKD12]. max-product [CG14d]. max-type [CK13a, LS12b, Ste10e, Ste11g, Ste12s, Ste12o, Ste13g, SAAS14a, SAAS14b, Ste14e, SHWX14]. maximal [Ou10, She14a, SLZ11]. Maximally [Mat14b]. maximization [BGRS11, FNV10, HJZC13, MLM^{+}14]. Maximum [ABGGS13, BA14b, GCG14, HT11a, LC14c, SPP12, YLAT14, Duk12, FKS13, FP13d, HLH12, Ilt11, JCW12, QZ10, Rud12, WLJ11, XYL10, ZDH14, FT11, Tag14]. Maxwell [Imo14, ANHA14, BC11a, EL13, CT14b, CYZ14b, LCH10, lLZ14a, Maa14, PM12, TPD10, WFL14b, YWW13, YLS13, YJ14a, dlRS14]. mBPSO [FYL13]. McDougall [Fuk14b]. mCH [gZrLfM10]. MCMC [SLM14]. mDP [gZrLfM10]. Mean [LL14a, RH14, ZZQ13, AEESMA10, ABK14a, AA12c, BSM14, CL11a, CDLZ11, CJRV12, DJ12, DZZZ12, EG11, EKK14, GK11, GKV12, GYL14, GC11d, HS14a, HS14b, Koy12, LW11c, LLZ14b, LP13d, PTW11, Ral12, SS12a, TK13, Wan14a, Wan14f, YLMZ13, ZZWF13, dlHV12, CCJ11, MS11]. mean-covariance [GYL14]. Mean-square [ZZQ13, DZZZ12, Wan14a]. mean-variance [YLMZ13]. meandering [ACG13]. meaning [Car12]. meaningful [GGM^{+}13]. Means [HH13a, Asl10a, BP10b, BP13, BEdAdSK14, CHC12, CS11c, CT13c, CK13b, EV14, HD14b, Kov10, MKSN13, PQW10, ROA11, RRRRTR12, SMDZG13, SS10e, Wan10e, Yüz13]. measure [AOTV14, AP13, AAH14, DM10b, Gra13, GSC14, RH14, Esm12]. measured [Yua11]. measurement [RVS10, ZL11c]. measurements [BWHQ14, CÁGGLP14, IG11, WLP^{+}14]. measures [BCM12, CDS13, DM10a, HMR12, KHP11, Mal14, MVRVSCMV14, Sza13, YLC14, ZSZ12, dACSS12]. Measuring [LDR10, MOZ11]. Mechanical [EHSM14, BS12b]. mechanics [JJL14, MSCK11, SSK10, SMR11b, XGG^{+}11, ZLX12]. mechanism [BZZ14, ES12c, VD14a, XBYF14]. mechanisms [HT14, YWLX11, YZZ^{+}14]. media [Alf10, Alf12, All11, ABK^{+}14b, BCWM10, qClLyW10, CRX12, CS12b, HFS12, HCCT14, JAG11, KPC14, MATA14, SRS^{+}13, SM14a, XLL11, YS10a, YLS13, ZLZY12]. Median [BSH13, qGyFzX13]. Median-oriented [BSH13]. mediated [HRH14, LKY14, WPKE14]. medical [CCJ10]. medium [Abb14, AAMADH11, AV14, CS13d, CMS11, DAA14, ITLA10, KM11a, Kha14a, Lot14, PC13a, PMV14, Pan15, PP13b, PMD14, Sha12b, SM12, SD13a, VS10, Vis14, Xia10a, ZCL11, YJ14a]. Mehrotra [YLD14]. Mehrotra-type [YLD14]. Meir [MAK13]. Mellin [BM12a]. mellitus [FGFRV11]. Melnikov [Sha14b]. melting [BMHM14]. member [B– D11]. Membership [HS13, YK14b]. membrane [SFF11]. memetic [VC12]. Memory [XPLZ14, ABKG^{+}13, ASB10, Ber14, DPP12, Gee12, GM13b, HZW12, Kan12b, KfM11b, Laz11, PID10, PDN11, SAL12, SGG12, TS12b, VL12, LLS11]. memoryless [LH11a]. MEMS [KCGH14]. Meng [Mar14]. Menger [CIVS14, GAV14, VDFV12]. Mercer [PPS13]. merge [CHK10, SLM14]. merging [XBYF14]. merit [PZ14b, TC14b]. meromorphic [BÖ11, CALR11, CY11b, DO11, Hus10b, SYW13, SÖCS12, WLX11, YWCH14]. meromorphically [SI11a]. Mesh [HL13c, BHG10, CLX13, CVP^{+}10, DN14a, HJVC14, Jha13, JHN11, LLS^{+}13a, SL14b, VT13, ZLX12]. mesh-free [BHG10]. meshes [Bra13, LMT11, MVM10, MGC12, Oss10, PP12b, PP15, TD14a, TVZ10, WJ11, ZCZ14]. meshfree [MR13b]. Meshless [DOV^{+}11, HBTU10, HWPY14, KM13a, MPS14a, SWL12, WQW^{+}10, AGM14, DZLW13, DWZW14, DS13b, HHU12, Kia14, LLT^{+}14, Reu11, uiH14, ZD11a]. mesoscopic [NKM^{+}10]. messages [AdC14a]. metaheuristics [CR11a, IS14]. metal [DKU13]. metallic [dCDNV^{+}13]. metapopulation [NN14]. Method [GH10d, KM13a, MV14, MCR14, Shi13a, YL14b, Aba12, eMA13a, eMBHA13, AdC14a, AK10a, ABJ10, AC10a, ARE14, AGM14, AMS14, AAAH12, ASRI12, AH10, AHRV11, AT10b, AB12b, AGHdD13, ACG14, AANA10, AAB^{+}13, AMR13, AHVR14, Amd10, AC10b, AL10a, AY11b, AHKS12, AHDD13, Ant11a, Ant12, Apa10, ASB10, AP14, ANS14, Ard13, ACH10b, AH12b, AG13, AH13a, AH13b, ACK14, AM14f, AG14f, AM14e, AHM14, AS12b, AS13a, AASS13, AE13, Asl10a, Asl10c, Asl10d, Asl12a, Asl12b, AZAK10, Bab12, BRS13, BDQR11, BDL13, BM14b, BP11a, BDR12b, BY11a, BRAB11, BP11b, BS12a, Bek10, BA10, BS12b, Ber12, BR14a, BTVC10b, BDG13, BES10b, BTS12, Bhr13, BY11b, BCC11a, Bie11, BC10b, BCW13, BHG10, BCEAV12, BCÁV14, BHP13, BBK13, BA14a, Bog12, Bot12, EL13]. method [BRM11a, BRM11b, BMR13, BR13, BA14b, BLR10, Bra13, BS14c, BCGV11, Bum12, Buo10, CHS14a, CSQ11, ÇE11, Cao10b, CLZmF11, CHX13, CBA^{+}14, CT12a, CK10, CC11b, ÇBS14, CSC10, Cha10, CLCL11, Cha14b, CC14a, CBO12, CS10a, Che10c, CS10b, CZZ^{+}11, CZ11b, CJ11b, Che11b, CTLL11, CX11, CLW11, CW12, CD12, CLF13, CS13d, CYM13, CLTA13, CF13, Che13d, Che14j, Che14f, CFZ14a, CZ14b, CFX14, CD10a, CFZ14b, CZL14, CCA11, CD13, CL13a, CS10c, CEXG12, CDMT12, CP14b, CC12c, CP12b, CSF12, Cov11, DGSB14, DRJ11, DW12, DWQG12, DZLW13, DWZW14, DW14, DWY14, DOV^{+}11, DGDA10, DNSM11, DB11a, DX10, DS13b, DS14e, DJ14b, Don12, DZ13, – Dor12a, Dos11, DFM10, Dri13, DHSZ10, DW11, DR11c, DCR12, DHQ14, Dum10, DAA13, DLS13, ESV12, EK11, EDAM10, EGMEM14, EEA11, EEEE05, EBR13, EZRMR12]. method [EZ14a, ECG14, Els12, EÖ10, EALS10, EH14b, ELIGMB11, Eto11, EZ13b, Fal14, Fan10a, FZLZ13, FL13, Fan10b, FH11a, FEGEA14, FC10, FcDDqZ10, FLW11, FFOC14, Fer14, FNVLPD^{+}13, FR11, Fuk12, Fuk14b, Fuy14a, GGP^{+}13, GK14, GG13a, qGyFzX13, GLS13, GC11a, Gen11, GQ14, GS10b, Gor10, GKW14, GHS11, GSN12, GSSS13, Gu11, GG14d, GT11c, GG14e, GÖS10, GZXJD10, GZ10c, GZ10d, GZZ10, GL10c, GMZL11, GML13, GYS13, GFZ13, GZL14, GK12b, GSRR13, Gür10, Gür12, GMSE13, GMR13, HJVC14, HB14, HXY12, HBTU10, HHU12, Har10, HKI12, HET12, HMIZ13, HBB11, Hay10, HZ10, HYD10, HCL^{+}13, Her11, HH13a, HOMA13, HSW12, HHM14, HWPY14, HDT11, HR13, Hou13, HL14b, Hua10, HKH10, lHyLfG10, Hua11a, HM12, HL13d, HL14d, HM14b, HBDC14, HMGC18, HG13b, Iid11, IU10, Inc12, IISA10, IJ11, JKN11]. method [Jan14d, Jat10, Jat12, JPB10a, JA14, Jha13, Ji14b, JC14, Jia13a, JHX14, bJLyZlL14, JC13, JJL14, JSSB14, JMS13, Jun10a, JBDC12, KPC14, gKW12, KPP11, KR13b, KP14b, KR12, KM14b, KW10, Kes13, KcZ14, Kha14a, KR11, Koç11, Koç13, Koh12, KV13, Kot14, Kov10, KC13b, KCC12, Kud10a, KR10, KSS12, KET10, La 10, LP12a, LJ12a, Lee12, LPB10, LH11a, LW13a, LWQ10, LCH10, LMMW10, LSX10, LZZ11, LH11b, LZQ11, LSJ11, LF11c, LYZ11, LF11a, LL11a, LWW11, LTZ12, LZ12b, LTL12, LWL12, LK12, xLhYlZ13, LL13a, lL13c, LWZZ14, LWYY14, LW14d, Li14a, LR14a, LL10d, LZ13b, LKS14, LW14f, LZ14a, LYD10, LLBW11, LW11c, LWF11, LY12a, LWFG13, LTW10, sL10a, LTM10, LLL10a, LLWW10, LZZ10, LHL11, LH11c, LX11a, LqTyCZ12, LCC12, Liu12a, Liu12b, LJ13, LW13c, LYXP13, LM13b, LLS14, LWN14a, LFLH14, LL14f, LW14g, LZX14, LP12b, LZ10f]. method [LZ10g, LXZ12, dALORPGC14, LLN10, LL10h, LZX10, LWM11, LLZC11, LV14, lL10g, LHL^{+}13, LLS13b, LD13c, LZ14d, LXLY14, Luo14, MG10a, ML10b, MXH10, ML12a, MFZ12, qMWjHpX13, MH13a, Maa12, MDR14, Mah13, MC13, MAN10a, MHA12, MCM10, MT10, MP10, MP12a, Mar11, MR13b, MMO13, MH10a, MP14c, Mat10, Mat12, MF13, MC12c, MA13c, Men14, MO12a, MNP13, MC14c, MMS11, MMS– D13, MO14, MG10b, MJ12, MB13b, MB14, dMbJmY11, MQ14, MESAN13, Moh14, MG14c, MAHAS14, MYL^{+}10b, Mos14, MZL07, MPW12, MS11, NTT10, Nar13, NAL12, NR14, NCS12, NKM^{+}10, NZW11, NY12, NMDWAS10, NK12, Occ11, Odi10, ONBN10, OPD12, ÖA10, Pak11a, Pak11b, hPrLgZ10, PG13, PV10, PSSS14, PN11b, Pan12b, Pan15, PT13, PR12a, PG11, Par10c, Par10b, PZ14b, PL10, PW10, PR11, PMC11, PTS11, PC14b, QZ10, QS11]. method [QQXY12, QMM13, QFCZ13, QS14, QLL12, ROK10, RMN14, RB14, Rah13, RDKU14, RK12, RAIM14, RSN11, RWB10, RA10b, RA10c, RCH12, RA12c, fRbLJyS13, Reu11, Reu14a, Reu14b, RCSH10, RGL12, RW14b, RB11b, Ros11b, RnZ10, RX14, RSK11, RC12, Sab13, Sab14, SNMD13, SWL12, SR14a, SB13, SHA12a, Sar12c, SIDR11, SS14c, SATdV13, SSTZ14, SSZ11, SSCA13, SJR10, SMBI14, Sha13a, SBS10, SS10c, Sha12b, She10, SX14, SWX10a, SZS13, SP13a, SWW14, SCC11, SDK10, SY10b, SAK10, SR12, SU12, SKD12, SSCT14, SSU13, SGA14, SHZ10, SP13b, SXDJ11, SXCD11, SA12b, SXC^{+}12, Suh13, SK10c, Tad12, TZZX11, THDF11, TDZF14, TW14b, TD14a, TD14b, Tem12, TA13, THNTNXR11, TKC12, Thu10, hTlCqL10, TXX14, TSJ13, Tom12, Tom14, Tom15, TQWZ10, TDT12, TÇ13, TS13, TT13, TN14, Tra12, Tri12]. method [Tsa11, TTQ11, Tua13, VMS10, VG14a, VZ11, VDK10, VDK13, VL12, Wan10e, WQW^{+}10, Wan10d, WLCW10, WHC10, WY10a, WT11b, WWLL11, WaZZW11, WCL11, WZ11f, WZZH11, WLZ11, WW12, WCL12, WLH12, WF12, WW13a, WaZW13, Wan13a, WW13e, WWG13, WW13c, WYT13, WD14b, WSCH14, WW14b, WMM14, WLXX14, WZ14b, WS14, WW14a, WJ11, Waz10a, WRD13, WTZ11, WWC12, Wei14, WWY10, WYW14, WFZ13, Wil10a, Wil10b, Wil11, Wil13, WZ11g, WL10b, WLHZ10, WT11c, WD14c, XYY12, XWWF10, XY10, XD11a, Xie11a, Xie12a, Xin10, XG13, wX13b, Xu14c, Xu14a, XDZ14, YAW10, YZL10, YH10, Yan10f, Yan10b, YGL11, YCTD11, YPSW11, Yan11c, Yan12b, YZK12, YY13a, YLLL13, Yan13e, Yan14a, Yao10, YLAT14, YHW13, YCK13, YH14d, YT13c, YM10c, YSW11, YLTB12, Yu12b, YLWH14, Yua11, Yua13, YZ14d, Yun11a, YS13b, YS13a, ZZ14a, Zay11, ZB11, ZOA10, ZFH14, Zha10b, Zha10c, ZZQ10]. method [gZrLfM10, Zha11c, ZX11a, ZYZ12, ZC12, ZM12, ZLX12, ZJ13, ZW13a, ZC13b, ZW13e, ZY13a, ZZL13, ZHY13, ZY13b, ZL14b, ZM14a, ZL14c, ZLL14, ZxHW14, ZDW10, ZLY11b, ZLLL12, ZHZ12, ZY14b, ZTX10, ZL10c, ZZZM10, ZYSQ11, ZD11a, ZW11c, Zhe11, ZYLS12, ZhGQ14, ZM14e, ZS14c, Zho13, Zho12b, ZZC13, ZS14d, ZZG10, ZL10e, ZSD10, ZC11, ZXZ11, ZZZL14, Zhu11c, ZZ13c, ZR14, ZL10g, Zit11, ZANA12, Zuo10, ZZ11c, ZFCW13, ARH^{+}12, BZ10a, CXG10, GYD14, MH14b]. methodology [JE11]. Methods [KI10b, AS12a, Abh14, ACZ12, Ach11, AISS13, AA12b, ABKG^{+}13, AG14c, APA13, AST12, AAS12, ASST11, Alz14, ABGGS13, AAB^{+}14, ACCT14, And11, Ard11, AH11c, AH11d, AH11e, AH12a, BDD^{+}12a, BDGP14b, BDD12b, BJ12, BBSC13, BS11c, BFXZ10, BX12, BJ11, BIT12a, BIT12b, BL11, CMM10, CMX14, CC13, CFGRV13, CAY11, CB10, CZ11a, CW11b, CSZ12, Che13a, CS14a, CCGT13, Cho11a, CLND12, CL13b, CN14c, CNKS14, CD11b, CT11c, CHMT11, CTV12, CTV13, CST14, Cor11a, DEP12, DR10a, DS12a, fDxZxChT12, DAEY12, DHZ11, DC14, – Dor12a, DJT10, DR11b, Duk12, DA14, DHMU14, DPP11, Džu11, DPP12, EES14, EVDG12, EHRV12, Fan13, FHWZ12, FLZ14, FG11, FG13b, FG14b, Gal11e, GK10a, GP12, GT11a, GK10b, GK10c, GKS12, GTPM14, GLV10, GSGN11, GSDB11, GHZ^{+}14, GW12, GZF12, GMV11, Han11, HW14b]. methods [Hu10, HZ12c, HG13a, JMY13, Ji14a, JLTZ13, Jun10b, KG10, Kal13, KBK13, KYH10, KAG12, KA11, KL14b, Kim10, Kim12, KI12b, Kos13, KS11d, KKS12, Kum14, KP11a, LP11, LL10c, LHZ^{+}11, LD13b, LDWY14, LC14a, LLK14, LSL10, LLY14b, LZ14b, LW10a, LL10f, Liu11c, LL12e, LW12b, LNPS10, LHTN13, LWN14b, LL12f, MDD12, Mag14a, MA14b, MA10a, MLG13, MN14a, MIHS14, MPM14, Mil14b, MJ11a, MX11, Mon12, MAD11, MJ13b, NLLC10, NAL12, NBR10, NN11a, NSC12b, NSC12a, NCS13, NS13a, NC13b, NCS14, NC14, NZ14, NK12, OSRFF14, OR11a, OO14, OP14b, Pat13a, Pat13b, PC11, PID10, PMM11, PDN11, Pet13, PNPD14, QCK10a, QW11a, QMZ14, RD11, RMY14, Ram10b, RP14a, RY10, RW14a, RP13, RA12b, RDFL13, RV10b, Sad11, SSK13, SNC11, SMIS14, SGG12, SA13, SAP14, She12c, SDK10, Sho14, uiH14]. methods [SCK13, SL10b, SKC10, SWX10b, SCK14b, TM13, TS12a, TS14, TS10c, VLMHK^{+}13, Wan10a, WYY11, WZ11e, WDZ12, WWZ13, WWT13, WR10, WYBT12, Wu12a, WS13, WHZ14, WTQU14, XH13a, XZY14a, XT14, XLY14, YH14a, Yan13d, YLS13, YJ14a, YCF13, YZC14, YXZ10, YDTW11, YWL13, YZ12b, Yun10b, Yun11b, ZD10, ZA11, ZS10a, ZL11b, ZF11, ZLS12a, ZZGL14a, ZL14e, ZZGL14b, ZQ14, ZCU14, ZZY12, Zha13d, ZL14f, ZZH11, ZY13c, ZCS13, ZS14e, Zhu12b, AGM12, AKH13]. Metric [SGJ10, AKR10a, ARN10, ANR10, AKN11, AN– D12, ANR12, ANR13, ARR12, ABR^{+}11, AKR11, BB12, Bor12, sCYW10, CIVS14, ĆASH11, ĆSAV11, DV12b, DV13, – Dor13b, EVR12, EVR13, GIR10, GS12a, GGM14b, GMM^{+}14, HKK12, KKVCS12, LAR13, LYD10, yLbLjH10, LYLS14, MAK13, NKR12, NA14a, OP12a, PP13a, RRR11, Rad14, RRG11, RPA14, SSRA12, SI14, VR12, YK12]. metrically [KI12a]. metrics [ĆARA12]. Metropolis [RCSH10]. MEW [WTFC13]. Meyers [DCR12]. MFR [QLL12]. MFS [CTLL11]. MHD [ASR^{+}14, Pan12b, ANHA14, ASRI12, AV14, AA12d, BTRZ14, EEEE05, Fer10a, HCCT14, KBB12, KS13d, LL13d, PC13a, PMD14, VPS12, VSS13]. MHSS [wX13b]. micelle [NW12]. Micro [BM14c]. microarray [PJFL14]. microbial [GW13, YZZ^{+}14]. microbubble [BYM14]. microchannel [KML14]. microelongated [SM12]. micropolar [ASM11, EEEE05, GG12a, GZW12, Mah13, Pan12b, SBS10]. microscope [RVS10]. microstretch [OADL14]. microtubes [Che10c]. microtubules [YA14, ZMD^{+}14, ZSMB14]. midpoint [ACH10b, Bab12, Hwa14, Liu11g]. midpoint-trapezoid [Liu11g]. migration [BDDK14, GWX12, NN14, Qiu11, Sun14]. migration-driven [Sun14]. mild [AH11c, KD13, LP13b, MN10, Mop11, RSRA14, WZ11b]. Miller [Hus10b]. Milne [HS10b]. Milovic [Kha11]. MIMO [HMH13, JWWY13]. Min [MSS14b, NKM^{+}10]. Min-max [MSS14b]. Min-protein [NKM^{+}10]. Minda [BAR11]. mine [Gol14]. Minesweeper [Gol14, Gol11]. minification [COPP14]. Minimal [KWS10, KKVCS12, ABG11, AKC12, AAI14, ASB10, BC12c, BB13b, Dim13b, FGIV12, JK13, KP12, WW10, Wan14c, ZT13]. minimal-memory [ASB10]. minimalist [GAK12]. minimality [ZL13e]. minimax [Ant11b, AZ14c, BM13b, HZ11, bJLyZlL14, LH11b, WW11b, WW13a, WWW13a, XY10, YCXL12]. minimisation [ZB14]. Minimization [KR13b, TP12, AM14a, BAO14, CKP^{+}13, EPMU10, GS13c, qMWjHpX13, Rud12, ZL11b, ZC12, ZGY14]. minimize [HLS10, LL12a, ZHC^{+}14]. minimizer [LWFG13]. Minimizing [EG11, JCW12, mL12b, YW11b, CHX13, JZ11, LH10b, XYL10]. Minimum [CP12c, ZX11b, BFH12, BR14a, BO13, DA11a, Duk12, Kam14a, KGB12, Kyr12, LWZD10a, LL11b, Sun12, WZZ14, WTZ11, YX12b]. minimum-norm [LWZD10a, LL11b]. mining [Sev11, SBS11, YG13]. Minkowski [AOM^{+}11, JK13, LKY10]. Minlos [KM14c]. minmax [PR14]. minor [SC11]. minors [JK14a]. Miodek [MGYQ12, RK14a, XTAJ12, yZqLwW12]. Miodekequation [BK10c]. Mirakjan [ÖD11a]. Mises [SLW12]. mismatched [WS12]. missing [CÁGGLP14, GRKP10, RMMM11, RGA14, YL10a, ZL14d]. Mittag [TS12b]. Mittag-Leffler [TS12b]. Miwa [TMXG11, TTM12, ZgJlK12]. Mixed [CM12c, HMD14, KP11a, Liu12c, LHCI12, Pal10, RSBK14, RnZ10, Sha13b, TR13, TD13, WC13, XHXH12, YL11c, ABEK11, ACZ12, AAS14b, AZAK10, BBS10, BH12b, CMM10, CGGMG12, CC12a, sCCLY10, CS14a, CRX12, CZZL13, DB11a, DNI^{+}10, Din11, Din12, DLLW13, EJ12, EZ12, EGH12, FFOC14, FR11, GWW14, GL10a, GS11a, GD11, GÖS10, Guo11, GZF12, GFZ13, GK12c, HZ10, Hu10, IG11, JE11, JJPW13, KKV12, KR12, Kim12, KD13, KPW11, LJR14, LWQ10, LLC12, LXK13, LR14a, LSW^{+}14a, LKH12, LTM10, LL12e, LFLH14, Mat14b, MO12a, MHB14, Pak11b, PK12, RRM12, SK11a, Sas12, SZ11, SZS13, SZ14b, SR13, Ste10j, Ste10k, Ste12w, TSG10, TJ14, WZ13a, WLCW10, WCL11, WWG13, WZZ13, WW14a, WY12c, WFZ13, WPSC12, YZK12, YY13a, YLS13, YJ14a, bZlD11, ZaFT11, ZQ14, ZCD10, ZCC13]. mixed-norm [Ste10j, Ste10k, Ste12w]. mixed-parity [EZ12]. Mixed-type [LHCI12, WC13, YL11c, DLLW13, WZZ13]. mixing [CS11b, IC11, SS11i]. Mixture [RAW13, SLM14, BGP^{+}13, CBLC12, KRMM14, LCCA11, SPT12]. mixtures [GG12a]. MKdV [ZZ11c, SHC11, AMW12, GZ10c, GMZL11, HLZ13, JZW12, Sal10b, TI10a, Waz12a, YDT^{+}14, ZCG10, ZgJhD12, ZL10g]. mKdV-type [AMW12]. Mkdv integrable [Xu10b]. ML [IS14]. mobile [JDD14, LLT^{+}14]. mobile/immobile [LLT^{+}14]. Mocanu [WZ13c, Hus10b, NM08]. Modal [NMM14, KTT12, LGKM13, SSP11, Yua11]. mode [AA12a, BRP13, fD12b, GLX13, GWW14, JQ14, KXW14, LR13a, Liu10c, LZL12, Pai14, RSPK13, TLZZ12, WS12, ZX10, ZLZ14b]. mode-dependent [BRP13, GWW14, KXW14, Liu10c, LZL12, RSPK13, TLZZ12]. Model [CFN12, SBM12, Abb14, AAP10, AM14b, AKP13, AB12b, AKS11, ADD12, Ama14b, ABK^{+}14b, And12a, AL10b, ALP11, AWHW11, BKA12, BC12a, BLZZ14, Bai14, BPP12b, BBDS14, Bel11a, BC14a, BBI13a, BBI13b, BH12a, BJSS14, BH12b, BWaY14, BLR13a, BI10, BRW13, Cai10b, CL10a, CGL12, CGW14, CFL14, CNTG12, CBTGWW14, CCK11, CJK12, CCG10, Che10d, CBLC12, CLD12, CXL12, CH13, CW14b, CYG14, CZ11c, CW11c, CC11c, CX14a, CX14b, CC10c, CS12b, COPP14, CO13, DL12a, DS14a, DMS11, DRSG14, DMGK12, DMGK13, DGS14c, DS14d, DIAQ13, Den14, DISS14a, DK14b, DJY13, DFGL14, DWP10, DRT11, DYL14, DL11, EB12, ELS11, EWS10, FCJ14, FW10, FD12a, FXD12, FW14, Fan14a, FZP12, FHWZ12, Fer10a, Fer12b, FGS14, FKH10a, FL11, GGYW11, GC12, GL13a, GYL^{+}13, GHC13, GZ13, GGM^{+}13, GLL13, GS12c]. model [GRKP10, GW13, GHZ^{+}14, GS14c, GD11, Gui14, GTX11, GZ11, GC11c, GXF13, GSC14, HL11a, HdCOL14, HY11a, HLM14, HYT13, HSLG10, HB11b, HRH14, HDZW11, HYZH14, HLW12, INE10, IG11, JM13, JKN11, JSWJ13, JXA11, JZY12, JSH10, JWZ^{+}12, JV13, JK14c, JM10b, Juk13, KHA12, Kal12, KG11, KJ12, KK12b, Kha14b, KB14, KGC11, KG13, KR14, KN12, KS11e, KLY12, LSS11, LOKB12, LS13a, LKP12, LG10a, LSJ10, LQY10, LG10b, LZLF10, LF11b, LWS11a, LTY11, Li12c, LCH12, LW13b, LXH13, LXC14, LX14c, LWB14, LKY14, LSSZ14, LSW^{+}14a, LLY14a, LS10c, LZWZ14, LWL14, lLZ14a, LGS^{+}11, LOD12, LJX14, LZG10, LWTC10, LWWL10, Liu11e, LY11b, LWW12, Liu13a, LWG13, Liu14b, LH14b, LB14c, LC14d, LB14d, LLT^{+}14, LB14b, LZLY14, LB14a, LD14, LDZ10, LL10i, LF14, LYP14, MYJ10, MYT10, MHTH12, MLY12, MCWC13, MH14a, MDR14, MC14b, MAB14, MTD10, wMwWsT11]. model [MS14a, Mat14b, MS10, zM10b, MHXyY14, MA11d, MG12b, MP11b, MSR14, MV13, MM13c, MM13d, MSS14b, MGGS12, MGSMRÁ14, Muk10, MEL13, NM14, NAHH11, NSL14, NTS11a, NTS11b, NBG10, NKM^{+}10, NCdCDM13, NN14, OP13, OR13a, PSC12, PM14a, PSJGLS14, PAL14, PR14, PLF13, PMP11, PLTH12, PM14b, PFG^{+}14, PS10, QaC13, lQZWC14, QS11, QS14, dejW14, Qiu10, Qiu11, QL12, RG14, RRLA14, RTV12, SG11, SGE13, SP14a, SP14b, SP14c, Sam11, SRS^{+}13, SS14a, SSL10, SSL12, SSL13, San11a, San11b, San11c, San12, SCN14, SM11, Sar12a, Sar12b, Sar13, SM13, SMS14, SGCG14, SVGC10, SSS14, SHAO14, SY10a, SZH13, SY14a, SP14d, SW13, SRA10, SKIH12, SM14b, Siu10, SS12b, SMR14, SHZ14, SDB14, STV12, SS13d, SS10g, SS11i, SHM13, SLYJ13, SZ14d, TES14, TGL14, TDSA14, TL14, TKC12, TW11a]. model [TX14, Tie12, TVZ10, TPD10, Tsa13a, VDL12b, VDL14, VJ11, VJBDV13, VDFV12, WDBM14, WWZ10b, WzH10, WT11a, WY11c, WF11, WZ11d, WQ11a, WES12, WY12b, WZZL12, WW13g, dWcFwWjZ13, Wan13e, WW13b, WWW13d, WWZ14a, WW14d, Wan14a, WL14b, WC14a, WPKE14, Wan14b, WJ14, WHWC14, WWK^{+}14, WC14c, WDS11, Wil12, WZ11g, Wu14b, XGZW13, XLS13a, Xie11a, Xie12a, XX10, XTL10, Xu12b, XZ13a, Xu14b, XYL14, YL11a, Yan11a, YLWU13, YS11b, YPSW11, Yan13c, YL14a, YLMZ13, YJ14b, YA14, YZX13, YLC12, Yua13, YLK14, ZSMB14, Zei11, ZA14b, ZJS10, ZZQ10, pZJ11, ZL11a, ZMZS12, Zha12a, ZJD12, ZX12, ZjHO13, ZY13a, ZX13, ZYZ14a, ZZ14c, ZHXM14, ZT14, ZJLO14, ZW14a, ZZ15, ZT11, ZY12, ZWZ12, ZWP13, ZW14c, ZY14a, ZJ14b, ZZ12a, ZL10d, ZZW11b, ZLZY12, ZCC13, ZZQ13, ZZL14, ZZY14b, ZMC10, ZL10f, ZGY^{+}10, ZWLW10, ZYYZ12, dlHV12, dlSAQ11]. model-based [DK14b]. Modeling [CBA^{+}14, CZL14, Gal11d, GAK12, KMD13, KMC12, KDS13, LWTC10, Mis11, MMP^{+}12, Mis14, MZWZ12, SSST14, AMM11, BAO14, CNTG12, CC11b, CCCRR13, CTLD12, GAC12, Hua11b, KMH^{+}13, LJ12b, LD13a, lLW14, LLT^{+}14, MCL14a, Mad14a, MTSN14, NA14b, PKD14, PM14a, PM12, PPC13, SZG12, SHN10, SHBN13, SFF11, SWD11, TSJ13, URR11, YLC12, ZCL14a]. modelled [ML11b]. Modelling [Ber10, BMH12, TKLH10, XZ13a, XH13b, YX12a, CC10c, CSSSV13, Gür10, Gür12, KPC14, SJR10, SM10, Svá11]. Models [Alf10, AAKB13, Alf12, AAA13b, Aub10, BFH12, BM14c, BCÁV14, BMI^{+}11, BCNP14a, BCNP14b, BR10c, CCJ10, CCS14, CDR13, CFGRV13, CS11b, CGO14, CLS12, CEP11, CS14b, Dol14, EMM^{+}12, EGO13, FO11, Far14b, FGRLSV14, FP13c, GLSJCBS10, GS12d, Gwi11, Gwi12, IG11, KBAvBW14, KMR14, LR11, LDWY14, Lin13b, LHT14b, LXGM14, LS14b, LCCA11, LY14e, MLM^{+}14, MZWZ12, MKP13, NRWF12, NDZ14, OSRPC12, Örk13, Pal13, PR10, PC14a, PW14, RMMM11, SLM14, uiH14, SKC14, Soo11a, Soo11b, SS13c, TFMB12, VDLEK14, WL12b, WM13b, WW13h, WYL10b, YHS13, ZA14a, ZL14a, ZLY11a, ZZC13, Zil12, Moh14]. modes [CC11b, LPP^{+}12, NZ14, TJZ14, ZW11b]. Modification [NBP10, OS14, CHYT12, DS14b, DR11c, Ere11, MK08, MK10a, VT13, bOYqGL^{+}14]. modification-based [CHYT12]. modifications [– Dor12a, HH13a, Tsi14]. Modified [sCCL11, EH14a, EH15, GW12, IRS14, JSWJ13, JPB10a, KT14, LWYY14, LZ14a, LLL12, MMS11, SB13, SS10c, SKC11, SR10, VDK10, YSW11, ZZY12, AAC11, AKS14, AAA13a, AG13, AS13a, AM12d, Asl10d, dCBdFM14, CQJ10, CSC10, CLCZ12, CLZ11, Che13a, CLW14, CC14c, CC12c, DK14a, Den11d, DR11b, DCR12, FYYT11, FYL13, FL13, FYM13, FDHK14, FMS11, FGO12, GGYW11, GLR12, HMIZ13, HML10, HjH12, lHyLfG10, HL14d, HMPVQ14, IU10, Izg11, Kle12, Kre13, KKS12, LY13a, LW14d, LZH10, LTM10, LLWW10, LP12b, LP13a, LCLW10, LWX13, ML12a, MM11a, Mer14, MMCD12, MJ12, MB14, MMGM13, NCS12, NWZ13, PJFL14, RY10, RA10b, RB11b, SSZ11, She10, SX14, SS10d, SW10a, SW14c, Sun10, TW14a, TMJ14, TW11a, TTQ11, WDLM10a, WDL12a, WDZ13, yWZhL14, Waz12b, Wei14, WH11, WX13d, WYL10b, Yad14, YPSW11, YT13c, ZZ14a, Zay11]. modified [bZjZS11, ZZ10c, ZL10c, Zhe12, ZS10b, ZGY^{+}10, Zhu11b, DAV14a, BD12c, RNZH13]. modular [Sah11b]. modulated [SHI13b]. modulating [FC10]. Modulation [GQ13, Rem10]. Module [NWNL10, yWZhL14, SBM12]. Module-phase [NWNL10, yWZhL14]. moduli [DM11]. modulus [Caz14, KM14b, KA12, SPP12, SK10b]. modulus-based [KM14b]. Moerbeke [YL13]. MOICA [EYAD13]. moisture [KS13c, SSS13b]. Molai [Shi10]. molecular [CGVR14, FOZX14]. molecules [WTH^{+}14]. mollification [LF11c]. Moment [MIT12, PTK14, PJ12, CLR12, FT11, GT10a, GT10b, SPA12, Tag11, Tag14, XLS13a, ZZ14d, HRX14]. Moment-based [PTK14]. Moments [dACSS12, GT10a, GT10b, GT12, Hwa11, Hwa14, Kwo12, MS13g, PKK10, QQXY12, SU12, Tag11]. momentum [ASM11]. Monge [Gor10]. monic [ZY14b]. monitored [WW11h]. monitoring [GLSV10]. monkey [Zhe13]. monkeypox [BMH12]. Monod [AE11, Cai10b]. monodromy [AGR14b, DA11b]. monogenic [ADÇ14b]. monoids [KAÇ^{+}11]. monomiality [Fau14]. monomials [BP10a, VD14b]. monopolies [MS14b]. monopolistic [CN14a]. monosplines [Che13d]. monostability [WL10d]. monostable [ZL12a]. Monotone [Bog11, HLX13, Jan14d, LSJ11, PV10, YZX13, AMA12a, BBI13a, BBI13b, Bog12, BK11b, CZ14a, HZ10, Hou14, La 13, yL14c, Li10b, pLjH10, MDR14, MM10a, NW10, SMNN14, She14a, Sta14d, SLZ11]. Monotonic [Dar13, ESH10, DLL11a, FW14, GQ12, Guo13a, Guo13b, MCL14b, Sal11c]. Monotonicity [AMA12a, AR14b, Che10a, Seg14, CB12b, ET11, Koh13, KSD13, Kra14a, Mor10b, VC14b]. Monotonicity-preserving [AMA12a, AR14b]. Monte [ÁB14b, DJT10, Sie14, XLY14, ZZ14a, ZA14a]. Moore [CW11b, DD13, GH10d, HLT13, JC14, KPP11, MD11a, Rad11, SSCT14, SSU13, SG10, TSP10b, TS13, USAF14, ZCZP14]. MORAP [FYL13]. Morioka [EDYA14]. Morsean [ZC14a]. mortality [ZT11]. MOSOA [FS12]. most [GSN12, GG14g, LL11a]. most-obtuse-angle [LL11a]. motion [AAAH12, BW14, GSF12, HRX14, LXWW14, LZZZ10, PRT11, SWD11, Sol12, ZHY11]. motivated [Van11a]. motors [Ver11]. mountain [GS13a]. Movement [RY14a, Hua11a]. moving [ARH^{+}12, AS11b, AS14c, CLX13, GSSS13, HLZ14, LXTC13, Oss10, RMN14, RCH12, RPW14, RLV10, RRP14, SJR10, SBS10, WSCH14, ZL14c, ZD11a, dCCL14]. MPDPGA [UJH14]. MPSO [BSH13]. MRA [LZ11b]. much [NNPH14]. Muhammad [WZ13c]. Mühlbach [Ger12]. Mulay [XC13]. Multi [CSZ12, HL12e, LW11b, LL12b, MGYQ12, PD14, Pet14, SXCD11, jWtW13, WS13, XBYF14, Yu13a, ZZW^{+}12, AKP13, ABD^{+}12, ÁD10, BC13b, BBC14, BBF12a, BDG13, BDDK14, BC12d, CÁGGLP14, CGTCF13, CBTGWW14, CW11a, CLC^{+}13, sCWW^{+}14, CT12c, CVP^{+}10, CBRTA11, ĆARA12, DS14d, DN14b, DZS^{+}14, DE11, dCDNRH12, Din13b, DLS13, EM12, Ema13d, EYAD13, FTAB14, FYL13, FH11a, FC14, FBGH12, FS12, GC14, GMST10, GK10b, GGTCdlF13, Haj11, HL12a, HL12b, HL12c, HL13b, HB13, HLZ14, HCD14, JS14a, JSWJ13, JAG11, JZZ13, lJCWpW12, Jia13b, JYF13, JLW13, KLW14a, KB13, KDTSN12, KSFG14, KUKU14, KBG11, LPX12, LSW^{+}14a, LZWZ14, LLK13, LSL^{+}14, LNPS10, Lu12d, LL10i, Mao12, MC12b, MF13, MM13a, OGJLCCB14, PQRR10, PR14, QTYS14, RM11, Reu14b, SWL12, Sal11c, SKB14, SDL10a, Sha10b, Soo11b, Su13, Sun14, TNB11]. multi [TA11, UJH14, Wan10f, WTL^{+}11, WD14a, WLLY14, XDG12, XLLL14, XYL14, YWK11, YWLX11, YLAT14, YG13, YDZH14, ZW11b, ZHY11, ZL13b, ZL14d, ZWZ12, ZPZ14, ZHGP14, FM14]. multi-agent [LZWZ14, LSL^{+}14, WLLY14, XLLL14, ZL14d, ZPZ14]. multi-asset [KLW14a]. multi-channel [YWK11]. Multi-choice [PD14, BBC14]. Multi-class [ZZW^{+}12]. multi-constraint [TNB11]. multi-controls [DN14b]. multi-core [UJH14]. multi-criteria [CW11a, DZS^{+}14, DE11, YG13]. Multi-degree [LW11b]. multi-delay [LL10i]. multi-dimensional [SWL12, ZHY11, ZL13b]. multi-domain [FBGH12]. multi-element [DLS13]. multi-facility [JS14a]. multi-failure [ZW11b]. multi-fields [XDG12]. multi-group [LSW^{+}14a, XYL14]. multi-holed [dCDNRH12]. multi-innovation [HLZ14]. multi-instance [QTYS14]. multi-item [AKP13]. multi-layer [BC13b]. multi-linear [Sha10b]. multi-link [Sun14]. multi-manufacturers [DS14d]. multi-nonlinearities [JZZ13]. multi-objective [Ema13d, EYAD13, FYL13, FS12, GGTCdlF13, KDTSN12, LPX12, MF13, SKB14, FM14]. multi-overlapping [PQRR10]. multi-pantograph [BC12d]. multi-parameter [BBF12a, GK10b, ZHGP14]. multi-parametric [KSFG14]. Multi-period [LL12b, KDTSN12]. multi-periods [DS14d]. multi-point [CT12c, Haj11, HL12a, HL12b, HL12c, HL13b, Jia13b, JLW13, Reu14b, TA11, Wan10f]. multi-predators [SDL10a]. multi-product [Din13b, Soo11b, TNB11]. multi-products [CBTGWW14, JAG11]. multi-projection [LNPS10]. multi-quadric [GC14]. multi-regressions [JSWJ13]. multi-resolution [CVP^{+}10]. Multi-rogue [Yu13a]. multi-sample [MC12b]. Multi-scale [SXCD11, KUKU14]. multi-scenario [PR14]. multi-scroll [CGTCF13, OGJLCCB14]. multi-section [KB13]. multi-sensor [CÁGGLP14, SKB14]. multi-shift [Su13]. multi-solitary [WD14a]. Multi-soliton [MGYQ12, GMST10, HCD14, MM13a, WTL^{+}11]. multi-solitons [ÁD10]. multi-source [lJCWpW12]. multi-species [ZWZ12]. multi-stability [Mao12]. Multi-stage [HL12e, JYF13, KBG11]. Multi-step [Pet14, WS13]. multi-stripe [YWLX11]. multi-swarm [BDDK14, YDZH14]. Multi-symplectic [CSZ12, jWtW13]. multi-team [EM12]. multi-term [BDG13, YLAT14]. multi-therapeutic [HB13]. multi-threaded [ABD^{+}12, FTAB14]. multi-valued [CLC^{+}13, sCWW^{+}14, ĆARA12, FH11a, FC14, LLK13]. multiasset [XLY14]. Multibump [Zha14d]. multicast [SXGR11]. Multiclass [ZDH14, Niu11]. multiclassification [MIM^{+}14]. Multicolor [RGL12]. multicriteria [BPDZ12, SGJ10]. multidelay [MPS14b]. multiderivative [RP14a]. multidimensional [Bra13, CHC12, PG13, RY13a, RY14b, RY14c]. multidomain [EVDG12]. multifold [HLeH12]. multifractal [WSD13]. multifunctions [YSB11]. Multigrid [DS13c, YCS11, lL13c, MGC12, RBMA12, RGL12, ZY13a]. multigrid-based [RBMA12]. multigroup [CS11b, SS11i]. multilattices [CCG^{+}12a]. multilayer [NTdCJL12]. multilevel [DGM^{+}14, FOZX14, Hor10, Oss10, SF11b, WWZ14b]. multilinear [DFTQOB14]. multimaps [O’R12b]. multimodal [VC12, WWW13b, Zil12]. multinomial [Lip13, XFW^{+}12]. Multiobjective [GK12c, Kas12a, MT12a, AC14, Ant12, CGVR14, GW11a, GHLF14, GS14d, MP14c, MEBYG11, PN11a, TD13, Tri14, Ver13]. multiparameter [EDGC14, GHPR12]. multiperiodicity [HZW12]. multiperspective [TDSA14]. multiphysics [JHN11]. Multiple [AAT^{+}13, CV12, DLL11a, FV14, KS11a, Lin13a, LWZD14, MTT14, MR11, SZZ^{+}11, Sit14, Sta14b, TNB11, WF11, WHF14, Waz10c, WT10, Waz10d, Wei11b, WX13d, YWT14, YYY10, ZZW11a, ZLZ12, Amm14, BB11c, Che11g, Che14j, CS13f, DP10, DW13b, EWS10, FZLZ13, FP13a, Gao10, GLHCLP12, HH12, JS12, JHN11, Kan11, KBK13, KH14a, KKS12, xLhYlZ13, LXH13, LZL14a, LT12, LSL^{+}14, LZ14d, MLH11, MZ12, wMwWsT11, Mat14a, MPM14, NCÁHCLP13, NP10a, NSC12a, NCS13, NC13b, NC14, NZ12, Örk13, QZ10, QHZ^{+}13, RA10b, SG11, SA11, SCN14, SSTZ14, SAL12, Sha14c, SS10c, SS11c, Tad10, TJZ14, Van10, Ver12, Vod13, WL12a, WL13a, WW13e, WSY14, Waz12a, Waz12b, XSX12, XQ13, Yan10f, YX11, Yua12, Yun10b, gZrLX10, ZLS12a, ZLW12b, ZT14, ZTZ14, ZZY12, ZZX14, ZCS13, ZCY14, ZS14e, ZW14d, Waz12a]. Multiple-buyer [TNB11]. multiple-sets [ZZY12]. multiple-species [EWS10]. multiple-stepped [DW13b, WW13e]. multiple-vendor [TNB11]. multiplication [SK10a, Ste10h, SSB11a, SSB11b]. multiplicative [Cao10b, – Dor12a, GWM10, HL13d, KEH14, KLW14b, LLW10a, MDD12, MMRS13, WZH13]. Multiplicity [Aou11a, CT13a, FW14, FKH14, JL14, WW11f, XC14, CT12b, CLW14, DH13b, Gal11c, HL12a, NWZ13, TJZ13, WWS10, XCH14]. multiplier [CY11b, ES12a, SMBI14, WXA10]. multipliers [MK10b, SX14]. multiply [NMIA11, SMN11, San14]. Multipoint [PNPD14, GK10c, KKS12, PDN11, Pet13]. Multipower [dV12a]. multiprocessor [CDIS10]. multiquadric [JW12]. multiresolution [AY14]. Multiscale [qClLyW10, DC14, SXDJ11, SXC^{+}12, CLZ11, CGLM^{+}12, LWN14b, Sie11, SHZ10, YLLL13, ZYLS12]. multisemilattices [CCG^{+}12a]. multiserver [JU11]. multisoliton [jZzFbL10]. multispecies [CHC11, CDK13, LDL13, ZLJZ14]. multisplitting [Che10b, WWY10, ZZGL14a, ZZGL14b]. Multistability [HZW12, HZW14, NP14]. multistage [PM14b]. Multistate [LY11a]. multistationarity [Jos13]. Multistep [CC13, Ach11, Jat10, MC12c, Ram10a, SL10b, XZY14a, YZC14]. Multisymplectic [Cai10a]. Multithreaded [UJH14]. multivalent [CALR11, CKS10, CY11b, Dzi13, WXA10, WLX11, YL11b]. multivalently [SB12b]. multivariable [AŠA11, Dua10a, HSX^{+}11, JWW14, ZL14a, ZJD12]. Multivariate [AB12a, Akt14a, AT11, DGS11, GMGV^{+}11, KK13a, LMZ13, LZ11b, NTdCJL12, NA14b, Öza14, SMDZG13, ZZ14c, ZZ15]. multiwalled [SK10c]. Multiwavelet [LP13c, Kwo12]. museum [YLC10].

MUSIC [JKP14b]. MUSIC-type [JKP14b]. Mustafa [PK18]. Mutation [JYYJ10, Che13c, GCLL10, HAA^{+}14, LWZY11, TPA10, WWW13b]. mutilated [CGJ12]. Mutual [BPDZ12, GC11c, LDR10, Liu13a, MCWC13, ZjHO13, ZMC10]. mutualism [LWTC10, WW14d, ZLJZ14]. mutualisms [GZ14a]. Mutually [STH10].

N [GXF13, DLP10, HN10b, Ste10b, Ste10c, Tie12, ZA14b]. N.T [Pan12b]. nabla [DBK14]. Naghdi [FA10]. Nagumo [Bhr13]. nano [AV14, EAAN^{+}14, FHAL14, GAC12, YA14]. nano-fluids [FHAL14]. nano-ionic [YA14]. nano-sized [GAC12]. nano-Timoshenko [EAAN^{+}14]. nanobeams [AHDD13, EEM12, EMAM13, EHSM14, EKSO14, ZAA^{+}14]. nanobiosciences [SSZ11]. nanofluid [Das13, KRMM14, KML14, SHAO14, TSS^{+}14]. nanofluids [AHMS14, GSF12, HMD14, MA13b, PMV14]. nanoparticle [KRMM14]. nanoparticle-water [KRMM14]. nanorods [Cha13b]. nanotube [Nar12, SRA10]. nanotubes [Dan14, Kia14, SK10c]. NARMAX [Tsa13a]. narrow [ZL10a]. Narumi [Aza14]. Nash [RK14b, Soo11a]. Nasser [Pan12b]. Natural [Che10f, SK14a, ZhGQ14, AR14c, CNTG12, DW13a, DSPZ13, JDN14, KM11a, KRMM14, MMM12]. Navier [AL10a, CFZ14a, Che14h, CRW11, DHQ14, DAA13, Gal11a, Liu14c, LMB13, NR13, QMZ14, RW14a, SYK10, SH12a, SZ14c, TD14b, Wan10d, WH12, WZ14a, WWW12, ZZL13, ZLZ14c, dFGAN12]. navigation [LXTC13]. NCP [CLW12, HM12, Yu14]. NCP-functions [CLW12]. NCPs [LWW11]. NDDEs [RnZ10]. Near [MC12b, RD11, ZTX10, ARG14, CWW14, DLSS11, KRMM14, WLL14, ZW13b, BR14a]. near-collision [ARG14]. Near-exact [MC12b]. Near-field [RD11, ZTX10]. near-optimal [ZW13b]. nearest [YZ13]. Necessary [LS14a, Rag12, XLLL14, ZW13b, ZZJ14, Bes10a, CH13, – Dor13b, ZL10b]. Negative [DGMR10, AAAH12, ANR14, CLD12, Cve14, Gat14, GZW12, Imo14, NOS10, STM10, WWW13d, XC14]. negative-energy [XC14]. negaton [HLZ13]. negatons [ZCG10]. Nehari [yCfW12]. neighbor [PLTH12, ZZR11]. Neighborhood [SB12b, Mat14b, Ste12t, Ste13l, Tsa14, ZBM10, ZANA12]. Nekrasov [CKD12, CDDL13]. nematic [ZCF13]. NEPS [Ste12a]. Nesterov [AG14a, WYT13]. Net [WY14, HZ14d]. Nets [CFN12]. network [AOTV12, AOTV14, AM12e, BH12b, Che11c, CS13c, CZX^{+}14a, CZX^{+}14b, CC11c, CK12b, DM10a, EGO13, FO11, GRKP10, HYZH14, JJP^{+}12, JFHZ14, KAY12, KGO10, LJPJ12, LPJ^{+}12, LWP12, LWY14, LDR10, LY11a, LC12, LX11b, MP11b, MVS14b, NL12b, PR10, RG14, Sal11a, Sev13, SHWT14a, SSST14, SWLP13, TP12, TPTW14, Tsa13b, WiN12, WC14a, WC14b, WT11d, YCXL12, YYL^{+}13, YY14a, YLC12, ZLL11a, ZCL14b, ZWL14a, ZHW10, LG10b, MC14b]. network-based [GRKP10, WC14b]. Network-on-Chip [LG10b]. network-virus [YLC12]. networked [DdGL10, KH14a, LDR10, WY11b, WY12a, ZDC12]. networks [ASMA12, BGV10, BC13b, BC12b, BBI14, Ber14, Che11f, CZD^{+}14, CM10, CDIS10, DZZ^{+}13, DZZ14, ES12c, EVE12, FWZW13, FP13b, FA12, FW12, FOL12, GLY14, GWZ14, GJ14, GR11, GMA13, GYG11, GL13c, GSDW14, HI14, HYJ14, HLTT14, HLC13, HYYG12, HZW12, HZW14, IHM14, JJPW13, JKW^{+}11, JHP11, Jos13, KS11a, KfM11b, Ko13, KJW10, KPLC11, KLPC12, KPP^{+}14, LPJB12, LPL^{+}13, LYJP13, Li10a, LZZL10, LS12a, LX12, LYWL12, LLC12, LYC12, LY14b, LX14b, LW14f, LGCX13, LTT^{+}13, LLJF12, LSJZ13, LYXP13, LLKP14, LP13a, LLS11, MWJW10, MI10, MK14a, Mao12, MPSA14, MHS12, NTdCJL12, Ou10, PKP^{+}12a, PKP^{+}12b, PM11b, QLCF10, RV10a, RSPK13, SAMA11, SAL12, Sha11b, SD10, SZZH13, SZ14b, Sie11, Son10, SPWZ12, SIV10, SCLC10, SC13, THP14, TZ11, TLZZ12, TXX14, TJ12, TJ13, TJ14, WZ13a, WHW10, WMJW11, WPSW13, WZWW13, WW13b, WXWY14]. networks [WQ13, WYL11, WLZ12b, WLY^{+}12, XWMW12, XCT10, Xu10a, YZW^{+}10, YCS11, Yan12e, YLQ13, YC14, YH14b, ZX10, ZXCL10, pZJ11, ZaFT11, pZJ12, ZS13b, ZLSW13, ZAY^{+}14, ZW10d, ZCY14, ZLY11c]. Neumann [ÁB14b, CLZmF11, CNRY10, Gao14, GP10, JPM12, LY13a, LLC11, LL13e, NMIA11, SK12a, XCH14, XD11b, YZ12a, Yao10]. Neural [MC14b, ASMA12, BC13b, BBI14, BH12b, Che11c, CZX^{+}14a, CZX^{+}14b, CZD^{+}14, DZZ^{+}13, DZZ14, EGO13, FP13b, FO11, FA12, FOL12, GLY14, GWZ14, GR11, GRKP10, GMA13, GSDW14, HYJ14, HLTT14, HYYG12, HZW12, HZW14, JJPW13, JKW^{+}11, JHP11, KS11a, KfM11b, Ko13, KGO10, KPLC11, KLPC12, KPP^{+}14, LPJB12, LPL^{+}13, Li10a, LZZL10, LX12, LYWL12, LLC12, LY14b, LGCX13, LX11b, LYXP13, LLKP14, LP13a, MI10, MK14a, MPSA14, MHS12, NL12b, PKP^{+}12a, PM11b, PR10, QLCF10, RSPK13, SAMA11, SAL12, Sev13, Sha11b, SD10, SZZH13, SZ14b, Son10, SIV10, SWLP13, SC13, THP14, TZ11, TLZZ12, TXX14, TJ12, TJ13, TJ14, WZ13a, WHW10, WiN12, WPSW13, WQ13, WLZ12b, XCT10, YZW^{+}10, Yan12e, YCXL12, YH14b, YY14a, ZX10, ZLL11a, ZS13b, ZLSW13, ZW10d, ZCY14, ZLY11c, ZHW10]. neuron [SHWT14a, SHWT14b, WS12]. neuronal [BCNP14a, BCNP14b]. neurons [MWJW10, WMJW11, YWW^{+}14]. Neutral [Yue14, AKR14, ABLZ13, Ala14, BD12a, BMR10, BRP13, CDMS14, CZZ12, CZZ14, CW14a, CZZL13, CY12, DMW13, DX14b, EAAN^{+}14, FP13a, FH13, GLX13, HR10, HPB11, JVK10, Jan13d, JS13b, Kar13b, KTS11, Li10a, LL12d, LY14b, LJKU11, LY11b, LL10i, MI10, MM14b, MAD14b, Mil14b, MF14, MN10, RCS14, RHSC14, RSRA14, RN10a, SMN12, Shi14b, Ste12m, Ste13a, Ste14c, Sun11a, SHM12, VSM14, WZ11d, Wan13d, WFZL14, Wei11a, XSC11, YYJ14, ZW10a, ZXCL10, ZWZ12, Zha13c, dS10]. neutral-type [GLX13, MI10, ZXCL10]. neutron [Mar11]. Nevanlinna [Zhu10e]. Neville [Ger12, ADGP10, ACQOR11, Peñ14]. Newell [SMM14, XZ13b]. newsboy [Din13b]. newsvendor [PSJGLS14]. Newton [AKH13, AT11, ABGGS13, AMR13, AAB^{+}14, Ard13, AH11e, AH12a, AH12b, AG13, AH13a, AH13b, AM14f, AG14f, AM14e, AHM14, BRS13, Bur12, CSC10, Che13d, CG14e, Dar11, EALS10, EGH12, Fal14, Fan10b, GPK13, GG14d, GMR13, HH13a, HM12, HM14b, KL14b, Koh12, LP11, LZ10a, LWW11, LTM10, LLL10a, LHL11, Liu12b, LdS10, jLLW12, LV14, MA14b, PMC11, RA10b, SA13, SDK10, THDF11, VG14a, WZ14b, XY10, Yan10f, YCTD11, ZCU14, ZH10b, ZS14c, ZL10e]. Newton-like [ABGGS13, AH12a, Koh12, MA14b, SDK10]. Newton-Product [BRS13]. Newton-type [AMR13, AAB^{+}14, Ard13, AH11e, ZH10b]. Newtonian [CS12b, FMSC13, GL13a, HCCT14, KM11a, KK11b, Lac14, NA10a, SFF11, SWD11, VO12, VO14, WGS10]. next [KAÇ^{+}11]. NFEM [GS14b]. niching [VC12]. Nicholson [WD13, Wan14a, ZZQ13]. Nicolson [KMR14, LKS14, LLS13b, SW10a]. Nilpotent [CHX14, LW14c, MM10b, Pan11, SK11b]. nilpotent-Poincarè [LW14c]. nine [SBDB13]. nine-point [SBDB13]. ninth [GMX10, ZB13]. ninth-order [GMX10]. Nirenberg [Bah13a]. Nizhnik [Pan12a, THZF12, Waz12b, ZX11a]. NLMS [cLYrL11]. NLS [LZaY14, XX11]. NLSE [DW14]. NLSE-based [DW14]. nnd [Liu14d, XQ11a]. NNV [ZYC14]. no [AA12c, Par12, Yan13c, YZ14d]. no-slip [Par12]. no-wait [AA12c]. Nodal [RV13a, BC11a, BCM12, BCDI14, DH13a, KCZK11, PRV13, She12b]. node [CHF14, GAC12, LC11, TZZX11, WY10b, YS12]. node-based [TZZX11]. node-disjoint [CHF14]. node-foci [WY10b]. noded [Kha14c]. nodes [AOTV12, AOTV14, FWZW13, GMGV^{+}11, GL13c, LYJP13, LW14f, ZXCL10]. Noether [FAM11, FT10, Shi12]. Noise [SLYJ13, XWWF10, BLZZ14, Cao10b, Che12b, DRSG14, Gal11e, GC10a, KLW14b, LLW10a, LW13b, LCL14, MDD12, SHWT14a, SHWT14b, SPWZ12, dWcFwWjZ13, YLZ13, ZTX10]. noises [CDLZ11, WLP^{+}14, WZH13, Zha14e]. noising [SCC12]. Noisy [Ben14b, BLR10, Che14g, LX11a]. Non [ANR14, ADBT13, CV11, CSSSV13, dCDNVS11, dCDNRH12, EASK10, FP13b, FHWZ12, GWW14, KK11c, Li14b, MNOB10, OC10, Par13, RZ12, Wah13, Waz10f, YYJ14, ABG11, AAMAS11, AAMADH11, ASM11, ABE^{+}12, AC10a, ARE14, AIAD12, AHDD13, AJPZ14, BBI13b, BDG13, BY11b, Bie11, BLR13b, aBS14, BNP11, BK11b, BT10c, BPX14, Cai11, CMX14, CDR13, CDK13, Cha13b, CDLZ11, Che14k, CD11a, Chu11, CT13b, CS12b, CFG13, Cor11b, DLL11a, DBK14, DI13a, DP13b, fDxZxChT12, DGM^{+}14, DWP10, DL11, DLS13, DHMU14, EDGC14, EZ14a, ELS13, Et13, FV14, FNV10, FHW14, FG14a, FGS14, FMV14, FMSC13, FPTS14, GL12, Gal11d, Gin12a, Gin12b, GNS14, Gu10, GZW13, GS12d, GW12, GGOS12, GMV12, GG14g, HVD^{+}14, HK11, HZ14a, HCCT14, ICKV14, JB10, Jos13, JV13, JK14c, KM11a, KEGH13, Kan13, KAF13]. non [KS11b, Khe13, KS10a, KS12a, KM13c, Kov10, KRC10, KLW14b, KSS12, LOKB12, Lat14, LKY10, LYJP13, LL10c, LLW13, LW14d, LH10b, LWD10, LGN10, LNG12, LL10e, LM12, LHTN13, LD14, LO14, LZ14d, Luo14, MC14a, MAB14, MM14a, MM14b, MVQDE12, MS14a, MWAM14, MR11, MVS13, MG11b, MG14c, MYL^{+}10b, NNG10, NW10, NWZ13, NA10b, PSJGLS14, Par11, PP13a, PAC13, PFS10, PB14a, PMD14, QZS13, Qui10, RK14b, RHSC14, RSRA14, SMI12, SSP11, SFF11, SL14b, Sta14c, Sta14d, Sun10, TDZF14, eT12a, THP14, TS13, VPS12, VSS13, VS10, Vir13, WGS10, WF11, WCZ13, WDT^{+}13, WFL14b, Wan14b, WWK^{+}14, Wen10a, WYW14, WL10d, WLY^{+}12, XC10b, XC10a, YAW10, YWW13, YTC13, YZX13, ZCG10, ZCT11, ZC12, ZL13a, ZHXM14, ZC14a, ZhGQ14, Zhu12b, ZZZL14, ZD11b, ZWH14, SQJ11]. non- [Vir13]. non-abelian [ABG11]. non-absolute [Et13]. non-acute [aBS14]. non-analytic [LLW13]. non-autonomous [AIAD12, CDR13, GZW13, JV13, JK14c, Kan13, KLW14b, LM12, LD14, MAB14, MS14a, PB14a, QZS13, RHSC14, RSRA14, THP14, WF11]. non-availability [ZL13a]. Non-Boussinesq [EASK10]. non-central [EDGC14]. non-classical [AHDD13, BT10c, KS12a, MVQDE12, WDT^{+}13]. non-coercive [NWZ13]. non-compact [LHTN13]. non-concave [FNV10]. non-conforming [BY11b, DLS13]. non-constant [DWP10, DL11]. non-convex [FGS14, ZC12]. non-cooperative [RK14b]. non-cylindrical [KS11b]. non-Darcian [PMD14]. non-Darcy [KM11a]. non-decreasingness [eT12a]. non-definite [KM13c]. non-degeneracy [PP13a, PFS10]. non-degenerate [Gin12a, Gin12b, LKY10]. non-delayed [WLY^{+}12]. non-differentiability [FPTS14]. non-differentiable [BNP11]. Non-equilibrium [CV11]. non-expansive [DI13a]. non-flow [Jos13]. Non-fragile [ADBT13, FP13b, GWW14, RZ12, LZ14d]. non-Gaussian [CDLZ11]. non-Hermitian [CMX14, DGM^{+}14, FG14a, GW12, LW14d, WYW14, YAW10, Zhu12b, ZZZL14]. non-homogeneous [AAMAS11, AAMADH11, Cha13b, CT13b, DBK14, FMV14, GNS14, GMV12, HK11, MR11, VS10]. non-ideal [KAF13]. non-identical [LYJP13]. non-implicit [Gu10]. non-increasing [PSJGLS14]. Non-integrable [Waz10f]. non-interior [TDZF14]. non-isentropic [WFL14b, YWW13]. non-isospectral [ELS13, Sun10]. non-isothermal [VPS12]. Non-iterative [FHWZ12, Chu11]. non-Kerr [Lat14]. Non-linear [dCDNVS11, dCDNRH12, MNOB10, AC10a, ARE14, AJPZ14, BDG13, Bie11, EZ14a, GL12, HVD^{+}14, KEGH13, Kov10, KSS12, LOKB12, MC14a, MWAM14, MG11b, NNG10, PAC13, SL14b, ZHXM14]. non-linearly [Cai11, Cor11b]. non-Lipschitz [MM14b, WCZ13, ZWH14]. non-local [Gal11d, KS10a, SMI12, WL10d]. non-modal [SSP11]. non-monotone [BBI13b, BK11b, NW10, Sta14d, YZX13]. non-monotonic [DLL11a]. non-Morsean [ZC14a]. Non-negative [ANR14]. non-Newtonian [CS12b, FMSC13, HCCT14, KM11a, SFF11, WGS10]. Non-oscillation [Par13, Par11]. Non-oscillatory [YYJ14, FHW14]. Non-parametric [CSSSV13]. non-permutation [LH10b]. Non-polynomial [KK11c, fDxZxChT12, MG14c]. non-possibility [GS12d]. non-reflecting [HZ14a, LGN10, LNG12]. non-regular [Khe13]. non-reliability [GGOS12]. Non-resonant [Li14b]. non-respondents [DP13b]. non-selective [CDK13, Wan14b]. non-self [ICKV14]. Non-self-adjoint [SQJ11]. Non-selfadjoint [OC10]. non-semisimple [MM14a]. non-separable [MYL^{+}10b]. non-separated [CD11a]. non-simultaneous [KRC10, LL10e]. non-smooth [BLR13b, Che14k, CFG13, DHMU14, JB10, YTC13]. non-spreading [ZD11b]. non-square [LL10c, TS13]. non-stability [BPX14]. non-standard [Qui10]. non-stationary [ABE^{+}12, FHWZ12, Luo14]. non-symmetric [MVS13, Sta14c]. non-synchronized [WWK^{+}14]. Non-systematic [Wah13]. non-traveling [LWD10, XC10b, XC10a]. non-travelling [Wen10a]. non-trivial [NWZ13]. non-uniform [ASM11, Cha13b, PMD14, VSS13, ZhGQ14]. non-uniformity [ZCG10]. non-viscous [NA10b]. non-zero [GG14g, ZCT11]. Nonautonomous [YXO14, Dia13, EGO13, Far14b, GH10e, Hu13a, KN12, LS12b, PSS11, Sam11, Ste11g, WLQ^{+}10, XGZW13]. noncentral [YK11b]. nonclassical [SW14d, WWQ14]. noncoercivity [NMR14]. Noncommutative [YL10c]. noncompactness [AAH14, CDS13, DM10b]. noncomputability [GZB12]. Nonconforming [SP13a, Bra13, CF13, MJ13b, SZ11, SZS13, SZW14, XYC11, XSL10, YLS13, YJ14a, ZCZ14]. nonconservative [SRA10]. nonconstant [PR12b, RS13]. nonconvex [Ant11a, EES14, ZL11b, ZT13]. Nondecreasing [DS13a, ZL13f]. nondegenerate [ZYP13]. nondifferentiable [Ant12, TD13]. nondiscrete [AH11d]. nonexistence [Ste12e, Xu13b]. nonexpansive [AT12, AK11, BL11, sCCLY10, sCLCY11, sCCL11, sCLCaL11, CWL^{+}12, CLCZ12, sCWT^{+}12, sCCYW13, CLC^{+}13, sCWW^{+}14, CWWC14, CS10c, CMW10, CMM12, DHZ11, Gu10, Iid11, Inc12, KudKL14, KPK14, NS11, OS11, QCK10a, QCK10b, SK11a, She12a, SLZ11, pYX10, YZK12, pYmK13, YK12, ZOS10, ZS14a]. nonholonomic [GY13]. nonhomogeneous [AY11b, CV12, JW10a, LL14e, NTT10, Par13, PW14, Tua13, WW14b]. noninteger [MÇ12a, MRE^{+}12]. Nonisospectral [GPW14, LTG^{+}12, Xu10b, pZLC10]. noniterative [Jun10a]. Nonlinear [ALNGMGMM14, dCBdFM14, CH14, CN14a, Den10, dCDNCFRH10, dCDNV^{+}13,

– Dor13b, FGS14, HCL^{+}13, HCLP11, JLLD13, KS10c, KS12d, Ma11c, NA14b, NSZ^{+}14, NCANCF12, SHAO14, VR12, VS13, YS11a, ZS13a, ZY13a, ZW14c, ZYG14, ABRT14, AB11, AG10, ABEK11, AMB^{+}14, AOAA10, AKR14, ABP13, Afu12, ABT^{+}12, AZL13, ABLZ14, AS10, AZ14a, ADÇY11, Ala14, AH10, AHVR14, AS14d, Ans13, Ard13, AG13, AM14e, AS12b, AP10, AM12c, AE13, Asl11b, Asl12a, BRS13, Bah13a, BZ10b, BZ11, Bai14, BBJ10, BC11b, Bek10, BA10, BB11c, BAFP12, BAP^{+}14, BBI14, BTVC10b, BK14, BTS12, Bhr14, BK10b, BSS12, BRM11b, BR13, BWHQ14, BR10c, CL10a, CC10a, CGL12, Çak13, CB12a, CLTY13, CLCL11, Cha14b, CCJT11, CLL10, COPR10, CS11b, CLAT11, CJ11b, CM12a, CYM13, Che13c, CLW14, CXD14, CM14, CL14c, CYY14]. nonlinear [CZZL13, CL10b, CCGT13, CD13, CZL11, CL13b, CT13b, CS14b, CHMT11, CTV12, CST14, Cor11a, CYZ13, Cve14, DW12, DW14, Dar11, DMS11, DGDA10, DK11b, DX10, DW13a, DZZZ12, DKU13, Don14, DK14c, Du11, DR11c, DLL11b, DCR12, DPP11, Džu11, DPP12, ESV12, EAA11, EN11, ESH10, EZRMR12, EB14, EAAN^{+}14, EMM^{+}12, EZ13a, EALS10, EGH12, EZ13b, EKZ11, FHLW11, FZ13, Fan10b, FDHK14, Feb11, FcDDqZ10, FG14b, GLP11, Gal10b, GYL^{+}13, GLZ10a, GY13, Gao14, GC14, GLZ10b, GC11a, GZ13, GK10c, GYD14, GB10, GK13, Goć14, Goo13, GMA13, GQ13, GSGN11, Gu11, GZ12a, Gu13, GZ14b, GG14d, GM14b, GADK14, GMSH13, Guo14, GM14c, Gür10, Gür12, GMSE13, GMV11, Han11, HLZ12, HHU12, HR10, HKI12, Has11, HET12, HZ14b, mHhL10, HM10b, HZ11, HSZ14, HL13b, HH13a, HSLG10, HRH14, HM10c]. nonlinear [HRAA12, Hu12a, HLX13, HLZ14, HM08, HMH13, HZW13, HX13, HYW14, HBDC14, Hus10a, IISA10, Iri10b, Jan14a, Jan14b, JKN11, Jan13c, JSWJ13, JPB10a, JPB10b, JM11, Jha13, JH10, JX10, mJ10, JLHW13, JYF13, JYL14, JM10b, Juk13, JBDC12, gKW12, KJ10, Kar14, Kas12a, KK12b, KSFG14, KL14b, KA10b, KS10b, Kim10, Koç14, Koh12, KW13, Kot14, KL10, KAB11, KS12c, KK12c, Kud13a, KKS12, KET10, La 10, yL14c, Lat14, LZ12a, LP10, LJKP11, Lee12, LKP12, LW13a, LSJ10, LMMW10, LLY11a, LYZ11, LLJ11, LZW12a, LL12c, LZW13a, LD13b, LFR14, LX14a, Li14a, LW14e, LZL14b, Lia11, LYAH13, LKS14, LYD10, LTOX11, lLZ14a, rL11, LY12a, LNC12, LLL13, Lin13c, LTW10, LW10a, LLW10b, LLX10, Liu11e, LL11e, Liu11f, LJKU11, LW12b, LXC12, LL13f, LTU13, LWG13, LM13b, LH14b, LLS14, LSW14b, LC14b, LZSD14, LZ10f, LZ10g, LGKM13, LTR11]. nonlinear [LLN10, LL10h, LZX10, Lu12b, ML10b, MLH11, qMWjHpX13, MDR14, MA10a, MZ13b, MAN10a, MCM10, MN12, MMMMMG^{+}11, MS14b, MPS14b, MEBYG11, Mel14, MH11b, MGYQ12, Mer14, MAD14b, MR11, Mil14a, Mil14b, MH14c, MHB14, MJ12, MSC^{+}13, MT12b, MEG10, MM10c, Mol14, MEL13, Mur14, NTS11a, NTS11b, NMR14, NR14, NP10b, NSC12b, NS13a, NC13b, NCS14, NC14, NW10, NZ12, NMDWAS10, NK12, OM11, ONBN10, OGO14, PSP10, PMB12, hPrLgZ10, PRT11, Pan10, PTY13, PR12a, Par10c, Par10b, PT14a, PC13b, PZ14b, PID10, PDN11, Pet13, PNPD14, PD13, PEMS13, QaC13, QTL^{+}12, Ram10b, RGL10, RGL11, RS10, RP13, RKA^{+}13, RWB10, RY12, RA12b, RWBA13, RD14, RM12b, RMJM12, RZ11, RnZ10, RV13b, RSK11, RC12, SuR13, SMR11a, SGA13, SDAN14, SKSK12, SSZ11, Sha10a, SQ11, SS11c, SGG12, SA13]. nonlinear [SAP14, She10, SDL10b, SXL12, SP13a, Shi14b, SZW14, SDK10, SY10b, SAK10, SF13, SMH14, SKD12, Son10, SL11b, Ste12c, Ste12l, Ste12f, Ste12g, Ste12h, Ste12m, Ste12r, Ste12u, Ste13a, Ste13k, Ste13d, SAAS13, Ste13l, Ste13o, SS13d, Suh13, SS11i, zSWZW12, SWLP13, SH13, SCK14b, SSM11, SK10c, TA11, Thu10, hTlCqL10, Tsa13a, Tur14b, Val11, VSV10, VV10, Van10, Van11a, VG14a, Ver10, VDK10, VD14b, Wan10f, WGZ10, WGS10, Wan10c, WLMJ10, WT11b, WZ11f, WZ11e, WDZ12, WKLW12, WZG12, WSD13, WWG13, WWZ13, Wan13e, WSL^{+}13, Wan13d, WFZL14, WCL14, WLP^{+}14, WL14b, WYTL14, Waz10a, Wei11a, WGW11, WG12, WW14e, WDS11, WR10, Wu10b, WD12, WS13, XC10c, XWC11, XP12, XZY14a, Xin10, XGM10, Xu11, XZ12, XZ13b, XC14, YKI10, Yan11a, YP14, YZD10, YZL10, YW11a, YCTD11, YL11e, YL11d, YSZ12, Yan12f, YZC13, dYyR13, Yan13e]. nonlinear [Yan14b, Yao10, YZ11, YYJ14, YCK13, yYyL10, YDT^{+}14, YJ14b, YuRI14, YM10c, Yu13a, YL14d, YY14b, YLC12, YHS13, YXO14, YZZ^{+}14, Yun10b, Yun11a, Zah11, ZA11, ZSMB14, ZB11, ZA14a, ZL14a, yZ14a, ZH10a, ZTW10, yZhLjMzC10, ZL10b, ZW10a, ZSL11, ZL11b, fZwXyZ11, Zha11c, ZLS12a, ZLW12b, ZYP13, ZJ13, ZLWL13, ZL14b, ZWS14, ZZW14, ZQ14, ZxHW14, ZDH10, Zha10e, ZSHZ11, ZHZ12, ZZR14, ZL14f, ZZY^{+}14a, ZZH11, ZLH11, Zhe11, ZS10b, jZzFbL10, ZCC13, ZCS13, ZCL13b, ZCZ13b, ZXF14, ZS14e, ZGW14b, ZZG10, ZLW10, ZYZ14c, ZR14, ZL10g, Zuf12]. nonlinearities [CM12a, CLD12, CFCLM13, Dua11b, JZZ13, KRC10, Lee11, OO11, OO12, ÖP12b, Sin11b, WL13a, YMZS10, ZL10e]. Nonlinearity [ZS13a, AMB10, BBDS14, BZR10, BK10c, BT10b, BR10b, CYY14, DAA14, GWW14, Hay10, Hay11, IB10b, KW13, Lat14, NC13a, hPnZrL12, S.14, SB11, TB14, YZ10, yZhLjMzC10]. nonlinearly [Hil12, LLR10]. nonlocal [AS10, AP11, Aou11a, AY11c, AO12, Bai10, BAJK10, BRW13, BR13, Cha13b, CSSN11, CFCLM13, CYZ13, DT14, DN14b, EB12, EMAM13, FXL14, GV14, GQ14, Goo13, GAC12, HR12, IKL11, JLW11, Ji14b, Kia14, LLY14a, lLW14, MTT14, MO12a, MSC^{+}13, Nar12, NSYA14, NRL12, RYÖ11, SGA14, TGL14, Web10, Web12, Xu13b, ZAA^{+}14, ZL11a]. nonmagnetic [KMD13]. nonmatching [BA14b]. Nonmonotone [Gu11, PM11b, Qj10, WW11b, CD10a, qDG10, SWX10a, SA12b, WW13a, ZZQ10, ZC12, ZSQ10, ZZC13]. Nonnegative [WY14, GL10b, LZDS12, Lin10, LL11d, SR13, WLY11, WCZ11, WH11]. nonobtuse [KK13c]. nonparallel [QTYS14]. nonparametric [cLYrL11]. nonperiodic [LSLL14]. nonpolynomial [JM11, SJPS14]. nonprincipal [EZ13a, ÖZ10]. nonradial [Con13]. nonself [sCWT^{+}12]. nonsingular [BR10a, LWYY14, Wan14d]. Nonsmooth [Ant11b, Aou11a, DHM14, GS14d, Ilt11, LWH14, Nar13]. nonstandard [EM13]. nonstationary [WWY10]. nonstiff [CD11b]. nonsymmetric [DSZ14, GL10b, HKH10, LC13, Liu12c, dLfX14, WFwC12, Yin10, ZW14d]. Nontrivial [FHLW11, GP10, LLW10b, WL12a, YZ12a, CLW14, Che14d, DH14, Li11b, LWZD14]. Nonuniform [PCM14]. nonviscous [JA14]. nonzero [FP13c]. Noor [WZ13c, ARC14]. Nordsieck [BJ12, BJ11]. Norm [Ste12i, BA14b, CKD12, CDDL13, CD14, DC13a, FF11, GS13c, bJLyZlL14, Koh13, Kyr12, LWZD10a, LGZ14a, LL11b, LDL11, LQY11, LC14c, MM11a, RS10, Sol14, Ste10d, Ste10j, Ste10k, SS11f, SSB11a, Ste12w, WY10c]. norm-bounded [LDL11]. norm-relaxed [bJLyZlL14]. normal [Ada11, CMX14, DAV12a, DAV12b, JCW12, KM13b, KM14a, LPP^{+}12, LCCA11, MD12b, PHB14, PPMCG14, RTPH14, Sah11b, DAV14b, Ste12d]. normal-exponential [LCCA11]. normalized [BC10a, KSM11, QTL^{+}12, RDA14, Zhu10c]. normed [SA11, Sav10]. Norms [Ste10h, Ste10i, Boz13, İpe11, KSM11, SMDZG13, SC10a, ZJ14c]. Northern [NA14b, MSGN^{+}14, NCdCDM13]. Note [DISS14b, Fan14c, KE11, SM06, ABBR12, AAC11, AÇT12, Akt14b, ADP11, Asl10c, BDQR11, BD14c, BTRZ14, BCF13, CFL14, CL11b, CJZ12, CC14a, CZZ14, CZ14b, Che14h, CD10a, Cho11a, CK14a, DA11a, Dem11, Dem14, Den11d, DWP10, Elo13, Far14b, Fen13, Gal11c, GV13, HR11, HL11d, HMGC18, Jor10, KKA12, Kre13, Kud10a, Kud10c, KRS10, LR13a, LYS10, LLCX10, LR13b, Lu12d, MN12, MCdSS11, MG11a, Men12b, MC14c, MS13g, NR12, NN11b, NCS12, Pak11b, PM11a, Par10a, Pet13, PS11, Qui10, Rad14, RV10a, RHSC14, RC12, She14c, Shi14a, Shi10, Shi13a, SY14b, Sul10, TH14b, TTQ10, WC14b, Woź13, Wri11, WL13c, XQ13, XYL10, Yan12a, Yay11, YRV10, Zay11, ZLL10, Zha10a, Zha11b, ZC12, ZT13, ZS15, ZY14a, Zho12b]. Notes [EMS10b, Has10, KS11f]. notice [AZ19, EH15, Rho14b, Rho19]. Novel [ABDKA12, AS14b, Ara14, CZZL13, FCZ^{+}10, LzJZ12, LCLW10, MP14b, Sin11b, WWW13c, BP11b, DMW13, EYAD13, GWX12, GxFX^{+}12, GjLxF^{+}13, GC11a, Gen11, GGM^{+}13, HYZH14, HZZ14, JM13, JK14b, KSFG14, KGB12, LZGZ11, LHS12, LWG13, zM10b, NR13, ONBN10, QHZ^{+}13, RSG14, SEED13, SNH10, WLXX14, XWZF11, YZC14, YJCC14, YDZH14, ZLG^{+}14, Zhu10a, Tsa14]. Novikov [Pan12a, THZF12, Waz12b, ZX11a, ZZ10c]. NP [Rud12]. NP-hardness [Rud12]. NPV [Bas13]. NRAM [AM12e]. NS [TZZX11, NTPVNHLX14]. NS-PIM [TZZX11]. NS/ [NTPVNHLX14]. NSBM [AM12e]. NSC [And12b, And13b]. NT [LS11a]. NU [HMIZ13]. nuclear [KMH^{+}13, MÇ12a]. null [AAA13b, CFCLM13, qGyFzX13, LR13c, Shk14]. null-controllability [Shk14]. number [BOR14a, Boz13, CLTL13, CFZ14a, CL10c, DMP13, GGM14b, GG14c, HS11b, IS10c, LL12a, LZM^{+}12, LW14f, LCL^{+}13b, dLfX14, LMT11, MVRVSCMV14, PG13, Par12, PJ10, RW14b, RBT14, SLM14, WZL14b, XBC^{+}14, XH10, Yan13c, YGZW14, ZHC^{+}14]. numbers [Ara14, AAS14a, BSS14, BE13, CL11c, Cho11c, CK14a, Cvi10, DGMR12, DRS11b, EDGC14, EMS10a, EMA13b, FPP11a, GCMP14, GTU13, GQ11, İpe11, IA14, Kek11, KR13c, Li12a, LLCX10, Liu12c, dLL13, LS11c, MÇ12a, MS13d, NPA11, OS14, PP10, PW14, SC10a, SCH11b, SK10a, SK11b, SS10d, Sta11, Wan10d, WT12, Wło13, Xia10c, YT13a, Yor14, ZJ14c, BT12]. Numeric [AYS13, LCF^{+}14]. Numerical [ANHA14, ASM11, AAV10, AMA12b, AH11a, AT10a, AC10b, AS13a, AAS14b, Asa10a, Asa10b, BRS13, BCGK14, BB11c, BAE13, BLK^{+}13, BRVR12, BCKL11, Boy12, BA13, CGGMG13, CRR^{+}13, CFGRV13, Cha10, Cha13c, Che11g, CLAT11, Che13a, CSL14, CC14c, CLLS14, CL10b, CC10c, CS11d, CE13, DZLW13, DL12b, DGSB12, DZ12, DJ14a, DAA14, EB12, EZ13b, FA10, FdOdSM14, Fra10, GK14, GLS13, GC14, GS11a, GO13, GÖ14, GZXJD10, GZL14, GK12b, Gür10, Gür12, HYX14, Hsi10, HG13a, IKH11, JWZ^{+}12, JW12, JT13, JM14b, KLBR13, KP14b, KK11b, KBB12, KSS12, KK13d, LCD12, LH11c, LPPG14, LSF^{+}13, LMB13, LYL11, LL12f, MH13a, Mah13, Mam13, Mao11b, MVQDE12, MA13c, MHB14, MSR14, MJ11b, MJ12, MB13b, MHA10, Mos14, NSL14, NAA13a, NAA13b, OR11a, OH12, PPC13, PR12a]. Numerical [PAC13, RS11a, RA12a, RAIM14, RSN11, Res11, RB11b, SR14a, SSL10, SZG12, SMI12, Set14, SLC^{+}12, She12c, SWW14, SP14d, uiH14, STV12, Svá11, TSN13, TSS^{+}14, TRE13, WWZ13, WWZ14b, XRBZ13, XW13, XX14c, Yüz13, YS13a, Zei11, ZLZ^{+}14a, ZL14f, ZLYL10, ZX14, ZK10a, ZSD10, ZR14, ZK14, AMIR14, Abh14, ASS13, Ach10a, Ada11, ABJ10, AHJ14, ASVV10, AV14, Amd10, AK13, AHKS12, Aou11b, AJPZ14, Bar10, BMD13, BBF12a, BS12a, BB14a, BR14a, BVTC13, Bho10, BCC11a, BIT12a, Bur12, Cai11, CMM10, CGGMG12, qClLyW10, CCJ14a, CCG10, CLA^{+}12, CSD12, CLTA13, CN14b, CTY11, Civ13, CFCLM13, CRC14, DLL11a, DT10, DGCFM11, DFGL14, DC14, DFM10, DG10b, Dri13, DR11b, FYM13, FH11b, Fuy14b, GVW^{+}14, GRV13, GS12c, GT11c, GADK14, HVD^{+}14, Haj11, HBTU10, HHU12, Har10, HLMW13, HS11a]. numerical [HTX^{+}12, HM12, ITLA10, JMS13, KG10, KM12b, KKV12, Kha14c, KKH14, KS10b, Koh12, KV13, Kos13, Kos11b, KR10, KCLC13, LPR10, LCH10, LZZ11, LWZY11, Li12b, LK12, LL10d, Lia14, LNC12, LZ11c, LXZ12, LJ11, LHL^{+}13, LXLY14, MM10b, MS13e, MIHS14, MC12c, ML11a, Men11, MMCD12, MNP13, Mil14b, MVGS11, MO12b, MB14, dMbJmY11, MEG10, Mol14, MSJ14, MPW12, MV11, NRWF12, NCANCF12, OSRFF14, PB13b, PCT^{+}10, QHL10, RMY14, RK12, Ray12, RP13, Res13, RLV10, RRM12, RWkPk10, RBT14, Ryo10, SSL12, SM14a, SKC11, SFF11, SF13, SY12, SKD12, SGA14, TS12a, TK14b, TKC12, Tur14a, VMS10, VO12, WY10a, WiN11, WD13, WJ14, WTZ11, Wei14, WL10c, Xie11a, Xie12a, XWZF11, XZZX14, YTCY12, Yan14a, YLAT14, ZR13, ZK10b, dLDP13, Kos11a, KS12b]. Numerically [HW13a, JSF14, MLM^{+}14, Ser13, WMM14]. Numerov [YCF13]. nutrient [Ban14, Cai10b, DBS12, FHW13, GW13, SSA14a]. nutrient-plankton [FHW13]. NWF [MAD11]. Nyström [AST12, ASST11, FG11, FG13b, MMO13, RJ12, Tsi14, Zho13]. Nyström-type [Zho13].

object [KCGH14]. objective [ARS14, DWY14, Ema13d, EYAD13, FYL13, FS12, GGTCdlF13, KHD12, KB13, KDTSN12, LPX12, Mae10, MF13, dejW14, SKB14, TIT13, Ver12, Xu12a, Zil12, FM14]. objects [CL13a, LXZ14, BR14a]. oblique [PPC13]. Obrechkoff [Ach11, YNBV11]. observability [LWZD10b]. observable [WZ11c]. observation [DYYL13, LL10b]. Observations [Par10c, Par10b, HCLP11, NHCLP10, NCÁHCLP13, ZZ13b]. Observer [HL10, LX14a, WZH13, ZB11, ZLMG12, GWW14, HZ14b, JJPW13, KV12, YCZ14, ZSWS12]. Observer-based [HL10, WZH13, ZLMG12, GWW14, HZ14b, YCZ14]. observers [Abd13, MZXZ10]. obsolescent [WT11a]. obstacle [Ryo10]. obstacles [LXTC13]. obtain [Dos11, EZ14a, XG13]. obtained [ADÇ14b, JKP14a, Sol12, WP14]. obtaining [GGM^{+}13, RnZ10, VDK13]. obtuse [LL11a]. occurring [FP13b, SS11a, SSK10]. ocean [AWHW11, LXZ14]. oceanic [HB11b, LZZZ10]. Odd [WW14e, DLL14, KB11b, Rim13a, Rim13c, Ste13h, VD14b, YWCH14]. Odd-soliton [WW14e]. ODE [SQ11, BNBV10, GL12, GLV10, HM08, LL10h, Wei14]. ODEs [BIT12a, BIT12b, PRT11, TS12a, ZZH11]. off [AK10b, MT12b, MS12b, SL14b]. off-step [MT12b, MS12b, SL14b]. offers [TC14a]. offs [KMC12]. Ohmic [FD12a]. oil [MS10, YLK14]. Olami [ZYZ14a]. older [LD13a]. Oldroyd [Lee12]. Oldroyd-B [Lee12]. oligopolies [CMS13]. oligopolists [SVV13]. oligopoly [Ask14]. Olver [SHY11, Zay11]. on-line [ZZYX12]. One [AC13a, Boy11, BV12b, HKK12, Yan13a, Yig11, ABJ10, ANR14, AKB13a, AKB13b, Asa10b, Bac14, BKA12, Bel12, BF12, BIT12b, CSC10, DR11a, DGSB12, DC13a, DR11b, ES12c, Fan10b, GG10a, GV14, GLM14, Gol14, GL11b, GQ12, GYS13, ITLA10, JKN11, JZG13, JW12, Jos13, KA10b, KB11a, Koç14, KB11b, Kos11b, KR14, LH11a, LTL11, LXWW14, LWF11, LTM10, LL10f, MÇ12a, MIM^{+}14, MAO10, MB13b, MG11b, MG14c, MK14b, MPB13, PM14a, PM11a, Reu11, RX14, RH14, SATdV13, SdAT14, SDL10a, SPP12, TSN13, TZJZ12, TK12, VT13, WaZZW11, WZ11e, WL12b, WN12b, Wu10a, YS13a, ZF13a, ZC14a, ZCC13, Zhu11c]. one-center [MÇ12a]. One-dimensional [Yig11, ABJ10, Asa10b, Bac14, Bel12, Boy11, DR11a, ES12c, GV14, GL11b, GYS13, ITLA10, JZG13, JW12, KA10b, LL10f, MAO10, Reu11, SATdV13, SdAT14, TSN13, TZJZ12, VT13, WaZZW11, WL12b, YS13a, Zhu11c]. one-go [Koç14]. One-local [HKK12]. one-norm [DC13a]. One-parameter [BV12b]. one-parametric [RX14]. One-point [Boy11, Koç14, PM11a]. one-predator [PM14a]. one-prey [BKA12, SDL10a]. one-shot [DGSB12]. one-sided [LXWW14]. one-space [MG11b, MG14c]. one-step [BIT12b, DR11b, Fan10b, LTM10]. One-switch [AC13a]. one-versus-all [MIM^{+}14]. online [Che13e, ZLL11a]. onlookers [LWX13]. only [Don12, WLLY14]. Ono [Esf11b, hXqXlC10]. onto [SMN11]. opacities [MS13e]. Open [MC13, BBS10, GI13, Jos13, LZ13a, ZB14]. Open-loop [MC13, BBS10]. open-shop [ZB14]. operating [SR12]. Operational [KAG12, AG14c, BDD^{+}12a, BDGP14b, DGSB12, GP12, KYH10, LZ10b, PKBD12, Ray12, RP13, YHW13, YH14d]. operations [AS12b, Aza14, BLR13b, DD12]. Operator [GT13, GT11a, AS11a, AG14d, AAB^{+}13, Aou10, AY11c, BP11b, BDAD10, CKS10, CG14d, DN12a, Dra11, DB12b, GG11a, GPK13, GS11d, Gwi11, HT11a, fHcXH11, HM10d, HSC14, HP12, HKK12, JKP14a, Kay11, Koh12, La 10, LS10a, LjH11, LZLW11, LZW12a, Liu11a, LDW14, MO12a, MMKD14, OM11, PN11b, PH10a, PR12b, Ren14a, SS10b, SI11a, SMI12, SKM11, SK12b, Ste10j, SU10a, Ste10k, Ste11b, Ste12d, Ste12w, SKW11, TSG10, VCNV13, WZW12, Waz12a, Waz13, WWZ14b, XXS14, YP14, YL11b, Yan12a, YK14b, YL11f, ZS15, Zhu11a, dlRS14]. Operator-splitting [GT13, GT11a]. operatorial [Tri12]. Operators [MKSN13, AAR14, ABK14a, AR11, AGK13, AGKK14, AKS14, AKR10b, AKZ14, Alt11, ADT13, AM12a, AM13b, Ans13, Aou11b, AG14f, AB12c, AS11c, BAR11, BM12a, BM11, Bir13, BCC11b, BD14d, CMGR13, CMGR14, CMG14, CFS11, CCLF13, CD12, CBS12, DB11b, Deo12, DR10b, DM10b, DJ14b, EGOM10, Ere11, EDD11, Fer10c, Fer13, GG14a, GjLxF^{+}13, GGS13, GA10, GA11b, GY12, Gup13, GS14e, GRY14, GS11c, Gwi12, HAA^{+}14, Izg11, Jan11, Jan13b, JS10, JDD14, KPT11, Kar14, Kat14, Kir11, Kre13, LS10b, LZ10d, LLZ14b, LNPS10, LHTN13, LWN14b, pLjH10, MMX10, Man11, Man12b, Mat14a, MP14d, Mor13, MKK14, OC10, ÖD11a, RDD14, Ren14b, Sad11, SS14d, She14a, SKC10, ST14, SDO10, Ste10b, Ste10c, Ste10d, Ste10g, Ste10h, Ste10i, SU10b, Ste10m, Ste10l, Ste10n, SS11e, SS11f]. operators [SSB11a, SSB11b, Ste11e, SS11h, SS12c, Ste12i, Ste12v, SLZ11, Suc14, TZS14, TKLH10, TZJZ12, Tun14, VGA12, VRMT14, Wal10, WYZ14, WZY14, Yad14, YM10a, Yan11e, Yan12d, YHWZ11, YK13b, Yu12a, YD14, YTS12b, ZH12, ZLL10, Zhu10e, ZX11b, dV12a]. Opial [LH14a, MAO10]. Optical [BCWM10, LGKM13, GQ13, SMC12, WLMJ10, Yu13a]. optics [WW14e]. Optimal [AVF10, ALP11, BFH12, BM12c, CC12a, CCK11, CT13c, DN14a, EÖ10, FOZX14, GKC11, GG14c, HKI12, HLWH12, HB13, HLL14, JAG11, JWWY13, Jun12, KP14a, KK14b, KFM11a, KEH14, KLY12, LWS11a, LL12c, LHT14a, LB14d, LH14c, LF14, PM14b, RZHH13, SO10, San10, SMBI14, TS10a, THCW11, WWWJ14, WW11h, YGX14, YX12b, ZYYZ12, dFGAN12, ACCT14, BLR13b, Bot12, Che11a, CYW^{+}14, CCO12, CL10b, Chu11, CLND12, CL13b, CN14c, CLS12, CT11c, DN14b, DS13c, DPP11, FR11, Gal10c, GS14b, GFZ13, HY11b, HYC^{+}10, Ilt11, IISA10, IJ11, Isa10, JMB12, Kam14b, KBK13, Kes13, KBYL14, KP10, KKR11, Kuj14, LL14b, LXTC13, LY11a, LOD12, LLZ14c, LYB12, LMT11, MC13, Mak14, MS10, Men14, NP10b, NCS14, OYCCB13, PM14a, PG13, PPD10, Pet13, RZHH14, SGC12, SZH13, SS10e, SSK12, SD14, TDSA14, TC14a, Tsa13a, WZ11a]. optimal [WRZ12, WiN12, WaZW13, WY13b, WTZ11, WWC12, WW13j, YTS12a, YS10b, YH13, YCJ^{+}14, ZW13b, ZLH11, ZZ12a, ZCD10, MH14b, NTT13, BM12b]. optimal-order [WRZ12]. Optimality [CMGR13, GS14d, WTQU14, ZL13e, DGS14c, DS14d, Ver12, WWW13a, WYBT12]. Optimally [Liu12a]. optimisation [HS11a]. Optimization [AHS14, Ant13, BD12c, Civ13, EdCVB^{+}14, EH15, KTW10, KTT12, WYL11, YG13, ASS13, AG14a, AT10b, AC14, And13a, AM14f, ARS14, BHJ13, BMD13, Bas12a, BSR10, BHP13, BS14c, CBA^{+}14, CMP13, CSD12, CYW^{+}14, CHC12, Cho11b, CTY11, CFSYPP10, CvW10, dSCAM14, CSF12, CK12b, DWY14, DM13a, DYYL13, DQL^{+}12, Duk12, DHDB12, ES12a, EEA11, EH14a, ESE13, EL14, ESE14, ET11, FYL13, FOZX14, FOL12, FS12, GPP^{+}10, GWX12, GWM10, GW11a, GVW^{+}14, GGFM14, Gee12, GLV10, GCLL10, Gu13, GZ14b, GGTCdlF13, GZ12b, Gwi11, Gwi12, HVD^{+}14, HLeH12, HCW11, HjH12, HW12, HjH14, Hor10, HL12e, Hsi14, HWQG11, HSD11, HL14d, ICS^{+}13, Iid11, JSWJ13, JLL12, JWG^{+}14, JHX14, JLHW10, JLHW13, JW10b, JLTZ13, JYL14, JS14b, Jun11a, KHD12, KRJ11, KB13, KK12a, KA13a, KHP11, KS13a, KcZ14, KL14b, KBYL14]. optimization [KGB12, KAY12, KH14a, KZS12, KL13a, LFLS14, LH11a, LWQ10, LH11b, LWZY11, LF11a, LL12b, LSZ12, LL13a, LL14d, jL12, LKH12, LCL^{+}13a, LK11, LWF11, LDW14, LL14f, LJ11, LWX13, MYL10a, Mae10, MMM12, MT10, MLGdSC11, MMCD12, MHS12, dMbJmY11, MMM13, NMM14, PKD14, Pak11b, PSTL10, PTK14, PZ14b, PAC13, Qj10, RY14a, RP14b, RCSH10, RAZES13, RDFL13, Sal11b, SGJ10, SEED13, SW14a, SYl14, SKC11, SA12b, SFP^{+}11, THDF11, TIT13, TMJ14, TPA10, TPAB11, TQWZ10, TS10c, VTM14, VFGP13, VC12, WW11b, WZZH11, WYT13, WWW13b, WSY14, WZ14b, WR10, WYBT12, WTQU14, XBC^{+}13, XBC^{+}14, XZF11, Xu12a, Xu13a, XZZX14, YCXL12, YWYY12, YTCY12, YLD14, YLWH14, YZ14b, YHL08, YYW12, YDZH14, ZZQ10, ZLZ14d, ZHW11, ZSQ10, ZZC13, ZK10b, ZYM11, ZL13e, ZP13, Zil12, BSH13, CFG13, DS14c, For11, RY14a, SK14b, SDN14]. optimization-based [KH14a]. optimizations [UJH14]. Optimize [MZXZ10]. Optimized [DKK^{+}13, Kha14c, Ped13, ZZW^{+}12]. optimizer [HSD11]. Optimizing [CGTCF13, LWL12, MA14b, SR12, SYB13, SSG13, CEM11, LXH13]. optimum [B– D11]. Option [ECS13, CZ13, FCJ14, GO10, HLL14, LW13a, NN11a, RA12a, ZW13c]. optional [YWK11]. options [BP11a, CLX13, HL11c, KLW14a, MPW12, SY14a, Tch11, WWWJ14, Xia10c, XLY14, ZY13c]. orbit [ARC14, GLM14]. Orbital [yOZ14, IBS11]. orbitals [MÇ12a]. orbits [AFSMRL14, BY11a, Bel11b, CT11a, EDYSA12, fGyLJ11, LMY12, LM12, LR13b, ZT12]. Order [HMSW14, ABRT14, AB14a, Abd13, Ach11, AOAA10, AS11a, AC10a, Afu12, ABT^{+}12, ABLZ13, AZL13, ABLZ14, AS10, AISS13, ADÇY11, AJF13, AMS14, AT10a, Ali12, Alz14, AAB^{+}13, AZ10, ACCT14, AHDD13, AZ14c, jAzBS14, ANS14, Ard13, AAS14b, AKB13a, AKB13b, APJ14, ADBT13, Asl10d, AS14e, Bab12, BD12a, BDR12a, Bai10, BDL13, BDD12b, Bal13, BEP13, BC10a, BKP14, BDG13, BK14, BES10b, BTS12, BC10b, BD10, EL13, Bra13, BK12, BPX14, CST12, CT13a, ÇDO11, Çak10, CLZmF11, CBTGWW14, Car13a, CCJ14a, CF14, CDMS14, CS10a, CZM11, CT11a, CW11b, CLAT11, CT12b, CMH12, CM12b, CT12c, Che13a, CS14a, CLW14, CFZ14a, CSL14, CC14c, CLLS14, Chi10, CK14a, CGK12, CLND12, CL13b, CN14c, CLS12, CP10, CG12b, CT11c, CHMT11, CZC11, CIP10, CIN14, DH13a, DHI13]. order [DAV12a, DAV12b, DSGB12, DGSB12, DS12a, DSX11, DN12a, Dia13, Dia14, DD13, DP10, DJ14b, DAEY12, Dol14, DK14c, DFM10, Du11, DR11b, DR11c, DLS13, ESEEKH12, ESAH10, EHSM14, EA14, EDD11, EZ13a, EVDG12, FHLW11, FNG11, FFB13, Fer12a, FLP11, FKH14, Fig11, Fra10, FG14b, FPTS14, Fre11, Gal11c, GS13a, GB12, Gao10, GMX10, GL11a, Gao14, GS14a, GWS13, GWA14, GM14a, GK10b, GK10c, Gin12a, GPCCA14, GG12b, GK12a, GMR11, Gro13, GV13, GL13b, GK12c, HT11a, Han11, HLZ12, HL11b, HZ13, HH14, HR10, Has11, HET12, HXZ12, HSZ14, HL12b, HL12c, HL13b, Her11, HLTU12, HLT13, HZ12c, HQ14, HX14, HM08, HM14a, IU10, Iri10a, Iri10b, ISG11b, Jan10, Jan14d, JMB12, Jat10, Jat12, JM11, Jha13, JH10, KBAvBW14, KC12, KBK13, Kar13a, Kar13b, Kas12a, KK12b]. order [KfM11b, Kha14a, KK11c, KS13b, KA11, KKBO12, KA10b, KA13b, Kim10, KK13a, Koç13, KI11b, KB11b, KV13, Kon14, Kos13, KRC10, KS12c, KS12d, KR10, KM12c, LP12a, LP14, LW06, LZ10b, LMMW10, LSJ11, LS12a, LTG^{+}12, LSZ12, LZW13a, LZ13a, LY14b, LC14a, LSLL14, LH14a, LLL11a, LTOX11, LW11c, rL11, LNC12, LYL13a, Lin13c, LGN10, LNG12, LW10a, LLW10b, LY10b, LZZ10, LJKU11, LLC11, LW12b, LTW12, LHCI12, LJ13, LSW14b, LZB14, LT14, LFLH14, LZSD14, LP13b, LM12, LD11, LL10h, lL10g, LLS13b, LD13c, Luo14, LL12f, MG10a, MXH12, MYW13, MCH14, Ma14a, Mal13a, MMMMMG^{+}11, MIHS14, MH11b, MJ11a, MB13b, MX11, MMS14, MT12b, MG14c, MSS14a, MK10b, MD11a, MD11b, MD12d, MD13b, NN13, NN11a, Ném10b, NP10b, NCS14, NZ14, Nie13, NSYA14, NRL12, OO11, OO12, OYCCB13, OP14b, PN11a, PSP10, hPnZrL12, PSJGLS14, Par11]. order [Par13, PT11a, PC11, PC13b, PPD10, PD13, lQZWC14, QMM13, RSBK14, RDD14, Ram10b, RGL10, RGL11, RP14a, RK12, Rao14, Ray12, RP13, RHSC14, Rim13a, Rim13c, Rim13b, RTPH14, RSK11, S.14, SHT14, SAO12, SMR11a, SMN12, Sal10a, SDAN14, SSTZ14, DAV14b, SMIS14, Sha14a, SQ11, SS11c, SAP14, SHBN13, She12b, SLC^{+}12, SGY^{+}14, SEL14, SDL10b, SZS13, Shi14b, SDK10, SU12, SL14b, SLW12, Ste12k, SAAS13, Ste13h, Ste13e, Ste13f, SAAS14c, SW10b, SH12b, SW12, TMM13, TMM14, THDF11, TDZF14, TD14a, TA11, TA13, Thu10, TT13, THAB11, Tri14, Tun13, Tur14a, Tur14b, UOK10, Van10, VMRA14, WLH10, Wan10f, Wan10c, WCZ11, WCL11, Wan11a, WH12, WL12a, WRZ12, jWtW13, WC13, WXWY14, WW14c, WLLY14, WW14b, WYTL14, WYK11, WT10, Waz10e, Wei12, WGW11, WL10b, Wu10a, WW11f, WZ11i, WE13, Wu14c, XZZK12, Xia13]. order [XLGW12, XCH14, XQ13, XM10, XGM10, Xu11, XZ12, XCS13, YNBV11, YWT14, YL10b, Yan10b, YWLX11, YL11c, YZ12a, Yan13e, YKL14, YCZ14, YZ11, YT13b, YZZ10, YCF13, Yu11a, Yu11b, Yu13a, YWCH14, Zah11, ZB13, ZH12, ZL14a, ZFH14, ZW10a, ZF11, ZLW12a, ZLW12b, ZSWS12, ZYP13, ZHY13, ZZZ14, Zha14d, ZLZ^{+}14a, Zha13d, ZPZ14, ZZH11, Zhe12, ZH14, ZCS13, ZGY14, Zhu10d, ZLW10, Zhu11c, ZR14, ZL10g, dLDP13, WL11]. order-size [CLS12]. ordered [ANR12, ARR12, BB12, Bor12, ĆASH11,

– Dor13b, EVR12, EVR13, NKR12, NA14a, Rad14, RPA14, SSRA12, SK13a, SY14d]. ordering [CCO12, JGM13, LOD12, THCW11, ZZ12a]. ordering/trade [ZZ12a]. ordering/trade-credit [ZZ12a]. orderings [TCI12]. orders [JWWY13]. ordinal [HL12e]. Ordinary [SI11a, AB11, AISS13, ADÇY11, Car13b, CD11a, DK11b, DX10, DCR12, JM11, JS13a, KL14b, Kos13, LL12g, MSC^{+}13, Ort14b, Ram10b, RGL10, RGL11, Ser13, TMM13, TS14, TFMB12, VMRA14, Wan11c, Wan11a, ZH14]. Ore [ZB12]. organism [ZHH^{+}13]. organization [LWWL10]. organizational [LDR10]. organizing [FS12]. Orientation [CHT14, KUKU14, MRCD12]. Orientation-based [CHT14]. oriented [BSH13]. originating [GL11b]. Orlicz [CV12, CFS11, CT13b, Fer10c, MR11, NMR14, Sav10, SS14d]. Ornstein [GNdC12]. orography [WJ11]. ortho [WQ11b]. ortho-triples [WQ11b]. Orthogonal [XBC^{+}13, AAGR13, BPR13, BR14c, CT11b, CSMT12, DSGB12, DGSB12, DSM^{+}12, ET11, HMR12, HMPVQ14, Kra14b, Kuj14, LZ11b, LLL11b, MS12a, MMMMB13, MR14a, Mor13, PKK10, Pet12, RSG14, SMDZG13, SIDR11, Sta14b, WZ11e, WY12b, WLXL10, XRBZ13, XBC^{+}14]. orthogonalized [Sza13]. orthotropic [AAMAS11, AS14c, JE11, VS10]. Osada [HMGC18, ZS14d]. oscillating [ÇÇD10, MS13e, Med12, Mur11, Pan10, Yig11]. Oscillation [ABT^{+}12, BD12a, BDR12a, GWH12, HL11b, HR10, HZ12b, HM08, LML12, Par11, SQ11, Shi14b, Sta14d, ZW10a, ZZW13, AZL13, BD14a, BEP13, BBI13b, BK11b, ÇT10b, DF11, Has11, Jan14b, Kar13b, LLLZ11, LZW13a, MH11b, NKM^{+}10, Par13, SDAN14, SH12b, Vít14, ZGW14b]. Oscillations [SY10b, AR14a, BYM14, BZZ14, PT11a]. oscillator [BAFP12, BAP^{+}14, BCKL11, dCBdFM14, Cve14, EK11, EZRMR12, EZ12, EZ14b, Feb11, GjLxF^{+}13, MGCFMM13, Ped13, WKLW12, ZGH10, ZK11, DAV14a]. oscillators [CGTCF13, CGGZ11, EZ14a, FG13b, GL10c, Kov10, KRC10, MSC^{+}13, Oni11, PRV13, SZ14a, TGTCCGRG13, WLMJ10, WW11e, WW13i, YZC14, ZZW11a, ZWD12]. Oscillatory [ZX12, ABLZ13, AISS13, AAS12, Che14e, Che14f, DK14c, FHW14, FG14b, HMPVQ14, KX10, KX11, KW10, Kel12, KC13a, LWW10, LZLB14, MX11, RP13, Sha13a, XH13a, XW13, Xia14, XX14c, YYJ14, YCF13]. Oseen [ZYLS12]. osmosis [ZTF10]. osmotic [MSCK11]. Ostroumov [AS14d]. Ostrovsky [Par10a, OSRFF14, Yaş10]. Ostrowski [Alo13, FPTS14, GSGN11, HN14b, HHT13, LG14, Nie11, THYC10, Tse12, WV14, WW13j, Xia12]. Ostrowski-Grüss [HN14b]. Ostrowski-like [WW13j, Xia12]. other [BDK^{+}14, GLSJCBS10, MWJW10, NBR10, Sri11b]. out-flux [DMGK12]. outbreak [SC14a]. outbreaks [SRS^{+}13]. outer [MMS11, SPKS12, SP13b]. outerplanar [HLS^{+}14]. outliers [JSWJ13, MSGN^{+}14]. outperforms [ADGP10]. Output [LZWZ14, Abd13, DGS14b, FXD12, GC10b, HLZ14, KWJ13, MMM11, RLW14, TPT12, XSX12, yZ14a, ZLM12]. output-affine [Abd13]. outputs [LT12, MTSN14, TPT12]. over-current [TPA10]. over-relaxed [yL14c]. overall [LT12, ZF10]. overdetermination [IKL11]. Overlapping [GPZ12, CS10b, DGS14a, PQRR10, Sab13, Sab14, SR14b]. overrelaxation [Yan13d, YT13c]. overshooting [MSR14]. Overtaking [LYB12]. Oviedo [NCdCDM13]. oxide [LKP12]. oxygen [Mis11]. Ozaki [NS13b].

p [DR10a, ZT12, MZWZ12]. p-linear [ZT12]. p53 [ZCL14a]. Pachpatte [AY11a]. package [CZL11, RRSPTR12, RRSPTR14, dYyR13]. packages [jZzFbL10]. packaging [WKLW12]. packed [QS11]. packed-bed [QS11]. packet [AS14a, CBRTA11, HdCOL14, WY11b, WY12a, WZH13]. packets [LLL11b]. packing [KMW12]. Padé [CZL11, CG14e, GRV13, GT11b, Kuj14, TW14b, WZZ11, ZCL11]. Padé-type [TW14b, GT11b]. PageRank [AOTV12]. Painlevé [GGYW11, LWL10, ZGY^{+}10, CGK12, GP11b, GMP13, GB13, HVC12, SPV10, SGK11, SMH14, Wan10b, pZLC10, ZYC14]. pair [BT13, CWS10, FYM13, HP12, MKeM13, PH10a, RRG11, TI10a, Wan10b, WLY11, WTL^{+}11]. pairs [Cvi12, HA11, HKK12, Kau12, KI12a, MPC14, Ola11, Tsi14, TÜ12]. pairwise [GRKP10, MPS14b]. Paley [Wu14c]. pan [AA12b]. Panconnectivity [LWY14]. pancyclicity [LWY14, Sze10, Sze11, Sze12]. panel [Bas12b]. pantograph [AYS13, BC12d, GWH12, HG13a, LJ13, LL12f, Mil14a, Tri12, Yus10, Yüz14]. pantograph-type [Tri12, Yüz14]. paper [AÇT12, HLSS10, Ibr14, JS14b, Kre13, Mar14, NTdCJL12, Pan15, PK18, SQ11, Shi10, SM06, Yan12a, ZS15]. papers [MK08, MK10a]. para [MR14a]. para-orthogonal [MR14a]. parabolic [ALP11, BZ10b, BP11b, Bog11, Bog12, BR13, BH11, Bra13, CFCLM13, CG12b, CG14c, CYZ13, DNSM11, DE12, DM13b, FXL14, GGP^{+}13, GT13, GO12, Guo11, GZF12, GL13b, JM14b, Khe13, KS12a, Kwo11, LLS13b, MJ11a, MS12b, NMR14, SS13b, SL14d, SXDJ11, SD13b, zSWZW12, VS13, WaZZW11, WL12b, WW14b, Xia13, Xu14c, Yan12b, YLZ13, YS13a, ZH10a, Zha13b, Zha14e, Zhi11, ZWT12, ZS14f]. paradox [UW10]. Parallel [AMV12, CGO14, Ery14, GG12b, HL14c, Niu11, NNK14, QFCZ13, UJH14, WW11d, ZL13a, ZHW11, AT10a, BTT13, CEM11, CR11a, CS10b, CYW^{+}14, CvW10, FG11, GGOS12, Hou13, KBB12, lL13c, PW10, QQXY12, SSC11, SSN^{+}12, XT14, YZL10, YYL14, YYW12, yZxGyM10]. Parallel-machine [ZL13a]. parallel-series [BTT13]. parallelism [GL13b, HVD^{+}14]. parallelization [BCGV11]. Parameter [AŞK10, CLD12, DL13, GjLxF^{+}13, GS10a, GC10a, HZ12a, LD13b, SSL13, WzH10, YXZ10, Aha14, Akt14b, BC13a, BBF12a, BKW11, BN10, BV12b, CB12a, CPKT13, CYZ14a, CZ14a, CXD14, CR12, DEP12, DNSM11, DDWW12, EL14, EMP12, EH14b, FHR13, FA12, Gal11c, Gao10, GLX13, GGFM14, Gar11, GK10b, HHM10, HSX^{+}11, HdCOL14, HXZ12, Hil12, JLHW10, JM10b, KKBO12, KR11, KG13, KT14, KR10, KGO10, LPB10, LY14a, LWF11, MZXZ10, MYL10a, Ma12, MG14a, MTD10, Mea13, NMM14, Rec13, Rec14, RRLA14, RLW14, SS13b, SLW12, SNH10, TYCZ14, TK13, Tun13, WN12b, Xu13a, YK14a, YY14a, ZWS14, ZLY11b, ZHGP14, ZZ12b, ZZS14]. Parameter-dependent [HZ12a]. parameter-robust [KR10]. Parameter-setting-free [GS10a]. parameter-space [CR12, Rec13]. parameterization [GY13, KN13, Lip13, LLS^{+}13a]. Parameterizations [SHSC13].

Parameterized

[Kol14a, CZ14b, JCY10, LWYY14, LZ13b, Lu14c, SCC12, WL13c, ZW13a]. Parameters [LL13g, MGGS12, Aja11, AGHdD13, BMR10, BBC14, BAA13, CkPqL12, CK13a, DAEY12, – Dor12a, FG13a, FG14a, Gal11b, GS12b, GS13a, GxFX^{+}12, Her11, HR13, HHA14, JSWJ13, Jea13, JWWY13, KK11a, Kot14, LH10a, LfXfC13, MSGN^{+}14, MLG13, NOS10, NSNEV14, Oni11, OD11b, PB13a, QHZ^{+}13, RRRRTR12, SvB10, TKL14, WS12, WY13b, WLHZ10, XwK11, Yan10a, Yan11b, YS11b, YSI11, ZWD12, ZCL14b, ZZ10b, ZWL14a, ZANA12, TS10a]. Parametric [KS13a, KS13b, Lin14a, CCS13, CSSSV13, DRT11, DA14, DZ14, DPP11, GK10c, HSD11, KSFG14, Kol14a, KZ14c, ÖP12b, ÖTC11, RX14, TP12, WLMJ10]. parametrical [ZC14b]. parametrically [BZZ14]. paranormed [SK10b]. paraxial [qSyGH11]. parental [Liu14a]. Pareto [DGS14c, FM14, MF13]. Parity [BX12, EZ12]. parliamentary [Tsi11a]. Parseval [LZ11b]. parsimony [WY11a]. Part [LYLS14, MSC^{+}13, GP10, MCR14, Boy13]. Part-metric [LYLS14]. Partial [CXL12, LfXfC13, AAKB13, AM14d, Ans13, AAGR13, BP11b, Bho10, BJSS14, BR13, BX12, CHX13, CG12a, CDK13, Cha11, CYM13, CZZ14, CSL14, ĆSAV11, DN12b, ESEEKH12, EH14b, FCJ14, FcDDqZ10, Gal11e, GGP^{+}13, GMX10, GKC11, GMSE13, Har10, HLTU12, HLT13, HHM14, HLL14, Hua11b, Ign13, KLBR13, Kau14, KI11b, LYZ11, LDWY14, LSL10, LLS14, LFLH14, LXZ12, LZX10, MXH10, Mal13a, Man12a, MIHS14, MJ11a, MJPB11, MS12b, MSS14a, NR14, hPrLgZ10, PM11a, PB14a, RMY14, RDD14, SS13a, San10, SGC12, SSZ11, She12c, SSRA12, SÖCS12, TCI12, TT13, VR12, VD14b, WQX^{+}10, WDT^{+}13, WMM14, WWS^{+}14, WHWC14, Wu12a, XD11b, YHW13, ZLA10, ZHY13, ZHZ12, ZYSQ11, ZLZY12, ZSD10, dS10]. partially [ANR12, ARR12, BRP13, BB12, Bor12,

– Dor13b, MK14a, MVM10, Rad14, SSRA12, TLZZ12, Wu14b]. participating [VD14b]. Particle [BD12c, BSH13, JLTZ13, Kho14, TPAB11, WWW13b, WYL11, eMBHA13, AC14, BHJ13, BMMS13, Ber10, CYW^{+}14, CHC12, CTY11, CV11, DGCFM11, FYL13, Gad11, GWX12, GGFM14, GSF12, GO13, HjH12, HjH14, HCCT14, HWQG11, HSD11, JSWJ13, JLHW10, JLHW13, Kam14a, KRJ11, KGB12, KZS12, LCH10, LDW14, MHS12, MMM13, NMM14, PAC13, RP14b, SKC11, SK14b, SMC12, SWD11, SFP^{+}11, TIT13, TTWL13, WiN11, Wan13b, Xu13a, YTCY12, YZ14b, ZLZ14d, dCCL14, CFG13, SK14b, SDN14]. particles [HdCOL14, Kos11a, YD13]. particular [EMS10b, JS13a, Kal12, NRAV14, Ort14a, Ort14b, Reu14b]. Partition [Mat14b, HL13c, YRV10]. partitioned [KZ14a, LJCI13, Mon12]. partitioning [PTS11, TXX14]. partitions [CGO14]. partly [Bho10]. parts [WT11c]. Pascal [BS11c]. pass [GS13a]. passage [RRSPTR12, RRSPTR14]. passenger [JHZC13]. passive [JXK14, KEGH13]. Passivity [JKW^{+}11, LLC12, WW13b, DZZ^{+}13, FB10, GS13b, KPLC11, NLLC10, SAMA11]. Passivity-based [JKW^{+}11]. past [Afz10, BD14b, CRP13, Das13, EEEE05, Fig11, KS13d, Pan12b, Pan15, PP11, SBS10, Sha12b, SGC10]. Pasternak [HA12, KAF13]. patch [DWC13, Far14a, Mur14, WY11c, XTL10]. patches [YHL14]. patchy [LZTZ13]. path [AMS14, B– D11, COPR10, DMP13, LWY14, MT13, PR14, VLMHK^{+}13, Wan14e, Zhe14, ZSQ10]. pathogenesis [WDS10]. pathogens [BC14b]. pathologic [FdOdSM14]. paths [BB13b, CHF14, LC11, SN12]. pathway [MH11a]. Pattern [Fan14a, KfM11b, RDFL13, SZS^{+}12, Ban14, BWaY14, CV11, CSF12, DBS12, Gaz12, KUTY14, KR14, Sev13, SLYJ13, TZYH14, WWZ10b]. patterns [Gui14, GTX11, Hua11a, KRMM14, Liu14a, ZWLW10]. pay [AABV12]. pay-as-you-go [AABV12]. payment [BJSS14, KGC11, TC14a]. payments [AC13a, Sar12a, SGCG14]. payoff [KI11b]. PBC [LP10]. PC [CD11b]. PCA [AAM13, GP13]. PCALDA [Hua10]. pd [Liu14d, XQ11a, RZ12]. PDE [BA11, CFGRV13, KK12a, STV12, ZS14f]. PDEs [Sab14, Bal14, BB14c, BA14b, MK10b, Sab13, VDK10, ZC13b]. peak [WTFC13]. Peakon [FYYT11, HML10, HM10a]. peakons [yOZ14]. Peano [Che13d]. Pearson [BSKL13]. pedestal [ZCT11]. pedestrian [JWZ^{+}12]. Pell [KA13b, Li12a, WT12, YB12, dF14a]. Pempinelli [WM10, qYqLtW10]. Penalized [AAM13, AM14e, DU13]. penalties [WW11d]. penalty [Ant11a, Ant12, CBA^{+}14, Cop14, CFG13, EVDG12, jL12, dFPF14, QC12, QMZ14, SMWR10, ZY12]. penalty-finite [Cop14]. penalty-free-type [QC12]. penalty-like [CBA^{+}14]. pencils [PR11]. pendents [GG14c]. pendulum [LLLZ11, PPMCG14, ZHX11]. penetrable [LXZ14]. Penrose [CW11b, DD13, GH10d, HLT13, JC14, KPP11, MD11a, SSCT14, SSU13, SG10, TSP10b, TS13, USAF14, ZCZP14]. pension [AABV12]. penta [DSZ14, GK10c, Nem10a]. penta-diagonal [DSZ14, Nem10a]. penta-parametric [GK10c]. pentadiagonal [AKB13a, AKB13b, Elo11, Elo13, IA14]. perceptron [NTdCJL12]. percolation [CLTA13, ISH12]. perfectly [CL13a, JKP14b]. perforated [DC14]. Performance [CHK10, HSX^{+}11, LP10, LT11, MAD11, AT10a, BPDZ12, Che13d, CSSSV13, dCDNVS11, DQL^{+}12, JPK12, Kay11, KHP11, KPBC14, KPP^{+}13, LLP14a, LT12, RH14, SK14b, TK13, WHL14, XZF11, YY13b, YAAS10]. peridynamic [GHZ^{+}14]. peridynamics [WM13b]. period [BX12, FP12, Fuk12, KDTSN12, LL12b, Öca14, Özk14, QK13, Ste11d, SXH11, TLW11, Vil10, XGZW13, Xu12b]. period- [SXH11]. period-two [Öca14]. Periodic [CL11a, DSX11, DDWW12, GLZ10a, KK14a, KC12, LZSD14, LM12, Ma14a, SL14c, SHM12, WGS10, qWH11, Wan13e, WLQ^{+}10, YT13b, ANR10, Ach11, AKR14, AIAD12, AAS14b, AKMM12, Asl10b, BKW11, BSM14, CHC11, Çak13, CYHL11, CW12, CMH12, Che14d, CLL14, lChH14, CTLD12, CK13a, CSS11, CHS14b, DLL11a, DPC10, DISS14b, DN12b, DSH12, DLX13, DP10, DX14b, DZ14, DL11, EMK13, FW14, FNG11, FK13, GL12, GWZ14, GLM14, GWA14, GX12, GO13, GC11c, Han11, HV12a, HM10a, HRL14, HPB11, HLTT14, HL11e, Hus10a, Ito13, JŞ11, KS11a, KRS10, Li10c, Li10d, LX12, LHC14, LS10c, LDL13, LWTC10, Liu10a, LZ10e, LY11b, LW14h, LR13b, LO12, LDZ10, LL10i, LDLL10, LW10b, LDLM13, MXH12, MYW13, MCH14, MZ13b, MTD10, MHLL11, MAHAS14, NCL11, NC13a, PSP10, PZ14a, PR10, RDKU14, Ram10b, RGL10, RCYL10, SSF10, SvB10, Sha11b, She12c, SD10, SS13c]. periodic [Ste12o, Ste13o, SDS14, SXCD11, TRE13, WDLM10b, WF10, Wan11d, WqXlC11, WF11, WZ11d, Wan11a, Wan14a, WFL14a, Wei11b, WQ13, Wu10c, WW11f, Wu12d, XHG11, XM10, hXqXlC10, XLW11, YWT14, YMG11, Yan12e, Zha12a, ZL12a, ZS13b, ZjHO13, ZT13, ZLJZ14, ZWZ12, ZZW11b, ZZ12b, ZZQ13]. periodic-soliton [Wei11b]. periodic-wave [DPC10]. Periodically [SM12, Wei14, Yig11]. Periodicity [DMGK13, Ste11g, BX12, SAAS14a, YZW^{+}10]. periods [DS14d, Vil12, qWH11, XH11]. perishable [GKC11, LXZX13, San11b]. Peristaltic [JHA14, TPD10, Gad11, Gad14, Lac14, NA10a, RNEA14, VSS13]. peritonitis [NM14]. Permanence [CCWM13, DWC13, Hu13a, KN12, LTY11, LZ10e, LZTZ13, PZ14a, SY10a, WF10, WH14, YL11e, Far14b, HL11e, ZZJ14]. permanent [Ver11]. permanents [SY14b]. permeability [KMH^{+}13, MNOB10, SD13a]. permeable [AHMS14, DNI^{+}10, Fit12, Ish10, Kha14a, LWZ13b, RRP14, VV11a]. permissible [BJSS14, KGC11]. permitting [KBK13, TV13]. permutable [MPS14b]. Permutation [XZZK12, LH10b, LYG14]. permutations [Man10b]. Perov [CR14]. perpendicular [XK14]. perpetual [ZC11]. Perrin [Li12a]. Perron [CD12, DR10b, WH11]. Perry [LP12b]. persevering [Yua13]. Persistence [FL10, KX14, LD14, LS14a, VDL14, WW11a, YLC11, ZY14a]. Persistent [GS14c, SS12b]. personal [TIT13]. perspective [Xia10b]. perspectives [TPAB11]. Persson [WZ11h]. persymmetric [AdFY13, GG11c, GG14g, Rim13a, Rim13c]. perturb [LWZY11]. Perturbation [Das12, LYS12, LQY11, PE10, BS12b, BCWM10, CJ11b, CZ11c, CL10b, DMGK13, DLL11b, Dum10, ESEEKH12, EES14, EB14, Fer10a, Fer14, GB10, GML13, Has10, HCL^{+}13, HSW12, HMZ14, HHR13, LS14a, LYS10, rL11, Liu11c, LD14, Mar11, NK12, SS11a, SY12, WiN12, WaZW13, WWZ14a, XSL10, Xie11a, Xie12a, ZJLO14]. Perturbations [CDGR13, CZZL13, DZZZ12, DX14b, GMR11, GWH12, GS11b, HLW12, HMR12, KZ14b, LFR14, LCL^{+}13a, Lin13c, LW14h, qMWjHpX13, RM12a, SW14d, YC14, YH14b, Zha13c, ZZY14b, ZMC10]. perturbative [Boy12]. Perturbed [BDE^{+}14, AC10b, AK10d, AK10c, BP11b, BC10a, BLK11, CDR13, CPKT13, CH12, CG10, CG12b, Dar12, DN14a, Dem14, DL14, FKH14, FG13b, Gen11, GC10b, GO12, HV12b, HM10c, JPB10c, JZW12, KG10, KKV12, KCG10, KR10, KSS12, Kum14, Lat14, LL14d, LC14c, MVM10, PMC11, RS11a, RK12, Rep11, SRP13, She10, SPWZ12, VKV14, WWW13c, XH14, YS13a, yZhLjMzC10, ZYC12, ZLW12b, ZC13b, ZCZ14, Zho11b, ZXZ11]. perturbing [SL11b, WFW10, ZSZ12]. pest [GHC13, GC10c, LZG10]. pests [JDN14]. Petri [CFN12]. Petrov [DZLW13, DWZW14, DAEY12, RB11b, Ros11b, RV10b]. Petrovskii [FLW11]. Petrovsky [MDP12, Wu13]. Petviashvili [MA13c, SYZ10, AM12d, BR10b, Elb12, GB13, KA10a, Kud10c, LWL10, Liu10a, MTZ10, Sun10, WD14a, Waz10b, Wei11b, Wu10c, XC10b, XC10a, XZT10, XLZ10, Zha10a]. Petviashvilli [Zha10d]. Pfaffian [AM12d, MGYQ12]. Pfaffianization [MTZ10]. PH [JHZC13, QHYW14]. Phase [dCCL14, Ach11, AISS13, BC12a, DR11a, DP13b, FYM13, HYT14a, HY11b, HJZC13, Lin12, MSR14, Mon12, NWNL10, OR13a, Rem10, SAuAH14, SHBN13, SZW14, SWD11, SMWR10, TKC12, TSJ13, yWZhL14, WW13h, YLK14, Zei11, ZCL13a, ZCL14a, dOL13]. phase-field [YLK14]. phase-fitted [AISS13, FYM13]. phase-lag [Ach11, Mon12]. phase-type [HJZC13, SMWR10, WW13h]. phenomena [BSZ13, HW13b, LzJZ12, SBR^{+}12, WX13d]. phenomenon [KS13a, MSR14, SCLC10]. pheromone [SEED13]. phi [ZS10b]. phi-four [ZS10b]. Phillips [GG14a]. Philos [ABLZ14]. Philos-type [ABLZ14]. phosphate [QS14]. Photograph [Udw10]. photonic [BRVR12]. Physical [CCG12b, Apa10, IS12, Shi12, Sta12b, TPTW14, WY12b, YHS13]. physically [GGM^{+}13]. Physics [CCG12b, Pal12, BBB12, XZF11, CG12a]. physiological [MSCK11]. physiologically [PM14b]. Phytoplankton [CRC13, CPC11, EB12, JBW14, MAB14]. Phytoplankton-zooplankton [CRC13, JBW14]. PI [LP10]. Picard [AM13b]. Picone [Yos13]. PID [PPMCG14]. piece [XZY14b]. piece-rate [XZY14b]. pieces [DLL14, PTY12]. Piecewise [CJ11b, DR10b, KN13, RTPH14, BTS12, CMGR13, CLTY13, CD12, HXY12, KP14a, LSL10, LL10d, LR14b, LYL11, NRL12, PRV13, Sch11a, Sha14b, SL10b, Wan13d, Wan14a, WLL14, XH14]. piecewise-constant [CLTY13]. Piecewise-quadratics [KN13]. piezoelectric [AA13, Bas12b, LZ10c, RA10a]. PIM [TZZX11]. pine [GSC14, LL14b]. Pinning [SHWT14a, WLY^{+}12, LW14f, WXWY14, XWMW12, ZL13c]. pipe [CGLM^{+}12, LL14e, SZZ^{+}11]. pipes [CGLM^{+}12, DWQG12, NZW11, SM14a]. PIRKN [CD11b]. PIRKN-type [CD11b]. Piscounov [MDP12]. Piskunov [FLW11]. Pitaevskii [JZ10, Yu13b]. PIU [GK10a]. pivoting [Peñ13]. placed [ZA14b]. plains [WDBM14]. planar [ABG14, AGR14b, ARG14, AA13, CNRY10, CR11a, CDST14, Den11c, HHR13, LZM^{+}12, MZWZ12, pSbL14, TZYH14, ZZ10d, Zho12c]. Planck [WL10c]. Plane [CMS11, AJR12, Che11g, CLW11, Che14j, CDMT12, CTS14, ECG14, GMV11, LSX10, Li13, Mag14b, Mar12, Meh11, Shi12, SK11b, Sin11a, Ste10b, Ste10c, SS11h, Vod13, XK14]. planes [Rec14]. planetary [Med12, TRE13]. plankton [Ban14, CPC11, DBS12, FHW13, Gaz12, LCH12, SSA14a]. plankton-nutrient [SSA14a]. planning [BFH12, JAG11, MT13, VLMHK^{+}13]. plant [BC14b, LWWL10, XGZW13]. plant-pathogens [BC14b]. plants [YH14e]. plasma [GGYW11, YLS13]. plasmas [AMB10, Awa12, BZR10, BZGK10, SB11, WFL14b]. plastic [ASR^{+}14, BHG10]. plate [AA12d, AM13c, BD14b, BH12a, CW12, DNI^{+}10, FFOC14, HCCT14, KAF13, Mey14, PC13a, PMD14, RRP14, SBS10, SGC10, YXZ10]. plates [AS14c, CCS13, DB11a, DKU13, EJ12, GAC12, HA12, JE11, RMN14, THNTNXR11, WW13f]. plating [LNdCJV14]. playas [WDBM14]. played [QGD14]. players [CN14a, DLP10, FXD12, PPK14]. plotting [PBP14]. plus [BS12c, Boy13]. ply [CALR11]. PM [GAV14]. PM-spaces [GAV14]. PMC [PLTH12, TL14]. Pochhammer [SÇK14a, SC14b, Zuo10]. POD [AHJ14, LLS13b, Luo14, SJ12]. Poincaré [ARG14, MPB13, PBG14, LW14c]. point [AKR10a, ANR10, AKN11, AN– D12, ANR12, ANR13, AG14a, ARR12, AS10, Alz14, AZ10, Aou11a, AKR11, AAS14b, Bab12, BZ11, Bar10, Bas12a, BCK14, BCC11a, Boy11, B– D11, Buo10, CT13a, CSQ11, CJZ12, CL14a, sCCLY10, CIVS14, CT12c, Che14c, CZ14b, Cho11b, ĆARA12, CJRV12, – Dor13b, DPP11, Džu11, DPP12, EGOM10, ECG14, EVR12, EVR13, EKZ11, FFB13, FHK13, FPTS14, GIR10, GS12a, GH10a, GK10a, GFMP13, Gar11, GC11a, Goć14, GAV14, Haj11, Hao10, HWL11, HL12a, HL12b, HL12c, HL13b, HZ14c, HZW13, HA11, HP12, Iid11, ICKV14, Inc12, Jha13, JCY10, Jia13b, JLW13, JS14b, KR12, KI12a, KS13b, Koç14, yL14c, LPAE12, LPB10, LZZ11, LSJ11, LWYY14, LZ13b, LZ14a, LZ14b, LNC12, LLW10b, LS11a, LLS^{+}13a, LLK13, LS12c, dALORPGC14, Lu14c, MAD14b, MNP13, MC14c, MT12b, MM10c, NKR12, NA14a]. point [Ols12, PMV14, Pan11, PM11a, PC13b, PL10, PPD10, PID10, PK12, PD13, Pre11a, Rad14, RS11a, Reu14b, RRP14, RPA14, SN12, SKB14, She14a, She12b, SY13a, She14c, SSRA12, SR10, SSA14b, SIR14, SR14c, SK13a, SY12, SFP^{+}11, SBDB13, TYCZ14, TZZX11, TZYH14, TA11, Ver10, Wan10f, WHC10, WCZ11, WHL11, WDZ12, WYT13, WZ14b, WZY14, WL10b, WZ12, WL13c, XL11, XG13, XL14, YZK12, Zah11, ZS14a, ZZ10a, ZS10a, ZL10b, ZLX12, ZW13a, ZGZ14, ZGZG10, Zha13c, ZC14a, ZM14d, ZM14e, ZH14, Zhu12b, ZZZL14]. point-forces [FHK13]. point-to-point [SN12]. points [ARN10, BS10c, CALR11, CHX14, CDMT12, ĆASH11, ĆSAV11, DMS12, DI13b, DV12b, DV13, DV14, DLSS11, Dos14, EHRV12, FZLZ13, fHcXH11, Kan11, KM12a, KM13c, KPK14, LAR13, yLbLjH10, MM11a, Mur11, Mur13, Ola11, PH10a, PBP14, RRR11, SRP13, TLDZ13, Wag10, Wei14, Wu12a, XM12, YK12, ZX11b, Zou12]. Pointwise [JDD14, Łen11, Len14b]. Poisson [CBO12, CY12, DAHAK10, DG14b, GS14a, GDH10, GGOS12, HY11a, Imo14, KAK11, LK14, RN10b, SPWZ12, TDT12, WYL10a, dWcFwWjZ13, WW11g, XC14, Yan11c, YWW13, YWL13, Yue14, ZFH14, ZL13b, Zit11, LGZ14b]. Pol [GjLxF^{+}13, KK14a, dCBdFM14, EK11, LXWW14, PS14a, RP13, ZZW11a]. Polak [Li14a]. polar [HMZ14, LNG12]. polarity [HLH12]. pole [OÖ13, Yan11d]. pole-clustering [Yan11d]. police [Mis14]. policies [JGM13]. policy [CCO12, JU11, LWS11a, LHT14a, LOD12, LB14d, MB10, MB11, OYCCB13, PM14a, SGC12, WZ11a, WW11h, ZW11b, ZZ12a]. pollutants [LPY11, SS11a, SS10g]. polluted [Cai10b, LC14e, LH14c, LF14]. pollution [LZG10, VJBDV13, WL12b, XLL11, ZW14c]. Polyak [Li14a]. polydisperse [NBG10]. polygon [LMT11]. polygonal [PQR14]. polygons [DLL14, SSV11]. Polyharmonic [BLR10]. polymers [MO12b]. Polynomial [AE13, BHP13, CCJT11, DMS12, GPCCA14, LLC11, SHT14, YLD14, AY14, AAGR13, BDE^{+}14, BC12c, BC12d, BS12c, CCCRR13, DG13, DG14a, DM11, fDxZxChT12, FF11, GMX10, GMGV^{+}11, GGM^{+}14a, GM14a, GLV14, GGÖS11, GS14f, HSW11, HVC12, Hu13b, Hu14, IB10c, JBL^{+}11, KK11c, KK13b, Kot14, LPH11, LL10d, LM11, MTER14, Mil14a, MPM14, MG14c, Pan11, PP10, PB13b, PMM11, PC14b, QHZ^{+}13, QFCZ13, SIDR11, Sas12, SGY^{+}14, TS10b, WQ11b, Wan11c, WW14a, WLLD13, WTQU14, XH10, ZTD10, ZLJ11, ZS12, ZZ10d, Zho11c, Zho12c, ZGY14]. Polynomials [Kud13a, MS13d, AAC11, AYS13, AŠA11, Akt14b, AG14c, AT11, ÁNA12, AG14e, Ara14, AGLR13, BC12c, BSS14, BB14b, BPR13, BR14c, Cao10a, CS13b, CLR12, CLLS14, CK14a, Chu10, CT11b, CSMT12, DGMR12, DS11, DAEY12, Dua10a, Dua10b, Dua11a, EDG14, EMS10b, EMA13b, EMA13c, Elo14, Els12, ET11, EDD11, Fan14d, Fau14, Fin14, GG10a, Gal10a, GI14, GÖS11, HXY12, HLPW14, HMR12, HMPVQ14, Içö12, Ign10, ISG11b, KÖ11, KK11a, KASY10, KYH10, KKA12, KAG12, KR13c, KP13, KK13a, KK14c, Kud12, Kuj14, KS11f, LW11a, LWN13, LMS11, LLLS11, LLSW12, LLLS13, LS11c, LM11, Ma11b, Mah14a, Mah14b, MS12a, MMMMB13, MS13c, MR14a, Men12a, MJ13a, Mor13, MKSN13, OÖ13, OO10, Öza11, Öza14, Ozd11, OS14, PE10, PSSS14, Pet12, PAA14, Ram14, RSG14, SMDZG13, SA10b]. polynomials [SI11b, Sim11, SC12, SC14b, Sta14b, Sza13, TKS11, WLZ11, WY12b, Woź12, XQ11b, YC12, Yüz13, ZZ11a]. polyphase [SHSC13]. polytopes [AT11, Gue14a, Gue14b, KKVCS12]. polytopic [ZFCW13]. polytropic [SJ13, WGS10]. Polytube [SLBS11]. Ponderomotive [dOS13]. Population [UJH14, AEESMA10, AAA13b, BBDS14, Bel11a, BC14a, CZL14, Far14b, GK11, GKV12, HS14a, HS14b, INE10, Koy12, KK13d, LS13a, LL14b, LWS11a, Li12c, LH14b, LB14d, LTSV14, LL10i, LH14c, MYJ10, MZW14, MCL14a, MDR14, MG12b, NNG10, PM14b, RWkPk10, RGA14, SVGC10, SL14a, SS12a, SGK14, SM14c, SS12b, WZ11d, YT11, YJ14b, YZX13, ZWZ12, ZY14a]. populations [CDK13, DIAQ13, Gal11d, RTVV11, WY11a]. pore [CS12b]. pore-scale [CS12b]. porosity [ABBC10, VDFV12]. porous [All11, AV14, ABK^{+}14b, AA12d, AM13c, BD14b, CRX12, CS12b, DAA14, Haj11, HFS12, HMA10, HCCT14, KM11a, Kha14a, PC13a, PMV14, Pan15, PMD14, SM14a, Sha12b, SZZ^{+}11, SD13a, XLL11]. portfolio [BO13, CFG13, JWG^{+}14, KKR11, LQY10, LL12b, RZHH13, RZHH14, TQWZ10, ZZYX12]. portfolios [OGO14, YLC14]. portraits [OR13a]. posed [AG13, FLMR14, GPK13, GHS11, gKW12, Liu12a, LW13d, SA10a, VG14a, YLLL13, ZM12]. posedness [AO12, BW14]. position [CWCW11, Cor11a, DJY13, DKK^{+}13, EAAN^{+}14, Rud13, WLLY14, XZY14b, Yan10e]. position-based [CWCW11]. position-dependent [XZY14b]. position-only [WLLY14]. Positioning [OSRPC12, XWWF10]. positions [CKH11]. Positive [AZ10, AKB13a, AKB13b, Bai10, CHC11, CMH12, CM12b, DSH12, DX14b, FCJ14, Goo13, GK12a, HL12c, HL13b, HL14a, Hus10a, Jan14e, JW10a, KB11b, KM12c, LZLW11, LLJ11, LS10c, LY11b, MYW13, MCH14, MCL14a, NCL11, ÖA13, SHN10, WZ11d, Wan11a, WLW14, WZ11i, XZ12, YP14, YMG11, Yan12f, Yan14b, fZwXyZ11, ZGZG10, ZSHZ11, Zha12c, ZZW11b, ZCL13b, ZLW10, AGT11, AMA14a, AS14a, BC13a, BBDS14, CC10a, CJZ12, CMX14, CGTCF13, CK13a, Cve14, DP10, DWP10, DKMM11, ESH10, Fen13, GX11, Goo11, GC11c, GW12, GZF12, GFZ13, mHhL10, HL12a, HL12b, HLTT14, JXA11, JLW13, Kar13a, KSM14, xLhYlZ13, LXK13, LW14d, LZL14b, LZKK11, LJKU11, LC14b, LDZ10, LL10i, MXH12, MCdSS11, MKK14, PEP14, RSK14, Sta12a, SAAS14b, Ste14e, Sza13, Wan11d, WD14a, WJ14, WWY10, WYW14, Wu10a, WHG13, XX14a, XM10, Xu13b, YAW10, YW11a, yYyL10]. positive [ZL10b, ZLW12b, ZG12, ZWS14, ZGZ14, ZF13a]. positive-definite [GW12, ZGZ14]. Positivity [AMA14a, SHA12a, AGKR10, GKS10, HS11b, JK14a, KV13, MDP12]. Positivity-preserving [AMA14a, SHA12a, MDP12]. positon [HLZ13]. positone [AZ10, Zha12c]. Positons [ZCG10]. positron [Awa12]. possessing [WLZ11]. possibility [GS12d]. post [RY12]. post-nonlinear [RY12]. posterior [SLM14]. posteriori [ABE^{+}12, Bac14, CL14c, CYZ14b, Kim12, LC14c, ZF11, ZZL13, ZCZ14]. postural [ZLC^{+}13]. Potential [AK10a, All11, DSX11, Fer12b, GW11b, GB13, HCD14, IS10a, IS10b, Kam14a, LG10b, Li13, LL13f, NWZ13, PP13b, RNZH13, Sin12, TW11b, WSD12, XC10b, XC10a, YT13b, YS10b, ZTZ14, dCCL14]. potential-Yu [HCD14]. potentially [ZWL14b]. potentials [AGM12, CT14b, GM10, HMIZ13, IS12, Kri14, ML10a, WW13d, ZD11b]. Potra [MA14b]. Potter [Rad11]. Poussin [BSM14, EKK14]. Power [AAR14, MP14a, NAHH11, WN11, ANR13, AMFM12, AMB10, BAO14, BZR10, BK10c, BT10b, BPM10, BR10b, Bor14b, Boy14, CCJ11, CPKT13, CYY14, CT13c, EMP12, FKH10a, GJGSNRÁ12, Hay10, Hay11, IB10b, JLLD13, Kek11, KKBO12, KRC10, KT14, LdS10, jLLW12, Nov14, PC13a, PCM14, PP11, Sta11, SB11, TL10, TLL13, Tom13, TB14, Van11b, VVK12, VKV14, WW13d, XC10b, XC10a]. power-law [FKH10a, Hay10, Hay11, PC13a, PP11, Tom13]. Powers [GG11c, AKB13a, AKB13b, BS11a, Fen13, GG13b, KI10a, KB11b, Kuj14, ÖA13, Rim13a, Rim13c, Rim13b, SMDZG13, Wu10a]. PR [Don12].

Prabhakar [GGPT14]. practical [Ach13, Don12, HD14a, Kel12, Men14]. Prandtl [Par12, RNEA14]. pre [NBYHV11, YSB11]. pre-analysis [NBYHV11]. pre-strong [YSB11]. Precise [TM11, JKP14a, TK14b]. precision [AR12b, BBB12, DAV12a, DAV12b, KPBC14, DAV14b]. Preclusion [FW12]. preconditioned [DGM^{+}14, GK10a, GPZ12, GHZ^{+}14, HL13d, Jia13a, KZ14a, LZZ11, LW14b, LZ13b, RW14a, WWT13, wX13b, YAW10, YZ12b, Yun11b, ZSQ10]. preconditioner [CJZ12, HH12, LV14, LHL^{+}13, WHW13a, ZGZ14]. preconditioners [GWLZ10, HN10a, JCY10, Lu14c, NL10, She14c, WCN11, WL13c, Yin10, ZZ10a]. preconditioning [CR11b, LHZ^{+}11, RW14b, YH14a]. predation [SKC14]. predator [AKS11, BLZZ14, BWaY14, BLR13a, CHC11, CCK11, CJK12, CXL12, CWLC13, CCWM13, CZ11c, CD10b, DLL11a, DRSG14, DMGK12, DMGK13, DWP14, DSH12, DWC13, DL11, FW14, Gui14, Hu13a, HL11e, Jan13a, JXA11, JSH10, KG11, KJ12, Kha14b, LR11, LG10a, LW13b, LL13b, LXC14, LS10c, LDL13, LZ10e, LY11b, LZTZ13, Liu13a, LC14e, Liu14a, LB14c, LLP14b, LZLY14, LB14a, LDZ10, LYP14, LY14e, MCWC13, MHXyY14, MB11, MM13d, PM14a, PLF13, QLWW12, SP14a, SP14c, SL13a, SKC14, SZS^{+}12, SHM13, TW11a, WF11, WQ11a, WW11a, WWW13d, WFL14a, Wan14b, WH14, WZSW14, WC14c, XTL10, XZ13a, YT11, YLWU13, YLC11, Yan13c, YTC13, YL14a, ZLY11a, ZL11a, ZT14, ZJLO14, ZW14c, ZMC10, ZWLW10, ZM10c, Zuo13, dlHV12]. predator-prey [BLZZ14, BWaY14, CXL12, CWLC13, CCWM13, CZ11c, DMGK12, DMGK13, DWP14, DSH12, DWC13, DL11, FW14, Gui14, Hu13a, HL11e, Jan13a, KG11, Kha14b, LW13b, LL13b, LXC14, LY11b, LZTZ13, Liu13a, LC14e, LB14c, LLP14b, LB14a, LYP14, LY14e, MCWC13, MHXyY14, MB11, MM13d, QLWW12, SP14c, SL13a, SKC14, SZS^{+}12, SHM13, TW11a, WF11, WQ11a, WWW13d, WFL14a, WH14, WZSW14, XZ13a, YT11, YLC11, YTC13, YL14a, ZLY11a, ZW14c, Zuo13, dlHV12]. predators [BKA12, Liu14a, SDL10a, ZMC10]. Predication [Wan13b]. predicting [Kam14a]. Prediction [HB11b, SFL^{+}11, CH13, GXF13, JFHZ14, LNdCJV14, Lin12, RRSS14, SMC12, Tie12, Tsa13a, UMMT14, WS13, YLD14, Jun10a]. prediction-correction [JFHZ14]. prediction-corrector [YLD14]. Prediction/Correction [Jun10a]. predictions [MNP13]. Predictive [LKP12, RGA14]. Predictor [Gal11e, CTV12, DGSB14, HZW13, LZL14a, LLL10a, LW13d, ZW12, ZZ11b]. Predictor-corrector [Gal11e, CTV12, DGSB14, HZW13, LW13d, ZW12]. predictors [Jat10]. Preface [BDLY10, CCG12b, CCC11, KS11c, KS12b]. Preference [LZLS14, HL14c, LIX14, XKH10, ZF10]. preferences [WHI12]. prefixed [CCJT11]. preinvex [Mat14c]. Preissmann [jWtW13]. preload [CLD12]. premium [Kar10]. prescribed [CLTL13, Cra12]. presence [AHMS14, BD14b, BM12d, BM14c, BCNP14a, BCNP14b, CRC13, Che11d, CTH14, Dol14, KK12b, PV10, PBP14, Sar12a, Sun11b, WiN12, ZW14b]. present [Fig11]. presentation [SÖCS12]. Preservation [DLSS11, TYCZ14]. Preserving [AGKR10, AMA12a, AMA14a, AR14b, AR12b, ADT13, HLW13, HHA14, KV13, LWZZ14, MDP12, SHA12a, Sok13, VC14b, WM13a, Xie11a, Xie12a]. pressure [ABBC10, Che14h, CE13, dCDNV^{+}13, LL13d, MWY13, SMR14]. prestarlike [Dzi13]. pretreatment [MZL07, Pak11a]. prevalence [Mis14, YLC12]. preventive [WZ11a, XH13b]. Prey [CD10b, AKS11, BKA12, BLZZ14, BWaY14, BLR13a, CCK11, CJK12, CXL12, CWLC13, CCWM13, CZ11c, DLL11a, DRSG14, DMGK12, DMGK13, DWP14, DSH12, DWC13, DL11, FW14, Gui14, Hu13a, HL11e, Jan13a, JXA11, JSH10, KG11, KJ12, Kha14b, LR11, LG10a, LW13b, LL13b, LXC14, LX14c, LS10c, LZ10e, LY11b, LZTZ13, Liu13a, LC14e, Liu14a, LB14c, LLP14b, LZLY14, LB14a, LDZ10, LYP14, LY14e, MHTH12, MCWC13, MHXyY14, MB11, MM13d, PM14a, PLF13, QLWW12, SP14a, SP14c, SDL10a, SL13a, SY10a, SKC14, SZS^{+}12, SHM13, TW11a, WF11, WQ11a, WW11a, WWW13d, WFL14a, Wan14b, WH14, WZSW14, WC14c, XLS13a, XTL10, XZ13a, YT11, YLC11, Yan13c, YTC13, YL14a, ZLY11a, ZL11a, ZT14, ZJLO14, ZW14c, ZMC10, ZWLW10, ZM10c, Zuo13, dlHV12]. prey-predator [BLR13a, CCK11, CJK12, DRSG14, JXA11, Liu14a, PLF13, Wan14b, WC14c, Yan13c, ZT14]. prey-predator-resource [LZLY14]. prey-taxis [MHTH12]. Price [San11b, BP11a, GKC11, GO10, MPW12, San10, San11a, San11c, SMS14, SS13c, TC14a, WLYC13]. price-discount [TC14a]. Price-sensitive [San11b]. prices [And12a, Xia10c]. Pricing [CBLC12, FGRLSV14, LWL14, SY14a, ACG14, CFGRV13, CLX13, CZ13, DWLL10, FCJ14, HL11c, KLW14a, LXZX13, MA11d, NN11a, Siu10, Soo11a, Soo11b, STV12, Tch11, WT11a, ZW13c, ZY13c, ZC11, Zim14]. Prim [WZZ14]. primal [Ach10a, AG14a, CPOdCN11, ZGY14]. primal-dual [Ach10a, AG14a, ZGY14]. prime [FP12]. primitive [LMB13, MGMM14]. Prince [JSSB14]. Principal [NAA13b, ÖZ10, ÖP12b, AMA12b, DE12, DF11, EGMPMRBC12, EZ13a, JK14a, Kel12, LWM10, LWW10, MÇ12a, MTR13, MTER14, NAA13a, ZLZ14d]. principle [BZGK10, CWL^{+}12, CXD14, Din11, Fau14, HSI13, HZ12b, HR12, Ilt11, LP11, WLZ11, YLAT14]. principles [FKS13, HT11a, MJPB11]. priori [AHRV11]. priority [Dim13a, Pet13]. prism [BS14a]. Prisoner [ES12b]. Privalov [Ste10l]. Privalov-type [Ste10l]. probabilistic [BDH^{+}14, CIVS14, CSSSV13,

– Dor13b, HS10b, LPL^{+}13, Liu10d, NBG10, Pal13, RSPK13]. probabilities [BRP13, CYG14, Gat14, HL13a, ZCZ^{+}13a]. probability [CZX^{+}14a, CZX^{+}14b, ESAH10, Gho14, Ign13, LW06, Mam11, MMRS13, MIT12, SO10, SDC10, SKIH12, SSS13a, YX12b]. probes [For11]. problem [AE11, eMBHA13, AAV10, ABJ10, Aja11, AGM13, ARL13, ARG14, AANA10, AC10b, Ami14, AK10b, ACCT14, BMHM14, BC11a, Bai10, BD11, BC13a, BU12, BPT12, BA11, BGRS11, BM13b, BKW11, Bel12, BH13, BMMS13, BCK14, BY11b, BRGP14, BD10, Bot12, BAJK10, BH11, BA13, BT10c, CT13a, CDS13, CV12, CNRY10, CRG11, Car13a, CC12a, CLR12, Cha10, sCCL11, CLC^{+}13, CMP13, Che10c, COPR10, Che11g, CLW11, CGRS12, CLF13, CS13d, CF13, CLL14, CT13b, CvW10, Cop11, Cop14, CFG13, CL14d, Cor10, Cov11, DA11a, DJ12, DNSM11, DJ14a, DWC11, DM13b, Din13b, DKU13, DGOR13, DHZ11, DX14a, DS13c, DG10b, Dri13, qDG10, DCL11, DCR12, DHQ14, DFG14, Dum10, EdCVB^{+}14, ES12a, EEA11, ECG14, Ema13d, EA14, EG06, ELS13, FYL13, FXL14, Fan10b, FF11, FFB13, FT11]. problem [FR11, FIK14, Gal10b, GS12b, Gal11d, GK10a, GHLF14, GI13, GS11a, GO12, GD11, GFZ13, GL14b, GSRR13, GAC12, Gus12, GT10a, GT10b, HMA10, HL12a, HH11, HW14b, HDT11, HY11b, HLS10, HYC^{+}10, HG13b, HS10b, HHDM13, Iid11, IBS11, JS14a, JMY13, JLL12, JWG^{+}14, JCW12, JLHW13, JJL14, JFHZ14, lJLpWxZ14, JSF14, JM14b, JM10b, Kan12a, KMW12, KAÇ^{+}11, KR12, KS12a, KR11, Kle12, KB11a, KC13b, KCBH14, KP10, Kuj14, KPW11, KZS12, KH14b, LCH13, LWM10, LL10b, LSX10, LSJ11, LF11c, LLJ11, Li13, LL13a, LXTC13, LW14e, LH10b, LY11a, rL11, Liu10d, LTM10, LLW10b, LLW11, Liu11a, Liu11b, LMS^{+}12, LW13c, LTU13, LYG14, LWZD14, LIC12, LPPG14, LfXfC13, LYB12, MFZ12, MZ13a, Maa14, MB12, MV14, MIM^{+}14, MP14c, MVM10, MEBYG11, MKSC13, MO12a, MR11, MO12b, MQ14, MMS14, NTT10, NB11, NH12, Ni13]. problem [NSYA14, NA10b, OO12, OH12, ÖBT10, OD11b, PTY12, PTY13, PN11b, PR14, PH10b, PT11b, Plo14, QW11a, QK13, QW14, Qui10, RS11a, RFK12, Rud12, RS12, SK11a, Sal11b, SK12a, SSC11, SSCA13, Sev11, She12b, SEL14, SWW14, Sit14, SM10, SK13b, SSV13, SSK12, SD14, Tag11, Tag14, TNB11, TDZF14, TSP10a, Tem12, TK14b, TDT12, TRE13, TTQ10, TTQ11, Tua13, Tun13, VFGP13, Wag10, WT11b, WW11e, WaZZW11, WL11, WYY11, WZZL12, WW13a, WaZW13, WSY14, WTH^{+}14, WTZ11, Wu12c, WCW^{+}13, XX11, XL11, XLL11, XYC11, Xie11a, XM11, Xie12a, XLGW12, XCH14, XYL10, XGM10, XY11, XZ12, YTS12a, YS10a, YGL11, YZK12, YS10b, YH13, YH12, YLC10, YCJ^{+}14, Yua13, Yun10a, ZD10, ZOS10, ZA14b, Zha11a, Zha11c, ZLW12a, ZLZ12, ZW13b, ZZW13, ZZL13, ZB14, ZZGL14b, ZCZ14, ZxHW14, ZW10b, ZGZG10, ZZY12, ZY14b]. problem [ZW11c, ZYLS12, CP12c, Mat14b]. problems [AMIR14, AB11, Abh14, ACZ12, Ach10a, Ach11, AS10, AP11, AMS14, ARH^{+}12, Ali12, AT10b, AAS12, AB12b, ASST11, AC14, AZ10, And11, And12b, And13b, Ant12, jASzZ11, jAzBS14, jAS14, AG13, AM14e, AAS14b, AY11c, AO12, AMA10, Bac14, BCGK14, BZ11, BDR12b, BBC14, Bar10, Bas12a, BC10a, BS12a, BS12b, BCC11a, BLR13b, BRAS12, BA14a, Bog11, Bog12, BDK^{+}14, BRW13, BR13, BV12b, Buo10, CCMV12, CS12a, CJZ12, CCJ14a, CT12a, CKSV14, CF14, sCCLY10, CW11a, CYZ14a, Cha14b, CSLL10, CTLL11, CT12b, CT12c, CTA13, CS14a, Che14d, CZ14b, CL14c, CW11c, CHC12, CZL14, CQ14, CS10c, CD11a, CXG10, CEXG12, CZL11, CTY11, Civ13, CG10, CG12b, CK12b, DH13a, DZLW13, DWZW14, DWY14, DR11a, DN14a, DR10a, DS12a, DH13b, Din11, Din12, DZ13, DW11, DR11c, DHMU14, EAA11]. problems [EDAM10, EGMEM14, EGOM10, EES14, ET11, Eto11, FTAB14, FHLW11, FL13, FKS13, FLMR14, FNV10, FNG11, Fer12a, FKH14, FNVLPD^{+}13, FBGH12, FOL12, FS12, Fra10, FGO12, GT13, GX11, Gao14, GP10, GWS13, GWA14, GC11a, Gen11, GQ14, GM13a, GYD14, Goo11, GPW14, GK12a, GKW14, GHS11, GS14b, GS11b, GK12b, HJVC14, Haj11, HN14a, Han11, HLZ12, Hao10, HH14, HAA^{+}14, fHcXH11, HZ11, HCW11, HW12, HCL^{+}13, HMMK11, HL12b, HL12c, HL13b, Her11, HS11a, HWPY14, HL12e, Hsi14, HLX13, HN10a, HLW13, HM14a, IS14, Ilt11, Inc12, ISG11b, JYCY10, Jan11, Jan13d, Jan14c, JMB12, Jat10, JSMA11, Jha13, JY12, JL14, Jia13a, bJLyZlL14, JCY10, JW10a, lJCWpW12, Jia13b, JKP14b, Jun10a, Jun11a, KG10, Kam14b, KCZK11, gKW12, Kar13a, KB13, KK12a, KA13a, KKV12, KM14b, Kes13, KDTSN12, KK11c]. problems [KS13b, KKBO12, KM13c, Kim12, Kim14b, KAY12, KZ12, KPK14, KR10, KSS12, Kum14, KCLC13, LCD12, Len14a, LWQ10, LTL11, LZZ11, LF11a, LPX12, LL14d, LWYY14, LZ13b, LZ14a, LZ14b, LKH12, LFC^{+}12, Lim12, LNC12, LZFR10, LLL10a, LHL11, LLC11, Liu12a, LW13d, LO12, LNPS10, LZ10f, LZ10g, LHTN13, LWN14b, LS12c, Lu14c, lL10g, MXH12, qMWjHpX13, Mae10, MT10, MVQDE12, MR13b, MS13e, MF13, MC14c, Mit13, dMbJmY11, MT12b, MAHAS14, MJ13b, NMR14, Ném10b, NTPVNHLX14, NW10, NL12b, NN19, OO11, OS11, OPD12, PN11a, Pak11b, PSTL10, PP10, PT14a, PC13b, PL10, PMC11, PK12, PZ11, QCK10b, Qj10, RSBK14, RY14a, Ram11b, RY10, RK12, RA10c, Reu11, Reu14a, Reu14b, RD14, RAZES13, RRM12, RRSPTR14, Rud13, RV13b, Ryo10, SK11a, SS11a, SMNN14, SGJ10, SSK13, SBR^{+}12, Sev13, SS13b, She12a]. problems [She14a, She14c, SYl14, SCC11, SY12, SCK13, SY14c, Sta12a, SXDJ11, SD13b, SNH10, SBDB13, SW12, Tad12, Tan14, TA11, TJ14, VTM14, VMS10, VV10, Van10, VG14a, Ver10, VT13, Wan10f, WLCW10, WHC10, WCZ11, WW11b, WCL11, WMH12, WCL12, WY13b, WY13a, WLW14, WY12c, Waz11b, Web10, Web12, Wei10b, Wil10a, Wil10b, Wil13, WR10, WL10b, WZ11i, WYBT12, WL13c, XY10, XZY14a, XWWC10, XSL10, XG10, XM10, XXS12, Xu13b, Yan10e, Yan10b, YW11a, YCXL12, YZ12a, Yao10, YXZ10, YHL08, YS13a, Zah11, ZS10a, Zha10c, ZF11, ZM12, ZS12, ZLX12, ZG12, ZW13a, ZHH^{+}13, ZL14c, ZLG^{+}14, Zha14b, ZGZ14, ZH10b, ZSHZ11, ZJ14a, ZM14d, ZM14e, ZS14c, ZH14, ZCD10, ZGY14, ZLW10, ZXZ11, Zhu12b, ZL13e, ZZZL14, Zit11, BD12c, FM14, ZZW13, JS14b]. procedure [ABCD11, CRX12, DK14a, GM14c, HB11b, Tag11, TZZX11, VMRA14, VC14b, WHL14, Zha10e, Zhu10d]. procedures [CLS12, Gue14b, Guo10, Guo11, Wah13]. Procesi [Den11d, LZH10]. Process [EH15, Agr14a, ALNGMGMM14, AŞK10, BJB11, B– D11, CCS14, CQ14, CMW10, COPP14, DWLL10, DZS^{+}14, EM11, EH14a, ECS13, GNdC12, Gu10, GJGSNRÁ12, HR11, Hu12b, Jea13, Kar10, Kim14a, KBG11, LNdCJV14, LHT14b, Lu12d, RM11, RRRRTR12, Sar12b, SVV13, SMWR10, Sun14, VG14a, WN14a]. processes [BDGIPG12, BCNP14a, BCNP14b, CZ13, FY12, Gat14, GR11, Gra11, JWW14, Kos11a, LMS10, OP12a, RRSPTR12, RRSPTR14, SO10, SMR14, SS13c, WL13b, ZBM10]. processing [AY14, LYZG13, LWW14, PM11b, Rud12, Rud13, WL14a, XZY14b]. processing-time [Rud12]. processor [Rud13]. procrust [WY13a]. procurement [HLL14]. produced [ÖKU12]. producing [CPC11, JBW14]. Product [GKV12, HN10a, ADÇ14b, CG14d, DE11, Din13b, DCK11, GT11a, GKC11, GP13, Jea13, Ker13, LW11b, LY10b, LHTN13, LS14e, ML11a, Mez13, PS11, SMNN14, SS14d, Soo11b, Ste11b, TNB11, TRA13, WLYC13, WLLD13, YK13a, YRV10, ZLC10, ZZR11, BRS13]. Product-type [HN10a, SS14d, Ste11b]. production [BM14a, CV13, CX14a, CX14b, DS14d, GLSJCBS10, GKC11, HRH14, HLL14, KBG11, PSC12, SM11, Sar12a, Sar12b, SYB13, VV11b, WWK^{+}14]. production-inventory [BM14a, DS14d, GLSJCBS10, PSC12]. Products [LS10b, LO14, SSB11b, BJB11, Bry11, CBTGWW14, CGM13, GCMP14, JAG11, KE11, LXZX13, LLLS13, San11a, SSB11a, WT11a, WHL14, XQ13]. Professor [Can11]. Profile [WZSW14, HLMW13, ZYZ14c]. profiles [CS14b, RH14]. profit [BGRS11]. profitless [CZ11c]. program [BK11a, CLW12, SC14a]. programming [Alz14, Ant11a, Ant11b, Ant12, AZ14c, AL10b, BBC14, DD12, EES14, Ema13d, Eto11, GWM10, GS14d, HMH13, HZW13, ICKV14, JSMA11, JLHW13, KBG11, LWQ10, LTL11, LWFG13, Mak14, MAN10a, MCR11, MP14c, Mat14b, MF13, MEBYG11, MZL07, NL12b, ÖB11, PN11a, Pak11a, Pak11b, PD14, PD12, RFK12, Sit14, SSTM14, TD13, Tri14, Ver12, Ver13, WZ11f, WWW13a, Xu12a, ZZG10, ZnXwY14, JYYJ10]. programs [Goy14, LTL12, SRS^{+}13, YDZ11]. progression [CFL14]. progressive [Dem14]. progressively [SSS13a]. Project [HZ14d, KMC12, KDTSN12]. Projected [WZ14b, Gu13, LL11d, PT11b, SB13, XDZ14, ZY13c]. Projection [bZlD11, CFZ14a, CFX14, DHZ11, DZ13, FH11a, lHyLfG10, JMY13, JFHZ14, KBAvBW14, LP11, LWS14, LNPS10, LZ10g, MM10a, Men12b, NL12b, dFPF14, QCK10a, Sab13, Sab14, WYY11, yWZhL14, YZL10, YZK12, YSW11, ZZY12, ZL10c, ZL10f]. projection-based [CFX14, JFHZ14, KBAvBW14]. projection-filter [LZ10g]. projection-proximal [MM10a]. projections [HH14, HS10a, HS13, PP14c]. Projective [ZX10, LLWW10, NBP10, NSNEV14, WJ12, XWMW12, Zhe12]. projectors [BT11a]. projects [Bas13]. Prolate [Dha13, KM12b]. proof [AZAK10, Pal12, PP12b, PP15, Ryo10, YB12]. proofs [SC11]. propagating [HLT14]. Propagation [AAMADH11, GMKV12, KGM14, Alf10, ÁD10, CCSM13, Dem11, GQ13, GM13b, GMV12, HYZH14, HLW13, KPC14, PPC13, Sha14a, SLZ12, Sin10a, Vis14, Wan11b]. propanediol [YZZ^{+}14]. Proper [XBC^{+}13, XBC^{+}14, XRBZ13, DAA13]. Properties [AR13, COPP14, LL10e, LOS13, MN14a, MJ13a, SKW11, WZW12, WJ14, Zha14a, ADÇ14b, ADT13, AB12c, BAR11, BTV12, CCJ11, CMGR14, Cha14c, Che10a, CB12b, CZ14a, CY11b, CBS12, CDCS13, Cur14, DLSS11, DK14c, EDD11, EKK14, GG10a, GPP11, Guo13b, GA11b, HdCOL14, HD14a, JM10a, JYF13, JPI^{+}13, KASY10, KI11b, Koj14a, KSD13, Kra14a, Kro11, KD13, LS13a, LTOX11, LWM11, Mal13a, MU12, MM13b, MMY11, NC13a, RS13, RYÖ11, RMJM12, SGJ10, Sar10, Sch11a, Seg14, Sha14a, SKM11, SB12b, Sta12a, TKC12, Tun13, UOK10, VGA12, VV11b, WES12, WW13d, WL13b, WSK12, WZW14, YL11b, YL11f, ZLL10, ZTY^{+}10, ZM10a, ZM14b, Zho12c, dCCL14]. property [AN– D12, ANR13, ASM11, AM12b, DW12, GGYW11, JS14c, Koh13, LWW11, LW14e, LCL^{+}13b, LW10b, Mon12, PP12b, PP15, PFS10, STH10, SMH14, Vod13, Wan10b, ZGY^{+}10]. proportion [RGA14]. Proportional [SLJ^{+}13, YJ10, Tsi11a, Xu14a, YX12b, ZLS12a, ZCY14]. proportionality [BCF13]. proportionally [LPAE12]. proposed [bOYqGL^{+}14]. protection [dCDNVS11, KMH^{+}13, ZWL14b]. protective [ZWL14b]. protein [LPP^{+}12, NKM^{+}10]. protein-ligand [LPP^{+}12]. proteins [SBS11]. Proving [ADT13, PP13a]. proximal [AM14e, CC14a, HD14a, yL14c, MM10a, Niu11, Ver10]. proximal-descent [HD14a]. proximity [Bas12a, DI13b, JS14b]. PRP [CD10a]. Pryce [NBYHV11]. PS [Udd13]. Pseudo [RAIM14, AOM^{+}11, AIAD12, ASR^{+}14, sCCLY10, CLC^{+}13, CMM12, CSS11, DLX13, DHS10, Guo10, Guo11, Jun10b, Jun11a, LSYS14, LL12e, lL10g, MB13a, Mop11, WZ12, Yan12b]. pseudo-almost [CSS11]. pseudo-contractions [DHS10]. pseudo-contractive [sCCLY10, CLC^{+}13, CMM12, Jun10b, Jun11a]. pseudo-hyperbolic [Guo10, LL12e]. pseudo-integral [AOM^{+}11]. pseudo-integrals [LSYS14]. pseudo-isochronous [WZ12]. pseudo-parabolic [Guo11, Yan12b]. pseudo-plastic [ASR^{+}14]. Pseudo-Spectral [RAIM14]. pseudo-transient [lL10g]. Pseudocomposition [CTV12]. pseudocontraction [SW14c]. pseudocontractions [CQ14, PK12]. pseudocontractive [DTP14, GFMP13]. pseudoconvexity [Kas12a]. pseudoelastic [KDS13]. pseudomonotone [Sad11, YSW11]. pseudoparabolic [Ami14]. pseudoplastic [AM14g]. Pseudospectral [BPPS14, Bhr14, Boy11, Don14, FFOC14, Gal11e, XYY12]. pseudospin [IS10b, IBS11]. psi [Che10a, CB12b, Che13b]. PSO [ICS^{+}13, LRM12, TS10c, WY13b, YD13]. PSS [She14c]. PSVM [ZZF12]. psychological [YLC12]. PTT [MZWZ12]. public [LSW^{+}14a, SDB14]. pull [GD11]. Pullback [Kan13, QZS13, SW14d]. Pulling [BB11a]. pulsatile [BMMS13, Gad11]. pulsating [ZL12a]. pulsations [WDS10]. pulse [FCZ^{+}10, LS14b, MHB14, SS14a, XLS13a, XHG11, YHW13]. pulsed [Wei14]. pulses [GHC13, GQ13, MYJ10, ZCL14a]. punches [CLW11]. pure [And12a, SWLP13, WY11a, YHWZ11, Yig11]. pure-feedback [SWLP13]. purely [CS12b]. pursuit [SGT10, SY14c]. push [GD11]. push-pull [GD11]. puts [ZZ11b, ZC11]. Puu [CG14a]. PWM [Kuj14]. pyroelectric [HYT14a].

Q [Pak11b, ZZ10b]. Q.I.F [RSK14].

QARIMA [RRSS14]. QLM [Kri14]. QoS [DXC^{+}14]. QoS-aware [DXC^{+}14]. QPQC [BDK^{+}14]. QRS [CBRTA11]. QSSA [LJ11]. quadrangulation [LMZ13]. quadrant [BKP14]. Quadratic [BC14a, Che11a, KY14, ABRT14, AG14a, Alz14, BAA13, BBDS14, BKP14, BAP^{+}14, Bel11b, Bel11a, BH13, Bog11, CLW12, Con13, DS13a, DS14b, DFG14, ESH10, EZ12, Eto11, HL13d, JLL12, KGC11, KH14b, LMZ13, LJR14, LWQ10, dLL13, LLZ14c, LfXfC13, Man12a, MZL07, NH12, NL12b, ÖTC11, Pak11a, Pak11b, RGL12, SGC12, WYT13, WYBT12, Yan12b, YDTW11, ZB14, ZY14b, ZC14a, Zho11c, ZL13f, CP12c, SS14b]. quadratics [KN13]. Quadrature [Wil11, AS13a, BV12a, Bur12, CM12d, DB12a, DB11a, DW11, ELIGMB11, FPP11b, JPM12, JMS13, LJT13, Maa12, Men11, MCS11, Mil12, MB14, Mur11, Mur13, Nar12, PM11a, PS12a, RMN14, RSG14, SN11, SPP12, UB13, WW13f, WW13e]. quadratures [HN10b]. quadric [GC14, Hu14]. quadrilateral [KP11a, SP13a]. quadropoly [EAE13]. quadruple [JPI^{+}13, Zha14c]. Qualitative [GSC14, LZLB14, Liu13a, SG11, SGE13, WQ11a, BGGH12, CL14b, KD13, FKH10a]. quality [LD11, MSGN^{+}14, NCdCDM13, NA14b, PB13a, SGJ10]. quantifying [Lin12, RS10]. quantiles [WN14a]. quantisation [Abb12]. Quantitative [AM12a, GG11a]. quantities [WHL11, WHW13b]. quantity [OYCCB13, WWK^{+}14]. Quantization [SGN13]. quantized [JT13, WY12b]. Quantum [LJ11, SFP^{+}11, WiN11, Abb12, BS12b, FG13a, GW11a, GHLF14, GK13, Goć14, IS12, IO12, JSWJ13, JJL14, Kam14a, LU12a, MÇ12a, SSK10, Wan13b, gZrLX10]. Quantum-behaved [SFP^{+}11, JSWJ13]. Quantum-inspired [LJ11, GW11a, GHLF14]. quarantine [SG11]. quarantine/isolation [SG11]. quark [EN11]. Quartic [LL10f, Boy14, CMD13, KZ14c, LKS14, TIT13, WWZ14b, WZL14b, WTQU14]. Quasi [CR14, KL14b, KK12c, Kud13b, LP11, LTU13, LXZ12, LXLY14, MMX10, SZW14, Wu12d, YL11d, AAI14, Bor14b, BH11, ÇÇD10, sCYW10, sCLCY11, CLCZ12, sCWT^{+}12, sCCYW13, CLC^{+}13, CWWC14, DHZ11, DCL11, Fau14, GIR10, GL13a, GB12, GC14, GX12, HZ10, bJLyZlL14, JW12, KPW11, LMZ13, yLbLjH10, LW13c, Maa12, MS12b, MG14c, MSS14a, MV11, NB11, NNK14, OS11, QCK10a, RSG14, RM12a, SK11a, She12a, SXCD11, SD13b, TL10, VR12, WWZ14b, XFW^{+}12, XLY14, YZK12, YK12, YY10, ZOS10, ZK10a, BGP^{+}13, sCCL11]. quasi- [sCLCY11, CLCZ12, sCWT^{+}12, sCCYW13, QCK10a, SK11a, YZK12, sCCL11]. quasi-asymptotically [CWWC14]. Quasi-contraction [CR14]. quasi-contractions [GIR10]. quasi-coprime [TL10]. Quasi-exact [KK12c, Kud13b]. quasi-geostrophic [RM12a]. quasi-interpolants [LMZ13, XFW^{+}12]. Quasi-interpolation [MMX10, GC14, JW12, MV11, WWZ14b, ZK10a]. quasi-linear [MS12b, MG14c, MSS14a, SD13b, YY10]. Quasi-Linearisation [BGP^{+}13]. quasi-Lipschitzian [sCYW10, yLbLjH10]. quasi-minimal [AAI14]. quasi-monomiality [Fau14]. quasi-Monte [XLY14]. Quasi-Newton [KL14b, LP11]. quasi-Newtonian [GL13a]. quasi-nonexpansive [CLC^{+}13, DHZ11, YK12, ZOS10]. Quasi-periodic [Wu12d, GX12, SXCD11]. quasi-polynomials [RSG14]. Quasi-rational [YL11d]. quasi-reversibility [LW13c]. quasi-singular [Maa12]. Quasi-solution [LTU13]. quasi-strongly [bJLyZlL14]. quasi-variational [BH11, HZ10]. quasi-variational-like [KPW11]. Quasi-Wilson [SZW14]. quasicrystals [GL11b, GYS13, LF11b, Li13]. Quasilinear [SW10b, Ami14, ÇT10a, Cov11, DH13b, FV14, Kum14, Laz11, Li11b, MK14b, SdAT14, TMM13, TMM14, YKL12b, Yos13, YYY10]. Quasilinearization [LW14g, SuR13]. quasinorm [CGJ12]. quasiperiodic [WW11e]. quaternion [AR14c, BAV11, BI14, Fal14, JJL14, KWS10, KM14c, Kyr10, Kyr14, LWZZ14, SW11, Son12, WM13a, WLLD13, YWD13, ZW11a, ZW13d]. quaternionic [BHAV10, BRAS12, KGH12]. Quenching [Cha13a, JZZ13, BS12a, Zhi11, ZMZ11]. queue [Amm14, BM14a, CK14b, Dim13a, GCG14, LT11, LWZ14, LLHY11, wMwWsT11, RRSS14, TLW11, YWK11]. queueing [GNS14, JU11, JS12, YCS11]. queues [CCJ14b, SLL14, WZ11c]. quick [ÁGVZ12]. quicker [Lu14b]. quickly [LW14f, ZYM11]. Quintic [ZHY13, Beh10, BJS14, EZ14b, GLV14, LL13f, NC13a, She10, WD12, XH10, Xu11, ZM14b]. Quintication [EZ14a]. quorum [JZY10]. quotient [OP14a]. quotients [Ant13, KP13].

R [Pan15, CZZ^{+}11, CT13b, RRSPTR12, RRSPTR14]. R.S [Ban14]. Racah [ÁNA12]. Rach [DCR12]. radar [AWS10]. Radau [PS12a]. radial [BYM14, BW10, Boy13, CSQ11, CJM14, Con13, EGJB13, FYM13, GA11a, KAY12, Ko13, KGO10, Lac14, LY10a, LLSY11, MÇ12a, Neu12b, OPD12, PR12a, Sar12c, Ste12v, Ste12w, WWC12, ZSMB14, RY14a]. Radially [KS11e, BMHM14, YY10]. radially-symmetric [BMHM14]. radiation [BD14b, Har14, PMV14, SHAO14]. radiative [HS12, KCBH14, MS13e, SWL12, SdAT14]. Radii [AJR12, XFS12]. radioactive [MP14a]. radiotherapy [BFH12, RDFL13]. radius [AÇ13b, BB13a, Das11, GD10a, HY14, HMGC18, Kol14b, LW14a, LZDS12, MDÇ13, PB13b, RA10b, SB12a, TH14b, ZS14d]. rail [JHZC13]. Rainbow [HLS^{+}14]. rainwater [NCANCF12]. Ramani [WT10]. Ramanujan [GBD^{+}12, Mor10e]. ramp [AAKB13, ZAA^{+}14]. ramp-type [ZAA^{+}14]. Random [CCJ10, CCJ11, LLW10a, YLZ13, AEESMA10, BSKL13, Cha14c, CCCRR13, CBLC12, CJRV12, Gho14, GPCCA14, GK11, GKV12, GX13, HY11a, Hwa11, Hwa14, IG11, KBG11, LS14a, mL12b, Ma12, MKP13, NCÁHCLP13, PJ10, RJ14, SSM13, San11c, SAuAH14, SDC10, SAA14, Sha13b, SS12a, SM14c, Tsa13a, WPSW13, WZH13, XD11b, ZR13, ZWP13, Zha14e, ZWL14b, LRM12]. random-local-optimal [Tsa13a]. Randomized [Goć14, GK13, KHP11, ZY14a]. randomly [FP13b, LJPJ12, NHCLP10, VKV14, WLP^{+}14]. randomness [DZ12, KCK08]. Range [Ram11a, SYl14, CRR^{+}13, Hut13, Kol14b, MM10b, MMSZ14, RY12, WN14a]. ranges [Ada11, CN14b, DGS14a]. Rank [ZF13a, AGM13, BBT10, CSC10, LH11a, LWWZ11, Sau11, SCIMM11, SPKS12, TW13, Wan14c, WFwC12, Wri11, ZW11a, Zho11b, ZL10f]. Rank-one [ZF13a, CSC10]. rank-type [Sau11]. ranking [AOTV12, AOTV14, Iç14, WY10c, WYL10b]. ranks [GH10c, HW12, KWS10, WLY11]. rare [SSCA13]. rarefied [YHH11]. Rashed [EM11]. ratchet [STV12]. ratchet-cap [STV12]. rate [AAKB13, BB11a, Bog11, BR10c, DMS11, DIAQ13, Deo12, DL11, FGRLSV14, GYL^{+}13, GGM^{+}13, HYT13, Hou13, Kar10, LTY11, Liu11e, LWG13, Liu14b, MS14a, PT13, QLL12, SSS14, SX14, SYB13, SL14d, STV12, SMPA13, WL14b, WZ11g, XZY14b, Zha11a, ZW13c, ZZL14, ZYZ14c]. Rates [BD14d, BSP11, CS11b, CW11c, EMM^{+}12, Gra11, HLM14, Kan12b, KXW14, KPT11, LSJ10, MEL13, MM12, Mus14, SS11i, TLZZ12]. ratio [AMM11, AKS11, BKA12, BLZZ14, CHMT11, DLL11a, DWP14, DWC13, EV14, FW14, LY11b, LZTZ13, LTSV14, LDZ10, MC12b, MCL14b, SAuAH14, SY10a, SS11d, Van11b, VG14b, WQ11a, WW11a, WFL14a, YK13a, ZLD11]. ratio-dependent [AKS11, BKA12, BLZZ14, DLL11a, DWP14, DWC13, FW14, LY11b, LZTZ13, SY10a, WQ11a, WW11a, WFL14a, ZLD11]. ratio-like [Van11b]. ratio-scale [SS11d]. ratio-type [VG14b]. Rational [CM11, Ech12, Fer11, AMA12a, AMA14a, AM12c, Ata11, BAA13, BB13b, DAV12a, DAV12b, DB12a, DP13a, Den11a, DW11, HHX12, Hu13b, HX14, HW11, ISG11b, JZ13, JDLL14, Kud10b, LZLF10, LDJZ13, LCL^{+}13b, Nov14, Rao14, Ren14b, Sas12, DAV14b, Ste12j, SAAS14c, TS10b, WW12, Wan13a, Wen10a, Xin10, XwK11, Yam14, YH10, YL11d, YZZP14a, YZZP14b, ZM14a]. rationality [DLGJ14, PMP11]. ratios [JLL12]. Raviart [BY11b]. ray [AP10, Zak10]. Rayleigh [GG13a, GLZ10a, HZ12b, HS12]. RBF [AGM14, EL14, LLT^{+}14, Udd13]. RBF-based [AGM14]. RBFNN [EL14, KPBC14]. Re [Liu14d, PJ10, HLWH12, XQ11a]. Re-nnd [Liu14d, XQ11a]. Re-pd [Liu14d, XQ11a]. Re-seeding [PJ10]. re-substitution [HLWH12]. reachable [FL14, ZFCW13]. reacting [MU12, SJ13]. Reaction [Alf12, AP14, BD14b, BC10a, BA13, CL14a, CG10, CG12b, CG14c, DN14a, DM13b, DWP10, AA14a, GS12c, GMZL11, Hay11, HVC12, IS10c, Jos13, JBDC12, KR10, LLW10a, LZQ11, Lia11, lLW14, LY14c, MC14a, MVQDE12, MR14b, MJ11b, NRWF12, NKM^{+}10, OP14b, Sar12c, Sha14a, SZ14b, VT13, WZ13b, WW13b, WZZ13, YWZ10, ZL12a, ZCZ14, ZLY11c]. Reaction-diffusion [Alf12, BA13, CL14a, CG10, CG12b, CG14c, DN14a, AA14a, HVC12, LLW10a, LZQ11, Lia11, lLW14, MC14a, MR14b, NRWF12, OP14b, Sha14a, SZ14b, VT13, WZ13b, WW13b, WZZ13, ZL12a, ZCZ14]. reactions [BSØ11]. Reactive [FMSC13, SS10g]. reactor [CBA^{+}14, CM12b, MVRVSCMV14, QS11]. Real [KV12, Len14a, SAA14, BU12, BS12c, BPR13, BR14c, Che11c, CZ14a, DHDB12, GCLL10, GMA13, GJGSNRÁ12, GTM13, Hla13, HZW14, JPI^{+}13, KK11a, KKS11, KZ11, KZ12, Kol14b, Mag14b, MMSZ14, Pet14, PKS10, QFCZ13, Rim13a, Rim13c, Rim13b, Ste12u, TMJ14, Van11b, YH14a, YK13b]. Real-coded [SAA14, DHDB12, GCLL10, YK13b]. real-imaginary-type [HZW14]. Real-time [KV12, Che11c]. reality [Sta12b].

Realization

[HYD10, BBF12a, BB14a, ESNE^{+}14, Sin11b]. realizations [DG14b]. Realizing [WWS^{+}14]. Reallocating [LT12]. really [KNS14]. rearrangement [AM14a]. reboot [KL14a]. recesses [dCDNRH12]. recessive [Dos14]. recipes [Res11, Res13]. reciprocal [CK13a, KA13b, WT12, WN12b]. reciprocally [LP14]. reciprocals [BS13]. reciprocity [CS10b]. recognition [qGyFzX13, Hua10, KB11a, QZ10, Sev13]. recombination [Hsi14]. recommendation [DXC^{+}14]. recommendations [CCCRR13]. recommender [GJ14]. Reconstructing [ZS14f]. Reconstruction [LX11a, AGM12, ADT13, CL13a, Cor11a, KK12b, LWS14, ZCZ14, ZL13d]. record [JS14c, WHL14]. recorded [KVP14]. Recovered [HLM14]. Recovering [YS11b, MS13g]. recovery [Dim13a, FGO12, MJ13b]. recovery-based [MJ13b]. rectangle [MRCD12]. rectangular [BB13b, DB11a, JE11, LR14a, LWZ13b, RNEA14, YHL14, Zha10f]. Recurrence [Dua10b, Dua11a, Ger12, QQXY12, qWH11]. recurrences [KA13b, PSWW14]. Recurrent [Che13e, Dim13a, HZW14, LYWL12, PM11b, SC13]. recursion [WHW13b, Waz12a, Waz13]. Recursive [Fuk13, Res13, Asa10b, CHK10, FC10, FWZW10, Mat12, ZLYL10]. red [KMAG12]. redefined [MJ11a]. Redheffer [Zhu11b]. redistribution [Ber10]. Reduced [KK12b, SHBN13, ZSWS12, Abd13, EES14, EEZ14, KBAvBW14, Kha14c, KN13, LWL14, LJ13, LLS13b, Luo14, MVRVSCMV14, Mis10, MT12b, VD14a, WZ11f, XC10a, Xia10b, YCZ14, ZHT13]. Reduced-order [KK12b, ZSWS12, Abd13, LLS13b, Luo14, YCZ14]. Reducing [Che11c, GZ12b, hPnZrL12]. Reduction [CN14b, AHRV11, BB11b, CHYT12, ELS13, GPP11, Kha14c, LW11b, Liu10b, Lu12d, RM11, Ram10b, SM13, SMPA13, XLZ10, YCS11]. Reductions [AM13c, BGGH12, CS11d, GMP13, GPW14, IKU10, MK13b, SGK11, TV13, WDL12b, WWWJ14, XC10b, YH14c, yZqLwW12]. reductions/analysis [MK13b]. redundancy [HY11b, SSTM14]. redundant [Yua12]. Reeves [LL13a]. Reexamination [CGAK10]. refinable [Ros11a]. refined [LMZ13, ZAA^{+}14]. Refinement [KM12a, AST12, KM11b, Mit13, PP12b, PP15, Wah13]. Refinements [HYT14b, CN12, KK13c]. Reflected [Hu12b, GNdC12, HR11]. reflecting [BCNP14a, BCNP14b, HZ14a, LGN10, LNG12, Zho11a]. Reflection [Sin11a]. reflections [AAC11]. reflective [ZZ12b]. reflexive [CWWC14, Din12, LLW11, PX13, XQ13, ZJ14a]. reformulation [CFG13, LHL11]. refuge [Jan13a, JXA11, LL13b, MCWC13]. refuges [Yan13c]. Regenerative [HSW11]. regime [ECS13, HYX14, HMPVQ14, LWL14, SL14a, Siu10]. regime-switching [ECS13, HYX14, Siu10]. regimes [CFGRV13]. Regińska [BBF12a]. region [Bes10a, CKP^{+}13, DG13, ES12a, GZ14b, KR13b, PZ14b, Qj10, QC12, Tan14, Van11b, WW13a, ZZQ10, ZZC13]. regional [AMPT14]. regions [AAP10, BJ12, CKN13, Dub13, lJLpWxZ14, ML11a, NMIA11, Oss10, Rec14, SMN11, San14, SN11, Yan11d]. Regression [GJNN14, AY14, AE13, COPP14, GYL14, LNdCJV14, LY14a, LPY11, NTdCJL12, NCdCDM13, NA14b, OSRPC12, Örk13, PDPGOR13, SKIH12, Wu14a]. regressions [Hua11b, JSWJ13]. regret [JWG^{+}14, PR14]. Regular [Zha14b, Che11f, KP14a, Khe13, Kri14, KCLC13, KM12c, LMY12, Lin13a, PR11, SW12, TMM13, TMM14, WMJW11, Zha10c, ZCZP14]. Regularity [FLP11, GZW12, Kol14b, Kum11, LL13d, LP13b, Zha14e, ABR^{+}11, BTRZ14, Che11e, Che14h, KS11b, WWW12, ZLZ14c]. Regularization [HHDM13, Sal11b, BBF12a, CPS14, CM12c, GPK13, LPB10, Liu12a, LW13d, LW13c, MQ14, PZ11, QW11a, VG14a, WT11b, YLLL13, ZxHW14, ZW11c, SJ12, TTQ10]. Regularized [Cor11a, Nie13, AS11a, Cai11, GZL14, HM12, HM14b, LPR10, Lin14a, LIC12, ÖK11, SXL13, Tua13, YWY^{+}13]. regularizing [GHS11, Mam13]. regulation [LZWZ14]. regulatory [CW14b, HLC13, ZaFT11, ZCL14b]. reinforced [BHG10, HA12, Vis14, ZA14b]. Reinforcement [Ko13]. reinstatements [AH11b]. reinsurance [AH11b, YJ10, YX12b]. rejection [ZL13a]. relapse [GZ13]. related [AAS12, Bot12, BH11, CST12, Che10a, Che11e, Che14a, CS13f, DR10a, DBG11, DCK11, DRS11b, Guo13b, HT14, IA14, KR13c, LDWY14, LO14, MS12a, MD13a, PS11, Pré11b, Ras14, RY14b, SFL^{+}11, SW14a, SI14, SÖCS12, SG10, SKW11, SV10a]. relating [Sza13]. relation [Ger12, KI11a, KI10a, WLZ11]. Relations [TW13, ÇM13b, LWN13, NA14a, WT11c]. Relationship [Liu13b, Elo14, FKH10b, Zho11c]. relationships [Aou10, BC11b]. relative [AHJ14, LLY11a, WFW10, WL13d, Wu14b]. relatively [sCLCaL11, CS10c, DHZ11, NS11, OS11, She12a, SLZ11, ZOS10, ZS14a]. Relativistic [IHF13, PM12, Woo20, HdCOL14]. relativity [PM12]. relaxation [aBS14, Fan10a, Fan13, MKSC13, Xia10a, ZL13b]. Relaxed [CAY11, AMR13, bJLyZlL14, yL14c, Liu12b, WYY11, ZZGL14a, ZGZ14]. relay [TPA10]. release [LWW14, PKD14, WCW^{+}13]. released [JDN14]. Reliability [BSR10, GGOS12, LC12, SLBS11, WW11g, Wu12b, Yua12, dLDP13, BTT13, CHK10, ET12b, GYG11, HY11b, KLBR13, LY11a, Lin12, Ou10, PKD14, RTW13, Sar12b, SMS14, UMMT14, ZZWF13, ZLY11b]. reliability-redundancy [HY11b]. reliable [CGVR14, HX13, JXK14, MPW12, Waz11b, XC10c, XWC11, XSC11]. remainder [Lin14b, PS12a, Tsi11a]. Remark [Asl10d, Den11c, Con13, Gal11a, Guo11, LB14b]. Remarks [BC12e, BPX14, EMA13c, Esf11b, GBA14, HJMY14, HLSS10, JS14b, Mar14, Ste14a, Web10, Web12, Xia12, Cur14, EGJB13, Tom12, Tom15]. remotely [GXSH14]. removable [YY12]. Removal [KMAG12, dWcFwWjZ13]. removals [SSS13a]. Removing [GRV13]. renders [Koç14]. renewable [PP14a, TKLH10]. renewal [ZY12]. Renormalized [NMR14]. Rényi [SMDZG13]. repair [CK14b, KR13a, MCPO12, WZ11a, WSY14, YX12a, YGX14]. repairable [KL14a, WZ11a, WW11g, XH13b]. Repairing [PHB14]. repairman [WX12, YX11]. repairmen [Yua12]. repairs [LWZ14, WZ11c]. Repeated [BDGP14a, Dim13a, IG11]. repetitive [BTV12]. replacement [WZ11a, ZW11b]. replenishment [Ban14, DBS12, OYCCB13, SGC12]. repopulation [PHB14, WDBM14]. repository [KMH^{+}13]. Representation [BZT11, LS11b, MD12c, Bez13, CS11c, DAHAK10, yDpY14, HSC14, JKW^{+}11, KAK11, LW11a, Li11a, LZ10d, Son12, SG14, TQ11, WZZ14, XRBZ13, YWY^{+}13, YK13b, ZCL11]. Representations [BFB12, GD10b, LJCI13, jAS14, GWA14, Ger12, Guo14, Kha14c, KT14, Kyr13, Kyr14, LLLS11, LXY10, LLLS13, SC11, Sht11, SCIMM11, SPKS12, TÜ12]. represented [AdC14a]. representing [WHI12]. reproducing [AASS13, Gen11, GQ14, JC13, LNC12, WDT^{+}13, WL10b]. reproduction [BDAD10]. reproductive [PW14]. repulsive [BHJ13]. required [CKH11, HS11b]. requires [BX12]. rescaling [dOL13]. Research [DLGJ14, Zho12c, DD12, GXF13, Ign10, Tie12]. Residual [DR10a, La 13, AAI14, DF14b, GHS11, GC11d, La 10, qMWjHpX13, Öke11]. Residue [GM14c, dlS10]. resistant [LSJZ13]. resistive [PP13b]. resistivity [KR11]. Resolution [Lim12, CVP^{+}10, MAD11, TKL14, XKH10, Zei11]. resolvent [HSC14, VRMT14]. Resolvents [Bec13, Fan14b]. resolving [AHKS12, KKVCS12]. resonance [CT12c, CvW10, Jan14d, Jia14, KHK^{+}11, WCZ11, WKLW12, WL12a, Web10, Web12, XGM10, ZLZ12]. resonances [Fer12b, ÖP12b, RA10a]. resonant [CGRS12, DGOR13, DFG14, Li14b]. resonator [GO13, JSF14]. Resource [KM10, DXC^{+}14, DLGJ14, FYL13, LZLY14, PP14a, YH14e]. Resource-saving [KM10]. resources [GZ11, Lin11c, TKLH10]. respect [BX12, CALR11, CDMS14, GH10c, LQY11, Mor13, SK13b, VD14b, Wan14c]. respective [ZLL13]. respondents [DP13b]. response [AS11b, AA13, CHC11, CLD12, DSM^{+}12, FW14, GC11c, HL11e, LZ10e, LZ10c, LY11b, LB14c, MEL13, PLF13, QLWW12, SDL10a, SL13a, SM12, SR12, TX14, WES12, WW14d, Wan14b, XWMW12, XDG12, ZjHO13, ZCL14b, ZM10c]. responses [BKA12, DLL11a, WPKE14, ZK14]. restart [QLL12]. restarting [LLBW11]. resting [DB11a, HA12]. restoration [GG14d, HL13d, LP11, LLP10, PS10, YWD13, ZP13]. restoration-free [ZP13]. restoring [Jan14b, ZLC^{+}13]. restrained [WaZW13]. restricted [ARG14, Bah13b, BGV10, CG14a, LY14a, LMY12, Ou10, SW11, Wu14a]. restrictions [LWY14, WYL10b]. result [Aou11a, BS14b, Bor14b, CW14a, DMW13, DRGR10, DK14b, FL14, Ge14, SHL10]. resultants [YZZ12]. results [AGT11, ANR10, AKN11, ABRT14, AR14a, AOAA10, ARR12, AHJ14, ASVV10, AKR11, APJ14, BMR10, CH12, CS13f, CG14d, Cov11, CIP10, CISM11, CI11, DB11b, DN12a, Dia14, Du11, EK10, FKH14, GS11c, HY14, HA11, HP12, IB11, Jan13c, KYH10, Khe13, KD13, KPP^{+}14, LZ12a, LYWL12, LWZ14, LDL10, LY12c, LZB14, LMA14, MI10, MA13d, MD11a, MD13b, NA14a, OP12a, PAA14, QCK10b, RPA14, SAL12, SHI13b, SK13a, TJZ13, VRMT14, WWW13c, WWT13, WF14, Wei11a, WW11f, XX14b, Yun11b, ZLW12b]. resurrection [WL13b]. resuscitation [Jun12]. retailer [CCO12, LOD12]. retailers [SCN14]. retardation [CT11a, SW10b]. retarded [ABT^{+}12, ABLZ14, CL14b, TFM13, WZG12, YG11]. retirement [BM12b, BM12c]. RETRACTED [Rho14a]. Retraction [AZ19, EH15, Rho14b, Rho19]. retracts [HKK12]. retrial [CK14b, GCG14]. returns [THCW11]. revenue [Lin11b]. reversal [RJ14]. Reverse [CIN14, MD11b, MD12d, CIP10, DD13, LY10b, LHCI12, LZB14, MD11a, MD13b, PHPD11, WC13, XQ13, YL11c, Yu11b]. reverse-flow [PHPD11]. reverse-order [LHCI12, WC13, YL11c]. reversed [LSJ11]. Reversibility [dRS11a, LW13c]. reversible [AM14b, CHYT12, GZW13, GX13, JZY12, KX14, LL10b, ZC14a]. review [MSS14b, SRP13, Soo11b, WR10]. Revisited [CTS14, BBI13a, MS13c, VA12]. revolution [Ros11a, VRZ11]. reweighted [dWcFwWjZ13]. rewiring [SSG13]. rework [WWK^{+}14]. reworkable [PSC12]. Reynolds [CFZ14a, KBB12, Wan10d]. Rhoades [KI12a]. rhombic [WW13f]. rhythms [FW10]. Ribaric [JKP14a]. ribbon [DFGL14]. Ribière [Li14a]. Ricardo [ZL10d]. Riccati [GL14a, Ech12, EBR13, GL10b, GMZL11, HKH10, KCG10, LC13, LLWW10, Liu12c, dLfX14, Lu12b, MH12, MH13b, Mis10, Rad11, Rao14, SS10a, SY10b, TV13, THP14, UT13, WFwC12, Yüz13, Zha10b]. Ricci [YTY^{+}14]. Richards [PZ14a]. Richardson [GGTCdlF13, MG10a]. Ricker [Bel11a, SvB10]. ridge [TK13, WQW^{+}10]. Riemann [ABRT14, Alo13, BRGP14, DHM14, GKW14, Jan13d, Mer14, PTY12, PTY13, PHGV13, lQZWC14, RY13a, WYK11, ZLWW14]. Riemannian [AAB^{+}13, AAB^{+}14, AM14f, CZSV12, MVS14a]. Riesz [FT10, RMY14, YLAT14]. right [EA14, GVT^{+}13, HH12, Ste12t, YH13, ZB12, ZW14d]. right-hand [HH12, ZW14d]. rigid [Che11g, PRT11, Vis14, Zen12]. Rigidity [VRZ11]. rigidly [CLW11]. Rimming [FKH10a, FMSC13]. ring [BCI13, dCDNCFRH10, ZX12, ZCZP14]. ring-type [dCDNCFRH10]. rings [LZB14, MD11a, MD12b, MD12d, MD13b, MD14, SGZC13]. Riordan [CCJ14b]. Risk [Kar10, BO13, CYG14, CW11c, DS14a, Gat14, HY11a, LQY10, LWL14, LC14d, PAL14, SMWR10, Tsa11, ZY12]. risks [DG10a]. risky [Tch11]. RISM [IG11]. Ritz [CZL14, EJ12, GG13a, GZL14, ZL14b, ZL14c]. river [PB13a, VJBDV13]. RKHS [And12b, And13b]. RKM [GC11a]. RLS [NHCLP10, Nak13, NCÁHCLP13]. RLS-FIR [Nak13]. RLW [ÖK11, MC12c, WP14]. road [NCANCF12]. Robin [CPS14, CRG11, DWC11, Khe13, Kle12]. robot [KPBC14, MT13]. robots [CC11b]. Robust [ABE^{+}12, AHM14, ASMA12, BN10, Che12b, EL14, GYL14, HL13a, HX13, JS14a, JQ14, JXK14, JYF13, KUKU14, LJKP11, LZLF10, LSSZ14, LCL^{+}13a, LTR11, MA10a, NZ12, SAMA11, SKSK12, SGT10, Tsa13a, WHW10, XC10c, XWC11, XSC11, YMZS10, Yan11d, YZZ^{+}14, ZXCL10, ZLM12, ZLS12b, AA12a, AS14b, Bas13, Che10e, CZZL13, CP14b, Das12, JPK12, KR10, KPP^{+}13, LLP14a, Li10a, LL12d, MI10, MPSA14, MB10, PR14, Sal11b, SAL12, TKC12, WS11, WPSW13, XSL10, ZW10d]. Robustness [ZZS14, ASSPA10, BTV12, ZCL14b]. rod [BB11a]. Roesser [Che10d]. Rogers [Cao10a, CS13b]. rogons [LGKM13]. rogue [Yu13a, YY14b]. Rojo [VA12]. Role [Ban14, DBS12, Ver12, BCGK14, BIT12a, DMGK12, GS14c, Gwi11, Gwi12, JHN11, NNG10]. rolling [CZX^{+}14b]. Romania [RG14]. romantic [BC11b]. roofs [dCDNV^{+}13]. root [BBSC13, DR10a, GK10b, GK12c, KK13b, Mag14a, MTER14, PTW11, PPD10, Pet14, PC14b, Tri14, WH11, Yan10f]. root-finding [Mag14a, PC14b]. root-mean-square [PTW11]. root-solvers [PPD10]. roots [BS12c, CL11c, CNKS14, GMA13, KBK13, KKS12, LG11, NSC12b, NSC12a, NCS13, NS13a, NC13b, NCS14, NC14, ONBN10, PE10, SSTZ14, SS10c, SS11c, Yun10b, ZCS13, ZS14e]. rootvectors [Tun14]. Rose [WS12]. Rosenau [HHX11, PZ12, WW11c, WP14, XZ13c]. Rosenthal [Hu11]. Rota [TZS14]. rotary [Nar12]. rotating [AL10a, Cha13c, CWW14, GG13a, MA13b, Nar12, Neu12b, SSM11, WS10]. Rotation [PDPGOR13, AAMAS11, Che11d, JK13, TZYH14]. rotational [ABG14, KMD13, LLLZ11, Liu14c]. Rothe [RB14]. rotor [MSJ14]. Rotorcraft [KGO10]. round [TL14]. rounding [Tsi11a]. roundoff [Pre11a]. Routh [AA12b]. routing [KZS12, LG10b, SSV13, SXGR11, YLC10]. row [LWWZ11, ZL10f]. row-rank [LWWZ11, ZL10f]. Ruin [AH11b, CW11c, Gat14, KCK08, YX12b]. rule [BDGP14a, BBF12a, CN14a, CM12d, HT11b, Kyr10, Kyr12, LL11a, LLX10, PVV11, SW11, WLZ12a]. Ruled [LKY10, Sag11, YLJ14]. rules [Bas13, DB12a, Liu11g, LJT13, ML11a, Men11, OCN11, SN11, SSC11, SDN14, SBS11]. run [KK14b, TES14]. Runge [JSSB14, BDD12b, DEP12, FYM13, FG11, FG13b, GTPM14, GSDB11, Kal13, LLY14b, LJ13, MDD12, Mon12, SMIS14, Sho14, TS14, Tsi14, ZQ14]. Runge-Kutta [Tsi14]. Russell [Esm12]. RVT [HS12].

S [EM11, Pan12b, PS11, DMS11, ZZ10b]. S-I [DMS11]. SAA [ZxHW14]. SABR [FGRLSV14]. SABR/LIBOR [FGRLSV14]. saddle [CJZ12, Che14c, CZ14b, GK10a, JCY10, LZZ11, LWYY14, LZ13b, LZ14a, LZ14b, Lu14c, MC14c, PL10, She14c, WZY14, WLL14, WL13c, ZZ10a, ZS10a, ZW13a, ZGZ14, ZM14d, ZM14e, ZLZY12, Zhu12b, ZZZL14]. saddle-point [Che14c, LZ14a, LZ14b]. saddle-spiral [ZLZY12]. Sadia [PK18]. Sahle [NSNEV14]. Said [ZTD10]. sale [GKC11]. sales [BM14a, San11c]. Salesman [FM14, Liu10d, TSP10a, WZZL12]. salesmen [San11a]. salt [CJM14, MATA14]. Samarskii [AO12]. same [Zho11a]. sample [KB11a, LTSV14, MC12b, SSS13a]. Sampled [RSPK13, KH14a, LWP12, ZL14a, ZL14d]. Sampled-data [RSPK13, KH14a, LWP12, ZL14a, ZL14d]. samples [SSCA13]. sampling [AEESMA10, BD14d, Dha13, DP13b, Gat14, GK11, GKV12, HB11b, SAuAH14, SS12a, SM14c, VG14b, WWC12, ZA14a]. sampling-type [BD14d]. SAMR [MMX10]. Sandor [CS11a]. Sandwich [CKS10, Bas12b, RA10a]. Sandwich-type [CKS10]. sandwiched [Vis14]. Sarmanov [HL11d]. SARS [NSL14]. Sasa [HCD14]. satellite [HSD11, SKB14, Wil11]. satellites [XBYF14]. satisfy [SV12]. satisfying [AN– D12, ANR13, ARR12, DI13b, NA14a, SSRA12, Ste13l, YX12a]. Satsuma [Aba12, GZ10d, Hon10a, eMA13a, GH10b, MKeM13, XD11a, ZZ11c]. saturable [KW13]. saturated [ABK^{+}14b, KM11a, KMH^{+}13, Liu14b, MNOB10, RBT14, WW14d, ZZL14]. saturation [Che10d, Lee11, LSSZ14, RKA^{+}13, Sin11b, Sun11b, TX14, Xu12b]. Saunders [CL14e]. saving [KM10]. savings [SSV13]. Sawada [QMM13, GS10b, MG12a, SGY^{+}14, Waz10e, XLZ10]. Saxon [RNZH13]. scalable [KC12]. Scalar [MMS– D13, Far14b, GC10a, Shk14, Sol12, dlS10]. scale [BBF12a, CR11b, Che11g, CS12b, DR10a, FOL12, Gu13, Hil12, KUKU14, LH11a, LH11b, Li14a, LSJZ13, NCdCDM13, NA14b, PM11a, SS11d, SXCD11, TZYH14, YLWH14, ZCŠ10, ZHW11]. scale-free [LSJZ13]. Scaled [LH11a, LLSY11, Liu12a, MH10a]. scales [AZ10, AM13b, BEP13, BFM13, CSSN11, DV14, EMP12, FZ12, GWZ14, GPK13, GM14b, Has11, HZ14c, KCLC13, LY14b, LH14a, LN10, MT10, MS13f, SH13, Tun14, Vít14, Vod13, Wan14a, ZH12, ZW10a, UK11]. scaling [CPOdCN11, GMP13, GM12, LU12a, NBP10, WZ14b, WLXL10, ZZ10b]. scanning [RVS10]. scatter [HJZC13, QZ10]. Scattered [ABCD11, BLR10, LGCX13, SHA12a]. scattering [Apa10, GPW14, HS12, JKP14b, Kim14b, LZ13a, SSK13, TSN13, WW11c]. scenario [PR14]. scenarios [MATA14]. schedule [CK14b]. Scheduling [GL10a, LYZG13, YK11a, Ber14, Che11c, CWCW11, EG06, HZ14b, HH11, HL12e, HLS10, HL14c, HZ14d, JCW12, KBYL14, LL12a, LCH13, mL12b, LPX12, LH10b, LYL13a, Liu11b, LYG14, LWW14, MKSC13, dejW14, Rud12, WWZ10a, WW11d, WL14a, XZZK12, XBYF14, XG10, XYL10, XY11, XZY14b, Yan10e, YW11b, YCC^{+}12, ZL13a, ZB14, ZHC^{+}14, ZDC12]. schema [GLB12]. scheme [AS14e, AJPZ14, BLK11, BK14, Bra13, CLX13, Che11c, CC14c, CG10, CG12b, CG14c, Cui14, DL10, fDxZxChT12, Dub10, DC13b, EM13, FC10, FHW14, FGO12, Haj11, HWL11, HB13, HHX11, Ito13, JMS13, Jun11a, KKV12, KK13b, KK13d, LZW12a, LFC^{+}12, LY12b, LZW13b, LSW14b, LQ13, MDP12, MMS14, Mol14, MS11, Nie13, NSYA14, NDZ14, PZ12, PP12b, PP15, lQZWC14, RW14b, SHA12a, SYK10, SXL13, SLZ12, SZ11, SSA14b, SIR14, SW10a, SW10b, zSWZW12, TYCZ14, Udd13, WG11, WH12, WRZ12, WZJ13, jWtW13, WFZ13, WP14, WL10c, XWZF11, XFW^{+}12, XZ13c, ZB13, ZW12, ZZ10b, ZZ11b, ZWH14]. schemes [ABDKA12, ADT13, AY11c, Bho10, BBFK12, Cai10a, Cai11, Che10b, CCG10, Chu11, CZC11, CK12c, Dub13, EA14, EVDG12, FCJ14, GRV13, GG12b, GL13b, HZ13, KCZK11, KKH14, LY10a, LY13a, Lia14, lLZ14a, LL13e, Maa14, MC14a, Mal13b, MAD11, NNK14, Ram10a, SGDK11, SR10, SR14c, SF13, TPA10, TW11a, TK12, WGZ10, XLGW12, Xie12b, ZFH14, ZLX12, ZHGP14, Zuf12, dFGAN12, dLDP13]. Schiff [MESAN13]. schistosomiasis [DISS14a]. Schmitendorf [DLP10, JMHZ10]. Scholes [Cha11, LW13a, TM13, TKL14]. school [CSSSV13]. Schröder [CL11c]. schrödinger [Cai10a, BT10b, AMB^{+}14, AAS12, Asa10b, BK14, BCW13, qClLyW10, CM12a, CM14, CT14b, Elb12, ES12d, EKZ11, FYM13, GLZ10b, HM10c, Kat14, KW13, KL10, LZW12a, LKS14, Lia14, Liu11f, LL13f, LGZ14b, LGKM13, Lu12b, MZ13b, MRE^{+}12, NWZ13, RNZH13, SSK10, She10, SDL10b, Sin12, SD14, hTlCqL10, TZJZ12, Üna13, WGZ10, WZ11e, WWZ13, WSL^{+}13, WCL14, WYTL14, WW14e, WD12, Xu11, XZ13b, XC14, YTS12a, YZD10, YZC13, YS10b, YYY10, Yu13a, YY14b, ZA11, yZhLjMzC10, ZL14b, ZYZ14c]. Schrödinger-like [Lu12b]. Schrödinger-Poisson [LGZ14b]. Schrödinger’s [Lat14]. Schur [CN12, DA11b, HB11b, tLpOjCW10, Li11a, LHZ10, LJCI13, XQ12]. Schurer [AGK13, AKS14]. Schwarz [CS10b, GPZ12, LS14e]. Schwarzian [WDLM10b, Asl11a, PQW10, TW11b]. science [Cha14b]. sciences [CNTG12, SF11b]. scoring [NDZ14]. Scott [GZW13, GX13, JZY12]. Scout [ZYM11]. screening [NTS11a, WWK^{+}14]. scroll [CGTCF13, OGJLCCB14]. SDEs [MM14b]. SDP [LS11a]. Search [Civ13, CAC14, DS14c, KKSS14, SK14b, Vir13, ABDKA12, ABKG^{+}13, Ala10, BH13, BDDK14, CkPqL12, CLD12, CD10a, CHC12, CSF12, EA13, FGIV12, GVW^{+}14, GS10a, Gee12, HJZC13, KEH14, KA13a, LH10b, LJ11, LWM11, Mat14b, MHS12, MKP13, NW10, PSTL10, PG13, PZ14b, QLL12, RY14a, RCSH10, RDFL13, SAA14, SWX10a, SW14b, TNB11, VTM14, WZ11f, WW13a, WY13b, WZ14b, bOYqGL^{+}14, YG13, YJCC14, ZC12, ZBM10, ZLZ14d, ZF13b, ZANA12, CEP11, HAA^{+}14]. search-space [RY14a]. searching [ABCD11, SXGR11]. seasonal [CCS14, DS14a]. secant [AHVR14, Cvi11, EHRV12, GSN12, LP13a, RWB10, RA10c, RA12b, ZZZM10, AH11d]. secant-like [EHRV12, RWB10]. secant-type [RA12b]. Second [AZ14c, Fig11, Koç11, LP14, LSZ12, Mal12b, NRL12, RK12, Ach11, AZL13, AISS13, Alz14, AZ10, ANS14, AAS14b, APJ14, AKÖ11, Bah13b, BDL13, BM14b, BEP13, BKP14, BES10b, Bie11, Bou10, Boy12, Bra13, CST12, CCJ14a, CK10, Çet13, CS10a, CJ11b, CT12b, CM12b, CLND12, DAV12b, DS12a, DSX11, Dia14, DJ14b, EA14, EZ13a, FPP11a, Fra10, FG14b, Gal11c, GS13a, Gao10, Gao14, GÖS11, GK12c, HT11a, Han11, HL11b, HH14, HR10, HKI12, Has11, HET12, HL12c, HL13b, HQ14, HX14, HM08, HM14a, Jat10, JM11, KC12, Kar13a, Kas12a, KfM11b, Kha14a, Kim10, KZ11, KV13, Kon14, Kos13, KS12d, LSJ11, LZW13a, LLK14, LH14a, LL10d, LNC12, LN10, LLC11, LLLS13, lL10g, LS11c, MXH12, MCH14, Ma14a, Mal12a, MMO13, MH11b, Men12b, MB13b, MV11, NSYA14, Par11, PT11a, Pre11a]. second [RP14a, RA12c, RHSC14, SHT14, SR14a, SAO12, SDAN14, SHMCM12, DAV14b, SQ11, SK14b, SAAS13, SAAS14c, TMM13, THDF11, TDZF14, TW14b, TA11, TT13, THAB11, UB13, VMRA14, WLH10, WCZ11, WL12a, WF12, WLLY14, XW13, XCH14, XM10, YWT14, Yan10b, YWK11, YYL14, YT13b, YCF13, ZW10a, Zha14d, ZPZ14, Zho13, ZH14, ZX14, Zhu10d, SGN13]. Second-order [Fig11, LP14, LSZ12, NRL12, AZL13, Alz14, AZ10, ANS14, AAS14b, APJ14, BDL13, BES10b, Bra13, CST12, CCJ14a, CS10a, CT12b, CM12b, Dia14, EZ13a, Gal11c, Gao10, Gao14, GK12c, Han11, HL12c, HL13b, HM08, HM14a, Kar13a, KfM11b, KV13, Kos13, LSJ11, LZW13a, LLC11, lL10g, MXH12, MCH14, MH11b, PT11a, RP14a, RHSC14, SHT14, SAO12, SQ11, SAAS14c, THDF11, TDZF14, VMRA14, WLLY14, XCH14, YWT14, Yan10b, YT13b, YCF13, ZW10a, Zha14d, ZPZ14, ZH14, Zhu10d]. secondary [Gra13, RA10a]. section [KB13, PP12b, PP15, SMPA13]. sections [CXG10]. sector [GAC12, LKP12, SDB14]. sectorial [HA12]. secular [VS10]. secure [Abd13, Che12b, Has14, HCC^{+}12, YCZ14, DAV14a]. Security [dejW14, Ach13, HZZ14]. sedimentation [CV11]. seeding [PJ10]. seek [FLW11]. Segal [SGN13]. Segalman [ARH^{+}12, ASRI12]. segmentation [LK12]. segmented [JAG11]. segments [ZT14]. segregation [BCGK14, BA13]. SEI [XYL14, SM14b, TGL14]. Seidel [YZ12b]. Seidelization [GMV11]. Seiffert [CT13c]. Seiffert-type [CT13c]. SEIR [SGE13, WWZ14a, ZW14a, dlSAQ11]. SEIR-epidemic [dlSAQ11]. SEIRI [Xu14b]. SEIRS [KN12, MS14a, QaC13].

SEIS [Xu12b]. seismic [KPC14, KGM14, TSN13]. Selecting [SSC11]. selection [ABDKA12, ABKG^{+}13, AKP13, Chu11, CV11, CFG13, DEP12, GYL14, HLWH12, HZ14d, JLTZ13, JHN11, Kay11, KDTSN12, LQY10, OSRPC12, Örk13, PJFL14, RMMM11, SNH10, ZZYX12]. selections [ZX11b]. selective [CDK13, PHB14, Wan14b]. Self [Fre11, IKV11, Ach13, AS11a, AY11c, Car13a, Car13b, dSCAM14, DI13a, DI13b, DV13, ESE14, ES12b, FS12, GB12, Gui14, HM10b, ICKV14, JFHZ14, KA13a, Kyc11, NW12, PSTL10, PT11a, PB14a, Shi14a, Sun10, SQJ11, Tsa14, VV10, Vis14, WX13a, WLHZ10, XZZX14, YSW11, ZZW^{+}12, ZL12b]. self- [Gui14]. self-adaptive [Ach13, dSCAM14, ESE14, HM10b, JFHZ14, KA13a, PSTL10, Tsa14, XZZX14, YSW11, ZZW^{+}12]. Self-adjoint [Fre11, AY11c, Car13a, Car13b, GB12, PT11a, PB14a]. Self-adjointness [IKV11]. self-affine [ZL12b]. self-assembling [NW12]. self-balanced [Shi14a]. self-consistent [Sun10, WX13a, WLHZ10]. self-inhibition [ES12b]. self-maps [DI13a, DI13b]. self-organizing [FS12]. self-reinforced [Vis14]. self-similar [Kyc11, VV10]. selfadjoint [OC10]. Selkov [AM14b]. selling [San10]. semanteme [LYGZ13]. semantics [GAK12]. Semi [AM14b, Gra11, JZG13, AG14a, AZ10, AZ14c, AHM14, BS12a, BJS14, BH12a, Cao10b, sCCLY10, sCCYW13, CYG14, CZC11, CSMT12, DNSM11, Dub10, DC13b, EKZ11, GYS13, HN14a, KGH12, KS13d, LZ11b, LL14e, LZ14b, LC14d, LDZ10, MMY11, MF14, NN19, POR13, RA12c, SMNN14, SO10, SBS10, TD14a, TD14b, VJBDV13, WFZ13, XG13, YLZ13, Zha12c, ZZF12, ZWH14]. semi-adaptive [BS12a]. Semi-analytical [AM14b, JZG13, XG13]. semi-cardinal [BJS14]. semi-conservative [Dub10]. semi-convergence [LZ14b]. semi-definite [AG14a]. semi-discrete [EKZ11]. semi-discretized [VJBDV13]. semi-Euclidean [KGH12]. semi-groups [sCCLY10, sCCYW13]. semi-implicit [Cao10b, CZC11, DC13b, TD14a, TD14b]. semi-infinite [AZ14c, BH12a, KS13d, LL14e, SBS10]. semi-inner [SMNN14]. semi-integer [CSMT12]. semi-inverse [GYS13]. semi-linear [DNSM11, MF14, POR13, WFZ13, YLZ13]. semi-local [AHM14, RA12c]. Semi-Markov [Gra11, CYG14, SO10]. semi-Markovian [LC14d]. semi-orthogonal [LZ11b]. semi-positone [AZ10, Zha12c]. semi-ratio-dependent [LDZ10]. semi-supervised [ZZF12]. semi-tamed [ZWH14]. semicircle [Sta14b]. Semiclassical [YHH11]. semiconductor [JBDC12, LY12b, Mam11, YY13a]. semiconductors [BK10b]. semicontinuity [GX13, QZS13, SW14d, WWQ14]. semidefinite [BK11a, DKMM11, ZF13a]. Semidiscrete [WG11, Zha11c, ZnZ14]. semigroup [DTP14, OD11b]. semigroups [ADÇ14b, BL11, RBT14, pYmK13]. Semilinear [Kan11, Bog11, Bog12, CB10, yCfW12, Chi10, Ji14b, KD13, Kyc11, MZ13a, MN10, Mop11, Mop14, NV12, zSWZW12, TTQ11, VRMT14, WL11, Zhi11]. Semilocal [PC14b, AH12b, GMR13]. seminormed [AY10, CS13a]. semiorthogonal [SR14a]. semiparametric [GYL14]. semipositone [WLW14, Yan14b, ZLW10]. semisimple [MM14a]. semismooth [LHL11]. sense [AKÖ11, FT10, PK12, Shi12]. sensed [GXSH14]. sensing [JZY10]. sensitive [DGS14c, San11a, San11b, TK14b]. Sensitivity [AM12e, DA11b, WiN12, AHJ14, BM13b, DHQ14, FHR13, FKH10b, FGD13, GGTCdlF13, RRLA14, WTZ11, XXZH10, ZCL14b, ZLY11b]. sensor [CÁGGLP14, CK12b, SKB14]. sensors [GLHCLP12, KCGH14]. separability [BSS12]. separable [Bec13, BA14a, BDK^{+}14, MYL^{+}10b]. separated [CD11a]. separation [CXD14, GM11b, YLGW14]. separators [KMAG12, PHPD11]. septic [KS13b, WZ12]. Sequence [LYL13a, Abe10, AR11, Bal13, Bil14, ÇT10b, CVP^{+}10, DS12b, DM12, DB12b, EVR13, Et13, EKK14, HMSW14, MD13a, Özk14, PM11b, SA11, SK10b, SK12b, YTS12b]. Sequence-dependent [LYL13a]. Sequences [Man10b, AK11, BM11, BR12, Bor14b, Boz13, Bra12, Bum12, CGVR14, DM10b, ES12c, Et13, FTAB14, Guo13b, yLbLjH10, Lu14b, Ota14, PS13, Pet12, PCM14, Sah11a, Sav10, Sol14, XC13, Yay11]. Sequential [KAY12, PQRR10, ZA14a, BMA11, BTT13, CT12c, CGO14, JHX14, PJFL14, VC12, WZZ14, WWC12, ZZZL14]. Serbia [KVP14]. Series [KM13a, Pré11b, WN14b, AB11, AQ11, AAR14, AM13a, BRAB11, BS13, BTT13, BTVC10b, BSR10, BS12c, Boy14, CCJ11, CPKT13, Cha14b, EK10, EMP12, ET12b, ELIGMB11, FY12, GC10a, GGOS12, HM10d, ISG11a, JZW12, Kek11, KKBO12, Kra14b, KT14, LLWW10, LS14c, MMKD14, Nov14, Oul11, PD13, ROA11, Sáf14, Sar10, SS11b, Sri11b, Sul10, TN14, Van11b, Ver11, WW13d, WN11, WHS14a, WHS14b, WS13, ZC13b, Zhi12, ADÇY11, BM14b]. series-plus-linear-polynomial [BS12c]. Serre [Kho14]. server [JS12, KHP11, LWZ14, LK11, Sal11a]. service [MSS14b]. Set [CS10d, LYXP13, Ahn13, BS14b, ĆARA12, DSGB12, DSM^{+}12, DV14, DHQ14, ES12a, FL14, FM14, GFMP13, HH13b, Iid11, KI10a, LZM^{+}12, LSZ12, LL14d, Lim12, LCL^{+}13b, LCF^{+}14, Mae10, Mal12a, Mal12b, MF13, Öza14, PDBRFER14, RBT14, WY12b, WH13, YK14b, YDZ11, Yor14, Yu14, ZS14a, ZLYL10, ZX11b, ZFCW13].

Set-valued [CS10d, LYXP13, LSZ12,

Mae10, Mal12b, ZX11b]. sets

[AT12, AK10d, AK10c, AK13, ARC14, BBCR13, ER13, HS10a, HS13, HW13a, KKVCS12, PBP14, SL13b, TZYH14, Xie11b, Yan10f, ZZY12]. setting [BB11a, BD14d, DKK^{+}13, GS10a, Lin11c, TES14]. settings [Ali12, GK13, TPA10]. settler [DGCFM11]. setup [EG06, SM13, WL14a, XY11]. seven [YS12]. seven-node [YS12]. seventh [Sal10a, SGY^{+}14]. seventh-order [Sal10a, SGY^{+}14]. Several [Bab12, GMSH13, MD13c, AAC11, AMFM12, BC10b, BK12, DRS12a, KKA12, LW11c, LfXfC13, MM13b, NSC12b, WJ12, XX10]. Sextic [Moh14, KK11c, WZ13d]. SFVE [Luo14]. SG [HSX^{+}11]. Shabat [HSI13]. shallow [DOV^{+}11, DC13b, EVDG12, FH11b, FK12, GLW12, LLL12, LLL10b, MGWG11, RYÖ11, RTV12, TD14a, TK12, WTL^{+}11, Waz11a, YH10, dLDP13]. Shape [BM13b, DHQ14, HHM10, HHA14, ZZ14d, AAP10, BAA13, CCLW10, EÖ10, EH14b, KK12a, QHZ^{+}13, VCNV13, WSD12, XBC^{+}13, XBC^{+}14]. Shape-preserving [HHA14]. shaped [AT12, CRP13, KRS10, Li10d, Sar12c, SSS14]. shapes [PP13a, TVZ10, WQX^{+}10]. shaping [LP10]. sharing [HW14b, SLJ^{+}13]. Sharma [Ban14, Pan15, SHY11, Zay11]. Sharp [Aza14, GQ11, KOP11, LW14a, MDÇ13, Mor14a, XCS13, Zhu12a, Hon10b, WS14]. sharp-edged [WS14]. Shattuck [XC13]. Shaw [Cur14]. Sheaf [Mat10]. Sheaf-theoretic [Mat10]. shear [AM14c, BSP11, CS12b, Nar12, THNTNXR11]. shear-thinning [BSP11]. shears [PQR14]. sheep [SZ14d]. sheet [ASM11, AM14g, Cor10, EEEE05, FZZ12, HAMA14, Pan12b, PK18, Van10]. sheets [Fit12]. Sheffer [KASY10]. shell [Dan14, DRS11b, FA10, Naz10]. shell-like [DRS11b]. shelving [dCDNVS11]. shift [BEGM12, Dar11, Kau12, LC13, Su13, TKLH10, Yu12a]. shifted [DBBEE13, GÖS11, GL10b, SGA14]. Shifting [Aub10]. shifts [AKR14]. Shimizu [EDYA14]. Shishkin [MVM10, VT13]. Shiu [LSS11, LC14d, SMWR10]. Shock [MCPO12, GO13, SLZ12, SJ13, THAB11, YA14, Zak10, ZA14b, ZLZY12]. shocks [WW11g, Wu12b]. shockwave [CTH14]. shooting [Cha10, Cov11]. shop [HH11, HL12e, mL12b, LPX12, LYG14, dejW14, XZZK12, YW11b, ZB14]. short [Ach10a, Bar13, CF14, DL12a, DZ12, FCZ^{+}10, LZLS14, TES14, Woź13, XHG11]. short-pulse [XHG11]. short-run [TES14]. short-step [Ach10a]. shortage [PSJGLS14]. shortages [JGM13, SGC12]. shortest [LC11, PR14, SN12]. Shortfall [OGO14]. shot [DGSB12]. Should [BO13, NNPH14]. shows [Koç13]. Shrinking [YD13, Cor10, DHZ11, PMV14, QS14]. shuffled [Aha14, LPX12, MLGdSC11]. Shulman [Sar13]. SI [Sha13b, SW13, TGL14, WWZ10b]. sickle [CCG10]. side [DG10a, EA14, SHA12a]. side-vertex [SHA12a]. sided [CS13f, LXWW14, SMPA13, ZCU14]. sides [BCC11b, HH12, SN11, ZW14d]. Sierpinski [BV12b]. Sieved [MR14a]. sight [LZLS14]. sigma [PD13]. Sigmoidal [RJ12]. sign [Bah13a, CI13, CT14b, DH13a, DSX11, FHLW11, GLZ10a, HT11a, KSM14, LR13a, LLW10b, YP14, YT13b, ZLZ12, ZTZ14, ZL10e]. sign-changing [DH13a, FHLW11, KSM14, LLW10b, YP14, ZLZ12, ZTZ14, ZL10e]. sign-variable [GLZ10a]. signal [GM12, LWS14, MPB14, RY12]. signalized [CTH14]. signals [BHAV10, Cha13a, Che14g, GV14, KUTY14, MMKD14, MRSS11, Rem10]. signed [CGO14]. significant [MG14c]. signless [LW14a, MDÇ13]. Signorini [Cop14]. signs [HT11b, WZ11i]. silicon [LWL12]. similar [FWZW13, HYT14b, Kyc11, San11a, VV10].

Similarity [DNI^{+}10, Ish10, Liu10b, BAE13, CS11d, Džu11, GPW14, IKU10, MKeM13, TPTW14, SGK11]. Simo [CTS14]. Simon [Gar11]. Simple [Woź12, AEESMA10, And12b, Bas13, BIT12b, Cha10, Che11b, CNKS14, DHI13, Fuk12, GK10b, GK11, GKV12, Han14, JPB10a, NSC12b, NS13a, NCS14, Ort14b, hPrLgZ10, hPnZrL12, RY14a, SN12, SH12a, Sid10, SS12a, SM14c, TZS14, TCBL11, hTlCqL10, VOL14, WS14, XDZ14, YW11b, Zay11, ZYZ14a, ZZC13]. simpler [ZW10c, ZW14d]. simplest [BC10b, BCW13, VDK10, VDK13]. simplex [aBS14, LL11a]. simplification [SMR14]. simplified [CF13, KCGH14, Tag11]. simplify [TVZ10]. simply [KAF13]. Simpson [PVV11]. Simulated [CSD12, AGZ12, AK10b, CEM11, KNS14, MT13, Örk13, SKC11, RRRRTR12]. Simulating [Xu10a, GVT^{+}13]. Simulation [CFH^{+}13, PS12b, AAAH12, BAE13, BSØ11, CMT11, Cha13c, CLAT11, DGCFM11, FdOdSM14, GXC^{+}11, GHZ^{+}14, GJGSNRÁ12, Har14, HLMW13, JWZ^{+}12, KP14b, KSS12, KK13d, LPR10, Li12b, Lia14, LFC^{+}12, Liu13b, LSF^{+}13, Mah13, MA13c, MZWZ12, NRWF12, NCANCF12, SATdV13, SVGC10, SP14d, Sie14, TSS^{+}14, Xie12b, XLY14, YAAS10, PHPD11]. simulations [BLK^{+}13, BBFK12, CRW11, CE13, MMX10, WJ14, ZZ13c]. Simultaneous [LR13c, AGHdD13, GCM10, GY12, Jan14b, KRC10, LL10e, MPM14, PC14b, TSN13]. Sinc [EGMEM14, AH11a, AB12b]. Sine [KRS10, BTVC10b, mJ10, LTZ12, ZA11, JCB11, Li10d, JW12, SP13a]. sine-cosine [ZA11]. Sine-Gordon [KRS10, LTZ12, JCB11, Li10d, JW12]. Sinelshchikov [Rya10]. Single [CYY14, LWW14, QK13, WWZ10a, WL14a, WTFC13, XG10, XZY14b, Yan10e, ARS14, ÇBS14, CGLM^{+}12, yDpY14, Dan14, lF12, FPTS14, GM12, GC10c, HLS10, Hu12a, lJCWpW12, JCW12, JHN11, KR13a, LCH13, LT11, LWZ14, LYZG13, PS13, RM12b, TLW11, WCB14, WX12, WS12, Wan13e, Wan13b, WCW^{+}13, Yan10a, Yan11b, YCC^{+}12, YH12, ZWD12, ZHC^{+}14, BM14b, JYYJ10]. single-arrival [WCB14]. single-degree-of-freedom [ZWD12]. single-input [RM12b]. Single-machine [WWZ10a, WL14a, XG10, XZY14b, Yan10e, HLS10, JCW12, LCH13, WCW^{+}13, YCC^{+}12]. Single-period [QK13]. single-pest [GC10c]. single-server [LWZ14]. single-source [lJCWpW12]. Single-Term [BM14b]. single-unit [KR13a]. Singular [CT14a, CLW11, GG13c, Lin13c, Rep11, SS11a, SK13b, SQJ11, YCZ14, AH11a, AM12a, BRS13, Bec13, Bie11, BRM11a, CS12a, Car13a, Cha14b, CJ11a, CZ11b, CJ11b, CC12b, CMH12, Che14j, CZ14b, CHX14, Cho11a, CXG10, CEXG12, Dar11, Das12, DBK14, DL12b, DZZZ12, DW13b, EGOM10, EZRMR12, ECG14, FHLW11, FZ13, FPC14, Fan14e, FFB13, FLZ14, GMR11, GG14e, GS11b, GÇ11b, GAC12, IJ11, ISG11a, JM11, JZZ13, JLW13, Kan11, KXW14, Kar14, KW10, KJW10, Kri14, Laz11, Li11a, LWZZ14, LWYY14, LZ13b, LZ14b, LWCS14, rL11, LZW13b, LZL12, Maa12, Mao11a, NY12, OO11, OO12, Öza14, Pan11, PV10, PKS10, SK12a, Set14, Shi14a, SCC12, Sid10, SY12, SNZL13, SW14d, WLCW10, WHL11, WCL11, Wan11a, WWW13c, WD14b, WLW14, Waz11b, Waz12a, WPSC11, WPSC12, WZ12, XSL10, XZ12, YKI10, YW11a, YSI11]. singular [YY10, ZL10b, ZLW12a, ZLM12, ZW13a, ZHY13, ZLWW14, Zha14b, ZMZ11, ZLW10, CXG10]. singularities [BK10a, KY14, MCH14, Ma14a, OC10]. singularity [CEXG12, ECG14, KX10, MX11, ZYP13]. singularly [AC10b, BP11b, BC10a, BLK11, CG10, CG12b, DN14a, Gen11, GC10b, GO12, HV12b, KG10, KKV12, KCG10, KR10, KSS12, Kum14, LC14c, MVM10, RS11a, RK12, SRP13, YS13a, ZCZ14, ZXZ11]. sinh [LTZ12, HRL14]. sinh-Gordon [LTZ12]. sink [AHMS14, KML14]. sinusoidal [YWW^{+}14]. SIQS [ZHXM14]. SIR [HLM14, CC10c, EMM^{+}12, HLM14, KK14b, LTY11, LLY14a, LJX14, ZMZS12, ZZY14b]. SIRA [Ama14b]. SIRI [GZ13]. SIRS [BR10c, LOKB12, LS14a, ZX13]. SIS [ZZL14, DIAQ13, INE10, KK14b, LS14b, OR13a, WY11c, ZJ14b]. Sisko [HMA10]. situations [FdOdSM14]. Sivashinsky [HBTU10, Kud13b, SJ12]. six [YS12]. six-node [YS12]. Sixteenth [LMMW10]. Sixteenth-order [LMMW10]. Sixth [KA11, AMS14, AHDD13, KS13b, Kim10, LZ13a, MT12b, WW14c, WT10, Waz10e, WGW11]. sixth-order [AMS14, AHDD13, KS13b, Kim10, LZ13a, WT10, Waz10e, WGW11]. size [BLR10, CBTGWW14, CLS12, EEM12, GLSJCBS10, GKC11, INE10, Jun12, LS13a, cLYrL11, LH14b, PZ11, SYB13, TCBL11, ZL10c]. size-dependent [EEM12]. sized [GAC12]. sizes [Lip13]. sizing [BJB11, KBG11]. Skan [ANHA14, Afz10, PP11]. Skew [SBDB13, CL11b, Elo11, GG11c, HYC^{+}10, Kyr14, Rim13a, Rim13c, Son12, WY13a, Yu11b, ZW11a, ZGZ14, Zhu12b]. skew- [WY13a]. Skew-Hermitian [SBDB13, ZGZ14, Zhu12b]. skew-persymmetric [GG11c, Rim13a, Rim13c]. skew-symmetric [Elo11]. skim [GXC^{+}11]. skull [GS12c]. slag [MWAM14]. slag-type [MWAM14]. Slant [OGYE13, KGH12]. slate [MIM^{+}14]. Slater [MÇ12a]. slave [GHG14, BYM14]. sliding [AA12a, GLX13, GWW14, JQ14, Pai14, WS12, YTC13, ZX10, ZLZ14b, ZT14]. slip [Che10c, CRP13, Kha14a, Par12, RRP14, VSS13, ZM14c]. slip-stick [CRP13]. slit [PQR14]. slits [SMN11, San14]. slope [IG11]. sloping [HLT14]. sloshing [KS13a]. slotted [HMD14]. slow [ÇT10b, GLM14, KCG10, SC14a, VO12, VO14, WW10, WWW13c]. slow-fast [KCG10, SC14a, WWW13c]. slowly [ÇÇD10, SZZ^{+}11]. SMA [KDS13]. Smale [AKH13]. small [BRVR12, CL14d, Cor11a, CO13, CDIS10, Fer10b, Her11, Kau14, LC14c, Rem10, eT12a, VCĆV12, WW11c, Xu10a, XC14, ZZ12b]. small-amplitude [Fer10b]. small-world [Xu10a]. smart [KSY12]. Smooth [JB10, AAM13, AM12a, BLR13b, Che14k, CFG13, DHMU14, HXY12, HM10a, KKS11, LR14b, Sha14b, Shk14, SY12, WFL14b, WLL14, XH14, YTC13, Zah11, Zit11]. smoothed [KP14b, NTPVNHLX14, TZZX11]. Smoothing [jL12, ZS14c, AAI14, Eto11, Fan10b, HM10b, HM12, HM14b, JLHW13, KCGH14, KFM11a, LH11b, LWW11, LTL12, LTM10, LLL10a, Nar13, NW10, RX14, THDF11, TQWZ10, XY10, XM11, YCTD11, ZZ13b]. smoothing-type [NW10]. SMS [MMS11]. Sobolev [AM14d, CHS14a, CV12, CGM13, CT13b, DN14b, GLB12, HMPVQ14, Kan11, KR13b, KA10b, MMMMB13, MR11, PPK11, RSN11, SZ11, SGA14]. Sobolev-type [SGA14]. SOCCPs [HM14b, RX14]. social [PPK14]. society [Mis14]. soft [ML10a]. Software [PKD14, Lin12, RTW13, SSST14, UMMT14]. soil [CTLD12, MNOB10]. soils [BHG10]. Sokolov [MG10b, MK13b, BD12b, Hu12a, SV10a, SV10b, SV12, ZTW10]. solar [KV12, SHAO14]. solid [AY11b, LKP12, LZ10c, NTPVNHLX14, OADL14, Sin10a, VO14, ZLX12]. solidification [Yig11]. solids [ZD11a]. Solitary [HA10, LZ11a, PCT^{+}10, SC10b, SHC11, TI10a, THAB11, VD14b, YA14, Asl10b, Awa12, DPC10, Esf11b, FYYT11, HML10, HSZ14, HM10c, Kud13a, Liu10a, LDLL10, NR14, SYZ10, SS10f, WD14a, WTFC13, hXqXlC10, gZrLX10, Zha10d, ZHZ12]. solitary-wave [hXqXlC10]. Soliton [AMB10, BI10, BK10d, BK10c, BK10b, BZGK10, BT10b, BT11b, BTHA11, GB10, HBDC14, IB10a, IPB10, IB10b, JPB10c, JPB10b, KZJ13, Kud10b, LTJ^{+}11, QTL^{+}12, TI10b, TW10b, TB14, WSL^{+}13, XGG^{+}11, Asl12b, Bai13, BCWM10, BPM10, BR10b, BSS12, CYY14, CC12c, DW14, Elb12, GMST10, GX12, GARK12, HCD14, Lat14, LMS10, LL13f, LDLM13, Ma13a, MM14a, MM13a, MGYQ12, Pin10, ROK12, RCYL10, SSF10, SF11a, SMM14, SB10, SB11, TW10a, WDLM10a, WTL^{+}11, WYTL14, WT10, Waz10d, Waz10e, Waz11a, Waz12a, Waz12b, Wei11b, WG12, WX13d, WW14e, WLHZ10, YDXC10, YZZP14a, YZZP14b, YZZ10, Yu12b, YL14d, YY14b, gZrLX10, ZlDxX11, ZCT11, ZgJhD12, ZGY^{+}10]. soliton-like [Bai13, DW14, RCYL10, ZGY^{+}10]. Solitonic [LGS^{+}11, MGWG11]. Solitons [JX10, ÁD10, BZR10, CP14b, DL12a, JCB11, MZ13b, Pan10, TW11b, WDLM10b, Waz10f, WD12, YA14]. Solow [FGS14]. Solution [ARH^{+}12, ÇBS14, EZ14b, GDH10, GSRR13, HG13b, HS12, LHT14b, MHA12, NMR14, PKBD12, SB13, SATdV13, AMIR14, ABEK11, ASM11, AAV10, ARE14, Aja11, ADÇY11, AT10a, AANA10, AS11b, AS14c, AS14d, Aou11b, AP14, AS13a, Asa10a, Asa10b, AMB10, ÁB14b, AB12c, BRS13, BP10b, Bar10, BBS10, Bas12a, BC10a, BGRS11, BS12a, BB11c, BKW11, BK11a, BMMS13, Bes10a, BI10, BK10d, BK10c, BK10b, BZGK10, BT10b, BPM10, BR10b, BT11b, BTHA11, BCKL11, Bot12, EL13, BDK^{+}14, CHC11, CMM10, CP12a, CK10, CWS10, Cha10, CB10, CJ11a, Che11g, CLA^{+}12, CSL14, CLLS14, Chi14, CZL11, CL13b, CLS12, DZLW13, DR11a, DVGY11, DJ12, DL12b, DSGB12, DGSB12, Dem14, DF11, DS13c, DSPZ13, DFG14, DZ14, EAA11, EM11, ESH10, EZ12, EB14, EALS10, EZ13b, FYM13, FF11, FLW11]. solution [FBGH12, FP12, Fuy14b, GLP11, GK14, GLS13, GC14, GI13, GYD14, Goo11, GG12b, GS11a, GT11c, GÖS10, GÖS11, GÖ14, GZXJD10, GL10b, GC11c, GL11b, GLW12, GM14c, GS11d, HN14a, HBTU10, HHU12, Har10, HBB11, mHhL10, HLTT14, HS10b, IJ11, ISG11a, IB10a, IPB10, IB10b, JM11, JS13a, JW12, JMS13, JS14b, Jum12, KP14a, KZJ13, KLBR13, KK11b, KK11c, KS13b, KS11b, KA10b, KS10b, Kia14, KB11a, Koh12, Kol14a, Kos11b, KD13, KBB12, KH14b, Lai10, Laz11, LWZD10a, LX12, LY14b, LL14d, LL10d, LL11b, LYAH13, LK14, LNC12, LDL13, Lin14a, Liu10d, LWZ13b, LTU13, LMB13, LS14d, MH13a, MR14b, MC12c, Mil14a, MHB14, MB13a, MSR14, Mis10, MVGS11, MJ11b, MB13b, MG11b, MT12b, MEG10, MYL^{+}10b, MV11, Naz10, NH12, Neu12b, NN19, OH12, OPD12, Ort14a, Ort14b, PZ14a, PR12a]. solution [PAC13, PK12, Pic12, PBP14, QFT^{+}14, RA12a, RSK14, RK14b, RAIM14, Ray12, RSN11, RLV10, RZ11, RnZ10, RS12, SS14b, Sar13, SSK10, Set14, SYK10, SWW14, SL14b, uiH14, SF13, SY12, Ste14f, SB10, STV12, SB11, SD14, SD13b, SNH10, TK14b, TJZ14, TN14, Tra12, TW10a, UT13, Üna13, VMS10, Wan10e, WDL10, Wan11d, WZ11d, WWH13, WXLW13, WFZL14, WWZ14a, WLQ^{+}10, WFwC12, Wu10c, WLH11, Wu12d, XG13, YDZS11, Yan12a, Yun10a, ZS14a, Zha10c, Zha11a, ZHX11, ZgJhD12, ZW13d, Zha14c, ZZZ14, ZL14b, ZLJZ14, Zha14a, ZS15, ZH14, ZLYL10, ZZW11b, ZZ12b, ZZQ13, ZX14, ZK10a, fuy13]. Solutions [BEdAdSK14, Cha11, DLL11b, DCR12, Han11, HW12, IS10b, KPK14, Ste12r, Ste12s, Ste12t, Yad13, YlTA10, ANHA14, eMBHA13, AG10, AMB^{+}14, ASS13, AK12, AKR14, ABP13, Afu12, Afz10, AYS13, AIAD12, AZ14b, AM14b, AAT^{+}13, AV14, ABF14, AZ10, Aou12, AGKR10, AAH14, AAGR13, AM12c, AM12d, Asl10b, Asl10d, Asl11a, Asl11b, Asl12b, Awa12, AMA10, AM13c, AZ19, BP10a, Bah13a, Bah13b, Bai10, BD11, Bai13, BC13a, BD12b, BMA11, BBJ10, BDE^{+}14, Bec13, BA10, BAP^{+}14, BS11a, BTVC10a, BTVC10b, BVTC13, BR14b, BI14, BC12d, Bou10, BAJK10, CST12, CI13, CT13a, CQJ10, CC10a, Cai11, CCJ11, CV12, CB12a, CYHL11, CL14a, CG12a, CCJ14a, Cel14, CDMS14, CJM14, CLL10, CSLL10, CPZ10, CT12b, CMH12, CM12b, CYM13, CTA13, CLW14, CM14, Che14d, CT14b, CYY14, lChH14, CS13e, CC12c, Con13]. solutions [CK13a, CO13, CSS11, CHS14b, CYZ13, DLFZ10, DLL11a, DH13a, yDpY14, DW14, Dar13, DS13a, DK11a, DGDA10, DK11b, DPC10, Den11d, DSX11, fD12b, DH13b, Den14, DNI^{+}10, Dia13, DRS12a, DRS12b, DS12b, DCR13, DV14, DISS14b, DN12b, Din12, DSH12, DP10, DWP10, DX14a, Dos14, DG10b, DX14b, DH14, Ech12, EGJB13, Elb12, EBR13, EZ14a, EZ13a, FHLW11, FYYT11, lF12, FW14, FV14, FcDDqZ10, FNG11, FLP11, Fit12, FKH10a, FCZ^{+}10, FH13, GL12, GGYW11, GLY14, Gal11a, Gal11c, GS13a, Gao10, GZ10a, GLZ10a, GX11, GWZ14, GP12, GP10, GS14a, GMST10, GZ10b, GH10b, GLZ10b, GX12, GHPR12, GG12a, GKS12, GG14b, Goo13, GARK12, GK12a, GW11b, GB13, GM11b, GH10e, GH11, HB11a, HLM14, Has10, HA10, HMIZ13, HV12b, Hay10, Hay11, HML10, HM10a, HRL14, HVC12, HSZ14, HL12a, HL12b, HL12c, HL13b]. solutions [HL14a, HSLG10, HPB11, HR12, HYX14, Hon10a, HL12d, HM10d, Hsi10, Hu12a, HLX13, HLZ13, HRX14, HCD14, HW13b, HYW14, HBDC14, Hus10a, Ign10, IS10a, ITLA10, IU10, IISA10, Ish10, IS10c, JS14a, Jan13b, Jan14e, JPB10c, JPB10b, JŞ11, JZZ13, JX10, JXA11, JZG13, JL14, JW10a, JB10, JLM^{+}12, JLW13, JLZB14, JZW12, JKB10, JL10, JBDC12, KK14a, Kam14b, KC12, Kar13a, KJ10, KS11a, KG14, KR12, KA10a, Kha11, KW13, Kon14, KL10, Kov10, KT14, KS11d, KAB11, Kud10c, KRS10, Kud10b, KS12c, KS12d, KK12c, Kud13b, KR14, KPW11, Kum11, KSM14, KM12c, Kyc11, Kyr12, Kyr13, Lat14, LMS10, Li10c, Li10d, LWL10, Li11b, LF11b, LSJ11, LYZ11, LLJ11, LZ11a, LTZ12, LS13b, xLhYlZ13, LDWY14, LHC14, LWB14, LKY14, LSLL14, LGZ14a, LZLB14, LZaY14, LZL14b, LS10c, Lia11, LWD10, LGS^{+}11, LLL13]. solutions [LZH10, Liu10a, LLW10b, LLWW10, Liu10b, LJW10, LZ11c, Liu11d, LZKK11, LL11e, Liu11f, LJKU11, LCH11, LH11c, LY11b, LTJ^{+}11, LTW12, LLL12, LL13f, LY14d, LZ14c, LLS14, Liu14c, Liu14d, LGZ14b, LXGM14, LC14b, LWZD14, LZSD14, LP13b, LLR10, LS12c, LLL10b, LCLW10, LZX10, LDZ10, LL10i, LDLL10, LDLM13, LM11, LL12g, LY14e, MAA11, ML12a, MXH12, MZTT12, MYW13, MCH14, MCL14a, Ma14a, MKeM13, MDR14, MCM10, Mal12b, MV14, MN12, Mao11b, MM14b, MCdSS11, MTT14, MV12, MM13a, Mel14, MHLL11, MGYQ12, Mer14, MZX10, MM13b, MR11, MPSD12, MMS– D13, MG10b, MJ12, MESAN13, MK13a, MN10, Mop11, Mos14, NV12, NCL11, NR14, NWZ13, NA10b, NMDWAS10, OM11, OO14, ÖZ10, PSP10, PP14a, PV10, PRS12, PEP14, Par11, Par12, Par10a, Par10c, Par10b, PD12, PC13b, PX13, POR13, Pin10, PR10, PPV14, PV11, QW10b, QW10a]. solutions [QS11, QS14, QW11b, QHL10, QTL^{+}12, ROK12, RV14, RNZH13, Ram10b, RGL10, RK14a, RP13, RM12a, Rep11, Reu14a, Reu14b, RSRA14, RD14, RRM12, RWkPk10, RB11b, RCYL10, Rya10, RSK11, Ryo10, SSF10, S10, S.14, SHT14, SvB10, SR14a, Sal10a, Sal10b, Sal10c, SF11a, Sal11c, SHMCM12, SPV10, SM14a, Sha10a, SHY11, She10, She12b, She12c, SDL10b, SD10, SXL12, SW13, SZZ^{+}11, Sie10, Sin12, SS12b, SS14e, SMH14, SYZ10, SS10f, SY10c, SC10b, SHC11, SL14c, Sta12a, Ste12c, Ste12l, Ste12e, Ste12f, Ste12g, Ste12h, Ste12m, Ste12u, Ste13a, Ste13j, Ste13b, Ste13k, Ste13c, Ste13d, Ste13l, Ste13o, Ste14b, Ste14c, Ste14d, SAAS14b, SDS14, Ste14e, SS13d, SHM12, SV10a, SV10b, SV12, SK10c, Tad10, TM11, TMM13, TMM14, THH10, TMXG11, THZF12, TSG10, TIT13, Tch11, TKC12, TV13, TZJZ12, TW13, TC14a]. solutions [TRE13, TI10b, TW10b, TI10a, THAB11, TB14, Tur14a, VSV10, VV10, Van10, VDOG13, VDK13, VD14b, WDLM10a, WDLM10b, WWS10, WLH10, WW10, Wan10f, WF10, WGS10, Wan10c, WLY11, WW11c, WZ11b, WCZ11, WW11e, WqXlC11, WTL^{+}11, WF11, Wan11a, qWH11, WCL12, WL12a, WW13g, WY13a, Wan13e, WCZ13, Wan13d, WFL14b, WCL14, WW14c, Wan14a, WFL14a, WYTL14, WWWJ14, WJ14, WLW14, Wan14f, Waz10c, WT10, Waz10d, Waz10e, WM10, Waz11a, Waz12a, Waz12b, Wei10b, Wei11b, WQ13, WTFC13, Wen10a, Wen10b, WGW11, WG12, WX13d, WW14e, WN11, WN14b, WZDW10, Wu10b, WFDL11, WW11f, WZ11i, WC11, WZL12, WE13, XL11, XC10b, XC10a, Xia10a, Xia10b, XW13, XP12, XZT10, XW10, XHG11, XZH13, Xin10, pXqLlZ10, XLZ10, XCH14, XQ11a, XM10, hXqXlC10, Xu11, XGG^{+}11, XZ12, Xu13b, XL14, XC14, Yan11a, YWT14, YP14, YZD10, YW11a, YMG11, YL11d, Yan12e, YZ12a, Yan12f]. solutions [Yan13b, YZZP14a, YZZP14b, YH14c, Yan14b, Yaş10, YZ10, YZ11, YT13b, YYJ14, yYyL10, YLGW14, YZZ10, YYY10, YuRI14, YM10b, qYqLtW10, YM10c, Yu12b, Yu13b, Yu14, YY14b, Yua10, YHS13, YWD13, YWCH14, YS13b, Yüz13, YS13a, Zah11, ZA11, ZOA10, ZH10a, ZTW10, Zha10b, pZLC10, Zha10a, yZhLjMzC10, gZrLX10, gZrLfM10, ZL10b, ZlDxX11, ZLGZ11, Zha11b, fZwXyZ11, ZCT11, ZW11a, ZYC12, ZLW12b, Zha12a, ZgJlK12, ZLZ12, ZG12, yZqLwW12, ZS13b, ZjHO13, ZR13, ZC13b, ZLWL13, ZT13, Zha13b, ZZ14b, ZTZ14, ZWS14, ZM14a, Zha14d, ZSL10, Zha10e, ZGZG10, ZZ10c, ZSHZ11, ZWZ12, ZHZ12, Zha12c, Zha13e, ZF13a, ZZR14, Zhe11, ZMG13, ZS10b, jZzFbL10, ZTF10, ZCL13b, ZnZ14, ZC14b, bZpZJ10, ZL10e, ZSD10, ZGY^{+}10, ZLW10, ZL13f, ZYZ14c, ZL10g, uRE13]. Solvability [BZ11, BDL13, CDS13, CT12c, CLL14, Jia14, KCBH14, LW14e, XGM10, Ami14, HM14a, SMNN14]. solvable [Ste12j, Ste12p, Ste13h, Ste13e, Ste13f, TYT14]. solve [AKP13, AHVR14, CFGRV13, IN10, JPM12, Liu12a, MQ14, SDK10, Wan11c, WTH^{+}14, WP14, YL11d, ZZ14a, Zuo10, ZZ11c]. solved [DM13b, DKU13, Ram10a, Ste13i]. solver [BNBV10, CKSV14, CLZ11, Koç11, Koç14, MAD11, YNBV11]. solvers [AT10a, MC14a, NP10b, NBYHV11, PPD10, Pet14, lQZWC14, SS11c]. Solving [ABD^{+}12, AASS13, BP13, BDG13, BCEAV12, BMR13, BRW13, BR13, COPR10, CX11, CD11a, CP12c, CJRV12, Eto11, FTAB14, FL13, GL10b, Jat10, JC13, Kuj14, MFZ12, MZ12, MIM^{+}14, MN14a, MVM10, ÖTC11, QW14, RWBA13, Ser13, VA12, WZZL12, WLHZ10, YS10a, Yun11a, ZHH^{+}13, ZL10c, eMA13a, eMBHA13, AISS13, AH11a, AMS14, AHRV11, ACCT14, And12b, And13b, Ard13, ACH10b, AG13, AS12b, AAS14b, BA11, BTS12, Bhr13, BFXZ10, BBK13, BA14a, Bog11, Bog12, BRM11a, BRM11b, BX12, CR11a, CW11a, CLCL11, Cha14b, CZ11b, CJ11b, CLW12, Che13a, CC14c, CHC12, CD13, CTY11, CAP14, DS13b, fDxZxChT12, DBBEE13, DJT10, Dri13, DSZ14, DR11c, DPP11, Džu11, DPP12, ESV12, EGMEM14, ESEEKH12, ESAH10, ES12a, EEA11, EH14b, FC14, FLZ14, FFOC14, Fra10, FG13b, FG14b, Fuy14a]. solving [Gen11, GK13, GKW14, GSGN11, GT11b, GGÖS11, GWML12, GK12b, GS14f, GMV11, HKI12, HSW11, HM10b, HH13a, HHM14, HRAA12, Hou13, Hsi14, Hu10, HKH10, lHyLfG10, HM12, HLW13, HM14b, Jan14a, JK14b, JLHW13, JSF14, KG10, KDTSN12, Kim10, Kle12, KC13b, KK13d, KZS12, La 10, La 13, LW06, LZ10b, Li10b, LZLW11, Li14a, LYD10, LWH14, Liu10d, LW10a, LTM10, LL10f, Liu11c, LLC11, LW12b, LL13e, LSW14b, LZ10f, LZ10g, Lu12b, MC14b, MH14b, Mar11, MP14c, MIHS14, MEBYG11, MG11a, Men12a, Men12b, MH14c, MJ11a, Moh14, MAHAS14, MS11, Nar13, Nem10a, NW10, NL12b, NK12, NNK14, PID10, PDN11, Pet13, PNPD14, PK12, PB14b, PEMS13, QFCZ13, RB14, Rah13, RP14a, RWB10, RA10c, RA12b, fRbLJyS13, RJ12, Sab13, Sab14, SWL12, SSK13, SSCA13, SMIS14, SGG12, SA13, SAP14, SPK14, SAK10]. solving [SGK14, SSV13, Suh13, TS14, TDZF14, TSP10a, TA11, Tem12, Thu10, Tom12, Tom15, TQWZ10, TT13, Udd13, WLCW10, WZZ11, WaZZW11, WCL11, WDZ12, WDL12a, WF12, WW13c, WDT^{+}13, WMM14, WRD13, Wil10a, Wil10b, Wil13, WR10, WLD11, Wu12a, XM11, Xu10c, XWZF11, XD11b, YKI10, YCTD11, YCXL12, Yan14a, YHW13, YCK13, YH14d, YSW11, Yus10, Yüz14, ZM12, ZY13a, ZLG^{+}14, ZZH11, ZLH11, ZW11c, ZGY14, dLDP13]. Some [AGT11, ANR10, ADÇ14b, Aou10, Bal13, BSS14, Bel14, Cai11, CS13b, CCCRR13, CB12b, CS13f, Cur14, DB11b, DFGL14, DCK11, Dra14, EGJB13, EKK14, FMF13, GAV14, Gu10, GM14b, GÇ11b, Guo13b, GA11b, GS11c, HMZ14, Hus10b, Hwa11, Hwa14, IISA10, Irm11, JVK10, KG14, KASY10, KYH10, KK13a, KK14c, KZ14c, KSD13, Kra14a, LLSY11, LH14a, LW14b, LN10, LLZ14b, LWM11, Lu14a, Lu14b, LS11c, LMA14, Mal13a, MS13f, Mez13, MMY11, Neu11, Öza11, PS11, PAA14, QCK10b, QW11a, RT14, RP14a, Rho14b, Rho14a, Rho19, RPA14, RCYL10, SDAN14, Sch11a, SGG12, SJPS14, SK13a, SDO10, Sri11b, Sri14, SC14b, Sta11, SH13, TC14a, Tom12, Tom15, UOK10, VGA12, Wan10c, WLX11, WZG12, WDL12b, WWT13, Wan14d, WSK12, Wło13, XXS14, YHWZ11, YL11f, ZCZ13b, AB11, AY11a, Afu12, AR13, AIAD12, ABGGS13]. some [AAA13b, Aou11a, Aou12, Aou11b, Bai10, Bar13, BS13, Bel11b, BS11b, Bir13, BCF13, Bry11, BC12e, BFB12, CGGMG12, CRR^{+}13, CL11b, CG12a, Cen11, CL14b, Chi10, CBS12, CLS12, CS14b, CMM12, CT13c, CIP10, CISM11, CK12c, DR10a, DI13b, DN12a, Dia13, Dia14, DN12b, DM12, Džu11, EK10, EMA13b, EMA13c, FMS11, FcDDqZ10, FGRLSV14, GLP11, GL12, GS12b, GGPT14, Gem11, GNS14, GSN12, GKS10, HMSW14, IS12, Iri10b, JX10, JLZB14, KOP11, Kek11, KWS10, KI11b, KKVCS12, KKS12, KET10, KPP^{+}13, Kyr10, Kyr12, Kyr13, Kyr14, LYZ11, LGMZ12, LJT13, Lu12b, LS14e, MDR14, MCM10, MD13a, MN12, MH10b, MN10, MK10b, MCL14b, NN13, NBYHV11, ÖTC11, PC11, RB14, RGL10, Ros11a, Ryo10, Sad11, Sah11b, SI11a, SS14d, SSZ11, SW11, SY14d, Ste10f, Ste10i, Ste11e]. some [Ste12c, Ste12f, Ste12i, Ste12r, Ste12n, Ste12o, Ste12p, Ste13b, Ste13c, Ste13d, SAAS14a, Ste14d, SDS14, Ste14e, TQ11, TRA13, TJ14, Tos14, VV11b, WV14, Wei10b, XP12, XQ11b, YL11d, YuRI14, ZCU14, Zho12c, ZS14e, ZX11b, dF14a]. Sommerfeld [Fuk14a]. SOR [Fan13, LZ14a, Men14, ZS10a, ZM14e]. SOR-Like [ZM14e]. Soret [HCCT14, PC13a]. source [AHMS14, CLAT11, CL10b, CYZ13, IKL11, lJCWpW12, KM11c, Pan15, Sha12b, SM12, SWW14, SL14d, SS10g, VJBDV13, WCL12, WL12b, WS10, XWWC10, YGL11, Yan11c, ZW13e]. source/sink [AHMS14]. sources [CSLL10, HMMK11, LL10e, Sun10, WGS10, WX13a, WLHZ10]. southern [WDBM14]. Space [STM10, AAAH12, AASS13, AA12d, BTRZ14, BDDK14, CHS14a, CNTG12, CL14a, CKK12, CS10c, CG14c, CR12, DM13b, DB12b, EKK14, For11, FKH10b, GxFX^{+}12, GM14a, GL14b, GMKV12, GMV12, GS11c, HdCOL14, fHcXH11, HMMK11, ITLA10, Jar10, JK13, JFHZ14, KGH12, KG14, KGM14, Kyc11, La 13, LMZ13, LW13b, LNC12, LZ10d, LH11c, LJ11, MLM^{+}14, MVS14a, MG11b, MS12b, MG14c, MK14b, MSS14a, Mos14, NRWF12, Ols12, PP13a, RMY14, RY14a, RDD14, Rec13, fRbLJyS13, RPW14, SGJ10, SN12, SLZ12, SEL14, Sin10a, Sin11b, SK10b, SK12b, Ste10a, Ste10b, Ste10c, Ste10k, SU10b, Ste10m, SS11e, Ste11e, Ste12d, Ste12w, TÜ12, Vis14, WG11, Wu14c, Yan12d, YTY^{+}14, YLAT14, YuRI14, YLTB12, YLJ14, Yue14, ZA14b, ZLA10, ZLZ^{+}14a, ZL12c, Zhu11a, ZZHZ14, dCCL14, dOL13]. space- [LH11c, Mos14]. space-dependent [HMMK11]. space-fractional [ZLA10]. Spaces [NMR14, AKR10a, ARN10, ANR10, AKN11, AN– D12, ANR12, ANR13, ARR12, AT12, AKZ14, ABR^{+}11, AM14d, AKR11, ACK14, AG14f, BT13, Bal13, BM11, BCK14, BS10c, BB12, Bor12, Buo10, CV12, CFS11, CCLF13, sCYW10, sCCL11, CLCL11, CWL^{+}12, sCCYW13, sCWW^{+}14, CWWC14, CIVS14, CCA11, CS13e, CT13b, CM12c, ĆASH11, ĆSAV11, CZSV12, DV12b, DV13, Din11, Din12, DM10b, DM12, DHS10, – Dor13b, DCK11, EVR12, EVR13, Et13, Fer10c, FLP11, Fer13, GIR10, GS12a, GM11a, GG11b, GZW12, HDK^{+}10, Hao10, HMSW14, HA11, HKK12, ICKV14, JS10, JW10a, Jun10b, KKS11, KAS12b, Ker13, KI12a, KudKL14, Kim14a, KD13, KPW11, LAR13, LZ12a, LKY10, LS10a, LS10b, LjH11, Lim12, yLbLjH10, LHCI12, pLjH10, Mal12a, MD13a, Man11, MR11, MMS– D13, MVS13, MAK13, NV12, NKR12, NA14a, NS11]. spaces [OP12a, PG11, PPK11, PS11, RRR11, Rad14, RRG11, Ren14a, RPA14, Sab13, Sab14, Sad11, SA11, SMNN14, Sav10, SS14d, She14a, She14b, SSRA12, Sin11a, SK13a, SKC10, Sta14c, Ste10a, Ste10d, Ste10j, Ste10g, SU10a, Ste10h, Ste10l, Ste10n, Ste11a, SS11e, SS11f, SSB11a, SS11g, SSB11b, Ste11b, Ste11e, SS11h, Ste12d, SS12c, Ste12i, Ste12v, TR13, VR12, WFZL14, WY12c, WWW12, XZ12, YM10a, Yan11e, YZK12, YK12, bZlD11, Zhu10e, ZX11b, GAV14]. Spain [NA14b, MSGN^{+}14, NCdCDM13]. Spanish [CSSSV13]. Spanning [CP12c, LTT^{+}13, WZZ14]. spare [WT11c]. sparse [LWS14, ZS12]. sparseness [MCM13]. sparsest [Zha13e]. Spatial [LWWL10, SK12a, ZL10a, BA13, Bra13, ES12b, FS12, Gal11d, Gui14, GTX11, Liu14b, LXLY14, MLY12, SZS^{+}12, Yan12b, YZ14c, ZLY11a, ZX13, ZWLW10]. spatially [HWPY14, IS10b, KLKA14]. spatially-dependent [IS10b]. spatio [YZ14c, ZS14b]. spatio-temporal [YZ14c, ZS14b]. Spatiotemporal [YLWU13]. Special [KS12b, Kud12, SS10a, Sta14c, ZCT11, BSS14, Boy13, Bry11, BZT11, Cen11, CBS12, DGMR10, Din13a, EMA13b, GBD^{+}12, Haj11, HW14a, IG11, Koh13, Kud10b, LMZ13, Pic12, RP14a, SMDZG13, SMIS14, Udw10, XQ12, ZLGZ11, CCG12b, DSZ12]. specially [XQ11b, Zho11b]. specie [DMGK13]. species [ADD12, DRSG14, DWC13, DX14b, EWS10, GZ14a, GGM^{+}13, LWTC10, LZ11c, LYB12, LF14, MLY12, PLF13, Sam10, SY10a, VDL12a, WW11a, Wan13e, Wei10a, YL11e, Zha13b, ZLD11, ZWZ12]. specific [GMX10, Mur11, Mur13, RY12]. specifications [CFN12]. spectra [Alt11, Aou11b, DB12b, Gus12, ZWZ13]. Spectral [AST12, ABBC10, Boy13, CPKT13, DGM^{+}14, EMP12, GHPR12, JPI^{+}13, KKBO12, LJ12b, LZX14, Rah13, RDKU14, RAIM14, ZW11c, AB12c, BK10a, BDR12b, BES10b, BX12, Boz13, CP14b, Das11, DAEY12, DBBEE13, DC13b, DAA13, GPZ12, GKW14, GD10a, Hil12, HY14, Hua11b, HG13b, İpe11, Ito13, KY14, KT14, La 13, LW14a, LZDS12, LGN10, LNG12, LLZC11, MDÇ13, Mah13, OC10, Sab13, Sab14, SMI12, Sol14, SB12a, TH14b, Tun13, WWZ13, XFS12, ZW10b, ZJ14c, ZC11, Zhu11c, ZR14, dV12a, DHMU14, EVDG12]. spectral-collocation [ZC11]. spectral-element [Zhu11c]. spectral-Petrov [DAEY12]. Spectral-plus [Boy13]. spectral-type [LLZC11]. spectrally [Gat14]. spectrogram [GV14]. Spectrum [AKZ14, jASzZ11, jAzBS14, MCR14, SK12b]. speed [CTH14, LLW13, WY10a, Xia10c]. speeds [CWW14, GM13b]. SPH [TDT12]. sphere [CP14a, LP13d, Reu14a, SF13, SY14c, VO14, dlRS14]. spheres [Sab13, Sab14]. Spherical [PPL12, Abb14, BDGP14b, BS14b, CT12a, CTLL11, Meh11, Meh13, MVGS11, Naz10, SA12a, SWD11, Tom13]. sphericity [MC12b]. spheroidal [Dha13, KM12b]. spill [YZ14d]. spill-over [YZ14d]. Spin [IBS11, CGGZ11, CO13, IS10b]. spin-coating [CO13]. spin-torque [CGGZ11]. spindle [WY10a]. Spiral [San14, Amd10, ACG13, MWJW10, WMJW11, ZLZY12]. spirallike [XLS13b]. Spline [PT14a, AMA12a, AMIR14, AMA14a, ABJ10, ARE14, AAM13, Apa10, AS13a, CF12, CF14, fDxZxChT12, GK12b, HWPY14, JM11, KK11c, KS10b, KS10a, KS12a, KS11e, LMZ13, LKS14, Lin14a, LL10f, LLC11, MMX10, MMSZ14, MB13b, MB14, Moh14, MG11b, MG14c, SR14a, SHN10, WZ11e, WLH12, WZ13d, XwK11, XFW^{+}12, Zah11, ZHY13, ZK10a]. splines [Beh10, BJS14, BLR10, CHS14a, Isa10, KFM11a, KS13b, MJ11a, MJ12, MYL^{+}10b, NTdCJL12, NA14b, QHYW14, RV10b, ZM10a, ZM14b, Zha10f]. Split [CLC^{+}13, LZ10a, WZ11e, YPSW11, GMR11, Guo10, Guo11, LWL12, SLM14, WYY11, WWZ13, ZLL14, ZZY12]. split-merge [SLM14]. Split-step [WZ11e, LWL12, ZLL14]. Splitting [SF13, YH14a, BS12a, CHX13, CMX14, DMR10, GT13, GT11a, GT11c, GZF12, GFZ13, Hou13, HD14a, KM14b, PW10, RW14b, Wan10a, WYW14, Yin10, ZGZ14, Zhu12b]. Splitting-based [SF13, YH14a]. splittings [Hou14]. Sponge [VDFV12]. spray [NBG10, NAL12]. Spread [MM13c, Goy14, WC14a, YY12]. spreading [PS12b, RG14, pZJ12, ZD11b]. spring [LL13g]. spring-mass [LL13g]. Sprott [HCC^{+}12]. Spurious [SBR^{+}12]. SQP [JW10b, WZZH11, ZZG10, ZP13]. SQP-filter [WZZH11]. square [CCJ11, CJRV12, DZZZ12, GK12c, Hua11b, LL10c, LG11, OR11a, PTW11, SK14a, Seg14, TK13, TS13, Tri14, Wan14a, YH14a, YWY^{+}13, YK11b, ZZQ13]. Squared [Tom13, GI14, ZDZ10]. Squares [JJL14, AM14e, AE13, BDR12b, CBO12, CF13, Cor11a, DR10b, FTAB14, Guo10, Guo11, JWW14, Juk13, Kyr12, Lee12, LWZD10a, LGZ14a, LL11b, Mea13, MMS– D13, PX13, PZ11, QW14, RMN14, RA10c, RCH12, RPW14, Sal11b, SZS13, TW13, WSCH14, Wu12a, XLY10, Yam14, Yua10, ZLGZ11, ZJD12, ZL14c, ZJ14a, ZZX14, Zit11, LIC12]. squaring [SCIMM11]. squeeze [KBB12]. Srivastava [Aou10, Can11, CS13b, CKS10, Cvi10, Deo12, EDD11, Kir11, LLSW12, LLLS13, Öza11, PPS13, SKW11, WZW12, XXS14, Yad14, YL11b, YL11f]. SRLW [XHXH12]. SSLE [dMbJmY11, ZHW11]. SSOR [RY10]. Stabilisation [KXW14]. stabilities [FO11]. Stability [AA12b, BBI13b, BC14b, Cel14, CKN13, CJ14, CZZ12, CCS13, DGS14a, DE12, DRGR10, Don14, EMM^{+}12, FS14, Ge14, GGM^{+}13, HL11a, HYT13, HSZ14, HLC13, KHA12, KJ12, KZ14b, KB14, LPL^{+}13, LL14b, LZZL10, LXK13, LL13b, LLY14b, LX11b, LXC12, LZ12c, LWG13, LZLY14, LLS11, MLY12, Mao12, zM10b, MHXyY14, MG12b, NC13a, OP12a, PP14a, PPMCG14, RS13, SY13a, SS12b, SL11b, SHZ14, Ste14f, SC13, TW11a, VDL12b, VVK12, VKV14, Wan13d, WL13d, Xie11b, XTL10, ZSZ12, ABRT14, Ach13, AC10a, AZ14a, All11, Amd10, AM13b, AM14d, ASMA12, BMR10, Bal14, BC12b, BJ12, BC14a, BBI14, BH12b, BCÁV14, BN10, BJ11, BK12, BC12e, BPX14, BR10c, CGL12, Cao10b, CDR13, Che10d, CS11b, CWLC13, CS13c, CZZ14, CW14a, CZZL13, CP10, Das12, DdGL10, DI13a, DMW13, DLSS11, DZZZ12, DG10b, DZZ14]. stability [DYL14, Dub13, DA11b, DZ14, EGO13, FXL14, FP13a, FA12, FA10, FdBCC14, GWZ14, GGFM14, GSH12, GYBR10, GS12d, GGXH12, GL13b, GSDW14, GH11, HN14a, HZ13, HLTT14, HX14, HRX14, HYYG12, Ign10, Ign13, Iri10b, JZ13, Jan14b, JPK12, JQ14, KAS12b, Kot14, KTS11, KLPC12, KPP^{+}13, KPP^{+}14, LOKB12, LLP14a, LP14, Li10a, LLY11a, LX12, LCH12, LYWL12, LAF13, LFR14, LSL10, LDL13, LB14c, LLP14b, LdS10, LYL11, LL12f, ML10a, MZW14, MH14a, MI10, MA10a, Mao11a, MPSA14, McC10, MPS14b, MA10b, MA13d, Mel14, MAD14b, Mey14, Mil14a, MEL13, Mur14, NM14, NP10a, Ngo12, NW12, NN19, Oni11, yOZ14, PKPL11, PJ12, PR10, PR12b, QaC13, QLCF10, QHL10, RM12a, RN10a, SvB10, SHL10, SKSK12, SSP11, SAL12, Sha11b, SD10, SZZH13, SZ14b, Sie14, SM14b, Sin10b, Sin11b, SHI13b, Sta14a, Ste10e]. stability [Ste10f, SD14, Sun11a, SS11i, SHM13, SZ14d, SSM11, TZ11, TLZZ12, TX14, TXX14, TJ12, TJ13, TJ14, Tom14, WL10a, WFW10, WS11, WY12a, WL13a, WFZL14, Wan14a, WC14b, WF14, WLQ^{+}10, WLZ12b, WHZ14, WD14c, XX14a, XSX12, XCH14, XCT10, XLW11, XYL14, YZW^{+}10, Yan13c, YTC13, YL14a, YH14b, ZPXX14, ZZC10, mZT11, ZaFT11, ZLJZ14, ZSL10, ZWZ12, Zha13c, ZYZ14b, ZZY^{+}14a, ZW10d, ZCY14, ZXF14, Zhu10d, ZLY11c, ZZS14, ZWH14, Zuo13]. stabilizability [BN10]. Stabilization [LLZY10, Liu10c, Sun11b, ABE^{+}12, AAA13b, AGH12, BNP11, CS13c, Che14k, GY13, GLX13, GM13b, LLX10, LHS12, LXC12, MA10a, Mao11a, NSZ^{+}14, SKSK12, THP14, WF14, XWC11, XSC11, yZ14a, ZSL11, ZYLS12, ZWZH13]. stabilize [BBF12b]. Stabilized [EJ11, CFZ14a, CFX14, HJVC14, ZZL13]. Stabilizing [GWLZ10]. Stable [AY11c, Bho10, DL11, LYXP13, CJZ12, CF13, CFZ14b, CDST14, DL13, GRV13, Jun10a, KBAvBW14, LK12, LD13c, MDR14, PVV11, SS10g, WY10b, Xie12b, YZC14]. stably [BBFK12]. stage [AE13, BBC14, BNBV10, CCK11, CJK12, CXL12, CWLC13, CCWM13, DWP14, EWS10, GC10c, HL12e, JDN14, JSH10, JYF13, KJ12, Kha14b, KBG11, Liu11b, LZLY14, MYJ10, MG14a, MHXyY14, SJL10, Sal11a, SL13a, SHM13, SLL14, WH14, YNBV11, YH12, ZL11a, ZW14c, ZZZL14]. stage-structure [GC10c, SHM13]. stage-structured [CWLC13, CCWM13, EWS10, JDN14, JSH10, WH14, ZW14c]. staged [CFL14, LH14b]. stages [SG11]. staggered [CYZ14b, DC13b, TD14a, TD14b]. stagnation [PMV14, RRP14]. Stalford [FNV10]. Stancu [AGK13, AKS14, GG10a, GG14a, Içö12, VGA12, WYZ14]. standard [AGNAL14, DAV12a, DAV12b, Len14a, NBR10, Qui10, RTPH14, SS10a, DAV14b, SL13b]. standby [ET12b, WZ11a, WW11g, Wu12b, YGX14]. star [AT12, DD11, HLTU12, JM14a, LX14b, PM12, Yan10c, Yan10d, Zho12a]. star-shaped [AT12]. Starlike [PP12a, BAR11, BÖ11, ÇDO11, CNRY10, CRG11, Kwo11, SKW11, UOK10, WLX11, XCS13]. starlikeness [AJR12]. stars [EN11]. start [TLW11, Yan10e]. start-time [Yan10e]. start-up [TLW11]. starting [EHRV12]. STATCOM [JLLD13]. State [LL10b, AZ14a, APJ14, BBDS14, CÁGGLP14, Che10d, CTS14, DGCFM11, DWP10, DHDB12, Fan14e, FH13, GG12a, GP11a, HHHS11, HMIZ13, HB11b, IS10a, JŞ11, KK12b, KV12, KPP^{+}13, LPJB12, LXC14, LLX10, LGZ14b, LLKP14, lL10g, MLM^{+}14, Mao11a, MCR14, PP13b, RNZH13, RSPK13, RZ12, Sin11b, SS10g, WCB14, WC14c, WW11h, XPLZ14, ZZY^{+}14a, ZYG14, dS10]. state-dependent [APJ14, FH13, LLX10, dS10]. state-space [Sin11b]. states [CR12, HL14b, IS12, MH14a, TPT12, Üna13, YY10]. Static [EKSO14, KWJ13, KPP^{+}14, LLKP14, SC13, WD14b, ZLM12]. Stationary [Kha11, SSK13, SL14a, ZWLW10, ABE^{+}12, FHWZ12, IS10c, LY10a, Liu14a, Luo14, Neu12b, TV13, WS10, YZD10, ZL13b, ZZY14b]. stations [JHZC13]. statistic [MC12b]. Statistical [EDD11, HdCOL14, MKK14, BM13a, Bel14, Bra12, BSM14, CCJ11, EMK13, GJGSNRÁ12, dALORPGC14, MH11a, MKEG12, MKSN13, ÖD11a, PB13a, PS13, SMR11b, Zil12]. statistically [LZZP10]. statistically-constrained [LZZP10]. Status [Ber12]. Staudt [AAS14a]. Stavroulakis [SM06]. STD [Qiu10]. Steady [DGCFM11, ARH^{+}12, BBDS14, CFX14, DWQG12, DWP10, DHDB12, GG12a, HJVC14, HHHS11, lL10g, MH14a, MMS14, PP13b, QMZ14, RY10, VO12, VO14, WCB14, ZZL13]. steady-state [DWP10, HHHS11, lL10g, PP13b, WCB14]. steepest [Boy14, CCA11, Jia13a, LZGZ11, ZF13b]. Stefan [BT10c, CT12a, Reu11, Reu14a, Tad12].

Steffensen [PC11, CT11c, LZZ10, PPS13, ZZZM10, ZZH11, ZLH11]. Steffensen-secant [ZZZM10]. Steffensen-type [ZZH11, ZLH11]. Steklov [BY11b]. stencil [AS14e]. stenosis [NAHH11]. Step [DS14e, Abh14, Ach10a, AG14a, AAS12, AMR13, BIT12b, Chu11, DEP12, DZ13, DR11b, Fan10b, FG14b, GK10b, HMD14, Jat12, Kar10, KAÇ^{+}11, KM14b, Kim10, LWL12, LTHG14, LTM10, LS11a, cLYrL11, MT12b, MS12b, MK14b, PCT^{+}10, PRK13, Pet14, RP14a, SL14b, TSS^{+}14, WZ11e, WWZ13, WYT13, WS13, YNBV11, ZLL14, ZL10c, dCCL14]. step-size [cLYrL11, ZL10c]. Stepanov [MB13a]. Stepanov-like [MB13a]. stepped [DW13b, MP10, WW13e]. stepping [CRX12, MC14a, OP14b, RV10b]. Steps [UMMT14, HS11b, Xia10c]. stepsize [MN14a]. Stević [Iri10a]. Stević-type [Iri10a]. stick [CRP13]. sticking [LXWW14]. Stieltjes [ABRT14, Alo13, DHM14, GT10b, Jan14e, ZCL11, ZLWW14]. Stiff [GG13a, Abh14, XZY14a]. stiffness [AS13b, DZ14]. Stiglitz [TES14]. still [AM14c]. Stirling [DGMR12, DP14, EDGC14, FPP11a, Lin14b, LS11c, Mor10a, Mor10d, Mor10e, PG11]. Stirling-like [PG11]. stirred [GZ11]. Stochastic [Bai13, BC11b, Cha14c, CZD^{+}14, DG10c, EES14, GR11, GL13c, LZaY14, LLZ14c, LZL12, Tsa14, VJ11, ZaFT11, ZW14a, ZZY^{+}14a, ZHC11, AZ14a, Alz14, Ama14b, And12a, AL10b, AGKR10, APJ14, ÁB14b, BPDZ12, BLZZ14, BC12b, BBC14, BM12b, BM12c, BSR10, BW14, CRR^{+}13, CYHL11, CBLC12, CZZ14, CW14a, CZ13, CY12, DMS11, DMW13, DZZ14, ELS11, Ery14, Fan10a, FZLZ13, Fan13, Gal11e, GY13, GGFM14, GX13, GSDW14, GJGSNRÁ12, HR11, HLW12, Hu12b, HG13a, HLZ14, HRX14, HS10b, HS12, Ign13, JVK10, JDJ11, JDJ12, JQ14, JXK14, JV13, JK14c, Kar10, KLBR13, KMC12, KLW14b, mL12b, Li10a, LWZD10b, LWS11a, LZLW11, LY14a, Lin11a, LJX14, LLWW10, LJW10, LZ11c, LHL11, LWW12, LZ12c, LM13b, LCL14, LC14e, LB14c, LLP14b, LB14d, LB14b, LB14a, LD11, LD14, MDD12, Ma12, qMWjHpX13, MZW14, MK14a, Mal13b, MAB14, Mao11b, zM10b, Mil14a]. stochastic [Mil14b, MGGS12, MGSMRÁ14, NHCLP10, Nak13, NCÁHCLP13, PS14a, PJ12, QLWW12, QHL10, RCS14, RHSC14, RSRA14, RWkPk10, Sab13, Sab14, San11c, SKSK12, SL14a, SZ14b, SY14a, SHWT14a, SHWT14b, Siu10, TNB11, TLZZ12, TJ12, TJ13, TJ14, VVK12, WZ13a, WzH10, WW12, WCZ13, Wan14a, Wan14e, Wil10a, Wil10b, Wil13, WZW14, WHZ14, WD14c, XL11, XPLZ14, XWC11, XM11, Xu13c, YT11, YMZS10, YSZ12, YLZ13, Yue14, ZLM12, ZSZ12, ZR13, ZW13c, ZLSW13, ZW13b, ZHXM14, ZJLO14, Zha14a, ZLL14, ZxHW14, ZSL10, ZY12, Zha13c, ZPZ14, ZYZ14b, ZL14f, ZJ14b, ZY13c, ZZQ13, ZXF14, ZZY14b, bZpZJ10, ZLY11c, ZC11, ZL12d, ZWH14, Zuf12, dlHV12, MH14c]. stock [CW11a, DW13a, San11a, Sar12a, WW13h]. stocking [WH14]. stocks [CC12a]. Stockwell [Hut13]. Stokes [AL10a, CF13, CFZ14a, Che14h, CRP13, CP14a, CRW11, DS13c, DHQ14, DAA13, FL13, FHWZ12, FHK13, Gal11a, HMA10, JMY13, LR14a, Liu14c, LPPG14, LMB13, Luo14, MV12, NR13, QMZ14, RW14a, SYK10, SH12a, SY13b, SZ14c, TD14b, Wan10d, WH12, WZ14a, WWW12, Yan13d, ZZL13, ZLZ14c, dFGAN12]. Stokes-type [MV12]. Stolarsky [BP10b]. Stoneley [VS10]. Stoner [Fuk14b]. stopped [LM13b]. stopping [Mak14]. storage [CTLD12, GW13, KS13a, MM10c, YY12]. storm [AWHW11]. straight [CEXG12]. Strain [HTX^{+}12, CMS11, Dan14, ZLX12]. strain-constructed [ZLX12]. strain-inertia [Dan14]. strains [MG14b]. Strang [WZ11h]. strategies [BM12d, CCJ10, CW11a, DG10a, ESE14, GS13b, GH10a, HH11, LWZ14, MCR11, MMRS13, PKK10, Peñ13, SLL14, WLYC13, YGX14, dlSAQ11]. strategy [BLR13a, CR11b, Che13c, CYW^{+}14, DE11, Gol14, HLeH12, HWQG11, KMW12, LHT14a, MS13e, SSTM14, SS11d, WT11a, Yu14]. stratified [BBFK12, SAuAH14]. Stream [CFN12, SH12a]. streamer [ZZ13c]. strength [AAMB14, BS14a]. strengthen [ZWL14b]. stress [AHDD13, Che14j, CTS14, CS12b, ECG14, OADL14]. stressed [AAMADH11, GMKV12, GMV12, KGM14, Sin10a]. Stressing [STM10]. stretch [BY14]. stretch-twist-fold [BY14]. stretching [ASM11, Afz10, AHMS14, AV14, AM14g, Cor11b, Das13, EEEE05, FZY10, FZZ12, Fit12, HAMA14, Kha14a, KAB11, Pal10, PMV14, Pan12b, Pan15, PK18, PP11, PMD14, Sha12b, VPS12, VSV10, VV11a]. stretching/shrinking [PMV14]. strict [BK11a, CLC^{+}13, CIVS14, CQ14, DV14, DHS10, JHX14, KI11b, PK12, SW14c, ZL13e]. strictly [sCCLY10, CMM12, Jun10b, Jun11a]. strike [GO10]. string [AM14a, GG13a]. string-rearrangement [AM14a]. strings [DZ12]. strip [ABCD11, AA13, ZL10f]. strip-based [ZL10f]. stripe [YWLX11]. stromal [LKY14]. Strong [AT12, AK11, Buo10, sCLCaL11, sCWT^{+}12, sCCYW13, CWWC14, DTP14, DHS10, Jun10b, KM13a, Kim14a, LM13b, NS11, She12a, She14b, SW14c, WHC10, ZLL14, BCGK14, BDK^{+}14, CQ14, CMW10, CRW11, DM12, Gal11a, KKVCS12, MD13a, Rud12, SW13, SJ13, Xu13b, pYX10, YSB11, Zil12]. Stronger [sCYW10]. Strongly [Fer12a, AG14d, AK11, BM11, Bat14, Cai10a, CWW14, DM10b, GS11d, HRAA12, bJLyZlL14, LY13b, dMbJmY11, RT14, VO14, Yan12a, Zha13b, ZS15, Sav10]. strongly-coupled [Zha13b]. Structural [Mey14, MMM13, HTX^{+}12, NMM14, SA12a, SBS11, Wan14e, YZ14d, ZLY11b]. structurally [BBSC13]. Structure [HLW13, LWN13, Par12, YLJ14, Bat14, BA14a, CCK11, CXL12, DZ12, DWP14, EÖ10, Far14a, FGD13, GC10c, JCB11, KSM11, LWZZ14, LZG10, LH14b, LZLY14, MYJ10, MXH12, MHXyY14, Mur14, Pal13, RV10a, SJL10, She12b, SKC14, SB12b, SHM13, WM13a, Xie11a, Xie12a, Xu10b, XTAJ12, YG13, Yua13, ZCŠ10, ZL11a, ZLGZ11, pZJ12, ZF10]. structure-persevering [Yua13]. Structure-preserving [HLW13, LWZZ14, WM13a, Xie11a, Xie12a]. structure-sensitivity [FGD13]. Structured [ZS12, ZJ14a, BBK13, CJK12, CWLC13, CCWM13, EWS10, Gem11, JDN14, JSH10, KJ12, Kha14b, KLY12, LZ11c, PM14b, RWkPk10, SGE13, SL13a, WH14, WJ11, Yan11d, ZW14c, ZMC10]. Structures [Waz12b, AZ14b, ASKA13, AZ19, FMSGS12, KP14b, LPAE12, LH10a, LDR10, MWJW10, MR13a, MMM12, Pan12a, SXCD11, SSS13b, WDL12b, WD14b, WHG13, ZGW14a]. Struve [BDGP14b]. Stuart [ASM14, SZ14a]. studied [BGGH12]. studies [CLL10, GGOS12, LdS10, SJPS14, ZJ13]. Study [AWHW11, CTLL11, DMS11, HYJ14, MAD14b, Pan10, PRS12, PMP11, ABLZ13, AAP10, ÁD10, CCS14, CP14a, dCDNCFRH10, DAA14, FL11, GVT^{+}13, Gra13, GADK14, GC10c, HM12, IKH11, JDN14, JS12, JT13, KS13a, KMAG12, KCC12, KZS12, LLKP14, LLL11b, Mag14b, Mam13, MAB14, Mar12, MV13, MB14, NSL14, NKM^{+}10, NCdCDM13, Odi10, PKK10, RB11a, SG11, SSL10, SSL12, SS14c, SGK14, Sol12, URR11, VZ11, VO12, WaZW13, Waz10g, WDS10, ZB14, ZCL14b]. Studying [NR12, She12c]. Sturm [SQJ11, ZZW13, AGM12, AMS14, And11, jASzZ11, jAS14, BD10, Dum10, EDAM10, EMP12, FHLW11, GWS13, GM10, KKBO12, KM13c, KY14, KT14, Liu11a, OC10, SW12, Tun14, WSD12, YW11a, ZH12, Zha10c, ZLZ12, ZWZ13, Zha14b, ZGZG10]. Sturmian [Yos13]. style [GO10, XLY14]. Sub [WZY14, CYM13, Fre11, bJLyZlL14, LYZ11, LLS14, LL10h, dMbJmY11]. Sub-band [WZY14]. sub-classes [Fre11]. sub-equation [CYM13, LLS14]. sub-equations [LYZ11]. sub-feasible [bJLyZlL14, dMbJmY11]. sub-ODE [LL10h]. subadditive [RT14]. subclass [BÖ11, ÇDO11, KOP11, Kwo11, Sek11, SYW13, SKM11, XXS12]. subclasses [WXA10, XCS13, YL11f]. subcritical [RM12a]. subdiffusion [HZ12c]. subdivision [AR12b, ADT13, BCR10, HWL11, KK13b, LY10a, LLS^{+}13a, LQ13, RLC11, SR10, SSA14b, SIR14, SR14c, TYCZ14, ZHGP14]. Subgradient [CAP14, FC14]. subgrid [ABE^{+}12]. subgroups [Sah11b]. Subharmonic [ZC14b, LZSD14]. subject [AY11c, AA12c, Che10d, Che12b, DF14b, KS10a, KS12a, LLYH10, LY11a, LC12, LW14h, LWW14, NRL12, NL12b, PP14a, YCZ14, ZCZ^{+}13a]. subjected [Cha14c, MWY13, PPMCG14]. sublattices [KP12]. sublinear [TM11]. submanifold [MVS14a]. submatrices [Cra12]. submatrix [PX13]. suboptimal [LTR11]. suborbital [DBG11]. Subordination [CY11b, CBS12, Sok13, XXS14]. subordination-preserving [Sok13]. subordinations [CK13b]. subparametric [NN13]. Subperiodic [BV12a]. subproblems [Tan14]. subquadratic [CT14b]. subsampling [DP13b]. Subset [Örk13]. Subsets [BT10a]. subspace [GP13]. subspaces [BT10a, KP12, Ker13, MZTT12, MVS13, SG10, ZMG13]. substance [lLZ14a]. substances [Sam10]. substitution [AMM11, HLWH12]. substrate [AE11]. subtly [Mor14b]. Successive [CY12, WCZ13, BCC11a, SCIMM11, Yan13d, YT13c]. Successively [Yao10]. such [AZ14b, AZ19, BR10a]. suction [Afz10, AA12d, SZZ^{+}11, VV11a]. Sudan [EMH10]. sufficiency [JSMA11]. Sufficient [Kar13b, CH13, DW12, FNV10, LS14a, LF11a, MPB13, Rag12, XLLL14, ZL10b, ZW13b, ZZJ14]. Sugeno [AMO^{+}10, AFFV11, CS11a, Hon10b, LSY14]. suggested [Öza14]. suitable [yZxGyM10]. sum [BFB12, CHX13, CFSYPP10, EVE12, GI14, JLL12, KA13b, Kuj14, LYZG13, Ras14, WT12, ZB14, ZZX14]. sum-of-processing-times-based [LYZG13]. summability [Bel14, Bor11, BSM14, ÇT10b, ÇE11, Kra14b, Pat13a, Pat13b, SS11b, Sul10, TÇ13]. summable [BM11, BRGP14, DM10b, Sav10]. Summation [Cvi11, Nov14, Cho11c, EK10, MM13b]. summer [NATQ14]. sums [Abe10, Ant13, Bry11, CM11, CL10c, Cvi11, FMF13, Mez13, Zhi12]. Sun [Mar14]. Super [Che11f, LMY12, Yan12c, BGV10, CM14, NCS12, PM12, SA12a, SGA14, WW11e, WX13a, WZ13d, Xia10b, YDXC10]. super-dense [PM12]. super-Halley [NCS12]. super-Hamiltonian [YDXC10]. super-KdV [Xia10b]. super-linear [WW11e]. Superattracting [Liu12b]. Superclose [SY13b]. supercomputing [SA10a, SA10b]. Superconvergence [DX10, HH14, JMY13, YLS13, ZCD10, CS14a, LZW13b, SY13b]. Superconvergent [ASST11, AST12]. Superlinear [Suh13, BEP13, CM12a, HL11b, MXH12, ZT13]. superlinearly [bJLyZlL14]. superordination [CBS12]. superposition [BD12b, CFS11, Fer13, WLZ11]. superquadratic [Bat14]. supersingular [WLZ12a]. superstatistics [MH11a]. supersymmetry [TZJZ12]. supervised [ZZF12]. Supplementary [PP14c, SLBS11]. supplier [AKP13, LOD12]. supplier-retailer [LOD12]. suppliers [SCN14]. supplies [LLY14a]. supply [HLL14, PSC12, PSJGLS14, RK14b, SCN14, SM13, SYB13, TNB11, TC14a, ZZ12a]. supplying [SP14c]. Support [MQ14, CF14, CYW^{+}14, LNdCJV14, LIX14, LPY11, NDZ14, PM11a, SFL^{+}11, ZZW^{+}12, dACSS12, LYGZ13, Niu11]. supported [KAF13]. Supporting [HVD^{+}14]. suppression [KEGH13, Koh13, Kos11a, XWWF10]. supremum [CDS13, DS13a]. sure [Mil14a, ZXF14]. surface [Afz10, AHKS12, AHDD13, BMHM14, BCGK14, Bar13, Che10f, CWW14, DOV^{+}11, Das13, EMAM13, FZY10, GMKV12, GMV12, HLW13, Ish10, KK12a, KHK^{+}11, Kha14a, KAB11, LWZ13a, Pal10, PMV14, Pan15, RRP14, SJR10, Sha12b, SR12, Van11a, VV11a, VCNV13, WW10]. surfaces [AM14a, ASRI12, AV14, BB13b, Cor11b, DA14, GA11a, GSSS13, Hu13b, Hu14, HHA14, JK13, KBB12, LKY10, LW11b, LWZ11, LR10, LP13d, QHZ^{+}13, Ros11a, Sag11, Sch11a, SIR14, Sol12, VRZ11, VCĆV12, VC14a, YYL14, YLJ14]. surfactant [PS12b, YLK14]. surge [AWHW11]. survey [BTVC10a, KG10, Lol12, PNPD14, TGTCCGRG13]. surveys [LTSV14]. Survival [CYG14, ZZY14b, CG14a, CXL12, Fuk12, SSS14]. susceptible [DIAQ13, HLM14]. Susceptible-Infected-Recovered [HLM14]. suspension [Gad11, Kan13, MWC11]. Sustainability [AABV12]. sustainable [DGCFM11]. Suzuki [ĆARA12]. SVIR [DYL14]. SVM [MIM^{+}14, NCdCDM13]. SVM-based [NCdCDM13]. SVR [GWML12, HDZW11]. swaps [LWL14, ZL12d]. Swarm [BD12c, BSH13, CFG13, LFLS14, SK14b, SDN14, WYL11, AC14, BHJ13, BDDK14, Che10e, CYW^{+}14, CHC12, CTY11, FYL13, GWX12, GGFM14, HjH12, HjH14, HWQG11, HSD11, JSWJ13, JLHW10, JLHW13, JLTZ13, Kam14a, KRJ11, KGB12, KZS12, LDW14, MHS12, MMM13, NMM14, PAC13, SKC11, SFP^{+}11, TIT13, TPAB11, TTWL13, WWW13b, Xu13a, YTCY12, YZ14b, YDZH14, ZLZ14d]. Swarm/flock [LFLS14]. Sweep [MS11]. Swift [ASVV10, VDOG13]. swinging [ALP11]. Switch [JA14, AC13a, FW12, WLL14]. switched [AZ14a, ASMA12, CS13c, DLZ14, HX13, JXK14, LXK13, LX14a, LCL^{+}13a, LDL11, Liu10c, LLX10, LXC12, LZ14d, NZ12, WF14, XC10c, XWC11, XSC11, XX14a, XX14b, ZWZH13]. switches [CJ14, Mao12]. Switching [LSL^{+}14, BRP13, BLR13a, DL13, DZZZ12, ECS13, GSDW14, HB13, HYX14, HX13, KP14a, LS13a, LJPJ12, LWL14, LLX10, LS14b, MZW14, MM14b, RWkPk10, SL14a, Siu10, XC10c, XWC11, XSC11, YMZS10, ZLSW13, ZLL14, ZZWY14, ZPZ14, ZCZ^{+}13a, ZWZH13]. Switchings [LRM12]. Sylvester [BPR13, CR11b, CCL12, Chu10, WDL12a, WZDW10, WZL12]. Sylvester-conjugate [WZDW10, WZL12]. symbol [DGM^{+}14, ES12c, SÇK14a]. Symbolic [LLL13, Sal10c, SSZ11, ZZ10c, DTN14, GMST10, La 10, LWL10, ÖA10, QTL^{+}12, WSL^{+}13]. symbols [SC14b]. Symmetric [Ach11, JLW13, MN11, ZG12, AAS12, AB12b, ARL13, AKB13a, AKB13b, BMHM14, BAP^{+}14, CALR11, CT14a, CPZ10, DH13b, EMS10b, EMA13c, Elo11, Elo13, FV14, FG13b, GM10, Gue14b, GV13, GK12c, HYC^{+}10, HA11, ICKV14, Jia13a, JPI^{+}13, Kas12a, KK13a, KZ12, KS11e, LH11a, LGZ14a, LW14e, LLL10a, MVS13, NB11, PN11a, RSK14, Rec13, RWBA13, Sta14c, Ste10f, Ste10h, Ste10i, Ste13g, Su13, SHSC13, VA12, Vir13, Wan10a, WY13a, WYT13, WWY10, wX13b, XZ13b, YLD14, YH13, YZC14, YY10, YL11f, ZLA10, ZW11a, Nie13]. Symmetries [CG12a, Wan14f, Yan10f, AMA10, CL14a, GB13, IHF13, RK14a, SS10a, VMRA14, Woo20, LLL12]. Symmetrization [GG14f, Gue14b]. Symmetry [MK13a, XC10b, XC10a, XLZ10, yZqLwW12, BX12, BGGH12, HYSZ14, IS10b, IBS11, JZW12, KG14, KA10a, KP13, KK14c, MESAN13, MM10c, WWWJ14, YZD10, YH14c, pZLC10, bZjZS11, ZC13b, ZYC14]. Symplectic [DG14b, MDD12, Mon12, BS10b, CSZ12, CZ10, DLSS11, Dos14, FG13b, Hil12, HZ12b, HZ14c, Kal13, jWtW13]. symplecticity [BIT12a]. Synchronization [Abd13, CC11c, FPC14, Fan14e, KJW10, LWP12, LYC12, MGCFMM13, MMMMMG^{+}11, PKP^{+}12a, PKP^{+}12b, SHWT14b, Son10, WLMJ10, ZC13a, ZLL11b, AA12a, ADBT13, Che11a, Che12b, FWZW13, FP13b, GHG14, GL13c, JJPW13, JJP^{+}12, LL10a, LJPJ12, LPJ^{+}12, LYJP13, LW14f, LLYH10, LLJF12, MZXZ10, MGMM14, NWNL10, NSNEV14, Pai14, SHWT14a, SPWZ12, SCLC10, SL13b, THX14, Ver11, WJ12, WS12, WZWW13, WXWY14, yWZhL14, WWS^{+}14, WLY^{+}12, XWMW12, Yan10a, Yan11b, Yan13a, YC14, YCZ14, ZX10, ZXCL10, ZWD12, ZSWS12, ZZ10b, ZWL14a, ZDB10, Zhe12]. synchronize [Zhu10a]. synchronized [WWK^{+}14]. Synchronizing [GM12]. Synchronous [JU11, Ota14]. synergistic [HT14]. synthesis [B– D11, Dan10, FL14, RKA^{+}13, SSM13, SL11a, ZB11]. system [AK12, AA12b, AGR14b, AABV12, ADBT13, Asl10a, AZAK10, BRS13, BM14a, Bai13, BRP13, Ban14, Bat14, BTT13, BK14, Bhr14, BSR10, BZZ14, BDAD10, BRAS12, Bot12, BAJK10, BF12, Cai10a, CHC11, CB12a, ÇM13a, ÇBS14, sCCLY10, CLCL11, Cha13a, Cha14c, CZZ^{+}11, CC12b, Che13c, CWLC13, CCWM13, CW14b, CW14a, Cho11a, CS10d, CFN12, CFCLM13, Con13, Dan10, DL13, DN14a, DGDA10, DWP14, DRGR10, DBS12, dCDNVS11, DRS12b, Din11, Din12, DWC13, DU13, DS14e, DGOR13, Dou10, DHSZ10, DX14b, DG14b, DH14, EDYA14, ESNE^{+}14, EEZ14, EM12, ET12b, EALS10, FHW13, FZ13, FS14, FMS14, FP13d, Fuy14a, GS13b, GHLF14, Gaz12, Ge14, GMST10, AA14a, GHPR12, GC10a, Goo11, Gu11, GZW13, GX13, GGXH12, GM11b, GH10e, HHHS11, HZ10, HM10b, HL12b, HL12c, HL13b, HL14a, HR12, HR13, Hon10a, Hu10, Hu12a]. system [HLZ13, HL11e, HH13b, JU11, Jan13a, Jan13b, JS12, JYF13, JLW13, JC13, JZY10, KR13a, Kan12b, KKS11, KAS12b, KR12, KHP11, KL14a, KSFG14, KW13, KX14, KPK14, KZS12, LS14a, LLLZ11, LzJZ12, LL10a, LCH10, LL10b, LWWZ11, Li12b, LL13b, LXWW14, Li14b, LTX11, LK11, LDL13, lLW14, Lis14, LZ10e, LLWW10, LZ11c, LZKK11, LLL12, LZTZ13, LW14h, LC14e, Liu14a, LLP14b, LL13g, LYB12, LH14c, Ma12, MCL14a, Mad14a, Mal14, MTD10, MTT14, MM13a, MGWG11, Mey14, MJ11b, MS12b, MPB13, Mur14, MM12, NSL14, NWNL10, NW12, NNK14, ÖP12b, Pan11, Pan10, PTY12, PTY13, PEP14, PLF13, PW10, PWZZ13, PK12, PBG14, QLWW12, RB11a, RK12, RP13, RS13, Rec14, RWBA13, RRLA14, RPA14, SR14a, SHL10, Sam10, San11a, San12, SM11, Sha14b, SDL10a, SL13a, SXL12, SHWT14b, SK13b, Ste11c, Ste11d, Ste12g, Ste12s, Ste12j, Ste12t]. system [Ste12k, Ste13a, Ste13j, Ste13k, SAAS13, Ste13m, Ste13n, Ste13h, Ste13i, Ste13e, Ste13f, Ste13g, Ste14b, SAAS14b, Sun10, Sun11a, SZS^{+}12, SV12, TW14a, TCBL11, TC14a, UJH14, Val11, VSV10, Vil12, Wan10e, WYL10a, Wan10b, WGZ10, WF10, Wan10c, WZ11a, Wan11d, WHL11, WW11a, WX12, WKLW12, WDL12b, WDZ13, WZ13b, WWW13c, WZZ13, WSL^{+}13, WLP^{+}14, WFL14a, WHF14, yWZhL14, WWWJ14, WH14, WY10b, WLQ^{+}10, Wen10a, WX13d, WZ12, Wu12b, Wu13, WZW14, XWWF10, XP12, XLGW12, Xin10, Xu10c, XXZH10, XH10, XH11, XLW11, XL14, Xu14a, Yad13, YZL10, YW11a, YWLX11, YL11e, YLC11, Yan12f, YTC13, Yan13a, Yan14b, YT13b, YL10c, YX11, YX12a, Yua12, YJCC14, YGX14, YLK14, ZLC^{+}13, ZTW10, ZZC10, ZLJ11, ZW11b, ZHX11, bZlD11, ZW11a, ZCZS12, ZR13, ZW13d, ZZ13a, ZT13, ZW13b, ZC13a, Zha13b, ZL13b, ZZ14b, Zha14c, ZLJZ14, ZLL14, ZW10b, ZLL11b]. system [Zha12b, ZW14b, ZZJ14, ZTX10, ZDB10, ZK11, Zho11c, ZHC11, ZnZ14, ZYG14, Zhu10b, Zhu10a, ZL10f, ZGY^{+}10, ZM10c, Zuo13, dLDP13, OSRPC12]. systematic [Wah13, WLZ11]. systems [Abd13, AKR14, AA12a, AZ14a, Ala14, AKR10b, AMV12, AFSMRL14, AS14b, AHVR14, AAH14, AP14, APJ14, AS13b, ÁB14b, AGH12, AC12, BDL13, BDD12b, BD12b, BM14b, BMR10, Bal14, BDE^{+}14, BAE13, Bel11b, Bel11a, Bel12, BBI14, BS11b, BCM12, BCDI14, BOR14a, Bie11, BCEAV12, BS10b, BI14, BN10, BNP11, BA13, BC12e, BWHQ14, CÁGGLP14, ÇT10a, Çak13, CGGMG13, CMX14, CKN13, CLTY13, CHK10, CCJT11, Che10e, Che10d, CT11a, Che11a, yCfW12, CH12, CLD12, CMH12, CM12b, CGRS12, CZZ12, CXD14, Che12b, CZZL13, Che14k, CD10b, CZ10, CG14c, CHMT11, CTV12, CST14, DLL11a, DGS14a, DGS14b, DBK14, DL12b, DdGL10, DI13a, DSX11, DMW13, DV14, DISS14b, DMR10, DLSS11, DZZZ12, DSH12, DLZ14, DXC^{+}14, Dos14, DDWW12, DCL11, Du11, DSZ14, DGH11, DZ14, ESV12, EDYSA12, Ery14, FP13a, FPC14]. systems [Fan14e, Far14a, FV14, FAM11, FL14, FB10, Fuy14b, GLSV10, fGyLJ11, GxFX^{+}12, GY13, GLX13, GWW14, GLM14, GHG14, GJ14, GLV14, GNS14, GC10b, GAK12, GMA13, GZB12, GM12, GSGN11, GZ12a, GM13b, GW12, GMSH13, GM14c, GGOS12, GC11d, HSX^{+}11, HV12a, Has14, HZ14b, HSW11, HL10, HXZ12, HZ12a, HL13a, HC14, HCLP11, Hil12, HZ12b, HZ14c, HSS10, HM11, HRAA12, HCC^{+}12, Hsi10, HZ12c, HHR13, Hu13a, HLZ14, HH12, HX13, IT12, Ilt11, JZ13, JBW14, JPK12, JLW11, Ji14b, JQ14, JK14b, JXK14, JLLD13, JWWY13, JYL14, KP14a, KNS14, Kan11, KHD12, KWJ13, KXW14, KZ14b, KK12b, KV12, KLKA14, KH14a, Koh13, Kol14a, KWM13, Kot14, KTS11, KPP^{+}13, LL14a, LFLS14, Lee11, LJKP11, LKP12, LLP14a, LP14, LR13a, LWZD10b, LLY11a, LZQ11, LZLW11, LL12d, LD13b, LXK13, LAF13, LW14c, LFR14, LZL14a, LW14d, LX14a, Li14a]. systems [LSSZ14, LZZP10, LZWZ14, LLY14b, LCL^{+}13a, LLYH10, LDL11, LLCX10, LLZY10, Liu10c, LLX10, Liu11c, LCC12, LHS12, LXC12, LZ12c, LYL13b, LCL14, LGZ14b, LR14b, LYLS14, LSL^{+}14, LLZ14c, LS13c, LZL12, LdS10, LTR11, LZ14d, LL12g, LY14e, MZXZ10, MLH11, MDD12, Ma14a, MC14a, MA10a, MS13a, MHA12, Mao11a, MGCFMM13, MMMMMG^{+}11, MA10b, MA13d, MF14, MCPO12, MK10b, Mus14, NHCLP10, Nak13, NCÁHCLP13, NP10a, Nar13, Nem10a, NZ12, OS11, Pai14, PQRR10, PRS12, PKPL11, PAC13, PB14a, PS14b, QFCZ13, Ram10a, RY10, RS10, RSK14, RKA^{+}13, RS11b, RZ12, RM12b, RLW14, SJ12, SMN12, SSM13, SS14c, SKSK12, SL14a, Sha10b, SA13, SAP14, She12a, SDK10, Shk14, SAK10, Sin10b, Sin11b, SDN14, SS10e, SL11b, Ste12c, Ste12l, Ste12f, Ste12h, Ste12m, Ste12r, Ste12o, Ste12p, Ste13c, Ste13d, Ste13l, Ste13o]. systems [Ste14c, SAAS14a, Ste14d, SDS14, Ste14e, SS13d, Sun11b, SWLP13, TKLH10, TA13, TPT12, TYT14, LM13a, TK12, TN14, Tsa13a, THX14, VDL12a, Vej13, VVK12, VKV14, VA12, Vil10, Vir13, WCB14, Wan10a, WFW10, WY11b, WZ11b, WS11, Wan11a, WJ12, WY12a, WS12, WL12a, WL13a, WXLW13, WLLY14, WF14, Wan14e, WZSW14, WWS^{+}14, WLL14, Wei14, WWY10, WZH13, WYW14, WT11d, WPSC11, WW11f, WW11h, WW11g, WPSC12, XPLZ14, XC10c, XWC11, XSC11, XX14a, XX14b, XLLL14, XH14, Xu10b, Xu12a, wX13b, XH13b, XC14, YKI10, YMZS10, YWT14, YAW10, YDXC10, Yan11d, YS11b, YSZ12, YKL12b, YKL12a, YCZ14, YHWZ11, Yin10, YZ14c, Yua11, YXO14, YZ14d, YZZ^{+}14, ZMD^{+}14, ZCŠ10, Zei11, ZB11, ZPXX14, ZA14a, ZL14a, yZ14a, ZL10b, ZSL11, mZT11, ZLMG12, ZT12, ZJD12, ZLM12, ZLS12b, ZSWS12, ZY13b, ZLZ14b, ZZGL14a, ZL14d, ZZW14, Zha14d, ZZ10b]. systems [ZZ13b, Zha13e, ZZWY14, ZPZ14, ZC14a, ZWL14a, ZZY^{+}14a, Zhe12, ZCZ^{+}13a, Zho11a, ZDC12, ZZ12b, ZLW10, ZZS14, ZWZH13, ZW14d, Zuf12, ZFCW13, fuy13]. systolic [CCSM13]. Szasz [Suc14, AGKK14, GA10, ÖD11a, YD14]. Szego [Cao10a, CS13b, Kan12a, SPP12]. Szynal [HJMY14].

T [PK18]. Tabu [LH10b, YG13]. Tackling [AGM13]. TAGE [JM11]. tails [GRV13]. Takagi [CT14a]. take [LYPC12]. Takens [XM12]. taking [Mis11]. tamed [ZWH14]. Tanaka [ZHY11]. Tang [AÇT12]. tangency [CDMT12]. tangent [Cvi11, KCC12, WKLW12]. tangle [IN10]. tanh [SSZ11, SAK10, ZA11]. tanh-function [SSZ11, ZA11]. tanks [KS13a]. Tanner [Liu13a]. tapered [AAP10, AE13, MP12a, NAHH11, Ngu14]. tardiness [LH10b, WCW^{+}13, XZZK12]. tardy [ZHC^{+}14]. target [KEH14, MWJW10]. targeting [BMMS13]. targets [AP13, Lin11c]. task [KHD12, XBYF14]. Tasso [SHY11, Zay11]. tau [fRbLJyS13, Tri12, EDAM10, SGA14]. Tauberian [Bel14, ÇE11]. tax [BM14c]. taxation [Ber10]. taxis [MHTH12]. Taylor [AB11, BRAB11, BNBV10, Cha14a, Cha14b, KM13a, LS12c, SI11b, TN14, WW14a, XD11b, Xu14a]. TB [GC12, PC14a]. TC [YL10c]. team [EM12]. technique [AAI14, BRM11b, BCGV11, BGP^{+}13, CTV12, DN14a, DGS14c, DF11, ESEEKH12, EB14, GSN12, HET12, HRAA12, HTX^{+}12, HS12, Jun12, KB13, Li10b, LZH10, Luo14, MH14c, MMGM13, Pic12, SuR13, SLBS11, SSK12, TA11, THNTNXR11, WT11b, ZL11b, ZZWF13]. techniques [AG14b, EL14, GX11, GSDB11, LHZ^{+}11, LC13, MPS14a, MN14a, MO12b, QC12, SLC^{+}12, XZZX14, dlSAQ11]. technology [FGS14]. telegraph [DVGY11, fDxZxChT12, JPM12, LD13c, MB13b, MB14, ZL12c]. telegraphic [LL10f, LL13e]. temperature [Abb14, CEM11, Che10f, FKH10b, KM11c, Lot14, NA10a, SHI13b, VSS13, VO12, VO14, Yig11]. temperature-dependent [Che10f, VO12, VO14]. temperatures [BSZ13]. tempered [AM12b]. templates [Kha14c]. Temporal [KVP14, PM11b, YZ14c, ZS14b]. Temporally [LPR10]. temporary [ZX13]. Tension [ARE14, BMHM14, HHA14, XwK11]. tensor [ADÇ14b, HOMA13, IS10a, IHF13, KE11, LW11b, DFTQOB14, Woo20, ZLZ14c]. tensors [CSC10]. term [And13a, BDG13, BGGH12, CCJ10, CLAT11, CXD14, CL10b, Con13, Cov11, FHLW11, FDHK14, GLZ10a, GSH12, GK12c, HRH14, HL12d, Iid11, JB10, KM11c, Lia14, LLW10b, LdS10, MB13a, MG14c, PS12a, Qui10, SHL10, Sal11c, SAAS14b, SS13d, TW11b, Tri14, Wan14f, XHXH12, YZ11, YLAT14, ZGW14b, BM14b]. terminal [NA10b]. termination [ZH10b]. terms [Bah13a, CM14, Che14d, Che14h, FW14, GWZ14, Ibr14, ISH12, LY14b, LL13d, LL10h, MT12a, Mor14a, Sah11a, Wan10c, WF11, WZ13b, Waz10g, XFS12, YZX13, YXO14, ZCG10, ZSZ12, ZS13b, Zha13a, ZL10g]. Ternary [SR14c, SR10, SSA14b, SIR14]. territory [WJ14]. test [GH11, HVC12, Lin12, LD11, MC12b, PTW11, RTW13, SSS13a, TL14, Van11b, YAAS10, ZWP13]. testing [Mit13, SEKR14]. Tests [KCK08, HS13, PJ10, TPTW14]. tethered [Wil11]. tetrahedral [KK13c]. text [JM13]. textural [OSAG13]. Texture [BB13a]. TFETI [CKSV14]. TGMCHS [KML14]. th [CL11c, DR10a, DB12b, FNG11, Fra10, MTER14, PC13b, Shi14b, Ste10a, Ste10b, Ste10c, Ste10n, Ste12w, Ste13f, Wan10f, WL11, XZ12, Yan11e]. th-order [FNG11, PC13b, Wan10f, XZ12]. Thailand [AWHW11]. Their [Dub13, AAC11, AAGR13, AS11c, BD14a, Bai14, Bal13, Bel12, BS11a, Bez13, CKN13, Cha14a, CDMS14, CK12a, CLND12, DSGB12, Den10, EG11, FZ12, GMSH13, HW14a, HMPVQ14, HYT14b, Irm11, KPT11, KI10b, KT14, La 10, LLLS13, LLZ14b, MSJ14, OO10, OS14, PMB12, PTS11, SS14e, SL14c, Sri11b, SC14b, Ste13d, TPA10, TFM13, Tur14b, WZG12, WDL12b, Wan13d, XT14, YG11, YWD13, ZTD10, ZGW14a, ZM10a, ZM14b, Zha12b, ZLL13, ZK14, dlS10]. theirs [Bai13]. them [ÇM13b]. theorem [AAA12, ABLZ14, AT12, AAS14a, Asl11b, BEP13, BCK14, BSM14, Buo10, Dha13, EVR12, EGH12, Fin14, FT10, GBD^{+}12, GS12d, GS11d, HL11b, HT11b, Hen11, KM14c, KP11b, Łen11, Len14b, LW12a, LHZ10, Lu12b, MA14b, Mez13, MKEG12, Pat13b, SS11b, She14b, Shi12, Sok13, WSD12, Yan12a, YD14, ZS15, dlS10]. theorems [AKR10a, AK11, BD12a, Bas12a, Bel14, BB12, Bor12, ÇE11, sCYW10, sCLCY11, sCLCaL11, CWL^{+}12, sCWT^{+}12, sCCYW13, sCWW^{+}14, CWWC14, CIVS14, CKS10, ĆARA12, DHZ11, – Dor13b, Dou10, EVR13, GAV14, Gu10, GG11b, HZ12b, Hou14, ICKV14, JS14b, KOP11, KI12a, LLK13, NKR12, NS11, Ols12, PE10, PK12, Rad14, SK11a, She12a, She14a, Wal10, pYX10, YK12, YZ12b, ZOS10, ZZGL14b]. theoretic [MTSN14, Mat10, ZLSW13]. Theoretical [CFCLM13, WY10a, WDS10, KRJ11]. theoretically [LMT11]. Theories [CCG12b, Lot14, OADL14]. theory [Abb14, AP11, AH11b, Aou11a, AKH13, AP10, BD14a, Bel13, BS10b, BX12, Cha13b, CZ10, Dan14, DZS^{+}14, FOZX14, GFMP13, GXC^{+}11, GB10, GAC12, Hil12, Isa10, Li13, LD11, MM11b, MB10, O’R12b, Pal12, PCT^{+}10, PPMCG14, RRSS14, SdAT14, SMBI14, SEL14, Sho14, WFwC12, Xie11a, Xie12a, Yos13, Zei11, ZAA^{+}14, Zha13c, ZM14c, NSV13]. therapeutic [HB13]. therapy [MGGS12, MGSMRÁ14, ZK14]. there [CL12]. Thermal [BD14b, FHR13, KML14, ZA14b, Aja11, BB11a, BSØ11, dCDNRH12, FD12a, HWPY14, JYCY10, KV12, MWAM14, MM10c, NBG10, PC13a, PMV14, SHAO14, TH14a]. thermionic [Mam11]. Thermo [Bas12b, HMMK11, OADL14, Sin11a]. Thermo-elastic [Bas12b, HMMK11, Sin11a]. thermoacoustic [SSP11]. thermodynamic [GP11a]. thermodynamically [Zei11]. thermoelastic [Abb14, DKK^{+}13, Lot14, Mey14, Sin10a, WS10]. thermoelasticity [BAJK10, Cop11, Cop14, Sar13, SEL14, ZAA^{+}14]. thermophoretic [HCCT14]. thermoviscoelastic [Mus14]. these [Sza13]. theta [ZT11]. thick [DB11a]. thickness [FZZ12, GMKV12, LNdCJV14]. Thiele [CG14e]. Thin [ASRI12, ASR^{+}14, ARH^{+}12, Bas12b, DKU13, LWCS14, PP13b, PS12b]. Thin-film [ASRI12]. thinning [BSP11, GSC14]. third [Abh14, AOAA10, AC10a, ABT^{+}12, ABLZ14, AAB^{+}13, Ard13, AA12d, Bab12, BDR12a, DGSB12, DAEY12, Du11, FKS13, FHAL14, FFB13, FLP11, Fre11, GB12, KS12d, LKY10, rL11, Lin13c, MIHS14, MK10b, Par13, PS12a, Ram10b, RGL10, RGL11, SMR11a, SSTZ14, SMIS14, SS11c, Sta14c, Ste12k, WL10b, WZ11i, XLGW12, ZYP13, ZCS13]. third- [DAEY12]. third-grade [FHAL14]. third-order [AC10a, ABT^{+}12, ABLZ14, Ard13, Du11, FFB13, rL11, Lin13c, MIHS14, MK10b, Ram10b, RGL10, RGL11, SMR11a, Ste12k, WZ11i, ZYP13]. Thomas [ABF14, Fer11, Oul11, ZLLL12]. Thompson [LYD10]. threaded [ABD^{+}12, FTAB14]. Three [AAP10, BD11, BC13a, DPP12, Har14, HYT14a, Iid11, LJW10, OPD12, PSP10, PSC12, SSN^{+}12, SCK13, WQX^{+}10, Yu13b, Abh14, AZ10, And13a, Aou12, AWHW11, AC12, BZ11, BPT12, BNBV10, CHX13, CRG11, CZM11, CTLL11, DLFZ10, DWQG12, DRSG14, DHMU14, DPP11, Džu11, FFB13, GK10b, GZW13, GX13, GL14b, GG14g, HHHS11, JZY12, JM10b, KÖ11, Kan11, KMW12, Kle12, LSJ11, LTX11, LNC12, LL13d, LZ11c, LWZ13b, LH14b, Lot14, MLY12, ML11a, MGC12, MHB14, MCPO12, MZWZ12, MPB13, OGYE13, PTY12, PTY13, PRS12, PC11, PLF13, PPD10, PB14b, Pre11a, Rec13, RNEA14, SCN14, SHAO14, SY10a, SFF11, WDL10, WZ11e, WDZ12, WLXX14, WP14, WL10b, Wu10b, WW11h, XYC11, XKH10, Yan12b, YY14a, YWW^{+}14, YSI11, ZD10, ZZW11a, Zha13b, ZDW10, ZLD11, ZZJ14, ZTX10, ZH14, ZHW10]. three-component [GZW13, GX13]. three-dimension [GL14b]. Three-dimensional [HYT14a, OPD12, WQX^{+}10, Yu13b, AC12, BPT12, CRG11, CTLL11, DWQG12, HHHS11, KMW12, LL13d, LWZ13b, MHB14, MZWZ12, PB14b, SHAO14, XYC11]. three-element [ZTX10]. Three-layer [PSC12]. Three-level [SSN^{+}12, WP14, XKH10]. three-parameter [JM10b, Rec13, YY14a]. Three-point [DPP12, AZ10, BZ11, DPP11, Džu11, FFB13, LSJ11, LNC12, PPD10, Pre11a, WDZ12, WL10b, ZH14]. three-species [MLY12, SY10a]. three-stage [BNBV10]. three-staged [LH14b]. three-state [WW11h]. three-step [Abh14, GK10b]. Three-term [Iid11, And13a]. three-unit [ZHW10]. three-variable [KÖ11]. three-wave [DLFZ10, WDL10, ZDW10]. threshold [Alf12, DMGK12, ELS11, LR11, MB10, MB11, SO10, WPKE14, ZJ14b]. thresholding [FOZX14, Hor10]. thymic [BPP12b]. tight [SHSC13]. tightly [JHN11]. tigrinum} [WDBM14]. Tikhonov [BBF12a, CPS14, CM12c, LPB10, LW13d, MQ14]. tile [MIM^{+}14]. tilings [GGM^{+}14a, Mar12]. Time [CLTY13, GLB12, HS11a, LXTC13, RV10b, TS12a, AZ14a, Ala14, AS14b, AK10b, AZ10, AM13b, AP14, ASMA12, Asa10b, AMB10, AGH12, AA12c, BMR10, BRP13, BP11a, BEP13, BK14, Bhr13, BCWM10, BFM13, BNP11, EL13, Bra13, CHC11, CNTG12, CMT11, CC12a, CFH^{+}13, Che10d, CDLZ11, Che11c, CS13d, CW14a, CRX12, CZZL13, CZD^{+}14, Che14k, Chu11, CYZ14b, CSSN11, CFN12, CG10, CTLD12, Cui14, DS14a, DVGY11, DGS14c, DSGB12, DIAQ13, DJ14a, DMW13, DM13b, DRGR10, DV14, DMR10, DZZZ12, DWC13, DKU13, DLP10, DDWW12, DZZ^{+}13, DZZ14, DZ14, DLS13, EM11, EWS10, EMP12, Fan10a, FWZW13, FP13a, FP13b, FF11, FZ12, FL14, FR11, Gal11d, GxFX^{+}12, GWZ14, GWW14, GL10a, Ge14, GHG14, GM13a, GC10a, GNS14, GCG14, GM14b, Gün10, GC10c, GL13c, GB13, HB11a, HL11a, Has11, HL10, HL13a, HYJ14]. time [HMMK11, Hil12, HZ14c, HM11, Hsi10, HLC13, HYSZ14, HYYG12, Hua11a, HX13, HHDM13, IKL11, JPK12, JJPW13, JS12, JQ14, JXK14, JCW12, JFHZ14, JKB10, JK10, JV13, JL10, KPC14, KJ12, KfM11b, KMC12, KGC11, KV12, KM11c, KJW10, KSY12, KPLC11, KLPC12, KPP^{+}13, KPP^{+}14, LL14a, LPL^{+}13, Lee11, LJKP11, LWP12, LLP14a, LP14, LR13a, LL10b, LWZD10b, LLY11a, LZQ11, LTY11, Li12c, LX12, LLC12, LYC12, LYPC12, LXK13, LY14b, LX14a, LSSZ14, LH14a, LS10c, LKS14, LCL^{+}13a, LK11, LDL11, LDL13, LJX14, LHT14b, LZG10, LLZY10, LN10, Liu10c, LH11c, LHS12, LXC12, LFLH14, LSL^{+}14, LLKP14, LLZ14c, LZL12, LWW14, LDZ10, LMT11, MLH11, MLY12, Maa14, MC14a, MA10a, MS13a, MHA12, MT10, MTD10, Mao12, MP11a, MA13d, MS13f, MG12b, Mil14b, MMP^{+}12, MSS14b, Mos14, NRWF12, NHCLP10, Nak13, NCÁHCLP13, NP10a, NTS11a, NSZ^{+}14, OP14b]. time [PKD14, Pai14, PT13, PKPL11, PKP^{+}12a, PKP^{+}12b, PWZZ13, PD13, QaC13, QHL10, RSBK14, RSPK13, Rec14, RRLA14, RRSPTR12, Rud12, RH14, Sáf14, SP14b, SAMA11, San10, SGC12, SMS14, SGCG14, SKSK12, SN12, SAL12, SYK10, Sha14c, Sha11b, SLZ12, SLC^{+}12, SZZH13, SZ14b, Sin10b, Sin11b, SSS13a, SS14e, SVV13, SSTM14, Son10, SYB13, SL14d, SHZ14, SDB14, SEKR14, STM10, Sun11b, SC13, SH13, TNB11, TKL14, TPT12, TZ11, TZJZ12, TLZZ12, TXX14, TJ12, TJ13, TJ14, TW10a, Tsa13a, Tun14, Vít14, WCB14, WZ13a, WHW10, WY11b, WS11, WW11a, WH12, WY12a, WL13a, WW13g, WXLW13, WZWW13, Wan14a, WF14, WWS^{+}14, Waz10g, WLYC13, WQ13, WZH13, WT11d, WPSC11, WPSC12, WLZ12b, WS13, WL13d, XPLZ14, XC10c, XX14a, XP12, XX14b, XCT10, Xu14b, Xu14a, YMZS10, YZW^{+}10, YlTA10, Yan10e, Yan10b, YGL11, YW11b, YL11e, YS11b]. time [Yan12e, Yan13d, YLS13, YJ14a, YLAT14, Yig11, YuRI14, YLTB12, ZH12, ZB11, ZPXX14, yZ14a, ZW10a, ZX10, ZLL11a, ZHX11, ZaFT11, ZLMG12, ZCZS12, ZX12, ZLM12, ZLS12b, ZL12a, ZS13b, ZZ13a, ZW13e, ZL13b, ZYZ14a, ZLJZ14, ZJLO14, ZLZ^{+}14a, ZL12c, ZZ13b, ZW14b, ZZWY14, ZW11c, ZCZ^{+}13a, ZW10d, ZMC10, ZLY11c, ZnXwY14, ZZS14, ZD11b, ZS14f, ZWZH13, Zuf12, ZFCW13, dCCL14, dlHV12, dlS10, dlRS14, CFN12, UK11]. time- [HMMK11]. time-accurate [SYK10]. time-cost [KMC12]. Time-delay [CLTY13, Ala14, DZZZ12, FWZW13, FP13a, FL14, GHG14, JXK14, KfM11b, MLH11, Sun11b, WWS^{+}14, WZH13, XPLZ14, XX14a, Xu14a, YMZS10, ZB11, ZW14b, ZWZH13]. time-delayed [HL10, WW13g, ZX10, ZLMG12, ZZ13a]. time-delays [GxFX^{+}12, LSL^{+}14, WZ13a]. time-dependent [AMB10, BP11a, Bhr13, BCWM10, CFH^{+}13, CFN12, DMR10, FR11, GL10a, HHDM13, IKL11, JCW12, JKB10, KMC12, LWW14, Mil14b, OP14b, RSBK14, SGC12, TKL14, TZJZ12, TW10a, WH12, Waz10g, YlTA10, YGL11, Yan13d, YLS13, YJ14a, ZS14f]. Time-discretization [GLB12]. time-domain [KPC14]. time-fractional [DVGY11, LH11c, LFLH14, Mos14, ZW13e, ZW11c]. time-harmonic [EL13, CS13d, CYZ14b, Maa14, dlRS14]. time-inconsecutive [Hua11a]. time-independent [Asa10b]. time-invariant [LR13a]. time-limited [GC10c]. time-quadratic [SGC12]. time-series [WS13]. time-space [YLAT14, YuRI14, ZL12c]. Time-stepping [RV10b, CRX12, MC14a, OP14b]. time-variable [HB11a]. time-variant [YS11b]. time-varying [AS14b, BMR10, BRP13, BNP11, Che10d, CW14a, CZZL13, CZD^{+}14, Che14k, DIAQ13, DMW13, DRGR10, DKU13, DZZ^{+}13, DZZ14, FP13b, GWZ14, Ge14, GL13c, HYJ14, HM11, Hsi10, HYYG12, JPK12, JQ14, JV13, KJW10, KPLC11, KLPC12, KPP^{+}13, KPP^{+}14, LPL^{+}13, LWP12, LLP14a, LP14, LX12, LLC12, LXK13, LX14a, LSSZ14, LCL^{+}13a, Liu10c, LHS12, LXC12, LLKP14, MS13a, MHA12, MTD10, PKPL11, PKP^{+}12a, PKP^{+}12b, RSPK13, SAMA11, Sha14c, Sha11b, SZZH13, Son10, SC13, TPT12, TZ11, TLZZ12, TXX14, TJ12, TJ13, WHW10, WL13a, WZWW13, WLYC13, WPSC11, WPSC12, XCT10, Yan12e, ZPXX14, yZ14a, ZLL11a, ZS13b, ZZ13b, ZZWY14, ZLY11c, dlS10]. times [CK14b, EG11, EG06, LYZG13, LWW14, Mam13, RAW13, RRSPTR14, Rud13, SMWR10, WL14a, WCW^{+}13, XY11, XZY14b, ZB14]. timestepping [KMR14]. Timoshenko [Ahn12, EKSO14, EAAN^{+}14, FMS14, GM13b, MM12, SHL10, TJZ14]. tissue [LX11a]. tissues [GSRR13]. Titchmarsh [BS10b, CZ10]. Tobacco [Goy14]. Toda [HCD14, HG13b, LW10b, Wu12d]. Todd [AG14a, WYT13]. Todd-step [AG14a]. Toeplitz [Alt11, BD14c, DGM^{+}14, DSZ14, Elo11, Elo13, Elo14, GS11c, HH12, LZFR10, MMS11, Nem10a, VA12, WCN11, Xu10c]. together [Ste13d]. Tolerable [DGS14b]. tolerance [BGV10, HZFZ14, LX14b, Sze11, Sze12]. tolerant [LU12a, LG10b, Sze10]. tollgates [Dol14]. tomography [CEM11, KR11, LWWZ11, ZL10f]. tongues [ER13]. tool [Mag14b, OSAG13]. tools [NSV13, Sal11b]. top [CPC11, HL13c, MD10]. topography [OP13, dLDP13]. Topological [BPM10, BR10b, SB10, AK10d, AK10c, BOR14a, FHK13, FGD13, GG14c, HI14, IHM14, JCB11, TR13]. topologies [Sha14c, ZPZ14]. topology [LJPJ12, ZX12]. TOPSIS [Iç14, Tsa11]. tori [KX14]. torque [CGGZ11]. Torrey [YLTB12]. torsion [MVS13]. Torsional [GMV12, Vis14, GMKV12]. Torvik [Ray12]. Total [BS14a, sCLCY11, CWL^{+}12, sCCYW13, CLC^{+}13, Dub13, GKS10, GT12, HG14, KHK^{+}11, LL12a, LH10b, SSS13a, WW11d, dWcFwWjZ13, WCW^{+}13, YW11b]. totally [CLCZ12, CWWC14]. Touchard [DGMR12, MS13c]. toxic [CRC13, Sam10]. toxicity [LX14c]. toxin [CPC11, JBW14]. Trace [ZCZ13b, AS11a, BBT10, Wri11]. tracing [NTS11b]. track [ABCD11]. tracking [KCGH14, KM10, KPBC14, SMC12, SWLP13, WLLY14, WY14, XPLZ14]. tractable [GD11]. trade [AK10b, CCO12, CLS12, KMC12, LHT14a, LOD12, OYCCB13, Tsa13b]. trade-credit [ZZ12a]. trade-off [AK10b]. trade-offs [KMC12]. tradeoffs [STM10]. trading [CW11a]. traffic [Aub10, COPR10, Dol14, HDZW11, lQZWC14]. traffics [CTH14]. trailing [GRV13]. trained [PM11b]. Training [MHS12, LP13a, Niu11]. trajectories [eMBHA13, DDWW12, HH13b, Kho14, RP14b]. trajectory [KPBC14, XLW11, YD13]. tranches [CBLC12]. transaction [KKR11, LW13a]. transcendental [Öke11]. transcription [ALNGMGMM14]. transdermal [RLW14]. Transfer [DSM^{+}12, KS12b, ASM11, AHMS14, All11, AM14g, Che10c, DWQG12, EEEE05, FHAL14, FMV14, Fit12, GSF12, GSRR13, HDK^{+}10, HMD14, HZZ14, HS12, Ish10, JHA14, JHN11, KM11a, Kha14a, KC13b, KCBH14, KS10c, MS13e, MWAM14, Pal10, PC13a, PMV14, Pan12b, Pan15, RRP14, SWL12, SSP11, SBS10, Sha12b, TSS^{+}14, VPS12, VV10, VV11a]. Transform [TM13, Abb12, ABJ10, AS11b, AS14c, AM12b, BHAV10, BAV11, CHT14, Cvi12, Els12, EÖ10, GTM13, HGN11, Jum12, KM14c, LH11c, LG14, MO14, Mos14, Occ11, RN10b, SS13a, SB13, SGN13, SD13b, TZYH14, TQ11, UY11, Waz10a, Boy14, ZW13c]. transformable [HV12a]. Transformation [Shi13a, Yun10b, AM13a, Ber12, Bes10a, CY11b, DTN14, ESG11, EB14, GGYW11, Gal10c, GMST10, GH10b, GKV12, GZ12b, GH10d, GAC12, HCLP11, HG14, IKU10, JYCY10, LTG^{+}12, LTJ^{+}11, jLLW12, MM13a, MGWG11, NZW11, PPK11, PCT^{+}10, QTL^{+}12, TLDZ13, TZJZ12, Wan10b, WTL^{+}11, WGW11, Wen12, WY10c, XGG^{+}11, XTAJ12, YZZP14a, YZZP14b, YWCH11, ZZR14]. transformational [Yam14]. transformations [BD12b, CC11a, HC14, HMSW14, HS10b, MG12a, MD13a, MH10b, NN13, SHMCM12, SPV10, SHY11, hTlCqL10, ZlDxX11, ZHT13, ZLL13]. transformed [JS14c, ZM14a]. transforms [Che14g, Hut13, LDS14, LOS13, SE12, SS11e, UY11, VOL14, WXA10, Xia14, XWWC10]. Transient [AM14g, LZ10c, Mal14, AMA10, Bac14, CCJ14b, DZLW13, HDK^{+}10, HMMK11, KDS13, Kos11b, lL10g, MVGS11, SHBN13, WCB14]. transit [JHZC13]. Transition [MWJW10, BRP13, ESAH10, HL13a, KXW14, LW06, MK14a, SLYJ13, TLZZ12, WLMJ10, ZCL13a, ZCZ^{+}13a, ZYG14]. transitions [Lin12]. transitive [Han14, LHL14]. translation [TZYH14]. Transmission [LSW^{+}14a, jASzZ11, BU12, BAO14, HBDC14, JM14b, LCD12, LL14b, Lim12, lLW14, LWW12, LSJZ13, MP11b, NN14, QaC13, Sin11a, YJ14b, YA14, Zha14b]. transport [ABE^{+}12, BPPS14, BSZ13, EJ11, Gad11, Gad14, KS13c, KMH^{+}13, Lac14, LZ12a, LLT^{+}14, Mar11, MU12, NA10a, OPD12, SDC10, SATdV13, SP14d, SSS13b, VSS13, WY11c, WFwC12, YZZ^{+}14, dCCL14]. transport-fragmentation [BPPS14]. transportation [Liu11b, YH12, ZAY^{+}14]. transpose [WLD11]. transversal [cLYrL11]. transverse [Mam13, Nar12, NZ14, TJZ14]. transversely [LWZ13b, Sin10a]. trapezoid [Liu11g]. trapezoidal [Hwa11, Lin13b, MS11, WLZ12a]. Traub [Pet14]. traveler [XMWD14]. Traveling [AMW12, ÁME^{+}12, FM14, LLY14a, MYT10, SW13, SV12, WL10d, YZ14c, AMB^{+}14, Asl11a, CQJ10, CL11a, yDpY14, DL11, lF12, FLW11, GZ10b, GLZ10b, GYBR10, Hay11, Hu12a, HYW14, JB10, JLM^{+}12, JLZB14, LZ11a, LZLB14, LWD10, LZZZ10, Liu10d, LY14e, ML12a, NMDWAS10, RC12, S10, She10, SDL10b, SY10c, TSP10a, TH14a, VDK13, WZZL12, Wu10b, XC10b, XC10a, XHG11, XZH13, Yan13b, YZX13, YHS13, ZC14b]. traveling-wave [Hay11, VDK13, XHG11, XZH13]. Travelling [CS14b, LY14c, LZX10, THZF12, Xia10b, YuRI14, ZX13, Zhe11, BTVC10a, CLL10, Den11d, Den14, FYYT11, FL10, Kud10c, LY14d, LLR10, MCM10, Par10a, QW10b, QW10a, She12c, THH10, WCL14, WM10, Wen10a, XZT10, Yaş10, ZTW10, Zha10a, ZZ10c, ZS10b, ZTF10]. traversing [CS12b]. treatment [AH11a, BFH12, CGW14, EB12, Fer10a, Fra10, GS12c, HB13, KKH14, SGE13, SGC10, Waz11b, ZZL14]. Tree [CP12c, WZZ14]. trees [Dim13b, FGIV12, GG14c, LGMZ12, MMRS13, Xia10c]. Trefftz [CTLL11, FL13]. tri [GQ12]. tri-gamma [GQ12]. trial [GMSE13]. triangle [AJR11, Akt14b, BCC11b, DCK11, Dua10b, MMO13, PS11, TVZ10]. triangles [DM10b, PFS10, SMPA13]. triangular [Alt11, ÁGVZ12, BD14c, CJZ12, GS11a, HM11, Hu13b, HSC14, JCY10, KML14, LY12b, LY12c, Lu14c, MP12b, MH10a, NN13, PP12b, PP13a, PP15, WL13c, XSL10, YHL14, ZLX12, ZY13c, ZCD10, ZH13]. Tribonacci [MS13d]. trichotomy [SS14c]. Tricomi [Lu12c]. tridiagonal [AdFY13, Chu10, Fen11, GG11c, GG13b, IS11, LV14, Neu13, ÖA13, Rim13a, Rim13c, Rim13b, SEM11, SY14b, VA12, Wu10a, YB12]. trigamma [Mor10b]. Trigonometric [SV10b, AMIR14, BAA13, BV12a, CSMT12, Don14, Gem11, HHM10, LZ12b, MMKD14, Neu14, XW13, ZMG13]. Trigonometrically [FG14b, Kal13]. trilinear [Cao12]. trim [GG12c]. trinomial [BDR12a]. triopoly [CG14a]. triple [CS11c, FL11, LLP14a]. Tripled [BB12, Bor12, Rad14]. triples [WQ11b]. triplet [LfXfC13]. triplets [NY12]. tripotent [ÖKU12]. trisection [PP13a, PFS10]. tritrophic [FL11]. trivial [FV14, NWZ13]. Troesch [Cha10, Tem12, TK14b]. Tropical [CCSM13]. true [Caz14]. truly [BAP^{+}14, Cve14, GTM13]. Truncated [Özk14, Fuk14a, LV14, Wal10, XY10]. Truncation [Pre11a, NTT10, Yan11c, ZW13e]. truss [MMM12]. Trust [Tan14, CKP^{+}13, ES12a, GZ14b, KR13b, PZ14b, Qj10, QC12, WW13a, ZZQ10, ZZC13]. trust-region [CKP^{+}13, ES12a, KR13b, PZ14b, WW13a]. TS [DK14a]. TS-iterative [DK14a]. TSK [Che13c]. TSK-type [Che13c]. tube [AHMS14, Bar13, Zak10]. tuberculosis [NM14]. tubular [QS11, SFF11]. tumor [GS12c, HHHS11, KB14, LKY14, Mad14a, ZK14]. tumor-stromal [LKY14]. tumour [HRH14, MGSMRÁ14, RRLA14]. tumour-immune [RRLA14]. tuning [Xu13a]. tunnel [HB14]. turbine [LSF^{+}13]. turbulent [EASK10, TSS^{+}14, VD14a]. Turing [AKS11, WZ13b, ZLY11a]. turning [KM13c, PBP14, RS11a, SRP13]. Tutte [GGM^{+}14a]. twist [BY14]. twisted [Han14, SS10d]. Twitter [AdC14a]. Two [AAA13a, BAV11, BBC14, Bil14, CMG14, CB10, CWCW11, CHF14, – Dor12a, ESG11, FAM11, GZF12, Ji14a, Kwo12, LLY11b, dLL13, LS12c, Lot14, LW10b, PTY12, PTY13, QZ10, SWD11, SMPA13, TTM12, XZY14a, XZH13, Yan10b, YCC^{+}12, ZGW14a, ZCU14, ZLL13, AKN11, AB14a, Abb14, AMA12b, AA12a, AJF13, AFSMRL14, AKP13, AHRV11, AMR13, AAH14, AAS14b, AM12c, AE13, Ata11, AJPZ14, BKA12, BHAV10, BC12a, BP10b, BAA13, BCK14, BH12a, BJSS14, BY11b, BCC11a, BFB12, CT13a, Çak13, CZM11, CKH11, CMD13, CLTA13, Che13a, Che14g, CZ13, CGO14, CS13f, ĆARA12, Cop14, CJRV12, DEP12, DGS14a, Den10, DC13a, Den14, DWP14, DP13b, DWC13, DLLW13, DPP11, EDYSA12, EGOM10, EMA13c, ET12b, Esf11b, FHAL14, Fer12a, FP13c, FP12, FG14b, GG10a, GLM14, GLZ10b, GC11a, GZ14a, Goć14, GZ11, GM11b, Gus12, HS14a]. two [Has10, HYT13, HC14, HSI13, Her11, HYT14a, HY11b, HLW12, HYW14, HA11, HD14b, Iri10a, Jat12, Jha13, JPM12, KCGH14, Kar10, KMR14, KM14b, KI12a, KS13b, KM13c, Kim10, KW13, KBB12, KCLC13, LCH13, LF11b, LWWZ11, LY14a, LL10d, LHT14a, LMY12, yLbLjH10, LTW10, LWTC10, LXY10, LY10b, LZ10c, Liu11b, LY12b, LYL13b, LTU13, LYZG13, LLL11b, MG14a, MCL14a, MS13b, Mal14, MÇ12a, MMO13, Men11, MWY13, MG14b, MM13c, MJ11b, MB14, MT12b, MS12b, MK08, MK10a, NTT10, NAA13b, Öca14, OYCCB13, PM14a, Par12, PC13b, PDBRFER14, PID10, RGL11, RP14a, RDA14, Rud13, Sal11a, SS13a, Sam10, SAuAH14, SR14b, SN11, SJR10, SH12a, SHBN13, She12b, SY13a, SC11, Sin10b, Sin11a, Sin12, SY12, SY14c, Sol12, SDB14, SMR11b, Ste11d, SCLC10, SLL14, Sza13, Tan14, TW14b, TD14a, TD14b, TKC12, TFM13]. two [TK13, TDT12, TCI12, TK12, Tua13, Üna13, VDL12a, VG14b, WLJ11, WDLM10a, WY11c, WZ11e, WW11a, WX12, WS12, WZ14a, Wan14d, Waz10f, WHWC14, WY10b, WFZ13, WW11g, WLY^{+}12, XL11, XMWD14, XTL10, XLW11, wX13c, XL14, YK14a, YTS12a, YH14a, YW11b, Yu12a, YL11f, YHS13, YWCH14, YGX14, Zah11, Zei11, ZFH14, ZGH10, Zha10b, ZSZ12, ZWZ13, ZL14c, ZLZ14c, ZD11a, Zhe11, ZYLS12, ZWT12, ZGY^{+}10, ZZZL14, Zil12, dFGAN12, KS11b]. Two-agent [CWCW11, YCC^{+}12]. two-body [Cop14]. two-by-two [Wan14d, wX13c, YH14a]. two-center [MÇ12a]. two-component [XZH13]. two-coupled [ZGH10]. two-degree-of-freedom [XLW11]. two-dependent-unit [ET12b]. Two-dimensional [BAV11, ESG11, FAM11, PTY12, PTY13, QZ10, AHRV11, BHAV10, CLTA13, Che13a, Che14g, Den14, Fer12a, GM11b, KCLC13, LF11b, LY12b, LYL13b, LS12c, MJ11b, NTT10, NAA13b, Par12, SN11, Sol12, TK12, WLJ11, WZ14a, YTS12a, ZFH14, ZL14c, ZD11a, ZWT12]. two-dimensions [HSI13]. Two-direction [Kwo12, LLL11b]. two-filter-smoothing [KCGH14]. Two-grid [CB10, BY11b, WFZ13, dFGAN12]. two-layer [FHAL14, ZGY^{+}10]. two-level [AKP13, MS12b, SH12a]. two-levels [LHT14a]. two-machine [YW11b]. two-norm [DC13a]. two-parameter [LY14a]. two-patch [DWC13, WY11c, XTL10]. Two-periodic [LW10b]. Two-phase [SWD11, BC12a, DP13b, HYT14a, HY11b, SAuAH14, SHBN13, TKC12, Zei11]. Two-point [LS12c, AAS14b, BCK14, BCC11a, CJRV12, EGOM10, GC11a, Goć14, KS13b, PC13b, PID10, She12b, SY12, XL11, XL14]. two-predators [BKA12]. two-prey [PM14a]. two-processor [Rud13]. Two-sided [SMPA13, ZCU14, CS13f]. two-soliton [WDLM10a]. two-species [DWC13, GZ14a, LWTC10, Sam10, VDL12a, WW11a]. Two-stage [BBC14, AE13, Liu11b, MG14a, Sal11a, SLL14, ZZZL14]. two-step [AMR13, DEP12, FG14b, Jat12, Kar10, KM14b, Kim10, RP14a]. two-temperature [Abb14]. Two-time [Yan10b]. two-unit [WW11g, YGX14]. two-variable [BP10b]. two-warehouse [BJSS14]. Twostep [CD11b]. Twostep-by-twostep [CD11b]. Type [Rho19, AY11a, ACZ12, AR14a, ABK14a, AMO^{+}10, AY10, AOM^{+}11, AZL13, ABLZ14, Agr14a, AGKK14, AKS14, AAKB13, AYS13, Alo13, Alo14, ÁNA12, AMR13, AAB^{+}14, AMW12, Aou12, AKR11, Ard13, AH11e, AKB13a, AKB13b, AM12c, AM12d, AO12, Ata11, AKÖ11, AS11c, BAR11, BZ10b, Bal14, BSS14, BWaY14, Bir13, BCC11b, BHP13, BFXZ10, BD14d, BD10, BB12, Bor12, BSM14, Bum12, CS13a, Cai10b, Çak10, ÇT10a, Çak13, CMGR14, CMG14, CK10, CLCZ12, CZ11b, CSZ12, Che13c, Che14f, CS13e, CD13, CKS10, CG14b, CT13b, ĆARA12, CD11b, CT11c, CTV13, CT13c, CK13a, CG14e, CR14, DLL11a, Dar12, DS14b, DN14b, DB12a, dCDNCFRH10, DSH12, DWC13, Dou10, Dra11, Dra14, DLLW13, DH14, Ere11, Et13, EGH12, FZ12, FMF13, Fer10c, FKH14, FGFRV11, FG13b, FPTS14, GG10a]. type [Gal10a, GLSV10, GMX10, GLZ10a, GLX13, GLZ10b, AA14a, GAV14, GO10, GGS13, GARK12, GSGN11, GKV12, GT11b, GM14b, GLB12, GS14d, GS14f, HL11a, HHU12, Has11, Hon10b, HJZC13, HYSZ14, HXQ14, HN10a, HL11e, HLW13, HZW14, HMPVQ14, HN10b, Içö12, IHF13, IJ11, Iri10a, IW14, Izg11, JS10, JW10b, JKP14b, JDD14, KM11b, Kar13b, Kar14, Kha14b, KI12a, KKH14, Khe13, KA10b, Koh12, KB11b, Kre13, KV14, Kud10c, KGM14, KS11f, KM12c, Lai10, LCD12, LJR14, Li10a, LS10a, LZZL10, LHZ^{+}11, LY14b, LSYS14, LSY14, LZaY14, LH14a, LGS^{+}11, Lin13b, LWH14, LZ10e, LLWW10, LN10, LJW10, Liu11g, LY11b, LS12b, LHCI12, LJ13, LC14d, LT14, LLR10, LLZC11, LS14e, LY14e, Ma11b, MI10, Mah14a, MM10a, Mal12a, Mal12b, MC14b, MÇ12a, Mao11b, MP12a, MCdSS11, MTT14, MV12, MS13f, MWAM14, Mis11, MX11, MAHAS14, MKEG12, MKK14]. type [MAK13, Mus14, Neu14, NW10, Nie11, O’R10, OM11, OO14, OP12a, ÖD11a, OS14, ÖA10, PKBD12, PEP14, PPK11, PC11, Pet14, PS14b, PKS10, QC12, RRR11, RY13a, RY14b, RY14c, RA10c, RA12b, Ren14a, Rep11, Rho14b, Rho14a, RRRRTR12, SPA12, SP14a, SS14c, Sau11, Sch11a, SS14d, SMIS14, Set14, SDL10a, She14b, Sim11, SS14e, SKD12, SGA14, SSTM14, SMWR10, SL14d, SÖCS12, Sta14c, Ste10a, Ste10b, Ste10c, Ste10d, Ste10e, Ste10j, Ste10g, SU10a, Ste10h, Ste10k, Ste10m, Ste10l, Ste10n, Ste11a, SS11e, SS11g, Ste11g, Ste11b, Ste11e, Ste12d, Ste12i, Ste12m, Ste12s, Ste12o, Ste12w, Ste13a, Ste13g, SAAS14a, SAAS14b, Ste14e, SHWX14, SV10b, TKL14, TS14, TSG10, TW11a, TS12b, Tri12, THAB11, TD13, Tri14, VG14a, VR12, VG14b, Vít14, VS13, WDLM10a, WDL10, WZ10, jWtW13]. type [WC13, WZZ13, WD14a, Wan14b, WV14, WYK11, WW13h, Woo20, Wu10a, Wu14b, XSL10, YM10a, YL10b, Yan11e, YL11c, Yan12d, YKL12b, YKL12a, YKL14, YLD14, YWCH11, YG11, YD14, Yüz13, Yüz14, ZAA^{+}14, ZTD10, Zha10a, ZXCL10, ZL11a, ZL12b, ZjHO13, ZTZ14, ZCU14, ZH10b, ZZH11, ZLH11, Zho13, ZHC11, bZpZJ10, Zhu10c, Zhu11b, ZXZ11, ZL13f, ZM10c, Zuo13, EGOM10, SSS13a, Ste10l, TW14b]. type-II [BWaY14, SSS13a]. type-III [Mis11]. type-IV [HL11e]. types [Bel11b, Bel11a, CS13b, KJ10, LJW10, MCPO12, PMB12, WYZ14, XZH13, Xu11]. Typhoon [AWHW11].

Uhlenbeck [GNdC12]. Ulam [AM13b, AM14d, BC12e, PR12b]. ultimate [DI13a]. ultrasonic [KEGH13]. umbral [Ara14, DS11]. un-stirred [GZ11]. unbalanced [MB12]. Unbounded [DCR13, Abb14, AS11a, FO11, Lim12, Lin13a, LGN10, LXLY14, Ols12, RLW14, SIV10, VSM14, YXO14, ZZZ11]. Uncertain [Din13b, Abd13, AS14b, BMR10, BNP11, Che11a, CZZ12, Che12b, DMW13, GGXH12, HCLP11, HX13, HZ14d, JJPW13, JJP^{+}12, JQ14, JXK14, KXW14, KM10, LSSZ14, LLZ14b, LZ14d, MS13a, Ni13, Pai14, QK13, RZ12, SZH13, SS13c, WHW10, WS11, XPLZ14, XC10c, XSC11, YMZS10, Yan11b, Yan11d, YC14, YCZ14, YK14b, ZSL11, ZP12, ZLS12b, ZLZ14b, ZZ13b]. uncertainties [AA12a, FA12, GLX13, HLL14, LLYH10, QLCF10, WPSW13, ZFCW13]. Uncertainty [KS13c, Ask14, CMS13, JS14a, KK12b, LXZX13, LIX14, Lin11d, OR14, PAL14, SSTM14, Wan11b]. Unconditional [GL13b, HZ13, Xie12b, Zhu10d]. unconditionally [LK12, LD13c]. unconstrained [And13a, BS14c, dSCAM14, GVW^{+}14, HL14d, KL14b, LF11a, LWF11, LL14f, MYL10a, MLGdSC11, Qj10, RAZES13, UJH14, ZZQ10, ZSQ10, ZZC13]. Unconventional [APA13]. uncoupled [HMMK11]. undamped [YZ14d]. undefinable [Ste13c]. under-balanced [SHBN13]. undercooling [BMHM14]. underdetermined [ESV12, Zha13e]. understanding [BVTC13]. underwater [PCT^{+}10]. undesirable [MTSN14]. unequal [FP13c, WCW^{+}13]. uniaxial [YS10a]. unicyclic [HLH12]. unidirectional [Alf10]. Unification [UK11, Ozd11, OS14]. Unified [Kri14, BCR10, Che12b, EDG14, FH11a, GAK12, Kan12a, LWZD10b, LJT13, NMM14, PS14b, SÖCS12, TS12a, ZSL11]. Uniform [BK12, GZW13, LAF13, MWC11, Oni11, Ros11a, eT12a, WW11a, YWZ10, YLC11, Zha13b, ASM11, BLK11, Boy13, Cha13b, CG10, PMD14, SAA14, VSS13, ZhGQ14, ZX11b]. uniformity [ZCG10]. uniformly [BCM12, sCYW10, CG14c, KKS11, KudKL14, Kwo11, yLbLjH10, RK12, SK11a, SKW11, ZXZ11]. Unifying [dF14a]. unilateral [SHWT14a]. union [Yan11d]. Unique [Ste12u, Ste13o, BKW11, WZSW14]. Uniqueness [Jan14b, ZL12a, AANA10, CP12a, CJM14, Du11, FP12, GLW12, Khe13, Lai10, Mil14a, PD12, PC13b, PV11, QFT^{+}14, TES14, WLH10, Zha13e, uRE13]. Unisolvency [GMGV^{+}11]. unit [AGLR13, CGM13, ET12b, GM13a, KR13a, MR14a, Ste10a, Ste10b, Ste10d, Ste10j, Ste10g, SU10a, Ste10k, SU10b, Ste10l, Ste10n, SS11e, Ste11b, Ste12v, Ste12w, WW11g, Yan11e, YGX14, ZHW10, dlRS14]. unit-time [GM13a]. units [Che12a, LT12, SGCG14]. unity [HL13c]. Univalence [DO11, OP14a, Kir11, SDO10]. Univalency [BDN^{+}14]. univalent [XXS12]. univariate [BCR10, KK13b]. Universal [ES12c, CL12]. univex [TD13, Tri14]. unknown [AA12a, BRP13, GxFX^{+}12, KP14a, KM11c, LYJP13, NSNEV14, OD11b, SLM14, THX14, Yan10a, YS11b, Yan11c, ZWD12, ZW13e, ZZ10b, ZWL14a, Zhe12, ZS14f]. unknowns [PB14b, SW10a]. unreliable [CK14b, KHP11]. unsaturated [CMT11]. unscented [Gro13, HCLP11]. unsolved [SK13b]. unstable [JWW14, Zou12]. unstaggered [TK12]. Unsteady [KS13d, NR13, PMD14, SS10g, AM13c, BD14b, CFZ14a, CP14a, FZY10, Pan15, RW14a, RNEA14, SBS10, Sha12b]. unstressed [CMS11]. unstructured [MGC12, TD14a]. unsymmetric [LZ14a]. unwittingly [Koç13]. updates [ASB10]. updating [KS13c, SEED13, Xie11a, Xie12a, Yua11, Yua13, YZ14d]. upon [Abb14]. Uppaal [CFN12]. Upper [Chi10, GX13, JZ10, QZS13, SW14d, WWQ14, WC11, YSB11, AG14d, ÁGVZ12, GD10a, HW14b, JL14, LSJ11, LLHY11, LXGM14, MDÇ13, PV10, RTPH14, SB12a, Ste10c, SS11h, TH14b, UT13, WHL14]. upside [SSS14]. upside-down [SSS14]. upstream [CTH14]. upward [SHBN13]. Upwind [SYK10, LW13a, WG11, ZY13b]. upwind-like [ZY13b]. upwinding [LFC^{+}12, SFF11]. urban [AOTV12, DGCFM11, JHZC13, MSGN^{+}14, NCdCDM13, NA14b, TES14]. Urysohn [Dar12]. Usage [VDFV12, YG13]. USAOR [ZZGL14a]. use [Apa10, DD12, DFTQOB14]. used [OSRPC12]. users [GJ14]. users-items [GJ14]. Using [CTH14, EGMPMRBC12, FLW11, GA11a, Liu11b, NTdCJL12, Sáf14, SMC12, SKIH12, SBS11, WDT^{+}13, Waz13, ZZ14a, Zou12, AEESMA10, ASS13, ABJ10, AHS14, AAI14, AH11a, AYS13, AA12b, AG14c, AZ14b, ARH^{+}12, ASRI12, AWS10, ABCD11, Ard11, AASS13, AWHW11, Asl11b, AMA10, AZ19, Bab12, BPDZ12, BB13a, BHJ13, BFH12, BYM14, BA10, BAE13, BHG10, CDS13, CMT11, CBA^{+}14, CR11a, CX11, Che12a, Che13c, CF13, Che14j, CXD14, CFZ14a, CSL14, CYW^{+}14, CZL14, CVP^{+}10, CK14a, CP14b, CFN12, CEP11, dSCAM14, CSSSV13, D’A12, DZLW13, Dan14, DN14a, DGS14b, DNSM11, DSGB12, DGSB12, DB11a, DGCFM11, DM13b, DAEY12, Don12, DJT10, DHQ14, DHDB12, DPP11, EM12, EZ14b, EB14, EL14, ESE14, EÖ10, Eto11, EZ13b, FOZX14, FC10, FcDDqZ10, FHK13, GC14, Gee12, GTPM14, GRKP10, GP13, GG12b, GMA13]. using [GSDB11, GK11, GZL14, GK12b, GS14d, GAC12, GH11, HB14, HXY12, HHU12, HS14a, HET12, Has14, HB13, Hor10, Hua11a, HBDC14, HS10b, HS12, IISA10, JJPW13, JKW^{+}11, JW12, JC13, JLTZ13, KHD12, KP14b, KS13a, KA10a, KYH10, Kim11, KMAG12, KAY12, KG13, KB11a, Koy12, KGO10, KSS12, KPBC14, La 10, LR13a, LYZ11, LGN10, LNG12, LWS11b, Liu11a, LZX10, Mah13, MC13, MMM12, MHA12, MS13b, Mam11, MÇ12a, MP10, MP12a, MVRVSCMV14, MSGN^{+}14, MIM^{+}14, MA13c, MAD14b, MT13, MMSZ14, MHS12, MG10b, MESAN13, MMM13, MT12b, MYL^{+}10b, MV11, MKP13, Nak13, NMM14, NAL12, NP10b, NA14b, NDZ14, NK12, Ola11, OPD12, ÖA10, PKD14, PG13, PKBD12, PSSS14, Pan10, Pan15, PB13b, PJFL14, PPMCG14, PAC13, QS11, RMN14, RVS10, SHMCM12, SFL^{+}11, SHN10, SHA12a, SEED13, SSZ11]. using [SSCA13, SMBI14, SLBS11, SBS10, Sha12b, She10, SY13a, Shi13a, SSA14b, SIR14, SR12, Sin10b, SS12a, SM14c, SMH14, SK10b, SGN13, SK10c, TSJ13, Tsa13a, UMMT14, VMS10, VLMHK^{+}13, VJBDV13, VG14b, WLJ11, WLLY14, WYL11, WW13h, WY10c, WYL10b, Wu12a, XRBZ13, XWWF10, XFW^{+}12, Xu14a, YK14a, YWY^{+}13, YB12, YM10c, ZA11, ZJD12, ZLX12, ZL14b, ZB14, Zhe11, Zhi12, ZLYL10, ZYG14, Zhu10b, Zim14, ZANA12, dlS10]. utility [AC13a, PAL14]. utilizing [AHMS14, MB11, Sin11b]. Uzawa [CZ14b, LWYY14, LZ13b, LZ14b, ZS10a, ZW13a, ZM14d]. Uzawa-SOR [ZS10a].

vacation [CK14b, JU11, KHP11, LT11, TLW11, WX12]. vacations [Amm14, JS12, wMwWsT11, SLL14, YX11, Yua12]. Vaccination [dlSAQ11, CCS14, DYL14, GYL^{+}13, LOKB12, LWG13, SS14a, SSL12, ZJ14b]. vaccine [CFL14]. Vakhnenko [LMS10, ML12a]. valent [SI11a]. valently [Irm11]. validating [AANA10]. validation [FLMR14, GS12c, SBM12]. Vallée [EKK14, BSM14]. Valuation [WW13h, ECS13, LW13a, Öke11, ZL12d]. Value [AGM12, AE11, AB11, Abh14, Ach11, AMA12b, AS10, Ali12, AANA10, Ami14, And12b, And13b, jAzBS14, AAS14b, AO12, AMA10, Bai10, BZ11, BD11, BC13a, BU12, BKW11, BCK14, BCC11a, BRAS12, BRGP14, BD10, BAJK10, BRW13, BR13, CCMV12, CS12a, CT13a, CH14, CDS13, Cha14b, CT14a, CSLL10, CZ11a, CT12c, CLL14, CD11a, Cor10, CJRV12, DN14a, Dri13, DR11c, DCR12, EGOM10, FHLW11, FNG11, FFB13, Fra10, GS12b, GX11, Gao14, GWA14, GC11a, Gen11, GQ14, GYD14, Goo11, GK12a, GKW14, GG13c, Haj11, HN14a, Han11, HLZ12, HL12a, HL12b, HL12c, Her11, HLX13, HD14b, Jan11, Jan13d, Jat10, Jha13, JY12, JL14, JW10a, Jia13b, Kar13a, Kel12, KK11c, KS13b, KSS12, LWW10, LSJ11, LLJ11, LXH13, LWZZ14, rL11, LNC12, LLW10b, LLC11, LS12c, MXH12, MV14, MVM10, MT12b]. value [MAHAS14, NAA13a, NA10b, NN19, PT14a, PC13b, Plo14, RK12, Reu14b, RV13b, She12b, SCC12, SK13b, SY12, Sta12a, SNZL13, SD13b, TA11, Tun13, VMS10, Van10, WW10, Wan10f, WLCW10, WCZ11, WCL11, WL11, WLW14, Waz11b, Web10, Web12, WL10b, WZ11i, XL11, XZY14a, XWWC10, XLGW12, XCH14, XM10, XZ12, YKI10, YS10a, YW11a, YZ12a, Yao10, Zah11, Zha11a, ZLZ12, ZG12, ZGZG10, ZSHZ11, ZH14, ZLW10, NAA13b]. valued [AR14c, AB12c, CLC^{+}13, sCWW^{+}14, CS10d, ĆARA12, DGM^{+}14, FH11a, FS14, FC14, GT11b, HMSW14, HZW14, JSMA11, KB13, LSZ12, LLK13, LYXP13, LLL11b, Mae10, Mal12b, Man11, SK10b, WS14, ZLL11a, ZX11b]. values [BU12, BOR14a, DJ12, FPP11a, GÇ11b, HW11, ISG11b, JWWY13, JS14c, KI12b, Koj13, Koj14b, Koj14a, MMSZ14, WRD13, WHL14]. valuing [ZZ11b]. Vandermonde [Res11, Res13]. vanishing [LMS11, dOL13]. VaR [KKR11, OGO14]. variability [Lac14]. Variable [LLP10, OSRPC12, PJFL14, YNBV11, YLGW14, ZX11a, Ali12, AY11b, Aou12, BP10b, BA10, BFM13, BC12d, CCJ14a, COPR10, CLAT11, Che13a, CSL14, CLLS14, CAP14, Cui14, DWZW14, Dem11, Dem14, DAA14, EB14, FZZ12, Fau14, FZP12, For11, GGYW11, GLZ10a, GKV12, GZ10c, GMZL11, Guo14, GYL14, GM11b, HT11a, HB11a, HL12d, Ibr14, INE10, JZG13, KÖ11, KAY12, LWL10, LGS^{+}11, LLWW10, Liu10b, Liu11f, cLYrL11, LLL10b, MU12, Mat14b, MN14a, MNOB10, MTZ10, MGWG11, MWAM14, MK13a, MSS14b, NSZ^{+}14, Oni11, PC13a, PZ14a, PPV14, RA12a, RCH12, Sal10b, SGCG14, SJR10, SLBS11, SLC^{+}12, SGK11, SL14b, SD13a, TNB11, WW11a, WJ14, Wei10b, Yan11a, YH14b, YH14d, YM10b, Yu13b, YY14b, ZA11, ZOA10, Zha13a, ZZW14, ZLZ^{+}14a, ZBM10, ZYC14, ZH14, ZGY^{+}10, ZANA12]. Variable-coefficient [ZX11a, GGYW11, GMZL11, Guo14, HL12d, LWL10, LGS^{+}11, LLWW10, Liu10b, MTZ10, MGWG11, PPV14, ZGY^{+}10]. variable-order [CLAT11, Che13a, YNBV11]. Variable-step [YNBV11]. variable-territory [WJ14]. variables [AB14a, AAC11, AAH14, Ata11, Den10, GG10a, GLM14, GGOS12, Hwa11, Hwa14, JM10b, KKA12, LMB13, MMO13, SS13a, SMR11b, TFM13, Van11b, WWS^{+}14, Zha14c]. variably [KMH^{+}13]. variance [BO13, ECS13, LTSV14, SM14c, YK14a, YLMZ13, ZL12d]. variant [AAI14, AGZ12, CAC14, EGH12, Içö12, KPT11, LGS^{+}11, LZZ10, LLZC11, Wen10b, YS11b]. Variants [ZZZM10, And13b, GGFM14, GSN12, Waz10f]. variate [Mat14a, ZZ14a]. variates [VG14b]. Variation [Oul11, VCNV13, Alo13, Dub13, JZ11, KHK^{+}11, KM12c, LG14, NM08, TMM13, TMM14, THYC10, Tse12, WZ13c, dWcFwWjZ13, WTZ11, WLHZ10, XG13, DWZ12]. Variational [AM14a, AP11, DF11, HSI13, KWM13, BDL13, BZGK10, BFXZ10, BBK13, BA14a, BH11, BA13, Buo10, CAY11, CLCL11, CS10a, CCA11, DHZ11, DZ13, DH14, FH11a, FC14, FV14, GS10b, GML13, GL14b, Han11, Hao10, HZ10, Hou13, lHyLfG10, JMB12, KKS11, KAS12b, KR12, KRC10, KPW11, KPK14, pLjH10, MFZ12, qMWjHpX13, MC13, MT12a, NNK14, OO12, PK12, PS10, Ram11b, Sad11, SK11a, SMNN14, SHZ10, SSK12, Ver10, WT11b, XL11, YZL10, YNL10, YSW11, ZOS10, ZS14a, bZlD11, ZH10b, ZL10c, ZYLS12]. variational-like [KAS12b]. variations [MT10, VC14a]. varieties [HW13a]. variety [BDH^{+}14, WM10]. various [ASMA12, BSP11, EÖ10, Gat14, KRMM14, NSC12a, SNC11, Set14, SMR11b, Yan10d]. Varley [Wan14b]. Varying [MMMMB13, AS14b, BMR10, BRP13, BNP11, Che10d, CW14a, CZZL13, CZD^{+}14, Che14k, DIAQ13, DMW13, DRGR10, DKU13, DZZ^{+}13, DZZ14, DZ14, FP13b, GWZ14, Ge14, GL13c, HYJ14, HWPY14, HM11, Hsi10, HYYG12, JPK12, JQ14, JV13, KJW10, KPLC11, KLPC12, KPP^{+}13, KPP^{+}14, LS13a, LPL^{+}13, LWP12, LLP14a, LP14, LX12, LLC12, LXK13, LX14a, LSSZ14, LCL^{+}13a, Liu10c, LHS12, LXC12, LLKP14, MS13a, MHA12, MTD10, PKPL11, PKP^{+}12a, PKP^{+}12b, RSPK13, SAMA11, San10, Sha14c, Sha11b, SM12, SZZH13, Son10, SC13, TPT12, TZ11, TLZZ12, TXX14, TJ12, TJ13, WHW10, WL13a, WZWW13, WLYC13, WPSC11, WPSC12, XCT10, Yan12e, ZPXX14, yZ14a, ZLL11a, ZS13b, ZZ13b, ZZWY14, ZLY11c, ZZS14, dlS10]. Vasicek [Siu10]. Vector [AZ19, LYGZ13, Niu11, AOTV12, ABG14, AZ14b, Ant12, BT13, CL10a, CZX^{+}14b, CYW^{+}14, DWZ12, DLC10, Hu10, KP12, LNdCJV14, LL14b, LZM^{+}12, LLW13, LWB14, LL11c, LPY11, rL11, Liu12a, LLL11b, Man11, MG14b, MQ14, NNG10, NDZ14, RW14b, SFL^{+}11, SATdV13, pSbL14, SK10b, VDLEK14, WL13d, YZD10, ZZW^{+}12, ZZ10d]. vector-form [DWZ12]. vector-host [CL10a]. vector-valued [LLL11b, Man11, SK10b]. vectors [Ahn13, CGJ12, LZW12b]. vegetation [GTX11, SLYJ13]. vehicle [Cha14c, KZS12, LL13g, SSV13]. vehicular [CTH14, Dol14]. Velocity [NTT13, AWS10, JZG13, LMB13, SMC12, VSS13, ZLZ14c]. velocity-vorticity [LMB13]. vendor [SM13, TNB11, TCBL11, WLYC13]. vendor-buyer [SM13, WLYC13]. Vergnaud [PT13]. verification [FdBCC14, PPL12, SKC11, SBM12]. verified [BJ11]. version [DMS11, MF13, RRSPTR14]. versions [AAA13a]. versus [AM12e, CN14a, MIM^{+}14, McC10, ZS13a]. vertex [BOR14a, SHA12a, Sze10, Sze11, Yan10d]. vertex-degree-based [BOR14a]. vertical [ARH^{+}12, ASRI12, ASR^{+}14, AWS10, FHAL14, HCCT14, KM11a, KRMM14, KS13d, MP11b, PMD14, QaC13, YJ14b, ZxHW14]. vertices [BOR14a, CKH11]. Verticum [GLSV10]. Verticum-type [GLSV10]. Veselov [Pan12a, THZF12, Waz12b]. Vesselov [ZX11a]. vessels [BMMS13, LJ12b]. vex [MMY11]. via [eMBHA13, AR14a, AT11, AM13b, Aou11a, AAH14, AM12b, AKÖ11, BDL13, BB14a, BM13b, BSR10, BEGM12, BRAS12, BSM14, CG12a, CD12, Chu11, Cov11, DW14, DSM^{+}12, EMK13, EMA13b, EZ13a, EJ11, Fen11, GT11a, GMST10, GM12, GMR11, GT11b, GAC12, Han11, HW13a, HM11, Hsi10, HL14c, IW14, JM13, JKP14a, Ji14b, JQ14, JYL14, KEGH13, Kuj14, KPP^{+}13, LPJ^{+}12, LWP12, LLP14a, LWZD10a, LLXZ12, LDR10, Lin11c, Lin11b, Lip13, LL10e, LLWW10, LLX10, LLLS13, LG14, LWZD14, MB12, MGMM14, MGYQ12, MHB14, NZ12, Nov14, Pai14, PTK14, PR12a, Pet12, RWBA13, SLM14, Sha14b, SHWT14b, SY10b, Siu10, Sta11, THP14, Tra12, VVK12, VRMT14, WFW10, WS12, WSL^{+}13, WLLY14, WWS^{+}14, WT11d, Yan10a, Yan11b, YLLL13, YT13a, YJCC14, ZA14b, ZLM12, ZSZ12, ZC13b, Zha13c, ZDH14]. viability [KI11a]. Vibration [DWQG12, EAAN^{+}14, KEGH13, AGM14, AHDD13, Cha13b, CTA13, Che14d, Dan14, DB11a, DW13b, EJ12, EEM12, EMAM13, GAC12, HA12, JE11, Koh13, MR13a, MP10, MP12a, Nar12, NZW11, TJZ14, WW13f, WW13e, WD14b, ZAA^{+}14]. vibration-suppression [Koh13]. Vibrational [SRA10]. Vibrations [GYBR10, Cha13c, CCS13, GG12a, Kia14, SK10c]. vibratory [MSJ14]. vibro [XLW11]. vibro-impact [XLW11]. VIE [CJ11b]. Viète [GCMP14]. Viète-like [GCMP14]. view [GCMP14, JKP14b]. views [TC14b]. VIM [HA10]. viral [FW10]. virose [WC14b]. Virtual [YTY^{+}14, TIT13]. virus [HT10, HYT13, RG14, WPKE14, YLC12, ZYYZ12]. viruses [Ama14b, GYL^{+}13, WC14b, YY12, YYL^{+}13]. visceral [EMH10]. Viscoelastic [WM13b, Ahn12, ASKA13, BSP11, KPC14, LPAE12, Lee12, LJ12b, MSCK11, MZWZ12, MM12, QZS13, TPD10, Vis14, WDS11]. viscoelasticity [eT12a]. Viscosity [Inc12, Che10f, DTP14, KS10c, Lac14, LDWY14, LPPG14, NA10a, VO12, VO14]. Viscous [All11, BD14b, CFH^{+}13, CS12b, Fer10a, JA14, KM11a, Kha14a, Neu12b, NA10b, Pan15, RY10, RAIM14, RTV12, SBS10, Sha12b, SGC10, SD13a, Svá11, TH14a, WZ14a, Zha11a]. viscous-capillary [TH14a]. visitor [YLC10]. visualization [VRZ11]. visualizing [HT14]. VMO [Rag12, Sha11a]. voids [Sin11a]. Voigt [AQ11]. volatility [CZ13, ELS11, GA11a, RA12a, SY14a, ZW13c, ZY13c, ZC11, ZL12d]. volcanoes [Moo12]. Volkenborn [AAC11, KP13]. voltage [Ach13]. Volterra [ABRT14, Tom15, AC12, BRS13, BZ10b, BM14b, Bec13, BB11c, Bie11, BRM11a, CGGMG13, CDR13, CC13, CZ11a, CJ11a, CGRS12, CDCS13, Dan10, Dia14, DS12b, DGOR13, DX14b, Far14a, FMF13, GAV14, GM14b, GADK14, GC11c, GWML12, HL11a, HSW12, JC13, KJ10, KG13, KD13, LTX11, LCL14, LXLY14, MCWC13, MN12, MN14a, MS13f, Men11, MM13b, MH14c, MHB14, MEG10, MAHAS14, Mur14, Ngo12, PR12a, RGL10, SNMD13, Sha10b, Sho14, Tom12, Tur14b, Vil10, Wan11d, WHL11, WWG13, WW13c, WFZL14, WW14a, Waz10a, WRD13, WLQ^{+}10, WN11, XLS13a, XH13a, XW13, Xia14, Yan13c, YZ14c, Yüz14, ZLS12a, ZjHO13, ZYZ14b, ZZJ14, ZL13f, dlHV12]. volume [AP14, BB13a, Bra13, CB10, FK12, KK13d, LLS13b, MYT10, MJ13b, Naj14, WaZZW11, Yan10b, Yan12b, YWL13, Zha11c]. volume-filling [MYT10]. volume-radius [BB13a]. Voronovskaja [BM12a, Fin14, YD14]. Voronovskaya [Kre13, Izg11]. Vortex [ZL13c, ASM14, IKH11, JT13, PHPD11, ZTX10]. Vortex-filaments [ZL13c]. vorticity [AWS10, LMB13]. Vries [eMA13a, Asl11a, Dem14, GGYW11, GMST10, GT11c, Kud10b, LJW10, ÖK11, QW10b, SB13, Sie10, THAB11, VDK13, WDLM10a, ZGY^{+}10]. vs [JHN11]. VSVO [BNBV10].

wage [LWM10]. wait [AA12c]. walk [BSKL13]. walks [XFS12]. wall [ASR^{+}14, HMA10, KS13d, RNEA14, SZZ^{+}11]. walled [Dan14, Kia14, SRA10]. walls [Gad14, JHA14, KRMM14]. Walras [Hen11]. Walsh [BM14b, AM13a]. Wang [ZTD10]. warehouse [BJSS14]. warranty [KR13a]. waste [KMH^{+}13]. Watanabe [ZHY11]. water [eMBHA13, CJM14, DOV^{+}11, Dem11, DC13b, EVDG12, FHW13, FH11b, FK12, GZZ10, GLW12, HSZ14, KRMM14, LGS^{+}11, LLL12, LLL10b, MGWG11, PB13a, RYÖ11, RTV12, TD14a, TK12, WTL^{+}11, Waz11a, WX13d, XWWC10, YH10, YLK14, dLDP13]. water-oil-surfactant [YLK14]. water-wave [WX13d]. waterborne [WC14a]. watermarking [MO14, PTK14, SNZL13]. Wave [AS14a, Sin10a, Xie12b, AAMADH11, AMB^{+}14, AM12c, Asl10b, Asl11a, Awa12, Bai13, Bai14, BTVC10a, BDG13, BK10b, CQJ10, Cai11, CLL10, CLA^{+}12, lChH14, CS14b, DLFZ10, DL12a, yDpY14, Dea13, Dem11, Dem14, DPC10, Den11d, Den14, Dha13, DC14, FYYT11, lF12, FLW11, GZ10b, GLZ10b, GSH12, GB10, GO13, GM13b, GZZ10, GZL14, GMV12, HdCOL14, HA10, Hay11, HML10, HM10a, HRL14, HSZ14, Hu12a, HZ12c, HLW13, HYW14, HBDC14, JPB10b, JB10, JLM^{+}12, JLZB14, KPC14, KM12b, KS10a, KE11, Kud10c, KRS10, Kyc11, Laz11, Li10c, Li10d, LZ11a, LZW12a, LY13b, LZLB14, LWD10, Lin14a, Liu10a, LLL12, LL13f, LY14d, LSW14b, LXZ14, LLR10, LLL10b, LZX10, LDLL10, LY14e, MWJW10, ML12a, MCM10, Meh11, MHLL11, Mis10, MG11b, NR14, NSZ^{+}14, NMDWAS10, ÖK11, PPC13, Par10a, PCT^{+}10, QW10b, QW10a, QTL^{+}12, RYÖ11, RRM12, RCYL10, SBR^{+}12]. wave [SXL13, She10, SLZ12, She12c, SDL10b, SXL12, SW13, SAK10, SJ13, SMH14, SYZ10, SS10f, SY10c, SC10b, SHC11, SL14c, SV12, THH10, THZF12, TS12b, TI10a, THAB11, Vis14, VDK13, VD14b, WDL10, WMJW11, WqXlC11, jWtW13, WCL14, WM10, WTFC13, Wen10a, WX13d, Wu10c, Wu10b, Wu12d, Wu13, XC10b, XC10a, Xia10b, XP12, XZT10, XW10, XHG11, XZH13, hXqXlC10, YH10, Yan13b, Yaş10, YZ11, YA14, YuRI14, YHS13, ZOA10, ZTW10, Zha10a, gZrLX10, ZL12a, ZDW10, ZZ10c, ZHZ12, Zhe11, ZS10b, ZTF10, ZLZY12, ZC14b, ZZHZ14, Nie13]. wave-penetrable [LXZ14]. waveblock [Alf12]. waveform [Fan10a, Fan13]. waveguide [FHR13, LWL12]. wavelength [LWL12]. Wavelet [AJR11, LXLY14, PN11b, PB13a, SGDK11, YH14d, BAV11, CK10, CX11, CVP^{+}10, CBRTA11, DFM10, GL11a, Hua11b, Jia12, LZ10b, LDWY14, LDS14, LXZ12, LLL11b, MR13a, MO14, MMM11, QFT^{+}14, Ray12, RP13, SuR13, SR14a, SD13b, SHSC13, WF12, WMM14, WT11d, ZYSQ11, ZLYL10]. wavelet-based [CVP^{+}10, MR13a]. wavelet-Galerkin [DFM10]. wavelet-quasilinearization [SuR13]. wavelets [CF12, CF14, HHM14, KSY12, LSX10, LZ11b, LZ12b, LLL11b, VMS10, WZY14, ZX14]. wavepackets [NL12a, ZD11b]. Waves [ASM14, eMBHA13, Amd10, ACG13, AMW12, ÁME^{+}12, CL11a, DL11, Esf11b, GMKV12, HM10c, HLT14, KS10c, Kud13a, KGM14, LLY14a, LGS^{+}11, LNG12, LZZZ10, LY14c, LW10b, NC13a, RC12, S10, SBR^{+}12, Sin11a, TH14a, VS10, WTL^{+}11, WD14a, Waz11a, WGW11, WG12, WL10d, YDT^{+}14, Yu13a, YZX13, YY14b, YZ14c, ZX13, Zha10d]. wavetrain [She12c]. wavy [KRMM14].

Wazewska [ZZW11b]. WDS [HS11b]. Weak [AH11e, CMW10, Ilt11, Üna13, Ver13, pYX10, DX14a, GLW12, HH14, ICKV14, JZ10, Lin11a, MCH14, MM13d, Neu12b]. weak-limitation [JZ10]. weakened [BH12a]. weakly [ANR12, AP10, Bie11, BRM11a, CJ11a, CJ11b, EVR12, GG14e, ISG11a, LAR13, Mam13, Ola11, SK13a, ZHY13]. weakly-singular [BRM11a]. wear [MCPO12]. web [PBG14, WW13b, Wei10a, YG13]. Weber [AS11c, JS14a, lJCWpW12, lJLpWxZ14, PT11b]. Weber-type [AS11c]. Website [YG13]. Webster [Tsi11b]. wedge [ECG14, PP11]. weed [BMD13]. Wei [DW12, HL14d, LWM11, ZJ13]. Weibull [AŞK10, GC11d, JM10b, RAW13, YY14a]. Weierstrass [AS11c, BM12a, GW11b, Mez13]. Weierstrass-type [AS11c]. weight [Bab12, CNKS14, DH13a, DC13a, Ger12, KSM14, PS12a, SPP12]. weight-functions [Ger12]. Weighted [DLX13, HHT13, Man11, MB13a, Mop11, MKEG12, ÖD11a, SU10b, Ste10m, Ste10l, Ste10n, SS11h, Ste12v, Ste12w, Zhu10e, AIAD12, BM13a, Bog11, CCLF13, CBO12, CCLW10, Das11, DH13b, EMK13, EG11, ELIGMB11, FHW14, Fer13, Gho14, Guo10, HLT13, HW11, Hwa11, Hwa14, JS10, JMS13, JM10b, Juk13, Koh13, KS11e, LL12a, Lee12, LZ10d, MLG13, MMKD14, MD12a, SS14d, SXL13, SA13, SSCT14, SSU13, SL14d, SB12a, Ste10a, Ste10b, Ste10c, Ste10d, SU10a, Ste10h, SS11f, SSB11a, SS11g, SSB11b, Ste11b, SS12c, Ste12i, TSP10b, TH14b, USAF14, XZW11, XZZK12, Xu10a, Yan11e, YW11b, Yan12d, ZHC^{+}14, Zit11]. weighted-EP [MD12a]. weighted-Newton [SA13]. weighted-type [JS10, SS14d, Ste10a, Ste10b, Ste10c, Ste10d, SU10a, Ste10h, Ste10m, Ste10l, Ste10n, Ste11b, Ste12i, Ste12w, Yan11e]. weighting [Che10b]. weights [CM12d, Kat14, RS12, SS11d, Ste12d]. Weiszfeld [PT11b]. welding [TSJ13]. welfare [PPK14]. Well [BW14, TK12, AP13, dCCL14, AO12]. Well-balanced [TK12]. well-defined [AP13]. Well-posedness [BW14, AO12]. wellposedness [HR12]. wells [FG13a]. WENO [CZC11, Ger12, LP12a, ZB13]. Werner [HKH10]. West [YAAS10]. Weyl [Ans13, BS10b, CZ10, ZL12b]. wheel [SI14]. where [Alo13, ADGP10, SSS13a]. which [BCM12, ÇT10b, Imo14, KL12, Ste13i, SV12]. white [EN11, Gal11e, WZ11g]. Whitham [yDpY14, GZ10d, HML10, JB10, Pin10]. whose [B– D11, DV14, DP10, FPTS14, HHT13, HD14b, LN10, MS13d]. Wick [LZaY14, LLWW10, LJW10, bZpZJ10]. Wick-type [LZaY14, LLWW10, LJW10, bZpZJ10]. Widder [UY11]. width [DS13b, lF12, Ros11b, Udd13]. Wiener [HGN11, HLH12, LGMZ12, NHCLP10, NCÁHCLP13, WN14a, Wu14c]. Wilker [Mor14b, Neu14]. Willmore [DA14]. Willmore-like [DA14]. Wilson [MG10b, MK13b, SZW14]. wilt [LL14b]. wind [AWS10, CWW14, dCDNV^{+}13]. window [GM13a]. Windowed [BHAV10]. Windschitl [Che14a]. windup [RKA^{+}13]. Winner [LYPC12]. Winner-take-all [LYPC12]. winning [KI11a]. wireless [CK12b]. wires [BRVR12]. withdrawal [LLHY11]. within [HMPVQ14, Mah13, OGO14, TFMB12]. within-hosts [TFMB12]. without [BBDS14, CSLL10, CT13b, CKP^{+}13, DPP12, Jat10, JS12, JHX14, LSLL14, LW13c, MVS13, PID10, RCYL10, SM14a, TCBL11, ZF13b]. Witness [HS10a, HW13a]. WLANs [YTY^{+}14]. WNLRT [AP10]. Womersley [BMMS13]. Wong [AKR11]. Wong-type [AKR11]. Woods [RNZH13]. word [KAÇ^{+}11]. words [FMSGS12]. work [BJB11, SFL^{+}11]. work-related [SFL^{+}11]. working [B– D11, JU11, LT11, SFL^{+}11, Ser13, SLL14, TLW11]. world [Xu10a]. worms [MP11b]. worst [JWG^{+}14]. worst-case [JWG^{+}14]. Wright [Kir11]. Wronskian [KZJ13, LGS^{+}11, LCH11, MAA11, TMXG11, TTM12, XGG^{+}11, ZgJlK12]. Wu [PPS13].

XII [WHL14].

Yakubovich [WFDL11]. Yakubovich-conjugate [WFDL11]. Yao [DW12, HL14d, LWM11, ZJ13]. Ye [SX14]. Ye-Yuan [SX14]. yields [AB12b, KBG11]. Young [MCS11, Caz14, Hu13c, KA12]. Yström [FdBCC14]. Yu [HCD14]. Yuan [SX14]. Yukawa [Woo20, IHF13, Wan13b]. Yukawa-type [Woo20, IHF13]. Yule [Gar11]. Yule-Simon [Gar11].

Zabczyk [SS14c]. Zagreb [HG14, RLC11, Ste12a]. Zakharov [Asl10b, Awa12, BTVC10a, Bhr14, BZGK10, HSI13, IB10b, LZLB14, SXL12, SY10c, SC10b, SB11, TB14, WZJ13, gZrLX10]. Zassenhaus [GT11a]. Zermelo [LXTC13]. Zernike [QQXY12, SU12]. Zero [AISS13, GSDB11, SA10b, ZL14a, ZL13b, Ach11, AJF13, Che13d, GLM14, GG14g, fHcXH11, Kos11b, LTL11, QFCZ13, Ste12r, Ste12t, Ste13l, SDS14, Ste14f, ZCT11]. Zero-clusters [SA10b]. Zero-dissipative [AISS13]. Zero-electron-mass [ZL13b]. Zero-finder [GSDB11]. zero-finding [Che13d]. zero-matching [QFCZ13]. zero-one [LTL11]. zerorelaxationtime [ZL13b]. Zeros [HMR12, AGLR13, DG13, DG14a, DM11, ET11, MMMMB13, MPM14, PB13b, PMM11, PC14b, Ras14, RA10b, SAO12, SKC10, ST14, WC11, WE13, WZL14b, ZZ11a]. zeta [Cvi12, FK13, Pré11b, RY13a, Ras14, SÖCS12, WYK11]. Zhang [Ibr14, LWN14a, ZLL11a]. ZK [Esf11a, SL14c]. zone [ZWL14b]. zones [eT12a]. zooplankton [CRC13, JBW14]. Zygmund [LS10a, LS10b, Ren14a]. Zygmund-type [Ren14a].

References

Aghababa:2012:CFR

[AA12a] Mohammad Pourmahmood Aghababa and Mohammad Esmaeel Akbari. A chattering-free robust adaptive sliding mode controller for synchronization of two different chaotic systems with unknown uncertainties and external disturbances. Applied Mathematics and Computation, 218(9):5757–5768, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                                        //www.sciencedirect.
                                                        com/science/article/
                                                        pii/S0096300311014263.

Al-Azzawi:2012:SBP

[AA12b] Saad Fawzi Al-Azzawi. Stability and bifurcation of pan chaotic system by using Routh– Hurwitz and Gardan methods. Applied Mathematics and Computation, 219(3):1144–1152, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                                        //www.sciencedirect.

com/science/article/
pii/S0096300312007229.

Aydilek:2012:HNW

[AA12c] Harun Aydilek and Ali Allahverdi. Heuristics for no-wait flowshops with makespan subject to mean completion time. Applied Mathematics and Computation, 219(1):351–359, September 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312006236.

Aziz:2012:MFT

[AA12d] Asim Aziz and Taha Aziz.

MHD flow of a third grade fluid in a porous half space with plate suction or injection: an analytical approach. Applied Mathematics and Computation, 218(21):10443– 10453, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031200389X.

Ang:2013:DRP

[AA13] W. T. Ang and L. Athanasius. Dynamic response of planar cracks in an infinitely long piezoelectric strip. Applied Mathematics and Computation, 219(14):7711–7724, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001008.

Ghafli:2014:IEB

[AA14a] Ahmed A. Al Ghafli and Hassan J. Al Salman. An improved error bound for a finite element approximation of a reactiondiffusion system of λ − ω type. Applied Mathematics and Computation, 246(??):491–501, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401159X.

Ansari:2014:FCF

[AA14b] Alireza Ansari and Hassan Askari. On fractional calculus of A_{∈\+∞}(§) function. Applied Mathematics and Computation, 232(??):487–497, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000952.

Abdulhadi:2012:LTF

[AAA12] Z. Abdulhadi, Y. Abumuhanna, and R. M. Ali. Landau’s theorem for functions with logharmonic Laplacian. Applied Mathematics and Computation, 218(12):6798–6802, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311015335.

Alizadegan:2013:TMV

[AAA13a] Amir Alizadegan, Babak Asady, and Mohammad Ahmadpour. Two modified versions of artificial bee colony algorithm. Applied Mathematics and Computation, 225(??):601–609, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009843.

Anita:2013:INS

[AAA13b] Laura-Iulia Anita, Sebastian Anita, and Viorel Arnautu. Internal null stabilization for some diffusive models of population dynamics. Applied Mathematics and Computation, 219(20):10231– 10244, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031300386X.

Alagoz:2012:SCM

[AAAH12] Baris Baykant Alagoz, Hafiz Z. Alisoy, Serkan Alagoz, and Fevzi Hansu. A space charge motion simulation with FDTD method and application in negative corona electrostatic field analysis. Applied Mathematics and Computation, 218(17):9007–9017, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002032.

Amat:2013:BOF

[AAB^{+}13] S. Amat, I. K. Argyros, S. Busquier, R. Castro, S. Hilout, and S. Plaza. On a bilinear operator free third order method on Riemannian manifolds. Applied Mathematics and Computation, 219(14):7429–7444, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313000696.

Amat:2014:NTM

[AAB^{+}14] S. Amat, I. K. Argyros, S. Busquier, R. Castro, S. Hilout, and S. Plaza. Newton-type methods on Riemannian manifolds under Kantorovich-type conditions. Applied Mathematics and Computation, 227(??):762–787, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012241.

Angrisani:2012:SPY

[AABV12] Massimo Angrisani, Anna Attias, Sergio Bianchi, and Zoltàn Varga. Sustainability of a pay-as-you-go pension system by dynamic immigration control. Applied Mathematics and Computation, 219(5):2442–2452, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008739.

Acikgoz:2011:NMB

[AAC11] Mehmet Açikgöz, Serkan Araci, and Ismail Naci Cangül. A note on the modified q-Bernstein polynomials for functions of several variables and their reflections on q-Volkenborn integration. Applied Mathematics and Computation, 218(3):707–712, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001263.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Area:2013:LPD

[AAGR13] I. Area, N. Atakishiyev, E. Godoy, and J. Rodal. Linear partial q-difference equations on q-linear lattices and their bivariate qorthogonal polynomial solutions. Applied Mathematics and Computation, 223(??):520–536, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008679.

Arab:2014:ESI

[AAH14] Reza Arab, Reza Allahyari, and Ali Shole Haghighi. Existence of solutions of infinite systems of integral equations in two variables via measure of noncompactness. Applied Mathematics and Computation, 246(??):283–291, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011175.

Aihara:2014:QMR

[AAI14] Kensuke Aihara, Kuniyoshi Abe, and Emiko Ishiwata. A quasiminimal residual variant of IDRstab using the residual smoothing technique. Applied Mathematics and Computation, 236(??):67– 77, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003877.

Ahmed:2013:IMR

[AAKB13] M. A. Ahmed, T. A. Al-Khamis, and L. Benkherouf. Inventory models with ramp type demand rate, partial backlogging and general deterioration rate. Applied Mathematics and Computation, 219(9):4288–4307, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009800.

Aguilera:2013:PPA

[AAM13] A. M. Aguilera and M. C. Aguilera-Morillo. Penalized PCA approaches for B-spline expansions of smooth functional data. Applied Mathematics and Computation, 219(14):7805–7819, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001252.

Abd-Alla:2011:PWN

[AAMADH11] A. M. Abd-Alla, S. R. Mahmoud, S. M. Abo-Dahab, and M. I. Helmy. Propagation of S-wave in a non-homogeneous anisotropic incompressible and initially stressed medium under influence of gravity field. Applied Mathematics and Computation, 217(9):4321–4332, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031001060X.

Abd-Alla:2011:ERN

[AAMAS11] A. M. Abd-Alla, S. R.

Mahmoud, and N. A. Al-Shehri. Effect of the rotation on a nonhomogeneous infinite cylinder of orthotropic material. Applied Mathematics and Computation, 217(22):8914–8922, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004450.

Ahmad:2014:EIS

[AAMB14] Ali Ahmad, Omar Bin Saeed Al-Mushayt, and Martin Baca. On edge irregularity strength of graphs. Applied Mathematics and Computation, 243(??):607–610, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008649.

Amairi:2010:CEM

[AANA10] M. Amairi, M. Aoun, S. Najar, and M. N. Abdelkrim. A constant enclosure method for validating existence and uniqueness of the solution of an initial value problem for a fractional differential equation. Applied Mathematics and Computation, 217(5):2162–2168, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007605.

Agrawal:2010:TDF

[AAP10] Mamta Agrawal, Neeru Adlakha, and K. R. Pardasani. Three dimensional finite element model to study heat flow in dermal regions of elliptical and tapered shape human limbs. Applied Mathematics and Computation, 217(8):4129–4140, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010635.

Acar:2014:PSB

[AAR14] Tuncer Acar, Ali Aral, and Ioan Rasa. Power series of Beta operators. Applied Mathematics and Computation, 247(??):815–823, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314012740.

Alolyan:2012:NFS

[AAS12] I. Alolyan, Z. A. Anastassi, and T. E. Simos. A new family of symmetric linear four-step methods for the efficient integration of the Schrödinger equation and related oscillatory problems. Applied Mathematics and Computation, 218(9):5370–5382, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013506.

Araci:2014:SCT

[AAS14a] Serkan Araci, Mehmet Acikgoz, and Erdogan Sen. On the von Staudt– Clausen’s theorem associated with q-Genocchi numbers. Applied Mathematics and Computation, 247(??):780–785, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012788.

Arqub:2014:NAS

[AAS14b] Omar Abu Arqub and Mohammed Al-Smadi. Numerical algorithm for solving two-point, secondorder periodic boundary value problems for mixed integro-differential equations. Applied Mathematics and Computation, 243(??):911–922, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009060.

Arqub:2013:SFI

[AASS13] Omar Abu Arqub, Mohammed Al-Smadi, and Nabil Shawagfeh. Solving Fredholm integrodifferential equations using reproducing kernel Hilbert space method. Applied Mathematics and Computation, 219(17): 8938–8948, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313002592.

Alomari:2013:MSF

[AAT^{+}13] A. K. Alomari, F. Awawdeh, N. Tahat, F. Bani Ahmad, and W. Shatanawi. Multiple solutions for fractional differential equations: Analytic approach. Applied Mathematics and Computation, 219(17):8893–8903, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003196.

Accion:2010:NSF

[AAV10] A. Acción, I. Arregui, and C. Vázquez. Numerical solution of a free boundary problem associated to investments with instantaneous irreversible environmental effects. Applied Mathematics and Computation, 215(9):3461–3472, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009473.

Abbasbandy:2011:ACT

[AB11] S. Abbasbandy and C. Bervillier. Analytic continuation of Taylor series and the boundary value problems of some nonlinear ordinary differential equations. Applied Mathematics and Computation, 218(5):2178–2199, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009684.

Allasia:2012:MHB

[AB12a] Giampietro Allasia and Cesare Bracco. Multivariate Hermite–Birkhoff interpolation by a class of cardinal basis functions. Applied Mathematics and Computation, 218(18):9248–9260, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002330.

Alonso:2012:SSG

[AB12b] Nicomedes Alonso III and Kenneth L. Bowers. The symmetric Sinc–Galerkin method yields ADI model problems. Applied Mathematics and Computation, 219(8):3847–3854, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010028.

Aygar:2012:JSS

[AB12c] Yelda Aygar and Elgiz Bairamov. Jost solution and the spectral properties of the matrix-valued difference operators. Applied Mathematics and Computation, 218(19):9676–9681, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002354.

Abbas:2014:FOI

[AB14a] Saïd Abbas and Mouffak Benchohra. Fractional order integral equations of two independent variables. Applied Mathematics and Computation, 227(??):755–761, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011533.

AviladaS:2014:EBM

[ÁB14b] Cláudio R. Ávila da S., Jr.

and André Teófilo Beck. Efficient bounds for the Monte Carlo–Neumann solution of stochastic systems. Applied Mathematics and Computation, 248(??):672–686, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013617.

Abazari:2012:CNM

[Aba12]  Reza Abazari.  Com-

ment on “A new method for a generalized Hirota– Satsuma coupled KdV equation” [Appl. Math. Comput. 217 (17) (2011) 7117–7125]”. Applied Mathematics and Computation, 218(9):5838–5839, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014123.

See [XD11a].

Abbott:2012:QQF

[Abb12] Alastair A. Abbott. Dequantisation of the quantum Fourier transform. Applied Mathematics and Computation, 219(1):3– 13, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311009027.

Abbas:2014:GMB

[Abb14] Ibrahim A. Abbas. A GN model based upon two-temperature generalized thermoelastic theory in an unbounded medium with a spherical cavity. Applied Mathematics and Computation, 245(??):108–115, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010169.

Azaiez:2010:SDD

[ABBC10] Mejdi Azaïez, Faker Ben Belgacem, Christine Bernardi, and Nejmeddine Chorfi. Spectral discretization of Darcy’s equations with pressure dependent porosity. Applied Mathematics and Computation, 217 (5):1838–1856, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006995.

Abreu-Blaya:2012:NBM

[ABBR12] Ricardo Abreu-Blaya and Juan Bory-Reyes. A note on the Bochner–Martinelli integral. Applied Mathematics and Computation, 218(17):9018–9023, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002093.

Allasia:2011:STD

[ABCD11] G. Allasia, R. Besenghi, R. Cavoretto, and A. De Rossi. Scattered and track data interpolation using an efficient strip searching procedure. Applied Mathematics and Computation, 217(12):5949–5966, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013275.

Aliaga:2012:SDG

[ABD^{+}12] José I. Aliaga, Paolo Bientinesi, Davor Davidović, Edoardo Di Napoli, Francisco D. Igual, and Enrique S. Quintana-Ortí. Solving dense generalized eigenproblems on multi-threaded architectures. Applied Mathematics and Computation, 218(22):11279–11289, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200505X.

Abdullah:2013:SSC

[Abd13] Ali Abdullah. Synchronization and secure communication of uncertain chaotic systems based on full-order and reducedorder output-affine observers. Applied Mathematics and Computation, 219(19):10000–10011, June 1, 2013. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313003950.

Al-Betar:2012:NSS

[ABDKA12] Mohammed Azmi Al-Betar, Iyad Abu Doush, Ahamad Tajudin Khader, and Mohammed A. Awadallah. Novel selection schemes for harmony search. Applied Mathematics and Computation, 218(10):6095–6117, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101441X.

Abel:2010:CAE

[Abe10] Ulrich Abel. A complete asymptotic expansion for a sequence of certain sums. Applied Mathematics and Computation, 217(8):4302–4305, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010568.

Achchab:2012:RPE

[ABE^{+}12] B. Achchab, A. Benjouad, M. El Fatini, A. Souissi, and G. Warnecke. Robust a posteriori error estimates for subgrid stabilization of non-stationary convection dominated diffusive transport. Applied Mathematics and Computation, 218(9):5276–5291, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013427.

Abdou:2011:SMN

[ABEK11] M. A. Abdou, A. A. Badr, and M. M. El-Kojok. On the solution of a mixed nonlinear integral equation. Applied Mathematics and Computation, 217(12):5466–5475, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012166.

Amore:2014:ACS

[ABF14] Paolo Amore, John P. Boyd, and Francisco M. Fernández. Accurate calculation of the solutions to the Thomas– Fermi equations. Applied Mathematics and Computation, 232(??):929–943, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001829.

Abbaspour:2011:EMN

[ABG11] Mohammad Hassan Abbaspour, Houshang Behravesh, and Ghodrat Ghaffarzadeh. Embeddings of minimal non-abelian p-groups. Applied Mathematics and Computation, 218(3):658– 661, October 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311000816.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Aimi:2014:DPV

[ABG14] A. Aimi, G. Buffoni, and M. Groppi. Decomposition of a planar vector field into irrotational and rotational components. Applied Mathematics and Computation, 244(??):63–90, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009242.

Amat:2013:MEF

[ABGGS13] Sergio Amat, Sonia Busquier, Àngela Grau, and Miquel Grau-Sánchez. Maximum efficiency for a family of Newton-like methods with frozen derivatives and some applications. Applied Mathematics and Computation, 219(15):7954–7963, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000878.

Abhulimen:2014:EFT

[Abh14] C. E. Abhulimen. Exponentially fitted third derivative three-step methods for numerical integration of stiff initial value problems. Applied Mathematics and Computation, 243(??):446–453, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007966.

Addam:2010:NMO

[ABJ10] M. Addam, A. Bouhamidi, and K. Jbilou. A numerical method for one-dimensional diffusion problem using Fourier transform and the B-spline Galerkin method. Applied Mathematics and Computation, 215(12):4067–4079, February 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031000024X.

Adiyasuren:2014:HDH

[ABK14a] Vandanjav Adiyasuren, Tserendorj Batbold, and Mario Krnić. Halfdiscrete Hilbert-type inequalities with mean operators, the best constants, and applications. Applied Mathematics and Computation, 231(??):148–159, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000411.

Amodio:2014:ACM

[ABK^{+}14b] P. Amodio, C. J. Budd, O. Koch, G. Settanni, and E. Weinmüller. Asymptotical computations for a model of flow in saturated porous media. Applied Mathematics and Computation, 237(??):155–167, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004251.

Al-Betar:2013:ASM

[ABKG^{+}13] Mohammed Azmi Al-Betar, Ahamad Tajudin Khader, Zong Woo Geem, Iyad Abu Doush, and Mohammed A. Awadallah. An analysis of selection methods in memory consideration for harmony search. Applied Mathematics and Computation, 219(22):10753–10767, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004888.

Agarwal:2013:NAS

[ABLZ13] Ravi P. Agarwal, Martin Bohner, Tongxing Li, and Chenghui Zhang. A new approach in the study of oscillatory behavior of even-order neutral delay differential equations. Applied Mathematics and Computation, 225(??):787–794, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010096.

Agarwal:2014:PTT

[ABLZ14] Ravi P. Agarwal, Martin Bohner, Tongxing Li, and Chenghui Zhang. A Philos-type theorem for third-order nonlinear retarded dynamic equations. Applied Mathematics and Computation, 249(??):527–531, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013514.

Aero:2013:FIS

[ABP13] E. L. Aero, A. N.

Bulygin, and Yu. V. Pavlov. Functionally invariant solutions of nonlinear Klein–Fock–Gordon equation. Applied Mathematics and Computation, 223(??):160–166, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008448.

Alimohammady:2011:CRC

[ABR^{+}11] M. Alimohammady, J. Balooee, S. Radojević, V. Rakočević, and M. Roohi. Conditions of regularity in cone metric spaces. Applied Mathematics and Computation, 217(13):6359–6363, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000245.

Abbas:2014:ESR

[ABRT14] S. Abbas, M. Benchohra, M. Rivero, and J. J. Trujillo. Existence and stability results for nonlinear fractional order Riemann– Liouville Volterra–Stieltjes quadratic integral equations. Applied Mathematics and Computation, 247(??):319–328, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012417.

Boonperm:2014:AFS

[aBS14] Aua aree Boonperm and Krung Sinapiromsaran. Artificial-free simplex algorithm based on the non-acute constraint relaxation. Applied Mathematics and Computation, 234(??):385–401, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002793.

Agarwal:2012:OAB

[ABT^{+}12] Ravi P. Agarwal, Martin Bohner, Shuhong Tang, Tongxing Li, and Chenghui Zhang. Oscillation and asymptotic behavior of third-order nonlinear retarded dynamic equations. Applied Mathematics and Computation, 219(8):3600–3609, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009538.

Afuwape:2010:AES

[AC10a] Anthony Uyi Afuwape and Jairo Eloy Castellanos. Asymptotic and exponential stability of certain third-order nonlinear delayed differential equations: Frequency domain method. Applied Mathematics and Computation, 216(3):940–950, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001335.

Amiraliyev:2010:NMS

[AC10b] Gabil M. Amiraliyev and Erkan Cimen. Numerical method for a singularly perturbed convection– diffusion problem with delay. Applied Mathematics and Computation, 216(8):2351–2359, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003346.

Aziz:2012:LIL

[AC12] Waleed Aziz and Colin Christopher. Local integrability and linearizability of three-dimensional Lotka–Volterra systems. Applied Mathematics and Computation, 219(8):4067– 4081, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010673.

Abbas:2013:OSU

[AC13a] A. E. Abbas and J. Chudziak. One-switch utility functions with annuity payments. Applied Mathematics and Computation, 219(14):7699–7710, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000994.

Akgunes:2013:NBR

[AÇ13b] Nihat Akgünes and A. Sinan Çevik. A new bound of radius with irregularity index. Applied Mathematics and Computation, 219(11):5750–5753, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012489.

Alves:2014:ABP

[AC14] Maria João Alves and João Paulo Costa. An algorithm based on particle swarm optimization for multiobjective bilevel linear problems. Applied Mathematics and Computation, 247(??):547–561, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012314.

Andreu:2014:COE

[ACCT14] Carlos Andreu, Noelia Cambil, Alicia Cordero, and Juan R. Torregrosa. A class of optimal eighth-order derivativefree methods for solving the Danchick–Gauss problem. Applied Mathematics and Computation, 232(??):237–246, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000939.

Amdjadi:2013:EEM

[ACG13] Faridon Amdjadi, Mark Carter, and Jaganathan Gomatam. The eikonal equation and meandering of spiral waves: a geometrical approach. Applied Mathematics and Computation, 225(??):1–6, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009995.

Aluigi:2014:CKL

[ACG14] Federico Aluigi, Massimiliano Corradini, and Andrea Gheno. Chapman– Kolmogorov lattice method for derivatives pricing. Applied Mathematics and Computation, 226(??): 606–614, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011521.

Achache:2010:CAN

[Ach10a] Mohamed Achache. Complexity analysis and numerical implementation of a short-step primaldual algorithm for linear complementarity problems. Applied Mathematics and Computation, 216(7):1889–1895, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002572.

Argyros:2010:MMS

[ACH10b] Ioannis K. Argyros, Yeol Je Cho, and Saïd Hilout. On the midpoint method for solving equations. Applied Mathematics and Computation, 216(8):2321–2332, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003309.

Achar:2011:SMO

[Ach11] Sujatha D. Achar. Symmetric multistep Obrechkoff methods with zero phaselag for periodic initial value problems of second order differential equations. Applied Mathematics and Computation, 218(5):2237–2248, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311009878.

Acharjee:2013:AES

[Ach13]  P. Acharjee.  Appli-

cation of efficient selfadaptive differential evolutionary algorithm for voltage stability analysis under practical security constraints. Applied Mathematics and Computation, 219(23):10882–10897, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300492X.

Argyros:2014:CBM

[ACK14] I. K. Argyros, Y. J. Cho, and S. K. Khattri. On the convergence of Broyden’s method in Hilbert spaces. Applied Mathematics and Computation, 242(??):945–951, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008352.

Alonso:2011:IDL

[ACQOR11] Pedro Alonso, Raquel Cortina, Enrique S. Quintana-Ortí, and José Ranilla. Increasing data locality and introducing Level-3 BLAS in the Neville elimination. Applied Mathematics and Computation, 218(7):3348–3358, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011179.

Aktas:2012:NTH

[AÇT12] Mustafa Fahri Aktas, Devrim Çakmak, and Aydin Tiryaki. A note on Tang and He’s paper. Applied Mathematics and Computation, 218(9):4867–4871, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012951.

Abide:2012:CMM

[ACZ12] Stéphane Abide, Xavier Chesneau, and Belkacem Zeghmati. Compact mixed methods for convection/diffusion type problems. Applied Mathematics and Computation, 218(10):5867–5876, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013579.

Alvarez:2010:EPW

[ÁD10] J. Álvarez and A. Durán. Error propagation when approximating multi-solitons: The KdV equation as a case study. Applied Mathematics and Computation, 217(4):1522–1539, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309005943.

Adam:2011:NRC

[Ada11] Maria Adam. On numerical ranges of the compressions of normal matrices. Applied Mathematics and Computation, 217(9):4699–4709, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011355.

Asheghan:2013:NFC

[ADBT13] Mohammad Mostafa Asheghan, Saleh S. Delshad, Mohammad Taghi Hamidi Beheshti, and Mohammad Saleh Tavazoei. Nonfragile control and synchronization of a new fractional order chaotic system. Applied Mathematics and Computation, 222(??):712–721, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007959.

Abilhoa:2014:KEM

[AdC14a] Willyan D. Abilhoa and Leandro N. de Castro. A keyword extraction method from Twitter messages represented as graphs. Applied Mathematics and Computation, 240(??):308–325, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006304.

Akgunes:2014:SPT

[ADÇ14b] Nihat Akgünes, Kinkar Ch.

Das, and A. Sinan Çevik. Some properties on the tensor product of graphs obtained by monogenic semigroups. Applied Mathematics and Computation, 235(??):352–357, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003609.

Akyuz-Dascioglu:2011:SHO

[ADÇY11] Ayşegül Akyüz-Daşcıoḡlu and Handan Çerdık-Yaslan. The solution of high-order nonlinear ordinary differential equations by Chebyshev Series. Applied Mathematics and Computation, 217(12):5658–5666, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012464.

Alzahrani:2012:SCM

[ADD12] Ebraheem O. Alzahrani, Fordyce A. Davidson, and Niall Dodds. A 3-species competition model for biocontrol. Applied Mathematics and Computation, 218(19):9690–9698, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002445.

Akbulak:2013:EFP

[AdFY13] Mehmet Akbulak, C. M. da Fonseca, and Fatih Yilmaz. The eigenvalues of a family of persymmetric anti-tridiagonal 2-Hankel matrices. Applied Mathematics and Computation, 225(??):352–357, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009867.

Alonso:2010:CEW

[ADGP10] Pedro Alonso, Jorge Delgado, Rafael Gallego, and Juan Manuel Peña. A collection of examples where Neville elimination outperforms Gaussian elimination. Applied Mathematics and Computation, 216(9):2525–2533, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003486.

Ashyralyev:2011:NFH

[ADP11]  Allaberen  Ashyralyev,

Fadime Dal, and Zehra Pınar. A note on the fractional hyperbolic differential and difference equations. Applied Mathematics and Computation, 217(9):4654–4664, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001129X.

Amat:2013:PCP

[ADT13] S. Amat, R. Donat, and J. C. Trillo. Proving convexity preserving properties of interpolatory subdivision schemes through reconstruction operators. Applied Mathematics and Computation, 219(14):7413–7421, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000635.

Abbas:2011:ASF

[AE11] Fazal Abbas and Hermann J. Eberl. Analytical substrate flux approximation for the Monod boundary value problem. Applied Mathematics and Computation, 218(4):1484–1494, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007831.

Asikgil:2013:PTT

[AE13] Baris Asikgil and Aydin Erar. Polynomial tapered two-stage least squares method in nonlinear regression. Applied Mathematics and Computation, 219(18):9743–9754, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003494.

Abd-Elfattah:2010:IEP

[AEESMA10] A. M. Abd-Elfattah, E. A.

El-Sherpieny, S. M. Mohamed, and O. F. Abdou. Improvement in estimating the population mean in simple random sampling using information on auxiliary attribute. Applied Mathematics and Computation, 215(12):4198–4202, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010881.

Agahi:2011:FLS

[AFFV11] Hamzeh Agahi, A. Flores-Franulič, and S. Mansour Vaezpour. Fatou’s lemma for Sugeno integral. Applied Mathematics and Computation, 217(13):6092–6096, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012968.

Algaba:2014:CEH

[AFSMRL14] Antonio Algaba, Fernando Fernández-Sánchez, Manuel Merino, and Alejandro J. Rodríguez-Luis. Comment on “Existence of heteroclinic and homoclinic orbits in two different chaotic dynamical systems” [appl. math. comput. 218 (2012) 11859– 11870]. Applied Mathematics and Computation, 244(??):49–56, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009291.

See [EDYSA12].

Afuwape:2012:CSS

[Afu12] Anthony Uyi Afuwape. Convergence of solutions of some fourth order nonlinear differential equations. Applied Mathematics and Computation, 219(8):3638–3645, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009599.

Afzal:2010:FSE

[Afz10]  Noor Afzal.  Falkner–

Skan equation for flow past a stretching surface with suction or blowing: Analytical solutions. Applied Mathematics and Computation, 217(6):2724–2736, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008593.

Abdel-Gawad:2010:ASN

[AG10] H. I. Abdel-Gawad. Approximate solutions of nonlinear fractional equations. Applied Mathematics and Computation, 215(12):4094–4100, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010303.

Argyros:2013:EAM

[AG13] Ioannis K. Argyros and Santhosh George. Expanding the applicability of a modified Gauss– Newton method for solving nonlinear ill-posed problems. Applied Mathematics and Computation, 219(21):10518–10526, July 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313004414.

Achache:2014:FNT

[AG14a] Mohamed Achache and Loubna Guerra. A full Nesterov–Todd-step feasible primal-dual interior point algorithm for convex quadratic semi-definite optimization. Applied Mathematics and Computation, 231(??):581–590, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013532.

Achchab:2014:ALF

[AG14b] Boujemâa Achchab and Allal Guessab. On approximation of linear functionals over convex functions: Construction techniques and new directions. Applied Mathematics and Computation, 247(??):417–430, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031401248X.

Al-Gonah:2014:GFI

[AG14c] Ahmed Ali Al-Gonah. Generating functions involving Hermite–Laguerre polynomials using operational methods. Applied Mathematics and Computation, 239(??):38–46, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005906.

Alabdali:2014:ASC

[AG14d] Osama Yousif Mohammed Alabdali and Allal Guessab. On the approximation of strongly convex functions by an upper or lower operator. Applied Mathematics and Computation, 247 (??):1129–1138, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012235.

Alzer:2014:IIC

[AG14e] Horst Alzer and Allal Guessab. An integral inequality for cosine polynomials. Applied Mathematics and Computation, 249(??):532–534, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014611.

Argyros:2014:EAN

[AG14f] I. K. Argyros and D. González. Extending the applicability of Newton’s method for k-Fréchet differentiable operators in Banach spaces. Applied Mathematics and Computation, 234(??):167–178, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002859.

Awwad:2012:BSF

[AGH12] Essam Awwad, István Győri, and Ferenc Hartung. BIBO stabilization of feedback control systems with time dependent delays. Applied Mathematics and Computation, 219(8):3664–3676, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009642.

Alonso:2013:ACM

[AGHdD13] M. M. Canedo Alonso, J. L. González-Hernández, and S. Encinar del Dedo. Application the computational method KIN-MODEL(AGDC) to the simultaneous determination of kinetic and analytical parameters. Applied Mathematics and Computation, 219(12):7089–7101, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000684.

Agrawal:2013:ABS

[AGK13] P. N. Agrawal, Vijay Gupta, and A. Sathish Kumar. On q-analogue of Bernstein–Schurer–Stancu operators. Applied Mathematics and Computation, 219(14):7754–7764, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001148.

Agrawal:2014:GBS

[AGKK14] P. N. Agrawal, Vijay Gupta, A. Sathish Kumar, and Arun Kajla. Generalized Baskakov–Szász type operators. Applied Mathematics and Computation, 236(??):311–324, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004469.

Appleby:2010:PPS

[AGKR10] John A. D. Appleby, Małgorzata Guzowska, Cónall Kelly, and Alexandra Rodkina. Preserving positivity in solutions of discretised stochastic differential equations. Applied Mathematics and Computation, 217(2):763– 774, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007009.

Area:2013:BHP

[AGLR13] I. Area, E. Godoy, R. L. Lamblém, and A. Sri Ranga. Basic hypergeometric polynomials with zeros on the unit circle. Applied Mathematics and Computation, 225(??):622–630, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010394.

Aceto:2012:BVM

[AGM12] Lidia Aceto, Paolo Ghelardoni, and Cecilia Magherini. Boundary Value Methods for the reconstruction of Sturm–Liouville potentials. Applied Mathematics and Computation, 219(6):2960–2974, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009204.

Aledo:2013:TRA

[AGM13] Juan A. Aledo, Jose A.

Gámez, and David Molina. Tackling the rank aggregation problem with evolutionary algorithms. Applied Mathematics and Computation, 222(??): 632–644, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031300831X.

Al-Gahtani:2014:RBM

[AGM14] Husain J. Al-Gahtani and Faisal M. Mukhtar. RBFbased meshless method for the free vibration of beams on elastic foundations. Applied Mathematics and Computation, 249(??):198–208, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013319.

Apolloni:2014:EED

[AGNAL14] Javier Apolloni, José García-Nieto, Enrique Alba, and Guillermo Leguizamón. Empirical evaluation of distributed Differential Evolution on standard benchmarks. Applied Mathematics and Computation, 236(??):351–366, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004457.

Agratini:2014:API

[Agr14a] Octavian Agratini. On an approximation process of integral type. Applied Mathematics and Computation, 236(??):195–201, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004147.

Algaba:2014:NAD

[AGR14b] A. Algaba, C. García, and M. Reyes. A new algorithm for determining the monodromy of a planar differential system. Applied Mathematics and Computation, 237(??):419–429, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005001.

Abad:2011:SRA

[AGT11] Manuel F. Abad, María T.

Gassó, and Juan R. Torregrosa. Some results about inverse-positive matrices. Applied Mathematics and Computation, 218(1):130–139, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007442.

Alvarez:2012:QEA

[ÁGVZ12] Rafael Álvarez, Clara Gallardo, Jose Vicent, and Antonio Zamora. A quick exponentiation algorithm for 3 × 3 block upper triangular matrices. Applied Mathematics and Computation, 219 (4):2004–2016, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008351.

Ali:2012:DFV

[AGZ12] M. M. Ali, M. N. Gabere, and Wenxing Zhu. A derivative-free variant called DFSA of Dekkers and Aarts’ continuous simulated annealing algorithm. Applied Mathematics and Computation, 219(2):605–616, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006571.

Ali:2010:GEF

[AH10] Ahmad T. Ali and Ezzat R. Hassan. General Exp _{a}-function method for nonlinear evolution equations. Applied Mathematics and Computation, 217(2):451–459, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007101.

Akel:2011:NTS

[AH11a] Mohamed S. Akel and Hussein S. Hussein. Numerical treatment of solving singular integral equations by using Sinc approximations. Applied Mathematics and Computation, 218(7):3565–3573, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031101143X.

Albrecher:2011:RTE

[AH11b] Hansjörg Albrecher and Sandra Haas. Ruin theory with excess of loss reinsurance and reinstatements. Applied Mathematics and Computation, 217(20):8031–8043, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100333X.

Argyros:2011:CDM

[AH11c] Ioannis K. Argyros and Saïd Hilout. Convergence of directional methods under mild differentiability and applications. Applied Mathematics and Computation, 217(21):8731–8746, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004905.

Argyros:2011:EAS

[AH11d] Ioannis K. Argyros and Saïd Hilout. Extending the applicability of Secant methods and nondiscrete induction. Applied Mathematics and Computation, 218(7):3238–3246, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011040.

Argyros:2011:WCC

[AH11e] Ioannis K. Argyros and Saïd Hilout. Weak convergence conditions for Inexact Newton-type methods. Applied Mathematics and Computation, 218(6):2800–2809, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010496.

Argyros:2012:NCC

[AH12a] Ioannis K. Argyros and Saïd Hilout. New conditions for the convergence of Newton-like methods and applications. Applied Mathematics and Computation, 219(6):3279–3289, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010053.

Argyros:2012:SCD

[AH12b] Ioannis K. Argyros and Saïd Hilout. On the semilocal convergence of damped Newton’s method. Applied Mathematics and Computation, 219(5):2808– 2824, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009101.

Argyros:2013:EAN

[AH13a] Ioannis K. Argyros and Saïd Hilout. Extending the applicability of Newton’s method on Lie groups. Applied Mathematics and Computation, 219(20):10355– 10365, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004062.

Argyros:2013:ICA

[AH13b] Ioannis K. Argyros and Saïd Hilout. On an improved convergence analysis of Newton’s method. Applied Mathematics and Computation, 225(??): 372–386, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031301028X.

Ahandani:2014:DSF

[Aha14] Morteza Alinia Ahandani. A diversified shuffled frog leaping: An application for parameter identification. Applied Mathematics and Computation, 239(??):1– 16, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005682.

Ansari:2013:SOC

[AHDD13] R. Ansari, K. Hosseini, A. Darvizeh, and B. Daneshian. A sixth-order compact finite difference method for non-classical vibration analysis of nanobeams including surface stress effects. Applied Mathematics and Computation, 219(10):4977–4991, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031201212X.

Akkari:2014:MNR

[AHJ14] N. Akkari, A. Hamdouni, and M. Jazar. Mathematical and numerical results on the sensitivity of the POD approximation relative to the Burgers equation. Applied Mathematics and Computation, 247(??):951–961, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012211.

Ansari:2012:NLR

[AHKS12] A. R. Ansari, B. Hossain, B. Koren, and G. I. Shishkin. A numerical layer resolving method for the flow over a curved surface. Applied Mathematics and Computation, 218(21):10656– 10668, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004316.

Argyros:2014:RSL

[AHM14] Ioannis K. Argyros, Saïd Hilout, and Ángel A. Magreñán. Robust semilocal convergence analysis for inexact Newton method. Applied Mathematics and Computation, 227(??):741–754, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012459.

Ahmed:2014:BLF

[AHMS14] Sameh E. Ahmed, Ahmed Kadhim Hussein, H. A. Mohammed, and S. Sivasankaran. Boundary layer flow and heat transfer due to permeable stretching tube in the presence of heat source/sink utilizing nanofluids. Applied Mathematics and Computation, 238(??):149–162, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004779.

Ahn:2012:VTB

[Ahn12] Jeongho Ahn. A viscoelastic Timoshenko beam with Coulomb law of friction. Applied Mathematics and Computation, 218(13):7078–7099, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015566.

Ahn:2013:ADL

[Ahn13]  Hyo-Sung Ahn.  An

algorithm to determine linear independence of a set of interval vectors. Applied Mathematics and Computation, 219(22):10822–10830, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004827.

Allery:2011:PRM

[AHRV11] C. Allery, A. Hamdouni, D. Ryckelynck, and N. Verdon. A priori reduction method for solving the two-dimensional Burgers’ equations. Applied Mathematics and Computation, 217(15):6671–6679, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001007.

Ahuja:2014:OHA

[AHS14] Vivek Ahuja, Roy J. Hartfield, and Andrew Shelton. Optimization of hypersonic aircraft using genetic algorithms. Applied Mathematics and Computation, 242(??):423–434, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314008200.

Amat:2014:IAS

[AHVR14] S. Amat, M. A. Hernández-Verón, and M. J. Rubio. Improving the applicability of the secant method to solve nonlinear systems of equations. Applied Mathematics and Computation, 247(??):741–752, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012910.

Al-Islam:2012:EWP

[AIAD12] Najja S. Al-Islam, Saud M. Alsulami, and Toka Diagana. Existence of weighted pseudo antiperiodic solutions to some non-autonomous differential equations. Applied Mathematics and Computation, 218(11):6536–6548, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015001.

Ahmad:2013:ZDP

[AISS13] S. Z. Ahmad, F. Ismail, N. Senu, and M. Suleiman. Zero-dissipative phasefitted hybrid methods for solving oscillatory second order ordinary differential equations. Applied Mathematics and Computation, 219(19):10096– 10104, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003779.

Ajadi:2011:AAS

[Aja11] Suraju Olusegun Ajadi.

Approximate analytic solution for critical parameters in a thermal explosion problem. Applied Mathematics and Computation, 218(5):2005–2010, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009313.

Al-Jaser:2013:IID

[AJF13] Asma Al-Jaser and Khaled M.

Furati. Inequalities for integro-differential equations involving derivatives of order between zero and two. Applied Mathematics and Computation, 221(??):535–546, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313007285.

Atanackovic:2014:CAN

[AJPZ14] Teodor M. Atanackovic, Marko Janev, Stevan Pilipovic, and Dusan Zorica. Convergence analysis of a numerical scheme for two classes of nonlinear fractional differential equations. Applied Mathematics and Computation, 243(??):611–623, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008832.

Ajmi:2011:WBT

[AJR11] Neyla Ajmi, Abdellatif Jouini, and Pierre Gilles Lemarié Rieusset. Wavelet bases on a triangle. Applied Mathematics and Computation, 217(22):9387–9396, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005844.

Ali:2012:RSA

[AJR12] Rosihan M. Ali, Naveen K. Jain, and V. Ravichandran. Radii of starlikeness associated with the lemniscate of Bernoulli and the left-half plane. Applied Mathematics and Computation, 218(11):6557–6565, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015074.

AbuElSeoud:2010:PMA

[AK10a] E. Y. Abu El Seoud and M. M. Kassem. Potential method applied to Boussinesq equation. Applied Mathematics and Computation, 215(11):3991–3997, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030901039X.

Anagnostopoulos:2010:EES

[AK10b] K. P. Anagnostopoulos and L. Kotsikas. Experimental evaluation of simulated annealing algorithms for the time–cost trade-off problem. Applied Mathematics and Computation, 217(1):260–270, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006156.

Andreadis:2010:TCPb

[AK10c] Ioannis Andreadis and Theodoros E. Karakasidis. On a topological closeness of perturbed Julia sets. Applied Mathematics and Computation, 217(6):2883–2890, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008787.

Andreadis:2010:TCPa

[AK10d] Ioannis Andreadis and Theodoros E. Karakasidis. On a topological closeness of perturbed Mandelbrot sets. Applied Mathematics and Computation, 215(10):3674–3683, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009801.

Aoyama:2011:SCT

[AK11] Koji Aoyama and Yasunori Kimura. Strong convergence theorems for strongly nonexpansive sequences. Applied Mathematics and Computation, 217(19):7537–7545, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001317.

Adem:2012:SCL

[AK12] Abdullahi Rashid Adem and Chaudry Masood Khalique. On the solutions and conservation laws of a coupled KdV system. Applied Mathematics and Computation, 219(3):959–969, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200700X.

Andreadis:2013:NAA

[AK13] Ioannis Andreadis and Theodoros E. Karakasidis. On numerical approximations of the area of the generalized Mandelbrot sets. Applied Mathematics and Computation, 219(23):10974–10982, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004876.

Arslan:2013:CPI

[AKB13a] Saadet Arslan, Fikri Köken, and Durmus Bozkurt. Corrigendum on “Positive integer powers and inverse for one type of even order symmetric pentadiagonal matrices” [Applied Mathematics and Computation 219 (2012) 5241– 5248]. Applied Mathematics and Computation, 219(12):6978, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000647.

See [AKB13b].

Arslan:2013:PIP

[AKB13b] Saadet Arslan, Fikri Köken, and Durmus Bozkurt. Positive integer powers and inverse for one type of even order symmetric pentadiagonal matrices. Applied Mathematics and Computation, 219(10):5241–5248, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012076.

See corrigendum [AKB13a].

Ahn:2012:IKF

[AKC12] Hyo-Sung Ahn, Young-Soo Kim, and YangQuan Chen. An interval Kalman filtering with minimal conservatism. Applied Mathematics and Computation, 218(18):9563–9570, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001907.

Argyros:2013:EAI

[AKH13] Ioannis K. Argyros, Sanjay K. Khattri, and Saïd Hilout. Expanding the applicability of Inexact Newton Methods under Smale’s (αγ)- theory. Applied Mathematics and Computation, 224(??):224–237, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009181.

Asada:2012:AEP

[AKMM12] T. Asada, V. Kalantonis, M. Markakis, and P. Markellos. Analytical expressions of periodic disequilibrium fluctuations generated by Hopf bifurcations in economic dynamics. Applied Mathematics and Computation, 218(13):7066–7077, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015505.

Abbas:2011:CCF

[AKN11] Mujahid Abbas, Abdul Rahim Khan, and Talat Nazir. Coupled common fixed point results in two generalized metric spaces. Applied Mathematics and Computation, 217(13):6328–6336, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000105.

Avci:2011:NIH

[AKÖ11] Merve Avci, Havva Kavurmaci, and M. Emin Özdemir. New inequalities of Hermite– Hadamard type via sconvex functions in the second sense with applications. Applied Mathematics and Computation, 217(12):5171–5176, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011768.

Aliabadi:2013:TLG

[AKP13] Danial Esmaeili Aliabadi, Abolfazl Kaazemi, and Behrooz Pourghannad. A two-level GA to solve an integrated multi-item supplier selection model. Applied Mathematics and Computation, 219(14): 7600–7615, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313000866.

Abbas:2010:CCF

[AKR10a] M. Abbas, M. Ali Khan, and S. Radenović. Common coupled fixed point theorems in cone metric spaces for w-compatible mappings. Applied Mathematics and Computation, 217(1):195–202, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005862.

Aldaz:2010:BOE

[AKR10b] J. M. Aldaz, O. Kounchev, and H. Render. Bernstein operators for extended Chebyshev systems. Applied Mathematics and Computation, 217(2):790–800, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007034.

Arandjelovic:2011:BWT

[AKR11] Ivan Arandjelović, Zoran Kadelburg, and Stojan Radenović. Boyd– Wong-type common fixed point results in cone metric spaces. Applied Mathematics and Computation, 217(17):7167–7171, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100155X.

Adivar:2014:EPS

[AKR14] Murat Adivar, Halis Can Koyuncuoglu, and Youssef N. Raffoul. Existence of periodic solutions in shifts δ_{±} for neutral nonlinear dynamic systems. Applied Mathematics and Computation, 242(??):328–339, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007279.

Aly:2011:TIR

[AKS11] Shaban Aly, Imbunm Kim, and Dongwoo Sheen. Turing instability for a ratio-dependent predator– prey model with diffusion. Applied Mathematics and Computation, 217(17):7265–7281, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001949.

Agrawal:2014:STG

[AKS14] P. N. Agrawal, A. Sathish Kumar, and T. A. K. Sinha. Stancu type generalization of modified schurer operators based on q-integers. Applied Mathematics and Computation, 226(??):765–776, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011351.

Aktas:2014:MLI

[Akt14a] Mustafa Fahri Aktas. On the multivariate Lyapunov inequalities. Applied Mathematics and Computation, 232(??):784–786, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000691.

Aktas:2014:NPD

[Akt14b] Rabia Aktas. A note on parameter derivatives of the Jacobi polynomials on the triangle. Applied Mathematics and Computation, 247(??):368–372, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012120.

Aleksic:2014:SLM

[AKZ14] J. Aleksić, V. Kostić, and M. Zigić. Spectrum localizations for matrix operators on ℓ^{p} spaces. Applied Mathematics and Computation, 249(??):541–553, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314014465.

An:2010:BIM

[AL10a] Rong An and Kaitai Li. The boundary integral method for the linearized rotating Navier– Stokes equations in exterior domain. Applied Mathematics and Computation, 216(9):2671–2678, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003668.

Aouni:2010:GSG

[AL10b] Belaid Aouni and Davide La Torre. A generalized stochastic goal programming model. Applied Mathematics and Computation, 215(12):4347–4357, February 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310000032.

Alatas:2010:CHS

[Ala10]  Bilal Alatas.  Chaotic

harmony search algorithms. Applied Mathematics and Computation, 216(9):2687–2699, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003681.

Alaviani:2014:CCN

[Ala14] S. Sh. Alaviani. Controllability of a class of nonlinear neutral timedelay systems. Applied Mathematics and Computation, 232(??):1235–1241, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000393.

Alford:2010:MUP

[Alf10] John G. Alford. Models of unidirectional propagation in heterogeneous excitable media. Applied Mathematics and Computation, 216(4):1337–1348, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001979.

Alford:2012:RDM

[Alf12] John G. Alford. Reactiondiffusion models of threshold and waveblock in heterogeneous excitable media. Applied Mathematics and Computation, 218(13):7212–7220, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000045.

Alikhanov:2012:BVP

[Ali12] Anatoly A. Alikhanov. Boundary value problems for the diffusion equation of the variable order in differential and difference settings. Applied Mathematics and Computation, 219(8):3938–3946, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010302.

Allah:2011:VPF

[All11] M. H. Obied Allah. Viscous potential flow analysis of interfacial stability with mass transfer through porous media. Applied Mathematics and Computation, 217(20):7920–7931, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002736.

Aguilar-Lopez:2014:NEC

[ALNGMGMM14] Ricardo Aguilar-López, M. Isabel Neria-González, Rafael Martínez-Guerra, and Juan L. Mata-Machuca. Nonlinear estimation in a class of gene transcription process. Applied Mathematics and Computation, 226(??):131–144, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010679.

Alomari:2013:COI

[Alo13] Mohammad Wajeeh Alomari. A companion of Ostrowski’s inequality for the Riemann–Stieltjes integral ∫_{a}^{b} f(t) du(t), where f is of bounded variation and u is of r-H-Hölder type and applications. Applied Mathematics and Computation, 219(9):4792–4799, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011356.

Alomari:2014:NGT

[Alo14] Mohammad W. Alomari. New Grüss type inequalities for double integrals. Applied Mathematics and Computation, 228(??):102–107, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012629.

Asano:2011:OCS

[ALP11] E. Asano, S. Lenhart, and V. Protopopescu. Optimal control of swinging alliances in a parabolic competition model. Applied Mathematics and Computation, 217(16):6838–6855, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000452.

Altun:2011:FST

[Alt11] Muhammed Altun. On the fine spectra of triangular Toeplitz operators. Applied Mathematics and Computation, 217(20):8044–8051, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003390.

Alzalg:2014:DBI

[Alz14] Baha Alzalg. Decompositionbased interior point methods for stochastic quadratic second-order cone programming. Applied Mathematics and Computation, 249(??):1–18, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013812.

Anastassiou:2012:QAF

[AM12a] George A. Anastassiou and Razvan A. Mezei. Quantitative approximation by fractional smooth general singular operators. Applied Mathematics and Computation, 218(11):6200–6213, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013695.

Arteaga:2012:HTT

[AM12b] Cristian Arteaga and Isabel Marrero. The Hankel transform of tempered Boehmians via the exchange property. Applied Mathematics and Computation, 219(3):810–818, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200642X.

Asaad:2012:EGT

[AM12c] Magdy G. Asaad and Wen-Xiu Ma. Extended Gram-type determinant, wave and rational solutions to two (3 + 1)-dimensional nonlinear evolution equations. Applied Mathematics and Computation, 219(1):213–225, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005905.

Asaad:2012:PSD

[AM12d] Magdy G. Asaad and Wen-Xiu Ma. Pfaffian solutions to a (3 + 1)-dimensional generalized B-type Kadomtsev– Petviashvili equation and its modified counterpart. Applied Mathematics and Computation, 218(9):5524–5542, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013725.

Avkiran:2012:SAN

[AM12e] Necmi K. Avkiran and Alan McCrystal. Sensitivity analysis of network DEA: NSBM versus NRAM. Applied Mathematics and Computation, 218(22):11226–11239, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004997.

Alghamdi:2013:HMT

[AM13a] Mohammed A. Alghamdi and M. Mursaleen. Hankel matrix transformation of the Walsh–Fourier series. Applied Mathematics and Computation, 224(??):278–282, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009430.

Andras:2013:UHS

[AM13b] Szilárd András and Alpár Richárd Mészáros. Ulam–Hyers stability of dynamic equations on time scales via Picard operators. Applied Mathematics and Computation, 219(9):4853–4864, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011514.

Aziz:2013:RSU

[AM13c] Taha Aziz and F. M. Mahomed. Reductions and solutions for the unsteady flow of a fourth grade fluid on a porous plate. Applied Mathematics and Computation, 219(17):9187–9195, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002841.

Ahmad:2014:VMS

[AM14a] Daud Ahmad and Bilal Masud. Variational minimization on stringrearrangement surfaces, illustrated by an analysis of the bilinear interpolation. Applied Mathematics and Computation, 233(??):72– 84, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002252.

AlNoufaey:2014:SAS

[AM14b] K. S. Al Noufaey and T. R. Marchant. Semianalytical solutions for the reversible Selkov model with feedback delay. Applied Mathematics and Computation, 232(??):49– 59, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000964.

Andersson:2014:BLD

[AM14c] Helge I. Andersson and Swati Mukhopadhyay. Boundary layers due to shear flow over a still fluid: a direct integration approach. Applied Mathematics and Computation, 242(??):856–862, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008893.

Andras:2014:UHS

[AM14d] Szilárd András and Alpár Richárd Mészáros. Ulam–Hyers stability of elliptic partial differential equations in Sobolev spaces. Applied Mathematics and Computation, 229(??):131–138, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313012976.

Argyros:2014:LCA

[AM14e] Ioannis K. Argyros and Á. Alberto Magreñán. Local convergence analysis of proximal Gauss–Newton method for penalized nonlinear least squares problems. Applied Mathematics and Computation, 241(??):401–408, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314006274.

Argyros:2014:EAG

[AM14f] Ioannis K. Argyros and Ángel Alberto Magreñán. Extending the applicability of Gauss–Newton method for convex composite optimization on Riemannian manifolds. Applied Mathematics and Computation, 249(??): 453–467, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031401354X.

Ashrafi:2014:TFH

[AM14g] Nariman Ashrafi and Meysam Mohamadali. Transient flow and heat transfer of pseudoplastic fluids on a stretching sheet. Applied Mathematics and Computation, 228(??): 153–161, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313012356.

Azad:2010:ASI

[AMA10] H. Azad, M. T. Mustafa, and A. F. M. Arif. Analytic solutions of initial– boundary-value problems of transient conduction using symmetries. Applied Mathematics and Computation, 215(12):4132–4140, February 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309010637.

Abbas:2012:MPR

[AMA12a] Muhammad Abbas, Ahmad Abd. Majid, and Jamaludin Md. Ali. Monotonicitypreserving C^{2} rational cubic spline for monotone data. Applied Mathematics and Computation, 219(6):2885–2895, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200906X.

Acharya:2012:NET

[AMA12b] M. Acharya, S. N. Mohapatra, and B. P. Acharya. Numerical evaluation of two dimensional complex Cauchy principal value integrals. Applied Mathematics and Computation, 218(19):9980–9982, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S009630031200344X.

Abbas:2014:PPR

[AMA14a] Muhammad Abbas, Ahmad Abd. Majid, and Jamaludin Md. Ali. Positivitypreserving rational bicubic spline interpolation for 3D positive data. Applied Mathematics and Computation, 234(??):460–476, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002707.

Amador:2014:SSM

[Ama14b]  Julia Amador.  The

stochastic SIRA model for computer viruses. Applied Mathematics and Computation, 232(??):1112–1124, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001702.

Atif:2010:SSC

[AMB10] Sami Atif, Daniela Milovic, and Anjan Biswas. 1soliton solution of the complex KdV equation in plasmas with power law nonlinearity and timedependent coefficients.

Applied Mathematics and Computation, 217(4):1785– 1789, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310001931.

Abdoulkary:2014:ETW

[AMB^{+}14] Saïdou Abdoulkary, Alidou Mohamadou, Tibi Beda, Betchewe Gambo, Serge Y. Doka, Alim, and Aboubakar Mahamoudou. Exact traveling wave solutions to the nonlinear Schrödinger equation. Applied Mathematics and Computation, 233(??):109–115, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001544.

Amdjadi:2010:NMD

[Amd10] Faridon Amdjadi. A numerical method for the dynamics and stability of spiral waves. Applied Mathematics and Computation, 217(7):3385–3391, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310009495.

Avila:2012:TWL

[ÁME^{+}12] Ricardo P. Ávila, Paulo F. A. Mancera, Lourdes Esteva, Marcio R. Pie, and Cláudia P. Ferreira. Traveling waves in the Lethargic Crab Disease. Applied Mathematics and Computation, 218(19):9898–9910, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003347.

Alonso-Meijide:2012:CSP

[AMFM12] J. M. Alonso-Meijide, J. Freixas, and X. Molinero. Computation of several power indices by generating functions. Applied Mathematics and Computation, 219(8):3395–3402, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010089.

Amirov:2014:SFI

[Ami14] Sherif Amirov. On solvability of the first initialboundary value problem for quasilinear pseudoparabolic equations.

Applied Mathematics and Computation, 246(??): 112–120, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314010625.

Abbas:2014:ACT

[AMIR14] Muhammad Abbas, Ahmad Abd. Majid, Ahmad Izani Md. Ismail, and Abdur Rashid. The application of cubic trigonometric B-spline to the numerical solution of the hyperbolic problems. Applied Mathematics and Computation, 239(??):74–88, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005645.

Adnan:2011:MME

[AMM11] R. Adnan, H. H. Masjuki, and T. M. I. Mahlia. Mathematical modeling on the effect of equivalence ratio in emission characteristics of compression ignition engine with hydrogen substitution. Applied Mathematics and Computation, 217(13):6144–6158, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013135.

Ammar:2014:AID

[Amm14] Sherif I. Ammar. Analysis of an M/M/1 driven fluid queue with multiple exponential vacations. Applied Mathematics and Computation, 227(??):329–334, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011417.

Agahi:2010:BTI

[AMO^{+}10] Hamzeh Agahi, Radko Mesiar, Yao Ouyang, Endre Pap, and Mirjana Štrboja. Berwald type inequality for Sugeno integral. Applied Mathematics and Computation, 217(8):4100–4108, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010581.

Arponen:2014:CAG

[AMPT14] Teijo Arponen, Andreas Müller, Samuli Piipponen, and Jukka Tuomela. Computational algebraic geometry and global analysis of regional manipulators. Applied Mathematics and Computation, 232(??):820–835, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001830.

Amat:2013:TSR

[AMR13] S. Amat, Á. A. Magreñán, and N. Romero. On a twostep relaxed Newton-type method. Applied Mathematics and Computation, 219(24):11341–11347, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005626.

Al-Mdallal:2014:CCP

[AMS14] Qasem M. Al-Mdallal and Muhammed I. Syam. The Chebyshev collocationpath following method for solving sixth-order Sturm– Liouville problems. Applied Mathematics and Computation, 232(??): 391–398, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314001209.

Aledo:2012:PDS

[AMV12] Juan A. Aledo, S. Martínez, and Jose C. Valverde. Parallel dynamical systems over directed dependency graphs. Applied Mathematics and Computation, 219(3):1114–1119, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007187.

Anco:2012:TWC

[AMW12] Stephen C. Anco, Mohammad Mohiuddin, and Thomas Wolf. Traveling waves and conservation laws for complex mKdVtype equations. Applied Mathematics and Computation, 219(2):679–698, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006698.

Abbas:2012:CFPa

[AN– D12] Mujahid Abbas, Talat Nazir, and Dragan – Dorić. Common fixed point of mappings satisfying (E.A) property in generalized metric spaces. Applied Mathematics and Computation, 218(14):7665–7670, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000732.

Alvarez-Nodarse:2012:RKT

[ÁNA12] R. Álvarez-Nodarse and R. Sevinik Adigüzel. The q-Racah–Krall-type polynomials. Applied Mathematics and Computation, 218(23):11362–11369, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005000.

Andrew:2011:FDM

[And11] Alan L. Andrew. Finite difference methods for half inverse Sturm–Liouville problems. Applied Mathematics and Computation, 218(2):445–457, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007661.

Andreyev:2012:DIP

[And12a] M. Yu. Andreyev. Defaults and infinite prices in a stochastic pure exchange model. Applied Mathematics and Computation, 218(16):8112–8119, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011805.

Andrzejczak:2012:SNR

[And12b] Grzegorz Andrzejczak. A simple NSC–RKHS algorithm for solving linear boundary value problems

— error bounds and implementations. Applied Mathematics and Computation, 218(22):11009–11020, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004614.

Andrei:2013:TTC

[And13a]  Neculai Andrei.  On

three-term conjugate gradient algorithms for unconstrained optimization. Applied Mathematics and Computation, 219(11): 6316–6327, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312012908.

Andrzejczak:2013:FVN

[And13b] G. Andrzejczak. Fast variants of the NSC–RKHS algorithm for solving linear boundary value problems. Applied Mathematics and Computation, 219(22):10706–10725, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004815.

Abbasbandy:2014:NAS

[ANHA14] Saeid Abbasbandy, Rahila Naz, Tasawar Hayat, and A. Alsaedi. Numerical and analytical solutions for Falkner–Skan flow of MHD Maxwell fluid. Applied Mathematics and Computation, 242(??):569–575, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006420.

Anonymous:2011:EB

[Ano11]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 218(1):??, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311008605.

Anonymous:2012:EB

[Ano12]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 219(1):??, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200745X.

Abbas:2010:SPP

[ANR10] Mujahid Abbas, Talat Nazir, and Stojan Radenović. Some periodic point results in generalized metric spaces. Applied Mathematics and Computation, 217(8):4094–4099, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001057X.

Abbas:2012:CFPb

[ANR12] Mujahid Abbas, Talat Nazir, and Stojan Radenović. Common fixed point of generalized weakly contractive maps in partially ordered G-metric spaces. Applied Mathematics and Computation, 218(18):9383–9395, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002548.

Abbas:2013:CFP

[ANR13] Mujahid Abbas, Talat Nazir, and Stojan Radenović. Common fixed point of power contraction mappings satisfying (E.A) property in generalized metric spaces. Applied Mathematics and Computation, 219(14):7663–7670, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000957.

Anbumalar:2014:NNM

[ANR14] S. Anbumalar, R. Ananda Natarajan, and P. Rameshbabu. Non-negative matrix factorization algorithm for the deconvolution of one dimensional chromatograms. Applied Mathematics and Computation, 241(??):242–258, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006870.

Ansari:2013:FEO

[Ans13]  Alireza Ansari.  Frac-

tional exponential operators and nonlinear partial fractional differential equations in the Weyl fractional derivatives. Applied Mathematics and Computation, 220(??):149–154, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006279.

Araujo:2014:ADI

[ANS14] Adérito Araújo, Cidália Neves, and Ercília Sousa. An alternating direction implicit method for a second-order hyperbolic diffusion equation with convection. Applied Mathematics and Computation, 239(??):17–28, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005840.

Antczak:2011:NEE

[Ant11a] Tadeusz Antczak. A new exact exponential penalty function method and nonconvex mathematical programming. Applied Mathematics and Computation, 217(15):6652–6662, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000671.

Antczak:2011:NMP

[Ant11b] Tadeusz Antczak. Nonsmooth minimax programming under locally Lipschitz (ϕ, ρ)-invexity. Applied Mathematics and Computation, 217(23): 9606–9624, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005947.

Antczak:2012:VEP

[Ant12] Tadeusz Antczak. The vector exact l_{1} penalty method for nondifferentiable convex multiobjective programming problems. Applied Mathematics and Computation, 218(18):9095–9106, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001968.

Antunes:2013:OSQ

[Ant13] Pedro R. S. Antunes. Optimization of sums and quotients of Dirichlet– Laplacian eigenvalues. Applied Mathematics and Computation, 219(9):4239– 4254, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312011253.

Ashyralyev:2012:BST

[AO12] Allaberen Ashyralyev and Elif Ozturk. On Bitsadze– Samarskii type nonlocal boundary value problems for elliptic differential and difference equations: Wellposedness. Applied Mathematics and Computation, 219(3):1093–1107, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007163.

Ademola:2010:BRC

[AOAA10] Timothy A. Ademola, Michael O. Ogundiran, Peter O. Arawomo, and Olufemi Adeyinka Adesina. Boundedness results for a certain third order nonlinear differential equation. Applied Mathematics and Computation, 216(10):3044–3049, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000442X.

Agahi:2011:HMT

[AOM^{+}11] Hamzeh Agahi, Yao Ouyang, Radko Mesiar, Endre Pap, and Mirjana Strboja. Hölder and Minkowski type inequalities for pseudointegral. Applied Mathematics and Computation, 217(21):8630–8639, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004681.

Agryzkov:2012:ARN

[AOTV12] Taras Agryzkov, Jose L. Oliver, Leandro Tortosa, and Jose F. Vicent. An algorithm for ranking the nodes of an urban network based on the concept of PageRank vector. Applied Mathematics and Computation, 219 (4):2186–2193, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008570.

Agryzkov:2014:NBC

[AOTV14] Taras Agryzkov, Jose L.

Oliver, Leandro Tortosa, and Jose Vicent. A new betweenness centrality measure based on an algorithm for ranking the nodes of a network. Applied Mathematics and Computation, 244(??):467–478, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009837.

Aouf:2010:SIR

[Aou10] M. K. Aouf. Some inclusion relationships associated with Dziok– Srivastava operator. Applied Mathematics and Computation, 216(2):431– 437, March 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310000470.

Aouaoui:2011:MRS

[Aou11a] Sami Aouaoui. Multiplicity result for some nonlocal anisotropic equation via nonsmooth critical point theory approach. Applied Mathematics and Computation, 218(2):532–541, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007788.

Aouicha:2011:CSS

[Aou11b] Hichem Ben Aouicha. Computation of the spectra of some integral operators and application to the numerical solution of some linear integral equations. Applied Mathematics and Computation, 218(7):3217–3229, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011027.

Aouaoui:2012:ETS

[Aou12]  Sami Aouaoui.  Exis-

tence of three solutions for some equation of Kirchhoff type involving variable exponents. Applied Mathematics and Computation, 218(13):7184–7192, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015773.

Arun:2010:EKC

[AP10] K. R. Arun and Phoolan Prasad. Eigenvalues of kinematical conservation laws (KCL) based 3D weakly nonlinear ray theory (WNLRT). Applied Mathematics and Computation, 217(5):2285–2288, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310007265.

Aksoylu:2011:VTD

[AP11] Burak Aksoylu and Michael L.

Parks. Variational theory and domain decomposition for nonlocal problems. Applied Mathematics and Computation, 217(14):6498–6515, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000439.

Aparicio:2013:WDE

[AP13] Juan Aparicio and Jesus T. Pastor. A welldefined efficiency measure for dealing with closest targets in DEA. Applied Mathematics and Computation, 219(17):9142–9154, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002956.

Arachchige:2014:FVM

[AP14] Jayantha Pasdunkorale Arachchige and Graeme J. Pettet. A finite volume method with linearisation in time for the solution of advection-reaction-diffusion systems. Applied Mathematics and Computation, 231(??):445–462, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000186.

Apaydin:2010:UPS

[Apa10] Gökhan Apaydin. On the use of physical spline finite element method for acoustic scattering. Applied Mathematics and Computation, 215(10):3576–3588, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009709.

Ali:2013:UIM

[APA13] Musrrat Ali, Millie Pant, and Ajith Abraham. Unconventional initialization methods for differential evolution. Applied Mathematics and Computation, 219(9):4474–4494, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010697.

Arthi:2014:ECR

[APJ14] G. Arthi, Ju H. Park, and H. Y. Jung. Existence and controllability results for second-order impulsive stochastic evolution systems with state-dependent delay. Applied Mathematics and Computation, 248(??):328–341, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013095.

Abrarov:2011:EAI

[AQ11] S. M. Abrarov and B. M.

Quine. Efficient algorithmic implementation of the Voigt /complex error function based on exponential series approximation. Applied Mathematics and Computation, 218(5):1894–1902, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311009179.

Agratini:2011:EBI

[AR11] Octavian Agratini and Cristina Radu. An extension based on qR-integral for a sequence of operators. Applied Mathematics and Computation, 218(1):140–147, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007545.

Akgun:2012:EGH

[AR12a] F. Aydin Akgun and B. E. Rhoades. The equivalence of generalized Hölder and Cesáro matrices. Applied Mathematics and Computation, 218(12):6677–6683, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015013.

Albrecht:2012:CPI

[AR12b] Gudrun Albrecht and Lucia Romani. Convexity preserving interpolatory subdivision with conic precision. Applied Mathematics and Computation, 219(8):4049–4066, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010648.

Akgun:2013:PSQ

[AR13] F. Aydin Akgun and B. E. Rhoades. Properties of some q-Hausdorff matrices. Applied Mathematics and Computation, 219(14):7392–7397, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000519.

Acu:2014:NRC

[AR14a] Ana-Maria Acu and Maria-Daniela Rusu. New results concerning Chebyshev– Grüss-type inequalities via discrete oscillations. Applied Mathematics and Computation, 243(??): 585–593, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314008443.

Akgun:2014:MPD

[AR14b] F. Aydin Akgun and B. E. Rhoades. Monotonicitypreserving doubly infinite matrices. Applied Mathematics and Computation, 235(??):105–109, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003208.

Akila:2014:NCQ

[AR14c] Lakshmanan Akila and Rajakumar Roopkumar. A natural convolution of quaternion valued functions and its applications. Applied Mathematics and Computation, 242(??):633–642, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008431.

Araci:2014:NII

[Ara14] Serkan Araci. Novel identities involving Genocchi numbers and polynomials arising from applications of umbral calculus. Applied Mathematics and Computation, 233(??):599–607, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000502.

Ashish:2014:JSM

[ARC14] Ashish, Mamta Rani, and Renu Chugh. Julia sets and Mandelbrot sets in Noor orbit. Applied Mathematics and Computation, 228(??):615–631, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012460.

Ardelean:2011:CBI

[Ard11] Gheorghe Ardelean. A comparison between iterative methods by using the basins of attraction. Applied Mathematics and Computation, 218(1):88–95, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007259.

Ardelean:2013:NTO

[Ard13] Gheorghe Ardelean. A new third-order Newtontype iterative method for solving nonlinear equations. Applied Mathematics and Computation, 219(18):9856–9864, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003937.

Aghamohamadi:2014:TSM

[ARE14] Masomeh Aghamohamadi, Jalil Rashidinia, and Reza Ezzati. Tension spline method for solution of non-linear Fisher equation. Applied Mathematics and Computation, 249(??):399–407, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012223.

Alvarez-Ramirez:2014:PMN

[ARG14] Martha Alvarez-Ramírez and Antonio García. Poincaré maps and near-collision dynamics for a restricted planar (n + 1)-body problem. Applied Mathematics and Computation, 233(??):328–337, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314002446.

Alam:2012:SST

[ARH^{+}12] M. Kamran Alam, M. T.

Rahim, T. Haroon, S. Islam, and A. M. Siddiqui. Solution of steady thin film flow of Johnson– Segalman fluid on a vertical moving belt for lifting and drainage problems using Adomian Decomposition Method. Applied Mathematics and Computation, 218(21):10413– 10428, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031200358X.

Alvarez-Ramirez:2013:SCC

[ARL13] Martha Alvarez-Ramírez and Jaume Llibre. The symmetric central configurations of the 4-body problem with masses m_{1} = m_{2} ≠ m_{3} = m_{4}. Applied Mathematics and Computation, 219(11):5996–6001, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013124.

Abbas:2010:CFP

[ARN10] Mujahid Abbas, B. E.

Rhoades, and Talat Nazir. Common fixed points for four maps in cone metric spaces. Applied Mathematics and Computation, 216(1):80–86, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000123.

Aghajani:2012:CFP

[ARR12] Asadollah Aghajani, Stojan Radenovíc, and Jamal Rezaei Roshan. Common fixed point results for four mappings satisfying almost generalized (ST)-contractive condition in partially ordered metric spaces. Applied Mathematics and Computation, 218(9):5665–5670, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014044.

Asafuddoula:2014:AHD

[ARS14] Md. Asafuddoula, Tapabrata Ray, and Ruhul Sarker. An adaptive hybrid differential evolution algorithm for single objective optimization. Applied Mathematics and Computation, 231(??):601–618, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000782.

Ahmad:2010:FPN

[AS10] Bashir Ahmad and S. Sivasundaram. On fourpoint nonlocal boundary value problems of nonlinear integro-differential equations of fractional order. Applied Mathematics and Computation, 217(2):480–487, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006399.

Adiguzelov:2011:RTS

[AS11a] Ehliman Adigüzelov and Yonca Sezer. The regularized trace of a self adjoint differential operator of higher order with unbounded operator coefficient. Applied Mathematics and Computation, 218(5):2113–2121, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009611.

An:2011:DRC

[AS11b] Chen An and Jian Su. Dynamic response of clamped axially moving beams: Integral transform solution. Applied Mathematics and Computation, 218(2):249–259, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007041.

Aygunes:2011:AHO

[AS11c] A. Ahmet Aygunes and Yilmaz Simsek. The action of Hecke operators to families of Weierstrass-type functions and Weber-type functions and their applications. Applied Mathematics and Computation, 218(3):678–682, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004589.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Abe:2012:HBC

[AS12a] Kuniyoshi Abe and Gerard L. G. Sleijpen. Hybrid Bi-CG methods with a Bi-CG formulation closer to the IDR approach. Applied Mathematics and Computation, 218(22):10889–10899, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004481.

Aristov:2012:MFC

[AS12b] Vladimir Aristov and Andrey Stroganov. A method of formalizing computer operations for solving nonlinear differential equations. Applied Mathematics and Computation, 218(16):8083–8098, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011994.

Arora:2013:NSB

[AS13a] Geeta Arora and Brajesh Kumar Singh. Numerical solution of Burgers’ equation with modified cubic B-spline differential quadrature method. Applied Mathematics and Computation, 224(??): 166–177, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313009399.

Assar:2013:IBH

[AS13b] Rodrigo Assar and David J. Sherman. Implementing biological hybrid systems: Allowing composition and avoiding stiffness. Applied Mathematics and Computation, 223(??):167–179, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008606.

Abdullah:2014:WPF

[AS14a] Abdullah and Firdous A. Shah. Wave packet frames on local fields of positive characteristic. Applied Mathematics and Computation, 249(??):133–141, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013654.

Ali:2014:NDD

[AS14b] M. Syed Ali and R. Saravanakumar. Novel delay-dependent robust H_{∞} control of uncertain systems with distributed time-varying delays. Applied Mathematics and Computation, 249(??):510–520, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014271.

An:2014:DAA

[AS14c] Chen An and Jian Su.

Dynamic analysis of axially moving orthotropic plates: Integral transform solution. Applied Mathematics and Computation, 228(??):489–507, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012848.

Andreev:2014:OBS

[AS14d] Viktor K. Andreev and Irina V. Stepanova. Ostroumov– Birikh solution of convection equations with nonlinear buoyancy force. Applied Mathematics and Computation, 228(??):59– 67, February 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313011624.

Asthana:2014:EHO

[AS14e] K. Asthana and T. K. Sengupta. An explicit higher order difference scheme on a compact stencil for elliptic equations on curvilinear geometries. Applied Mathematics and Computation, 242(??):143–158, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007140.

Asaithambi:2010:NSB

[Asa10a] Asai Asaithambi. Numerical solution of the Burgers’ equation by automatic differentiation. Applied Mathematics and Computation, 216(9):2700–2708, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003796.

Asaithambi:2010:NSO

[Asa10b] Asai Asaithambi. Numerical solution of the onedimensional time-independent Schrödinger’s equation by recursive evaluation of derivatives. Applied Mathematics and Computation, 215(12):4400–4405, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000111.

Aktas:2011:MEH

[AŠA11] Rabia Aktaş, Recep Šahin, and Abdullah Altin. On a multivariable extension of the Humbert polynomials. Applied Mathematics and Computation, 218(3):662–666, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001172.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Apostolopoulou:2010:CMB

[ASB10] M. S. Apostolopoulou, D. G. Sotiropoulos, and C. A. Botsaris. A curvilinear method based on minimal-memory BFGS updates. Applied Mathematics and Computation, 217(2):882–892, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007174.

Aydogdu:2010:PEG

[AŞK10] Halil Aydoḡdu, Birdal Şenoḡlu, and Mahmut Kara. Parameter estimation in geometric process with Weibull distribution. Applied Mathematics and Computation, 217(6):2657–2665, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008532.

Askar:2014:ICU

[Ask14] S. S. Askar. The impact of cost uncertainty on Cournot oligopoly game with concave demand function. Applied Mathematics and Computation, 232(??):144–149, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001349.

Ashrafi:2013:BEF

[ASKA13] H. Ashrafi, M. Shariyat, S. M. R. Khalili, and K. Asemi. A boundary element formulation for the heterogeneous functionally graded viscoelastic structures. Applied Mathematics and Computation, 225(??):246–262, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009971.

Aslan:2010:ALL

[Asl10a]  İsmail Aslan.  The

Ablowitz–Ladik lattice system by means of the extended (\frac{G′}{G})-expansion method. Applied Mathematics and Computation, 216(9):2778–2782, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003899.

Aslan:2010:GSP

[Asl10b] İsmail Aslan. Generalized solitary and periodic wave solutions to a (2 + 1)-dimensional Zakharov– Kuznetsov equation. Applied Mathematics and Computation, 217(4):1421– 1429, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309005396.

Aslan:2010:NEM

[Asl10c] İsmail Aslan. A note on the (\frac{G′}{G})-expansion method again. Applied Mathematics and Computation, 217(2):937–938, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310006703.

Aslan:2010:REF

[Asl10d] İsmail Aslan. Remark on: “Exp-function method for the exact solutions of fifth order KdV equation and modified Burgers equation” [Appl. Math. Comput. (2009) doi:10.1016/j.amc.2009.07.009]. Applied Mathematics and Computation, 217(7):2912, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310009628.

See [IU10].

Aslan:2011:AID

[Asl11a] İsmail Aslan. Analytic investigation of the (2 + 1)-dimensional Schwarzian Korteweg–de Vries equation for traveling wave solutions. Applied Mathematics and Computation, 217(12):6013–6017, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001341X.

Aslan:2011:EES

[Asl11b] Ismail Aslan. Exact and explicit solutions to nonlinear evolution equations using the division theorem. Applied Mathematics and Computation, 217(20):8134–8139, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003213.

Aslan:2012:AEFa

[Asl12a] Ismail Aslan. On the application of the Expfunction method to nonlinear differential-difference equations. Applied Mathematics and Computation, 218(19):9594–9595, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200330X.

Aslan:2012:AEFb

[Asl12b] Ismail Aslan. On the application of the Expfunction method to the KP equation for N-soliton solutions. Applied Mathematics and Computation, 219(6):2825–2828, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009459.

Abel:2011:NSM

[ASM11] M. Subhas Abel, P. G. Siddheshwar, and N. Mahesha. Numerical solution of the momentum and heat transfer equations for a hydromagnetic flow due to a stretching sheet of a nonuniform property micropolar liquid. Applied Mathematics and Computation, 217(12):5895–5909, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012981.

Alobaidi:2014:WSV

[ASM14] G. Alobaidi, C. J. Smith, and R. Mallier. Waves on a Stuart vortex. Applied Mathematics and Computation, 227(??):370–383, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011880.

Arunkumar:2012:RSC

[ASMA12] A. Arunkumar, R. Sakthivel, K. Mathiyalagan, and S. Marshal Anthoni. Robust stability criteria for discrete-time switched neural networks with various activation functions. Applied Mathematics and Computation, 218(22):10803–10816, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004043.

Alam:2014:TFF

[ASR^{+}14] M. K. Alam, A. M. Siddiqui, M. T. Rahim, S. Islam, E. J. Avital, and J. J. R. Williams. Thin film flow of magnetohydrodynamic (MHD) pseudoplastic fluid on vertical wall. Applied Mathematics and Computation, 245(??):544–556, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010042.

Alam:2012:TFF

[ASRI12] M. K. Alam, A. M.

Siddiqui, M. T. Rahim, and S. Islam. Thin-film flow of magnetohydrodynamic (MHD) Johnson– Segalman fluid on vertical surfaces using the Adomian decomposition method. Applied Mathematics and Computation, 219(8):3956–3974, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010338.

Abraham:2013:FNS

[ASS13] Siby Abraham, Sugata Sanyal, and Mukund Sanglikar. Finding numerical solutions of Diophantine equations using ant colony optimization. Applied Mathematics and Computation, 219(24):11376–11387, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005985.

Ahrari:2010:LCB

[ASSPA10] Ali Ahrari, Mohammad R. Saadatmand, Masoud Shariat-Panahi, and Ali A. Atai. On the limitations of classical benchmark functions for evaluating robustness of evolutionary algorithms. Applied Mathematics and Computation, 215(9):3222–3229, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009096.

Alouch:2011:SND

[ASST11] C. Alouch, P. Sablonnière, D. Sbibih, and M. Tahrichi. Superconvergent Nyström and degenerate kernel methods for eigenvalue problems. Applied Mathematics and Computation, 217(20):7851–7866, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001378.

Allouch:2012:SRB

[AST12] C. Allouch, D. Sbibih, and M. Tahrichi. Spectral refinement based on superconvergent Nyström and degenerate kernel methods. Applied Mathematics and Computation, 218(22):10777–10790, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200450X.

Akyildiz:2010:ANR

[ASVV10] F. Talay Akyildiz, Dennis A. Siginer, K. Vajravelu, and Robert A. Van Gorder. Analytical and numerical results for the Swift–Hohenberg equation. Applied Mathematics and Computation, 216(1): 221–226, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000548.

Ali:2010:NPP

[AT10a] Norhashidah Hj. Mohd. Ali and Khoo Kok Teong. Numerical performance of parallel group explicit solvers for the solution of fourth order elliptic equations. Applied Mathematics and Computation, 217(6):2737–2749, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008635.

Almeida:2010:LDM

[AT10b] Ricardo Almeida and Delfim F. M. Torres. Leitmann’s direct method for fractional optimization problems. Applied Mathematics and Computation, 217(3):956–962, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003395.

Allem:2011:GMP

[AT11] Luiz Emilio Allem and Vilmar Trevisan. GCD of multivariate polynomials via Newton polytopes. Applied Mathematics and Computation, 217(21):8377–8386, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004012.

Akashi:2012:SCT

[AT12] Shigeo Akashi and Wataru Takahashi. Strong convergence theorem for nonexpansive mappings on star-shaped sets in Hilbert spaces. Applied Mathematics and Computation, 219(4):2035–2040, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008399.

Atakut:2011:DBT

[Ata11]  Çigdem Atakut.  On

derivatives of Bernstein type rational functions of two variables. Applied Mathematics and Computation, 218(3):673–677, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004590.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Aubin:2010:MTM

[Aub10] Jean-Pierre Aubin. Macroscopic traffic models: Shifting from densities to “Celerities”. Applied Mathematics and Computation, 217(3):963–971, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002006.

Aly:2014:ENS

[AV14] E. H. Aly and K. Vajravelu. Exact and numerical solutions of MHD nano boundary-layer flows over stretching surfaces in a porous medium. Applied Mathematics and Computation, 232(??):191–204, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014379.

Alvarez-Vazquez:2010:OCB

[AVF10] Lino J. Alvarez-Vázquez and Francisco J. Fernández. Optimal control of a bioreactor. Applied Mathematics and Computation, 216(9):2559–2575, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003516.

Awawdeh:2012:NES

[Awa12] Fadi Awawdeh. New exact solitary wave solutions of the Zakharov– Kuznetsov equation in the electron-positron-ion plasmas. Applied Mathematics and Computation, 218(13):7139–7143, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015669.

Aschariyaphotha:2011:SSS

[AWHW11] N. Aschariyaphotha, P. Wongwises, U. W. Humphries, and S. Wongwises. Study of storm surge due to Typhoon Linda (1997) in the Gulf of Thailand using a three dimensional ocean model. Applied Mathematics and Computation, 217(21):8640–8654, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004735.

Alaoui:2010:EVW

[AWS10] Abdelhamid Hamidi Alaoui, Luther White, and Alan Shapiro. Estimation of vertical wind velocity using Doppler radar data with vorticity constraints. Applied Mathematics and Computation, 215(12):4119–4131, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010625.

Agahi:2010:GHT

[AY10] Hamzeh Agahi and M. A. Yaghoobi. General Hardy type inequality for seminormed fuzzy integrals. Applied Mathematics and Computation, 216(7):1972– 1977, June 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300310002699.

Abdeldaim:2011:SNI

[AY11a] A. Abdeldaim and M. Yakout. On some new integral inequalities of Gronwall– Bellman–Pachpatte type. Applied Mathematics and Computation, 217(20): 7887–7899, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031100316X.

Ang:2011:CVB

[AY11b] W. T. Ang and B. I.

Yun. A complex variable boundary element method for axisymmetric heat conduction in a nonhomogeneous solid. Applied Mathematics and Computation, 218(5):2225–2236, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009726.

Ashyralyev:2011:SDS

[AY11c] Allaberen Ashyralyev and Ozgur Yildirim. Stable difference schemes for the hyperbolic problems subject to nonlocal boundary conditions with selfadjoint operator. Applied Mathematics and Computation, 218(3):1124–1131, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005352.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Arandiga:2014:CAM

[AY14] Francesc Aràndiga and Dionisio F. Yáñez. Cellaverage multiresolution based on local polynomial regression. application to image processing. Applied Mathematics and Computation, 245(??):1–16, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010364.

Akkaya:2013:NSP

[AYS13] Tugçe Akkaya, Salih Yalçinbas, and Mehmet Sezer. Numeric solutions for the pantograph type delay differential equation using First Boubaker polynomials. Applied Mathematics and Computation, 219(17):9484–9492, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002749.

Anderson:2010:PSS

[AZ10] Douglas R. Anderson and Chengbo Zhai. Positive solutions to semi-positone second-order three-point problems on time scales. Applied Mathematics and Computation, 215(10): 3713–3720, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300309009990.

Ai:2014:FTS

[AZ14a] Zidong Ai and Guangdeng Zong. Finite-time stochastic input-to-state stability of impulsive switched stochastic nonlinear systems. Applied Mathematics and Computation, 245(??):462–473, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010595.

Al-Zhour:2014:GVS

[AZ14b] Zeyad Abdel Aziz Al-Zhour. The general (vector) solutions of such linear (coupled) matrix fractional differential equations by using Kronecker structures. Applied Mathematics and Computation, 232(??):498–510, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001167.

See retraction notice [AZ19].

Antczak:2014:SOI

[AZ14c] Tadeusz Antczak and G. J. Zalmai. Second order (Φ, ρ)–V -invexity and duality for semi-infinite minimax fractional programming. Applied Mathematics and Computation, 227(??):831–856, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011077.

AzizAl-Zhour:2019:RNG

[AZ19] Zeyad Abdel AzizAl-Zhour. Retraction notice to “The general (Vector) solutions of such linear (coupled) matrix fractional differential equations by using Kronecker structures” [Applied Mathematics and Computation 232 (2014) 498–510]. Applied Mathematics and Computation, 361(??):889, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304357.

See [AZ14b].

Azari:2014:SLB

[Aza14] Mahdieh Azari. Sharp lower bounds on the Narumi–Katayama index of graph operations. Applied Mathematics and Computation, 239(??):409–421, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006286.

Az-Zobi:2010:NCP

[AZAK10] Emad A. Az-Zo’bi and Kamel Al-Khaled. A new convergence proof of the Adomian decomposition method for a mixed hyperbolic elliptic system of conservation laws. Applied Mathematics and Computation, 217(8):4248–4256, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310011008.

Agarwal:2013:NKT

[AZL13] Ravi P. Agarwal, Chenghui Zhang, and Tongxing Li. New Kamenevtype oscillation criteria for second-order nonlinear advanced dynamic equations. Applied Mathematics and Computation, 225(??):822–828, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010515.

Bulatovic:2011:COS

[B– D11] Radovan R. Bulatović and Stevan R. – Dordević. Control of the optimum synthesis process of a four-bar linkage whose point on the working member generates the given path. Applied Mathematics and Computation, 217(23):9765–9778, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006345.

Bekir:2010:ESN

[BA10] Ahmet Bekir and Esin Aksoy. Exact solutions of nonlinear evolution equations with variable coefficients using exp-function method. Applied Mathematics and Computation, 217(1):430–436, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005904.

Baranger:2011:CLG

[BA11] T. N. Baranger and S. Andrieux. Constitutive law gap functionals for solving the Cauchy problem for linear elliptic PDE. Applied Mathematics and Computation, 218(5):1970–1989, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009283.

Bozorgnia:2013:NAV

[BA13] Farid Bozorgnia and Avetik Arakelyan. Numerical algorithms for a variational problem of the spatial segregation of reactiondiffusion systems. Applied Mathematics and Computation, 219(17):8863–8875, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003354.

Bnouhachem:2014:DLA

[BA14a] Abdellah Bnouhachem and Qamrul Hasan Ansari. A descent LQP alternating direction method for solving variational inequality problems with separable structure. Applied Mathematics and Computation, 246(??):519–532, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011631.

Boulbrachene:2014:MNE

[BA14b] Messaoud Boulbrachene and Qais Al Farei. Maximum norm error analysis of a nonmatching grids finite element method for linear elliptic PDEs. Applied Mathematics and Computation, 238(??):21– 29, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005177.

Bashir:2013:RQT

[BAA13] Uzma Bashir, Muhammad Abbas, and Jamaludin Md Ali. The G^{2} and C^{2} rational quadratic trigonometric Bézier curve with two shape parameters with applications. Applied Mathematics and Computation, 219(20):10183– 10197, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003718.

Babajee:2012:SIP

[Bab12] Diyashvir Kreetee Rajiv Babajee. Several improvements of the 2-point third order midpoint iterative method using weight functions. Applied Mathematics and Computation, 218(15):7958–7966, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001373.

Baccouch:2014:AEP

[Bac14] Mahboub Baccouch. Asymptotically exact a posteriori LDG error estimates for one-dimensional transient convection-diffusion problems. Applied Mathematics and Computation, 226(??):455–483, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010837.

Bellamine:2013:NSD

[BAE13] F. H. Bellamine, A. Almansoori, and A. Elkamel. Numerical simulation of distributed dynamic systems using hybrid intelligent computing combined with generalized similarity analysis. Applied Mathematics and Computation, 223(??):88–100, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008515.

Belendez:2012:CFE

[BAFP12] A. Beléndez, E. Arribas, J. Francés, and I. Pascual. Comments on “A finite extensibility nonlinear oscillator”. Applied Mathematics and Computation, 218(10):6168–6175, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014706.

See [Feb11].

Bahrouni:2013:ISC

[Bah13a] Anouar Bahrouni. Infinite solutions for a class of Brézis–Nirenberg equations with an indefinite linear and nonlinear terms in sign. Applied Mathematics and Computation, 219(24):11198–11205, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005638.

Bahyrycz:2013:SSG

[Bah13b]  Anna Bahyrycz.  On

solutions of the second generalization of d’Alembert’s functional equation on a restricted domain. Applied Mathematics and Computation, 223(??):209–215, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008618.

Bai:2010:PSS

[Bai10] Zhanbing Bai. Positive solutions of some nonlocal fourth-order boundary value problem. Applied Mathematics and Computation, 215(12):4191–4197, February 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630030901087X.

Bai:2013:SSL

[Bai13] Cheng-Lin Bai. Stochastic soliton-like solutions and theirs stochastic excitations under a (2 + 1)-dimensional stochastic dispersive long wave system. Applied Mathematics and Computation, 219(14):7795–7804, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001215.

Bai:2014:CWE

[Bai14] Cheng-Lin Bai. Complex wave excitations and their interaction behaviors for higher-dimensional nonlinear model. Applied Mathematics and Computation, 240(??):115–121, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006365.

Bougoffa:2010:SBV

[BAJK10] Lazhar Bougoffa, Hind K. Al-Jeaid, and Ammar Khanfer. On the solutions of a boundary value problem of linear thermoelasticity system with nonlocal conditions. Applied Mathematics and Computation, 217(8):4227–4233, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010921.

Baliarsingh:2013:SND

[Bal13] P. Baliarsingh. Some new difference sequence spaces of fractional order and their dual spaces. Applied Mathematics and Computation, 219(18):9737–9742, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003342.

Balbus:2014:ASC

[Bal14]  Joanna Balbus.  At-

tractivity and stability in the competitive systems of PDEs of Kolmogorov type. Applied Mathematics and Computation, 237(??):69–76, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314004871.

Banerjee:2014:CSD

[Ban14] Malay Banerjee. Comments on “J. Dhar, R.S. Baghel, A.K. Sharma, Role of instant nutrient replenishment on plankton dynamics with diffusion in a closed system: a pattern formation, Appl. Math. Comput. 218 (2012) 8925–8936”. Applied Mathematics and Computation, 232(??):771–774, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400068X.

See [DBS12].

Bamigbola:2014:MME

[BAO14] O. M. Bamigbola, M. M. Ali, and M. O. Oke. Mathematical modeling of electric power flow and the minimization of power losses on transmission lines. Applied Mathematics and Computation, 241(??):214–221, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007048.

Belendez:2014:EAS

[BAP^{+}14] A. Beléndez, E. Arribas, C. Pascual, T. Beléndez, M. L. Alvarez, and A. Hernández. Exact and approximate solutions for the anti-symmetric quadratic truly nonlinear oscillator. Applied Mathematics and Computation, 246(??):355–364, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010996.

Bartoszewski:2010:NAN

[Bar10] Z. Bartoszewski. A new approach to numerical solution of fixed-point problems and its application to delay differential equations. Applied Mathematics and Computation, 215(12):4320–4331, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309011151.

Badghaish:2011:CPO

[BAR11] Abeer O. Badghaish, Rosihan M. Ali, and V. Ravichandran. Closure properties of operators on the Ma–Minda type starlike and convex functions. Applied Mathematics and Computation, 218(3):667–672, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000828.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Barbu:2013:SEF

[Bar13]  Luminita Barbu.  On

some estimates for a fluid surface in a short capillary tube. Applied Mathematics and Computation, 219(15):8192–8197, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001379.

Basha:2012:BPP

[Bas12a] S. Sadiq Basha. Best proximity point theorems: an exploration of a common solution to approximation and optimization problems. Applied Mathematics and Computation, 218(19):9773–9780, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002652.

See remarks [JS14b].

Bassiouny:2012:TEB

[Bas12b] E. Bassiouny. Thermoelastic behavior of thin sandwich panel made of piezoelectric layers. Applied Mathematics and Computation, 218(20): 10009–10021, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312002615.

Bas:2013:RAD

[Bas13] Esra Bas. A robust approach to the decision rules of NPV and IRR for simple projects. Applied Mathematics and Computation, 219(11):5901–5908, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013070.

Batkam:2014:SES

[Bat14] Cyril Joel Batkam. On a superquadratic elliptic system with strongly indefinite structure. Applied Mathematics and Computation, 233(??):243–251, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002392.

Bahri:2011:TDQ

[BAV11] Mawardi Bahri, Ryuichi Ashino, and Rémi Vaillancourt. Two-dimensional quaternion wavelet transform. Applied Mathematics and Computation, 218(1):10–21, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006990.

Balint:2011:PRT

[BB11a] Agneta M. Balint and Stefan Balint. Pulling rate, thermal and capillary conditions setting for rod growth. Applied Mathematics and Computation, 218(5):1634–1649, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008769.

Bardet:2011:CRC

[BB11b] Magali Bardet and Islam Boussaada. Complexity reduction of C-Algorithm. Applied Mathematics and Computation, 217(17):7318–7323, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001998.

Belbas:2011:NSM

[BB11c] S. A. Belbas and Yuriy Bulka. Numerical solution of multiple nonlinear Volterra integral equations. Applied Mathematics and Computation, 217(9):4791–4804, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001146X.

Borcut:2012:TCTa

[BB12] Marin Borcut and Vasile Berinde. Tripled coincidence theorems for contractive type mappings in partially ordered metric spaces. Applied Mathematics and Computation, 218(10):5929–5936, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013920.

Backes:2013:TAU

[BB13a] André R. Backes and Odemir M. Bruno. Texture analysis using volumeradius fractal dimension. Applied Mathematics and Computation, 219(11):5870–5875, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012854.

Bez:2013:NMB

[BB13b] H. E. Bez and N. Bez. New minimal bounds for the derivatives of rational Bézier paths and rational rectangular Bézier surfaces. Applied Mathematics and Computation, 225(??):475–479, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010187.

Bedin:2014:BEC

[BB14a] Luciano Bedin and Fermín S. Viloche Bazán. On the 2D bioheat equation with convective boundary conditions and its numerical realization via a highly accurate approach. Applied Mathematics and Computation, 236(??):422–436, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004330.

Belbachir:2014:AAC

[BB14b] Hacène Belbachir and Athmane Benmezai. An alternative approach to Cigler’s q-lucas polynomials. Applied Mathematics and Computation, 226(??):691–698, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010643.

Berntsson:2014:CIP

[BB14c] Fredrik Berntsson and George Baravdish. Coefficient identification in PDEs applied to image inpainting. Applied Mathematics and Computation, 242(??):227–235, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007164.

Bailey:2012:HPC

[BBB12] David H. Bailey, Roberto Barrio, and Jonathan M.

Borwein. High-precision computation: Mathematical physics and dynamics. Applied Mathematics and Computation, 218(20): 10106–10121, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://docserver.
carma.newcastle.edu.
au/775/;   http://www.
sciencedirect.com/science/
article/pii/S0096300312003505.

Barik:2014:TSS

[BBC14] S. K. Barik, M. P. Biswal, and D. Chakravarty. Twostage stochastic programming problems involving multi-choice parameters. Applied Mathematics and Computation, 240(??):109–114, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003981.

Bailey:2013:EFS

[BBCR13] David H. Bailey, Jonathan M. Borwein, Richard E. Crandall, and Michael G. Rose. Expectations on fractal sets. Applied Mathematics and Computation, 220(??):695–721, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007273.

Bastinec:2014:DPM

[BBDS14] Jaromír Bastinec, Leonid Berezansky, Josef Diblík, and Zdenek Smarda. On a delay population model with a quadratic nonlinearity without positive steady state. Applied Mathematics and Computation, 227(??):622–629, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012307.

Bazan:2012:GRP

[BBF12a] Fermín S. Viloche Bazán, Leonardo S. Borges, and Juliano B. Francisco. On a generalization of Regińska’s parameter choice rule and its numerical realization in large-scale multi-parameter Tikhonov regularization. Applied Mathematics and Computation, 219(4):2100–2113, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312008478.

Bielczyk:2012:DCS

[BBF12b] Natalia Bielczyk, Marek Bodnar, and Urszula Foryś. Delay can stabilize: Love affairs dynamics. Applied Mathematics and Computation, 219(8):3923–3937, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010296.

Bodnar:2012:ACF

[BBFK12] T. Bodnár, L. Benes, Ph. Fraunié, and K. Kozel. Application of compact finite-difference schemes to simulations of stably stratified fluid flows. Applied Mathematics and Computation, 219(7):3336– 3353, December 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010988.

Berezansky:2013:MGMa

[BBI13a] Leonid Berezansky, Elena Braverman, and Lev Idels. Mackey–Glass model of hematopoiesis with monotone feedback revisited. Applied Mathematics and Computation, 219(9):4892–4907, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010685.

Berezansky:2013:MGMb

[BBI13b] Leonid Berezansky, Elena Braverman, and Lev Idels. Mackey–Glass model of hematopoiesis with nonmonotone feedback: Stability, oscillation and control. Applied Mathematics and Computation, 219(11):6268–6283, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013197.

Berezansky:2014:NGE

[BBI14] Leonid Berezansky, Elena Braverman, and Lev Idels. New global exponential stability criteria for nonlinear delay differential systems with applications to BAM neural networks. Applied Mathematics and Computation, 243(??):899–910, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009035.

Banas:2010:EGA

[BBJ10] Józef Banaś, Krishnan Balachandran, and Diana Julie. Existence and global attractivity of solutions of a nonlinear functional integral equation. Applied Mathematics and Computation, 216(1):261–268, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000627.

Bnouhachem:2013:IAD

[BBK13] Abdellah Bnouhachem, Hafida Benazza, and Mohamed Khalfaoui. An inexact alternating direction method for solving a class of structured variational inequalities. Applied Mathematics and Computation, 219(14):7837–7846, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300129X.

Basar:2010:DGM

[BBS10] Tamer Başar, Alain Bensoussan, and Suresh P. Sethi. Differential games with mixed leadership: The open-loop solution. Applied Mathematics and Computation, 217(3):972– 979, October 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310000615.

Basto:2013:CBA

[BBSC13] Mário Basto, Luís P.

Basto, Viriato Semiao, and Francisco L. Calheiros. Contrasts in the basins of attraction of structurally identical iterative root finding methods. Applied Mathematics and Computation, 219(15):7997–8008, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002014.

Baksalary:2010:EBR

[BBT10] Oskar Maria Baksalary, Dennis S. Bernstein, and Götz Trenkler. On the equality between rank and trace of an idempotent matrix. Applied Mathematics and Computation, 217(8):4076–4080, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310010520.

Bawa:2010:HOG

[BC10a] R. K. Bawa and C. Clavero. Higher order global solution and normalized flux for singularly perturbed reaction–diffusion problems. Applied Mathematics and Computation, 216(7):2058–2068, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000278X.

Bilige:2010:ESE

[BC10b] Sudao Bilige and Temuer Chaolu. An extended simplest equation method and its application to several forms of the fifth-order KdV equation. Applied Mathematics and Computation, 216(11):3146–3153, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004492.

Badia:2011:CNC

[BC11a] Santiago Badia and Ramon Codina. A combined nodal continuousdiscontinuous finite element formulation for the Maxwell problem. Applied Mathematics and Computation, 218(8):4276–4294, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311012306.

Barley:2011:SND

[BC11b] Kamal Barley and Alhaji Cherif. Stochastic nonlinear dynamics of interpersonal and romantic relationships. Applied Mathematics and Computation, 217(13):6273–6281, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000038.

Bai:2012:ATP

[BC12a] Meng Bai and Shangbin Cui. Analysis of a two-phase cell division model. Applied Mathematics and Computation, 218(9):4849–4859, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012938.

Bao:2012:ESS

[BC12b] Haibo Bao and Jinde Cao. Exponential stability for stochastic BAM networks with discrete and distributed delays. Applied Mathematics and Computation, 218(11):6188–6199, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013658.

Bayad:2012:MPD

[BC12c] Abdelmejid Bayad and Ismail Naci Cangul. The minimal polynomial of 2 cos(π/q) and Dickson polynomials. Applied Mathematics and Computation, 218(13):7014–7022, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015189.

Bota:2012:APS

[BC12d] Constantin Bota and Bogdan Caruntu. ϵ-Approximate polynomial solutions for the multipantograph equation with variable coefficients. Applied Mathematics and Computation, 219(4):1785– 1792, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312008107.

Brzdek:2012:RHU

[BC12e] Janusz Brzdek and Krzysztof Ciepliński. Remarks on the Hyers–Ulam stability of some systems of functional equations. Applied Mathematics and Computation, 219(8):4096–4105, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312010739.

Bai:2013:TPS

[BC13a] Dingyong Bai and Yuming Chen. Three positive solutions for a generalized Laplacian boundary value problem with a parameter. Applied Mathematics and Computation, 219(9):4782–4788, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011307.

Ban:2013:DML

[BC13b] Jung-Chao Ban and Chih-Hung Chang. Diamond in multi-layer cellular neural networks. Applied Mathematics and Computation, 222(??):1–12, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007601.

Belozyorov:2014:QMI

[BC14a] V. Ye. Belozyorov and S. V. Chernyshenko. Quadratic model of inter-population interaction: Investigation of stability areas. Applied Mathematics and Computation, 230(??):43–56, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013854.

Buonomo:2014:SBP

[BC14b] Bruno Buonomo and Marianna Cerasuolo. Stability and bifurcation in plant-pathogens interactions. Applied Mathematics and Computation, 232(??):858–871, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001726.

Blanco-Cocom:2014:CSA

[BCÁV14] Luis Blanco-Cocom and Eric Ávila-Vales. Convergence and stability analysis of the θ-method for delayed diffusion mathematical models. Applied Mathematics and Computation, 231(??):16–25, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000277.

Bica:2011:ANM

[BCC11a] A. M. Bica, M. Curila, and S. Curila. About a numerical method of successive interpolations for two-point boundary value problems with deviating argument. Applied Mathematics and Computation, 217(19):7772–7789, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003080.

Blaga:2011:BTO

[BCC11b] Petru Blaga, Teodora Catinas, and Gheorghe Coman. Bernstein-type operators on a triangle with all curved sides. Applied Mathematics and Computation, 218(7):3072– 3082, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311010538.

Berriochoa:2014:ACH

[BCDI14] E. Berriochoa, A. Cachafeiro, J. Díaz, and J. Illán. Algorithms and convergence for Hermite interpolation based on extended Chebyshev nodal systems. Applied Mathematics and Computation, 234(??):223–236, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002665.

Blanco-Cocom:2012:SDD

[BCEAV12] Luis Blanco-Cocom, Angel G. Estrella, and Eric Avila-Vales. Solving delay differential systems with history functions by the Adomian decomposition method. Applied Mathematics and Computation, 218(10):5994–6011, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311014287.

Brunelli:2013:NPB

[BCF13] Matteo Brunelli, Andrew Critch, and Michele Fedrizzi. A note on the proportionality between some consistency indices in the AHP. Applied Mathematics and Computation, 219(14):7901–7906, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000751.

Baginski:2014:NIR

[BCGK14] Frank E. Baginski, Roberto Croce, Susan Gillmor, and Rolf Krause. Numerical investigations of the role of curvature in strong segregation problems on a given surface. Applied Mathematics and Computation, 227(??):399–411, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011685.

Breuss:2011:ADD

[BCGV11] Michael Breuß, Emiliano Cristiani, Pascal Gwosdek, and Oliver Vogel. An adaptive domaindecomposition technique for parallelization of the fast marching method. Applied Mathematics and Computation, 218(1):32– 44, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311007107.

Benitez:2013:ER

[BCI13] Julio Benítez and Dragana Cvetković-Ilić. On the elements aa^{†} and a^{†}a in a ring. Applied Mathematics and Computation, 222(??):478–489, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007650.

Berzig:2014:GKF

[BCK14] Maher Berzig, Sumit Chandok, and Mohammad Saeed Khan. Generalized Krasnoselskii fixed point theorem involving auxiliary functions in bimetric spaces and application to two-point boundary value problem. Applied Mathematics and Computation, 248(??):323–327, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013307.

Blaszczyk:2011:NSF

[BCKL11] T. Blaszczyk, M. Ciesielski, M. Klimek, and J. Leszczynski. Numerical solution of fractional oscillator equation. Applied Mathematics and Computation, 218(6):2480–2488, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010101.

Berriochoa:2012:AMN

[BCM12] E. Berriochoa, A. Cachafeiro, and E. Martínez. About measures and nodal systems for which the Hermite interpolants uniformly converge to continuous functions on the circle and interval. Applied Mathematics and Computation, 218(9):4813–4824, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012872.

Buonocore:2014:CG

[BCNP14a] A. Buonocore, L. Caputo, A. G. Nobile, and E. Pirozzi. Corrigendum to: “Gauss–Markov processes in the presence of a reflecting boundary and applications in neuronal models” [Applied Mathematics and Computation 232 (2014) 799–809]. Applied Mathematics and Computation, 241(??):11–12, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006614.

See [BCNP14b].

Buonocore:2014:GMP

[BCNP14b] A. Buonocore, L. Caputo, A. G. Nobile, and E. Pirozzi. Gauss–Markov processes in the presence of a reflecting boundary and applications in neuronal models. Applied Mathematics and Computation, 232(??):799–809, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001891.

See corrigendum [BCNP14a].

Beccari:2010:UFI

[BCR10] C. V. Beccari, G. Casciola, and L. Romani. A unified framework for interpolating and approximating univariate subdivision. Applied Mathematics and Computation, 216(4):1169–1180, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000175X.

Bilige:2013:AES

[BCW13] Sudao Bilige, Temuer Chaolu, and Xiaomin Wang. Application of the extended simplest equation method to the coupled Schrödinger–Boussinesq equation. Applied Mathematics and Computation, 224(??):517–523, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300951X.

Biswas:2010:OSP

[BCWM10] Anjan Biswas, Carl Cleary, James E. Watson, Jr., and Daniela Milovic. Optical soliton perturbation with time-dependent coefficients in a log law media. Applied Mathematics and Computation, 217(6):2891–2894, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007770.

Bonanno:2010:FOB

[BD10] Gabriele Bonanno and Beatrice Di Bella. A fourth-order boundary value problem for a Sturm–Liouville type equation. Applied Mathematics and Computation, 217(8):3635–3640, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010490.

Bai:2011:TSL

[BD11] Liang Bai and Binxiang Dai. Three solutions for a p-Laplacian boundary value problem with impulsive effects. Applied Mathematics and Computation, 217(24):9895–9904, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004656.

Baculikova:2012:OTH

[BD12a] B. Baculíková and J. Dzurina. Oscillation theorems for higher order neutral differential equations. Applied Mathematics and Computation, 219(8):3769–3778, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009939.

Balakhnev:2012:ABT

[BD12b] M. Yu. Balakhnev and D. K. Demskoi. Auto– Bäcklund transformations and superposition formulas for solutions of Drinfeld–Sokolov systems. Applied Mathematics and Computation, 219(8):3625– 3637, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009563.

Bansal:2012:MBP

[BD12c] Jagdish Chand Bansal and Kusum Deep. A Modified Binary Particle Swarm Optimization for Knapsack Problems. Applied Mathematics and Computation, 218(22):11042–11061, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004766.

Baculikova:2014:FIT

[BD14a] B. Baculíková and J. Dzurina. On functional inequalities and their applications in the oscillation theory. Applied Mathematics and Computation, 226(??):266–273, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011144.

Barik:2014:TRE

[BD14b] R. N. Barik and G. C. Dash. Thermal radiation effect on an unsteady magnetohydrodynamic flow past inclined porous heated plate in the presence of chemical reaction and viscous dissipation. Applied Mathematics and Computation, 226(??):423–434, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010783.

Belhaj:2014:NCI

[BD14c] Skander Belhaj and Marwa Dridi. A note on computing the inverse of a triangular Toeplitz matrix. Applied Mathematics and Computation, 236(??):512–523, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004391.

Boccuto:2014:RAG

[BD14d] Antonio Boccuto and Xenofon Dimitriou. Rates of approximation for general sampling-type operators in the setting of filter convergence. Applied Mathematics and Computation, 229(??):214–226, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313013271.

Bozhkov:2014:GCG

[BD14e] Y. Bozhkov and S. Dimas. Group classification of a generalization of the Heath equation. Applied Mathematics and Computation, 243(??):121–131, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008005.

Biswas:2010:ARO

[BDAD10] Arijit Biswas, Swagatam Das, Ajith Abraham, and Sambarta Dasgupta. Analysis of the reproduction operator in an artificial bacterial foraging system. Applied Mathematics and Computation, 215(9):3343–3355, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030900928X.

Babusci:2012:DIO

[BDD^{+}12a] D. Babusci, G. Dattoli, G. H. E. Duchamp, K. Górska, and K. A. Penson. Definite integrals and operational methods. Applied Mathematics and Computation, 219(6):3017–3021, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009289.

Baier:2012:HOR

[BDD12b] R. Baier, Q. Din, and T. Donchev. Higher order Runge–Kutta methods for impulsive differential systems. Applied Mathematics and Computation, 218(24):11790–11798, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200522X.

Biswas:2014:CEB

[BDDK14] Subhodip Biswas, Swagatam Das, Shantanab Debchoudhury, and Souvik Kundu. Co-evolving bee colonies by forager migration: a multiswarm based Artificial Bee Colony algorithm for global search space. Applied Mathematics and Computation, 232(??): 216–234, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031301299X.

Bates:2014:PHF

[BDE^{+}14] Daniel J. Bates, Brent Davis, David Eklund, Eric Hanson, and Chris Peterson. Perturbed homotopies for finding all isolated solutions of polynomial systems. Applied Mathematics and Computation, 247(??):301–311, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012107.

Bhalekar:2013:CSM

[BDG13] Sachin Bhalekar and Varsha Daftardar-Gejji. Corrigendum to “Solving multi-term linear and nonlinear diffusion-wave equations of fractional order by Adomian decomposition method” [Applied Mathematics and Computation 202 (2008) 113-120]. Applied Mathematics and Computation, 219(16):8413–8415, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002105.

Benitez:2012:ICA

[BDGIPG12] J. Benítez, X. Delgado-Galván, J. Izquierdo, and R. Pérez-García. Improving consistency in AHP decision-making processes. Applied Mathematics and Computation, 219(5):2432–2441, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008727.

Babusci:2014:RDC

[BDGP14a] D. Babusci, G. Dattoli, K. Górska, and K. A. Penson. Repeated derivatives of composite functions and generalizations of the Leibniz rule. Applied Mathematics and Computation, 241(??):193–199, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006109.

Babusci:2014:SBS

[BDGP14b] D. Babusci, G. Dattoli, K. Górska, and K. A. Penson. The spherical Bessel and Struve functions and operational methods. Applied Mathematics and Computation, 238(??):1– 6, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005086.

Bates:2014:CPA

[BDH^{+}14] Daniel J. Bates, Wolfram Decker, Jonathan D.

Hauenstein, Chris Peterson, Gerhard Pfister, Frank-Olaf Schreyer, Andrew J. Sommese, and Charles W. Wampler. Comparison of probabilistic algorithms for analyzing the components of an affine algebraic variety. Applied Mathematics and Computation, 231(??):619–633, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000034.

Bouchala:2014:SCQ

[BDK^{+}14] Jirí Bouchala, Zdenek Dostál, Tomás Kozubek, Lukás Pospísil, and Petr Vodstrcil. On the solution of convex QPQC problems with elliptic and other separable constraints with strong curvature. Applied Mathematics and Computation, 247(??):848–864, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012697.

Bai:2013:SSO

[BDL13] Liang Bai, Binxiang Dai, and Feng Li. Solvability of second-order Hamiltonian systems with impulses via variational method. Applied Mathematics and Computation, 219(14):7542–7555, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000805.

Bai:2010:P

[BDLY10] Zhong-Zhi Bai, Hua Dai, Xing-Ping Liu, and Jun-Feng Yin. Preface. Applied Mathematics and Computation, 216(6):1681, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010832.

Boyd:2014:UCH

[BDN^{+}14] Z. Boyd, M. Dorff, M. Nowak, M. Romney, and M. Woloszkiewicz. Univalency of convolutions of harmonic mappings. Applied Mathematics and Computation, 234(??):326–332, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400215X.

Babusci:2011:NED

[BDQR11] D. Babusci, G. Dattoli, M. Quattromini, and P. E. Ricci. A note on the extension of the Dirac method. Applied Mathematics and Computation, 218(4):1495–1497, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008010.

Baculikova:2012:OTO

[BDR12a] Blanka Baculíková, Jozef Dzurina, and Yuri V. Rogovchenko. Oscillation of third order trinomial delay differential equations. Applied Mathematics and Computation, 218(13):7023–7033, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015359.

Balyan:2012:LSS

[BDR12b] Lokendra K. Balyan, Pravir K. Dutt, and R. K. S. Rathore. Least squares h-p spectral element method for elliptic eigenvalue problems. Applied Mathematics and Computation, 218(19):9596–9613, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001543.

Buric:2013:AEM

[BE13] Tomislav Burić and Neven Elezović. Approximants of the Euler–Mascheroni constant and harmonic numbers. Applied Mathematics and Computation, 222(??):604–611, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008412.

Becker:2013:RSS

[Bec13] Leigh C. Becker. Resolvents and solutions of singular Volterra integral equations with separable kernels. Applied Mathematics and Computation, 219(24):11265–11277, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005729.

Bertoli:2014:SEI

[BEdAdSK14] Sávio L. Bertoli, Laércio Ender, Juliano de Almeida, and Pollyana M. da Silva Kalvelage. Solutions of elliptic integrals and generalizations by means of Bessel functions. Applied Mathematics and Computation, 243(??):33–43, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400784X.

Biezuner:2012:EEL

[BEGM12] Rodney Josué Biezuner, Grey Ercole, Breno Loureiro Giacchini, and Eder Marinho Martins. Eigenvalues and eigenfunctions of the Laplacian via inverse iteration with shift. Applied Mathematics and Computation, 219(1):360–375, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006248.

Behforooz:2010:IIQ

[Beh10] Hossein Behforooz. Interpolation by integro quintic splines. Applied Mathematics and Computation, 216(2):364–367, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000226.

Bekir:2010:AEF

[Bek10] Ahmet Bekir. Application of the Exp-function method for nonlinear differential-difference equations. Applied Mathematics and Computation, 215(11):4049–4053, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010376.

Belozyorov:2011:NTD

[Bel11a] Vasiliy Ye. Belozyorov.

New types of 3-D systems of quadratic differential equations with chaotic dynamics based on Ricker discrete population model. Applied Mathematics and Computation, 218(8):4546–4566, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012823.

Belozyorov:2011:EHO

[Bel11b] Vasiliy Ye. Belozyorov. On existence of homoclinic orbits for some types of autonomous quadratic systems of differential equations. Applied Mathematics and Computation, 217(9):4582–4595, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011227.

Belozyorov:2012:IOD

[Bel12] Vasiliy Ye. Belozyorov. Implicit one-dimensional discrete maps and their connection with existence problem of chaotic dynamics in 3-D systems of differential equations. Applied Mathematics and Computation, 218(17):8869–8886, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001737.

Belbas:2013:FTI

[Bel13] S. A. Belbas. Floquet theory for integral and integro-differential equations. Applied Mathematics and Computation, 223(??):327–345, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S009630031300845X.

Belen:2014:STT

[Bel14]  Cemal Belen.  Some

Tauberian theorems for statistical (C, 1) summability. Applied Mathematics and Computation, 240(??):252–258, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006511.

Bengochea:2014:AAL

[Ben14a] Gabriel Bengochea. Algebraic approach to the Lane–Emden equation. Applied Mathematics and Computation, 232(??): 424–430, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314001337.

BenSaida:2014:NCI

[Ben14b] Ahmed BenSaïda. Noisy chaos in intraday financial data: Evidence from the American index. Applied Mathematics and Computation, 226(??):258–265, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011211.

Baoguo:2013:OTS

[BEP13] Jia Baoguo, Lynn Erbe, and Allan Peterson. An oscillation theorem for second order superlinear dynamic equations on time scales. Applied Mathematics and Computation, 219 (20):10333–10342, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004037.

Bertotti:2010:MTR

[Ber10] Maria Letizia Bertotti. Modelling taxation and redistribution: a discrete active particle kinetic approach. Applied Mathematics and Computation, 217(2):752–762, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006983.

Bervillier:2012:SDT

[Ber12] C. Bervillier. Status of the differential transformation method. Applied Mathematics and Computation, 218(20):10158– 10170, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003578.

Berlinska:2014:CSD

[Ber14] Joanna Berlińska. Communication scheduling in data gathering networks with limited memory. Applied Mathematics and Computation, 235(??):530–537, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003774.

Besong:2010:NTB

[Bes10a] D. O. Besong. A new transformation of Burger’s equation for an exact solution in a bounded region necessary for certain boundary conditions. Applied Mathematics and Computation, 215(9):3455–3460, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009461.

Bhrawy:2010:JSG

[BES10b] A. H. Bhrawy and S. I. El-Soubhy. Jacobi spectral Galerkin method for the integrated forms of second-order differential equations. Applied Mathematics and Computation, 217(6):2684–2697, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008568.

Bezerra:2013:ECB

[Bez13] Licio Hernanes Bezerra. Efficient computation of Bézier curves from their Bernstein–Fourier representation. Applied Mathematics and Computation, 220(??):235–238, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006590.

Bozhkov:2012:LES

[BF12] Yuri Bozhkov and Igor Leite Freire. On the Lane– Emden system in dimension one. Applied Mathematics and Computation, 218(21):10762– 10766, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004328.

Bu:2012:RDI

[BFB12] Changjiang Bu, Chengcheng Feng, and Shuyan Bai. Representations for the Drazin inverses of the sum of two matrices and some block matrices. Applied Mathematics and Computation, 218 (20):10226–10237, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003657.

Barnard:2012:ORT

[BFH12] Richard Barnard, Martin Frank, and Michael Herty. Optimal radiotherapy treatment planning using minimum entropy models. Applied Mathematics and Computation, 219(5):2668–2679, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008922.

Bohner:2013:CDI

[BFM13] Martin Bohner, Márcia Federson, and Jaqueline Godoy Mesquita. Continuous dependence for impulsive functional dynamic equations involving variable time scales. Applied Mathematics and Computation, 221(??):383–393, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300605X.

Bnouhachem:2010:NET

[BFXZ10] Abdellah Bnouhachem, Xiao-Ling Fu, M. H. Xu, and Sheng Zhaohan. New extragradient-type methods for solving variational inequalities. Applied Mathematics and Computation, 216(8):2430–2440, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003425.

Bruzon:2012:EGE

[BGGH12] M. S. Bruzón, M. L. Gandarias, G. A. González, and R. Hansen. The K(mn) equation with generalized evolution term studied by symmetry reductions and qualitative analysis. Applied Mathematics and Computation, 218(20):10094– 10105, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003426.

Bykov:2013:AGQ

[BGP^{+}13] V. Bykov, J. F. Griffiths, R. Piazzesi, S. S. Sazhin, and E. M. Sazhina. The application of the Global Quasi-Linearisation technique to the analysis of the cyclohexane/air mixture autoignition. Applied Mathematics and Computation, 219(14):7338–7347, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013185.

Bayon:2011:ESP

[BGRS11] L. Bayón, J. M. Grau, M. M. Ruiz, and P. M. Suárez. The explicit solution of the profit maximization problem with box-constrained inputs. Applied Mathematics and Computation, 217(21):8705–8715, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100484X.

Balbuena:2010:EFT

[BGV10] C. Balbuena and P. García-Vázquez. Edge fault tolerance analysis of super krestricted connected networks. Applied Mathematics and Computation, 216(2):506–513, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000731.

Boulaaras:2011:ABF

[BH11] Salah Boulaaras and Mohamed Haiour. L^{∞}- asymptotic behavior for a finite element approximation in parabolic quasivariational inequalities related to impulse control problem. Applied Mathematics and Computation, 217(14):6443–6450, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000415.

Bhargava:2012:MMC

[BH12a] R. R. Bhargava and S. Hasan. Mathematical model for crack arrest of an infinite plate weakened by a finite and two semiinfinite cracks. Applied Mathematics and Computation, 218(11):6576–6584, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015414.

Bi:2012:HBS

[BH12b] Ping Bi and Zhixing Hu.

Hopf bifurcation and stability for a neural network model with mixed delays. Applied Mathematics and Computation, 218(12):6748–6761, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015165.

Benlic:2013:BLS

[BH13] Una Benlic and Jin-Kao Hao. Breakout local search for the quadratic assignment problem. Applied Mathematics and Computation, 219(9):4800–4815, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011368.

Bahri:2010:WFT

[BHAV10] Mawardi Bahri, Eckhard S. M. Hitzer, Ryuichi Ashino, and Rémi Vaillancourt. Windowed Fourier transform of twodimensional quaternionic signals. Applied Mathematics and Computation, 216(8):2366–2379, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000336X.

Binesh:2010:EPA

[BHG10] S. M. Binesh, N. Hataf, and A. Ghahramani. Elasto-plastic analysis of reinforced soils using mesh-free method. Applied Mathematics and Computation, 215(12):4406–4421, February 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000135.

Badyrka:2013:ADO

[BHJ13] Jeffrey M. Badyrka, Roy J. Hartfield, and Rhonald M. Jenkins. Aerospace design optimization using a compound repulsive particle swarm. Applied Mathematics and Computation, 219(15):8311–8331, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011435.

Bhowmik:2010:SNS

[Bho10] Samir Kumar Bhowmik.

Stable numerical schemes for a partly convolutional partial integro-differential equation. Applied Mathematics and Computation, 217(8):4217–4226, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010799.

Bleylevens:2013:POJ

[BHP13] Ivo W. M. Bleylevens, Michiel E. Hochstenbach, and Ralf L. M. Peeters. Polynomial optimization and a Jacobi– Davidson type method for commuting matrices. Applied Mathematics and Computation, 224(??):564–580, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008643.

Bhrawy:2013:JGL

[Bhr13] A. H. Bhrawy. A Jacobi– Gauss–Lobatto collocation method for solving generalized Fitzhugh– Nagumo equation with time-dependent coefficients. Applied Mathematics and Computation, 222(??):255–264, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008060.

Bhrawy:2014:EJP

[Bhr14] A. H. Bhrawy. An efficient Jacobi pseudospectral approximation for nonlinear complex generalized Zakharov system. Applied Mathematics and Computation, 247(??):30–46, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011722.

Biswas:2010:SSCa

[BI10] Anjan Biswas and M. S. Ismail. 1-soliton solution of the coupled KdV equation and Gear–Grimshaw model. Applied Mathematics and Computation, 216(12):3662–3670, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005369.

Bolat:2014:SQI

[BI14] Cennet Bolat and Ahmet Ipek. On the solutions of the quaternion interval systems [x] = [A][x] + [b]. Applied Mathematics and Computation, 244(??):375–381, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009503.

Bieniasz:2011:AHM

[Bie11]  L. K. Bieniasz.  An

adaptive Huber method for non-linear systems of weakly singular second kind Volterra integral equations. Applied Mathematics and Computation, 217(12):5622–5631, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012427.

Bilgici:2014:TGL

[Bil14] Goksal Bilgici. Two generalizations of Lucas sequence. Applied Mathematics and Computation, 245(??):526–538, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010807.

Birou:2013:BTO

[Bir13] Marius Mihai Birou. Bernstein type operators with a better approximation for some functions. Applied Mathematics and Computation, 219(17):9493–9499, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003135.

Brugnano:2012:LCR

[BIT12a] Luigi Brugnano, Felice Iavernaro, and Donato Trigiante. The lack of continuity and the role of infinite and infinitesimal in numerical methods for ODEs: the case of symplecticity. Applied Mathematics and Computation, 218(16):8056–8063, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003766.

Brugnano:2012:SFD

[BIT12b] Luigi Brugnano, Felice Iavernaro, and Donato Trigiante. A simple framework for the derivation and analysis of effective one-step methods for ODEs. Applied Mathematics and Computation, 218(17):8475–8485, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001269.

Bras:2011:NMC

[BJ11] M. Braś and Z. Jackiewicz.

Nordsieck methods with computationally verified algebraic stability. Applied Mathematics and Computation, 217(21):8598–8610, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004644.

Bartoszewski:2012:ENM

[BJ12] Z. Bartoszewski and Z. Jackiewicz. Explicit Nordsieck methods with extended stability regions. Applied Mathematics and Computation, 218(10):6056–6066, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014342.

Barzoki:2011:EIP

[BJB11] Morteza Rasti Barzoki, Morteza Jahanbazi, and Mehdi Bijari. Effects of imperfect products on lot sizing with work in process inventory. Applied Mathematics and Computation, 217(21):8328–8336, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003778.

Jian:2014:SCN

[bJLyZlL14] Jin bao Jian, Jie Li, Hai yan Zheng, and Jian ling Li. A superlinearly convergent norm-relaxed method of quasi-strongly subfeasible direction for inequality constrained minimax problems. Applied Mathematics and Computation, 226(??):673–690, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313011399.

Bejancu:2014:FDE

[BJS14] Aurelian Bejancu, Michael J.

Johnson, and Hamid Said. Finite difference end conditions and semi-cardinal interpolation with quintic B-splines. Applied Mathematics and Computation, 237(??):318–329, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004883.

Bhunia:2014:TWI

[BJSS14] A. K. Bhunia, Chandra K.

Jaggi, Anuj Sharma, and Ritu Sharma. A two-warehouse inventory model for deteriorating items under permissible delay in payment with partial backlogging. Applied Mathematics and Computation, 232(??):1125–1137, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400160X.

Bairamov:2010:EDD

[BK10a] Elgiz Bairamov and Turhan Koprubasi. Eigenparameter dependent discrete Dirac equations with spectral singularities. Applied Mathematics and Computation, 215(12):4216–4220, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030901090X.

Biswas:2010:SSCd

[BK10b] Anjan Biswas and A. H. Kara. 1-soliton solution and conservation laws for nonlinear wave equation in semiconductors. Applied Mathematics and Computation, 217(8):4289–4292, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010143.

Biswas:2010:SSCc

[BK10c] Anjan Biswas and A. H. Kara. 1-soliton solution and conservation laws for the Jaulent– Miodekequation with power law nonlinearity. Applied Mathematics and Computation, 217(2):944–948, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031000706X.

Biswas:2010:SSCb

[BK10d] Anjan Biswas and A. H.

Kara. 1-soliton solution and conservation laws of the generalizedDullin– Gottwald–Holm equation. Applied Mathematics and Computation, 217(2):929– 932, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310006545.

Benterki:2011:FSF

[BK11a] Djamel Benterki and Abdelkrim Keraghel. Finding a strict feasible solution of a linear semidefinite program. Applied Mathematics and Computation, 217(13):6437–6440, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013007.

Braverman:2011:ODD

[BK11b] Elena Braverman and Basak Karpuz. On oscillation of differential and difference equations with non-monotone delays. Applied Mathematics and Computation, 218(7):3880–3887, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012070.

Braverman:2012:UES

[BK12] Elena Braverman and Basak Karpuz. Uniform exponential stability of first-order dynamic equations with several delays. Applied Mathematics and Computation, 218(21): 10468–10485, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003931.

Bhatt:2014:HOE

[BK14] Harish P. Bhatt and Abdul Q. M. Khaliq. Higher order exponential time differencing scheme for system of coupled nonlinear Schrödinger equations. Applied Mathematics and Computation, 228(??):271–291, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012587.

Baek:2012:COP

[BKA12] Seunghyeon Baek, Wonlyul Ko, and Inkyung Ahn. Coexistence of a one-prey two-predators model with ratio-dependent functional responses. Applied Mathematics and Computation, 219(4):1897–1908, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008259.

Bektesevic:2014:GDQ

[BKP14] J. Bektesević, M. R. S.

Kulenović, and E. Pilav. Global dynamics of quadratic second order difference equation in the first quadrant. Applied Mathematics and Computation, 227(??):50–65, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011053.

Bellout:2011:DUS

[BKW11] Hamid Bellout, Qingkai Kong, and Min Wang. Dependence of the unique solution of a periodic boundary value problem on the parameter. Applied Mathematics and Computation, 217(19):7838–7844, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003341.

Buong:2011:HMH

[BL11] Nguyen Buong and Nguyen Duc Lang. Hybrid Mann– Halpern iteration methods for nonexpansive mappings and semigroups. Applied Mathematics and Computation, 218(6):2459– 2466, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010071.

Bawa:2011:UHS

[BLK11] Rajesh K. Bawa, A. K. Lal, and Vinod Kumar. An ϵ-uniform hybrid scheme for singularly perturbed delay differential equations. Applied Mathematics and Computation, 217(21):8216–8222, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003122.

Benes:2013:NSF

[BLK^{+}13] Ludek Benes, Petr Louda, Karel Kozel, Radka Keslerová, and Jaroslav Stigler. Numerical simulations of flow through channels with T - junction. Applied Mathematics and Computation, 219(13):7225–7235, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006448.

Bozzini:2010:PSA

[BLR10] Mira Bozzini, Licia Lenarduzzi, and Milvia Rossini. Polyharmonic splines: an approximation method for noisy scattered data of extra-large size. Applied Mathematics and Computation, 216(1):317–331, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000780.

Bischi:2013:PPF

[BLR13a] Gian-Italo Bischi, Fabio Lamantia, and Davide Radi. A prey-predator fishery model with endogenous switching of harvesting strategy. Applied Mathematics and Computation, 219(20):10123– 10142, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001999.

Blanchard:2013:CAC

[BLR13b] E. A. Blanchard, R. Loxton, and V. Rehbock. A computational algorithm for a class of non-smooth optimal control problems arising in aquaculture operations. Applied Mathematics and Computation, 219(16):8738–8746, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002087.

Bai:2014:ASR

[BLZZ14] Ling Bai, Jingshi Li, Kai Zhang, and Wenju Zhao. Analysis of a stochastic ratio-dependent predatorprey model driven by Lévy noise. Applied Mathematics and Computation, 233(??):480–493, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000265.

Basar:2011:CCO

[BM11] Feyzı Başar and Eberhard Malkowsky. The characterization of compact operators on spaces of strongly summable and bounded sequences. Applied Mathematics and Computation, 217(12): 5199–5207, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310012075.

Bardaro:2012:VFL

[BM12a] Carlo Bardaro and Ilaria Mantellini. On Voronovskaja formula for linear combinations of Mellin–Gauss– Weierstrass operators. Applied Mathematics and Computation, 218(20): 10171–10179, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312003608.

Barucci:2012:COI

[BM12b] Emilio Barucci and Daniele Marazzina. Corrigendum to ‘Optimal investment, stochastic labor income and retirement’ Appl. Math. Comput. 218 (2012) 5588–5604. Applied Mathematics and Computation, 218(11):6627, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015141.

See [BM12c].

Barucci:2012:OIS

[BM12c] Emilio Barucci and Daniele Marazzina. Optimal investment, stochastic labor income and retirement. Applied Mathematics and Computation, 218(9):5588–5604, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013956.

See corrigendum [BM12b].

Bernard:2012:BSE

[BM12d] C. Bernard and S. Martin. Building strategies to ensure language coexistence in presence of bilingualism. Applied Mathematics and Computation, 218(17):8825–8841, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001683.

Belen:2013:GWS

[BM13a] C. Belen and S. A. Mohiuddine. Generalized weighted statistical convergence and application. Applied Mathematics and Computation, 219(18):9821–9826, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003767.

Belhachmi:2013:SSA

[BM13b] Z. Belhachmi and H. Meftahi. Shape sensitivity analysis for an interface problem via minimax differentiability. Applied Mathematics and Computation, 219 (12):6828–6842, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000623.

Baek:2014:MMQ

[BM14a] Jung Woo Baek and Seung Ki Moon. The M/M/1 queue with a production-inventory system and lost sales. Applied Mathematics and Computation, 233(??):534–544, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314002720.

Balakumar:2014:STW

[BM14b] V. Balakumar and K. Murugesan. Single-Term Walsh Series method for systems of linear Volterra integral equations of the second kind. Applied Mathematics and Computation, 228(??):371–376, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012216.

Bertotti:2014:MMM

[BM14c] Maria Letizia Bertotti and Giovanni Modanese. Micro to macro models for income distribution in the absence and in the presence of tax evasion. Applied Mathematics and Computation, 244(??):836–846, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010121.

Baleanu:2011:SCS

[BMA11] Dumitru Baleanu, Octavian G. Mustafa, and Ravi P. Agarwal. On L^{p}- solutions for a class of sequential fractional differential equations. Applied Mathematics and Computation, 218(5):2074–2081, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031100957X.

Basak:2013:DIW

[BMD13] Aniruddha Basak, Dipankar Maity, and Swagatam Das. A differential invasive weed optimization algorithm for improved global numerical optimization. Applied Mathematics and Computation, 219(12):6645–6668, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013331.

Bhunu:2012:MHA

[BMH12] C. P. Bhunu, S. Mushayabasa, and J. M. Hyman. Modelling HIV/AIDS and monkeypox co-infection. Applied Mathematics and Computation, 218(18): 9504–9518, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312002743.

Back:2014:EST

[BMHM14] Julian M. Back, Scott W. McCue, Mike H.-N. Hsieh, and Timothy J. Moroney. The effect of surface tension and kinetic undercooling on a radiallysymmetric melting problem. Applied Mathematics and Computation, 229(??):41–52, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012770.

Botta:2011:FFA

[BMI^{+}11] N. Botta, A. Mandel, C. Ionescu, M. Hofmann, D. Lincke, S. Schupp, and C. Jaeger. A functional framework for agent-based models of exchange. Applied Mathematics and Computation, 218(8):4025–4040, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010915.

Berselli:2013:ESI

[BMMS13] Luigi C. Berselli, Piero Miloro, Arianna Menciassi, and Edoardo Sinibaldi. Exact solution to the inverse Womersley problem for pulsatile flows in cylindrical vessels, with application to magnetic particle targeting. Applied Mathematics and Computation, 219(10):5717–5729, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012386.

Balasubramaniam:2010:ESR

[BMR10] P. Balasubramaniam, A. Manivannan, and R. Rakkiyappan. Exponential stability results for uncertain neutral systems with interval time-varying delays and Markovian jumping parameters. Applied Mathematics and Computation, 216(11):3396–3407, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005175.

Bougoffa:2013:SCI

[BMR13] Lazhar Bougoffa, Abdelaziz Mennouni, and Randolph C. Rach. Solving Cauchy integral equations of the first kind by the Adomian decomposition method. Applied Mathematics and Computation, 219(9):4423–4433, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010624.

Botmart:2010:RES

[BN10] T. Botmart and P. Niamsup. Robust exponential stability and stabilizability of linear parameter dependent systems with delays. Applied Mathematics and Computation, 217(6):2551–2566, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008271.

Bozic:2010:TSV

[BNBV10] Vladan Bozic, Truong Nguyen-Ba, and Rémi Vaillancourt. A threestage, VSVO, Hermite– Birkhoff–Taylor, ODE solver. Applied Mathematics and Computation, 216(2):598–610, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300310001062.

Botmart:2011:DDE

[BNP11] T. Botmart, P. Niamsup, and V. N. Phat. Delaydependent exponential stabilization for uncertain linear systems with interval non-differentiable time-varying delays. Applied Mathematics and Computation, 217(21): 8236–8247, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311003201.

Bostanci:2011:NSM

[BÖ11] Hakan Bostanci and Metin Öztürk. A new subclass of the meromorphic harmonic γ-starlike functions. Applied Mathematics and Computation, 218(3):683–688, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005297.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Bodnar:2013:BRA

[BO13] Taras Bodnar and Yarema Okhrin. Boundaries of the risk aversion coefficient: Should we invest in the global minimum variance portfolio? Applied Mathematics and Computation, 219(10):5440–5448, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012167.

Boglaev:2011:MIQ

[Bog11] Igor Boglaev. Monotone iterates with quadratic convergence rate for solving weighted average approximations to semilinear parabolic problems. Applied Mathematics and Computation, 217(13):6390–6400, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000300.

Boglaev:2012:IMM

[Bog12] Igor Boglaev. An inexact monotone method for solving semilinear parabolic problems. Applied Mathematics and Computation, 219(6):3253–3263, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009794.

Bor:2011:GAC

[Bor11] Hüseyin Bor. On generalized absolute Cesàro summability. Applied Mathematics and Computation, 217(22):8923–8926, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004462.

Borcut:2012:TCTb

[Bor12] Marin Borcut. Tripled coincidence theorems for contractive type mappings in partially ordered metric spaces. Applied Mathematics and Computation, 218(14):7339–7346, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000665.

Berrocal:2014:EVV

[BOR14a] Lilia Berrocal, Aurora Olivieri, and Juan Rada. Extremal values of vertexdegree-based topological indices over hexagonal systems with fixed number of vertices. Applied Mathematics and Computation, 243(??):176–183, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008121.

Bor:2014:NRQ

[Bor14b] Hüseyin Bor. A new result on the quasi power increasing sequences. Applied Mathematics and Computation, 248(??): 426–429, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314013022.

Botelho:2012:ESG

[Bot12]  Fabio Botelho.  Ex-

istence of solution for the Ginzburg–Landau system, a related optimal control problem and its computation by the generalized method of lines. Applied Mathematics and Computation, 218(24):11976–11989, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005887.

Bougoffa:2010:NEG

[Bou10] Lazhar Bougoffa. New exact general solutions of Abel equation of the second kind. Applied Mathematics and Computation, 216(2):689–691, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001384.

Boyd:2010:LBE

[Boy10]  John P. Boyd.  The

Legendre–Burgers equation: When artificial dissipation fails. Applied Mathematics and Computation, 217(5):1949–1964, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007368.

Boyd:2011:OPP

[Boy11]  John P. Boyd.  One-

point pseudospectral collocation for the onedimensional Bratu equation. Applied Mathematics and Computation, 217(12):5553–5565, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310012312.

Boyd:2012:NPC

[Boy12] John P. Boyd. Numerical, perturbative and Chebyshev inversion of the incomplete elliptic integral of the second kind. Applied Mathematics and Computation, 218(13):7005–7013, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014809.

Boyd:2013:FEA

[Boy13] John P. Boyd. A Fourier error analysis for radial basis functions on an infinite uniform grid. Part 2: Spectral-plus is special. Applied Mathematics and Computation, 225(??):695–707, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010527.

Boyd:2014:FTQ

[Boy14] John P. Boyd. The Fourier Transform of the quartic Gaussian exp(−Ax^{4}): Hypergeometric functions, power series, steepest descent asymptotics and hyperasymptotics and extensions to exp(−Ax^{2n}). Applied Mathematics and Computation, 241(??):75– 87, August 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314006572.

Yang:2014:ICH

[bOYqGL^{+}14] Hai bin Ou Yang, Li qun Gao, Steven Li, Xiangyong Kong, and De xuan Zou. On the iterative convergence of harmony search algorithm and a proposed modification. Applied Mathematics and Computation, 247 (??):1064–1095, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401296X.

Bozkurt:2013:SNM

[Boz13]  Durmus Bozkurt.  On

the spectral norms of the matrices connected to integer number sequences. Applied Mathematics and Computation, 219(12):6576–6579, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000490.

Bae:2010:FEH

[BP10a] Jae-Hyeong Bae and Won-Gil Park. A functional equation having monomials as solutions. Applied Mathematics and Computation, 216(1):87–94, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000196.

Bajak:2010:CAS

[BP10b] Szabolcs Baják and Zsolt Páles. Computer aided solution of the invariance equation for two-variable Stolarsky means. Applied Mathematics and Computation, 216(11):3219–3227, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004741.

Ballestra:2011:BEM

[BP11a] Luca Vincenzo Ballestra and Graziella Pacelli. A boundary element method to price time-dependent double barrier options. Applied Mathematics and Computation, 218(8):4192– 4210, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311012227.

Bashier:2011:NFO

[BP11b] E. B. M. Bashier and K. C. Patidar. A novel fitted operator finite difference method for a singularly perturbed delay parabolic partial differential equation. Applied Mathematics and Computation, 217(9):4728–4739, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011409.

Bajak:2013:SIE

[BP13] Szabolcs Baják and Zsolt Páles. Solving invariance equations involving homogeneous means with the help of computer. Applied Mathematics and Computation, 219(11):6297–6315, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201199X.

Babalos:2012:MFP

[BPDZ12] Vassilios Babalos, Nikolaos Philippas, Michael Doumpos, and Constantin Zopounidis. Mutual funds performance appraisal using stochastic multicriteria acceptability analysis. Applied Mathematics and Computation, 218(9):5693–5703, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014093.

Biswas:2010:TES

[BPM10] Anjan Biswas, Marko D. Petković, and Daniela Milović. Topological exact soliton solution of the power law KdV equation. Applied Mathematics and Computation, 217(4):1780–1784, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010364.

Bakula:2012:CJI

[BPP12a] M. Klaricić Bakula, J. Pecarić, and J. Perić. On the converse Jensen inequality. Applied Mathematics and Computation, 218 (11):6566–6575, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015116.

Balasubramaniam:2012:HBA

[BPP12b] P. Balasubramaniam, M. Prakash, and Ju H. Park. Hopf bifurcation analysis of delayed model of thymic infection with HIV-1. Applied Mathematics and Computation, 218(23): 11505–11517, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312005425.

Banasiak:2014:PLA

[BPPS14] J. Banasiak, N. Parumasur, W. D. Poka, and S. Shindin. Pseudospectral Laguerre approximation of transport-fragmentation equations. Applied Mathematics and Computation, 239(??):107–125, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005992.

Branquinho:2013:SEL

[BPR13] A. Branquinho, A. Paiva, and M. N. Rebocho. Sylvester equations for Laguerre–Hahn orthogonal polynomials on the real line. Applied Mathematics and Computation, 219(17):9118–9131, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003445.

Baldi:2012:TDK

[BPT12] Mauro Maria Baldi, Guido Perboli, and Roberto Tadei. The threedimensional knapsack problem with balancing constraints. Applied Mathematics and Computation, 218(19):9802–9818, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002913.

Brzdek:2014:RSN

[BPX14] Janusz Brzdek, Dorian Popa, and Bing Xu. Remarks on stability and non-stability of the linear functional equation of the first order. Applied Mathematics and Computation, 238(??):141–148, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005311.

Benitez:2010:MSN

[BR10a] Julio Benítez and Vladimir Rakočević. Matrices A such that AA^{†} − A^{†}A are nonsingular. Applied Mathematics and Computation, 217(7):3493–3503, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310009690.

Biswas:2010:TSS

[BR10b] Anjan Biswas and Arjuna Ranasinghe. Topological 1-soliton solution of Kadomtsev–Petviashvili equation with power law nonlinearity. Applied Mathematics and Computation, 217(4):1771–1773, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309008674.

Buonomo:2010:LSS

[BR10c] Bruno Buonomo and Salvatore Rionero. On the Lyapunov stability for SIRS epidemic models with general nonlinear incidence rate. Applied Mathematics and Computation, 217(8):4010–4016, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310010374.

Benitez:2012:CAL

[BR12] Julio Benítez and Vladimir Rakocević. Canonical angles and limits of sequences of EP and co-EP matrices. Applied Mathematics and Computation, 218(17):8503–8512, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001300.

Bougoffa:2013:SNIb

[BR13] Lazhar Bougoffa and Randolph C. Rach. Solving nonlocal initial-boundary value problems for linear and nonlinear parabolic and hyperbolic partial differential equations by the Adomian decomposition method. Applied Mathematics and Computation, 225(??):50–61, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009831.

Besse:2014:NMC

[BR14a] I. M. Besse and N. H.

Rhee. A numerical method for calculating minimum distance to Near Earth Objects. Applied Mathematics and Computation, 237(??):274–281, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400486X.

Bisci:2014:ESL

[BR14b] Giovanni Molica Bisci and Dusan Repovs. Existence of solutions for p-Laplacian discrete equations. Applied Mathematics and Computation, 242(??):454–461, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008182.

Branquinho:2014:DLH

[BR14c] A. Branquinho and M. N.

Rebocho. Deformed Laguerre-Hahn orthogonal polynomials on the real line. Applied Mathematics and Computation, 233(??):359–368, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002586.

Braha:2012:ALS

[Bra12] N. L. Braha. On asymptotically Δ^{m} lacunary statistical equivalent sequences. Applied Mathematics and Computation, 219(1):280–288, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005991.

Bradji:2013:ASO

[Bra13]  Abdallah Bradji.  An

analysis of a second-order time accurate scheme for a finite volume method for parabolic equations on general nonconforming multidimensional spatial meshes. Applied Mathematics and Computation, 219(11):6354–6371, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312013264.

Barrio:2011:BLT

[BRAB11] R. Barrio, M. Rodríguez, A. Abad, and F. Blesa. Breaking the limits: The Taylor series method. Applied Mathematics and Computation, 217(20): 7940–7954, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311003031.

Blaya:2012:BVP

[BRAS12] Ricardo Abreu Blaya, Juan Bory Reyes, Alí Guzmán Adán, and Baruch Schneider. Boundary value problems for the Cimmino system via quaternionic analysis. Applied Mathematics and Computation, 219(8):3872–3881, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010090.

Blaya:2014:ARB

[BRGP14] Ricardo Abreu Blaya, Juan Bory Reyes, Tania Moreno García, and Yudier Peña Pérez. Analytic Riemann boundary value problem on h-summable closed curves. Applied Mathematics and Computation, 227(??):593–600, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012228.

Bougoffa:2011:AMS

[BRM11a] Lazhar Bougoffa, Randolph C. Rach, and Abdelaziz Mennouni. An approximate method for solving a class of weaklysingular Volterra integrodifferential equations. Applied Mathematics and Computation, 217(22): 8907–8913, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311003250.

Bougoffa:2011:CTS

[BRM11b] Lazhar Bougoffa, Randolph C. Rach, and Abdelaziz Mennouni. A convenient technique for solving linear and nonlinear Abel integral equations by the Adomian decomposition method. Applied Mathematics and Computation, 218(5):1785–1793, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009076.

Balasubramaniam:2013:FNM

[BRP13] P. Balasubramaniam, V. M.

Revathi, and Ju H. Park. L_{2}–L_{∞} filtering for neutral Markovian switching system with modedependent time-varying delays and partially unknown transition probabilities. Applied Mathematics and Computation, 219(17):9524–9542, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002907.

Babayar-Razlighi:2013:NSS

[BRS13] B. Babayar-Razlighi and B. Soltanalizadeh. Numerical solution for system of singular nonlinear Volterra integrodifferential equations by Newton-Product method. Applied Mathematics and Computation, 219(15): 8375–8383, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031300043X.

Blanco-Rodriguez:2012:NAE

[BRVR12] Francisco J. Blanco-Rodríguez, Francisco R. Villatoro, and Juan I. Ramos. Numerical analysis of the effect of small geometrical imperfections on photonic crystal wires. Applied Mathematics and Computation, 218(10):5989–5993, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014275.

Bougoffa:2013:SNIa

[BRW13] Lazhar Bougoffa, Randolph C. Rach, and Abdul-Majid Wazwaz. Solving nonlocal initialboundary value problems for the Lotka–von Foerster model. Applied Mathematics and Computation, 225(??):7–15, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009739.

Brychkov:2011:SDI

[Bry11] Yu. A. Brychkov. On some definite integrals and infinite sums containing products of special functions. Applied Mathematics and Computation, 218(3):689–692, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100378X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Baricz:2010:BGM

[BS10a] Árpád Baricz and Yin Sun. Bounds for the generalized Marcum Qfunction. Applied Mathematics and Computation, 217(5):2238–2250, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007691.

Bohner:2010:WTT

[BS10b] Martin Bohner and Shurong Sun. Weyl–Titchmarsh theory for symplectic difference systems. Applied Mathematics and Computation, 216(10):2855–2864, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004029.

Boonchari:2010:CCF

[BS10c] Daruni Boonchari and Satit Saejung. Construction of common fixed points of a countable family of λ-demicontractive mappings in arbitrary Banach spaces. Applied Mathematics and Computation, 216(1):173–178, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000408.

Berg:2011:DEP

[BS11a] Lothar Berg and Stevo Stević. On difference equations with powers as solutions and their connection with invariant curves. Applied Mathematics and Computation, 217(17):7191–7196, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001639.

Berg:2011:SSD

[BS11b] Lothar Berg and Stevo Stević. On some systems of difference equations. Applied Mathematics and Computation, 218(5):1713–1718, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008836.

Bezerra:2011:CBC

[BS11c] Licio Hernanes Bezerra and Leonardo Koller Sacht. On computing Bézier curves by Pascal matrix methods. Applied Mathematics and Computation, 217(24):10118–10128, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100662X.

Beauregard:2012:SAC

[BS12a] Matthew A. Beauregard and Qin Sheng. A semiadaptive compact splitting method for the numerical solution of 2dimensional quenching problems. Applied Mathematics and Computation, 218(22):11240–11254, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005012.

Bera:2012:HPM

[BS12b] P. K. Bera and Tapas Sil. Homotopy perturbation method in quantum mechanical problems. Applied Mathematics and Computation, 219(6):3272–3278, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009915.

Boyd:2012:CRR

[BS12c] John P. Boyd and Burhan A. Sadiq. Computing the real roots of a Fourier seriesplus-linear-polynomial: a Chebyshev companion matrix approach. Applied Mathematics and Computation, 219(3):819–826, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312006443.

Batir:2013:SSI

[BS13] Necdet Batir and Anthony Sofo. On some series involving reciprocals of binomial coefficients. Applied Mathematics and Computation, 220(??):331–338, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006000.

Baca:2014:TEI

[BS14a] Martin Baca and Muhammad Kamran Siddiqui. Total edge irregularity strength of generalized prism. Applied Mathematics and Computation, 235(??):168–173, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003543.

Baladram:2014:NSI

[BS14b] Mohammad Samy Baladram and Djoko Suprijanto. New spherical 2s + 1-designs from Kuperberg’s set: an experimental result. Applied Mathematics and Computation, 249(??):45–52, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014118.

Bravi:2014:IDM

[BS14c] Luca Bravi and Marco Sciandrone. An incremental decomposition method for unconstrained optimization. Applied Mathematics and Computation, 235(??):80–86, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003361.

Beheshti:2013:MMO

[BSH13] Zahra Beheshti, Siti Mariyam Hj. Shamsuddin, and Shafaatunnur Hasan. MPSO: Median-oriented Particle Swarm Optimization. Applied Mathematics and Computation, 219(11):5817–5836, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012805.

Bojarski:2013:PRW

[BSKL13] Jacek Bojarski, Robert Smolenski, Adam Kempski, and Piotr Lezynski. Pearson’s random walk approach to evaluating interference generated by a group of converters. Applied Mathematics and Computation, 219(12):6437–6444, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S009630031300057X.

Braha:2014:KTA

[BSM14] Naim L. Braha, H. M. Srivastava, and S. A. Mohiuddine. A Korovkin’s type approximation theorem for periodic functions via the statistical summability of the generalized de la Vallée poussin mean. Applied Mathematics and Computation, 228(??):162–169, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012642.

Britz:2011:DST

[BSØ11] Dieter Britz, Jörg Strutwolf, and Ole Østerby. Digital simulation of thermal reactions. Applied Mathematics and Computation, 218(4):1280–1290, October 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311008228.

Bodnar:2011:STV

[BSP11] T. Bodnár, A. Sequeira, and M. Prosi. On the shear-thinning and viscoelastic effects of blood flow under various flow rates. Applied Mathematics and Computation, 217(11):5055–5067, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000799X.

Bhunia:2010:RSO

[BSR10] A. K. Bhunia, L. Sahoo, and D. Roy. Reliability stochastic optimization for a series system with interval component reliability via genetic algorithm. Applied Mathematics and Computation, 216(3):929–939, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000130X.

Blaszak:2012:CSN

[BSS12] Maciej Blaszak, Blazej M. Szablikowski, and Burcu Silindir. Construction and separability of nonlinear soliton integrable couplings. Applied Mathematics and Computation, 219(4):1866–1873, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008211.

Bayad:2014:SAT

[BSS14] Abdelmejid Bayad, Yilmaz Simsek, and H. M. Srivastava. Some array type polynomials associated with special numbers and polynomials. Applied Mathematics and Computation, 244(??):149–157, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009308.

Benes:2013:ACT

[BSZ13] Michal Benes, Radek Stefan, and Jan Zeman. Analysis of coupled transport phenomena in concrete at elevated temperatures. Applied Mathematics and Computation, 219(13):7262–7274, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002682.

Baksalary:2010:SS

[BT10a] Oskar Maria Baksalary and Götz Trenkler. Subsets and subspaces. Applied Mathematics and Computation, 216(2):484– 491, March 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310000639.

Biswas:2010:SSK

[BT10b] Anjan Biswas and Houria Triki. 1-soliton solution of the Klein–Gordon– Schrodinger’s equation with power law nonlinearity. Applied Mathematics and Computation, 217(8):3869–3874, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010064.

Briozzo:2010:SPN

[BT10c] Adriana C. Briozzo and Domingo A. Tarzia. A Stefan problem for a nonclassical heat equation with a convective condition. Applied Mathematics and Computation, 217(8):4051–4060, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010453.

Baksalary:2011:P

[BT11a] Oskar Maria Baksalary and Götz Trenkler. On the projectors F F † and F † F. Applied Mathematics and Computation, 217(24):10213–10223, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006734.

Biswas:2011:SSE

[BT11b] Anjan Biswas and Houria Triki. 1-soliton solution of the D(m, n) equation with generalized evolution. Applied Mathematics and Computation, 217(21):8482–8488, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004176.

Bozkurt:2012:DIC

[BT12] Durmus Bozkurt and Tin-Yau Tam. Determinants and inverses of circulant matrices with Jacobsthal and Jacobsthal– Lucas Numbers. Applied Mathematics and Computation, 219(2):544–551, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006388.

Baksalary:2013:PVS

[BT13] Oskar Maria Baksalary and Götz Trenkler. On a pair of vector spaces. Applied Mathematics and Computation, 219(17): 9572–9580, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003639.

Baksalary:2014:GCI

[BT14] Oskar Maria Baksalary and Götz Trenkler. On a generalized core inverse. Applied Mathematics and Computation, 236(??):450–457, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400410X.

Biswas:2011:SSG

[BTHA11] Anjan Biswas, Houria Triki, T. Hayat, and Omar M. Aldossary. 1soliton solution of the generalized Burgers equation with generalized evolution. Applied Mathematics and Computation, 217(24):10289–10294, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007004.

Benbernou:2014:NRC

[BTRZ14] Samia Benbernou, Mekki Terbeche, Maria Alessandra Ragusa, and Zujin Zhang. A note on the regularity criterion for 3D MHD equations in Ḃ_{∞,∞}^{−1} space. Applied Mathematics and Computation, 238(??):245–249, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004573.

Bhrawy:2012:NBM

[BTS12] A. H. Bhrawy, E. Tohidi, and F. Soleymani. A new Bernoulli matrix method for solving high-order linear and nonlinear Fredholm integro-differential equations with piecewise intervals. Applied Mathematics and Computation, 219(2):482–497, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312006194.

Benkamra:2013:BSE

[BTT13] Zohra Benkamra, Mekki Terbeche, and Mounir Tlemcani. Bayesian sequential estimation of the reliability of a parallelseries system. Applied Mathematics and Computation, 219(23):10842– 10852, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313005298.

Bifaretti:2012:EIC

[BTV12] S. Bifaretti, P. Tomei, and C. M. Verrelli. Establishing improved convergence and robustness properties for the repetitive learning control. Applied Mathematics and Computation, 218(22):11311–11322, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005103.

Betchewe:2010:DSG

[BTVC10a] Gambo Betchewe, Bouetou Bouetou Thomas, Kuetche Kamgang Victor, and Kofane Timoleon Crepin. Dynamical survey of a generalized-Zakharov equation and its exact travelling wave solutions. Applied Mathematics and Computation, 217(1):203–211, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005886.

Betchewe:2010:ESS

[BTVC10b] Gambo Betchewe, Bouetou Bouetou Thomas, Kuetche Kamgang Victor, and Kofane Timoleon Crepin. Explicit series solutions to nonlinear evolution equations: The sine–cosine method. Applied Mathematics and Computation, 215(12):4239–4247, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010935.

Bairamov:2012:CVR

[BU12] Elgiz Bairamov and Ekin Ugurlu. On the characteristic values of the real component of a dissipative boundary value transmission problem. Applied Mathematics and Computation, 218(19):9657–9663, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002305.

Bumbariu:2012:NAT

[Bum12] Oana Bumbariu. A new Aitken type method for accelerating iterative sequences. Applied Mathematics and Computation, 219(1):78–82, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005929.

Buong:2010:SCT

[Buo10] Nguyen Buong. Strong convergence theorem of an iterative method for variational inequalities and fixed point problems in Hilbert spaces. Applied Mathematics and Computation, 217(1):322–329, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006235.

Burg:2012:DBC

[Bur12] Clarence O. E. Burg. Derivative-based closed Newton–Cotes numerical quadrature. Applied Mathematics and Computation, 218(13):7052–7065, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015475.

Bos:2012:STI

[BV12a] Len Bos and Marco Vianello. Subperiodic trigonometric interpolation and quadrature. Applied Mathematics and Computation, 218(21): 10630–10638, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004237.

Breckner:2012:OPD

[BV12b] Brigitte E. Breckner and Csaba Varga. Oneparameter Dirichlet problems on the Sierpinski gasket. Applied Mathematics and Computation, 219(4):1813–1820, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008132.

Betchewe:2013:NSG

[BVTC13] Gambo Betchewe, Kuetche Kamgang Victor, Bouetou Bouetou Thomas, and Kofane Timoleon Crepin. New solutions of the Gardner equation: Analytical and numerical analysis of its dynamical understanding. Applied Mathematics and Computation, 223(??):377–388, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008965.

Boyd:2010:ACG

[BW10] John P. Boyd and Lei Wang. Asymptotic coefficients for Gaussian radial basis function interpolants. Applied Mathematics and Computation, 216(8):2394–2407, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003383.

Bian:2014:WPS

[BW14] Baojun Bian and Guolian Wang. Well-posedness of stochastic KdV–BO equation driven by fractional Brownian motion. Applied Mathematics and Computation, 243(??):657–669, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314008145.

Bie:2014:CDI

[BWaY14] Qunyi Bie, Qiru Wang, and Zheng an Yao. Crossdiffusion induced instability and pattern formation for a Holling type-II predator-prey model. Applied Mathematics and Computation, 247(??):1– 12, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314011989.

Bu:2014:ICC

[BWHQ14] Xuhui Bu, Hongqi Wang, Zhongsheng Hou, and Wei Qian. H_{∞} control for a class of 2-D nonlinear systems with intermittent measurements. Applied Mathematics and Computation, 247(??):651–662, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012260.

Boyd:2012:PSR

[BX12] John P. Boyd and Yang (Chris) Xiu. Parity symmetry with respect to both x = 0 and x = L requires periodicity with period 4L: Connections between computer graphics, group theory and spectral methods for solving partial differential equations. Applied Mathematics and Computation, 219(8):3461–3468, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009034.

Bao:2011:NMF

[BY11a] Jianghong Bao and Qigui Yang. A new method to find homoclinic and heteroclinic orbits. Applied Mathematics and Computation, 217(14):6526–6540, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000488.

Bi:2011:TGM

[BY11b] Hai Bi and Yidu Yang. A two-grid method of the non-conforming Crouzeix– Raviart element for the Steklov eigenvalue problem. Applied Mathematics and Computation, 217(23):9669–9678, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006217.

Bao:2014:BAG

[BY14] Jianghong Bao and Qigui Yang. Bifurcation analysis of the generalized stretchtwist-fold flow. Applied Mathematics and Computation, 229(??):16–26, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313013131.

Behnia:2014:ICM

[BYM14] S. Behnia, M. Yahyavi, and F. Mobadersani. Intelligent controlling microbubble radial oscillations by using Slave– Master Feedback control. Applied Mathematics and Computation, 245(??):404–415, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314010716.

Bai:2010:NGA

[BZ10a] Cheng-Lin Bai and Hong Zhao. A new General Algebraic Method and its applications to the (2+1)-dimensional Broer– Kaup–Kupershmidt equations. Applied Mathematics and Computation, 217(4):1719–1732, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009394.

Bai:2010:CVN

[BZ10b] Yuzhen Bai and Pingping Zhang. On a class of Volterra nonlinear equations of parabolic type. Applied Mathematics and Computation, 216(1):236– 240, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000573.

Bai:2011:SFT

[BZ11] Zhanbing Bai and Yinghan Zhang. Solvability of fractional threepoint boundary value problems with nonlinear growth. Applied Mathematics and Computation, 218(5):1719–1725, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008848.

Biswas:2010:SSG

[BZGK10] Anjan Biswas, Essaid Zerrad, Jude Gwanmesia, and Ramzi Khouri. 1-soliton solution of the generalized Zakharov equation in plasmas by He’s variational principle. Applied Mathematics and Computation, 215(12):4462–4466, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000093.

Zhang:2011:BAS

[bZjZS11] Jian bing Zhang, Da jun Zhang, and Qing Shen. Bilinear approach for a symmetry constraint of the modified KdV equation. Applied Mathematics and Computation, 218(8):4494–4500, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311012604.

Zhang:2011:PAS

[bZlD11] Qing bang Zhang and Ru liang Deng. Projection algorithms for the system of generalized mixed variational inequalities in Banach spaces. Applied Mathematics and Computation, 217(19):7718–7724, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002979.

Zhu:2010:ESW

[bZpZJ10] Fu bo Zhu, Qing pei Zang, and Jie Ji. Exact solutions for a Wick-type stochastic 2D KdV equation. Applied Mathematics and Computation, 216(9):2766–2770, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003541.

Biswas:2010:DSP

[BZR10] Anjan Biswas, Essaid Zerrad, and Arjuna Ranasinghe. Dynamics of solitons in plasmas for the complex KdV equation with power law nonlinearity. Applied Mathematics and Computation, 217(4):1491–1496, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030900589X.

Bu:2011:RDI

[BZT11] Changjiang Bu, Jiemei Zhao, and Jiapei Tang. Representation of the Drazin inverse for special block matrix. Applied Mathematics and Computation, 217(10):4935–4943, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011719.

Bi:2014:BMB

[BZZ14] Qinsheng Bi, Ran Zhang, and Zhengdi Zhang. Bifurcation mechanism of bursting oscillations in parametrically excited dynamical system. Applied Mathematics and Computation, 243(??):482–491, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314007085.

Chen:2014:EII

[CA14] Ruyun Chen and Congpei An. On the evaluation of infinite integrals involving Bessel functions. Applied Mathematics and Computation, 235(??):212–220, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003695.

Contreras:2014:IVC

[CAC14] Jhonatan Contreras, Iván Amaya, and Rodrigo Correa. An improved variant of the conventional harmony search algorithm. Applied Mathematics and Computation, 227(??):821–830, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012198.

Caballero-Aguila:2014:IFA

[CÁGGLP14] R. Caballero-Águila, I. García-Garrido, and J. Linares-Pérez. Information fusion algorithms for state estimation in multi-sensor systems with correlated missing measurements. Applied Mathematics and Computation, 226(??):

548–563, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313011259.

Cai:2010:MSS

[Cai10a]  Jiaxiang Cai.  Multi-

symplectic schemes for strongly coupled schrödinger system. Applied Mathematics and Computation, 216(8):2417–2429, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003413.

Cai:2010:ADM

[Cai10b] Shaohong Cai. Analysis of a delayed Monod type chemostat model with impulsive input the polluted nutrient. Applied Mathematics and Computation, 217(6):2320–2326, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007757.

Cai:2011:SLN

[Cai11] Jiaxiang Cai. Some linearly and non-linearly implicit schemes for the numerical solutions of the regularized long-wave equation. Applied Mathematics and Computation, 217(24):9948–9955, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005984.

Cakmak:2010:LTIa

[Çak10] Devrim Çakmak. Lyapunovtype integral inequalities for certain higher order differential equations. Applied Mathematics and Computation, 216(2):368–373, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000238.

Cakmak:2013:LTI

[Çak13] Devrim Çakmak. Lyapunovtype inequalities for two classes of nonlinear systems with anti-periodic boundary conditions. Applied Mathematics and Computation, 223(??): 237–242, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

com/science/article/
pii/S009630031300862X.

Chandrashekar:2011:CMM

[CALR11] R. Chandrashekar, Rosihan M. Ali, See Keong Lee, and V. Ravichandran. Convolutions of meromorphic multivalent functions with respect to nply points and symmetric conjugate points. Applied Mathematics and Computation, 218(3):723–728, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004371.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Cangul:2011:PHM

[Can11] Ismail Naci Cangül. Professor H. M. Srivastava: Man and mathematician. Applied Mathematics and Computation, 218(3):655–657, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009465.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Cao:2010:BGF

[Cao10a]  Jian Cao.  Bivariate

generating functions for Rogers–Szegő polynomials. Applied Mathematics and Computation, 217 (5):2209–2216, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007666.

Cao:2010:SSI

[Cao10b] Wanrong Cao. T -stability of the semi-implicit Euler method for delay differential equations with multiplicative noise. Applied Mathematics and Computation, 216(3):999–1006, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001426.

Cao:2012:CTG

[Cao12] Jian Cao. On Carlitz’s trilinear generating functions. Applied Mathematics and Computation, 218(19):9839–9847, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300312002986.

Cruz:2014:SAS

[CAP14] J. Y. Bello Cruz, G. Bouza Allende, and L. R. Lucambio Pérez. Subgradient algorithms for solving variable inequalities. Applied Mathematics and Computation, 247 (??):1052–1063, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012983.

Carsetti:2012:EMC

[Car12]  Arturo Carsetti.  The

emergence of meaning at the co-evolutionary level: an epistemological approach. Applied Mathematics and Computation, 219(1):14–23, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010794.

Caruntu:2013:ESP

[Car13a] Dumitru I. Caruntu. Eigenvalue singular problem of factorized fourth-order self-adjoint differential equations. Applied Mathematics and Computation, 224(??):603–610, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009557.

Caruntu:2013:FSA

[Car13b] Dumitru I. Caruntu. Factorization of self-adjoint ordinary differential equations. Applied Mathematics and Computation, 219(14):7622–7631, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000891.

Ciric:2012:STF

[ĆARA12] Ljubomir Ćirić, Mujahid Abbas, Miloje Rajović, and Basit Ali. Suzuki type fixed point theorems for generalized multivalued mappings on a set endowed with two bmetrics. Applied Mathematics and Computation, 219(4):1712–1723, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200803X.

Ciric:2011:CFPa

[ĆASH11] Ljubomir Ćirić, Mujahid Abbas, Reza Saadati, and Nawab Hussain. Common fixed points of almost generalized contractive mappings in ordered metric spaces. Applied Mathematics and Computation, 217(12):5784–5789, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012622.

Ceng:2011:REI

[CAY11] Lu-Chuan Ceng, Qamrul Hasan Ansari, and Jen-Chih Yao. Relaxed extragradient iterative methods for variational inequalities. Applied Mathematics and Computation, 218(3):1112–1123, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000889.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Cazzani:2014:TEY

[Caz14] Antonio Cazzani. On the true extrema of Young’s modulus in hexagonal materials. Applied Mathematics and Computation, 238(??):397–407, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005360.

Chen:2010:TGM

[CB10] Chuanjun Chen and Chunjia Bi. Two-grid methods for characteristic finite volume element solution of semilinear convection–diffusion equations. Applied Mathematics and Computation, 217(5):1896–1906, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007290.

Candito:2012:ESN

[CB12a] Pasquale Candito and Giovanni Molica Bisci. Existence of solutions for a nonlinear algebraic system with a parameter. Applied Mathematics and Computation, 218(23):11700–11707, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312005735.

Chen:2012:SIM

[CB12b] Chao-Ping Chen and Necdet Batir. Some inequalities and monotonicity properties associated with the gamma and psi functions. Applied Mathematics and Computation, 218(17):8217–8225, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001257.

Carvalho:2014:MOA

[CBA^{+}14] Esdras P. Carvalho, Carolina Borges, Doherty Andrade, Jin Yun Yuan, and Mauro A. S. S. Ravagnani. Modeling and optimization of an ammonia reactor using a penaltylike method. Applied Mathematics and Computation, 237(??):330–339, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004615.

Chen:2012:PCT

[CBLC12] Zhe Chen, Qunfang Bao, Shenghong Li, and Jianli Chen. Pricing CDO tranches with stochastic correlation and random factor loadings in a mixture copula model. Applied Mathematics and Computation, 219(6):2909–2916, November 25, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009137.

Chaudhry:2012:WAL

[CBO12] Jehanzeb Hameed Chaudhry, Stephen D. Bond, and Luke N. Olson. A weighted adaptive least-squares finite element method for the Poisson–Boltzmann equation. Applied Mathematics and Computation, 218(9):4892–4902, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013002.

Chouakri:2011:QCD

[CBRTA11] S. A. Chouakri, F. Bereksi-Reguig, and A. Taleb-Ahmed. QRS complex detection based on multi wavelet packet decomposition. Applied Mathematics and Computation, 217(23):9508–9525, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003377.

Cho:2012:GFI

[CBS12] Nak Eun Cho, Teodor Bulboaca, and H. M. Srivastava. A General family of integral operators and associated subordination and superordination properties of some special analytic function classes. Applied Mathematics and Computation, 219(4):2278–2288, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008685.

Cevik:2014:SDS

[ÇBS14] Mehmet Çevik, M. Mustafa Bahsi, and Mehmet Sezer. Solution of the delayed single degree of freedom system equation by exponential matrix method. Applied Mathematics and Computation, 242(??):444–453, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400811X.

Cardenas-Barron:2014:CMP

[CBTGWW14] Leopoldo Eduardo Cárdenas-Barrón, Gerardo Treviño-Garza, Gede Agus Widyadana, and Hui-Ming Wee. A constrained multi-products EPQ inventory model with discrete delivery order and lot size. Applied Mathematics and Computation, 230(??):359–370, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031301360X.

Cai:2010:HPD

[CC10a] Jing Cai and Guoliang Chen. On the Hermitian positive definite solutions of nonlinear matrix equation X^{s} + A ∗ X^{−t}A = Q. Applied Mathematics and Computation, 217(1):117–123, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310005679.

Chen:2010:GEA

[CC10b] Zhixiang Chen and Feilong Cao. Global errors for approximate approximations with Gaussian kernels on compact intervals. Applied Mathematics and Computation, 217(2):725–734, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006843.

Chinviriyasit:2010:NMS

[CC10c] Settapat Chinviriyasit and Wirawan Chinviriyasit. Numerical modelling of an SIR epidemic model with diffusion. Applied Mathematics and Computation, 216(2):395–409, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000041X.

Cao:2011:CBT

[CC11a] Xifang Cao and Ying Chen. On classification of Bäcklund transformations. Applied Mathematics and Computation, 217(21):8552–8559, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004504.

Celentano:2011:CEM

[CC11b] Laura Celentano and Alessandro Coppola. A computationally efficient method for modeling flexible robots based on the assumed modes method. Applied Mathematics and Computation, 218(8):4483– 4493, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311012598.

Cheng:2011:SGC

[CC11c] Zunshui Cheng and Jinde Cao. Synchronization of a growing chaotic network model. Applied Mathematics and Computation, 218(5):2122–2127, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009623.

Castellano:2012:OCI

[CC12a] Rosella Castellano and Roy Cerqueti. Optimal consumption/investment problem with light stocks: a mixed continuous-discrete time approach. Applied Mathematics and Computation, 218(12):6887–6898, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311015529.

Chen:2012:BCB

[CC12b] Boshan Chen and Jiejie Chen. Bifurcation and chaotic behavior of a discrete singular biological economic system. Applied Mathematics and Computation, 219(5):2371–2386, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200759X.

Chung:2012:EFS

[CC12c] K. W. Chung and Y. Y.

Cao. Exact front, soliton and hole solutions for a modified complex Ginzburg–Landau equation from the harmonic balance method. Applied Mathematics and Computation, 218(9):5140–5145, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013269.

Cardone:2013:MCM

[CC13] Angelamaria Cardone and Dajana Conte. Multistep collocation methods for Volterra integrodifferential equations. Applied Mathematics and Computation, 221(??): 770–785, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313007625.

Chao:2014:NCA

[CC14a] Miantao Chao and Caozong Cheng. A note on the convergence of alternating proximal gradient method. Applied Mathematics and Computation, 228(??):258–263, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012708.

Chen:2014:IAE

[CC14b] Chao-Ping Chen and Junesang Choi. Inequalities and asymptotic expansions for the constants of Landau and Lebesgue. Applied Mathematics and Computation, 248(??):610–624, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314013836.

Chen:2014:NSH

[CC14c] Y. Chen and Chang-Ming Chen. Numerical scheme with high order accuracy for solving a modified fractional diffusion equation. Applied Mathematics and Computation, 244(??):772–782, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010017.

Chidume:2011:CHS

[CCA11] C. E. Chidume, C. O.

Chidume, and Bashir Ali. Convergence of hybrid steepest descent method for variational inequalities in Banach spaces. Applied Mathematics and Computation, 217(23):9499–9507, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003225.

Cangul:2011:P

[CCC11] Ismail Naci Cangül, Nak Eun Cho, and Junesang Choi. Preface. Applied Mathematics and Computation, 218(3):653–654, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009453.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Chen-Charpentier:2013:SRA

[CCCRR13] Benito M. Chen-Charpentier, J.-C. Cortés, J.-V. Romero, and M.-D. Roselló. Some recommendations for applying gPC (generalized polynomial chaos) to modeling: an analysis through the Airy random differential equation. Applied Mathematics and Computation, 219(9):4208–4218, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011605.

Cakalli:2010:QSO

[ÇÇD10] Hüseyin Çakallı, İbrahim Çanak, and Mehmet Dik. Δ-quasi-slowly oscillating continuity. Applied Mathematics and Computation, 216(10):2865–2868, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310004030.

Chen:2010:ANS

[CCG10] Mingxiang Chen, Dominic P. Clemence, and Gregory Gibson. Analysis of numerical schemes of a mathematical model for sickle cell depolymerization. Applied Mathematics and Computation, 216(5):1489–1500, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002390.

Cabrera:2012:FCM

[CCG^{+}12a] I. P. Cabrera, P. Cordero, G. Gutiérrez, J. Martínez, and M. Ojeda-Aciego. Finitary coalgebraic multisemilattices and multilattices. Applied Mathematics and Computation, 219(1):31–44, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013270.

Calude:2012:PSI

[CCG12b] Cristian S. Calude, José Félix Costa, and Hélia Guerra. Preface to the Special Issue on Physics and Computation “Towards a Computational Interpretation of Physical Theories”. Applied Mathematics and Computation, 219(1):1–2, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007412.

Chicharro:2013:CDD

[CCGT13] Francisco Chicharro, Alicia Cordero, José M. Gutiérrez, and Juan R. Torregrosa. Complex dynamics of derivativefree methods for nonlinear equations. Applied Mathematics and Computation, 219(12):7023–7035, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000167.

Calbo:2010:RMD

[CCJ10] G. Calbo, J.-C. Cortés, and L. Jódar. Random matrix difference models arising in longterm medical drug strategies. Applied Mathematics and Computation, 217(5):2149–2161, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007599.

Calbo:2011:RHD

[CCJ11] G. Calbo, J.-C. Cortés, and L. Jódar. Random Hermite differential equations: Mean square power series solutions and statistical properties. Applied Mathematics and Computation, 218(7):3654–3666, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101174X.

Casaban:2014:CFN

[CCJ14a] M.-C. Casabán, R. Company, and L. Jódar. Closed form numerical solutions of variable coefficient linear second-order elliptic problems. Applied Mathematics and Computation, 238(??):266–280, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400558X.

Cheon:2014:ARA

[CCJ14b] Gi-Sang Cheon, Bong Dae Choi, and Sung-Tae Jin. An application of Riordan arrays to the transient analysis of M/M/1 queues. Applied Mathematics and Computation, 237(??):659–671, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400513X.

Chen-Charpentier:2011:PAN

[CCJT11] Benito Chen-Charpentier, Lucas Jódar, and Aleksey S. Telyakovskiy. Polynomial approximation of nonlinear differential systems with prefixed accuracy. Applied Mathematics and Computation, 218(5):1650–1657, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008770.

Chakraborty:2011:OCH

[CCK11] Kunal Chakraborty, Milon Chakraborty, and T. K. Kar. Optimal control of harvest and bifurcation of a prey-predator model with stage structure. Applied Mathematics and Computation, 217(21):8778–8792, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005194.

Chiang:2012:SEA

[CCL12] Chun-Yueh Chiang, Eric King-Wah Chu, and Wen-Wei Lin. On the ⋆- Sylvester equation AX ± X^{⋆}B^{⋆} = C. Applied Mathematics and Computation, 218(17):8393–8407, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001142.

Castillo:2013:COL

[CCLF13] René E. Castillo, Dana D.

Clahane, Juan F. Farías López, and Julio C. Ramos Fernández. Composition operators from logarithmic Bloch spaces to weighted Bloch spaces. Applied Mathematics and Computation, 219(12): 6692–6706, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312012581.

Chen:2010:WCF

[CCLW10] Sheng-Gwo Chen, Mei-Hsiu Chi, Ying-Jen Lin, and Jyh-Yang Wu. A weighted curvature flow for shape deformation. Applied Mathematics and Computation, 217(5):2097– 2117, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000754X.

Cabada:2012:CGF

[CCMV12] Alberto Cabada, José Ángel Cid, and Beatriz Máquez-Villamarín. Computation of Green’s functions for boundary value problems with Mathematica. Applied Mathematics and Computation, 219(4):1919–1936, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008284.

Cheng:2012:ROO

[CCO12] Mei-Chuan Cheng, Chun-Tao Chang, and Liang-Yuh Ouyang. The retailer’s optimal ordering policy with trade credit in different financial environments. Applied Mathematics and Computation, 218(19):9623–9634, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002068.

Chen:2013:SPV

[CCS13] Wei-Ren Chen, Chun-Sheng Chen, and Jenq-Huey Shyu. Stability of parametric vibrations of laminated composite plates. Applied Mathematics and Computation, 223(??):127–138, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008564.

Canto:2014:SVM

[CCS14] B. Cantó, C. Coll, and E. Sánchez. A study on vaccination models for a seasonal epidemic process. Applied Mathematics and Computation, 243(??):152–160, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008042.

Ciric:2013:TAB

[CCSM13] Vladimir Ciric, Aleksandar Cvetković, Vladimir Simić, and Ivan Milentijević. Tropical algebra based framework for error propagation analysis in systolic arrays. Applied Mathematics and Computation, 225(??):512–525, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010382.

Chen:2013:PSS

[CCWM13] Fengde Chen, Wanlin Chen, Yumin Wu, and Zhaozhi Ma. Permanence of a stagestructured predator-prey system. Applied Mathematics and Computation, 219(17):8856–8862, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003081.

Chen:2011:LCA

[CCX11] Xiaohong Chen, Songcan Chen, and Hui Xue. Large correlation analysis. Applied Mathematics and Computation, 217(22):9041–9052, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004851.

Cheng:2010:NCP

[CD10a] Liqing Cheng and Yunda Dong. A note on convergence of PRP method with new nonmonotone line search. Applied Mathematics and Computation, 216(7):2232–2234, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003139.

Chu:2010:PPS

[CD10b] Yuming Chu and Changming Ding. Prey– predator systems with intertwined basins of attraction. Applied Mathematics and Computation, 215(12):4422–4425, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000147.

Chompuvised:2011:SBV

[CD11a] Kanittha Chompuvised and Ampon Dhamacharoen.

Solving boundary value problems of ordinary differential equations with non-separated boundary conditions. Applied Mathematics and Computation, 217(24):10355–10360, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007144.

Cong:2011:TTC

[CD11b] Nguyen Huu Cong and La Tri Dung. Twostepby-twostep continuous PIRKN-type PC methods for nonstiff IVPs. Applied Mathematics and Computation, 217(20):8010–8019, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003298.

Chen:2012:PCM

[CD12] Suanrong Chen and Jiu Ding. A piecewise constant method for Frobenius– Perron operators via delta function approximations. Applied Mathematics and Computation, 219(3):1047– 1052, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312007096.

Chidume:2013:IMS

[CD13] C. E. Chidume and N. Djitté. An iterative method for solving nonlinear integral equations of Hammerstein type. Applied Mathematics and Computation, 219(10):5613–5621, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012180.

Cvetkovic:2014:MNE

[CD14] Ljiljana Cvetković and Ksenija Doroslovacki. Max norm estimation for the inverse of block matrices. Applied Mathematics and Computation, 242(??):694–706, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008716.

Cuevas:2013:BPV

[CDCS13] Claudio Cuevas, Filipe Dantas, Mario Choquehuanca, and Herme Soto. l^{p}-boundedness properties for Volterra difference equations. Applied Mathematics and Computation, 219(12):6986–6999, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201329X.

Cvetkovic:2013:INB

[CDDL13] Ljiljana Cvetković, Ping-Fan Dai, Ksenija Doroslovacki, and Yao-Tang Li. Infinity norm bounds for the inverse of Nekrasov matrices. Applied Mathematics and Computation, 219(10):5020–5024, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012234.

Chen:2014:EBD

[CDESS14] Zengqiang Chen, Matthias Dehmer, Frank Emmert-Streib, and Yongtang Shi. Entropy bounds for dendrimers. Applied Mathematics and Computation, 242(??):462–472, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008054.

Castillo:2013:PAH

[CDGR13] K. Castillo, D. K. Dimitrov, L. E. Garza, and F. R. Rafaeli. Perturbations on the antidiagonals of Hankel matrices. Applied Mathematics and Computation, 221(??):444–452, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007522.

Cvetkovic:2010:GSM

[CDIS10] Dragoš Cvetković, Tatjana Davidović, Aleksandar Ilić, and Slobodan K. Simić. Graphs for small multiprocessor interconnection networks. Applied Mathematics and Computation, 217(6):2468–2480, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008039.

Chakraborty:2013:NSH

[CDK13] Kunal Chakraborty, Sanjoy Das, and T. K. Kar. On non-selective harvesting of a multispecies fishery incorporating partial closure for the populations. Applied Mathematics and Computation, 221(??):581–597, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007042.

Chen:2011:CAM

[CDLZ11] Huiqin Chen, Jinqiao Duan, Xiaofan Li, and Chengjian Zhang. A computational analysis for mean exit time under non-Gaussian Lévy noises. Applied Mathematics and Computation, 218(5):1845–1856, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009131.

Chatzarakis:2014:CND

[CDMS14] G. E. Chatzarakis, J. Diblík, G. N. Miliaras, and I. P. Stavroulakis. Classification of neutral difference equations of any order with respect to the asymptotic behavior of their solutions. Applied Mathematics and Computation, 228(??):77–90, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012484.

Christou:2012:HMC

[CDMT12] D. Christou, A. Danelakis, M. Mitrouli, and D. Triantafyllou. A hybrid method for computing the intersection and tangency points of plane curves. Applied Mathematics and Computation, 219(4):1975–1987, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008326.

Caglar:2011:CBS

[ÇDO11] Murat Çaglar, Erhan Deniz, and Halit Orhan. Coefficient bounds for a subclass of starlike functions of complex order. Applied Mathematics and Computation, 218(3):693–698, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001202.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Capone:2013:SNA

[CDR13] F. Capone, R. De Luca, and S. Rionero. On the stability of nonautonomous perturbed Lotka–Volterra models. Applied Mathematics and Computation, 219(12): 6868–6881, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000374.

Caballero:2013:SFH

[CDS13] Josefa Caballero, Mohamed Abdalla Darwish, and Kishin Sadarangani. Solvability of a fractional hybrid initial value problem with supremum by using measures of noncompactness in Banach algebras. Applied Mathematics and Computation, 224(??):553–563, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009284.

Cong:2014:SMP

[CDST14] N. D. Cong, T. S. Doan, S. Siegmund, and H. T. Tuan. On stable manifolds for planar fractional differential equations. Applied Mathematics and Computation, 226(??):157–168, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010655.

Canak:2011:TTS

[ÇE11] Ibrahim Çanak and Yilmaz Erdem. On Tauberian theorems for (A)(C, α) summability method. Applied Mathematics and Computation, 218(6):2829– 2836, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311010526.

Crowder:2013:NSF

[CE13] T. R. Crowder and V. J.

Ervin. Numerical simulations of fluid pressure in the human eye. Applied Mathematics and Computation, 219(24):11119– 11133, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313005614.

Celinski:2014:SSA

[Cel14] Rafał Celiński. Stability of solutions to aggregation equation in bounded domains. Applied Mathematics and Computation, 228(??):49–58, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011892.

Cai:2011:ITP

[CEM11] Weiwei Cai, David J. Ewing, and Lin Ma. Investigation of temperature parallel simulated annealing for optimizing continuous functions with application to hyperspectral tomography. Applied Mathematics and Computation, 217(12):5754–5767, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012567.

Cengellenmis:2011:SSC

[Cen11]  Y. Cengellenmis.  On

some special codes over F_{p} + uF_{p} + u^{2}F_{p}. Applied Mathematics and Computation, 218(3):720–722, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/

pii/S0096300311004280.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Cobos:2011:GLG

[CEP11] Carlos Cobos, Dario Estupiñán, and José Pérez. GHS + LEM: Globalbest Harmony Search using learnable evolution models. Applied Mathematics and Computation, 218(6):2558–2578, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010228.

Cetinkaya:2013:ISA

[Çet13] Aysegül Çetinkaya. The incomplete second Appell hypergeometric functions. Applied Mathematics and Computation, 219(15):8332–8337, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012179.

Christodoulou:2012:SFB

[CEXG12] E. Christodoulou, M. Elliotis, C. Xenophontos, and G. C. Georgiou. The singular function boundary integral method for 3-D Laplacian problems with a boundary straight edge singularity. Applied Mathematics and Computation, 219(3):1073–1081, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007138.

Cerna:2012:CSW

[CF12] Dana Cerná and Václav Finek. Cubic spline wavelets with complementary boundary conditions. Applied Mathematics and Computation, 219(4):1853–1865, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200820X.

Chen:2013:NAS

[CF13] Gang Chen and Minfu Feng. A new absolutely stable simplified Galerkin least-squares finite element method using nonconforming element for the Stokes problem. Applied Mathematics and Computation, 219(10):5356–5366, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012143.

Cerna:2014:CSW

[CF14] Dana Cerná and Václav Finek. Cubic spline wavelets with short support for fourth-order problems. Applied Mathematics and Computation, 243(??):44–56, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007309.

Clark:2013:TNL

[CFCLM13] H. R. Clark, E. Fernández-Cara, J. Limaco, and L. A. Medeiros. Theoretical and numerical local null controllability for a parabolic system with local and nonlocal nonlinearities. Applied Mathematics and Computation, 223(??):483–505, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300903X.

Corazza:2013:PSO

[CFG13] Marco Corazza, Giovanni Fasano, and Riccardo Gusso. Particle Swarm Optimization with nonsmooth penalty reformulation, for a complex portfolio selection problem. Applied Mathematics and Computation, 224(??):611–624, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008473.

Castillo:2013:NMS

[CFGRV13] D. Castillo, A. M. Ferreiro, J. A. García-Rodríguez, and C. Vázquez. Numerical methods to solve PDE models for pricing business companies in different regimes and implementation in GPUs. Applied Mathematics and Computation, 219(24):11233– 11257, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313005651.

Cesenek:2013:SCV

[CFH^{+}13] Jan Cesenek, Miloslav Feistauer, Jaromír Horácek, Václav Kucera, and Jaroslava Prokopová. Simulation of compressible viscous flow in time-dependent domains. Applied Mathematics and Computation, 219(13):7139–7150, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011180.

Cai:2014:NSP

[CFL14] Liming Cai, Bin Fang, and Xuezhi Li. A note of a staged progression HIV model with imperfect vaccine. Applied Mathematics and Computation, 234(??):412–416, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002483.

Cicirelli:2012:MCT

[CFN12] Franco Cicirelli, Angelo Furfaro, and Libero Nigro. Model checking time-dependent system specifications using Time Stream Petri Nets and Uppaal. Applied Mathematics and Computation, 218(16):8160–8186, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001397.

Castillo:2011:BSO

[CFS11] René Erlin Castillo, Julio C. Ramos Fernández, and Miguel Salazar. Bounded superposition operators between Bloch–Orlicz and α-Bloch spaces. Applied Mathematics and Computation, 218(7):3441–3450, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011301.

Chun-Feng:2010:GOS

[CFSYPP10] Wang Chun-Feng, Liu San-Yang, and Shen Pei-Ping. Global optimization for sum of geometric fractional functions. Applied Mathematics and Computation, 216(8):2263–2270, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003152.

Chen:2014:NPB

[CFX14] Gang Chen, Minfu Feng, and Chunmei Xie. A new projection-based stabilized method for steady convection-dominated convectiondiffusion equations. Applied Mathematics and Computation, 239(??):89– 106, July 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314005426.

Chen:2014:LPS

[CFZ14a] Gang Chen, Minfu Feng, and Hong Zhou. Local projection stabilized method on unsteady Navier– Stokes equations with high Reynolds number using equal order interpolation. Applied Mathematics and Computation, 243(??):465–481, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007863.

Cheng:2014:IMS

[CFZ14b] Hao Cheng, Chu-Li Fu, and Yuan-Xiang Zhang. An iteration method for stable analytic continuation. Applied Mathematics and Computation, 233(??):203–213, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000903.

Clavero:2010:UCF

[CG10] C. Clavero and J. L. Gracia. On the uniform convergence of a finite difference scheme for time dependent singularly perturbed reaction-diffusion problems. Applied Mathematics and Computation, 216(5):1478–1488, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002389.

Caraffini:2012:SES

[CG12a] G. L. Caraffini and M. Galvani. Symmetries and exact solutions via conservation laws for some partial differential equations of Mathematical Physics. Applied Mathematics and Computation, 219(4):1474–1484, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007667.

Clavero:2012:HOH

[CG12b] C. Clavero and J. L. Gracia. A high order HODIE finite difference scheme for 1D parabolic singularly perturbed reaction-diffusion problems. Applied Mathematics and Computation, 218(9):5067–5080, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101318X.

Canovas:2014:DRC

[CG14a] Jose S. Cánovas and María Muñoz Guillermo. On the dynamics of a restricted Cournot–Puu triopoly: Firms survival and complexity. Applied Mathematics and Computation, 237(??):356–372, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004317.

Choudhury:2014:DEM

[CG14b] A. Ghose Choudhury and Partha Guha. Damped equations of Mathieu type. Applied Mathematics and Computation, 229(??):85–93, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012812.

Clavero:2014:IUC

[CG14c] C. Clavero and J. L. Gracia. An improved uniformly convergent scheme in space for 1D parabolic reaction-diffusion systems. Applied Mathematics and Computation, 243(??):57– 73, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007693.

Coroianu:2014:LRB

[CG14d] Lucian Coroianu and Sorin G. Gal. Localization results for the Bernstein max-product operator. Applied Mathematics and Computation, 231(??):73– 78, March 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000290.

Cui:2014:BGI

[CG14e] Rongrong Cui and Chuanqing Gu. Bivariate generalized inverse Newton– Thiele type matrix Padé approximation. Applied Mathematics and Computation, 236(??):202–214, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003841.

Cafaro:2010:RIG

[CGAK10] C. Cafaro, A. Giffin, S. A. Ali, and D.-H. Kim. Reexamination of an information geometric construction of entropic indicators of complexity. Applied Mathematics and Computation, 217(7):2944–2951, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008829.

Calio:2012:ASN

[CGGMG12] Franca Caliò, A. I. Garralda-Guillem, E. Marchetti, and M. Ruiz Galán. About some numerical approaches for mixed integral equations. Applied Mathematics and Computation, 219(2):464–474, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005966.

Calio:2013:NAS

[CGGMG13] F. Caliò, A. I. Garralda-Guillem, E. Marchetti, and M. Ruiz Galán. Numerical approaches for systems of Volterra–Fredholm integral equations. Applied Mathematics and Computation, 225(??):811–821, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313010606.

Consolo:2011:LFL

[CGGZ11] G. Consolo, G. Gubbiotti, L. Giovannini, and R. Zivieri. Lagrangian formulation of the linear autonomous magnetization dynamics in spin-torque auto-oscillators. Applied Mathematics and Computation, 217(21):8204–8215, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002335.

Cen:2012:CLM

[CGJ12] Yi Cen, Rentao Gu, and Yuefeng Ji. Concentration of S-largest mutilated vectors with l_{p}-quasinorm for 0 < p ≤ 1 and its applications. Applied Mathematics and Computation, 219(3):1087–1092, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007151.

Choudhury:2012:LDH

[CGK12] A. Ghose Choudhury, Partha Guha, and Nikolay A. Kudryashov. A Lagrangian description of the higher-order Painlevé equations. Applied Mathematics and Computation, 218(11):6612–6619, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014664.

Cai:2012:GSD

[CGL12] Li-Ming Cai, Bao-Zhu Guo, and Xue-Zhi Li. Global stability for a delayed HIV-1 infection model with nonlinear incidence of infection. Applied Mathematics and Computation, 219(2):617–623, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006595.

Corli:2012:MAL

[CGLM^{+}12] Andrea Corli, Ingenuin Gasser, Mária Lukácová-Medvid’ová, Arne Roggensack, and Ulf Teschke. A multiscale approach to liquid flows in pipes I: the single pipe. Applied Mathematics and Computation, 219(3):856–874, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006625.

Castillo:2013:CAD

[CGM13] Kenier Castillo, Lino G. Garza, and Francisco Marcellán. On computational aspects of discrete Sobolev inner products on the unit circle. Applied Mathematics and Computation, 223(??):452–460, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008989.

Chiaselotti:2014:PSD

[CGO14] G. Chiaselotti, T. Gentile, and P. A. Oliverio. Parallel and sequential dynamics of two discrete models of signed integer partitions. Applied Mathematics and Computation, 232(??):1249–1261, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001635.

Chen:2012:RCP

[CGRS12] Xingwu Chen, Jaume Giné, Valery G. Romanovski, and Douglas S. Shafer. The 1: −q resonant center problem for certain cubic Lotka–Volterra systems. Applied Mathematics and Computation, 218(23):11620–11633, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005607.

Carbajal-Gomez:2013:OPL

[CGTCF13] V. H. Carbajal-Gómez, E. Tlelo-Cuautle, and F. V. Fernández. Optimizing the positive Lyapunov exponent in multiscroll chaotic oscillators with differential evolution algorithm. Applied Mathematics and Computation, 219(15):8163–8168, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001793.

Chaves-Gonzalez:2014:MAB

[CGVR14] José M. Chaves-González and Miguel A. Vega-Rodríguez. A multiobjective approach based on the behavior of fireflies to generate reliable DNA sequences for molecular computing. Applied Mathematics and Computation, 227(??):291–308, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012009.

Cai:2014:AEH

[CGW14] Liming Cai, Shuli Guo, and Shuping Wang. Analysis of an extended HIV/ AIDS epidemic model with treatment. Applied Mathematics and Computation, 236(??):621–627, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003269.

Chen:2012:NRP

[CH12] Huiwen Chen and Zhimin He. New results for perturbed Hamiltonian systems with impulses. Applied Mathematics and Computation, 218(18): 9489–9497, May 15, 2012.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312002706.

Chen:2013:NSC

[CH13] Chun-I Chen and Shou-Jen Huang. The necessary and sufficient condition for GM(1, 1) grey prediction model. Applied Mathematics and Computation, 219(11):6152–6162, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012829.

Cabada:2014:NFD

[CH14] Alberto Cabada and Zakaria Hamdi. Nonlinear fractional differential equations with integral boundary value conditions. Applied Mathematics and Computation, 228(??):251–257, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012265.

Chang:2010:NST

[Cha10] Shih-Hsiang Chang. Numerical solution of Troesch’s problem by simple shooting method. Applied Mathematics and Computation, 216(11):3303–3306, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004844.

Chanane:2011:SCP

[Cha11] Bilal Chanane. Solutions of a class of partial differential equations with application to the Black– Scholes equation. Applied Mathematics and Computation, 217(19):7845–7850, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003365.

Chang:2013:QCM

[Cha13a] Shun-Chang Chang. Quenching chaos of a magnetically levitated system based on dither signals. Applied Mathematics and Computation, 222(??):539–547, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008126.

Chang:2013:AVN

[Cha13b] Tai-Ping Chang. Axial vibration of non-uniform and non-homogeneous nanorods based on nonlocal elasticity theory. Applied Mathematics and Computation, 219(10):4933–4941, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201226X.

Chatjigeorgiou:2013:NSC

[Cha13c] Ioannis K. Chatjigeorgiou.

Numerical simulation of the chaotic lateral vibrations of long rotating beams. Applied Mathematics and Computation, 219(10):5592–5612, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201243X.

Chang:2014:NAT

[Cha14a] Shih-Hsiang Chang. New algorithms for Taylor coefficients of indefinite integrals and their applications. Applied Mathematics and Computation, 247(??):872–879, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012429.

Chang:2014:TSM

[Cha14b] Shih-Hsiang Chang. Taylor series method for solving a class of nonlinear singular boundary value problems arising in applied science. Applied Mathematics and Computation, 235(??):110–117, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003427.

Chang:2014:SDF

[Cha14c] T.-P. Chang. Stochastic dynamic finite element analysis of bridge-vehicle system subjected to random material properties and loadings. Applied Mathematics and Computation, 242(??):20–35, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007036.

Cai:2011:PPS

[CHC11] Zuowei Cai, Lihong Huang, and Haibo Chen. Positive periodic solution for a multispecies competitionpredator system with Holling III functional response and time delays. Applied Mathematics and Computation, 217(10):4866–4878, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010441.

Cheng:2012:MPS

[CHC12] Min-Yuan Cheng, Kuo-Yu Huang, and Hung-Ming Chen. K-means particle swarm optimization with embedded chaotic search for solving multidimensional problems. Applied Mathematics and Computation, 219(6):3091–3099, November 25, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009381.

Chen:2010:MPF

[Che10a] Chao-Ping Chen. Monotonicity properties of functions related to the psi function. Applied Mathematics and Computation, 217(7):2905–2911, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009604.

Chen:2010:AMI

[Che10b] Fang Chen. Asynchronous multisplitting iteration with different weighting schemes. Applied Mathematics and Computation, 216(6):1771–1776, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010741.

Chen:2010:LBM

[Che10c]  Sheng Chen.  Lattice

Boltzmann method for slip flow heat transfer in circular microtubes: Extended Graetz problem. Applied Mathematics and Computation, 217(7):3314–3320, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009276.

Chen:2010:DDS

[Che10d] Shyh-Feng Chen. Delaydependent stability for 2D systems with timevarying delay subject to state saturation in the Roesser model. Applied Mathematics and Computation, 216(9):2613–2622, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003589.

Chen:2010:ARC

[Che10e] Ye-Hwa Chen. Adaptive robust control of artificial swarm systems. Applied Mathematics and Computation, 217(3):980–987, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005060.

Cheng:2010:NCB

[Che10f] Ching-Yang Cheng. Natural convection boundary layer flow of fluid with temperature-dependent viscosity from a horizontal elliptical cylinder with constant surface heat flux. Applied Mathematics and Computation, 217(1):83–91, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005308.

Chen:2011:QOS

[Che11a] Chaio-Shiung Chen. Quadratic optimal synchronization of uncertain chaotic systems. Applied Mathematics and Computation, 217(24):10377–10386, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100720X.

Chen:2011:SAF

[Che11b] Guoxin Chen. A simple adaptive feedback control method for chaos and hyper-chaos control. Applied Mathematics and Computation, 217(17):7258–7264, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001937.

Chen:2011:RNC

[Che11c] Ruey-Maw Chen. Reducing network and computation complexities in neural based real-time scheduling scheme. Applied Mathematics and Computation, 217(13):6379–6389, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000282.

Chen:2011:EGD

[Che11d] Sheng Chen. Entropy generation of double-diffusive convection in the presence of rotation. Applied Mathematics and Computation, 217(21):8575–8597, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004632.

Chen:2011:LIR

[Che11e] Xiaochun Chen. Logarithmically improved regularity criterion for the harmonic heat flow and related equations. Applied Mathematics and Computation, 217(13):6260–6263, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013317.

Chen:2011:SCR

[Che11f]  Y-Chuang Chen.  Su-

per connectivity of kregular interconnection networks. Applied Mathematics and Computation, 217(21):8489–8494, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100419X.

Chen:2011:NSD

[Che11g] Y. Z. Chen. Numerical solution for degenerate scale problem arising from multiple rigid lines in plane elasticity. Applied Mathematics and Computation, 218(1):96–106, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007272.

Chen:2012:CDE

[Che12a] Jin-Xiao Chen. A comment on DEA efficiency assessment using ideal and anti-ideal decision making units. Applied Mathematics and Computation, 219(2):583–591, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006534.

Cheng:2012:RSU

[Che12b] Chao-Jung Cheng. Robust synchronization of uncertain unified chaotic systems subject to noise and its application to secure communication. Applied Mathematics and Computation, 219(5):2698–2712, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312008946.

Chen:2013:NMS

[Che13a] Chang-Ming Chen. Numerical methods for solving a two-dimensional variable-order modified diffusion equation. Applied Mathematics and Computation, 225(??):62–78, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009326.

Chen:2013:CFE

[Che13b] Chao-Ping Chen. Continued fraction estimates for the psi function. Applied Mathematics and Computation, 219(19):9865–9871, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003962.

Chen:2013:DTT

[Che13c] Cheng-Hung Chen. Design of TSK-type fuzzy controllers using differential evolution with adaptive mutation strategy for nonlinear system control. Applied Mathematics and Computation, 219(15):8277–8294, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001483.

Chen:2013:PCZ

[Che13d] Chin-Yun Chen. A performance comparison of the zero-finding by extended interval Newton method for Peano monosplines. Applied Mathematics and Computation, 219(12):6919–6930, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012751.

Chen:2013:RID

[Che13e] Ke Chen. Recurrent implicit dynamics for online matrix inversion. Applied Mathematics and Computation, 219(20):10218– 10224, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003780.

Chen:2014:AEG

[Che14a] Chao-Ping Chen. Asymptotic expansions of the gamma function related to Windschitl’s formula. Applied Mathematics and Computation, 245(??): 174–180, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314010376.

Chen:2014:NBA

[Che14b] Chao-Ping Chen. New bounds and asymptotic expansions for the constants of Landau and Lebesgue. Applied Mathematics and Computation, 242(??):790–799, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009229.

Chen:2014:CSP

[Che14c] Fang Chen. On conditioning of saddle-point matrices with Lagrangian augmentation. Applied Mathematics and Computation, 248(??):4–7, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013125.

Chen:2014:IMN

[Che14d]  Guanwei Chen.  In-

finitely many nontrivial periodic solutions for damped vibration problems with asymptotically linear terms. Applied Mathematics and Computation, 245(??):438–446, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010923.

Chen:2014:FCC

[Che14e] Ruyun Chen. Fast computation of a class of highly oscillatory integrals. Applied Mathematics and Computation, 227(??):494–501, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031301237X.

Chen:2014:IAF

[Che14f] Ruyun Chen. On the implementation of the asymptotic Filon-type method for infinite integrals with oscillatory Bessel kernels. Applied Mathematics and Computation, 228(??):477–488, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012939.

Chen:2014:CTD

[Che14g] Weidong Chen. Computation of two-dimensional Fourier transforms for noisy band-limited signals. Applied Mathematics and Computation, 246(??):199–209, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006651.

Chen:2014:NRC

[Che14h] Wenying Chen. A note on regularity criterion for the Navier–Stokes equations in terms of the pressure. Applied Mathematics and Computation, 248(??):1–3, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012570.

Chen:2014:FAC

[Che14i] Xie-Bin Chen. A fast algorithm for computing the determinants of banded circulant matrices. Applied Mathematics and Computation, 229(??):201–207, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013313.

Chen:2014:ESM

[Che14j] Y. Z. Chen. Evaluation of the T -stress for multiple cracks in an elastic halfplane using singular integral equation and Green’s function method. Applied Mathematics and Computation, 228(??):17–30, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012630.

Chentouf:2014:BSH

[Che14k] Boumediène Chentouf. On boundary stabilization of heat exchangers systems with nonsmooth time-varying coefficients. Applied Mathematics and Computation, 233(??):398–403, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000848.

Cheng:2014:TND

[CHF14] Dongqin Cheng, Rong-Xia Hao, and Yan-Quan Feng. Two node-disjoint paths in balanced hypercubes. Applied Mathematics and Computation, 242(??):127–142, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007024.

Chiappinelli:2010:ULB

[Chi10] Raffaele Chiappinelli. Upper and lower bounds for higher order eigenvalues of some semilinear elliptic equations. Applied Mathematics and Computation, 216(12):3772–3777, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006090.

Chiang:2014:SLM

[Chi14] Chun-Yueh Chiang. On the solution of the linear matrix equation X = Af(X)B + C. Applied Mathematics and Computation, 244(??):925–935, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010182.

Chang:2010:PAR

[CHK10] Yung-Ruei Chang, Chin-Yu Huang, and Sy-Yen Kuo. Performance assessment and reliability analysis of dependable and distributed computing systems based on BDD and recursive merge. Applied Mathematics and Computation, 217(1):403–413, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000634X.

Caceres:2013:LDC

[CHM^{+}13] José Cáceres, Carmen Hernando, Mercè Mora, Ignacio M. Pelayo, and María Luz Puertas. Locatingdominating codes: Bounds and extremal cardinalities. Applied Mathematics and Computation, 220(??):38–45, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006085.

Cordero:2011:EHO

[CHMT11] Alicia Cordero, José L.

Hueso, Eulalia Martínez, and Juan R. Torregrosa. Efficient high-order methods based on golden ratio for nonlinear systems. Applied Mathematics and Computation, 217(9):4548–4556, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011185.

Cho:2011:NSL

[Cho11a] Durkbin Cho. A note on the singular linear system of the generalized finite element methods. Applied Mathematics and Computation, 217(15):6691–6699, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001056.

Cho:2011:IPA

[Cho11b]  Gyeong-Mi Cho.  An

interior-point algorithm for linear optimization based on a new barrier function. Applied Mathematics and Computation, 218(2):386–395, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007569.

Choi:2011:CSF

[Cho11c] Junesang Choi. Certain summation formulas involving harmonic numbers and generalized harmonic numbers. Applied Mathematics and Computation, 218(3):734–740, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000890.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Cabada:2014:CIS

[CHS14a] A. Cabada, A. R. Hayotov, and Kh. M. Shadimetov. Construction of D^{m}-splines in L_{2}^{(m)}(0, 1) space by Sobolev method. Applied Mathematics and Computation, 244(??):

542–551, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314009904.

Cuevas:2014:APS

[CHS14b] Claudio Cuevas, Hernán R.

Henríquez, and Herme Soto. Asymptotically periodic solutions of fractional differential equations. Applied Mathematics and Computation, 236(??):524–545, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003993.

Chung:2014:OBD

[CHT14] Kuo-Liang Chung, Yong-Huai Huang, and Shiang-Ren Tsai. Orientationbased discrete Hough transform for line detection with low computational complexity. Applied Mathematics and Computation, 237(??):430–437, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004998.

Chu:2010:FPS

[Chu10] Wenchang Chu. Fibonacci polynomials and Sylvester determinant of tridiagonal matrix. Applied Mathematics and Computation, 216(3):1018–1023, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000113X.

Chu:2011:BAF

[Chu11] Kevin T. Chu. Boosting the accuracy of finite difference schemes via optimal time step selection and non-iterative defect correction. Applied Mathematics and Computation, 218(7):3596–3614, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011490.

Cao:2013:NPS

[CHX13] Cuixia Cao, Deren Han, and Lingling Xu. A new partial splitting augmented Lagrangian method for minimizing the sum of three convex functions. Applied Mathematics and Computation, 219(10):5449–5457, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011447.

Chen:2014:NSP

[CHX14] Aiyong Chen, Wentao Huang, and Yongan Xie. Nilpotent singular points and compactons. Applied Mathematics and Computation, 236(??):300–310, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400407X.

Chung:2012:DRH

[CHYT12] Kuo-Liang Chung, Yong-Huai Huang, Wen-Ming Yan, and Wei-Chung Teng. Distortion reduction for histogram modification-based reversible data hiding. Applied Mathematics and Computation, 218(9):5819– 5826, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311013026.

Cvetkovic-Ilic:2011:NAR

[CI11] Dragana S. Cvetković-Ilić.

New additive results on Drazin inverse and its applications. Applied Mathematics and Computation, 218(7):3019–3024, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011246.

Cabada:2013:EHC

[CI13] Alberto Cabada and Antonio Iannizzotto. Existence of homoclinic constant sign solutions for a difference equation on the integers. Applied Mathematics and Computation, 224(??):216–223, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008667.

Cvetkovic-Ilic:2014:ROL

[CIN14] Dragana Cvetković-Ilić and Jovana Nikolov. Reverse order laws for {1, 2, 3}-generalized inverses. Applied Mathematics and Computation, 234(??):114–117, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314002161.

Cvetkovic-Ilic:2010:CSR

[CIP10] Dragana S. Cvetković-Ilić and Vladimir Pavlović. A comment on some recent results concerning the reverse order law for {1, 3, 4}-inverses. Applied Mathematics and Computation, 217(1):105–109, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310005321.

Cvetkovic-Ilic:2011:CSR

[CISM11] Dragana S. Cvetković-Ilić, Predrag S. Stanimirović, and Marko Miladinović. Comments on some recent results concerning {2, 3} and {2, 4}-generalized inverses. Applied Mathematics and Computation, 218(5):1512–1514, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009957.

Civicioglu:2013:BSO

[Civ13] Pinar Civicioglu. Backtracking Search Optimization Algorithm for numerical optimization problems. Applied Mathematics and Computation, 219(15):8121–8144, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001380.

Chauhan:2014:HCC

[CIVS14] Sunny Chauhan, Mohammad Imdad, Calogero Vetro, and Wutiphol Sintunavarat. Hybrid coincidence and common fixed point theorems in Menger probabilistic metric spaces under a strict contractive condition with an application. Applied Mathematics and Computation, 239(??):422–433, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006195.

Chen:2011:ESC

[CJ11a] Zhong Chen and Wei Jiang. The exact solution of a class of Volterra integral equation with weakly singular kernel. Applied Mathematics and Computation, 217(18):7515–7519, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002505.

Chen:2011:PHP

[CJ11b] Zhong Chen and Wei Jiang. Piecewise homotopy perturbation method for solving linear and nonlinear weakly singular VIE of second kind. Applied Mathematics and Computation, 217(19):7790–7798, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003092.

Cermak:2014:SSL

[CJ14] Jan Cermák and Jirí Jánský. Stability switches in linear delay difference equations. Applied Mathematics and Computation, 243(??):755–766, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009084.

Chakraborty:2012:GDB

[CJK12] Kunal Chakraborty, Soovoojeet Jana, and T. K. Kar. Global dynamics and bifurcation in a stage structured prey-predator fishery model with harvesting. Applied Mathematics and Computation, 218(18):9271–9290, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002378.

Chehade:2014:EUR

[CJM14] H. Hajj Chehade, M. Jazar, and A. Mourad. Existence and uniqueness of radial solutions for fresh/salt water interface equation in a confined aquifer. Applied Mathematics and Computation, 244(??):101–117, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008807.

Cortes:2012:SIT

[CJRV12] J.-C. Cortés, L. Jódar, M.- D. Roselló, and L. Villafuerte. Solving initial and two-point boundary value linear random differential equations: a mean square approach. Applied Mathematics and Computation, 219(4):2204–2211, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008594.

Cao:2012:NPS

[CJZ12] Yang Cao, Mei-Qun Jiang, and Ying-Long Zheng. A note on the positive stable block triangular preconditioner for generalized saddle point problems. Applied Mathematics and Computation, 218(22):11075–11082, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200478X.

Cattani:2010:HWM

[CK10] C. Cattani and A. Kudreyko. Harmonic wavelet method towards solution of the Fredholm type integral equations of the second kind. Applied Mathematics and Computation, 215(12):4164–4171, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010844.

Choi:2012:ACF

[CK12a] Jin Ho Choi and Young Ho Kim. Associated curves of a Frenet curve and their applications. Applied Mathematics and Computation, 218(18):9116–9124, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002044.

Cvetkovic:2012:AGD

[CK12b] Lj. Cvetkovíc and V. Kostíc. Application of generalized diagonal dominance in wireless sensor network optimization problems. Applied Mathematics and Computation, 218(9):4798–4805, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012537.

Czernous:2012:CBS

[CK12c] W. Czernous and Z. Kamont. Comparison between some explicit and implicit difference schemes for Hamilton Jacobi functional differential equations. Applied Mathematics and Computation, 218(17):8758–8772, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312001610.

Cranston:2013:BPS

[CK13a] Daniel W. Cranston and Candace M. Kent. On the boundedness of positive solutions of the reciprocal max-type difference equation x_{n} = max{A_{n−1}^{1}/x_{n−1}, A_{n−2}^{2}/x_{n2}, …, A_{n−l}^{l}/x_{n−l}} with periodic parameters. Applied Mathematics and Computation, 221(??):144–151, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006796.

Crisan:2013:DSI

[CK13b] O. Crisan and S. Kanas.

Differential subordinations involving arithmetic and geometric means. Applied Mathematics and Computation, 222(??): 123–131, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313008011.

Choi:2014:NHO

[CK14a] Jongsung Choi and Young-Hee Kim. A note on high order Bernoulli numbers and polynomials using differential equations. Applied Mathematics and Computation, 249(??):480–486, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014490.

Choudhury:2014:URQ

[CK14b] Gautam Choudhury and Jau-Chuan Ke. An unreliable retrial queue with delaying repair and general retrial times under Bernoulli vacation schedule. Applied Mathematics and Computation, 230(??):436–450, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013982.

Cvetkovic:2012:MNB

[CKD12] Ljiljana Cvetković, Vladimir Kostić, and Ksenija Doroslovacki. Max-norm bounds for the inverse of S-Nekrasov matrices. Applied Mathematics and Computation, 218(18): 9498–9503, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S009630031200272X.

Chen:2011:EHC

[CKH11] Hon-Chan Chen, Tzu-Liang Kung, and Lih-Hsing Hsu. Embedding a Hamiltonian cycle in the crossed cube with two required vertices in the fixed positions. Applied Mathematics and Computation, 217(24):10058–10065, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006539.

Choi:2012:BCD

[CKK12] Jin Ho Choi, Tae Ho Kang, and Young Ho Kim. Bertrand curves in 3-dimensional space forms. Applied Mathematics and Computation, 219(3):1040–1046, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007084.

Cermak:2013:SRL

[CKN13] Jan Cermák, Tomás Kisela, and Ludek Nechvátal. Stability regions for linear fractional differential systems and their discretizations. Applied Mathematics and Computation, 219(12):7012–7022, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012957.

Conejo:2013:GCT

[CKP^{+}13] P. D. Conejo, E. W. Karas, L. G. Pedroso, A. A. Ribeiro, and M. Sachine. Global convergence of trust-region algorithms for convex constrained minimization without derivatives. Applied Mathematics and Computation, 220(??):324–330, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006802.

Chen:2012:HSA

[CkPqL12] Jing Chen, Quan ke Pan, and Jun qing Li. Harmony search algorithm with dynamic control parameters. Applied Mathematics and Computation, 219(2):592–604, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312006558.

Cho:2010:STT

[CKS10] Nak Eun Cho, In Hwa Kim, and H. M. Srivastava. Sandwich-type theorems for multivalent functions associated with the Srivastava–Attiya operator. Applied Mathematics and Computation, 217(2):918–928, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007216.

Cermak:2014:TDD

[CKSV14] M. Cermák, T. Kozubek, S. Sysala, and J. Valdman. A TFETI domain decomposition solver for elastoplastic problems. Applied Mathematics and Computation, 231(??):634–653, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000253.

Cai:2010:GAV

[CL10a] Li-Ming Cai and Xue-Zhi Li. Global analysis of a vector-host epidemic model with nonlinear incidences. Applied Mathematics and Computation, 217(7):3531–3541, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009756.

Chi:2010:NIN

[CL10b] Guangsheng Chi and Gongsheng Li. Numerical inversions for a nonlinear source term in the heat equation by optimal perturbation algorithm. Applied Mathematics and Computation, 216(8):2408–2416, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003401.

Coffey:2010:GHN

[CL10c] Mark W. Coffey and Nicholas Lubbers. On generalized harmonic number sums. Applied Mathematics and Computation, 217(2):689–698, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006806.

Cai:2011:PTW

[CL11a] Jingjing Cai and Bendong Lou. Periodic traveling waves of a mean curvature equation in high dimensional cylinders. Applied Mathematics and Computation, 217(22):9267–9277, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005455.

Cao:2011:NGI

[CL11b] Chongguang Cao and Jiamei Li. A note on the group inverse of some 2×2 block matrices over skew fields. Applied Mathematics and Computation, 217(24):10271–10277, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006965.

Cardoso:2011:CSI

[CL11c] João R. Cardoso and Ana F. Loureiro. On the convergence of Schröder iteration functions for p th roots of complex numbers. Applied Mathematics and Computation, 217(21):8833–8839, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004152.

Calude:2012:TUI

[CL12] Cristian S. Calude and J. P. Lewis. Is there a universal image generator? Applied Mathematics and Computation, 218(16):8151–8159, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100868X.

Chiu:2013:IRP

[CL13a] Chien-Ching Chiu and Jhuo-Ru Li. Image reconstruction of perfectly conducting objects by a cascaded method. Applied Mathematics and Computation, 219(11):6093–6105, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012490.

Chun:2013:NOE

[CL13b] Changbum Chun and Mi Young Lee. A new optimal eighth-order family of iterative methods for the solution of nonlinear equations. Applied Mathematics and Computation, 223(??):506–519, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009016.

Cao:2014:LPS

[CL14a] Zhijie Cao and Yiping Lin.

Lie point symmetries, conservation laws, and solutions of a space dependent reaction-diffusion equation. Applied Mathematics and Computation, 248(??):386–398, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013277.

Chang:2014:SGR

[CL14b] Jung-Chan Chang and Dah-Chin Luor. On some generalized retarded integral inequalities and the qualitative analysis of integral equations. Applied Mathematics and Computation, 244(??):324–334, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314009515.

Chen:2014:PEE

[CL14c] Yanping Chen and Yijie Lin. A posteriori error estimates for control problems governed by nonlinear elliptic equations in hp-FEM. Applied Mathematics and Computation, 238(??):163–176, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004159.

Corbera:2014:CCB

[CL14d] Montserrat Corbera and Jaume Llibre. Central configurations of the 4body problem with masses m_{1} = m_{2} > m_{3} = m_{4} = m > 0 and m small. Applied Mathematics and Computation, 246(??):121–147, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010789.

Cordeiro:2014:EGB

[CL14e] Gauss M. Cordeiro and Artur J. Lemonte. The exponentiated generalized Birnbaum–Saunders distribution. Applied Mathematics and Computation, 247(??):762–779, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401279X.

Chen:2012:ASN

[CLA^{+}12] J. Chen, F. Liu, V. Anh, S. Shen, Q. Liu, and C. Liao. The analytical solution and numerical solution of the fractional diffusion-wave equation with damping. Applied Mathematics and Computation, 219(4):1737–1748, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008065.

Chen:2011:NSV

[CLAT11] Chang-Ming Chen, F. Liu, V. Anh, and I. Turner. Numerical simulation for the variable-order Galilei invariant advection diffusion equation with a nonlinear source term. Applied Mathematics and Computation, 217(12):5729–5742, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012518.

Chang:2013:SFP

[CLC^{+}13] S. S. Chang, H. W. Joseph Lee, C. K. Chan, L. Wang, and L. J. Qin. Split feasibility problem for quasi-nonexpansive multivalued mappings and total asymptotically strict pseudo-contractive mapping. Applied Mathematics and Computation, 219 (20):10416–10424, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004207.

Chang:2011:NMS

[CLCL11] S. S. Chang, H. W. Joseph Lee, Chi Kin Chan, and J. A. Liu. A new method for solving a system of generalized nonlinear variational inequalities in Banach spaces. Applied Mathematics and Computation, 217(16):6830–6837, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000373.

Chang:2012:MHT

[CLCZ12] S. S. Chang, H. W. Joseph Lee, Chi Kin Chan, and W. B. Zhang. A modified halpern-type iteration algorithm for totally quasiϕ-asymptotically nonexpansive mappings with applications. Applied Mathematics and Computation, 218(11):6489–6497, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014780.

Chen:2012:PIS

[CLD12] Jing Chen, Xianling Lu, and Rui Ding. Parameter identification of systems with preload nonlinearities based on the finite impulse response model and negative gradient search. Applied Mathematics and Computation, 219(5):2498–2505, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008788.

Chen:2013:ADG

[CLF13] Yumei Chen, Yan Luo, and Minfu Feng. Analysis of a discontinuous Galerkin method for the Biot’s consolidation problem. Applied Mathematics and Computation, 219(17):9043–9056, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003652.

Chen:2010:BST

[CLL10] Aiyong Chen, Yiping Lin, and Shumin Li. Bifurcation studies on travelling wave solutions for an integrable nonlinear wave equation. Applied Mathematics and Computation, 216(2):611–622, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001074.

Chen:2014:SPB

[CLL14] Taiyong Chen, Wenbin Liu, and Jiaying Liu. Solvability of periodic boundary value problem for fractional p-Laplacian equation. Applied Mathematics and Computation, 244(??):422–431, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314009497.

Chen:2014:NSV

[CLLS14] Yiming Chen, Liqing Liu, Baofeng Li, and Yannan Sun. Numerical solution for the variable order linear cable equation with Bernstein polynomials. Applied Mathematics and Computation, 238(??):329–341, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004287.

Chun:2012:OFO

[CLND12] Changbum Chun, Mi Young Lee, Beny Neta, and Jovana Dzunić. On optimal fourth-order iterative methods free from second derivative and their dynamics. Applied Mathematics and Computation, 218(11):6427–6438, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014718.

Castillo:2012:MPA

[CLR12] K. Castillo, R. L. Lamblém, and A. Sri Ranga. On a moment problem associated with Chebyshev polynomials. Applied Mathematics and Computation, 218(18):9571–9574, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002718.

Chung:2012:CSP

[CLS12] Kun-Jen Chung, Shy-Der Lin, and H. M. Srivastava. The complete solution procedures for the mathematical analysis of some families of optimal inventory models with order-size dependent trade credit and deterministic and constant demand. Applied Mathematics and Computation, 219(1):142–157, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005747.

Chen:2013:INM

[CLTA13] S. Chen, F. Liu, I. Turner, and V. Anh. An implicit numerical method for the two-dimensional fractional percolation equation. Applied Mathematics and Computation, 219(9):4322–4331, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009903.

Chen:2013:LBN

[CLTL13] Jheng-Cheng Chen, Chia-Jui Lai, Chang-Hsiung Tsai, and Pao-Lien Lai. A lower bound on the number of Hamiltonian cycles through a prescribed edge in a crossed cube. Applied Mathematics and Computation, 219(19):9885–9892, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003901.

Chai:2013:TDE

[CLTY13] Qinqin Chai, Ryan Loxton, Kok Lay Teo, and Chunhua Yang. Timedelay estimation for nonlinear systems with piecewiseconstant input. Applied Mathematics and Computation, 219(17):9543–9560, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002683.

Chen:2011:SIE

[CLW11] Y. Z. Chen, X. Y. Lin, and Z. X. Wang. Singular integral equation method for contact problem for rigidly connected punches on elastic halfplane. Applied Mathematics and Computation, 217(12):5680–5694, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012488.

Chen:2012:CAS

[CLW12] Jein-Shan Chen, Jing-Fan Li, and Jia Wu. A continuation approach for solving binary quadratic program based on a class of NCPfunctions. Applied Mathematics and Computation, 219(8):3975–3992, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031201034X.

Chen:2014:EMN

[CLW14] Shaoxiong Chen, Jiu Liu, and Xian Wu. Existence and multiplicity of nontrivial solutions for a class of modified nonlinear fourthorder elliptic equations on R^{n}. Applied Mathematics and Computation, 248(??):593–601, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013873.

Cen:2013:FDS

[CLX13] Zhongdi Cen, Anbo Le, and Aimin Xu. Finite difference scheme with a moving mesh for pricing Asian options. Applied Mathematics and Computation, 219(16):8667–8675, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002038.

Liu:2011:NVS

[cLYrL11] Jian chang Liu, Xia Yu, and Hong ru Li. A nonparametric variable stepsize NLMS algorithm for transversal filters. Applied Mathematics and Computation, 217(17):7365–7371, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002025.

Chen:2011:FMS

[CLZ11] Zhongying Chen, Jianfei Li, and Yongdong Zhang. A fast multiscale solver for modified Hammerstein equations. Applied Mathematics and Computation, 218(7):3057–3067, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010484.

Cao:2011:FOM

[CLZmF11] Huai-Huo Cao, Li-Bin Liu, Yong Zhang, and Sheng mao Fu. A fourth-order method of the convectiondiffusion equations with Neumann boundary conditions. Applied Mathematics and Computation, 217(22):9133–9141, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005212.

Comellas:2010:FEA

[CM10] Francesc Comellas and Alicia Miralles. A fast and efficient algorithm to identify clusters in networks. Applied Mathematics and Computation, 217(5):2007–2014, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007459.

Cheon:2011:RCS

[CM11] Gi-Sang Cheon and Toufik Mansour. Rational combinations for the sums involving inverse binomial coefficients. Applied Mathematics and Computation, 218(6):2641–2646, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010290.

Chen:2012:DNS

[CM12a] Guanwei Chen and Shiwang Ma. Discrete nonlinear Schrödinger equations with superlinear nonlinearities. Applied Mathematics and Computation, 218(9):5496–5507, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013683.

Chen:2012:PSS

[CM12b] Ruipeng Chen and Ruyun Ma. Positive solutions of the second-order differential systems in reactor dynamics. Applied Mathematics and Computation, 219(8):3882–3892, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010107.

Cimrak:2012:MTR

[CM12c] Ivan Cimrák and Valdemar Melicher. Mixed Tikhonov regularization in Banach spaces based on domain decomposition. Applied Mathematics and Computation, 218(23):11583–11596, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005553.

Cvetkovic:2012:GIQ

[CM12d] Aleksandar S. Cvetković and Gradimir V. Milovanović. Gaussian interval quadrature rule for exponential weights. Applied Mathematics and Computation, 218(18):9332–9341, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312002482.

Celik:2013:HBA

[ÇM13a] C. Çelik and H. Merdan.

Hopf bifurcation analysis of a system of coupled delayed-differential equations. Applied Mathematics and Computation, 219(12):6605–6617, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000040.

Cevik:2013:NDB

[ÇM13b] Ahmet Sinan Çevik and Ayse Dilek Maden. New distance-based graph invariants and relations among them. Applied Mathematics and Computation, 219(24):11171– 11177, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313005444.

Chen:2013:LIA

[CM13c] Chao-Ping Chen and Cristinel Mortici. Limits and inequalities associated with the Euler– Mascheroni constant. Applied Mathematics and Computation, 219(18):

9755–9761, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003500.

Chen:2014:HSD

[CM14] Guanwei Chen and Shiwang Ma. Homoclinic solutions of discrete nonlinear Schrödinger equations with asymptotically or super linear terms. Applied Mathematics and Computation, 232(??):787–798, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002008.

Chen:2013:CCT

[CMD13] Xiao-Diao Chen, Weiyin Ma, and Chongyang Deng. Conditions for the coincidence of two quartic Bézier curves. Applied Mathematics and Computation, 225(??):731–736, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010321.

Cardenas-Morales:2014:TFB

[CMG14] Daniel Cárdenas-Morales and Vijay Gupta. Two families of Bernstein– Durrmeyer type operators. Applied Mathematics and Computation, 248(??):342–353, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013289.

Cardenas-Morales:2013:OPI

[CMGR13] D. Cárdenas-Morales, P. Garrancho, and I. Rasa. Optimality of piecewise τlinear interpolating operators. Applied Mathematics and Computation, 219(12):6445–6448, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000118.

Cardenas-Morales:2014:APB

[CMGR14] D. Cárdenas-Morales, P. Garrancho, and I. Rasa. Approximation properties of Bernstein–Durrmeyer type operators. Applied Mathematics and Computation, 232(??):1–8, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314000836.

Chen:2012:PPS

[CMH12] Ruipeng Chen, Ruyun Ma, and Zhiqian He. Positive periodic solutions of first-order singular systems. Applied Mathematics and Computation, 218(23):11421–11428, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005164.

Calio:2010:DIM

[CMM10] F. Caliò, M. V. Fernández Muñoz, and E. Marchetti. Direct and iterative methods for the numerical solution of mixed integral equations. Applied Mathematics and Computation, 216(12):3739–3746, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000576X.

Colao:2012:SAM

[CMM12] Vittorio Colao, Giuseppe Marino, and Luigi Muglia. On some auxiliary mappings generated by nonexpansive and strictly pseudo-contractive mappings. Applied Mathematics and Computation, 218(11):6232–6241, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014160.

Changdar:2013:ACO

[CMP13] C. Changdar, G. S. Mahapatra, and R. K. Pal. An ant colony optimization approach for binary knapsack problem under fuzziness. Applied Mathematics and Computation, 223(??):243–253, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008278.

Chugh:2011:PSD

[CMS11] Shamta Chugh, Dinesh Kumar Madan, and Kuldip Singh. Plane strain deformation of an initially unstressed elastic medium. Applied Mathematics and Computation, 217(21):8683–8692, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311004802.

Chiarella:2013:IOU

[CMS13] Carl Chiarella, Akio Matsumoto, and Ferenc Szidarovszky. Isoelastic oligopolies under uncertainty. Applied Mathematics and Computation, 219 (21):10475–10486, July 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004645.

Carr:2011:ESU

[CMT11] E. J. Carr, T. J. Moroney, and I. W. Turner. Efficient simulation of unsaturated flow using exponential time integration. Applied Mathematics and Computation, 217(14):6587–6596, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000579.

Cianciaruso:2010:WSC

[CMW10]  Filomena  Cianciaruso,

Giuseppe Marino, and Xuewu Wang. Weak and strong convergence of the Ishikawa iterative process for a finite family of asymptotically nonexpansive mappings. Applied Mathematics and Computation, 216(12):3558–3567, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005187.

Cao:2014:HNS

[CMX14] Yang Cao, Lin Mao, and Xun-Qian Xu. Hermitian and normal splitting methods for non-Hermitian positive definite linear systems. Applied Mathematics and Computation, 243(??):690–702, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314008674.

Cvetkovic:2012:ELR

[CN12] Ljiljana Cvetković and Maja Nedović. Eigenvalue localization refinements for the Schur complement. Applied Mathematics and Computation, 218(17):8341–8346, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200104X.

Cavalli:2014:CDG

[CN14a] Fausto Cavalli and Ahmad Naimzada. A Cournot duopoly game with heterogeneous players: Nonlinear dynamics of the gradient rule versus local monopolistic approach. Applied Mathematics and Computation, 249(??):382–388, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013976.

Chien:2014:RJN

[CN14b] M. T. Chien and H. Nakazato. Reduction of joint cnumerical ranges. Applied Mathematics and Computation, 232(??):178–182, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001040.

Chun:2014:ANF

[CN14c] Changbum Chun and Beny Neta. An analysis of a new family of eighth-order optimal methods. Applied Mathematics and Computation, 245(??):86–107, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401025X.

Chun:2014:CWF

[CNKS14] Changbum Chun, Beny Neta, Jeremy Kozdon, and Melvin Scott. Choosing weight functions in iterative methods for simple roots. Applied Mathematics and Computation, 227(??):788–800, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012538.

Caratelli:2010:NPH

[CNRY10] D. Caratelli, P. Natalini, P. E. Ricci, and A. Yarovoy. The Neumann problem for the Helmholtz equation in a starlike planar domain. Applied Mathematics and Computation, 216(2):556–564, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001013.

Calidonna:2012:CIS

[CNTG12] C. R. Calidonna, A. Naddeo, G. A. Trunfio, and S. Di Gregorio. From classical infinite space– time CA to a hybrid CA model for natural sciences modeling. Applied Mathematics and Computation, 218(16):8137–8150, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009386.

Cregan:2013:EAS

[CO13] Vincent Cregan and Stephen B. G. O’Brien. Extended asymptotic solutions to the spin-coating model with small evaporation. Applied Mathematics and Computation, 223(??):76–87, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008217.

Constantin:2013:RRN

[Con13] Bogdan Constantin. A remark on the radial and nonradial solutions for an elliptic system with a quadratic gradient term. Applied Mathematics and Computation, 219(15):8302–8310, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001707.

Copetti:2011:CPG

[Cop11] M. I. M. Copetti. A contact problem in generalized thermoelasticity. Applied Mathematics and Computation, 218 (5):2128–2145, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009635.

Copetti:2014:PFE

[Cop14] M. I. M. Copetti. A penalty-finite element approximation to a Signorini two-body contact problem in thermoelasticity. Applied Mathematics and Computation, 234(??):293–308, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002896.

Cordeiro:2014:LGD

[COPP14] Gauss M. Cordeiro, Edwin M. M. Ortega, Bozidar V. Popović, and Rodrigo R. Pescim. The Lomax generator of distributions: Properties, minification process and regression model. Applied Mathematics and Computation, 247(??):465–486, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401220X.

Chen:2010:SBT

[COPR10] Anthony Chen, Jun-Seok Oh, Dongjoo Park, and Will Recker. Solving the bicriteria traffic equilibrium problem with variable demand and nonlinear path costs. Applied Mathematics and Computation, 217(7):3020–3031, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310008891.

Cortell:2010:CBV

[Cor10] Rafael Cortell. On a certain boundary value problem arising in shrinking sheet flows. Applied Mathematics and Computation, 217(8):4086–4093, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010556.

Cornelio:2011:RNL

[Cor11a] Anastasia Cornelio. Regularized nonlinear least squares methods for hit position reconstruction in small gamma cameras. Applied Mathematics and Computation, 217(12):5589–5595, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012373.

Cortell:2011:HFF

[Cor11b] Rafael Cortell. Heat and fluid flow due to nonlinearly stretching surfaces. Applied Mathematics and Computation, 217(19):7564–7572, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002050.

Covei:2011:ERQ

[Cov11] Dragos-Patru Covei. Existence results for a quasilinear elliptic problem with a gradient term via shooting method. Applied Mathematics and Computation, 218(8):4161–4168, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012197.

Cimpean:2010:SLD

[CP10] Dalia Sabina Cîmpean and Dorian Popa. On the stability of the linear differential equation of higher order with constant coefficients. Applied Mathematics and Computation, 217(8):4141–4146, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010660.

Carvajal:2012:UDS

[CP12a] Xavier Carvajal and Mahendra Panthee. On uniqueness and decay of solution for Hirota equation. Applied Mathematics and Computation, 218(9):4928–4943, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101304X.

Constantine:2012:LMA

[CP12b] Paul G. Constantine and Eric T. Phipps. A Lanczos method for approximating composite functions. Applied Mathematics and Computation, 218(24):11751–11762, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004948.

Cordone:2012:SQM

[CP12c] Roberto Cordone and Gianluca Passeri. Solving the Quadratic Minimum Spanning Tree Problem. Applied Mathematics and Computation, 218(23):11597–11612, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005577.

Choudhuri:2014:SAU

[CP14a] D. Choudhuri and B. Sri Padmavati. A study of an arbitrary unsteady Stokes flow in and around a liquid sphere. Applied Mathematics and Computation, 243(??):644–656, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007346.

Christou:2014:KSB

[CP14b] M. A. Christou and N. C. Papanicolaou. Kawahara solitons in Boussinesq equations using a robust Christov–Galerkin spectral method. Applied Mathematics and Computation, 243(??):245–257, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007632.

Chatterjee:2011:BTE

[CPC11] A. Chatterjee, S. Pal, and S. Chatterjee. Bottom up and top down effect on toxin producing phytoplankton and its consequence on the formation of plankton bloom. Applied Mathematics and Computation, 218(7):3387–3398, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011234.

Castillo-Perez:2013:SPP

[CPKT13] Raúl Castillo-Pérez, Vladislav V.

Kravchenko, and Sergii M. Torba. Spectral parameter power series for perturbed Bessel equations. Applied Mathematics and Computation, 220(??):676–694, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007856.

Cunha:2011:CPA

[CPOdCN11] F. G. M. Cunha, A. W. M.

Pinto, P. R. Oliveira, and J. X. da Cruz Neto. A class of primal affine scaling algorithms. Applied Mathematics and Computation, 218(8):4523–4532, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012793.

Cao:2014:DTR

[CPS14] Hui Cao, Sergei V.

Pereverzev, and Eva Sincich. Discretized Tikhonov regularization for Robin boundaries localization. Applied Mathematics and Computation, 226(??):374–385, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031301093X.

Chen:2010:LIC

[CPZ10] Yabo Chen, Zhenyun Peng, and Tiejun Zhou. LSQR iterative common symmetric solutions to matrix equations AXB = E and CXD = F. Applied Mathematics and Computation, 217(1):230–236, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006120.

Cho:2014:SCI

[CQ14] Sun Young Cho and Xiaolong Qin. On the strong convergence of an iterative process for asymptotically strict pseudocontractions and equilibrium problems. Applied Mathematics and Computation, 235(??):430–438, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400366X.

Cai:2010:ETW

[CQJ10] Jionghui Cai, Wen Qiu, and Pizhu Jia. Exact traveling wave solutions for a modified Camassa–Holm equation. Applied Mathematics and Computation, 217(2):607–611, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006697.

Chandra:2011:SFK

[CR11a] Rohitash Chandra and Luc Rolland. On solving the forward kinematics of 3RPR planar parallel manipulator using hybrid metaheuristics. Applied Mathematics and Computation, 217(22):8997–9008, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004747.

Chehab:2011:IPS

[CR11b] Jean-Paul Chehab and Marcos Raydan. An implicit preconditioning strategy for large-scale generalized Sylvester equations. Applied Mathematics and Computation, 217(21):8793–8803, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005285.

Correia:2012:HSP

[CR12] Marcos J. Correia and Paulo C. Rech. Hyperchaotic states in the parameter-space. Applied Mathematics and Computation, 218(12):6711–6715, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015098.

Cvetkovic:2014:QCP

[CR14] Marija Cvetković and Vladimir Rakocević. Quasicontraction of Perov type. Applied Mathematics and Computation, 237(??): 712–722, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400304X.

Cravo:2012:MPE

[Cra12] Glória Cravo. Matrices with prescribed eigenvalues and prescribed submatrices II. Applied Mathematics and Computation, 219(1):180–187, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005784.

Chaudhuri:2013:PZD

[CRC13] Sanjay Chaudhuri, Shovonlal Roy, and J. Chattopadhyay. Phytoplanktonzooplankton dynamics in the ‘presence’ or ‘absence’ of toxic phytoplankton. Applied Mathematics and Computation, 225(??):102–116, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009508.

Cruceanu:2014:ANC

[CRC14] Stefan Gicu Cruceanu, Eleonora Rapeanu, and Adrian Carabineanu. Arclength numerical continuation in free-boundary flow. Applied Mathematics and Computation, 230(??):138–153, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013702.

Caratelli:2011:RPL

[CRG11] D. Caratelli, P. E. Ricci, and J. Gielis. The Robin problem for the Laplace equation in a three-dimensional starlike domain. Applied Mathematics and Computation, 218(3):713–719, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005261.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Choudhuri:2013:SFP

[CRP13] D. Choudhuri, R. Radha, and B. Sri Padmavati. Stokes flow past an arbitrary shaped body with slip-stick boundary conditions. Applied Mathematics and Computation, 219(10):5367–5375, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011538.

Camenga:2013:NRS

[CRR^{+}13] Kristin A. Camenga, Patrick X. Rault, Daniel J. Rossi, Tsvetanka Sendova, and Ilya M. Spitkovsky. Numerical range of some doubly stochastic matrices. Applied Mathematics and Computation, 221(??):40–47, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006231.

Cousins:2011:EEH

[CRW11] Benjamin R. Cousins, Leo G. Rebholz, and Nicholas E. Wilson. Enforcing energy, helicity and strong mass conservation in finite element computations for incompressible Navier–Stokes simulations. Applied Mathematics and Computation, 218(4):1208–1221, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007946.

Cheng:2012:TSP

[CRX12] Aijie Cheng, Yongqiang Ren, and Kaihua Xi. A time-stepping procedure based on convolution for the mixed finite element approximation for porous media flow. Applied Mathematics and Computation, 218(9):5319–5326, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013452.

Chen:2010:AVM

[CS10a] Haibo Chen and Juntao Sun. An application of variational method to second-order impulsive differential equation on the half-line. Applied Mathematics and Computation, 217(5):1863–1869, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007253.

Chen:2010:PIO

[CS10b] Ke Chen and Kamal Shanazari. A parallel implementation of overlapping Schwarz method for the dual reciprocity method. Applied Mathematics and Computation, 216(11):3355–3364, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004911.

Cholamjiak:2010:IME

[CS10c] Prasit Cholamjiak and Suthep Suantai. An iterative method for equilibrium problems and a finite family of relatively nonexpansive mappings in a Banach space. Applied Mathematics and Computation, 217(8):3825–3831, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010015.

Cichon:2010:SVS

[CS10d] Mieczysław Cichoń and Hussein A. H. Salem. Setvalued system of fractional differential equations with hysteresis. Applied Mathematics and Computation, 215(11):3824–3829, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010121.

Caballero:2011:SIS

[CS11a] J. Caballero and K. Sadarangani.

Sandor’s inequality for Sugeno integrals. Applied Mathematics and Computation, 218(5):1617–1622, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311008745.

Chen:2011:GSD

[CS11b] Hao Chen and Jitao Sun. Global stability of delay multigroup epidemic models with group mixing and nonlinear incidence rates. Applied Mathematics and Computation, 218(8):4391–4400, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012458.

Choi:2011:AFT

[CS11c] Junesang Choi and H. M. Srivastava. Asymptotic formulas for the triple Gamma function Γ_{3} by means of its integral representation. Applied Mathematics and Computation, 218(6):2631–2640, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010289.

Christou:2011:NSR

[CS11d] M. A. Christou and C. Sophocleous. Numerical similarity reductions of the (1+3)-dimensional Burgers equation. Applied Mathematics and Computation, 217(18):7455–7461, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002323.

Cabada:2012:FFB

[CS12a] Alberto Cabada and Svatoslav Stanek. Functional fractional boundary value problems with singular ϕ-Laplacian. Applied Mathematics and Computation, 219(4):1383–1390, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007928.

Cloete:2012:APS

[CS12b] M. Cloete and G. J. F.

Smit. An analytical pore-scale, shear stress model for purely viscous non-Newtonian fluids traversing porous media. Applied Mathematics and Computation, 219(7):3367–3384, December 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011003.

Caballero:2013:MTI

[CS13a] J. Caballero and K. Sadarangani.

A markov-type inequality for seminormed fuzzy integrals. Applied Mathematics and Computation, 219(22):10746–10752, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004864.

Cao:2013:SGF

[CS13b] Jian Cao and H. M. Srivastava. Some q-generating functions of the Carlitz and Srivastava–Agarwal types associated with the generalized Hahn polynomials and the generalized Rogers–Szegő polynomials. Applied Mathematics and Computation, 219(15):8398–8406, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001161.

Chen:2013:GSSb

[CS13c] Hao Chen and Jitao Sun. Global stability and stabilization of switched Boolean network with impulsive effects. Applied Mathematics and Computation, 224(??):625–634, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009442.

Chen:2013:CCD

[CS13d] Xiaoming Chen and Keran Shao. Complex coefficients domain decomposition method for heterogeneous medium time-harmonic problem. Applied Mathematics and Computation, 219(17):9331–9339, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002737.

Chidume:2013:IAS

[CS13e] C. E. Chidume and Y. Shehu. Iterative approximation of solutions of equations of Hammerstein type in certain Banach spaces. Applied Mathematics and Computation, 219(10):5657–5667, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012088.

Choi:2013:STS

[CS13f] Junesang Choi and H. M.

Srivastava. Some twosided inequalities for multiple Gamma functions and related results. Applied Mathematics and Computation, 219(20): 10343–10354, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313004050.

Chen:2014:EES

[CS14a] Yanping Chen and Chunmei Sun. Error estimates and superconvergence of mixed finite element methods for fourth order hyperbolic control problems. Applied Mathematics and Computation, 244(??):642–653, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008583.

Coelho:2014:TWP

[CS14b] Isabel Coelho and Luís Sanchez. Travelling wave profiles in some models with nonlinear diffusion. Applied Mathematics and Computation, 235(??):469–481, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400352X.

Ciric:2011:CFPb

[ĆSAV11] Ljubomir Ćirić, Bessem Samet, Hassen Aydi, and Calogero Vetro. Common fixed points of generalized contractions on partial metric spaces and an application. Applied Mathematics and Computation, 218(6):2398–2406, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009246.

Chang:2010:MNM

[CSC10] Jingya Chang, Wenyu Sun, and Yannan Chen. A modified Newton’s method for best rankone approximation to tensors. Applied Mathematics and Computation, 216(6):1859–1867, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010662.

Chen:2012:SAB

[CSD12] Shi-Ming Chen, Ali Sarosh, and Yun-Feng Dong. Simulated annealing based artificial bee colony algorithm for global numerical optimization. Applied Mathematics and Computation, 219(8):3575–3589, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009514.

Costa:2012:HGP

[CSF12] Lino Costa, Isabel A. C. P. Espírito Santo, and Edite M. G. P. Fernandes. A hybrid genetic pattern search augmented Lagrangian method for constrained global optimization. Applied Mathematics and Computation, 218(18):9415–9426, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002573.

Chen:2014:NSF

[CSL14] Yiming Chen, Yannan Sun, and Liqing Liu. Numerical solution of fractional partial differential equations with variable coefficients using generalized fractional-order Legendre functions. Applied Mathematics and Computation, 244(??):847–858, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010078.

Chen:2010:ISB

[CSLL10] Jeng-Tzong Chen, Hung-Chih Shieh, Ying-Te Lee, and Jia-Wei Lee. Image solutions for boundary value problems without sources. Applied Mathematics and Computation, 216(5):1453–1468, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002201.

Cvetkovic:2012:ABO

[CSMT12] Aleksandar S. Cvetković, Marija P. Stanić, Zvezdan M. Marjanović, and Tatjana V. Tomović. Asymptotic behavior of orthogonal trigonometric polynomials of semiinteger degree. Applied Mathematics and Computation, 218(23):11528– 11533, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312005449.

Cai:2011:LRP

[CSQ11] Xinghui Cai, G. H. Su, and Suizheng Qiu. Local radial point interpolation method for the fully developed magnetohydrodynamic flow. Applied Mathematics and Computation, 217(9):4529–4539, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011161.

Cuevas:2011:APP

[CSS11] Claudio Cuevas, Alex Sepúlveda, and Herme Soto. Almost periodic and pseudo-almost periodic solutions to fractional differential and integro-differential equations. Applied Mathematics and Computation, 218(5):1735–1745, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008873.

Cichon:2011:IND

[CSSN11] Mieczysław Cichoń, Bianca Satco, and Aneta Sikorska-Nowak. Impulsive nonlocal differential equations through differential equations on time scales. Applied Mathematics and Computation, 218(6):2449–2458, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010058.

Cortes:2013:NPP

[CSSSV13] J.-C. Cortés, A. Sánchez-Sánchez, F.-J. Santonja, and R.-J. Villanueva. Nonparametric probabilistic forecasting of academic performance in Spanish high school using an epidemiological modelling approach. Applied Mathematics and Computation, 221(??):648–661, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007091.

Cabada:2012:HSS

[CST12] Alberto Cabada, Diko Souroujon, and Stepan Tersian. Heteroclinic solutions of a second-order difference equation related to the Fisher–Kolmogorov’s equation. Applied Mathematics and Computation, 218(18):9442–9450, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002640.

Cordero:2014:DAI

[CST14] Alicia Cordero, Fazlollah Soleymani, and Juan R. Torregrosa. Dynamical analysis of iterative methods for nonlinear systems or how to deal with the dimension? Applied Mathematics and Computation, 244(??):398–412, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009679.

Chen:2012:MSM

[CSZ12] Yaming Chen, Songhe Song, and Huajun Zhu. Multi-symplectic methods for the Ito-type coupled KdV equation. Applied Mathematics and Computation, 218(9):5552–5561, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013889.

Cakmak:2010:LTIb

[ÇT10a] Devrim Çakmak and Aydın Tiryaki. On Lyapunov-type inequality for quasilinear systems. Applied Mathematics and Computation, 216(12):3584–3591, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005217.

Canak:2010:CUW

[ÇT10b] İbrahim Çanak and Ümit Totur. A condition under which slow oscillation of a sequence follows from Cesàro summability of its generator sequence. Applied Mathematics and Computation, 216(5):1618–1623, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002596.

Chen:2011:EIM

[CT11a] Peng Chen and X. H.

Tang. Existence of infinitely many homoclinic orbits for fourth-order difference systems containing both advance and retardation. Applied Mathematics and Computation, 217(9):4408–4415, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010726.

Chu:2011:FCF

[CT11b] Wenchang Chu and Peipei Tang. Five convolution formulae of orthogonal polynomials. Applied Mathematics and Computation, 217(21):8825–8832, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003833.

Cordero:2011:CST

[CT11c] Alicia Cordero and Juan R. Torregrosa. A class of Steffensen type methods with optimal order of convergence. Applied Mathematics and Computation, 217(19):7653–7659, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002748.

Case:2012:IEM

[CT12a] Elizabeth Case and Johannes Tausch. An integral equation method for spherical Stefan problems. Applied Mathematics and Computation, 218(23):11451–11460, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005188.

Chen:2012:EMS

[CT12b] Peng Chen and Xianhua Tang. Existence and multiplicity of solutions for second-order impulsive differential equations with Dirichlet problems. Applied Mathematics and Computation, 218(24):11775–11789, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005127.

Chen:2012:SSF

[CT12c] Yi Chen and Xianhua Tang. Solvability of sequential fractional order multi-point boundary value problems at resonance. Applied Mathematics and Computation, 218(14):7638–7648, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000690.

Cabada:2013:MST

[CT13a] Alberto Cabada and Stepan Tersian. Multiplicity of solutions of a two point boundary value problem for a fourth-order equation. Applied Mathematics and Computation, 219(10):5261–5267, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012337.

Chung:2013:NNH

[CT13b] N. T. Chung and H. Q. Toan. On a nonlinear and non-homogeneous problem without (A–R) type condition in Orlicz– Sobolev spaces. Applied Mathematics and Computation, 219(14):7820–7829, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001276.

Costin:2013:OES

[CT13c] Iulia Costin and Gheorghe Toader. Optimal evaluations of some seifferttype means by power means. Applied Mathematics and Computation, 219(9):4745–4754, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011071.

Chebotarev:2014:SVD

[CT14a] Alexander M. Chebotarev and Alexander E. Teretenkov. Singular value decomposition for the Takagi factorization of symmetric matrices. Applied Mathematics and Computation, 234(??):380–384, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002239.

Chen:2014:IMS

[CT14b] Peng Chen and Cai Tian.

Infinitely many solutions for Schrödinger–Maxwell equations with indefinite sign subquadratic potentials. Applied Mathematics and Computation, 226(??):492–502, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313011260.

Chen:2013:FHS

[CTA13] Peng Chen, Xianhua Tang, and Ravi P. Agarwal. Fast homoclinic solutions for a class of damped vibration problems. Applied Mathematics and Computation, 219(11):6053–6065, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011332.

Cho:2014:UDV

[CTH14] Hsun-Jung Cho, Ming-Te Tseng, and Ming-Chorng Hwang. Using detection of vehicular presence to estimate shockwave speed and upstream traffics for a signalized intersection. Applied Mathematics and Computation, 232(??):1151–1165, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000198.

Cordier:2012:LTB

[CTLD12] Stephane Cordier, Le Xuan Truong, Nguyen Thanh Long, and Alain Pham Ngoc Dinh. Large time behavior of differential equations with drifted periodic coefficients modeling carbon storage in soil. Applied Mathematics and Computation, 218(9):5641–5654, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014019.

Chen:2011:SCM

[CTLL11] Jeng-Tzong Chen, Jhen-Jyun Tsai, Ying-Te Lee, and Jia-Wei Lee. Study on connections of the MFS, Trefftz method, indirect BIEM and invariant MFS in the threedimensional Laplace problems containing spherical boundaries. Applied Mathematics and Computation, 218(8):4056–4074, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101201X.

Cleja-Tigoiu:2014:RSA

[CTS14] Sanda Cleja-Tigoiu and Nadia Elena Stoicuta. Revisited Simo algorithm for the plane stress state. Applied Mathematics and Computation, 237(??):

730–751, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004974.

Cordero:2012:PTD

[CTV12] Alicia Cordero, Juan R. Torregrosa, and María P. Vassileva. Pseudocomposition: a technique to design predictor-corrector methods for systems of nonlinear equations. Applied Mathematics and Computation, 218(23):11496– 11504, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005413.

Cordero:2013:DFC

[CTV13] Alicia Cordero, Juan R. Torregrosa, and Pura Vindel. Dynamics of a family of Chebyshev–Halley type methods. Applied Mathematics and Computation, 219(16):8568–8583, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001732.

Chuang:2011:CCP

[CTY11] Li-Yeh Chuang, Sheng-Wei Tsai, and Cheng-Hong Yang. Chaotic catfish particle swarm optimization for solving global numerical optimization problems. Applied Mathematics and Computation, 217(16):6900–6916, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001160.

Cui:2014:CCD

[Cui14] Mingrong Cui. Combined compact difference scheme for the time fractional convection-diffusion equation with variable coefficients. Applied Mathematics and Computation, 246(??):464–473, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011199.

Curt:2014:SRC

[Cur14] Paula Curt. Some remarks on certain invariant geometric properties in Hele–Shaw flows. Applied Mathematics and Computation, 236(??):384–390, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004032.

Chung:2011:NEP

[CV11] Bong Jae Chung and Ashwin Vaidya. Nonequilibrium pattern selection in particle sedimentation. Applied Mathematics and Computation, 218(7):3451–3465, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011313.

Cammaroto:2012:MSN

[CV12] F. Cammaroto and L. Vilasi. Multiple solutions for a nonhomogeneous Dirichlet problem in Orlicz–Sobolev spaces. Applied Mathematics and Computation, 218(23):11518–11527, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005437.

Chen:2013:GCH

[CV13] Bang-Yen Chen and Gabriel Eduard Vîlcu. Geometric classifications of homogeneous production functions. Applied Mathematics and Computation, 225(??):345–351, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301031X.

Cveticanin:2014:TNO

[Cve14] L. Cveticanin. On the truly nonlinear oscillator with positive and negative damping. Applied Mathematics and Computation, 243(??):433–445, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008455.

Cvijovic:2010:DSC

[Cvi10] Djurdje Cvijović. The Dattoli–Srivastava conjectures concerning generating functions involving the harmonic numbers. Applied Mathematics and Computation, 215(11):4040–4043, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300309010583.

Cvijovic:2011:SFF

[Cvi11] Djurdje Cvijović. Summation formulae for finite tangent and secant sums. Applied Mathematics and Computation, 218(3):741–745, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001147.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Cvijovic:2012:ADF

[Cvi12] Djurdje Cvijović. Another discrete Fourier transform pairs associated with the Lipschitz–Lerch zeta function. Applied Mathematics and Computation, 218(12):6744–6747, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015153.

Cho:2010:DMS

[CVP^{+}10] Jae W. Cho, S. Valette, Ju H. Park, Ho Y. Jung, and R. Prost. 3-D mesh sequence compression using wavelet-based multi-resolution analysis. Applied Mathematics and Computation, 216(2):410–425, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000457.

Clason:2010:BOP

[CvW10] Christian Clason and Gregory von Winckel. On a bilinear optimization problem in parallel magnetic resonance imaging. Applied Mathematics and Computation, 216(5):1443–1452, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002195.

Chang:2011:DMC

[CW11a] Ying-Hua Chang and Tz-Ting Wu. Dynamic multicriteria evaluation of coevolution strategies for solving stock trading problems. Applied Mathematics and Computation, 218(8):4075–4089, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311012033.

Chen:2011:FHO

[CW11b] Haibin Chen and Yiju Wang. A family of higherorder convergent iterative methods for computing the Moore–Penrose inverse. Applied Mathematics and Computation, 218(8):4012–4016, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007478.

Cheng:2011:RPA

[CW11c] Jianhua Cheng and Dehui Wang. Ruin problems for an autoregressive risk model with dependent rates of interest. Applied Mathematics and Computation, 218(7):3822–3833, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012008.

Chen:2012:BIE

[CW12] Y. Z. Chen and Z. X. Wang. Boundary integral equation method for periodic dissimilar elastic inclusions in an infinite plate. Applied Mathematics and Computation, 218(17):8578–8591, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001439.

Chen:2014:NRE

[CW14a] Huabin Chen and Liu Wang. New result on exponential stability for neutral stochastic linear system with timevarying delay. Applied Mathematics and Computation, 239(??):320–325, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005876.

Chen:2014:GAM

[CW14b] Shanshan Chen and Junjie Wei. Global attractivity in a model of genetic regulatory system with delay. Applied Mathematics and Computation, 232(??):411–415, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400071X.

Cheng:2011:TAS

[CWCW11] T. C. E. Cheng, Wen-Hsiang Wu, Shuenn-Ren Cheng, and Chin-Chia Wu. Two-agent scheduling with position-based deteriorating jobs and learning effects. Applied Mathematics and Computation, 217(21):8804–8824, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005467.

Chang:2012:DPC

[CWL^{+}12] S. S. Chang, L. Wang, H. W. Joseph Lee, C. K. Chan, and L. Yang. Demiclosed principle and Δconvergence theorems for total asymptotically nonexpansive mappings in CAT(0) spaces. Applied Mathematics and Computation, 219(5):2611–2617, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008880.

Chen:2013:GSSa

[CWLC13] Fengde Chen, Haina Wang, Yuhua Lin, and Wanlin Chen. Global stability of a stagestructured predator-prey system. Applied Mathematics and Computation, 223(??):45–53, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008382.

Chang:2010:CSP

[CWS10] Hai-Xia Chang, Qing-Wen Wang, and Guang-Jing Song. (RS)-conjugate solution to a pair of linear matrix equations. Applied Mathematics and Computation, 217(1):73– 82, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310004819.

Crowell:2014:ENS

[CWW14] Sean Crowell, Luther White, and Louis Wicker. Estimation of near surface wind speeds in strongly rotating flows. Applied Mathematics and Computation, 235(??):201–211, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400040X.

Chang:2014:SCT

[CWWC14] S. S. Chang, L. Wang, X. R. Wang, and C. K. Chan. Strong convergence theorems for Bregman totally quasiasymptotically nonexpansive mappings in reflexive Banach spaces. Applied Mathematics and Computation, 228(??):38–48, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012435.

Chen:2011:SDE

[CX11] Xuefeng Chen and Jiawei Xiang. Solving diffusion equation using wavelet method. Applied Mathematics and Computation, 217(13):6426–6432, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000580.

Cheng:2014:ACP

[CX14a] Maolin Cheng and Minyin Xiang. Application of a combination production function model. Applied Mathematics and Computation, 236(??):33–40, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003245.

See corrigendum [CX14b].

Cheng:2014:CAC

[CX14b] Maolin Cheng and Minyin Xiang. Corrigendum to “Application of a combination production function model” [Applied Mathematics and Computation 236 (2014) 33– 40]. Applied Mathematics and Computation, 242(??):952, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009138.

See [CX14a].

Chen:2014:HGP

[CXD14] Huibo Chen, Yongsong Xiao, and Feng Ding. Hierarchical gradient parameter estimation algorithm for Hammerstein nonlinear systems using the key term separation principle. Applied Mathematics and Computation, 247 (??):1202–1210, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012958.

Christodoulou:2010:SFB

[CXG10] Evgenia Christodoulou, Christos Xenophontos, and Georgios C. Georgiou. The Singular Function Boundary Integral Method for singular Laplacian problems over circular sections. Applied Mathematics and Computation, 217(6):2773–2787, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008660.

Chen:2012:PSE

[CXL12] Fengde Chen, Xiangdong Xie, and Zhong Li. Partial survival and extinction of a delayed predatorprey model with stage structure. Applied Mathematics and Computation, 219(8):4157–4162, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010715.

Chen:2011:NAC

[CY11a] Yue-Hui Chen and Cheng-Yi Yu. A new algorithm for computing the inverse and the determinant of a Hessenberg matrix. Applied Mathematics and Computation, 218(8):4433–4436, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012525.

Cho:2011:SPM

[CY11b] N. E. Cho and M. Yoon.

Subordination properties for meromorphic multivalent functions associated with the multiplier transformation. Applied Mathematics and Computation, 218(3):729–733, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001664.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Cui:2012:SAN

[CY12] Jing Cui and Litan Yan.

Successive approximation of neutral stochastic evolution equations with infinite delay and Poisson jumps. Applied Mathematics and Computation, 218(12):6776–6784, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015311.

Chen:2014:SPD

[CYG14] Mi Chen, Kam Chuen Yuen, and Junyi Guo. Survival probabilities in a discrete semi-Markov risk model. Applied Mathematics and Computation, 232(??):205–215, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000940.

Cao:2011:AAP

[CYHL11] Junfei Cao, Qigui Yang, Zaitang Huang, and Qing Liu. Asymptotically almost periodic solutions of stochastic functional differential equations. Applied Mathematics and Computation, 218(5):1499–1511, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008666.

Chen:2013:DSE

[CYM13] Huai-Tang Chen, Shu-Huan Yang, and Wen-Xiu Ma. Double subequation method for complexiton solutions of nonlinear partial differential equations. Applied Mathematics and Computation, 219(9):4775–4781, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011241.

See comment [LLS14].

Chen:2014:TOS

[CYW^{+}14] Hui Ling Chen, Bo Yang, Su Jing Wang, Gang Wang, Da You Liu, Huai Zhong Li, and Wen Bin Liu. Towards an optimal support vector machine classifier using a parallel particle swarm optimization strategy. Applied Mathematics and Computation, 239(??):180–197, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005724.

Chen:2014:SSS

[CYY14] Jing Chen, Jie Yang, and Hongxiang Yang. Single soliton solutions of the coupled nonlinear Klein–Gordon equations with power law nonlinearity. Applied Mathematics and Computation, 246(??):184–191, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011060.

Cui:2013:BSN

[CYZ13] Zhoujin Cui, Zuodong Yang, and Rui Zhang. Blow-up of solutions for nonlinear parabolic equation with nonlocal source and nonlocal boundary condition. Applied Mathematics and Computation, 224(??):1–8, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009120.

Chang:2014:AFE

[CYZ14a] Yanzhen Chang, Danping Yang, and Zhijuan Zhang. Adaptive finite element approximation for a class of parameter estimation problems. Applied Mathematics and Computation, 231(??):284–298, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014318.

Chung:2014:PEA

[CYZ14b] Eric T. Chung, Man Chun Yuen, and Liuqiang Zhong. A-posteriori error analysis for a staggered discontinuous Galerkin discretization of the timeharmonic Maxwell’s equations. Applied Mathematics and Computation, 237(??):613–631, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005050.

Clark:2010:WTT

[CZ10] Stephen Clark and Petr Zemánek. On a Weyl– Titchmarsh theory for discrete symplectic systems on a half line. Applied Mathematics and Computation, 217(7):2952–2976, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310008830.

Chen:2011:BVM

[CZ11a] Hao Chen and Chengjian Zhang. Boundary value methods for Volterra integral and integro-differential equations. Applied Mathematics and Computation, 218(6):2619–2630, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010277.

Chen:2011:NMS

[CZ11b] Zhong Chen and Yong-Fang Zhou. A new method for solving Hilbert type singular integral equations. Applied Mathematics and Computation, 218(2):406–412, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007594.

Cheng:2011:NPP

[CZ11c] Huidong Cheng and Tongqian Zhang. A new predatorprey model with a profitless delay of digestion and impulsive perturbation on the prey. Applied Mathematics and Computation, 217(22):9198–9208, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100539X.

Chiarella:2013:AOP

[CZ13] Carl Chiarella and Jonathan Ziveyi. American option pricing under two stochastic volatility processes. Applied Mathematics and Computation, 224(??):283–310, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009156.

Chen:2014:CMA

[CZ14a] Bin Chen and Haigang Zhou. On completely monotone [properties] of an arbitrary real parameter function involving the gamma function. Applied Mathematics and Computation, 242(??):658–663, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006997.

Chen:2014:NGP

[CZ14b] Yuan Chen and Naimin Zhang. A note on the generalization of parameterized inexact Uzawa method for singular saddle point problems. Applied Mathematics and Computation, 235(??):318–322, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003373.

Crnjaric-Zic:2011:HOA

[CZC11] Nelida Crnjarić-Zic and Bojan Crnković. High order accurate semi-implicit WENO schemes for hyperbolic balance laws. Applied Mathematics and Computation, 217(21):8611–8629, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004668.

Cheng:2014:SFT

[CZD^{+}14] Jun Cheng, Hong Zhu, Yucai Ding, Shouming Zhong, and Qishui Zhong. Stochastic finite-time boundedness for Markovian jumping neural networks with time-varying delays. Applied Mathematics and Computation, 242(??):281–295, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400736X.

Chu:2011:MPN

[CZL11] Hongmei Chu, Yinlong Zhao, and Yinping Liu. A MAPLE package of new ADM-Padé approximate solution for nonlinear problems. Applied Mathematics and Computation, 217(17):7074–7091, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013020.

Cheng:2014:MBP

[CZL14] R. J. Cheng, L. W.

Zhang, and K. M. Liew. Modeling of biological population problems using the element-free kp-Ritz method. Applied Mathematics and Computation, 227(??):274–290, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012010.

Chen:2011:AEB

[CZM11] Shaochun Chen, Yanjun Zheng, and Shipeng Mao. Anisotropic error bounds of Lagrange interpolation with any order in two and three dimensions. Applied Mathematics and Computation, 217(22):9313–9321, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005716.

Ciric:2012:GME

[CZSV12] Marija S. ’irić, Milan Lj. Zlatanović, Mića S. Stanković, and Ljubica S. Velimirović. On geodesic mappings of equidistant generalized Riemannian spaces. Applied Mathematics and Computation, 218(12): 6648–6655, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300311014512.

Chen:2014:FDBa

[CZX^{+}14a] Xiaoyue Chen, Jianzhong Zhou, Han Xiao, Ercheng Wang, Jian Xiao, and Huifeng Zhang. Fault diagnosis based on comprehensive geometric characteristic and probability neural network. Applied Mathematics and Computation, 230(??):542–554, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014124.

Chen:2014:FDBb

[CZX^{+}14b] Xiaoyue Chen, Jianzhong Zhou, Jian Xiao, Xinxin Zhang, Han Xiao, Wenlong Zhu, and Wenlong Fu. Fault diagnosis based on dependent feature vector and probability neural network for rolling element bearings. Applied Mathematics and Computation, 247(??):835–847, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012879.

Chen:2011:NEE

[CZZ^{+}11] Yiming Chen, Zhiquan Zhou, Jing Zhang, Yulian Li, and Xiaojuan Wang. A new error estimation approach based on fuzzy logic system for H-R adaptive boundary element method. Applied Mathematics and Computation, 218(5):2167–2177, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311009660.

Chen:2012:SAU

[CZZ12] Huabin Chen, Yong Zhang, and Yang Zhao. Stability analysis for uncertain neutral systems with discrete and distributed delays. Applied Mathematics and Computation, 218(23):11351–11361, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004420.

Chen:2014:NES

[CZZ14] Huabin Chen, Chuanxi Zhu, and Yingying Zhang. A note on exponential stability for impulsive neutral stochastic partial functional differential equations. Applied Mathematics and Computation, 227(??):139–147, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011156.

Cheng:2013:NDD

[CZZL13] Jun Cheng, Hong Zhu, Shouming Zhong, and Guihua Li. Novel delaydependent robust stability criteria for neutral systems with mixed timevarying delays and nonlinear perturbations. Applied Mathematics and Computation, 219(14):7741–7753, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001136.

Daneshpajouh:2011:NMI

[DA11a] Hamed Daneshpajouh and Mohsen Alimomeni. A note on the minimum interval cost flow problem. Applied Mathematics and Computation, 217(17):7051–7052, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001950.

Duman:2011:SSS

[DA11b] Ahmet Duman and Kemal Aydın. Sensitivity of Schur stability of monodromy matrix. Applied Mathematics and Computation, 217(15):6663–6670, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311000683.

DAlotto:2012:CAU

[D’A12] Louis D’Alotto. Cellular automata using infinite computations. Applied Mathematics and Computation, 218(16):8077–8082, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013117.

Duldul:2014:WLM

[DA14] Mustafa Düldül and Özlem Akbaba. Willmore-like methods for the intersection of parametric (hyper)surfaces. Applied Mathematics and Computation, 226(??):516–527, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011405.

Dumon:2013:PGD

[DAA13] A. Dumon, C. Allery, and A. Ammar. Proper Generalized Decomposition method for incompressible Navier–Stokes equations with a spectral discretization. Applied Mathematics and Computation, 219(15):8145–8162, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001434.

Duque:2014:NSP

[DAA14] José C. M. Duque, Rui M. P. Almeida, and Stanislav N. Antontsev. Numerical study of the porous medium equation with absorption, variable exponents of nonlinearity and free boundary. Applied Mathematics and Computation, 235(??):137–147, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003488.

deAmo:2012:MAM

[dACSS12] Enrique de Amo, Manuel Díaz Carrillo, Juan Fernández Sánchez, and Antonio Salmerón. Moments and associated measures of copulas with fractal support. Applied Mathematics and Computation, 218(17):8634–8644, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312001476.

Doha:2012:ESP

[DAEY12] E. H. Doha, W. M. Abd-Elhameed, and Y. H. Youssri. Efficient spectral-Petrov–Galerkin methods for the integrated forms of third- and fifth-order elliptic differential equations using general parameters generalized Jacobi polynomials. Applied Mathematics and Computation, 218(15):7727–7740, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000677.

Dafa-Alla:2010:NFD

[DAHAK10] Anour F. A. Dafa-Alla, Kyung-Won Hwang, Soyoung Ahn, and Philsu Kim. A new finite difference representation for Poisson’s equation on R^{3} from a contour integral. Applied Mathematics and Computation, 217(8):3624–3634, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010477.

Lopez:2014:SMG

[dALORPGC14] Francisco de Asís López, Celestino Ordóñez, Javier Roca-Pardiñas, and Silverio García-Cortés. A statistical method for geometry inspection from point clouds. Applied Mathematics and Computation, 242(??):562–568, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008303.

Danca:2010:ASL

[Dan10] Marius-F. Danca. Attractors synthesis for a Lotka–Volterra like system. Applied Mathematics and Computation, 216(7):2107–2117, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002869.

Daneshmand:2014:CSI

[Dan14] Farhang Daneshmand.

Combined strain-inertia gradient elasticity in free vibration shell analysis of single walled carbon nanotubes using shell theory. Applied Mathematics and Computation, 243(??):856–869, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007942.

Dardery:2011:NKA

[Dar11] Samah M. Dardery. Newton– Kantorovich approximations to nonlinear singular integral equation with shift. Applied Mathematics and Computation, 217(22):8873–8882, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005388.

Darwish:2012:PFI

[Dar12] Mohamed Abdalla Darwish. On a perturbed functional integral equation of Urysohn type. Applied Mathematics and Computation, 218(17): 8800–8805, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001646.

Darwish:2013:MSC

[Dar13] Mohamed Abdalla Darwish. Monotonic solutions of a convolution functional-integral equation. Applied Mathematics and Computation, 219(22):10777–10782, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004906.

Das:2011:EGC

[Das11] Kinkar Ch. Das. Extremal graph characterization from the bounds of the spectral radius of weighted graphs. Applied Mathematics and Computation, 217(18):7420–7426, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002232.

Dassios:2012:PRS

[Das12] Ioannis K. Dassios. Perturbation and robust stability of autonomous singular linear matrix difference equations. Applied Mathematics and Computation, 218(12):6912–6920, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311015542.

Das:2013:LGA

[Das13]  K. Das.  Lie group

analysis for nanofluid flow past a convectively heated stretching surface. Applied Mathematics and Computation, 221(??):547–557, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007224.

DeSchrijver:2012:DPRa

[DAV12a] Steven K. De Schrijver, El-Houssaine Aghezzaf, and Hendrik Vanmaele. Double precision rational approximation algorithm for the inverse standard normal first order loss function. Applied Mathematics and Computation, 219(3):1375–1382, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007114.

DeSchrijver:2012:DPRb

[DAV12b] Steven K. De Schrijver, El-Houssaine Aghezzaf, and Hendrik Vanmaele. Double precision rational approximation algorithms for the standard normal first and second order loss functions. Applied Mathematics and Computation, 219(4):2320–2330, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008041.

laHoz:2014:MCC

[DAV14a] Mauricio Zapateiro De la Hoz, Leonardo Acho, and Yolanda Vidal. A modified Chua chaotic oscillator and its application to secure communications. Applied Mathematics and Computation, 247(??):712–722, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012491.

Schrijver:2014:DPR

[DAV14b] Steven K. De Schrijver, El-Houssaine Aghezzaf, and Hendrik Vanmaele. Double precision rational approximation algorithm for the inverse standard normal second order loss function. Applied Mathematics and Computation, 232(??):247–253, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000319.

Dehghan:2011:BFV

[DB11a] Mehdi Dehghan and Gholam Hosein Baradaran. Buckling and free vibration analysis of thick rectangular plates resting on elastic foundation using mixed finite element and differential quadrature method. Applied Mathematics and Computation, 218(6):2772–2784, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010460.

Deo:2011:SAR

[DB11b] Naokant Deo and Neha Bhardwaj. Some approximation results for Durrmeyer operators. Applied Mathematics and Computation, 217(12):5531–5536, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012282.

Deckers:2012:ECR

[DB12a] Karl Deckers and Adhemar Bultheel. The existence and construction of rational Gauss-type quadrature rules. Applied Mathematics and Computation, 218(20):10299– 10320, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003918.

Dutta:2012:FSG

[DB12b] S. Dutta and P. Baliarsingh. On the fine spectra of the generalized r th difference operator Δ_{ν}^{r} on the sequence space ℓ_{1}. Applied Mathematics and Computation, 219(4):1776–1784, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008090.

Doha:2013:SJS

[DBBEE13] E. H. Doha, A. H. Bhrawy, D. Baleanu, and S. S. Ezz-Eldien. On shifted Jacobi spectral approximations for solving fractional differential equations. Applied Mathematics and Computation, 219(15):8042–8056, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300091X.

Deger:2011:SGR

[DBG11] Ali H. Deger, Murat Besenk, and Bahadir Ö. Güler. On suborbital graphs and related continued fractions. Applied Mathematics and Computation, 218(3):746–750, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004334.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Dassios:2014:NHS

[DBK14] Ioannis K. Dassios, Dumitru I. Baleanu, and Grigoris I. Kalogeropoulos. On non-homogeneous singular systems of fractional nabla difference equations. Applied Mathematics and Computation, 227(??):112–131, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313011570.

Dhar:2012:RIN

[DBS12] Joydip Dhar, Randhir Singh Baghel, and Anuj Kumar Sharma. Role of instant nutrient replenishment on plankton dynamics with diffusion in a closed system: a pattern formation. Applied Mathematics and Computation, 218(17):8925–8936, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001920.

See comments [Ban14].

Deng:2013:ACD

[DC13a] Peter Shaohua Deng and Henry C. J. Chao. Analysis on comparison of distances derived by onenorm and two-norm with weight functions. Applied Mathematics and Computation, 219(17):9093–9098, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300341X.

Dumbser:2013:SSI

[DC13b] Michael Dumbser and Vincenzo Casulli. A staggered semi-implicit spectral discontinuous Galerkin scheme for the shallow water equations. Applied Mathematics and Computation, 219(15):8057–8077, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001720.

Dong:2014:MAE

[DC14] Qiao-Li Dong and Li-Qun Cao. Multiscale asymptotic expansions methods and numerical algorithms for the wave equations in perforated domains. Applied Mathematics and Computation, 232(??):872–887, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014021.

Braga:2014:NAM

[dCBdFM14] Denis de Carvalho Braga, Nivaldo Gonçalves de Faria, and Luis Fernando Mello. Nonlinear analysis in a modified van der Pol oscillator. Applied Mathematics and Computation, 245(??):474–486, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010728.

daCosta:2014:PSP

[dCCL14] Diogo Ricardo da Costa, Iberê L. Caldas, and Edson D. Leonel. Phase space properties and chaotic transport for a particle moving in a time dependent step potential well. Applied Mathematics and Computation, 236(??):215–228, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003749.

Diaz:2010:NEA

[dCDNCFRH10] J. J. del Coz Díaz, P. J. García Nieto, D. Castro-Fresno, and J. Rodríguez-Hernández. Nonlinear explicit analysis and study of the behaviour of a new ring-type brake energy dissipator by FEM and experimental comparison. Applied Mathematics and Computation, 216(5):1571–1582, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310002511.

Diaz:2012:NLT

[dCDNRH12] J. J. del Coz Díaz, P. J. García Nieto, F. P. Alvarez Rabanal, and J. Domínguez Hernández. Non-linear thermal analysis of the efficiency of light concrete multi-holed bricks with large recesses by FEM. Applied Mathematics and Computation, 218 (20):10040–10049, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312003591.

Diaz:2013:NAP

[dCDNV^{+}13] J. J. del Coz Díaz, P. J. García Nieto, J. A. Vilán Vilán, F. P. Alvarez Rabanal, A. Navarro-Manso, and M. Alonso-Martínez. Nonlinear analysis of the pressure field in industrial buildings with curved metallic roofs due to the wind effect by FEM. Applied Mathematics and Computation, 221(??):202–213, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300313007066.

Diaz:2011:NLA

[dCDNVS11] J. J. del Coz Díaz, P. J. García Nieto, J. A. Vilán Vilán, and J. L. Suárez Sierra. Non-linear analysis and calculation of the performance of a shelving protection system by FEM. Applied Mathematics and Computation, 218(6):2365–2376, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010150.

Dragomir:2011:SII

[DCK11] S. S. Dragomir, Y. J.

Cho, and S. S. Kim. Some inequalities in inner product spaces related to the generalized triangle inequality. Applied Mathematics and Computation, 217(18):7462–7468, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002372.

See note [PS11].

Du:2011:CPB

[DCL11] Chaoxiong Du, Haibo Chen, and Yirong Liu.

Center problem and bifurcation behavior for a class of quasi analytic systems. Applied Mathematics and Computation, 217(9):4665–4675, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011306.

Duan:2012:SIV

[DCR12] Jun-Sheng Duan, Temuer Chaolu, and Randolph Rach. Solutions of the initial value problem for nonlinear fractional ordinary differential equations by the Rach–Adomian– Meyers modified decomposition method. Applied Mathematics and Computation, 218(17):8370–8392, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001129.

Diblik:2013:USE

[DCR13] Josef Diblík, Radoslav Chupác, and Miroslava Ruzicková. Unbounded solutions of the equation ẏ(t) = ∑_{i=1}^{n} β_{i}(t)[y(t − δ_{i}) − y(t − τ_{i})]. Applied Mathematics and Computation, 221(??):610–619, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313007492.

Dongqin:2011:CES

[DD11] Cheng Dongqin and Guo Dachang. Cycle embedding in star graphs with more conditional faulty edges. Applied Mathematics and Computation, 218(7):3856–3867, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012057.

DeCosmis:2012:UGM

[DD12] Sonia De Cosmis and Renato De Leone. The use of grossone in mathematical programming and operations research. Applied Mathematics and Computation, 218(16):8029–8038, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009891.

Dincic:2013:ICR

[DD13] Nebojsa C. Dincić and Dragan S. Djordjević.

Identities concerning the reverse order law for the Moore–Penrose inverse. Applied Mathematics and Computation, 220(??):439–445, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006565.

DeLellis:2010:ASC

[DdGL10] Pietro DeLellis, Mario diBernardo, Franco Garofalo, and Davide Liuzza. Analysis and stability of consensus in networked control systems. Applied Mathematics and Computation, 217(3):988–1000, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001505.

Drame:2012:PTD

[DDWW12] A. K. Drame, D. Dochain, J. J. Winkin, and P. R. Wolenski. Periodic trajectories of distributed parameter biochemical systems with time delay. Applied Mathematics and Computation, 218(14):7395–7405, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015761.

Dhouib:2011:NMC

[DE11] Diala Dhouib and Sonda Elloumi. A new multicriteria approach dealing with dependent and heterogeneous criteria for end-of-life product strategy. Applied Mathematics and Computation, 218(5):1668–1681, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008794.

Demir:2012:SDP

[DE12] Ali Demir and Arzu Erdem. Stability for determining the principal coefficient of parabolic equation. Applied Mathematics and Computation, 219(8):3826–3830, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009988.

Dea:2013:ABC

[Dea13] John R. Dea. Absorbing boundary conditions for the fractional wave equation. Applied Mathematics and Computation, 219(18):9810–9820, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003743.

Qing-dao-er-ji:2014:SBB

[dejW14] Ren Qing dao-er ji and Yuping Wang. Security based bi-objective flow shop scheduling model and its hybrid genetic algorithm. Applied Mathematics and Computation, 243(??):637–643, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008558.

Demiray:2011:NWP

[Dem11] Hilmi Demiray. A note on the wave propagation in water of variable depth. Applied Mathematics and Computation, 218(5):2294–2299, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009969.

Demiray:2014:NPW

[Dem14] Hilmi Demiray. A note on the progressive wave solution of the perturbed Korteweg–de Vries equation with variable dissipation. Applied Mathematics and Computation, 248(??):562–566, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013861.

Deng:2010:NDI

[Den10] Shengfu Deng. Nonlinear discrete inequalities with two variables and their applications. Applied Mathematics and Computation, 217(5):2217–2225, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007678.

Deng:2011:IBM

[Den11a] Chongyang Deng. Improved bounds on the magnitude of the derivative of rational Bézier curves. Applied Mathematics and Computation, 218(1):204–206, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007491.

Deng:2011:LCB

[Den11b] Qin Deng. On the logarithmic coefficients of Bazilevič functions. Applied Mathematics and Computation, 217(12): 5889–5894, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012774.

Deng:2011:RIC

[Den11c] Shengfu Deng. Remark on invariant curves for planar mappings. Applied Mathematics and Computation, 217(21):8419–8424, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004048.

Deng:2011:NET

[Den11d]  Xijun Deng.  A note

on exact travelling wave solutions for the modified Camassa–Holm and Degasperis–Procesi equations. Applied Mathematics and Computation, 218(5):2269–2276, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100991X.

Deng:2013:CCA

[Den13] Chun Yuan Deng. Characterizations of the commutators and the anticommutator involving idempotents. Applied Mathematics and Computation, 221(??):351–359, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006954.

Deng:2014:ETW

[Den14]  Shengfu Deng.  Exis-

tence of travelling wave front solutions of a twodimensional anisotropic model. Applied Mathematics and Computation, 226(??):503–515, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010497.

Deo:2012:FRC

[Deo12]  Naokant Deo.  Faster

rate of convergence on Srivastava–Gupta operators. Applied Mathematics and Computation, 218 (21):10486–10491, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312003955.

DAmbrosio:2012:EFT

[DEP12] R. D’Ambrosio, E. Esposito, and B. Paternoster. Exponentially fitted twostep Runge–Kutta methods: Construction and parameter selection. Applied Mathematics and Computation, 218(14):7468–7480, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000471.

DiRocco:2010:AAI

[DESW10] Sandra Di Rocco, David Eklund, Andrew J. Sommese, and Charles W. Wampler. Algebraic C^{⋆}-actions and the inverse kinematics of a general 6R manipulator. Applied Mathematics and Computation, 216(9):2512–2524, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300309010613.

Dosly:2011:VTP

[DF11] Ondřej Došlý and Simona Fišnarová. Variational technique and principal solution in halflinear oscillation criteria. Applied Mathematics and Computation, 217 (12):5385–5391, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012063.

daFonseca:2014:USP

[dF14a] Carlos M. da Fonseca.

Unifying some Pell and Fibonacci identities. Applied Mathematics and Computation, 236(??):41–42, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004263.

Durante:2014:DER

[DF14b] Fabrizio Durante and Rachele Foschi. Dependence of exchangeable residual lifetimes subject to failure. Applied Mathematics and Computation, 235(??):502–511, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003178.

Dukaric:2014:SRC

[DFG14] Masa Dukarić, Brigita Fercec, and Jaume Giné. The solution of the 1:- 3 resonant center problem in the quadratic case. Applied Mathematics and Computation, 237(??):501–511, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005189.

deFrutos:2012:OEB

[dFGAN12] Javier de Frutos, Bosco García-Archilla, and Julia Novo. Optimal error bounds for twogrid schemes applied to the Navier–Stokes equations. Applied Mathematics and Computation, 218(13):7034–7051, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015384.

Dobry:2014:SNE

[DFGL14] A. Dobry, O. Fojón, M. Gadella, and L. P. Lara. Some numerical estimations of energy levels on a model for a graphene ribbon in a magnetic field. Applied Mathematics and Computation, 235(??):8– 16, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002999.

Dou:2010:WGM

[DFM10] Fang-Fang Dou, Chu-Li Fu, and Yun-Jie Ma. A wavelet-Galerkin method for high order numerical differentiation. Applied Mathematics and Computation, 215(10):3702–3712, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009977.

Pinto:2014:EPF

[dFPF14] Ricardo Luiz Utsch de Freitas Pinto and Ricardo Poley Martins Ferreira. An exact penalty function based on the projection matrix. Applied Mathematics and Computation, 245(??):66–73, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010650.

Napoli:2014:TEU

[DFTQOB14] Edoardo Di Napoli, Diego Fabregat-Traver, Gregorio Quintana-Ortí, and Paolo Bientinesi. Towards an efficient use of the BLAS library for multilinear tensor contractions. Applied Mathematics and Computation, 235(??):454–468, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002902.

Dockner:2010:DIS

[DG10a] Engelbert J. Dockner and Andrea Gaunersdorfer. Dynamic investment strategies with demand-side and cost-side risks. Applied Mathematics and Computation, 217(3):1001–1009, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004935.

Dragomirescu:2010:ANS

[DG10b] F. I. Dragomirescu and C. I. Gheorghiu. Analytical and numerical solutions to an electrohydrodynamic stability problem. Applied Mathematics and Computation, 216(12):3718–3727, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005722.

Dudzinski:2010:SEC

[DG10c] Marcin Dudziński and Przemysław Górka. Stochastic evolution of 2D crystals. Applied Mathematics and Computation, 216(1): 205–212, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310000482.

Dalal:2013:RCA

[DG13] Aseem Dalal and N. K.

Govil. On region containing all the zeros of a polynomial. Applied Mathematics and Computation, 219(17):9609–9614, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313002117.

Dalal:2014:ACA

[DG14a] Aseem Dalal and N. K. Govil. Annulus containing all the zeros of a polynomial. Applied Mathematics and Computation, 249(??):429–435, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314014040.

Du:2014:SRA

[DG14b] Dianlou Du and Xue Geng. Symplectic realizations and action-angle coordinates for the Lie– Poisson system of Dirac hierarchy. Applied Mathematics and Computation, 244(??):222–234, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009473.

DelCozDiaz:2011:SSN

[DGCFM11] J. J. Del Coz Díaz, P. J. García Nieto, D. Castro-Fresno, and P. Menéndez Rodríguez. Steady state numerical simulation of the particle collection efficiency of a new urban sustainable gravity settler using design of experiments by FVM. Applied Mathematics and Computation, 217(21):8166–8178, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004103.

Davodi:2010:FGE

[DGDA10] Amin G. Davodi, D. D.

Ganji, Arash G. Davodi, and A. Asgari. Finding general and explicit solutions (2 + 1) dimensional Broer–Kaup– Kupershmidt system nonlinear equation by expfunction method. Applied Mathematics and Computation, 217(4):1415–1420, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309005359.

Dun:2011:DCL

[DGH11] Ao Dun, Zhiyong Geng, and Lin Huang. Design of controller for Lur’e systems guaranteeing dichotomy. Applied Mathematics and Computation, 217(22):8927–8935, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004620.

Donatelli:2014:SBP

[DGM^{+}14] Marco Donatelli, Carlo Garoni, Mariarosa Mazza, Stefano Serra-Capizzano, and Debora Sesana. Spectral behavior of preconditioned non-Hermitian multilevel block Toeplitz matrices with matrixvalued symbol. Applied Mathematics and Computation, 245(??):158–173, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009801.

Dattoli:2010:NDS

[DGMR10] G. Dattoli, B. Germano, M. R. Martinelli, and P. E. Ricci. Negative derivatives and special functions. Applied Mathematics and Computation, 217(8):3924–3928, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010167.

Dattoli:2012:TLP

[DGMR12] G. Dattoli, B. Germano, M. R. Martinelli, and P. E. Ricci. Touchard like polynomials and generalized Stirling numbers. Applied Mathematics and Computation, 218(12):6661–6665, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014743.

Dolicanin:2013:CPI

[DGOR13] Diana Dolićanin, Jaume Giné, Regilene Oliveira, and Valery G. Romanovski. The center problem for a 2: −3 resonant cubic Lotka–Volterra system. Applied Mathematics and Computation, 220(??):12–19, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006140.

DiNardo:2011:NAC

[DGS11] E. Di Nardo, G. Guarino, and D. Senato. A new algorithm for computing the multivariate Faà di Bruno’s formula. Applied Mathematics and Computation, 217(13):6286–6295, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000051.

Das:2014:SAL

[DGS14a] Dushmanta Kumar Das, Sandip Ghosh, and Bidyadhar Subudhi. Stability analysis of linear systems with two delays of overlapping ranges. Applied Mathematics and Computation, 243(??):83–90, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007255.

Das:2014:TDM

[DGS14b] Dushmanta Kumar Das, Sandip Ghosh, and Bidyadhar Subudhi. Tolerable delay-margin improvement for systems with input-output delays using dynamic delayed feedback controllers. Applied Mathematics and Computation, 230(??):57–64, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313013568.

De:2014:IPP

[DGS14c] Sujit Kumar De, Adrijit Goswami, and Shib Sankar Sana. An interpolating bypass to Pareto optimality in intuitionistic fuzzy technique for a EOQ model with time sensitive backlogging. Applied Mathematics and Computation, 230(??):664–674, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014276.

Deb:2012:NST

[DGSB12] Anish Deb, Anindita Ganguly, Gautam Sarkar, and Amitava Biswas. Numerical solution of third order linear differential equations using generalized one-shot operational matrices in orthogonal hybrid function domain. Applied Mathematics and Computation, 219(4):1485–1514, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007679.

Daftardar-Gejji:2014:NPC

[DGSB14] Varsha Daftardar-Gejji, Yogita Sukale, and Sachin Bhalekar. A new predictorcorrector method for fractional differential equations. Applied Mathematics and Computation, 244(??):158–182, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009412.

Dai:2013:GBN

[DH13a] Guowei Dai and Xiaoling Han. Global bifurcation and nodal solutions for fourth-order problems with sign-changing weight. Applied Mathematics and Computation, 219(17):9399–9407, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003640.

Deng:2013:EMS

[DH13b] Zhiying Deng and Yisheng Huang. Existence and multiplicity of symmetric solutions for the weighted critical quasilinear problems. Applied Mathematics and Computation, 219(9):4836–4846, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011393.

Duan:2014:ENS

[DH14] Lian Duan and Lihong Huang. Existence of nontrivial solutions for Kirchhoff-type variational inclusion system in R^{N}. Applied Mathematics and Computation, 235(??): 174–186, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031400318X.

Dhaouadi:2013:PSW

[Dha13]  Lazhar Dhaouadi.  q-

prolate spheroidal wave functions and q-sampling theorem. Applied Mathematics and Computation, 222(??):453–461, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008151.

Dyer:2012:ADO

[DHDB12] John D. Dyer, Roy J.

Hartfield, Gerry V. Dozier, and John E. Burkhalter. Aerospace design optimization using a steady state real-coded genetic algorithm. Applied Mathematics and Computation, 218(9):4710–4730, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009714.

Dargazany:2013:SAF

[DHI13] Roozbeh Dargazany, Karl Hörnes, and Mikhail Itskov. A simple algorithm for the fast calculation of higher order derivatives of the inverse function. Applied Mathematics and Computation, 221(??):833–838, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006747.

Dragomir:2014:EBA

[DHM14] S. S. Dragomir, Charis Harley, and E. Momoniat. Error bounds in approximating the Riemann–Stieltjes integral of C^{n+1}-class integrands and nonsmooth integrators. Applied Mathematics and Computation, 249(??):237–246, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013885.

Dutt:2014:ISE

[DHMU14] P. Dutt, Akhlaq Husain, A. S. Vasudeva Murthy, and C. S. Upadhyay. h– p Spectral element methods for three dimensional elliptic problems on nonsmooth domains. Applied Mathematics and Computation, 234(??):13–35, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002641.

Duan:2014:SIN

[DHQ14] Xian-Bao Duan, Nan-Nan He, and Xin-Qiang Qin. Shape identification for Navier–Stokes problem using shape sensitivity analysis and level set method. Applied Mathematics and Computation, 231(??):1–8, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000058.

Dong:2010:SCI

[DHS10] Qiao-Li Dong, Songnian He, and Fang Su. Strong convergence of an iterative algorithm for an infinite family of strict pseudo-contractions in Banach spaces. Applied Mathematics and Computation, 216(3):959–969, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001359.

Du:2010:NMC

[DHSZ10] Jianguo Du, Tingwen Huang, Zhaohan Sheng, and Haibin Zhang. A new method to control chaos in an economic system. Applied Mathematics and Computation, 217(6):2370–2380, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007812.

Dong:2011:CTS

[DHZ11] Qiao-Li Dong, Songnian He, and Jing Zhao. Convergence theorems of shrinking projection methods for equilibrium problem, variational inequality problem and a finite family of relatively quasi-nonexpansive mappings. Applied Mathematics and Computation, 217(24):10256–10266, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006941.

DelaSen:2013:ANE

[DI13a] M. De la Sen and A. Ibeas.

Asymptotically non-expansive self-maps and global stability with ultimate boundedness of dynamic systems. Applied Mathematics and Computation, 219(22):10655–10667, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004086.

DelaSen:2013:FPB

[DI13b] M. De la Sen and A. Ibeas.

Fixed points and best proximity points in contractive cyclic self-maps satisfying constraints in closed integral form with some applications. Applied Mathematics and Computation, 219(10):5410–5426, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011691.

Diagana:2013:EGA

[Dia13]  Toka Diagana.  Ex-

istence of globally attracting almost automorphic solutions to some nonautonomous higherorder difference equations. Applied Mathematics and Computation, 219(12):6510–6519, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000088.

Diagana:2014:ERS

[Dia14] Toka Diagana. Existence results for some damped second-order Volterra integrodifferential equations. Applied Mathematics and Computation, 237(??): 304–317, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004767.

DelaSen:2013:TVS

[DIAQ13] M. De la Sen, A. Ibeas, and S. Alonso-Quesada. A time-varying SIS epidemic model with incidence rate depending on the susceptible and infective populations with eventual impulsive effects. Applied Mathematics and Computation, 219(10):5516–5536, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201154X.

Dimitriou:2013:BAP

[Dim13a]  Ioannis Dimitriou.  A

batch arrival priority queue with recurrent repeated demands, admission control and hybrid failure recovery discipline. Applied Mathematics and Computation, 219(24):11327–11340, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300595X.

Dimitrov:2013:ECT

[Dim13b]  Darko Dimitrov.  Ef-

ficient computation of trees with minimal atombond connectivity index. Applied Mathematics and Computation, 224(??):663–670, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009570.

Ding:2011:API

[Din11] Xie Ping Ding. Auxiliary principle and iterative algorithm for a new system of generalized mixed equilibrium problems in Banach spaces. Applied Mathematics and Computation, 218(7):3507–3514, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011386.

Ding:2012:IAS

[Din12] Xie Ping Ding. Iterative algorithm of solutions for a system of generalized mixed implicity equilibrium problems in reflexive Banach spaces. Applied Mathematics and Computation, 218(9):4953–4961, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013063.

Ding:2013:CME

[Din13a]  Feng Ding.  Compu-

tation of matrix exponentials of special matrices. Applied Mathematics and Computation, 223(??):311–326, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008291.

Ding:2013:UMP

[Din13b]  Sibo Ding.  Uncertain

multi-product newsboy problem with chance constraint. Applied Mathematics and Computation, 223(??):139–146, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008333.

Diaby:2014:GAS

[DISS14a] Mouhamadou Diaby, Abderrahman Iggidr, Mamadou Sy, and Abdou Sène. Global analysis of a schistosomiasis infection model with biological control. Applied Mathematics and Computation, 246(??):731–742, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314011710.

Diblik:2014:NEP

[DISS14b] Josef Diblík, Bratislav Iricanin, Stevo Stević, and Zdenek Smarda. Note on the existence of periodic solutions of a class of systems of differentialdifference equations. Applied Mathematics and Computation, 232(??): 922–928, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001416.

Dascal:2012:CAS

[DJ12] Judita Dascal and Justyna Jarczyk. Computer assisted solution of an equality problem of mean values. Applied Mathematics and Computation, 219(2):475–481, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006170.

Delic:2014:NAI

[DJ14a] Aleksandra Delić and Bosko S. Jovanović. Numerical approximation of an interface problem for fractional in time diffusion equation. Applied Mathematics and Computation, 229(??):467–479, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301343X.

Dobrogowska:2014:FMA

[DJ14b] Alina Dobrogowska and Grzegorz Jakimowicz. Factorization method applied to the second order q-difference operators. Applied Mathematics and Computation, 228(??):147–152, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301271X.

Doucet:2010:SIE

[DJT10] Arnaud Doucet, Adam M.

Johansen, and Vladislav B. Tadić. On solving integral equations using Markov chain Monte Carlo methods. Applied Mathematics and Computation, 216(10):2869–2880, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310004042.

Ding:2013:DHB

[DJY13] Yuting Ding, Weihua Jiang, and Pei Yu. Double Hopf bifurcation in a container crane model with delayed position feedback. Applied Mathematics and Computation, 219(17):9270–9281, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002762.

Das:2011:AASc

[DK11a] S. Das and R. Kumar. Approximate analytical solutions of fractional gas dynamic equations. Applied Mathematics and Computation, 217(24):9905–9915, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005248.

Demina:2011:ESN

[DK11b] Maria V. Demina and Nikolai A. Kudryashov. On elliptic solutions of nonlinear ordinary differential equations. Applied Mathematics and Computation, 217(23):9849–9853, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006011.

DelaSen:2014:CJM

[DK14a] M. De la Sen and E. Karapinar. On a cyclic Jungck modified TS-iterative procedure with application examples. Applied Mathematics and Computation, 233(??):383–397, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002409.

Dimitrova:2014:NRM

[DK14b] Neli S. Dimitrova and Mikhail I. Krastanov. New result on the modelbased biological control of the chemostat. Applied Mathematics and Computation, 237(??):686–694, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004986.

Dosla:2014:AOP

[DK14c] Zuzana Doslá and Jana Krejcová. Asymptotic and oscillatory properties of the fourth-order nonlinear difference equations. Applied Mathematics and Computation, 249(??):164–173, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014088.

Dolezel:2013:OCF

[DKK^{+}13] Ivo Dolezel, Pavel Karban, Petr Kropík, Václav Kotlan, and David Pánek. Optimized control of field current in thermoelastic actuator for accurate setting of position. Applied Mathematics and Computation, 219(13):7187–7193, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008216.

Dostal:2011:CDP

[DKMM11] Zdeněk Dostál, Tomáš Kozubek, Alexandros Markopoulos, and Martin Menšík. Cholesky decomposition of a positive semidefinite matrix with known kernel. Applied Mathematics and Computation, 217(13):6067–6077, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012713.

Darwish:2014:FIE

[DKS14] Mohamed Abdalla Darwish, Bothayna S. Kashkari, and Kishin Sadarangani. On a functional-integral equation with deviating arguments. Applied Mathematics and Computation, 246(??):64–70, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401114X.

Dolezel:2013:IHT

[DKU13] Ivo Dolezel, Petr Kropík, and Bohus Ulrych. Induction heating of thin metal plates in time-varying external magnetic field solved as nonlinear hardcoupled problem. Applied Mathematics and Computation, 219(13):7159–7169, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010824.

Ding:2010:ESC

[DL10] Xiaohua Ding and Kaining Wu Mingzhu Liu. The Euler scheme and its convergence for impulsive delay differential equations. Applied Mathematics and Computation, 216(5):1566–1570, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002493.

Duque:2011:SPT

[DL11] Cosme Duque and Marcos Lizana. Stable periodic traveling waves for a predator-prey model with non-constant death rate and delay. Applied Mathematics and Computation, 217(23):9717–9722, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006266.

Dai:2012:IBB

[DL12a] Chao-Qing Dai and Cui-Yun Liu. Interaction behaviors between solitons for the (2 + 1)- dimensional long wave short wave interaction model. Applied Mathematics and Computation, 219(5):2658–2667, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008910.

DeBonis:2012:NSS

[DL12b] Maria Carmela De Bonis and Concetta Laurita. Numerical solution of systems of Cauchy singular integral equations with constant coefficients. Applied Mathematics and Computation, 219(4):1391–1410, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008156.

Danca:2013:PSG

[DL13] Marius-F. Danca and Nicolae Lung. Parameter switching in a generalized Duffing system: Finding the stable attractors. Applied Mathematics and Computation, 223(??):101–114, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008436.

Deng:2014:DPL

[DL14] Guifeng Deng and Xianyi Li. Dichotomy of a perturbed Lyness difference equation. Applied Mathematics and Computation, 236(??):229–234, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004408.

Du:2010:BLC

[DLC10] Chaoxiong Du, Yirong Liu, and Haibo Chen. The bifurcation of limit cycles in Z_{n}-equivariant vector fields. Applied Mathematics and Computation, 217(5):2041–2056, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007502.

deLuna:2013:RFO

[dLDP13] T. Morales de Luna, M. J. Castro Díaz, and C. Parés. Reliability of first order numerical schemes for solving shallow water system over abrupt topography. Applied Mathematics and Computation, 219(17):9012–9032, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002865.

Liu:2014:CNB

[dLfX14] Lan dong Liu and Shu fang Xu. Condition number and backward errors of nonsymmetric algebraic Riccati equation. Applied Mathematics and Computation, 242(??):716–728, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008728.

Dai:2010:ETW

[DLFZ10] Zhengde Dai, Songqing Lin, Haiming Fu, and Xiping Zeng. Exact threewave solutions for the KP equation. Applied Mathematics and Computation, 216(5):1599–1604, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002559.

Ding:2014:RDR

[DLGJ14] Zhanwen Ding, Qiang Li, Dongliang Ge, and Shumin Jiang. Research on dynamics in a resource extraction game with bounded rationality. Applied Mathematics and Computation, 236(??):628–634, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004378.

delaHoz:2012:MET

[dlHV12] F. de la Hoz and F. Vadillo. A mean extinction-time estimate for a stochastic Lotka–Volterra predatorprey model. Applied Mathematics and Computation, 219(1):170–179, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005760.

Dai:2011:MPS

[DLL11a] Binxiang Dai, Ying Li, and Zhenguo Luo. Multiple periodic solutions for impulsive Gauss-type ratio-dependent predator– prey systems with nonmonotonic numerical responses. Applied Mathematics and Computation, 217(18):7478–7487, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002402.

Duan:2011:SPA

[DLL11b] Xuefeng Duan, Chunmei Li, and Anping Liao. Solutions and perturbation analysis for the nonlinear matrix equation X + ∑_{i=1}^{m} A_{i}^{∗}X^{−1}X_{a} = I. Applied Mathematics and Computation, 218 (8):4458–4466, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012562.

Liu:2013:TKC

[dLL13] Lan dong Liu and Xin Lu.

Two kinds of condition numbers for the quadratic matrix equation. Applied Mathematics and Computation, 219(16):8759–8769, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300180X.

Deng:2014:AOD

[DLL14] Chongyang Deng, Jianzhen Liu, and Yajuan Li. Approximating odd degree Bézier pieces by degreeelevation cutdown polygons. Applied Mathematics and Computation, 239(??):56–59, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005967.

Duan:2013:TCM

[DLLW13] Xuefeng Duan, Chunmei Li, Anping Liao, and Ronghao Wang. On two classes of mixed-type Lyapunov equations. Applied Mathematics and Computation, 219(16):8486–8495, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001768.

Dragone:2010:LSA

[DLP10] Davide Dragone, Luca Lambertini, and Arsen Palestini. The Leitmann– Schmitendorf advertising game with n players and time discounting. Applied Mathematics and Computation, 217(3):1010–1016, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001992.

laRosa:2014:HOH

[dlRS14] M. A. Pérez de la Rosa and M. Shapiro. On the Hilbert operator and the Hilbert formulas on the unit sphere for the timeharmonic Maxwell equations. Applied Mathematics and Computation, 248(??):480–493, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013423.

delaSen:2010:AET

[dlS10] M. de la Sen. Asymptotic expansions for timevarying scalar differential equations using the residue theorem for their discretized difference equations. Applied Mathematics and Computation, 216(1):149–160, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310000378.

Durochat:2013:HON

[DLS13] Clément Durochat, Stéphane Lanteri, and Claire Scheid. High order non-conforming multi-element discontinuous Galerkin method for time domain electromagnetics. Applied Mathematics and Computation, 224(??):681–704, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009375.

delaSen:2011:VSB

[dlSAQ11] M. de la Sen and S. Alonso-Quesada. Vaccination strategies based on feedback control techniques for a general SEIR-epidemic model. Applied Mathematics and Computation, 218(7):3888–3904, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012082.

Ding:2011:PSP

[DLSS11] Xiaohua Ding, Hongyu Liu, Zaijiu Shang, and Geng Sun. Preservation of stability properties near fixed points of linear Hamiltonian systems by symplectic integrators. Applied Mathematics and Computation, 217(13):6105–6114, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013056.

Ding:2013:WPA

[DLX13] Hui-Sheng Ding, Jin Liang, and Ti-Jun Xiao. Weighted pseudo almost periodic functions and applications to evolution equations with delay. Applied Mathematics and Computation, 219(17):8949–8958, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003469.

Ding:2014:DIC

[DLZ14] Xiuyong Ding, Xiu Liu, and Shouming Zhong. Delay-independent criteria for exponential admissibility of switched descriptor delayed systems. Applied Mathematics and Computation, 228(??):432–445, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012824.

Dehmer:2010:IEB

[DM10a] Matthias Dehmer and Abbe Mowshowitz. Inequalities for entropybased measures of network information content. Applied Mathematics and Computation, 215(12):4263–4271, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010984.

Djolovic:2010:HMN

[DM10b] Ivana Djolović and Eberhard Malkowsky. The Hausdorff measure of noncompactness of operators on the matrix domains of triangles in the spaces of strongly C_{1} summable and bounded sequences. Applied Mathematics and Computation, 216(4):1122–1130, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001700.

Dehmer:2011:BMP

[DM11] Matthias Dehmer and Abbe Mowshowitz. Bounds on the moduli of polynomial zeros. Applied Mathematics and Computation, 218(8):4128–4137, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101215X.

Djolovic:2012:MMS

[DM12] Ivana Djolovíc and Eberhard Malkowsky. On matrix mappings into some strong Cesàro sequence spaces. Applied Mathematics and Computation, 218(10):6155–6163, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014500.

Das:2013:CIG

[DM13a] Kedar Nath Das and Rajashree Mishra. Chemoinspired genetic algorithm for function optimization. Applied Mathematics and Computation, 220(??):394–404, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006541.

Deolmi:2013:PIC

[DM13b] G. Deolmi and F. Marcuzzi. A parabolic inverse convection-diffusion-reaction problem solved using space–time localization and adaptivity. Applied Mathematics and Computation, 219(16): 8435–8454, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313001719.

Mo:2011:ISS

[dMbJmY11] Xing de Mo, Jin bao Jian, and Su min Yang. An improved strongly subfeasible SSLE method for optimization problems and numerical experiments. Applied Mathematics and Computation, 217(17):7226–7237, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001706.

Ddumba:2012:RPF

[DMGK12] H. Ddumba, J. Y. T. Mugisha, J. W. Gonsalves, and G. I. H. Kerley. The role of predator fertility and prey threshold bounds on the global and local dynamics of a predator-prey model with a prey out-flux dilution effect. Applied Mathematics and Computation, 218(18):9169–9186, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200224X.

Ddumba:2013:PLC

[DMGK13] H. Ddumba, J. Y. T.

Mugisha, J. W. Gonsalves, and G. I. H. Kerley. Periodicity and limit cycle perturbation analysis of a predator-prey model with interspecific species’ interference, predator additional food and dispersal. Applied Mathematics and Computation, 219(15):8338–8357, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012301.

Detti:2013:LBH

[DMP13] Paolo Detti, Carlo Meloni, and Marco Pranzo. A lower bound on the Hamiltonian path completion number of a line graph. Applied Mathematics and Computation, 220(??):296–304, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006334.

Diele:2010:SAI

[DMR10] Fasma Diele, Carmela Marangi, and Stefania Ragni. SB^{3}A splitting for approximation of invariants in timedependent Hamiltonian systems. Applied Mathematics and Computation, 217(6):2798–2807, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008684.

Das:2011:SEM

[DMS11] Prasenjit Das, Debasis Mukherjee, and A. K. Sarkar. Study of an S-I epidemic model with nonlinear incidence rate: Discrete and stochastic version. Applied Mathematics and Computation, 218(6):2509–2515, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010149.

DeMarchi:2012:PAC

[DMS12] Stefano De Marchi, Martina Marchioro, and Alvise Sommariva. Polynomial approximation and cubature at approximate Fekete and Leja points of the cylinder. Applied Mathematics and Computation, 218(21):10617– 10629, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004225.

Deng:2013:NRS

[DMW13] Feiqi Deng, Weihua Mao, and Anhua Wan. A novel result on stability analysis for uncertain neutral stochastic time-varying delay systems. Applied Mathematics and Computation, 221(??):132–143, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006486.

Diagana:2012:ERS

[DN12a] Toka Diagana and Valerie Nelson. Existence results for some higherorder evolution equations with operator coefficients. Applied Mathematics and Computation, 219(6):2923–2931, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009150.

Dimbour:2012:APS

[DN12b] William Dimbour and Gaston M. N’Guérékata. S-asymptotically ω-periodic solutions to some classes of partial evolution equations. Applied Mathematics and Computation, 218(14):7622–7628, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000653.

Das:2014:OEE

[DN14a] Pratibhamoy Das and Srinivasan Natesan. Optimal error estimate using mesh equidistribution technique for singularly perturbed system of reaction-diffusion boundary-value problems. Applied Mathematics and Computation, 249(??): 265–277, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013897.

Debbouche:2014:STF

[DN14b] Amar Debbouche and Juan J. Nieto. Sobolev type fractional abstract evolution equations with nonlocal conditions and optimal multi-controls. Applied Mathematics and Computation, 245(??):74– 85, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314010303.

Deswita:2010:SSM

[DNI^{+}10] L. Deswita, R. Nazar, A. Ishak, R. Ahmad, and I. Pop. Similarity solutions for mixed convection boundary layer flow over a permeable horizontal flat plate. Applied Mathematics and Computation, 217(6):2619–2630, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008350.

Davoodi:2011:CPE

[DNSM11] F. Davoodi, A. Abbas Nejad, A. Shahrezaee, and M. J. Maghrebi. Control parameter estimation in a semi-linear parabolic inverse problem using a high accurate method. Applied Mathematics and Computation, 218(5):1798–1804, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100909X.

Deniz:2011:UCM

[DO11] Erhan Deniz and Halit Orhan. Univalence criterion for meromorphic functions and Loewner chains. Applied Mathematics and Computation, 218(3):751–755, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002360.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

deOliveira:2013:RPS

[dOL13] Juliano A. de Oliveira and Edson D. Leonel. A rescaling of the phase space for Hamiltonian map: Applications on the Kepler map and mappings with diverging angles in the limit of vanishing action. Applied Mathematics and Computation, 221(??):32–39, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300684X.

Dolfin:2014:BCF

[Dol14]  M. Dolfin.  Boundary

conditions for first order macroscopic models of vehicular traffic in the presence of tollgates. Applied Mathematics and Computation, 234(??):260–266, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400277X.

Dong:2012:PPC

[Don12] Yunda Dong. A practical PR+ conjugate gradient method only using gradient. Applied Mathematics and Computation, 219(4):2041–2052, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008405.

Dong:2014:SCT

[Don14] Xuanchun Dong. Stability and convergence of trigonometric integrator pseudospectral discretization for N-coupled nonlinear Klein–Gordon equations. Applied Mathematics and Computation, 232(??):752–765, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001908.

Doschoris:2011:AMO

[Dos11] Michael Doschoris. An alternative method to obtain integral kernels for Lamé integral equations. Applied Mathematics and Computation, 218(5):1596–1604, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008708.

deOliveira:2013:PFC

[dOS13] Luiz P. L. de Oliveira and Jean Schmith. Ponderomotive force as a cause for chaos in loudspeakers. Applied Mathematics and Computation, 219(12):6449–6456, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312012817.

Dosly:2014:FPR

[Dos14] Ondrej Doslý. Focal points and recessive solutions of discrete symplectic systems. Applied Mathematics and Computation, 243(??):963–968, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008510.

Dou:2010:LTT

[Dou10] Jingbo Dou. Liouville type theorems for the system of integral equations. Applied Mathematics and Computation, 217(6):2586–2594, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008301.

Darbani:2011:MMS

[DOV^{+}11] M. Darbani, A. Ouahsine, P. Villon, H. Naceur, and H. Smaoui. Meshless method for shallow water equations with free surface flow. Applied Mathematics and Computation, 217(11):5113–5124, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007939.

Dix:2010:EMP

[DP10] Julio G. Dix and Seshadev Padhi. Existence of multiple positive periodic solutions for delay differential equation whose order is a multiple of 4. Applied Mathematics and Computation, 216(9):2709–2717, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003802.

Delgado:2013:ACC

[DP13a] J. Delgado and J. M. Peña. Accurate computations with collocation matrices of rational bases. Applied Mathematics and Computation, 219(9):4354–4364, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010065.

Diana:2013:CET

[DP13b] Giancarlo Diana and Pier Francesco Perri. A class of estimators in two-phase sampling with subsampling the nonrespondents. Applied Mathematics and Computation, 219(19):10033– 10043, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003676.

Delgado:2014:FAA

[DP14] Jorge Delgado and Juan Manuel Peña. Fast and accurate algorithms for Jacobi– Stirling matrices. Applied Mathematics and Computation, 236(??):253–259, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004093.

Deng:2010:ESP

[DPC10] Xijun Deng, E. J. Parkes, and Jinlong Cao. Exact solitary and periodicwave solutions of the K(2, 2) equation (defocusing branch). Applied Mathematics and Computation, 217(4):1566–1576, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300309006249.

Dzunic:2011:FOT

[DPP11] Jovana Dzunić, Miodrag S. Petković, and Ljiljana D. Petković. A family of optimal threepoint methods for solving nonlinear equations using two parametric functions. Applied Mathematics and Computation, 217(19):7612–7619, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002463.

Dzunic:2012:TPM

[DPP12] J. Dzuníc, M. S. Petkovíc, and L. D. Petkovíc. Three-point methods with and without memory for solving nonlinear equations. Applied Mathematics and Computation, 218(9):4917–4927, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013038.

Ding:2012:CAP

[DQL^{+}12] Dongsheng Ding, Donglian Qi, Xiaoping Luo, Jinfei Chen, Xuejie Wang, and Pengyin Du. Convergence analysis and performance of an extended central force optimization algorithm. Applied Mathematics and Computation, 219(4):2246–2259, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008648.

DeAbreu:2010:RML

[DR10a] Braulio De Abreu and Marcos Raydan. Residual methods for the largescale matrix p th root and some related problems. Applied Mathematics and Computation, 217(2):650–660, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006764.

Ding:2010:PLL

[DR10b] J. Ding and N. H. Rhee.

Piecewise linear least squares approximations of Frobenius–Perron operators. Applied Mathematics and Computation, 217(7):3257–3262, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009239.

Das:2011:AASa

[DR11a] S. Das and Rajeev. An approximate analytical solution of one-dimensional phase change problems in a finite domain. Applied Mathematics and Computation, 217(13):6040–6046, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010659.

Duan:2011:NHO

[DR11b] Jun-Sheng Duan and Randolph Rach. New higherorder numerical one-step methods based on the Adomian and the modified decomposition methods. Applied Mathematics and Computation, 218(6):2810–2828, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010502.

Duan:2011:NMA

[DR11c] Jun-Sheng Duan and Randolph Rach. A new modification of the Adomian decomposition method for solving boundary value problems for higher order nonlinear differential equations. Applied Mathematics and Computation, 218(8):4090–4118, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012094.

Dragomir:2011:HHT

[Dra11] S. S. Dragomir. Hermite– Hadamard’s type inequalities for operator convex functions. Applied Mathematics and Computation, 218(3):766–772, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100083X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Dragomir:2014:SLT

[Dra14] S. S. Dragomir. Some Lipschitz type inequalities for complex functions. Applied Mathematics and Computation, 230(??):516–529, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013477.

Dey:2010:SAC

[DRGR10] Rajeeb Dey, G. Ray, Sandip Ghosh, and A. Rakshit. Stability analysis for continuous system with additive time-varying delays: a less conservative result. Applied Mathematics and Computation, 215(10):3740–3745, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010030.

See comments [Ge14].

Drignei:2013:NMS

[Dri13] Mihaela-Cristina Drignei. A numerical method for solving a Goursat–Cauchy boundary value problem. Applied Mathematics and Computation, 220(??):123–141, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313005882.

Dai:2011:AMM

[DRJ11] Hong-Liang Dai, Yan-Ni Rao, and Hao-Jie Jiang.

An analytical method for magnetothermoelastic analysis of functionally graded hollow cylinders. Applied Mathematics and Computation, 218(4):1467–1477, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008575.

delRey:2011:RLC

[dRS11a] A. Martín del Rey and G. Rodríguez Sánchez. Reversibility of linear cellular automata. Applied Mathematics and Computation, 217(21):8360–8366, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003997.

Dziok:2011:CFR

[DRS11b] Jacek Dziok, Ravinder Krishna Raina, and Janusz Sokól. On α-convex functions related to shelllike functions connected with Fibonacci numbers. Applied Mathematics and Computation, 218(3):996–1002, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000865.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Diblik:2012:ACSa

[DRS12a] J. Diblík, M. Ruzicková, and Z. Sutá. Asymptotic convergence of the solutions of a discrete equation with several delays. Applied Mathematics and Computation, 218(9):5391–5401, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013531.

Diblik:2012:ACSb

[DRS12b] J. Diblík, M. Ruzicková, and Z. Sutá. Asymptotic convergence of the solutions of a discrete system with delays. Applied Mathematics and Computation, 219(8):4036–4044, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010417.

delRey:2013:ICA

[dRS13] Ángel Martín del Rey and Gerardo Rodríguez Sánchez. On the invertible cellular automata 150 over F_{p}. Applied Mathematics and Computation, 219(10):5427–5432, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012039.

Das:2014:CDT

[DRSG14] Kalyan Das, K. Shiva Reddy, M. N. Srinivas, and N. H. Gazi. Chaotic dynamics of a three species prey-predator competition model with noise in ecology. Applied Mathematics and Computation, 231(??):117–133, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000216.

Dopazo:2011:PGM

[DRT11] E. Dopazo and M. Ruiz-Tagle. A parametric GP model dealing with incomplete information for group decisionmaking. Applied Mathematics and Computation, 218(2):514–519, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007752.

dosSantos:2010:SDD

[dS10] José P. C. dos Santos. On state-dependent delay partial neutral functional integro–differential equations. Applied Mathematics and Computation, 216(5):1637–1644, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002614.

Dere:2011:GPA

[DS11] Rahime Dere and Yilmaz Simsek. Genocchi polynomials associated with the Umbral algebra. Applied Mathematics and Computation, 218(3):756–761, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001135.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Deka:2012:FEM

[DS12a] Bhupen Deka and Rajen Kumar Sinha. Finite element methods for second order linear hyperbolic interface problems. Applied Mathematics and Computation, 218(22):10922–10933, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004547.

Diblik:2012:ESL

[DS12b] Josef Diblík and Ewa Schmeidel. On the existence of solutions of linear Volterra difference equations asymptotically equivalent to a given sequence. Applied Mathematics and Computation, 218(18):9310–9320, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002421.

Darwish:2013:NSQ

[DS13a] Mohamed Abdalla Darwish and Kishin Sadarangani. Nondecreasing solutions of a quadratic Abel equation with supremum in the kernel. Applied Mathematics and Computation, 219(14):7830–7836, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001288.

Dereli:2013:MKB

[DS13b] Yilmaz Dereli and Robert Schaback. The meshless kernel-based method of lines for solving the equal width equation. Applied Mathematics and Computation, 219(10):5224–5232, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011022.

Draganescu:2013:MSD

[DS13c] Andrei Draganescu and Ana Maria Soane. Multigrid solution of a distributed optimal control problem constrained by the Stokes equations. Applied Mathematics and Computation, 219(10):5622–5634, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012374.

Damarackas:2014:BSD

[DS14a] Julius Damarackas and Jonas Siaulys. Biseasonal discrete time risk model. Applied Mathematics and Computation, 247(??):930–940, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401265X.

Darwish:2014:EKT

[DS14b] Mohamed Abdalla Darwish and Kishin Sadarangani. On Erdélyi–Kober type quadratic integral equation with linear modification of the argument. Applied Mathematics and Computation, 238(??):30– 42, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005256.

Das:2014:DFS

[DS14c] Kedar Nath Das and Tapan Kumar Singh. Drosophila Food-Search Optimization. Applied Mathematics and Computation, 231(??):566–580, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000770.

De:2014:MPP

[DS14d] Sujit Kumar De and Shib Sankar Sana. A multi-periods productioninventory model with capacity constraints for multi-manufacturers — a global optimality in intuitionistic fuzzy environment. Applied Mathematics and Computation, 242(??):825–841, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009199.

Dobritoiu:2014:SMS

[DS14e] Maria Dobritoiu and Marcel-Adrian Serban. Step method for a system of integral equations from biomathematics. Applied Mathematics and Computation, 227(??):412–421, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031301206X.

Coelho:2014:SAC

[dSCAM14] Leandro dos Santos Coelho, Helon Vicente Hultmann Ayala, and Viviana Cocco Mariani. A self-adaptive chaotic differential evolution algorithm using gamma distribution for unconstrained global optimization. Applied Mathematics and Computation, 234(??):452–459, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002124.

Deb:2012:AID

[DSGB12] Anish Deb, Gautam Sarkar, Anindita Ganguly, and Amitava Biswas. Approximation, integration and differentiation of time functions using a set of orthogonal hybrid functions (HF) and their application to solution of first order differential equations. Applied Mathematics and Computation, 218(9):4731–4759, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101006X.

Ding:2012:PPS

[DSH12] Xiaoquan Ding, Bentang Su, and Jianmin Hao. Positive periodic solutions for impulsive Gausetype predator-prey systems. Applied Mathematics and Computation, 218(12):6785–6797, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015323.

Deb:2012:TFI

[DSM^{+}12] Anish Deb, Gautam Sarkar, Priyaranjan Mandal, Amitava Biswas, Anindita Ganguly, and Debasish Biswas. Transfer function identification from impulse response via a new set of orthogonal hybrid functions (HF). Applied Mathematics and Computation, 218(9):4760–4787, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011295.

Duan:2013:NGD

[DSPZ13] Xiaomin Duan, Huafei Sun, Linyu Peng, and Xinyu Zhao. A natural gradient descent algorithm for the solution of discrete algebraic Lyapunov equations based on the geodesic distance. Applied Mathematics and Computation, 219(19):9899–9905, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003809.

Deng:2011:PSS

[DSX11] Xiaoqing Deng, Haiping Shi, and Xiaoliang Xie. Periodic solutions of second order discrete Hamiltonian systems with potential indefinite in sign. Applied Mathematics and Computation, 218(1):148–156, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007582.

DeLeone:2012:FSI

[DSZ12] Renato De Leone, Yaroslav D.

Sergeyev, and Anatoly Zhigljavsky. Foreword: Special issue INFINITY 2010. Applied Mathematics and Computation, 218(16):7977–7978, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002731.

Du:2014:ASN

[DSZ14] Lei Du, Tomohiro Sogabe, and Shao-Liang Zhang. An algorithm for solving nonsymmetric pentadiagonal Toeplitz linear systems. Applied Mathematics and Computation, 244(??):10–15, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009485.

Deano:2010:ANA

[DT10] Alfredo Deaño and Nico M. Temme. Analytical and numerical aspects of a generalization of the complementary error function. Applied Mathematics and Computation, 216(12):3680–3693, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005692.

Debbouche:2014:ACF

[DT14] Amar Debbouche and Delfim F. M. Torres. Approximate controllability of fractional delay dynamic inclusions with nonlocal control conditions. Applied Mathematics and Computation, 243(??):161–175, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007875.

Diaz-Toca:2014:DST

[DTN14] Gema María Díaz-Toca and Ioana Necula. Direct symbolic transformation from 3D cartesian into hyperboloidal coordinates. Applied Mathematics and Computation, 228(??):349–365, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301268X.

Dewangan:2014:SCA

[DTP14] Rajshree Dewangan, Balwant Singh Thakur, and Mihai Postolache. Strong convergence of asymptotically pseudocontractive semigroup by viscosity iteration. Applied Mathematics and Computation, 248(??):160–168, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013496.

Du:2011:EUR

[Du11]  Zengji Du.  Existence

and uniqueness results for third-order nonlinear differential systems. Applied Mathematics and Computation, 218(7):2981–2987, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311010241.

Dione:2013:FEA

[DU13] Ibrahima Dione and José. M. Urquiza. Finite element approximations of the Lamé system with penalized ideal contact boundary conditions. Applied Mathematics and Computation, 223(??):115–126, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007765.

Duan:2010:EAM

[Dua10a]  Jun-Sheng Duan.  An

efficient algorithm for the multivariable Adomian polynomials. Applied Mathematics and Computation, 217(6):2456–2467, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300310007915.

Duan:2010:RTA

[Dua10b] Jun-Sheng Duan. Recurrence triangle for Adomian polynomials. Applied Mathematics and Computation, 216(4):1235–1241, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001815.

Duan:2011:CAR

[Dua11a] Jun-Sheng Duan. Convenient analytic recurrence algorithms for the Adomian polynomials. Applied Mathematics and Computation, 217(13):6337–6348, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000117.

Duan:2011:NID

[Dua11b] Jun-Sheng Duan. New ideas for decomposing nonlinearities in differential equations. Applied Mathematics and Computation, 218(5):1774–1784, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311009064.

Dubey:2010:CSC

[Dub10] Ritesh Kumar Dubey. A composite semi-conservative scheme for hyperbolic conservation laws. Applied Mathematics and Computation, 215(9):3335–3342, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009278.

Dubey:2013:FLS

[Dub13] Ritesh Kumar Dubey. Flux limited schemes: Their classification and accuracy based on total variation stability regions. Applied Mathematics and Computation, 224(??):325–336, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008953.

Dukkipati:2012:MEM

[Duk12] Ambedkar Dukkipati. On maximum entropy and minimum KL-divergence optimization by Gröbner basis methods. Applied Mathematics and Computation, 218(23):11674– 11687, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031200567X.

Duman:2010:AES

[Dum10] Melda Duman. Asymptotic expansions for the Sturm–Liouville problem by homotopy perturbation method. Applied Mathematics and Computation, 216(2):492–496, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000688.

dAlmeida:2012:CMD

[dV12a] Filomena D. d’Almeida and Paulo B. Vasconcelos. Convergence of Multipower Defect-Correction for spectral computations of integral operators. Applied Mathematics and Computation, 219(4):1601–1606, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200793X.

DiBari:2012:CFP

[DV12b] Cristina Di Bari and Pasquale Vetro. Common fixed points in generalized metric spaces. Applied Mathematics and Computation, 218(13):7322–7325, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200029X.

DiBari:2013:CFP

[DV13] Cristina Di Bari and Pasquale Vetro. Common fixed points for self mappings on compact metric spaces. Applied Mathematics and Computation, 219(12):6804–6808, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000611.

Diblik:2014:ABS

[DV14] Josef Diblík and Jirí Vítovec. Asymptotic behavior of solutions of systems of dynamic equations on time scales in a set whose boundary is a combination of strict egress and strict ingress points. Applied Mathematics and Computation, 238(??):289–299, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005451.

Das:2011:AASb

[DVGY11] S. Das, K. Vishal, P. K.

Gupta, and A. Yildirim. An approximate analytical solution of timefractional telegraph equation. Applied Mathematics and Computation, 217(18):7405–7411, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002062.

Du:2011:RDQ

[DW11] Liangliang Du and Xionghua Wu. On a rational differential quadrature method in irregular domains for problems with boundary layers. Applied Mathematics and Computation, 218(4):1379–1388, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008472.

Dai:2012:AIW

[DW12] Zhifeng Dai and Fenghua Wen. Another improved Wei–Yao–Liu nonlinear conjugate gradient method with sufficient descent property. Applied Mathematics and Computation, 218(14):7421–7430, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000033.

Dieci:2013:III

[DW13a] Roberto Dieci and Frank Westerhoff. On the inherent instability of international financial markets: natural nonlinear interactions between stock and foreign exchange markets. Applied Mathematics and Computation, 221(??):306–328, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006814.

Duan:2013:FVA

[DW13b] Guohui Duan and Xinwei Wang. Free vibration analysis of multiplestepped beams by the discrete singular convolution. Applied Mathematics and Computation, 219(24):11096–11109, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005432.

Dai:2014:IGS

[DW14] Chao-Qing Dai and Yue-Yue Wang. Infinite generation of soliton-like solutions for complex nonlinear evolution differential equations via the NLSE-based constructive method. Applied Mathematics and Computation, 236(??):606–612, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400438X.

Deng:2011:LRE

[DWC11] Shao-Gao Deng, Qin Wang, and Shijuan Cheng. On the p(x)-Laplacian Robin eigenvalue problem. Applied Mathematics and Computation, 217(12):5643–5649, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310012440.

Ding:2013:PTS

[DWC13] Xiaoquan Ding, Chunwei Wang, and Peng Chen. Permanence for a twospecies Gauss-type ratiodependent predator-prey system with time delay in a two-patch environment. Applied Mathematics and Computation, 219(17):9099–9105, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003627.

Wang:2013:IRT

[dWcFwWjZ13] Xu dong Wang, Xiang chu Feng, Wei wei Wang, and Wen juan Zhang. Iterative reweighted total generalized variation based Poisson noise removal model. Applied Mathematics and Computation, 223(??):264–277, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008461.

Dai:2010:EAL

[DWLL10] Tian-Shyr Dai, Chuan-Ju Wang, Yuh-Dauh Lyuu, and Yen-Chun Liu. An efficient and accurate lattice for pricing derivatives under a jump-diffusion process. Applied Mathematics and Computation, 217(7):3174–3189, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009148.

Dong:2010:NPN

[DWP10] Wei Dong, Zhongli Wei, and Changci Pang. A note on positive nonconstant steady-state solutions to a reaction– diffusion model. Applied Mathematics and Computation, 217(8):4234–4238, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310010933.

Deng:2014:HBA

[DWP14] Lianwang Deng, Xuedi Wang, and Miao Peng. Hopf bifurcation analysis for a ratio-dependent predator-prey system with two delays and stage structure for the predator. Applied Mathematics and Computation, 231(??):214–230, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000629.

Dai:2012:VAT

[DWQG12] H. L. Dai, L. Wang, Q. Qian, and J. Gan. Vibration analysis of threedimensional pipes conveying fluid with consideration of steady combined force by transfer matrix method. Applied Mathematics and Computation, 219(5):2453–2464, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008740.

Dai:2014:NEA

[DWY14] Cai Dai, Yuping Wang, and Miao Ye. A new evolutionary algorithm based on contraction method for many-objective optimization problems. Applied Mathematics and Computation, 245(??):191–205, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010261.

Dong:2012:NDI

[DWZ12] Huanhe Dong, Xiangrong Wang, and Wencai Zhao. A new 4-dimensional implicit vector-form loop algebra with arbitrary constants and the corresponding computing formula of constant γ in the Variation identity. Applied Mathematics and Computation, 218(22):10998–11008, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004602.

Dai:2014:CVM

[DWZW14] Baodong Dai, Qifang Wang, Weiwei Zhang, and Linghui Wang. The complex variable meshless local Petrov–Galerkin method for elastodynamic problems. Applied Mathematics and Computation, 243(??):311–321, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008236.

Deng:2010:SDF

[DX10] Kang Deng and Zhiguang Xiong. Superconvergence of a discontinuous finite element method for a nonlinear ordinary differential equation. Applied Mathematics and Computation, 217(7):3511–3515, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009719.

Dong:2014:EWS

[DX14a] Wei Dong and Jiafa Xu. Existence of weak solutions for a p-Laplacian problem involving Dirichlet boundary condition. Applied Mathematics and Computation, 248(??): 511–518, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314013411.

Du:2014:PPS

[DX14b] Zengji Du and Min Xu. Positive periodic solutions of n-species neutral delayed Lotka– Volterra competition system with impulsive perturbations. Applied Mathematics and Computation, 243(??):379–391, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S009630031400705X.

Ding:2014:QAR

[DXC^{+}14] Shuai Ding, Chengyi Xia, Qiong Cai, Kaile Zhou, and Shanlin Yang. QoSaware resource matching and recommendation for cloud computing systems. Applied Mathematics and Computation, 247(??):941–950, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012831.

Duan:2014:GSS

[DYL14] Xichao Duan, Sanling Yuan, and Xuezhi Li. Global stability of an svir model with age of vaccination. Applied Mathematics and Computation, 226(??):528–540, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011302.

Deng:2013:IDC

[DYYL13] Zui-Cha Deng, Liu Yang, Jian-Ning Yu, and Guan-Wei Luo. Identifying the diffusion coefficient by optimization from the final observation. Applied Mathematics and Computation, 219(9):4410–4422, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010612.

Yang:2013:HMP

[dYyR13] Xu dong Yang and Hang yu Ruan. HBFGen: a Maple package to construct the Hirota bilinear form for nonlinear equations. Applied Mathematics and Computation, 219(15):8018–8025, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001598.

Delahaye:2012:NEA

[DZ12] Jean-Paul Delahaye and Hector Zenil. Numerical evaluation of algorithmic complexity for short strings: a glance into the innermost structure of randomness. Applied Mathematics and Computation, 219(1):63–77, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012367.

Dong:2013:NSL

[DZ13] Yunda Dong and Xue Zhang. New step lengths in projection method for variational inequality problems. Applied Mathematics and Computation, 220(??):239–245, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006280.

Dupal:2014:APS

[DZ14] Jan Dupal and Martin Zajíček. Analytical periodic solution and stability assessment of 1 DOF parametric systems with time varying stiffness. Applied Mathematics and Computation, 243(??):138–151, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007899.

Dziok:2013:AMP

[Dzi13] Jacek Dziok. Applications of multivalent prestarlike functions. Applied Mathematics and Computation, 221(??):230–238, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006930.

Dai:2013:NST

[DZLW13] Baodong Dai, Baojing Zheng, Qingxiang Liang, and Linghui Wang. Numerical solution of transient heat conduction problems using improved meshless local Petrov– Galerkin method. Applied Mathematics and Computation, 219(19):10044– 10052, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004244.

Deng:2014:EGT

[DZS^{+}14] Xinyang Deng, Xi Zheng, Xiaoyan Su, Felix T. S. Chan, Yong Hu, Rehan Sadiq, and Yong Deng. An evidential game theory framework in multicriteria decision making process. Applied Mathematics and Computation, 244(??):783–793, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010224.

Dzunic:2011:SST

[Džu11]  Jovana Džunić.  On

the similarity of some three-point methods for solving nonlinear equations. Applied Mathematics and Computation, 217(14):6633–6635, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000592.

Du:2013:EPB

[DZZ^{+}13] Yuanhua Du, Shouming Zhong, Nan Zhou, Lei Nie, and Wenqin Wang. Exponential passivity of BAM neural networks with time-varying delays. Applied Mathematics and Computation, 221(??):727–740, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300708X.

Du:2014:GAS

[DZZ14] Yuanhua Du, Shouming Zhong, and Nan Zhou. Global asymptotic stability of Markovian jumping stochastic Cohen– Grossberg BAM neural networks with discrete and distributed time-varying delays. Applied Mathematics and Computation, 243(??):624–636, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008571.

Ding:2012:EMS

[DZZZ12] Yucai Ding, Hong Zhu, Shouming Zhong, and Yong Zeng. Exponential mean-square stability of time-delay singular systems with Markovian switching and nonlinear perturbations. Applied Mathematics and Computation, 219(4):2350–2359, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312008600.

El-Abd:2013:IGB

[EA13] Mohammed El-Abd. An improved global-best harmony search algorithm. Applied Mathematics and Computation, 222(??):94– 106, October 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313007704.

Erdogan:2014:SOI

[EA14] A. S. Erdogan and A. Ashyralyev. On the second order implicit difference schemes for a right hand side identification problem. Applied Mathematics and Computation, 226(??):212–228, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010928.

El-Azab:2011:FES

[EAA11] M. S. El-Azab and K. M.

Abdelgaber. Finite element solution of nonlinear diffusion problems. Applied Mathematics and Computation, 217(13):6198–6205, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013226.

Eltaher:2014:VNG

[EAAN^{+}14] M. A. Eltaher, A. A.

Abdelrahman, A. Al-Nabawy, M. Khater, and A. Mansour. Vibration of nonlinear graduation of nano-Timoshenko beam considering the neutral axis position. Applied Mathematics and Computation, 235(??):512–529, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003816.

Elsadany:2013:CDC

[EAE13] A. A. Elsadany, H. N. Agiza, and E. M. Elabbasy. Complex dynamics and chaos control of heterogeneous quadropoly game. Applied Mathematics and Computation, 219(24):11110–11118, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005596.

Eshkuvatov:2010:ASS

[EALS10] Z. K. Eshkuvatov, Anvarjon Ahmedov, N. M. A. Nik Long, and O. Shafiq. Approximate solution of the system of nonlinear integral equation by Newton–Kantorovich method. Applied Mathematics and Computation, 217(8):3717–3725, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001074X.

El-Amin:2010:NBT

[EASK10] M. F. El-Amin, Shuyu Sun, and H. Kanayama. Non-Boussinesq turbulent buoyant jet of a low-density gas leaks into high-density ambient. Applied Mathematics and Computation, 217(8):3764–3778, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009951.

ElSaadi:2012:NTN

[EB12] Nadjia El Saadi and Alassane Bah. Numerical treatment of a nonlocal model for phytoplankton aggregation. Applied Mathematics and Computation, 218(17):8279–8287, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000963.

Elmas:2014:NPT

[EB14] N. Elmas and H. Boyaci.

A new perturbation technique in solution of nonlinear differential equations by using variable transformation. Applied Mathematics and Computation, 227(??):422–427, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012034.

Eleuch:2013:FCM

[EBR13] H. Eleuch, H. Bahlouli, and Y. V. Rostovtsev. A fast converging method for generating solutions to the Riccati equation. Applied Mathematics and Computation, 222(??):548–552, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008394.

Elliotis:2014:SFB

[ECG14] Miltiades C. Elliotis, Dimos C. Charmpis, and Georgios C. Georgiou. The singular function boundary integral method for an elastic plane stress wedge beam problem with a point boundary singularity. Applied Mathematics and Computation, 248(??):93–100, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013150.

Echi:2012:RSR

[Ech12] Nadhem Echi. Rational solutions of Riccati differential equation with coefficients rational. Applied Mathematics and Computation, 218(21):10341– 10366, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004304.

Elliott:2013:OVU

[ECS13] Robert J. Elliott, Leunglung Chan, and Tak Kuen Siu. Option valuation under a regime-switching constant elasticity of variance process. Applied Mathematics and Computation, 219(9):4434–4443, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010636.

El-Daou:2010:ITM

[EDAM10] Mohamed K. El-Daou and Nadia R. Al-Matar. An improved Tau method for a class of Sturm–Liouville problems. Applied Mathematics and Computation, 216(7):1923–1937, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000264X.

Echevarria:2014:FDI

[EdCVB^{+}14] Lídice Camps Echevarría, Haroldo Fraga de Campos Velho, José Carlos Becceneri, Antônio José da Silva Neto, and Orestes Llanes Santiago. The fault diagnosis inverse problem with ant colony optimization and ant colony optimization with dispersion. Applied Mathematics and Computation, 227(??):687–700, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012319.

Erkus-Duman:2011:SAP

[EDD11] Esra Erkus-Duman and Oktay Duman. Statistical approximation properties of high order operators constructed with the Chan–Chyan–Srivastava polynomials. Applied Mathematics and Computation, 218(5):1927–1933, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311009210.

El-Desouky:2014:NUF

[EDG14] B. S. El-Desouky and R. S.

Gomaa. A new unified family of generalized Apostol–Euler, Bernoulli and Genocchi polynomials. Applied Mathematics and Computation, 247(??):695–702, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012181.

El-Desouky:2014:AMN

[EDGC14] B. S. El-Desouky, R. S.

Gomaa, and Nenad P. Cakić. q-analogues of multiparameter non-central Stirling and generalized harmonic numbers. Applied Mathematics and Computation, 232(??): 132–143, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314000630.

El-Dessoky:2014:BAC

[EDYA14] M. M. El-Dessoky, M. T. Yassen, and E. S. Aly. Bifurcation analysis and chaos control in Shimizu– Morioka chaotic system with delayed feedback. Applied Mathematics and Computation, 243(??): 283–297, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007371.

El-Dessoky:2012:EHH

[EDYSA12] M. M. El-Dessoky, M. T. Yassen, E. Saleh, and E. S. Aly. Existence of heteroclinic and homoclinic orbits in two different chaotic dynamical systems. Applied Mathematics and Computation, 218(24):11859–11870, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005632.

See comment [AFSMRL14].

ElMoumen:2011:NHM

[EEA11] Samira El Moumen, Rachid Ellaia, and Rajae Aboulaich. A new hybrid method for solving global optimization problem. Applied Mathematics and Computation, 218(7):3265–3276, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311011064.

Eldabe:2005:CFD

[EEEE05] N. T. Eldabe, E. F. Elshehawey, Elsayed M. E. Elbarbary, and Nasser S. Elgazery. Chebyshev finite difference method for MHD flow of a micropolar fluid past a stretching sheet with heat transfer. Applied Mathematics and Computation, 160(2):437–450, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030301172X.

See comment [Pan12b].

Eltaher:2012:FVA

[EEM12] M. A. Eltaher, Samir A.

Emam, and F. F. Mahmoud. Free vibration analysis of functionally graded size-dependent nanobeams. Applied Mathematics and Computation, 218(14):7406–7420, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000021.

ElMouatasim:2014:SPR

[EES14] Abdelkrim El Mouatasim, Rachid Ellaia, and Eduardo Souza de Cursi. Stochastic perturbation of reduced gradient & grg methods for nonconvex programming problems. Applied Mathematics and Computation, 226(??):198–211, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010813.

Elabbasy:2014:BAC

[EEZ14] E. M. Elabbasy, A. A. Elsadany, and Yue Zhang. Bifurcation analysis and chaos in a discrete reduced Lorenz system. Applied Mathematics and Computation, 228(??):184–194, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012575.

Eren:2006:BFS

[EG06] Tamer Eren and Ertan Güner. A bicriteria flowshop scheduling problem with setup times. Applied Mathematics and Computation, 183(2):1292–1300, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030600676X.

See comments [XY11].

Erel:2011:MWM

[EG11] Erdal Erel and Jay B.

Ghosh. Minimizing weighted mean absolute deviation of job completion times from their weighted mean. Applied Mathematics and Computation, 217(22):9340–9350, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005777.

Ezquerro:2012:VNK

[EGH12] J. A. Ezquerro, D. González, and M. A. Hernández. A variant of the Newton– Kantorovich theorem for nonlinear integral equations of mixed Hammerstein type. Applied Mathematics and Computation, 218(18):9536–9546, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002810.

ElDoussouki:2013:SRR

[EGJB13] A. El Doussouki, M. Guedda, M. Jazar, and M. Benlahsen. Some remarks on radial solutions of Föppl–von Kármán equations. Applied Mathematics and Computation, 219(9):4340–4345, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010016.

El-Gamel:2014:SGM

[EGMEM14] Mohamed El-Gamel, Adel Mohsen, and Amgad Abd El-Mohsen. Sinc–Galerkin method for solving biharmonic problems. Applied Mathematics and Computation, 247(??):386–396, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011965.

Enguix-Gonzalez:2012:UCB

[EGMPMRBC12] A. Enguix-González, J. M. Muñoz-Pichardo, J. L. Moreno-Rebollo, and I. Barranco-Chamorro. Using conditional bias in principal component analysis for the evaluation of joint influence on the eigenvalues of the covariance matrix. Applied Mathematics and Computation, 218(17):8937–8950, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001944.

Esteves:2013:GES

[EGO13] Salete Esteves, Elçin Gökmen, and José J. Oliveira. Global exponential stability of nonautonomous neural network models with continuous distributed delays. Applied Mathematics and Computation, 219(17):9296–9307, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002889.

El-Gebeily:2010:TOA

[EGOM10] Mohamed A. El-Gebeily, Donal O’Regan, and Salim Messauodi. Type I operators and the approximation of singular two-point boundary value problems. Applied Mathematics and Computation, 216(12):3433–3438, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310005643.

Elhosseini:2014:MHA

[EH14a] Mostafa A. Elhosseini and Amira Y. Haikal. Modified hybrid algorithm for process optimization. Applied Mathematics and Computation, 232(??):1209–1224, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001623.

See retraction notice [EH15].

Esmaeilbeigi:2014:NAB

[EH14b] M. Esmaeilbeigi and M. M. Hosseini. A new approach based on the genetic algorithm for finding a good shape parameter in solving partial differential equations by Kansa’s method. Applied Mathematics and Computation, 249(??):419–428, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300314013782.

Elhosseini:2015:RNM

[EH15] Mostafa A. Elhosseini and Amira Y. Haikal. Retraction notice to “Modified Hybrid Algorithm for Process Optimization” [AMC 232 (2014) 1209– 1224]. Applied Mathematics and Computation, 266(??):1177, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300315009467.

See [EH14a].

Ezquerro:2012:IDS

[EHRV12] J. A. Ezquerro, M. A.

Hernández, N. Romero, and A. I. Velasco. Improving the domain of starting points for secant-like methods. Applied Mathematics and Computation, 219(8):3677–3692, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009824.

Eltaher:2014:MAH

[EHSM14] M. A. Eltaher, M. A.

Hamed, A. M. Sadoun, and A. Mansour. Mechanical analysis of higher order gradient nanobeams. Applied Mathematics and Computation, 229(??): 260–272, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313013593.

Ervin:2011:SAD

[EJ11] V. J. Ervin and E. W. Jenkins. Stabilized approximation to degenerate transport equations via filtering. Applied Mathematics and Computation, 217(17):7282–7294, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001962.

Eftekhari:2012:HAM

[EJ12] S. A. Eftekhari and A. A. Jafari. High accuracy mixed finite element-Ritz formulation for free vibration analysis of plates with general boundary conditions. Applied Mathematics and Computation, 219(3):1312–1344, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312007394.

El-Kalla:2010:NRA

[EK10] I. L. El-Kalla. New results on the analytic summation of Adomian series for some classes of differential and integral equations. Applied Mathematics and Computation, 217(8):3756–3763, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000994X.

Eigoli:2011:HAM

[EK11] Ali Kamali Eigoli and Mohammad Khodabakhsh. A homotopy analysis method for limit cycle of the van der Pol oscillator with delayed amplitude limiting. Applied Mathematics and Computation, 217(22):9404–9411, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005868.

Et:2014:SGP

[EKK14] Mikail Et, Murat Karakas, and Vatan Karakaya. Some geometric properties of a new difference sequence space defined by de la Vallée Poussin mean. Applied Mathematics and Computation, 234(??):237–244, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001672.

Eltaher:2014:SBA

[EKSO14] M. A. Eltaher, A. Khairy, A. M. Sadoun, and Fatema-Alzahraa Omar. Static and buckling analysis of functionally graded Timoshenko nanobeams. Applied Mathematics and Computation, 229(??): 283–295, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313013556.

Ezzoug:2011:FDG

[EKZ11] Emna Ezzoug, Wided Kechiche, and Ezzeddine Zahrouni. Finite dimensional global attractor for a semi-discrete nonlinear Schrödinger equation with a point defect. Applied Mathematics and Computation, 217(19):7818–7830, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003195.

Bouajaji:2013:HOD

[EL13] Mohamed El Bouajaji and Stéphane Lanteri. High order discontinuous Galerkin method for the solution of 2D timeharmonic Maxwell’s equations. Applied Mathematics and Computation, 219(13):7241–7251, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005200.

Elsayed:2014:RPD

[EL14] Khairy Elsayed and Chris Lacor. Robust parameter design optimization using kriging, RBF and RBFNN with gradient-based and evolutionary optimization techniques. Applied Mathematics and Computation, 236(??):325–344, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004445.

Elboree:2012:NSS

[Elb12] Mohammed K. Elboree.

New soliton solutions for a Kadomtsev–Petviashvili (KP) like equation coupled to a Schrödinger equation. Applied Mathematics and Computation, 218(10):5966–5973, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014172.

Esnaola:2011:CSM

[ELIGMB11] E. Berriochoa Esnaola, A. Cachafeiro López, J. R. Illán-González, and E. Martínez-Brey. Chebyshev series method for computing weighted quadrature formulas. Applied Mathematics and Computation, 218(8):4437–4447, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311012549.

Elouafi:2011:EFD

[Elo11] Mohamed Elouafi. An explicit formula for the determinant of a skewsymmetric pentadiagonal Toeplitz matrix. Applied Mathematics and Computation, 218(7):3466–3469, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311011325.

Elouafi:2013:NEF

[Elo13] Mohamed Elouafi. A note for an explicit formula for the determinant of pentadiagonal and heptadiagonal symmetric Toeplitz matrices. Applied Mathematics and Computation, 219(9):4789–4791, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011344.

Elouafi:2014:RBC

[Elo14] Mohamed Elouafi. On a relationship between Chebyshev polynomials and Toeplitz determinants. Applied Mathematics and Computation, 229(??):27–33, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013052.

Elliott:2011:FET

[ELS11] Robert J. Elliott, Chuin Ching Liew, and Tak Kuen Siu. On filtering and estimation of a threshold stochastic volatility model. Applied Mathematics and Computation, 218(1):61–75, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007223.

Elsaid:2012:FDT

[Els12] A. Elsaid. Fractional differential transform method combined with the Adomian polynomials. Applied Mathematics and Computation, 218(12):6899–6911, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311015530.

Estevez:2013:IID

[ELS13] P. G. Estévez, J. D. Lejarreta, and C. Sardón. Integrable 1 + 1 dimensional hierarchies arising from reduction of a nonisospectral problem in 2+1 dimensions. Applied Mathematics and Computation, 224(??):311–324, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313009107.

El-Mistikawy:2011:CGS

[EM11] Tarek M. A. El-Mistikawy.

Comment on “Group solution of a time dependent chemical convective process” by M. M. Kassem and A. S. Rashed, Applied Mathematics and Computation, 215 (2009) 1671– 1684. Applied Mathematics and Computation, 217(12):5999–6003, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012609.

See [?].

Elettreby:2012:CDM

[EM12] M. F. Elettreby and M. Mansour. On Cournot dynamic multi-team game using incomplete information dynamical system. Applied Mathematics and Computation, 218(21): 10691–10696, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004377.

Ehrhardt:2013:NFD

[EM13] Matthias Ehrhardt and Ronald E. Mickens. A nonstandard finite difference scheme for convectiondiffusion equations having constant coefficients. Applied Mathematics and Computation, 219(12): 6591–6604, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031300009X.

Abd-el-Malek:2013:LGM

[eMA13a] Mina B. Abd el Malek and Amr M. Amin. Lie group method for solving generalized Hirota–Satsuma coupled Korteweg–de Vries (KdV) equations. Applied Mathematics and Computation, 224(??):501–516, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313009387.

El-Mikkawy:2013:DII

[EMA13b] Moawwad El-Mikkawy and Faiz Atlan. Derivation of identities involving some special polynomials and numbers via generating functions with applications. Applied Mathematics and Computation, 220(??):518–535, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006267.

El-Mikkawy:2013:RTS

[EMA13c] Moawwad El-Mikkawy and Faiz Atlan. Remarks on two symmetric polynomials and some matrices. Applied Mathematics and Computation, 219(16):8770–8778, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002063.

Emam:2013:IAB

[Ema13d] O. E. Emam. Interactive approach to bi-level integer multi-objective fractional programming problem. Applied Mathematics and Computation, 223(??):17–24, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008357.

Eltaher:2013:CEN

[EMAM13] M. A. Eltaher, F. F.

Mahmoud, A. E. Assie, and E. I. Meletis. Coupling effects of nonlocal and surface energy on vibration analysis of nanobeams. Applied Mathematics and Computation, 224(??):760–774, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313009600.

Abd-el-Malek:2013:NSS

[eMBHA13] Mina B. Abd el Malek, Nagwa A. Badran, Hossam S. Hassan, and Heba H. Abbas. New solutions for solving problem of particle trajectories in linear deepwater waves via Lie-group method. Applied Mathematics and Computation, 219(24):11365–11375, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006061.

Elmojtaba:2010:MAD

[EMH10] Ibrahim M. Elmojtaba, J. Y. T. Mugisha, and Mohsin H. A. Hashim. Mathematical analysis of the dynamics of visceral leishmaniasis in the Sudan. Applied Mathematics and Computation, 217(6):2567–2578, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008283.

Edely:2013:APF

[EMK13] Osama H. H. Edely, M. Mursaleen, and Asif Khan. Approximation for periodic functions via weighted statistical convergence. Applied Mathematics and Computation, 219(15):8231–8236, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001458.

Enatsu:2012:SAD

[EMM^{+}12] Yoichi Enatsu, Eleonora Messina, Yoshiaki Muroya, Yukihiko Nakata, Elvira Russo, and Antonia Vecchio. Stability analysis of delayed SIR epidemic models with a class of nonlinear incidence rates. Applied Mathematics and Computation, 218(9):5327–5336, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311013464.

Erbe:2012:SPP

[EMP12] Lynn Erbe, Raziye Mert, and Allan Peterson. Spectral parameter power series for Sturm–Liouville equations on time scales. Applied Mathematics and Computation, 218(14): 7671–7678, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000756.

El-Mikkawy:2010:NFF

[EMS10a] Moawwad El-Mikkawy and Tomohiro Sogabe. A new family of k-Fibonacci numbers. Applied Mathematics and Computation, 215(12):4456–4461, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000007X.

El-Mikkawy:2010:NPS

[EMS10b] Moawwad El-Mikkawy and Tomohiro Sogabe. Notes on particular symmetric polynomials with applications. Applied Mathematics and Computation, 215(9):3311–3317, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009242.

El-Nabulsi:2011:FWD

[EN11] Ahmad Rami El-Nabulsi.

The fractional white dwarf hydrodynamical nonlinear differential equation and emergence of quark stars. Applied Mathematics and Computation, 218(6):2837–2849, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101054X.

Erdonmez:2010:OSA

[EÖ10] Ülker Erdönmez and İbrahim Özkol. Optimal shape analysis of a column structure under various loading conditions by using differential transform method (DTM). Applied Mathematics and Computation, 216(11):3172–3183, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004625.

Easwaran:2010:MMM

[EPMU10] Gopalakrishnan Easwaran, Larry E. Parten, Rafael Moras, and Paul X. Uhlig. Makespan minimization in machine dominated flowshop. Applied Mathematics and Computation, 217(1):110–116, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310005667.

Endler:2013:MLS

[ER13] Antonio Endler and Paulo C. Rech. From Mandelbrotlike sets to Arnold tongues. Applied Mathematics and Computation, 222(??): 559–563, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008369.

Erencin:2011:DTM

[Ere11] Aysegül Erençin. Durrmeyer type modification of generalized Baskakov operators. Applied Mathematics and Computation, 218(8):4384–4390, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300311012446.

Eryilmaz:2014:PCF

[Ery14] Serkan Eryilmaz. Parallel and consecutive-k-out-of-n:F systems under stochastic deterioration. Applied Mathematics and Computation, 227(??):19–26, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011387.

El-Sobky:2012:MAS

[ES12a]  Bothina El-Sobky.  A

multiplier active-set trustregion algorithm for solving constrained optimization problem. Applied Mathematics and Computation, 219(3):928–946, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312006959.

Emmert-Streib:2012:EDS

[ES12b]  Frank  Emmert-Streib.

Evolutionary dynamics of the spatial Prisoner’s Dilemma with self-inhibition. Applied Mathematics and Computation, 218(11): 6482–6488, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014779.

Emmert-Streib:2012:UCM

[ES12c]  Frank  Emmert-Streib.

Universal construction mechanism for networks from one-dimensional symbol sequences. Applied Mathematics and Computation, 219(3):1020–1030, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007060.

Essman:2012:BDC

[ES12d] Michael Essman and Junping Shi. Bifurcation diagrams of coupled Schrödinger equations. Applied Mathematics and Computation, 219(8):3646–3654, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009605.

El-Shehawey:2010:ECA

[ESAH10] M. A. El-Shehawey and A. S. Al-Henawy. Erratum to “Computer aided solving the higher-order transition probability matrix of the finite Markov chain” [Appl. Math. Comput. 172 (2006) 267– 285]. Applied Mathematics and Computation, 217(8):4311–4312, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000963X.

See [LW06].

Elsayed:2013:ACE

[ESE13] Saber M. Elsayed, Ruhul A.

Sarker, and Daryl L. Essam. Adaptive Configuration of evolutionary algorithms for constrained optimization. Applied Mathematics and Computation, 222(??):680–711, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008187.

Elsayed:2014:SAC

[ESE14] Saber M. Elsayed, Ruhul A.

Sarker, and Daryl L. Essam. A self-adaptive combined strategies algorithm for constrained optimization using differential evolution. Applied Mathematics and Computation, 241(??):267–282, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006742.

El-Sayed:2012:HPT

[ESEEKH12] A. M. A. El-Sayed, A. Elsaid, I. L. El-Kalla, and D. Hammad. A homotopy perturbation technique for solving partial differential equations of fractional order in finite domains. Applied Mathematics and Computation, 218(17):8329–8340, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001038.

Esfahani:2011:ZED

[Esf11a] Amin Esfahani. On the ZK equation with a directional dissipation. Applied Mathematics and Computation, 217(10):4911–4927, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011689.

Esfahani:2011:RSW

[Esf11b]  Amin Esfahani.  Re-

marks on solitary waves of the generalized two dimensional Benjamin–Ono equation. Applied Mathematics and Computation, 218(2):308–323, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007466.

El-Shahed:2011:TDD

[ESG11] Moustafa El-Shahed and Mohammed Gaber. Twodimensional q-differential transformation and its application. Applied Mathematics and Computation, 217(22):9165–9172, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005327.

El-Sayed:2010:MPS

[ESH10] A. M. A. El-Sayed and H. H. G. Hashem. Monotonic positive solution of a nonlinear quadratic functional integral equation. Applied Mathematics and Computation, 216(9):2576–2580, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000353X.

Esmaeili:2012:ERM

[Esm12]  Mehrab Esmaeili.  An

Enhanced Russell Measure in DEA with interval data. Applied Mathematics and Computation, 219(4):1589–1593, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007904.

El-Sayed:2014:CRB

[ESNE^{+}14] A. M. A. El-Sayed, H. M. Nour, A. Elsaid, A. E. Matouk, and A. Elsonbaty. Circuit realization, bifurcations, chaos and hyperchaos in a new 4D system. Applied Mathematics and Computation, 239(??):333–345, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006560.

Echebest:2012:DFM

[ESV12] N. Echebest, M. L. Schuverdt, and R. P. Vignau. A derivative-free method for solving box-constrained underdetermined nonlinear systems of equations. Applied Mathematics and Computation, 219(6):3198–3208, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009551.

Erb:2011:AME

[ET11] Wolfgang Erb and Ferenc Toókos. Applications of the monotonicity of extremal zeros of orthogonal polynomials in interlacing and optimization problems. Applied Mathematics and Computation, 217(9):4771–4780, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011446.

Tatar:2012:UDV

[eT12a] Nasser eddine Tatar. Uniform decay in viscoelasticity for kernels with small non-decreasingness zones. Applied Mathematics and Computation, 218(15):7939–7946, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001336.

Eryilmaz:2012:RAT

[ET12b] Serkan Eryilmaz and Fatih Tank. On reliability analysis of a two-dependent-unit series system with a standby unit. Applied Mathematics and Computation, 218(15):7792–7797, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000847.

Et:2013:GCD

[Et13] Mikail Et. Generalized Cesàro difference sequence spaces of non-absolute type involving lacunary sequences. Applied Mathematics and Computation, 219(17):9372–9376, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002920.

Etoa:2011:SQC

[Eto11] Jean Bosco Etoa Etoa. Solving quadratic convex bilevel programming problems using a smoothing method. Applied Mathematics and Computation, 217(15):6680–6690, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001019.

Elezovic:2014:AEI

[EV14] Neven Elezović and Lenka Vuksić. Asymptotic expansions of integral means and applications to the ratio of gamma functions. Applied Mathematics and Computation, 235(??):187–200, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002653.

Escobar-Vargas:2012:HOD

[EVDG12] J. A. Escobar-Vargas, P. J.

Diamessis, and F. X. Giraldo. High-order discontinuous element-based schemes for the inviscid shallow water equations: Spectral multidomain penalty and discontinuous Galerkin methods. Applied Mathematics and Computation, 218(9):4825–4848, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012914.

Estrada:2012:DSH

[EVE12] Ernesto Estrada and Eusebio Vargas-Estrada. Distance-sum heterogeneity in graphs and complex networks. Applied Mathematics and Computation, 218(21):10393– 10405, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312003542.

Esmaily:2012:CPT

[EVR12] Jafar Esmaily, S. Mansour Vaezpour, and B. E. Rhoades. Coincidence point theorem for generalized weakly contractions in ordered metric spaces. Applied Mathematics and Computation, 219(4):1536–1548, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007849.

Esmaily:2013:CCF

[EVR13] Jafar Esmaily, S. Mansour Vaezpour, and B. E.

Rhoades. Coincidence and common fixed point theorems for a sequence of mappings in ordered metric spaces. Applied Mathematics and Computation, 219(10):5684–5692, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201168X.

Emmert:2010:CTS

[EWS10] Keith E. Emmert, Peter White, and Katy Sims. Continuous-time, stagestructured, multiple-species model with applications to amphibians. Applied Mathematics and Computation, 216(4):1109–1121, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001694.

Enayatifar:2013:MNM

[EYAD13] Rasul Enayatifar, Moslem Yousefi, Abdul Hanan Abdullah, and Amer Nordin Darus. MOICA: a novel multi-objective approach based on imperialist competitive algorithm. Applied Mathematics and Computation, 219(17):8829–8841, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003603.

Elias-Zuniga:2012:ESQ

[EZ12] Alex Elías-Zúñiga. Exact solution of the quadratic mixed-parity Helmholtz– Duffing oscillator. Applied Mathematics and Computation, 218(14):7590–7594, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000604.

Ertem:2013:AIS

[EZ13a] T. Ertem and A. Zafer. Asymptotic integration of second-order nonlinear differential equations via principal and nonprincipal solutions. Applied Mathematics and Computation, 219(11):5876–5886, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012878.

Ezzati:2013:NSN

[EZ13b] R. Ezzati and S. Ziari. Numerical solution of nonlinear fuzzy Fredholm integral equations using iterative method. Applied Mathematics and Computation, 225(??):33–42, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009922.

Elias-Zuniga:2014:QMO

[EZ14a] Alex Elías-Zúñiga. “Quintication” method to obtain approximate analytical solutions of non-linear oscillators. Applied Mathematics and Computation, 243(??):849–855, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007851.

Elias-Zuniga:2014:SDC

[EZ14b] Alex Elías-Zúñiga. Solution of the damped cubicquintic Duffing oscillator by using Jacobi elliptic functions. Applied Mathematics and Computation, 246(??):474–481, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314010790.

Elias-Zuniga:2012:AEB

[EZRMR12] Alex Elías-Zúñiga, Ciro A. Rodríguez, and Oscar Martínez-Romero. Application of the elliptic balance method to a nonlinear singular oscillator. Applied Mathematics and Computation, 218(22):11112–11117, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004821.

Ferchichi:2010:NSD

[FA10] Hanen Ferchichi and Saloua Mani Aouadi. Numerical stability for a dynamic Naghdi shell. Applied Mathematics and Computation, 216(9):2483– 2506, July 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003528.

Faydasicok:2012:ESA

[FA12] Ozlem Faydasicok and Sabri Arik. Equilibrium and stability analysis of delayed neural networks under parameter uncertainties. Applied Mathematics and Computation, 218(12):6716–6726, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015104.

Falcao:2014:NMC

[Fal14] Maria Irene Falcão. Newton method in the context of quaternion analysis. Applied Mathematics and Computation, 236(??):458–470, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004123.

Farooq:2011:TDS

[FAM11] M. Umar Farooq, S. Ali, and F. M. Mahomed. Twodimensional systems that arise from the Noether classification of Lagrangians on the line. Applied Mathematics and Computation, 217(16):6959–6973, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001433.

Fan:2010:DTW

[Fan10a]  Zhencheng Fan.  Dis-

crete time waveform relaxation method for stochastic delay differential equations. Applied Mathematics and Computation, 217(8):3903–3909, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031001012X.

Fang:2010:NOS

[Fan10b]  Liang Fang.  A new

one-step smoothing Newton method for nonlinear complementarity problem with P_{0}-function. Applied Mathematics and Computation, 216(4):1087–1095, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001670.

Fan:2013:SWR

[Fan13]  Zhencheng Fan.  SOR

waveform relaxation methods for stochastic differential equations. Applied Mathematics and Computation, 219(10):4992–5003, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012222.

Fan:2014:PFE

[Fan14a] Ye Fan. Pattern formation of an epidemic model with cross diffusion. Applied Mathematics and Computation, 228(??):311–319, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012599.

Fan:2014:CCR

[Fan14b]  Zhenbin Fan.  Char-

acterization of compactness for resolvents and its applications. Applied Mathematics and Computation, 232(??):60–67, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000885.

Fang:2014:NQC

[Fan14c] Jian-Ping Fang. Note on a q-contour integral formula. Applied Mathematics and Computation, 233(??):292–297, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400201X.

Fang:2014:DEP

[Fan14d]  Jian-Ping Fang.  q-

difference equation and q-polynomials. Applied Mathematics and Computation, 248(??):550–561, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013769.

Fang:2014:SSF

[Fan14e] Qingxiang Fang. Synchronization and state feedback control of linearly coupled singular systems. Applied Mathematics and Computation, 232(??):381–390, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001295.

Faria:2014:GDL

[Far14a] Teresa Faria. Global dynamics for Lotka–Volterra systems with infinite delay and patch structure. Applied Mathematics and Computation, 245(??):575–590, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011035.

Faria:2014:NPN

[Far14b] Teresa Faria. A note on permanence of nonautonomous cooperative scalar population models with delays. Applied Mathematics and Computation, 240(??):82–90, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005803.

Faustino:2014:CHP

[Fau14] N. Faustino. Classes of hypercomplex polynomials of discrete variable based on the quasi-monomiality principle. Applied Mathematics and Computation, 247(??):607–622, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012454.

Frost:2010:ESA

[FB10] S. A. Frost and M. J.

Balas. Evolving systems:

Adaptive key component control and inheritance of passivity and dissipativity. Applied Mathematics and Computation, 217(3):1034–1044, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004534.

Florez:2012:LMC

[FBGH12] W. F. Florez, C. A. Bustamante, M. Giraldo, and A. F. Hill. Local mass conservative Hermite interpolation for the solution of flow problems by a multi-domain boundary element approach. Applied Mathematics and Computation, 218(11):6446–6457, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014731.

Fedele:2010:RSF

[FC10] Giuseppe Fedele and Loredana Coluccio. A recursive scheme for frequency estimation using the modulating functions method. Applied Mathematics and Computation, 216(5):1393–1400, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002110.

Fang:2014:SEA

[FC14] Changjie Fang and Shenglan Chen. A subgradient extragradient algorithm for solving multivalued variational inequality. Applied Mathematics and Computation, 229(??):123–130, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013155.

Feng:2010:HFS

[FcDDqZ10] Yang Feng, Yan cheng Dong, Qi Ding, and Hong qing Zhang. Hyperelliptic functions solutions of some nonlinear partial differential equations using the direct method. Applied Mathematics and Computation, 215(11):3868–3873, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010194.

Fakharany:2014:PFD

[FCJ14] M. Fakharany, R. Company, and L. Jódar. Positive finite difference schemes for a partial integro-differential option pricing model. Applied Mathematics and Computation, 249(??):320–332, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314014398.

Fu:2010:NES

[FCZ^{+}10] Zuntao Fu, Zhe Chen, Linna Zhang, Jiangyu Mao, and Shikuo Liu. Novel exact solutions to the short pulse equation. Applied Mathematics and Computation, 215(11):3899–3905, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010273.

Fan:2012:ABO

[FD12a] Mingshu Fan and Lili Du. Asymptotic behavior for an Ohmic heating model in thermal electricity. Applied Mathematics and Computation, 218(22):10906–10913, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004523.

Deng:2012:DMS

[fD12b] Shu fang Deng. The decay mode solutions for the cylindrical KP equation. Applied Mathematics and Computation, 218(10):5974–5981, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014226.

Fullmer:2014:ASV

[FdBCC14] William D. Fullmer, Martin A. Lopez de Bertodano, Min Chen, and Alejandro Clausse. Analysis of stability, verification and chaos with the Kreiss– Yström equations. Applied Mathematics and Computation, 248(??):28–46, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012995.

Fang:2014:FID

[FDHK14] Yong Fang, Huanhe Dong, Yijun Hou, and Yuan Kong. Frobenius integrable decompositions of nonlinear evolution equations with modified term. Applied Mathematics and Computation, 226(??):435–440, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011041.

Ferreira:2014:NSA

[FdOdSM14] J. A. Ferreira, P. de Oliveira, P. M. da Silva, and J. N. Murta. Numerical simulation of aqueous humor flow: From healthy to pathologic situations. Applied Mathematics and Computation, 226(??):777–792, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011272.

Ding:2012:CDS

[fDxZxChT12] Heng fei Ding, Yu xin Zhang, Jian xiong Cao, and Jun hong Tian. A class of difference scheme for solving telegraph equation by new non-polynomial spline methods. Applied Mathematics and Computation, 218(9):4671–4683, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013245.

Febbo:2011:FEN

[Feb11] M. Febbo. A finite extensibility nonlinear oscillator. Applied Mathematics and Computation, 217(14):6464–6475, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000257.

See [BAFP12].

Fathy:2014:LGM

[FEGEA14] Mohamed Fathy, Mohamed El-Gamel, and M. S. El-Azab. Legendre– Galerkin method for the linear Fredholm integrodifferential equations. Applied Mathematics and Computation, 243(??): 789–800, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031400900X.

Feng:2011:FID

[Fen11] Jishe Feng. Fibonacci identities via the determinant of tridiagonal matrix. Applied Mathematics and Computation, 217(12):5978–5981, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012270.

Feng:2013:NCP

[Fen13] Jishe Feng. A note on computing of positive integer powers for circulant matrices. Applied Mathematics and Computation, 223(??):472–475, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008655.

Fernandez:2010:PTM

[Fer10a] Francisco M. Fernández. On a perturbation treatment of a model for MHD viscous flow. Applied Mathematics and Computation, 217(6):2307–2310, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310007733.

Fernandez:2010:SAA

[Fer10b] Francisco M. Fernández.

On the small-amplitude approximation to the differential equation ẍ + (1 + ẋ^{2})x = 0. Applied Mathematics and Computation, 216(10):3111–3113, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003929.

Fernandez:2010:COB

[Fer10c] Julio C. Ramos Fernández.

Composition operators on Bloch–Orlicz type spaces. Applied Mathematics and Computation, 217(7):3392–3402, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009513.

Fernandez:2011:RAT

[Fer11] Francisco M. Fernández.

Rational approximation to the Thomas–Fermi equations. Applied Mathematics and Computation, 217(13):6433–6436, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311000658.

Fericean:2012:SLF

[Fer12a] Denisa Fericean. Strongly Φ-like functions of order α in two-dimensional free boundary problems. Applied Mathematics and Computation, 218(15): 7856–7863, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001063.

Fernandez:2012:ACR

[Fer12b] Francisco M. Fernández. Accurate calculation of resonances for a centralfield model potential. Applied Mathematics and Computation, 218(10): 5961–5965, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014068.

Fernandez:2013:BSO

[Fer13] Julio C. Ramos Fernández. Bounded superposition operators between weighted Banach spaces of analytic functions. Applied Mathematics and Computation, 219(10):4942–4949, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012210.

Fernandez:2014:HPM

[Fer14] Francisco M. Fernández.

On the homotopy perturbation method for Boussinesq-like equations. Applied Mathematics and Computation, 230(??): 208–210, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313014446.

Fedele:2011:ESF

[FF11] Giuseppe Fedele and Andrea Ferrise. Explicit solution of the finite time L_{2}- norm polynomial approximation problem. Applied Mathematics and Computation, 217(21):8354–8359, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003985.

Feng:2013:EST

[FFB13] Xingfang Feng, Hanying Feng, and Donglong Bai. Eigenvalue for a singular third-order threepoint boundary value problem. Applied Mathematics and Computation, 219(18):9783–9790, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003688.

Fernandez-Feria:2014:PBM

[FFOC14] R. Fernandez-Feria and J. Ortega-Casanova. A pseudospectral based method of lines for solving integrodifferential boundary-layer equations. Application to the mixed convection over a heated horizontal plate. Applied Mathematics and Computation, 242(??):388–396, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007681.

Franco:2011:FEP

[FG11] J. M. Franco and I. Gómez. A family of explicit parallel Runge–Kutta–Nyström methods. Applied Mathematics and Computation, 218(8):4177–4191, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311012215.

Fernandez:2013:LAC

[FG13a] Francisco M. Fernández and Javier Garcia. Local approximation to the critical parameters of quantum wells. Applied Mathematics and Computation, 220(??):580–592, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006887.

Franco:2013:CES

[FG13b] J. M. Franco and I. Gómez.

Construction of explicit symmetric and symplectic methods of Runge– Kutta–Nyström type for solving perturbed oscillators. Applied Mathematics and Computation, 219(9):4637–4649, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010910.

Fernandez:2014:CPN

[FG14a] Francisco M. Fernández and Javier Garcia. Critical parameters for non-Hermitian Hamiltonians. Applied Mathematics and Computation, 247(??): 141–151, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031401193X.

Franco:2014:TFN

[FG14b] J. M. Franco and I. Gómez. Trigonometrically fitted nonlinear two-step methods for solving second order oscillatory IVPs. Applied Mathematics and Computation, 232(??): 643–657, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001155.

Furtula:2013:SSD

[FGD13] Boris Furtula, Ivan Gutman, and Matthias Dehmer. On structure-sensitivity of degree-based topological indices. Applied Mathematics and Computation, 219(17):8973–8978, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003330.

Flores-Gutierrez:2011:HLI

[FGFRV11] C. P. Flores-Gutiérrez, R. Femat, and E. Ruiz-Velázquez. On hypoglycemic levels induced by H_{∞} control on type I diabetes mellitus. Applied Mathematics and Computation, 218(2):376–385, September 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311007557.

Furtula:2012:CST

[FGIV12] Boris Furtula, Ivan Gutman, Milos Ivanović, and Damir Vukicević. Computer search for trees with minimal ABC index. Applied Mathematics and Computation, 219(2):767–772, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006996.

French:2012:EAM

[FGO12] Donald A. French, Marshall C. Galbraith, and Mauricio Osorio. Error analysis of a modified discontinuous Galerkin recovery scheme for diffusion problems. Applied Mathematics and Computation, 218(13):7144–7154, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015724.

Ferreiro:2014:SLM

[FGRLSV14] Ana M. Ferreiro, José A. García-Rodríguez, José G. López-Salas, and Carlos Vázquez. SABR/LIBOR market models: Pricing and calibration for some interest rate derivatives. Applied Mathematics and Computation, 242(??):65–89, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006730.

Ferrara:2014:NDS

[FGS14] Massimiliano Ferrara, Luca Guerrini, and Mauro Sodini. Nonlinear dynamics in a solow model with delay and non-convex technology. Applied Mathematics and Computation, 228(??):1–12, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012514.

Gao:2011:IAB

[fGyLJ11] Wei feng Gao, San yang Liu, and Fei Jiang. An improved artificial bee colony algorithm for directing orbits of chaotic systems. Applied Mathematics and Computation, 218(7):3868–3879, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012069.

Fang:2011:DPA

[FH11a] Changjie Fang and Yiran He. A double projection algorithm for multi-valued variational inequalities and a unified framework of the method. Applied Mathematics and Computation, 217(23):9543–9551, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005509.

Felcman:2011:NFS

[FH11b] J. Felcman and O. Havle.

On a numerical flux for the shallow water equations. Applied Mathematics and Computation, 217(11):5160–5170, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007927.

Fu:2013:ESN

[FH13] Xianlong Fu and Rong Huang. Existence of solutions for neutral integrodifferential equations with state-dependent delay. Applied Mathematics and Computation, 224(??): 743–759, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009818.

Farooq:2014:HMT

[FHAL14] U. Farooq, T. Hayat, A. Alsaedi, and Shijun Liao. Heat and mass transfer of twolayer flows of third-grade nano-fluids in a vertical channel. Applied Mathematics and Computation, 242(??):528–540, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008261.

He:2011:IAZ

[fHcXH11] Xin feng He, Yong chun Xu, and Zhen He. Iterative approximation for a zero of accretive operator and fixed points problems in Banach space. Applied Mathematics and Computation, 217(9):4620–4626, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011264.

Ferchichi:2013:DPF

[FHK13] J. Ferchichi, M. Hassine, and H. Khenous. Detection of point-forces location using topological algorithm in Stokes flows. Applied Mathematics and Computation, 219(12):7056–7074, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201288X.

Fan:2011:NSS

[FHLW11] Wenxi Fan, Xinan Hao, Lishan Liu, and Yonghong Wu. Nontrivial solutions of singular fourth-order Sturm–Liouville boundary value problems with a sign-changing nonlinear term. Applied Mathematics and Computation, 217(15):6700–6708, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001068.

Farina:2013:TEP

[FHR13] D. Farina, P. Hammes, and S. Reitzinger. Thermal expansion as a parameter for sensitivity analysis of waveguide filters. Applied Mathematics and Computation, 219(13):7181–7186, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311009350.

Fan:2013:GDN

[FHW13] Aijun Fan, Peng Han, and Kaifa Wang. Global dynamics of a nutrientplankton system in the water ecosystem. Applied Mathematics and Computation, 219(15):8269–8276, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001835.

Feng:2014:IMW

[FHW14] Hui Feng, Cong Huang, and Rong Wang. An improved mapped weighted essentially non-oscillatory scheme. Applied Mathematics and Computation, 232(??):453–468, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000988.

Feng:2012:NID

[FHWZ12] Wenqiang Feng, Xiaoming He, Zhu Wang, and Xu Zhang. Noniterative domain decomposition methods for a nonstationary Stokes–Darcy model with Beavers– Joseph interface condition. Applied Mathematics and Computation, 219(2):453–463, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004973.

Figueroa:2011:SOF

[Fig11] Rubén Figueroa. Secondorder functional differential equations with past, present and future dependence. Applied Mathematics and Computation, 217(18):7448–7454, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002311.

Figueroa:2012:DFD

[Fig12] Rubén Figueroa. Discontinuous functional differential equations with delayed or advanced arguments. Applied Mathematics and Computation, 218(19):9882–9889, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312003256.

Furati:2014:IPG

[FIK14] Khaled M. Furati, Olaniyi S. Iyiola, and Mokhtar Kirane. An inverse problem for a generalized fractional diffusion. Applied Mathematics and Computation, 249(??):24–31, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031401412X.

Finta:2014:GVT

[Fin14] Zoltán Finta. Generalized Voronovskaja theorem for q-Bernstein polynomials. Applied Mathematics and Computation, 246(??):619–627, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011576.

Fitt:2012:GES

[Fit12] A. D. Fitt. Generalized exact solutions for boundary layer flow and heat transfer over permeable stretching sheets. Applied Mathematics and Computation, 219(4):1468–1473, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007655.

Felcman:2012:AFV

[FK12] J. Felcman and L. Kadrnka.

Adaptive finite volume approximation of the shallow water equations. Applied Mathematics and Computation, 219(7):3354–3366, December 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031100600X.

Fucci:2013:ACD

[FK13] Guglielmo Fucci and Klaus Kirsten. Analytic continuation of the doubly-periodic Barnes zeta function. Applied Mathematics and Computation, 221(??):598–609, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007418.

Fomin:2010:RFP

[FKH10a] Sergei Fomin, Kris Kilpatrick, and Richard Hubbard. Rimming flow of a power-law fluid: Qualitative analysis of the mathematical model and analytical solutions. Applied Mathematics and Computation, 216(7):2169–2176, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002948.

Frankel:2010:LSA

[FKH10b] J. I. Frankel, M. Keyhani, and M. G. Huang. Local sensitivity analysis for the heat flux-temperature integral relationship in the half-space. Applied Mathematics and Computation, 217(1):363–375, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006296.

Ferrara:2014:MRP

[FKH14] Massimiliano Ferrara, Somaye Khademloo, and Shapour Heidarkhani. Multiplicity results for perturbed fourth-order Kirchhoff type elliptic problems. Applied Mathematics and Computation, 234(??):316–325, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400280X.

Farago:2013:CDM

[FKS13] István Faragó, Sergey Korotov, and Tamás Szabó. On continuous and discrete maximum principles for elliptic problems with the third boundary condition. Applied Mathematics and Computation, 219(13):7215–7224, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311007120.

Fu:2010:PTF

[FL10] Yanggeng Fu and Zhengrong Liu. Persistence of travelling fronts of KdV–Burgers–Kuramoto equation. Applied Mathematics and Computation, 216(7):2199–2206, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003115.

Francoise:2011:AST

[FL11] Jean-Pierre Françoise and Jaume Llibre. Analytical study of a triple Hopf bifurcation in a tritrophic food chain model. Applied Mathematics and Computation, 217(17):7146–7154, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001512.

Fan:2013:SIS

[FL13] Chia-Ming Fan and Hong-Huei Li. Solving the inverse Stokes problems by the modified collocation Trefftz method and Laplacian decomposition. Applied Mathematics and Computation, 219(12):6520–6535, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000222.

Feng:2014:IRR

[FL14] Zhiguang Feng and James Lam. An improved result on reachable set estimation and synthesis of timedelay systems. Applied Mathematics and Computation, 249(??):89–97, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013691.

Favati:2014:GCV

[FLMR14] P. Favati, G. Lotti, O. Menchi, and F. Romani. Generalized crossvalidation applied to conjugate gradient for discrete ill-posed problems. Applied Mathematics and Computation, 243(??):258–268, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008091.

Fernandez:2011:RST

[FLP11] Claudio Fernández, Carlos Lizama, and Verónica Poblete. Regularity of solutions for a third order differential equation in Hilbert spaces. Applied Mathematics and Computation, 217(21):8522–8533, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004243.

Feng:2011:UEM

[FLW11] Jishe Feng, Wanjun Li, and Qiaoling Wan. Using \frac{G′}{G} -expansion method to seek the traveling wave solution of Kolmogorov– Petrovskii–Piskunov equation. Applied Mathematics and Computation, 217(12):5860–5865, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012737.

Feng:2014:IEE

[FLZ14] Hui Feng, Yang Liu, and Xiaoping Zhang. L_{2} error estimates of collocation methods for solving certain singular integral equations. Applied Mathematics and Computation, 229(??):396–413, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013301.

Florios:2014:GEP

[FM14] Kostas Florios and George Mavrotas. Generation of the exact Pareto set in Multi-Objective Traveling Salesman and Set Covering Problems. Applied Mathematics and Computation, 237(??):1–19, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004810.

Feng:2013:SNG

[FMF13] Qinghua Feng, Fanwei Meng, and Bosheng Fu. Some new generalized Volterra–Fredholm type finite difference inequalities involving four iterated sums. Applied Mathematics and Computation, 219(15):8247–8258, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001331.

Fatehi:2011:MBF

[FMS11] Ali Fatehi, Jonathan Murley, and Nasser Saad. On a modified beta function and some applications. Applied Mathematics and Computation, 218(3):1012–1018, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005108.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Fatori:2014:TSH

[FMS14] Luci Harue Fatori, Rodrigo Nunes Monteiro, and Hugo D. Fernández Sare. The Timoshenko system with history and Cattaneo law. Applied Mathematics and Computation, 228(??):128–140, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031301223X.

Fomin:2013:RRF

[FMSC13] Sergei Fomin, Anton Mazurenko, Ravi Shankar, and Vladimir Chugunov. Reactive rimming flow of non-Newtonian fluids. Applied Mathematics and Computation, 219(14):7671–7684, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000969.

Fernandez-Martinez:2012:FDF

[FMSGS12] M. Fernández-Martínez, M. A. Sánchez-Granero, and J. E. Trinidad Segovia. Fractal dimension for fractal structures: Applications to the domain of words. Applied Mathematics and Computation, 219(3):1193–1199, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007291.

Fiordilino:2014:HTF

[FMV14] J. Fiordilino, M. Massoudi, and A. Vaidya. On the heat transfer and flow of a non-homogeneous fluid. Applied Mathematics and Computation, 243(??):184–196, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008170.

Feng:2011:ESO

[FNG11] Xiaojing Feng, Pengcheng Niu, and Qianqiao Guo. Existence of solutions for 2m th-order periodic boundary value problems. Applied Mathematics and Computation, 218(5):2343–2352, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100885X.

Feichtinger:2010:ALS

[FNV10] G. Feichtinger, A. J. Novak, and V. M. Veliov. Applying the Leitmann–Stalford sufficient conditions to maximization control problems with non-concave Hamiltonian. Applied Mathematics and Computation, 217(3):1017–1022, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003243.

Filobello-Nino:2013:GBM

[FNVLPD^{+}13] U. Filobello-Nino, H. Vazquez-Leal, D. Pereyra-Diaz, A. Yildirim, A. Perez-Sesma, R. Castaneda-Sheissa, J. Sanchez-Orea, and C. Hoyos-Reyes.

A generalization of the Bernoulli’s method applied to brachistochronelike problems. Applied Mathematics and Computation, 219(12):6707–6718, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313000520.

Faria:2011:GCA

[FO11] Teresa Faria and José J.

Oliveira. General criteria for asymptotic and exponential stabilities of neural network models with unbounded delays. Applied Mathematics and Computation, 217(23):9646–9658, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006199.

Fontova:2012:HNN

[FOL12] Marta I. Velazco Fontova, Aurelio R. L. Oliveira, and Christiano Lyra. Hopfield neural networks in largescale linear optimization problems. Applied Mathematics and Computation, 218(12):6851–6859, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015463.

Formato:2011:CFO

[For11] Richard A. Formato. Central Force Optimization with variable initial probes and adaptive decision space. Applied Mathematics and Computation, 217(21):8866–8872, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005315.

Fan:2014:OMT

[FOZX14] Chaodong Fan, Honglin Ouyang, Yingjie Zhang, and Leyi Xiao. Optimal multilevel thresholding using molecular kinetic theory optimization algorithm. Applied Mathematics and Computation, 239(??):391–408, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006432.

Fotiades:2012:EUA

[FP12] N. Fotiades and G. Papaschinopoulos. Existence, uniqueness and attractivity of prime period two solution for a difference equation of exponential form. Applied Mathematics and Computation, 218(23):11648–11653, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005620.

Fang:2013:MIA

[FP13a] Mei Fang and Ju H.

Park. A multiple integral approach to stability of neutral time-delay systems. Applied Mathematics and Computation, 224(??):714–718, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009466.

Fang:2013:NFS

[FP13b] Mei Fang and Ju H.

Park. Non-fragile synchronization of neural networks with time-varying delay and randomly occurring controller gain fluctuation. Applied Mathematics and Computation, 219(15):8009–8017, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001525.

Forys:2013:AHB

[FP13c] Urszula Foryś and Monika J. Piotrowska. Analysis of the Hopf bifurcation for the family of angiogenesis models II: the case of two nonzero unequal delays. Applied Mathematics and Computation, 220(??):277–295, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006577.

Fotiades:2013:SDE

[FP13d] N. Fotiades and G. Papaschinopoulos. On a system of difference equations with maximum. Applied Mathematics and Computation, 221(??):684–690, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007649.

Fang:2014:SAL

[FPC14] Qingxiang Fang, Jigen Peng, and Feilong Cao. Synchronization analysis of linearly coupled singular systems. Applied Mathematics and Computation, 248(??):536–549, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013733.

Ferenczik:2011:EVS

[FPP11a] J. Ferenczik, Á. Pintér, and B. Porvázsnyik. On equal values of Stirling numbers of the second kind. Applied Mathematics and Computation, 218(3):980–984, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001238.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Franjic:2011:FQF

[FPP11b] Iva Franjić, Josip Pečarić, and Ivan Perić. On families of quadrature formulas based on Euler identities. Applied Mathematics and Computation, 217(9):4516–4528, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011148.

Franjic:2014:OTI

[FPTS14] Iva Franjić, Josip Pecarić, and Sanja Tipurić-Spuzević. Ostrowski type inequalities for functions whose higher order derivatives have a single point of nondifferentiability. Applied Mathematics and Computation, 245(??):557–565, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011138.

Fu:2011:CMF

[FR11] Hongfei Fu and Hongxing Rui. A characteristicmixed finite element method for time-dependent convectiondiffusion optimal control problem. Applied Mathematics and Computation, 218(7):3430–3440, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011283.

Frammartino:2010:NTO

[Fra10] Carmelina Frammartino. Numerical treatment of sth order boundary value problems by solving second kind Fredholm integral equations. Applied Mathematics and Computation, 215(12):4141–4153, February 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309010649.

Ren:2013:ECT

[fRbLJyS13] Rong fen Ren, Hou biao Li, Wei Jiang, and Ming yan Song. An efficient Chebyshev-tau method for solving the space fractional diffusion equations. Applied Mathematics and Computation, 224(??):259–267, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009417.

Freire:2011:SAS

[Fre11] Igor Leite Freire. Selfadjoint sub-classes of third and fourth-order evolution equations. Applied Mathematics and Computation, 217(22):9467–9473, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311005996.

Fotakis:2012:NMO

[FS12] Dimitris Fotakis and Epameinondas Sidiropoulos. A new multi-objective self-organizing optimization algorithm (MOSOA) for spatial optimization problems. Applied Mathematics and Computation, 218(9):5168–5180, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013300.

Fang:2014:SCV

[FS14] Tao Fang and Jitao Sun. Stability of complexvalued impulsive system with delay. Applied Mathematics and Computation, 240(??):102–108, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400602X.

Frederico:2010:FNT

[FT10] Gastão S. F. Frederico and Delfim F. M. Torres. Fractional Noether’s theorem in the Riesz–Caputo sense. Applied Mathematics and Computation, 217(3):1023–1033, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001244.

Frontini:2011:HMP

[FT11] M. Frontini and A. Tagliani.

Hausdorff moment problem and Maximum Entropy: On the existence conditions. Applied Mathematics and Computation, 218(2):430–433, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007636.

Fabregat-Traver:2014:SSG

[FTAB14] Diego Fabregat-Traver, Yurii S. Aulchenko, and Paolo Bientinesi. Solving sequences of generalized least-squares problems on multi-threaded architectures. Applied Mathematics and Computation, 234(??):606–617, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002951.

Fukuda:2012:SMA

[Fuk12] Kosei Fukuda. A simple method for age-period-cohort decomposition of firm survival data. Applied Mathematics and Computation, 219(2):741–747, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006753.

Fukushima:2013:RCD

[Fuk13] Toshio Fukushima. Recursive computation of derivatives of elliptic functions and of incomplete elliptic integrals. Applied Mathematics and Computation, 221(??):21–31, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006152.

Fukushima:2014:ACG

[Fuk14a] Toshio Fukushima. Analytical computation of generalized Fermi–Dirac integrals by truncated Sommerfeld expansions. Applied Mathematics and Computation, 234(??): 417–433, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314002926.

Fukushima:2014:CGI

[Fuk14b] Toshio Fukushima. Computation of a general integral of Fermi–Dirac distribution by McDougall– Stoner method. Applied Mathematics and Computation, 238(??):485–510, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400561X.

Fuyong:2011:ICM

[Fuy11] Lin Fuyong. The inverse of circulant matrix. Applied Mathematics and Computation, 217(21):8495–8503, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004206.

fuyong:2013:SLS

[fuy13] Lin fuyong. The solution of linear systems equations with circulant-like coefficient matrices. Applied Mathematics and Computation, 219(15):8259–8268, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001422.

Fuyong:2014:DMS

[Fuy14a] Lin Fuyong. A direct method for solving block circulant banded system of linear equations. Applied Mathematics and Computation, 232(??):1269–1276, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001362.

Fuyong:2014:NKS

[Fuy14b] Lin Fuyong. A numerical kernel solution of beam systems. Applied Mathematics and Computation, 243(??):594–600, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008662.

Farkas:2014:MSI

[FV14] Csaba Farkas and Csaba Varga. Multiple symmetric invariant non trivial solutions for a class of quasilinear elliptic variational systems. Applied Mathematics and Computation, 241(??):347–355, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006699.

Fan:2010:VIM

[FW10] Aijun Fan and Kaifa Wang. A viral infection model with immune circadian rhythms. Applied Mathematics and Computation, 215(9):3369–3374, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009333.

Feliu:2012:PSB

[FW12] Elisenda Feliu and Carsten Wiuf. Preclusion of switch behavior in networks with mass-action kinetics. Applied Mathematics and Computation, 219(4):1449–1467, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007643.

Fan:2014:MPS

[FW14] Yong-Hong Fan and Lin-Lin Wang. Multiplicity of periodic solutions for a delayed ratio-dependent predator-prey model with monotonic functional response and harvesting terms. Applied Mathematics and Computation, 244(??):878–894, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010030.

Fu:2010:RAC

[FWZW10] Hongguang Fu, Ying Wang, Shizhong Zhao, and Qingxian Wang. A recursive algorithm for constructing complicated Dixon matrices. Applied Mathematics and Computation, 217 (6):2595–2601, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008313.

Fan:2013:SCD

[FWZW13] Yong-Qing Fan, Yin-He Wang, Yun Zhang, and Qin-Ruo Wang. The synchronization of complex dynamical networks with similar nodes and coupling time-delay. Applied Mathematics and Computation, 219(12):6719–6728, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000489.

Fan:2012:CDD

[FXD12] Yueqian Fan, Tao Xie, and Jianguo Du. Complex dynamics of duopoly game with heterogeneous players: a further analysis of the output model. Applied Mathematics and Computation, 218(15):7829–7838, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001014.

Fan:2014:ASN

[FXL14] Mingshu Fan, Anyin Xia, and Shan Li. Asymptotic stability for a nonlocal parabolic problem. Applied Mathematics and Computation, 243(??):740–745, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006249.

Faydaoglu:2012:AFI

[FY12] S. Faydaoglu and V. G. Yakhno. Applications of finite and infinite Fourier series in heat processes with impulse data. Applied Mathematics and Computation, 218(16):8120–8130, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011349.

Fan:2013:EMB

[FYL13] Kun Fan, Weijia You, and Yuanyuan Li. An effective modified binary particle swarm optimization (mbpso) algorithm for multi-objective resource allocation problem (morap). Applied Mathematics and Computation, 221(??):257–267, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006784.

Fang:2013:NPF

[FYM13] Yonglei Fang, Xiong You, and Qinghe Ming. A new phase-fitted modified Runge–Kutta pair for the numerical solution of the radial Schrödinger equation. Applied Mathematics and Computation, 224(??):432–441, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009491.

Fan:2011:PLS

[FYYT11] Xinghua Fan, Shouxiang Yang, Jiuli Yin, and Lixin Tian. Peakon, loop and solitary travelling wave solutions for the general modified CH–DP equation. Applied Mathematics and Computation, 217(9):4812–4818, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011483.

Feng:2012:GGB

[FZ12] Qinghua Feng and Bin Zheng. Generalized Gronwall–Bellman-type delay dynamic inequalities on time scales and their applications. Applied Mathematics and Computation, 218(15):7880–7892, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312001191.

Fan:2013:LMA

[FZ13] Jinyan Fan and Jinlong Zeng. A Levenberg– Marquardt algorithm with correction for singular system of nonlinear equations. Applied Mathematics and Computation, 219(17):9438–9446, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002798.

Fan:2013:MSM

[FZLZ13] Zhi-Ping Fan, Xiao Zhang, Yang Liu, and Yao Zhang. A method for stochastic multiple attribute decision making based on concepts of ideal and anti-ideal points. Applied Mathematics and Computation, 219(24):11438–11450, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313005869.

Feng:2012:LBM

[FZP12] Heying Feng, Xiaoqing Zhang, and Yehui Peng. A lattice Boltzmann model for elliptic equations with variable coefficient. Applied Mathematics and Computation, 219(5):2798– 2807, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009095.

Zhang:2011:PDS

[fZwXyZ11] Guo feng Zhang, Wei wei Xie, and Jing yu Zhao. Positive definite solutions of the nonlinear matrix equation X + A^{∗}X^{q}A = Q(q > 0). Applied Mathematics and Computation, 217(22):9182–9188, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005340.

Fang:2010:NFU

[FZY10] Tiegang Fang, Ji Zhang, and Shanshan Yao. A new family of unsteady boundary layers over a stretching surface. Applied Mathematics and Computation, 217(8):3747–3755, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009781.

Fang:2012:BLF

[FZZ12] Tiegang Fang, Ji Zhang, and Yongfang Zhong. Boundary layer flow over a stretching sheet with variable thickness. Applied Mathematics and Computation, 218(13):7241–7252, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000203.

Gupta:2010:CAS

[GA10] Vijay Gupta and Ali Aral. Convergence of the q analogue of Szász– Beta operators. Applied Mathematics and Computation, 216(2):374–380, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000317.

Glover:2011:URB

[GA11a] J. Glover and M. M. Ali. Using radial basis functions to construct local volatility surfaces. Applied Mathematics and Computation, 217(9):4834–4839, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300310011069.

Gupta:2011:SAP

[GA11b] Vijay Gupta and Ali Aral. Some approximation properties of q-Baskakov–Durrmeyer operators. Applied Mathematics and Computation, 218(3):783–788, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000841.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Gurses:2012:MMV

[GAC12] Murat Gürses, Bekir Akgöz, and Ömer Civalek. Mathematical modeling of vibration problem of nano-sized annular sector plates using the nonlocal continuum theory via eight-node discrete singular convolution transformation. Applied Mathematics and Computation, 219(6):3226–3240, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009617.

Gad:2011:EHC

[Gad11] N. S. Gad. Effect of Hall currents on interaction of pulsatile and peristaltic transport induced flows of a particle– fluid suspension. Applied Mathematics and Computation, 217(9):4313–4320, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008702.

Gad:2014:EHC

[Gad14] N. S. Gad. Effects of Hall currents on peristaltic transport with compliant walls. Applied Mathematics and Computation, 235(??):546–554, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003294.

Guebbai:2014:ANS

[GADK14] H. Guebbai, M. Z. Aissaoui, I. Debbar, and B. Khalla. Analytical and numerical study for an integro-differential nonlinear Volterra equation. Applied Mathematics and Computation, 229(??):367–373, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013295.

Golden:2012:MCS

[GAK12] Boris Golden, Marc Aiguier, and Daniel Krob. Modeling of complex systems II: a minimalist and unified semantics for heterogeneous integrated systems. Applied Mathematics and Computation, 218(16):8039–8055, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000860.

Gal:2010:ACG

[Gal10a]  Sorin G. Gal.  Ap-

proximation by complex genuine Durrmeyer type polynomials in compact disks. Applied Mathematics and Computation, 217(5):1913–1920, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007319.

Galewski:2010:DPN

[Gal10b]  Marek Galewski.  On

the Dirichlet problem for a nonlinear elastic beam equation. Applied Mathematics and Computation, 217(8):4295–4301, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001043X.

Galperin:2010:GLI

[Gal10c] E. A. Galperin. G. Leitmann’s integral transformation for global optimal control. Applied Mathematics and Computation, 217(3):1045–1052, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004364.

Gala:2011:RBC

[Gal11a] Sadek Gala. A remark on the blow-up criterion of strong solutions to the Navier–Stokes equations. Applied Mathematics and Computation, 217(22):9488–9491, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005364.

Galewski:2011:DPD

[Gal11b] Marek Galewski. Dependence on parameters for a discrete Emden–Fowler equation. Applied Mathematics and Computation, 218(4):1247–1253, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008186.

Galewski:2011:NMS

[Gal11c] Marek Galewski. A note on the multiplicity of solutions for a second-order difference equation with a parameter. Applied Mathematics and Computation, 218(7):3954–3956, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011556.

Galiano:2011:MSA

[Gal11d] Gonzalo Galiano. Modeling spatial adaptation of populations by a time non-local convection crossdiffusion evolution problem. Applied Mathematics and Computation, 218(8):4587–4594, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012860.

Gallego:2011:PCP

[Gal11e] Rafael Gallego. Predictorcorrector pseudospectral methods for stochastic partial differential equations with additive white noise. Applied Mathematics and Computation, 218(7):3905–3917, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012100.

Gao:2010:EMS

[Gao10] Liming Gao. Existence of multiple solutions for a second-order difference equation with a parameter. Applied Mathematics and Computation, 216(5):1592–1598, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002547.

Gao:2014:LND

[Gao14]  Chenghua Gao.  On

the linear and nonlinear discrete secondorder Neumann boundary value problems. Applied Mathematics and Computation, 233(??):62–71, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400157X.

Garcia:2011:FPA

[Gar11] Juan Manuel Garcia Garcia. A fixed-point algorithm to estimate the Yule-Simon distribution parameter. Applied Mathematics and Computation, 217(21):8560–8566, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004607.

Goyal:2012:LTS

[GARK12] Amit Goyal, Alka, Thokala Soloman Raju, and C. Nagaraja Kumar. Lorentziantype soliton solutions of ac-driven complex Ginzburg–Landau equation. Applied Mathematics and Computation, 218(24):11931–11937, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312005796.

Gatto:2014:ISA

[Gat14] Riccardo Gatto. Importance sampling approximations to various probabilities of ruin of spectrally negative Lévy risk processes. Applied Mathematics and Computation, 243(??):91–104, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007644.

Gopal:2014:SNF

[GAV14] Dhananjay Gopal, Mujahid Abbas, and Calogero Vetro. Some new fixed point theorems in Menger PM-spaces with application to Volterra type integral equation. Applied Mathematics and Computation, 232(??):955–967, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001805.

Gazi:2012:DMP

[Gaz12] Nurul Huda Gazi. Dynamics of a marine plankton system: Diffusive instability and pattern formation. Applied Mathematics and Computation, 218(17):8895–8905, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001750.

Girgis:2010:SPT

[GB10] Laila Girgis and Anjan Biswas. Soliton perturbation theory for nonlinear wave equations. Applied Mathematics and Computation, 216(7):2226–2231, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002985.

Gandarias:2012:CLC

[GB12] M. L. Gandarias and M. S.

Bruzon. Conservation laws for a class of quasi selfadjoint third order equations. Applied Mathematics and Computation, 219(2):668–678, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006674.

Gupta:2013:PAL

[GB13] R. K. Gupta and Anupma Bansal. Painlevé analysis, Lie symmetries and invariant solutions of potential Kadomstev–Petviashvili equation with time dependent coefficients. Applied Mathematics and Computation, 219(10):5290–5302, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012118.

Ghorbani:2014:RCC

[GBA14] Modjtaba Ghorbani and Elaheh Bani-Asadi. Remarks on characteristic coefficients of fullerene graphs. Applied Mathematics and Computation, 230(??):428–435, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031301357X.

Gorska:2012:RMT

[GBD^{+}12] K. Górska, D. Babusci, G. Dattoli, G. H. E. Duchamp, and K. A. Penson. The Ramanujan master theorem and its implications for special functions. Applied Mathematics and Computation, 218(23):11466–11471, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005218.

Ghosh:2010:PED

[GC10a] Dibakar Ghosh and A. Roy Chowdhury. Parameter estimation of delay dynamical system from a scalar time series under external noise. Applied Mathematics and Computation, 216(7):2069–2076, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002833.

Glielmo:2010:OFC

[GC10b] Luigi Glielmo and Martin Corless. On output feedback control of singularly perturbed systems. Applied Mathematics and Computation, 217(3):1053–1070, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002031.

Guo:2010:STL

[GC10c] Hongjian Guo and Lansun Chen. A study on time-limited control of single-pest with stagestructure. Applied Mathematics and Computation, 217(2):677–684, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006788.

Geng:2011:NMN

[GC11a] Fazhan Geng and Minggen Cui. A novel method for nonlinear two-point boundary value problems: Combination of ADM and RKM. Applied Mathematics and Computation, 217(9):4676–4681, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031001132X.

Gungor:2011:SBE

[GÇ11b] A. Dilek Maden (Güngör) and A. Sinan Çevik. Some bounds for extreme singular values of a complex matrix. Applied Mathematics and Computation, 218(3):762–765, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004413.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Guo:2011:EGA

[GC11c] Haijun Guo and Xiaoxing Chen. Existence and global attractivity of positive periodic solution for a Volterra model with mutual interference and Beddington– DeAngelis functional response. Applied Mathematics and Computation, 217(12):5830–5837, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012671.

Gurler:2011:AMR

[GC11d] Selma Gurler and Sedat Capar. An algorithm for mean residual life computation of (n−k+1)-out-of-n systems: an application of exponentiated Weibull distribution. Applied Mathematics and Computation, 217(19):7806–7811, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003171.

Gakkhar:2012:DMH

[GC12] Sunita Gakkhar and Nareshkumar Chavda. A dynamical model for HIV-TB co-infection. Applied Mathematics and Computation, 218(18):9261–9270, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002366.

Gao:2014:NSN

[GC14] Feng Gao and Chunmei Chi. Numerical solution of nonlinear Burgers’ equation using high accuracy multi-quadric quasiinterpolation. Applied Mathematics and Computation, 229(??):414–421, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013118.

Gomez-Corral:2014:MQL

[GCG14] A. Gómez-Corral and M. López García. Maximum queue lengths during a fixed time interval in the M/M/c retrial queue. Applied Mathematics and Computation, 235(??):124–136, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003221.

Gong:2010:RCB

[GCLL10] Wenyin Gong, Zhihua Cai, Charles X. Ling, and Hui Li. A realcoded biogeography-based optimization with mutation. Applied Mathematics and Computation, 216(9):2749–2758, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003887.

Garrancho:2010:CAF

[GCM10] P. Garrancho and D. Cárdenas-Morales. A converse of asymptotic formulae in simultaneous approximation. Applied Mathematics and Computation, 217(6):2676–2683, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008556.

Garcia-Caballero:2014:CVV

[GCMP14] Esther M. García-Caballero, Samuel G. Moreno, and Michael P. Prophet. A complete view of Viète-like infinite products with Fibonacci and Lucas numbers. Applied Mathematics and Computation, 247(??):703–711, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012363.

Gungor:2010:IUL

[GD10a] A. Dilek Güngör and Kinkar Ch. Das. Improved upper and lower bounds for the spectral radius of digraphs. Applied Mathematics and Computation, 216(3):791–799, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001049.

Guo:2010:RDI

[GD10b] Li Guo and Xiankun Du. Representations for the Drazin inverses of 2 × 2 block matrices. Applied Mathematics and Computation, 217 (6):2833–2842, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008738.

Guerry:2011:ETA

[GD11] Marie-Anne Guerry and Tim De Feyter. An extended and tractable approach on the convergence problem of the mixed push-pull manpower model. Applied Mathematics and Computation, 217(22):9062–9071, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004917.

Ghadimi:2010:SPE

[GDH10] Parviz Ghadimi, Abbas Dashtimanesh, and Hossein Hosseinzadeh. Solution of Poisson’s equation by analytical boundary element integration. Applied Mathematics and Computation, 217(1):152–163, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005783.

Ge:2014:CIR

[Ge14] Xiaohua Ge. Comments and an improved result on “Stability analysis for continuous system with additive time-varying delays: A less conservative result”. Applied Mathematics and Computation, 241(??):42–46, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006225.

See [DRGR10].

Geem:2012:EIM

[Gee12] Zong Woo Geem. Effects of initial memory and identical harmony in global optimization using harmony search algorithm. Applied Mathematics and Computation, 218(22):11337–11343, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004869.

Gemignani:2011:SCS

[Gem11]  L. Gemignani.  On

some classes of structured matrices with algebraic trigonometric eigenvalues. Applied Mathematics and Computation, 217(19):7573–7578, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002220.

Geng:2011:NMS

[Gen11] Fazhan Geng. A novel method for solving a class of singularly perturbed boundary value problems based on reproducing kernel method. Applied Mathematics and Computation, 218(8):4211–4215, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012240.

Gerolymos:2012:GRR

[Ger12] G. A. Gerolymos. A general recurrence relation for the weight-functions in Mühlbach–Neville–Aitken representations with application to WENO interpolation and differentiation. Applied Mathematics and Computation, 219(8):4133–4142, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312009435.

Garcia-Falset:2013:FPT

[GFMP13] J. Garcia-Falset and O. Muñiz-Pérez. Fixed point theory for 1-set contractive and pseudocontractive mappings. Applied Mathematics and Computation, 219(12):6843–6855, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000209.

Guo:2013:SPD

[GFZ13] Hui Guo, Hongfei Fu, and Jiansong Zhang. A splitting positive definite mixed finite element method for elliptic optimal control problem. Applied Mathematics and Computation, 219(24):11178–11190, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313005407.

Gadjiev:2010:APN

[GG10a] A. D. Gadjiev and A. M. Ghorbanalizadeh. Approximation properties of a new type Bernstein– Stancu polynomials of one and two variables. Applied Mathematics and Computation, 216(3):890–901, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001232.

Gutman:2010:DGG

[GG10b] Shaul Gutman and Orly Goldan. 3D differential game guidance. Applied Mathematics and Computation, 217(3):1077–1084, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004388.

Gal:2011:QEN

[GG11a] Sorin G. Gal and Vijay Gupta. Quantitative estimates for a new complex Durrmeyer operator in compact disks. Applied Mathematics and Computation, 218(6):2944–2951, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010848.

Guo:2011:MTC

[GG11b] Weiping Guo and Wei Guo. The matching theorems in G-convex spaces with applications. Applied Mathematics and Computation, 218(7):3746–3752, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011854.

Gutierrez-Gutierrez:2011:PCP

[GG11c] Jesús Gutiérrez-Gutiérrez. Powers of complex persymmetric or skew-persymmetric anti-tridiagonal matrices with constant antidiagonals. Applied Mathematics and Computation, 217(13):6125–6132, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310013081.

Ghiba:2012:FSM

[GG12a] Ionel-Dumitrel Ghiba and Catalin Gales. On the fundamental solutions for micropolar fluid-fluid mixtures under steady state vibrations. Applied Mathematics and Computation, 219(5):2749–2759, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009009.

Gordon:2012:PSH

[GG12b] Dan Gordon and Rachel Gordon. Parallel solution of high frequency Helmholtz equations using high order finite difference schemes. Applied Mathematics and Computation, 218(21):10737– 10754, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004511.

Gouravaraju:2012:EHB

[GG12c] Saipraneeth Gouravaraju and Ranjan Ganguli. Estimating the Hausdorff– Besicovitch dimension of boundary of basin of attraction in helicopter trim. Applied Mathematics and Computation, 218(21): 10435–10442, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003670.

Ganesh:2013:SSA

[GG13a] R. Ganesh and Ranjan Ganguli. Stiff string approximations in Rayleigh– Ritz method for rotating beams. Applied Mathematics and Computation, 219(17):9282–9295, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002701.

Gutierrez-Gutierrez:2013:BCP

[GG13b] Jesús Gutiérrez-Gutiérrez. Binomial coefficients and powers of large tridiagonal matrices with constant diagonals. Applied Mathematics and Computation, 219(17):9219–9222, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003512.

Gutierrez-Gutierrez:2013:SVD

[GG13c] Jesús Gutiérrez-Gutiérrez. Singular value decomposition for comb filter matrices. Applied Mathematics and Computation, 222(??):472–477, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S009630031300800X.

Gal:2014:ACP

[GG14a] Sorin G. Gal and Vijay Gupta. Approximation by complex Phillips–Stancu operators in compact disks under exponential growth conditions. Applied Mathematics and Computation, 234(??):309–315, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002884.

Gong:2014:ESH

[GG14b] Dong Gong and Xianguo Geng. Explicit solutions for a hierarchy of differential-difference equations. Applied Mathematics and Computation, 247(??):898–917, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012661.

Goubko:2014:DBT

[GG14c] Mikhail Goubko and Ivan Gutman. Degree-based topological indices: Optimal trees with given number of pendents. Applied Mathematics and Computation, 240(??):387–398, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006213.

Gu:2014:NNF

[GG14d] Xiaojuan Gu and Li Gao. A new nonlinear filter constructed from the Newton method and EPR in image restoration. Applied Mathematics and Computation, 234(??):520–530, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002434.

Guebbai:2014:NDK

[GG14e] Hamza Guebbai and Laurence Grammont. A new degenerate kernel method for a weakly singular integral equation. Applied Mathematics and Computation, 230(??):414–427, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013921.

Guessab:2014:SCA

[GG14f] Allal Guessab and Florian Guessab. Symmetrization, convexity and applications. Applied Mathematics and Computation, 240(??):149–160, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006031.

Gutierrez-Gutierrez:2014:EDP

[GG14g] Jesús Gutiérrez-Gutiérrez.

Eigenvalue decomposition for persymmetric Hankel matrices with at most three non-zero anti-diagonals. Applied Mathematics and Computation, 234(??):333–338, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002227.

Garcia-Gonzalo:2014:CSS

[GGFM14] Esperanza García-Gonzalo and Juan Luis Fernández-Martínez. Convergence and stochastic stability analysis of particle swarm optimization variants with generic parameter distributions. Applied Mathematics and Computation, 249(??):286–302, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314014416.

Ghag:2013:SAS

[GGM^{+}13] G. Ghag, P. Ghosh, A. Mauro, V. Rangachari, and A. Vaidya. Stability analysis of 4species A^{β} aggregation model: a novel approach to obtaining physically meaningful rate constants. Applied Mathematics and Computation, 224(??):205–215, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009211.

Garijo:2014:CTP

[GGM^{+}14a] D. Garijo, M. E. Gegúndez, A. Márquez, M. P. Revuelta, and F. Sagols. Computing the Tutte polynomial of Archimedean tilings. Applied Mathematics and Computation, 242(??):842–855, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009102.

Garijo:2014:DBM

[GGM14b] Delia Garijo, Antonio González, and Alberto Márquez. The difference between the metric dimension and the determining number of a graph. Applied Mathematics and Computation, 249(??):487–501, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014003.

Gulsu:2011:LPA

[GGÖS11] Mustafa Gülsu, Burcu Gürbüz, Yalçın Öztürk, and Mehmet Sezer. Laguerre polynomial approach for solving linear delay difference equations. Applied Mathematics and Computation, 217(15):6765–6776, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001548.

Gupta:2012:RNR

[GGOS12] Pushpa L. Gupta, Ramesh C.

Gupta, Seng-Huat Ong, and H. M. Srivastava. Reliability and non-reliability studies of Poisson variables in series and parallel systems. Applied Mathematics and Computation, 218(9):5112–5120, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013221.

Ganaie:2013:ACC

[GGP^{+}13] Ishfaq Ahmad Ganaie, Bharti Gupta, N. Parumasur, P. Singh, and V. K. Kukreja. Asymptotic convergence of cubic Hermite collocation method for parabolic partial differential equation. Applied Mathematics and Computation, 220(??):560–567, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300653X.

Garra:2014:HPD

[GGPT14] Roberto Garra, Rudolf Gorenflo, Federico Polito, and Zivorad Tomovski. Hilfer–Prabhakar derivatives and some applications. Applied Mathematics and Computation, 242(??):576–589, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008297.

Govil:2013:CNC

[GGS13] N. K. Govil, Vijay Gupta, and Danyal Soybas. Certain new classes of Durrmeyer type operators. Applied Mathematics and Computation, 225(??): 195–203, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313010023.

Guerra-Gomez:2013:REB

[GGTCdlF13] I. Guerra-Gómez, E. Tlelo-Cuautle, and Luis G. de la Fraga. Richardson extrapolation-based sensitivity analysis in the multi-objective optimization of analog circuits. Applied Mathematics and Computation, 222(??):167–176, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008096.

Guo:2012:AES

[GGXH12] Liangdong Guo, Hong Gu, Jun Xing, and Xiqin He.

Asymptotic and exponential stability of uncertain system with interval delay. Applied Mathematics and Computation, 218(19):9997–10006, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003323.

Gai:2011:PPA

[GGYW11] Xiao-Ling Gai, Yi-Tian Gao, Xin Yu, and Lei Wang. Painlevé property, auto-Bäcklund transformation and analytic solutions of a variablecoefficient modified Korteweg– de Vries model in a hot magnetized dusty plasma with charge fluctuations. Applied Mathematics and Computation, 218(2):271–279, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007193.

Gao:2010:AES

[GH10a] Zhen Gao and Jan S. Hesthaven. On ANOVA expansions and strategies for choosing the anchor point. Applied Mathematics and Computation, 217(7):3274–3285, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009252.

Geng:2010:DTE

[GH10b] Xianguo Geng and Guoliang He. Darboux transformation and explicit solutions for the Satsuma–Hirota coupled equation. Applied Mathematics and Computation, 216(9):2628–2634, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003619.

Guo:2010:ERM

[GH10c] Wenbin Guo and Tingzhu Huang. Extremal ranks of matrix expression of A − BXC with respect to Hermitian matrix. Applied Mathematics and Computation, 217(6):2381–2389, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007836.

Guo:2010:MET

[GH10d] Wenbin Guo and Tingzhu Huang. Method of elementary transformation to compute Moore–Penrose inverse. Applied Mathematics and Computation, 216(5):1614–1617, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002584.

Gyori:2010:ABS

[GH10e] I. Győri and L. Horváth. Asymptotic behaviour of the solutions of a nonautonomous linear delay difference system. Applied Mathematics and Computation, 217(8):4205–4216, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010787.

Gyori:2011:SSA

[GH11] I. Györi and L. Horváth. l^{p}-solutions and stability analysis of difference equations using the Kummer’s test. Applied Mathematics and Computation, 217(24):10129–10145, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006631.

Gao:2013:EMP

[GHC13] Shujing Gao, Yuying He, and Lansun Chen. An epidemic model with pulses for pest management. Applied Mathematics and Computation, 219(9):4308–4321, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009873.

Ge:2014:MSS

[GHG14] Chao Ge, Changchun Hua, and Xinping Guan. Master-slave synchronization criteria of Lur’e systems with time-delay feedback control. Applied Mathematics and Computation, 244(??):895–902, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010029.

Gao:2014:QIA

[GHLF14] Jiaquan Gao, Guixia He, Ronghua Liang, and Zhilin Feng. A quantum-inspired artificial immune system for the multiobjective 0–1 knapsack problem. Applied Mathematics and Computation, 230(??):120–137, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013787.

Ghosal:2014:GWR

[Gho14]  Sanjoy Ghosal.  Gen-

eralized weighted random convergence in probability. Applied Mathematics and Computation, 249(??):502–509, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014313.

Gheorghiu:2012:SCS

[GHPR12] C. I. Gheorghiu, M. E. Hochstenbach, B. Plestenjak, and J. Rommes. Spectral collocation solutions to multiparameter Mathieu’s system. Applied Mathematics and Computation, 218(24):11990– 12000, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300312005899.

Grasmair:2011:RMR

[GHS11] Markus Grasmair, Markus Haltmeier, and Otmar Scherzer. The residual method for regularizing ill-posed problems. Applied Mathematics and Computation, 218(6):2693–2710, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010356.

Gu:2014:CPI

[GHZ^{+}14] Xian-Ming Gu, Ting-Zhu Huang, Xi-Le Zhao, Wei-Ru Xu, Hou-Biao Li, and Liang Li. Circulant preconditioned iterative methods for peridynamic model simulation. Applied Mathematics and Computation, 248(??):470–479, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013472.

Gavrea:2013:SOP

[GI13] Ioan Gavrea and Mircea Ivan. A solution to an open problem on the Euler–Mascheroni constant. Applied Mathematics and Computation, 224(??):54–57, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009144.

Gavrea:2014:CCS

[GI14] Ioan Gavrea and Mircea Ivan. On a conjecture concerning the sum of the squared Bernstein polynomials. Applied Mathematics and Computation, 241(??):70–74, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006596.

Gine:2012:HOL

[Gin12a] Jaume Giné. Higher order limit cycle bifurcations from non-degenerate centers. Applied Mathematics and Computation, 218(17):8853–8860, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001713.

Gine:2012:LCB

[Gin12b] Jaume Giné. Limit cycle bifurcations from a nondegenerate center. Applied Mathematics and Computation, 218(9):4703–4709, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009720.

Gajic:2010:CMG

[GIR10] Ljiljana Gajić, Dejan Ilić, and Vladimir Rakočević. On Ćirić maps with a generalized contractive iterate at a point and Fisher’s quasi-contractions in cone metric spaces. Applied Mathematics and Computation, 216(8):2240–2247, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002523.

Gharibshah:2014:CUI

[GJ14] Joobin Gharibshah and Mahdi Jalili. Connectedness of users-items networks and recommender systems. Applied Mathematics and Computation, 243(??):578–584, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008601.

Gutierrez-Jaimez:2012:PSG

[GJGSNRÁ12] R. Gutiérrez-Jáimez, R. Gutiérrez-Sánchez, A. Nafidi, and E. Ramos-Ábalos. A τpower stochastic gamma diffusion process: Computational statistical inference and simulation aspects. A real example. Applied Mathematics and Computation, 219 (4):1576–1588, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007886.

Gao:2013:PIV

[GjLxF^{+}13] Fei Gao, Xue jing Lee, Feng xia Fei, Heng qing Tong, Yi bo Qi, Yan fang Deng, Ilangko Balasingham, and Hua ling Zhao. Parameter identification for van der Pol–Duffing oscillator by a novel artificial bee colony algorithm with differential evolution operators. Applied Mathematics and Computation, 222(??):132–144, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008035.

Guyader:2014:GAI

[GJNN14] Arnaud Guyader, Nicolas Jégou, Alexander B. Németh, and Sándor Z. Németh. A geometrical approach to iterative isotone regression. Applied Mathematics and Computation, 227(??):359–369, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012162.

Gao:2010:BDP

[GK10a] Ning Gao and Xu Kong.

Block diagonally preconditioned PIU methods of saddle point problem. Applied Mathematics and Computation, 216(6):1880–1887, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010224.

Geum:2010:MPF

[GK10b] Young Hee Geum and Young Ik Kim. A multiparameter family of threestep eighth-order iterative methods locating a simple root. Applied Mathematics and Computation, 215(9):3375–3382, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009357.

Geum:2010:PPF

[GK10c] Young Hee Geum and Young Ik Kim. A penta-parametric family of fifteenth-order multipoint methods for nonlinear equations. Applied Mathematics and Computation, 217(6):2311–2319, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007745.

Grover:2011:IEE

[GK11] Lovleen Kumar Grover and Parmdeep Kaur. An improved exponential estimator of finite population mean in simple random sampling using an auxiliary attribute. Applied Mathematics and Computation, 218(7):3093–3099, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010617.

Graef:2012:PSC

[GK12a] John R. Graef and Lingju Kong. Positive solutions for a class of higher order boundary value problems with fractional qderivatives. Applied Mathematics and Computation, 218(19):9682–9689, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200238X.

Gupta:2012:NAS

[GK12b] Bharti Gupta and V. K.

Kukreja. Numerical approach for solving diffusion problems using cubic B-spline collocation method. Applied Mathematics and Computation, 219(4):2087–2099, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008466.

Gupta:2012:MSO

[GK12c] S. K. Gupta and N. Kailey. Multiobjective secondorder mixed symmetric duality with a square root term. Applied Mathematics and Computation, 218(14):7602–7613, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200063X.

Gocwin:2013:CSN

[GK13] Maciej Goćwin and Bolesław Kacewicz. Complexity of solving nonlinear equations in the deterministic, randomized and quantum settings. Applied Mathematics and Computation, 224(??):652–662, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009788.

Ganaie:2014:NSB

[GK14] I. A. Ganaie and V. K. Kukreja. Numerical solution of Burgers’ equation by cubic Hermite collocation method. Applied Mathematics and Computation, 237(??):571–581, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004731.

Ghosh:2011:OPL

[GKC11] S. K. Ghosh, S. Khanra, and K. S. Chaudhuri. Optimal price and lot size determination for a perishable product under conditions of finite production, partial backordering and lost sale. Applied Mathematics and Computation, 217(13):6047–6053, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001252X.

Gupta:2010:LDF

[GKS10] Ramesh C. Gupta, S. N.

U. A. Kirmani, and H. M. Srivastava. Local dependence functions for some families of bivariate distributions and total positivity. Applied Mathematics and Computation, 216(4):1267–1279, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001876.

Gnecco:2012:AAS

[GKS12] Giorgio Gnecco, Vera Kurková, and Marcello Sanguineti. Accuracy of approximations of solutions to Fredholm equations by kernel methods. Applied Mathematics and Computation, 218(14):7481–7497, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000483.

Grover:2012:PTE

[GKV12] Lovleen Kumar Grover, Parmdeep Kaur, and Gajendra K. Vishawkarma. Product type exponential estimators of population mean under linear transformation of auxiliary variable in simple random sampling. Applied Mathematics and Computation, 219(4):1937–1946, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008296.

Kang:2012:CAN

[gKW12] Chuan gang Kang and Zhitao Wu. A convergence analysis of nonlinear implicit iterative method for nonlinear illposed problems. Applied Mathematics and Computation, 218(9):5265–5275, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013415.

Graef:2014:CSM

[GKW14] John R. Graef, Lingju Kong, and Min Wang. A Chebyshev spectral method for solving Riemann– Liouville fractional boundary value problems. Applied Mathematics and Computation, 241(??): 140–150, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314006687.

Gawiejnowicz:2010:STD

[GL10a] Stanisław Gawiejnowicz and Bertrand M. T. Lin. Scheduling timedependent jobs under mixed deterioration. Applied Mathematics and Computation, 216(2):438– 447, March 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310000500.

Guo:2010:SNS

[GL10b] Xiaoxia Guo and Chuanxing Li. Solving the nonnegative solution for a (shifted) nonsymmetric algebraic Riccati equation in the critical case. Applied Mathematics and Computation, 216(6):1682–1686, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010996.

Guo:2010:IHH

[GL10c] Zhongjin Guo and A. Y. T. Leung. The iterative homotopy harmonic balance method for conservative Helmholtz–Duffing oscillators. Applied Mathematics and Computation, 215(9):3163–3169, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309008121.

Gao:2011:DAF

[GL11a] Zhe Gao and Xiaozhong Liao. Discretization algorithm for fractional order integral by Haar wavelet approximation. Applied Mathematics and Computation, 218(5):1917–1926, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300311009209.

Guo:2011:ESF

[GL11b] Jun-Hong Guo and Zi-Xing Lu. Exact solution of four cracks originating from an elliptical hole in one-dimensional hexagonal quasicrystals. Applied Mathematics and Computation, 217(22):9397–9403, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005856.

Gadella:2012:DAP

[GL12] M. Gadella and L. P. Lara.

On the determination of approximate periodic solutions of some non-linear ODE. Applied Mathematics and Computation, 218(10):6038–6044, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014317.

Galvin:2013:AAC

[GL13a] Keith Galvin and Hyesuk Lee. Analysis and approximation of the Cross model for quasi-Newtonian flows with defective boundary conditions. Applied Mathematics and Computation, 222(??):244–254, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300756X.

Guo:2013:USA

[GL13b] Geyang Guo and Bo Liu. Unconditional stability of alternating difference schemes with intrinsic parallelism for the fourthorder parabolic equation. Applied Mathematics and Computation, 219(14):7319–7328, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012775.

Guo:2013:SAS

[GL13c] Xiaoyong Guo and Junmin Li. Stochastic adaptive synchronization for time-varying complex delayed dynamical networks with heterogeneous nodes. Applied Mathematics and Computation, 222(??):381–390, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313007807.

Garrett:2014:GIM

[GL14a] C. Kristopher Garrett and Ren-Cang Li. GIP integrators for Matrix Riccati Differential Equations. Applied Mathematics and Computation, 241(??):283–297, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006705.

Guo:2014:HVP

[GL14b] Cuiping Guo and Weiyu Liu. Homogenization of a variational problem in three-dimension space. Applied Mathematics and Computation, 232(??):261–271, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400109X.

Guezane-Lakoud:2012:TDS

[GLB12] A. Guezane-Lakoud and D. Belakroum. Timediscretization schema for an integrodifferential Sobolev type equation with integral conditions. Applied Mathematics and Computation, 218(9):4695–4702, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014238.

Garcia-Ligero:2012:DCF

[GLHCLP12] M. J. García-Ligero, A. Hermoso-Carazo, and J. Linares-Pérez. Distributed and centralized fusion estimation from multiple sensors with Markovian delays. Applied Mathematics and Computation, 219(6):2932–2948, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009162.

Ghosh:2013:BCM

[GLL13] Mini Ghosh, Abid Ali Lashari, and Xue-Zhi Li. Biological control of malaria: a mathematical model. Applied Mathematics and Computation, 219(15):7923–7939, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001859.

Garcia:2014:POB

[GLM14] Isaac A. García, Jaume Llibre, and Susanna Maza. On the periodic orbit bifurcating from a zero Hopf bifurcation in systems with two slow and one fast variables. Applied Mathematics and Computation, 232(??):84–90, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400023X.

Gadella:2011:ISS

[GLP11] M. Gadella, L. P. Lara, and G. P. Pronko. Iterative solution of some nonlinear differential equations. Applied Mathematics and Computation, 217(22):9480–9487, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100628X.

Garcia-Lopez:2012:ECM

[GLR12] C. M. García-López and J. I. Ramos. Effects of convection on a modified GRLW equation. Applied Mathematics and Computation, 219(8):4118–4132, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201082X.

Gao:2013:NSB

[GLS13] Yun Gao, Li-Hua Le, and Bao-Chang Shi. Numerical solution of Burgers’ equation by lattice Boltzmann method. Applied Mathematics and Computation, 219(14):7685–7692, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000970.

Garcia-Laguna:2010:ILS

[GLSJCBS10] Juan García-Laguna, Luis A. San-José, Leopoldo E. Cárdenas-Barrón, and Joaquín Sicilia. The integrality of the lot size in the basic EOQ and EPQ models: Applications to other productioninventory models. Applied Mathematics and Computation, 216(5):1660–1672, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002146.

Gamez:2010:VTS

[GLSV10] M. Gámez, I. López, I. Szabó, and Z. Varga. Verticum-type systems applied to ecological monitoring. Applied Mathematics and Computation, 215(9):3230–3238, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009102.

Goh:2010:OMO

[GLV10] B. S. Goh, C. S. Lee, and T. L. Vincent. ODE methods for optimization with equality constraints. Applied Mathematics and Computation, 217(3):1071–1076, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000263.

Gine:2014:CCG

[GLV14] Jaume Giné, Jaume Llibre, and Claudia Valls. Centers for a class of generalized quintic polynomial differential systems. Applied Mathematics and Computation, 242(??):187–195, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007127.

Guo:2012:EUG

[GLW12] Yunxi Guo, Shaoyong Lai, and Yonghong Wu. On existence and uniqueness of the global weak solution for a shallow water equation. Applied Mathematics and Computation, 218(23):11410–11420, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005140.

Gao:2013:ESC

[GLX13] Cunchen Gao, Zhen Liu, and Ruiping Xu. On exponential stabilization for a class of neutraltype systems with parameter uncertainties: an integral sliding mode approach. Applied Mathematics and Computation, 219(23):11044–11055, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004530.

Gai:2014:ELS

[GLY14] Mingjiu Gai, Xiaolei Liu, and Jian Yuan. Existence of local solutions for a class of delayed neural networks with discontinuous activations. Applied Mathematics and Computation, 245(??):520–525, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008613.

Gao:2010:PSR

[GLZ10a] F. B. Gao, S. P. Lu, and W. Zhang. Periodic solutions for a Rayleigh type p-Laplacian equation with sign-variable coefficient of nonlinear term. Applied Mathematics and Computation, 216(7):2010–2015, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002730.

Geng:2010:EET

[GLZ10b] Yixiang Geng, Jibin Li, and Lixiang Zhang. Exact explicit traveling wave solutions for two nonlinear Schrödinger type equations. Applied Mathematics and Computation, 217(4):1509–1521, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309005931.

Ghelardoni:2010:BCS

[GM10] Paolo Ghelardoni and Cecilia Magherini. BVMs for computing Sturm– Liouville symmetric potentials. Applied Mathematics and Computation, 217(7):3032–3045, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008908.

Georgiou:2011:CDF

[GM11a] D. N. Georgiou and A. C. Megaritis. Covering dimension and finite spaces. Applied Mathematics and Computation, 218(7):3122–3130, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010800.

Gurefe:2011:NVS

[GM11b] Yusuf Gurefe and Emine Misirli. New variable separation solutions of two-dimensional Burgers system. Applied Mathematics and Computation, 217(22):9189–9197, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005376.

Grassi:2012:SCS

[GM12] Giuseppe Grassi and Damon A. Miller. Synchronizing chaotic systems up to an arbitrary scaling matrix via a single signal. Applied Mathematics and Computation, 218(10):6118–6124, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014433.

Gerstl:2013:DWA

[GM13a] Enrique Gerstl and Gur Mosheiov. Due-window assignment problems with unit-time jobs. Applied Mathematics and Computation, 220(??):487–495, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313005924.

Guesmia:2013:STS

[GM13b] Aissa Guesmia and Salim A. Messaoudi. On the stabilization of Timoshenko systems with memory and different speeds of wave propagation. Applied Mathematics and Computation, 219(17):9424–9437, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003664.

Georgiou:2014:APO

[GM14a] D. N. Georgiou and A. C. Megaritis. An algorithm of polynomial order for computing the covering dimension of a finite space. Applied Mathematics and Computation, 231(??):276–283, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000241.

Gu:2014:SNN

[GM14b] Juan Gu and Fanwei Meng. Some new nonlinear Volterra–Fredholm type dynamic integral inequalities on time scales.

Applied Mathematics and Computation, 245(??): 235–242, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314010133.

Guo:2014:RHB

[GM14c] Zhongjin Guo and Xiaoyan Ma. Residue harmonic balance solution procedure to nonlinear delay differential systems. Applied Mathematics and Computation, 237(??):20–30, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004524.

Goulianas:2013:FAR

[GMA13] Konstantinos Goulianas, Athanasios Margaris, and Miltiades Adamopoulos. Finding all real roots of 3 × 3 nonlinear algebraic systems using neural networks. Applied Mathematics and Computation, 219(9):4444–4464, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031201065X.

Garcia-March:2011:UMP

[GMGV^{+}11] Miguel Ángel García-March, Fernando Giménez, Francisco R. Villatoro, Jezabel Pérez, and Pedro Fernández de Córdoba. Unisolvency for multivariate polynomial interpolation in Coatmèlec configurations of nodes. Applied Mathematics and Computation, 217(18):7427–7431, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002244.

Gupta:2012:PTS

[GMKV12] S. Gupta, D. K. Majhi, S. Kundu, and S. K. Vishwakarma. Propagation of torsional surface waves in a homogeneous layer of finite thickness over an initially stressed heterogeneous half-space. Applied Mathematics and Computation, 218(9):5655–5664, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014032.

Guo:2013:FVH

[GML13] Shimin Guo, Liquan Mei, and Ying Li. Fractional variational homotopy perturbation iteration method and its application to a fractional diffusion equation. Applied Mathematics and Computation, 219(11):5909–5917, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012611.

Grigorious:2014:MDC

[GMM^{+}14] Cyriac Grigorious, Paul Manuel, Mirka Miller, Bharati Rajan, and Sudeep Stephen. On the metric dimension of circulant and Harary graphs. Applied Mathematics and Computation, 248(??):47–54, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012703.

Gonzalez-Morales:2013:NII

[GMMIDM13] M. J. González-Morales, R. Mahillo-Isla, and C. Dehesa-Martínez. A new integral identity involving the elliptic integral E(m). Applied Mathematics and Computation, 221(??):568–570, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007303.

Gordoa:2013:GSR

[GMP13] P. R. Gordoa, U. Mugan, and A. Pickering. Generalized scaling reductions and Painlevé hierarchies. Applied Mathematics and Computation, 219(15):8104–8111, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001744.

Guo:2014:CCG

[GMP14] Songbai Guo, Wanbiao Ma, and B. G. Sampath Aruna Pradeep. Complete characterizations of the gamma function. Applied Mathematics and Computation, 244(??):912–916, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009795.

Grossmann:2011:DHO

[GMR11] Christian Grossmann, Ranjan K. Mohanty, and Hans-Goerg Roos. A direct higher order discretization in singular perturbations via domain split — a computational approach. Applied Mathematics and Computation, 217(22):9302–9312, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005686.

Gutierrez:2013:SCN

[GMR13] J. M. Gutiérrez, Á. A. Magreñán, and N. Romero. On the semilocal convergence of Newton– Kantorovich method under center-lipschitz conditions. Applied Mathematics and Computation, 221(??):79–88, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006589.

Gurefe:2013:ETE

[GMSE13] Yusuf Gurefe, Emine Misirli, Abdullah Sonmezoglu, and Mehmet Ekici. Extended trial equation method to generalized nonlinear partial differential equations. Applied Mathematics and Computation, 219(10):5253–5260, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012131.

Guo:2013:SII

[GMSH13] Shuli Guo, Irene Moroz, Ligeng Si, and Lina Han. Several integral inequalities and their applications in nonlinear differential systems. Applied Mathematics and Computation, 219(9):4266–4277, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011320.

Geng:2010:BTM

[GMST10] Tao Geng, Xiang-Hua Meng, Wen-Rui Shan, and Bo Tian. Bäcklund transformation and multisoliton solutions for a (2 + 1)-dimensional Korteweg– de Vries system via symbolic computation. Applied Mathematics and Computation, 217(4):1470– 1475, October 15, 2010.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309005645.

Gutierrez:2011:GSI

[GMV11] José Manuel Gutiérrez, Ángel Alberto Magreñán, and Juan Luis Varona. The “Gauss–Seidelization” of iterative methods for solving nonlinear equations in the complex plane. Applied Mathematics and Computation, 218(6):2467–2479, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010095.

Gupta:2012:TSW

[GMV12] Shishir Gupta, Dinesh Kumar Majhi, and Sumit Kumar Vishwakarma. Torsional surface wave propagation in an initially stressed non-homogeneous layer over a non-homogeneous half-space. Applied Mathematics and Computation, 219(6):3209–3218, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312009575.

Gao:2010:FID

[GMX10] Liang Gao, Wen-Xiu Ma, and Wei Xu. Frobenius integrable decompositions for ninth-order partial differential equations of specific polynomial type. Applied Mathematics and Computation, 216(9):2728–2733, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003851.

Guo:2011:ERE

[GMZL11] Shimin Guo, Liquan Mei, Yubin Zhou, and Chao Li. The extended Riccati equation mapping method for variablecoefficient diffusion–reaction and mKdV equations. Applied Mathematics and Computation, 217(13): 6264–6272, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000026.

Giorno:2012:ROU

[GNdC12] V. Giorno, A. G. Nobile, and R. di Cesare. On the reflected Ornstein– Uhlenbeck process with catastrophes. Applied Mathematics and Computation, 218(23):11570– 11582, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312005498.

Giorno:2014:STN

[GNS14] V. Giorno, A. G. Nobile, and S. Spina. On some time non-homogeneous queueing systems with catastrophes. Applied Mathematics and Computation, 245(??):220–234, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314010339.

Gounden:2010:AFP

[GO10] S. Gounden and J. G.

O’Hara. An analytic formula for the price of an American-style Asian option of floating strike type. Applied Mathematics and Computation, 217(7):2923–2936, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310008799.

Gracia:2012:SPP

[GO12] J. L. Gracia and E. O’Riordan. A singularly perturbed parabolic problem with a layer in the initial condition. Applied Mathematics and Computation, 219(2):498–510, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006273.

Gubaidullin:2013:NIP

[GO13] D. A. Gubaidullin and P. P. Ossipov. Numerical investigation of particle drift in acoustic resonator with periodic shock wave. Applied Mathematics and Computation, 219(9):4535–4544, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010776.

Gulsu:2014:NAS

[GÖ14] Mustafa Gülsu and Yalçin Öztürk. Numerical approach for the solution of hypersingular integro-differential equations. Applied Mathematics and Computation, 230(??):701–710, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014537.

Gocwin:2014:RQC

[Goć14] Maciej Goćwin. Randomized and quantum complexity of nonlinear twopoint BVPs. Applied Mathematics and Computation, 245(??):357–371, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031401073X.

Golan:2011:MG

[Gol11] Shahar Golan. Minesweeper on graphs. Applied Mathematics and Computation, 217(14):6616–6623, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000622.

Golan:2014:MSO

[Gol14] Shahar Golan. Minesweeper strategy for one mine. Applied Mathematics and Computation, 232(??):

292–302, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000824.

Goodrich:2011:EPS

[Goo11] Christopher S. Goodrich. Existence of a positive solution to a system of discrete fractional boundary value problems. Applied Mathematics and Computation, 217(9):4740–4753, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011410.

Goodrich:2013:PSD

[Goo13] Christopher S. Goodrich. Positive solutions to differential inclusions with nonlocal, nonlinear boundary conditions. Applied Mathematics and Computation, 219(24):11071–11081, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313005420.

Gorguis:2010:MMI

[Gor10] Alice Gorguis. Monge’s method for intermediate integrals. Applied Mathematics and Computation, 216(8):2310–2316, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003218.

Gulsu:2010:NCM

[GÖS10] Mustafa Gülsu, Yalçın Öztürk, and Mehmet Sezer. A new collocation method for solution of mixed linear integro-differential-difference equations. Applied Mathematics and Computation, 216(7):2183– 2198, June 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310002961.

Gulsu:2011:SAE

[GÖS11] Mustafa Gülsu, Yalçın Öztürk, and Mehmet Sezer. On the solution of the Abel equation of the second kind by the shifted Chebyshev polynomials. Applied Mathematics and Computation, 217(9):4827–4833, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011732.

Goyal:2014:TEE

[Goy14] Ashish Goyal. Tobacco epidemics: Effect of marketing bans and awareness programs on its spread. Applied Mathematics and Computation, 247(??):1030–1051, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013010.

Gasinski:2010:NSN

[GP10] Leszek Gasiński and Nikolaos S. Papageorgiou. Nontrivial solutions for Neumann problems with an indefinite linear part. Applied Mathematics and Computation, 217(6):2666– 2675, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008544.

Goncalves:2011:CES

[GP11a] Eric Goncalvès and Regiane Fortes Patella. Constraints on equation of state for cavitating flows with thermodynamic effects. Applied Mathematics and Computation, 217(11):5095–5102, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008015.

Gordoa:2011:NDP

[GP11b] P. R. Gordoa and A. Pickering. A new derivation of Painlevé hierarchies. Applied Mathematics and Computation, 218(7):3942–3949, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011258.

Garra:2012:ASF

[GP12] Roberto Garra and Federico Polito. Analytic solutions of fractional differential equations by operational methods. Applied Mathematics and Computation, 218(21):10642– 10646, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004274.

Gopi:2013:FCP

[GP13] E. S. Gopi and P. Palanisamy.

Fast computation of PCA bases of image subspace using its inner-product subspace. Applied Mathematics and Computation, 219(12):6729–6732, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000179.

Gonzalez-Parra:2014:PCR

[GPCCA14] Gilberto González-Parra, Benito Chen-Charpentier, and Abraham J. Arenas. Polynomial chaos for random fractional order differential equations. Applied Mathematics and Computation, 226(??):123–130, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011089.

George:2013:NLR

[GPK13] Santhosh George, Suresan Pareth, and M. Kunhanandan. Newton Lavrentiev regularization for illposed operator equations in Hilbert scales. Applied Mathematics and Computation, 219(24):11191– 11197, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313005419.

Gan:2010:ADM

[GPP^{+}10] Min Gan, Hui Peng, Xiaoyan Peng, Xiaohong Chen, and Garba Inoussa. An adaptive decision maker for constrained evolutionary optimization. Applied Mathematics and Computation, 215(12):4172–4184, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010856.

Gordoa:2011:IPG

[GPP11] P. R. Gordoa, A. Pickering, and J. Prada. Integrability properties of a generalized reduction of the KdV6 equation. Applied Mathematics and Computation, 218(6):2968–2972, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010332.

Gordoa:2014:NSP

[GPW14] P. R. Gordoa, A. Pickering, and J. A. D. Wattis. Nonisospectral scattering problems and similarity reductions. Applied Mathematics and Computation, 237(??):77–84, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004780.

Ghezzi:2012:OSP

[GPZ12] L. Ghezzi, L. F. Pavarino, and E. Zampieri. Overlapping Schwarz preconditioned eigensolvers for spectral element discretizations. Applied Mathematics and Computation, 218(15):7700–7710, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000525.

Guo:2011:SBH

[GQ11] Bai-Ni Guo and Feng Qi. Sharp bounds for harmonic numbers. Applied Mathematics and Computation, 218(3):991–995, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/

pii/S009630031100124X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Guo:2012:CMF

[GQ12] Bai-Ni Guo and Feng Qi.

A completely monotonic function involving the trigamma function and with degree one. Applied Mathematics and Computation, 218(19):9890–9897, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003335.

Grajales:2013:MIN

[GQ13] Juan Carlos Muñoz Grajales and Julio César Quiceno. Modulation instability in nonlinear propagation of pulses in optical fibers. Applied Mathematics and Computation, 221(??):177–191, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006978.

Geng:2014:NRK

[GQ14] F. Z. Geng and S. P.

Qian. A new reproducing kernel method for linear nonlocal boundary value problems. Applied Mathematics and Computation, 248(??):421–425, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013678.

Giebel:2011:SPA

[GR11] Stefan Giebel and Martin Rainer. Stochastic processes adapted by neural networks with application to climate, energy, and finance. Applied Mathematics and Computation, 218(3):1003–1007, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004899.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Grabski:2011:SMF

[Gra11] Franciszek Grabski. Semi-Markov failure rates processes. Applied Mathematics and Computation, 217(24):9956–9965, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311006254.

Graham:2013:MCS

[Gra13] Jason M. Graham. A measure of control for secondary cytokine-induced injury of articular cartilage: a computational study. Applied Mathematics and Computation, 223(??):225–236, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008229.

Gomez-Ruiz:2010:EMJ

[GRKP10] Jose Antonio Gomez-Ruiz, Marcelo Karanik, and José Ignacio Peláez. Estimation of missing judgments in AHP pairwise matrices using a neural network-based model. Applied Mathematics and Computation, 216(10): 2959–2975, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310004297.

Grothe:2013:HOC

[Gro13] Oliver Grothe. A higher order correlation unscented Kalman filter. Applied Mathematics and Computation, 219(17): 9033–9042, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313002725.

Garralon:2013:RTT

[GRV13] Julio Garralón, Francisco Rus, and Francisco R. Villatoro. Removing trailing tails and delays induced by artificial dissipation in Padé numerical schemes for stable compacton collisions. Applied Mathematics and Computation, 220(??):185–192, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006309.

Gupta:2014:ALB

[GRY14] Vijay Gupta, Themistocles M. Rassias, and Rani Yadav. Approximation by Lupas–Beta integral operators. Applied Mathematics and Computation, 236(??):19–26, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003865.

Geem:2010:PSF

[GS10a] Zong Woo Geem and Kwee-Bo Sim. Parametersetting-free harmony search algorithm. Applied Mathematics and Computation, 217(8):3881–3889, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001009X.

Gomez:2010:VIM

[GS10b] Cesar A. Gómez and Alvaro H. Salas. The variational iteration method combined with improved generalized tanh–coth method applied to Sawada–Kotera equation. Applied Mathematics and Computation, 217(4):1408–1414, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309005347.

Gou:2011:NSG

[GS11a] Kun Gou and Bo Sun.

Numerical solution of the Goursat problem on a triangular domain with mixed boundary conditions. Applied Mathematics and Computation, 217(21):8765–8777, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004978.

Guesmia:2011:ASP

[GS11b] Senoussi Guesmia and Abdelmouhcene Sengouga. Anisotropic singular perturbations of hyperbolic problems. Applied Mathematics and Computation, 217(22):8983–8996, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004723.

Gurdal:2011:SRT

[GS11c] M. Gürdal and F. Söhret. Some results for Toeplitz operators on the Bergman space. Applied Mathematics and Computation, 218(3):789–793, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001044.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Gurudwan:2011:CTC

[GS11d] Niyati Gurudwan and B. K. Sharma. Convergence theorem for the common solution for a finite family of ϕ-strongly accretive operator equations. Applied Mathematics and Computation, 217(15):6748–6754, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001329.

See note [ZS15].

Gajic:2012:CGM

[GS12a] Lj. Gajić and M. Stojaković. On Ćirić generalization of mappings with a contractive iterate at a point in G-metric spaces. Applied Mathematics and Computation, 219(1):435–441, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006406.

Galewski:2012:SDB

[GS12b] Marek Galewski and Joanna Smejda. On some discrete boundary value problem with parameters. Applied Mathematics and Computation, 218(21):10708– 10716, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004390.

Giatili:2012:DNT

[GS12c] Stavroula G. Giatili and Georgios S. Stamatakos. A detailed numerical treatment of the boundary conditions imposed by the skull on a diffusionreaction model of glioma tumor growth. Clinical validation aspects. Applied Mathematics and Computation, 218(17):8779–8799, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001634.

Guerra:2012:NPT

[GS12d] Ana-Isabel Guerra and Ferran Sancho. A nonpossibility theorem for joint stability in interindustry models. Applied Mathematics and Computation, 219(8):4045– 4048, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010442.

Galewski:2013:DPM

[GS13a] Marek Galewski and Joanna Smejda. On the dependence on parameters for mountain pass solutions of second order discrete BVP’s. Applied Mathematics and Computation, 219(11):5963–5971, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013045.

Gambino:2013:IPB

[GS13b] G. Gambino and V. Sciacca. Intermittent and passivity based control strategies for a hyperchaotic system. Applied Mathematics and Computation, 221(??):367–382, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313007388.

Gonzalez:2013:IAI

[GS13c] Luis González and Antonio Suárez. Improving approximate inverses based on Frobenius norm minimization. Applied Mathematics and Computation, 219(17):9363–9371, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300310X.

Gaur:2014:GIS

[GS14a] M. Gaur and K. Singh. On group invariant solutions of fractional order Burgers–Poisson equation. Applied Mathematics and Computation, 244(??):870–877, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010108.

Guan:2014:HAN

[GS14b] Hong-Bo Guan and Dong-Yang Shi. A high accuracy NFEM for constrained optimal control problems governed by elliptic equations. Applied Mathematics and Computation, 245(??):382–390, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010340.

Guerrini:2014:PFD

[GS14c] Luca Guerrini and Mauro Sodini. Persistent fluctuations in a dual model with frictions: The role of delays. Applied Mathematics and Computation, 241(??):371–379, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006936.

Gupta:2014:ODN

[GS14d] Rekha Gupta and M. Srivastava. Optimality and duality for nonsmooth multiobjective programming using G-type I functions. Applied Mathematics and Computation, 240(??):294–307, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006067.

Gupta:2014:ACG

[GS14e] Vijay Gupta and Danyal Soybas. Approximation by complex genuine hybrid operators. Applied Mathematics and Computation, 244(??):526–532, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314009825.

Gurbuz:2014:LPA

[GS14f] Burcu Gürbüz and Mehmet Sezer. Laguerre polynomial approach for solving Lane–Emden type functional differential equations. Applied Mathematics and Computation, 242(??):255–264, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007231.

Guo:2014:QAK

[GSC14] Hongjian Guo, Xinyu Song, and Lansun Chen. Qualitative analysis of a Korean pine forest model with impulsive thinning measure. Applied Mathematics and Computation, 234(??):203–213, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002732.

Grau-Sanchez:2011:ZFM

[GSDB11] Miquel Grau-Sánchez and José Luis Dı´ az-Barrero. Zero-finder methods derived using Runge–Kutta techniques. Applied Mathematics and Computation, 217(12):5366–5376, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011926.

Guo:2014:GSS

[GSDW14] Ying Guo, Huan Su, Xiaohua Ding, and Ke Wang. Global stochastic stability analysis for stochastic neural networks with infinite delay and Markovian switching. Applied Mathematics and Computation, 245(??):53–65, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009382.

Giraldo:2012:HTN

[GSF12] Mauricio Giraldo, Daniel Sanín, and Whady F. Flórez. Heat transfer in nanofluids: a computational evaluation of the effects of particle motion. Applied Mathematics and Computation, 219(7):3308–3315, December 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311010964.

Grau-Sanchez:2011:OTM

[GSGN11] Miquel Grau-Sánchez, Àngela Grau, and Miquel Noguera. Ostrowski type methods for solving systems of nonlinear equations. Applied Mathematics and Computation, 218(6):2377–2385, November 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031101037X.

Gerbi:2012:EES

[GSH12] Stéphane Gerbi and Belkacem Said-Houari. Existence and exponential stability of a damped wave equation with dynamic boundary conditions and a delay term. Applied Mathematics and Computation, 218(24):11900–11910, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200570X.

Grau-Sanchez:2012:TCM

[GSN12] Miquel Grau-Sánchez and Miquel Noguera. A technique to choose the most efficient method between secant method and some variants. Applied Mathematics and Computation, 218(11):6415–6426, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101469X.

Gupta:2013:SHT

[GSRR13] Praveen Kumar Gupta, Jitendra Singh, K. N. Rai, and S. K. Rai. Solution of the heat transfer problem in tissues during hyperthermia by finite difference-decomposition method. Applied Mathematics and Computation, 219(12):6882–6892, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000581.

Grbic:2013:ASM

[GSSS13] Ratko Grbić, Klaudija Scitovski, Kristian Sabo, and Rudolf Scitovski. Approximating surfaces by the moving least absolute deviations method. Applied Mathematics and Computation, 219(9):4387–4399, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010430.

Gzyl:2010:HMP

[GT10a] H. Gzyl and A. Tagliani. Hausdorff moment problem and fractional moments. Applied Mathematics and Computation, 216(11):3319–3328, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000487X.

Gzyl:2010:SMP

[GT10b] H. Gzyl and A. Tagliani. Stieltjes moment problem and fractional moments. Applied Mathematics and Computation, 216(11):3307–3318, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004856.

Geiser:2011:OSM

[GT11a] Jürgen Geiser and Gamze Tanoglu. Operatorsplitting methods via the Zassenhaus product formula. Applied Mathematics and Computation, 217(9):4557–4575, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011197.

Gu:2011:FVP

[GT11b] Chuanqing Gu and Youtian Tao. Function-valued Padé-type approximant via E-algorithm and its applications in solving integral equations. Applied Mathematics and Computation, 217(20):7975–7984, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003134.

Gucuyenen:2011:NSK

[GT11c] Nurcan Gücüyenen and Gamze Tanoğlu. On the numerical solution of Korteweg–de Vries equation by the iterative splitting method. Applied Mathematics and Computation, 218(3):777–782, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004528.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Gzyl:2012:DDT

[GT12] Henryk Gzyl and Aldo Tagliani. Determination of the distribution of total loss from the fractional moments of its exponential. Applied Mathematics and Computation, 219(4):2124–2133, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008508.

Ganesan:2013:OSF

[GT13] Sashikumaar Ganesan and Lutz Tobiska. Operatorsplitting finite element algorithms for computations of high-dimensional parabolic problems. Applied Mathematics and Computation, 219(11): 6182–6196, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312013033.

Gzyl:2013:LTI

[GTM13] H. Gzyl, A. Tagliani, and M. Milev. Laplace transform inversion on the real line is truly ill-conditioned. Applied Mathematics and Computation, 219(18):9805–9809, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003731.

Gobithaasan:2014:GLA

[GTPM14] R. U. Gobithaasan, Y. M.

Teh, A. R. M. Piah, and Kenjiro T. Miura. Generation of log-aesthetic curves using adaptive Runge– Kutta methods. Applied Mathematics and Computation, 246(??):257–262, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314011266.

Gulec:2013:NAG

[GTU13] H. H. Gulec, N. Taskara, and K. Uslu. A new approach to generalized Fibonacci and Lucas numbers with binomial coefficients. Applied Mathematics and Computation, 220(??):482–486, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005900.

Guo:2011:ASP

[GTX11] Tao Guo, Qiulin Tan, and Jijun Xiong. Analysis of spatial patterns in a vegetation model. Applied Mathematics and Computation, 217(21):8303–8310, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003511.

Gu:2010:SCT

[Gu10] Feng Gu. Some convergence theorems of nonimplicit iteration process with errors for a finite families of I-asymptotically nonexpansive mappings. Applied Mathematics and Computation, 216(1):161– 172, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031000038X.

Gu:2011:NFD

[Gu11]  Chao Gu.  Nonmono-

tone filter DQMM method for the system of nonlinear equations. Applied Mathematics and Computation, 217(22):9351–9357, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311005807.

Gu:2013:DFI

[Gu13]  Chao Gu.  A dwin-

dling filter inexact projected Hessian algorithm for large scale nonlinear constrained optimization. Applied Mathematics and Computation, 219(23):10898–10908, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005304.

Guessab:2014:ADC

[Gue14a]  Allal Guessab.  Ap-

proximations of differentiable convex functions on arbitrary convex polytopes. Applied Mathematics and Computation, 240(??):326–338, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006158.

Guessab:2014:SPC

[Gue14b]  Allal Guessab.  Sym-

metrization procedures and convexity in centrally symmetric polytopes. Applied Mathematics and Computation, 243(??):74–82, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007711.

Guin:2014:ESP

[Gui14] Lakshmi Narayan Guin. Existence of spatial patterns in a predator-prey model with self- and cross-diffusion. Applied Mathematics and Computation, 226(??):320–335, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031301059X.

Gungor:2010:CGA

[Gün10] F. Güngör. Comment on “Group analysis of KdV equation with time dependent coefficients”. Applied Mathematics and Computation, 217(8):4293–4294, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010192.

See [JK10].

Guo:2010:ASW

[Guo10]  Hui Guo.  Analysis

of split weighted leastsquares procedures for pseudo-hyperbolic equations. Applied Mathematics and Computation, 217(8):4109–4121, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010593.

Guo:2011:RSL

[Guo11] Hui Guo. A remark on split least-squares mixed element procedures for pseudo-parabolic equations. Applied Mathematics and Computation, 217(9):4682–4690, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011331.

Guo:2013:LCM

[Guo13a] Senlin Guo. Logarithmically completely monotonic functions and applications. Applied Mathematics and Computation, 221(??):169–176, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313006760.

Guo:2013:SPC

[Guo13b] Senlin Guo. Some properties of completely monotonic sequences and related interpolation. Applied Mathematics and Computation, 219(10): 4958–4962, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012404.

Guo:2014:BRI

[Guo14]  Xiurong Guo.  On

bilinear representations and infinite conservation laws of a nonlinear variable-coefficient equation. Applied Mathematics and Computation, 248(??):531–535, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013630.

Gupta:2013:CIO

[Gup13]  Vijay Gupta.  Combi-

nations of integral operators. Applied Mathematics and Computation, 224(??):876–881, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009934.

Gurarslan:2010:NML

[Gür10] Gürhan Gürarslan. Numerical modelling of linear and nonlinear diffusion equations by compact finite difference method. Applied Mathematics and Computation, 216(8):2472– 2478, June 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310003474.

See erratum [Gür12].

Gurarslan:2012:ENM

[Gür12] Gürhan Gürarslan. Erratum to: “Numerical modelling of linear and nonlinear diffusion equations by compact finite difference method” [Applied Mathematics and Computation, 216 (8) (2010) 2472–2478]. Applied Mathematics and Computation, 218(9):5840, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012239.

See [Gür10].

Guseinov:2012:IPT

[Gus12] Gusein Sh. Guseinov. On an inverse problem for two spectra of finite Jacobi matrices. Applied Mathematics and Computation, 218(14):7573–7589, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000598.

Gulati:2013:NHO

[GV13] T. R. Gulati and Khushboo Verma. A note on higher-order symmetric duality. Applied Mathematics and Computation, 222(??):553–558, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300814X.

Galiano:2014:NSD

[GV14] G. Galiano and J. Velasco. On a nonlocal spectrogram for denoising one-dimensional signals. Applied Mathematics and Computation, 244(??):859–869, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300314009606.

Gomez-Villegas:2013:AEK

[GVMNS13] M. A. Gómez-Villegas, P. Main, H. Navarro, and R. Susi. Assessing the effect of kurtosis deviations from Gaussianity on conditional distributions. Applied Mathematics and Computation, 219(21): 10499–10505, July 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004463.

Ghostine:2013:CSA

[GVT^{+}13] R. Ghostine, J. Vazquez, A. Terfous, N. Rivière, A. Ghenaim, and R. Mosé. A comparative study of 1D and 2D approaches for simulating flows at right angled dividing junctions. Applied Mathematics and Computation, 219(10):5070–5082, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012155.

Gao:2014:GSA

[GVW^{+}14] Shangce Gao, Catherine Vairappan, Yan Wang, Qiping Cao, and Zheng Tang. Gravitational search algorithm combined with chaos for unconstrained numerical optimization. Applied Mathematics and Computation, 231(??):48–62, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000149.

Gao:2011:HQI

[GW11a] Jiaquan Gao and Jun Wang. A hybrid quantuminspired immune algorithm for multiobjective optimization. Applied Mathematics and Computation, 217(9):4754–4770, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011422.

Guo:2011:WEF

[GW11b] Yunxi Guo and Ying Wang. On Weierstrass elliptic function solutions for a (N + 1) dimensional potential KdV equation. Applied Mathematics and Computation, 217(20):8080–8092, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003468.

Guo:2012:MHI

[GW12] Xiao-Xia Guo and Shen Wang. Modified HSS iteration methods for a class of non-Hermitian positive-definite linear systems. Applied Mathematics and Computation, 218 (20):10122–10128, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003517.

Grover:2013:DMM

[GW13] James P. Grover and Feng-Bin Wang. Dynamics of a model of microbial competition with internal nutrient storage in a flowing habitat. Applied Mathematics and Computation, 225(??):747–764, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010333.

Ge:2014:MRF

[GWA14] Suqin Ge, Wanyi Wang, and Jijun Ao. Matrix representations of fourth order boundary value problems with periodic boundary conditions. Applied Mathematics and Computation, 227(??):601–609, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012289.

Guan:2012:OPD

[GWH12] Kaizhong Guan, Qisheng Wang, and Xiaobao He. Oscillation of a pantograph differential equation with impulsive perturbations. Applied Mathematics and Computation, 219(6):3147–3153, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009472.

Gwizdalla:2011:RCO

[Gwi11] Tomasz M. Gwizdalla. The role of crossover operator in the genetic optimization of magnetic models. Applied Mathematics and Computation, 217(22):9368–9379, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311005820.

Gwizdalla:2012:RDG

[Gwi12] Tomasz M. Gwizdalla.

The role of different genetic operators in the optimization of magnetic models. Applied Mathematics and Computation, 218(18):9220–9233, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002299.

Geng:2010:SBD

[GWLZ10] Xinyu Geng, Yin Wang, Jeonghwa Lee, and Jun Zhang. Stabilizing block diagonal preconditioners for complex dense matrices in electromagnetics. Applied Mathematics and Computation, 217 (5):1983–1990, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007411.

Gao:2010:NGO

[GWM10] Yuelin Gao, Guorong Wu, and Weimin Ma. A new global optimization approach for convex multiplicative programming. Applied Mathematics and Computation, 216(4):1206–1218, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001785.

Guo:2012:LSB

[GWML12] X. C. Guo, C. G. Wu, M. Marchese, and Y. C. Liang. LS-SVR-based solving Volterra integral equations. Applied Mathematics and Computation, 218(23):11404–11409, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005139.

Ge:2013:DEC

[GWS13] Suqin Ge, Wanyi Wang, and Jianqing Suo. Dependence of eigenvalues of a class of fourth-order Sturm–Liouville problems on the boundary. Applied Mathematics and Computation, 220(??):268–276, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313006437.

Gao:2014:NFO

[GWW14] Lijun Gao, Dandan Wang, and Yuqiang Wu. Nonfragile observer-based sliding mode control for Markovian jump systems with mixed modedependent time delays and input nonlinearity. Applied Mathematics and Computation, 229(??):374–395, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313012885.

Gang:2012:NPS

[GWX12] Ma Gang, Zhou Wei, and Chang Xiaolin. A novel particle swarm optimization algorithm based on particle migration. Applied Mathematics and Computation, 218(11): 6620–6626, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300311015062.

Gao:2014:ESA

[GWZ14] Jin Gao, Qi-Ru Wang, and Liu-Wei Zhang. Existence and stability of almost-periodic solutions for cellular neural networks with time-varying delays in leakage terms on time scales. Applied Mathematics and Computation, 237(??):639–649, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004135.

Gao:2011:BTP

[GX11] Chenghua Gao and Jia Xu. Bifurcation techniques and positive solutions of discrete Lidstone boundary value problems. Applied Mathematics and Computation, 218(2):434–444, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007648.

Geng:2012:SQP

[GX12] Xianguo Geng and Bo Xue. N-soliton and quasi-periodic solutions of the KdV6 equations. Applied Mathematics and Computation, 219(8):3504–3510, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009241.

Gu:2013:USR

[GX13] Anhui Gu and Hongjun Xiang. Upper semicontinuity of random attractors for stochastic three-component reversible Gray–Scott system. Applied Mathematics and Computation, 225(??):387–400, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010205.

Ge:2011:SBT

[GXC^{+}11] Lin Ge, Luo Xiaonan, Li Chunjing, Shi Xiquan, Li Yi, and Wang Ruomei. The skim of balance theory of 3D garment simulation. Applied Mathematics and Computation, 218(2):492–501, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007727.

Guo:2013:RCA

[GXF13] Huan Guo, Xinping Xiao, and Jeffrey Forrest. A research on a comprehensive adaptive grey prediction model CAGM(1,N). Applied Mathematics and Computation, 225(??):

216–227, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031300982X.

Gao:2012:NAB

[GxFX^{+}12] Fei Gao, Feng xia Fei, Qian Xu, Yan fang Deng, Yi bo Qi, and Ilangko Balasingham. A novel artificial bee colony algorithm with space contraction for unknown parameters identification and time-delays of chaotic systems. Applied Mathematics and Computation, 219(2):552–568, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200639X.

Gao:2014:KBB

[GXSH14] Jianqiang Gao, Lizhong Xu, Aiye Shi, and Fengchen Huang. A kernel-based block matrix decomposition approach for the classification of remotely sensed images. Applied Mathematics and Computation, 228(??):531–545, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313012757.

Gupta:2012:DES

[GY12] Vijay Gupta and Rani Yadav. Direct estimates in simultaneous approximation for BBS operators. Applied Mathematics and Computation, 218(22):11290–11296, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005061.

Gao:2013:ASS

[GY13] Fangzheng Gao and Fushun Yuan. Adaptive stabilization of stochastic nonholonomic systems with nonlinear parameterization. Applied Mathematics and Computation, 219(16):8676–8686, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002622.

Ghayesh:2010:VSA

[GYBR10] Mergen H. Ghayesh, Mostafa Yourdkhani, Sara Balar, and Tyler Reid. Vibrations and stability of axially traveling laminated beams. Applied Mathematics and Computation, 217(2):545–556, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006570.

Girgin:2014:CMS

[GYD14] Zekeriya Girgin, Yasin Yilmaz, and Ersin Demir. A Combining Method for solution of nonlinear boundary value problems. Applied Mathematics and Computation, 232(??):1037–1045, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001787.

Guo:2011:KRA

[GYG11] Litao Guo, Weihua Yang, and Xiaofeng Guo. On a kind of reliability analysis of networks. Applied Mathematics and Computation, 218(6):2711–2715, November 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311010368.

Gan:2013:EMC

[GYL^{+}13] Chenquan Gan, Xiaofan Yang, Wanping Liu, Qingyi Zhu, and Xulong Zhang. An epidemic model of computer viruses with vaccination and generalized nonlinear incidence rate. Applied Mathematics and Computation, 222(??):265–274, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008059.

Guo:2014:RVS

[GYL14] Chaohui Guo, Hu Yang, and Jing Lv. Robust variable selection in semiparametric meancovariance regression for longitudinal data analysis. Applied Mathematics and Computation, 245(??):343–356, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010509.

Guo:2013:SIM

[GYS13] Jun-Hong Guo, Jing Yu, and Riguleng Si. A semi-inverse method of a Griffith crack in onedimensional hexagonal quasicrystals. Applied Mathematics and Computation, 219(14):7445–7449, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000702.

Gao:2010:NES

[GZ10a] Hua Gao and Rong-Xia Zhao. New exact solutions to the generalized Burgers–Huxley equation. Applied Mathematics and Computation, 217(4):1598–1603, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309006705.

Geng:2010:BTW

[GZ10b] Yixiang Geng and Lixiang Zhang. Bifurcations of traveling wave solutions for the magma equation. Applied Mathematics and Computation, 217(4):1741–1748, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010248.

Guo:2010:AEM

[GZ10c] Shimin Guo and Yubin Zhou. Auxiliary equation method for the mKdV equation with variable coefficients. Applied Mathematics and Computation, 217(4):1476–1483, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309005840.

Guo:2010:EEM

[GZ10d] Shimin Guo and Yubin Zhou. The extended (\frac{G′}{G})- expansion method and its applications to the Whitham–Broer–Kaup– Like equations and coupled Hirota–Satsuma KdV equations. Applied Mathematics and Computation, 215(9):3214–3221, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009084.

Guo:2011:CMT

[GZ11] Haojie Guo and Sining Zheng. A competition model for two resources in un-stirred chemostat. Applied Mathematics and Computation, 217(16):6934–6949, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100141X.

Gu:2012:FAN

[GZ12a] Chao Gu and Detong Zhu. A filter algorithm for nonlinear systems of equalities and inequalities. Applied Mathematics and Computation, 218 (20):10289–10298, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312003906.

Guettal:2012:RTG

[GZ12b] Djaouida Guettal and Abdelkader Ziadi. Reducing transformation and global optimization. Applied Mathematics and Computation, 218(10):5848–5860, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013968.

Georgescu:2013:LFS

[GZ13] Paul Georgescu and Hong Zhang. A Lyapunov functional for a SIRI model with nonlinear incidence of infection and relapse. Applied Mathematics and Computation, 219(16):8496–8507, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001756.

Georgescu:2014:LFT

[GZ14a] Paul Georgescu and Hong Zhang. Lyapunov functionals for two-species mutualisms. Applied Mathematics and Computation, 226(??):754–764, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011181.

Gu:2014:DFT

[GZ14b] Chao Gu and Detong Zhu.

A dwindling filter trust region algorithm for nonlinear optimization. Applied Mathematics and Computation, 240(??):72–81, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006419.

Graca:2012:CNH

[GZB12] Daniel S. Graça, Ning Zhong, and Jorge Buescu. Computability, noncomputability, and hyperbolic systems. Applied Mathematics and Computation, 219(6):3039–3054, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009307.

Guo:2012:TSP

[GZF12] Hui Guo, Jiansong Zhang, and Hongfei Fu. Two splitting positive definite mixed finite element methods for parabolic integro-differential equations. Applied Mathematics and Computation, 218(22):11255–11268, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005036.

Guo:2014:NAG

[GZL14] P. F. Guo, L. W. Zhang, and K. M. Liew. Numerical analysis of generalized regularized long wave equation using the element-free kp-Ritz method. Applied Mathematics and Computation, 240(??):91– 101, August 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314005475.

Zhang:2010:NESc

[gZrLfM10] Ben gong Zhang, Zheng rong Liu, and Jian feng Mao. New exact solutions for mCH and mDP equations by auxiliary equation method. Applied Mathematics and Computation, 217(4):1306–1314, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030900441X.

Zhang:2010:NESb

[gZrLX10] Ben gong Zhang, Zheng rong Liu, and Qing Xiao. New exact solitary wave and multiple soliton solutions of quantum Zakharov–Kuznetsov equation. Applied Mathematics and Computation, 217(1):392–402, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006338.

Guo:2012:RCM

[GZW12] Congchong Guo, Zujin Zhang, and Jialin Wang. Regularity criteria for the 3D magnetomicropolar fluid equations in Besov spaces with negative indices. Applied Mathematics and Computation, 218(21):10755– 10758, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004833.

Gu:2013:UAN

[GZW13] Anhui Gu, Shengfan Zhou, and Zhaojuan Wang. Uniform attractor of non-autonomous threecomponent reversible Gray– Scott system. Applied Mathematics and Computation, 219(16):8718–8729, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300194X.

Guo-Zhong:2010:NSB

[GZXJD10] Zhao Guo-Zhong, Yu Xi-Jun, and Wu Di. Numerical solution of the Burgers’ equation by local discontinuous Galerkin method. Applied Mathematics and Computation, 216(12):3671–3679, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005655.

Guo:2010:IEM

[GZZ10] Shimin Guo, Yubin Zhou, and Chenxia Zhao. The improved (\frac{G′}{G})-expansion method and its applications to the Broer–Kaup equations and approximate long water wave equations. Applied Mathematics and Computation, 216(7):1965–1971, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002687.

Hassan:2010:SWS

[HA10] Saleh M. Hassan and Naif M. Alotaibi. Solitary wave solutions of the improved KdV equation by VIM. Applied Mathematics and Computation, 217(6):2397–2403, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000785X.

Hussain:2011:CFP

[HA11] N. Hussain and M. Abbas. Common fixed point results for two new classes of hybrid pairs in symmetric spaces. Applied Mathematics and Computation, 218(2):542–547, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100779X.

Hedayati:2012:IGA

[HA12] H. Hedayati and B. Sobhani Aragh. Influence of graded agglomerated CNTs on vibration of CNT-reinforced annular sectorial plates resting on Pasternak foundation. Applied Mathematics and Computation, 218(17):8715–8735, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001579.

Hasan:2014:HHS

[HAA^{+}14] Basima Hani F. Hasan, Iyad Abu Doush, Eslam Al Maghayreh, Faisal Alkhateeb, and Mohammad Hamdan. Hybridizing Harmony Search algorithm with different mutation operators for continuous problems. Applied Mathematics and Computation, 232(??):1166–1182, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301429X.

Hajji:2011:NSM

[Haj11] M. A. Hajji. A numerical scheme for multi-point special boundary-value problems and application to fluid flow through porous layers. Applied Mathematics and Computation, 217(12):5632–5642, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310012439.

Hayat:2014:BLF

[HAMA14] T. Hayat, Sadia Asad, M. Mustafa, and A. Alsaedi. Boundary layer flow of Carreau fluid over a convectively heated stretching sheet. Applied Mathematics and Computation, 246(??):12–22, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314010406.

See comment [PK18].

Han:2011:SPB

[Han11]  Zhiqing Han.  Solu-

tions of periodic boundary value problems for secondorder nonlinear differential equations via variational methods. Applied Mathematics and Computation, 217(14):6516–6525, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000440.

Han:2014:BTI

[Han14] Guangguo Han. Blocktransitive 2 − (ν, k, 1) designs and the twisted simple groups ^{2}E_{6}(q). Applied Mathematics and Computation, 242(??):109–115, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006924.

Hao:2010:VIC

[Hao10]  Yan Hao.  On varia-

tional inclusion and common fixed point problems in Hilbert spaces with applications. Applied Mathematics and Computation, 217(7):3000–3010, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008878.

Harley:2010:HMN

[Har10]  C. Harley.  Hopscotch

method: The numerical solution of the Frank– Kamenetskii partial differential equation. Applied Mathematics and Computation, 217(8):4065–4075, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310010507.

Harfash:2014:TDS

[Har14] A. J. Harfash. Three dimensional simulation of radiation induced convection. Applied Mathematics and Computation, 227(??):92–101, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011673.

Hasanov:2010:NTP

[Has10] Vejdi I. Hasanov. Notes on two perturbation estimates of the extreme solutions to the equations X ± A^{⋆}X^{−1}A = Q. Applied Mathematics and Computation, 216(5):1355–1362, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000216X.

Hassan:2011:KTO

[Has11] Taher S. Hassan. Kamenevtype oscillation criteria for second order nonlinear dynamic equations on time scales. Applied Mathematics and Computation, 217(12):5285–5297, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011811.

Hassan:2014:NAS

[Has14] Mohamed F. Hassan. A new approach for secure communication using constrained hyperchaotic systems. Applied Mathematics and Computation, 246(??):711–730, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314011230.

Hayek:2010:CES

[Hay10] Mohamed Hayek. Constructing of exact solutions to the KdV and Burgers equations with power-law nonlinearity by the extended (\frac{G′}{G})- expansion method. Applied Mathematics and Computation, 217(1):212– 221, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310005898.

Hayek:2011:ETW

[Hay11] Mohamed Hayek. Exact and traveling-wave solutions for convectiondiffusion-reaction equation with power-law nonlinearity. Applied Mathematics and Computation, 218(6):2407–2420, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009672.

Hammond:2011:ASF

[HB11a] Jason F. Hammond and David M. Bortz. Analytical solutions to Fisher’s equation with time-variable coefficients. Applied Mathematics and Computation, 218(6):2497–2508, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010137.

Hoang:2011:PES

[HB11b] Hong Son Hoang and Remy Baraille. Prediction error sampling procedure based on dominant Schur decomposition. Application to state estimation in high dimensional oceanic model. Applied Mathematics and Computation, 218(7):3689–3709, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011787.

Heydari:2013:OMT

[HB13] Ali Heydari and S. N. Balakrishnan. Optimal multitherapeutic HIV treatment using a global optimal switching scheme. Applied Mathematics and Computation, 219(14): 7872–7881, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

com/science/article/
pii/S009630031300132X.

Hajiazizi:2014:EAT

[HB14] M. Hajiazizi and P. Bastan. The elastoplastic analysis of a tunnel using the EFG method: a comparison of the EFGM with FEM and FDM. Applied Mathematics and Computation, 234(??):82–113, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400263X.

Hasseine:2011:ASD

[HBB11] A. Hasseine, A. Bellagoun, and H.-J. Bart. Analytical solution of the droplet breakup equation by the Adomian decomposition method. Applied Mathematics and Computation, 218(5):2249–2258, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100988X.

Hubert:2014:SWS

[HBDC14] Malwe Boudoue Hubert, Gambo Betchewe, Serge Y. Doka, and Kofane Timoleon Crepin. Soliton wave solutions for the nonlinear transmission line using the Kudryashov method and the G′/G-expansion method. Applied Mathematics and Computation, 239(??):299–309, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006055.

Haq:2010:MML

[HBTU10] Sirajul Haq, Nagina Bibi, S. I. A. Tirmizi, and M. Usman. Meshless method of lines for the numerical solution of generalized Kuramoto– Sivashinsky equation. Applied Mathematics and Computation, 217(6):2404– 2413, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310007861.

He:2014:HFT

[HC14] Baiying He and Liangyun Chen. Hamiltonian forms of the two new integrable systems and two kinds of Darboux transformations. Applied Mathematics and Computation, 244(??):261–273, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009631.

Hou:2012:DIS

[HCC^{+}12] Yi-You Hou, Hsin-Chieh Chen, Jen-Fuh Chang, Jun-Juh Yan, and Teh-Lu Liao. Design and implementation of the Sprott chaotic secure digital communication systems. Applied Mathematics and Computation, 218(24):11799–11805, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200536X.

Hsiao:2014:ITP

[HCCT14] Chao-Yin Hsiao, Wen-Jeng Chang, Ming-I Char, and Bo-Chen Tai. Influence of thermophoretic particle deposition on MHD free convection flow of non-Newtonian fluids from a vertical plate embedded in porous media considering Soret and Dufour effects. Applied Mathematics and Computation, 244(??):390–397, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009643.

Hu:2014:NKM

[HCD14] Yangjie Hu, Hanlin Chen, and Zhengde Dai. New kink multi-soliton solutions for the (3 + 1)-dimensional potential-Yu–Toda–Sasa–Fukuyama equation. Applied Mathematics and Computation, 234(??):548–556, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002835.

He:2013:NLD

[HCL^{+}13] Xiao-Ting He, Liang Cao, Zheng-Ying Li, Xing-Jian Hu, and Jun-Yi Sun. Nonlinear large deflection problems of beams with gradient: a biparametric perturbation method. Applied Mathematics and Computation, 219(14):7493–7513, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000763.

Hermoso-Carazo:2011:NEA

[HCLP11] A. Hermoso-Carazo and J. Linares-Pérez. Nonlinear estimation applying an unscented transformation in systems with correlated uncertain observations. Applied Mathematics and Computation, 217(20):7998–8009, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003286.

He:2011:NFF

[HCW11] Suxiang He, Weilai Chen, and Hui Wang. A new filled function algorithm for constrained global optimization problems. Applied Mathematics and Computation, 217(12):5853–5859, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012725.

Huang:2014:NPF

[HD14a] Yuanyuan Huang and Yunda Dong. New properties of forward-backward splitting and a practical proximal-descent algorithm. Applied Mathematics and Computation, 237(??):60–68, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400424X.

Hwang:2014:CTI

[HD14b] Dah-Yan Hwang and Silvestru Sever Dragomir. Comparing two integral means for absolutely continuous functions whose absolute value of the derivative are convex and applications. Applied Mathematics and Computation, 230(??):259–266, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013416.

Hansen:2014:SPD

[HdCOL14] Matheus Hansen, Diogo R. da Costa, Diego F. M. Oliveira, and Edson D. Leonel. Statistical properties for a dissipative model of relativistic particles in a wave packet: a parameter space investigation. Applied Mathematics and Computation, 238(??):387–392, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300314004858.

Hachem:2010:EFE

[HDK^{+}10] E. Hachem, H. Digonnet, N. Kosseifi, E. Massoni, and T. Coupez. Enriched finite element spaces for transient conduction heat transfer. Applied Mathematics and Computation, 217(8):3929–3943, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010179.

Hladik:2011:FMI

[HDT11] Milan Hladík, David Daney, and Elias Tsigaridas. A filtering method for the interval eigenvalue problem. Applied Mathematics and Computation, 217(12):5236–5242, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010702.

Hong:2011:HEA

[HDZW11] Wei-Chiang Hong, Yucheng Dong, Feifeng Zheng, and Shih Yung Wei. Hybrid evolutionary algorithms in a SVR traffic flow forecasting model. Applied Mathematics and Computation, 217(15):6733–6747, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001081.

Hendtlass:2011:CCW

[Hen11] Matthew Hendtlass. The computational content of Walras’ existence theorem. Applied Mathematics and Computation, 217(13):6185–6191, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310013196.

Herceg:2011:FOF

[Her11] Djordje Herceg. Fourthorder finite-difference method for boundary value problems with two small parameters. Applied Mathematics and Computation, 218(2):616–627, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100796X.

Hassan:2012:NTU

[HET12] Hany N. Hassan and Magdy A. El-Tawil. A new technique of using homotopy analysis method for second order nonlinear differential equations. Applied Mathematics and Computation, 219(2):708–728, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200673X.

Hasan:2012:FCF

[HFS12] Agus Hasan, Bjarne Foss, and Svein Sagatun. Flow control of fluids through porous media. Applied Mathematics and Computation, 219(7):3323–3335, December 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311009180.

Hu:2013:NAB

[HG13a] Lin Hu and Siqing Gan.

Numerical analysis of the balanced implicit methods for stochastic pantograph equations with jumps. Applied Mathematics and Computation, 223(??):281–297, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008709.

Huseynov:2013:SFC

[HG13b] Aydin Huseynov and Gusein Sh. Guseinov. Solution of the finite complex Toda lattice by the method of inverse spectral problem. Applied Mathematics and Computation, 219(10):5550–5563, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012398.

Hosamani:2014:ZIT

[HG14] Sunilkumar M. Hosamani and Ivan Gutman. Zagreb indices of transformation graphs and total transformation graphs. Applied Mathematics and Computation, 247(??):1156–1160, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013058.

Hayek:2011:MWT

[HGN11] N. Hayek, B. J. González, and E. R. Negrin. Matrix Wiener transform. Applied Mathematics and Computation, 218(3):773–776, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001196.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Hidri:2011:BSH

[HH11] Lotfi Hidri and Mohamed Haouari. Bounding strategies for the hybrid flow shop scheduling problem. Applied Mathematics and Computation, 217(21):8248–8263, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003328.

Huang:2012:AIP

[HH12] Jie Huang and Ting-Zhu Huang. An approximate inverse preconditioner for Toeplitz systems with multiple righthand sides. Applied Mathematics and Computation, 218(23):11370–11379, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312005024.

Herceg:2013:MBM

[HH13a] Dragoslav Herceg and Djordje Herceg. Means based modifications of Newton’s method for solving nonlinear equations. Applied Mathematics and Computation, 219(11):6126–6133, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012830.

Huseyin:2013:CST

[HH13b] Nesir Huseyin and Anar Huseyin. Compactness of the set of trajectories of the controllable system described by an affine integral equation. Applied Mathematics and Computation, 219(16):8416–8424, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002580.

Harris:2014:SWG

[HH14] Anna Harris and Stephen Harris. Superconvergence of weak Galerkin finite element approximation for second order elliptic problems by L ^{2}-projections. Applied Mathematics and Computation, 227(??):610–621, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012344.

Hussain:2014:SPS

[HHA14] Malik Zawwar Hussain, Maria Hussain, and Beenish Aqeel. Shape-preserving surfaces with constraints on tension parameters. Applied Mathematics and Computation, 247(??): 442–464, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314012259.

Huy:2013:BHP

[HHDM13] Tuan Nguyen Huy, Quan Pham Hoang, Trong Dang Duc, and Triet Le Minh. On a backward heat problem with time-dependent coefficient: Regularization and error estimates. Applied Mathematics and Computation, 219(11):6066–6073, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312012362.

Hao:2011:TDS

[HHHS11] Wenrui Hao, Jonathan D. Hauenstein, Bei Hu, and Andrew J. Sommese. A three-dimensional steadystate tumor system. Applied Mathematics and Computation, 218(6):2661– 2669, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010320.

Han:2010:SAC

[HHM10] Xi-An Han, XiLi Huang, and YiChen Ma. Shape analysis of cubic trigonometric Bézier curves with a shape parameter. Applied Mathematics and Computation, 217(6):2527– 2533, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008106.

Heydari:2014:LWM

[HHM14] M. H. Heydari, M. R. Hooshmandasl, and F. Mohammadi. Legendre wavelets method for solving fractional partial differential equations with Dirichlet boundary conditions. Applied Mathematics and Computation, 234(??):267–276, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002860.

Hu:2013:BPH

[HHR13] Zhaoping Hu, Maoan Han, and Valery G. Romanovski. Bifurcations of planar Hamiltonian systems with impulsive perturbation. Applied Mathematics and Computation, 219(12):6733–6742, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000416.

Hwang:2013:WOI

[HHT13] Shiow-Ru Hwang, Kai-Chen Hsu, and Kuei-Lin Tseng. Weighted Ostrowski integral inequalities for mappings whose derivatives belong to L_{p}(a, b). Applied Mathematics and Computation, 219(17):9516–9523, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031300355X.

Haq:2012:NSN

[HHU12] Sirajul Haq, Arshad Hussain, and Marjan Uddin. On the numerical solution of nonlinear Burgers’-type equations using meshless method of lines. Applied Mathematics and Computation, 218(11):6280–6290, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014524.

Hu:2011:AIL

[HHX11] Jinsong Hu, Bing Hu, and Youcai Xu. Average implicit linear difference scheme for generalized Rosenau–Burgers equation. Applied Mathematics and Computation, 217(19):7557–7563, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001743.

Hu:2012:GAB

[HHX12] Lin-Xia Hu, Wan-Sheng He, and Hong-Ming Xia. Global asymptotic behavior of a rational difference equation. Applied Mathematics and Computation, 218(15):7818–7828, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000999.

Hayat:2014:CTI

[HI14] Sakander Hayat and Muhammad Imran. Computation of topological indices of certain networks. Applied Mathematics and Computation, 240(??):213–228, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006316.

Hilscher:2012:ETT

[Hil12] Roman Simon Hilscher.

Eigenvalue theory for time scale symplectic systems depending nonlinearly on spectral parameter. Applied Mathematics and Computation, 219(6):2839–2860, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008193.

He:2012:MPS

[HjH12] Guang He and Nan jing Huang. A modified particle swarm optimization algorithm with applications. Applied Mathematics and Computation, 219(3):1053–1060, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007102.

He:2014:NPS

[HjH14] Guang He and Nan jing Huang. A new particle swarm optimization algorithm with an application. Applied Mathematics and Computation, 232(??):521–528, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000654.

Hamedani:2014:RCM

[HJMY14] G. G. Hamedani, Z. Javanshiri, M. Maadooliat, and A. Yazdani. Remarks on characterizations of Malinowska and Szynal. Applied Mathematics and Computation, 246(??):377–388, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011242.

Hachem:2014:SFE

[HJVC14] E. Hachem, G. Jannoun, J. Veysset, and T. Coupez. On the stabilized finite element method for steady convection-dominated problems with anisotropic mesh adaptation. Applied Mathematics and Computation, 232(??):581–594, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000046.

Huang:2011:ABC

[HJW11] Wei-Xian Huang, Cong-Jian Jin, and Guo-Jin Wang. Approximate Bézier curves by cubic LN curves. Applied Mathematics and Computation, 218(7):3083–3092, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010587.

Hu:2013:HSS

[HJZC13] Lu Hu, Yangsheng Jiang, Juanxiu Zhu, and Yanru Chen. Hybrid of the scatter search, improved adaptive genetic, and expectation maximization algorithms for phase-type distribution fitting. Applied Mathematics and Computation, 219(10):5495–5515, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011721.

Hamdan:2011:IFA

[HK11] M. H. Hamdan and M. T. Kamel. On the Ni(x) integral function and its application to the Airy’s non-homogeneous equation. Applied Mathematics and Computation, 217(17):7349–7360, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002013.

Huang:2010:KWM

[HKH10] Zhengda Huang, Xiangyin Kong, and Weidong Hu. The King–Werner method for solving nonsymmetric algebraic Riccati equation. Applied Mathematics and Computation, 216(6):1790–1804, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010728.

Hashmi:2012:OHA

[HKI12] M. S. Hashmi, N. Khan, and S. Iqbal. Optimal homotopy asymptotic method for solving nonlinear Fredholm integral equations of second kind. Applied Mathematics and Computation, 218(22):10982–10989, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004584.

Hussain:2012:OLR

[HKK12] N. Hussain, M. A. Khamsi, and W. A. Kirk. Onelocal retracts and Banach operator pairs in metric spaces. Applied Mathematics and Computation, 218(20):10072– 10081, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003244.

He:2010:OBF

[HL10] Shuping He and Fei Liu. Observer-based finite-time control of time-delayed jump systems. Applied Mathematics and Computation, 217(6):2327–2338, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007769.

See comment [ZLMG12].

Han:2011:SBA

[HL11a] Wei Han and Maoxing Liu. Stability and bifurcation analysis for a discrete-time model of Lotka–Volterra type with delay. Applied Mathematics and Computation, 217(12):5449–5457, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012142.

Hao:2011:OTS

[HL11b] Qing-Hai Hao and Fang Lu. Oscillation theorem for superlinear second order damped differential equations. Applied Mathematics and Computation, 217(17):7126–7131, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001020.

Hsu:2011:EPD

[HL11c] William W. Y. Hsu and Yuh-Dauh Lyuu. Efficient pricing of discrete Asian options. Applied Mathematics and Computation, 217(24):9875–9894, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000294.

Huang:2011:NSB

[HL11d] J. S. Huang and G. D.

Lin. A note on the Sarmanov bivariate distributions. Applied Mathematics and Computation, 218(3):919–923, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001226.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Huang:2011:DPP

[HL11e] Meihua Huang and Xuepeng Li. Dispersal permanence of a periodic predatorprey system with Holling type-IV functional response. Applied Mathematics and Computation, 218(2):502–513, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007739.

Henderson:2012:EMP

[HL12a] Johnny Henderson and Rodica Luca. Existence and multiplicity for positive solutions of a multi-point boundary value problem. Applied Mathematics and Computation, 218(21):10572– 10585, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004031.

Henderson:2012:EPS

[HL12b] Johnny Henderson and Rodica Luca. Existence of positive solutions for a system of higher-order multipoint boundary value problems. Applied Mathematics and Computation, 219(8):3709–3720, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200985X.

Henderson:2012:PSS

[HL12c] Johnny Henderson and Rodica Luca. Positive solutions for a system of second-order multi-point boundary value problems. Applied Mathematics and Computation, 218(10):6083–6094, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101438X.

Hong:2012:NES

[HL12d] Baojian Hong and Dianchen Lu. New exact solutions for the generalized variable-coefficient Gardner equation with forcing term. Applied Mathematics and Computation, 219(5):2732–2738, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008971.

Horng:2012:MSO

[HL12e] Shih-Cheng Horng and Shin-Yeu Lin. Multistage ordinal optimization based approach for job shop scheduling problems. Applied Mathematics and Computation, 219(3):1125–1134, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007205.

He:2013:RFT

[HL13a] Shuping He and Fei Liu. Robust finite-time estimation of Markovian jumping systems with bounded transition probabilities. Applied Mathematics and Computation, 222(??):297–306, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007820.

Henderson:2013:PSS

[HL13b] Johnny Henderson and Rodica Luca. Positive solutions for a system of second-order nonlinear multi-point eigenvalue problems. Applied Mathematics and Computation, 223(??):197–208, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008400.

Hong:2013:MBC

[HL13c] Won-Tak Hong and Phill-Seung Lee. Mesh based construction of flat-top partition of unity functions. Applied Mathematics and Computation, 219(16):8687–8704, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001938.

Huang:2013:PCG

[HL13d] Yu-Mei Huang and De-Yong Lu. A preconditioned conjugate gradient method for multiplicative half-quadratic image restoration. Applied Mathematics and Computation, 219(12):6556–6564, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000428.

Henderson:2014:PSS

[HL14a] Johnny Henderson and Rodica Luca. Positive solutions for a system of fractional differential equations with coupled integral boundary conditions. Applied Mathematics and Computation, 249(??):182–197, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013940.

Hua:2014:FEM

[HL14b] Dong-Ying Hua and Xiang-Gui Li. The finite element method for computing the ground states of the dipolar Bose–Einstein condensates. Applied Mathematics and Computation, 234(??):214–222, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001222.

Huang:2014:PMS

[HL14c] Ching-Jen Huang and Li-Man Liao. Parallel machines scheduling with machine preference via agentbased approach. Applied Mathematics and Computation, 233(??):298–309, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314001520.

Huang:2014:MWY

[HL14d] Hai Huang and Suihua Lin. A modified Wei–Yao– Liu conjugate gradient method for unconstrained optimization. Applied Mathematics and Computation, 231(??):179–186, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000496.

Hladik:2013:BER

[Hla13] Milan Hladík. Bounds on eigenvalues of real and complex interval matrices. Applied Mathematics and Computation, 219(10):5584–5591, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012428.

Hu:2013:SAG

[HLC13] Jianqiang Hu, Jinling Liang, and Jinde Cao. Stability analysis for genetic regulatory networks with delays: the continuoustime case and the discretetime case. Applied Mathematics and Computation, 220(??):507–517, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006103.

Hamaizia:2012:FCO

[HLeH12] Tayeb Hamaizia, René Lozi, and Nasr eddine Hamri. Fast chaotic optimization algorithm based on locally averaged strategy and multifold chaotic attractor. Applied Mathematics and Computation, 219(1):188–196, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005802.

Hou:2012:MWP

[HLH12] Huoquan Hou, Bolian Liu, and Yufei Huang. The maximum Wiener polarity index of unicyclic graphs. Applied Mathematics and Computation, 218(20):10149– 10157, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003530.

Hu:2014:OPP

[HLL14] Fei Hu, Cheng-Chew Lim, and Zudi Lu. Optimal production and procurement decisions in a supply chain with an option contract and partial backordering under uncertainties. Applied Mathematics and Computation, 232(??):1225–1234, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001957.

Harko:2014:EAS

[HLM14] Tiberiu Harko, Francisco S. N. Lobo, and M. K. Mak. Exact analytical solutions of the Susceptible-Infected-Recovered (SIR) epidemic model and of the SIR model with equal death and birth rates. Applied Mathematics and Computation, 236(??):184–194, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400383X.

Hastermann:2013:DPE

[HLMW13] G. Hastermann, P. M.

Lima, M. L. Morgado, and E. B. Weinmüller.

Density profile equation with p-Laplacian: Analysis and numerical simulation. Applied Mathematics and Computation, 225(??):550–561, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301045X.

Hetmaniok:2014:IIP

[HLPW14] Edyta Hetmaniok, Piotr Lorenc, Mariusz Pleszczyński, and Roman Witula. Iterated integrals of polynomials. Applied Mathematics and Computation, 249(??):389–398, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014325.

Hsu:2010:SMS

[HLS10] Chou-Jung Hsu, Chinyao Low, and Chwen-Tzeng Su. A single-machine scheduling problem with maintenance activities to minimize makespan. Applied Mathematics and Computation, 215(11): 3929–3935, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300309010297.

Huang:2014:RCO

[HLS^{+}14] Xiaolong Huang, Xueliang Li, Yongtang Shi, Jun Yue, and Yan Zhao. Rainbow connections for outerplanar graphs with diameter 2 and 3. Applied Mathematics and Computation, 242(??):277–280, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007310.

Han:2010:RPA

[HLSS10] Zhenlai Han, Tongxing Li, Shurong Sun, and Yibing Sun. Remarks on the paper [Appl. Math. Comput. 207 (2009) 388– 396]. Applied Mathematics and Computation, 215(11):3998–4007, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010406.

See [?].

Hernandez:2013:POD

[HLT13] A. Hernández, M. Lattanzi, and N. Thome. On a partial order defined by the weighted Moore– Penrose inverse. Applied Mathematics and Computation, 219(14):7310–7318, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001264.

Hsu:2014:LDG

[HLT14] Hung-Chu Hsu, Meng-Syue Li, and Li-Hung Tsai. Lagrangian description of gravity-capillary waves propagating on a sloping bottom. Applied Mathematics and Computation, 233(??):494–504, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002173.

Hien:2014:EGA

[HLTT14] L. V. Hien, T. T. Loan, B. T. Huyen Trang, and H. Trinh. Existence and global asymptotic stability of positive periodic solution of delayed Cohen–Grossberg neural networks. Applied Mathematics and Computation, 240(??):200–212, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006183.

Hernandez:2012:SPO

[HLTU12] A. Hernández, M. Lattanzi, N. Thome, and F. Urquiza. The star partial order and the eigenprojection at 0 on EP matrices. Applied Mathematics and Computation, 218(21):10669– 10678, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031200433X.

Hu:2012:ABE

[HLW12] Guixin Hu, Meng Liu, and Ke Wang. The asymptotic behaviours of an epidemic model with two correlated stochastic perturbations. Applied Mathematics and Computation, 218(21):10520– 10532, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003980.

Huang:2013:SPA

[HLW13] Tsung-Ming Huang, Wen-Wei Lin, and Chin-Tien Wu. Structure-preserving arnoldi-type algorithm for solving eigenvalue problems in leaky surface wave propagation. Applied Mathematics and Computation, 219(19):9947–9958, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003810.

He:2012:OBS

[HLWH12] Yu-Lin He, James N. K. Liu, Xi-Zhao Wang, and Yan-Xing Hu. Optimal bandwidth selection for resubstitution entropy estimation. Applied Mathematics and Computation, 219(8):3425–3460, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008491.

Hu:2013:MIS

[HLX13] Chaozhu Hu, Bin Liu, and Songfa Xie. Monotone iterative solutions for nonlinear boundary value problems of fractional differential equation with deviating arguments. Applied Mathematics and Computation, 222(??):72–81, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007984.

Han:2012:IBV

[HLZ12] Jiangfeng Han, Yiliang Liu, and Jing Zhao. Integral boundary value problems for first order nonlinear impulsive functional integro-differential differential equations. Applied Mathematics and Computation, 218(9):5002–5009, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013130.

Hu:2013:NAP

[HLZ13] Hengchun Hu, Lei Liu, and Ling Zhang. New analytical positon, negaton and complexiton solutions of a coupled KdV– mKdV system. Applied Mathematics and Computation, 219(11):5743–5749, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013306.

Hu:2014:MIG

[HLZ14] Yuanbiao Hu, Baolin Liu, and Qin Zhou. A multiinnovation generalized extended stochastic gradient algorithm for output nonlinear autoregressive moving average systems. Applied Mathematics and Computation, 247(??):218–224, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012065.

Huang:2008:OSO

[HM08] Yan Huang and Fanwei Meng. Oscillation of second-order nonlinear ODE with damping. Applied Mathematics and Computation, 199(2):644– 652, June 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300307010521.

See comments [SQ11].

He:2010:NEE

[HM10a] Bin He and Qing Meng. New exact explicit peakon and smooth periodic wave solutions of the K(3, 2) equation. Applied Mathematics and Computation, 217(4):1697–1703, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009072.

He:2010:SSA

[HM10b] Chan He and Changfeng Ma. A smoothing self-adaptive Levenberg– Marquardt algorithm for solving system of nonlinear inequalities. Applied Mathematics and Computation, 216(10):3056–3063, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004455.

Hoseini:2010:ESW

[HM10c] S. M. Hoseini and T. R.

Marchant. Evolution of solitary waves for a perturbed nonlinear Schrödinger equation. Applied Mathematics and Computation, 216(12): 3642–3651, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310005345.

Hou:2010:HSS

[HM10d] Qing-Hu Hou and Yan-Ping Mu. Hypergeometric series solutions of linear operator equations. Applied Mathematics and Computation, 215(10):3731–3739, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010029.

Hoseini:2011:ATV

[HM11] S. M. Hoseini and H. R. Marzban. Analysis of time-varying delay systems via triangular functions. Applied Mathematics and Computation, 217(18):7432–7441, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002256.

Huang:2012:NSR

[HM12] Na Huang and Changfeng Ma. The numerical study of a regularized smoothing Newton method for solving P_{0}-NCP based on the generalized smoothing Fischer–Burmeister function. Applied Mathematics and Computation, 218(13):7253–7269, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000215.

Huang:2014:EFS

[HM14a] Na Huang and Changfeng Ma. Exceptional family and solvability of the second-order cone complementarity problems. Applied Mathematics and Computation, 244(??):561–566, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009850.

Huang:2014:RSN

[HM14b] Na Huang and Changfeng Ma. A regularized smoothing Newton method for solving SOCCPs based on a new smoothing Cfunction. Applied Mathematics and Computation, 230(??):315–329, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014069.

Hayat:2010:SFP

[HMA10] T. Hayat, R. J. Moitsheki, and S. Abelman. Stokes’ first problem for Sisko fluid over a porous wall. Applied Mathematics and Computation, 217(2):622–628, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006727.

Heshmati:2014:MCH

[HMD14] A. Heshmati, H. A. Mohammed, and A. N. Darus. Mixed convection heat transfer of nanofluids over backward facing step having a slotted baffle. Applied Mathematics and Computation, 240(??):368–386, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005980.

Hueso:2018:NO

[HMGC18] José L. Hueso, Eulalia Martínez, D. K. Gupta, and Fabricio Cevallos. A note on “Convergence radius of Osada’s method under Hölder continuous condition”. Applied Mathematics and Computation, 321(??):689–699, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300317307762.

See [ZS14d].

Huang:2013:DIC

[HMH13] Zhijian Huang, Jie Ma, and He Huang. The derivation of iterative convergence calculation for a nonlinear MIMO approximate dynamic programming approach. Applied Mathematics and Computation, 219(9):4495–4502, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010703.

Hassanabadi:2013:AAS

[HMIZ13] H. Hassanabadi, E. Maghsoodi, Akpan N. Ikot, and S. Zarrinkamar. Approximate arbitrary-state solutions of Dirac equation for modified deformed Hylleraas and modified Eckart potentials by the NU method. Applied Mathematics and Computation, 219(17):9388–9398, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002646.

He:2010:EPS

[HML10] Bin He, Qing Meng, and Shaolin Li. Explicit peakon and solitary wave solutions for the modified Fornberg–Whitham equation. Applied Mathematics and Computation, 217(5):1976–1982, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000740X.

Hematiyan:2011:BEA

[HMMK11] M. R. Hematiyan, M. Mohammadi, L. Marin, and A. Khosravifard. Boundary element analysis of uncoupled transient thermoelastic problems with time- and space-dependent heat sources. Applied Mathematics and Computation, 218(5):1862–1882, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311009155.

Huertas:2014:ALS

[HMPVQ14] Edmundo J. Huertas, Francisco Marcellán, M. Francisca Pérez-Valero, and Yamilet Quintana. Asymptotics for Laguerre–Sobolev type orthogonal polynomials modified within their oscillatory regime. Applied Mathematics and Computation, 236(??):260–272, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004007.

Huertas:2012:ZOP

[HMR12] Edmundo J. Huertas, Francisco Marcellán, and Fernando R. Rafaeli. Zeros of orthogonal polynomials generated by canonical perturbations of measures. Applied Mathematics and Computation, 218(13):7109–7127, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015621.

Herawati:2014:OMT

[HMSW14] E. Herawati, M. Mursaleen, Supama, and I. E. Wijayanti. Order matrix transformations on some Banach lattice valued sequence spaces. Applied Mathematics and Computation, 247(??):1122–1128, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012375.

Hong:2014:SNP

[HMZ14] Xiaoli Hong, Lingsheng Meng, and Bing Zheng. Some new perturbation bounds of generalized polar decomposition. Applied Mathematics and Computation, 233(??):430–438, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002525.

Huang:2010:PTB

[HN10a] Yu-Mei Huang and Michael K. Ng. Product-type block preconditioners for electromagnetic cavity problems. Applied Mathematics and Computation, 216(6):1741–1751, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010777.

Huy:2010:ITI

[HN10b] Vu Nhat Huy and Quô´ c-Anh Ngô. On an Iyengartype inequality involving quadratures in n knots. Applied Mathematics and Computation, 217(1):289–294, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006193.

Hajmohammadi:2014:SCV

[HN14a] M. R. Hajmohammadi and S. S. Nourazar. On the solution of characteristic value problems arising in linear stability analysis; semi analytical approach. Applied Mathematics and Computation, 239(??):126–132, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006006.

See erratum [NN19].

Huy:2014:NOG

[HN14b] Vu Nhat Huy and Qu´ ôc-Anh Ngô. A new Ostrowski-Grüss inequality involving 3n knots. Applied Mathematics and Computation, 235(??): 272–282, May 25, 2014.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003385.

Hernandez:2012:NCA

[HOB12] Eduardo Hernández, Donal O’Regan, and Maria Aparecida Bená. On a new class of abstract integral equations and applications. Applied Mathematics and Computation, 219(4):2271–2277, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008673.

Hernandez:2013:ATA

[HOMA13] C. Hernandez, R. B. B. Ovando-Martinez, and M. A. Arjona. Application of tensor analysis to the finite element method. Applied Mathematics and Computation, 219(9):4625–4636, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010909.

Hong:2010:NEJ

[Hon10a] Baojian Hong. New exact Jacobi elliptic functions solutions for the generalized coupled Hirota– Satsuma KdV system. Applied Mathematics and Computation, 217(2):472– 479, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310006387.

Hong:2010:SHT

[Hon10b]  Dug Hun Hong.  A

sharp Hardy-type inequality of Sugeno integrals. Applied Mathematics and Computation, 217(1):437–440, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006302.

Horng:2010:MIT

[Hor10] Ming-Huwi Horng. A multilevel image thresholding using the honey bee mating optimization. Applied Mathematics and Computation, 215(9):3302–3310, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009230.

Hou:2013:CRP

[Hou13]  L. S. Hou.  On the

O(1/t) convergence rate of the parallel descentlike method and parallel splitting augmented Lagrangian method for solving a class of variational inequalities. Applied Mathematics and Computation, 219(11):5862–5869, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012866.

Hou:2014:CTD

[Hou14] Guoliang Hou. Comparison theorems for double splittings of K-monotone matrices. Applied Mathematics and Computation, 244(??):382–389, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400945X.

Hussain:2012:CFP

[HP12] N. Hussain and H. K. Pathak. Common fixed point and approximation results for H-operator pair with applications. Applied Mathematics and Computation, 218(22):11217– 11225, July 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004985.

Hernandez:2011:AAP

[HPB11] Eduardo Hernández, Michelle Pierri, and Maria Aparecida Bená. Asymptotically almost periodic solutions for abstract neutral integro-differential equations. Applied Mathematics and Computation, 217(22):8963–8972, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100470X.

Pan:2012:SAR

[hPnZrL12] Chao hong Pan, Xue na Zhu, and Zheng rong Liu. A simple approach for reducing the order of equations with higher order nonlinearity. Applied Mathematics and Computation, 218(17):8702–8714, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001567.

Pan:2010:SMA

[hPrLgZ10] Chao hong Pan, Zheng rong Liu, and Ben gong Zhang. A simple method and its applications to nonlinear partial differential equations. Applied Mathematics and Computation, 217(8):4197–4204, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010775.

Hu:2014:CSO

[HQ14] Shenglong Hu and Liqun Qi. Convergence of a second order Markov chain. Applied Mathematics and Computation, 241(??):183–192, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006675.

Hasanbulli:2010:OCS

[HR10] Mustafa Hasanbulli and Yuri V. Rogovchenko. Oscillation criteria for second order nonlinear neutral differential equations. Applied Mathematics and Computation, 215(12):4392–4399, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000010X.

Hu:2011:NRB

[HR11] Lanying Hu and Yong Ren. A note on the reflected backward stochastic differential equations driven by a Lévy process with stochastic Lipschitz condition. Applied Mathematics and Computation, 218(8):4325–4332, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012379.

Hinds:2012:DPW

[HR12] Brittney Hinds and Petronela Radu. Dirichlet’s principle and wellposedness of solutions for a nonlocal p-Laplacian system. Applied Mathematics and Computation, 219(4):1411–1419, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007618.

Holder:2013:IBM

[HR13] Andrew B. Holder and Marianito R. Rodrigo. An integration-based method for estimating parameters in a system of differential equations. Applied Mathematics and Computation, 219(18):9700–9708, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003056.

Hosen:2012:ATS

[HRAA12] Md. Alal Hosen, M. Saifur Rahman, M. Shamsul Alam, and M. Ruhul Amin. An analytical technique for solving a class of strongly nonlinear conservative systems. Applied Mathematics and Computation, 218(9):5474–5486, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031101366X.

Holder:2014:MAM

[HRH14] Andrew B. Holder, Marianito R. Rodrigo, and Miguel A. Herrero. A model for acid-mediated tumour growth with nonlinear acid production term. Applied Mathematics and Computation, 227(??):176–198, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011855.

He:2014:NED

[HRL14] Bin He, Weiguo Rui, and Yao Long. New exact double periodic wave and complex wave solutions for a generalized sinh–Gordon equation. Applied Mathematics and Computation, 229(??):159–172, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013167.

Hu:2014:MSS

[HRX14] Lanying Hu, Yong Ren, and Tianbao Xu. p-Moment stability of solutions to stochastic differential equations driven by G-Brownian motion. Applied Mathematics and Computation, 230(??): 231–237, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031301401X.

Hauenstein:2010:WSP

[HS10a] Jonathan D. Hauenstein and Andrew J. Sommese. Witness sets of projections. Applied Mathematics and Computation, 217(7):3349–3354, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009318.

Hussein:2010:DSS

[HS10b] A. Hussein and M. M. Selim. A developed solution of the stochastic Milne problem using probabilistic transformations. Applied Mathematics and Computation, 216(10):2910–2919, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004091.

Herty:2011:TDN

[HS11a] Michael Herty and Veronika Schleper. Time discretizations for numerical optimisation of hyperbolic problems. Applied Mathematics and Computation, 218(1):183–194, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007995.

Hou:2011:BNS

[HS11b] Xiaorong Hou and Junwei Shao. Bounds on the number of steps of WDS required for checking the positivity of integral forms. Applied Mathematics and Computation, 217(24):9978–9984, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006370.

Hussein:2012:SSR

[HS12] A. Hussein and M. M.

Selim. Solution of the stochastic radiative transfer equation with Rayleigh scattering using RVT technique. Applied Mathematics and Computation, 218(13):7193–7203, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015785.

Hauenstein:2013:MTI

[HS13] Jonathan D. Hauenstein and Andrew J. Sommese. Membership tests for images of algebraic sets by linear projections. Applied Mathematics and Computation, 219(12):6809–6818, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312013367.

Haq:2014:IEF

[HS14a] Abdul Haq and Javid Shabbir. An improved estimator of finite population mean when using two auxiliary attributes. Applied Mathematics and Computation, 241(??):14–24, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006092.

Haq:2014:IFE

[HS14b] Abdul Haq and Javid Shabbir. An improved family of estimators of finite population mean based on the auxiliary attribute. Applied Mathematics and Computation, 230(??):336–341, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014033.

Huang:2014:RDI

[HSC14] Junjie Huang, Yunfeng Shi, and Alatancang Chen. The representation of the Drazin inverse of antitriangular operator matrices based on resolvent expansions. Applied Mathematics and Computation, 242(??):196–201, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007188.

Hu:2011:IPS

[HSD11] Di Hu, Ali Sarosh, and Yun-Feng Dong. An improved particle swarm optimizer for parametric optimization of flexible satellite controller. Applied Mathematics and Computation, 217(21):8512–8521, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004231.

Hsiao:2010:NSL

[Hsi10]  Chun-Hui Hsiao.  Nu-

merical solutions of linear time-varying descriptor systems via hybrid functions. Applied Mathematics and Computation, 216(5):1363–1374, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002468.

Helal:2013:VPZ

[HSI13] M. A. Helal, A. R. Seadawy, and R. S. Ibrahim. Variational principle for Zakharov– Shabat equations in twodimensions. Applied Mathematics and Computation, 219(10):5635–5648, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010958.

Hsieh:2014:BGR

[Hsi14]  Tsung-Jung Hsieh.  A

bacterial gene recombination algorithm for solving constrained optimization problems. Applied Mathematics and Computation, 231(??):187–204, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000174.

Hernandez:2010:ESN

[HSLG10] Jairo Ernesto Castillo Hernández, Alvaro H. Salas, José Gonzalo Escobar Lugo, and Grupo CIBAVIR. Exact solutions for a nonlinear model. Applied Mathematics and Computation, 217(4):1646–1651, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309008091.

Hokayem:2010:CCA

[HSS10] Peter F. Hokayem, Dušan M.

Stipanović, and Mark W. Spong. Coordination and collision avoidance for Lagrangian systems with disturbances. Applied Mathematics and Computation, 217(3):1085–1094, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003280.

Hauenstein:2011:RCH

[HSW11] Jonathan D. Hauenstein, Andrew J. Sommese, and Charles W. Wampler. Regenerative cascade homotopies for solving polynomial systems. Applied Mathematics and Computation, 218(4):1240–1246, October 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008150.

Hetmaniok:2012:CEE

[HSW12] Edyta Hetmaniok, Damian Slota, and Roman Witula. Convergence and error estimation of homotopy perturbation method for Fredholm and Volterra integral equations. Applied Mathematics and Computation, 218(21):10717– 10725, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004407.

Hong:2010:DDE

[HSX10] Liang Hong, Taixiang Sun, and Hongjian Xi. On the dynamics of the difference equation x_{n+1} = \frac{1}{B_{n}x_{n}+x_{n−1}} . Applied Mathematics and Computation, 216(1):337–340, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000949.

Han:2011:PAS

[HSX^{+}11] Heqiang Han, Guanglei Song, Yongsong Xiao, Yuwu Liao, and Ruifeng Ding. Performance analysis of the AM-SG parameter estimation for multivariable systems. Applied Mathematics and Computation, 217(12):5566–5572, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310012324.

Helal:2014:SAS

[HSZ14] M. A. Helal, A. R.

Seadawy, and M. H. Zekry. Stability analysis of solitary wave solutions for the fourth-order nonlinear Boussinesq water wave equation. Applied Mathematics and Computation, 232(??):1094–1103, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001039.

Han:2010:DBC

[HT10] Xie Han and Qiulin Tan. Dynamical behavior of computer virus on Internet. Applied Mathematics and Computation, 217(6):2520–2526, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000809X.

Hakl:2011:MAP

[HT11a] Robert Hakl and Pedro J. Torres. Maximum and antimaximum principles for a second order differential operator with variable coefficients of indefinite sign. Applied Mathematics and Computation, 217(19):7599–7611, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100244X.

Haukkanen:2011:GDR

[HT11b] Pentti Haukkanen and Timo Tossavainen. A generalization of Descartes’ rule of signs and fundamental theorem of algebra. Applied Mathematics and Computation, 218(4):1203–1207, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007909.

Halkos:2014:AVS

[HT14] George E. Halkos and Kyriaki D. Tsilika. Analyzing and visualizing the synergistic impact mechanisms of climate change related costs. Applied Mathematics and Computation, 246(??):586–596, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011539.

Tian:2010:SDM

[hTlCqL10] Ying hui Tian, Han lin Chen, and Xi qiang Liu. A simple direct method to find equivalence transformations of a generalized nonlinear Schrödinger equation and a generalized KdV equation. Applied Mathematics and Computation, 215(10):3509–3514, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030900962X.

Hu:2012:SEN

[HTX^{+}12] Ming-Hui Hu, Shan-Tung Tu, Fu-Zhen Xuan, Chun-Ming Xia, and Hui-He Shao. Strain energy numerical technique for structural damage detection. Applied Mathematics and Computation, 219(5):2424–2431, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008715.

Hu:2010:DSM

[Hu10] Xiaolin Hu. Dynamic system methods for solving mixed linear matrix inequalities and linear vector inequalities and equalities. Applied Mathematics and Computation, 216(4):1181–1193, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001761.

Hu:2011:ERI

[Hu11] Xiao-Li Hu. An extension of Rosenthal’s inequality. Applied Mathematics and Computation, 218(8):4638–4640, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011520.

Hu:2012:CST

[Hu12a]  Jin-Yan Hu.  Classifi-

cation of single traveling wave solutions to the nonlinear dispersion Drinfel’d–Sokolov system. Applied Mathematics and Computation, 219(4):2017–2025, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008363.

Hu:2012:RBD

[Hu12b] Lanying Hu. Reflected backward doubly stochastic differential equations driven by a Lévy process with stochastic Lipschitz condition. Applied Mathematics and Computation, 219(3):1153–1157, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007230.

Hu:2013:PNP

[Hu13a] Hongxiao Hu. Permanence for nonautonomous predator-prey Kolmogorov systems with impulses and its applications. Applied Mathematics and Computation, 223(??):54–75, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008497.

Hu:2013:IAP

[Hu13b] Qianqian Hu. An iterative algorithm for polynomial approximation of rational triangular Bézier surfaces. Applied Mathematics and Computation, 219(17):9308–9316, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003068.

Hu:2013:EYI

[Hu13c]  Xiao-Li Hu.  An ex-

tension of Young’s inequality and its application. Applied Mathematics and Computation, 219(12):6393–6399, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S009630031300026X.

Hu:2014:CPA

[Hu14]  Qianqian Hu.  Con-

strained polynomial approximation of quadric surfaces. Applied Mathematics and Computation, 248(??):354–362, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013575.

Huang:2010:FPN

[Hua10] Guohong Huang. Fusion (2D)^{2} PCALDA: a new method for face recognition. Applied Mathematics and Computation, 216(11):3195–3199, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004686.

Huang:2011:DTI

[Hua11a] Qing-Hua Huang. Discovery of time-inconsecutive co-movement patterns of foreign currencies using an evolutionary biclustering method. Applied Mathematics and Computation, 218(8):4353–4364, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012410.

Huang:2011:ISC

[Hua11b] Shian-Chang Huang. Integrating spectral clustering with wavelet based kernel partial least square regressions for financial modeling and forecasting. Applied Mathematics and Computation, 217(15):6755–6764, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001354.

Huseynov:2010:PSN

[Hus10a] Aydin Huseynov. Positive solutions of a nonlinear impulsive equation with periodic boundary conditions. Applied Mathematics and Computation, 217(1):247–259, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006144.

Hussain:2010:SAM

[Hus10b]  Saqib Hussain.  Some

applications of Miller– Mocanu lemma on certain classes of meromorphic functions. Applied Mathematics and Computation, 216(10):3016–3021, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004339.

Hutnikova:2013:RST

[Hut13] Mária Hutníková. On the range of Stockwell transforms. Applied Mathematics and Computation, 219(17):8904–8909, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002816.

Hasil:2012:APT

[HV12a] Petr Hasil and Michal Veselý. Almost periodic transformable difference systems. Applied Mathematics and Computation, 218(9):5562–5579, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013932.

Haussermann:2012:ASS

[HV12b] John Haussermann and Robert A. Van Gorder. Asymptotic solutions for singularly perturbed Boussinesq equations. Applied Mathematics and Computation, 218(20):10238– 10243, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312003682.

Hearns:2012:PTI

[HVC12] Jessica Hearns, Robert A. Van Gorder, and S. Roy Choudhury. Painlevé test, integrability, and exact solutions for densitydependent reaction-diffusion equations with polynomial reaction functions. Applied Mathematics and Computation, 219(6):3055–3064, November 25, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009319.

Hadjidoukas:2014:SAI

[HVD^{+}14] P. E. Hadjidoukas, C. Voglis, V. V. Dimakopoulos, I. E. Lagaris, and D. G. Papageorgiou. Supporting adaptive and irregular parallelism for nonlinear numerical optimization. Applied Mathematics and Computation, 231(??):544–559, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013829.

Huang:2011:WBB

[HW11] Wei-Xian Huang and Guo-Jin Wang. A weighted bivariate blending rational interpolation based on function values. Applied Mathematics and Computation, 217(9):4644–4653, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011288.

He:2012:SOP

[HW12] Zhuo-Heng He and Qing-Wen Wang. Solutions to optimization problems on ranks and inertias of a matrix function with applications. Applied Mathematics and Computation, 219(6):2989–3001, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200923X.

Hauenstein:2013:NIA

[HW13a] Jonathan D. Hauenstein and Charles W. Wampler. Numerically intersecting algebraic varieties via witness sets. Applied Mathematics and Computation, 219(10):5730–5742, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006339.

Huang:2013:BPS

[HW13b] Shou-Jun Huang and Rui Wang. On blowup phenomena of solutions to the Euler equations for Chaplygin gases. Applied Mathematics and Computation, 219(9):4365–4370, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010272.

He:2014:DSC

[HW14a] Chengyuan He and Xiaoye Wang. The discriminance for a special class of circulant matrices and their diagonalization. Applied Mathematics and Computation, 238(??):7– 12, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400530X.

Hifi:2014:NUB

[HW14b] Mhand Hifi and Lei Wu.

New upper bounds and exact methods for the knapsack sharing problem. Applied Mathematics and Computation, 227(??):518–530, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012083.

Hwang:2011:SID

[Hwa11] Dah-Yan Hwang. Some inequalities for differentiable convex mapping with application to weighted trapezoidal formula and higher moments of random variables. Applied Mathematics and Computation, 217(23):9598–9605, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005935.

Hwang:2014:SID

[Hwa14] Dah-Yan Hwang. Some inequalities for differentiable convex mapping with application to weighted midpoint formula and higher moments of random variables. Applied Mathematics and Computation, 232(??):68–75, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000873.

Hao:2011:APB

[HWL11] Yong-Xia Hao, Ren-Hong Wang, and Chong-Jun Li. Analysis of a 6point binary subdivision scheme. Applied Mathematics and Computation, 218(7):3209–3216, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011015.

Hidayat:2014:MLS

[HWPY14] Mas Irfan P. Hidayat, Bambang A. Wahjoedi, Setyamartana Parman, and Puteri S. M. Megat Yusoff. Meshless local B-spline-FD method and its application for 2D heat conduction problems with spatially varying thermal conductivity. Applied Mathematics and Computation, 242(??):236–254, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006961.

Hu:2011:FES

[HWQG11] Jian Hu, Zhiqiang Wang, Shaojie Qiao, and Jian-Chao Gan. The fitness evaluation strategy in particle swarm optimization. Applied Mathematics and Computation, 217(21):8655–8670, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004760.

Huang:2013:RRC

[HX13] Shipei Huang and Zhengrong Xiang. Robust L_{∞} reliable control for uncertain switched nonlinear systems with time delay under asynchronous switching. Applied Mathematics and Computation, 222(??):658–670, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008345.

Hu:2014:GAS

[HX14] Lin-Xia Hu and Hong-Ming Xia. Global asymptotic stability of a second order rational difference equation. Applied Mathematics and Computation, 233(??):377–382, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400246X.

Hua:2014:IHH

[HXQ14] Jü Hua, Bo-Yan Xi, and Feng Qi. Inequalities of Hermite–Hadamard type involving an s-convex function with applications. Applied Mathematics and Computation, 246(??):752–760, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011515.

Xu:2010:NPS

[hXqXlC10] Zhen hui Xu, Da quan Xian, and Han lin Chen. New periodic solitarywave solutions for the Benjiamin Ono equation. Applied Mathematics and Computation, 215(12):4439–4442, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009989.

Hao:2012:GMK

[HXY12] Shu-Yan Hao, Shu-Sen Xie, and Su-Cheol Yi. The Galerkin method for the KdV equation using a new basis of smooth piecewise cubic polynomials. Applied Mathematics and Computation, 218(17):8659–8671, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200149X.

He:2012:HOF

[HXZ12] Tieshan He, Yuantong Xu, and Yanshan Zeng. Higher order functional differential systems with parameter dependence. Applied Mathematics and Computation, 218(15):7783–7791, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000811.

Hao:2011:CPR

[HY11a] Yuanyuan Hao and Hu Yang.

On a compound Poisson risk model with delayed claims and random incomes. Applied Mathematics and Computation, 217(24):10195–10204, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006710.

Hsieh:2011:EIB

[HY11b] Y.-C. Hsieh and P.-S. You. An effective immune based two-phase approach for the optimal reliability-redundancy allocation problem. Applied Mathematics and Computation, 218(4):1297–1307, October 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311008253.

Hong:2014:FRS

[HY14] Wenxi Hong and Lihua You. Further results on the spectral radius of matrices and graphs. Applied Mathematics and Computation, 239(??):326–332, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314006547.

Huang:2010:MIP

[HYC^{+}10] Guang-Xin Huang, Feng Yin, Hua-Fu Chen, Ling Chen, and Ke Guo. Matrix inverse problem and its optimal approximation problem for R-skew symmetric matrices. Applied Mathematics and Computation, 216(12):3515–3521, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005114.

He:2010:RHM

[HYD10] Songnian He, Caiping Yang, and Peichao Duan. Realization of the hybrid method for Mann iterations. Applied Mathematics and Computation, 217(8):4239–4247, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010945.

He:2014:SDA

[HYJ14] Hanlin He, Lu Yan, and Mei Jiang. Study on disturbance attenuation of cellular neural networks with time-varying delays. Applied Mathematics and Computation, 244(??):533–541, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009941.

Hu:2014:LSA

[HYSZ14] Juan Hu, Yujian Ye, Shoufeng Shen, and Jun Zhang. Lie symmetry analysis of the time fractional KdV-type equation. Applied Mathematics and Computation, 233(??):439–444, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002495.

Hattaf:2013:SAV

[HYT13] Khalid Hattaf, Noura Yousfi, and Abdessamad Tridane. Stability analysis of a virus dynamics model with general incidence rate and two delays. Applied Mathematics and Computation, 221(??):514–521, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313007534.

Hou:2014:TDG

[HYT14a] Peng-Fei Hou, Ke Yuan, and Wei Tian. Threedimensional Green’s functions for a fluid and pyroelectric two-phase material. Applied Mathematics and Computation, 249(??):303–319, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013964.

Hwang:2014:RSE

[HYT14b] Shiow-Ru Hwang, Shu-Ying Yeh, and Kuei-Lin Tseng. Refinements and similar extensions of Hermite-Hadamard inequality for fractional integrals and their applications. Applied Mathematics and Computation, 249(??):103–113, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013988.

Huang:2014:ATW

[HYW14] Yong Huang, Wenjun Yuan, and Yonghong Wu. All traveling wave exact solutions of two kinds of nonlinear evolution equations. Applied Mathematics and Computation, 235(??):148–156, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003191.

Hoang:2014:NSR

[HYX14] Tuan Anh Hoang, George Yin, and Fubao Xi. Numerical solutions of regime-switching jump diffusions. Applied Mathematics and Computation, 244(??):822–835, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010091.

Hua:2012:NSC

[HYYG12] Changchun Hua, Xian Yang, Jing Yan, and Xinping Guan. New stability criteria for neural networks with timevarying delays. Applied Mathematics and Computation, 218(9):5035–5042, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311013166.

Hong:2014:NDM

[HYZH14] Sheng Hong, Hongqi Yang, Enrico Zio, and Ning Huang. A novel dynamics model of fault propagation and equilibrium analysis in complex dynamical communication network. Applied Mathematics and Computation, 247 (??):1021–1029, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012478.

He:2010:IMS

[HZ10] Xiahui He and Yunshan Zhou. An iterative method for a system of generalized mixed quasi-variational inclusions with (Aη)-monotone mappings. Applied Mathematics and Computation, 215(12):4256–4262, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010960.

He:2011:NAL

[HZ11] Suxiang He and Shumin Zhou. A nonlinear augmented Lagrangian for constrained minimax problems. Applied Mathematics and Computation, 218(8):4567–4579, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012847.

He:2012:PDF

[HZ12a] Xu He and Jun Zhao. Parameter-dependent H_{∞} filter design for LPV systems and an autopilot application. Applied Mathematics and Computation, 218(9):5508–5517, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013701.

Hilscher:2012:OTR

[HZ12b] Roman Simon Hilscher and Vera Zeidan. Oscillation theorems and Rayleigh principle for linear Hamiltonian and symplectic systems with general boundary conditions. Applied Mathematics and Computation, 218(17):8309–8328, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001026.

Hu:2012:FDM

[HZ12c] Xiuling Hu and Luming Zhang. On finite difference methods for fourthorder fractional diffusionwave and subdiffusion systems. Applied Mathematics and Computation, 218(9):5019–5034, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013154.

Hao:2013:DDS

[HZ13] Wenrui Hao and Shaohong Zhu. Domain decomposition schemes with highorder accuracy and unconditional stability. Applied Mathematics and Computation, 219(11):6170–6181, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012593.

Hashemi:2014:ENR

[HZ14a] M. Y. Hashemi and Kamiar Zamzamian. Efficient and non-reflecting far-field boundary conditions for incompressible flow calculations. Applied Mathematics and Computation, 230(??):248–258, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013799.

Hassan:2014:OBC

[HZ14b] Mohamed Fahim Hassan and Mohamed Zribi. An observer-based controller for nonlinear systems: a gain scheduling approach. Applied Mathematics and Computation, 237(??):695–711, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005013.

Hilscher:2014:LPL

[HZ14c] Roman Simon Hilscher and Petr Zemánek. Limit point and limit circle classification for symplectic systems on time scales. Applied Mathematics and Computation, 233(??):623–646, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313014252.

Huang:2014:PSS

[HZ14d] Xiaoxia Huang and Tianyi Zhao. Project selection and scheduling with uncertain net income and investment cost. Applied Mathematics and Computation, 247(??):61–71, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314011928.

Hao:2014:HCE

[HZFZ14] Rong-Xia Hao, Ru Zhang, Yan-Quan Feng, and Jin-Xin Zhou. Hamiltonian cycle embedding for fault tolerance in balanced hypercubes. Applied Mathematics and Computation, 244(??):447–456, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009722.

Huang:2012:MMD

[HZW12] Yujiao Huang, Huaguang Zhang, and Zhanshan Wang. Multistability and multiperiodicity of delayed bidirectional associative memory neural networks with discontinuous activation functions. Applied Mathematics and Computation, 219(3):899–910, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006765.

Huang:2013:PCA

[HZW13] Qingqun Huang, Zhibin Zhu, and Xiangling Wang. A predictor-corrector algorithm combined conjugate gradient with homotopy interior point for general nonlinear programming. Applied Mathematics and Computation, 219(9):4379–4386, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010375.

Huang:2014:MCV

[HZW14] Yujiao Huang, Huaguang Zhang, and Zhanshan Wang. Multistability of complex-valued recurrent neural networks with realimaginary-type activation functions. Applied Mathematics and Computation, 229(??):187–200, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013039.

Hsu:2014:NGK

[HZZ14] Chingfang Hsu, Bing Zeng, and Maoyuan Zhang. A novel group key transfer for big data security. Applied Mathematics and Computation, 249(??):436–443, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401426X.

Ipek:2014:HPD

[IA14] Ahmet Ipek and Kamil Ari. On Hessenberg and pentadiagonal determinants related with Fibonacci and Fibonacci-like numbers. Applied Mathematics and Computation, 229(??):433–439, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013544.

Ismail:2010:SSGa

[IB10a] M. S. Ismail and Anjan Biswas. 1-soliton solution of the generalized KdV equation with generalized evolution. Applied Mathematics and Computation, 216(5):1673–1679, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002171.

Ismail:2010:SSK

[IB10b] M. S. Ismail and Anjan Biswas. 1-soliton solution of the Klein– Gordon–Zakharov equation with power law nonlinearity. Applied Mathematics and Computation, 217(8):4186–4196, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010763.

Itik:2010:CKB

[IB10c] Mehmet Itik and Stephen P. Banks. On the calculation of the Kauffman bracket polynomial. Applied Mathematics and Computation, 216(2):655–661, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310001268.

Ilic:2011:NRE

[IB11] Aleksandar Ilić and Milan Basić. New results on the energy of integral circulant graphs. Applied Mathematics and Computation, 218(7):3470–3482, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011350.

Ibragimov:2014:CPS

[Ibr14] Nail H. Ibragimov. Comments on the paper “Conservation laws of the (2 + 1)-dimensional KP equation and Burgers equation with variable coefficients and cross terms” by Li-Hua Zhang. Applied Mathematics and Computation, 226(??):1–2, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010904.

See [Zha13a].

Ikhdair:2011:SPS

[IBS11] Sameer M. Ikhdair, Cüneyt Berkdemir, and Ramazan Sever. Spin and pseudospin symmetry along with orbital dependency of the Dirac-Hulthén problem. Applied Mathematics and Computation, 217(22):9019–9032, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004772.

Ionescu:2011:NAC

[IC11] Adela Ionescu and Daniela Coman. New approaches in computational dynamics of the mixing flow. Applied Mathematics and Computation, 218(3):809–816, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005224.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ic:2014:TBD

[Iç14] Yusuf Tansel Iç. A TOP-SIS based design of experiment approach to assess company ranking. Applied Mathematics and Computation, 227(??):630–647, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012113.

Imdad:2014:FPT

[ICKV14] Mohammad Imdad, Sunny Chauhan, Zoran Kadelburg, and Calogero Vetro. Fixed point theorems for non-self mappings in symmetric spaces under ϕweak contractive conditions and an application to functional equations in dynamic programming. Applied Mathematics and Computation, 227(??):469–479, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011740.

Icoz:2012:KVN

[Içö12] Gürhan Içöz. A Kantorovich variant of a new type Bernstein–Stancu polynomials. Applied Mathematics and Computation, 218(17):8552–8560, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001385.

Idoumghar:2013:HIP

[ICS^{+}13] Lhassane Idoumghar, Nicolas Chérin, Patrick Siarry, Robin Roche, and Abdellatif Miraoui. Hybrid ICA–PSO algorithm for continuous optimization. Applied Mathematics and Computation, 219(24):11149–11170, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300547X.

Iyit:2011:CRI

[IG11] Neslihan Iyit and Asir Genc. Constitution of random intercept and slope model (RISM) as a special case of linear mixed models (LMMs) for repeated measurements data. Applied Mathematics and Computation, 218(3):827–831, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004292.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ignatyev:2010:HPL

[Ign10] Oleksiy A. Ignatyev. Homogeneous polynomials as Lyapunov functions in the stability research of solutions of difference equations. Applied Mathematics and Computation, 216(2):388–394, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000391.

Ignatyev:2013:NCP

[Ign13] Oleksiy Ignatyev. New criterion of partial asymptotic stability in probability of stochastic differential equations. Applied Mathematics and Computation, 219(23):10961– 10966, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313005158.

Ikhdair:2013:RSY

[IHF13] Sameer M. Ikhdair, Majid Hamzavi, and Babatunde J. Falaye. Relativistic symmetries in Yukawa-type interactions with Coulomb-like tensor. Applied Mathematics and Computation, 225(??):775–786, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313010849.

See comments [Woo20].

Imran:2014:TIC

[IHM14] Muhammad Imran, Sakander Hayat, and Muhammad Yasir Hayat Mailk. On topological indices of certain interconnection networks. Applied Mathematics and Computation, 244(??):936–951, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010212.

Iiduka:2011:TTC

[Iid11] Hideaki Iiduka. Threeterm conjugate gradient method for the convex optimization problem over the fixed point set of a nonexpansive mapping. Applied Mathematics and Computation, 217(13):6315–6327, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000099.

Iqbal:2010:SSL

[IISA10] S. Iqbal, M. Idrees, A. M. Siddiqui, and A. R. Ansari. Some solutions of the linear and nonlinear Klein–Gordon equations using the optimal homotopy asymptotic method. Applied Mathematics and Computation, 216(10):2898–2909, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004078.

Iqbal:2011:AOH

[IJ11] S. Iqbal and A. Javed. Application of optimal homotopy asymptotic method for the analytic solution of singular Lane–Emden type equation. Applied Mathematics and Computation, 217(19):7753–7761, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003067.

Ilea:2011:NSD

[IKH11] Catalin G. Ilea, Paweł Kosinski, and Alex C. Hoffmann. Numerical study on the dynamics of a jet-vortex interaction. Applied Mathematics and Computation, 217(11):5103–5112, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000860X.

Ismailov:2011:DTD

[IKL11] M. I. Ismailov, F. Kanca, and D. Lesnic. Determination of a time-dependent heat source under nonlocal boundary and integral overdetermination conditions. Applied Mathematics and Computation, 218(8):4138–4146, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012161.

Inan:2010:ABT

[IKU10] Ibrahim E. Inan, Doḡan Kaya, and Yavuz Ugurlu. Auto-Bäcklund transformation and similarity reductions for coupled Burger’s equation. Applied Mathematics and Computation, 216(9):2507–2511, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000384X.

Ibragimov:2011:SAG

[IKV11] N. H. Ibragimov, R. S.

Khamitova, and A. Valenti. Self-adjointness of a generalized Camassa–Holm equation. Applied Mathematics and Computation, 218(6):2579–2583, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101023X.

Ilter:2011:WMP

[Ilt11] Serkan Ilter. Weak maximum principle for optimal control problems of nonsmooth systems. Applied Mathematics and Computation, 218(3):805–808, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004322.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Imoto:2014:GCM

[Imo14] Tomoaki Imoto. A generalized Conway–Maxwell– Poisson distribution which includes the negative binomial distribution. Applied Mathematics and Computation, 247(??):824–834, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012776.

Ibarra:2010:ABB

[IN10] Hugo Cabrera Ibarra and David A. Lizárraga Navarro. An algorithm based on 3-braids to solve tangle equations arising in the action of Gin DNA invertase. Applied Mathematics and Computation, 216(1):95–106, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000202.

Inchan:2012:VIM

[Inc12] Issara Inchan. Viscosity iteration method for generalized equilibrium problems and fixed point problems of finite family of nonexpansive mappings. Applied Mathematics and Computation, 219(6):2949–2959, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009198.

Iggidr:2010:FSE

[INE10] A. Iggidr, K. Niri, and E. Ould Moulay Ely. Fluctuations in a SIS epidemic model with variable size population. Applied Mathematics and Computation, 217(1):55–64, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002821.

Iriyama:2012:CCA

[IO12] S. Iriyama and M. Ohya.

Computational complexity and applications of quantum algorithm. Applied Mathematics and Computation, 218(16): 8019–8028, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031100693X.

Ismail:2010:SSGb

[IPB10] M. S. Ismail, Marko D.

Petković, and Anjan Biswas. 1-soliton solution of the generalized KP equation with generalized evolution. Applied Mathematics and Computation, 216(7):2220–2225, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310002973.

Ipek:2011:SNC

[İpe11] Ahmet İpek. On the spectral norms of circulant matrices with classical Fibonacci and Lucas numbers entries. Applied Mathematics and Computation, 217(12):6011–6012, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310013111.

Iricanin:2010:BCT

[Iri10a] Bratislav D. Iričanin. The boundedness character of two Stević-type fourthorder difference equations. Applied Mathematics and Computation, 217(5):1857–1862, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007228.

Iricanin:2010:GSS

[Iri10b] Bratislav D. Iričanin. Global stability of some classes of higher-order nonlinear difference equations. Applied Mathematics and Computation, 216(4):1325–1328, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001955.

Irmak:2011:SFV

[Irm11] Hüseyin Irmak. Some families of p-valently analytic functions established by fractional derivatives and their certain consequences. Applied Mathematics and Computation, 218(3):822–826, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000853.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ismail:2014:MAK

[IRS14] Hassan N. A. Ismail, Tamer M. Rageh, and Ghada S. E. Salem. Modified approximation for the KdV-Burgers equation. Applied Mathematics and Computation, 234(??):58– 62, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002203.

Ikhdair:2010:ABS

[IS10a] Sameer M. Ikhdair and Ramazan Sever. Approximate bound state solutions of Dirac equation with Hulthén potential including Coulomb-like tensor potential. Applied Mathematics and Computation, 216(3):911–923, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001281.

Ikhdair:2010:SSD

[IS10b] Sameer M. Ikhdair and Ramazan Sever. Solutions of the spatiallydependent mass Dirac equation with the spin and pseudospin symmetry for the Coulomblike potential. Applied Mathematics and Computation, 216(2):545–555, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000962.

Iturriaga:2010:ENS

[IS10c] Leonelo Iturriaga and Justino Sánchez. Exact number of solutions of stationary reaction–diffusion equations. Applied Mathematics and Computation, 216(4):1250–1258, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001839.

Igelnik:2011:ETM

[IS11] Boris Igelnik and Dan Simon. The eigenvalues of a tridiagonal matrix in biogeography. Applied Mathematics and Computation, 218(1):195–201, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007247.

Ikhdair:2012:ABS

[IS12] Sameer M. Ikhdair and Ramazan Sever. Approximate bound states of the Dirac equation with some physical quantum potentials. Applied Mathematics and Computation, 218 (20):10082–10093, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003311.

Ijyas:2014:LCM

[IS14] V. P. Thafasal Ijyas and S. M. Sameer. Low complexity metaheuristics for joint ML estimation problems. Applied Mathematics and Computation, 230(??):342–358, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013933.

Isaev:2010:TOS

[Isa10] Vyacheslav K. Isaev. To the theory of optimal splines. Applied Mathematics and Computation, 217(3):1095–1109, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006107.

Isik:2011:BSS

[ISG11a] Osman Raşit Işik, Mehmet Sezer, and Zekeriya Güney. Bernstein series solution of a class of linear integro-differential equations with weakly singular kernel. Applied Mathematics and Computation, 217(16):7009–7020, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001561.

Isik:2011:RAB

[ISG11b] Osman Rasit Isik, Mehmet Sezer, and Zekeriya Güney. A rational approximation based on Bernstein polynomials for high order initial and boundary values problems. Applied Mathematics and Computation, 217(22):9438–9450, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005959.

Ishak:2010:SSF

[Ish10] Anuar Ishak. Similarity solutions for flow and heat transfer over a permeable surface with convective boundary condition. Applied Mathematics and Computation, 217(2):837–842, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007113.

Iudin:2012:IPT

[ISH12] D. I. Iudin, Ya. D.

Sergeyev, and M. Hayakawa.

Interpretation of percolation in terms of infinity computations. Applied Mathematics and Computation, 218(16):8099–8111, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013877.

Iscan:2014:NGI

[ISÖ14] Imdat Iscan, Erhan Set, and M. Emin Özdemir. On new general integral inequalities for s-convex functions. Applied Mathematics and Computation, 246(??):306–315, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011503.

Iavernaro:2012:CFD

[IT12] Felice Iavernaro and Donato Trigiante. Continued fractions as dynamical systems. Applied Mathematics and Computation, 218(16):8203–8216, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003158.

Ilic:2010:ANS

[ITLA10] M. Ilic, I. W. Turner, F. Liu, and V. Anh. Analytical and numerical solutions of a one-dimensional fractional-in-space diffusion equation in a composite medium. Applied Mathematics and Computation, 216(8):2248–2262, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003140.

Ito:2013:ECS

[Ito13]  Kazuho Ito.  An en-

ergy conserving spectral scheme for the periodic dynamic elastica. Applied Mathematics and Computation, 224(??):463–472, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031300948X.

Inan:2010:EFM

[IU10] Ibrahim E. Inan and Yavuz Ugurlu. Expfunction method for the exact solutions of fifth order KdV equation and modified Burgers equation. Applied Mathematics and Computation, 217(4):1294–1299, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309006560.

See remark [Asl10d].

Iscan:2014:HHT

[IW14] Imdat Iscan and Shanhe Wu. Hermite–Hadamard type inequalities for harmonically convex functions via fractional integrals. Applied Mathematics and Computation, 238(??):237–244, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400544X.

Izgi:2011:VTA

[Izg11] Aydin Izgi. Voronovskaya type asymptotic approximation by modified Gamma operators. Applied Mathematics and Computation, 217(20):8061–8067, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003419.

See note [Kre13].

Jerez:2014:SFL

[JA14] Silvia Jerez and Mario Arciga. Switch flux limiter method for viscous and nonviscous conservation laws. Applied Mathematics and Computation, 246(??):292–305, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011059.

Jha:2011:OMP

[JAG11] P. C. Jha, Remica Aggarwal, and Anshu Gupta. Optimal media planning for multi-products in segmented market. Applied Mathematics and Computation, 217(16):6802–6818, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013287.

Jankowski:2010:FOA

[Jan10] Tadeusz Jankowski. Firstorder advanced difference equations. Applied Mathematics and Computation, 216(4):1242–1249, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310001827.

Jankowski:2011:BVP

[Jan11] Tadeusz Jankowski. Boundary value problems for difference equations with causal operators. Applied Mathematics and Computation, 218(6):2549–2557, November 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010216.

Jana:2013:CDD

[Jan13a] Debaldev Jana. Chaotic dynamics of a discrete predator-prey system with prey refuge. Applied Mathematics and Computation, 224(??):848–865, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009582.

Jankowski:2013:ESC

[Jan13b] Tadeusz Jankowski. Existence of solutions for a coupled system of difference equations with causal operators. Applied Mathematics and Computation, 219(17):9348–9355, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003615.

Jankowski:2013:ERD

[Jan13c] Tadeusz Jankowski. Existence results to delay fractional differential equations with nonlinear boundary conditions. Applied Mathematics and Computation, 219(17):9155–9164, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002981.

Jankowski:2013:IVP

[Jan13d] Tadeusz Jankowski. Initial value problems for neutral fractional differential equations involving a Riemann–Liouville derivative. Applied Mathematics and Computation, 219(14):7772–7776, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001173.

Jang:2014:IEF

[Jan14a] T. S. Jang. An integral equation formalism for solving the nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 243(??):322–338, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008388.

Jang:2014:USS

[Jan14b] T. S. Jang. Uniqueness and stability of the simultaneous detection of the nonlinear restoring and excitation of a forced nonlinear oscillation. Applied Mathematics and Computation, 228(??):234–239, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012745.

Jankowski:2014:FPA

[Jan14c] Tadeusz Jankowski. Fractional problems with advanced arguments. Applied Mathematics and Computation, 230(??): 371–382, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031301309X.

Jankowski:2014:MIM

[Jan14d] Tadeusz Jankowski. Monotone iterative method for first-order differential equations at resonance. Applied Mathematics and Computation, 233(??):20– 28, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001684.

Jankowski:2014:PSF

[Jan14e] Tadeusz Jankowski. Positive solutions to fractional differential equations involving Stieltjes integral conditions. Applied Mathematics and Computation, 241(??):200–213, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006201.

Jaradat:2010:FGC

[Jar10] M. M. M. Jaradat. Fundamental group and cycle space of dual graphs and applications. Applied Mathematics and Computation, 217(2):580–582, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006600.

Ao:2014:MRS

[jAS14] Ji jun Ao and Jiong Sun. Matrix representations of Sturm– Liouville problems with coupled eigenparameterdependent boundary conditions. Applied Mathematics and Computation, 244(??):142–148, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009400.

Ao:2011:FSS

[jASzZ11] Ji jun Ao, Jiong Sun, and Mao zhu Zhang. The finite spectrum of Sturm–Liouville problems with transmission conditions. Applied Mathematics and Computation, 218(4):1166–1173, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007028.

Jator:2010:SSO

[Jat10] S. N. Jator. Solving second order initial value problems by a hybrid multistep method without predictors. Applied Mathematics and Computation, 217(8):4036–4046, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010404.

Jator:2012:CTS

[Jat12] S. N. Jator. A continuous two-step method of order 8 with a block extension for y′′ = f(x, y, y′). Applied Mathematics and Computation, 219(3):781–791, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006261.

Ao:2014:FOB

[jAzBS14] Ji jun Ao, Fang zhen Bo, and Jiong Sun. Fourth order boundary value problems with finite spectrum. Applied Mathematics and Computation, 244(??):952–958, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401011X.

Jiang:2010:SNS

[JB10] Bo Jiang and Qinsheng Bi. Smooth and non-smooth traveling wave solutions of the Fornberg–Whitham equation with linear dispersion term. Applied Mathematics and Computation, 216(7):2155–2162, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002924.

Justin:2012:ESS

[JBDC12] Mibaile Justin, Gambo Betchewe, Serge Y. Doka, and Kofane Timoleon Crepin. Exact solutions of a semiconductor nonlinear reaction diffusion equation through factorization method. Applied Mathematics and Computation, 219(6):2917–2922, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009149.

Jiang:2011:AEP

[JBL^{+}11] Hao Jiang, Roberto Barrio, Housen Li, Xiangke Liao, Lizhi Cheng, and Fang Su. Accurate evaluation of a polynomial in Chebyshev form. Applied Mathematics and Computation, 217(23):9702–9716, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006242.

Jang:2014:DPZ

[JBW14] Sophia Jang, James Baglama, and Li Wu. Dynamics of phytoplanktonzooplankton systems with toxin producing phytoplankton. Applied Mathematics and Computation, 227(??):717–740, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012204.

Jiang:2013:SSL

[JC13] Wei Jiang and Zhong Chen. Solving a system of linear Volterra integral equations using the new reproducing kernel method. Applied Mathematics and Computation, 219(20):10225– 10230, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003846.

Ji:2014:NMC

[JC14] Jun Ji and Xuzhou Chen. A new method for computing Moore– Penrose inverse through Gauss–Jordan elimination. Applied Mathematics and Computation, 245(??):271–278, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031401039X.

Johnson:2011:MST

[JCB11] Stephen Johnson, Feiyu Chen, and Anjan Biswas. Mathematical structure of topological solitons due to the Sine-Gordon Equation. Applied Mathematics and Computation, 217(13):6372–6378, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000270.

Jiang:2012:MML

[JCW12] Zhongyi Jiang, Fangfang Chen, and Chunqing Wu. Minimizing the maximum lateness in a single-machine scheduling problem with the normal time-dependent and job-dependent learning effect. Applied Mathematics and Computation, 218(18):9438–9441, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002603.

Jiang:2010:PBT

[JCY10] Mei-Qun Jiang, Yang Cao, and Lin-Quan Yao. On parameterized block triangular preconditioners for generalized saddle point problems. Applied Mathematics and Computation, 216(6):1777–1789, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030901073X.

Jung:2014:PAB

[JDD14] Hee Sun Jung, Naokant Deo, and Minakshi Dhamija. Pointwise approximation by Bernstein type operators in mobile interval. Applied Mathematics and Computation, 244(??):683–694, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314009916.

Jankovic:2011:CBD

[JDJ11] Svetlana Janković, Jasmina Djordjević, and Miljana Jovanović. On a class of backward doubly stochastic differential equations. Applied Mathematics and Computation, 217(21):8754–8764, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004966.

See corrigendum [JDJ12].

Jankovic:2012:CCB

[JDJ12] Svetlana Janković, Jasmina Djordjević, and Miljana Jovanović. Corrigendum to “On a class of backward doubly stochastic differential equations” [Applied Mathematics and Computation 217 (2011) 8754–8764]. Applied Mathematics and Computation, 218(17):9033–9034, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001002.

See [JDJ11].

Jin:2014:DBE

[JDLL14] Wenbiao Jin, Chongyang Deng, Yajuan Li, and Jianzhen Liu. Derivative bound estimations on rational conic Bézier curves. Applied Mathematics and Computation, 248(??):113–117, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013563.

Jatav:2014:MSS

[JDN14] Kunwer Singh Jatav, Joydip Dhar, and Atulya K. Nagar. Mathematical study of stage-structured pests control through impulsively released natural enemies with discrete and distributed delays. Applied Mathematics and Computation, 238(??):511–526, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005621.

Jafari:2011:EMM

[JE11] A. A. Jafari and S. A.

Eftekhari. An efficient mixed methodology for free vibration and buckling analysis of orthotropic rectangular plates. Applied Mathematics and Computation, 218(6):2670–2692, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010344.

Jeang:2013:CPP

[Jea13] Angus Jeang. Concurrent product and process parameters determination under deterioration process. Applied Mathematics and Computation, 219(17):9132–9141, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003470.

Jiang:2014:DSA

[JFHZ14] Yiping Jiang, Xiaoling Fu, Kai Huang, and Lindu Zhao. A decomposable self-adaptive projectionbased prediction-correction algorithm for convex time space network flow problem. Applied Mathematics and Computation, 231(??):422–434, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314000708.

Jaggi:2013:CFE

[JGM13] Chandra K. Jaggi, Satish K.

Goel, and Mandeep Mittal. Credit financing in economic ordering policies for defective items with allowable shortages. Applied Mathematics and Computation, 219(10):5268–5282, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011940.

Jia:2010:GAH

[JH10] Xiu-Mei Jia and Lin-Xia Hu. Global attractivity of a higherorder nonlinear difference equation. Applied Mathematics and Computation, 216(3):857–861, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001165.

Jha:2013:FOA

[Jha13]  Navnit Jha.  A fifth

order accurate geometric mesh finite difference method for general nonlinear two point boundary value problems. Applied Mathematics and Computation, 219(16):8425–8434, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002075.

Javed:2014:PFB

[JHA14] Maryiam Javed, T. Hayat, and A. Alsaedi. Peristaltic flow of Burgers’ fluid with compliant walls and heat transfer. Applied Mathematics and Computation, 244(??):654–671, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009667.

Johnson:2011:RDT

[JHN11] Richard W. Johnson, Glen Hansen, and Chris Newman. The role of data transfer on the selection of a single vs. multiple mesh architecture for tightly coupled multiphysics applications. Applied Mathematics and Computation, 217(22):8943–8962, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311004693.

Jin:2011:AAD

[JHP11] Dequan Jin, Zhili Huang, and Jigen Peng. Approximation on attraction domain of Cohen–Grossberg neural networks. Applied Mathematics and Computation, 217(21):8747–8753, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004954.

Jian:2014:FCS

[JHX14] Jinbao Jian, Daolan Han, and Qingjuan Xu. A fast convergent sequential linear equation method for inequality constrained optimization without strict complementarity. Applied Mathematics and Computation, 236(??):93–108, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003798.

Jiang:2013:PFA

[JHZC13] Yangsheng Jiang, Lu Hu, Juanxiu Zhu, and Yanru Chen. PH fitting of the arrival interval distribution of the passenger flow on urban rail transit stations. Applied Mathematics and Computation, 225(??):158–170, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009752.

Ji:2014:TIF

[Ji14a] Jun Ji. Two inverse-of-N-free methods for A_{M,N}^{†}. Applied Mathematics and Computation, 232(??):39– 48, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000861.

Ji:2014:ACS

[Ji14b] Shaochun Ji. Approximate controllability of semilinear nonlocal fractional differential systems via an approximating method. Applied Mathematics and Computation, 236(??):43– 53, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003804.

Jiang:2012:IEA

[Jia12]  Yingchun Jiang.  Im-

proved error analysis for adaptive wavelet algorithms. Applied Mathematics and Computation, 219(4):2134–2141, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200851X.

Jian:2013:BPS

[Jia13a] Shuai Jian. A block preconditioned steepest descent method for symmetric eigenvalue problems. Applied Mathematics and Computation, 219 (20):10198–10217, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003755.

Jiang:2013:EIM

[Jia13b] Weihua Jiang. Eigenvalue interval for multi-point boundary value problems of fractional differential equations. Applied Mathematics and Computation, 219(9):4570–4575, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312010843.

Jiang:2014:SFD

[Jia14]  Weihua Jiang.  Solv-

ability for fractional differential equations at resonance on the half line. Applied Mathematics and Computation, 247(??):90–99, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011771.

Jiang:2014:AMQ

[JJL14] Tongsong Jiang, Ziwu Jiang, and Sitao Ling. An algebraic method for quaternion and complex least squares coneigenproblem in quantum mechanics. Applied Mathematics and Computation, 249(??):222–228, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314014507.

Ji:2012:ALS

[JJP^{+}12] D. H. Ji, S. C. Jeong, Ju H. Park, S. M. Lee, and S. C. Won. Adaptive lag synchronization for uncertain complex dynamical network with delayed coupling. Applied Mathematics and Computation, 218(9):4872–4880, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012975.

Jeong:2013:ASU

[JJPW13] S. C. Jeong, D. H. Ji, Ju H. Park, and S. C. Won. Adaptive synchronization for uncertain chaotic neural networks with mixed time delays using fuzzy disturbance observer. Applied Mathematics and Computation, 219(11):5984–5995, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012842.

Johnpillai:2010:GAK

[JK10] A. G. Johnpillai and C. M.

Khalique. Group analysis of KdV equation with time dependent coefficients. Applied Mathematics and Computation, 216(12):3761–3771, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310005874.

See comment [Gün10].

Ji:2013:IBM

[JK13] Fenghui Ji and Young Ho Kim. Isometries between minimal helicoidal surfaces and rotation surfaces in Minkowski space. Applied Mathematics and Computation, 220(??):1–11, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313005997.

Jaklic:2014:PPM

[JK14a] Gasper Jaklic and Tadej Kanduc. On positivity of principal minors of bivariate Bézier collocation matrix. Applied Mathematics and Computation, 227(??):320–328, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012022.

Jia:2014:NEA

[JK14b] Jiteng Jia and Qiongxiang Kong. A novel elementary algorithm for solving comrade linear systems. Applied Mathematics and Computation, 236(??):642–646, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004585.

Jovanovic:2014:ANA

[JK14c] Miljana Jovanović and Marija Krstić. Analysis of non-autonomous stochastic Gompertz model with delay. Applied Mathematics and Computation, 242(??):101–108, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007115.

Johnpillai:2010:ESK

[JKB10] A. G. Johnpillai, C. M.

Khalique, and Anjan Biswas. Exact solutions of KdV equation with time-dependent coefficients. Applied Mathematics and Computation, 216(10):3114–3119, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003991.

Jangveladze:2011:GFE

[JKN11] Temur Jangveladze, Zurab Kiguradze, and Beny Neta. Galerkin finite element method for one nonlinear integro-differential model. Applied Mathematics and Computation, 217(16):6883–6892, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000695.

Jaksic:2014:MPE

[JKP14a] Rozarija Jaksić, Mario Krnić, and Josip Pecarić. More precise estimates for the Jensen operator inequality obtained via the Lah–Ribaric inequality. Applied Mathematics and Computation, 249(??):346–355, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314014015.

Joh:2014:MTI

[JKP14b] Young-Deuk Joh, Young Mi Kwon, and Won-Kwang Park. MUSIC-type imaging of perfectly conducting cracks in limited-view inverse scattering problems. Applied Mathematics and Computation, 240(??):273–280, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006377.

Ji:2011:PBC

[JKW^{+}11] D. H. Ji, J. H. Koo, S. C.

Won, S. M. Lee, and Ju H. Park. Passivity-based control for Hopfield neural networks using convex representation. Applied Mathematics and Computation, 217(13):6168–6175, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013172.

Jue:2010:LTB

[JL10] Wang Jue and Zhang Lei.

Long-time behavior of the difference solutions to the 2D GKS equation. Applied Mathematics and Computation, 216(5):1418–1427, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002158.

Lian:2012:SAE

[jL12] Shu jun Lian. Smoothing approximation to ℓ_{1} exact penalty function for inequality constrained optimization. Applied Mathematics and Computation, 219(6):3113–3121, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009411.

Jia:2014:MSI

[JL14] Mei Jia and Xiping Liu. Multiplicity of solutions for integral boundary value problems of fractional differential equations with upper and lower solutions. Applied Mathematics and Computation, 232(??):313–323, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001106.

Jiang:2010:IPS

[JLHW10] Yan Jiang, Changmin Liu, Chongchao Huang, and Xianing Wu. Improved particle swarm algorithm for hydrological parameter optimization. Applied Mathematics and Computation, 217(7):3207–3215, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009173.

Jiang:2013:APS

[JLHW13] Yan Jiang, Xuyong Li, Chongchao Huang, and Xianing Wu. Application of particle swarm optimization based on CHKS smoothing function for solving nonlinear bilevel programming problem. Applied Mathematics and Computation, 219(9):4332–4339, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009976.

Ji:2012:GOA

[JLL12] Ying Ji, Yijun Li, and Pengyu Lu. A global optimization algorithm for sum of quadratic ratios problem with coefficients. Applied Mathematics and Computation, 218(19):9965–9973, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312003414.

Jiang:2013:NLD

[JLLD13] Nan Jiang, Shengtao Li, Ting Liu, and Xiujuan Dong. Nonlinear large disturbance attenuation controller design for the power systems with STAT-COM. Applied Mathematics and Computation, 219 (20):10378–10386, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004104.

Lu:2012:FAL

[jLLW12] Dian jun Lu, Keh-Shin Lii, and Yu Wang. The formulae and algorithms for Lagrange-power basis transformation and Lagrange–Newton transformation. Applied Mathematics and Computation, 218(10):5861–5866, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012331.

Jiang:2012:BTW

[JLM^{+}12] Bo Jiang, Yi Lu, Qiang Ma, Yi Cao, and Yunzhi Zhang. Bifurcation of traveling wave solutions for the BBM-like B(2, 2) equation. Applied Mathematics and Computation, 218(14):7375–7381, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015487.

Jin:2013:PSO

[JLTZ13] Xin Jin, Yongquan Liang, Dongping Tian, and Fuzhen Zhuang. Particle swarm optimization using dimension selection methods. Applied Mathematics and Computation, 219(10):5185–5197, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011733.

Ji:2011:CID

[JLW11] Shaochun Ji, Gang Li, and Min Wang. Controllability of impulsive differential systems with nonlocal conditions. Applied Mathematics and Computation, 217(16):6981–6989, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311001494.

Jiang:2013:SPS

[JLW13] Jiqiang Jiang, Lishan Liu, and Yonghong Wu. Symmetric positive solutions to singular system with multi-point coupled boundary conditions. Applied Mathematics and Computation, 220(??):536–548, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006772.

Jiang:2014:BSN

[JLZB14] Bo Jiang, Yi Lu, Jianhao Zhang, and Qinsheng Bi. Bifurcations and some new traveling wave solutions for the CH-γ equation. Applied Mathematics and Computation, 228(??):220–233, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012253.

Jaklic:2010:PCM

[JM10a] Gašper Jaklič and Jolanda Modic. On properties of cell matrices. Applied Mathematics and Computation, 216(7):2016–2023, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002742.

Jukic:2010:NWE

[JM10b] Dragan Jukić and Darija Marković. On nonlinear weighted errors-in-variables parameter estimation problem in the three-parameter Weibull model. Applied Mathematics and Computation, 215(10):3599–3609, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009722.

Jha:2011:TIA

[JM11] Navnit Jha and R. K.

Mohanty. TAGE iterative algorithm and nonpolynomial spline basis for the solution of nonlinear singular second order ordinary differential equations. Applied Mathematics and Computation, 218(7):3289–3296, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311011088.

Jacobs:2013:NAT

[JM13] B. A. Jacobs and E. Momoniat. A novel approach to text binarization via a diffusion-based model. Applied Mathematics and Computation, 225(??):446–460, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010278.

Jaklic:2014:EGD

[JM14a] Gasper Jaklic and Jolanda Modic. Euclidean graph distance matrices of generalizations of the star graph. Applied Mathematics and Computation, 230(??):650–663, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313014483.

Jovanovic:2014:NAP

[JM14b] Bosko S. Jovanović and Zorica D. Milovanović. Numerical approximation of a 2D parabolic transmission problem in disjoint domains. Applied Mathematics and Computation, 228(??):508–519, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012769.

Jarad:2012:HOF

[JMB12] Fahd Jarad, Thabet Abdeljawad (Maraaba), and Dumitru Baleanu. Higher order fractional variational optimal control problems with delayed arguments. Applied Mathematics and Computation, 218(18):9234–9240, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002317.

Jorgensen:2010:LSA

[JMHZ10] Steffen Jørgensen, Guiomar Martín-Herrán, and Georges Zaccour. The Leitmann– Schmitendorf advertising differential game. Applied Mathematics and Computation, 217(3):1110–1116, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000603.

Jiwari:2013:NSB

[JMS13] Ram Jiwari, R. C. Mittal, and Kapil K. Sharma. A numerical scheme based on weighted average differential quadrature method for the numerical solution of Burgers’ equation. Applied Mathematics and Computation, 219(12):6680–6691, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312013112.

Jari:2013:SIF

[JMY13] Rabeea Jari, Lin Mu, and Xiu Ye. Superconvergence of H(div) finite element approximations for the Stokes problem by L^{2}-projection methods. Applied Mathematics and Computation, 219(10):5649–5656, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011770.

Jordan:2010:NCE

[Jor10]  P. M. Jordan.  A

note on Chrystal’s equation. Applied Mathematics and Computation, 217(2):933–936, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006685.

Joshi:2013:CCM

[Jos13] Badal Joshi. Complete characterization by multistationarity of fully open networks with one nonflow reaction. Applied Mathematics and Computation, 219(12):6931–6945, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313000660.

Jawad:2010:MSE

[JPB10a] Anwar Ja’afar Mohamad Jawad, Marko D. Petković, and Anjan Biswas. Modified simple equation method for nonlinear evolution equations. Applied Mathematics and Computation, 217(2):869–877, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310007150.

Jawad:2010:SSF

[JPB10b] Anwar Ja’afar Mohamad Jawad, Marko D. Petković, and Anjan Biswas. Soliton solutions of a few nonlinear wave equations. Applied Mathematics and Computation, 216(9):2649–2658, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003644.

Jawad:2010:SSB

[JPB10c] Anwar Ja’afar Mohamad Jawad, Marko D. Petković, and Anjan Biswas. Soliton solutions of Burgers equations and perturbed Burgers equation. Applied Mathematics and Computation, 216(11):3370–3377, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310005059.

Jivkov:2013:SPQ

[JPI^{+}13] Venelin Jivkov, Philip Philipoff, Anastas Ivanov, Mario Muñoz, Galerida Raikova, Mikhail Tatur, and Philip Michaylov. Spectral properties of quadruple symmetric real functions. Applied Mathematics and Computation, 221(??):344–350, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007261.

Jeong:2012:IAR

[JPK12] Changki Jeong, PooGyeon Park, and Sung Hyun Kim. Improved approach to robust stability and H_{∞} performance analysis for systems with an interval time-varying delay. Applied Mathematics and Computation, 218(21): 10533–10541, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003992.

Jiwari:2012:DQA

[JPM12] Ram Jiwari, Sapna Pandit, and R. C. Mittal. A differential quadrature algorithm to solve the two dimensional linear hyperbolic telegraph equation with Dirichlet and Neumann boundary conditions. Applied Mathematics and Computation, 218(13):7279–7294, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000240.

Ji:2014:RSC

[JQ14] Yude Ji and Jiqing Qiu. Robust stability criterion for uncertain stochastic systems with timevarying delay via sliding mode control. Applied Mathematics and Computation, 238(??):70–81, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004913.

Jiang:2010:CDW

[JS10] Zhi Jie Jiang and Stevo Stević. Compact differences of weighted composition operators from weighted Bergman spaces to weighted-type spaces. Applied Mathematics and Computation, 217(7):3522– 3530, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009744.

Jebelean:2011:GSP

[JŞ11] Petru Jebelean and Călin Şerban. Ground state periodic solutions for difference equations with discrete p-Laplacian. Applied Mathematics and Computation, 217(23):9820–9827, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006461.

Jeyakumar:2012:SBS

[JS12] S. Jeyakumar and B. Senthilnathan.

A study on the behaviour of the server breakdown without interruption in a M^{x}/G(a, b)/1 queueing system with multiple vacations and closedown time. Applied Mathematics and Computation, 219(5):2618–2633, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008892.

Jia:2013:PSO

[JS13a] Jiteng Jia and Tomohiro Sogabe. On particular solution of ordinary differential equations with constant coefficients. Applied Mathematics and Computation, 219(12):6761–6767, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000210.

Jiang:2013:ABN

[JS13b] Fangfang Jiang and Jitao Sun. Asymptotic behavior of neutral delay differential equation of Euler form with constant impulsive jumps. Applied Mathematics and Computation, 219(19):9906–9913, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004220.

Jamalian:2014:RSM

[JS14a] Ali Jamalian and Maziar Salahi. Robust solutions to multi-facility Weber location problem under interval and ellipsoidal uncertainty. Applied Mathematics and Computation, 242(??):179–186, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007176.

Jleli:2014:RPB

[JS14b] Mohamed Jleli and Bessem Samet. Remarks on the paper: Best proximity point theorems: an exploration of a common solution to approximation and optimization problems. Applied Mathematics and Computation, 228(??):366–370, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012472.

See [Bas12a].

Juhas:2014:CGC

[JS14c] Matej Juhás and Valéria Skrivánková. Characterization of general classes of distributions based on independent property of transformed record values. Applied Mathematics and Computation, 226(??):44–50, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010941.

Joubert:2014:NSE

[JSF14] Stephan V. Joubert, Michael Y.

Shatalov, and Temple H. Fay. On numerically solving an eigenvalue problem arising in a resonator gyroscope. Applied Mathematics and Computation, 246(??):561–571, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011655.

Jiang:2010:DBD

[JSH10] Xiaowu Jiang, Qiang Song, and Meiying Hao. Dynamics behaviors of a delayed stage-structured predator–prey model with impulsive effect. Applied Mathematics and Computation, 215(12):4221–4229, February 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010911.

Jayswal:2011:SDC

[JSMA11] Anurag Jayswal, Ioan Stancu-Minasian, and I. Ahmad. On sufficiency and duality for a class of interval-valued programming problems. Applied Mathematics and Computation, 218(8):4119–4127, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012136.

Jimenez:2014:LLR

[JSSB14] J. C. Jimenez, A. Sotolongo, and J. M. Sanchez-Bornot. Locally Linearized Runge Kutta method of Dormand and Prince. Applied Mathematics and Computation, 247(??):589–606, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401217X.

Jau:2013:MQB

[JSWJ13] You-Min Jau, Kuo-Lan Su, Chia-Ju Wu, and Jin-Tsong Jeng. Modified quantum-behaved particle swarm optimization for parameters estimation of generalized nonlinear multi-regressions model based on Choquet integral with outliers. Applied Mathematics and Computation, 221(??):282–295, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006899.

Jiang:2013:NSQ

[JT13] Wei Jiang and Qinglin Tang. Numerical study of quantized vortex interaction in complex Ginzburg– Landau equation on bounded domains. Applied Mathematics and Computation, 222(??):210–230, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007935.

Jain:2011:SWV

[JU11] M. Jain and Shweta Upadhyaya. Synchronous working vacation policy for finite-buffer multiserver queueing system. Applied Mathematics and Computation, 217(24):9916–9932, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005492.

Jukic:2013:NWL

[Juk13] Dragan Jukić. On nonlinear weighted least squares estimation of Bass diffusion model. Applied Mathematics and Computation, 219(14):7891–7900, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313001392.

Jumarie:2012:FSE

[Jum12] Guy Jumarie. On the fractional solution of the equation f(x + y) = f(x)f(y) and its application to fractional Laplace’s transform. Applied Mathematics and Computation, 219(4):1625–1643, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007965.

Jun:2010:SNP

[Jun10a] Younbae Jun. A stable noniterative Prediction/ Correction domain decomposition method for hyperbolic problems. Applied Mathematics and Computation, 216(8):2286–2292, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003188.

Jung:2010:SCI

[Jun10b] Jong Soo Jung. Strong convergence of iterative methods for k-strictly pseudo-contractive mappings in Hilbert spaces. Applied Mathematics and Computation, 215(10): 3746–3753, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300309010042.

Jung:2011:GIS

[Jun11a] Jong Soo Jung. A general iterative scheme for kstrictly pseudo-contractive mappings and optimization problems. Applied Mathematics and Computation, 217(12):5581–5588, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310012361.

Jung:2011:AAF

[Jun11b] Soon-Mo Jung. Approximation of analytic functions by Laguerre functions. Applied Mathematics and Computation, 218(3):832–835, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001214.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Jung:2012:OCR

[Jun12] Eunok Jung. Optimal cardiopulmonary resuscitation technique depending on body size. Applied Mathematics and Computation, 218(19):9614–9622, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001555.

Jovanovic:2013:DNA

[JV13] Miljana Jovanović and Maja Vasilova. Dynamics of non-autonomous stochastic Gilpin–Ayala competition model with time-varying delays. Applied Mathematics and Computation, 219(12): 6946–6964, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313000143.

Jankovic:2010:SAA

[JVK10] Svetlana Janković, Maja Vasilova, and Marija Krstić. Some analytic approximations for neutral stochastic functional differential equations. Applied Mathematics and Computation, 217(8):3615–3623, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000980X.

Jiang:2010:PSN

[JW10a] Weihua Jiang and Bin Wang. Positive solutions for nonhomogeneous boundary value problems in Banach spaces. Applied Mathematics and Computation, 217(7):3504–3510, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009707.

Jin:2010:TEF

[JW10b] Zhong Jin and Yuqing Wang. A type of efficient feasible SQP algorithms for inequality constrained optimization. Applied Mathematics and Computation, 215(10):3589–3598, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009710.

Jiang:2012:NSO

[JW12] Zi-Wu Jiang and Ren-Hong Wang. Numerical solution of onedimensional sine-Gordon equation using high accuracy multiquadric quasiinterpolation. Applied Mathematics and Computation, 218(15):7711–7716, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000574.

Ji:2014:WCD

[JWG^{+}14] Ying Ji, Tienan Wang, Mark Goh, Yong Zhou, and Bo Zou. The worst-case discounted regret portfolio optimization problem. Applied Mathematics and Computation, 239(??):310–319, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006122.

Wang:2013:MSP

[jWtW13] Jun jie Wang and Lian tang Wang. Multisymplectic preissmann scheme for a high order wave equation of KdV type. Applied Mathematics and Computation, 219(9):4400–4409, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010600.

Jin:2014:LSB

[JWW14] Qibing Jin, Zhu Wang, and Jing Wang. Least squares based iterative identification for multivariable integrating and unstable processes in closed loop. Applied Mathematics and Computation, 242(??):10–19, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007243.

Jin:2013:OID

[JWWY13] Qibing Jin, Zhu Wang, Qi Wang, and Ruigeng Yang. Optimal input design for identifying parameters and orders of MIMO systems with initial values. Applied Mathematics and Computation, 224(??):735–742, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009715.

Jiang:2012:NSC

[JWZ^{+}12] Yanqun Jiang, S. C.

Wong, Peng Zhang, Ruxun Liu, Yali Duan, and Keechoo Choi. Numerical simulation of a continuum model for bi-directional pedestrian flow. Applied Mathematics and Computation, 218(10):6135–6143, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014457.

Jia:2010:SSS

[JX10] Huabing Jia and Wei Xu.

Solitons solutions for some nonlinear evolution equations. Applied Mathematics and Computation, 217(4):1678–1687, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030900890X.

Jia:2011:EPS

[JXA11] Yunfeng Jia, Hong-Kun Xu, and Ravi P. Agarwal. Existence of positive solutions for a prey-predator model with refuge and diffusion. Applied Mathematics and Computation, 217(21):8264–8276, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003353.

Jia:2014:RRP

[JXK14] Huimei Jia, Zhengrong Xiang, and Hamid Reza Karimi. Robust reliable passive control of uncertain stochastic switched time-delay systems. Applied Mathematics and Computation, 231(??):254–267, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000575.

Ji:2012:ECC

[JY12] Jun Ji and Bo Yang. Eigenvalue comparisons for a class of boundary value problems of discrete beam equation. Applied Mathematics and Computation, 218(9):5402–5408, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013543.

Jang:2010:DTA

[JYCY10] Ming-Jyi Jang, Yen-Liang Yeh, Chieh-Li Chen, and Wei-Chih Yeh. Differential transformation approach to thermal conductive problems with discontinuous boundary condition. Applied Mathematics and Computation, 216(8):2339–2350, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003334.

Jiang:2013:RIP

[JYF13] Zhigang Jiang, Jinlong Yuan, and Enmin Feng. Robust identification and its properties of nonlinear bilevel multi-stage dynamic system. Applied Mathematics and Computation, 219(12):6979–6985, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313000234.

Jin:2014:DSC

[JYL14] Li Jin, Hongxia Yu, and Zhisong Liu. Differential systems for constrained optimization via a nonlinear augmented Lagrangian. Applied Mathematics and Computation, 235(??):482–491, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003452.

Ji:2010:SCM

[JYYJ10] Mingjun Ji, Hualong Yang, Yongzhi Yang, and Zhihong Jin. A Single Component Mutation Evolutionary Programming. Applied Mathematics and Computation, 215(10):3759–3768, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009655.

Jiang:2010:UBW

[JZ10] Yi Jiang and Shihui Zhu. Upper bound of weak-limitation for the Gross–Pitaevskii equation. Applied Mathematics and Computation, 217(2):488–494, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006405.

Jaklic:2011:CVM

[JZ11] Gasper Jaklic and Emil Zagar. Curvature variation minimizing cubic Hermite interpolants. Applied Mathematics and Computation, 218(7):3918–3924, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311012112.

Jammazi:2013:RSA

[JZ13] Chaker Jammazi and Maâli Zaghdoudi. On the rational stability of autonomous dynamical systems. Applications to control chained systems. Applied Mathematics and Computation, 219(20): 10158–10171, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313003573.

Jia:2013:SAS

[JZG13] Xinfeng Jia, Fanhua Zeng, and Yongan Gu. Semianalytical solutions to onedimensional advectiondiffusion equations with variable diffusion coefficient and variable flow velocity. Applied Mathematics and Computation, 221(??):268–281, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006917.

Jiao:2012:AHS

[JZW12] Xiaoyu Jiao, Ying Zheng, and Bo Wu. Approximate homotopy symmetry and infinite series solutions to the perturbed mKdV equation. Applied Mathematics and Computation, 218(17):8486–8491, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001270.

Juan:2010:HBB

[JZY10] Zhang Juan, Zhang Zhonghua, and Suo Yaohong. Hopf bifurcation of a bacteriaimmunity system with delayed quorum sensing. Applied Mathematics and Computation, 215(11):3936–3949, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010315.

Jia:2012:GAK

[JZY12] Xiaolin Jia, Caidi Zhao, and Xinbo Yang. Global attractor and Kolmogorov entropy of three component reversible Gray– Scott model on infinite lattices. Applied Mathematics and Computation, 218(19):9781–9789, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002688.

Ji:2013:QBS

[JZZ13] Ruihong Ji, Shuangshuang Zhou, and Sining Zheng. Quenching behavior of solutions in coupled heat equations with singular multinonlinearities. Applied Mathematics and Computation, 223(??):401–410, October 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031300859X.

Zhou:2010:MPC

[jZzFbL10] Zhen jiang Zhou, Jing zhi Fu, and Zhi bin Li. Maple packages for computing Hirota’s bilinear equation and multisoliton solutions of nonlinear evolution equations. Applied Mathematics and Computation, 217(1):92–104, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000531X.

Khalique:2010:ESG

[KA10a] Chaudry Masood Khalique and Abdullahi R. Adem. Exact solutions of a generalized (3 + 1)-dimensional Kadomtsev–Petviashvili equation using Lie symmetry analysis. Applied Mathematics and Computation, 216(10):2849–2854, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004017.

Khudaverdiyev:2010:GES

[KA10b] K. I. Khudaverdiyev and A. G. Alieva. On the global existence of solution to one-dimensional fourth order nonlinear Sobolev type equations. Applied Mathematics and Computation, 217(1):347–354, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006260.

Khattri:2011:SOD

[KA11] Sanjay K. Khattri and Ioannis K. Argyros. Sixth order derivative free family of iterative methods. Applied Mathematics and Computation, 217(12):5500–5507, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001221X.

Khan:2012:EYM

[KA12] R. A. Khan and F. Ahmad. Extrema of Young’s modulus in hexagonal materials. Applied Mathematics and Computation, 219(4):2260–2266, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200865X.

Kattan:2013:DSA

[KA13a] Ali Kattan and Rosni Abdullah. A dynamic selfadaptive harmony search algorithm for continuous optimization problems. Applied Mathematics and Computation, 219(16):8542–8567, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313002129.

Kilic:2013:MIS

[KA13b] Emrah Kiliç and Talha Arikan. More on the infinite sum of reciprocal Fibonacci, Pell and higher order recurrences. Applied Mathematics and Computation, 219(14):7783–7788, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001197.

Kudenatti:2011:AAS

[KAB11] Ramesh B. Kudenatti, Vishwanath B. Awati, and N. M. Bujurke. Approximate analytical solutions of a class of boundary layer equations over nonlinear stretching surface. Applied Mathematics and Computation, 218(6):2952–2959, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010897.

Karpuz:2011:NSW

[KAÇ^{+}11] E. Güzel Karpuz, Firat Ates, A. Sinan Çevik, I. Naci Cangül, and A. Dilek Maden (Güngör).

The next step of the word problem over monoids. Applied Mathematics and Computation, 218(3):794– 798, October 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311004449.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Khalili:2013:BNI

[KAF13] S. M. R. Khalili, P. Abbaspour, and K. Malekzadeh Fard. Buckling of nonideal simply supported laminated plate on Pasternak foundation. Applied Mathematics and Computation, 219(12):6420–6430, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031201332X.

Khan:2012:OML

[KAG12] Subuhi Khan and Ahmed Ali Al-Gonah. Operational methods and Laguerre– Gould Hopper polynomials. Applied Mathematics and Computation, 218(19):9930–9942, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312003384.

Kim:2011:LBE

[KAK11] Sangdong Kim, Soyoung Ahn, and Philsu Kim. Local boundary element based a new finite difference representation for Poisson equations. Applied Mathematics and Computation, 217(12):5186–5198, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011884.

Kalicinska:2012:ALM

[Kal12] Barbora Volná Kalicinská. Augmented IS-LM model based on particular functions. Applied Mathematics and Computation, 219(3):1244–1262, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007333.

Kalogiratou:2013:DIT

[Kal13] Z. Kalogiratou. Diagonally implicit trigonometrically fitted symplectic Runge–Kutta methods. Applied Mathematics and Computation, 219(14):7406–7412, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300060X.

Kamberaj:2014:GSQ

[Kam14a] Hiqmet Kamberaj. q-Gaussian swarm quantum particle intelligence on predicting global minimum of potential energy function. Applied Mathematics and Computation, 229(??):94–106, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301312X.

Kamocki:2014:EOS

[Kam14b] Rafał Kamocki. On the existence of optimal solutions to fractional optimal control problems. Applied Mathematics and Computation, 235(??):94–104, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003348.

Kang:2011:SSI

[Kan11] Dongsheng Kang. Semilinear systems involving multiple critical Hardy– Sobolev exponents and three singular points. Applied Mathematics and Computation, 218(8):4514– 4522, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031101263X.

Kanas:2012:UAF

[Kan12a] S. Kanas. An unified approach to the Fekete– Szegő problem. Applied Mathematics and Computation, 218(17):8453–8461, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200121X.

Kang:2012:EDR

[Kan12b]  Jum-Ran Kang.  En-

ergy decay rates for von Kármán system with memory and boundary feedback. Applied Mathematics and Computation, 218(18):9085–9094, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300312001932.

Kang:2013:PAN

[Kan13] Jum-Ran Kang. Pullback attractors for the non-autonomous coupled suspension bridge equations. Applied Mathematics and Computation, 219(16):8747–8758, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001781.

Karnaukh:2010:RPS

[Kar10] Ie. Karnaukh. Risk process with stochastic income and two-step premium rate. Applied Mathematics and Computation, 217(2):775–781, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007010.

Karaca:2013:PSS

[Kar13a] Ilkay Yaslan Karaca. On positive solutions for second-order boundary value problems of functional differential equations. Applied Mathematics and Computation, 219(10):5433–5439, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011265.

Karpuz:2013:SCO

[Kar13b] Basak Karpuz. Sufficient conditions for the oscillation and asymptotic behaviour of higher-order dynamic equations of neutral type. Applied Mathematics and Computation, 221(??):453–462, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300739X.

Karsli:2014:FTC

[Kar14] Harun Karsli. Fatou type convergence of nonlinear m-singular integral operators. Applied Mathematics and Computation, 246(??):221–228, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011114.

Kassem:2012:MNS

[Kas12a] Mohamed Abd El-Hady Kassem. Multiobjective nonlinear second order symmetric duality with (K, F)-pseudoconvexity. Applied Mathematics and Computation, 219(4):2142– 2148, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312008533.

Kazmi:2012:CSI

[KAS12b] K. R. Kazmi, Naeem Ahmad, and Mohammad Shahzad. Convergence and stability of an iterative algorithm for a system of generalized implicit variational-like inclusions in Banach spaces. Applied Mathematics and Computation, 218(18):9208–9219, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002287.

Khan:2010:SPH

[KASY10] Subuhi Khan, Mustafa Walid Al-Saad, and Ghazala Yasmin. Some properties of Hermite-based Sheffer polynomials. Applied Mathematics and Computation, 217(5):2169–2183, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007617.

Katugampola:2011:NAG

[Kat11] Udita N. Katugampola. New approach to a generalized fractional integral. Applied Mathematics and Computation, 218(3):860–865, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004309.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Katatbeh:2014:ESO

[Kat14] Qutaibeh D. Katatbeh. Eigenvalues of Schrödinger operators with definite and indefinite weights. Applied Mathematics and Computation, 247(??): 125–140, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011837.

Kauffman:2012:CPI

[Kau12] Louis H. Kauffman. Categorical pairs and the indicative shift. Applied Mathematics and Computation, 218(16):7989–8004, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200080X.

Kaushik:2014:EEC

[Kau14]  Aditya Kaushik.  Er-

ror estimates for a class of partial functional differential equation with small dissipation. Applied Mathematics and Computation, 226(??):250–257, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010977.

Kaya:2011:ENS

[Kay11] Mustafa Kaya. The effects of a new selection operator on the performance of a genetic algorithm. Applied Mathematics and Computation, 217(19):7669–7678, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002773.

Kitayama:2012:SAO

[KAY12] Satoshi Kitayama, Masao Arakawa, and Koetsu Yamazaki. Sequential approximate optimization for discrete design variable problems using radial basis function network. Applied Mathematics and Computation, 219(8):4143–4156, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010314.

Koc:2011:NSO

[KB11a] Mehmet Koç and Atalay Barkana. A new solution to one sample problem in face recognition using FLDA. Applied Mathematics and Computation, 217(24):10368–10376, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007181.

Koken:2011:PIP

[KB11b] Fikri Köken and Durmuş Bozkurt. Positive integer powers for one type of odd order circulant matrices. Applied Mathematics and Computation, 217(9):4377–4381, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010611.

Karmakar:2013:NMS

[KB13] Samiran Karmakar and Asoke Kumar Bhunia. A new multi-section based technique for constrained optimization problems with interval-valued objective function. Applied Mathematics and Computation, 225(??):487–502, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313010217.

Khajanchi:2014:SBA

[KB14] Subhas Khajanchi and Sandip Banerjee. Stability and bifurcation analysis of delay induced tumor immune interaction model. Applied Mathematics and Computation, 248(??):652–671, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013757.

Kalashnikova:2014:CES

[KBAvBW14] Irina Kalashnikova, Matthew F.

Barone, Srinivasan Arunajatesan, and Bart G. van Bloemen Waanders. Construction of energy-stable projection-based reduced order models. Applied Mathematics and Computation, 249(??):569–596, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314014489.

Kudenatti:2012:NSM

[KBB12] Ramesh B. Kudenatti, D. P. Basti, and N. M. Bujurke. Numerical solution of the MHD Reynolds equation for squeeze film lubrication between two parallel surfaces. Applied Mathematics and Computation, 218(18):9372–9382, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002512.

Konak:2011:DPA

[KBG11] Abdullah Konak, Michael R. Bartolacci, and Bezalel Gavish. A dynamic programming approach for batch sizing in a multi-stage production process with random yields. Applied Mathematics and Computation, 218(4):1399–1406, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008496.

Kanwar:2013:NOC

[KBK13] V. Kanwar, Saurabh Bhatia, and Munish Kansal. New optimal class of higher-order methods for multiple roots, permitting f′(x_{n}) = 0. Applied Mathematics and Computation, 222(??):564–574, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007467.

Kim:2014:BBO

[KBYL14] Sung-Soo Kim, Ji-Hwan Byeon, Hong Yu, and Hongbo Liu. Biogeographybased optimization for optimal job scheduling in cloud computing. Applied Mathematics and Computation, 247(??):266–280, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012247.

Kang:2012:PSS

[KC12] Shugui Kang and Sui Sun Cheng. Periodic solutions for second order periodic differential equations under scalable control. Applied Mathematics and Computation, 218(18):9138–9146, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002202.

Kim:2013:EEF

[KC13a] Kyung Joong Kim and Ronald Cools. Extended exponentially fitted interpolation formulas for oscillatory functions. Applied Mathematics and Computation, 224(??):178–195, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009077.

Kovtanyuk:2013:IMS

[KC13b] Andrey E. Kovtanyuk and Alexander Yu. Chebotarev. An iterative method for solving a complex heat transfer problem. Applied Mathematics and Computation, 219(17):9356–9362, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003524.

Kovtanyuk:2014:SAC

[KCBH14] Andrey E. Kovtanyuk, Alexander Yu. Chebotarev, Nikolai D. Botkin, and Karl-Heinz Hoffmann. Solvability of P_{1} approximation of a conductiveradiative heat transfer problem. Applied Mathematics and Computation, 249(??):247–252, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014295.

Krisnangkura:2012:ASG

[KCC12] Moke Krisnangkura, Settapat Chinviriyasit, and Wirawan Chinviriyasit. Analytic study of the generalized Burger’s–Huxley equation by hyperbolic tangent method. Applied Mathematics and Computation, 218(22):10843– 10847, July 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312004432.

Koskie:2010:ESF

[KCG10] S. Koskie, C. Coumarbatch, and Z. Gajic. Exact slow-fast decomposition of the singularly perturbed matrix differential Riccati equation. Applied Mathematics and Computation, 216(5):1401–1411, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002122.

Kamil:2014:LCO

[KCGH14] Mustafa Kamil, Thitipun Chobtrong, Ersan Günes, and Markus Haid. Low-cost object tracking with MEMS sensors, Kalman filtering and simplified twofilter-smoothing. Applied Mathematics and Computation, 235(??):323–331, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003683.

Kim:2008:TRG

[KCK08] Chihurn Kim, Geon Ho Choe, and Dong Han Kim. Tests of randomness by the gambler’s ruin algorithm. Applied Mathematics and Computation, 199(1):195–210, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300307009873.

See critical remarks [PJ10].

Kuo:2013:ADN

[KCLC13] Shyh-Rong Kuo, Jeng-Tzong Chen, Jia-Wei Lee, and Yi-Wei Chen. Analytical derivation and numerical experiments of degenerate scales for regular N-gon domains in two-dimensional Laplace problems. Applied Mathematics and Computation, 219(10):5668–5683, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011617.

Ketfi-cherif:2014:GDM

[KcZ14] Amine Ketfi-cherif and Abdelkader Ziadi. Global descent method for constrained continuous global optimization. Applied Mathematics and Computation, 244(??):209–221, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009333.

Kang:2011:INF

[KCZK11] Tong Kang, Tao Chen, Huai Zhang, and Kwang Ik Kim. Improved T -ψ nodal finite element schemes for eddy current problems. Applied Mathematics and Computation, 218(2):287–302, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007326.

Kucche:2013:ERQ

[KD13] Kishor D. Kucche and M. B. Dhakne. On existence results and qualitative properties of mild solution of semilinear mixed Volterra–Fredholm functional integrodifferential equations in Banach spaces. Applied Mathematics and Computation, 219(22):10806–10816, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004943.

Khalili:2013:MTD

[KDS13] S. M. R. Khalili, M. Botshekanan Dehkordi, and M. Shariyat. Modeling and transient dynamic analysis of pseudoelastic SMA hybrid composite beam. Applied Mathematics and Computation, 219(18):9762–9782, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003536.

Khalili-Damghani:2012:IMO

[KDTSN12] Kaveh Khalili-Damghani, Madjid Tavana, and Soheil Sadi-Nezhad. An integrated multi-objective framework for solving multi-period project selection problems. Applied Mathematics and Computation, 219(6):3122–3138, November 25, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009423.

Kilicman:2011:NWE

[KE11] Adem Kiliçman and Hassan Eltayeb. Note on the wave equation and tensor products. Applied Mathematics and Computation, 218(3):871–877, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000877.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Kamel:2013:VSU

[KEGH13] M. M. Kamel, W. A. A. El-Ganaini, and Y. S. Hamed. Vibration suppression in ultrasonic machining described by nonlinear differential equations via passive controller. Applied Mathematics and Computation, 219(9):4692–4701, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011009.

Kassem:2014:OMB

[KEH14] Mohamed Abd El-Hady Kassem and Mohamed Abd Allah El-Hadidy. Optimal multiplicative Bayesian search for a lost target. Applied Mathematics and Computation, 247(??):795–802, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012648.

Kekec:2011:SLP

[Kek11]  Gülcan Kekeç.  On

some lacunary power series with algebraic coefficients and Mahler’s Unumbers. Applied Mathematics and Computation, 218(3):866–870, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004310.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Keller:2012:PAC

[Kel12] Paweł Keller. A practical algorithm for computing Cauchy principal value integrals of oscillatory functions. Applied Mathematics and Computation, 218(9):4988–5001, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013105.

Kersey:2013:DBF

[Ker13] Scott N. Kersey. Dual basis functions in subspaces of inner product spaces. Applied Mathematics and Computation, 219(19):10012– 10024, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004153.

Kesri:2013:EMO

[Kes13] M’hamed Kesri. An explicit method for optimal control problems. Applied Mathematics and Computation, 219(15):8213–8221, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001227.

Kutluay:2010:EMS

[KET10] S. Kutluay, A. Esen, and O. Tasbozan. The (\frac{G′}{G})- expansion method for some nonlinear evolution equations. Applied Mathematics and Computation, 217(1):384–391, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006326.

Kano:2011:OSI

[KFM11a] Hiroyuki Kano, Hiroyuki Fujioka, and Clyde F. Martin. Optimal smoothing and interpolating splines with constraints. Applied Mathematics and Computation, 218(5):1831– 1844, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031100912X.

Ke:2011:PMA

[KfM11b] Yunquan Ke and Chun fang Miao. Pattern memory analysis of secondorder neural networks with time-delay. Applied Mathematics and Computation, 217(22):8883–8898, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005704.

Kadalbajoo:2010:BSN

[KG10] Mohan K. Kadalbajoo and Vikas Gupta. A brief survey on numerical methods for solving singularly perturbed problems. Applied Mathematics and Computation, 217(8):3641–3716, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010532.

Kar:2011:DBD

[KG11] T. K. Kar and Abhijit Ghorai. Dynamic behaviour of a delayed predator-prey model with harvesting. Applied Mathematics and Computation, 217(22):9085–9104, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004942.

Kloppers:2013:LVM

[KG13] P. H. Kloppers and J. C. Greeff. Lotka–Volterra model parameter estimation using experiential data. Applied Mathematics and Computation, 224(??):817–825, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009806.

Kaur:2014:SIS

[KG14] Lakhveer Kaur and R. K. Gupta. Some invariant solutions of field equations with axial symmetry for empty space containing an electrostatic field. Applied Mathematics and Computation, 231(??):560–565, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014100.

Kiran:2012:NHA

[KGB12] Mustafa Servet Kiran, Mesut Gündüz, and Ömer Kaan Baykan. A novel hybrid algorithm based on particle swarm and ant colony optimization for finding the global minimum. Applied Mathematics and Computation, 219(4):1515–1521, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007813.

Khanra:2011:EMD

[KGC11] S. Khanra, S. K. Ghosh, and K. S. Chaudhuri. An EOQ model for a deteriorating item with time dependent quadratic demand under permissible delay in payment. Applied Mathematics and Computation, 218(1):1–9, September 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006321.

Kahraman:2012:QSH

[KGH12] Ferdag Kahraman, Ismaïl Gök, and H. Hïlmï Hacisalïhoglu. On the quaternionic B_{2} slant helices in the semi-Euclidean space E_{2}^{4}. Applied Mathematics and Computation, 218(11): 6391–6400, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014652.

Kundu:2014:PGT

[KGM14] S. Kundu, S. Gupta, and S. Manna. Propagation of G-type seismic waves in heterogeneous layer lying over an initially stressed heterogeneous half-space. Applied Mathematics and Computation, 234(??):1– 12, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002197.

Kumar:2010:RPE

[KGO10] Rajan Kumar, Ranjan Ganguli, and S. N. Omkar. Rotorcraft parameter estimation using radial basis function neural network. Applied Mathematics and Computation, 216(2):584–597, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001050.

Kobayashi:2014:OBA

[KH14a] Koichi Kobayashi and Kunihiko Hiraishi. An optimization-based approach to sampled-data control of networked control systems with multiple delays. Applied Mathematics and Computation, 247(??):786–794, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012892.

Kurina:2014:ASL

[KH14b] Galina Kurina and Nguyen Thi Hoai. Asymptotic solution of a linear-quadratic problem with discontinuous coefficients and cheap control. Applied Mathematics and Computation, 232(??):347–364, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313013878.

Khalique:2011:SSB

[Kha11] Chaudry Masood Khalique. Stationary solutions for the Biswas–Milovic equation. Applied Mathematics and Computation, 217(18):7400–7404, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002049.

Kacha:2012:SHB

[KHA12] Abdelaziz Kacha, My Lhassan Hbid, and Pierre Auger. Stability and Hopf bifurcation of a mathematical model describing bacteria-fish interaction in marine environment. Applied Mathematics and Computation, 218(17):8226–8241, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310013019.

Khader:2014:LCM

[Kha14a] M. M. Khader. Laguerre collocation method for the flow and heat transfer due to a permeable stretching surface embedded in a porous medium with a second order slip and viscous dissipation. Applied Mathematics and Computation, 243(??):503–513, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008479.

Khajanchi:2014:DBB

[Kha14b] Subhas Khajanchi. Dynamic behavior of a Beddington–DeAngelis type stage structured predatorprey model. Applied Mathematics and Computation, 244(??):344–360, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009539.

Khajavi:2014:GTN

[Kha14c] R. Khajavi. General templates for n-noded bar elements based on reduced representations and numerical dispersion reduction by optimized finite elements. Applied Mathematics and Computation, 233(??):445–462, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002616.

Kang:2012:BOT

[KHD12] Qinma Kang, Hong He, and Rong Deng. Biobjective task assignment in heterogeneous distributed systems using honeybee mating optimization. Applied Mathematics and Computation, 219(5):2589–2600, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008867.

Kheloufi:2013:EUR

[Khe13] Arezki Kheloufi. Existence and uniqueness results for parabolic equations with Robin type boundary conditions in a non-regular domain of R^{3}. Applied Mathematics and Computation, 220(??):756–769, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007777.

Keeling:2011:TVB

[KHK^{+}11] Stephen L. Keeling, Michael Hintermüller, Florian Knoll, Daniel Kraft, and Antoine Laurain. A total variation based approach to correcting surface coil magnetic resonance images. Applied Mathematics and Computation, 218(2):219–232, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003389.

Khorsand:2014:PTS

[Kho14] Zahra Khorsand. Particle trajectories in the Serre equations. Applied Mathematics and Computation, 230(??):35–42, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012940.

Ke:2011:PMR

[KHP11] Jau-Chuan Ke, Kai-Bin Huang, and Wen Lea Pearn. The performance measures and randomized optimization for an unreliable server M^{[x]}/G/1 vacation system. Applied Mathematics and Computation, 217(21):8277–8290, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100347X.

Kojima:2010:ARC

[KI10a] Kentaro Kojima and Takehiro Inohara. An acyclic relation for comparison of bargaining powers of coalitions and its interrelationship with bargaining set. Applied Mathematics and Computation, 215(10):3665–3668, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009783.

Kojima:2010:MCC

[KI10b] Kentaro Kojima and Takehiro Inohara. Methods for comparison of coalition influence on games in characteristic function form and their interrelationships. Applied Mathematics and Computation, 217(8):4047–4050, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010428.

Kitamura:2011:NBR

[KI11a] Masahito Kitamura and Takehiro Inohara. A new binary relation to compare viability of winning coalitions and its interrelationships to desirability relation and blockability relation. Applied Mathematics and Computation, 217(13):6176–6184, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013184.

Kojima:2011:SPO

[KI11b] Kentaro Kojima and Takehiro Inohara. A strict partial order on payoff configurations and its some properties. Applied Mathematics and Computation, 218(5):2108–2112, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100960X.

Khan:2012:RTF

[KI12a] Ladlay Khan and M. Imdad. Rhoades type fixed point theorems for two hybrid pairs of mappings in metrically convex spaces. Applied Mathematics and Computation, 218(17):8861–8868, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001725.

Kojima:2012:CVD

[KI12b] Kentaro Kojima and Takehiro Inohara. Coalition values derived from methods for comparison of coalition influence for games in characteristic function form. Applied Mathematics and Computation, 219(3):1345–1353, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007400.

Kiani:2014:NMS

[Kia14] Keivan Kiani. A nonlocal meshless solution for flexural vibrations of double-walled carbon nanotubes. Applied Mathematics and Computation, 234(??):557–578, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000526.

Kim:2010:NTS

[Kim10] Young Ik Kim. A new twostep biparametric family of sixth-order iterative methods free from second derivatives for solving nonlinear algebraic equations. Applied Mathematics and Computation, 215(9):3418–3424, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009412.

Kim:2011:EAF

[Kim11] Kyung Joong Kim. Error analysis for frequencydependent interpolation formulas using first derivatives. Applied Mathematics and Computation, 217(19):7703–7717, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002967.

Kim:2012:GPE

[Kim12] Kwang-Yeon Kim. Guaranteed a posteriori error estimator for mixed finite element methods of elliptic problems. Applied Mathematics and Computation, 218(24):11820–11831, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005462.

Kim:2014:SCI

[Kim14a] Gang Eun Kim. Strong convergence for an implicit iteration process in Hilbert spaces. Applied Mathematics and Computation, 239(??):434–437, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004792.

Kim:2014:AAC

[Kim14b] Seungil Kim. Analysis of an approximate cloaking for acoustic scattering problems in R^{3}. Applied Mathematics and Computation, 232(??):117–131, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000101.

Kiryakova:2011:CUD

[Kir11]  V. Kiryakova.  Crite-

ria for univalence of the Dziok–Srivastava and the Srivastava–Wright operators in the class A. Applied Mathematics and Computation, 218(3):883–892, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001111.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Karoui:2010:EAC

[KJ10] Abderrazek Karoui and Adel Jawahdou. Existence and approximate L^{p} and continuous solutions of nonlinear integral equations of the Hammerstein and Volterra types. Applied Mathematics and Computation, 216(7):2077–2091, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002845.

Kar:2012:SBA

[KJ12] T. K. Kar and Soovoojeet Jana. Stability and bifurcation analysis of a stage structured predator prey model with time delay. Applied Mathematics and Computation, 219(8):3779–3792, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009940.

Koo:2010:SSC

[KJW10] J. H. Koo, D. H. Ji, and S. C. Won. Synchronization of singular complex dynamical networks with time-varying delays. Applied Mathematics and Computation, 217(8):3916–3923, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010155.

Kargin:2011:GEP

[KK11a] Levent Kargin and Veli Kurt. On the generalization of the Euler polynomials with the real parameters. Applied Mathematics and Computation, 218(3):856–859, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004565.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Keslerova:2011:NSL

[KK11b] R. Keslerová and K. Kozel.

Numerical solution of laminar incompressible generalized Newtonian fluids flow. Applied Mathematics and Computation, 217(11):5125–5133, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007940.

Khan:2011:NPS

[KK11c] Arshad Khan and Pooja Khandelwal. Non-polynomial sextic spline approach for the solution of fourthorder boundary value problems. Applied Mathematics and Computation, 218(7):3320–3329, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011143.

Kasumba:2012:FSP

[KK12a] H. Kasumba and K. Kunisch. On free surface PDE constrained shape optimization problems. Applied Mathematics and Computation, 218(23):11429–11450, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005176.

Kazantzis:2012:ROS

[KK12b] Nikolaos Kazantzis and Vasiliki Kazantzi. Reducedorder state reconstruction for nonlinear dynamical systems in the presence of model uncertainty. Applied Mathematics and Computation, 218(23):11708–11718, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005814.

Kudryashov:2012:QES

[KK12c] Nikolay A. Kudryashov and Mark B. Kochanov. Quasi-exact solutions of nonlinear differential equations. Applied Mathematics and Computation, 219(4):1793–1804, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008119.

Kim:2013:SSI

[KK13a] Daeyeoul Kim and Min-Soo Kim. Some symmetric identities on higher order q-euler polynomials and multivariate fermionic p-adic q-integral on Z_{p}. Applied Mathematics and Computation, 221(??):558–567, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007376.

Ko:2013:ISS

[KK13b] K. H. Ko and K. Kim. Improved subdivision scheme for the root computation of univariate polynomial equations. Applied Mathematics and Computation, 219(14):7450–7464, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000714.

Korotov:2013:LNT

[KK13c] Sergey Korotov and Michal Krízek. Local nonobtuse tetrahedral refinements around an edge. Applied Mathematics and Computation, 219(13):7236–7240, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005236.

Kumar:2013:NSC

[KK13d] Rajesh Kumar and Jitendra Kumar. Numerical simulation and convergence analysis of a finite volume scheme for solving general breakage population balance equations. Applied Mathematics and Computation, 219(10):5140–5151, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011289.

Kalas:2014:PSG

[KK14a] J. Kalas and Z. Kaderábek. Periodic solutions of a generalized Van der Pol– Mathieu differential equation. Applied Mathematics and Computation, 234(??):192–202, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002148.

Kandhway:2014:HRC

[KK14b] Kundan Kandhway and Joy Kuri. How to run a campaign: Optimal control of SIS and SIR information epidemics. Applied Mathematics and Computation, 231(??):79–92, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000022.

Kim:2014:SIS

[KK14c] Dae San Kim and Taekyun Kim. Some identities of symmetry for the generalized q-euler polynomials. Applied Mathematics and Computation, 235(??):408–411, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002975.

Khan:2012:NHP

[KKA12] Mumtaz Ahmad Khan, Abdul Hakim Khan, and Naeem Ahmad. A note on Hermite polynomials of several variables. Applied Mathematics and Computation, 218(11):6385–6390, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014640.

Khmelnytskaya:2012:SPP

[KKBO12] Kira V. Khmelnytskaya, Vladislav V. Kravchenko, and Jesús A. Baldenebro-Obeso. Spectral parameter power series for fourthorder Sturm–Liouville problems. Applied Mathematics and Computation, 219(8):3610–3624, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200954X.

Khan:2014:ANT

[KKH14] Abdul Rahim Khan, Vivek Kumar, and Nawab Hussain. Analytical and numerical treatment of Jungck-type iterative schemes. Applied Mathematics and Computation, 231(??): 521–535, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313014409.

Krejic:2011:VOP

[KKR11] Natasa Krejíc, Miles Kumaresan, and Andrea Roznjik. VaR optimal portfolio with transaction costs. Applied Mathematics and Computation, 218(8):4626–4637, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012926.

Kazmi:2011:SGV

[KKS11] K. R. Kazmi, F. A. Khan, and Mohammad Shahzad. A system of generalized variational inclusions involving generalized H(·, ·)- accretive mapping in real q-uniformly smooth Banach spaces. Applied Mathematics and Computation, 217(23):9679–9688, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006229.

Kumar:2012:SMF

[KKS12] Sanjeev Kumar, V. Kanwar, and Sukhjit Singh. On some modified families of multipoint iterative methods for multiple roots of nonlinear equations. Applied Mathematics and Computation, 218(14):7382–7394, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015712.

Khalili:2014:GDH

[KKSS14] Mohammad Khalili, Riyaz Kharrat, Karim Salahshoor, and Morteza Haghighat Sefat. Global dynamic harmony search algorithm: Gdhs. Applied Mathematics and Computation, 228(??):195–219, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012277.

Kaushik:2012:EMA

[KKV12] Aditya Kaushik, Vijayant Kumar, and Anil K. Vashishth. An efficient mixed asymptoticnumerical scheme for singularly perturbed convection diffusion problems. Applied Mathematics and Computation, 218(17):8645–8658, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001488.

Kratica:2012:MDR

[KKVCS12] Jozef Kratica, Vera Kovacević-Vujcić, Mirjana Cangalović, and Milica Stojanović. Minimal doubly resolving sets and the strong metric dimension of some convex polytopes. Applied Mathematics and Computation, 218(19):9790–9801, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002792.

Kougias:2010:EAS

[KL10] I. E. Kougias and S. Louvros. Exact analytic solutions of Schrödinger linear and nonlinear equations. Applied Mathematics and Computation, 216(8):2457–2462, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003450.

Kanovei:2012:IWD

[KL12] Vladimir Kanovei and Vassily Lyubetsky. An infinity which depends on the axiom of choice. Applied Mathematics and Computation, 218(16): 8196–8202, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311006588.

Kuo:2013:DGA

[KL13a] Hsin-Chuan Kuo and Ching-Hai Lin. A Directed Genetic Algorithm for global optimization. Applied Mathematics and Computation, 219(14):7348–7364, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312013227.

Kwiatkowski:2013:GFD

[KL13b]  Grzegorz  Kwiatkowski

and Sergey Leble. Green function diagonal for a class of heat equations. Applied Mathematics and Computation, 219(11):6084–6092, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012325.

Ke:2014:RSI

[KL14a] Jau-Chuan Ke and Tzu-Hsin Liu. A repairable system with imperfect coverage and reboot. Applied Mathematics and Computation, 246(??):148–158, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314010571.

Khiyabani:2014:QNM

[KL14b] Farzin Modarres Khiyabani and Wah June Leong. Quasi-Newton methods based on ordinary differential equation approach for unconstrained nonlinear optimization. Applied Mathematics and Computation, 233(??):272–291, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002240.

Kaur:2013:NSS

[KLBR13] Manwinder Kaur, Arvind Kumar Lal, Satvinder Singh Bhatia, and Akepati Sivarami Reddy. Numerical solution of stochastic partial differential difference equation arising in reliability engineering. Applied Mathematics and Computation, 219(14):7645–7652, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000921.

Kleefeld:2012:MBI

[Kle12] A. Kleefeld. A modified boundary integral equation for solving the exterior Robin problem for the Helmholtz equation in three dimensions. Applied Mathematics and Computation, 219(4):2114–2123, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200848X.

Kim:2014:ILC

[KLKA14] Byeong-Yeon Kim, Taekyung Lee, Young-Soo Kim, and Hyo-Sung Ahn. Iterative learning control for spatially interconnected systems. Applied Mathematics and Computation, 237(??):438–445, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004949.

Kwon:2012:NAS

[KLPC12] O. M. Kwon, S. M. Lee, Ju H. Park, and E. J. Cha. New approaches on stability criteria for neural networks with interval timevarying delays. Applied Mathematics and Computation, 218(19):9953–9964, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003402.

Kao:2014:HLP

[KLW14a] Wen-Hung Kao, Yuh-Dauh Lyuu, and Kuo-Wei Wen. The hexanomial lattice for pricing multiasset options. Applied Mathematics and Computation, 233(??):463–479, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002410.

Krause:2014:DNA

[KLW14b] Andrew Krause, Michael Lewis, and Bixiang Wang. Dynamics of the nonautonomous stochastic p-Laplace equation driven by multiplicative noise. Applied Mathematics and Computation, 246(??): 365–376, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314011278.

Kwon:2012:OCA

[KLY12] Hee-Dae Kwon, Jeehyun Lee, and Sung-Dae Yang.

Optimal control of an agestructured model of HIV infection. Applied Mathematics and Computation, 219(5):2766–2779, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009022.

Kryazhimskiy:2010:RSI

[KM10] Arkady Kryazhimskiy and Vyacheslav Maksimov. Resource-saving infinitehorizon tracking under uncertain input. Applied Mathematics and Computation, 217(3):1135–1140, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000275.

Kairi:2011:EVD

[KM11a] R. R. Kairi and P. V. S. N. Murthy. Effect of viscous dissipation on natural convection heat and mass transfer from vertical cone in a non-Newtonian fluid saturated non-Darcy porous medium. Applied Mathematics and Computation, 217(20):8100–8114, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003523.

Kapica:2011:MRT

[KM11b] Rafał Kapica and Janusz Morawiec. Matrix refinement type equations. Applied Mathematics and Computation, 217(21): 8311–8317, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311003547.

Kirane:2011:DUS

[KM11c] Mokhtar Kirane and Salman A.

Malik. Determination of an unknown source term and the temperature distribution for the linear heat equation involving fractional derivative in time. Applied Mathematics and Computation, 218(1):163–170, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100765X.

Kapica:2012:RED

[KM12a] Rafał Kapica and Janusz Morawiec. Refinement equations and distributional fixed points. Applied Mathematics and Computation, 218(15):7741–7746, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000719.

Karoui:2012:ABN

[KM12b] Abderrazek Karoui and Issam Mehrzi. Asymptotic behaviors and numerical computations of the eigenfunctions and eigenvalues associated with the classical and circular prolate spheroidal wave functions. Applied Mathematics and Computation, 218(22):10871–10888, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200447X.

Kusano:2012:PSF

[KM12c] Takasi Kusano and Jelena V. Manojlović. Positive solutions of fourth order Emden–Fowler type differential equations in the framework of regular variation. Applied Mathematics and Computation, 218(12):6684–6701, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015037.

Karamanli:2013:SFM

[KM13a] A. Karamanli and A. Mugan. Strong form meshless implementation of Taylor series method. Applied Mathematics and Computation, 219(17):9069–9080, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002919.

Katsouleas:2013:DIN

[KM13b] Georgios Katsouleas and John Maroulas. Diagonal imbeddings in a normal matrix. Applied Mathematics and Computation, 219(12):7048–7055, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300012X.

Kikonko:2013:NDS

[KM13c] Mervis Kikonko and Angelo B. Mingarelli. On non-definite Sturm–Liouville problems with two turning points. Applied Mathematics and Computation, 219(17):9508–9515, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002786.

Katsouleas:2014:IBH

[KM14a] Georgios Katsouleas and John Maroulas. Interaction between Hermitian and normal imbeddings. Applied Mathematics and Computation, 233(??):557–564, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002550.

Ke:2014:CAT

[KM14b] Yi-Fen Ke and Chang-Feng Ma. On the convergence analysis of two-step modulus-based matrix splitting iteration method for linear complementarity problems. Applied Mathematics and Computation, 243(??):413–418, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008194.

Kou:2014:ABQ

[KM14c] Kit Ian Kou and João Morais. Asymptotic behaviour of the quaternion linear canonical transform and the Bochner–Minlos theorem. Applied Mathematics and Computation, 247(??):675–688, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012004.

Kim:2012:RMI

[KMAG12] Jeongho Kim, Mehrdad Massoudi, James F. Antaki, and Alberto Gandini. Removal of malariainfected red blood cells using magnetic cell separators: a computational study. Applied Mathematics and Computation, 218(12):6841–6850, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101544X.

Ke:2012:MSP

[KMC12] Hua Ke, Weimin Ma, and Xiaowei Chen. Modeling stochastic project timecost trade-offs with timedependent activity durations. Applied Mathematics and Computation, 218(18):9462–9469, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002676.

Karban:2013:MRI

[KMD13] Pavel Karban, Frantisek Mach, and Ivo Dolezel. Modeling of rotational induction heating of nonmagnetic cylindrical billets. Applied Mathematics and Computation, 219(13):7170–7180, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010083.

Kuraz:2013:DPV

[KMH^{+}13] Michal Kuráz, Petr Mayer, Vojtech Havlícek, Pavel Pech, and Jirka Pavlásek. Dual permeability variably saturated flow and contaminant transport modeling of a nuclear waste repository with capillary barrier protection. Applied Mathematics and Computation, 219(13):7127–7138, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011507.

Kuppusamy:2014:THC

[KML14] Navin Raja Kuppusamy, H. A. Mohammed, and C. W. Lim. Thermal and hydraulic characteristics of nanofluid in a triangular grooved microchannel heat sink (TGM-CHS). Applied Mathematics and Computation, 246(??):168–183, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010546.

Kaya:2014:CNA

[KMR14] Songul Kaya, Carolina C.

Manica, and Leo G. Rebholz. On Crank–Nicolson Adams–Bashforth timestepping for approximate deconvolution models in two dimensions. Applied Mathematics and Computation, 246(??):23–38, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010698.

Kang:2012:HGA

[KMW12] Kyungdaw Kang, Ilkyeong Moon, and Hongfeng Wang. A hybrid genetic algorithm with a new packing strategy for the threedimensional bin packing problem. Applied Mathematics and Computation, 219(3):1287–1299, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007369.

Kuniya:2012:PEN

[KN12] Toshikazu Kuniya and Yukihiko Nakata. Permanence and extinction for a nonautonomous SEIRS epidemic model. Applied Mathematics and Computation, 218(18):9321–9331, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002433.

Kozera:2013:PQE

[KN13] R. Kozera and L. Noakes. Piecewise-quadratics and exponential parameterization for reduced data. Applied Mathematics and Computation, 221(??): 620–638, September 15, 2013. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313006991.

Kalmar-Nagy:2014:CCS

[KNS14] Tamas Kalmar-Nagy and Ilinca Stanciulescu. Can complex systems really be simulated? Applied Mathematics and Computation, 227(??):199–211, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012058.

Kaanoglu:2011:NFG

[KÖ11] Cem Kaanoğlu and Mehmet Ali Özarslan. New families of generating functions for certain class of three-variable polynomials. Applied Mathematics and Computation, 218(3):836–842, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001184.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ko:2013:RRB

[Ko13]  Chia-Nan Ko.  Re-

inforcement radial basis function neural networks with an adaptive annealing learning algorithm. Applied Mathematics and Computation, 221(??):503–513, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007443.

Kocak:2011:SDI

[Koç11] Mehmet Çetin Koçak. Second derivative of an iterative solver boosts its acceleration by Koçak’s method. Applied Mathematics and Computation, 218(3):893–898, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006163.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Kocak:2013:KMS

[Koç13] Mehmet Çetin Koçak. Koçak’s method shows that an unwittingly exaggerated convergence order is in fact 2. Applied Mathematics and Computation, 219(19):10053– 10058, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313002877.

Kocak:2014:KDL

[Koç14] Mehmet Çetin Koçak.

Koçak’s double linearization loop renders a onepoint, one-go nonlinear equation solver. Applied Mathematics and Computation, 248(??):494–502, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314013551.

Kohaupt:2012:NLM

[Koh12] L. Kohaupt. A Newtonlike method for the numerical solution of nonlinear Fredholm-type operator equations. Applied Mathematics and Computation, 218(20):10129– 10148, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003529.

Kohaupt:2013:VSP

[Koh13]  L. Kohaupt.  On the

vibration-suppression property and monotonicity behavior of a special weighted norm for dynamical systems ẋ = Ax, x(t_{0}) = x_{0}. Applied Mathematics and Computation, 222(??):307–330, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007406.

Kojima:2013:NAI

[Koj13]  Kentaro Kojima.  A

new axiom and interrelationships among coalition values for games in characteristic function form. Applied Mathematics and Computation, 224(??):705–713, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009594.

Kojima:2014:CPC

[Koj14a] Kentaro Kojima. Collusion properties of coalition values for games in characteristic function form. Applied Mathematics and Computation, 238(??):57– 69, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S009630031400527X.

Kojima:2014:CBI

[Koj14b]  Kentaro Kojima.  On

connections between individual values and coalition values for games in characteristic function form. Applied Mathematics and Computation, 229(??):60–69, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013143.

Kolev:2014:PSL

[Kol14a] Lubomir V. Kolev. Parameterized solution of linear interval parametric systems. Applied Mathematics and Computation, 246(??):229–246, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401131X.

Kolev:2014:RRR

[Kol14b] Lubomir V. Kolev. Regularity radius and real eigenvalue range. Applied Mathematics and Computation, 233(??):404–412, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001581.

Kong:2014:HSS

[Kon14] Lingju Kong. Homoclinic solutions for a second order difference equation with p-Laplacian. Applied Mathematics and Computation, 247(??):1113–1121, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012946.

Kanas:2011:SDT

[KOP11] S. Kanas and J. Orlowska-Puzio. Sharp distortion theorems for some subclass of close-to-convex functions. Applied Mathematics and Computation, 218(3):843–850, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012580.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Kosinski:2011:ESC

[Kos11a] Paweł Kosinski. Explosion suppression by a cloud of particles: Numerical analysis of the initial processes. Applied Mathematics and Computation, 217(11):5087–5094, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008337.

Kostoglou:2011:EAN

[Kos11b] M. Kostoglou. The equimaterial approach for the numerical solution of the one dimensional transient diffusion equation with zero flux boundary conditions. Applied Mathematics and Computation, 218(4):1353–1359, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008319.

Kostek:2013:DNM

[Kos13]  Robert Kostek.  Di-

rect numerical methods dedicated to secondorder ordinary differential equations. Applied Mathematics and Computation, 219(19):10082– 10095, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001409.

Kotsios:2014:AMC

[Kot14] Stelios Kotsios. An algorithmic method for checking global asymptotic stability of nonlinear polynomial systems with parameters. Applied Mathematics and Computation, 240(??):358–367, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006080.

Kovacic:2010:IAS

[Kov10] Ivana Kovacic. Invariants and approximate solutions for certain non-linear oscillators by means of the field method. Applied Mathematics and Computation, 215(9):3482–3487, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009308.

Koyuncu:2012:EEP

[Koy12] Nursel Koyuncu. Efficient estimators of population mean using auxiliary attributes. Applied Mathematics and Computation, 218(22):10900–10905, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004493.

Krabs:2010:OCP

[KP10] Werner Krabs and Stefan Pickl. An optimal control problem in cancer chemotherapy. Applied Mathematics and Computation, 217(3):1117–1124, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005278.

Kwak:2011:MFE

[KP11a] Do Y. Kwak and Hyun Chan Pyo. Mixed finite element methods for general quadrilateral grids. Applied Mathematics and Computation, 217(14): 6556–6565, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031100052X.

Kyritsi:2011:BTL

[KP11b] Sophia Th. Kyritsi and Nikolaos S. Papageorgiou. A bifurcation theorem for the p-logistic equation. Applied Mathematics and Computation, 217(18):7504–7508, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002475.

Katsikis:2012:CVS

[KP12] Vasilios N. Katsikis and Ioannis A. Polyrakis. Computation of vector sublattices and minimal lattice-subspaces of R^{k}: Applications in finance. Applied Mathematics and Computation, 218(12):6860–6873, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015499.

Kim:2013:ISB

[KP13] Dae San Kim and Kyoung Ho Park. Identities of symmetry for Bernoulli polynomials arising from quotients of Volkenborn integrals invariant under S_{3}. Applied Mathematics and Computation, 219(10):5096–5104, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012283.

Kacewicz:2014:OAS

[KP14a] Bolesław Kacewicz and Paweł Przybyłowicz. Optimal adaptive solution of piecewise regular systems of IVPs with unknown switching hypersurface. Applied Mathematics and Computation, 228(??):116–127, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012551.

Kazemzadeh-Parsi:2014:NFS

[KP14b] Mohammad Javad Kazemzadeh-Parsi. Numerical flow simulation in gated hydraulic structures using smoothed fixed grid finite element method. Applied Mathematics and Computation, 246(??):447–459, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011102.

Kumar:2014:EPP

[KPBC14] Naveen Kumar, Vikas Panwar, Jin-Hwan Borm, and Jangbom Chai. Enhancing precision performance of trajectory tracking controller for robot manipulators using RBFNN and adaptive bound. Applied Mathematics and Computation, 231(??):320–328, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013726.

Kalyani:2014:FDT

[KPC14] V. K. Kalyani, Pallavika, and S. K. Chakraborty. Finite-difference timedomain method for modelling of seismic wave propagation in viscoelastic media. Applied Mathematics and Computation, 237(??):133–145, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003828.

Kumam:2014:SSE

[KPK14] Wiyada Kumam, Hossein Piri, and Poom Kumam. Solutions of system of equilibrium and variational inequality problems on fixed points of infinite family of nonexpansive mappings. Applied Mathematics and Computation, 248(??):441–455, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013538.

Kwon:2011:NAL

[KPLC11] O. M. Kwon, Ju H.

Park, S. M. Lee, and E. J. Cha. A new augmented Lyapunov– Krasovskii functional approach to exponential passivity for neural networks with time-varying delays. Applied Mathematics and Computation, 217(24):10231–10238, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100676X.

Katsikis:2011:IMC

[KPP11] Vasilios N. Katsikis, Dimitrios Pappas, and Athanassios Petralias. An improved method for the computation of the Moore–Penrose inverse matrix. Applied Mathematics and Computation, 217(23):9828–9834, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006503.

Kwon:2013:ARI

[KPP^{+}13] O. M. Kwon, M. J. Park, Ju H. Park, S. M. Lee, and E. J. Cha. Analysis on robust H_{∞} performance and stability for linear systems with interval time-varying state delays via some new augmented Lyapunov– Krasovskii functional. Applied Mathematics and Computation, 224(??): 108–122, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009363.

Kwon:2014:NIR

[KPP^{+}14] O. M. Kwon, M. J. Park, Ju H. Park, S. M. Lee, and E. J. Cha. New and improved results on stability of static neural networks with interval timevarying delays. Applied Mathematics and Computation, 239(??):346–357, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006298.

Karsli:2011:RCB

[KPT11] Harun Karsli and Paulina Pych-Taberska. On the rates of convergence of BBH-Kantorovich operators and their Bézier variant. Applied Mathematics and Computation, 218(6):2960–2967, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101099X.

Kumam:2011:EIA

[KPW11] Poom Kumam, Narin Petrot, and Rabian Wangkeeree. Existence and iterative approximation of solutions of generalized mixed quasi-variational-like inequality problem in Banach spaces. Applied Mathematics and Computation, 217(18):7496–7503, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002451.

Kumar:2010:HOP

[KR10] Mukesh Kumar and S. Chandra Sekhara Rao. High order parameter-robust numerical method for singularly perturbed reaction– diffusion problems. Applied Mathematics and Computation, 216(4):1036– 1046, April 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310001451.

Kleefeld:2011:LMM

[KR11] A. Kleefeld and M. Reißel. The Levenberg–Marquardt method applied to a parameter estimation problem arising from electrical resistivity tomography. Applied Mathematics and Computation, 217(9):4490–4501, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011124.

Kazmi:2012:HEM

[KR12] K. R. Kazmi and S. H. Rizvi. A hybrid extragradient method for approximating the common solutions of a variational inequality, a system of variational inequalities, a mixed equilibrium problem and a fixed point problem. Applied Mathematics and Computation, 218(9):5439–5452, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013622.

Kadyan:2013:CBA

[KR13a] M. S. Kadyan and Ramniwas. Cost benefit analysis of a single-unit system with warranty for repair. Applied Mathematics and Computation, 223(??):346–353, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008539.

Kazemi:2013:MGL

[KR13b] P. Kazemi and R. J.

Renka. Minimization of the Ginzburg–Landau energy functional by a Sobolev gradient trustregion method. Applied Mathematics and Computation, 219(11):5936–5942, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013148.

Khan:2013:IAP

[KR13c] Subuhi Khan and Nusrat Raza. 2-iterated Appell polynomials and related numbers. Applied Mathematics and Computation, 219(17):9469–9483, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003433.

Kudryashov:2014:ESO

[KR14] Nikolay A. Kudryashov and Pavel N. Ryabov. Exact solutions of one pattern formation model. Applied Mathematics and Computation, 232(??): 1090–1093, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314001179.

Krasniqi:2014:CSC

[Kra14a] Valmir Krasniqi. Corrigendum to “Some complete monotonicity properties for the (p, q)-gamma function”. Applied Mathematics and Computation, 227(??):662, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012526.

See [KSD13].

Krasniqi:2014:AMS

[Kra14b] Xhevat Z. Krasniqi. On absolute matrix summability of orthogonal series. Applied Mathematics and Computation, 230(??):711–715, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014458.

Kovacic:2010:NSV

[KRC10] Ivana Kovacic, Zvonko Rakaric, and Livija Cveticanin. A non-simultaneous variational approach for the oscillators with fractionalorder power nonlinearities. Applied Mathematics and Computation, 217(8):3944–3954, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010180.

Krech:2013:NPS

[Kre13] Grazyna Krech. A note on the paper “Voronovskaya type asymptotic approximation by modified Gamma operators”. Applied Mathematics and Computation, 219(11):5787–5791, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012933.

See [Izg11].

Krivec:2014:UQR

[Kri14]  R. Krivec.  Unified

QLM for regular and arbitrary singular potentials. Applied Mathematics and Computation, 246(??):440–446, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011606.

Kanovic:2011:GPS

[KRJ11] Zeljko Kanović, Milan R. Rapaić, and Zoran D. Jelicić. Generalized particle swarm optimization algorithm — Theoretical and empirical analysis with application in fault detection. Applied Mathematics and Computation, 217(24):10175–10186, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006680.

Kashani:2014:EGN

[KRMM14] S. Kashani, A. A. Ranjbar, M. Mastiani, and H. Mirzaei. Entropy generation and natural convection of nanoparticlewater mixture (nanofluid) near water density inversion in an enclosure with various patterns of vertical wavy walls. Applied Mathematics and Computation, 226(??):180–193, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011119.

Krol:2011:APF

[Kro11] Katja Krol. Asymptotic properties of fractional delay differential equations. Applied Mathematics and Computation, 218(5):1515–1532, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006291.

Kudryashov:2010:NNK

[KRS10] Nikolay A. Kudryashov, Pavel N. Ryabov, and Dmitry I. Sinelshchikov. A note on “New kinkshaped solutions and periodic wave solutions for the (2 + 1)-dimensional Sine-Gordon equation”. Applied Mathematics and Computation, 216(8):2479– 2481, June 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031000322X.

See [Li10d].

Khuri:2010:SCAb

[KS10a] S. A. Khuri and A. Sayfy. A spline collocation approach for a generalized wave equation subject to non-local conservation condition. Applied Mathematics and Computation, 217(8):3993–4001, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010234.

Khuri:2010:SCAa

[KS10b] S. A. Khuri and A. Sayfy. A spline collocation approach for the numerical solution of a generalized nonlinear Klein– Gordon equation. Applied Mathematics and Computation, 216(4):1047–1056, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001463.

Kudryashov:2010:NEE

[KS10c] Nikolai A. Kudryashov and Dmitry I. Sinelshchikov. Nonlinear evolution equations for describing waves in bubbly liquids with viscosity and heat transfer consideration. Applied Mathematics and Computation, 217(1):414–421, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310005771.

Kaslik:2011:MPS

[KS11a] Eva Kaslik and Seenith Sivasundaram. Multiple periodic solutions in impulsive hybrid neural networks with delays. Applied Mathematics and Computation, 217(10):4890–4899, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011379.

Kheloufi:2011:RHE

[KS11b] Arezki Kheloufi and Boubaker-Khaled Sadallah. On the regularity of the heat equation solution in noncylindrical domains: Two approaches. Applied Mathematics and Computation, 218(5):1623–1633, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008757.

Kosinski:2011:P

[KS11c] P. Kosinski and T. Simos. Preface. Applied Mathematics and Computation, 217(11):5015, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012932.

Kropielnicka:2011:ESD

[KS11d] K. Kropielnicka and L. Sapa. Estimate of solutions for differential and difference functional equations with applications to difference methods. Applied Mathematics and Computation, 217(13):6206–6218, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013238.

Kunter:2011:RSW

[KS11e] Fulya C. Kunter and S. Selim Seker. Radially symmetric weighted extended B-spline model. Applied Mathematics and Computation, 217(24):10305– 10316, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031100703X.

Kurt:2011:NGB

[KS11f] Burak Kurt and Yilmaz Simsek. Notes on generalization of the Bernoulli type polynomials. Applied Mathematics and Computation, 218(3):906–911, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004541.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Khuri:2012:SCA

[KS12a] S. A. Khuri and A. Sayfy.

A spline collocation approach for a generalized parabolic problem subject to non-classical conditions. Applied Mathematics and Computation, 218(18):9187–9196, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002251.

Kosinski:2012:SIN

[KS12b] Paweł Kosinski and Theodore E. Simos. Special issue on Numerical Analysis of Fluid Flow and Heat Transfer: Preface. Applied Mathematics and Computation, 219(7):3291, December 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008958.

Kudryashov:2012:ESF

[KS12c] Nikolai A. Kudryashov and Dmitry I. Sinelshchikov. Elliptic solutions for a family of fifth order nonlinear evolution equations. Applied Mathematics and Computation, 218(12):6991–6997, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015451.

Kudryashov:2012:NDE

[KS12d] Nikolai A. Kudryashov and Dmitry I. Sinelshchikov.

Nonlinear differential equations of the second, third and fourth order with exact solutions. Applied Mathematics and Computation, 218(21):10454– 10467, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031200392X.

Ketabdari:2013:PSO

[KS13a] M. J. Ketabdari and H. Saghi. Parametric study for optimization of storage tanks considering sloshing phenomenon using coupled BEM– FEM. Applied Mathematics and Computation, 224(??):123–139, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009041.

Khandelwal:2013:PSS

[KS13b] Pooja Khandelwal and Talat Sultana. Parametric septic splines approach for the solution of linear sixth-order two-point boundary value problems. Applied Mathematics and Computation, 219(12):6856–6867, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000465.

Kucerova:2013:UUD

[KS13c] Anna Kucerová and Jan Sýkora. Uncertainty updating in the description of coupled heat and moisture transport in heterogeneous materials. Applied Mathematics and Computation, 219(13):7252–7261, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003018.

Kumar:2013:UMF

[KS13d] Anand Kumar and A. K. Singh. Unsteady MHD free convective flow past a semi-infinite vertical wall with induced magnetic field. Applied Mathematics and Computation, 222(??):462–471, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007947.

Krasniqi:2013:SCM

[KSD13] V. B. Krasniqi, H. M.

Srivastava, and S. S. Dragomir. Some complete monotonicity properties for the (p, q)- gamma function. Applied Mathematics and Computation, 219(21):10538– 10547, July 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004499.

See corrigendum [Kra14a].

Khan:2014:NCM

[KSFG14] Yasir Khan, K. Sayevand, M. Fardi, and M. Ghasemi. A novel computing multiparametric homotopy approach for system of linear and nonlinear Fredholm integral equations. Applied Mathematics and Computation, 249(??):229–236, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314014453.

Komuro:2011:ESA

[KSM11] Naoto Komuro, Kichi-Suke Saito, and Ken-Ichi Mitani. Extremal structure of absolute normalized norms on R^{2} and the James constant. Applied Mathematics and Computation, 217(24):10035–10048, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006497.

Kumar:2014:FEA

[KSM14] Manoj Kumar, Akanksha Srivastava, and Garima Mishra. A finite element approach for finding positive solutions of a logistic equation with a sign-changing weight function. Applied Mathematics and Computation, 241(??):112–124, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005839.

Kumar:2012:NSS

[KSS12] Manoj Kumar, Akanksha Srivastava, and Akhilesh Kumar Singh. Numerical simulation of singularly perturbed non-linear elliptic boundary value problems using finite element method. Applied Mathematics and Computation, 219(1):226–236, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005942.

Kucuk:2012:ACS

[KSY12] Ismail Kucuk, Ibrahim Sadek, and Yalcin Yilmaz. Active control of a smart beam with time delay by Legendre wavelets. Applied Mathematics and Computation, 218(17):8968–8977, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200197X.

Kravchenko:2014:MSP

[KT14] Vladislav V. Kravchenko and Sergii M. Torba. Modified spectral parameter power series representations for solutions of Sturm–Liouville equations and their applications. Applied Mathematics and Computation, 238(??):82– 105, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005347.

Kuang:2011:ASN

[KTS11] Jiaoxun Kuang, Hongjiong Tian, and Kaiting Shan. Asymptotic stability of neutral differential systems with many delays. Applied Mathematics and Computation, 217(24):10087–10094, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100659X.

Kuzmanovic:2012:OMM

[KTT12] Ivana Kuzmanović, Zoran Tomljanović, and Ninoslav Truhar. Optimization of material with modal damping. Applied Mathematics and Computation, 218(13):7326–7338, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000306.

Krasovskii:2010:OFD

[KTW10] Andrey A. Krasovskii, Alexander M. Tarasyev, and Chihiro Watanabe. Optimization of functionality development. Applied Mathematics and Computation, 217(3):1125–1134, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310003231.

Kapcak:2011:GAF

[KU11] Sinan Kapcak and Ünal Ufuktepe. Generalization of analytic functions. Applied Mathematics and Computation, 218(3):851–855, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003791.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Kudryashov:2010:NEM

[Kud10a] Nikolai A. Kudryashov. A note on the \frac{G′}{G} -expansion method. Applied Mathematics and Computation, 217(4):1755–1758, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003255.

Kudryashov:2010:SRS

[Kud10b] Nikolai A. Kudryashov.

Soliton, rational and special solutions of the Korteweg–de Vries hierarchy. Applied Mathematics and Computation, 217(4):1774–1779, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010133.

Kudryashov:2010:NNE

[Kud10c] Nikolay A. Kudryashov. A new note on exact complex travelling wave solutions for (2 + 1)-dimensional B-type Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 217(5):2282–2284, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007241.

Kudryashov:2012:SPA

[Kud12] Nikolay A. Kudryashov. Special polynomials associated with the Burgers hierarchy. Applied Mathematics and Computation, 218(15):7972–7976, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001580.

Kudryashov:2013:PLF

[Kud13a] Nikolay A. Kudryashov.

Polynomials in logistic function and solitary waves of nonlinear differential equations. Applied Mathematics and Computation, 219(17):9245–9253, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002853.

Kudryashov:2013:QES

[Kud13b] Nikolay A. Kudryashov.

Quasi-exact solutions of the dissipative Kuramoto– Sivashinsky equation. Applied Mathematics and Computation, 219(17): 9213–9218, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003159.

Khan:2014:CGA

[KudKL14] A. R. Khan, H. Fukhar ud din, A. Kalsoom, and B. S. Lee. Convergence of a general algorithm of asymptotically nonexpansive maps in uniformly convex hyperbolic spaces. Applied Mathematics and Computation, 238(??):547–556, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005281.

Kujan:2014:SCS

[Kuj14]  Petr Kujan.  Solving

composite sum of powers via Padé approximation and orthogonal polynomials with application to optimal PWM problem. Applied Mathematics and Computation, 236(??):593–605, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004433.

Khan:2014:RMS

[KUKU14] Mohammad Asmat Ullah Khan, Khalil Ullah, Asif Khan, and Ihtesham Ul Islam. Robust multi-scale orientation estimation: Directional filter bank based approach. Applied Mathematics and Computation, 242(??):814–824, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009151.

Kumar:2011:RSE

[Kum11] R. Ravi Kumar. Regularity of solutions of evolution integrodifferential equations with deviating argument. Applied Mathematics and Computation, 217(22):9111–9121, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005169.

Kumar:2014:LAM

[Kum14]  Sunil Kumar.  Layer-

adapted methods for quasilinear singularly perturbed delay differential problems. Applied Mathematics and Computation, 233(??):214–221, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002343.

Kaya:2014:LBP

[KUTY14] Yilmaz Kaya, Murat Uyar, Ramazan Tekin, and Selçuk Yildirim. 1D-local binary pattern based feature extraction for classification of epileptic EEG signals. Applied Mathematics and Computation, 243(??):209–219, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008285.

Kicsiny:2012:RTS

[KV12] R. Kicsiny and Z. Varga. Real-time state observer design for solar thermal heating systems. Applied Mathematics and Computation, 218(23):11558– 11569, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005486.

Koleva:2013:SOP

[KV13] Miglena N. Koleva and Lubin G. Vulkov. A secondorder positivity preserving numerical method for Gamma equation. Applied Mathematics and Computation, 220(??):722–734, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007315.

Krnic:2014:MAC

[KV14] Mario Krnić and Predrag Vuković. On a more accurate class of discrete Hilbert-type inequalities. Applied Mathematics and Computation, 234(??):543–547, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002987.

Kostic:2014:TDR

[KVP14] Srdan Kostić, Nebojsa Vasović, and Matjaz Perc. Temporal distribution of recorded magnitudes in Serbia earthquake catalog. Applied Mathematics and Computation, 244(??):917–924, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010200.

Keller:2010:CMI

[KW10] Paweł Keller and Paweł Woźny. On the convergence of the method for indefinite integration of oscillatory and singular functions. Applied Mathematics and Computation, 216(3):989–998, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310001414.

Kol:2013:ESS

[KW13] Guy Richard Kol and Paul Woafo. Exact solutions for a system of two coupled discrete nonlinear Schrödinger equations with a saturable nonlinearity. Applied Mathematics and Computation, 219(11):5956–5962, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012040.

Kang:2013:GCC

[KWJ13] Dongyeop Kang, Sangchul Won, and Yu Jin Jang. Guaranteed cost control for bilinear systems by static output feedback. Applied Mathematics and Computation, 219(14):7398–7405, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000556.

Kong:2013:VIF

[KWM13] Xinlei Kong, Huibin Wu, and Fengxiang Mei. Variational integrators for forced Birkhoffian systems. Applied Mathematics and Computation, 225(??):326–332, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010242.

Kwon:2011:SPS

[Kwo11] Oh Sang Kwon. A subclass of parabolic starlike and uniformly convex functions. Applied Mathematics and Computation, 218(3):912–918, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001676.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Kwon:2012:TDM

[Kwo12] Soon-Geol Kwon. Twodirection multiwavelet moments. Applied Mathematics and Computation, 219(8):3530–3540, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009332.

Khan:2010:MRS

[KWS10] Israr Ali Khan, Qing-Wen Wang, and Guang-Jing Song. Minimal ranks of some quaternion matrix expressions with applications. Applied Mathematics and Computation, 217(5):2031–2040, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007496.

Kang:2010:CHO

[KX10] Hongchao Kang and Shuhuang Xiang. On the calculation of highly oscillatory integrals with an algebraic singularity. Applied Mathematics and Computation, 217(8):3890–3897, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010106.

Kang:2011:EIC

[KX11] Hongchao Kang and Shuhuang Xiang. Efficient integration for a class of highly oscillatory integrals. Applied Mathematics and Computation, 218(7):3553–3564, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011428.

Kong:2014:PLD

[KX14] Yuedong Kong and Junxiang Xu. Persistence of lower dimensional hyperbolic tori for reversible system. Applied Mathematics and Computation, 236(??):408–421, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004184.

Kao:2014:SMD

[KXW14] Y. G. Kao, J. Xie, and C. H. Wang. Stabilisation of mode-dependent singular Markovian jump systems with generally uncertain transition rates. Applied Mathematics and Computation, 245(??): 243–254, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314009072.

Koprubasi:2014:QED

[KY14] Turhan Koprubasi and Nihal Yokus. Quadratic eigenparameter dependent discrete Sturm–Liouville equations with spectral singularities. Applied Mathematics and Computation, 244(??):57–62, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009163.

Kycia:2011:SSS

[Kyc11]  Radosław Kycia.  On

self-similar solutions of semilinear wave equations in higher space dimensions. Applied Mathematics and Computation, 217(22):9451–9466, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005960.

Khan:2010:SRI

[KYH10] Subuhi Khan, Ghazala Yasmin, and Nader Ali Makboul Hassan. Some results involving Hermite-base polynomials and functions using operational methods. Applied Mathematics and Computation, 215(11):3769–3776, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009618.

Kyrchei:2010:CRS

[Kyr10] I. I. Kyrchei. Cramer’s rule for some quaternion matrix equations. Applied Mathematics and Computation, 217(5):2024–2030, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310007484.

Kyrchei:2012:ACR

[Kyr12] Ivan Kyrchei. Analogs of Cramer’s rule for the minimum norm least squares solutions of some matrix equations. Applied Mathematics and Computation, 218(11):6375–6384, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014627.

Kyrchei:2013:EFD

[Kyr13] Ivan Kyrchei. Explicit formulas for determinantal representations of the Drazin inverse solutions of some matrix and differential matrix equations. Applied Mathematics and Computation, 219(14):7632–7644, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000908.

Kyrchei:2014:DRD

[Kyr14] Ivan Kyrchei. Determinantal representations of the Drazin inverse over the quaternion skew field with applications to some matrix equations. Applied Mathematics and Computation, 238(??):193–207, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004962.

Koczan:2011:FCD

[KZ11] Leopold Koczan and Paweł Zaprawa. On functions convex in the direction of the real axis with a fixed second coefficient. Applied Mathematics and Computation, 217(15):6644–6651, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000543.

Koczan:2012:CPF

[KZ12] Leopold Koczan and Paweł Zaprawa. Covering problems for functions n-fold symmetric and convex in the direction of the real axis. Applied Mathematics and Computation, 219(3):947–958, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006984.

Karimi:2014:BPL

[KZ14a] S. Karimi and B. Zali.

The block preconditioned LSQR and GL–LSQR algorithms for the block partitioned matrices. Applied Mathematics and Computation, 227(??):811–820, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012071.

Kayar:2014:SCL

[KZ14b] Z. Kayar and A. Zafer.

Stability criteria for linear Hamiltonian systems under impulsive perturbations. Applied Mathematics and Computation, 230(??):680–686, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313014185.

Kovac:2014:SNQ

[KZ14c] Bostjan Kovac and Emil Zagar. Some new G^{1} quartic parametric approximants of circular arcs. Applied Mathematics and Computation, 239(??):254–264, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006407.

Kang:2013:SSB

[KZJ13] Yingli Kang, Yi Zhang, and Ligang Jin. Soliton solution to BKP equation in Wronskian form. Applied Mathematics and Computation, 224(??):250–258, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009533.

Kuo:2012:HPS

[KZS12] R. J. Kuo, Ferani E. Zulvia, and Kadarsah Suryadi. Hybrid particle swarm optimization with genetic algorithm for solving capacitated vehicle routing problem with fuzzy demand — a case study on garbage collection system. Applied Mathematics and Computation, 219(5):2574–2588, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008855.

LaCruz:2010:RMS

[La 10] William La Cruz. A residual method for solving nonlinear operator equations and their application to nonlinear integral equations using symbolic computation. Applied Mathematics and Computation, 217(1):11–24, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000473X.

LaCruz:2013:RSA

[La 13] William La Cruz. Residual spectral algorithm for solving monotone equations on a Hilbert space. Applied Mathematics and Computation, 219(12):6633–6644, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312013355.

Lachiheb:2014:ECR

[Lac14] Mehdi Lachiheb. Effect of coupled radial and axial variability of viscosity on the peristaltic transport of Newtonian fluid. Applied Mathematics and Computation, 244(??):761–771, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010194.

Li:2013:USI

[LAF13] Xiaodi Li, Haydar Akca, and Xillin Fu. Uniform stability of impulsive infinite delay differential equations with applications to systems with integral impulsive conditions. Applied Mathematics and Computation, 219(14):7329–7337, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312013094.

Lai:2010:EUL

[Lai10] Shaoyong Lai. The existence and uniqueness of the local solution for a Camassa–Holm type equation. Applied Mathematics and Computation, 216(4):1287–1298, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000189X.

Lakzian:2013:FPW

[LAR13] Hossein Lakzian, Hassen Aydi, and B. E. Rhoades. Fixed points for (ϕ, ψ, p)- weakly contractive mappings in metric spaces with w-distance. Applied Mathematics and Computation, 219(12):6777–6782, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011794.

Latif:2014:BDS

[Lat14] M. S. Abdel Latif. Bright and dark soliton solutions for the perturbed nonlinear Schrödinger’s equation with Kerr law and non-Kerr law nonlinearity. Applied Mathematics and Computation, 247(??):501–510, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012089.

Lazo:2011:GSQ

[Laz11] Pedro Pablo Durand Lazo. Global solution for a quasilinear wave equation with singular memory. Applied Mathematics and Computation, 217(19):7660–7668, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002761.

Liu:2014:SDP

[LB14a] Meng Liu and Chuan Zhi Bai. On a stochastic delayed predator-prey model with Lévy jumps. Applied Mathematics and Computation, 228(??):563–570, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013027.

Liu:2014:RSL

[LB14b] Meng Liu and Chuan Zhi Bai. A remark on stochastic logistic model with diffusion. Applied Mathematics and Computation, 228(??):141–146, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012368.

Liu:2014:GAS

[LB14c] Meng Liu and Chuanzhi Bai. Global asymptotic stability of a stochastic delayed predator-prey model with Beddington– DeAngelis functional response. Applied Mathematics and Computation, 226(??):581–588, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011090.

Liu:2014:OHP

[LB14d] Meng Liu and Chuanzhi Bai. Optimal harvesting policy of a stochastic food chain population model. Applied Mathematics and Computation, 245(??):265–270, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314010704.

Lu:2011:FSP

[LC11] Xin Lu and Martin Camitz. Finding the shortest paths by node combination. Applied Mathematics and Computation, 217(13):6401–6408, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000336.

Lin:2012:REC

[LC12] Yi-Kuei Lin and Ping-Chen Chang. Reliability evaluation of a computer network in cloud computing environment subject to maintenance budget. Applied Mathematics and Computation, 219(8):3893–3902, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010259.

Lin:2013:STN

[LC13] Matthew M. Lin and Chun-Yueh Chiang. The shift techniques for a nonsymmetric algebraic Riccati equation. Applied Mathematics and Computation, 219(10):5083–5095, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201171X.

Li:2014:HOC

[LC14a] Maojun Li and Aimin Chen. High order central discontinuous Galerkinfinite element methods for the Camassa–Holm equation. Applied Mathematics and Computation, 227(??):237–245, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011831.

Liu:2014:HPD

[LC14b] Aijing Liu and Guoliang Chen. On the Hermitian positive definite solutions of nonlinear matrix equation X^{s} + ∑_{i=1}^{m} A^{i}^{∗}X^{t_{i}}A_{i} = Q. Applied Mathematics and Computation, 243 (??):950–959, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007905.

Liu:2014:MNP

[LC14c] Li-Bin Liu and Yanping Chen. Maximum norm a posteriori error estimates for a singularly perturbed differential difference equation with small delay. Applied Mathematics and Computation, 227(??):801–810, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011429.

Liu:2014:GST

[LC14d] Luyin Liu and Eric C. K. Cheung. On a Gerber– Shiu type function and its applications in a dual semi-Markovian risk model. Applied Mathematics and Computation, 247 (??):1183–1201, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012843.

Liu:2014:ASDb

[LC14e] Qun Liu and Qingmei Chen. Analysis of a stochastic delay predatorprey system with jumps in a polluted environment. Applied Mathematics and Computation, 242(??):90–100, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006985.

Liu:2012:BAA

[LCC12] J. K. Liu, F. X. Chen, and Y. M. Chen. Bifurcation analysis of aeroelastic systems with hysteresis by incremental harmonic balance method. Applied Mathematics and Computation, 219(5):2398–2411, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312008272.

Lopera:2011:HNE

[LCCA11] Liliana Garrido Lopera, Edilberto Cepeda-Cuervo, and Jorge Alberto Achcar. Heteroscedastic normalexponential mixture models: Bayesian and classical approaches. Applied Mathematics and Computation, 218(7):3635–3648, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011568.

Lancia:2012:NAT

[LCD12] Maria Rosaria Lancia, Massimo Cefalo, and Guido Dell’Acqua. Numerical approximation of transmission problems across Koch-type highly conductive layers. Applied Mathematics and Computation, 218(9):5453–5473, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013634.

Liu:2014:NCG

[LCF^{+}14] Shuai Liu, Xiaochun Cheng, Weina Fu, Yunpeng Zhou, and Qianzhong Li. Numeric characteristics of generalized M-set with its asymptote. Applied Mathematics and Computation, 243(??):767–774, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008522.

Li:2010:NMP

[LCH10] Xiang-Gui Li, C. K. Chan, and Y. Hou. A numerical method with particle conservation for the Maxwell– Dirac system. Applied Mathematics and Computation, 216(4):1096–1108, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001682.

Liu:2011:GWS

[LCH11] YuQing Liu, DengYuan Chen, and Chao Hu. The generalized Wronskian solutions of a inverse KdV hierarchy. Applied Mathematics and Computation, 218(5):2025–2035, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009337.

Li:2012:GSD

[LCH12] Zhong Li, Fengde Chen, and Mengxin He. Global stability of a delay differential equations model of plankton allelopathy. Applied Mathematics and Computation, 218(13): 7155–7163, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311015736.

Lee:2013:SMB

[LCH13] Wen-Chiung Lee, Yu-Hsiang Chung, and Zong-Ren Huang. A singlemachine bi-criterion scheduling problem with two agents. Applied Mathematics and Computation, 219(23):10831–10841, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313005456.

Cheng:2014:PWS

[lChH14] Zhi long Cheng and Xiao hong Hao. The periodic wave solutions for a (2 + 1)-dimensional AKNS equation. Applied Mathematics and Computation, 234(??):118–126, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001192.

Lien:2013:RFD

[LCL^{+}13a] Chang-Hua Lien, Jenq-Der Chen, Chin-Tan Lee, Ruey-Shin Chen, and Chyi-Da Yang. Robust H_{∞} filter design for discrete-time switched systems with interval timevarying delay and linear fractional perturbations: LMI optimization approach. Applied Mathematics and Computation, 219(24):11395–11407, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005705.

Liu:2013:FPG

[LCL^{+}13b] Shuai Liu, Xiaochun Cheng, Caihe Lan, Weina Fu, Jiantao Zhou, Qianzhong Li, and Guanglai Gao. Fractal property of generalized M-set with rational number exponent. Applied Mathematics and Computation, 220(??):668–675, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313007455.

Liu:2014:ASDa

[LCL14] Qun Liu, Qingmei Chen, and Zhenhai Liu. Analysis on stochastic delay Lotka– Volterra systems driven by Lévy noise. Applied Mathematics and Computation, 235(??):261–271, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314003646.

Lu:2010:NCF

[LCLW10] DaZhao Lü, YanYing Cui, Cheng Lü, and ChuYuan Wei. Novel composite function solutions of the modified KdV equation. Applied Mathematics and Computation, 217(1):283–288, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006181.

Lopez-Diaz:2011:TBS

[LD11] Miguel López-Díaz. A test for the bidirectional stochastic order with an application to quality control theory. Applied Mathematics and Computation, 217(19):7762–7771, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003079.

Lee:2013:MMA

[LD13a] Young S. Lee and Tae Sug Do. A mathematical modeling approach to gambling among older adults. Applied Mathematics and Computation, 221(??):403–410, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006553.

Li:2013:PEM

[LD13b] Junhong Li and Rui Ding. Parameter estimation methods for nonlinear systems. Applied Mathematics and Computation, 219(9):4278–4287, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009447.

Luo:2013:USF

[LD13c] Xuqiong Luo and Qikui Du. An unconditionally stable fourth-order method for telegraph equation based on Hermite interpolation. Applied Mathematics and Computation, 219(15):8237–8246, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001847.

Lu:2014:PEG

[LD14] Chun Lu and Xiaohua Ding. Persistence and extinction in general non-autonomous logistic model with delays and stochastic perturbation. Applied Mathematics and Computation, 229(??):1–15, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013180.

Li:2013:BDR

[LDJZ13] Yajuan Li, Chongyang Deng, Wenbiao Jin, and Nailiang Zhao. On the bounds of the derivative of rational Bézier curves. Applied Mathematics and Computation, 219(21): 10425–10433, July 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004773.

Liu:2010:CCD

[LDL10] Yan Liu, Yi Du, and Changhao Lin. Convergence and continuous dependence results for the Brinkman equations. Applied Mathematics and Computation, 215(12):4443–4455, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010947.

Lin:2011:FTB

[LDL11] Xiangze Lin, Haibo Du, and Shihua Li. Finitetime boundedness and L_{2}- gain analysis for switched delay systems with normbounded disturbance. Applied Mathematics and Computation, 217(12): 5982–5993, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310012348.

Lin:2013:GAS

[LDL13] Xiaojie Lin, Zengji Du, and Yansen Lv. Global asymptotic stability of almost periodic solution for a multispecies competition-predator system with time delays. Applied Mathematics and Computation, 219(9):4908– 4923, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312010995.

Luo:2010:EPS

[LDLL10] Hongying Luo, Mianyi Duan, Xi Liu, and Jun Liu. Explicit periodic solitary wave solutions for the (2 + 1)- dimensional Boussinesq equation. Applied Mathematics and Computation, 217(2):826–829, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007083.

Luo:2013:EDP

[LDLM13] Hongying Luo, Zhengde Dai, Jun Liu, and Gui Mu. Explicit doubly periodic soliton solutions for the (2 + 1)- dimensional Boussinesq equation. Applied Mathematics and Computation, 219(12):6618–6621, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000076.

Lin:2010:MAN

[LDR10] Yuan Lin, Kevin C. Desouza, and Sumit Roy. Measuring agility of networked organizational structures via network entropy and mutual information. Applied Mathematics and Computation, 216(10):2824–2836, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000398X.

Lopes:2010:NAL

[LdS10] B. I. Lima Lopes and A. C. Zambroni de Souza. A Newton approach for long term stability studies in power systems. Applied Mathematics and Computation, 215(9):3327–3334, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009266.

Li:2014:III

[LDS14] Shasha Li, Caixia Deng, and Wei Sun. The isometric identities and inversion formulas of complex continuous wavelet transforms. Applied Mathematics and Computation, 233(??):116–126, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013696.

Liu:2014:BBP

[LDW14] Hao Liu, Guiyan Ding, and Bing Wang. Barebones particle swarm optimization with disruption operator. Applied Mathematics and Computation, 238(??):106–122, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005232.

Li:2014:API

[LDWY14] Hua Li, Lan Di, Antony Ware, and George Yuan. The applications of partial integro-differential equations related to adaptive wavelet collocation methods for viscosity solutions to jump-diffusion models. Applied Mathematics and Computation, 246(??):316–335, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010960.

Luo:2010:EPPa

[LDZ10] Zhenguo Luo, Binxiang Dai, and Qianhong Zhang. Existence of positive periodic solutions for an impulsive semi-ratio-dependent predator–prey model with dispersion and time delays. Applied Mathematics and Computation, 215(9):3390–3398, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009370.

Lee:2011:ALF

[Lee11] S. M. Lee. Augmented Lyapunov function approach to L_{2} gain analysis for discrete-time systems with saturation nonlinearities. Applied Mathematics and Computation, 217(24):10205–10212, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006722.

Lee:2012:NWL

[Lee12]  Hsueh-Chen Lee.  A

nonlinear weighted leastsquares finite element method for the Oldroyd-B viscoelastic flow. Applied Mathematics and Computation, 219(1):421–434, September 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312006352.

Lenski:2011:PAJ

[Łen11] Włodzimierz Łenski. Pointwise analog of the Jackson approximation theorem. Applied Mathematics and Computation, 218(5):1587–1595, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008691.

See corrigendum [Len14b].

Leng:2014:REB

[Len14a] Huinan Leng. Real eigenvalue bounds of standard and generalized real interval eigenvalue problems. Applied Mathematics and Computation, 232(??):164–171, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001076.

Lenski:2014:CPA

[Len14b] Wlodzimierz Lenski. Corrigendum to “Pointwise analog of the Jackson approximation theorem”. Applied Mathematics and Computation, 244(??): 998–999, October 1, 2014.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314010613.

See [Łen11].

Li:2011:SDL

[LF11a] Min Li and Heying Feng.

A sufficient descent LS conjugate gradient method for unconstrained optimization problems. Applied Mathematics and Computation, 218(5):1577– 1586, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311008678.

Li:2011:ESG

[LF11b] Wu Li and Tianyou Fan.

Exact solutions of the generalized Dugdale model of two-dimensional decagonal quasicrystals. Applied Mathematics and Computation, 218(7):3068–3071, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311010514.

Li:2011:MMC

[LF11c] Zhenping Li and Chuli Fu.

A mollification method for a Cauchy problem for the Laplace equation. Applied Mathematics and Computation, 217(22):9209–9218, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005406.

Fan:2012:CST

[lF12] Hui ling Fan. The classification of the single traveling wave solutions to the generalized equal width equation. Applied Mathematics and Computation, 219(2):748–754, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006893.

Luo:2014:OCAb

[LF14] Zhixue Luo and Xueliang Fan. Optimal control for an age-dependent competitive species model in a polluted environment. Applied Mathematics and Computation, 228(??):91–101, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313012381.

Lima:2012:CDU

[LFC^{+}12] G. A. B. Lima, V. G.

Ferreira, E. R. Cirilo, A. Castelo, M. A. C. Candezano, I. V. M. Tasso, D. M. C. Sano, and L. V. A. Scalvi. A continuously differentiable upwinding scheme for the simulation of fluid flow problems. Applied Mathematics and Computation, 218(17):8614–8633, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001464.

Liu:2014:MFE

[LFLH14] Yang Liu, Zhichao Fang, Hong Li, and Siriguleng He. A mixed finite element method for a time-fractional fourthorder partial differential equation. Applied Mathematics and Computation, 243(??):703–717, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008595.

Laudani:2014:SFO

[LFLS14] Antonino Laudani, Francesco Riganti Fulginei, Gabriele Maria Lozito, and Alessandro Salvini. Swarm/flock optimization algorithms as continuous dynamic systems. Applied Mathematics and Computation, 243(??):670–683, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008820.

Li:2014:DDS

[LFR14] Xiaodi Li, Xilin Fu, and R. Rakkiyappan. Delaydependent stability analysis for a class of dynamical systems with leakage delay and nonlinear perturbations. Applied Mathematics and Computation, 226(??):10–19, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010588.

Lu:2013:PDE

[LfXfC13] Xin Lu, Shu fang Xu, and Yun feng Cai. Partial derivatives of the eigentriplet of the quadratic eigenvalue problem depending on several parameters. Applied Mathematics and Computation, 219(24):11348–11357, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006115.

Li:2010:APP

[LG10a] Jianjun Li and Wenjie Gao. Analysis of a prey– predator model with disease in prey. Applied Mathematics and Computation, 217(8):4024–4035, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310010398.

Li:2010:FTR

[LG10b] Yonghui Li and Huaxi Gu. Fault tolerant routing algorithm based on the artificial potential field model in Networkon-Chip. Applied Mathematics and Computation, 217(7):3226–3235, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009197.

Lu:2011:CSR

[LG11] Chengbo Lu and Chuanqing Gu. The computation of the square roots of circulant matrices. Applied Mathematics and Computation, 217(16):6819–6829, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000324.

Liu:2014:AFH

[LG14] Wenjun Liu and Xingyue Gao. Approximating the finite Hilbert transform via a companion of Ostrowski’s inequality for function of bounded variation and applications. Applied Mathematics and Computation, 247(??):373–385, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012090.

Lin:2013:ANN

[LGCX13] Shaobo Lin, Xiaofei Guo, Feilong Cao, and Zongben Xu. Approximation by neural networks with scattered data. Applied Mathematics and Computation, 224(??):29–35, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008631.

Loomba:2013:ORI

[LGKM13] Shally Loomba, Rama Gupta, C. N. Kumar, and Daniela Milovic. Optical rogons for inhomogeneous nonlinear Schrödinger equation with inter modal dispersion. Applied Mathematics and Computation, 225(??): 318–325, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313010229.

Liu:2012:WIS

[LGMZ12] Jiaguo Liu, Ivan Gutman, Zengchao Mu, and Yusen Zhang. q-Wiener index of some compound trees. Applied Mathematics and Computation, 218(18):9528–9535, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002809.

Lindquist:2010:KGE

[LGN10] Joseph M. Lindquist, Francis X. Giraldo, and Beny Neta. Klein– Gordon equation with advection on unbounded domains using spectral elements and high-order nonreflecting boundary conditions. Applied Mathematics and Computation, 217(6):2710–2723, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008581.

Lin:2011:SID

[LGS^{+}11] Guo-Dong Lin, Yi-Tian Gao, Zhi-Yuan Sun, Xin Yu, and De-Xin Meng. Solitonic interactions and double-Wronskian-type solutions for a variablecoefficient variant Boussinesq model in the long gravity water waves. Applied Mathematics and Computation, 217(9):4805– 4811, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310011471.

Li:2014:LSS

[LGZ14a] Hongyi Li, Zongsheng Gao, and Di Zhao. Least squares solutions of the matrix equation AXB + CY D = D with the least norm for symmetric arrowhead matrices. Applied Mathematics and Computation, 226(??):719–724, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011223.

Liu:2014:EGS

[LGZ14b] Zhisu Liu, Shangjiang Guo, and Ziheng Zhang. Existence of ground state solutions for the Schrödinger-Poisson systems. Applied Mathematics and Computation, 244(??):312–323, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009655.

Leng:2010:CBE

[LH10a] Huinan Leng and Zhiqing He. Computation of bounds for eigenvalues of structures with interval parameters. Applied Mathematics and Computation, 216(9):2734–2739, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310003863.

Liao:2010:TSN

[LH10b] Li-Man Liao and Ching-Jen Huang. Tabu search for non-permutation flowshop scheduling problem with minimizing total tardiness. Applied Mathematics and Computation, 217(2):557–567, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006582.

Leong:2011:SMS

[LH11a] Wah June Leong and Malik Abu Hassan. Scaled memoryless symmetric rank one method for largescale optimization. Applied Mathematics and Computation, 218(2):413– 418, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007612.

Li:2011:ISM

[LH11b] Junxiang Li and Jiazhen Huo. Inexact smoothing method for large scale minimax optimization. Applied Mathematics and Computation, 218(6):2750–2760, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010435.

Liu:2011:NSS

[LH11c] Jincun Liu and Guolin Hou. Numerical solutions of the space- and timefractional coupled Burgers equations by generalized differential transform method. Applied Mathematics and Computation, 217(16):7001–7008, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001536.

Li:2014:SND

[LH14a] Lianzhong Li and Maoan Han. Some new dynamic Opial type inequalities and applications for second order integro-differential dynamic equations on time scales. Applied Mathematics and Computation, 232(??):542–547, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314001817.

Liu:2014:BAN

[LH14b] Yan Liu and Ze-Rong He. Behavioral analysis of a nonlinear three-staged population model with age-size-structure. Applied Mathematics and Computation, 227(??):437–448, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012332.

Luo:2014:OCAa

[LH14c] Zhixue Luo and Ze-Rong He. Optimal control for age-dependent population hybrid system in a polluted environment. Applied Mathematics and Computation, 228(??):68–76, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012393.

Li:2014:APS

[LHC14] Zhong Li, Maoan Han, and Fengde Chen. Almost periodic solutions of a discrete almost periodic logistic equation with delay. Applied Mathematics and Computation, 232(??):743–751, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001945.

Liu:2012:MTR

[LHCI12] Xiaoji Liu, Shaowu Huang, and Dragana S. Cvetković-Ilić. Mixed-type reverseorder laws for {1, 3, 4}- generalized inverses over Hilbert spaces. Applied Mathematics and Computation, 218(17):8570–8577, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001427.

Liu:2011:NRF

[LHL11] Hongwei Liu, Yakui Huang, and Xiangli Li. New reformulation and feasible semismooth Newton method for a class of stochastic linear complementarity problems. Applied Mathematics and Computation, 217(23): 9723–9740, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311006308.

Luo:2013:NME

[LHL^{+}13] Wei-Hua Luo, Ting-Zhu Huang, Hou-Biao Li, Yong Zhang, and Guo-Cheng Wu. A numerical method and efficient preconditioner for generalized airfoil equations. Applied Mathematics and Computation, 219(24):11451–11459, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313005699.

Li:2014:BTI

[LHL14] Shangzhao Li, Guangguo Han, and Weijun Liu. Block-transitive 2 − (ν, k, 1) designs and the Chevalley groups F_{4}(q). Applied Mathematics and Computation, 248(??):380–385, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031401337X.

Liu:2012:NAS

[LHS12] Yun Liu, Li-Sheng Hu, and Peng Shi. A novel approach on stabilization for linear systems with time-varying input delay. Applied Mathematics and Computation, 218(10):5937–5947, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013993.

Liao:2014:OSD

[LHT14a] Jui-Jung Liao, Kuo-Nan Huang, and Pin-Shou Ting. Optimal strategy of deteriorating items with capacity constraints under two-levels of trade credit policy. Applied Mathematics and Computation, 233(??):647–658, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001143.

Lin:2014:SPI

[LHT14b] Jennifer Lin, Kuo-Chen Hung, and Didos Tang. Solution process for inventory models with crashable lead time. Applied Mathematics and Computation, 227(??):335–340, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011715.

Long:2013:DPI

[LHTN13] Guangqing Long, Xiaoyuan Huang, Aimei Tan, and Gnaneshwar Nelakanti. Discrete product integration methods for eigen-problems of a class of non-compact integral operators. Applied Mathematics and Computation, 219(15):7964–7972, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001562.

Huang:2010:MPM

[lHyLfG10] Ling ling Huang, San yang Liu, and Wei feng Gao. A modified projection method for solving co-coercive variational inequalities. Applied Mathematics and Computation, 217(7):3216–3225, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009185.

Liu:2010:DDD

[LHZ10] Jianzhou Liu, Zejun Huang, and Juan Zhang. The dominant degree and disc theorem for the Schur complement of matrix. Applied Mathematics and Computation, 215(12):4055–4066, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309011205.

Li:2011:CTM

[LHZ^{+}11] Hou-Biao Li, Ting-Zhu Huang, Yong Zhang, Xing-Ping Liu, and Tong-Xiang Gu. Chebyshevtype methods and preconditioning techniques. Applied Mathematics and Computation, 218(2):260– 270, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311007053.

Li:2010:GRS

[Li10a] Xiaodi Li. Global robust stability for stochastic interval neural networks with continuously distributed delays of neutral type. Applied Mathematics and Computation, 215(12):4370–4384, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000068.

Li:2010:MIT

[Li10b] Yongxiang Li. A monotone iterative technique for solving the bending elastic beam equations. Applied Mathematics and Computation, 217(5):2200–2208, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007654.

Li:2010:NEH

[Li10c] Zitian Li. New exact homoclinic wave and periodic wave solutions for the Ginzburg–Landau equation. Applied Mathematics and Computation, 217(4):1549–1554, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630030900616X.

Li:2010:NKS

[Li10d] Zitian Li. New kinkshaped solutions and periodic wave solutions for the (2 + 1)-dimensional Sine-Gordon equation. Applied Mathematics and Computation, 215(11):3777–3781, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010017.

See note [KRS10].

Li:2011:RDI

[Li11a] Xiezhang Li. A representation for the Drazin inverse of block matrices with a singular generalized Schur complement. Applied Mathematics and Computation, 217(18):7531–7536, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002657.

Li:2011:ENS

[Li11b]  Zhouxin Li.  Exis-

tence of nontrivial solutions for quasilinear elliptic equations at critical growth. Applied Mathematics and Computation, 218(1):76–87, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007235.

Li:2012:FHM

[Li12a]  Hsuan-Chu Li.  On

Fibonacci–Hessenberg matrices and the Pell and Perrin numbers. Applied Mathematics and Computation, 218(17):8353–8358, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001087.

Li:2012:CAN

[Li12b] Huanrong Li. Convergence analysis and numerical simulation for a hyperbolic system. Applied Mathematics and Computation, 218(22):11209–11216, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004961.

Li:2012:DDF

[Li12c] Yingguo Li. Dynamics of a discrete food-limited population model with time delay. Applied Mathematics and Computation, 218(12):6954–6962, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300311015657.

Li:2013:CPT

[Li13]  Lian He Li.  Com-

plex potential theory for the plane elasticity problem of decagonal quasicrystals and its application. Applied Mathematics and Computation, 219(19):10105– 10111, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003366.

Li:2014:PRP

[Li14a] Min Li. A Polak–Ribière– Polyak method for solving large-scale nonlinear systems of equations and its global convergence. Applied Mathematics and Computation, 248(??):314–322, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013460.

Li:2014:NRD

[Li14b] Yanqiu Li. Non-resonant double Hopf bifurcation of a class-B laser system. Applied Mathematics and Computation, 226(??):564–574, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301134X.

Liang:2011:BGS

[Lia11] Fei Liang. Blow-up and global solutions for nonlinear reaction-diffusion equations with nonlinear boundary condition. Applied Mathematics and Computation, 218(8):3993– 3999, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311012513.

Liang:2014:LIC

[Lia14] Hua Liang. Linearly implicit conservative schemes for long-term numerical simulation of Klein– Gordon–Schrödinger equations. Applied Mathematics and Computation, 238(??):475–484, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005657.

Llinares:2012:ARA

[LIC12] Raul Llinares, Jorge Igual, and Andres Camacho. Application of regularized Alternating Least Squares to an astrophysical problem. Applied Mathematics and Computation, 219(3):1367–1374, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007606.

Limam:2012:RST

[Lim12] Kheira Limam. Resolution, in L^{p}-spaces, of transmission problems set in an unbounded domains. Applied Mathematics and Computation, 218(9):5605–5619, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101397X.

Lin:2010:AGA

[Lin10]  Lu Lin.  Alternative

gradient algorithms with applications to nonnegative matrix factorizations. Applied Mathematics and Computation, 216(6):1763–1770, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010753.

Lin:2011:BDS

[Lin11a] Qian Lin. Backward doubly stochastic differential equations with weak assumptions on the coefficients. Applied Mathematics and Computation, 217(22):9322–9333, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005728.

Lin:2011:AFCb

[Lin11b] Ruiyue Lin. Allocating fixed costs and common revenue via data envelopment analysis. Applied Mathematics and Computation, 218(7):3680–3688, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011775.

Lin:2011:AFCa

[Lin11c]  Ruiyue Lin.  Allocat-

ing fixed costs or resources and setting targets via data envelopment analysis. Applied Mathematics and Computation, 217(13):6349–6358, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000129.

Lin:2011:JEU

[Lin11d] Shi-Woei Lin. Jackknife evaluation of uncertainty judgments aggregated by the Kullback–Leibler distance. Applied Mathematics and Computation, 218(2):469–479, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007685.

Lin:2012:EAS

[Lin12]  Chu-Ti Lin.  Enhanc-

ing the accuracy of software reliability prediction through quantifying the effect of test phase transitions. Applied Mathematics and Computation, 219(5):2478–2492, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008764.

Lin:2013:MRG

[Lin13a]  Jen-Chih Lin.  Multi-

ple regular graph embeddings into a hypercube with unbounded expansion. Applied Mathematics and Computation, 220(??):429–438, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006516.

Lin:2013:EIM

[Lin13b] Kuo-Ping Lin. An extended inventory models with trapezoidal type demands. Applied Mathematics and Computation, 219(24):11414–11419, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006036.

Lin:2013:SPT

[Lin13c]  Xiaojie Lin.  Singular

perturbations of thirdorder nonlinear differential equations with full nonlinear boundary conditions. Applied Mathematics and Computation, 224(??):88–95, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009168.

Lin:2014:PSS

[Lin14a]  Bin Lin.  Parametric

spline solution of the regularized long wave equation. Applied Mathematics and Computation, 243(??):358–367, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008339.

Lin:2014:SFR

[Lin14b]  Long Lin.  On Stir-

ling’s formula remainder. Applied Mathematics and Computation, 247(??):494–500, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012302.

Lipovetsky:2013:FCC

[Lip13] Stan Lipovetsky. Finding cluster centers and sizes via multinomial parameterization. Applied Mathematics and Computation, 221(??):571–580, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007479.

Lisena:2014:GDL

[Lis14] Benedetta Lisena. On the global dynamics of the lengyel-Epstein system. Applied Mathematics and Computation, 249(??):67–75, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013708.

Liu:2010:NEP

[Liu10a] Changfu Liu. New exact periodic solitary wave solutions for Kadomtsev– Petviashvili equation. Applied Mathematics and Computation, 217(4):1350– 1354, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309004512.

Liu:2010:SRE

[Liu10b] Na Liu. Similarity reduction and explicit solutions for the variable-coefficient coupled Burger’s equations. Applied Mathematics and Computation, 217(8):4178–4185, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010738.

Liu:2010:SSL

[Liu10c]  Xiuhong Liu.  Stabi-

lization of switched linear systems with modedependent time-varying delays. Applied Mathematics and Computation, 216(9):2581–2586, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003553.

Liu:2010:DIS

[Liu10d] Yu-Hsin Liu. Different initial solution generators in genetic algorithms for solving the probabilistic traveling salesman problem. Applied Mathematics and Computation, 216(1):125–137, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000342.

Liu:2011:IPC

[Liu11a] Chein-Shan Liu. An inverse problem for computing a leading coefficient in the Sturm–Liouville operator by using the boundary data. Applied Mathematics and Computation, 218(8):4245–4259, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012276.

Liu:2011:UGA

[Liu11b] Cheng-Hsiang Liu. Using genetic algorithms for the coordinated scheduling problem of a batching machine and two-stage transportation. Applied Mathematics and Computation, 217(24):10095–10104, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006606.

Liu:2011:AHP

[Liu11c] Hsuan-Ku Liu. Application of homotopy perturbation methods for solving systems of linear equations. Applied Mathematics and Computation, 217(12):5259–5264, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011367.

Liu:2011:EAS

[Liu11d] Lingxia Liu. Existence of analytic solutions of an iterative functional equation. Applied Mathematics and Computation, 217(17):7245–7257, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001731.

Liu:2011:BEE

[Liu11e]  Xuanliang Liu.  Bi-

furcation of an ecoepidemiological model with a nonlinear incidence rate. Applied Mathematics and Computation, 218(5):2300–2309, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009970.

Liu:2011:ESN

[Liu11f] Yang Liu. Exact solutions to nonlinear Schrödinger equation with variable coefficients. Applied Mathematics and Computation, 217(12):5866–5869, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012749.

Liu:2011:MAM

[Liu11g]  Zheng Liu.  More on

the averaged midpointtrapezoid type rules. Applied Mathematics and Computation, 218(4):1389– 1398, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008484.

Liu:2012:OSV

[Liu12a] Chein-Shan Liu. Optimally scaled vector regularization method to solve ill-posed linear problems. Applied Mathematics and Computation, 218(21): 10602–10616, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004213.

Liu:2012:SCR

[Liu12b] Gang Liu. Superattracting cycles of the relaxed Newton’s method for entire functions. Applied Mathematics and Computation, 218(24):12008–12012, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006005.

Liu:2012:MCC

[Liu12c]  Landong Liu.  Mixed

and componentwise condition numbers of nonsymmetric algebraic Riccati equation. Applied Mathematics and Computation, 218(14):7595–7601, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000616.

Liu:2013:QAH

[Liu13a] Jia Liu. Qualitative analysis of a Hölling–Tanner predator-prey model with mutual interference. Applied Mathematics and Computation, 219(10): 5283–5289, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312012350.

Liu:2013:RBI

[Liu13b]  Yong Liu.  Relation-

ship between industrial firms, high-carbon and low-carbon energy: an agent-based simulation approach. Applied Mathematics and Computation, 219(14):7472–7479, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000738.

Liu:2014:CDI

[Liu14a] Jia Liu. Cross-diffusion induced stationary patterns in a prey-predator system with parental care for predators. Applied Mathematics and Computation, 237(??):176–189, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004226.

Liu:2014:ASE

[Liu14b] Pan-Ping Liu. Analysis of a spatial epidemic model with saturated incidence rate. Applied Mathematics and Computation, 243(??):426–432, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008625.

Liu:2014:CRS

[Liu14c] Ping Liu. A class of rotational solutions for the Ndimensional incompressible Navier–Stokes equations. Applied Mathematics and Computation, 245(??):504–508, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010583.

Liu:2014:CSC

[Liu14d] Xifu Liu. Comments on “The common Re-nnd and Re-pd solutions to the matrix equations AX + C and XB = D”. Applied Mathematics and Computation, 236(??):663–668, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004366.

See [XQ11a].

Li:2014:CAP

[LIX14] Kevin W. Li, Takehiro Inohara, and Haiyan Xu. Coalition analysis with preference uncertainty in group decision support. Applied Mathematics and Computation, 231(??):307–319, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014331.

Lu:2011:QIS

[LJ11] Tzyy-Chyang Lu and Jyh-Ching Juang. Quantuminspired space search algorithm (QSSA) for global numerical optimization. Applied Mathematics and Computation, 218(6):2516–2532, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010162.

Layeni:2012:HHB

[LJ12a] O. P. Layeni and J. V. Johnson. Hybrids of the heat balance integral method. Applied Mathematics and Computation, 218(14):7431–7444, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000069.

Lee:2012:SEM

[LJ12b] Usik Lee and Injoon Jang.

Spectral element modeling and analysis of the blood flows in viscoelastic vessels. Applied Mathematics and Computation, 218(13):7295–7307, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000264.

Liu:2013:CAA

[LJ13] Jun Liu and Yao-Lin Jiang. Convergence analysis of an Arnoldi order reduced Runge– Kutta method for integrodifferential equations of pantograph type. Applied Mathematics and Computation, 219(24):11460– 11470, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313005948.

Liu:2012:ALG

[LJCI12] Xiaoji Liu, Hongwei Jin, and Dragana S. Cvetković-Ilić. The absorption laws for the generalized inverses. Applied Mathematics and Computation, 219(4):2053–2059, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008429.

Liu:2013:RGI

[LJCI13] Xiaoji Liu, Hongwei Jin, and Dragana S. Cvetković-Ilić. Representations of generalized inverses of partitioned matrix involving Schur complement. Applied Mathematics and Computation, 219(18):9615–9629, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012659.

Jiang:2012:ACS

[lJCWpW12] Jian lin Jiang, Kun Cheng, Can-Can Wang, and Li ping Wang. Accelerating the convergence in the single-source and multi-source Weber problems. Applied Mathematics and Computation, 218(12):6814–6824, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015372.

Li:2011:GCA

[LjH11] Xi Li and Nan jing Huang.

Graph convergence for the H(·, ·)-accretive operator in Banach spaces with an application. Applied Mathematics and Computation, 217(22):9053–9061, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004875.

Lee:2011:RFC

[LJKP11] S. M. Lee, D. H. Ji, O. M.

Kwon, and Ju H. Park. Robust H_{∞} filtering for a class of discrete-time nonlinear systems. Applied Mathematics and Computation, 217(20):7991–7997, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003262.

Liu:2011:GEM

[LJKU11] Zeqing Liu, Anshu Jiang, Shin Min Kang, and Jeong Sheok Ume. Global existence and Mann iterative algorithms of positive solutions for first order nonlinear neutral delay differential equations. Applied Mathematics and Computation, 217(22):9424–9437, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005911.

Jiang:2014:GWP

[lJLpWxZ14] Jian lin Jiang, Fei Luan, Li ping Wang, and Xiao xing Zhu. A generalized Weber problem with different gauges for different regions. Applied Mathematics and Computation, 244(??):244–260, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400962X.

Lee:2012:DGC

[LJPJ12] Tae H. Lee, D. H. Ji, Ju H. Park, and H. Y. Jung. Decentralized guaranteed cost dynamic control for synchronization of a complex dynamical network with randomly switching topology. Applied Mathematics and Computation, 219(3):996–1010, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312007047.

Lee:2014:DMT

[LJR14] Yang-Hi Lee, Soon-Mo Jung, and Michael Th. Rassias. On an ndimensional mixed type additive and quadratic functional equation. Applied Mathematics and Computation, 228(??):13– 16, February 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313012605.

Liu:2013:UGS

[LJT13] Wenjun Liu, Yong Jiang, and Adnan Tuna. A unified generalization of some quadrature rules and error bounds. Applied Mathematics and Computation, 219(9):4765–4774, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201123X.

Liu:2010:TTE

[LJW10] Qing Liu, Dong-Li Jia, and Zi-Hua Wang. Three types of exact solutions to Wick-type generalized stochastic Korteweg–de Vries equation. Applied Mathematics and Computation, 215(10):3495–3500, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030900914X.

Lin:2014:LTB

[LJX14] Yuguo Lin, Daqing Jiang, and Peiyan Xia. Longtime behavior of a stochastic SIR model. Applied Mathematics and Computation, 236(??):1–9, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400397X.

Lin:2011:DTS

[LK11] Chuen-Horng Lin and Jau-Chuan Ke. On the discrete-time system with server breakdowns: Computational algorithm and optimization algorithm. Applied Mathematics and Computation, 218(7):3624–3634, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011544.

Li:2012:USN

[LK12] Yibao Li and Junseok Kim. An unconditionally stable numerical method for bimodal image segmentation. Applied Mathematics and Computation, 219(6):3083–3090, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200937X.

Liemert:2014:ESP

[LK14] André Liemert and Alwin Kienle. Exact solution of Poisson’s equation with an elliptical boundary. Applied Mathematics and Computation, 238(??):123–131, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005268.

Liao:2012:HAC

[LKH12] T. Warren Liao, R. J. Kuo, and J. T. L. Hu. Hybrid ant colony optimization algorithms for mixed discrete-continuous optimization problems. Applied Mathematics and Computation, 219(6):3241– 3252, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009630.

Lee:2012:PCS

[LKP12] S. M. Lee, O. M. Kwon, and Ju H. Park. Predictive control for sector bounded nonlinear model and its application to solid oxide fuel cell systems. Applied Mathematics and Computation, 218(18):9296–9304, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002408.

Liang:2014:ETD

[LKS14] Xiao Liang, Abdul Q. M. Khaliq, and Qin Sheng. Exponential time differencing Crank–Nicolson method with a quartic spline approximation for nonlinear Schrödinger equations. Applied Mathematics and Computation, 235(??):235–252, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003026.

Lee:2010:RSN

[LKY10] Chul Woo Lee, Young Ho Kim, and Dae Won Yoon. Ruled surfaces of non-degenerate third fundamental forms in Minkowski 3-spaces. Applied Mathematics and Computation, 216(11): 3200–3208, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310004716.

Li:2014:GEC

[LKY14] Chenglin Li, Aditya Kaushik, and Guocheng Yin. Global existence of classical solutions to an acid-mediated invasion model for tumorstromal interactions. Applied Mathematics and Computation, 234(??): 599–605, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314002963.

Lan:2010:CSN

[LL10a] Yaoyao Lan and Qingguo Li. Chaos synchronization of a new hyperchaotic system. Applied Mathematics and Computation, 217(5):2125–2132, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007563.

Li:2010:SOP

[LL10b] Jing Li and Qi Lü. State observation problem for general time reversible system and applications. Applied Mathematics and Computation, 217(6):2843–2856, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000874X.

Li:2010:FIM

[LL10c] Weiguo Li and Zhi Li. A family of iterative methods for computing the approximate inverse of a square matrix and inner inverse of a non-square matrix. Applied Mathematics and Computation, 215(9):3433–3442, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009436.

Liang:2010:FNS

[LL10d] Fen Liang and Fu-Rong Lin. A fast numerical solution method for two dimensional Fredholm integral equations of the second kind based on piecewise polynomial interpolation. Applied Mathematics and Computation, 216(10):3073–3088, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004479.

Liu:2010:PNS

[LL10e] Bingchen Liu and Fengjie Li. Properties of nonsimultaneous blow-up in heat equations coupled via different localized sources. Applied Mathematics and Computation, 217(7):3403–3411, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009537.

Liu:2010:QSM

[LL10f] Li-Bin Liu and Huan-Wen Liu. Quartic spline methods for solving one-dimensional telegraphic equations. Applied Mathematics and Computation, 216(3):951–958, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001347.

Luo:2010:SOP

[lL10g] Xin long Luo. A secondorder pseudo-transient method for steady-state problems. Applied Mathematics and Computation, 216(6):1752–1762, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010765.

Lu:2010:SOM

[LL10h] Dianchen Lu and CuiLian Liu. A sub-ODE method for generalized Gardner and BBM equation with nonlinear terms of any order. Applied Mathematics and Computation, 217(4):1404–1407, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309005323.

Luo:2010:EPPb

[LL10i] Yan Luo and Zhiguo Luo. Existence of positive periodic solutions for neutral multi-delay logarithmic population model. Applied Mathematics and Computation, 216(4):1310–1315, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000192X.

Li:2011:SMM

[LL11a] Wei Li and Haohao Li.

On simplex method with most-obtuse-angle rule and cosine rule. Applied Mathematics and Computation, 217(20):7867–7873, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002670.

Liang:2011:IAM

[LL11b] Kaifu Liang and Jianzhou Liu. Iterative algorithms for the minimum-norm solution and the leastsquares solution of the linear matrix equations A_{1}XB_{1} + C_{1}X^{T} D_{1} = M_{1}, A_{2}XB_{2} + C_{2}X^{T} D_{2} = M_{2}. Applied Mathematics and Computation, 218(7):3166–3175, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010927.

Lima:2011:LCI

[LL11c] Mauricio Firmino Silva Lima and Jaume Llibre. Limit cycles and invariant cylinders for a class of continuous and discontinuous vector field in dimension 2n. Applied Mathematics and Computation, 217(24):9985–9996, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006394.

Lin:2011:APG

[LL11d] Lu Lin and Zhong-Yun Liu. An alternating projected gradient algorithm for nonnegative matrix factorization. Applied Mathematics and Computation, 217(24):9997–10002, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006400.

Liu:2011:EES

[LL11e] Hanze Liu and Lei Liu. Exact and explicit solutions for a nonlinear extended iterative differential equation. Applied Mathematics and Computation, 218(2):548–562, September 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311007806.

Lee:2012:GSD

[LL12a] Wen-Chiung Lee and Zon-Syun Lu. Group scheduling with deteriorating jobs to minimize the total weighted number of late jobs. Applied Mathematics and Computation, 218(17):8750–8757, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001609.

Li:2012:MPP

[LL12b] Chanjuan Li and Zhongfei Li. Multi-period portfolio optimization for asset-liability management with bankrupt control. Applied Mathematics and Computation, 218(22):11196–11208, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031200495X.

Li:2012:OAC

[LL12c] Huanrong Li and Yukun Li. Optimal approximation to a class of nonlinear evolution equations. Applied Mathematics and Computation, 218(17):8842–8852, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001701.

Li:2012:DRC

[LL12d] Ming Li and Long Liu. Decentralized robust control for interconnected neutral delay systems. Applied Mathematics and Computation, 218(14):7453–7458, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000318.

Liu:2012:CGM

[LL12e] Yang Liu and Hong Li. Corrigendum to “H^{1}-Galerkin mixed finite element methods for pseudo-hyperbolic equations” [Appl. Math. Comput. 212 (2009) 446– 457]. Applied Mathematics and Computation, 218 (19):10008, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003268.

See [?].

Lv:2012:NSH

[LL12f] Wanjin Lv and Hui Li. Numerical stability of higherorder derivative methods for the pantograph equation. Applied Mathematics and Computation, 218(9):5739–5745, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014147.

Lv:2012:HSO

[LL12g] Xiang Lv and Shiping Lu.

Homoclinic solutions for ordinary p-Laplacian systems. Applied Mathematics and Computation, 218(9):5682–5692, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014081.

Li:2013:EFR

[LL13a] Can Li and Dong-Hui Li. An extension of the Fletcher–Reeves method to linear equality constrained optimization problem. Applied Mathematics and Computation, 219(23):10909–10914, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005134.

Li:2013:SHB

[LL13b] Yongkun Li and Changzhao Li. Stability and Hopf bifurcation analysis on a delayed Leslie–Gower predator-prey system incorporating a prey refuge. Applied Mathematics and Computation, 219(9):4576– 4589, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312010855.

Li:2013:NPC

[lL13c] Chen liang Li. A new parallel cascadic multigrid method. Applied Mathematics and Computation, 219(20):10150– 10157, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300313003561.

Lin:2013:RCT

[LL13d] Hongxia Lin and Shan Li. Regularity criteria in terms of the pressure for the threedimensional MHD equations. Applied Mathematics and Computation, 221(??):164–168, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006942.

Liu:2013:CDS

[LL13e] Li-Bin Liu and Huan-Wen Liu. Compact difference schemes for solving telegraphic equations with Neumann boundary conditions. Applied Mathematics and Computation, 219(19):10112– 10121, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004219.

Liu:2013:MWS

[LL13f] Yifang Liu and Guo-Rong Li. Matter wave soliton solutions of the cubic-quintic nonlinear Schrödinger equation with an anharmonic potential. Applied Mathematics and Computation, 219(9):4847–4852, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201140X.

Lu:2013:PIC

[LL13g] Z. R. Lu and J. K. Liu. Parameters identification for a coupled bridge-vehicle system with spring-mass attachments. Applied Mathematics and Computation, 219(17):9174–9186, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003007.

Lai:2014:MTF

[LL14a] Chin-Diew Lai and Gwo Dong Lin. Mean time to failure of systems with dependent components. Applied Mathematics and Computation, 246(??):103–111, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314010601.

Lee:2014:SAO

[LL14b] Kwang Sung Lee and Abid Ali Lashari. Stability analysis and optimal control of pine wilt disease with horizontal transmission in vector population. Applied Mathematics and Computation, 226(??):793–804, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010400.

Li:2014:DDI

[LL14c] Chenkuan Li and Changpin Li. On defining the distributions δ^{k} and (δ′)^{k} by fractional derivatives. Applied Mathematics and Computation, 246(??):502–513, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011783.

Li:2014:HCP

[LL14d] X. B. Li and S. J.

Li. Hölder continuity of perturbed solution set for convex optimization problems. Applied Mathematics and Computation, 232(??):908–918, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001325.

Li:2014:CDN

[LL14e] Yuanfei Li and Changhao Lin. Continuous dependence for the nonhomogeneous Brinkman– Forchheimer equations in a semi-infinite pipe. Applied Mathematics and Computation, 244(??): 201–208, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009266.

Liu:2014:NHC

[LL14f] J. K. Liu and S. J. Li. New hybrid conjugate gradient method for unconstrained optimization. Applied Mathematics and Computation, 245(??):36–43, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010637.

Lin:2011:BGM

[LLBW11] Yiqin Lin, Wenbo Li, Liang Bao, and Qinghua Wu. On the block GM-RES method with deflated restarting. Applied Mathematics and Computation, 218(5):2146–2154, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009647.

Liu:2011:PSA

[LLC11] Li-Bin Liu, Huan-Wen Liu, and Yanping Chen. Polynomial spline approach for solving secondorder boundary-value problems with Neumann conditions. Applied Mathematics and Computation, 217(16):6872–6882, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000634.

Li:2012:PCC

[LLC12] Hongyi Li, James Lam, and K. C. Cheung. Passivity criteria for continuoustime neural networks with mixed time-varying delays. Applied Mathematics and Computation, 218(22):11062–11074, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004778.

Liu:2010:NCN

[LLCX10] Gang Liu, Shengqi Lu, Juping Chen, and Wei Xu. A note on condition numbers for generalized inverse A_{T,S}^{(2)} and constrained linear systems. Applied Mathematics and Computation, 217(7):3199–3206, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009161.

Ling:2011:EUL

[LLHY11] Liuyi Ling, Liang Liang, Zhiyong Hu, and Feng Yang. Effect of upper and lower withdrawal limits on queue length. Applied Mathematics and Computation, 218(7):3834–3846, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012021.

Li:2011:PSB

[LLJ11] Xiaoyan Li, Song Liu, and Wei Jiang. Positive solutions for boundary value problem of nonlinear fractional functional differential equations. Applied Mathematics and Computation, 217(22):9278–9285, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005479.

Liu:2012:ASC

[LLJF12] Song Liu, Xiaoyan Li, Wei Jiang, and Yi-Zheng Fan. Adaptive synchronization in complex dynamical networks with coupling delays for general graphs. Applied Mathematics and Computation, 219(1):83–87, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007715.

Liu:2013:FPT

[LLK13] Zeqing Liu, Yan Lu, and Shin Min Kang. Fixed point theorems for multivalued contractions with w-distance. Applied Mathematics and Computation, 224(??):535–552, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009296.

Li:2014:ICH

[LLK14] Dingfang Li, Ping Liu, and Jisheng Kou. An improvement of Chebyshev– Halley methods free from second derivative. Applied Mathematics and Computation, 235(??):221–225, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003312.

Liu:2014:SIS

[LLKP14] Yajuan Liu, S. M. Lee, O. M. Kwon, and Ju H. Park. A study on H_{∞} state estimation of static neural networks with timevarying delays. Applied Mathematics and Computation, 226(??):589–597, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011326.

Liu:2010:PCS

[LLL10a] Lixia Liu, Sanyang Liu, and Hongwei Liu. A predictor–corrector smoothing Newton method for symmetric cone complementarity problems. Applied Mathematics and Computation, 217(7):2989– 2999, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310008866.

Lu:2010:ESD

[LLL10b] Hailing Lü, Xiqiang Liu, and Na Liu. Explicit solutions of the (2 + 1)- dimensional AKNS shallow water wave equation with variable coefficients. Applied Mathematics and Computation, 217(4):1287–1293, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309004469.

Liang:2011:CBF

[LLL11a] Jitai Liang, Yiliang Liu, and Zhenhai Liu. A class of BVPS for first order impulsive integro-differential equations. Applied Mathematics and Computation, 218(7):3667–3672, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311011751.

Luo:2011:SOT

[LLL11b] Lisuo Luo, Wanshe Li, and Qiao Li. A study on orthogonal two-direction vector-valued wavelets and two-direction wavelet packets. Applied Mathematics and Computation, 217(24):10146–10157, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006643.

Liu:2012:MDD

[LLL12] Ping Liu, Zi-Liang Li, and Ren-Ze Luo. Modified (2 + 1)-dimensional displacement shallow water wave system: Symmetries and exact solutions. Applied Mathematics and Computation, 219(4):2149–2157, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008545.

Lin:2013:SCA

[LLL13] Yezhi Lin, Yinping Liu, and Zhibin Li. Symbolic computation of analytic approximate solutions for nonlinear differential equations with boundary conditions. Applied Mathematics and Computation, 222(??): 145–166, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031300725X.

Lin:2011:IRG

[LLLS11] Shy-Der Lin, Shuoh-Jung Liu, Han-Chun Lu, and H. M. Srivastava. Integral representations for the generalized Bedient polynomials and the generalized Cesàro polynomials. Applied Mathematics and Computation, 218(4):1330–1341, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008290.

Liu:2013:LPC

[LLLS13] Shuoh-Jung Liu, Shy-Der Lin, Han-Chun Lu, and H. M. Srivastava. Linearization of the products of the Carlitz–Srivastava polynomials of the first and second kinds via their integral representations. Applied Mathematics and Computation, 219(9):4545–4550, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010788.

Lai:2011:AAL

[LLLZ11] S. K. Lai, C. W. Lim, Zhang Lin, and W. Zhang. Analytical analysis for large-amplitude oscillation of a rotational pendulum system. Applied Mathematics and Computation, 217(13):6115–6124, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310013068.

Lu:2010:GEM

[LLN10] Hai-Ling Lü, Xi-Qiang Liu, and Lei Niu. A generalized (\frac{G′}{G})-expansion method and its applications to nonlinear evolution equations. Applied Mathematics and Computation, 215(11):3811–3816, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010108.

Li:2010:VEF

[LLP10] Fang Li, Zhibin Li, and Ling Pi. Variable exponent functionals in image restoration. Applied Mathematics and Computation, 216(3):870–882, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001189.

Lee:2014:ICR

[LLP14a] Won Il Lee, Seok Young Lee, and PooGyeon Park. Improved criteria on robust stability and H_{∞} performance for linear systems with interval timevarying delays via new triple integral functionals. Applied Mathematics and Computation, 243(??):570–577, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008169.

Liu:2014:GSS

[LLP14b] Qun Liu, Yiliang Liu, and Xue Pan. Global stability of a stochastic predator-prey system with infinite delays. Applied Mathematics and Computation, 235(??):1–7, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003397.

Long:2010:NTW

[LLR10] Yao Long, Zhenyang Li, and Weiguo Rui. New travelling wave solutions for a nonlinearly dispersive wave equation of Camassa–Holm equation type. Applied Mathematics and Computation, 217(4):1315–1320, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309004421.

Luo:2011:SBA

[LLS11] Zhiguo Luo, Jianli Li, and Jianhua Shen. Stability of Bidirectional Associative Memory networks with impulses. Applied Mathematics and Computation, 218(5):1658–1667, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008782.

Liu:2013:FMP

[LLS^{+}13a] Chengming Liu, Zhongxuan Luo, Xiquan Shi, Fengshan Liu, and Xiaonan Luo. A fast mesh parameterization algorithm based on 4-point interpolatory subdivision. Applied Mathematics and Computation, 219(10):5339–5344, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012106.

Luo:2013:ROC

[LLS13b] Zhendong Luo, Hong Li, and Ping Sun. A reducedorder Crank–Nicolson finite volume element formulation based on POD method for parabolic equations. Applied Mathematics and Computation, 219(11):5887–5900, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012507.

Liu:2014:CDS

[LLS14] Hong-Zhun Liu, Sen Lin, and Xiao-Quan Sun. Comment on: “Double subequation method for complexiton solutions of nonlinear partial differential equations”. Applied Mathematics and Computation, 246(??):597–598, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011187.

See [LLS14].

Liu:2012:BGF

[LLSW12] Shuoh-Jung Liu, Shy-Der Lin, H. M. Srivastava, and Mu-Ming Wong. Bilateral generating functions for the Erkus–Srivastava polynomials and the generalized Lauricella functions. Applied Mathematics and Computation, 218(15):7685–7693, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000239.

Lee:2011:SII

[LLSY11] Mun Bae Lee, Yeon Ju Lee, Hasik Sunwoo, and Jungho Yoon. Some issues on interpolation matrices of locally scaled radial basis functions. Applied Mathematics and Computation, 217(10):5011–5014, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011690.

Liu:2014:RMA

[LLT^{+}14] Q. Liu, F. Liu, I. Turner, V. Anh, and Y. T. Gu. A RBF meshless approach for modeling a fractal mobile/immobile transport model. Applied Mathematics and Computation, 226(??):336–347, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010631.

Li:2010:RAR

[LLW10a] Jia Li, Yangrong Li, and Bin Wang. Random attractors of reactiondiffusion equations with multiplicative noise in L^{p}. Applied Mathematics and Computation, 215(9):3399–3407, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009382.

Liu:2010:NSH

[LLW10b] Lishan Liu, Bingmei Liu, and Yonghong Wu. Nontrivial solutions for higherorder m-point boundary value problem with a sign-changing nonlinear term. Applied Mathematics and Computation, 217(8):3792–3800, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009987.

Liu:2011:IEP

[LLW11] Zhibing Liu, Linzhang Lu, and Kanmin Wang. The inverse eigenvalue problem of generalized reflexive matrices and its approximation. Applied Mathematics and Computation, 218(5):1611–1616, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008721.

Li:2013:CCC

[LLW13] Feng Li, Jing Li, and Yusen Wu. Computing center conditions for non-analytic vector fields with constant angular speed. Applied Mathematics and Computation, 220(??):200–207, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006164.

Lin:2014:RDS

[lLW14] Huei li Lin and Feng-Bin Wang. On a reactiondiffusion system modeling the dengue transmission with nonlocal infections and crowding effects. Applied Mathematics and Computation, 248(??):184–194, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013356.

Liu:2010:SES

[LLWW10] Qing Liu, Jia-Cheng Lan, Hong-Yu Wu, and Zi-Hua Wang. A series of exact solutions for (2 + 1)-dimensional Wicktype stochastic generalized Broer–Kaup system via a modified variablecoefficient projective Riccati equation mapping method. Applied Mathematics and Computation, 217(2):629–638, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310006739.

Liu:2010:SSN

[LLX10] Jun Liu, Xinzhi Liu, and Wei-Chau Xie. On the (h_{0}, h)-stabilization of switched nonlinear systems via state-dependent switching rule. Applied Mathematics and Computation, 217(5):2067–2083, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007526.

Li:2012:FAE

[LLXZ12] Yutian Li, Saiyu Liu, Shuaixia Xu, and Yuqiu Zhao. Full asymptotic expansions of the Landau constants via a difference equation approach. Applied Mathematics and Computation, 219(3):988–995, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007035.

Li:2011:EGS

[LLY11a] An Li, Wei Lv, and Jianxiong Ye. Exponential and global stability of nonlinear dynamical systems relative to initial time difference. Applied Mathematics and Computation, 217(12):5923–5929, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013147.

Li:2011:TFL

[LLY11b] Xin-Yue Li, Yu-Xia Li, and Hong-Xiang Yang. Two families of Liouville integrable lattice equations. Applied Mathematics and Computation, 217(21):8671–8682, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004796.

Li:2014:TWN

[LLY14a] Yan Li, Wan-Tong Li, and Fei-Ying Yang. Traveling waves for a nonlocal dispersal SIR model with delay and external supplies. Applied Mathematics and Computation, 247(??):723–740, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314012971.

Liang:2014:SAR

[LLY14b] Hui Liang, M. Z. Liu, and Z. W. Yang. Stability analysis of Runge– Kutta methods for systems u′(t) = Lu(t) + Mu([t]). Applied Mathematics and Computation, 228(??):463–476, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012897.

Lin:2010:ESC

[LLYH10] Jui-Sheng Lin, Teh-Lu Liao, Jun-Juh Yan, and Meei-Ling Hung. Exponential synchronization of chaotic systems subject to uncertainties in the control input. Applied Mathematics and Computation, 216(8):2441–2449, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003437.

Liao:2014:LLD

[lLZ14a] Hong lin Liao and Ying Zhao. Linearly localized difference schemes for the nonlinear Maxwell model of a magnetic field into a substance. Applied Mathematics and Computation, 233(??):608–622, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002604.

Liu:2014:SIU

[LLZ14b] Peide Liu, Zhengmin Liu, and Xin Zhang. Some intuitionistic uncertain linguistic Heronian mean operators and their application to group decision making. Applied Mathematics and Computation, 230(??):570–586, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014239.

Liu:2014:SLQ

[LLZ14c] Xikui Liu, Yan Li, and Weihai Zhang. Stochastic linear quadratic optimal control with constraint for discrete-time systems. Applied Mathematics and Computation, 228(??):264–270, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010084.

Lu:2011:VST

[LLZC11] Aiguo Lu, Hongwei Liu, Xiuyun Zheng, and Weijie Cong. A variant spectral-type FR conjugate gradient method and its global convergence. Applied Mathematics and Computation, 217(12):5547–5552, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012300.

Liu:2010:SAD

[LLZY10] Zixin Liu, Shu Lü, Shouming Zhong, and Mao Ye. Stabilization analysis for discrete-time systems with time delay. Applied Mathematics and Computation, 216(7):2024–2035, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002754.

Luzon:2011:BDE

[LM11] Ana Luzón and Manuel A. Morón. Bivariate deltaevolution equations and convolution polynomials: Computing polynomial expansions of solutions. Applied Mathematics and Computation, 218(4):1417–1435, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008514.

Llibre:2012:POF

[LM12] Jaume Llibre and Amar Makhlouf. Periodic orbits of the fourth-order non-autonomous differential equation u′′′ + qu′′ + pu = ϵF (t, u, u′, u′′, u′′′). Applied Mathematics and Computation, 219(3):827– 836, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312006546.

Torre:2013:CGA

[LM13a] D. La Torre and F. Mendivil. A chaos game algorithm for generalized iterated function systems. Applied Mathematics and Computation, 224(??):238–249, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031300917X.

Liu:2013:SCS

[LM13b] Wei Liu and Xuerong Mao. Strong convergence of the stopped Euler– Maruyama method for nonlinear stochastic differential equations. Applied Mathematics and Computation, 223(??):389–400, October 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008722.

Luo:2014:SRE

[LMA14] Min-Jie Luo, Gradimir V. Milovanovic, and Praveen Agarwal. Some results on the extended beta and extended hypergeometric functions. Applied Mathematics and Computation, 248(??):631–651, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013447.

Loukopoulos:2013:NSI

[LMB13] V. C. Loukopoulos, G. T. Messaris, and G. C. Bourantas. Numerical solution of the incompressible Navier–Stokes equations in primitive variables and velocity-vorticity formulation. Applied Mathematics and Computation, 222(??):575–588, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008370.

Liu:2012:OAA

[LML12] Haidong Liu, Fanwei Meng, and Puchen Liu. Oscillation and asymptotic analysis on a new generalized Emden–Fowler equation. Applied Mathematics and Computation, 219(5):2739–2748, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008995.

Li:2010:SOM

[LMMW10] Xiaowu Li, Chunlai Mu, Jinwen Ma, and Chan Wang. Sixteenth-order method for nonlinear equations. Applied Mathematics and Computation, 215(10):3754–3758, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300309010054.

Li:2010:IPS

[LMS10] Bangqing Li, Yulan Ma, and Jianzhi Sun. The interaction processes of the N-soliton solutions for an extended generalization of Vakhnenko equation. Applied Mathematics and Computation, 216(12):3522–3535, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310005126.

Liman:2011:IPV

[LMS11] A. Liman, R. N. Mohapatra, and W. M. Shah. Inequalities for polynomials not vanishing in a disk. Applied Mathematics and Computation, 218(3):949–955, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001123.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Liu:2012:EGA

[LMS^{+}12] Linzhong Liu, Haibo Mu, Yubo Song, Haiyan Luo, Xiaojing Li, and Fang Wu. The equilibrium generalized assignment problem and genetic algorithm. Applied Mathematics and Computation, 218(11):6526–6535, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014846.

Lyuu:2011:LTC

[LMT11] Yuh-Dauh Lyuu, Tak-Man Ma, and Yen-Wu Ti. Linear-time compression of 2-manifold polygon meshes into informationtheoretically optimal number of bits. Applied Mathematics and Computation, 217(21):8432–8437, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004061.

Lin:2012:SRE

[LMY12] Huiqiu Lin, Jixiang Meng, and Weihua Yang. Super restricted edge connectivity of regular graphs with two orbits. Applied Mathematics and Computation, 218(12):6656–6660, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014573.

Lamnii:2013:LQI

[LMZ13] M. Lamnii, H. Mraoui, and A. Zidna. Local quasiinterpolants based on special multivariate quadratic spline space over a refined quadrangulation. Applied Mathematics and Computation, 219(21):10456– 10467, July 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004426.

Liu:2010:SIT

[LN10] Wenjun Liu and Quô´ c-Anh Ngô. Some Iyengartype inequalities on time scales for functions whose second derivatives are bounded. Applied Mathematics and Computation, 216(11):3244–3251, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004777.

Lin:2012:NSN

[LNC12] Yingzhen Lin, Jing Niu, and Minggen Cui. A numerical solution to nonlinear second order threepoint boundary value problems in the reproducing kernel space. Applied Mathematics and Computation, 218(14):7362–7368, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013397.

Lasheras:2014:ESV

[LNdCJV14] F. Sánchez Lasheras, P. J. García Nieto, F. J. de Cos Juez, and J. A. Vilán Vilán. Evolutionary support vector regression algorithm applied to the prediction of the thickness of the chromium layer in a hard chromium plating process. Applied Mathematics and Computation, 227(??):164–170, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011995.

Lindquist:2012:HON

[LNG12] Joseph M. Lindquist, Beny Neta, and Francis X. Giraldo. High-order non-reflecting boundary conditions for dispersive waves in polar coordinates using spectral elements. Applied Mathematics and Computation, 218(12):6666–6676, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014822.

Long:2010:DMP

[LNPS10] Guangqing Long, Gnaneshwar Nelakanti, Bijaya Laxmi Panigrahi, and Mitali Madhumita Sahani. Discrete multi-projection methods for eigen-problems of compact integral operators. Applied Mathematics and Computation, 217(8):3974–3984, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010210.

Loi:2012:GFG

[LO12] Nguyen Van Loi and Valeri Obukhovskii. Guiding functions for generalized periodic problems and applications. Applied Mathematics and Computation, 218(24):11719–11726, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005590.

Loi:2013:CDE

[LO13] Nguyen Van Loi and Valeri Obukhovskii. On controllability of Duffing equation. Applied Mathematics and Computation, 219(21):10468– 10474, July 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004517.

Luis:2014:PIM

[LO14] Rafael Luís and Henrique M. Oliveira. Products of 2 × 2 matrices related to non autonomous Fibonacci difference equations. Applied Mathematics and Computation, 226(??):101–116, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010552.

Lin:2012:JOO

[LOD12] Yu-Jen Lin, Liang-Yuh Ouyang, and Ya-Fang Dang. A joint optimal ordering and delivery policy for an integrated supplierretailer inventory model with trade credit and defective items. Applied Mathematics and Computation, 218(14):7498–7514, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000495.

Lahrouz:2012:CGS

[LOKB12] Aadil Lahrouz, Lahcen Omari, Driss Kiouach, and Aziza Belmaâti. Complete global stability for an SIRS epidemic model with generalized nonlinear incidence and vaccination. Applied Mathematics and Computation, 218(11):6519–6525, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014834.

Lolli:2012:IIH

[Lol12] Gabriele Lolli. Infinitesimals and infinites in the history of mathematics: a brief survey. Applied Mathematics and Computation, 218(16):7979–7988, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311011337.

Liu:2013:PCT

[LOS13] Ming-Sheng Liu, Shigeyoshi Owa, and Nian-Sheng Song. Properties of certain transforms defined by convolution of analytic functions. Applied Mathematics and Computation, 219(9):4702–4709, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011010.

Lotfy:2014:TTG

[Lot14] Kh. Lotfy. Two temperature generalized magnetothermoelastic interactions in an elastic medium under three theories. Applied Mathematics and Computation, 227(??):871–888, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012320.

Lee:2010:PIP

[LP10] Sang-Choel Lee and Ju H.

Park. Performance improvement of PI controller with nonlinear error shaping function: IDA-PBC approach. Applied Mathematics and Computation, 215(10):3620–3630, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009746.

Landi:2011:QNP

[LP11] G. Landi and E. Loli Piccolomini. Quasi-Newton projection methods and the discrepancy principle in image restoration. Applied Mathematics and Computation, 218(5):2091–2107, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009593.

Lahooti:2012:NFO

[LP12a] Mohsen Lahooti and Ahmadreza Pishevar. A new fourth order central WENO method for 3D hyperbolic conservation laws. Applied Mathematics and Computation, 218 (20):10258–10270, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003864.

Livieris:2012:GCM

[LP12b] Ioannis E. Livieris and Panagiotis Pintelas. Globally convergent modified Perry’s conjugate gradient method. Applied Mathematics and Computation, 218(18):9197–9207, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002263.

Livieris:2013:NCG

[LP13a] Ioannis E. Livieris and Panagiotis Pintelas. A new conjugate gradient algorithm for training neural networks based on a modified secant equation. Applied Mathematics and Computation, 221(??):491–502, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007558.

Lizama:2013:RMS

[LP13b] Carlos Lizama and Felipe Poblete. Regularity of mild solutions for a class of fractional order differential equations. Applied Mathematics and Computation, 224(??):803–816, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300979X.

Locke:2013:MDE

[LP13c] Judson B. Locke and Adrian M. Peter. Multiwavelet density estimation. Applied Mathematics and Computation, 219(11):6002–6015, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012921.

Lopez:2013:CMC

[LP13d] Rafael López and Juncheol Pyo. Constant mean curvature surfaces with boundary on a sphere. Applied Mathematics and Computation, 220(??):316–323, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006498.

Lee:2014:SOR

[LP14] Won Il Lee and PooGyeon Park. Second-order reciprocally convex approach to stability of systems with interval time-varying delays. Applied Mathematics and Computation, 229(??):245–253, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013015.

Lazaro:2012:CEP

[LPAE12] Mario Lázaro, José L.

Pérez-Aparicio, and Marcelo Epstein. Computation of eigenvalues in proportionally damped viscoelastic structures based on the fixed-point iteration. Applied Mathematics and Computation, 219(8):3511–3529, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009253.

Leem:2010:FPI

[LPB10] Koung Hee Leem, George Pelekanos, and Fermín S. Viloche Bazán. Fixedpoint iterations in determining a Tikhonov regularization parameter in Kirsch’s factorization method. Applied Mathematics and Computation, 216(12):3747–3753, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310005801.

Lee:2011:EHD

[LPH11] Moon Ho Lee, Subash Shree Pokhrel, and Jia Hou. Euler Hadamard/ DCT polynomial matrix. Applied Mathematics and Computation, 217(23):9845–9848, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005789.

Lee:2012:GCS

[LPJ^{+}12] Tae H. Lee, Ju H. Park, D. H. Ji, O. M. Kwon, and S. M. Lee. Guaranteed cost synchronization of a complex dynamical network via dynamic feedback control. Applied Mathematics and Computation, 218(11):6469–6481, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014767.

Lakshmanan:2012:DSE

[LPJB12] S. Lakshmanan, Ju H.

Park, H. Y. Jung, and P. Balasubramaniam. Design of state estimator for neural networks with leakage, discrete and distributed delays. Applied Mathematics and Computation, 218(22):11297– 11310, July 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312005073.

Lakshmanan:2013:SCB

[LPL^{+}13] S. Lakshmanan, Ju H.

Park, Tae H. Lee, H. Y. Jung, and R. Rakkiyappan. Stability criteria for BAM neural networks with leakage delays and probabilistic time-varying delays. Applied Mathematics and Computation, 219(17):9408–9423, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003317.

Lima:2012:GPL

[LPP^{+}12] Angélica Nakagawa Lima, Eric Allison Philot, David Perahia, Antonio Sérgio Kimus Braz, and Luis P. B. Scott. GANM: a protein-ligand docking approach based on genetic algorithm and normal modes. Applied Mathematics and Computation, 219(2):511–520, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006297.

Lobanov:2014:NAS

[LPPG14] I. S. Lobanov, I. Yu. Popov, A. I. Popov, and T. V. Gerya. Numerical approach to the Stokes problem with high contrasts in viscosity. Applied Mathematics and Computation, 235(??):17–25, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003324.

Layton:2010:TRD

[LPR10] William J. Layton, C. David Pruett, and Leo G. Rebholz. Temporally regularized direct numerical simulation. Applied Mathematics and Computation, 216(12):3728–3738, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310005758.

Li:2012:ESF

[LPX12] Junqing Li, Quanke Pan, and Shengxian Xie. An effective shuffled frogleaping algorithm for multi-objective flexible job shop scheduling problems. Applied Mathematics and Computation, 218(18):9353–9371, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002500.

Lin:2011:FCA

[LPY11] Kuo-Ping Lin, Ping-Feng Pai, and Shun-Ling Yang. Forecasting concentrations of air pollutants by logarithm support vector regression with immune algorithms. Applied Mathematics and Computation, 217(12):5318–5327, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011859.

Luo:2013:ISA

[LQ13] Zhongxuan Luo and Wanfeng Qi. On interpolatory subdivision from approximating subdivision scheme. Applied Mathematics and Computation, 220(??):339–349, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006383.

Liu:2012:AMC

[LqTyCZ12] Zhi Liu, Jie qing Tan, Xiao yan Chen, and Li Zhang. An approximation method to circular arcs. Applied Mathematics and Computation, 219(3):1306–1311, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007382.

Li:2010:CCP

[LQY10] Xiang Li, Zhongfeng Qin, and Lixing Yang. A chance-constrained portfolio selection model with risk constraints. Applied Mathematics and Computation, 217(2):949–951, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310007204.

Liu:2011:PBG

[LQY11] Xiaoji Liu, Yonghui Qin, and Yaoming Yu. Perturbation bounds for the generalized inverse A_{T,S}^{(2)} with respect to the Frobenius norm. Applied Mathematics and Computation, 218(8):4595–4604, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012884.

Qiao:2014:DGF

[lQZWC14] Dian liang Qiao, Peng Zhang, S. C. Wong, and Keechoo Choi. Discontinuous Galerkin finite element scheme for a conserved higher-order traffic flow model by exploring Riemann solvers. Applied Mathematics and Computation, 244(??):567–576, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400959X.

Lin:2010:BII

[LR10] Qun Lin and J. G.

Rokne. Blending implicit interval curves and surfaces. Applied Mathematics and Computation, 215(9):3249–3264, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009126.

Leard:2011:APP

[LR11] Benjamin Leard and Jorge Rebaza. Analysis of predator–prey models with continuous threshold harvesting. Applied Mathematics and Computation, 217(12):5265–5278, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011793.

Leyva-Ramos:2013:NMD

[LR13a] J. Leyva-Ramos. A note on mode decoupling of linear time-invariant systems using the generalized sign matrix. Applied Mathematics and Computation, 219(22):10817–10821, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004803.

Llibre:2013:NPO

[LR13b] Jaume Llibre and Ana Rodrigues. A note on the periodic orbits of a kind of Duffing equations. Applied Mathematics and Computation, 219(15):8358–8365, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012544.

Louis-Rose:2013:SNC

[LR13c] Carole Louis-Rose. Simultaneous null controllability with constraint on the control. Applied Mathematics and Computation, 219(11):6372–6392, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012982.

Li:2014:RME

[LR14a] Xindong Li and Hongxing Rui. A rectangular mixed element method with continuous flux approximation for coupling Stokes and Darcy flows. Applied Mathematics and Computation, 246(??):39– 53, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314009461.

Liu:2014:LCB

[LR14b] Yuanyuan Liu and Valery G. Romanovski. Limit cycle bifurcations in a class of piecewise smooth systems with a double homoclinic loop. Applied Mathematics and Computation, 248(??):235–245, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013605.

Liu:2012:NPA

[LRM12] Jun Liu, Xuemei Ren, and Hongbin Ma. A new PSO algorithm with Random C/D Switchings. Applied Mathematics and Computation, 218(19):9579–9593, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001993.

Li:2010:ITO

[LS10a] Songxiao Li and Stevo Stević. On an integraltype operator from ω-Bloch spaces to μ-Zygmund spaces. Applied Mathematics and Computation, 215(12):4385–4391, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000081.

Li:2010:PCD

[LS10b] Songxiao Li and Stevo Stević. Products of composition and differentiation operators from Zygmund spaces to Bloch spaces and Bers spaces. Applied Mathematics and Computation, 217(7):3144– 3154, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310009112.

Liang:2010:PPS

[LS10c] Ruixi Liang and Jianhua Shen. Positive periodic solutions for impulsive predator–prey model with dispersion and time delays. Applied Mathematics and Computation, 217(2):661–676, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006776.

Liu:2011:FNS

[LS11a] Zhongyi Liu and Wenyu Sun. A full-NT-step infeasible interior-point algorithm for SDP based on kernel functions. Applied Mathematics and Computation, 217(10):4990–4999, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011781.

Llanas:2011:RCA

[LS11b] Bernardo Llanas and Francisco J. Sáinz. Representation complexity of adaptive 3D distance fields. Applied Mathematics and Computation, 217(15):6709–6732, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100107X.

Luo:2011:SGA

[LS11c] Qiu-Ming Luo and H. M. Srivastava. Some generalizations of the Apostol– Genocchi polynomials and the Stirling numbers of the second kind. Applied Mathematics and Computation, 217(12):5702–5728, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310012506.

Li:2012:CHO

[LS12a] Fangfei Li and Jitao Sun.

Controllability of higher order Boolean control networks. Applied Mathematics and Computation, 219(1):158–169, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005759.

Liu:2012:GAF

[LS12b] Wanping Liu and Stevo Stević. Global attractivity of a family of nonautonomous max-type difference equations. Applied Mathematics and Computation, 218(11):6297–6303, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014548.

Lopez:2012:TPT

[LS12c] J. L. López and Ester Pérez Sinusía. Twopoint Taylor approximations of the solutions of two-dimensional boundary value problems. Applied Mathematics and Computation, 218(18):9107–9115, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002007.

Lahrouz:2013:APS

[LS13a] Aadil Lahrouz and Adel Settati. Asymptotic properties of switching diffusion epidemic model with varying population size. Applied Mathematics and Computation, 219(24):11134–11148, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005389.

Li:2013:BES

[LS13b] Jibin Li and Jianping Shi. Bifurcations and exact solutions of ac-driven complex Ginzburg–Landau equation. Applied Mathematics and Computation, 221(??):102–110, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006425.

Llibre:2013:DCK

[LS13c] Jaume Llibre and Tayeb Salhi. On the dynamics of a class of Kolmogorov systems. Applied Mathematics and Computation, 225(??):242–245, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009892.

Lahrouz:2014:NSC

[LS14a] Aadil Lahrouz and Adel Settati. Necessary and sufficient condition for extinction and persistence of SIRS system with random perturbation. Applied Mathematics and Computation, 233(??):10–19, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002045.

Liu:2014:SMS

[LS14b] Xinzhi Liu and Peter Stechlinski. SIS models with switching and pulse control. Applied Mathematics and Computation, 232(??):727–742, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013908.

Lopez:2014:NSE

[LS14c] José L. López and Ester Pérez Sinusía. New series expansions for the confluent hypergeometric function M(a, b, z). Applied Mathematics and Computation, 235(??):26– 31, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003476.

Luo:2014:EHA

[LS14d] Zhikun Luo and Huafei Sun. Extended Hamiltonian algorithm for the solution of discrete algebraic Lyapunov equations. Applied Mathematics and Computation, 234(??):245–252, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002768.

Lupu:2014:ALS

[LS14e] Cezar Lupu and Dan Schwarz. Another look at some new Cauchy– Schwarz type inner product inequalities. Applied Mathematics and Computation, 231(??):463–477, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012150.

Louda:2013:NST

[LSF^{+}13] Petr Louda, Petr Svácek, Jaroslav Fort, Jirí Fürst, Jan Halama, and Karel Kozel. Numerical simulation of turbine cascade flow with blade-fluid heat exchange. Applied Mathematics and Computation, 219(13):7206–7214, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007880.

Li:2010:BCE

[LSJ10] Li Li, Gui-Quan Sun, and Zhen Jin. Bifurcation and chaos in an epidemic model with nonlinear incidence rates. Applied Mathematics and Computation, 216(4):1226–1234, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310001803.

Li:2011:MIM

[LSJ11] Fangfei Li, Jitao Sun, and Mei Jia. Monotone iterative method for the second-order three-point boundary value problem with upper and lower solutions in the reversed order. Applied Mathematics and Computation, 217(9):4840–4847, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031001115X.

Liu:2013:ATD

[LSJZ13] Maoxing Liu, Guiquan Sun, Zhen Jin, and Tao Zhou. An analysis of transmission dynamics of drug-resistant disease on scale-free networks. Applied Mathematics and Computation, 222(??):177–189, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300773X.

Liang:2010:CAS

[LSL10] Hui Liang, Dongyang Shi, and Wanjin Lv. Convergence and asymptotic stability of Galerkin methods for a partial differential equation with piecewise constant argument. Applied Mathematics and Computation, 217(2):854–860, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007137.

Liu:2014:SCD

[LSL^{+}14] Bo Liu, Housheng Su, Rong Li, Dehui Sun, and Weina Hu. Switching controllability of discrete-time multi-agent systems with multiple leaders and timedelays. Applied Mathematics and Computation, 228(??):571–588, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012964.

Li:2014:IMH

[LSLL14] Feng Li, Juntao Sun, Gangfu Lu, and Chengjun Lv. Infinitely many homoclinic solutions for a nonperiodic fourth-order differential equation without (AR)-condition. Applied Mathematics and Computation, 241(??):36–41, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006079.

Labbe:2011:GSF

[LSS11] Chantal Labbé, Hristo S. Sendov, and Kristina P. Sendova. The Gerber– Shiu function and the generalized Cramér–Lundberg model. Applied Mathematics and Computation, 218(7):3035–3056, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006977.

Li:2014:REI

[LSSZ14] Bo Li, Hao Shen, Xiaona Song, and Junjie Zhao. Robust exponential H_{∞} control for uncertain time-varying delay systems with input saturation: a Markov jump model approach. Applied Mathematics and Computation, 237(??):190–202, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300314004500.

Li:2014:TDM

[LSW^{+}14a] Ming-Tao Li, Gui-Quan Sun, Yan-Fang Wu, Juan Zhang, and Zhen Jin. Transmission dynamics of a multi-group brucellosis model with mixed cross infection in public farm. Applied Mathematics and Computation, 237(??):582–594, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004561.

Liu:2014:EHO

[LSW14b] Changying Liu, Wei Shi, and Xinyuan Wu. An efficient high-order explicit scheme for solving Hamiltonian nonlinear wave equations. Applied Mathematics and Computation, 246(??):696–710, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011874.

Li:2010:WMP

[LSX10] Song-Hua Li, Ming-Bao Sun, and Jun Xian. A wavelets method for plane elasticity problem. Applied Mathematics and Computation, 216(10):3035–3043, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004418.

Li:2014:HHT

[LSY14] Dong-Qing Li, Xiao-Qiu Song, and Tian Yue. Hermite-Hadamard type inequality for Sugeno integrals. Applied Mathematics and Computation, 237(??):632–638, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005153.

Li:2014:GLT

[LSYS14] Dong-Qing Li, Xiao-Qiu Song, Tian Yue, and Ya-Zhi Song. Generalization of the Lyapunov type inequality for pseudointegrals. Applied Mathematics and Computation, 241(??):64–69, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006626.

Li:2012:SOC

[LSZ12] S. J. Li, X. K. Sun, and J. Zhai. Second-order contingent derivatives of setvalued mappings with application to set-valued optimization. Applied Mathematics and Computation, 218(12):6874–6886, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015517.

Li:2011:PAG

[LT11] Jihong Li and Naishuo Tian. Performance analysis of a GI/M/1 queue with single working vacation. Applied Mathematics and Computation, 217(10):4960–4971, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011744.

Liu:2012:RMI

[LT12] Fuh-Hwa Franklin Liu and Ling-Chuan Tsai. Reallocating multiple inputs and outputs of units to improve overall performance. Applied Mathematics and Computation, 219(1):237–247, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005954.

Liu:2014:LTI

[LT14] Xin-Ge Liu and Mei-Lan Tang. Lyapunovtype inequality for higher order difference equations. Applied Mathematics and Computation, 232(??):666–669, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001775.

Li:2012:GTB

[LTG^{+}12] Hong-Zhe Li, Bo Tian, Rui Guo, Yu-Shan Xue, and Feng-Hua Qi. Gauge transformation between the first-order nonisospectral and isospectral Heisenberg hierarchies. Applied Mathematics and Computation, 218(15):7694–7699, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000513.

Li:2014:IKH

[LTHG14] Junpeng Li, Yinggan Tang, Changchun Hua, and Xinping Guan. An improved krill herd algorithm: Krill herd with linear decreasing step. Applied Mathematics and Computation, 234(??):356–367, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001921.

Liu:2011:SSB

[LTJ^{+}11] Wen-Jun Liu, Bo Tian, Yan Jiang, Kun Sun, Pan Wang, Min Li, and Qi-Xing Qu. Soliton solutions and Bäcklund transformation for the complex Ginzburg–Landau equation. Applied Mathematics and Computation, 217(9):4369–4376, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010519.

Li:2011:CCA

[LTL11] Yanyan Li, Tao Tan, and Xingsi Li. Convergence of a continuous approach for zero-one programming problems. Applied Mathematics and Computation, 217(9):4691–4698, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011343.

Li:2012:LES

[LTL12] Yanyan Li, Tao Tan, and Xingsi Li. A logexponential smoothing method for mathematical programs with complementarity constraints. Applied Mathematics and Computation, 218(10): 5900–5909, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013890.

Liu:2010:NMO

[LTM10] Sanyang Liu, Jia Tang, and Changfeng Ma. A new modified one-step smoothing Newton method for solving the general mixed complementarity problem. Applied Mathematics and Computation, 216(4):1140–1149, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001724.

Liao:2011:DPC

[LTOX11] Maoxin Liao, Xianhua Tang, Zigen Ouyang, and Changjin Xu. Dynamical properties of a class of higher-order nonlinear difference equations. Applied Mathematics and Computation, 217(12):5476–5479, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012178.

Loxton:2011:RSC

[LTR11] R. Loxton, K. L. Teo, and V. Rehbock. Robust suboptimal control of nonlinear systems. Applied Mathematics and Computation, 217(14):6566–6576, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000555.

Lone:2014:NAR

[LTSV14] Hilal A. Lone, Rajesh Tailor, Housila P. Singh, and Med Ram Verma. New alternatives to ratio estimators of population variance in sample surveys. Applied Mathematics and Computation, 247(??):255–265, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401203X.

Lin:2013:SLF

[LTT^{+}13] Cheng-Kuan Lin, Yuan-Hsiang Teng, Jimmy J. M. Tan, Lih-Hsing Hsu, and Dragan Marusic. The spanning laceability on the faulty bipartite hypercube-like networks. Applied Mathematics and Computation, 219(15):8095–8103, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001495.

Liu:2013:QSA

[LTU13] Yiliang Liu, Salih Tatar, and Suleyman Ulusoy. Quasi-solution approach for a two dimensional nonlinear inverse diffusion problem. Applied Mathematics and Computation, 219(23):10956–10960, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313005328.

Liu:2010:AEM

[LTW10] Xun Liu, Lixin Tian, and Yuhai Wu. Application of \frac{G′}{G} -expansion method to two nonlinear evolution equations. Applied Mathematics and Computation, 217(4):1376–1384, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309005086.

Liu:2012:ESF

[LTW12] Xun Liu, Lixin Tian, and Yuhai Wu. Exact solutions of four generalized Benjamin–Bona– Mahony equations with any order. Applied Mathematics and Computation, 218(17):8602–8613, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001452.

Liao:2011:DCL

[LTX11] Maoxin Liao, Xianhua Tang, and Changjin Xu. Dynamics of a competitive Lotka–Volterra system with three delays. Applied Mathematics and Computation, 217(24):10024–10034, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006485.

Li:2011:PSE

[LTY11] Chun-Hsien Li, Chiung-Chiou Tsai, and Suh-Yuh Yang. Permanence of an SIR epidemic model with density dependent birth rate and distributed time delay. Applied Mathematics and Computation, 218(5):1682–1693, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008800.

Li:2012:EMA

[LTZ12] Wen-Wu Li, Ye Tian, and Zhe Zhang. F -expansion method and its application for finding new exact solutions to the sine-Gordon and sinh-Gordon equations. Applied Mathematics and Computation, 219(3):1135–1143, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312007217.

Lanzagorta:2012:ESF

[LU12a] Marco Lanzagorta and Jeffrey Uhlmann. Error scaling in fault tolerant quantum computation. Applied Mathematics and Computation, 219(1):24–30, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010198.

Lu:2012:DTC

[Lu12b] Bin Lu. Division theorem combined with the Riccati equation for solving some nonlinear Schrödinger-like equations. Applied Mathematics and Computation, 219(4):1686–1694, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008016.

Lu:2012:GTH

[Lu12c] Da-Qian Lu. Generalized Tricomi and Hermite– Tricomi functions. Applied Mathematics and Computation, 218(9):5090–5098, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013208.

Lu:2012:NIP

[Lu12d] Lizheng Lu. A note on iterative process for G^{2}- multi degree reduction of Bézier curves. Applied Mathematics and Computation, 218(12):6987–6990, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015426.

Lu:2014:SNI

[Lu14a] Dawei Lu. Some new improved classes of convergence towards Euler’s constant. Applied Mathematics and Computation, 243(??):24–32, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400798X.

Lu:2014:SQC

[Lu14b] Dawei Lu. Some quicker classes of sequences convergent to Euler’s constant. Applied Mathematics and Computation, 232(??):172–177, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001052.

Lu:2014:GPB

[Lu14c] Junfeng Lu. A generalization of parameterized block triangular preconditioners for generalized saddle point problems. Applied Mathematics and Computation, 241(??):25–35, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006353.

Luo:2014:ROE

[Luo14]  Zhendong Luo.  A

reduced-order extrapolation algorithm based on SFVE method and POD technique for nonstationary Stokes equations. Applied Mathematics and Computation, 247(??):976–995, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401282X.

Luksan:2014:ETP

[LV14] Ladislav Luksan and Jan Vlcek. Efficient tridiagonal preconditioner for the matrix-free truncated Newton method. Applied Mathematics and Computation, 235(??):394–407, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003592.

Li:2006:CAS

[LW06] Zhenqing Li and Weiming Wang. Computer aided solving the highorder transition probability matrix of the finite Markov chain. Applied Mathematics and Computation, 172(1):267–285, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300305002031.

See erratum [ESAH10].

Liu:2010:EOM

[LW10a] Liping Liu and Xia Wang. Eighth-order methods with high efficiency index for solving nonlinear equations. Applied Mathematics and Computation, 215(9):3449–3454, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030900945X.

Luo:2010:TPW

[LW10b] Lin Luo and Luo Wu.

Two-periodic waves and asymptotic property for generalized 2D Toda lattice equation. Applied Mathematics and Computation, 217(6):2821–2826, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310008714.

Lewanowicz:2011:BRC

[LW11a] Stanisław Lewanowicz and Paweł Woźny. Bézier representation of the constrained dual Bernstein polynomials. Applied Mathematics and Computation, 218(8):4580–4586, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311012859.

Lewanowicz:2011:MDR

[LW11b] Stanisław Lewanowicz and Paweł Woźny. Multidegree reduction of tensor product Bézier surfaces with general boundary constraints. Applied Mathematics and Computation, 217(9):4596–4611, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011239.

Lin:2011:CMS

[LW11c] S. H. Lin and I. M. Wu. On the common mean of several inverse Gaussian distributions based on a higher order likelihood method. Applied Mathematics and Computation, 217(12):5480–5490, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012191.

Li:2012:LTL

[LW12a] Peijin Li and Xiantao Wang. Landau’s theorem for log-p-harmonic mappings. Applied Mathematics and Computation, 218(9):4806–4812, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311012835.

Liu:2012:CIF

[LW12b] Xi-Lan Liu and Xiao-Rui Wang. A convergence improvement factor and higher-order methods for solving nonlinear equations. Applied Mathematics and Computation, 218(15):7871–7875, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001130.

Lesmana:2013:UFD

[LW13a] Donny C. Lesmana and Song Wang. An upwind finite difference method for a nonlinear Black– Scholes equation governing European option valuation under transaction costs. Applied Mathematics and Computation, 219(16):8811–8828, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000180.

Li:2013:DPPa

[LW13b] Li Li and Zhi-Jun Wang.

Dynamics in a predatorprey model with space and noise. Applied Mathematics and Computation, 219(12):6542–6547, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000362.

Liu:2013:QRR

[LW13c] Ji-Chuan Liu and Ting Wei. A quasi-reversibility regularization method for an inverse heat conduction problem without initial data. Applied Mathematics and Computation, 219(23):10866–10881, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005286.

Liu:2013:PCI

[LW13d] Wen Liu and Chuansheng Wu. A predictor-corrector iterated Tikhonov regularization for linear ill-posed inverse problems. Applied Mathematics and Computation, 221(??):802–818, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313007297.

Lang:2014:SBS

[LW14a] Weiwei Lang and Ligong Wang. Sharp bounds for the signless Laplacian spectral radius of digraphs. Applied Mathematics and Computation, 238(??):43–49, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314005244.

Li:2014:SNE

[LW14b] Cui-Xia Li and Shi-Liang Wu. Some new estimates on the complex eigenvalues of the HSS preconditioned matrix. Applied Mathematics and Computation, 248(??):519–524, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013666.

Li:2014:CCL

[LW14c] Feng Li and Yusen Wu.

Center conditions and limit cycles for a class of nilpotent-Poincarè systems. Applied Mathematics and Computation, 243(??):114–120, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008212.

Li:2014:MGM

[LW14d] Wen-Wei Li and Xiang Wang. A modified GPSS method for non-Hermitian positive definite linear systems. Applied Mathematics and Computation, 234(??):253–259, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002847.

Li:2014:SBP

[LW14e] Yuan-Min Li and Deyun Wei. Solvability based on E-property for the nonlinear symmetric cone complementarity problem. Applied Mathematics and Computation, 236(??):437–449, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004111.

Liang:2014:MQC

[LW14f] Yi Liang and Xingyuan Wang. A method of quickly calculating the number of pinning nodes on pinning synchronization in complex networks. Applied Mathematics and Computation, 246(??):743–751, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011758.

Liu:2014:QMF

[LW14g] Zhenhai Liu and Rui Wang. Quasilinearization method for fractional differential equations with delayed arguments. Applied Mathematics and Computation, 248(??):301–308, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013629.

Liu:2014:APC

[LW14h] Zhijun Liu and Qinglong Wang. An almost periodic competitive system subject to impulsive perturbations. Applied Mathematics and Computation, 231(??):377–385, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314000538.

Li:2014:ESL

[LWB14] Zunxian Li, Peixuan Weng, and Jianxia Bai. Entire solutions of a lattice vector disease model. Applied Mathematics and Computation, 241(??):129–139, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006754.

Liang:2014:TFE

[LWCS14] Bo Liang, Meishan Wang, Yang Cao, and Huiying Shen. A thin film equation with a singular diffusion. Applied Mathematics and Computation, 227(??):1–10, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011545.

Lin:2010:NET

[LWD10] Songqing Lin, Chuanjian Wang, and Zhengde Dai. New exact traveling and non-traveling wave solutions for (2 + 1)- dimensional Burgers equation. Applied Mathematics and Computation, 216(10):3105–3110, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003917.

Lin:2011:FFM

[LWF11] Hongwei Lin, Yuping Wang, and Lei Fan. A filled function method with one parameter for unconstrained global optimization. Applied Mathematics and Computation, 218(7):3776–3785, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101188X.

Lin:2013:NDF

[LWFG13] Hongwei Lin, Yuping Wang, Lei Fan, and Yuelin Gao. A new discrete filled function method for finding global minimizer of the integer programming. Applied Mathematics and Computation, 219(9):4371–4378, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312010363.

Liu:2013:SANb

[LWG13] Defang Liu, Bochu Wang, and Songtao Guo. Stability analysis of a novel epidemics model with vaccination and nonlinear infectious rate. Applied Mathematics and Computation, 221(??):786–801, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007662.

Ling:2014:NLM

[LWH14] Chen Ling, Guifeng Wang, and Hongjin He. A new Levenberg–Marquardt type algorithm for solving nonsmooth constrained equations. Applied Mathematics and Computation, 229(??):107–122, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012915.

Li:2010:PAN

[LWL10] Xiao-Nan Li, Guang-Mei Wei, and Yue-Qian Liang. Painlevé analysis and new analytic solutions for variablecoefficient Kadomtsev– Petviashvili equation with symbolic computation. Applied Mathematics and Computation, 216(12): 3568–3577, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310005199.

Li:2012:OIC

[LWL12] Xuefeng Li, Zhaolu Wang, and Hongjun Liu. Optimizing initial chirp for efficient femtosecond wavelength conversion in silicon waveguide by split-step Fourier method. Applied Mathematics and Computation, 218(24):11970– 11975, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312005863.

Liang:2014:PCD

[LWL14] Xue Liang, Guojing Wang, and Hong Li. Pricing credit default swaps with bilateral counterparty risk in a reduced form model with Markov regime switching. Applied Mathematics and Computation, 230(??):290–302, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313014057.

Leopold-Wildburger:2010:BFW

[LWM10] Ulrike Leopold-Wildburger and Arleta Mietek. Bonus or flat wage? An experiment into the principal– agent problem. Applied Mathematics and Computation, 217(3):1141–1149, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006119.

Lu:2011:SGC

[LWM11] Sha Lu, Zengxin Wei, and Liliu Mo. Some global convergence properties of the Wei–Yao–Liu conjugate gradient method with inexact line search. Applied Mathematics and Computation, 217(17):7132–7137, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311001366.

Lewanowicz:2013:SRB

[LWN13] Stanisław Lewanowicz, Paweł Woźny, and Rafal Nowak. Structure relations for the bivariate big q-jacobi polynomials. Applied Mathematics and Computation, 219(16):8790–8802, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001975.

Liu:2014:HZC

[LWN14a] Hao Liu, Haijun Wang, and Qin Ni. On Hager and Zhang’s conjugate gradient method with guaranteed descent. Applied Mathematics and Computation, 236(??):400–407, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003282.

Long:2014:IFM

[LWN14b] Guangqing Long, Weifen Wu, and Gnaneshwar Nelakanti. Iterated fast multiscale Galerkin methods for eigen-problems of compact integral operators. Applied Mathematics and Computation, 246(??):638–647, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011679.

Lee:2012:SCD

[LWP12] Tae H. Lee, Zheng-Guang Wu, and Ju H. Park. Synchronization of a complex dynamical network with coupling time-varying delays via sampled-data control. Applied Mathematics and Computation, 219(3):1354–1366, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007588.

Li:2010:NLG

[LWQ10] G. Q. Li, Z. Y. Wu, and J. Quan. A new local and global optimization method for mixed integer quadratic programming problems. Applied Mathematics and Computation, 217(6):2501–2512, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008076.

See note [Pak11b].

Li:2011:OHP

[LWS11a] Wenxue Li, Ke Wang, and Huan Su. Optimal harvesting policy for stochastic Logistic population model. Applied Mathematics and Computation, 218(1):157–162, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007600.

Liu:2011:ACU

[LWS11b] Yongguo Liu, Xindong Wu, and Yidong Shen. Automatic clustering using genetic algorithms. Applied Mathematics and Computation, 218(4):1267– 1279, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311008204.

Liu:2014:IAG

[LWS14] Zhongyi Liu, Zhihui Wei, and Wenyu Sun. An iteratively approximated gradient projection algorithm for sparse signal reconstruction. Applied Mathematics and Computation, 228(??):454–462, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031301120X.

Liu:2010:MAP

[LWTC10] Zhijun Liu, Jianhua Wu, Ronghua Tan, and Yiping Chen. Modeling and analysis of a periodic delayed two-species model of facultative mutualism. Applied Mathematics and Computation, 217(2):893–903, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007186.

Li:2010:ECP

[LWW10] Jianbing Li, Xuesong Wang, and Tao Wang. Evaluation of Cauchy principal value integrals of oscillatory kind. Applied Mathematics and Computation, 217(6):2390–2396, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007848.

Li:2011:SNM

[LWW11] Yuan Min Li, Xing Tao Wang, and De Yun Wei. A smoothing Newton method for NCPs with the P_{0}-property. Applied Mathematics and Computation, 217(16):6917–6925, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100138X.

Liu:2012:AIE

[LWW12] Meng Liu, Qiong Wu, and Ke Wang. Analysis of an improved epidemic model with stochastic disease transmission. Applied Mathematics and Computation, 218(19):9750–9758, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002585.

Lu:2014:SMG

[LWW14] Yuan-Yuan Lu, Jian-Jun Wang, and Ji-Bo Wang. Single machine group scheduling with decreasing time-dependent processing times subject to release dates. Applied Mathematics and Computation, 234(??):286–292, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002215.

Liu:2010:SOP

[LWWL10] Zhao-Hua Liu, Ai-Ling Wang, Biao Wang, and Yong-Jiang Liu. Spatial organization in a plant model. Applied Mathematics and Computation, 216(3):983–988, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001402.

Li:2011:FRR

[LWWZ11] Xiezhang Li, Hua Wang, Yan Wu, and Jiehua Zhu. A full row-rank system matrix generated along two directions in discrete tomography. Applied Mathematics and Computation, 218(1):107–114, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311007284.

Luo:2013:MAB

[LWX13] Jun Luo, Qian Wang, and Xianghai Xiao. A modified artificial bee colony algorithm based on converge-onlookers approach for global optimization. Applied Mathematics and Computation, 219(20):10253– 10262, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003895.

Li:2014:PPA

[LWY14] Jing Li, Shiying Wang, and Yuxing Yang. Panconnectivity and pancyclicity of the 3-ary n-cube network under the path restrictions. Applied Mathematics and Computation, 243(??):339–348, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400825X.

Li:2014:MAP

[LWYY14] Xu Li, Yu-Jiang Wu, Ai-Li Yang, and Jin-Yun Yuan. Modified accelerated parameterized inexact Uzawa method for singular and nonsingular saddle point problems. Applied Mathematics and Computation, 244(??):552–560, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314009886.

Li:2011:DCD

[LWZ11] Cai-Yun Li, Ren-Hong Wang, and Chun-Gang Zhu. Design and G^{1} connection of developable surfaces through Bézier geodesics. Applied Mathematics and Computation, 218(7):3199–3208, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010976.

Li:2013:GSF

[LWZ13a] Cai-Yun Li, Ren-Hong Wang, and Chun-Gang Zhu. A generalization of surface family with common line of curvature. Applied Mathematics and Computation, 219(17):9500–9507, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300338X.

Liu:2013:BSF

[LWZ13b] Hai-Tao Liu, Wen-Juan Wu, and Zhen-Gong Zhou. Basic solution to four three-dimensional rectangular limited-permeable cracks in transversely isotropic magneto-electro-elastic material. Applied Mathematics and Computation, 225(??):117–141, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313010047.

Li:2014:NRE

[LWZ14] Xiangyu Li, Jinting Wang, and Feng Zhang. New results on equilibrium balking strategies in the singleserver queue with breakdowns and repairs. Applied Mathematics and Computation, 241(??):380–388, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006900.

Li:2010:LSS

[LWZD10a] Zhao-Yan Li, Yong Wang, Bin Zhou, and Guang-Ren Duan. Least squares solution with the minimumnorm to general matrix equations via iteration. Applied Mathematics and Computation, 215(10):3547–3562, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009680.

Li:2010:UCD

[LWZD10b] Zhao-Yan Li, Yong Wang, Bin Zhou, and Guang-Ren Duan. On unified concepts of detectability and observability for continuous-time stochastic systems. Applied Mathematics and Computation, 217(2):521–536, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006557.

Liu:2014:MNS

[LWZD14] Jinsheng Liu, Shuli Wang, Jie Zhao, and Chunya Duan. Multiple nontrivial solutions for asymptotically linear discrete problem via computations of the critical groups. Applied Mathematics and Computation, 247(??): 431–441, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012338.

Li:2011:PBB

[LWZY11] Xiangtao Li, Jinyan Wang, Junping Zhou, and Minghao Yin. A perturb biogeography based optimization with mutation for global numerical optimization. Applied Mathematics and Computation, 218(2):598–609, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007934.

Li:2014:FSP

[LWZZ14] Ying Li, Musheng Wei, Fengxia Zhang, and Jianli Zhao. A fast structurepreserving method for computing the singular value decomposition of quaternion matrices. Applied Mathematics and Computation, 235(??): 157–167, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314003166.

Liu:2011:RBT

[LX11a] Jijun Liu and Huilin Xu.

Reconstruction of biologic tissue conductivity from noisy magnetic field by integral equation method. Applied Mathematics and Computation, 218(6):2647–2660, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010307.

Liu:2011:SBC

[LX11b] Qiming Liu and Rui Xu. Stability and bifurcation of a Cohen– Grossberg neural network with discrete delays. Applied Mathematics and Computation, 218(6):2850–2862, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010551.

Li:2012:EES

[LX12] Bing Li and Daoyi Xu. Existence and exponential stability of periodic solution for impulsive Cohen– Grossberg neural networks with time-varying delays. Applied Mathematics and Computation, 219(5):2506–2520, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200879X.

Li:2014:ODD

[LX14a] Xia Li and Zhengrong Xiang. Observer design of discrete-time impulsive switched nonlinear systems with time-varying delays. Applied Mathematics and Computation, 229(??):327–339, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013362.

Li:2014:FTI

[LX14b] Xiang-Jun Li and Jun-Ming Xu. Fault-tolerance of (n, k)-star networks. Applied Mathematics and Computation, 248(??): 525–530, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031401371X.

Li:2014:DBF

[LX14c] Ya Li and Yumei Xue. Dynamical behavior of a food chain model with prey toxicity. Applied Mathematics and Computation, 242(??):551–561, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314007887.

Liu:2012:SSC

[LXC12] Shilong Liu, Zhengrong Xiang, and Qingwei Chen. Stability and stabilization of a class of switched nonlinear systems with timevarying delay. Applied Mathematics and Computation, 218(23):11534– 11546, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005450.

Li:2014:CDP

[LXC14] Yongfeng Li, Dongliang Xie, and Jingan Cui. Complex dynamics of a predator-prey model with impulsive state feedback control. Applied Mathematics and Computation, 230(??):395–405, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013970.

Liu:2014:EUB

[LXGM14] Jun Liu, Xinping Xiao, Jinhai Guo, and Shuhua Mao. Error and its upper bound estimation between the solutions of GM(1, 1) grey forecasting models. Applied Mathematics and Computation, 246(??):648–660, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011588.

Li:2013:EVM

[LXH13] Yanhong Li, Wangtu Xu, and Shiwei He. Expected value model for optimizing the multiple bus headways. Applied Mathematics and Computation, 219(11):5849–5861, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201291X.

Li:2013:SGC

[LXK13] Shuo Li, Zhengrong Xiang, and Hamid Reza Karimi. Stability and L_{1}-gain controller design for positive switched systems with mixed timevarying delays. Applied Mathematics and Computation, 222(??):507–518, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313008205.

Luo:2014:QWB

[LXLY14] Man Luo, Da Xu, Limei Li, and Xuehua Yang. Quasi wavelet based numerical method for Volterra integro-differential equations on unbounded spatial domains. Applied Mathematics and Computation, 227(??):509–517, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011971.

Li:2013:TOZ

[LXTC13] Bin Li, Chao Xu, Kok Lay Teo, and Jian Chu. Time optimal Zermelo’s navigation problem with moving and fixed obstacles. Applied Mathematics and Computation, 224(??):866–875, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009727.

Li:2014:ICS

[LXWW14] Chao Li, Wei Xu, Liang Wang, and Zhenpei Wang. Impulsive control of sticking motion in van der Pol one-sided constraint system. Applied Mathematics and Computation, 248(??):363–370, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013265.

Liu:2010:RDI

[LXY10] Xiaoji Liu, Liang Xu, and Yaoming Yu. The representations of the Drazin inverse of differences of two matrices. Applied Mathematics and Computation, 216(12):3652–3661, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005357.

Long:2012:QWB

[LXZ12] Wenting Long, Da Xu, and Xueying Zeng. Quasi wavelet based numerical method for a class of partial integro-differential equation. Applied Mathematics and Computation, 218(24):11842–11850, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312005541.

Liu:2014:IWP

[LXZ14] Keji Liu, Yongzhi Xu, and Jun Zou. Imaging wavepenetrable objects in a finite depth ocean. Applied Mathematics and Computation, 235(??):364–376, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002872.

Li:2013:DPPb

[LXZX13] Gendao Li, Zhongkai Xiong, Yu Zhou, and Yu Xiong. Dynamic pricing for perishable products with hybrid uncertainty in demand. Applied Mathematics and Computation, 219(20):10366– 10377, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004074.

Lee:2010:ASS

[LY10a] Yeon Ju Lee and Jungho Yoon. Analysis of stationary subdivision schemes for curve design based on radial basis function interpolation. Applied Mathematics and Computation, 215(11):3851–3859, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010170.

Liu:2010:ROL

[LY10b] Deqiang Liu and Hu Yang.

The reverse order law for {1, 3, 4}-inverse of the product of two matrices. Applied Mathematics and Computation, 215 (12):4293–4303, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309011138.

Lin:2011:MCA

[LY11a] Yi-Kuei Lin and Cheng-Ta Yeh. Multistate components assignment problem with optimal network reliability subject to assignment budget. Applied Mathematics and Computation, 217(24):10074–10086, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006564.

Liu:2011:PPS

[LY11b] Guirong Liu and Jurang Yan. Positive periodic solutions for neutral delay ratio-dependent predatorprey model with Holling type III functional response. Applied Mathematics and Computation, 218(8):4341–4348, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012392.

Lin:2012:NFF

[LY12a] Youjiang Lin and Yongjian Yang. A new filled function method for constrained nonlinear equations. Applied Mathematics and Computation, 219(6):3100–3112, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200940X.

Liu:2012:FDS

[LY12b] Wei Liu and Yirang Yuan. A finite difference scheme for two-dimensional semiconductor device of heat conduction on composite triangular grids. Applied Mathematics and Computation, 218(11):6458–6468, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014755.

Liu:2012:FRG

[LY12c] Xifu Liu and Hu Yang.

Further results on the group inverses and Drazin inverses of anti-triangular block matrices. Applied Mathematics and Computation, 218(17):8978–8986, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001981.

Li:2013:NAM

[LY13a] Jiequan Li and Zhicheng Yang. The von Neumann analysis and modified equation approach for finite difference schemes. Applied Mathematics and Computation, 225(??): 610–621, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313010254.

Li:2013:EAS

[LY13b] Ke Li and Zhijian Yang.

Exponential attractors for the strongly damped wave equation. Applied Mathematics and Computation, 220(??):155–165, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005912.

Li:2014:ESR

[LY14a] Yalian Li and Hu Yang. Efficiency of a stochastic restricted two-parameter estimator in linear regression. Applied Mathematics and Computation, 249(??):371–381, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013770.

Li:2014:AAS

[LY14b] Yongkun Li and Li Yang. Almost automorphic solution for neutral type high-order Hopfield neural networks with delays in leakage terms on time scales. Applied Mathematics and Computation, 242(??):679–693, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314008881.

Liu:2014:TWI

[LY14c] Guirong Liu and Weiping Yan. Travelling waves of iodate-arsenous-acid reaction with a constant electric field. Applied Mathematics and Computation, 232(??):27–38, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000356.

Liu:2014:BET

[LY14d] Haihong Liu and Fang Yan. Bifurcation and exact travelling wave solutions for Gardner–KP equation. Applied Mathematics and Computation, 228(??):384–394, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012794.

Lv:2014:ETW

[LY14e] Yunfei Lv and Rong Yuan. Existence of traveling wave solutions for Gauss-type models of predator-prey systems. Applied Mathematics and Computation, 229(??):70–84, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013076.

Liang:2013:ISC

[LYAH13] Jin Liang, Sheng-Hua Yan, Ravi P. Agarwal, and Ting-Wen Huang. Integral solution of a class of nonlinear integral equations. Applied Mathematics and Computation, 219(10):4950–4957, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011290.

Luo:2012:OOC

[LYB12] Zhi-Xue Luo, Xiao-Di Yu, and Zheng-Gang Ba. Overtaking optimal control problem for an age-dependent competition system of n species. Applied Mathematics and Computation, 218(17):8561–8569, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001415.

Li:2012:SBN

[LYC12] Rui Li, Meng Yang, and Tianguang Chu. Synchronization of Boolean networks with time delays. Applied Mathematics and Computation, 219(3):917–927, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200690X.

Liao:2010:TMM

[LYD10] An-Ping Liao, Guo-Zhu Yao, and Xue-Feng Duan. Thompson metric method for solving a class of nonlinear matrix equation. Applied Mathematics and Computation, 216(6):1831–1836, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010698.

Liu:2014:EDE

[LYG14] Ying Liu, Minghao Yin, and Wenxiang Gu. An effective differential evolution algorithm for permutation flow shop scheduling problem. Applied Mathematics and Computation, 248(??):143–159, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012284.

Li:2013:KIS

[LYGZ13] Zhihua Li, Xue Yang, Wenqiang Gu, and Haitao Zhang. Kernel-improved Support Vector Machine for semanteme data. Applied Mathematics and Computation, 219(17): 8876–8880, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003305.

Lee:2013:ICS

[LYJP13] D. W. Lee, W. J. Yoo, D. H. Ji, and Ju H. Park. Integral control for synchronization of complex dynamical networks with unknown non-identical nodes. Applied Mathematics and Computation, 224(??):140–149, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009272.

Lv:2011:NSA

[LYL11] W. J. Lv, Z. W. Yang, and M. Z. Liu. Numerical stability analysis of differential equations with piecewise constant arguments with complex coefficients. Applied Mathematics and Computation, 218(1):45–54, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007119.

Lin:2013:SDS

[LYL13a] B. M. T. Lin, P. Y. Yin, and Y. S. Liu. Sequencedependent scheduling with order deliveries. Applied Mathematics and Computation, 222(??):58–71, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007364.

Liu:2013:DFT

[LYL13b] Wanping Liu, Xiaofan Yang, and Xinzhi Liu. Dynamics of a family of two-dimensional difference systems. Applied Mathematics and Computation, 219(11):5949–5955, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012258.

Liu:2014:PMA

[LYLS14] Wanping Liu, Xiaofan Yang, Xinzhi Liu, and Stevo Stević. Part-metric and its applications in discrete systems. Applied Mathematics and Computation, 228(??):320–328, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012423.

Lv:2014:EHD

[LYP14] Yunfei Lv, Rong Yuan, and Yongzhen Pei. Effect of harvesting, delay and diffusion in a generalist predator-prey model. Applied Mathematics and Computation, 226(??):348–366, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011284.

Li:2012:WTA

[LYPC12] Shuai Li, Jiguo Yu, Mingming Pan, and Sanfeng Chen. Winner-take-all based on discrete-time dynamic feedback. Applied Mathematics and Computation, 219(4):1569–1575, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312007874.

Li:2010:NPA

[LYS10] Hanyu Li, Hu Yang, and Hua Shao. A note on the perturbation analysis for the generalized Cholesky factorization. Applied Mathematics and Computation, 215(11):4022–4027, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010431.

Li:2012:PAB

[LYS12] Hanyu Li, Hu Yang, and Hua Shao. Perturbation analysis for block downdating of the generalized Cholesky factorization. Applied Mathematics and Computation, 218(18):9451–9461, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312002664.

Li:2012:IDD

[LYWL12] Tao Li, Xiuming Yao, Lingyao Wu, and Jianqing Li. Improved delaydependent stability results of recurrent neural networks. Applied Mathematics and Computation, 218(19):9983–9991, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002457.

Liu:2013:SANa

[LYXP13] Zixin Liu, Jian Yu, Daoyun Xu, and Dingtao Peng. Stable analysis for neural networks: Setvalued mapping method. Applied Mathematics and Computation, 220(??):46– 52, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006073.

Li:2011:NES

[LYZ11] Wei Li, Emmanuel Yomba, and Hongqing Zhang. New exact solutions for some coupled nonlinear partial differential equations using extended coupled sub-equations expansion method. Applied Mathematics and Computation, 217(21):8468–8481, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004140.

Liu:2013:STA

[LYZG13] Peng Liu, Na Yi, Xiaoye Zhou, and Hua Gong. Scheduling two agents with sum-of-processing-times-based deterioration on a single machine. Applied Mathematics and Computation, 219(17):8848–8855, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002932.

Li:2010:SNI

[LZ10a] Dongfang Li and Chengjian Zhang. Split Newton iterative algorithm and its application. Applied Mathematics and Computation, 217(5):2260–2265, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031000771X.

Li:2010:HWO

[LZ10b] Yuanlu Li and Weiwei Zhao. Haar wavelet operational matrix of fractional order integration and its applications in solving the fractional order differential equations. Applied Mathematics and Computation, 216(8):2276–2285, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003176.

Liu:2010:TRT

[LZ10c] Fang Liu and Xian-Ci Zhong. Transient response of two collinear dielectric cracks in a piezoelectric solid under inplane impacts. Applied Mathematics and Computation, 217(8):3779–3791, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009963.

Liu:2010:IRW

[LZ10d] Xiaoji Liu and Jin Zhong. Integral representation of the W -weighted Drazin inverse for Hilbert space operators. Applied Mathematics and Computation, 216(11):3228–3233, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004753.

Liu:2010:PEA

[LZ10e] Zijian Liu and Shouming Zhong. Permanence and extinction analysis for a delayed periodic predator– prey system with Holling type II response function and diffusion. Applied Mathematics and Computation, 216(10):3002–3015, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004327.

Long:2010:NFL

[LZ10f] Jun Long and Sanyun Zeng. A new Filter-Levenberg–Marquardt method with disturbance for solving nonlinear complementarity problems. Applied Mathematics and Computation, 216(2):677–688, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310001220.

Long:2010:PFM

[LZ10g] Jun Long and Sanyun Zeng. A projectionfilter method for solving nonlinear complementarity problems. Applied Mathematics and Computation, 216(1):300–307, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000767.

Li:2011:SWC

[LZ11a] Jibin Li and Yi Zhang. Solitary wave and chaotic behavior of traveling wave solutions for the coupled KdV equations. Applied Mathematics and Computation, 218(5):1794–1797, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311009088.

Li:2011:CCM

[LZ11b] Yun-Zhang Li and Feng-Ying Zhou. The characterization of a class of multivariate MRA and semi-orthogonal Parseval frame wavelets. Applied Mathematics and Computation, 217(22):9151–9164, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005303.

Liu:2011:CNS

[LZ11c] Wei Liu and Qimin Zhang.

Convergence of numerical solutions to stochastic agestructured system of three species. Applied Mathematics and Computation, 218(8):3973–3980, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011453.

Latrach:2012:ERN

[LZ12a] K. Latrach and A. Zeghal.

Existence results for a nonlinear transport equation in bounded geometry on L^{1}-spaces. Applied Mathematics and Computation, 219(3):1163–1172, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312007266.

Li:2012:GBE

[LZ12b] Maojun Li and Jialin Zhu. A Galerkin boundary element method based on interpolatory Hermite trigonometric wavelets. Applied Mathematics and Computation, 218(9):5357– 5369, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031101349X.

Liu:2012:SSD

[LZ12c] Shutang Liu and Yongping Zhang. Stability of stochastic 2-D systems. Applied Mathematics and Computation, 219(1):197–212, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005875.

Li:2013:SOF

[LZ13a] Chenliang Li and Jingwen Zou. A sixthorder fast algorithm for the electromagnetic scattering from large open cavities. Applied Mathematics and Computation, 219(16):8656–8666, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313001355.

Liang:2013:BDP

[LZ13b] Zhao-Zheng Liang and Guo-Feng Zhang. On block-diagonally preconditioned accelerated parameterized inexact Uzawa method for singular saddle point problems. Applied Mathematics and Computation, 221(??):89–101, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313006097.

Liang:2014:MUS

[LZ14a] Zhao-Zheng Liang and Guo-Feng Zhang. Modified unsymmetric SOR method for saddle-point problems. Applied Mathematics and Computation, 234(??):584–598, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003002.

Liang:2014:SCC

[LZ14b] Zhao-Zheng Liang and Guo-Feng Zhang. On semi-convergence of a class of Uzawa methods for singular saddle-point problems. Applied Mathematics and Computation, 247(??):397–409, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012132.

Liu:2014:BIS

[LZ14c] Gongwei Liu and Hongwei Zhang. Blow up at infinity of solutions for integro-differential equation. Applied Mathematics and Computation, 230(??):303–314, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013957.

Luo:2014:GML

[LZ14d] Jinming Luo and Jun Zhao. A generalized multiple Lyapunov functions method based non-fragile guaranteed cost control for uncertain switched fuzzy systems. Applied Mathematics and Computation, 237(??):494–500, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314005207.

Li:2014:SES

[LZaY14] Yin Li, Yulin Zhao, and Zheng an Yao. Stochastic exact solutions of the Wick-type stochastic NLS equation. Applied Mathematics and Computation, 249(??):209–221, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013083.

Liu:2014:FRR

[LZB14] Xiaoji Liu, Miao Zhang, and Julio Benítez. Further results on the reverse order law for the group inverse in rings. Applied Mathematics and Computation, 229(??):316–326, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013064.

Li:2012:SRN

[LZDS12] Hongyi Li, Di Zhao, Fei Dai, and Donglin Su. On the spectral radius of a nonnegative centrosymmetric matrix. Applied Mathematics and Computation, 218(9):4962–4966, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013075.

Liu:2010:IEP

[LZFR10] Zhongyun Liu, Yulin Zhang, C. Ferreira, and Rui Ralha. On inverse eigenvalue problems for block Toeplitz matrices with Toeplitz blocks. Applied Mathematics and Computation, 216(6):1819–1830, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010704.

Liu:2010:DPC

[LZG10] Bing Liu, Qian Zhang, and Yinghui Gao. The dynamics of pest control pollution model with age structure and time delay. Applied Mathematics and Computation, 216(10):2814–2823, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003966.

Liu:2011:NHI

[LZGZ11] X. Y. Liu, A. L. Zhang, Y. L. Gao, and W. Zhao. A novel hybrid immune algorithm and its convergence based on the steepest descent algorithm. Applied Mathematics and Computation, 218(4):1291–1296, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100823X.

Liu:2010:FTN

[LZH10] Yifang Liu, Xiangyue Zhu, and Juxiang He. Factorization technique and new exact solutions for the modified Camassa–Holm and Degasperis–Procesi equations. Applied Mathematics and Computation, 217(4):1658–1665, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309008583.

Lai:2012:NIP

[LzJZ12] Xian-Jing Lai, Mei zhen Jin, and Jie-Fang Zhang. Novel interacting phenomena in (2 + 1) dimensional AKNS system. Applied Mathematics and Computation, 218 (20):10056–10062, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003219.

Liu:2011:EBP

[LZKK11] Zeqing Liu, Liangshi Zhao, Shin Min Kang, and Young Chel Kwun. Existence of bounded positive solutions for a system of difference equations. Applied Mathematics and Computation, 218(6):2889–2912, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010599.

Long:2012:SAC

[LZL12] Shaohua Long, Shouming Zhong, and Zijian Liu. Stochastic admissibility for a class of singular Markovian jump systems with modedependent time delays. Applied Mathematics and Computation, 219(8):4106– 4117, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300312010740.

Li:2014:LKF

[LZL14a] Zhao-Yan Li, Bin Zhou, and James Lam. Lyapunov– Krasovskii functionals for predictor feedback control of linear systems with multiple input delays. Applied Mathematics and Computation, 244(??):303–311, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009746.

Li:2014:TPD

[LZL14b] Zhao-Yan Li, Bin Zhou, and James Lam. Towards positive definite solutions of a class of nonlinear matrix equations. Applied Mathematics and Computation, 237(??):546–559, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005062.

Li:2014:QAT

[LZLB14] Xiang Li, Weiguo Zhang, Zhengming Li, and Lanyun Bian. Qualitative analysis to traveling wave solutions of Zakharov– Kuznetsov–Burgers equation and its damped oscillatory solutions. Applied Mathematics and Computation, 236(??):169–183, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003300.

Li:2010:RCG

[LZLF10] Minghao Li, Wuneng Zhou, Hongqian Lu, and Jian’an Fang. Robust H_{∞} control for a generic linear rational expectations model of economy. Applied Mathematics and Computation, 216(7):2145–2154, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002912.

Liu:2014:PDG

[LZLS14] Chanjuan Liu, Enqiang Zhu, Fenrong Liu, and Kaile Su. Preference dynamics in games with short sight. Applied Mathematics and Computation, 244(??):493–501, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314009709.

Li:2011:POB

[LZLW11] Zhao-Yan Li, Bin Zhou, James Lam, and Yong Wang. Positive operator based iterative algorithms for solving Lyapunov equations for Itô stochastic systems with Markovian jumps. Applied Mathematics and Computation, 217(21):8179–8195, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000476.

Liu:2014:SAD

[LZLY14] Chao Liu, Qingling Zhang, Jinna Li, and Wenquan Yue. Stability analysis in a delayed preypredator-resource model with harvest effort and stage structure. Applied Mathematics and Computation, 238(??):177–192, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005396.

Li:2012:BSN

[LZM^{+}12] J. Li, W. Zhang, S. F.

Miao, M. Sun, and Y. Tian. Bifurcation set and number of limit cycles in Z_{2}-equivariant planar vector fields of degree 5. Applied Mathematics and Computation, 218(9):4944–4952, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013051.

Li:2011:LMR

[LZQ11] Dongfang Li, Chengjian Zhang, and Hongyu Qin. LDG method for reactiondiffusion dynamical systems with time delay. Applied Mathematics and Computation, 217(22): 9173–9181, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311005339.

Liu:2014:PSS

[LZSD14] Xia Liu, Yuanbiao Zhang, Haiping Shi, and Xiaoqing Deng. Periodic and subharmonic solutions for fourth-order nonlinear difference equations. Applied Mathematics and Computation, 236(??):613–620, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004482.

Liu:2013:PGA

[LZTZ13] Zijian Liu, Shouming Zhong, Zhidong Teng, and Long Zhang. Permanence and global attractivity of an impulsive ratiodependent predator-prey system in a patchy environment. Applied Mathematics and Computation, 219(18):9791–9804, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300369X.

Li:2012:CFD

[LZW12a] Xin Li, Luming Zhang, and Shanshan Wang. A compact finite difference scheme for the nonlinear Schrödinger equation with wave operator. Applied Mathematics and Computation, 219 (6):3187–3197, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009502.

Lu:2012:ILD

[LZW12b] Gui-Fu Lu, Jian Zou, and Yong Wang. Incremental learning of discriminant common vectors for feature extraction. Applied Mathematics and Computation, 218(22):11269–11278, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005048.

Li:2013:IOG

[LZW13a] Xiaopei Li, Xiaoliang Zhou, and Wu-Sheng Wang. Interval oscillation of a general class of second-order nonlinear impulsive differential equations with nonlinear damping. Applied Mathematics and Computation, 219(16):8803–8810, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001471.

Liu:2013:CSC

[LZW13b] Dongjie Liu, Xiaoping Zhang, and Jiming Wu. A collocation scheme for a certain Cauchy singular integral equation based on the superconvergence analysis. Applied Mathematics and Computation, 219(10):5198–5209, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012015.

Liang:2014:ORH

[LZWZ14] Hongjing Liang, Huaguang Zhang, Zhanshan Wang, and Jilie Zhang. Output regulation for heterogeneous linear multiagent systems based on distributed internal model compensator. Applied Mathematics and Computation, 242(??):736–747, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314008777.

Lu:2010:TWS

[LZX10] Bin Lu, HongQing Zhang, and FuDing Xie. Travelling wave solutions of nonlinear partial equations by using the first integral method. Applied Mathematics and Computation, 216(4):1329–1336, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001967.

Liu:2014:SMA

[LZX14] Dongyi Liu, Liping Zhang, and Genqi Xu. Spectral method and its application to the conjugate gradient method. Applied Mathematics and Computation, 240(??):339–347, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013842.

Liu:2010:VSM

[LZZ10] Zhongli Liu, Quan Zheng, and Peng Zhao. A variant of Steffensen’s method of fourth-order convergence and its applications. Applied Mathematics and Computation, 216(7):1978–1983, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002705.

Li:2011:ENM

[LZZ11] Dongping Li, Jingyu Zhao, and Guofeng Zhang. An efficient numerical method for preconditioned saddle point problems. Applied Mathematics and Computation, 217(12):5596–5602, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012397.

Li:2010:SIC

[LZZL10] Kelin Li, Liping Zhang, Xinhua Zhang, and Zuoan Li. Stability in impulsive Cohen–Grossbergtype BAM neural networks with distributed delays. Applied Mathematics and Computation, 215(11):3970–3984, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010352.

Liang:2010:ESS

[LZZP10] Yan Liang, Lei Zhang, Donghua Zhou, and Quan Pan. Estimation of systems with statisticallyconstrained inputs. Applied Mathematics and Computation, 217(6):2644– 2656, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310008386.

Liu:2010:BTW

[LZZZ10] Qian Liu, Yuqian Zhou, Weinian Zhang, and Shouming Zhong. Bounded traveling waves of the oceanic currents motion equations. Applied Mathematics and Computation, 216(1):107–113, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000305.

Mahmoud:2010:RSS

[MA10a] Magdi S. Mahmoud and Naif B. Almutairi. Robust stability and stabilization methods for a class of nonlinear discrete-time delay systems. Applied Mathematics and Computation, 215(12):4280–4292, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309011114.

Melchor-Aguilar:2010:SID

[MA10b] Daniel Melchor-Aguilar. On stability of integral delay systems. Applied Mathematics and Computation, 217(7):3578–3584, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310009227.

Ma:2011:CDM

[Ma11a] N.-Y. Ma. A comment on “On the distribution of Ma and King”. Applied Mathematics and Computation, 218(1):202–203, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007296.

See [Nad07].

Ma:2011:IIG

[Ma11b]  Shi-Mei Ma.  Identi-

ties involving generalized Fibonacci-type polynomials. Applied Mathematics and Computation, 217(22):9297–9301, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005674.

Ma:2011:NCI

[Ma11c]  Wen-Xiu Ma.  Non-

linear continuous integrable Hamiltonian couplings. Applied Mathematics and Computation, 217(17):7238–7244, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100172X.

Merdan:2011:MMA

[MA11d] H. Merdan and M. Alisen. A mathematical model for asset pricing. Applied Mathematics and Computation, 218(4):1449–1456, October 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008551.

Ma:2012:SHBa

[Ma12]  Shaojuan Ma.  The

stochastic Hopf bifurcation analysis in Brusselator system with random parameter. Applied Mathematics and Computation, 219(1):306–319, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006200.

Ma:2013:SHA

[Ma13a] Wen-Xiu Ma. A soliton hierarchy associated with so(3, R). Applied Mathematics and Computation, 220(??):117–122, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005687.

Mahajan:2013:CRM

[MA13b] Amit Mahajan and Monika Arora. Convection in rotating magnetic nanofluids. Applied Mathematics and Computation, 219 (11):6284–6296, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012799.

Mekki:2013:NSK

[MA13c] Ayadi Mekki and Maâtoug Mohamed Ali. Numerical simulation of Kadomtsev– Petviashvili–Benjamin–Bona– Mahony equations using finite difference method. Applied Mathematics and Computation, 219(24): 11214–11222, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313004633.

Melchor-Aguilar:2013:FRE

[MA13d] Daniel Melchor-Aguilar. Further results on exponential stability of linear continuous time difference systems. Applied Mathematics and Computation, 219(19):10025– 10032, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003044.

Ma:2014:PSS

[Ma14a] Tiantian Ma. Periodic solutions of the second order damped systems with singularities and deviating arguments. Applied Mathematics and Computation, 227(??):148–155, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011867.

Magrenan:2014:OAT

[MA14b] Á. Alberto Magreñán and Ioannis K. Argyros. Optimizing the applicability of a theorem by F. Potra for Newton-like methods. Applied Mathematics and Computation, 242(??):612–623, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007656.

Ma:2011:WGS

[MAA11] Wen-Xiu Ma, Alrazi Abdeljabbar, and Magdy Gamil Asaad. Wronskian and Grammian solutions to a (3 + 1)-dimensional generalized KP equation. Applied Mathematics and Computation, 217(24):10016–10023, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006473.

Maatouk:2012:EQE

[Maa12] Khalil Maatouk. Estimating quadrature errors for an efficient method for quasi-singular boundary integral. Applied Mathematics and Computation, 218(9):4658–4670, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013129.

Maatouk:2014:DBF

[Maa14]  Khalil Maatouk.  Dis-

crete boundary finite element schemes for an exterior problem for the time-harmonic Maxwell’s equation. Applied Mathematics and Computation, 242(??):361–387, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400719X.

Mandal:2014:CSD

[MAB14] Partha Sarathi Mandal, Syed Abbas, and Malay Banerjee. A comparative study of deterministic and stochastic dynamics for a non-autonomous allelopathic phytoplankton model. Applied Mathematics and Computation, 238(??):300–318, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005323.

Moukalled:2011:PCN

[MAD11] F. Moukalled, A. Abdel Aziz, and M. Darwish. Performance comparison of the NWF and DC methods for implementing highresolution schemes in a fully coupled incompressible flow solver. Applied Mathematics and Computation, 217(11):5041–5054, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007976.

Maddalena:2014:AID

[Mad14a] Lucia Maddalena. Analysis of an integro-differential system modeling tumor growth. Applied Mathematics and Computation, 245(??):152–157, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010388.

Mesmouli:2014:SSN

[MAD14b] Mouataz Billah Mesmouli, Abdelouaheb Ardjouni, and Ahcene Djoudi. Study of the stability in nonlinear neutral differential equations with functional delay using Krasnoselskii– Burton’s fixed-point. Applied Mathematics and Computation, 243(??): 492–502, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031400839X.

Maeda:2010:OPS

[Mae10]  Takashi Maeda.  On

optimization problems with set-valued objective maps. Applied Mathematics and Computation, 217(3):1150–1157, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004285.

Magrenan:2014:DAJ

[Mag14a] Á. Alberto Magreñán. Different anomalies in a Jarratt family of iterative root-finding methods. Applied Mathematics and Computation, 233(??):29– 38, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000745.

Magrenan:2014:NTS

[Mag14b] Ángel Alberto Magreñán. A new tool to study real dynamics: the convergence plane. Applied Mathematics and Computation, 248(??):215–224, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/
pii/S0096300314012867.

Mahfouz:2013:NSF

[Mah13] F. M. Mahfouz. Numerical simulation of free convection within an eccentric annulus filled with micropolar fluid using spectral method. Applied Mathematics and Computation, 219(10):5397–5409, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012052.

Mahmudov:2014:ADT

[Mah14a] Nazim I. Mahmudov.

Approximation by q-Durrmeyer type polynomials in compact disks in the case q < 1. Applied Mathematics and Computation, 237(??):293–303, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004901.

Mahmudov:2014:DEA

[Mah14b] Nazim I. Mahmudov. Difference equations of q-Appell polynomials. Applied Mathematics and Computation, 245(??): 539–543, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314010741.

Momani:2014:CMS

[MAHAS14] Shaher Momani, Omar Abu Arqub, Tasawar Hayat, and Hamed Al-Sulami. A computational method for solving periodic boundary value problems for integrodifferential equations of Fredholm–Volterra type. Applied Mathematics and Computation, 240(??): 229–239, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314005979.

Mustafa:2013:GMK

[MAK13] Zead Mustafa, Hassen Aydi, and Erdal Karapinar. Generalized Meir– Keeler type contractions on G-metric spaces. Applied Mathematics and Computation, 219(21): 10441–10447, July 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004475.

Makasu:2014:BPA

[Mak14]  Cloud Makasu.  A

bilevel programming approach to double optimal stopping. Applied Mathematics and Computation, 238(??):393–396, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005487.

Malinowski:2012:SDE

[Mal12a] Marek T. Malinowski. On set differential equations in Banach spaces — a second type Hukuhara differentiability approach. Applied Mathematics and Computation, 219(1):289– 305, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312006182.

Malinowski:2012:STH

[Mal12b] Marek T. Malinowski. Second type Hukuhara differentiable solutions to the delay set-valued differential equations. Applied Mathematics and Computation, 218(18):9427–9437, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002597.

Malik:2013:SMP

[Mal13a] Saroj B. Malik. Some more properties of core partial order. Applied Mathematics and Computation, 221(??):192–201, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006243.

Malinowski:2013:ASF

[Mal13b] Marek T. Malinowski. Approximation schemes for fuzzy stochastic integral equations. Applied Mathematics and Computation, 219(24):11278–11290, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313005870.

Malligarjunan:2014:TBE

[Mal14] Rajkumar Malligarjunan. Transient behaviour and entropy measures on an inventory system with two demand classes. Applied Mathematics and Computation, 226(??):738–753, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011065.

Mamedov:2011:AET

[Mam11] B. A. Mamedov. Analytical evaluation of the thermionic emission probability for semiconductor– insulator interfaces using incomplete gamma functions. Applied Mathematics and Computation, 217(13):6282–6285, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100004X.

Mammeri:2013:NSR

[Mam13] Y. Mammeri. Numerical study of the regularizing effect of the 3D weakly transverse BBM equations for long times. Applied Mathematics and Computation, 219(10):5162–5173, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011423.

Malakooti:2010:CIB

[MAN10a] Behnam Malakooti and Camelia Al-Najjar. The Complex Interior-Boundary method for linear and nonlinear programming with linear constraints. Applied Mathematics and Computation, 216(7):1903–1917, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001372.

Manev:2010:SGP

[Man10b] Nikolai L. Manev. Sequences generating permutations. Applied Mathematics and Computation, 216(3):708–718, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310000718.

Manhas:2011:WCO

[Man11] J. S. Manhas. Weighted composition operators between weighted spaces of vector-valued holomorphic functions on Banach spaces. Applied Mathematics and Computation, 218(3):929–934, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/

pii/S009630031100511X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Man:2012:CAP

[Man12a] Yiu-Kwong Man. A coverup approach to partial fractions with linear or irreducible quadratic factors in the denominators. Applied Mathematics and Computation, 219(8):3855–3862, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201003X.

Mante:2012:AIB

[Man12b] Claude Manté. Application of iterated Bernstein operators to distribution function and density approximation. Applied Mathematics and Computation, 218(18):9156–9168, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002238.

Mashiyev:2010:LOB

[MAO10] R. A. Mashiyev, G. Alisoy, and S. Ogras. Lyapunov, Opial and Beesack inequalities for onedimensional p(t)-Laplacian equations. Applied Mathematics and Computation, 216(12):3459–3467, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004480.

Mao:2011:LAS

[Mao11a] Wei-Jie Mao. An LMI approach to D-stability and D-stabilization of linear discrete singular systems with state delay. Applied Mathematics and Computation, 218 (5):1694–1704, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008812.

Mao:2011:NSS

[Mao11b] Xuerong Mao. Numerical solutions of stochastic differential delay equations under the generalized Khasminskii-type conditions. Applied Mathematics and Computation, 217(12):5512–5524, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310012233.

Mao:2012:SSB

[Mao12] Xiaochen Mao. Stability switches, bifurcation, and multi-stability of coupled networks with time delays. Applied Mathematics and Computation, 218(11):6263–6274, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014421.

Martin:2011:HPM

[Mar11] Olga Martin. A homotopy perturbation method for solving a neutron transport equation. Applied Mathematics and Computation, 217(21):8567–8574, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004619.

Margenstern:2012:AGS

[Mar12] Maurice Margenstern. An application of Grossone to the study of a family of tilings of the hyperbolic plane. Applied Mathematics and Computation, 218(16):8005–8018, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005698.

Marik:2014:RPS

[Mar14] Robert Marík. Remarks on the paper by Sun and Meng, Appl. Math. Comput. 174 (2006). Applied Mathematics and Computation, 248(??):309–313, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013344.

See [SM06].

Matsutani:2010:STI

[Mat10] Shigeki Matsutani. Sheaftheoretic investigation of CIP-method. Applied Mathematics and Computation, 217(2):568–579, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006594.

Matsuura:2012:RMI

[Mat12] Masaya Matsuura. On a recursive method including both CG and Burg’s algorithms. Applied Mathematics and Computation, 219(3):773–780, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007242.

Mathai:2014:EMV

[Mat14a] A. M. Mathai. Evaluation of matrix-variate gamma and beta integrals as multiple integrals and Kober fractional integral operators in the complex matrix variate case. Applied Mathematics and Computation, 247(??):312–318, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012077.

Matic:2014:MIL

[Mat14b] Dragan Matić. A mixed integer linear programming model and variable neighborhood search for Maximally Balanced Connected Partition Problem. Applied Mathematics and Computation, 237(??):85–97, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314004603.

Matloka:2014:IPF

[Mat14c] Marian Matloka. Inequalities for h-preinvex functions. Applied Mathematics and Computation, 234(??):52–57, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002690.

Matloka:2014:HIC

[Mat14d]  Marian Matloka.  On

Hadamard’s inequality for h-convex function on a disk. Applied Mathematics and Computation, 235(??):118–123, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003336.

Manteca:2014:DCS

[MATA14] I. Alhama Manteca, M. Alcaraz, E. Trigueros, and F. Alhama. Dimensionless characterization of salt intrusion benchmark scenarios in anisotropic media. Applied Mathematics and Computation, 247 (??):1173–1182, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012582.

Meza:2010:CTM

[MB10] Magno Enrique Mendoza Meza and Amit Bhaya. Control theory and the management of ecosystems: a threshold policy with hysteresis is robust. Applied Mathematics and Computation, 216(11):3133–3145, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004352.

Meza:2011:CPP

[MB11] Magno Enrique Mendoza Meza and Amit Bhaya. Controlling predator-prey discrete dynamics utilizing a threshold policy with hysteresis. Applied Mathematics and Computation, 217(20):7874–7886, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003055.

Majumdar:2012:AAU

[MB12] Jayanta Majumdar and Asoke Kumar Bhunia. An alternative approach for unbalanced assignment problem via genetic algorithm. Applied Mathematics and Computation, 218(12):6934–6941, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015578.

Mishra:2013:WPA

[MB13a] Indira Mishra and D. Bahuguna. Weighted pseudo almost automorphic solution of an integro-differential equation, with weighted Stepanovlike pseudo almost automorphic forcing term. Applied Mathematics and Computation, 219(10): 5345–5355, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011642.

Mittal:2013:NSS

[MB13b] R. C. Mittal and Rachna Bhatia. Numerical solution of second order one dimensional hyperbolic telegraph equation by cubic B-spline collocation method. Applied Mathematics and Computation, 220(??):496–506, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006619.

Mittal:2014:NST

[MB14] R. C. Mittal and Rachna Bhatia. A numerical study of two dimensional hyperbolic telegraph equation by modified B-spline differential quadrature method. Applied Mathematics and Computation, 244(??):976–997, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010170.

Mutyalarao:2010:LDC

[MBR10] M. Mutyalarao, D. Bharathi, and B. Nageswara Rao. Large deflections of a cantilever beam under an inclined end load. Applied Mathematics and Computation, 217(7):3607–3613, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310009689.

Mamedov:2012:CTC

[MÇ12a] Bahtiyar A. Mamedov and Ebru Çopuroğlu. Calculation of two-center nuclear attraction integrals of Slater type orbitals with noninteger principal quantum numbers using Guseinov’s one-center expansion formulas and Löwdinα radial function. Applied Mathematics and Computation, 218(15):7893–7897, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001208.

Marques:2012:NED

[MC12b] Filipe J. Marques and Carlos A. Coelho. Near-exact distributions for the likelihood ratio test statistic of the multi-sample block-matrix sphericity test. Applied Mathematics and Computation, 219(6):2861–2874, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008521.

Mei:2012:EMM

[MC12c] Liquan Mei and Yaping Chen. Explicit multistep method for the numerical solution of RLW equation. Applied Mathematics and Computation, 218(18):9547–9554, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002822.

Maidi:2013:OLO

[MC13] Ahmed Maidi and J. P.

Corriou. Open-loop optimal controller design using variational iteration method. Applied Mathematics and Computation, 219(16):8632–8645, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002130.

Madzvamuse:2014:FIT

[MC14a] Anotida Madzvamuse and Andy H. W. Chung. Fully implicit time-stepping schemes and non-linear solvers for systems of reactiondiffusion equations. Applied Mathematics and Computation, 244(??): 361–374, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009618.

Mall:2014:CNN

[MC14b] Susmita Mall and S. Chakraverty. Chebyshev Neural Network based model for solving Lane–Emden type equations. Applied Mathematics and Computation, 247(??):100–114, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011953.

Miao:2014:NGM

[MC14c] Shu-Xin Miao and Yang Cao. A note on GPIU method for generalized saddle point problems. Applied Mathematics and Computation, 230(??):27– 34, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013507.

McCluskey:2010:DVA

[McC10] C. Connell McCluskey. Delay versus age-of-infection – Global stability. Applied Mathematics and Computation, 217(7):3046–3049, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031000891X.

Martinez:2011:NAP

[MCdSS11] André L. M. Martinez, Emerson V. Castelani, Jair da Silva, and Wesley V. I. Shirabayashi. A note about positive solutions for an equation of Kirchhoff type. Applied Mathematics and Computation, 218(5):2082–2090, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009581.

Ma:2014:PPS

[MCH14] Ruyun Ma, Ruipeng Chen, and Zhiqian He. Positive periodic solutions of second-order differential equations with weak singularities. Applied Mathematics and Computation, 232(??):97–103, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301432X.

Ma:2014:PSE

[MCL14a] Ruyun Ma, Ruipeng Chen, and Yanqiong Lu. Positive solutions of an elliptic system modeling a population with two age groups. Applied Mathematics and Computation, 249(??):114–120, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314014404.

Mortici:2014:CMF

[MCL14b] Cristinel Mortici, Valentin Gabriel Cristea, and Dawei Lu. Completely monotonic functions and inequalities associated to some ratio of gamma function. Applied Mathematics and Computation, 240(??):168–174, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006250.

Malik:2010:ETW

[MCM10] Anand Malik, Fakir Chand, and S. C. Mishra. Exact travelling wave solutions of some nonlinear equations by (\frac{G′}{G})-expansion method. Applied Mathematics and Computation, 216(9):2596–2612, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003577.

Menniti:2013:EGD

[MCM13] Salvatore Menniti, Emanuele Castagna, and Tommaso Mazza. Estimating the global density of graphs by a sparseness index. Applied Mathematics and Computation, 224(??):346–357, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009089.

Montoro-Cazorla:2012:SWD

[MCPO12] Delia Montoro-Cazorla and Rafael Pérez-Ocón. Shock and wear degradating systems under three types of repair. Applied Mathematics and Computation, 218(24):11727– 11737, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312005681.

Martin:2011:CEF

[MCR11] M. A. Martín, M. L.

Cuadrado, and C. Romero. Computing efficient financial strategies: an extended compromise programming approach. Applied Mathematics and Computation, 217(19): 7831–7837, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311003304.

Milanovic:2014:MGD

[MCR14] V. Milanović, N. Cakić, and J. Radovanović. Method for generating a discrete state in the continuum part of the spectrum. Applied Mathematics and Computation, 246(??):514–518, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011850.

Milovanovic:2011:GBY

[MCS11] Gradimir V. Milovanović, Aleksandar S. Cvetković, and Marija P. Stanić. A generalized Birkhoff– Young–Chebyshev quadrature formula for analytic functions. Applied Mathematics and Computation, 218(3):944–948, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001652.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ma:2013:DBL

[MCWC13] Zhaozhi Ma, Fengde Chen, Chengqiang Wu, and Wanlin Chen. Dynamic behaviors of a Lotka–Volterra predatorprey model incorporating a prey refuge and predator mutual interference. Applied Mathematics and Computation, 219(15):7945–7953, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001550.

Mukherjee:2010:IDC

[MD10] Supriya Mukherjee and Sourav Dutta. Integrability of the discretization of the controlled Euler Top. Applied Mathematics and Computation, 215(12):4185–4190, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300309010868.

Mosic:2011:FRR

[MD11a] Dijana Mosić and Dragan S. Djordjević. Further results on the reverse order law for the Moore–Penrose inverse in rings with involution. Applied Mathematics and Computation, 218(4):1478–1483, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008733.

Mosic:2011:ROL

[MD11b] Dijana Mosić and Dragan S. Djordjević. Reverse order law in C^{∗}- algebras. Applied Mathematics and Computation, 218(7):3934–3941, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012148.

Mosic:2012:FWE

[MD12a] Dijana Mosíc and Dragan S. Djordjevíc. Factorization of weighted-EP elements in C^{∗}- algebras. Applied Mathematics and Computation, 218(9):5383–5390, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013518.

Mosic:2012:NCE

[MD12b] Dijana Mosić and Dragan S. Djordjević. New characterizations of EP, generalized normal and generalized Hermitian elements in rings. Applied Mathematics and Computation, 218(12):6702–6710, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311015049.

Mosic:2012:RGD

[MD12c] Dijana Mosić and Dragan S. Djordjević. Representation for the generalized Drazin inverse of block matrices in Banach algebras. Applied Mathematics and Computation, 218(24):12001–12007, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005917.

Mosic:2012:ROL

[MD12d] Dijana Mosić and Dragan S. Djordjević. Reverse order law for the group inverse in rings. Applied Mathematics and Computation, 219(5):2526–2534, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008818.

Malkowsky:2013:BAM

[MD13a] Eberhard Malkowsky and Ivana Djolović. Banach algebras of matrix transformations between some sequence spaces related to Λ-strong convergence and boundedness. Applied Mathematics and Computation, 219(16):8779–8789, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002051.

Mosic:2013:FRR

[MD13b] Dijana Mosić and Dragan S. Djordjević. Further results on the reverse order law for the group inverse in rings. Applied Mathematics and Computation, 219(19):9971–9977, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004049.

Mosic:2013:SEG

[MD13c] Dijana Mosić and Dragan S. Djordjević. Several expressions for the generalized Drazin inverse of a block matrix in a Banach algebra. Applied Mathematics and Computation, 220(??):374–381, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006310.

Mosic:2014:IIK

[MD14] Dijana Mosić and Dragan S. Djordjević. Inner image-kernel (p, q)- inverses in rings. Applied Mathematics and Computation, 239(??):144–152, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400592X.

Maden:2013:SUB

[MDÇ13] A. Dilek (Güngör) Maden, Kinkar Ch. Das, and A. Sinan Çevik. Sharp upper bounds on the spectral radius of the signless Laplacian matrix of a graph. Applied Mathematics and Computation, 219(10):5025–5032, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012064.

Ma:2012:SCS

[MDD12] Qiang Ma, Deqiong Ding, and Xiaohua Ding. Symplectic conditions and stochastic generating functions of stochastic Runge– Kutta methods for stochastic Hamiltonian systems with multiplicative noise. Applied Mathematics and Computation, 219(2):635– 643, October 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312006613.

Macias-Diaz:2012:EPP

[MDP12] J. E. Macías-Díaz and A. Puri. An explicit positivity-preserving finitedifference scheme for the classical Fisher–Kolmogorov– Petrovsky–Piscounov equation. Applied Mathematics and Computation, 218(9):5829–5837, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101407X.

Macias-Diaz:2014:CSN

[MDR14] J. E. Macías-Díaz and K. A. Rejniak. On a conditionally stable nonlinear method to approximate some monotone and bounded solutions of a generalized population model. Applied Mathematics and Computation, 229(??):273–282, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013350.

Mead:2013:DPE

[Mea13] J. L. Mead. Discontinuous parameter estimates with least squares estimators. Applied Mathematics and Computation, 219(10):5210–5223, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012349.

Megahed:2011:CIA

[MEBYG11] A. A. Megahed, Abou-Zaid El-Banna, E. A. Youness, and H. G. Gomaa. A combined interactive approach for solving E-convex multiobjective nonlinear programming problem. Applied Mathematics and Computation, 217(16):6777–6784, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310013032.

Medjo:2012:APG

[Med12] T. Tachim Medjo. Averaging of the planetary 3D geostrophic equations with oscillating external forces. Applied Mathematics and Computation, 218(10):5910–5928, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013919.

Mohsen:2010:NSL

[MEG10] Adel Mohsen and Mohamed El-Gamel. On the numerical solution of linear and nonlinear Volterra integral and integro-differential equations. Applied Mathematics and Computation, 217(7):3330–3337, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000929X.

Mehrem:2011:PWE

[Meh11] R. Mehrem. The plane wave expansion, infinite integrals and identities involving spherical Bessel functions. Applied Mathematics and Computation, 217(12):5360–5365, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011914.

Mehrem:2013:AEI

[Meh13] Rami Mehrem. Analytical evaluation of an infinite integral over four spherical Bessel functions. Applied Mathematics and Computation, 219(21):10506– 10509, July 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004402.

Muroya:2013:GSD

[MEL13] Yoshiaki Muroya, Yoichi Enatsu, and Huaixing Li. Global stability of a delayed HTLV-I infection model with a class of nonlinear incidence rates and CTLs immune response. Applied Mathematics and Computation, 219(21): 10559–10573, July 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003421.

Melo:2014:ESE

[Mel14] César A. Hernández Melo. Existence and stability of equilibrium solutions of a nonlinear heat equation. Applied Mathematics and Computation, 232(??):1025–1036, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001866.

Mennouni:2011:NNA

[Men11] Abdelaziz Mennouni. A new numerical approximation for Volterra integral equations combining two quadrature rules. Applied Mathematics and Computation, 218(5):1962–1969, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009271.

Mennouni:2012:APS

[Men12a] Abdelaziz Mennouni. Airfoil polynomials for solving integro-differential equations with logarithmic kernel. Applied Mathematics and Computation, 218(24):11947–11951, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005838.

Mennouni:2012:NSC

[Men12b] Abdelaziz Mennouni. A note on solving Cauchy integral equations of the second kind by projection. Applied Mathematics and Computation, 219(3):1108–1113, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007175.

Meng:2014:PAO

[Men14] Guo-Yan Meng. A practical asymptotical optimal SOR method. Applied Mathematics and Computation, 242(??):707–715, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314008704.

Merdan:2014:SNF

[Mer14]  Mehmet Merdan.  On

the solutions of nonlinear fractional Klein–Gordon equation with modified Riemann–Liouville derivative. Applied Mathematics and Computation, 242(??):877–888, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008492.

Moatimid:2013:ESC

[MESAN13] G. M. Moatimid, Rehab M. El-Shiekh, and Abdul-Ghani A. A. H. Al-Nowehy. Exact solutions for Calogero– Bogoyavlenskii–Schiff equation using symmetry method. Applied Mathematics and Computation, 220(??): 455–462, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313006735.

Meyvaci:2014:SSL

[Mey14] M. Meyvaci. Structural stability for a linear system of thermoelastic plate equations. Applied Mathematics and Computation, 242(??):664–668, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008753.

Mezo:2013:SIS

[Mez13] István Mező. Some infinite sums arising from the Weierstrass product theorem. Applied Mathematics and Computation, 219(18):9838–9846, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003834.

Mavrotas:2013:IVA

[MF13] George Mavrotas and Kostas Florios. An improved version of the augmented ϵ-constraint method (AUGMECON2) for finding the exact Pareto set in multiobjective integer programming problems. Applied Mathematics and Computation, 219(18):9652–9669, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002166.

Mokkedem:2014:ACS

[MF14] Fatima Zahra Mokkedem and Xianlong Fu. Approximate controllability of semi-linear neutral integro-differential systems with finite delay. Applied Mathematics and Computation, 242(??):202–215, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007206.

Ma:2012:SBH

[MFZ12] Yun-Jie Ma, Chu-Li Fu, and Yuan-Xiang Zhang. Solving a backward heat conduction problem by variational method. Applied Mathematics and Computation, 219(2):624– 634, October 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006601.

Ma:2010:HOF

[MG10a] Yuezhen Ma and Yongbin Ge. A high order finite difference method with Richardson extrapolation for 3D convection diffusion equation. Applied Mathematics and Computation, 215(9):3408–3417, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009400.

Misirli:2010:ESD

[MG10b] Emine Misirli and Yusuf Gurefe. Exact solutions of the Drinfel’d– Sokolov–Wilson equation using the Exp-function method. Applied Mathematics and Computation, 216(9):2623–2627, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003590.

Mennouni:2011:NSC

[MG11a] Abdelaziz Mennouni and Said Guedjiba. A note on solving Cauchy integral equations of the first kind by iterations. Applied Mathematics and Computation, 217(18):7442–7447, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100230X.

Mohanty:2011:HAC

[MG11b] R. K. Mohanty and Venu Gopal. High accuracy cubic spline finite difference approximation for the solution of one-space dimensional non-linear wave equations. Applied Mathematics and Computation, 218(8):4234–4244, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012264.

Ma:2012:DBT

[MG12a] Yunling Ma and Xianguo Geng. Darboux and Bäcklund transformations of the bidirectional Sawada–Kotera equation. Applied Mathematics and Computation, 218(12):6963–6965, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311015670.

Merdan:2012:SAG

[MG12b] Hüseyin Merdan and Özlem Ak Gümüs. Stability analysis of a general discrete-time population model involving delay and Allee effects. Applied Mathematics and Computation, 219(4):1821–1832, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008144.

Ma:2014:PBC

[MG14a] Junhai Ma and Zhanbing Guo. The parameter basin and complex of dynamic game with estimation and two-stage consideration. Applied Mathematics and Computation, 248(??):131–142, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013034.

Mishra:2014:EAV

[MG14b] Arti Mishra and Sunita Gakkhar. The effects of awareness and vector control on two strains dengue dynamics. Applied Mathematics and Computation, 246(??):159–167, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010947.

Mohanty:2014:HAN

[MG14c] R. K. Mohanty and Venu Gopal. High accuracy nonpolynomial spline in compression method for onespace dimensional quasilinear hyperbolic equations with significant first order space derivative term. Applied Mathematics and Computation, 238(??):250–265, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005359.

Mesri:2012:ACT

[MGC12] Youssef Mesri, Hervé Guillard, and Thierry Coupez. Automatic coarsening of three dimensional anisotropic unstructured meshes for multigrid applications. Applied Mathematics and Computation, 218(21):10500– 10519, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312003979.

Martinez-Guerra:2013:SCL

[MGCFMM13] Rafael Martínez-Guerra, Dulce M. G. Corona-Fortunio, and Juan L. Mata-Machuca. Synchronization of chaotic Liouvillian systems: an application to Chua’s oscillator. Applied Mathematics and Computation, 219(23):10934–10944, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005146.

Moummou:2012:SGM

[MGGS12] El Kettani Moummou, R. Gutiérrez, and R. Gutiérrez-Sanchez. A stochastic Gompertz model with logarithmic therapy functions: Parameters estimation. Applied Mathematics and Computation, 219(8):3729–3739, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009885.

Martinez-Guerra:2014:GSD

[MGMM14] Rafael Martínez-Guerra and Juan L. Mata-Machuca. Generalized synchronization via the differential primitive element. Applied Mathematics and Computation, 232(??):848–857, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400188X.

Moummou:2014:SGM

[MGSMRÁ14] El Kettani Moummou, R. Gutiérrez-Sanchez, M. C. Melchor, and E. Ramos-Ábalos. A stochastic Gompertz model highlighting internal and external therapy function for tumour growth. Applied Mathematics and Computation, 246(??):1– 11, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314011023.

Meng:2011:FDT

[MGWG11] De-Xin Meng, Yi-Tian Gao, Lei Wang, and Xiao-Ling Gai. N-fold Darboux transformation and solitonic interactions of a variable-coefficient generalized Boussinesq system in shallow water. Applied Mathematics and Computation, 218(8):4049–4055, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011210.

Meng:2012:MSP

[MGYQ12] Gao-Qing Meng, Yi-Tian Gao, Xin Yu, and Yi Qin. Multi-soliton and Pfaffian solutions of a (2 + 1)-dimensional nonlinear evolution equation via the Jaulent–Miodek hierarchy. Applied Mathematics and Computation, 218(22):10791–10802, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002937.

Matejas:2010:AKM

[MH10a] J. Matejaš and V. Hari.

Accuracy of the Kogbetliantz method for scaled diagonally dominant triangular matrices. Applied Mathematics and Computation, 217(8):3726–3746, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310009677.

Mohammadi:2010:SAB

[MH10b] A. Mohammadi and E. Hesameddini. On some aspects of Bäcklund transformations. Applied Mathematics and Computation, 215(11):3985–3990, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010388.

Mathai:2011:PBS

[MH11a] A. M. Mathai and H. J. Haubold. A pathway from Bayesian statistical analysis to superstatistics. Applied Mathematics and Computation, 218(3):799–804, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100381X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Meng:2011:IOC

[MH11b] Fanwei Meng and Yan Huang. Interval oscillation criteria for a forced second-order nonlinear differential equations with damping. Applied Mathematics and Computation, 218(5):1857–1861, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009143.

Mak:2012:NIC

[MH12] M. K. Mak and T. Harko.

New integrability case for the Riccati equation. Applied Mathematics and Computation, 218(22):10974–10981, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004572.

Ma:2013:NSF

[MH13a] Xiaohua Ma and Chengming Huang. Numerical solution of fractional integro-differential equations by a hybrid collocation method. Applied Mathematics and Computation, 219(12):6750–6760, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313000131.

Mak:2013:NFI

[MH13b] M. K. Mak and T. Harko.

New further integrability cases for the Riccati equation. Applied Mathematics and Computation, 219(14):7465–7471, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000726.

Ma:2014:BSA

[MH14a] Manjun Ma and Jiajia Hu. Bifurcation and stability analysis of steady states to a Brusselator model. Applied Mathematics and Computation, 236(??):580–592, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003233.

Marinca:2014:OHA

[MH14b] Vasile Marinca and Nicolae Herisanu. The Optimal Homotopy Asymptotic Method for solving Blasius equation. Applied Mathematics and Computation, 231(??):134–139, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014112.

Mirzaee:2014:CTS

[MH14c] Farshid Mirzaee and Elham Hadadiyan. A collocation technique for solving nonlinear Stochastic Itô–Volterra integral equations. Applied Mathematics and Computation, 247 (??):1011–1020, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012727.

Momoniat:2010:NIG

[MHA10] E. Momoniat, C. Harley, and E. Adlem. Numerical investigation of the generalized lubrication equation. Applied Mathematics and Computation, 217(6):2631–2638, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008362.

Maleki:2012:STV

[MHA12] Mohammad Maleki, I. Hashim, and S. Abbasbandy. Solution of time-varying delay systems using an adaptive collocation method. Applied Mathematics and Computation, 219(4):1434– 1448, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312007631.

Mirzaee:2014:NST

[MHB14] Farshid Mirzaee, Elham Hadadiyan, and Saeed Bimesl. Numerical solution for three-dimensional nonlinear mixed Volterra– Fredholm integral equations via three-dimensional block-pulse functions. Applied Mathematics and Computation, 237(??): 168–175, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314004494.

He:2010:HPD

[mHhL10] You mei He and Jian hui Long. On the Hermitian positive definite solution of the nonlinear matrix equation X + ∑_{i=1}^{m} A_{i}^{⋆}X^{−1}A_{i} = I. Applied Mathematics and Computation, 216 (12):3480–3485, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004674.

Meng:2011:NEP

[MHLL11] Qing Meng, Bin He, Yao Long, and Zhenyang Li. New exact periodic wave solutions for the Dullin–Gottwald– Holm equation. Applied Mathematics and Computation, 218(8):4533–4537, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031101280X.

Mirjalili:2012:TFN

[MHS12] SeyedAli Mirjalili, Siti Zaiton Mohd Hashim, and Hossein Moradian Sardroudi. Training feedforward neural networks using hybrid particle swarm optimization and gravitational search algorithm. Applied Mathematics and Computation, 218(22):11125–11137, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004857.

Ma:2012:FCL

[MHTH12] Manjun Ma, Jianchao Hu, Jicheng Tao, and Yazhou Han. Food chain length of a ecological model with prey-taxis. Applied Mathematics and Computation, 219(8):4013–4030, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010399.

Meng:2014:SPP

[MHXyY14] Xin-You Meng, Hai-Feng Huo, Hong Xiang, and Qi yu Yin. Stability in a predator-prey model with Crowley– Martin function and stage structure for prey. Applied Mathematics and Computation, 232(??):810–819, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001854.

Mahmoud:2010:IRR

[MI10] Magdi S. Mahmoud and Abdulla Ismail. Improved results on robust exponential stability criteria for neutraltype delayed neural networks. Applied Mathematics and Computation, 217(7):3011–3019, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000888X.

Mechee:2014:DNM

[MIHS14] M. Mechee, F. Ismail, Z. M. Hussain, and Z. Siri. Direct numerical methods for solving a class of thirdorder partial differential equations. Applied Mathematics and Computation, 247(??):663–674, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012399.

Milovanovic:2012:GQF

[Mil12] Gradimir V. Milovanović.

Generalized quadrature formulae for analytic functions. Applied Mathematics and Computation, 218(17):8537–8551, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001361.

Milosevic:2014:EUA

[Mil14a] Marija Milosević. Existence, uniqueness, almost sure polynomial stability of solution to a class of highly nonlinear pantograph stochastic differential equations and the Euler–Maruyama approximation. Applied Mathematics and Computation, 237(??):672–685, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005037.

Milosevic:2014:INM

[Mil14b] Marija Milosević. Implicit numerical methods for highly nonlinear neutral stochastic differential equations with timedependent delay. Applied Mathematics and Computation, 244(??):741–760, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009990.

Martinez:2014:SST

[MIM^{+}14] J. Martínez, C. Iglesias, J. M. Matías, J. Taboada, and M. Araújo. Solving the slate tile classification problem using a DAGSVM multiclassification algorithm based on SVM binary classifiers with a one-versus-all approach. Applied Mathematics and Computation, 230(??):464–472, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013775.

Misir:2010:SCR

[Mis10] Adil Misir. On the solution of a class of the reduced wave equation and the Riccati differential equation. Applied Mathematics and Computation, 216(5):1375–1382, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000250X.

Misra:2011:MDD

[Mis11] A. K. Misra. Modeling the depletion of dissolved oxygen due to algal bloom in a lake by taking Holling type-III interaction. Applied Mathematics and Computation, 217(21):8367–8376, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004000.

Misra:2014:MEP

[Mis14] A. K. Misra. Modeling the effect of police deterrence on the prevalence of crime in the society. Applied Mathematics and Computation, 237(??):531–545, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005074.

Milev:2012:MIE

[MIT12] Mariyan Milev, Pierluigi Novi Inverardi, and Aldo Tagliani. Moment information and entropy evaluation for probability densities. Applied Mathematics and Computation, 218(9):5782–5795, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014391.

Mitchell:2013:CEP

[Mit13] William F. Mitchell. A collection of 2D elliptic problems for testing adaptive grid refinement algorithms. Applied Mathematics and Computation, 220(??):350–364, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313006449.

Jiang:2010:AFS

[mJ10] Dong mei Jiang. Application of Fibonacci sine and cosine function to a nonlinear differential– difference equation. Applied Mathematics and Computation, 217(8):3898– 3902, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310010118.

Mittal:2011:BSM

[MJ11a] R. C. Mittal and R. K.

Jain. B-splines methods with redefined basis functions for solving fourth order parabolic partial differential equations. Applied Mathematics and Computation, 217(23):9741–9755, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100631X.

Mittal:2011:NST

[MJ11b] R. C. Mittal and Ram Jiwari. Numerical solution of two-dimensional reaction–diffusion Brusselator system. Applied Mathematics and Computation, 217(12):5404–5415, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310012105.

Mittal:2012:NSN

[MJ12] R. C. Mittal and R. K. Jain. Numerical solutions of nonlinear Burgers’ equation with modified cubic B-splines collocation method. Applied Mathematics and Computation, 218(15):7839–7855, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001051.

Milovanovic:2013:PBP

[MJ13a] Gradimir V. Milovanović and Dusan Joksimović. Properties of Boubaker polynomials and an application to Love’s integral equation. Applied Mathematics and Computation, 224(??):74–87, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009235.

Mu:2013:RBE

[MJ13b] Lin Mu and Rabeea Jari.

A recovery-based error estimate for nonconforming finite volume methods of interface problems. Applied Mathematics and Computation, 220(??):63–74, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005675.

Mohamad-Jawad:2011:AHP

[MJPB11] Anwar Ja’afar Mohamad-Jawad, Marko D. Petković, and Anjan Biswas. Applications of He’s principles to partial differential equations. Applied Mathematics and Computation, 217(16):7039–7047, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001718.

Molai:2008:AMT

[MK08] Ali Abbasi Molai and Esmaile Khorram. Another modification from two papers of Ghodousian and Khorram and Khorram et al. Applied Mathematics and Computation, 197(2):559–565, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300307008028.

See erratum [MK10a].

Molai:2010:EAM

[MK10a] Ali Abbasi Molai and Esmaile Khorram. Erratum to: “Another modification from two papers of Ghodousian and Khorram and Khorram et al.” [Appl. Math. Comput. 197 (2008) 559–565]. Applied Mathematics and Computation, 215(9):3488, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009291.

See [MK08].

Morris:2010:NCL

[MK10b] R. Morris and A. H. Kara. New conservation laws of some third-order systems of PDEs arising from higher-order multipliers. Applied Mathematics and Computation, 217(6):2639–2643, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310008374.

Molati:2013:SCI

[MK13a] Motlatsi Molati and Chaudry Masood Khalique. Symmetry classification and invariant solutions of the variable coefficient BBM equation. Applied Mathematics and Computation, 219(15):7917–7922, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001586.

Morris:2013:DRA

[MK13b] R. Morris and A. H.

Kara. Double reductions/ analysis of the Drinfeld– Sokolov–Wilson equation. Applied Mathematics and Computation, 219(12): 6473–6483, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313000507.

Mahmoud:2014:DAD

[MK14a] Magdi S. Mahmoud and Gulam Dastagir Khan. Dissipativity analysis for discrete stochastic neural networks with Markovian delays and partially known transition matrix. Applied Mathematics and Computation, 228(??):292–310, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012563.

Mohanty:2014:NFA

[MK14b] R. K. Mohanty and Ravindra Kumar. A new fast algorithm based on halfstep discretization for one space dimensional quasilinear hyperbolic equations. Applied Mathematics and Computation, 244(??):624–641, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009771.

Mursaleen:2012:WSC

[MKEG12] Mohammad Mursaleen, Vatan Karakaya, Müzeyyen Ertürk, and Faik Gürsoy. Weighted statistical convergence and its application to Korovkin type approximation theorem. Applied Mathematics and Computation, 218(18):9132–9137, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002081.

Mabrouk:2013:GSS

[MKeM13] S. Mabrouk, M. Kassem, and M. Abd el Malek. Group similarity solutions of the Lax pair for a generalized Hirota–Satsuma equation. Applied Mathematics and Computation, 219(14):7882–7890, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001343.

Mursaleen:2014:SAN

[MKK14] M. Mursaleen, Faisal Khan, and Asif Khan. Statistical approximation for new positive linear operators of Lagrange type. Applied Mathematics and Computation, 232(??):548–558, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001301.

Muller:2013:CGM

[MKP13] Juliane Müller, Juho Kanniainen, and Robert Piché. Calibration of GARCH models using concurrent accelerated random search. Applied Mathematics and Computation, 221(??):522–534, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007509.

Mellouli:2013:LRC

[MKSC13] Racem Mellouli, Imed Kacem, Chérif Sadfi, and Chengbin Chu. Lagrangian relaxation and column generation-based lower bounds for the P m, H_{j1}// ∑ w_{i}C_{i} scheduling problem. Applied Mathematics and Computation, 219(22):10783– 10805, July 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004931.

Mursaleen:2013:OCM

[MKSN13] M. Mursaleen, Asif Khan, H. M. Srivastava, and K. S. Nisar. Operators constructed by means of q-lagrange polynomials and A-statistical approximation. Applied Mathematics and Computation, 219(12):6911–6918, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000672.

Ma:2010:SBE

[ML10a] Xuan Ma and Shuangqian Liu. H_{x,ν}^{s} stability of the Boltzmann equation and Landau equation with soft potentials. Applied Mathematics and Computation, 215(10):3721–3730, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010005.

Ma:2010:NAE

[ML10b] Yulan Ma and Bangqing Li. New application of (\frac{G′}{G})-expansion method to a nonlinear evolution equation. Applied Mathematics and Computation, 216(7):2137–2144, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002900.

Meng:2011:CNI

[ML11a] Zhao-Liang Meng and Zhong-Xuan Luo. The construction of numerical integration rules of degree three for product regions. Applied Mathematics and Computation, 218(5):2036–2043, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009349.

Mukhopadhyay:2011:ABF

[ML11b] S. Mukhopadhyay and G. C. Layek. Analysis of blood flow through a modelled artery with an aneurysm. Applied Mathematics and Computation, 217(16):6792–6801, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010416.

Ma:2012:DMC

[ML12a] Yu-Lan Ma and Bang-Qing Li. A direct method for constructing the traveling wave solutions of a modified generalized Vakhnenko equation. Applied Mathematics and Computation, 219(4):2212–2219, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312008612.

Lei:2012:MMS

[mL12b]  De ming Lei.  Min-

imizing makespan for scheduling stochastic job shop with random breakdown. Applied Mathematics and Computation, 218(24):11851–11858, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005565.

Mazucheli:2013:CEM

[MLG13] J. Mazucheli, F. Louzada, and M. E. Ghitany. Comparison of estimation methods for the parameters of the weighted Lindley distribution. Applied Mathematics and Computation, 220(??):463–471, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313006620.

Mariani:2011:HSC

[MLGdSC11] Viviana Cocco Mariani, Luiz Guilherme Justi Luvizotto, Fábio Alessandro Guerra, and Leandro dos Santos Coelho. A hybrid shuffled complex evolution approach based on differential evolution for unconstrained optimization. Applied Mathematics and Computation, 217(12):5822–5829, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001266X.

Ma:2011:ILC

[MLH11] Fali Ma, Chuandong Li, and Tingwen Huang. Iterative learning control design of nonlinear multiple time-delay systems. Applied Mathematics and Computation, 218(8):4333–4340, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012380.

Mader:2014:NEI

[MLM^{+}14] Wolfgang Mader, Yannick Linke, Malenka Mader, Linda Sommerlade, Jens Timmer, and Björn Schelter. A numerically efficient implementation of the expectation maximization algorithm for state space models. Applied Mathematics and Computation, 241(??):222–232, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006869.

Ma:2012:SHBb

[MLY12] Zhan-Ping Ma, Wan-Tong Li, and Xiang-Ping Yan. Stability and Hopf bifurcation for a three-species food chain model with time delay and spatial diffusion. Applied Mathematics and Computation, 219(5):2713–2731, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200896X.

Mainge:2010:NIT

[MM10a] Paul-Emile Maingé and Nora Merabet. A new inertial-type hybrid projectionproximal algorithm for monotone inclusions. Applied Mathematics and Computation, 215(9):3149– 3162, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630030900931X.

Matache:2010:WNR

[MM10b] Valentin Matache and Mihaela T. Matache. When is the numerical range of a nilpotent matrix circular? Applied Mathematics and Computation, 216(1):269–275, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000652.

Moitsheki:2010:CLP

[MM10c] R. J. Moitsheki and O. D. Makinde. Classical Lie point symmetry analysis of nonlinear diffusion equations describing thermal energy storage. Applied Mathematics and Computation, 216(1):251–260, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000597.

Mainge:2011:CNM

[MM11a] Paul-Emile Maingé and Stefan Maruster. Convergence in norm of modified Krasnoselski–Mann iterations for fixed points of demicontractive mappings. Applied Mathematics and Computation, 217(24):9864–9874, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006382.

Merkle:2011:KTD

[MM11b] Milan Merkle and Monica Moulin Ribeiro Merkle. Krull’s theory for the double gamma function. Applied Mathematics and Computation, 218(3):935– 943, October 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311001251.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Mustafa:2012:EDR

[MM12] Muhammad I. Mustafa and Salim A. Messaoudi. Energy decay rates for a Timoshenko system with viscoelastic boundary conditions. Applied Mathematics and Computation, 218(18):9125–9131, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002056.

Mei:2013:FDT

[MM13a] Jianqin Mei and Zhangyun Ma. N-fold Darboux transformation and multisoliton solutions for the classical Boussinesq–Burgers system. Applied Mathematics and Computation, 219(11):6163–6169, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312013203.

Migda:2013:APS

[MM13b] Małgorzata Migda and Jarosław Morchało. Asymptotic properties of solutions of difference equations with several delays and Volterra summation equations. Applied Mathematics and Computation, 220(??):365–373, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006504.

Misra:2013:SCI

[MM13c] O. P. Misra and D. K. Mishra. Spread and control of influenza in two groups: a model. Applied Mathematics and Computation, 219(15):7982–7996, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001823.

Mohammadi:2013:EWP

[MM13d] Hossein Mohammadi and Mojtaba Mahzoon. Effect of weak prey in Leslie–Gower predatorprey model. Applied Mathematics and Computation, 224(??):196–204, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009259.

Manukure:2014:BIC

[MM14a] Solomon Manukure and Wen-Xiu Ma. Bi-integrable couplings of a new soliton hierarchy associated with a non-semisimple Lie algebra. Applied Mathematics and Computation, 245(??):44–52, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010649.

Mao:2014:ASN

[MM14b] Wei Mao and Xuerong Mao. On the approximations of solutions to neutral SDEs with Markovian switching and jumps under non-Lipschitz conditions. Applied Mathematics and Computation, 230(??):104–119, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013830.

Mezura-Montes:2012:EAM

[MMCD12] Efrén Mezura-Montes and Omar Cetina-Domínguez. Empirical analysis of a modified Artificial Bee Colony for constrained numerical optimization. Applied Mathematics and Computation, 218(22): 10943–10973, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312004560.

Mohamed:2013:MDC

[MMGM13] S. A. Mohamed, N. A. Mohamed, A. F. Abdel Gawad, and M. S. Matbuly. A modified diffusion coefficient technique for the convection diffusion equation. Applied Mathematics and Computation, 219(17):9317–9330, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002671.

Mishra:2014:TAS

[MMKD14] Lakshmi Narayan Mishra, Vishnu Narayan Mishra, Kejal Khatri, and Deepmala. On the trigonometric approximation of signals belonging to generalized weighted Lipschitz W (L^{r}ξ(t)) (r ≥ 1)-class by matrix (C^{1} · N_{p}) operator of conjugate series of its Fourier series. Applied Mathematics and Computation, 237(??):252–263, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004470.

Mitra:2011:WFB

[MMM11] Sharmishtha Mitra, Vidit Maheswari, and Amit Mitra. A wavelet filtering based estimation of output gap. Applied Mathematics and Computation, 218(7):3710–3722, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311011799.

Majumdar:2012:DAT

[MMM12] Aditi Majumdar, Dipak Kumar Maiti, and Damodar Maity. Damage assessment of truss structures from changes in natural frequencies using ant colony optimization. Applied Mathematics and Computation, 218(19):9759–9772, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002639.

Mohan:2013:SDA

[MMM13] S. C. Mohan, D. K. Maiti, and D. Maity. Structural damage assessment using FRF employing particle swarm optimization. Applied Mathematics and Computation, 219 (20):10387–10400, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004165.

Manas-Manas:2013:VDL

[MMMMB13] Juan F. Mañas-Mañas, Francisco Marcellán, and Juan J. Moreno-Balcázar.

Varying discrete Laguerre-Sobolev orthogonal polynomials: Asymptotic behavior and zeros. Applied Mathematics and Computation, 222(??):612–618, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008242.

Martinez-Martinez:2011:SNF

[MMMMMG^{+}11] Rafael Martínez-Martínez, Juan L. Mata-Machuca, Rafael Martínez-Guerra, Jorge A. León, and Guillermo Fernández-Anaya. Synchronization of nonlinear fractional order systems. Applied Mathematics and Computation, 218(7):3338–3347, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011167.

Mastroianni:2013:NMF

[MMO13] Giuseppe Mastroianni, Gradimir V. Milovanović, and Donatella Occorsio. Nyström method for Fredholm integral equations of the second kind in two variables on a triangle. Applied Mathematics and Computation, 219(14):7653–7662, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000945.

Misra:2012:MET

[MMP^{+}12] A. K. Misra, S. N. Mishra, A. L. Pathak, Peeyush Misra, and Ram Naresh. Modeling the effect of time delay in controlling the carrier dependent infectious disease — Cholera. Applied Mathematics and Computation, 218(23):11547–11557, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005474.

Martinez:2013:NSF

[MMRS13] Irene Martínez, Serafín Moral, Carmelo Rodríguez, and Antonio Salmerón. New strategies for finding multiplicative decompositions of probability trees. Applied Mathematics and Computation, 225(??):573–589, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300313010801.

Miladinovic:2011:MSM

[MMS11] Marko Miladinović, Sladjana Miljković, and Predrag Stanimirović. Modified SMS method for computing outer inverses of Toeplitz matrices. Applied Mathematics and Computation, 218(7):3131–3143, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311010861.

Miljkovic:2013:SCM

[MMS– D13] Sladjana Miljković, Marko Miladinović, Predrag S. Stanimirović, and Dragan S. – Dordević. Scalar correction method for finding least-squares solutions on Hilbert spaces and its applications. Applied Mathematics and Computation, 219(18): 9639–9651, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313002154.

Mohamed:2014:EHO

[MMS14] N. Mohamed, S. A. Mohamed, and L. F. Seddek. Exponential higherorder compact scheme for 3D steady convectiondiffusion problem. Applied Mathematics and Computation, 232(??):1046–1061, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001994.

Michel:2014:CRV

[MMSZ14] D. Michel, H. Mraoui, D. Sbibih, and A. Zidna. Computing the range of values of real functions using B-spline form. Applied Mathematics and Computation, 233(??):85–102, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001593.

Ma:2010:QIO

[MMX10] Libin Ma, Zeyao Mo, and Xiaowen Xu. Quasiinterpolation operators based on a cubic spline and applications in SAMR simulations. Applied Mathematics and Computation, 217(8):3853–3868, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300310010052.

Mishra:2011:SPS

[MMY11] S. K. Mishra, R. N.

Mohapatra, and E. A. Youness. Some properties of semi E − v-vex functions. Applied Mathematics and Computation, 217(12):5525–5530, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012269.

Mophou:2010:EMS

[MN10] G. M. Mophou and G. M.

N’Guérékata. Existence of mild solutions of some semilinear neutral fractional functional evolution equations with infinite delay. Applied Mathematics and Computation, 216(1):61–69, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309011199.

Mandal:2011:SDI

[MN11] P. Mandal and C. Nahak. Symmetric duality with (p, r) – ρ – (η, θ)- invexity. Applied Mathematics and Computation, 217(21):8141–8148, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100275X.

Malolepszy:2012:NBS

[MN12] Tomasz Malolepszy and Maciej Niedziela. A note on blow-up solutions to some nonlinear Volterra integral equations. Applied Mathematics and Computation, 218(11):6401–6406, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014676.

Mazzia:2014:SVI

[MN14a] F. Mazzia and A. M. Nagy. Solving Volterra integro-differential equations by variable stepsize block BS methods: Properties and implementation techniques. Applied Mathematics and Computation, 239(??):198–210, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005633.

Morais:2014:LDM

[MN14b] J. Morais and C. A.

Nolder. Local distortion of M-conformal mappings. Applied Mathematics and Computation, 249(??):554–568, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013587.

Menendez:2010:NLA

[MNOB10] C. Menéndez, P. J. G. Nieto, F. A. Ortega, and A. Bello. Non-linear analysis of the consolidation of an elastic saturated soil with incompressible fluid and variable permeability by FEM. Applied Mathematics and Computation, 216(2):458–476, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000524.

Mhaskar:2013:FLE

[MNP13] Hrushikesh N. Mhaskar, V. Naumova, and S. V. Pereverzyev. Filtered Legendre expansion method for numerical differentiation at the boundary point with application to blood glucose predictions. Applied Mathematics and Computation, 224(??):835–847, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009879.

Mesloub:2012:AHA

[MO12a] Said Mesloub and Saleem Obaidat. On the application of the homotopy analysis method for a nonlocal mixed problem with Bessel operator. Applied Mathematics and Computation, 219(8):3477–3485, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009174.

Mitchell:2012:ANA

[MO12b] S. L. Mitchell and S. B. G. O’Brien. Asymptotic, numerical and approximate techniques for a free boundary problem arising in the diffusion of glassy polymers. Applied Mathematics and Computation, 219(1):376–388, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200625X.

Minamoto:2014:BDI

[MO14] Teruya Minamoto and Ryuji Ohura. A blind digital image watermarking method based on the dyadic wavelet transform and interval arithmetic. Applied Mathematics and Computation, 226(??):306–319, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010850.

Mohammadi:2014:SBS

[Moh14] Reza Mohammadi. Sextic B-spline collocation method for solving Euler– Bernoulli Beam Models. Applied Mathematics and Computation, 241(??):151–166, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400664X.

Mollapourasl:2014:ENS

[Mol14] R. Mollapourasl. An efficient numerical scheme for a nonlinear integrodifferential equations with an integral boundary condition. Applied Mathematics and Computation, 248(??):8–17, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013137.

Monovasilis:2012:SPR

[Mon12] Th. Monovasilis. Symplectic partitioned Runge– Kutta methods with the phase-lag property. Applied Mathematics and Computation, 218(18): 9075–9084, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001695.

Moody:2012:CIV

[Moo12] Dustin Moody. Computing isogeny volcanoes of composite degree. Applied Mathematics and Computation, 218(9):5249–5258, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013385.

Mophou:2011:WPA

[Mop11] Gisèle M. Mophou. Weighted pseudo almost automorphic mild solutions to semilinear fractional differential equations. Applied Mathematics and Computation, 217(19): 7579–7587, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311002396.

Mophou:2014:CBF

[Mop14]  Gisèle Mophou.  Con-

trollability of a backward fractional semilinear differential equation. Applied Mathematics and Computation, 242(??):168–178, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007073.

Mortici:2010:BEG

[Mor10a] Cristinel Mortici. Best estimates of the generalized Stirling formula. Applied Mathematics and Computation, 215(11):4044–4048, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010601.

Mortici:2010:EDT

[Mor10b] Cristinel Mortici. Estimating the digamma and trigamma functions by completely monotonicity arguments. Applied Mathematics and Computation, 217(8):4081–4085, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010544.

Mortici:2010:ICT

[Mor10c] Cristinel Mortici. Improved convergence towards generalized Euler– Mascheroni constant. Applied Mathematics and Computation, 215(9):3443– 3448, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009448.

Mortici:2010:NIS

[Mor10d] Cristinel Mortici. New improvements of the Stirling formula. Applied Mathematics and Computation, 217(2):699–704, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300310006818.

Mortici:2010:RFG

[Mor10e] Cristinel Mortici. Ramanujan formula for the generalized Stirling approximation. Applied Mathematics and Computation, 217(6):2579–2585, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310008295.

Morell:2013:OPR

[Mor13] Jorge Alberto Borrego Morell. On orthogonal polynomials with respect to a class of differential operators. Applied Mathematics and Computation, 219(14):7853–7871, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001318.

Mortici:2014:SBG

[Mor14a] Cristinel Mortici. Sharp bounds for gamma function in terms of x^{x−1}. Applied Mathematics and Computation, 249(??): 278–285, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013939.

Mortici:2014:SAW

[Mor14b] Cristinel Mortici. A subtly analysis of Wilker inequality. Applied Mathematics and Computation, 231(??):516–520, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400054X.

Mosayebidorcheh:2014:CNS

[Mos14] Sobhan Mosayebidorcheh. Comments on “Numerical solutions of the spaceand time-fractional coupled Burgers equations by generalized differential transform method, Applied Mathematics and Computation 217 (2011) 7001-7008”. Applied Mathematics and Computation, 243(??):960–962, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008984.

Martinez-Ortiz:2011:MCL

[MOZ11] Carlos Martinez-Ortiz and Jovisa Zunić. Measuring cubeness in the limit cases. Applied Mathematics and Computation, 217(21):8860–8865, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004826.

Mao:2010:FVA

[MP10] Qibo Mao and Stanisław Pietrzko. Free vibration analysis of stepped beams by using Adomian decomposition method. Applied Mathematics and Computation, 217(7):3429–3441, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310009574.

Massey:2011:EAL

[MP11a] Frank Massey and Jeffrey Prentis. Error analysis of large-time approximations for decay chains. Applied Mathematics and Computation, 218(7):3311–3319, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011131.

Mishra:2011:DMW

[MP11b] Bimal Kumar Mishra and Samir Kumar Pandey. Dynamic model of worms with vertical transmission in computer network. Applied Mathematics and Computation, 217(21):8438–8446, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004073.

Mao:2012:FVA

[MP12a] Qibo Mao and Stanisław Pietrzko. Free vibration analysis of a type of tapered beams by using Adomian decomposition method. Applied Mathematics and Computation, 219(6):3264–3271, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009812.

Mary:2012:IAL

[MP12b] Xavier Mary and Pedro Patrício. The inverse along a lower triangular matrix. Applied Mathematics and Computation, 219(3):886–891, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006686.

Massey:2014:PLA

[MP14a] Frank Massey and Jeffrey Prentis. Power law approximations for radioactive decay chains. Applied Mathematics and Computation, 245(??):135–144, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314010054.

Matcovschi:2014:NEE

[MP14b] Mihaela-Hanako Matcovschi and Octavian Pastravanu. Novel estimations for the eigenvalue bounds of complex interval matrices. Applied Mathematics and Computation, 234(??):645–666, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002938.

Matejas:2014:NIM

[MP14c] Josip Matejas and Tunjo Perić. A new iterative method for solving multiobjective linear programming problem. Applied Mathematics and Computation, 243(??):746–754, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008868.

Mishra:2014:GIB

[MP14d] Vishnu Narayan Mishra and Prashantkumar Patel. On generalized integral Bernstein operators based on q-integers. Applied Mathematics and Computation, 242(??):931–944, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314008340.

Mukherjee:2013:ODP

[MPB13] Sayan Mukherjee, Sanjay Kumar Palit, and D. K. Bhattacharya. Is one dimensional Poincaré map sufficient to describe the chaotic dynamics of a three dimensional system? Applied Mathematics and Computation, 219(23):11056–11064, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004785.

Mukherjee:2014:ADD

[MPB14] Sayan Mukherjee, Sanjay Kumar Palit, and Dilip Kumar Bhattacharya. Approximate discrete dynamics of EMG signal. Applied Mathematics and Computation, 243(??):879–888, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009023.

Marcellan:2014:CPL

[MPC14] Francisco Marcellán and Natalia Camila Pinzón-Cortés. (M, N)-coherent pairs of linear functionals and Jacobi matrices. Applied Mathematics and Computation, 232(??):76– 83, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000927.

Milosevic:2014:IMS

[MPM14] D. M. Milosević, M. S.

Petković, and M. R. Milosević. Improved methods for the simultaneous inclusion of multiple polynomial zeros. Applied Mathematics and Computation, 241(??):332–346, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400695X.

Marusic-Paloka:2012:CBD

[MPPM12] Eduard Marusić-Paloka, Igor Pazanin, and Sanja Marusić. Comparison between Darcy and Brinkman laws in a fracture. Applied Mathematics and Computation, 218(14):7538–7545, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000549.

Mazzia:2014:MTA

[MPS14a] Annamaria Mazzia, Giorgio Pini, and Flavio Sartoretto. Meshless techniques for anisotropic diffusion. Applied Mathematics and Computation, 236 (??):54–66, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314003853.

Medved:2014:ESN

[MPS14b] M. Medved’, M. Pospísil, and L. Skripková. On exponential stability of nonlinear fractional multidelay integro-differential equations defined by pairwise permutable matrices. Applied Mathematics and Computation, 227(??):456–468, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011727.

Mathiyalagan:2014:DFA

[MPSA14] K. Mathiyalagan, Ju H.

Park, R. Sakthivel, and S. Marshal Anthoni. Delay fractioning approach to robust exponential stability of fuzzy Cohen–Grossberg neural networks. Applied Mathematics and Computation, 230(??):451–463, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013465.

Mijajlovic:2012:ASF

[MPSD12] Z. Mijajlović, N. Pejović, S. Segan, and G. Damljanović. On asymptotic solutions of Friedmann equations. Applied Mathematics and Computation, 219(3):1273–1286, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007357.

Mudzimbabwe:2012:RNM

[MPW12] W. Mudzimbabwe, K. C. Patidar, and P. J. Witbooi. A reliable numerical method to price arithmetic Asian options. Applied Mathematics and Computation, 218(22):10934–10942, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004559.

Mo:2014:SVM

[MQ14] Yan Mo and Tao Qian. Support vector machine adapted Tikhonov regularization method to solve Dirichlet problem. Applied Mathematics and Computation, 245(??):509–519, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031401056X.

Mihailescu:2011:MSN

[MR11] Mihai Mihăilescu and Dušan Repovš. Multiple solutions for a nonlinear and non-homogeneous problem in Orlicz–Sobolev spaces. Applied Mathematics and Computation, 217(14):6624–6632, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100066X.

Mahdavi:2013:WBA

[MR13a] S. H. Mahdavi and H. Abdul Razak. A waveletbased approach for vibration analysis of framed structures. Applied Mathematics and Computation, 220(??):414–428, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006395.

Martins:2013:MME

[MR13b] Nuno F. M. Martins and Magda Rebelo. A meshfree method for elasticity problems with interfaces. Applied Mathematics and Computation, 219(22):10732–10745, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004852.

Marcellan:2014:SPO

[MR14a] F. Marcellán and A. Sri Ranga. Sieved paraorthogonal polynomials on the unit circle. Applied Mathematics and Computation, 244(??):335–343, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009710.

Matculevich:2014:CED

[MR14b] S. Matculevich and S. Repin. Computable estimates of the distance to the exact solution of the evolutionary reactiondiffusion equation. Applied Mathematics and Computation, 247(??):329–347, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011643.

Molano:2012:FLA

[MRCD12] Rubén Molano, Pablo G. Rodríguez, Andrés Caro, and M. Luisa Durán.

Finding the largest area rectangle of arbitrary orientation in a closed contour. Applied Mathematics and Computation, 218(19):9866–9874, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003207.

Martins:2012:FSE

[MRE^{+}12] J. Martins, H. V. Ribeiro, L. R. Evangelista, L. R. da Silva, and E. K. Lenzi. Fractional Schrödinger equation with noninteger dimensions. Applied Mathematics and Computation, 219(4):2313–2319, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007837.

Mittal:2011:ASC

[MRSS11] M. L. Mittal, B. E.

Rhoades, Smita Sonker, and U. Singh. Approximation of signals of class Lip. Applied Mathematics and Computation, 217(9):4483–4489, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011112.

Maurer:2010:OCM

[MS10] Helmut Maurer and Willi Semmler. An optimal control model of oil discovery and extraction. Applied Mathematics and Computation, 217(3):1163–1169, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003292.

Muthuvalu:2011:HSA

[MS11] Mohana Sundaram Muthuvalu and Jumat Sulaiman. Half-sweep Arithmetic Mean method with composite trapezoidal scheme for solving linear Fredholm integral equations. Applied Mathematics and Computation, 217(12):5442–5448, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012130.

Malik:2012:DFC

[MS12a] Pradeep Malik and A. Swaminathan. Derivatives of a finite class of orthogonal polynomials related to inverse Gamma distribution. Applied Mathematics and Computation, 218(11):6251–6262, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101424X.

Mohanty:2012:NHA

[MS12b] R. K. Mohanty and Nikita Setia. A new high accuracy two-level implicit off-step discretization for the system of two space dimensional quasi-linear parabolic partial differential equations. Applied Mathematics and Computation, 219(5):2680–2697, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008934.

Mahmoud:2013:DAD

[MS13a] Magdi S. Mahmoud and Abdul-Wahid A. Saif. Dissipativity analysis and design for uncertain Markovian jump systems with time-varying delays. Applied Mathematics and Computation, 219(18): 9681–9695, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002774.

Malik:2013:IEU

[MS13b] Sachin Malik and Rajesh Singh. An improved estimator using two auxiliary attributes. Applied Mathematics and Computation, 219(23):10983–10986, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300533X.

Mansour:2013:GTP

[MS13c] Toufik Mansour and Matthias Schork. The generalized Touchard polynomials revisited. Applied Mathematics and Computation, 219(19):9978–9991, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004098.

Mansour:2013:PWC

[MS13d] Toufik Mansour and Mark Shattuck. Polynomials whose coefficients are generalized Tribonacci numbers. Applied Mathematics and Computation, 219(15):8366–8374, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013288.

Mathiaud:2013:NSR

[MS13e] Julien Mathiaud and Francesco Salvarani. A numerical strategy for radiative transfer problems with highly oscillating opacities. Applied Mathematics and Computation, 221(??):249–256, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006966.

Meng:2013:SNV

[MS13f] Fanwei Meng and Jing Shao. Some new Volterra– Fredholm type dynamic integral inequalities on time scales. Applied Mathematics and Computation, 223(??):444–451, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300893X.

Mnatsakanov:2013:NRD

[MS13g] Robert M. Mnatsakanov and Khachatur Sarkisian. A note on recovering the distributions from exponential moments. Applied Mathematics and Computation, 219(16):8730–8737, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001951.

Mateus:2014:NAS

[MS14a] Joaquim P. Mateus and César M. Silva. A nonautonomous SEIRS model with general incidence rate. Applied Mathematics and Computation, 247(??):169–189, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011886.

Matsumoto:2014:DCD

[MS14b] Akio Matsumoto and Ferenc Szidarovszky. Discrete and continuous dynamics in nonlinear monopolies. Applied Mathematics and Computation, 232(??):632–642, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314001386.

Mohanasubha:2013:NCB

[MSC^{+}13] R. Mohanasubha, Jane H.

Sheeba, V. K. Chandrasekar, M. Senthilvelan, and M. Lakshmanan. A nonlocal connection between certain linear and nonlinear ordinary differential equations — part ii: Complex nonlinear oscillators. Applied Mathematics and Computation, 224(??):593–602, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009521.

Misra:2011:EOF

[MSCK11] J. C. Misra, G. C. Shit, S. Chandra, and P. K. Kundu. Electro-osmotic flow of a viscoelastic fluid in a channel: Applications to physiological fluid mechanics. Applied Mathematics and Computation, 217(20):7932–7939, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002980.

Martinez:2014:AQP

[MSGN^{+}14] J. Martínez, Á. Saavedra, P. J. García-Nieto, J. I. Piñeiro, C. Iglesias, J. Taboada, J. Sancho, and J. Pastor. Air quality parameters outliers detection using functional data analysis in the Langreo urban area (Northern Spain). Applied Mathematics and Computation, 241(??):1–10, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006602.

Motsepe:2014:AHF

[MSJ14] K. A. Motsepe, M. Y. Shatalov, and S. V. Joubert. Application of the Heun functions to rotor vibratory gyroscope, their numerical computation and accuracy estimation. Applied Mathematics and Computation, 239(??):47– 55, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005414.

Mishra:2014:NSD

[MSR14] T. N. Mishra, Saikat Sarkar, and K. N. Rai. Numerical solution of dualphase-lagging heat conduction model for analyzing overshooting phenomenon. Applied Mathematics and Computation, 236(??):693–708, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004019.

Mohanty:2014:NHO

[MSS14a] R. K. Mohanty, Suruchi Singh, and Swarn Singh. A new high order space derivative discretization for 3D quasi-linear hyperbolic partial differential equations. Applied Mathematics and Computation, 232(??):529–541, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001015.

Moon:2014:MMD

[MSS14b] Ilkyeong Moon, Eunjoo Shin, and Biswajit Sarkar. Min-max distribution free continuous-review model with a service level constraint and variable lead time. Applied Mathematics and Computation, 229(??):310–315, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013106.

Malinowska:2010:LDM

[MT10] Agnieszka B. Malinowska and Delfim F. M. Torres. Leitmann’s direct method of optimization for absolute extrema of certain problems of the calculus of variations on time scales. Applied Mathematics and Computation, 217(3):1158–1162, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000287.

Malinowska:2012:MFV

[MT12a] Agnieszka B. Malinowska and Delfim F. M. Torres. Multiobjective fractional variational calculus in terms of a combined Caputo derivative. Applied Mathematics and Computation, 218(9):5099–5111, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031101321X.

Mohanty:2012:CAU

[MT12b] R. K. Mohanty and Jyoti Talwar. A combined approach using coupled reduced alternating group explicit (CRAGE) algorithm and sixth order off-step discretization for the solution of two point nonlinear boundary value problems. Applied Mathematics and Computation, 219(1):248–259, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005978.

Miao:2013:DRP

[MT13] Hui Miao and Yu-Chu Tian. Dynamic robot path planning using an enhanced simulated annealing approach. Applied Mathematics and Computation, 222(??):420–437, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007728.

Malik:2014:NGI

[MT14] Saroj B. Malik and Néstor Thome. On a new generalized inverse for matrices of an arbitrary index. Applied Mathematics and Computation, 226(??):575–580, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301117X.

Mao:2010:AMC

[MTD10] Yu Mao, Wallace K. S. Tang, and Marius-F. Danca. An averaging model for chaotic system with periodic time-varying parameter. Applied Mathematics and Computation, 217(1):355–362, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006272.

Marrero:2014:EFP

[MTER14] J. Abderramán Marrero, R. Ben Taher, Y. El Khatabi, and M. Rachidi. On explicit formulas of the principal matrix p th root by polynomial decompositions. Applied Mathematics and Computation, 242(??):435–443, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314008108.

Marrero:2013:EFP

[MTR13] J. Abderramán Marrero, R. Ben Taher, and M. Rachidi. On explicit formulas for the principal matrix logarithm. Applied Mathematics and Computation, 220(??):142–148, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006127.

Mahdiloo:2014:GTA

[MTSN14] Mahdi Mahdiloo, Madjid Tavana, Reza Farzipoor Saen, and Abdollah Noorizadeh. A game theoretic approach to modeling undesirable outputs and efficiency decomposition in data envelopment analysis. Applied Mathematics and Computation, 244(??):479–492, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009928.

Massar:2014:MSN

[MTT14] M. Massar, M. Talbi, and N. Tsouli. Multiple solutions for nonlocal system of (p(x), q(x))-kirchhoff type. Applied Mathematics and Computation, 242(??):216–226, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400722X.

Meng:2010:PGV

[MTZ10] Xiang-Hua Meng, Bo Tian, and Hai-Qiang Zhang. Pfaffianization of the generalized variable-coefficient Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 217(4):1300–1305, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309004408.

Massoudi:2012:CRF

[MU12] Mehrdad Massoudi and A. K. Uguz. Chemicallyreacting fluids with variable transport properties. Applied Mathematics and Computation, 219(4):1761–1775, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312008089.

Mukherjee:2010:HBE

[Muk10] Debasis Mukherjee. Hopf bifurcation in an ecoepidemic model. Applied Mathematics and Computation, 217(5):2118–2124, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310007551.

Muravskii:2011:CSI

[Mur11] G. B. Muravskii. Choice of specific interpolation points in quadrature formula for fast oscillating integrals. Applied Mathematics and Computation, 218(6):2973–2978, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010423.

Muravskii:2013:CSI

[Mur13] G. B. Muravskii. Choice of specific interpolation points in quadrature formula for integrals containing Bessel functions. Applied Mathematics and Computation, 224(??): 724–734, November 1, 2013. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009119.

Muroya:2014:GSD

[Mur14] Yoshiaki Muroya. Global stability of a delayed nonlinear Lotka–Volterra system with feedback controls and patch structure. Applied Mathematics and Computation, 239(??):60– 73, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005694.

Mustafa:2014:DRT

[Mus14] Muhammad I. Mustafa. On the decay rates for thermoviscoelastic systems of type III. Applied Mathematics and Computation, 239(??):29–37, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005827.

Muller:2011:ANS

[MV11] Frank Müller and Werner Varnhorn. On approximation and numerical solution of Fredholm integral equations of second kind using quasiinterpolation. Applied Mathematics and Computation, 217(13):6409–6416, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000385.

Massoudi:2012:ASS

[MV12] Mehrdad Massoudi and Ashwin Vaidya. Analytical solutions to Stokes-type flows of inhomogeneous fluids. Applied Mathematics and Computation, 218(11):6314–6329, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014561.

Misra:2013:MMS

[MV13] A. K. Misra and Maitri Verma. A mathematical model to study the dynamics of carbon dioxide gas in the atmosphere. Applied Mathematics and Computation, 219(16):8595–8609, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313001963.

Mallory:2014:MCA

[MV14] Kristina Mallory and Robert A. Van Gorder. Method for constructing analytical solutions to the dym initial value problem. Applied Mathematics and Computation, 226(??):67–82, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010746.

Mitchell:2011:ANS

[MVGS11] S. L. Mitchell, M. Vynnycky, I. G. Gusev, and S. S. Sazhin. An accurate numerical solution for the transient heating of an evaporating spherical droplet. Applied Mathematics and Computation, 217(22):9219–9233, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005418.

Meenakshi:2010:SPS

[MVM10] P. Maragatha Meenakshi, S. Valarmathi, and J. J. H. Miller. Solving a partially singularly perturbed initial value problem on Shishkin meshes. Applied Mathematics and Computation, 215(9):3170–3180, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309008637.

Martin-Vaquero:2012:NAD

[MVQDE12] J. Martín-Vaquero, A. Queiruga-Dios, and A. H. Encinas. Numerical algorithms for diffusion-reaction problems with non-classical conditions. Applied Mathematics and Computation, 218(9):5487–5495, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013671.

Marquez-Vera:2014:FIL

[MVRVSCMV14] M. A. Márquez-Vera, L. E.

Ramos-Velasco, J. Suárez-Cansino, and C. A. Márquez-Vera. Fuzzy iterative learning control applied in a biological reactor using a reduced number of measures. Applied Mathematics and Computation, 246(??):608–618, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011825.

Mincic:2013:SNS

[MVS13] Svetislav M. Mincić, Ljubica S. Velimirović, and Mića S. Stanković. On spaces with non-symmetric affine connection, containing subspaces without torsion. Applied Mathematics and Computation, 219(9):4346–4353, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010041.

Mincic:2014:ICD

[MVS14a] Svetislav M. Mincić, Ljubica S. Velimirović, and Mića S. Stanković. Integrability conditions of derivational equations of a submanifold of a generalized Riemannian space. Applied Mathematics and Computation, 226(??):3–9, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010710.

Misra:2014:CIB

[MVS14b] A. K. Misra, Maitri Verma, and Anupama Sharma. Capturing the interplay between malware and anti-malware in a computer network. Applied Mathematics and Computation, 229(??): 340–349, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313013428.

Miao:2014:HTF

[MWAM14] Ling Miao, Wei-Tao Wu, Nadine Aubry, and Mehrdad Massoudi. Heat transfer and flow of a slag-type non-linear fluid: Effects of variable thermal conductivity. Applied Mathematics and Computation, 227(??):77–91, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011703.

Ma:2011:UCA

[MWC11] Qiaozhen Ma, Suping Wang, and Xiaobao Chen. Uniform compact attractors for the coupled suspension bridge equations. Applied Mathematics and Computation, 217(14):6604–6615, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000610.

Ma:2010:TSW

[MWJW10] Jun Ma, Chun-Ni Wang, Wu-Yin Jin, and Ying Wu. Transition from spiral wave to target wave and other coherent structures in the networks of Hodgkin–Huxley neurons. Applied Mathematics and Computation, 217(8):3844–3852, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010039.

Miao:2013:ITC

[MWY13] Changqing Miao, Yintao Wei, and Xiangqiao Yan. Interactions of two collinear circular hole cracks subjected to internal pressure. Applied Mathematics and Computation, 223(??):216–224, October 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300313008527.

Mo:2011:AOF

[MX11] Hongmin Mo and Shuhuang Xiang. On the asymptotic order of Filon-type methods for highly oscillatory integrals with an algebraic singularity. Applied Mathematics and Computation, 217(22):9105–9110, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005157.

Ma:2010:NIM

[MXH10] Zhongyan Ma, Huiqing Xie, and Yaping Hu. A new iterative method for many eigenpair partial derivatives. Applied Mathematics and Computation, 217(5):1991–1996, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007423.

Ma:2012:GSP

[MXH12] Ruyun Ma, Jia Xu, and Xiaoling Han. Global structure of positive solutions for superlinear second-order periodic boundary value problems. Applied Mathematics and Computation, 218(10): 5982–5988, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300311014251.

Ma:2010:ICS

[MYJ10] Zhongjun Ma, Jian Yang, and Guirong Jiang. Impulsive control in a stage structure population model with birth pulses. Applied Mathematics and Computation, 217(7):3453–3460, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009598.

Ma:2010:PFF

[MYL10a] Suzhen Ma, Yongjian Yang, and Huaqun Liu. A parameter free filled function for unconstrained global optimization. Applied Mathematics and Computation, 215(10): 3610–3619, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300309009734.

Moon:2010:NSS

[MYL^{+}10b] B. S. Moon, D. S. Yoo, Y. H. Lee, I. S. Oh, J. W. Lee, D. Y. Lee, and K. C. Kwon. A non-separable solution of the diffusion equation based on the Galerkin’s method using cubic splines. Applied Mathematics and Computation, 217(5):1831–1837, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031000562X.

Ma:2010:TFV

[MYT10] Manjun Ma, Daoming Yang, and Hengsheng Tang. Traveling fronts of the volume-filling chemotaxis model with general kinetics. Applied Mathematics and Computation, 216(11):3162–3171, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004546.

Ma:2013:PPS

[MYW13] Ruyun Ma, Bianxia Yang, and Zhenyan Wang. Positive periodic solutions of first-order delay differential equations with impulses. Applied Mathematics and Computation, 219(11):6074–6083, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012969.

Ma:2012:SDG

[MZ12] Wen-Xiu Ma and Zuonong Zhu. Solving the (3 + 1)-dimensional generalized KP and BKP equations by the multiple exp-function algorithm. Applied Mathematics and Computation, 218(24):11871–11879, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005644.

Maagli:2013:SFD

[MZ13a] Habib Mâagli and Malek Zribi. On a semilinear fractional Dirichlet problem on a bounded domain. Applied Mathematics and Computation, 222(??):331–342, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007911.

Mai:2013:DSP

[MZ13b] Ali Mai and Zhan Zhou.

Discrete solitons for periodic discrete nonlinear Schrödinger equations. Applied Mathematics and Computation, 222(??):34–41, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007923.

Mu:2007:NBB

[MZL07] Xuewen Mu, Yaling Zhang, and Sanyang Liu. A new branch and bound method with pretreatment for the binary quadratic programming. Applied Mathematics and Computation, 192(1):252–259, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300307003293.

See erratum [Pak11a].

Zhang:2011:LIS

[mZT11] Qi ming Zhang and X. H.

Tang. Lyapunov inequalities and stability for discrete linear Hamiltonian systems. Applied Mathematics and Computation, 218(2):574–582, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100782X.

Ma:2012:HBE

[MZTT12] Wen-Xiu Ma, Yi Zhang, Yaning Tang, and Junyi Tu. Hirota bilinear equations with linear subspaces of solutions. Applied Mathematics and Computation, 218(13):7174–7183, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031101575X.

Ma:2014:ASS

[MZW14] Weijun Ma, Qimin Zhang, and Zhanping Wang. Asymptotic stability of stochastic age-dependent population equations with Markovian switching. Applied Mathematics and Computation, 227(??): 309–319, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011661.

Mu:2012:MST

[MZWZ12] Yue Mu, Guoqun Zhao, Xianghong Wu, and Jiqiang Zhai. Modeling and simulation of three-dimensional planar contraction flow of viscoelastic fluids with PTT, Giesekus and FENE-P constitutive models. Applied Mathematics and Computation, 218(17): 8429–8443, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312001166.

Mi:2010:BSA

[MZX10] Lufang Mi, Pingping Zhang, and Huazhong Xu. Boundedness of solutions for asymmetric Duffing equations. Applied Mathematics and Computation, 217(7):3102–3112, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008969.

Ma:2010:ODA

[MZXZ10] Jun Ma, Ai-Hua Zhang, Ya-Feng Xia, and Li-Ping Zhang. Optimize design of adaptive synchronization controllers and parameter observers in different hyperchaotic systems. Applied Mathematics and Computation, 215(9):3318–3326, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009254.

Nadeem:2010:ITD

[NA10a] S. Nadeem and Noreen Sher Akbar. Influence of temperature dependent viscosity on peristaltic transport of a Newtonian fluid: Application of an endoscope. Applied Mathematics and Computation, 216(12):3606–3619, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310005242.

Nieto:2010:ASI

[NA10b] Juan J. Nieto and Bashir Ahmad. Approximation of solutions for an initial and terminal value problem for the forced Duffing equation with nonviscous damping. Applied Mathematics and Computation, 216(7):2129–2136, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002882.

Nashine:2014:NFP

[NA14a] Hemant Kumar Nashine and Ishak Altun. New fixed point results for maps satisfying implicit relations on ordered metric spaces and application. Applied Mathematics and Computation, 240(??):259–272, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006134.

Nieto:2014:NAQ

[NA14b] P. J. García Nieto and J. C. Álvarez Antón. Nonlinear air quality modeling using multivariate adaptive regression splines in Gijón urban area (northern spain) at local scale. Applied Mathematics and Computation, 235(??):50– 65, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003440.

Nayak:2013:NCC

[NAA13a] M. M. Nayak, M. Acharya, and B. P. Acharya. Numerical computation of complex Cauchy principal value integrals. Applied Mathematics and Computation, 219(18):9696–9699, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002993.

Nayak:2013:NET

[NAA13b] M. M. Nayak, M. Acharya, and B. P. Acharya. Numerical evaluation of two-dimensional Cauchy Principal Value integrals. Applied Mathematics and Computation, 225(??):807–810, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010722.

Nadarajah:2007:DMK

[Nad07] Saralees Nadarajah. On the distribution of Ma and King. Applied Mathematics and Computation, 189(1):732–733, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030601650X.

See comment [Ma11a].

Nadeem:2011:PLF

[NAHH11] S. Nadeem, Noreen Sher Akbar, Awatif A. Hendi, and T. Hayat. Power law fluid model for blood flow through a tapered artery with a stenosis. Applied Mathematics and Computation, 217(17):7108–7116, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000427.

Najdanovic:2014:IBI

[Naj14] Marija S. Najdanović. Infinitesimal bending influence on the volume change. Applied Mathematics and Computation, 243(??):801–808, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008686.

Nakamori:2013:DRF

[Nak13] Seiichi Nakamori. Design of RLS-FIR filter using covariance information in linear continuoustime stochastic systems. Applied Mathematics and Computation, 219(17): 9598–9608, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313002750.

Nave:2012:ADF

[NAL12] Ophir Nave, Suraj Ajadi, and Yaron Lehavi. Analysis of the dynamics of fuel spray using asymptotic methods: The method of integral invariant manifolds. Applied Mathematics and Computation, 218(10):5877–5890, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013609.

Narendar:2012:DQB

[Nar12] S. Narendar. Differential quadrature based nonlocal flapwise bending vibration analysis of rotating nanotube with consideration of transverse shear deformation and rotary inertia. Applied Mathematics and Computation, 219(3):1232–1243, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007321.

Narushima:2013:SCG

[Nar13] Yasushi Narushima. A smoothing conjugate gradient method for solving systems of nonsmooth equations. Applied Mathematics and Computation, 219(16):8646–8655, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001987.

Nam:2014:IFA

[NATQ14] Phan Thanh Nam, Pham Ngoc Dinh Alain, Dang Duc Trong, and Pham Hoang Quan. Ice formation in the Arctic during summer: False-bottoms. Applied Mathematics and Computation, 227(??):857–870, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011636.

Naz:2010:GIS

[Naz10] R. Naz. Group invariant solution for a free jet on a hemi-spherical shell. Applied Mathematics and Computation, 215(9):3265–3270, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009138.

Nazari:2011:IEP

[NB11] A. M. Nazari and Z. Beiranvand. The inverse eigenvalue problem for symmetric quasi anti-bidiagonal matrices. Applied Mathematics and Computation, 217(23):9526–9531, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003973.

Nave:2010:PMT

[NBG10] Ophir Nave, Vitcheslav Bykov, and Vladimir Gol’dshtein. A probabilistic model of thermal explosion in polydisperse fuel spray. Applied Mathematics and Computation, 217(6):2698–2709, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000857X.

Nayak:2010:MKP

[NBP10] R. K. Nayak, M. P. Biswal, and S. Padhy. Modification of Karmarkar’s projective scaling algorithm.

Applied Mathematics and Computation, 216(1):227– 235, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310000561.

Navickas:2010:GEF

[NBR10] Zenonas Navickas, Liepa Bikulciene, and Minvydas Ragulskis. Generalization of Exp-function and other standard function methods. Applied Mathematics and Computation, 216(8):2380–2393, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003371.

Nguyen-Ba:2011:PPA

[NBYHV11] T. Nguyen-Ba, H. Yagoub, H. Hao, and R. Vaillancourt. Pryce pre-analysis adapted to some DAE solvers. Applied Mathematics and Computation, 217(21):8403–8418, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004036.

Natali:2013:SPP

[NC13a] Fábio Natali and Eleomar Cardoso, Jr. Stability properties of periodic waves for the Klein–Gordon equation with quintic nonlinearity. Applied Mathematics and Computation, 224(??):581–592, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009314.

Neta:2013:FLM

[NC13b] Beny Neta and Changbum Chun. On a family of Laguerre methods to find multiple roots of nonlinear equations. Applied Mathematics and Computation, 219(23):10987–11004, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004918.

See corrigendum [NC14].

Neta:2014:CFL

[NC14] Beny Neta and Changbum Chun. Corrigendum to “On a family of Laguerre methods to find multiple roots of nonlinear equations”. Applied Mathematics and Computation, 248(??):693–696, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013642.

See [NC13b].

Nakamori:2013:RWE

[NCÁHCLP13] Seiichi Nakamori, Raquel Caballero-Águila, Aurora Hermoso-Carazo, and Josefa Linares-Pérez. RLS Wiener estimators from observations with multiple and random delays in linear discrete-time stochastic systems. Applied Mathematics and Computation, 225(??):184–194, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009855.

Norambuena-Contreras:2012:NNS

[NCANCF12] J. Norambuena-Contreras, G. Arbat, P. J. García Nieto, and D. Castro-Fresno. Nonlinear numerical simulation of rainwater infiltration through road embankments by FEM. Applied Mathematics and Computation, 219(4):1843–1852, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008181.

Nieto:2013:SBR

[NCdCDM13] P. J. García Nieto, E. F. Combarro, J. J. del Coz Díaz, and E. Montañés. A SVM-based regression model to study the air quality at local scale in Oviedo urban area (Northern Spain): a case study. Applied Mathematics and Computation, 219(17):8923–8937, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002713.

Nan:2011:PPS

[NCL11] Zhijie Nan, Weijun Chen, and Lin Li. Positive periodic solutions for a class of delay differential equations. Applied Mathematics and Computation, 218(8):4647–4650, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012434.

Neta:2012:NMS

[NCS12] Beny Neta, Changbum Chun, and Melvin Scott. A note on the modified super-Halley method. Applied Mathematics and Computation, 218(18): 9575–9577, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312002780.

Neta:2013:DIM

[NCS13] Beny Neta, Changbum Chun, and Melvin Scott. On the development of iterative methods for multiple roots. Applied Mathematics and Computation, 224(??):358–361, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009454.

Neta:2014:BAO

[NCS14] Beny Neta, Changbum Chun, and Melvin Scott. Basins of attraction for optimal eighth order methods to find simple roots of nonlinear equations. Applied Mathematics and Computation, 227(??): 567–592, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011843.

Niklis:2014:CMA

[NDZ14] Dimitrios Niklis, Michael Doumpos, and Constantin Zopounidis. Combining market and accountingbased models for credit scoring using a classification scheme based on support vector machines. Applied Mathematics and Computation, 234(??):69– 81, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002677.

Nemani:2010:FAS

[Nem10a] S. S. Nemani. A fast algorithm for solving Toeplitz penta-diagonal systems. Applied Mathematics and Computation, 215(11):3830–3838, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010157.

Nemeth:2010:EFE

[Ném10b]  S. Z. Németh.  Ex-

ceptional family of elements for general order complementarity problems. Applied Mathematics and Computation, 217(5):2184–2190, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007630.

Neuman:2011:SIG

[Neu11] Edward Neuman. Some inequalities for the gamma function. Applied Mathematics and Computation, 218(8):4349–4352, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012409.

Neuman:2012:IJE

[Neu12a] Edward Neuman. Inequalities for Jacobian elliptic functions and Gauss lemniscate functions. Applied Mathematics and Computation, 218(15):7774–7782, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000793.

Neustupa:2012:ASV

[Neu12b] Tomás Neustupa. The analysis of stationary viscous incompressible flow through a rotating radial blade machine, existence of a weak solution. Applied Mathematics and Computation, 219(7):3316–3322, December 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006758.

Neuman:2013:BEI

[Neu13] Edward Neuman. Bounds for entries of the inverse matrix of a tridiagonal matrix. Applied Mathematics and Computation, 219(17):9210–9212, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003457.

Neuman:2014:WHT

[Neu14] Edward Neuman. Wilker and Huygens-type inequalities for the generalized trigonometric and for the generalized hyperbolic functions. Applied Mathematics and Computation, 230(??):211–217, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014264.

Ngoc:2012:ESV

[Ngo12] Pham Huu Anh Ngoc.

On exponential stability of Volterra differential equations with delay. Applied Mathematics and Computation, 218(24):11922–11930, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005772.

Nguyen:2014:LDB

[Ngu14] Dinh Kien Nguyen. Large displacement behaviour of tapered cantilever Euler– Bernoulli beams made of functionally graded material. Applied Mathematics and Computation, 237(??):340–355, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004755.

Nehi:2012:CBC

[NH12] Hasan Mishmast Nehi and Farhad Hamidi. A comment on “Biological computation of the solution to the quadratic assignment problem”. Applied Mathematics and Computation, 218(21):10759– 10761, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004298.

Nakamori:2010:DRW

[NHCLP10] Seiichi Nakamori, Aurora Hermoso-Carazo, and Josefa Linares-Pérez. Design of RLS Wiener estimators from randomly delayed observations in linear discrete-time stochastic systems. Applied Mathematics and Computation, 217(8):3801–3815, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009999.

Ni:2013:ECP

[Ni13] Yaodong Ni. Edge covering problem under hybrid uncertain environments. Applied Mathematics and Computation, 219(11):6044–6052, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312012891.

Niezgoda:2011:GOT

[Nie11] Marek Niezgoda. Grüss and Ostrowski type inequalities. Applied Mathematics and Computation, 217(23):9779–9789, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006369.

Nie:2013:DCD

[Nie13] Tao Nie. A decoupled and conservative difference scheme with fourth-order accuracy for the Symmetric Regularized Long Wave equations. Applied Mathematics and Computation, 219(17):9461–9468, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003378.

Niu:2011:PAT

[Niu11] Lingfeng Niu. Parallel algorithm for training multiclass proximal Support Vector Machines. Applied Mathematics and Computation, 217(12):5328–5337, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310011860.

Noor:2012:NIM

[NK12] Muhammad Aslam Noor and Waseem Asghar Khan. New iterative methods for solving nonlinear equation by using homotopy perturbation method. Applied Mathematics and Computation, 219(8):3565–3574, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009393.

Ngamsaad:2010:EBC

[NKM^{+}10] Waipot Ngamsaad, Paisan Kanthang, Charin Modchang, Somchai Sriyab, and Wannapong Triampo. The effect of boundary conditions on the mesoscopic lattice Boltzmann method: Case study of a reaction–diffusion based model for Min-protein oscillation. Applied Mathematics and Computation, 217(6):2339–2347, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310007782.

Nashine:2012:CCF

[NKR12] Hemant Kumar Nashine, Zoran Kadelburg, and Stojan Radenovíc. Coupled common fixed point theorems for w^{∗}-compatible mappings in ordered cone metric spaces. Applied Mathematics and Computation, 218(9):5422–5432, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013592.

Niu:2010:FAF

[NL10] Qiang Niu and Lin-Zhang Lu. Fourier analysis of frequency filtering decomposition preconditioners. Applied Mathematics and Computation, 216(6):1805–1818, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010716.

Nguyen:2012:CAW

[NL12a] Huyen Nguyen and John Lekner. Construction of accelerating wavepackets. Applied Mathematics and Computation, 218 (22):10990–10997, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004596.

Niu:2012:NDP

[NL12b] Jiuxiao Niu and Deyou Liu. A new delayed projection neural network for solving quadratic programming problems subject to linear constraints. Applied Mathematics and Computation, 219(6):3139–3146, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009460.

Navarro-Lopez:2010:CPC

[NLLC10] Eva M. Navarro-López and Eduardo Licéaga-Castro. Combining passivity and classical frequencydomain methods: an insight into decentralised control. Applied Mathematics and Computation, 215(12):4426–4438, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310000251.

Noor:2008:AFG

[NM08] Khalida Inayat Noor and Ali Muhammad. On analytic functions with generalized bounded Mocanu variation. Applied Mathematics and Computation, 196(2):802–811, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300307007321.

See comments [WZ13c].

Nadeau:2014:GSE

[NM14] Julie Nadeau and C. Connell McCluskey. Global stability for an epidemic model with applications to feline infectious peritonitis and tuberculosis. Applied Mathematics and Computation, 230(??):473–483, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014148.

Noor:2010:EFM

[NMDWAS10] Muhammad Aslam Noor, Syed Tauseef Mohyud-Din, Asif Waheed, and Eisa A. Al-Said. Expfunction method for traveling wave solutions of nonlinear evolution equations. Applied Mathematics and Computation, 216(2):477–483, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000055X.

See note [RC12].

Nasser:2011:BIE

[NMIA11] M. M. S. Nasser, A. H. M. Murid, M. Ismail, and E. M. A. Alejaily. Boundary integral equations with the generalized Neumann kernel for Laplace’s equation in multiply connected regions. Applied Mathematics and Computation, 217(9):4710–4727, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011392.

Nanda:2014:MPB

[NMM14] Bharadwaj Nanda, Damodar Maity, and Dipak K. Maiti. Modal parameter based inverse approach for structural joint damage assessment using unified particle swarm optimization. Applied Mathematics and Computation, 242(??):407–422, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008157.

Nassar:2014:RSN

[NMR14] S. Hadj Nassar, H. Moussa, and M. Rhoudaf. Renormalized solution for a nonlinear parabolic problems with noncoercivity in divergence form in Orlicz spaces. Applied Mathematics and Computation, 249(??):253–264, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013927.

Ndogmo:2011:HOA

[NN11a] J. C. Ndogmo and D. B.

Ntwiga. High-order accurate implicit methods for barrier option pricing. Applied Mathematics and Computation, 218(5):2210–2224, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009702.

Nemes:2011:NLC

[NN11b] Gergö Nemes and Antal Nemes. A note on the Landau constants. Applied Mathematics and Computation, 217(21):8543–8546, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004267.

Naidu:2013:ACA

[NN13] V. Kesavulu Naidu and K. V. Nagaraja. Advantages of cubic arcs for approximating curved boundaries by subparametric transformations for some higher order triangular elements. Applied Mathematics and Computation, 219(12):6893–6910, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000386.

Njagarah:2014:MMC

[NN14] J. B. H. Njagarah and F. Nyabadza. A metapopulation model for cholera transmission dynamics between communities linked by migration. Applied Mathematics and Computation, 241(??):317–331, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007012.

Nowruzi:2019:ESC

[NN19] Hashem Nowruzi and S. Salman Nourazar. Erratum to “On the solution of characteristic value problems arising in linear stability analysis; semi analytical approach”. Applied Mathematics and Computation, 350(??):462, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031930044X.

See [HN14a].

Ngwa:2010:MAR

[NNG10] Gideon A. Ngwa, Ashrafi M.

Niger, and Abba B. Gumel. Mathematical assessment of the role of non-linear birth and maturation delay in the population dynamics of the malaria vector. Applied Mathematics and Computation, 217(7):3286–3313, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310009264.

Noor:2014:PSS

[NNK14] Muhammad Aslam Noor, Khalida Inayat Noor, and Awais Gul Khan. Parallel schemes for solving a system of extended general quasi variational inequalities. Applied Mathematics and Computation, 245(??):566–574, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011527.

Nave:2014:HMI

[NNPH14] Ophir Nave, Yair Neuman, Leonid Perlovsky, and Newton Howard. How much information should we drop to become intelligent? Applied Mathematics and Computation, 245(??):261–264, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009862.

Ng:2010:CBN

[NOS10] Choung Min Ng, Seng-Huat Ong, and H. M. Srivastava. A class of bivariate negative binomial distributions with different index parameters in the marginals. Applied Mathematics and Computation, 217(7):3069–3087, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008945.

Novac:2014:SPS

[Nov14] Adela Novac. Summation of power series with rational coefficients via divided differences. Applied Mathematics and Computation, 236(??):345–350, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004044.

Nam:2010:IDD

[NP10a] Phan T. Nam and Pubudu N. Pathirana. Improvement on delay dependent absolute stability of Lurie control systems with multiple time delays. Applied Mathematics and Computation, 216(3):1024–1027, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310001141.

Neta:2010:COO

[NP10b] B. Neta and M. S. Petković. Construction of optimal order nonlinear solvers using inverse interpolation. Applied Mathematics and Computation, 217(6):2448–2455, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007903.

Naimzada:2014:ADD

[NP14] Ahmad Naimzada and Marina Pireddu. Adaptive decision dynamics: Bifurcations, multistability and chaos. Applied Mathematics and Computation, 239(??):375–390, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006043.

Nandi:2011:DNC

[NPA11] Mrinal Nandi, Subrata Parui, and Avishek Adhikari. The domination numbers of cylindrical grid graphs. Applied Mathematics and Computation, 217(10):4879–4889, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011318.

Naimzada:2012:SHA

[NR12] Ahmad K. Naimzada and Giorgio Ricchiuti. Studying heterogeneity among fundamentalists in financial markets: a note. Applied Mathematics and Computation, 219(3):792–799, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006315.

Narayan:2013:UFC

[NR13] S. P. Atul Narayan and K. R. Rajagopal. Unsteady flows of a class of novel generalizations of the Navier–Stokes fluid. Applied Mathematics and Computation, 219(19): 9935–9946, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313003020.

Navickas:2014:CNA

[NR14] Zenonas Navickas and Minvydas Ragulskis. Comments on “A new algorithm for automatic computation of solitary wave solutions to nonlinear partial differential equations based on the Exp-function method”. Applied Mathematics and Computation, 243(??):419–425, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008650.

Nunez:2014:CPF

[NRAV14] Juan Núñez, María Luisa Rodríguez-Arévalo, and María Trinidad Villar. Certain particular families of graphicable algebras. Applied Mathematics and Computation, 246(??):416–425, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011011.

Nieto:2012:SOL

[NRL12] Juan J. Nieto and Rosana Rodri’guez-López. Secondorder linear differential equations with piecewise constant arguments subject to nonlocal boundary conditions. Applied Mathematics and Computation, 218(19):9647–9656, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002214.

Nagaiah:2012:AST

[NRWF12] Ch. Nagaiah, S. Rüdiger, G. Warnecke, and M. Falcke. Adaptive space and time numerical simulation of reaction-diffusion models for intracellular calcium dynamics. Applied Mathematics and Computation, 218(20):10194– 10210, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003633.

Nilsrakoo:2011:SCT

[NS11] Weerayuth Nilsrakoo and Satit Saejung. Strong convergence theorems by Halpern–Mann iterations for relatively nonexpansive mappings in Banach spaces. Applied Mathematics and Computation, 217(14):6577–6586, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311000567.

Neuman:2012:IHF

[NS12] Edward Neuman and József Sándor. Inequalities for hyperbolic functions. Applied Mathematics and Computation, 218(18):9291–9295, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002391.

Neta:2013:FHL

[NS13a] Beny Neta and Melvin Scott. On a family of Halley-like methods to find simple roots of nonlinear equations. Applied Mathematics and Computation, 219(15):7940–7944, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001574.

Nunokawa:2013:IOC

[NS13b] Mamoru Nunokawa and Janusz Sokól. An improvement of Ozaki’s condition. Applied Mathematics and Computation, 219(22):10768–10776, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300489X.

Neta:2012:BAV

[NSC12a] Beny Neta, Melvin Scott, and Changbum Chun. Basin attractors for various methods for multiple roots. Applied Mathematics and Computation, 218(9):5043–5066, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013178.

Neta:2012:BAS

[NSC12b] Beny Neta, Melvin Scott, and Changbum Chun. Basins of attraction for several methods to find simple roots of nonlinear equations. Applied Mathematics and Computation, 218(21):10548– 10556, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004018.

Naheed:2014:NSS

[NSL14] Afia Naheed, Manmohan Singh, and David Lucy. Numerical study of SARS epidemic model with the inclusion of diffusion in the system. Applied Mathematics and Computation, 229(??):480–498, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013453.

Nik:2014:HPS

[NSNEV14] H. Saberi Nik, J. Saberi-Nadjafi, S. Effati, and R. A. Van Gorder. Hybrid projective synchronization and control of the Baier–Sahle hyperchaotic flow in arbitrary dimensions with unknown parameters. Applied Mathematics and Computation, 248(??):55–69, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012272.

Nunez:2013:MTF

[NSV13] Juan Núñez, Marithania Silvero, and María Trinidad Villar. Mathematical tools for the future: Graph Theory and graphicable algebras. Applied Mathematics and Computation, 219(11):6113–6125, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012623.

Nie:2014:HOC

[NSYA14] Cunyun Nie, Shi Shu, Haiyuan Yu, and Qianjiang An. A high order composite scheme for the second order elliptic problem with nonlocal boundary and its fast algorithm. Applied Mathematics and Computation, 227(??):212–221, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011235.

Ning:2014:NBS

[NSZ^{+}14] Zhen-Hu Ning, Chang-Xiang Shen, Xiaopeng Zhao, Hao Li, Changsong Lin, and Yanmei Zhang. Nonlinear boundary stabilization of the wave equations with variable coefficients and time dependent delay. Applied Mathematics and Computation, 232(??):511–520, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314001878.

Nieto:2012:UMA

[NTdCJL12] P. J. García Nieto, J. Martínez Torres, F. J. de Cos Juez, and F. Sánchez Lasheras. Using multivariate adaptive regression splines and multilayer perceptron networks to evaluate paper manufactured using Eucalyptus globulus. Applied Mathematics and Computation, 219(2):755–763, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006960.

Nguyen-Thoi:2014:SCN

[NTPVNHLX14] T. Nguyen-Thoi, P. Phung-Van, S. Nguyen-Hoang, and Q. Lieu-Xuan. A smoothed coupled NS/nES– FEM for dynamic analysis of 2D fluid-solid interaction problems. Applied Mathematics and Computation, 232(??):324–346, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000897.

Naresh:2011:NAE

[NTS11a] Ram Naresh, Agraj Tripathi, and Dileep Sharma. A nonlinear AIDS epidemic model with screening and time delay. Applied Mathematics and Computation, 217(9):4416– 4426, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310010805.

Naresh:2011:NHA

[NTS11b] Ram Naresh, Agraj Tripathi, and Dileep Sharma. A nonlinear HIV/AIDS model with contact tracing. Applied Mathematics and Computation, 217(23):9575–9591, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100590X.

Nam:2010:TMT

[NTT10] Phan Thanh Nam, Dang Duc Trong, and Nguyen Huy Tuan. The truncation method for a twodimensional nonhomogeneous backward heat problem. Applied Mathematics and Computation, 216(12):3423–3432, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002808.

Nakic:2013:ODV

[NTT13] Ivica Nakić, Zoran Tomljanović, and Ninoslav Truhar. Optimal Direct Velocity Feedback. Applied Mathematics and Computation, 225(??): 590–600, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010576.

NGuerekata:2012:ASS

[NV12] Gaston M. N’Guérékata and Vincent Valmorin. Antiperiodic solutions of semilinear integrodifferential equations in Banach spaces. Applied Mathematics and Computation, 218(22):11118–11124, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004845.

Ni:2010:STA

[NW10] Tie Ni and Ping Wang. A smoothing-type algorithm for solving nonlinear complementarity problems with a non-monotone line search. Applied Mathematics and Computation, 216(7):2207–2214, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003127.

Niu:2012:AAS

[NW12] Wei Niu and Dongming Wang. Algebraic analysis of stability and bifurcation of a self-assembling micelle system. Applied Mathematics and Computation, 219(1):108–121, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005504.

Nian:2010:MPS

[NWNL10] Fuzhong Nian, Xingyuan Wang, Yujun Niu, and Da Lin. Module-phase synchronization in complex dynamic system. Applied Mathematics and Computation, 217(6):2481– 2489, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310008040.

Nie:2013:EMN

[NWZ13] Jianjun Nie, Xian Wu, and Minfang Zhu. Existence and multiplicity of non-trivial solutions for a class of modified Schrödinger equation with non-coercive potential. Applied Mathematics and Computation, 225(??):677–694, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010539.

Niu:2012:HLB

[NY12] Datian Niu and Xuegang Yuan. A harmonic Lanczos bidiagonalization method for computing interior singular triplets of large matrices. Applied Mathematics and Computation, 218(14):7459–7467, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200046X.

Niu:2012:RCC

[NZ12] Ben Niu and Jun Zhao. Robust H_{∞} control for a class of switched nonlinear cascade systems via multiple Lyapunov functions approach. Applied Mathematics and Computation, 218(11):6330–6339, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014585.

Nguyen:2014:HOF

[NZ14] Duc Duy Nguyen and Shan Zhao. High order FDTD methods for transverse magnetic modes with dispersive interfaces. Applied Mathematics and Computation, 226(??):699–707, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011594.

Ni:2011:ADT

[NZW11] Q. Ni, Z. L. Zhang, and L. Wang. Application of the differential transformation method to vibration analysis of pipes conveying fluid. Applied Mathematics and Computation, 217(16):7028–7038, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001627.

Ozis:2010:AEM

[ÖA10] Turgut Öziş and İsmail Aslan. Application of the (\frac{G′}{G})-expansion method to Kawahara type equations using symbolic computation. Applied Mathematics and Computation, 216(8):2360–2365, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003358.

Oteles:2013:PIP

[ÖA13] Ahmet Öteles and Mehmet Akbulak. Positive integer powers of certain complex tridiagonal matrices. Applied Mathematics and Computation, 219(21): 10448–10455, July 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004451.

Othman:2014:GEI

[OADL14] M. I. A. Othman, S. M. Abo-Dahab, and Kh. Lotfy. Gravitational effect and initial stress on generalized magnetothermo-microstretch elastic solid for different theories. Applied Mathematics and Computation, 230(??):597–615, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014380.

Orkcu:2011:GPA

[ÖB11] H. Hasan Örkcü and Hasan Bal. Goal programming approaches for data envelopment analysis cross efficiency evaluation. Applied Mathematics and Computation, 218(2):346–356, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100751X.

Ozbakir:2010:BAG

[ÖBT10] Lale Özbakir, Adil Baykasoḡlu, and Pınar Tapkan. Bees algorithm for generalized assignment problem. Applied Mathematics and Computation, 215(11):3782–3795, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010078.

Olgun:2010:NSM

[OC10] Murat Olgun and Cafer Coskun. Non-selfadjoint matrix Sturm–Liouville operators with spectral singularities. Applied Mathematics and Computation, 216(8):2271–2275, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003164.

Ocalan:2014:DDE

[Öca14]  Özkan Öcalan.  Dy-

namics of the difference equation x_{n+1} = p_{n} + x_{n−k}/x_{n} with a periodtwo coefficient. Applied Mathematics and Computation, 228(??):31–37, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011879.

Occorsio:2011:MEH

[Occ11] Donatella Occorsio. A method to evaluate the Hilbert transform on (0, +∞). Applied Mathematics and Computation, 217(12):5667–5679, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012476.

Osipov:2011:ELC

[OCN11] Nikolay N. Osipov, Ronald Cools, and Michael V. Noskov. Extremal lattices and the construction of lattice rules. Applied Mathematics and Computation, 217(9):4397–4407, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010696.

Orkcu:2011:WSA

[ÖD11a] Mediha Örkcü and Ogün Dogru. Weighted statistical approximation by Kantorovich type q-Szász– Mirakjan operators. Applied Mathematics and Computation, 217(20): 7913–7919, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311003481.

Ozbilge:2011:ASA

[OD11b] Ebru Ozbilge and Ali Demir. Analysis of a semigroup approach in the inverse problem of identifying an unknown parameters. Applied Mathematics and Computation, 218(3):965–969, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001159.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Odibat:2010:SCH

[Odi10]  Zaid M. Odibat.  A

study on the convergence of homotopy analysis method. Applied Mathematics and Computation, 217(2):782–789, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007022.

Ontanon-Garcia:2014:FHM

[OGJLCCB14] L. J. Ontañón-García, E. Jiménez-López, E. Campos-Cantón, and M. Basin. A family of hyperchaotic multi-scroll attractors in R^{n}. Applied Mathematics and Computation, 233(??):522–533, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314001799.

Ortiz-Gracia:2014:EVE

[OGO14] Luis Ortiz-Gracia and Cornelis W. Oosterlee. Efficient VaR and Expected Shortfall computations for nonlinear portfolios within the delta– gamma approach. Applied Mathematics and Computation, 244(??):16–31, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009540.

Okuyucu:2013:SHT

[OGYE13] O. Zeki Okuyucu, I. Gök, Y. Yayli, and N. Ekmekci. Slant helices in three dimensional Lie groups. Applied Mathematics and Computation, 221(??):672–683, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007583.

Omar:2012:NSF

[OH12] A. H. Alsonosi Omar and Y. Abu Hasan. Numerical solution of fuzzy differential equations and the dependency problem. Applied Mathematics and Computation, 219(3):1263–1272, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007345.

Ozer:2011:DKV

[ÖK11] Mehmet Naci Özer and Murat Koparan. Derivation of Korteweg–de Vries flow equations from the regularized long-wave (RLW) equation. Applied Mathematics and Computation, 218(5):2277–2280, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009921.

Oke:2011:RTE

[Öke11] Figen Öke. A residual transcendental extension of a valuation on K to K(x_{1}, …, x_{n}). Applied Mathematics and Computation, 218(3):956–958, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005145.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ozdemir:2012:DID

[ÖKU12] Halim Özdemir, Emre Kisi, and Mahmut Uç. On a disjoint idempotent decomposition for linear combinations produced from n commutative tripotent matrices. Applied Mathematics and Computation, 219 (4):1909–1918, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008260.

Olaleru:2011:ACF

[Ola11] J. O. Olaleru. Approximation of common fixed points of weakly compatible pairs using the Jungck iteration. Applied Mathematics and Computation, 217(21):8425–8431, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100405X.

Olszowy:2012:FPT

[Ols12] Leszek Olszowy. Fixed point theorems in the Fréchet space C(R_{+}) and functional integral equations on an unbounded interval. Applied Mathematics and Computation, 218(18):9066–9074, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002767.

Ofoedu:2011:HAS

[OM11] Eric U. Ofoedu and David M. Malonza. Hybrid approximation of solutions of nonlinear operator equations and application to equation of Hammerstein-type. Applied Mathematics and Computation, 217(13): 6019–6030, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310012750.

Oftadeh:2010:NCC

[ONBN10] R. Oftadeh, M. Nikkhah-Bahrami, and A. Najafi. A novel cubically convergent iterative method for computing complex roots of nonlinear equations. Applied Mathematics and Computation, 217(6):2608–2618, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008349.

Onitsuka:2011:UAS

[Oni11] Masakazu Onitsuka. Uniform asymptotic stability for damped linear oscillators with variable parameters. Applied Mathematics and Computation, 218(4):1436–1442, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008526.

Ostrovskii:2010:AFA

[OO10] Iossif Ostrovskii and Sofiya Ostrovska. On the analyticity of functions approximated by their q-Bernstein polynomials when q > 1. Applied Mathematics and Computation, 217(1):65–72, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004406.

ORegan:2011:EIH

[OO11] Donal O’Regan and Aleksandra Orpel. Eigenvalue intervals for higher order problems with singular nonlinearities. Applied Mathematics and Computation, 218(4):1233–1239, October 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311008022.

ORegan:2012:VAE

[OO12] Donal O’Regan and Aleksandra Orpel. A variational approach to the eigenvalue problem for higher order BVPs with singular nonlinearities. Applied Mathematics and Computation, 218(20): 10063–10071, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312003232.

Ostrovska:2013:BPC

[OÖ13] Sofiya Ostrovska and Ahmet Yasar Özban. The q-Bernstein polynomials of the Cauchy kernel with a pole on [0, 1] in the case q > 1. Applied Mathematics and Computation, 220(??):735–747, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313007844.

Ofoedu:2014:NIE

[OO14] Eric U. Ofoedu and Charles E. Onyi. New implicit and explicit approximation methods for solutions of integral equations of Hammerstein type. Applied Mathematics and Computation, 246(??):628–637, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011667.

Olatinwo:2012:SRJ

[OP12a] Memudu Olaposi Olatinwo and Mihai Postolache. Stability results for Jungck-type iterative processes in convex metric spaces. Applied Mathematics and Computation, 218(12):6727–6732, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015128.

Ozhan:2012:PPR

[ÖP12b] B. Burak Özhan and Mehmet Pakdemirli. Principal parametric resonances of a general continuous system with cubic nonlinearities. Applied Mathematics and Computation, 219(5):2412–2423, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312008417.

Okrasinski:2013:BFM

[OP13] Wojciech Okrasiński and Lukasz Plociniczak. Bessel function model of corneal topography. Applied Mathematics and Computation, 223(??):436–443, October 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313008588.

Obradovic:2014:UQA

[OP14a] Milutin Obradović and Saminathan Ponnusamy. Univalence of quotient of analytic functions. Applied Mathematics and Computation, 247(??):689–694, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012466.

Owolabi:2014:HOT

[OP14b] Kolade M. Owolabi and Kailash C. Patidar. Higherorder time-stepping methods for time-dependent reaction-diffusion equations arising in biology. Applied Mathematics and Computation, 240(??):30–50, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005955.

Ooi:2012:TDS

[OPD12] E. H. Ooi, V. Popov, and H. Dogan. Threedimensional solution for acoustic and transport problems using the radial basis integral equation method. Applied Mathematics and Computation, 218(18):9470–9488, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200269X.

ORegan:2010:EET

[O’R10] Donal O’Regan. Essential extension type maps for general classes of maps. Applied Mathematics and Computation, 216(11):3391–3395, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310005151.

Occorsio:2011:NMF

[OR11a] Donatella Occorsio and Maria Grazia Russo. Numerical methods for Fredholm integral equations on the square. Applied Mathematics and Computation, 218(5):2318–2333, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010010.

ORegan:2011:ALH

[O’R11b] Donal O’Regan. Applications of linking to Hammerstein integral equations. Applied Mathematics and Computation, 217(12):5279–5284, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031001180X.

Dordevic:2012:TMM

[– Dor12a] Snezana S. – Dordević. Two modifications of the method of the multiplicative parameters in descent gradient methods. Applied Mathematics and Computation, 218(17):8672–8683, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001518.

ORegan:2012:CTM

[O’R12b] Donal O’Regan. Coincidence theory for multimaps. Applied Mathematics and Computation, 219(4):2026–2034, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008375.

Oliveira:2013:GPP

[OR13a] Regilene D. S. Oliveira and Alex C. Rezende. Global phase portraits of a SIS model. Applied Mathematics and Computation, 219(9):4924–4930, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031201106X.

Doric:2013:NCC

[– Dor13b] Dragan – Dorić. Nonlinear coupled coincidence and coupled fixed point theorems for not necessary commutative contractive mappings in partially ordered probabilistic metric spaces. Applied Mathematics and Computation, 219(11):5926–5935, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312013239.

Ossandon:2014:CAE

[OR14] Sebastián Ossandón and Camilo Reyes. On the calculation of the acoustic eigenfrequencies for anisotropic materials with uncertainty. Applied Mathematics and Computation, 230(??):330–335, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313014070.

Orkcu:2013:SSM

[Örk13] H. Hasan Örkcü. Subset selection in multiple linear regression models: a hybrid of genetic and simulated annealing algorithms. Applied Mathematics and Computation, 219(23):11018–11028, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005353.

Ortigueira:2014:PSC

[Ort14a] Manuel D. Ortigueira. On the particular solution of constant coefficient fractional differential equations. Applied Mathematics and Computation, 245(??):255–260, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010273.

Ortigueira:2014:SAP

[Ort14b] Manuel D. Ortigueira.

A simple approach to the particular solution of constant coefficient ordinary differential equations. Applied Mathematics and Computation, 232(??):254–260, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000733.

Ofoedu:2011:CAF

[OS11] Eric U. Ofoedu and Y. Shehu. Convergence analysis for finite family of relatively quasi nonexpansive mappings and systems of equilibrium problems. Applied Mathematics and Computation, 217(22):9142–9150, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005273.

Ozden:2014:MUA

[OS14] Hacer Ozden and Yilmaz Simsek. Modification and unification of the Apostol-type numbers and polynomials and their applications. Applied Mathematics and Computation, 235(??):338–351, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003579.

Ordonez:2013:FDA

[OSAG13] C. Ordóñez, C. Sierra, T. Albuquerque, and J. R. Gallego. Functional data analysis as a tool to correlate textural and geochemical data. Applied Mathematics and Computation, 223(??):476–482, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009004.

Obregon:2014:HAN

[OSRFF14] M. A. Obregon, E. Sanmiguel-Rojas, and R. Fernandez-Feria. High accuracy numerical methods for the Gardner–Ostrovsky equation. Applied Mathematics and Computation, 240(??):140–148, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002355.

Ordonez:2012:VSR

[OSRPC12] Celestino Ordóñez, Marta Sestelo, Javier Roca-Pardiñas, and Enrique Covián. Variable selection in regression models used to analyse Global Positioning System accuracy in forest environments. Applied Mathematics and Computation, 219(4):2220–2230, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312008624.

Ossipov:2010:HAG

[Oss10]  P. Ossipov.  Heuris-

tic algorithm for generating multilevel Cartesian meshes inmultidimensional regions with moving boundaries. Applied Mathematics and Computation, 215(10):3684–3695, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009813.

Otachel:2014:SSI

[Ota14] Zdzislaw Otachel. Synchronous sequences and inequalities for convex functions. Applied Mathematics and Computation, 247(??):865–871, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012624.

Ozkoc:2011:SSP

[ÖTC11] Arzu Özkoç, Ahmet Tekcan, and Ismail Naci Cangül. Solving some parametric quadratic Diophantine equation over Z and F_{p}. Applied Mathematics and Computation, 218(3):703–706, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004395.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ou:2010:MRE

[Ou10] Jianping Ou. On maximal 3-restricted edge connectivity and reliability analysis of hypercube networks. Applied Mathematics and Computation, 217(6):2602–2607, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008325.

Oulne:2011:VSA

[Oul11]  M. Oulne.  Variation

and series approach to the Thomas–Fermi equation. Applied Mathematics and Computation, 218(2):303–307, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007454.

Ouyang:2013:CEO

[OYCCB13] Liang-Yuh Ouyang, Chih-Te Yang, Ya-Lan Chan, and Leopoldo Eduardo Cárdenas-Barrón. A comprehensive extension of the optimal replenishment decisions under two levels of trade credit policy depending on the order quantity. Applied Mathematics and Computation, 224(??):268–277, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009302.

Ozbekler:2010:PNS

[ÖZ10] A. Özbekler and A. Zafer. Principal and nonprincipal solutions of impulsive differential equations with applications. Applied Mathematics and Computation, 216(4):1158–1168, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001748.

Ozarslan:2011:SFG

[Öza11] M. Ali Özarslan. Some families of generating functions for the extended Srivastava polynomials. Applied Mathematics and Computation, 218(3):959–964, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000919.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ozarslan:2014:SIE

[Öza14] Mehmet Ali Özarslan. On a singular integral equation including a set of multivariate polynomials suggested by Laguerre polynomials. Applied Mathematics and Computation, 229(??):350–358, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013337.

Ozden:2011:ADU

[Ozd11] Hacer Ozden. p-Adic distribution of the unification of the Bernoulli, Euler and Genocchi polynomials. Applied Mathematics and Computation, 218(3):970–973, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100110X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ozkan:2014:TLS

[Özk14]  Engin Özkan.  Trun-

cated Lucas sequence and its period. Applied Mathematics and Computation, 232(??):285–291, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000514.

Prajapati:2014:SRD

[PAA14] Jyotindra C. Prajapati, Naresh K. Ajudia, and Praveen Agarwal. Some results due to Konhauser polynomials of first kind and Laguerre polynomials. Applied Mathematics and Computation, 247(??):639–650, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012387.

Perez:2013:NSC

[PAC13] Oscar Pérez, Iván Amaya, and Rodrigo Correa. Numerical solution of certain exponential and nonlinear Diophantine systems of equations by using a discrete particle swarm optimization algorithm. Applied Mathematics and Computation, 225(??):737–746, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301062X.

Pai:2014:GSU

[Pai14] Ming-Chang Pai. Global synchronization of uncertain chaotic systems via discrete-time sliding mode control. Applied Mathematics and Computation, 227(??):663–671, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012447.

Pakdaman:2011:ENB

[Pak11a] Morteza Pakdaman. Erratum to “A new branch and bound method with pretreatment for the binary quadratic programming” [Appl. Math. Comput. 192 (2007) 252–259]. Applied Mathematics and Computation, 217(13):6441, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000312.

See [MZL07].

Pakdaman:2011:NNL

[Pak11b] Morteza Pakdaman. A note on “A new local and global optimization method for mixed integer quadratic programming problems” by G. Q. Li et al. Applied Mathematics and Computation, 217(12):5998, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012592.

See [LWQ10].

Pal:2010:MCH

[Pal10] Dulal Pal. Mixed convection heat transfer in the boundary layers on an exponentially stretching surface with magnetic field. Applied Mathematics and Computation, 217(6):2356–2369, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007800.

Paleo:2012:PPT

[Pal12] Bruno Woltzenlogel Paleo. Physics and proof theory. Applied Mathematics and Computation, 219(1):45–53, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009039.

Paladi:2013:PAH

[Pal13]  Florentin Paladi.  On

the probabilistic approach to heterogeneous structure interactions in agentbased computational models. Applied Mathematics and Computation, 219(24):11430–11437, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005894.

Park:2014:BDM

[PAL14] Changsoon Park, Suneung Ahn, and Sangwon Lee. A Bayesian decision model based on expected utility and uncertainty risk. Applied Mathematics and Computation, 242(??):643–648, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008418.

Pang:2010:SBO

[Pan10] QingLe Pang. Study on the behavior of oscillating solitons using the (2 + 1)-dimensional nonlinear system. Applied Mathematics and Computation, 217(5):2015–2023, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007472.

Pan:2011:BLC

[Pan11] Xuejun Pan. Bifurcation of limit cycles for a class of cubic polynomial system having a nilpotent singular point. Applied Mathematics and Computation, 218(4):1161–1165, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008241.

Pang:2012:CLC

[Pan12a] QingLe Pang. Construction of localized coherent structures based on the (3 + 1)-dimensional Nizhnik–Novikov–Veselov equation. Applied Mathematics and Computation, 218(22):11333–11336, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005516.

Pantokratoras:2012:CCF

[Pan12b] Asterios Pantokratoras.

Comment on “Chebyshev finite difference method for MHD flow of a micropolar fluid past a stretching sheet with heat transfer”, by N.T. Eldabe, E.F. Elshehawey, Elsayed M.E. Elbarbary and Nasser S. Elgazery [Applied Mathematics and Computation, 160, 2005, 437– 450]. Applied Mathematics and Computation, 218(9):5827–5828, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013555.

See [EEEE05].

Pantokratoras:2015:CPE

[Pan15] Asterios Pantokratoras.

Comment on the paper “Effect of viscous dissipation and heat source on unsteady boundary layer flow and heat transfer past a stretching surface embedded in a porous medium using element free Galerkin method, by R. Sharma, Applied Mathematics and Computation (2012) 219:976– 987”. Applied Mathematics and Computation, 266(??):70–71, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031500661X.

See [Sha12b].

Parkes:2010:NNT

[Par10a] E. J. Parkes. A note on “New travelling wave solutions to the Ostrovsky equation”. Applied Mathematics and Computation, 217(7):3575–3577, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009082.

See [Yaş10].

Parkes:2010:OEM

[Par10b] E. J. Parkes. Observations on the tanh– coth expansion method for finding solutions to nonlinear evolution equations. Applied Mathematics and Computation, 217(4):1749–1754, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010261.

Parkes:2010:OBE

[Par10c] E. J. Parkes. Observations on the basic (\frac{G′}{G})- expansion method for finding solutions to nonlinear evolution equations. Applied Mathematics and Computation, 217(4):1759–1763, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003279.

Parhi:2011:ONO

[Par11]  N. Parhi.  Oscillation

and non-oscillation of solutions of second order difference equations involving generalized difference. Applied Mathematics and Computation, 218(2):458–468, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007673.

Park:2012:SBS

[Par12] Jungho Park. Structure of bifurcated solutions of two-dimensional infinite Prandtl number convection with no-slip boundary conditions. Applied Mathematics and Computation, 218(21):10586– 10601, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004201.

Parhi:2013:NOC

[Par13] N. Parhi. Non-oscillation of a class of third order nonhomogeneous difference equations. Applied Mathematics and Computation, 219(15):8384–8397, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000532.

Patterson:2013:FDM

[Pat13a] Richard F. Patterson. Four dimensional matrix characterization P - convergence fields of summability methods. Applied Mathematics and Computation, 219(12):6783–6791, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313000106.

Patterson:2013:TEF

[Pat13b] Richard F. Patterson. A theorem on entire four dimensional summability methods. Applied Mathematics and Computation, 219(14):7777–7782, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001185.

Parmar:2013:WSA

[PB13a] Kulwinder Singh Parmar and Rashmi Bhardwaj. Wavelet and statistical analysis of river water quality parameters. Applied Mathematics and Computation, 219 (20):10172–10182, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003706.

Paul:2013:EBZ

[PB13b] Kallol Paul and Santanu Bag. Estimation of bounds for the zeros of a polynomial using numerical radius. Applied Mathematics and Computation, 222(??):231–243, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007960.

Ponsoda:2014:EIC

[PB14a] Enrique Ponsoda and Sergio Blanes. Exponential integrators for coupled self-adjoint nonautonomous partial differential systems. Applied Mathematics and Computation, 243(??):1– 11, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314007152.

Poullet:2014:EBI

[PB14b] Pascal Poullet and Amir Boag. Equation-based interpolation and incremental unknowns for solving the three-dimensional Helmholtz equation. Applied Mathematics and Computation, 232(??): 1200–1208, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300314001210.

Priyadarshi:2014:GPC

[PBG14] A. Priyadarshi, Sandip Banerjee, and S. Gakkhar. Geometry of the Poincaré compactification of a fourdimensional food-web system. Applied Mathematics and Computation, 226(??):229–237, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011247.

Pollack:2014:AAP

[PBP14] Steven Pollack, Daniel S.

Badali, and Jonathan Pollack. An algorithm for autonomously plotting solution sets in the presence of turning points. Applied Mathematics and Computation, 232(??):375–380, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000162.

Pavaloiu:2011:BAS

[PC11] I. Păvăloiu and E. Cătinaş.

Bilateral approximations for some Aitken–Steffensen– Hermite type methods of order three. Applied Mathematics and Computation, 217(12):5838–5846, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012695.

Pal:2013:SDE

[PC13a] Dulal Pal and Sewli Chatterjee. Soret and Dufour effects on MHD convective heat and mass transfer of a power-law fluid over an inclined plate with variable thermal conductivity in a porous medium. Applied Mathematics and Computation, 219(14):7556–7574, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000829.

Pei:2013:EUS

[PC13b] Minghe Pei and Sung Kag Chang. Existence and uniqueness of solutions for n th-order nonlinear two-point boundary value problems. Applied Mathematics and Computation, 219(23):11005–11017, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313005262.

Pinto:2014:NFD

[PC14a] Carla M. A. Pinto and Ana R. M. Carvalho. New findings on the dynamics of HIV and TB coinfection models. Applied Mathematics and Computation, 242(??):36–46, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007267.

Proinov:2014:SCC

[PC14b] Petko D. Proinov and Slav I. Cholakov. Semilocal convergence of Chebyshevlike root-finding method for simultaneous approximation of polynomial zeros. Applied Mathematics and Computation, 236(??):669–682, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004548.

Popa:2014:NPI

[PCM14] Ioan-Lucian Popa, Traian Ceausu, and Mihail Megan. Nonuniform power instability and Lyapunov sequences. Applied Mathematics and Computation, 247(??):969–975, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012764.

Pelinovsky:2010:SWT

[PCT^{+}10] Efim Pelinovsky, Byung Ho Choi, Tatiana Talipova, Seung Buhm Woo, and Dong Chule Kim. Solitary wave transformation on the underwater step: Asymptotic theory and numerical experiments. Applied Mathematics and Computation, 217(4):1704–1718, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009345.

Pathak:2012:EUS

[PD12] H. K. Pathak and Deepmala. Existence and uniqueness of solutions of functional equations arising in dynamic programming. Applied Mathematics and Computation, 218(13):7221–7230, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000057.

Ponomareva:2013:HOS

[PD13] Ksenia Ponomareva and Paresh Date. Higher order sigma point filter: a new heuristic for nonlinear time series filtering. Applied Mathematics and Computation, 221(??):662–671, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007339.

Panda:2014:MCL

[PD14] Arpita Panda and Chandan Bikash Das. Multichoice linear programming for matrix game. Applied Mathematics and Computation, 237(??):411–418, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003701.

Perez-Demydenko:2014:CSH

[PDBRFER14] C. Pérez-Demydenko, I. Brito-Reyes, B. Aragón Fernández, and E. Estevez-Rams. The complete set of homogeneous Hilbert curves in two dimensions. Applied Mathematics and Computation, 234(??):531–542, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001842.

Petkovic:2011:IMM

[PDN11] Miodrag S. Petković, Jovana Dzunić, and Beny Neta. Interpolatory multipoint methods with memory for solving nonlinear equations. Applied Mathematics and Computation, 218(6):2533–2541, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010174.

Pardo:2013:RFR

[PDPGOR13] Carlos Pardo, José F. Diez-Pastor, César García-Osorio, and Juan J. Rodríguez. Rotation forests for regression. Applied Mathematics and Computation, 219(19): 9914–9924, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004128.

Pakdemirli:2010:PTE

[PE10] M. Pakdemirli and N. Elmas. Perturbation theorems for estimating magnitudes of roots of polynomials. Applied Mathematics and Computation, 216(5):1645–1651, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002626.

Pedram:2013:AOO

[Ped13]  Pouria Pedram.  An-

harmonic oscillator and the optimized basis expansion. Applied Mathematics and Computation, 219(9):4655–4666, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010946.

Pourrajabian:2013:AGA

[PEMS13] Abolfazl Pourrajabian, Reza Ebrahimi, Masoud Mirzaei, and Mehrzad Shams. Applying genetic algorithms for solving nonlinear algebraic equations. Applied Mathematics and Computation, 219(24):11483–11494, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313006048.

Pena:2013:ELP

[Peñ13] J. M. Peña. Eigenvalue localization and pivoting strategies for Gaussian elimination. Applied Mathematics and Computation, 219(14):7725–7729, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001100.

Pena:2014:ELN

[Peñ14] J. M. Peña. Eigenvalue localization and Neville elimination. Applied Mathematics and Computation, 242(??):340–345, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007280.

Papaschinopoulos:2014:ABP

[PEP14] G. Papaschinopoulos, G. Ellina, and K. B. Papadopoulos. Asymptotic behavior of the positive solutions of an exponential type system of difference equations. Applied Mathematics and Computation, 245(??):181–190, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010315.

Petronilho:2012:GFS

[Pet12] J. Petronilho. Generalized Fibonacci sequences via orthogonal polynomials. Applied Mathematics and Computation, 218(19):9819–9824, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002925.

Petkovic:2013:NPO

[Pet13] M. S. Petković. A note on the priority of optimal multipoint methods for solving nonlinear equations. Applied Mathematics and Computation, 219(10):5249–5252, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012313.

Petkovic:2014:MSR

[Pet14] Miodrag S. Petković. Multi-step root solvers of Traub’s type in real interval arithmetic. Applied Mathematics and Computation, 248(??):430–440, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013599.

Pfeiffer:2010:IF

[Pfe10] Friedrich Pfeiffer. On impacts with friction. Applied Mathematics and Computation, 217(3):1184– 1192, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006065.

Pittavino:2014:CEM

[PFG^{+}14] Marta Pittavino, Luca Ferreri, Mario Giacobini, Luigi Bertolotti, Sergio Rosati, and Ezio Venturino. A caev epidemiological model for goat breeding. Applied Mathematics and Computation, 227(??):156–163, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300313011983.

Plaza:2010:NDP

[PFS10] Ángel Plaza, Sergio Falcón, and José P. Suárez. On the non-degeneracy property of the longest-edge trisection of triangles. Applied Mathematics and Computation, 216(3):862–869, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001177.

Parhi:2011:SLM

[PG11] S. K. Parhi and D. K.

Gupta. A Stirling-like method with Hölder continuous first derivative in Banach spaces. Applied Mathematics and Computation, 217(23):9567–9574, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005893.

Panda:2013:DFM

[PG13] Geetanjali Panda and Suman Ghosh. A derivative free multidimensional optimal search method using Lucas number. Applied Mathematics and Computation, 219(12):6536–6541, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010922.

Pathak:2010:CFP

[PH10a] H. K. Pathak and N. Hussain. Common fixed points for P-operator pair with applications. Applied Mathematics and Computation, 217(7):3137–3143, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009100.

Patricio:2010:GIP

[PH10b] P. Patrício and R. E. Hartwig. The (2, 2, 0) group inverse problem. Applied Mathematics and Computation, 217(2):516– 520, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006533.

P:2014:RNE

[PHB14] S. Ivvan Valdez P., Arturo Hernández, and Salvador Botello. Repairing normal EDAs with selective repopulation. Applied Mathematics and Computation, 230(??):65–77, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013714.

Parra-Hinojosa:2013:FIE

[PHGV13] Alfredo Parra-Hinojosa and Julio C. Gutiérrez-Vega. Fractional Ince equation with a Riemann– Liouville fractional derivative. Applied Mathematics and Computation, 219(22):10695–10705, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004797.

Pisarev:2011:LES

[PHPD11] Gleb I. Pisarev, Alex C.

Hoffmann, Weiming Peng, and Henk A. Dijkstra. Large Eddy Simulation of the vortex end in reverseflow centrifugal separators. Applied Mathematics and Computation, 217(11):5016–5022, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310007952.

Pickering:2012:SST

[Pic12]  Andrew Pickering.  A

special solution technique: Further extensions. Applied Mathematics and Computation, 218(22):11177–11181, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004924.

Petkovic:2010:DFT

[PID10] M. S. Petković, S. Ilić, and J. Džunić. Derivative free two-point methods with and without memory for solving nonlinear equations. Applied Mathematics and Computation, 217(5):1887–1895, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007289.

Ping:2010:NES

[Pin10] Zhang Ping. New exact solutions to breaking soliton equations and Whitham– Broer–Kaup equations. Applied Mathematics and Computation, 217(4):1688– 1696, October 15, 2010.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309008911.

Plesser:2010:RSI

[PJ10] Hans Ekkehard Plesser and Anders Grønvik Jahnsen. Re-seeding invalidates tests of random number generators. Applied Mathematics and Computation, 217(1):339–346, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006259.

See [KCK08].

Pavlovic:2012:MES

[PJ12] Gorica Pavlovíc and Svetlana Jankovíc. Moment exponential stability and integrability of stochastic functional differential equations. Applied Mathematics and Computation, 218(10):6125–6134, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014445.

Peng:2014:VSF

[PJFL14] Hong-Yi Peng, Chun-Fu Jiang, Xiang Fang, and Jin-Shan Liu. Variable selection for Fisher linear discriminant analysis using the modified sequential backward selection algorithm for the microarray data. Applied Mathematics and Computation, 238(??):132–140, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005128.

Phuangphoo:2012:ATS

[PK12] Pongrus Phuangphoo and Poom Kumam. Approximation theorems for solving common solution of a system of mixed equilibrium problems and variational inequality problems and fixed point problems for asymptotically strict pseudocontractions in the intermediate sense. Applied Mathematics and Computation, 219(3):837–855, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200656X.

Pantokratoras:2018:CPB

[PK18] Asterios Pantokratoras and Muhammad Kamran. Comment on the paper “Boundary layer flow of Carreau fluid over a convectively heated stretching sheet, T. Hayat, Sadia Asad, M. Mustafa, A. Alsaedi, Applied Mathematics and Computation 246 (2014) 12–22”. Applied Mathematics and Computation, 329(??):227–229, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300318301206.

See [HAMA14].

Pandey:2012:SLE

[PKBD12] Rajesh K. Pandey, Narayan Kumar, Abhinav Bhardwaj, and Goutam Dutta. Solution of Lane–Emden type equations using Legendre operational matrix of differentiation. Applied Mathematics and Computation, 218(14):7629–7637, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000689.

Pachauri:2014:SRG

[PKD14]  Bhoopendra  Pachauri,

Ajay Kumar, and Joydip Dhar. Software reliability growth modeling with dynamic faults and release time optimization using GA and MAUT. Applied Mathematics and Computation, 242(??):500–509, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008078.

Papakostas:2010:CSO

[PKK10] G. A. Papakostas, D. E. Koulouriotis, and E. G. Karakasis. Computation strategies of orthogonal image moments: a comparative study. Applied Mathematics and Computation, 216(1):1–17, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000640.

Park:2012:SCC

[PKP^{+}12a] M. J. Park, O. M. Kwon, Ju H. Park, S. M. Lee, and E. J. Cha. Synchronization criteria for coupled neural networks with interval time-varying delays and leakage delay. Applied Mathematics and Computation, 218 (12):6762–6775, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015177.

Park:2012:SCF

[PKP^{+}12b] M. J. Park, O. M. Kwon, Ju H. Park, S. M. Lee, and E. J. Cha. Synchronization criteria of fuzzy complex dynamical networks with interval time-varying delays. Applied Mathematics and Computation, 218(23):11634–11647, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005619.

Park:2011:NAL

[PKPL11] M. J. Park, O. M.

Kwon, Ju H. Park, and S. M. Lee. A new augmented Lyapunov– Krasovskii functional approach for stability of linear systems with timevarying delays. Applied Mathematics and Computation, 217(17):7197–7209, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311001640.

Pylak:2010:CTS

[PKS10] Dorota Pylak, Paweł Karczmarek, and Michail A. Sheshko. Cauchy-type singular integral equation with constant coefficients on the real line. Applied Mathematics and Computation, 217(7):2977–2988, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008854.

Peng:2010:ADI

[PL10] Xiao-Fei Peng and Wen Li. The alternatingdirection iterative method for saddle point problems. Applied Mathematics and Computation, 216(6):1845–1858, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010674.

Pei:2013:MMT

[PLF13] Yongzhen Pei, Changguo Li, and Shunhou Fan. A mathematical model of a three species prey-predator system with impulsive control and Holling functional response. Applied Mathematics and Computation, 219(23):10945–10955, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005316.

Luo:2010:MOB

[pLjH10] Xue ping Luo and Nan jing Huang. (H, ϕ)-η-monotone operators in Banach spaces with an application to variational inclusions. Applied Mathematics and Computation, 216(4):1131–1139, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001712.

Plociniczak:2014:EAF

[Plo14] Lukasz Plociniczak. Eigenvalue asymptotics for a fractional boundary-value problem. Applied Mathematics and Computation, 241(??):125–128, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314006948.

Peng:2012:GNC

[PLTH12] Shao-Lun Peng, Cheng-Kuan Lin, Jimmy J. M. Tan, and Lih-Hsing Hsu. The g-good-neighbor conditional diagnosability of hypercube under PMC model. Applied Mathematics and Computation, 218 (21):10406–10412, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312003554.

Panja:2011:NOP

[PM11a] M. M. Panja and B. N. Mandal. A note on onepoint quadrature formula for Daubechies scale function with partial support. Applied Mathematics and Computation, 218(8):4147–4151, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012173.

Peng:2011:NBT

[PM11b] Chun-Cheng Peng and George D. Magoulas. Nonmonotone BFGS-trained recurrent neural networks for temporal sequence processing. Applied Mathematics and Computation, 217(12):5421–5441, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012129.

Pant:2012:RMC

[PM12] Neeraj Pant and S. K.

Maurya. Relativistic modeling of charged superdense star with Einstein– Maxwell equations in general relativity. Applied Mathematics and Computation, 218(17):8260–8268, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000823.

Pal:2014:BMT

[PM14a] D. Pal and G. S. Mahapatra. A bioeconomic modeling of twoprey and one-predator fishery model with optimal harvesting policy through hybridization approach. Applied Mathematics and Computation, 242(??):748–763, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008546.

Picart:2014:OCM

[PM14b] Delphine Picart and Fabio Milner. Optimal control in a multistage physiologically structured insect population model. Applied Mathematics and Computation, 247(??):573–588, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012326.

Palit:2012:NTN

[PMB12] Sanjay Kumar Palit, Sayan Mukherjee, and D. K. Bhattacharya. New types of nonlinear autocorrelations of bivariate data and their applications. Applied Mathematics and Computation, 218(17):8951–8967, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001956.

Peris:2011:CPN

[PMC11] R. Peris, A. Marquina, and V. Candela. The convergence of the perturbed Newton method and its application for ill-conditioned problems. Applied Mathematics and Computation, 218(7):2988– 3001, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311010459.

Prakash:2014:UMN

[PMD14] D. Prakash, M. Muthtamilselvan, and Deog-Hee Doh. Unsteady MHD non-Darcian flow over a vertical stretching plate embedded in a porous medium with non-uniform heat generation. Applied Mathematics and Computation, 236(??):480–492, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004342.

Petkovic:2011:EMI

[PMM11] Miodrag S. Petković, Mimica R. Milosević, and Dusan M. Milosević. Efficient methods for the inclusion of polynomial zeros. Applied Mathematics and Computation, 217(19):7636–7652, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002694.

Peng:2011:SDG

[PMP11] Jing Peng, Zehua Miao, and Fang Peng. Study on a 3-dimensional game model with delayed bounded rationality. Applied Mathematics and Computation, 218(5):1568–1576, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008654.

Pal:2014:FHT

[PMV14] Dulal Pal, Gopinath Mandal, and K. Vajravelu. Flow and heat transfer of nanofluids at a stagnation point flow over a stretching/shrinking surface in a porous medium with thermal radiation. Applied Mathematics and Computation, 238(??):208–224, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005165.

Padhan:2011:HOS

[PN11a] S. K. Padhan and C. Nahak. Higher-order symmetric duality in multiobjective programming problems under higher-order invexity. Applied Mathematics and Computation, 218(5):1705–1712, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008824.

Panigrahi:2011:WGM

[PN11b] Bijaya Laxmi Panigrahi and Gnaneshwar Nelakanti. Wavelet Galerkin method for eigenvalue problem of a compact integral operator. Applied Mathematics and Computation, 218(4):1222–1232, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007971.

Petkovic:2014:MMS

[PNPD14] Miodrag S. Petković, Beny Neta, Ljiljana D. Petković, and Jovana Dzunić. Multipoint methods for solving nonlinear equations: a survey. Applied Mathematics and Computation, 226(??):635–660, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011296.

Polat:2011:LE

[Pol11] Necat Polat. Letter to the Editor. Applied Mathematics and Computation, 217(12):6018, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011380.

Pierri:2013:ESS

[POR13] Michelle Pierri, Donal O’Regan, and Vanessa Rolnik. Existence of solutions for semi-linear abstract differential equations with not instantaneous impulses. Applied Mathematics and Computation, 219(12):6743–6749, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000258.

Papathanasiou:2010:CNP

[PP10] Nikolaos Papathanasiou and Panayiotis Psarrakos. On condition numbers of polynomial eigenvalue problems. Applied Mathematics and Computation, 216(4):1194–1205, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001773.

Postelnicu:2011:FSB

[PP11] A. Postelnicu and I. Pop.

Falkner–Skan boundary layer flow of a powerlaw fluid past a stretching wedge. Applied Mathematics and Computation, 217(9):4359–4368, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009975.

Pascu:2012:SAA

[PP12a] Mihai N. Pascu and Nicolae R. Pascu. Starlike approximations of analytic functions. Applied Mathematics and Computation, 218(12):6825–6832, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015396.

Perdomo:2012:NPD

[PP12b] Francisco Perdomo and Ángel Plaza. A new proof of the degeneracy property of the longestedge n-section refinement scheme for triangular meshes. Applied Mathematics and Computation, 219(4):2342–2344, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008223.

See corrigendum [PP15].

Perdomo:2013:PND

[PP13a] Francisco Perdomo and Ángel Plaza. Proving the non-degeneracy of the longest-edge trisection by a space of triangular shapes with hyperbolic metric. Applied Mathematics and Computation, 221(??):424–432, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007248.

Perrussel:2013:AES

[PP13b] Ronan Perrussel and Clair Poignard. Asymptotic expansion of steady-state potential in a high contrast medium with a thin resistive layer. Applied Mathematics and Computation, 221(??):48–65, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006863.

Padhi:2014:SSE

[PP14a] Seshadev Padhi and Smita Pati. Stability of solutions of an equation governing dynamics of a renewable resource subject to additive Allee effect. Applied Mathematics and Computation, 247(??):631–638, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012405.

Pascu:2014:CAA

[PP14b] Mihai N. Pascu and Nicolae R. Pascu. Convex approximations of analytic functions. Applied Mathematics and Computation, 232(??):559–567, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400126X.

Pomparau:2014:SPA

[PP14c] I. Pomparau and C. Popa.

Supplementary projections for the acceleration of Kaczmarz algorithm. Applied Mathematics and Computation, 232(??):104–116, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001350.

Perdomo:2015:CNP

[PP15] Francisco Perdomo and Ángel Plaza. Corrigendum to “A new proof of the degeneracy property of the longest-edge n-section refinement scheme for triangular meshes” [Applied Mathematics and Computation 219 (4) (2012) 2342–2344]. Applied Mathematics and Computation, 260(??):412–413, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300315003975.

See [PP12b].

Papanicolaou:2013:NME

[PPC13] N. C. Papanicolaou, A. C. Polycarpou, and M. A. Christou. Numerical modeling of electromagnetic wave propagation in a liquid crystal cell at oblique incidence. Applied Mathematics and Computation, 219(22):10643–10654, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003986.

Petkovic:2010:CTP

[PPD10] Ljiljana D. Petković, Miodrag S. Petković, and Jovana Džunić. A class of three-point root-solvers of optimal order of convergence. Applied Mathematics and Computation, 216(2):671–676, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001475.

Pathak:2011:DST

[PPK11] R. S. Pathak, Akhilesh Prasad, and Manish Kumar. n-Dimensional Sobolev type spaces involving Hankel transformation. Applied Mathematics and Computation, 218(3):899–905, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/

                                        pii/S0096300311003730.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Perc:2014:IDI

[PPK14] Matjaz Perc, Otilija Petek, and Salahuddin M. Kamal. Impact of density and interconnectedness of influential players on social welfare. Applied Mathematics and Computation, 249(??):19–23, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314014106.

Petkovic:2012:SCV

[PPL12] Marko D. Petković, Dragoljub Pokrajac, and Longin Jan Latecki. Spherical coverage verification. Applied Mathematics and Computation, 218(19):9699–9715, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002469.

Perez-Polo:2014:SCB

[PPMCG14] Manuel F. Pérez-Polo, Manuel Pérez Molina, Javier Gil Chica, and José A. Berna Galiano. Stability and chaotic behavior of a PID controlled inverted pendulum subjected to harmonic base excitations by using the normal form theory. Applied Mathematics and Computation, 232(??):698–718, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001398.

Pecaric:2013:MWS

[PPS13] Josip Pecarić, Anamarija Perusić, and Ksenija Smoljak. Mercer and Wu– Srivastava generalisations of Steffensen’s inequality. Applied Mathematics and Computation, 219(21): 10548–10558, July 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004438.

Pocheketa:2014:GCE

[PPV14] Oleksandr A. Pocheketa, Roman O. Popovych, and Olena O. Vaneeva. Group classification and exact solutions of variablecoefficient generalized Burgers equations with linear damping. Applied Mathematics and Computation, 243(??):232–244, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007991.

Ponnusamy:2014:HSS

[PQR14] Saminathan Ponnusamy, Tri Quach, and Antti Rasila. Harmonic shears of slit and polygonal mappings. Applied Mathematics and Computation, 233(??):588–598, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001131.

Palacios-Quinonero:2010:SDM

[PQRR10] Francisco Palacios-Quiñonero, José Rodellar, and Josep M. Rossell. Sequential design of multi-overlapping controllers for longitudinal multi-overlapping systems. Applied Mathematics and Computation, 217(3):1170–1183, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001542.

Ponnusamy:2010:SDI

[PQW10] S. Ponnusamy, J. Qiao, and X. Wang. Schwarzian derivative, integral means, and the affine and linear invariant families of biharmonic mappings. Applied Mathematics and Computation, 216(12):3468–3479, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004510.

Pinto:2010:ESA

[PR10] Manuel Pinto and Gonzalo Robledo. Existence and stability of almost periodic solutions in impulsive neural network models. Applied Mathematics and Computation, 217(8):4167–4177, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010714.

Pereira:2011:DMR

[PR11] Edgar Pereira and Cecilia Rosa. A deflation method for regular matrix pencils. Applied Mathematics and Computation, 218(6):2913–2920, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010605.

Parand:2012:NSN

[PR12a] K. Parand and J. A. Rad. Numerical solution of nonlinear Volterra– Fredholm–Hammerstein integral equations via collocation method based on radial basis functions. Applied Mathematics and Computation, 218(9):5292–5309, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013439.

Popa:2012:HUS

[PR12b] Dorian Popa and Ioan Rasa. Hyers–Ulam stability of the linear differential operator with nonconstant coefficients. Applied Mathematics and Computation, 219(4):1562–1568, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007862.

Pascoal:2014:MRR

[PR14] Marta M. B. Pascoal and Marisa Resende. The minmax regret robust shortest path problem in a finite multi-scenario model. Applied Mathematics and Computation, 241(??):88–111, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400616X.

Prentice:2011:TRE

[Pre11a] J. S. C. Prentice. Truncation and roundoff errors in three-point approximations of first and second derivatives. Applied Mathematics and Computation, 217(9):4576–4581, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011203.

Prevost:2011:SIZ

[Pré11b] Marc Prévost. Series involving zeta and related functions. Applied Mathematics and Computation, 217(12):5307–5317, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011847.

Pericaro:2013:GCG

[PRK13] Gislaine A. Periçaro, Ademir A. Ribeiro, and Elizabeth W. Karas. Global convergence of a general filter algorithm based on an efficiency condition of the step. Applied Mathematics and Computation, 219(17):9581–9597, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313002658.

Papaschinopoulos:2012:SAB

[PRS12] G. Papaschinopoulos, M. Radin, and C. J. Schinas. Study of the asymptotic behavior of the solutions of three systems of difference equations of exponential form. Applied Mathematics and Computation, 218(9):5310–5318, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013440.

Panayotounakos:2011:NMC

[PRT11] Dimitrios E. Panayotounakos, Ioanna Rizou, and Efstathios E. Theotokoglou. A new mathematical construction of the general nonlinear ODEs of motion in rigid body dynamics (Euler’s equations). Applied Mathematics and Computation, 217(21):8534–8542, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004255.

Ponce:2013:ACP

[PRV13] Enrique Ponce, Javier Ros, and Elísabet Vela. Algebraically computable piecewise linear nodal oscillators. Applied Mathematics and Computation, 219(9):4194–4207, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011630.

Prasath:2010:HCV

[PS10] V. B. Surya Prasath and Arindama Singh. A hybrid convex variational model for image restoration. Applied Mathematics and Computation, 215(10):3655–3664, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300309009771.

Plaza:2011:NSI

[PS11] Ángel Plaza and Kishin Sadarangani. A note on “Some inequalities in inner product spaces related to the generalized triangle inequality” by S. S. Dragomir et al. Applied Mathematics and Computation, 217(22):9497–9498, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005765.

See [DCK11].

Pejcev:2012:RTG

[PS12a] Aleksandar V. Pejcev and Miodrag M. Spalević. On the remainder term of Gauss–Radau quadrature with Chebyshev weight of the third kind for analytic functions. Applied Mathematics and Computation, 219(5):2760–2765, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009010.

Peterson:2012:SSS

[PS12b] Ellen R. Peterson and Michael Shearer. Simulation of spreading surfactant on a thin liquid film. Applied Mathematics and Computation, 218(9):5157–5167, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013294.

Patterson:2013:CAS

[PS13] Richard F. Patterson and Ekrem Savas. Characterization of asymptotic statistical equivalent double and single sequences. Applied Mathematics and Computation, 224(??):484–491, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009429.

Patel:2014:FDV

[PS14a] Hiren G. Patel and Shambhu N.

Sharma. Filtering for a duffing-van der Pol stochastic differential equation. Applied Mathematics and Computation, 226(??):386–397, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313010953.

Popa:2014:HTA

[PS14b] C. Popa and C. Serban. Han-type algorithms for inconsistent systems of linear inequalities — a unified approach. Applied Mathematics and Computation, 246(??):247–256, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011126.

Shi:2014:BLC

[pSbL14] Jian ping Shi and Ji bin Li. Bifurcations of limit cycles in a Z_{6}-equivariant planar vector field of degree 7. Applied Mathematics and Computation, 244(??):191–200, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009357.

Pal:2012:TLS

[PSC12] Brojeswar Pal, Shib Sankar Sana, and Kripasindhu Chaudhuri. Three-layer supply chain — a productioninventory model for reworkable items. Applied Mathematics and Computation, 219(2):530–543, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006376.

Pando:2014:NIM

[PSJGLS14] Valentín Pando, Luis A.

San-José, Juan García-Laguna, and Joaquín Sicilia. A newsvendor inventory model with an emergency order to supply a non-increasing fraction of shortage. Applied Mathematics and Computation, 237(??):560–570, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004937.

Padhi:2010:TPS

[PSP10] Seshadev Padhi, Shilpee Srivastava, and Smita Pati. Three periodic solutions for a nonlinear first order functional differential equation. Applied Mathematics and Computation, 216(8):2450–2456, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310003449.

Papaschinopoulos:2011:NDE

[PSS11] G. Papaschinopoulos, C. J. Schinas, and G. Stefanidou. On the nonautonomous difference equation x_{n+1} = A_{n} + x_{n−1}^{p}/x_{n}^{q}. Applied Mathematics and Computation, 217(12):5573–5580, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012336.

Pandey:2014:AMA

[PSSS14] Rajesh K. Pandey, Suraj Suman, Koushlendra K. Singh, and Om P. Singh. An approximate method for Abel inversion using Chebyshev polynomials. Applied Mathematics and Computation, 237(??):120–132, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004056.

Pan:2010:SAG

[PSTL10] Quan-Ke Pan, P. N. Suganthan, M. Fatih Tasgetiren, and J. J. Liang. A self-adaptive global best harmony search algorithm for continuous optimization problems. Applied Mathematics and Computation, 216(3):830–848, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001128.

Panario:2014:GCR

[PSWW14] Daniel Panario, Murat Sahin, Qiang Wang, and William Webb. General conditional recurrences. Applied Mathematics and Computation, 243(??):220–231, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400808X.

Pasic:2011:FOS

[PT11a] Mervan Pasić and Satoshi Tanaka. Fractal oscillations of self-adjoint and damped linear differential equations of secondorder. Applied Mathematics and Computation, 218(5):2281–2293, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311009945.

Pilotta:2011:PWA

[PT11b] Elvio A. Pilotta and Germán A. Torres. A projected Weiszfeld algorithm for the box-constrained Weber location problem. Applied Mathematics and Computation, 218(6):2932–2943, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010812.

Papathanasiou:2013:VTI

[PT13] T. K. Papathanasiou and G. J. Tsamasphyros. On Vergnaud’s time integration method for autocatalytic cure rate equations. Applied Mathematics and Computation, 220(??):748–755, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007716.

Pedas:2014:SCN

[PT14a] Arvet Pedas and Enn Tamme. Spline collocation for nonlinear fractional boundary value problems. Applied Mathematics and Computation, 244(??):502–513, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009734.

Psarrakos:2014:GEM

[PT14b] Panayiotis J. Psarrakos and Michael J. Tsatsomeros. On the geometry of the envelope of a matrix. Applied Mathematics and Computation, 244(??):132–141, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009217.

Papakostas:2014:MBL

[PTK14] G. A. Papakostas, E. D.

Tsougenis, and D. E. Koulouriotis. Momentbased local image watermarking via genetic optimization. Applied Mathematics and Computation, 227(??):222–236, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012046.

Petkovic:2011:EPM

[PTS11] Marko D. Petković, Milan B. Tasić, and Predrag S. Stanimirović. Effective partitioning method for computing generalized inverses and their gradients. Applied Mathematics and Computation, 217(19):7588–7598, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002426.

Perkins:2011:CCL

[PTW11] William Perkins, Mark Tygert, and Rachel Ward. Computing the confidence levels for a root-mean-square test of goodnessof-fit. Applied Mathematics and Computation, 217(22):9072–9084, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004929.

Pang:2012:TDR

[PTY12] Yicheng Pang, Jianjun Paul Tian, and Hanchun Yang. Two-dimensional Riemann problem for a hyperbolic system of conservation laws in three pieces. Applied Mathematics and Computation, 219(4):1695–1711, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008028.

Pang:2013:TDR

[PTY13] Yicheng Pang, Jianjun Paul Tian, and Hanchun Yang. Two-dimensional Riemann problem involving three J’s for a hyperbolic system of nonlinear conservation laws. Applied Mathematics and Computation, 219(9):4614–4624, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010880.

Pandey:2010:MMS

[PV10] R. K. Pandey and Amit K.

Verma. Monotone method for singular BVP in the presence of upper and lower solutions. Applied Mathematics and Computation, 215(11):3860–3867, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300309010182.

Portman:2011:EUS

[PV11] N. Portman and E. R. Vrscay. Existence and uniqueness of solutions to the GRID macroscopic growth equation. Applied Mathematics and Computation, 217(21):8318–8327, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003754.

Pandey:2011:SSR

[PVV11] K. Pandey, Lajja Verma, and Amit K. Verma. L-stable Simpson’s 3/8 rule and Burgers’ equation. Applied Mathematics and Computation, 218(4):1342–1352, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311008307.

Peng:2010:IPS

[PW10] Zheng Peng and DongHua Wu. An inexact parallel splitting augmented Lagrangian method for large system of linear equations. Applied Mathematics and Computation, 216(5):1624–1636, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002602.

Posny:2014:CBR

[PW14] Drew Posny and Jin Wang. Computing the basic reproductive numbers for epidemiological models in nonhomogeneous environments. Applied Mathematics and Computation, 242(??):473–490, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400767X.

Peng:2013:BAA

[PWZZ13] Jian Peng, Lianhua Wang, Yueyu Zhao, and Yaobing Zhao. Bifurcation analysis in active control system with time delay feedback. Applied Mathematics and Computation, 219(19):10073– 10081, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004141.

Peng:2013:RLS

[PX13] Zhuohua Peng and Huimin Xin. The reflexive least squares solutions of the general coupled matrix equations with a submatrix constraint. Applied Mathematics and Computation, 225(??):425–445, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010412.

Xin:2010:ESB

[pXqLlZ10] Xiang peng Xin, Xi qiang Liu, and Lin lin Zhang. Explicit solutions of the Bogoyavlensky–Konoplechenko equation. Applied Mathematics and Computation, 215(10):3669–3673, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009795.

Yang:2013:GIA

[pYmK13] Li ping Yang and Wei ming Kong. A general iterative algorithm for semigroups of nonexpansive mappings with generalized contractive mapping. Applied Mathematics and Computation, 222(??):671–679, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008047.

Yang:2010:WSC

[pYX10] Li ping Yang and Xiangsheng Xie. Weak and strong convergence theorems for a finite family of I-asymptotically nonexpansive mappings. Applied Mathematics and Computation, 216(4):1057–1064, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001499.

Piccolomini:2011:IAL

[PZ11] E. Loli Piccolomini and F. Zama. An iterative algorithm for large size least-squares constrained regularization problems. Applied Mathematics and Computation, 217(24):10343–10354, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007132.

Pan:2012:NFD

[PZ12] Xintian Pan and Luming Zhang. A new finite difference scheme for the Rosenau–Burgers equation. Applied Mathematics and Computation, 218(17):8917–8924, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001919.

Pang:2014:PAP

[PZ14a] Liyan Pang and Tianwei Zhang. Permanence and almost periodic solution of a discrete impulsive Richards growth equation with variable delays and feedback control. Applied Mathematics and Computation, 248(??):399–420, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013459.

Pei:2014:TRA

[PZ14b] Yonggang Pei and Detong Zhu. A trust-region algorithm combining line search filter method with Lagrange merit function for nonlinear constrained optimization. Applied Mathematics and Computation, 247(??):281–300, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012193.

Zhang:2011:AEM

[pZJ11] Ju ping Zhang and Zhen Jin. The analysis of an epidemic model on networks. Applied Mathematics and Computation, 217(17):7053–7064, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010672.

Zhang:2012:ESC

[pZJ12] Ju ping Zhang and Zhen Jin. Epidemic spreading on complex networks with community structure. Applied Mathematics and Computation, 219(6):2829–2838, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006650.

Zhang:2010:FSG

[pZLC10] Huan ping Zhang, Biao Li, and Yong Chen. Full symmetry groups, Painlevé integrability and exact solutions of the nonisospectral BKP equation. Applied Mathematics and Computation, 217(4):1555–1560, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309006195.

Qi:2013:SSM

[QaC13] Longxing Qi and Jing an Cui. The stability of an SEIRS model with nonlinear incidence, vertical transmission and time delay. Applied Mathematics and Computation, 221(??):360–366, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300636X.

Qiu:2012:NPF

[QC12] Songqiang Qiu and Zhongwen Chen. A new penaltyfree-type algorithm based on trust region techniques. Applied Mathematics and Computation, 218(22):11089–11099, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004808.

Qin:2010:HPM

[QCK10a] Xiaolong Qin, Sun Young Cho, and Shin Min Kang. On hybrid projection methods for asymptotically quasi-ϕ-nonexpansive mappings. Applied Mathematics and Computation, 215(11):3874–3883, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010200.

Qin:2010:SRG

[QCK10b] Xiaolong Qin, Sun Young Cho, and Shin Min Kang. Some results on generalized equilibrium problems involving a family of nonexpansive mappings. Applied Mathematics and Computation, 217(7):3113–3126, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008970.

Cao:2010:MAN

[qClLyW10] Li qun Cao, Jian lan Luo, and Chong yu Wang. Multiscale analysis and numerical algorithm for the Schrödinger equations in heterogeneous media. Applied Mathematics and Computation, 217(8):3955–3973, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010209.

Du:2010:CAN

[qDG10] Shou qiang Du and Yan Gao. Convergence analysis of nonmonotone Levenberg–Marquardt algorithms for complementarity problem. Applied Mathematics and Computation, 216(5):1652–1659, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002638.

Qin:2013:PCR

[QFCZ13] Xiaolin Qin, Yong Feng, Jingwei Chen, and Jingzhong Zhang. Parallel computation of real solving bivariate polynomial systems by zero-matching method. Applied Mathematics and Computation, 219(14):7533–7541, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000787.

Qin:2014:EUS

[QFT^{+}14] Xinqiang Qin, Baoyan Fang, Shuangliang Tian, Xiaohong Tong, Zhigang Wang, and Lijun Su. The existence and uniqueness of solution to wavelet collocation. Applied Mathematics and Computation, 231(??):63–72, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013969.

Qiu:2014:ADG

[QGD14] Zhifeng Qiu, Ning Gui, and Geert Deconinck. Analysis of dynamic game played with inaccurate demand beliefs. Applied Mathematics and Computation, 230(??):530–541, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014215.

Gao:2013:MNB

[qGyFzX13] Jian qiang Gao, Li ya Fan, and Li zhong Xu. Median null(S_{w})-based method for face feature recognition. Applied Mathematics and Computation, 219(12):6410–6419, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000404.

Qu:2010:CSN

[QHL10] Xiaomei Qu, Chengming Huang, and Chao Liu. Convergence and stability of numerical solutions to a class of index 1 stochastic differential algebraic equations with time delay. Applied Mathematics and Computation, 215(11):4008–4021, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010418.

Qin:2014:CPS

[QHYW14] Xinqiang Qin, Gang Hu, Yang Yang, and Guo Wei. Construction of PH splines based on H-Bézier curves. Applied Mathematics and Computation, 238(??):460–467, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005669.

Qin:2013:NEP

[QHZ^{+}13] Xinqiang Qin, Gang Hu, Nianjuan Zhang, Xiaoli Shen, and Yang Yang. A novel extension to the polynomial basis functions describing Bézier curves and surfaces of degree n with multiple shape parameters. Applied Mathematics and Computation, 223(??):1–16, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008230.

Qiu:2010:CDS

[Qiu10] Zhipeng Qiu. The classification for dynamics of an STD model. Applied Mathematics and Computation, 216(5):1553–1565, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002481.

Qiu:2011:DEM

[Qiu11]  Zhipeng Qiu.  Dy-

namics of an epidemic model with host migration. Applied Mathematics and Computation, 218(8):4614–4625, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012902.

Qing-jun:2010:NTR

[Qj10] Wu Qing-jun. Nonmonotone trust region algorithm for unconstrained optimization problems. Applied Mathematics and Computation, 217(8):4274– 4281, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010040.

Qin:2013:SPI

[QK13] Zhongfeng Qin and Samarjit Kar. Single-period inventory problem under uncertain environment. Applied Mathematics and Computation, 219(18):9630–9638, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300313001367.

Qu:2012:HBC

[QL12] Rongning Qu and Xiaofang Li. Hopf bifurcation of a chemostat model. Applied Mathematics and Computation, 219(8):3541–3552, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009356.

Qian:2010:ISA

[QLCF10] Wei Qian, Tao Li, Shen Cong, and Shumin Fei. Improved stability analysis on delayed neural networks with linear fractional uncertainties. Applied Mathematics and Computation, 217(7):3596–3606, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009616.

Qu:2012:CRR

[QLL12] Aiping Qu, Donghui Li, and Min Li. The convergence rate of a restart MFR conjugate gradient method with inexact line search. Applied Mathematics and Computation, 218(23):11380–11390, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005085.

Qiu:2012:DSP

[QLWW12] Hong Qiu, Meng Liu, Ke Wang, and Yang Wang. Dynamics of a stochastic predator-prey system with Beddington– DeAngelis functional response. Applied Mathematics and Computation, 219(4):2303–2312, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008703.

Qin:2013:IEM

[QMM13] Zhenyun Qin, Gui Mu, and Hongcai Ma. G′/Gexpansion method for the fifth-order forms of KdV– Sawada–Kotera equation. Applied Mathematics and Computation, 222(??):29– 33, October 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031300742X.

Ma:2013:ERM

[qMWjHpX13] Hui qiang Ma, Meng Wu, Nan jing Huang, and Jiu ping Xu. Expected residual minimization method for stochastic variational inequality problems with nonlinear perturbations. Applied Mathematics and Computation, 219(11):6256–6267, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012416.

Qiu:2014:IPM

[QMZ14] Hailong Qiu, Liquan Mei, and Yamiao Zhang. Iterative penalty methods for the steady Navier– Stokes equations. Applied Mathematics and Computation, 237(??):110–119, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004834.

Qin:2012:PRM

[QQXY12] Huafeng Qin, Lan Qin, Lian Xue, and Chengbo Yu. A parallel recurrence method for the fast computation of Zernike moments. Applied Mathematics and Computation, 219(4):1549–1561, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007850.

Qin:2011:AAS

[QS11] Xiao-Ying Qin and Yan-Ping Sun. Approximate analytical solutions for a mathematical model of a tubular packed-bed catalytic reactor using an Adomian decomposition method. Applied Mathematics and Computation, 218(5):1990–1996, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009295.

Qin:2014:AAS

[QS14] Xiao-Ying Qin and Yan-Ping Sun. Approximate analytical solutions for a shrinking core model for the discharge of a lithium iron-phosphate electrode by the Adomian decomposition method. Applied Mathematics and Computation, 230(??):267–275, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014197.

Sun:2011:GIP

[qSyGH11] Jian qiang Sun, Xiao yan Gu, and Xun-Feng Hu. Geometric integration of the paraxial equation. Applied Mathematics and Computation, 218(7):3149–3157, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010885.

Qu:2012:SSB

[QTL^{+}12] Qi-Xing Qu, Bo Tian, Wen-Jun Liu, Pan Wang, and Yan Jiang. Soliton solutions and Bäcklund transformation for the normalized linearly coupled nonlinear wave equations with symbolic computation. Applied Mathematics and Computation, 218(21):10386–10392, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312003359.

Qi:2014:MIL

[QTYS14] Zhiquan Qi, Yingjie Tian, Xiaodan Yu, and Yong Shi. A multi-instance learning algorithm based on nonparallel classifier. Applied Mathematics and Computation, 241(??):233–241, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006729.

Quintanilla:2010:NNS

[Qui10] R. Quintanilla. A note on a non-standard problem for an equation with a delay term. Applied Mathematics and Computation, 216(9):2759–2765, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003905.

Qiang:2010:ETWb

[QW10a] Liu Qiang and Zhang Weiguo. Exact travelling wave solutions for the dissipative (2 + 1)- dimensional AKNS equation. Applied Mathematics and Computation, 217(2):735–744, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310006855.

Qiang:2010:ETWa

[QW10b] Liu Qiang and Zhang Weiguo. Exact travelling wave solutions of the dissipative coupled Korteweg–de Vries equation. Applied Mathematics and Computation, 215(12):4203–4215, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010893.

Qin:2011:SFR

[QW11a] Hai-Hua Qin and Ting Wei. Some filter regularization methods for a backward heat conduction problem. Applied Mathematics and Computation, 217(24):10317–10327, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007077.

Qiu:2011:EFS

[QW11b] Yuyang Qiu and Anding Wang. Eigenvector-free solutions to AX = B with P X = XP and XH = sX constraints. Applied Mathematics and Computation, 217(12):5650–5657, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012452.

Qiu:2014:SLC

[QW14] Yuyang Qiu and Anding Wang. Solving linearly constrained matrix least squares problem by LSQR. Applied Mathematics and Computation, 236(??):273–286, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400321X.

Wang:2011:PSB

[qWH11] Gen qiang Wang and Bo Han. Periodic solutions of a bang bang recurrence equation with least periods 1 through 9. Applied Mathematics and Computation, 217(9):4333–4349, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011057.

Yu:2010:ESCb

[qYqLtW10] Jin qian Yu, Xi qiang Liu, and Ting ting Wang. Exact solutions and conservation laws of (2 + 1)- dimensional Boiti–Leon– Pempinelli equation. Applied Mathematics and Computation, 216(8):2293– 2300, June 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031000319X.

Qi:2010:TDM

[QZ10] Yongfeng Qi and Jiashu Zhang. Two-dimensional multiple maximum scatter difference method for face recognition. Applied Mathematics and Computation, 216(12):3550–3557, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005163.

Qin:2013:USP

[QZS13] Yuming Qin, Jianpeng Zhang, and Linlin Sun. Upper semicontinuity of pullback attractors for a non-autonomous viscoelastic equation. Applied Mathematics and Computation, 223(??):362–376, October 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009028.

Rechdaoui:2010:ACS

[RA10a] M. S. Rechdaoui and L. Azrar. Active control of secondary resonances piezoelectric sandwich beams. Applied Mathematics and Computation, 216(11):3283–3302, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004832.

Ren:2010:CRM

[RA10b] Hongmin Ren and Ioannis K. Argyros. Convergence radius of the modified Newton method for multiple zeros under Hölder continuous derivative. Applied Mathematics and Computation, 217(2):612–621, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006715.

Ren:2010:LCS

[RA10c] Hongmin Ren and Ioannis K. Argyros. Local convergence of a secant type method for solving least squares problems. Applied Mathematics and Computation, 217(8):3816–3824, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010003.

Rana:2012:NSE

[RA12a] U. S. Rana and Asad Ahmad. Numerical solution of European call option with dividends and variable volatility. Applied Mathematics and Computation, 218(11):6242–6250, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014184.

Ren:2012:LCE

[RA12b] Hongmin Ren and I. K.

Argyros. Local convergence of efficient secanttype methods for solving nonlinear equations. Applied Mathematics and Computation, 218(14): 7655–7664, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312000720.

Ren:2012:SLC

[RA12c] Hongmin Ren and Ioannis K. Argyros. On the semi-local convergence of Halley’s method under a center-Lipschitz condition on the second Fréchet derivative. Applied Mathematics and Computation, 218(23):11488–11495, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312005383.

Radisavljevic:2011:IPA

[Rad11] Verica Radisavljevic. Improved Potter–Anderson– Moore algorithm for the differential Riccati equation. Applied Mathematics and Computation, 218(8):4641–4646, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011581.

Radenovic:2014:NTC

[Rad14]  Stojan Radenović.  A

note on tripled coincidence and tripled common fixed point theorems in partially ordered metric spaces. Applied Mathematics and Computation, 236(??):367–372, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004214.

Ragusa:2012:NSC

[Rag12] Maria Alessandra Ragusa.

Necessary and sufficient condition for a VMO function. Applied Mathematics and Computation, 218(24):11952–11958, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200584X.

Rahmoune:2013:SCM

[Rah13] Azedine Rahmoune. Spectral collocation method for solving Fredholm integral equations on the half-line. Applied Mathematics and Computation, 219(17):9254–9260, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002968.

Rashid:2014:NSC

[RAIM14] Abdur Rashid, Muhammad Abbas, Ahmad Izani Md. Ismail, and Ahmad Abd Majid. Numerical solution of the coupled viscous Burgers equations by Chebyshev– Legendre pseudo-spectral method. Applied Mathematics and Computation, 245(??):372–381, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010248.

Ralha:2012:GMA

[Ral12] Rui Ralha. The geometric mean algorithm. Applied Mathematics and Computation, 219(4):1607–1615, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312007941.

Ramos:2010:CDD

[Ram10a] Higinio Ramos. Contributions to the development of differential systems exactly solved by multistep finite-difference schemes. Applied Mathematics and Computation, 217(2):639–649, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006740.

Ramos:2010:AMB

[Ram10b] J. I. Ramos. Approximate methods based on order reduction for the periodic solutions of nonlinear third-order ordinary differential equations. Applied Mathematics and Computation, 215(12):4304–4319, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030901114X.

Ramadan:2011:RLB

[Ram11a] Ayad M. Ramadan. Range of lower bounds. Applied Mathematics and Computation, 218(3):1008–1011, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004346.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Rami:2011:FVP

[Ram11b] El-Nabulsi Ahmad Rami.

Fractional variational problems from extended exponentially fractional integral. Applied Mathematics and Computation, 217(22):9492–9496, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005480.

Ramirez:2014:CGF

[Ram14] José L. Ramírez. On convolved generalized Fibonacci and Lucas polynomials. Applied Mathematics and Computation, 229(??):208–213, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013325.

Raouf:2014:GBH

[Rao14] Azizi Raouf. Global behavior of the higher order rational Riccati difference equation. Applied Mathematics and Computation, 230(??):1–8, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013386.

Rassias:2014:CSR

[Ras14] Michael Th. Rassias. A cotangent sum related to zeros of the Estermann zeta function. Applied Mathematics and Computation, 240(??):161–167, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006262.

Razali:2013:MWD

[RAW13] Ahmad Mahir Razali and Ali A. Al-Wakeel. Mixture Weibull distributions for fitting failure times data. Applied Mathematics and Computation, 219(24):11358–11364, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006176.

Ray:2012:HWO

[Ray12] S. Saha Ray. On Haar wavelet operational matrix of general order and its application for the numerical solution of fractional Bagley Torvik equation. Applied Mathematics and Computation, 218(9):5239–5248, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013373.

Rizk-Allah:2013:HAC

[RAZES13] R. M. Rizk-Allah, Elsayed M. Zaki, and Ahmed Ahmed El-Sawy. Hybridizing ant colony optimization with firefly algorithm for unconstrained optimization problems. Applied Mathematics and Computation, 224(??):473–483, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008485.

Raheem:2011:SDC

[RB11a] Abdur Raheem and Dhirendra Bahuguna. A study of delayed cooperation diffusion system with Dirichlet boundary conditions. Applied Mathematics and Computation, 218(8):4169–4176, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012203.

Roshan:2011:NSM

[RB11b] Thoudam Roshan and K. S. Bhamra. Numerical solutions of the modified Burgers’ equation by Petrov–Galerkin method. Applied Mathematics and Computation, 218(7):3673–3679, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011763.

Raheem:2014:RMS

[RB14] Abdur Raheem and Dhirendra Bahuguna. Rothe’s method for solving some fractional integral diffusion equation. Applied Mathematics and Computation, 236(??):161–168, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003786.

Rakai:2012:AMB

[RBMA12] Logan Rakai, Laleh Behjat, Sebastian Martin, and Jose Aguado. An algebraic multigrid-based algorithm for circuit clustering. Applied Mathematics and Computation, 218(9):5202–5216, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013336.

Rosales:2014:ESS

[RBT14] J. C. Rosales, M. B. Branco, and D. Torrão. On the enumeration of the set of saturated numerical semigroups with fixed Frobenius number. Applied Mathematics and Computation, 236(??):471–479, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004202.

Ryabov:2012:NEF

[RC12] Pavel N. Ryabov and Stanislav A. Chesnokov. A note on the “Exp-function method for traveling waves of nonlinear evolution equations”. Applied Mathematics and Computation, 218(17):9024–9026, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001282.

See [NMDWAS10].

Ren:2012:CVI

[RCH12] Hongping Ren, Jing Cheng, and Aixiang Huang. The complex variable interpolating moving leastsquares method. Applied Mathematics and Computation, 219(4):1724–1736, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312008053.

Ren:2014:INS

[RCS14] Yong Ren, Xing Cheng, and R. Sakthivel. Impulsive neutral stochastic functional integrodifferential equations with infinite delay driven by fBm. Applied Mathematics and Computation, 247(??):205–212, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314012053.

Rios-Coelho:2010:MAC

[RCSH10] A. C. Rios-Coelho, W. F.

Sacco, and N. Henderson. A Metropolis algorithm combined with Hooke– Jeeves local search method applied to global optimization. Applied Mathematics and Computation, 217(2):843–853, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007125.

Rui:2010:SNS

[RCYL10] Weiguo Rui, Can Chen, Xinsong Yang, and Yao Long. Some new solitonlike solutions and periodic wave solutions with loop or without loop to a generalized KdV equation. Applied Mathematics and Computation, 217(4):1666–1677, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309008613.

Rach:2011:NFF

[RD11] Randolph Rach and Jun-Sheng Duan. Near-field and far-field approximations by the Adomian and asymptotic decomposition methods. Applied Mathematics and Computation, 217(12):5910–5922, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S009630031001310X.

Rhouma:2014:LEL

[RD14] Nedra Belhaj Rhouma and Amor Drissi. Large and entire large solutions for a class of nonlinear problems. Applied Mathematics and Computation, 232(??):272–284, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000381.

Rooin:2014:CNJ

[RDA14] Jamal Rooin, Hossein Dehghan, and Akram Alikhani. Comparison of the normalized Jensen functionals of two convex functions with applications. Applied Mathematics and Computation, 243(??):528–536, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008315.

Rakic:2014:CIC

[RDD14] Dragan S. Rakić, Nebojsa C. Dincić, and Dragan S. Djordjević. Core inverse and core partial order of Hilbert space operators. Applied Mathematics and Computation, 244(??):283–302, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009564.

Rocha:2013:PSM

[RDFL13] H. Rocha, J. M. Dias, B. C. Ferreira, and M. C. Lopes. Pattern search methods framework for beam angle optimization in radiotherapy design. Applied Mathematics and Computation, 219(23):10853–10865, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004955.

Raju:2014:SEM

[RDKU14] G. Naga Raju, P. Dutt, N. Kishore Kumar, and C. S. Upadhyay. Spectral element method for elliptic equations with periodic boundary conditions. Applied Mathematics and Computation, 246(??):426–439, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011321.

Rech:2013:PST

[Rec13]  Paulo C. Rech.  The

parameter-space of the three-parameter A2 symmetric flow. Applied Mathematics and Computation, 220(??):208–212, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006838.

Rech:2014:DHR

[Rec14] Paulo C. Rech. Delimiting hyperchaotic regions in parameter planes of a 5D continuous-time dynamical system. Applied Mathematics and Computation, 247(??):13–17, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011941.

Rembe:2010:EAM

[Rem10] Christian Rembe. Employing applied mathematics to expand the bandwidth of heterodyne carrier signals with a small phase modulation index. Applied Mathematics and Computation, 217(3):1202–1212, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004261.

Ren:2014:ITO

[Ren14a] Yong Ren. An integraltype operator from Q_{k}(p, q) spaces to Zygmund-type spaces. Applied Mathematics and Computation, 236(??):27–32, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003671.

Render:2014:CRB

[Ren14b] Hermann Render. Convergence of rational Bernstein operators. Applied Mathematics and Computation, 232(??):1076–1089, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001982.

Repovs:2011:SSP

[Rep11]  Dusan Repovs.  Sin-

gular solutions of perturbed logistic-type equations. Applied Mathematics and Computation, 218(8):4414–4422, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012483.

Respondek:2011:NRH

[Res11] Jerzy Stefan Respondek.

Numerical recipes for the high efficient inverse of the confluent Vandermonde matrices. Applied Mathematics and Computation, 218(5):2044–2054, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009362.

Respondek:2013:RNR

[Res13] Jerzy S. Respondek. Recursive numerical recipes for the high efficient inversion of the confluent Vandermonde matrices. Applied Mathematics and Computation, 225(??):718–730, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010734.

Reutskiy:2011:MMO

[Reu11]  S. Y. Reutskiy.  A

meshless method for onedimensional Stefan problems. Applied Mathematics and Computation, 217(23):9689–9701, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006230.

Reutskiy:2014:MAF

[Reu14a]  S. Y. Reutskiy.  The

method of approximate fundamental solutions (mafs) for Stefan problems for the sphere. Applied Mathematics and Computation, 227(??):648–655, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012101.

Reutskiy:2014:MPS

[Reu14b]  S. Yu. Reutskiy.  A

method of particular solutions for multi-point boundary value problems. Applied Mathematics and Computation, 243(??):559–569, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008509.

Rong:2012:DPB

[RFK12] Aiying Rong, José Rui Figueira, and Kathrin Klamroth. Dynamic programming based algorithms for the discounted 0-1 knapsack problem. Applied Mathematics and Computation, 218(12):6921–6933, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015554.

Rausanu:2014:HNM

[RG14] Silvia Rausanu and Crina Grosan. A hierarchical network model for epidemic spreading. analysis of A/H1N1 virus spreading in Romania. Applied Mathematics and Computation, 233(??):39–54, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000150.

Rueda:2014:PEF

[RGA14] M. Rueda, S. González, and A. Arcos. A predictive estimator of finite population proportion despite missing data. Applied Mathematics and Computation, 233(??):1–9, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001738.

Ramos:2010:VIF

[RGL10] J. I. Ramos and C. M. Garcı´ a-López. A Volterra integral formulation for determining the periodic solutions of some autonomous, nonlinear, third-order ordinary differential equations. Applied Mathematics and Computation, 216(9):2635– 2644, July 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310003620.

Ramos:2011:TCA

[RGL11] J. I. Ramos and C. M. García-López. On two classes of autonomous third-order nonlinear ordinary differential equations. Applied Mathematics and Computation, 218(5):2011–2024, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009325.

Rodrigo:2012:MFA

[RGL12] C. Rodrigo, F. J. Gaspar, and F. J. Lisbona. Multicolor Fourier analysis of the multigrid method for quadratic FEM discretizations. Applied Mathematics and Computation, 218(22):11182–11195, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004936.

Rusagara:2014:MAT

[RH14] I. Rusagara and C. Harley.

Mean action time as a measure for fin performance in one dimensional fins of exponential profiles. Applied Mathematics and Computation, 238(??):319–328, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005372.

Rhoades:2014:RSC

[Rho14a]  B. E. Rhoades.  RE-

TRACTED: Some comments concerning cyclic type contractions. Applied Mathematics and Computation, 233(??):505–506, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002021.

See retraction notices [Rho14b, Rho19].

Rhoades:2014:RNT

[Rho14b] B. E. Rhoades. Retraction notice to “Some comments concerning cyclic type contractions [AMC 233 (2014) 505–506]”. Applied Mathematics and Computation, 233(??):R1, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006766.

See [Rho14a]. From the Web page: “This article has been retracted at the request of the Author and Editor in Chief due to serious errors in the data.”.

Rhoades:2019:RNS

[Rho19] B. E. Rhoades. Retraction notice to “Some Comments Concerning Cyclic Type Contractions” [Applied Mathematics and Computation 233 (2014) 505–506]. Applied Mathematics and Computation, 341(??):452, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031830818X.

See [Rho14a].

Ren:2014:NSO

[RHSC14] Yong Ren, Tingting Hou, R. Sakthivel, and Xing Cheng. A note on the second-order nonautonomous neutral stochastic evolution equations with infinite delay under Carathéodory conditions. Applied Mathematics and Computation, 232(??):658–665, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001283.

Rimas:2013:CSP

[Rim13a] Jonas Rimas. Corrigendum to “Integer powers of real odd order skew-persymmetric antitridiagonal matrices with constant anti-diagonals (antitridiag_{n} (a, c, −a), a ∈ R[??]{0}, c ∈ R )” [Appl. Math. Comput. 219 (2013) 7075– 7088]. Applied Mathematics and Computation, 219 (17):9347, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003032.

See [Rim13c].

Rimas:2013:IPRb

[Rim13b]  Jonas Rimas.  Integer

powers of real even order anti-tridiagonal Hankel matrices of the form antitridiag_{n}(a, c, −a). Applied Mathematics and Computation, 225(??): 204–215, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010011.

Rimas:2013:IPRa

[Rim13c]  Jonas Rimas.  Integer

powers of real odd order skew-persymmetric antitridiagonal matrices with constant anti-diagonals (antitridiag_{n} (a, c, −a), a ∈ R[??]{0}, c ∈ R). Applied Mathematics and Computation, 219 (12):7075–7088, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/

pii/S0096300312013276.

See corrigendum [Rim13a].

Rostamy:2012:SNS

[RJ12] D. Rostamy and M. Jabbari. Sigmoidal–Nyström for solving an aerofoil equation. Applied Mathematics and Computation, 218(17):8492–8502, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001294.

Reidys:2014:ERR

[RJ14] Christian M. Reidys and Emma Yu Jin. The evolution of the random reversal graph. Applied Mathematics and Computation, 227(??):347–358, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012149.

Rao:2012:SOG

[RK12] S. Chandra Sekhara Rao and Sunil Kumar. Second order global uniformly convergent numerical method for a coupled system of singularly perturbed initial value problems. Applied Mathematics and Computation, 219(8):3740–3753, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009897.

Rashed:2014:HSE

[RK14a] A. S. Rashed and M. M. Kassem. Hidden symmetries and exact solutions of integro-differential Jaulent–Miodek evolution equation. Applied Mathematics and Computation, 247(??):1141–1155, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012430.

Rashedi:2014:CNC

[RK14b] Navid Rashedi and Hamed Kebriaei. Cooperative and non-cooperative Nash solution for linear supply function equilibrium game. Applied Mathematics and Computation, 244(??):794–808, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009989.

Rehan:2013:AWB

[RKA^{+}13] Muhammad Rehan, Abdul Qayyum Khan, Muhammad Abid, Naeem Iqbal, and Babar Hussain. Antiwindup-based dynamic controller synthesis for nonlinear systems under input saturation. Applied Mathematics and Computation, 220(??):382–393, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313006607.

Lin:2011:SPB

[rL11]  Su rong Lin.  The

singular perturbation of boundary value problem for third-order nonlinear vector integro-differential equation and its application. Applied Mathematics and Computation, 218(5):1746–1751, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008897.

Ranjini:2011:ZIL

[RLC11] P. S. Ranjini, V. Lokesha, and I. N. Cangül. On the Zagreb indices of the line graphs of the subdivision graphs. Applied Mathematics and Computation, 218(3):699–702, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004930.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Rincon:2010:ANS

[RLV10] M. A. Rincon, J. Límaco, and R. Vale. Analysis and numerical solution of Benjamin–Bona– Mahony equation with moving boundary. Applied Mathematics and Computation, 216(1):138–148, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000366.

Rosen:2014:BDD

[RLW14] I. G. Rosen, Susan E. Luczak, and Jordan Weiss. Blind deconvolution for distributed parameter systems with unbounded input and output and determining blood alcohol concentration from transdermal biosensor data. Applied Mathematics and Computation, 231(??): 357–376, March 15, 2014.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313013891.

Rababah:2011:IPM

[RM11] Abedallah Rababah and Stephen Mann. Iterative process for G^{2}-multi degree reduction of Bézier curves. Applied Mathematics and Computation, 217(20):8126–8133, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003559.

Ren:2012:SSS

[RM12a] Junbai Ren and Xuan Ma. L^{2} stability for solutions of subcritical quasi-geostrophic equation with large perturbations. Applied Mathematics and Computation, 218(9):5337–5345, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013476.

Rivadeneira:2012:ICS

[RM12b] P. S. Rivadeneira and C. H. Moog. Impulsive control of single-input nonlinear systems with application to HIV dynamics. Applied Mathematics and Computation, 218(17):8462–8474, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001221.

Rivaie:2012:NCN

[RMJM12] Mohd Rivaie, Mustafa Mamat, Leong Wah June, and Ismail Mohd. A new class of nonlinear conjugate gradient coefficients with global convergence properties. Applied Mathematics and Computation, 218(22):11323–11332, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005152.

Raya-Miranda:2011:DDL

[RMMM11] R. Raya-Miranda and M. D. Martínez-Miranda. Data-driven local bandwidth selection for additive models with missing data. Applied Mathematics and Computation, 217(24):10328–10342, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007090.

Ragb:2014:ACP

[RMN14] Ola Ragb, M. S. Matbuly, and M. Nassar. Analysis of composite plates using moving least squares differential quadrature method. Applied Mathematics and Computation, 238(??):225–236, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400349X.

Rahman:2014:IMF

[RMY14] Muhammad Rahman, Afzal Mahmood, and Muhammad Younis. Improved and more feasible numerical methods for Riesz space fractional partial differential equations. Applied Mathematics and Computation, 237(??): 264–273, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314004743.

Rojsiraphisal:2010:ESC

[RN10a] Thaned Rojsiraphisal and Piyapong Niamsup. Exponential stability of certain neutral differential equations. Applied Mathematics and Computation, 217(8):3875–3880, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010076.

Roopkumar:2010:PTB

[RN10b] R. Roopkumar and E. R. Negrin. Poisson transform on Boehmians. Applied Mathematics and Computation, 216(9):2740–2748, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003875.

Riaz:2014:IWF

[RNEA14] Arshad Riaz, S. Nadeem, R. Ellahi, and Noreen Sher Akbar. The influence of wall flexibility on unsteady peristaltic flow of Prandtl fluid in a three dimensional rectangular duct. Applied Mathematics and Computation, 241(??):389–400, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005864.

Rui:2010:MFM

[RnZ10] Weiguo Rui and Zuo nong Zhu. Mixed function method for obtaining exact solution of nonlinear differential-difference equations and coupled NDDEs. Applied Mathematics and Computation, 216(5):1501–1512, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002407.

Rahimov:2013:SSS

[RNZH13] H. Rahimov, H. Nikoofard, S. Zarrinkamar, and H. Hassanabadi. Any l-state solutions of the Schrödinger equation for the Modified Woods– Saxon potential in arbitrary dimensions. Applied Mathematics and Computation, 219(9):4710–4717, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011034.

Rhoades:2011:DAF

[ROA11] B. E. Rhoades, Kevser Ozkoklu, and Inci Albayrak. On the degree of approximation of functions belonging to a Lipschitz class by Hausdorff means of its Fourier series. Applied Mathematics and Computation, 217(16):6868–6871, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000506.

Rady:2010:HBM

[ROK10] A. S. Abdel Rady, E. S. Osman, and Mohammed Khalfallah. The homogeneous balance method and its application to the Benjamin–Bona–Mahoney (BBM) equation. Applied Mathematics and Computation, 217(4):1385–1390, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309005141.

Rady:2012:SSD

[ROK12] A. S. Abdel Rady, E. S. Osman, and Mohammed Khalfallah. On soliton solutions of the (2 + 1) dimensional Boussinesq equation. Applied Mathematics and Computation, 219(8):3414–3419, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309005153.

Rosca:2011:URG

[Ros11a]  Daniela Rosca.  Uni-

form and refinable grids on elliptic domains and on some surfaces of revolution. Applied Mathematics and Computation, 217(19):7812–7817, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003183.

Roshan:2011:PGM

[Ros11b]  Thoudam Roshan.  A

Petrov–Galerkin method for equal width equation. Applied Mathematics and Computation, 218(6):2730–2739, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010393.

Ray:2013:HWO

[RP13] S. Saha Ray and A. Patra. Haar wavelet operational methods for the numerical solutions of fractional order nonlinear oscillatory Van der Pol system. Applied Mathematics and Computation, 220(??):659–667, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007868.

Ramos:2014:SNI

[RP14a] Higinio Ramos and M. F. Patricio. Some new implicit two-step multiderivative methods for solving special secondorder IVP’s. Applied Mathematics and Computation, 239(??):227–241, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005815.

Resendiz:2014:ABO

[RP14b] Edgar Reséndiz and René Pinnau. Adjoint-based optimization of particle trajectories in laminar flows. Applied Mathematics and Computation, 248(??):567–583, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013721.

Roshan:2014:SCP

[RPA14] J. R. Roshan, V. Parvaneh, and I. Altun. Some coincidence point results in ordered b-metric spaces and applications in a system of integral equations. Applied Mathematics and Computation, 226(??):725–737, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011004.

Ren:2014:EAM

[RPW14] Hongping Ren, Kaiyan Pei, and Liping Wang. Error analysis for moving least squares approximation in 2D space. Applied Mathematics and Computation, 238(??):527–546, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005700.

Razani:2011:GCP

[RRG11] Abdolrahman Razani, Vladimir Rakocević, and Zahra Goodarzi. Generalized ϕcontraction for a pair of mappings on cone metric spaces. Applied Mathematics and Computation, 217(22):8899–8906, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002293.

Rihan:2014:TDM

[RRLA14] F. A. Rihan, D. H. Abdel Rahman, S. Lakshmanan, and A. S. Alkhajeh. A time delay model of tumourimmune system interactions: Global dynamics, parameter estimation, sensitivity analysis. Applied Mathematics and Computation, 232(??):606–623, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001568.

Rodriguez:2012:ESN

[RRM12] F. Rodríguez, M. Roales, and J. A. Martín. Exact solutions and numerical approximations of mixed problems for the wave equation with delay. Applied Mathematics and Computation, 219(6):3178–3186, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009496.

Rosca:2014:ASP

[RRP14] Alin V. Rosca, Natalia C.

Rosca, and Ioan Pop. Axisymmetric stagnation point flow and heat transfer towards a permeable moving flat plate with surface slip condition. Applied Mathematics and Computation, 233(??):139–151, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001751.

Radenovic:2011:CFP

[RRR11] Stojan Radenović, Vladimir Rakocević, and Shahram Resapour. Common fixed points for (g, f) type maps in cone metric spaces. Applied Mathematics and Computation, 218(2):480–491, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007697.

Roman-Roman:2012:EPG

[RRRRTR12] P. Román-Román, D. Romero, M. A. Rubio, and F. Torres-Ruiz. Estimating the parameters of a Gompertztype diffusion process by means of Simulated Annealing. Applied Mathematics and Computation, 218(9):5121–5131, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013233.

Roman-Roman:2012:RPE

[RRSPTR12] P. Román-Román, J. J. Serrano-Pérez, and F. Torres-Ruiz. An R package for an efficient approximation of first-passage-time densities for diffusion processes based on the FPTL function. Applied Mathematics and Computation, 218(17):8408–8428, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001154.

Roman-Roman:2014:MGP

[RRSPTR14] P. Román-Román, J. J. Serrano-Pérez, and F. Torres-Ruiz. More general problems on first-passage times for diffusion processes: a new version of the fptdApprox R package. Applied Mathematics and Computation, 244(??):432–446, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009552.

Rafsanjani:2014:QNA

[RRSS14] Marjan Kuchaki Rafsanjani, Atieh Rezaei, Amin Shahraki, and Arsham Borumand Saeid. QARIMA: a new approach to prediction in queue theory. Applied Mathematics and Computation, 244(??):514–525, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009527.

Rangarajan:2010:INQ

[RS10] Govindan Rangarajan and Shrihari Sridharan. Invariant norm quantifying nonlinear content of Hamiltonian systems. Applied Mathematics and Computation, 217(6):2495–2500, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310008052.

Rai:2011:NAS

[RS11a] Pratima Rai and Kapil K. Sharma. Numerical analysis of singularly perturbed delay differential turning point problem. Applied Mathematics and Computation, 218(7):3483–3498, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011362.

Ren:2011:DID

[RS11b] Guojing Ren and Yuming Shi. Defect indices and definiteness conditions for a class of discrete linear Hamiltonian systems. Applied Mathematics and Computation, 218(7):3414–3429, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011271.

Rutka:2012:CSE

[RS12] Przemysław Rutka and Ryszard Smarzewski. Complete solution of the electrostatic equilibrium problem for classical weights.

Applied Mathematics and Computation, 218(10): 6027–6037, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300311014305.

Rebenda:2013:SAP

[RS13] Josef Rebenda and Zdenek Smarda. Stability and asymptotic properties of a system of functional differential equations with nonconstant delays. Applied Mathematics and Computation, 219(12):6622–6632, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313000027.

Rademacher:2014:MFH

[RSBK14] Andreas Rademacher, Andreas Schröder, Heribert Blum, and Heiko Kleemann. Mixed FEM of higher-order for timedependent contact problems. Applied Mathematics and Computation, 233(??):165–186, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314001003.

Rapaic:2014:NCF

[RSG14] Milan R. Rapaić, Tomislav B. Sekara, and Vidan Govedarica. A novel class of fractionally orthogonal quasi-polynomials and new fractional quadrature formulas. Applied Mathematics and Computation, 245(??):206–219, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314010480.

Ryabov:2011:AKM

[RSK11] Pavel N. Ryabov, Dmitry I. Sinelshchikov, and Mark B. Kochanov. Application of the Kudryashov method for finding exact solutions of the high order nonlinear evolution equations. Applied Mathematics and Computation, 218(7):3965–3972, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011969.

Rao:2014:BEA

[RSK14] S. Chandra Sekhara Rao, Sarita, and Rabia Kamra.

Backward error analysis of Choleski q.i.f. for the solution of symmetric positive definite linear systems. Applied Mathematics and Computation, 226(??):398–405, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010990.

Raza:2011:NSB

[RSN11] N. Raza, S. Sial, and J. W.

Neuberger. Numerical solution of Burgers’ equation by the Sobolev gradient method. Applied Mathematics and Computation, 218(8):4017–4024, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010319.

Rakkiyappan:2013:SDS

[RSPK13] R. Rakkiyappan, N. Sakthivel, Ju H. Park, and O. M. Kwon. Sampleddata state estimation for Markovian jumping fuzzy cellular neural networks with modedependent probabilistic time-varying delays. Applied Mathematics and Computation, 221(??):

741–769, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007571.

Revathi:2014:EAA

[RSRA14] P. Revathi, R. Sakthivel, Yong Ren, and S. Marshal Anthoni. Existence of almost automorphic mild solutions to non-autonomous neutral stochastic differential equations. Applied Mathematics and Computation, 230(??):639–649, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013581.

Rambaud:2014:SCS

[RT14] Salvador Cruz Rambaud and María José Muñoz Torrecillas. Some characterizations of (strongly) subadditive discounting functions. Applied Mathematics and Computation, 243(??):368–378, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007954.

Rossi:2014:PLL

[RTPH14] Roberto Rossi, S. Armagan Tarim, Steven Prestwich, and Brahim Hnich. Piecewise linear lower and upper bounds for the standard normal first order loss function. Applied Mathematics and Computation, 231(??):489–502, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000563.

Rodriguez:2012:DNA

[RTV12] J. M. Rodríguez and R. Taboada-Vázquez. Derivation of a new asymptotic viscous shallow water model with dependence on depth. Applied Mathematics and Computation, 219(7):3292–3307, December 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010939.

Rivero:2011:FDP

[RTVV11] Margarita Rivero, Juan J.

Trujillo, Luis Vázquez, and M. Pilar Velasco. Fractional dynamics of populations. Applied Mathematics and Computation, 218(3):1089–1095, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003717.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Rekab:2013:ETA

[RTW13] Kamel Rekab, Herbert Thompson, and Wei Wu. An efficient test allocation for software reliability estimation. Applied Mathematics and Computation, 220(??):94–103, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006255.

Rudek:2012:SNH

[Rud12] Radosław Rudek. The strong NP-hardness of the maximum lateness minimization scheduling problem with the processingtime based aging effect. Applied Mathematics and Computation, 218(11):6498–6510, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311014792.

Rudek:2013:CCA

[Rud13] Radosław Rudek. The computational complexity analysis of the twoprocessor flowshop problems with position dependent job processing times. Applied Mathematics and Computation, 221(??):819–832, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007352.

Rajtmajer:2010:NEH

[RV10a] Sarah Michele Rajtmajer and Damir Vukičević. A note on the Estrada– Hatano communicability algorithm for detecting community structure in complex networks. Applied Mathematics and Computation, 217(7):3516–3521, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009732.

Rus:2010:TSP

[RV10b] Francisco Rus and Francisco R. Villatoro. Timestepping in Petrov–Galerkin methods based on cubic B-splines for compactons. Applied Mathematics and Computation, 217(6):2788–2797, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008672.

Rahman:2013:FMA

[RV13a] Talal Rahman and Jan Valdman. Fast MAT-LAB assembly of FEM matrices in 2D and 3D: Nodal elements. Applied Mathematics and Computation, 219(13):7151–7158, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010836.

Russo:2013:CEH

[RV13b] Matthew Russo and Robert A. Van Gorder. Control of error in the homotopy analysis of nonlinear Klein– Gordon initial value problems. Applied Mathematics and Computation, 219(12):6494–6509, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013252.

Rafei:2014:CSF

[RV14] M. Rafei and W. T. Van Horssen. On constructing solutions for the functional equation Z(x, y, n) = Z(a_{11}x + a_{12}y, a_{21}x + a_{22}y, n + 1). Applied Mathematics and Computation, 237(??):373–385, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005049.

Reithmeier:2010:MUS

[RVS10] Eduard Reithmeier, Taras Vynnyk, and Thanin Schultheis. 3D-measurement using a scanning electron microscope. Applied Mathematics and Computation, 217(3):1193–1201, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001311.

Ran:2014:PIM

[RW14a] Yu-Hong Ran and Jun-Gang Wang. On preconditioned iterative methods for unsteady incompressible Navier–Stokes equations. Applied Mathematics and Computation, 234(??):477–485, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002689.

Rong:2014:FVS

[RW14b] Yisheng Rong and Yuechuan Wei. A flux vector splitting scheme for low Mach number flows in preconditioning method. Applied Mathematics and Computation, 242(??):296–308, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314007322.

Ren:2010:CNS

[RWB10] Hongmin Ren, Qingbiao Wu, and Weihong Bi. On convergence of a new secant-like method for solving nonlinear equations. Applied Mathematics and Computation, 217(2):583–589, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006612.

Ren:2013:SNE

[RWBA13] Hongmin Ren, Long Wu, Weihong Bi, and Ioannis K. Argyros. Solving nonlinear equations system via an efficient genetic algorithm with symmetric and harmonious individuals. Applied Mathematics and Computation, 219(23):10967–10973, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004657.

Ronghua:2010:CNS

[RWkPk10] Li Ronghua, Pang Wankai, and Leung Ping-kei. Convergence of numerical solutions to stochastic age-structured population equations with diffusions and Markovian switching. Applied Mathematics and Computation, 216(3):744–752, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000974.

Rui:2014:ISM

[RX14] Shaoping Rui and Chengxian Xu. An inexact smoothing method for SOCCPs based on a one-parametric class of smoothing function. Applied Mathematics and Computation, 241(??): 167–182, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314006638.

Ran:2010:MBS

[RY10] Yu-Hong Ran and Li Yuan. On modified block SSOR iteration methods for linear systems from steady incompressible viscous flow problems. Applied Mathematics and Computation, 217(7):3050–3068, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310008933.

Ren:2012:EPN

[RY12] Dongxiao Ren and Mao Ye. Extracting postnonlinear signal with specific kurtosis range. Applied Mathematics and Computation, 218(9):5726– 5738, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311014135.

Rassias:2013:MHD

[RY13a] Michael Th. Rassias and Bicheng Yang. A multidimensional half-discrete Hilbert-type inequality and the Riemann zeta function. Applied Mathematics and Computation, 225(??):263–277, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010199.

Rassias:2013:HDH

[RY13b] Michael Th. Rassias and Bicheng Yang. On halfdiscrete Hilbert’s inequality. Applied Mathematics and Computation, 220(??):75–93, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300622X.

Rahmani:2014:NSF

[RY14a] Rasoul Rahmani and Rubiyah Yusof. A new simple, fast and efficient algorithm for global optimization over continuous search-space problems: Radial Movement Optimization. Applied Mathematics and Computation, 248(??):287–300, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013368.

Rassias:2014:MHD

[RY14b] Michael Th. Rassias and Bicheng Yang. On a multidimensional halfdiscrete Hilbert-type inequality related to the hyperbolic cotangent function. Applied Mathematics and Computation, 242(??):800–813, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008996.

Rassias:2014:MHT

[RY14c] Michael Th. Rassias and Bicheng Yang. On a multidimensional Hilberttype integral inequality associated to the gamma function. Applied Mathematics and Computation, 249(??):408–418, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014039.

Ryabov:2010:ESK

[Rya10] Pavel N. Ryabov. Exact solutions of the Kudryashov–Sinelshchikov equation. Applied Mathematics and Computation, 217(7):3585–3590, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009501.

Ryoo:2010:NAP

[Ryo10] Cheon Seoung Ryoo. A numerical approach to the proof of existence of solutions for some generalized obstacle problems. Applied Mathematics and Computation, 216(11):3365–3369, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004923.

Rezvan:2011:GPC

[RYÖ11] Farshad Rezvan, Emrullah Yasar, and Teoman Özer. Group properties and conservation laws for nonlocal shallow water wave equation. Applied Mathematics and Computation, 218(3):974–979, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003742.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ruan:2011:GDS

[RZ11] Lizhi Ruan and Zhiyong Zhang. Global decaying solution to dissipative nonlinear evolution equations with ellipticity. Applied Mathematics and Computation, 217(13):6054–6066, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012683.

Ren:2012:NFP

[RZ12] Junchao Ren and Qingling Zhang. Non-fragile PD state H_{∞} control for a class of uncertain descriptor systems. Applied Mathematics and Computation, 218(17):8806–8815, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200166X.

Ruan:2013:OPC

[RZHH13] Xinfeng Ruan, Wenli Zhu, Jin Hu, and Jiexiang Huang. Optimal portfolio and consumption with habit formation in a jump diffusion market. Applied Mathematics and Computation, 222(??):391–401, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008138.

See errata [RZHH14].

Ruan:2014:ECO

[RZHH14] Xinfeng Ruan, Wenli Zhu, Jin Hu, and Jiexiang Huang. Errata corrige optimal portfolio and consumption with habit formation in a jump diffusion market. Applied Mathematics and Computation, 232(??):235–236, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000976.

See [RZHH13].

S:2010:NTW

[S10] Cesar A. Gomez S. New traveling waves solutions to generalized Kaup– Kupershmidt and Ito equations. Applied Mathematics and Computation, 216(1):241–250, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000585.

S:2014:CFS

[S.14] Cesar A. Gómez S. Closed form solutions for a generalized Benjamin–Bona– Mahony–Burgers equation with higher-order nonlinearity. Applied Mathematics and Computation, 234(??):618–622, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003038.

Sen:2010:CDL

[SA10a] S. K. Sen and Ravi P. Agarwal. Central difference limit for derivatives of ill-posed functions: Best approach in supercomputing era. Applied Mathematics and Computation, 217(7):3338–3348, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009306.

Sen:2010:ZCP

[SA10b] S. K. Sen and Ravi P.

Agarwal. Zero-clusters of polynomials: Best approach in supercomputing era. Applied Mathematics and Computation, 215(12):4080–4093, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010066.

Sahiner:2011:NSS

[SA11] Ahmet Sahiner and H. Seda Akin. New sequence spaces in multiple normed spaces. Applied Mathematics and Computation, 218(3):1030–1035, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100525X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Sarvi:2012:DIN

[SA12a] Masoud Nasiri Sarvi and M. T. Ahmadian. Design and implementation of a new spherical super element in structural analysis. Applied Mathematics and Computation, 218(14):7546–7561, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000550.

Su:2012:GCN

[SA12b] Ke Su and Hui An. Global convergence of a nonmonotone filter method for equality constrained optimization. Applied Mathematics and Computation, 218(18):9396–9404, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200255X.

Sharma:2013:EWN

[SA13] Janak Raj Sharma and Himani Arora. On efficient weighted-Newton methods for solving systems of nonlinear equations. Applied Mathematics and Computation, 222(??):497–506, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313008163.

Sawyerr:2014:RCG

[SAA14] B. A. Sawyerr, A. O. Adewumi, and M. M. Ali. Real-coded genetic algorithm with uniform random local search. Applied Mathematics and Computation, 228(??):589–597, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012666.

Stevic:2013:NSO

[SAAS13] Stevo Stević, Mohammed A.

Alghamdi, Abdullah Alotaibi, and Naseer Shahzad. On a nonlinear second order system of difference equations. Applied Mathematics and Computation, 219(24):11388–11394, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005341.

Stevic:2014:EPS

[SAAS14a] Stevo Stević, Mohammed A.

Alghamdi, Abdullah Alotaibi, and Naseer Shahzad. Eventual periodicity of some systems of maxtype difference equations. Applied Mathematics and Computation, 236(??):635–641, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313014392.

Stevic:2014:LTB

[SAAS14b] Stevo Stević, Mohammed A. Alghamdi, Abdullah Alotaibi, and Naseer Shahzad. Long-term behavior of positive solutions of a system of max-type difference equations. Applied Mathematics and Computation, 235(??):567–574, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012137.

Stevic:2014:MRS

[SAAS14c] Stevo Stević, Mohammed A. Alghamdi, Abdullah Alotaibi, and Naseer Shahzad. More on a rational secondorder difference equation. Applied Mathematics and Computation, 236(??):565–571, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012496.

Sabelfeld:2013:SSP

[Sab13]  K. K. Sabelfeld.  A

stochastic spectral projection method for solving PDEs in domains composed by overlapping discs, spheres, and halfspaces. Applied Mathematics and Computation, 219(10):5123–5139, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011745.

See corrigendum [Sab14].

Sabelfeld:2014:CSS

[Sab14] K. K. Sabelfeld. Corrigendum to “A stochastic spectral projection method for solving PDEs in domains composed by overlapping discs, spheres, and halfspaces”. Applied Mathematics and Computation, 248(??):697–701, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401340X.

See [Sab13].

Saddeek:2011:CAG

[Sad11] A. M. Saddeek. Convergence analysis of generalized iterative methods for some variational inequalities involving pseudomonotone operators in Banach spaces. Applied Mathematics and Computation, 217(10):4856–4865, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011501.

Safadi:2014:UIC

[Sáf14] Thelma Sáfadi. Using independent component for clustering of time series data. Applied Mathematics and Computation, 243(??):522–527, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314008637.

Sagiroglu:2011:ECA

[Sag11] Yasemin Sagiroğlu. The equivalence of curves in SL(n, R) and its application to ruled surfaces. Applied Mathematics and Computation, 218(3):1019–1024, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004383.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Sahin:2011:GFF

[Sah11a] Murat Sahin. The generating function of a family of the sequences in terms of the continuant. Applied Mathematics and Computation, 217(12):5416–5420, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012117.

Sahin:2011:SNS

[Sah11b] Recep Sahin. On the some normal subgroups of the extended modular group. Applied Mathematics and Computation, 218(3):1025–1029, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004425.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Srivastava:2014:GFG

[SAJ14] H. M. Srivastava, Praveen Agarwal, and Shilpi Jain. Generating functions for the generalized Gauss hypergeometric functions. Applied Mathematics and Computation, 247(??): 348–352, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012156.

Shukri:2010:ETM

[SAK10] Sami Shukri and Kamel Al-Khaled. The extended tanh method for solving systems of nonlinear wave equations. Applied Mathematics and Computation, 217(5):1997–2006, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007435.

Salas:2010:CES

[Sal10a] Alvaro H. Salas. Computing exact solutions to a generalized Lax seventhorder forced KdV equation (KdV7). Applied Mathematics and Computation, 216(8):2333–2338, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003322.

Salas:2010:ESM

[Sal10b] Alvaro H. Salas. Exact solutions to mKdV equation with variable coefficients. Applied Mathematics and Computation, 216(10):2792–2798, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003942.

Salas:2010:SCS

[Sal10c] Alvaro H. Salas. Symbolic computation of solutions for a forced Burgers equation. Applied Mathematics and Computation, 216(1):18–26, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630030901042X.

Salah:2011:ATS

[Sal11a]  K. Salah.  Analysis

of a two-stage network server. Applied Mathematics and Computation, 217(23):9635–9645, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006187.

Salahi:2011:RTR

[Sal11b]  Maziar Salahi.  Regu-

larization tools and robust optimization for illconditioned least squares problem: a computational comparison. Applied Mathematics and Computation, 217(20):7985–7990, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003237.

Salem:2011:GMS

[Sal11c] Hussein A. H. Salem. Global monotonic solutions of multi term fractional differential equations. Applied Mathematics and Computation, 217(14):6597–6603, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000609.

Senan:2012:NRS

[SAL12] Sibel Senan, Sabri Arik, and Derong Liu. New robust stability results for bidirectional associative memory neural networks with multiple time delays. Applied Mathematics and Computation, 218(23):11472–11482, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005358.

Samanta:2010:TSC

[Sam10] G. P. Samanta. A twospecies competitive system under the influence of toxic substances. Applied Mathematics and Computation, 216(1):291–299, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000743.

Samanta:2011:AND

[Sam11] G. P. Samanta. Analysis of a nonautonomous dynamical model of diseases through droplet infection and direct contact. Applied Mathematics and Computation, 217(12):5870–5888, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012762.

Sakthivel:2011:RPA

[SAMA11] R. Sakthivel, A. Arunkumar, K. Mathiyalagan, and S. Marshal Anthoni. Robust passivity analysis of fuzzy Cohen–Grossberg BAM neural networks with time-varying delays. Applied Mathematics and Computation, 218(7):3799–3809, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011908.

Sana:2010:OSP

[San10] Shib Sankar Sana. Optimal selling price and lotsize with time varying deterioration and partial backlogging. Applied Mathematics and Computation, 217(1):185–194, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310005849.

Sana:2011:EMS

[San11a] Shib Sankar Sana. An EOQ model for salesmen’s initiatives, stock and price sensitive demand of similar products — a dynamical system. Applied Mathematics and Computation, 218(7):3277–3288, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311011076.

Sana:2011:PSD

[San11b] Shib Sankar Sana. Pricesensitive demand for perishable items — an EOQ model. Applied Mathematics and Computation, 217(13):6248–6259, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013305.

Sana:2011:SEM

[San11c] Shib Sankar Sana. The stochastic EOQ model with random sales price. Applied Mathematics and Computation, 218(2):239– 248, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311006709.

Sana:2012:EMD

[San12] Shib Sankar Sana. The EOQ model — a dynamical system. Applied Mathematics and Computation, 218(17):8736–8749, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312001592.

Sangawi:2014:SSM

[San14] Ali W. K. Sangawi. Spiral slits map and its inverse of bounded multiply connected regions. Applied Mathematics and Computation, 228(??):520–530, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012836.

Saker:2012:GBZ

[SAO12] S. H. Saker, R. P. Agarwal, and Donal O’Regan. Gaps between zeros of secondorder half-linear differential equations. Applied Mathematics and Computation, 219(3):875–885, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006662.

Sharma:2014:EDF

[SAP14] Janak Raj Sharma, Himani Arora, and Miodrag S. Petković. An efficient derivative free family of fourth order methods for solving systems of nonlinear equations. Applied Mathematics and Computation, 235(??):383–393, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003518.

Sarigol:2010:LPF

[Sar10] Mehmet Ali Sarıgöl. On the local properties of factored Fourier series. Applied Mathematics and Computation, 216(11): 3386–3390, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310005102.

Sarkar:2012:EMD

[Sar12a] Biswajit Sarkar. An EOQ model with delay in payments and stock dependent demand in the presence of imperfect production. Applied Mathematics and Computation, 218(17):8295–8308, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000987.

Sarkar:2012:IMR

[Sar12b] Biswajit Sarkar. An inventory model with reliability in an imperfect production process. Applied Mathematics and Computation, 218(9):4881–4891, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012999.

Sarra:2012:LRB

[Sar12c] Scott A. Sarra. A local radial basis function method for advectiondiffusion-reaction equations on complexly shaped domains. Applied Mathematics and Computation, 218(19):9853–9865, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003013.

Sarkar:2013:DSL

[Sar13] Nantu Sarkar. On the discontinuity solution of the Lord–Shulman model in generalized thermoelasticity. Applied Mathematics and Computation, 219 (20):10245–10252, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313003883.

Sastre:2012:EMR

[Sas12]  Jorge Sastre.  Effi-

cient mixed rational and polynomial approximation of matrix functions. Applied Mathematics and Computation, 218(24):11938–11946, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005826.

Sauter:2013:SOD

[SATdV13] Esequia Sauter, Fabio Souto Azevedo, Mark Thompson, and Marco Tullio Menna Barreto de Vilhena. Solution of the one-dimensional transport equation by the vector Green function method: Error bounds and simulation. Applied Mathematics and Computation, 219(24):11291–11301, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005584.

Sauer:2011:CRT

[Sau11] Timothy Sauer. Convergence of rank-type equations. Applied Mathematics and Computation, 217(9):4540–4547, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011173.

Sanaullah:2014:GEC

[SAuAH14] Aamir Sanaullah, Humera Ameer Ali, Muhammad Noor ul Amin, and Muhammad Hanif. Generalized exponential chain ratio estimators under stratified two-phase random sampling. Applied Mathematics and Computation, 226(??):541–547, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011557.

Savas:2010:SSS

[Sav10] E. Savaş. Δ^{m}-Strongly summable sequences spaces in 2-normed spaces defined by ideal convergence and an Orlicz function. Applied Mathematics and Computation, 217(1):271–276, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006168.

Savas:2011:IML

[Sav11] Ekrem Savas. On infinite matrices and lacunary σ-convergence. Applied Mathematics and Computation, 218(3):1036–1040, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004516.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Sturdevant:2010:TSS

[SB10] Benjamin J. M. Sturdevant and Anjan Biswas. Topological 1-soliton solution of the generalized KdV equation with generalized evolution. Applied Mathematics and Computation, 217(5):2289–2294, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310007393.

Suarez:2011:ESS

[SB11] Pablo Suarez and Anjan Biswas. Exact 1soliton solution of the Zakharov equation in plasmas with power law nonlinearity. Applied Mathematics and Computation, 217(17):7372–7375, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002268.

Sorgun:2012:NUB

[SB12a] Sezer Sorgun and Serife Büyükköse. The new upper bounds on the spectral radius of weighted graphs. Applied Mathematics and Computation, 218(9):5231–5238, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013361.

Srivastava:2012:NPC

[SB12b] H. M. Srivastava and Serap Bulut. Neighborhood properties of certain classes of multivalently analytic functions associated with the convolution structure. Applied Mathematics and Computation, 218(11):6511–6518, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300311014810.

Salkuyeh:2013:SCM

[SB13] Davod Khojasteh Salkuyeh and Mehdi Bastani. Solution of the Complex Modified Korteweg–de Vries equation by the projected differential transform method. Applied Mathematics and Computation, 219(10):5105–5112, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012295.

Sunhaloo:2013:SHB

[SBDB13] Mohammad Sameer Sunhaloo, Ravindra Boojhawon, Kumar Dookhitram, and Muddun Bhuruth. Skew-Hermitian based iterations for nine-point approximations of convectiondiffusion problems. Applied Mathematics and Computation, 219(10): 5384–5396, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312011587.

Somasundaram:2012:MBL

[SBM12] Karthik Somasundaram, Erik Birgersson, and Arun Sadashiv Mujumdar. Model for a bipolar Li-ion battery module: Automated model generation, validation and verification. Applied Mathematics and Computation, 219(4):2231–2245, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008636.

Sengupta:2012:SWD

[SBR^{+}12] Tapan K. Sengupta, Yogesh G. Bhumkar, Manoj K. Rajpoot, V. K. Suman, and Shakti Saurabh. Spurious waves in discrete computation of wave phenomena and flow problems. Applied Mathematics and Computation, 218(18):9035–9065, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002627.

Sharma:2010:CEM

[SBS10] Rajesh Sharma, R. Bhargava, and I. V. Singh. Combined effect of magnetic field and heat absorption on unsteady free convection and heat transfer flow in a micropolar fluid past a semiinfinite moving plate with viscous dissipation using element free Galerkin method. Applied Mathematics and Computation, 217(1):308–321, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006211.

Stelle:2011:UDM

[SBS11] Diogo Stelle, Maria C.

Barioni, and Luis P. Scott. Using data mining to identify structural rules in proteins. Applied Mathematics and Computation, 218(5):1997–2004, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009301.

Shen:2010:BNC

[SC10a] Shouqiang Shen and Jianmiao Cen. On the bounds for the norms of r-circulant matrices with the Fibonacci and Lucas numbers. Applied Mathematics and Computation, 216(10):2891–2897, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310004066.

Song:2010:SWS

[SC10b] Ming Song and Jionghui Cai. Solitary wave solutions and kink wave solutions for a generalized Zakharov–Kuznetsov equation. Applied Mathematics and Computation, 217(4):1455–1462, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630030900558X.

Sheng:2011:NPT

[SC11] Xingping Sheng and Guoliang Chen. New proofs of two representations and minor of generalized inverse A_{T,S}^{(2)}. Applied Mathematics and Computation, 217(13):6309–6314, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000075.

Srivastava:2012:GFC

[SC12] Rekha Srivastava and Nak Eun Cho. Generating functions for a certain class of incomplete hypergeometric polynomials. Applied Mathematics and Computation, 219(6):3219–3225, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009587.

Sun:2013:SAS

[SC13] Jian Sun and Jie Chen.

Stability analysis of static recurrent neural networks with interval time-varying delay. Applied Mathematics and Computation, 221(??):111–120, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006413.

Samanta:2014:EAP

[SC14a] S. Samanta and J. Chattopadhyay. Effect of awareness program in disease outbreak — a slowfast dynamics. Applied Mathematics and Computation, 237(??):98–109, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004809.

Srivastava:2014:SEP

[SC14b] Rekha Srivastava and Nak Eun Cho. Some extended Pochhammer symbols and their applications involving generalized hypergeometric polynomials. Applied Mathematics and Computation, 234(??):277–285, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002756.

Shih:2011:EFF

[SCC11] Yintzer Shih, Jun-Yong Cheng, and Kuen-Tsann Chen. An exponentialfitting finite element method for convection–diffusion problems. Applied Mathematics and Computation, 217(12):5798–5809, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012646.

Shih:2012:APB

[SCC12] Y.-T. Shih, C.-S. Chien, and C.-Y. Chuang. An adaptive parameterized block-based singular value decomposition for image de-noising and compression. Applied Mathematics and Computation, 218 (21):10370–10385, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300312003281.

Chang:2011:MBI

[sCCL11] Shih sen Chang, Chi Kin Chan, and H. W. Joseph Lee. Modified block iterative algorithm for Quasi-ϕ-asymptotically nonexpansive mappings and equilibrium problem in Banach spaces. Applied Mathematics and Computation, 217(18):7520–7530, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002645.

Chang:2010:SME

[sCCLY10] Shih sen Chang, Chi Kin Chan, H. W. Joseph Lee, and Li Yang. A system of mixed equilibrium problems, fixed point problems of strictly pseudo-contractive mappings and nonexpansive semi-groups. Applied Mathematics and Computation, 216(1):51–60, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309011175.

Chang:2013:SCT

[sCCYW13] Shih sen Chang, Chi Kin Chan, Li Yang, and Xiong Rui Wang. Strong convergence theorems of total quasi-φ-asymptotically nonexpansive semi-groups in Banach spaces. Applied Mathematics and Computation, 224(??):398–404, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009260.

Schneider:2011:SPC

[Sch11a] Baruch Schneider. Some properties of the Clifford Cauchy type integrals associated to Helmholtz equation on a piecewise Lyapunov surfaces in R^{m}. Applied Mathematics and Computation, 218(8):4268–4275, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031101229X.

Shen:2011:DIC

[SCH11b] Shou-Qiang Shen, Jian-Miao Cen, and Yong Hao. On the determinants and inverses of circulant matrices with Fibonacci and Lucas numbers. Applied Mathematics and Computation, 217(23):9790–9797, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006424.

Stanimirovic:2011:FRR

[SCIMM11] Predrag S. Stanimirović, Dragana S. Cvetković-Ilić, Sladjana Miljković, and Marko Miladinović. Fullrank representations of {2, 4}, {2, 3}-inverses and successive matrix squaring algorithm. Applied Mathematics and Computation, 217(22):9358–9367, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005819.

Solin:2013:TAB

[SCK13] Pavel Solin, Ondrej Certik, and Lukas Korous. Three anisotropic benchmark problems for adaptive finite element methods. Applied Mathematics and Computation, 219(13):7286–7295, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031001297X.

Srivastava:2014:CGP

[SÇK14a] H. M. Srivastava, Aysegül Çetinkaya, and I. Onur Kiymaz. A certain generalized Pochhammer symbol and its applications to hypergeometric functions. Applied Mathematics and Computation, 226(??):484–491, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010898.

Sun:2014:ABN

[SCK14b] Leping Sun, Yuhao Cong, and Jiaoxun Kuang. Asymptotic behavior of nonlinear delay differential-algebraic equations and implicit Euler methods. Applied Mathematics and Computation, 228(??):395–403, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012678.

Sun:2010:IHS

[SCLC10] Wen Sun, Zhong Chen, Yibing Lü, and Shihua Chen. An intriguing hybrid synchronization phenomenon of two coupled complex networks. Applied Mathematics and Computation, 216(8):2301–2309, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003206.

Chang:2011:SCT

[sCLCaL11] Shih sen Chang, H. W. Joseph Lee, Chi Kin Chan, and Jing ai Liu. Strong convergence theorems for countable families of asymptotically relatively nonexpansive mappings with applications. Applied Mathematics and Computation, 218(7):3187–3198, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010952.

Chang:2011:ATT

[sCLCY11] Shih sen Chang, H. W. Joseph Lee, Chi Kin Chan, and Li Yang. Approximation theorems for total quasi-ϕ-asymptotically nonexpansive mappings with applications. Applied Mathematics and Computation, 218(6):2921–2931, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010629.

Sana:2014:TLS

[SCN14] Shib Sankar Sana, Jaime Acevedo Chedid, and Katherinne Salas Navarro. A three layer supply chain model with multiple suppliers, manufacturers and retailers for multiple items. Applied Mathematics and Computation, 229(??):139–150, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313012800.

Chang:2012:SCT

[sCWT^{+}12] Shih sen Chang, Lin Wang, Yong-Kun Tang, Bin Wang, and Li-Juan Qin. Strong convergence theorems for a countable family of quasiϕ-asymptotically nonexpansive nonself mappings. Applied Mathematics and Computation, 218(15):7864–7870, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001099.

Chang:2014:ICT

[sCWW^{+}14] Shih sen Chang, Gang Wang, Lin Wang, Yong Kun Tang, and Zhao Li Ma. Δconvergence theorems for multi-valued nonexpansive mappings in hyperbolic spaces. Applied Mathematics and Computation, 249(??):535–540, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014519.

Chang:2010:SCT

[sCYW10] Shih sen Chang, Li Yang, and Xiong Rui Wang. Stronger convergence theorems for an infinite family of uniformly quasi-Lipschitzian mappings in convex metric spaces. Applied Mathematics and Computation, 217(1):277– 282, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031000617X.

Shi:2010:EES

[SD10] Peilin Shi and Lingzhen Dong. Existence and exponential stability of anti-periodic solutions of Hopfield neural networks with impulses. Applied Mathematics and Computation, 216(2):623–630, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001190.

Srivastava:2013:EMF

[SD13a] Bal Govind Srivastava and Satya Deo. Effect of magnetic field on the viscous fluid flow in a channel filled with porous medium of variable permeability. Applied Mathematics and Computation, 219(17):8959–8964, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003184.

Suleimanov:2013:ADW

[SD13b] B. A. Suleimanov and O. A. Dyshin. Application of discrete wavelet transform to the solution of boundary value problems for quasi-linear parabolic equations. Applied Mathematics and Computation, 219(12):7036–7047, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012003.

Subasi:2014:SSO

[SD14] Murat Subasi and Hülya Durur. On the stability of the solution in an optimal control problem for a Schrödinger equation. Applied Mathematics and Computation, 249(??):521–526, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014441.

Salhin:2014:SOC

[SDAN14] Ambarka Abdalla Salhin, Ummul Khair Salma Din, Rokiah Rozita Ahmad, and Mohd Salmi Md Noorani. Some oscillation criteria for a class of second order nonlinear damped differential equations. Applied Mathematics and Computation, 247(??):962–968, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012855.

Sauter:2014:ETO

[SdAT14] Esequia Sauter, Fabio Souto de Azevedo, and Mark Thompson. Existence theory for one-dimensional quasilinear coupled conductiveradiative flows. Applied Mathematics and Computation, 233(??):545–556, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002574.

Sportelli:2014:DLM

[SDB14] Mario Sportelli, Luigi De Cesare, and Maria T. Binetti. A dynamic IS– LM model with two time delays in the public sector. Applied Mathematics and Computation, 243(??):728–739, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400887X.

Santos:2010:PDF

[SDC10] L. T. Santos, F. A. Dorini, and M. C. C. Cunha. The probability density function to the random linear transport equation. Applied Mathematics and Computation, 216(5):1524–1530, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002432.

Shin:2010:CNK

[SDK10] Byeong-Chun Shin, M. T.

Darvishi, and Chang-Hyun Kim. A comparison of the Newton– Krylov method with high order Newton-like methods to solve nonlinear systems. Applied Mathematics and Computation, 217(7):3190–3198, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000915X.

Shao:2010:DIO

[SDL10a] Yuanfu Shao, Binxiang Dai, and Zhenguo Luo. The dynamics of an impulsive one-prey multi-predators system with delay and Hollingtype II functional response. Applied Mathematics and Computation, 217(6):2414–2424, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007873.

Shi:2010:CTW

[SDL10b] Yeqiong Shi, Zhengde Dai, and Donglong Li. The correct traveling wave solutions for the highorder dispersive nonlinear Schrödinger equation. Applied Mathematics and Computation, 216(5):1583–1591, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002535.

Singh:2014:CLP

[SDN14] Garima Singh, Kusum Deep, and Atulya K. Nagar. Cell-like P -systems based on rules of Particle Swarm Optimization. Applied Mathematics and Computation, 246(??):546–560, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314011217.

Srivastava:2010:SGU

[SDO10] H. M. Srivastava, Erhan Deniz, and Halit Orhan. Some general univalence criteria for a family of integral operators. Applied Mathematics and Computation, 215(10):3696–3701, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009825.

Stevic:2014:PSC

[SDS14] Stevo Stević, Josef Diblík, and Zdenek Smarda. On periodic and solutions converging to zero of some systems of differential-difference equations. Applied Mathematics and Computation, 227(??):43–49, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011120.

Sangwine:2012:CHD

[SE12] Stephen J. Sangwine and Todd A. Ell. Complex and hypercomplex discrete Fourier transforms based on matrix exponential form of Euler’s formula. Applied Mathematics and Computation, 219(2):644–655, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006637.

Seckiner:2013:ACO

[SEED13] Serap Ulusam Seçkiner, Yunus Eroglu, Merve Emrullah, and Türkay Dereli. Ant colony optimization for continuous functions by using novel pheromone updating. Applied Mathematics and Computation, 219(9):4163–4175, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011277.

Segura:2014:MPB

[Seg14] Javier Segura. Monotonicity properties and bounds for the chi-square and gamma distributions. Applied Mathematics and Computation, 246(??):399–415, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031401128X.

Seker:2011:CNS

[Sek11] Bilal Seker. On certain new subclass of close-to-convex functions. Applied Mathematics and Computation, 218(3):1041–1045, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003729.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Stehlik:2014:KLL

[SEKR14] M. Stehlík, P. Economou, J. Kiselák, and W.-D. Richter. Kullback–Leibler life time testing. Applied Mathematics and Computation, 240(??):122–139, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005608.

Sherief:2014:AFO

[SEL14] Hany H. Sherief and A. M. Abd El-Latief. Application of fractional order theory of thermoelasticity to a 2D problem for a half-space. Applied Mathematics and Computation, 248(??):584–592, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031401385X.

Sogabe:2011:FBD

[SEM11] Tomohiro Sogabe and Moawwad El-Mikkawy. Fast block diagonalization of k-tridiagonal matrices. Applied Mathematics and Computation, 218(6):2740–2743, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031101040X.

Sergeyev:2013:SOD

[Ser13] Yaroslav D. Sergeyev. Solving ordinary differential equations on the infinity computer by working with infinitesimals numerically. Applied Mathematics and Computation, 219(22):10668–10681, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004190.

Setia:2014:NSV

[Set14] Amit Setia. Numerical solution of various cases of Cauchy type singular integral equation. Applied Mathematics and Computation, 230(??):200–207, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014045.

Sever:2011:PDM

[Sev11] Ali Sever. A problem of data mining in Ecommerce. Applied Mathematics and Computation, 217(24):9966–9970, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006278.

Sever:2013:NNA

[Sev13] Ali Sever. A neural network algorithm to pattern recognition in inverse problems. Applied Mathematics and Computation, 221(??):484–490, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007431.

Salas:2011:CSS

[SF11a] Alvaro H. Salas and FIZ-MAKO Research Group.

Construction of N-soliton solutions to (2 + 1)- dimensional Konopelchenko– Dubrovsky (KD) equations. Applied Mathematics and Computation, 217(18):7391–7399, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002037.

Simsek:2011:AMA

[SF11b] Burcin Simsek and Mehmet Ziya Firat. Application of multilevel analysis in animal sciences. Applied Mathematics and Computation, 218(3):1067–1071, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004577.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Skiba:2013:SBS

[SF13] Yuri N. Skiba and Denis M. Filatov. Splittingbased schemes for numerical solution of nonlinear diffusion equations on a sphere. Applied Mathematics and Computation, 219(16):8467–8485, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300204X.

Silva:2011:ETU

[SFF11] J. M. Silva, V. G. Ferreira, and S. R. Fontes. An evaluation of three upwinding approximations for numerical modeling the flow in tubular membrane of Newtonian and non-Newtonian fluids. Applied Mathematics and Computation, 217(20):7955–7965, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003043.

Sanchez:2011:PWR

[SFL^{+}11] A. Suárez Sánchez, P. Riesgo Fernández, F. Sánchez Lasheras, F. J. de Cos Juez, and P. J. García Nieto. Prediction of workrelated accidents according to working conditions using support vector machines. Applied Mathematics and Computation, 218(7):3539–3552, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311011416.

Sun:2011:QBP

[SFP^{+}11] Jun Sun, Wei Fang, Vasile Palade, Xiaojun Wu, and Wenbo Xu. Quantum-behaved particle swarm optimization with Gaussian distributed local attractor point. Applied Mathematics and Computation, 218(7):3763–3775, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011878.

Suarez:2010:GMP

[SG10] Antonio Suárez and Luis González. A generalization of the Moore– Penrose inverse related to matrix subspaces of C^{n×m}. Applied Mathematics and Computation, 216(2):514–522, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000755.

Safi:2011:QSQ

[SG11] Mohammad A. Safi and Abba B. Gumel. Qualitative study of a quarantine/isolation model with multiple disease stages. Applied Mathematics and Computation, 218(5):1941–1961, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100926X.

Srivastava:2014:NR

[SG14] Shwetabh Srivastava and Dharmendra K. Gupta. A new representation for A_{T,S}^{(2,3)}. Applied Mathematics and Computation, 243(??):514–521, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008467.

Sakthivel:2013:ACF

[SGA13] R. Sakthivel, R. Ganesh, and S. M. Anthoni. Approximate controllability of fractional nonlinear differential inclusions. Applied Mathematics and Computation, 225(??):708–717, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313010473.

Soltanalizadeh:2014:SAS

[SGA14] B. Soltanalizadeh, H. Roohani Ghehsareh, and S. Abbasbandy. A super accurate shifted Tau method for numerical computation of the Sobolevtype differential equation with nonlocal boundary conditions. Applied Mathematics and Computation, 236(??):683–692, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004068.

Singh:2010:CTF

[SGC10] J. Singh, S. K. Gupta, and Srinivasan Chandrasekaran. Computational treatment of free convection effect on flow of elastico-viscous fluid past an accelerated plate with constant heat flux. Applied Mathematics and Computation, 217(2):685–688, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000679X.

Sarkar:2012:OIR

[SGC12] T. Sarkar, S. K. Ghosh, and K. S. Chaudhuri. An optimal inventory replenishment policy for a deteriorating item with timequadratic demand and time-dependent partial backlogging with shortages in all cycles. Applied Mathematics and Computation, 218(18):9147–9155, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002226.

Sarkar:2014:IIM

[SGCG14] Biswajit Sarkar, Hiranmoy Gupta, Kripasindhu Chaudhuri, and Suresh Kumar Goyal. An integrated inventory model with variable lead time, defective units and delay in payments. Applied Mathematics and Computation, 237(??):650–658, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004238.

Sandeep:2011:WBS

[SGDK11] K. Sandeep, Shikha Gaur, D. Dutta, and H. S. Kushwaha. Wavelet based schemes for linear advection-dispersion equation. Applied Mathematics and Computation, 218(7):3786–3798, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011891.

Safi:2013:QAA

[SGE13] Mohammad A. Safi, Abba B. Gumel, and Elamin H. Elbasha. Qualitative analysis of an age-structured SEIR epidemic model with treatment. Applied Mathematics and Computation, 219(22):10627–10642, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003871.

Sharma:2012:SED

[SGG12] Janak Raj Sharma, Rangan K. Guha, and Puneet Gupta. Some efficient derivative free methods with memory for solving nonlinear equations. Applied Mathematics and Computation, 219(2):699–707, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006704.

Salukvadze:2010:MPQ

[SGJ10] M. Salukvadze, R. Gogsadze, and N. Jibladze. Metric properties of quality criteria space for multicriteria optimization problems. Applied Mathematics and Computation, 217(3):1220–1230, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004522.

Singh:2011:BBM

[SGK11] K. Singh, R. K. Gupta, and Sachin Kumar. Benjamin– Bona–Mahony (BBM) equation with variable coefficients: Similarity reductions and Painlevé analysis. Applied Mathematics and Computation, 217(16):7021–7027, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001603.

Singh:2014:CSD

[SGK14] Mehakpreet Singh, Dhrubajyoti Ghosh, and Jitendra Kumar. A comparative study of different discretizations for solving bivariate aggregation population balance equation. Applied Mathematics and Computation, 234(??):434–451, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002914.

Srivastava:2013:CSQ

[SGN13] H. M. Srivastava, B. J. González, and E. R. Negrin. A characterization of the Second Quantization by using the Segal duality transform. Applied Mathematics and Computation, 219(11):6236–6240, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012477.

Shinar:2010:RPH

[SGT10] Josef Shinar, Valery Y. Glizer, and Vladimir Turetsky. Robust pursuit of a hybrid evader. Applied Mathematics and Computation, 217(3):1231–1245, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000439X.

Shen:2014:BPA

[SGY^{+}14] Yu-Jia Shen, Yi-Tian Gao, Xin Yu, Gao-Qing Meng, and Yi Qin. Bellpolynomial approach applied to the seventh-order Sawada–Kotera–Ito equation. Applied Mathematics and Computation, 227(??):502–508, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031301165X.

Sheng:2013:GIC

[SGZC13] Yuqiu Sheng, Yanling Ge, Hanyu Zhang, and Chongguang Cao. Group inverses for a class of 2 × 2 block matrices over rings. Applied Mathematics and Computation, 219(17):9340–9346, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300177X.

Shao:2012:EES

[SH12a] Xinping Shao and Danfu Han. Error estimate for a simple two-level discretization of stream function form of the Navier– Stokes equations. Applied Mathematics and Computation, 218(19):9943–9952, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003396.

Sun:2012:FOH

[SH12b] Yuangong Sun and Zhenlai Han. On forced oscillation of higher order functional differential equations. Applied Mathematics and Computation, 218(12):6966–6971, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015682.

Sun:2013:SND

[SH13] Yuangong Sun and Taher Hassan. Some nonlinear dynamic integral inequalities on time scales. Applied Mathematics and Computation, 220(??):221–225, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300313006759.

Shang:2010:ESC

[Sha10a] Desheng Shang. Exact solutions of coupled nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 217(4):1577–1583, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309006390.

Sharma:2010:CBM

[Sha10b] Shambhu N. Sharma. A connection between multilinear and Volterra systems. Applied Mathematics and Computation, 216(7):1918–1922, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000144X.

Shakhmurov:2011:AEE

[Sha11a] Veli B. Shakhmurov.

Anisotropic elliptic equations with VMO coefficients. Applied Mathematics and Computation, 218(3):1057–1062, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004553.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Shao:2011:ESP

[Sha11b] Yuanfu Shao. Exponential stability of periodic neural networks with impulsive effects and timevarying delays. Applied Mathematics and Computation, 217(16):6893–6899, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001032.

Sarfraz:2012:PPS

[SHA12a] Muhammad Sarfraz, Malik Zawwar Hussain, and Muhammad Arfan Ali. Positivity-preserving scattered data interpolation scheme using the sidevertex method. Applied Mathematics and Computation, 218(15):7898–7910, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001233.

Sharma:2012:EVD

[Sha12b]  Rajesh Sharma.  Ef-

fect of viscous dissipation and heat source on unsteady boundary layer flow and heat transfer past a stretching surface embedded in a porous medium using element free Galerkin method. Applied Mathematics and Computation, 219(3):976–987, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007023.

See comment [Pan15].

Shampine:2013:EFM

[Sha13a] L. F. Shampine. Efficient Filon method for oscillatory integrals. Applied Mathematics and Computation, 221(??):691–702, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313007340.

Shang:2013:MSR

[Sha13b]  Yilun Shang.  Mixed

SI(R) epidemic dynamics in random graphs with general degree distributions. Applied Mathematics and Computation, 219(10):5042–5048, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312011939.

Shamseldeen:2014:IPP

[Sha14a] Samir A. Shamseldeen. Instabilities and propagation properties in a fourthorder reaction-diffusion equation. Applied Mathematics and Computation, 231(??):395–406, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314000721.

Shang:2014:BPS

[Sha14b] Desheng Shang. Bifurcation of a piecewise smooth cubic system via expansion of Melnikov function. Applied Mathematics and Computation, 242(??):510–517, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007929.

Shang:2014:CTA

[Sha14c] Yilun Shang. Continuoustime average consensus under dynamically changing topologies and multiple time-varying delays. Applied Mathematics and Computation, 244(??):457–466, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400976X.

Shehzad:2014:NTR

[SHAO14] Sabir Ali Shehzad, Tasawar Hayat, Ahmed Alsaedi, and Mustafa Ali Obid. Nonlinear thermal radiation in threedimensional flow of Jeffrey nanofluid: a model for solar energy. Applied Mathematics and Computation, 248(??):273–286, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013162.

Shekari:2013:ROM

[SHBN13] Younes Shekari, Ebrahim Hajidavalloo, and Morteza Behbahani-Nejad. Reduced order modeling of transient two-phase flows and its application to upward two-phase flows in the under-balanced drilling. Applied Mathematics and Computation, 224(??):775–790, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009958.

Song:2011:SWS

[SHC11] Ming Song, Xuejiong Hou, and Jun Cao. Solitary wave solutions and kink wave solutions for a generalized KDV–mKDV equation. Applied Mathematics and Computation, 217(12):5942–5948, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310013263.

Shehata:2010:TWS

[She10]  A. R. Shehata.  The

traveling wave solutions of the perturbed nonlinear Schrödinger equation and the cubic– quintic Ginzburg Landau equation using the modified (\frac{G′}{G})-expansion method. Applied Mathematics and Computation, 217(1):1–10, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002894.

Shehu:2012:SCT

[She12a] Yekini Shehu. Strong convergence theorems for infinite family of relatively quasi nonexpansive mappings and systems of equilibrium problems. Applied Mathematics and Computation, 218(9):5146–5156, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013282.

Shen:2012:GSN

[She12b] Wenguo Shen. Global structure of nodal solutions for a fourth-order two-point boundary value problem. Applied Mathematics and Computation, 219(1):88–98, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015700.

Sherratt:2012:NCM

[She12c] Jonathan A. Sherratt.

Numerical continuation methods for studying periodic travelling wave (wavetrain) solutions of partial differential equations. Applied Mathematics and Computation, 218(9):4684–4694, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101335X.

Shehu:2014:CTM

[She14a] Yekini Shehu. Convergence theorems for maximal monotone operators and fixed point problems in Banach spaces. Applied Mathematics and Computation, 239(??):285–298, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006237.

Shehu:2014:SCT

[She14b] Yekini Shehu. Strong convergence theorem for integral equations of Hammerstein type in Hilbert spaces. Applied Mathematics and Computation, 231(??):140–147, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313014471.

Shen:2014:NPP

[She14c] Shu-Qian Shen. A note on PSS preconditioners for generalized saddle point problems. Applied Mathematics and Computation, 237(??):723–729, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005220.

Shieh:2010:NPM

[Shi10] Bih-Sheue Shieh. A note on a paper by Molai and Khorram. Applied Mathematics and Computation, 216(11):3419–3422, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004790.

Shi:2012:CIS

[Shi12] Weichen Shi. Conservation integrals in the sense of Noether’s theorem for an analytic function on a physical plane and application. Applied Mathematics and Computation, 219(6):3009–3016, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009277.

Shiri:2013:NUD

[Shi13a] B. Shiri. A note on using the Differential Transformation Method for the Integro–Differential Equations. Applied Mathematics and Computation, 219(14):7306–7309, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001124.

Singh:2013:GSR

[SHI13b] Jitender Singh, Evor Hines, and Daciana Iliescu. Global stability results for temperature modulated convection in ferrofluids. Applied Mathematics and Computation, 219(11):6204–6211, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012441.

Shi:2014:BNS

[Shi14a] Weichen Shi. A brief note on the self-balanced singular elastic field caused by an edge dislocation. Applied Mathematics and Computation, 226(??): 194–197, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313011168.

Shi:2014:OCT

[Shi14b] Yanxiang Shi. Oscillation criteria for n th order nonlinear neutral differential equations. Applied Mathematics and Computation, 235(??):423–429, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003713.

Shklyar:2014:ENC

[Shk14] B. Shklyar. Exact nullcontrollability of evolution equations by smooth scalar distributed controls and applications to controllability of interconnected systems. Applied Mathematics and Computation, 238(??):444–459, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400455X.

Said-Houari:2010:SRT

[SHL10] Belkacem Said-Houari and Yamina Laskri. A stability result of a Timoshenko system with a delay term in the internal feedback. Applied Mathematics and Computation, 217(6):2857–2869, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008751.

Sun:2012:PSC

[SHM12] Xiao-Ke Sun, Hai-Feng Huo, and Cao-Chuan Ma. Periodic solutions of a class of impulsive neutral delay differential equation. Applied Mathematics and Computation, 219(8):3947–3955, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010326.

Sun:2013:HBS

[SHM13] Xiao-Ke Sun, Hai-Feng Huo, and Cao-Chuan Ma. Hopf bifurcation and stability in predatorprey model with a stagestructure for prey. Applied Mathematics and Computation, 219(20):10313– 10324, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313004013.

Salinas-Hernandez:2012:NGS

[SHMCM12] E. Salinas-Hernández, J. Martínez-Castro, and R. Muñoz. New general solutions to the Abel equation of the second kind using functional transformations. Applied Mathematics and Computation, 218(17):8359–8362, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312001105.

Sarfraz:2010:PDM

[SHN10] Muhammad Sarfraz, Malik Zawwar Hussain, and Asfar Nisar. Positive data modeling using spline function. Applied Mathematics and Computation, 216(7):2036–2049, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002766.

Shoufu:2014:CTR

[Sho14] Li Shoufu. Classical theory of Runge–Kutta methods for Volterra functional differential equations. Applied Mathematics and Computation, 230 (??):78–95, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313013805.

Sun:2013:PMB

[SHSC13] Jianjun Sun, Yan Huang, Shuyao Sun, and Lihong Cui. Parameterizations of masks for 3-band tight wavelet frames by symmetric extension of polyphase matrix. Applied Mathematics and Computation, 225(??):461–474, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010448.

Shtern:2011:HKL

[Sht11] Alexander I. Shtern. Hochschild kernel for locally bounded finite-dimensional representations of a connected Lie group. Applied Mathematics and Computation, 218(3):1063–1066, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311000907.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Saad:2014:PSC

[SHT14] Nasser Saad, Richard L.

Hall, and Victoria A. Trenton. Polynomial solutions for a class of second-order linear differential equations. Applied Mathematics and Computation, 226(??):615–634, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011132.

Shi:2014:PSU

[SHWT14a] Xuerong Shi, Lixin Han, Zuolei Wang, and Keming Tang. Pinning synchronization of unilateral coupling neuron network with stochastic noise. Applied Mathematics and Computation, 232(??):1242–1248, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001714.

Shi:2014:SDB

[SHWT14b] Xuerong Shi, Lixin Han, Zuolei Wang, and Keming Tang. Synchronization of delay bursting neuron system with stochastic noise via linear controllers. Applied Mathematics and Computation, 233(??):232–242, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002422.

Sun:2014:GBM

[SHWX14] Taixiang Sun, Qiuli He, Xin Wu, and Hongjian Xi. Global behavior of the max-type difference equation x_{n} = max\{\frac{1}{x_{n−m}} , \frac{A_{n}}{x_{n−r}} \}. Applied Mathematics and Computation, 248(??): 687–692, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013848.

Shang:2011:BTA

[SHY11] Yadong Shang, Yong Huang, and Wenjun Yuan. Bäcklund transformations and abundant exact explicit solutions of the Sharma–Tasso–Olver equation. Applied Mathematics and Computation, 217(17):7172–7183, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001573.

Song:2010:VMM

[SHZ10] Lina Song, Yanren Hou, and Haibiao Zheng. A variational multiscale method based on bubble functions for convection-dominated convection–diffusion equation. Applied Mathematics and Computation, 217(5):2226–2237, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000768X.

Song:2014:SHB

[SHZ14] Yongli Song, Yanyan Han, and Tonghua Zhang. Stability and Hopf bifurcation in a model of gene expression with distributed time delays. Applied Mathematics and Computation, 243(??):398–412, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314008224.

San:2011:ODO

[SI11a] M. San and H. Irmak. Ordinary differential operator and some of its applications to certain meromorphically p-valent functions. Applied Mathematics and Computation, 218(3):817–821, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100227X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Serenbay:2011:ACT

[SI11b] S. Kirci Serenbay and E. Ibikli. Approximation by Chlodowsky–Taylor polynomials. Applied Mathematics and Computation, 218(3):878–882, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100453X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Siddiqui:2014:CMD

[SI14] Hafiz Muhammad Afzal Siddiqui and Muhammad Imran. Computing the metric dimension of wheel related graphs. Applied Mathematics and Computation, 242(??):624–632, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400842X.

Sidi:2010:SAA

[Sid10] Avram Sidi. A simple approach to asymptotic expansions for Fourier integrals of singular functions. Applied Mathematics and Computation, 216(11):3378–3385, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005072.

Sastre:2011:EOM

[SIDR11] J. Sastre, J. Ibáñez, E. Defez, and P. Ruiz. Efficient orthogonal matrix polynomial based method for computing matrix exponential. Applied Mathematics and Computation, 217(14):6451–6463, March 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311000087.

Sierra:2010:NES

[Sie10] Cesar A. Gomez Sierra. New exact solutions for a generalization of the Korteweg–de Vries equation (KdV6). Applied Mathematics and Computation, 216(1):357–360, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000536.

Siettos:2011:EFM

[Sie11] Constantinos I. Siettos. Equation-free multiscale computational analysis of individual-based epidemic dynamics on networks. Applied Mathematics and Computation, 218(2):324–336, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100748X.

Siettos:2014:CGC

[Sie14] Constantinos Siettos. Coarsegrained computational stability analysis and acceleration of the collective dynamics of a Monte Carlo simulation of bacterial locomotion. Applied Mathematics and Computation, 232(??):836–847, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001970.

Simsek:2011:CNG

[Sim11]  Yilmaz Simsek.  Con-

struction a new generating function of Bernstein type polynomials. Applied Mathematics and Computation, 218(3):1072–1076, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001093.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Singh:2010:WPI

[Sin10a]  Baljeet Singh.  Wave

propagation in an initially stressed transversely isotropic thermoelastic solid half-space. Applied Mathematics and Computation, 217(2):705–715, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000682X.

Singh:2010:LAS

[Sin10b] Vimal Singh. LMI approach to stability of a class of discrete-time systems using two’s complement arithmetic. Applied Mathematics and Computation, 217(6):2877–2882, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310008775.

Singh:2011:RTC

[Sin11a] S. S. Singh. Reflection and transmission of couple longitudinal waves at a plane interface between two dissimilar half-spaces of thermo-elastic materials with voids. Applied Mathematics and Computation, 218(7):3359–3371, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011192.

Singh:2011:NCS

[Sin11b] Vimal Singh. Novel criterion for stability of discrete-time systems in a state-space realization utilizing saturation nonlinearities. Applied Mathematics and Computation, 218(8):4305–4311, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101232X.

Singh:2012:ESS

[Sin12] Ram Mehar Singh. Exact solutions of the Schrödinger equation for an anharmonic potential in two dimensions. Applied Mathematics and Computation, 218(11):6439–6445, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101472X.

Siddiqi:2014:GFC

[SIR14] Shahid S. Siddiqi, Usama Idrees, and Kashif Rehan. Generation of fractal curves and surfaces using ternary 4-point interpolatory subdivision scheme. Applied Mathematics and Computation, 246(??):210–220, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314010352.

Sastre:2013:ECM

[SIRD13] Jorge Sastre, Javier Ibáñez, Pedro Ruiz, and Emilio Defez. Efficient computation of the matrix cosine. Applied Mathematics and Computation, 219(14):7575–7585, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000830.

Sitarz:2014:MCD

[Sit14] Sebastian Sitarz. Multiple criteria dynamic programming and multiple knapsack problem. Applied Mathematics and Computation, 228(??):598–605, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313012988.

Siu:2010:BPU

[Siu10] Tak Kuen Siu. Bond pricing under a Markovian regime-switching jumpaugmented Vasicek model via stochastic flows. Applied Mathematics and Computation, 216(11): 3184–3190, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310004637.

Stamova:2010:ICC

[SIV10] Ivanka M. Stamova, Rajcho Ilarionov, and Rositsa Vaneva. Impulsive control for a class of neural networks with bounded and unbounded delays. Applied Mathematics and Computation, 216(1):285– 290, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031000072X.

Sabetghadam:2012:RPG

[SJ12] Feriedoun Sabetghadam and Alireza Jafarpour. α Regularization of the POD–Galerkin dynamical systems of the Kuramoto– Sivashinsky equation. Applied Mathematics and Computation, 218(10): 6012–6026, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300311014299.

Singh:2013:IAW

[SJ13] Randheer Singh and J. Jena.

Interaction of an acceleration wave with a strong shock in reacting polytropic gases. Applied Mathematics and Computation, 225(??):638–644, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313010540.

Saito:2010:ECB

[SJL10] Yasuhisa Saito, Changdo Jung, and Yongdo Lim. Effects of cannibalism on a basic stage structure. Applied Mathematics and Computation, 217(5):2133–2141, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007575.

Singh:2014:SSN

[SJPS14] Pushpendra Singh, Shiv Dutt Joshi, Rakesh Kumar Patney, and Kaushik Saha. Some studies on nonpolynomial interpolation and error analysis. Applied Mathematics and Computation, 244(??):809–821, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010066.

Senthil:2010:VBM

[SJR10] G. Senthil, Girija Jayaraman, and A. D. Rao. A variable boundary method for modelling two dimensional free surface flows with moving boundaries. Applied Mathematics and Computation, 216(9):2544–2558, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003504.

Shuster:2010:ECN

[SK10a] John A. Shuster and Jens Köplinger. Elliptic complex numbers with dual multiplication. Applied Mathematics and Computation, 216(12):3497–3514, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005096.

Srivastava:2010:GVV

[SK10b] P. D. Srivastava and Sudhanshu Kumar. Generalized vector-valued paranormed sequence space using modulus function.

Applied Mathematics and Computation, 215(12): 4110–4118, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300309010595.

Sweilam:2010:ASN

[SK10c] N. H. Sweilam and M. M. Khader. Approximate solutions to the nonlinear vibrations of multiwalled carbon nanotubes using Adomian decomposition method. Applied Mathematics and Computation, 217(2):495–505, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006417.

Saewan:2011:CTM

[SK11a] Siwaporn Saewan and Poom Kumam. Convergence theorems for mixed equilibrium problems, variational inequality problem and uniformly quasi-ϕ-asymptotically nonexpansive mappings. Applied Mathematics and Computation, 218(7):3522– 3538, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311011404.

Shuster:2011:DNN

[SK11b] John A. Shuster and Jens Köplinger. Doubly nilpotent numbers in the 2D plane. Applied Mathematics and Computation, 217(17):7295–7310, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311001974.

Salakhitdinov:2012:SBP

[SK12a] M. S. Salakhitdinov and E. T. Karimov. Spatial boundary problem with the Dirichlet–Neumann condition for a singular elliptic equation. Applied Mathematics and Computation, 219(8):3469–3476, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009125.

Srivastava:2012:FSG

[SK12b] P. D. Srivastava and Sudhanshu Kumar. Fine spectrum of the generalized difference operator Δ_{uv} on sequence space l_{1}. Applied Mathematics and Computation, 218(11):6407–6414, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014688.

Sintunavarat:2013:SFP

[SK13a] Wutiphol Sintunavarat and Poom Kumam. Some fixed point results for weakly isotone mappings in ordered Banach spaces. Applied Mathematics and Computation, 224(??): 826–834, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313009740.

Smarda:2013:SIV

[SK13b] Zdenek Smarda and Yasir Khan. Singular initial value problem for a system of integro-differential equations unsolved with respect to the derivative. Applied Mathematics and Computation, 222(??):290–296, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007881.

Saravanan:2014:NCS

[SK14a] S. Saravanan and A. R. Vidhya Kumar. Natural convection in square cavity with heat generating baffles. Applied Mathematics and Computation, 244(??):1–9, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009369.

Shin:2014:SPI

[SK14b] Young-Bin Shin and Eisuke Kita. Search performance improvement of Particle Swarm Optimization by second best particle information. Applied Mathematics and Computation, 246(??):346–354, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011072.

Senthilnath:2014:APM

[SKB14] J. Senthilnath, Naveen P.

Kalro, and J. A. Benediktsson. Accurate point matching based on multiobjective Genetic Algorithm for multi-sensor satellite imagery. Applied Mathematics and Computation, 236(??):546–564, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004329.

Song:2010:IMZ

[SKC10] Yisheng Song, Jung Im Kang, and Yeol Je Cho. On iterations methods for zeros of accretive operators in Banach spaces. Applied Mathematics and Computation, 216(3):1007– 1017, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310001487.

Shieh:2011:MPS

[SKC11] Horng-Lin Shieh, Cheng-Chien Kuo, and Chin-Ming Chiang. Modified particle swarm optimization algorithm with simulated annealing behavior and its numerical verification. Applied Mathematics and Computation, 218(8):4365–4383, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012422.

Solis:2014:GPA

[SKC14] Francisco J. Solis and Roberto A. Ku-Carrillo. Generic predation in age structure predator-prey models. Applied Mathematics and Computation, 231(??):205–213, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014367.

Soheili:2012:ANM

[SKD12] Ali R. Soheili, Asghar Kerayechian, and Noshin Davoodi. Adaptive numerical method for Burgerstype nonlinear equations. Applied Mathematics and Computation, 219(8):3486– 3495, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312009186.

Shyur:2012:UEH

[SKIH12] Huan-Jyh Shyur, Hwa Keng, Ia-Ka I., and Cheng-Lung Huang. Using extended hazard regression model to assess the probability of aviation event. Applied Mathematics and Computation, 218(21):10647– 10655, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004286.

Srivastava:2011:IPS

[SKM11] H. M. Srivastava, S. M. Khairnar, and Meena More. Inclusion properties of a subclass of analytic functions defined by an integral operator involving the Gauss hypergeometric function. Applied Mathematics and Computation, 218(7):3810–3821, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011957.

Sathananthan:2012:RSS

[SKSK12] S. Sathananthan, M. J. Knap, A. Strong, and L. H. Keel. Robust stability and stabilization of a class of nonlinear discrete time stochastic systems: an LMI approach. Applied Mathematics and Computation, 219(4):1988–1997, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300312008338.

Sun:2011:PUS

[SKW11] Yong Sun, Wei-Ping Kuang, and Zhi-Gang Wang. Properties for uniformly starlike and related functions under the Srivastava– Attiya operator. Applied Mathematics and Computation, 218(7):3615–3623, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311011532.

Liu:2010:EHA

[sL10a]  Cheng shi Liu.  The

essence of the homotopy analysis method. Applied Mathematics and Computation, 216(4):1299–1303, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001906.

Song:2010:ILM

[SL10b] M. H. Song and X. Liu.

The improved linear multistep methods for differential equations with piecewise continuous arguments. Applied Mathematics and Computation, 217(8):4002–4009, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010246.

Sanchez-Lopez:2011:ASC

[SL11a] C. Sánchez-López. Automatic synthesis of chaotic attractors. Applied Mathematics and Computation, 217(9):4350–4358, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310011215.

Song:2011:SBC

[SL11b] Xinyu Song and An Li. Stability and boundedness criteria of nonlinear impulsive systems employing perturbing Lyapunov functions. Applied Mathematics and Computation, 217(24):10166–10174, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311006667.

Shao:2013:DAS

[SL13a] Yuanfu Shao and Ying Li. Dynamical analysis of a stage structured predator-prey system with impulsive diffusion and generic functional response. Applied Mathematics and Computation, 220(??):472–481, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313006188.

Sun:2013:DFC

[SL13b] Weihua Sun and Haiyan Lv. Dissipative feedback control and synchronization of Julia sets of the complex standard family. Applied Mathematics and Computation, 219(15):8026–8032, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001549.

Settati:2014:SDS

[SL14a] Adel Settati and Aadil Lahrouz. Stationary distribution of stochastic population systems under regime switching. Applied Mathematics and Computation, 244(??):235–243, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009692.

Singh:2014:HOV

[SL14b] Swarn Singh and Ping Lin. High order variable mesh off-step discretization for the solution of 1-D non-linear hyperbolic equation. Applied Mathematics and Computation, 230(??):629–638, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313014343.

Song:2014:PWS

[SL14c] Ming Song and Zhengrong Liu. Periodic wave solutions and their limits for the ZK–BBM equation. Applied Mathematics and Computation, 232(??):9– 26, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031400085X.

Song:2014:BBT

[SL14d] Xianfa Song and Xiaoshuang Lv. Bounds for the blowup time and blowup rate estimates for a type of parabolic equations with weighted source. Applied Mathematics and Computation, 236(??):78–92, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003762.

Shakuntla:2011:RAP

[SLBS11] S. Shakuntla, A. K. Lal, S. S. Bhatia, and Jai Singh. Reliability analysis of Polytube industry using supplementary variable technique. Applied Mathematics and Computation, 218(8):3981–3992, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101246X.

Shen:2012:NTV

[SLC^{+}12] S. Shen, F. Liu, J. Chen, I. Turner, and V. Anh. Numerical techniques for the variable order time fractional diffusion equation. Applied Mathematics and Computation, 218(22):10861–10870, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312004468.

Si:2013:PSD

[SLJ^{+}13] Xiaoling Si, Liang Liang, Guozhu Jia, Li Yang, Huaqing Wu, and Yongjun Li. Proportional sharing and DEA in allocating the fixed cost. Applied Mathematics and Computation, 219(12):6580–6590, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000398.

Sun:2014:EBS

[SLL14] Wei Sun, Shiyong Li, and Quan-Lin Li. Equilibrium balking strategies of customers in Markovian queues with two-stage working vacations. Applied Mathematics and Computation, 248(??):195–214, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013502.

Saraiva:2014:MMU

[SLM14] Erlandson F. Saraiva, Francisco Louzada, and Luis Milan. Mixture models with an unknown number of components via a new posterior split-merge MCMC algorithm. Applied Mathematics and Computation, 244(??):959–975, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009898.

Song:2012:HOP

[SLW12] Heping Song, Jun Liu, and Guoli Wang. High-order parameter approximation for von Mises–Fisher distributions. Applied Mathematics and Computation, 218(24):11880–11890, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005656.

Sun:2013:NIP

[SLYJ13] Gui-Quan Sun, Jing Li, Benguo Yu, and Zhen Jin. Noise induced pattern transition in a vegetation model. Applied Mathematics and Computation, 221(??):463–468, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313007212.

Su:2011:NMH

[SLZ11] Yongfu Su, Mengqin Li, and Hong Zhang. New monotone hybrid algorithm for hemi-relatively nonexpansive mappings and maximal monotone operators. Applied Mathematics and Computation, 217(12):5458–5465, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012154.

Shen:2012:LST

[SLZ12] Hua Shen, Kaixin Liu, and Deliang Zhang. A local space–time conservation scheme and its application in shock wave propagation. Applied Mathematics and Computation, 219(4):1958–1974, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008314.

Sun:2006:NPD

[SM06] Yuan Gong Sun and Fan Wei Meng. Note on the paper of Džurina and Stavroulakis. Applied Mathematics and Computation, 174(2):1634–1641, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300305005849.

See remarks [?].

Slodicka:2010:IAC

[SM10] Marián Slodička and Valdemar Melicher. An iterative algorithm for a Cauchy problem in eddycurrent modelling. Applied Mathematics and Computation, 217(1):237–246, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310006132.

Sarkar:2011:EMI

[SM11] Biswajit Sarkar and Ilkyeong Moon. An EPQ model with inflation in an imperfect production system. Applied Mathematics and Computation, 217(13):6159–6167, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013159.

Shaw:2012:PVH

[SM12] Soumen Shaw and Basudeb Mukhopadhyay. Periodically varying heat source response in a functionally graded microelongated medium. Applied Mathematics and Computation, 218(11):6304–6313, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031101455X.

Sarkar:2013:IVB

[SM13] Biswajit Sarkar and Arunava Majumder. Integrated vendor-buyer supply chain model with vendor’s setup cost reduction. Applied Mathematics and Computation, 224(??):362–371, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009405.

Shaikh:2014:ANS

[SM14a] Abdul Wasim Shaikh and Ghulam Qadir Memon. Analytical and numerical solutions of fluid flow filled with and without porous media in circular pipes. Applied Mathematics and Computation, 232(??):983–999, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000757.

Sigdel:2014:GSS

[SM14b] Ram P. Sigdel and C. Connell McCluskey. Global stability for an SEI model of infectious disease with immigration. Applied Mathematics and Computation, 243(??):684–689, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031400856X.

Singh:2014:IEP

[SM14c] Rajesh Singh and Sachin Malik. Improved estimation of population variance using information on auxiliary attribute in simple random sampling. Applied Mathematics and Computation, 235(??):43–49, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003555.

Shahriari:2014:OML

[SMBI14] A. Shahriari, H. Mokhlis, A. H. A. Bakar, and H. A. Illias. Optimal multiplier load flow method using concavity theory. Applied Mathematics and Computation, 245(??):487–503, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010820.

Shindler:2012:UOF

[SMC12] Luca Shindler, Monica Moroni, and Antonio Cenedese. Using optical flow equation for particle detection and velocity prediction in particle tracking. Applied Mathematics and Computation, 218(17):8684–8694, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200152X.

Sanchez-Moreno:2013:REI

[SMDZG13] P. Sánchez-Moreno, J. S. Dehesa, A. Zarzo, and A. Guerrero. Rényi entropies, L_{q} norms and linearization of powers of hypergeometric orthogonal polynomials by means of multivariate special functions. Applied Mathematics and Computation, 223(??):25–33, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008266.

Sohail:2014:PPA

[SMH14] Ayesha Sohail, Khadija Maqbool, and Tasawar Hayat. Painlevé property and approximate solutions using Adomian decomposition for a nonlinear KdV-like wave equation. Applied Mathematics and Computation, 229(??):359–366, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013398.

Sapagovas:2012:NSA

[SMI12] M. Sapagovas, T. Meskauskas, and F. Ivanauskas. Numerical spectral analysis of a difference operator with non-local boundary conditions. Applied Mathematics and Computation, 218(14):7515–7527, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312000501.

Senu:2014:EER

[SMIS14] N. Senu, M. Mechee, F. Ismail, and Z. Siri. Embedded explicit Runge–Kutta type methods for directly solving special third order differential equations y′′ = f(x, y). Applied Mathematics and Computation, 240(??):281–293, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314006341.

Shi:2014:ICD

[SMM14] Chang-Guang Shi, Wen-Xiu Ma, and Morgan McAnally. Integrable counterparts of the D– Kaup–Newell soliton hierarchy. Applied Mathematics and Computation, 248(??):463–469, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314013393.

Sangawi:2011:LIE

[SMN11] Ali W. K. Sangawi, Ali H. M. Murid, and M. M. S. Nasser. Linear integral equations for conformal mapping of bounded multiply connected regions onto a disk with circular slits. Applied Mathematics and Computation, 218(5):2055–2068, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009374.

Sakthivel:2012:CCF

[SMN12] R. Sakthivel, N. I. Mahmudov, and Juan. J. Nieto. Controllability for a class of fractional-order neutral evolution control systems. Applied Mathematics and Computation, 218(20):10334– 10340, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003566.

Sahu:2014:ASC

[SMNN14] N. K. Sahu, R. N. Mohapatra, C. Nahak, and S. Nanda. Approximation solvability of a class of A-monotone implicit variational inclusion problems in semi-inner product spaces. Applied Mathematics and Computation, 236(??):109–117, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314003439.

Suarez:2013:TSE

[SMPA13] José P. Suárez, Tania Moreno, Ángel Plaza, and Pilar Abad. Twosided estimation of diameters reduction rate for the longest edge nsection of triangles with n ≥ 4. Applied Mathematics and Computation, 224(??):492–500, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009090.

Sakthivel:2011:ACN

[SMR11a] R. Sakthivel, N. I. Mahmudov, and Yong Ren. Approximate controllability of the nonlinear thirdorder dispersion equation. Applied Mathematics and Computation, 217(21):8507–8511, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100422X.

Stankovic:2011:DEF

[SMR11b] Miomir S. Stanković, Sladjana D. Marinković, and Predrag M. Rajković. The deformed exponential functions of two variables in the context of various statistical mechanics. Applied Mathematics and Computation, 218 (6):2439–2448, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010046.

Smith:2014:ASM

[SMR14] N. A. S. Smith, S. L.

Mitchell, and A. M. Ramos. Analysis and simplification of a mathematical model for highpressure food processes. Applied Mathematics and Computation, 226(??):20– 37, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313010874.

Sarkar:2014:EMP

[SMS14] Biswajit Sarkar, Papiya Mandal, and Sumon Sarkar. An EMQ model with price and time dependent demand under the effect of reliability and inflation. Applied Mathematics and Computation, 231(??):414–421, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000344.

Song:2010:GSD

[SMWR10] Min Song, Qingbin Meng, Rong Wu, and Jiandong Ren. The Gerber–Shiu discounted penalty function in the risk process with phase-type interclaim times. Applied Mathematics and Computation, 216(2):523–531, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310000913.

Sarada:2011:GGQ

[SN11] Jayan Sarada and K. V. Nagaraja. Generalized Gaussian quadrature rules over twodimensional regions with linear sides. Applied Mathematics and Computation, 217(12):5612–5621, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300310012415.

Sedeno-Noda:2012:ETS

[SN12] Antonio Sedeño-Noda. An efficient time and space K point-to-point shortest simple paths algorithm. Applied Mathematics and Computation, 218 (20):10244–10257, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003852.

Scott:2011:BAV

[SNC11] Melvin Scott, Beny Neta, and Changbum Chun. Basin attractors for various methods. Applied Mathematics and Computation, 218(6):2584–2599, November 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311010253.

Sun:2010:ANE

[SNH10] Daim-Yuang Sun, Sheng-Chung Ni, and Tzu-Chen Huang. Apply a novel evolutionary algorithm to the solution of parameter selection problems. Applied Mathematics and Computation, 216(11):3343–3354, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000490X.

Saberi-Nadjafi:2013:CGM

[SNMD13] J. Saberi-Nadjafi, M. Mehrabinezhad, and T. Diogo. The Coiflet–Galerkin method for linear Volterra integral equations. Applied Mathematics and Computation, 221(??):469–483, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007546.

Su:2013:BDC

[SNZL13] Qingtang Su, Yugang Niu, Hailin Zou, and Xianxi Liu. A blind dual color images watermarking based on singular value decomposition. Applied Mathematics and Computation, 219(16):8455–8466, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300266X.

Sakaguchi:2010:OTP

[SO10] Masahiko Sakaguchi and Yoshio Ohtsubo. Optimal threshold probability and expectation in semi-Markov decision processes. Applied Mathematics and Computation, 216(10):2947–2958, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004273.

Srivastava:2012:UPC

[SÖCS12] H. M. Srivastava, Hacer Özden, Ismail Naci Cangül, and Yilmaz Simsek. A unified presentation of certain meromorphic functions related to the families of the partial zeta type functions and the Lfunctions. Applied Mathematics and Computation, 219(8):3903–3913, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010260.

Sokol:2013:SPT

[Sok13] Janusz Sokól. On subordinationpreserving theorem. Applied Mathematics and Computation, 219(14): 7847–7852, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001306.

Solouma:2012:LSS

[Sol12] E. M. Solouma. Local study of scalar curvature of two-dimensional surfaces obtained by the motion of circle. Applied Mathematics and Computation, 219(8):3385–3394, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009654.

Solak:2014:SNM

[Sol14]  Süleyman Solak.  On

the spectral norm of the matrix with integer sequences. Applied Mathematics and Computation, 232(??):919–921, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314001696.

Song:2010:SAA

[Son10] Qiankun Song. Synchronization analysis in an array of asymmetric neural networks with timevarying delays and nonlinear coupling. Applied Mathematics and Computation, 216(5):1605–1613, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002560.

Song:2012:DRG

[Son12] Guang-Jing Song. Determinantal representation of the generalized inverses over the quaternion skew field with applications. Applied Mathematics and Computation, 219(2):656–667, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006649.

Soon:2011:INE

[Soo11a] Wanmei Soon. Investigation of Nash Equilibrium existence involving complementarity-constrained pricing models. Applied Mathematics and Computation, 218(5):1553–1567, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311008587.

Soon:2011:RMP

[Soo11b] Wanmei Soon. A review of multi-product pricing models. Applied Mathematics and Computation, 217(21):8149–8165, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311004085.

Saxena:2011:FIF

[SP11] Ram K. Saxena and Tibor K. Pogány. On fractional integration formulae for Aleph functions. Applied Mathematics and Computation, 218(3):985–990, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003808.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Shi:2013:NQF

[SP13a] Dongyang Shi and Lifang Pei. Nonconforming quadrilateral finite element method for a class of nonlinear sine–Gordon equations. Applied Mathematics and Computation, 219(17):9447–9460, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002610.

Stanimirovic:2013:GJE

[SP13b] Predrag S. Stanimirović and Marko D. Petković. Gauss–Jordan elimination method for computing outer inverses. Applied Mathematics and Computation, 219(9):4667–4679, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010971.

Sahoo:2014:DPP

[SP14a] Banshidhar Sahoo and Swarup Poria. Diseased prey predator model with general Holling type interactions. Applied Mathematics and Computation, 226(??):83–100, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010680.

Sahoo:2014:EAF

[SP14b] Banshidhar Sahoo and Swarup Poria. Effects of additional food on an ecoepidemic model with time delay on infection. Applied Mathematics and Computation, 245(??):17–35, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010236.

Sahoo:2014:ESA

[SP14c] Banshidhar Sahoo and Swarup Poria. Effects of supplying alternative food in a predator-prey model with harvesting. Applied Mathematics and Computation, 234(??):150–166, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002781.

Shindin:2014:NST

[SP14d] Sergey Shindin and Nabendra Parumasur. Numerical simulation of a transport fragmentation coagulation model. Applied Mathematics and Computation, 246(??):192–198, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314011151.

Saboor:2012:MGF

[SPA12] Abdus Saboor, Serge B.

Provost, and Munir Ahmad. The moment generating function of a bivariate gamma-type distribution. Applied Mathematics and Computation, 218(24):11911–11921, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005723.

Shin:2014:HFS

[SPK14] Jaemin Shin, Seong-Kwan Park, and Junseok Kim. A hybrid FEM for solving the Allen–Cahn equation. Applied Mathematics and Computation, 244(??):606–612, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009977.

Stanimirovic:2012:FRR

[SPKS12] Predrag S. Stanimirović, Dimitrios Pappas, Vasilios N. Katsikis, and Ivan P. Stanimirović. Full-rank representations of outer inverses based on the QR decomposition. Applied Mathematics and Computation, 218 (20):10321–10333, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003943.

Spalevic:2012:MMK

[SPP12] Miodrag M. Spalevíc, Miroslav S. Praníc, and Aleksandar V. Pejcev. Maximum of the modulus of kernels of Gaussian quadrature formulae for one class of Bernstein–Szegő weight functions. Applied Mathematics and Computation, 218(9):5746–5756, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014159.

Sarabia:2012:FCF

[SPT12] José María Sarabia, Faustino Prieto, and Carmen Trueba. The n-fold convolution of a finite mixture of densities. Applied Mathematics and Computation, 218(19):9992–9996, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002998.

Senthilkumaran:2010:EES

[SPV10] M. Senthilkumaran, D. Pandiaraja, and B. Mayil Vaganan. Exact and explicit solutions of Euler– Painlevé equations through generalized Cole–Hopf transformations. Applied Mathematics and Computation, 217(7):3412–3416, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310009549.

Song:2012:GSC

[SPWZ12] Bo Song, Ju H. Park, Zheng-Guang Wu, and Ya Zhang. Global synchronization of complex networks perturbed by the Poisson noise. Applied Mathematics and Computation, 219(8):3831–3839, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031200999X.

Shang:2011:CPO

[SQ11] Nina Shang and Huizeng Qin. Comments on the paper: “Oscillation of second-order nonlinear ODE with damping” [Applied Mathematics and Computation 199 (2008) 644–652]. Applied Mathematics and Computation, 218(6):2979–2980, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010125.

See [HM08].

Sun:2011:CNS

[SQJ11] Huaqing Sun, Jiangang Qi, and Haibin Jing. Classification of Non-self-adjoint Singular Sturm– Liouville Difference Equations. Applied Mathematics and Computation, 217(20):8020–8030, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003316.

Siddiqi:2010:MFB

[SR10] Shahid S. Siddiqi and Kashif Rehan. Modified form of binary and ternary 3-point subdivision schemes. Applied Mathematics and Computation, 216(3):970–982, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001396.

Silva:2012:ODO

[SR12] Valter Bruno Silva and Abel Rouboa. Optimizing the DMFC operating conditions using a response surface method. Applied Mathematics and Computation, 218(12):6733–6743, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101513X.

Soltuz:2013:MIN

[SR13] Stefan M. Soltuz and B. E.

Rhoades. A mixed iteration for nonnegative matrix factorizations. Applied Mathematics and Computation, 219(18):9847–9855, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003858.

Sahu:2014:NSS

[SR14a] P. K. Sahu and S. Saha Ray. Numerical solutions for the system of Fredholm integral equations of second kind by a new approach involving semiorthogonal Bspline wavelet collocation method. Applied Mathematics and Computation, 234(??):368–379, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002823.

Sanchez-Reyes:2014:CCO

[SR14b] J. Sánchez-Reyes. The conditions for the coincidence or overlapping of two Bézier curves. Applied Mathematics and Computation, 248(??):625–630, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314013745.

Siddiqi:2014:TPL

[SR14c] Shahid S. Siddiqi and Kashif Rehan. Ternary 2 N-point Lagrange subdivision schemes. Applied Mathematics and Computation, 249(??):444–452, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300314014374.

Shubov:2010:VFD

[SRA10] Marianna A. Shubov and Miriam Rojas-Arenaza. Vibrational frequency distribution for nonconservative model of doublewalled carbon nanotube. Applied Mathematics and Computation, 217(3):1246– 1252, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310004376.

Srivastava:2011:F

[Sri11a] H. M. Srivastava. Foreword. Applied Mathematics and Computation, 218(3):651–652, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311009441.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Srivastava:2011:SFC

[Sri11b] Rekha Srivastava. Some families of combinatorial and other series identities and their applications. Applied Mathematics and Computation, 218(3):1077–1083, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310012531.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Srivastava:2014:SCG

[Sri14] Rekha Srivastava. Some classes of generating functions associated with a certain family of extended and generalized hypergeometric functions. Applied Mathematics and Computation, 243(??):132–137, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314007395.

Sharma:2013:RSP

[SRP13] Kapil K. Sharma, Pratima Rai, and Kailash C. Patidar. A review on singularly perturbed differential equations with turning points and interior layers. Applied Mathematics and Computation, 219(22):10575–10609, July 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313004840.

Samanta:2013:EAP

[SRS^{+}13] S. Samanta, S. Rana, A. Sharma, A. K. Misra, and J. Chattopadhyay. Effect of awareness programs by media on the epidemic outbreaks: a mathematical model. Applied Mathematics and Computation, 219(12):6965–6977, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000441.

S:2010:SSS

[SS10a] Cesar A. Gómez S. and Alvaro Salas. Special symmetries to standard Riccati equations and applications. Applied Mathematics and Computation, 216(10):3089–3096, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004650.

Saad:2010:AHD

[SS10b] Husam L. Saad and Abbas A. Sukhi. Another homogeneous q-difference operator. Applied Mathematics and Computation, 215(12):4332–4339, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309011187.

Sharma:2010:MJM

[SS10c] Janak Raj Sharma and Rajni Sharma. Modified Jarratt method for computing multiple roots. Applied Mathematics and Computation, 217(2):878– 881, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310007162.

Simsek:2010:FAT

[SS10d] Yilmaz Simsek and H. M. Srivastava. A family of padic twisted interpolation functions associated with the modified Bernoulli numbers. Applied Mathematics and Computation, 216(10):2976–2987, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004303.

Solis:2010:AID

[SS10e] Francisco J. Solis and Rosa I. Sánchez. Attractor information for discrete dynamical systems by means of optimal discrete Galerkin bases. Applied Mathematics and Computation, 216(12):3578–3583, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005205.

Song:2010:ESWb

[SS10f] Ming Song and Shuguang Shao. Exact solitary wave solutions of the generalized (2 + 1) dimensional Boussinesq equation. Applied Mathematics and Computation, 217(7):3557–3563, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000977X.

Sulochana:2010:USM

[SS10g] C. Sulochana and Moka Shekhu. Unsteady state mathematical model of chemically reactive pollutants from an instantaneous line source into a stable atmospheric boundary layer. Applied Mathematics and Computation, 217(6):2757–2767, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008611.

Sahmurova:2011:SPD

[SS11a] Aida Sahmurova and Veli B. Shakhmurov. Singular perturbation degenerate problems occurring in atmospheric dispersion of pollutants. Applied Mathematics and Computation, 218(3):1052–1056, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311005121.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Sevli:2011:GSF

[SS11b] Hamdullah Sevli and Ekrem Savas. On general summability factor theorem of infinite series. Applied Mathematics and Computation, 218(3):1046–1051, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S009630031100436X.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Sharma:2011:NTF

[SS11c] Janak Raj Sharma and Rajni Sharma. New third and fourth order nonlinear solvers for computing multiple roots. Applied Mathematics and Computation, 217(23):9756–9764, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006333.

Srdjevic:2011:BCE

[SS11d] Bojan Srdjevic and Zorica Srdjevic. Bi-criteria evolution strategy in estimating weights from the AHP ratio-scale matrices. Applied Mathematics and Computation, 218(4):1254–1266, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008198.

Stevic:2011:COS

[SS11e] Stevo Stević and Ajay K.

Sharma. Composition operators from the space of Cauchy transforms to Bloch and the little Blochtype spaces on the unit disk. Applied Mathematics and Computation, 217(24):10187–10194, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006692.

Stevic:2011:ENC

[SS11f] Stevo Stević and Ajay K. Sharma. Essential norm of composition operators between weighted Hardy spaces. Applied Mathematics and Computation, 217(13):6192–6197, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013202.

Stevic:2011:IDF

[SS11g] Stevo Stević and Ajay K. Sharma. Iterated differentiation followed by composition from Bloch-type spaces to weighted BMOA spaces. Applied Mathematics and Computation, 218(7):3574–3580, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011441.

Stevic:2011:WCO

[SS11h] Stevo Stević and Ajay K.

Sharma. Weighted composition operators between Hardy and growth spaces on the upper halfplane. Applied Mathematics and Computation, 217(10):4928–4934, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310011707.

Sun:2011:GSM

[SS11i] Ruoyan Sun and Junping Shi. Global stability of multigroup epidemic model with group mixing and nonlinear incidence rates. Applied Mathematics and Computation, 218(2):280–286, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007260.

Singh:2012:IEP

[SS12a] Housila P. Singh and Ramkrishna S. Solanki. Improved estimation of population mean in simple random sampling using information on auxiliary attribute. Applied Mathematics and Computation, 218(15):7798–7812, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000859.

Skakauskas:2012:SPS

[SS12b] Vladas Skakauskas and Olga Stikoniene. Stability of persistent solutions to a population dynamics model. Applied Mathematics and Computation, 218(17):8987–8996, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002019.

Stevic:2012:GCO

[SS12c] Stevo Stević and Ajay K. Sharma. Generalized composition operators on weighted Hardy spaces. Applied Mathematics and Computation, 218(17): 8347–8352, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312001075.

Salehbhai:2013:LTT

[SS13a] I. A. Salehbhai and A. K.

Shukla. The Laguerre transform in two variables and its application in partial differential equations. Applied Mathematics and Computation, 219(9):4650–4654, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010934.

Shakhmurov:2013:APP

[SS13b] Veli B. Shakhmurov and Aida Sahmurova. Abstract parabolic problems with parameter and application. Applied Mathematics and Computation, 219(17):9561–9571, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002804.

Stamov:2013:APP

[SS13c] Gani Tr. Stamov and Alexander G. Stamov. On almost periodic processes in uncertain impulsive delay models of price fluctuations in commodity markets. Applied Mathematics and Computation, 219(10):5376–5383, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011411.

Sugie:2013:CSN

[SS13d] Jitsuro Sugie and Yasuhisa Saito. Convergence of solutions of nonlinear systems with integrable forcing term and its applications to a biological model. Applied Mathematics and Computation, 219(15):8169–8177, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001446.

Samanta:2014:ADC

[SS14a] G. P. Samanta and Swarnali Sharma. Analysis of a delayed Chlamydia epidemic model with pulse vaccination. Applied Mathematics and Computation, 230(??):555–569, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014136.

Sanches:2014:CDS

[SS14b] Carlos Alberto Alonso Sanches and Nei Yoshihiro Soma. A computational DNA solution approach for the Quadratic Diophantine Equation. Applied Mathematics and Computation, 238(??):436–443, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005463.

Sasu:2014:ZTM

[SS14c] Adina Luminita Sasu and Bogdan Sasu. A Zabczyk type method for the study of the exponential trichotomy of discrete dynamical systems. Applied Mathematics and Computation, 245(??):447–461, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314010753.

Sehba:2014:SPT

[SS14d] Benoît Sehba and Stevo Stević. On some producttype operators from Hardy-Orlicz and Bergman-Orlicz spaces to weightedtype spaces. Applied Mathematics and Computation, 233(??):565–581, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000320.

Slavik:2014:ESD

[SS14e] Antonín Slavík and Petr Stehlík. Explicit solutions to dynamic diffusiontype equations and their time integrals. Applied Mathematics and Computation, 234(??):486–505, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002458.

Sharma:2014:DPN

[SSA14a] Amit Sharma, Anuj Kumar Sharma, and Kulbhushan Agnihotri. The dynamic of planktonnutrient interaction with delay. Applied Mathematics and Computation, 231(??):503–515, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000794.

Siddiqi:2014:FGU

[SSA14b] Shahid S. Siddiqi, Saima Siddiqui, and Nadeem Ahmad. Fractal generation using ternary 5point interpolatory subdivision scheme. Applied Mathematics and Computation, 234(??):402–411, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002549.

Stevic:2011:ENP

[SSB11a] Stevo Stević, Ajay K.

Sharma, and Ambika Bhat. Essential norm of products of multiplication composition and differentiation operators on weighted Bergman spaces. Applied Mathematics and Computation, 218(6):2386–2397, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008885.

Stevic:2011:PMC

[SSB11b] Stevo Stević, Ajay K.

Sharma, and Ambika Bhat. Products of multiplication composition and differentiation operators on weighted Bergman spaces. Applied Mathematics and Computation, 217(20):8115–8125, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311003535.

Sarkar:2011:SIR

[SSC11] Bikash Kanti Sarkar, Shib Sankar Sana, and Kripasindhu Chaudhuri. Selecting informative rules with parallel genetic algorithm in classification problem. Applied Mathematics and Computation, 218(7):3247–3264, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011052.

Selvan:2013:BLS

[SSCA13] S. Easter Selvan, M. S. P. Subathra, A. Hepzibah Christinal, and Umberto Amato. On the benefits of Laplace samples in solving a rare event problem using cross-entropy method. Applied Mathematics and Computation, 225(??):843–859, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313010667.

Soleimani:2014:EHM

[SSCT14] Farahnaz Soleimani, Fazlollah Soleymani, Alicia Cordero, and Juan R. Torregrosa. On the extension of Householder’s method for weighted Moore–Penrose inverse. Applied Mathematics and Computation, 231(??):407–413, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000587.

S:2010:NPS

[SSF10] Cesar A. Gómez S., Alvaro H. Salas, and Bernardo Acevedo Frias. New periodic and soliton solutions for the Generalized BBM and Burgers– BBM equations. Applied Mathematics and Computation, 217(4):1430–1434, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309005517.

Sydney:2013:OAC

[SSG13] Ali Sydney, Caterina Scoglio, and Don Gruenbacher. Optimizing algebraic connectivity by edge rewiring. Applied Mathematics and Computation, 219(10):5465–5479, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011551.

Saxena:2010:CSF

[SSK10] R. K. Saxena, Ravi Saxena, and S. L. Kalla. Computational solution of a fractional generalization of the Schrödinger equation occurring in quantum mechanics. Applied Mathematics and Computation, 216(5):1412–1417, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002134.

Subasi:2012:VTO

[SSK12] Murat Subasi, Yesim Saraç, and Ahmet Kaçar. A variational technique for optimal boundary control in a hyperbolic problem. Applied Mathematics and Computation, 218(12):6629–6636, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311015402.

Samokhin:2013:SIM

[SSK13] Alexander Samokhin, Yury Shestopalov, and Kazuya Kobayashi. Stationary iteration methods for solving 3D electromagnetic scattering problems. Applied Mathematics and Computation, 222(??):107–122, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007698.

Samsuzzoha:2010:NSI

[SSL10] Md. Samsuzzoha, Manmohan Singh, and David Lucy. Numerical study of an influenza epidemic model with diffusion. Applied Mathematics and Computation, 217(7):3461– 3479, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300310009641.

Samsuzzoha:2012:NSI

[SSL12] Md. Samsuzzoha, Manmohan Singh, and David Lucy. A numerical study on an influenza epidemic model with vaccination and diffusion. Applied Mathematics and Computation, 219(1):122–141, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005528.

Samsuzzoha:2013:PEI

[SSL13] Md. Samsuzzoha, Manmohan Singh, and David Lucy. Parameter estimation of influenza epidemic model. Applied Mathematics and Computation, 220(??):616–629, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300790X.

Sunil:2011:NSA

[SSM11] Sunil, Poonam Sharma, and Amit Mahajan. A nonlinear stability analysis of a rotating doublediffusive magnetized ferrofluid. Applied Mathematics and Computation, 218(6):2785–2799, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010472.

Sakthivel:2013:AAC

[SSM13] R. Sakthivel, Srimanta Santra, and K. Mathiyalagan. Admissibility analysis and control synthesis for descriptor systems with random abrupt changes. Applied Mathematics and Computation, 219(18):9717–9730, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003111.

Singer:2012:TLP

[SSN^{+}12] Sanja Singer, Saša Singer, Vedran Novakovíc, Davor Davidovíc, Krešimir Bokulíc, and Aleksandar Ušćumlíc. Threelevel parallel J-Jacobi algorithms for Hermitian matrices. Applied Mathematics and Computation, 218(9):5704–5725, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031101410X.

Sheng:2011:CIC

[SSO11] Changtao Sheng, Jiuliang Shi, and Yao Ouyang. Chebyshev’s inequality for Choquet-like integral. Applied Mathematics and Computation, 217(22):8936–8942, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031100467X.

Selimefendigil:2011:IHT

[SSP11] F. Selimefendigil, R. I. Sujith, and W. Polifke. Identification of heat transfer dynamics for non-modal analysis of thermoacoustic stability. Applied Mathematics and Computation, 217(11):5134–5150, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007964.

Shobkolaei:2012:CFP

[SSRA12] N. Shobkolaei, S. Sedghi, J. R. Roshan, and I. Altun. Common fixed point of mappings satisfying almost generalized (ST)- contractive condition in partially ordered partial metric spaces. Applied Mathematics and Computation, 219(2):443–452, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006716.

Singh:2013:ETT

[SSS13a] Sanjay Kumar Singh, Umesh Singh, and Vikas Kumar Sharma. Expected total test time and Bayesian estimation for generalized Lindley distribution under progressively Type-II censored sample where removals follow the Beta-binomial probability law. Applied Mathematics and Computation, 222(??):402–419, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008084.

Sykora:2013:HCH

[SSS13b] Jan Sýkora, Michal Sejnoha, and Jirí Sejnoha. Homogenization of coupled heat and moisture transport in masonry structures including interfaces. Applied Mathematics and Computation, 219(13):7275–7285, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311002414.

Sharma:2014:NUB

[SSS14] Vikas Kumar Sharma, Sanjay Kumar Singh, and Umesh Singh. A new upside-down bathtub shaped hazard rate model for survival data analysis. Applied Mathematics and Computation, 239(??):242–253, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005888.

Shukla:2014:MAE

[SSST14] J. B. Shukla, Gaurav Singh, Poonam Shukla, and Agraj Tripathi. Modeling and analysis of the effects of antivirus software on an infected computer network. Applied Mathematics and Computation, 227(??):11–18, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313011582.

Soltani:2014:IPR

[SSTM14] Roya Soltani, Seyed Jafar Sadjadi, and Reza Tavakkoli-Moghaddam. Interval programming for the redundancy allocation with choices of redundancy strategy and component type under uncertainty: Erlang time to failure distribution. Applied Mathematics and Computation, 244(??):413–421, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009576.

Sbibih:2014:GFT

[SSTZ14] D. Sbibih, A. Serghini, A. Tijini, and A. Zidna. A general family of third order method for finding multiple roots. Applied Mathematics and Computation, 233(??):338–350, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001532.

Soleymani:2013:AIM

[SSU13] F. Soleymani, Predrag S.

Stanimirović, and Malik Zaka Ullah. An accelerated iterative method for computing weighted Moore–Penrose inverse. Applied Mathematics and Computation, 222(??): 365–371, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313007893.

Santin:2011:ACF

[SSV11] G. Santin, A. Sommariva, and M. Vianello. An algebraic cubature formula on curvilinear polygons. Applied Mathematics and Computation, 217(24):10003–10015, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006412.

Stanojevic:2013:ESC

[SSV13] Milan Stanojević, Bogdana Stanojević, and Mirko Vujosević. Enhanced savings calculation and its applications for solving capacitated vehicle routing problem. Applied Mathematics and Computation, 219 (20):10302–10312, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300313004001.

Sekulic:2011:SCS

[SSZ11] Dalibor L. Sekulić, Miljko V.

Satarić, and Milos B. Zivanov. Symbolic computation of some new nonlinear partial differential equations of nanobiosciences using modified extended tanh-function method. Applied Mathematics and Computation, 218(7):3499–3506, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011374.

Song:2014:EIZ

[ST14] Yisheng Song and ChangAn Tian. An explicit iteration for zeros of accretive operators. Applied Mathematics and Computation, 233(??):369–376, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314001258.

Stanimirovic:2011:SIC

[Sta11] Stefan Stanimirović. Some identities on Catalan numbers and hypergeometric functions via Catalan matrix power. Applied Mathematics and Computation, 217(22):9122–9132, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005182.

Stanek:2012:LPP

[Sta12a] Svatoslav Stanek. Limit properties of positive solutions of fractional boundary value problems. Applied Mathematics and Computation, 219(5):2361–2370, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312009071.

Stannett:2012:CAP

[Sta12b]  Mike Stannett.  Com-

puting the appearance of physical reality. Applied Mathematics and Computation, 219(1):54–62, September 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010204.

Stamova:2014:GSI

[Sta14a] Ivanka Stamova. Global stability of impulsive fractional differential equations. Applied Mathematics and Computation, 237(??):605–612, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004299.

Stanic:2014:MOP

[Sta14b] Marija P. Stanić. Multiple orthogonal polynomials on the semicircle and applications. Applied Mathematics and Computation, 243(??):269–282, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007917.

Stankovic:2014:SEA

[Sta14c] Mića S. Stanković. Special equitorsion almost geodesic mappings of the third type of nonsymmetric affine connection spaces. Applied Mathematics and Computation, 244(??):695–701, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300314009783.

Stavroulakis:2014:OCD

[Sta14d] I. P. Stavroulakis. Oscillation criteria for delay and difference equations with non-monotone arguments. Applied Mathematics and Computation, 226(??):661–672, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313010989.

Stevic:2010:CFD

[Ste10a] Stevo Stević. Composition followed by differentiation from H^{∞} and the Bloch space to n th weighted-type spaces on the unit disk. Applied Mathematics and Computation, 216(12):3450–3458, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003814.

Stevic:2010:COH

[Ste10b] Stevo Stević. Composition operators from the Hardy space to the n th weighted-type space on the unit disk and the halfplane. Applied Mathematics and Computation, 215(11):3950–3955, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309010327.

Stevic:2010:COW

[Ste10c] Stevo Stević. Composition operators from the weighted Bergman space to the n th weighted-type space on the upper halfplane. Applied Mathematics and Computation, 217(7):3379–3384, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310009355.

Stevic:2010:END

[Ste10d] Stevo Stević. Essential norm of differences of weighted composition operators between weightedtype spaces on the unit ball. Applied Mathematics and Computation, 217(5):1811–1824, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031000202X.

Stevic:2010:GSM

[Ste10e] Stevo Stević. Global stability of a max-type difference equation. Applied Mathematics and Computation, 216(1):354–356, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000330.

Stevic:2010:GSS

[Ste10f] Stevo Stević. Global stability of some symmetric difference equations. Applied Mathematics and Computation, 216(1):179– 186, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310000421.

Stevic:2010:ITOb

[Ste10g] Stevo Stević. Integraltype operators between α-Bloch spaces and Besov spaces on the unit ball. Applied Mathematics and Computation, 216(12): 3541–3549, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031000514X.

Stevic:2010:NMO

[Ste10h]  Stevo Stević.  Norms

of multiplication operators on Hardy spaces and weighted composition operators from Hardy spaces to weighted-type spaces on bounded symmetric domains. Applied Mathematics and Computation, 217(6):2870–2876, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008763.

Stevic:2010:NSO

[Ste10i]  Stevo Stević.  Norms

of some operators on bounded symmetric domains. Applied Mathematics and Computation, 216(1):187–191, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310000433.

Stevic:2010:ITOa

[Ste10j]  Stevo Stević.  On an

integral-type operator from logarithmic Bloch-type spaces to mixed-norm spaces on the unit ball. Applied Mathematics and Computation, 215(11): 3817–3823, February 1, 2010. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630030901011X.

Stevic:2010:OLB

[Ste10k] Stevo Stević. On operator P_{ϕ}^{g} from the logarithmic Bloch-type space to the mixed-norm space on the unit ball. Applied Mathematics and Computation, 215(12):4248–4255, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309010959.

Stevic:2010:WCOc

[Ste10l] Stevo Stević. Weighted composition operators from Bergman–Privalovtype spaces to weightedtype spaces on the unit ball. Applied Mathematics and Computation, 217(5):1939–1943, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007344.

Stevic:2010:WCOb

[Ste10m] Stevo Stević. Weighted composition operators from the logarithmic weighted-type space to the weighted Bergman space in C^{n}. Applied Mathematics and Computation, 216(3):924–928, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310001293.

Stevic:2010:WDC

[Ste10n] Stevo Stević. Weighted differentiation composition operators from H^{∞} and Bloch spaces to n th weighted-type spaces on the unit disk. Applied Mathematics and Computation, 216(12):3634–3641, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005333.

Stevic:2011:CCF

[Ste11a] Stevo Stevíc. Characterizations of composition followed by differentiation between Bloch-type spaces. Applied Mathematics and Computation, 218(8):4312–4316, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300311012343.

Stevic:2011:PTO

[Ste11b]  Stevo Stević.  On a

product-type operator from Bloch spaces to weighted-type spaces on the unit ball. Applied Mathematics and Computation, 217(12):5930–5935, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310013160.

Stevic:2011:SDEa

[Ste11c] Stevo Stević. On a system of difference equations. Applied Mathematics and Computation, 218(7):3372–3378, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011209.

Stevic:2011:SDEb

[Ste11d] Stevo Stevíc. On a system of difference equations with period two coefficients. Applied Mathematics and Computation, 218(8):4317–4324, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012355.

Stevic:2011:SIT

[Ste11e] Stevo Stević. On some integral-type operators between a general space and Bloch-type spaces. Applied Mathematics and Computation, 218(6):2600–2618, November 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311010265.

Stevic:2011:DE

[Ste11f] Stevo Stevíc. On the difference equation x_{n} = x_{n−2}/(b_{n} + c_{n}x_{n−1}x_{n−2}). Applied Mathematics and Computation, 218(8):4507– 4513, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300311012628.

Stevic:2011:PCN

[Ste11g] Stevo Stević. Periodicity of a class of nonautonomous max-type difference equations. Applied Mathematics and Computation, 217(23):9562–9566, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311005790.

Stevanovic:2012:CZI

[Ste12a] Dragan Stevanović. Comparing the Zagreb indices of the NEPS of graphs. Applied Mathematics and Computation, 219(3):1082–1086, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200714X.

Stevic:2012:ADE

[Ste12b] Stevo Stević. Application of difference equations in divisibility. Applied Mathematics and Computation, 219(4):2060–2066, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008442.

Stevic:2012:BSS

[Ste12c] Stevo Stević. Bounded solutions of some systems of nonlinear functional differential equations with iterated deviating argument. Applied Mathematics and Computation, 218(21):10429– 10434, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312003669.

Stevic:2012:BCI

[Ste12d] Stevo Stevíc. Boundedness and compactness of an integral-type operator from Bloch-type spaces with normal weights to F (pqs) space. Applied Mathematics and Computation, 218(9):5414–5421, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013580.

Stevic:2012:DEN

[Ste12e]  Stevo Stević.  Dif-

ference equations and nonexistence of solutions of an integral equation. Applied Mathematics and Computation, 219(4):1998–2003, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031200834X.

Stevic:2012:EBS

[Ste12f] Stevo Stević. Existence of bounded solutions of some systems of nonlinear functional differential equations with complicated deviating argument. Applied Mathematics and Computation, 218(19):9974–9979, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312003438.

Stevic:2012:GBS

[Ste12g]  Stevo Stević.  Glob-

ally bounded solutions of a system of nonlinear functional differential equations with iterated deviating argument. Applied Mathematics and Computation, 219(4):2180–2185, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008569.

Stevic:2012:LEL

[Ste12h]  Stevo Stević.  Local

existence of Lipschitzcontinuous solutions of systems of nonlinear functional equations with iterated deviations. Applied Mathematics and Computation, 218(21):10542– 10547, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312004006.

Stevic:2012:NSO

[Ste12i]  Stevo Stević.  Norm

of some operators from logarithmic Bloch-type spaces to weighted-type spaces. Applied Mathematics and Computation, 218(22):11163–11170, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312004900.

Stevic:2012:SRS

[Ste12j]  Stevo Stević.  On a

solvable rational system of difference equations. Applied Mathematics and Computation, 219(6):2896–2908, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009113.

Stevic:2012:TOS

[Ste12k] Stevo Stević. On a thirdorder system of difference equations. Applied Mathematics and Computation, 218(14):7649–7654, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000707.

Stevic:2012:CSC

[Ste12l] Stevo Stević. On continuous solutions of a class of systems of nonlinear functional difference equations with deviating argument. Applied Mathematics and Computation, 218 (20):10188–10193, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003621.

Stevic:2012:SCS

[Ste12m] Stevo Stević. On solutions of a class of systems of nonlinear functional differential equations of neutral type with complicated deviations of an argument. Applied Mathematics and Computation, 219(8):3693–3700, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312009836.

Stevic:2012:SLD

[Ste12n] Stevo Stević. On some linear difference equations and inequalities with continuous argument. Applied Mathematics and Computation, 218(19):9831–9838, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002974.

Stevic:2012:SPS

[Ste12o] Stevo Stević. On some periodic systems of maxtype difference equations. Applied Mathematics and Computation, 218(23):11483–11487, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312005371.

Stevic:2012:SSS

[Ste12p] Stevo Stevíc. On some solvable systems of difference equations. Applied Mathematics and Computation, 218(9):5010–5018, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311013142.

Stevic:2012:DEI

[Ste12q] Stevo Stević. On the difference equation x_{n} = \frac{x_{n−k}}{b+cx_{n−1}···x_{n−k}}. Applied Mathematics and Computation, 218(11):6291–6296, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014536.

Stevic:2012:SCZ

[Ste12r] Stevo Stević. Solutions converging to zero of some systems of nonlinear functional differential equations with iterated deviating argument. Applied Mathematics and Computation, 219(8):4031–4035, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010405.

Stevic:2012:SMT

[Ste12s] Stevo Stević. Solutions of a max-type system of difference equations. Applied Mathematics and Computation, 218(19):9825–9830, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312002962.

Stevic:2012:SSF

[Ste12t] Stevo Stević. Solutions of a system of functional differential equations in a right neighborhood of zero. Applied Mathematics and Computation, 219(3):1011–1019, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312007059.

Stevic:2012:UEB

[Ste12u] Stevo Stević. Unique existence of bounded continuous solutions on the real line of a class of nonlinear functional equations with complicated deviations. Applied Mathematics and Computation, 218(15):7813–7817, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312000951.

Stevic:2012:WIR

[Ste12v] Stevo Stević. Weighted iterated radial operators between different weighted Bergman spaces on the unit ball. Applied Mathematics and Computation, 218(17):8288–8294, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000975.

Stevic:2012:WRO

[Ste12w] Stevo Stević. Weighted radial operator from the mixed-norm space to the n th weighted-type space on the unit ball. Applied Mathematics and Computation, 218(18):9241–9247, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002329.

Stevic:2013:ACS

[Ste13a] Stevo Stević. Asymptotically convergent solutions of a system of nonlinear functional differential equations of neutral type with iterated deviating arguments. Applied Mathematics and Computation, 219(11):6197–6203, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312013069.

Stevic:2013:BSS

[Ste13b] Stevo Stević. Behaviour of solutions of some linear functional equations at infinity. Applied Mathematics and Computation, 219(11):6134–6141, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012519.

Stevic:2013:DUS

[Ste13c] Stevo Stević. Domains of undefinable solutions of some equations and systems of difference equations. Applied Mathematics and Computation, 219(24):11206–11213, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313005365.

Stevic:2013:ESB

[Ste13d] Stevo Stević. Existence of solutions bounded together with their first derivatives of some systems of nonlinear functional differential equations. Applied Mathematics and Computation, 219(12):6457–6465, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312012520.

Stevic:2013:SSDa

[Ste13e] Stevo Stević. On a solvable system of difference equations of fourth order. Applied Mathematics and Computation, 219(10):5706–5716, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011782.

Stevic:2013:SSDb

[Ste13f] Stevo Stević. On a solvable system of difference equations of k th order. Applied Mathematics and Computation, 219(14):7765–7771, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031300115X.

Stevic:2013:SSM

[Ste13g] Stevo Stević. On a symmetric system of maxtype difference equations. Applied Mathematics and Computation, 219(15):8407–8412, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001240.

Stevic:2013:SDEb

[Ste13h] Stevo Stević. On a system of difference equations of odd order solvable in closed form. Applied Mathematics and Computation, 219(15):8222–8230, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001811.

Stevic:2013:SDEc

[Ste13i] Stevo Stević. On a system of difference equations which can be solved in closed form. Applied Mathematics and Computation, 219(17):9223–9228, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300313003391.

Stevic:2013:BCD

[Ste13j] Stevo Stević. On bounded continuously differentiable solutions with Lipschitz first derivatives of a system of functional differential equations. Applied Mathematics and Computation, 219(11):6344–6353, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031201315X.

Stevic:2013:DAS

[Ste13k]  Stevo Stević.  On q-

difference asymptotic solutions of a system of nonlinear functional differential equations. Applied Mathematics and Computation, 219(15):8295–8301, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313001513.

Stevic:2013:SCS

[Ste13l] Stevo Stević. On solutions of a class of systems of nonlinear functional equations in a neighborhood of zero satisfying Lipschitz condition. Applied Mathematics and Computation, 219(19):9893–9898, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003172.

Stevic:2013:S

[Ste13m] Stevo Stević. On the system x_{n+1} = y_{n}x_{n−k}/(y_{n−k+1}(a_{n}+ b_{n}y_{n}x_{n−k})), y_{n+1} = x_{n}y_{n−k}/(x_{n−k+1}(c_{n}+d_{n}x_{n}y_{n−k})). Applied Mathematics and Computation, 219(9):4526– 4534, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312010764.

Stevic:2013:SDEa

[Ste13n] Stevo Stević. On the system of difference equations x_{n} = c_{n}y_{n−1}/(a_{n} + b_{n}y_{n−1}x_{n−2}y_{n−3}), y_{n} = γ_{n}x_{n−3}/(α_{n}+β_{n}x_{n−2}x_{n−3}). Applied Mathematics and Computation, 219(9):4755– 4764, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300312011083.

Stevic:2013:UEP

[Ste13o] Stevo Stević. Unique existence of periodic solutions of a class of systems of nonlinear difference equations with continuous argument. Applied Mathematics and Computation, 219(14):7422–7428, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000659.

Stevanovic:2014:RDL

[Ste14a] Dragan Stevanović. Remarks on dynamic load balancing of integer loads and integral graphs. Applied Mathematics and Computation, 226(??):38– 43, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313011107.

Stevic:2014:ABS

[Ste14b] Stevo Stević. Asymptotic behavior of solutions of a system of qfunctional-difference equations. Applied Mathematics and Computation, 235(??):538–545, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0096300314003567.

Stevic:2014:EBS

[Ste14c] Stevo Stević. Existence of bounded solutions of a class of neutral systems of functional differential equations. Applied Mathematics and Computation, 231(??):478–488, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014410.

Stevic:2014:FBC

[Ste14d] Stevo Stević. On families of bounded continuous solutions of some systems of functionaldifference equations on half-intervals. Applied Mathematics and Computation, 231(??):169–178, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313014288.

Stevic:2014:PSS

[Ste14e] Stevo Stević. On positive solutions of some classes of max-type systems of difference equations. Applied Mathematics and Computation, 232(??):445–452, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313014161.

Stevic:2014:SZS

[Ste14f] Stevo Stević. Stability of the zero solution of a family of functionaldifferential equations. Applied Mathematics and Computation, 238(??):50– 56, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314005335.

Shih:2010:MIB

[STH10] Yuan-Kang Shih, Jimmy J. M. Tan, and Lih-Hsing Hsu. Mutually independent bipanconnected property of hypercube. Applied Mathematics and Computation, 217(8):4017–4023, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010386.

Subramani:2010:STT

[STM10] K. Subramani, C. Tauras, and K. Madduri. Space– time tradeoffs in negative cycle detection — an empirical analysis of the Stressing Algorithm. Applied Mathematics and Computation, 215(10):3563–3575, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009692.

Suarez-Taboada:2012:NSP

[STV12] M. Suárez-Taboada and C. Vázquez. Numerical solution of a PDE model for a ratchet-cap pricing with BGM interest rate dynamics. Applied Mathematics and Computation, 218(9):5217–5230, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013348.

Stevic:2010:ITOc

[SU10a] Stevo Stević and Sei-Ichiro Ueki. On an integraltype operator between weighted-type spaces and Bloch-type spaces on the unit ball. Applied Mathematics and Computation, 217(7):3127–3136, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009094.

Stevic:2010:WCOa

[SU10b] Stevo Stević and Sei-Ichiro Ueki. Weighted composition operators from the weighted Bergman space to the weighted Hardy space on the unit ball. Applied Mathematics and Computation, 215(10):3526–3533, January 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309009643.

Singh:2012:FAM

[SU12] Chandan Singh and Rahul Upneja. Fast and accurate method for high order Zernike moments computation. Applied Mathematics and Computation, 218(15):7759–7773, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000781.

Su:2013:CMS

[Su13] Qifang Su. The convergence of multi-shift QR algorithm for symmetric matrices. Applied Mathematics and Computation, 222(??):343–355, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313008114.

Sucu:2014:DAS

[Suc14]  Sezgin Sucu.  Dunkl

analogue of Szasz operators. Applied Mathematics and Computation, 244(??):42–48, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009321.

Suhadolnik:2013:SBM

[Suh13] Alojz Suhadolnik. Superlinear bracketing method for solving nonlinear equations. Applied Mathematics and Computation, 219(14):7369–7376, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000052.

Sulaiman:2010:NSF

[Sul10] W. T. Sulaiman. A note on ||A||_{k} summability factors of infinite series. Applied Mathematics and Computation, 216(9):2645–2648, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310003632.

Sun:2010:NIM

[Sun10]  Ye-Peng Sun.  The

non-isospectral modified Kadomtsev–Petviashvili equation with self-consistent sources and its coupled system. Applied Mathematics and Computation, 217(4):1639–1645, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309007930.

Sun:2011:ASS

[Sun11a] Leping Sun. Asymptotic stability for the system of neutral delay differential equations. Applied Mathematics and Computation, 218(2):337–345, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311007508.

Sun:2011:SAT

[Sun11b] Wei Wei Sun. Stabilization analysis of timedelay Hamiltonian systems in the presence of saturation. Applied Mathematics and Computation, 217(23):9625–9634, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311006023.

Sun:2012:BFH

[Sun12] Chao-Ming Sun. Bipanconnectivity of faulty hypercubes with minimum degree. Applied Mathematics and Computation, 218(9):5518–5523, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013713.

Sun:2014:KJM

[Sun14]  Ruoyan Sun.  Kinet-

ics of jobs in multi-link cities with generalized migration-driven aggregation process. Applied Mathematics and Computation, 247(??):1003–1010, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012818.

Saeed:2013:HWQ

[SuR13] Umer Saeed and Mujeeb ur Rehman. Haar waveletquasilinearization technique for fractional nonlinear differential equations. Applied Mathematics and Computation, 220(??):630–648, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007686.

Sweet:2010:ASG

[SV10a] Erik Sweet and Robert A. Van Gorder. Analytical solutions to a generalized Drinfel’d–Sokolov equation related to DSSH and KdV6. Applied Mathematics and Computation, 216(10):2783–2791, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310003930.

Sweet:2010:THT

[SV10b] Erik Sweet and Robert A. Van Gorder. Trigonometric and hyperbolic type solutions to a generalized Drinfel’d–Sokolov equation. Applied Mathematics and Computation, 217(8):4147–4166, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010684.

Sweet:2012:TWS

[SV12] Erik Sweet and Robert A.

Van Gorder. Traveling wave solutions (u, v) to a generalized Drinfel’d– Sokolov system which satisfy u = a_{1}v^{m}+a_{0}. Applied Mathematics and Computation, 218(19):9911–9921, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003360.

Svacek:2011:NMA

[Svá11]  P. Sváček.  Numerical

modelling of aeroelastic behaviour of an airfoil in viscous incompressible flow. Applied Mathematics and Computation, 217(11):5078–5086, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300310008027.

Sacker:2010:CSP

[SvB10] Robert J. Sacker and Hubertus F. von Bremen. A conjecture on the stability of the periodic solutions of Ricker’s equation with periodic parameters. Applied Mathematics and Computation, 217(3):1213–1219, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006089.

Sebestyen:2010:DMS

[SVGC10] Zoltán Sebestyén, Zoltán Varga, József Garay, and Roberta Cimmaruta. Dynamic model and simulation analysis of the genetic impact of population harvesting. Applied Mathematics and Computation, 216(2):565–575, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310001025.

Snyder:2013:CTD

[SVV13] Brooke C. Snyder, Robert A. Van Gorder, and K. Vajravelu. Continuoustime dynamic games for the Cournot adjustment process for competing oligopolists. Applied Mathematics and Computation, 219(12):6400–6409, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313000192.

Song:2010:MCN

[SW10a] Lun-Ji Song and Yu-Jiang Wu. A modified Crank– Nicolson scheme with incremental unknowns for convection dominated diffusion equations. Applied Mathematics and Computation, 215(9):3293–3301, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309009229.

Sun:2010:QIS

[SW10b] Yufeng Sun and Peiguang Wang. Quasilinear iterative scheme for a fourthorder differential equation with retardation and anticipation. Applied Mathematics and Computation, 217(7):3442–3452, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310009586.

Song:2011:CCR

[SW11] Guang-Jing Song and Qing-Wen Wang. Condensed Cramer rule for some restricted quaternion linear equations. Applied Mathematics and Computation, 218(7):3110–3121, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311010782.

Suo:2012:ECR

[SW12] Jianqing Suo and Wanyi Wang. Eigenvalues of a class of regular fourthorder Sturm–Liouville problems. Applied Mathematics and Computation, 218(19):9716–9729, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312002470.

Shu:2013:TWS

[SW13] Li Shu and Peixuan Weng. Traveling wave solutions of a diffusive SI model with strong Allee effect. Applied Mathematics and Computation, 222(??):190–200, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007078.

Shen:2014:AOD

[SW14a] Qi Shen and Xin Wang.

An analysis of the optimization disaggregation algorithm in the estimation related to energy consumption of appliances in buildings. Applied Mathematics and Computation, 234(??):506–519, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400294X.

Song:2014:DDC

[SW14b] M. L. Song and S. H.

Wang. DEA decomposition of China’s environmental efficiency based on search algorithm. Applied Mathematics and Computation, 247(??):562–572, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314012752.

Song:2014:SCM

[SW14c] Yisheng Song and Hongjun Wang. Strong convergence for the modified Mann’s iteration of λ-strict pseudocontraction. Applied Mathematics and Computation, 237(??):405–410, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004597.

Sui:2014:USP

[SW14d] Meiyu Sui and Yejuan Wang. Upper semicontinuity of pullback attractors for lattice nonclassical diffusion delay equations under singular perturbations. Applied Mathematics and Computation, 242(??):315–327, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314007103.

Smit:2011:TPF

[SWD11] G. J. F. Smit, J. M. Wilms, and G. P. J. Diedericks. Two-phase flow modeling for low concentration spherical particle motion through a Newtonian fluid. Applied Mathematics and Computation, 217(11):5068–5077, February 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310008003.

Sadat:2012:MMS

[SWL12] Hamou Sadat, Cheng-An Wang, and Vital Le Dez. Meshless method for solving coupled radiative and conductive heat transfer in complex multi-dimensional geometries. Applied Mathematics and Computation, 218 (20):10211–10225, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003645.

Sun:2013:DAA

[SWLP13] Gang Sun, Dan Wang, Xiaoqiang Li, and Zhouhua Peng. A DSC approach to adaptive neural network tracking control for pure-feedback nonlinear systems. Applied Mathematics and Computation, 219(11):6224–6235, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300312013100.

Shi:2014:NSI

[SWW14] Cong Shi, Chen Wang, and Ting Wei. Numerical solution for an inverse heat source problem by an iterative method. Applied Mathematics and Computation, 244(??):577–597, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314009953.

Shi:2010:CCG

[SWX10a] Zhen-Jun Shi, Shengquan Wang, and Zhiwei Xu. The convergence of conjugate gradient method with nonmonotone line search. Applied Mathematics and Computation, 217(5):1921–1932, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310007320.

Su:2010:FDM

[SWX10b] Lijuan Su, Wenqia Wang, and Qiuyan Xu. Finite difference methods for fractional dispersion equations. Applied Mathematics and Computation, 216(11):3329–3334, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004881.

Shen:2014:ICR

[SX14] Yuan Shen and Minghua Xu. On the O(1/t) convergence rate of Ye-Yuan’s modified alternating direction method of multipliers. Applied Mathematics and Computation, 226(??):367–373, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011028.

Su:2012:MCM

[SXC^{+}12] Fang Su, Zhan Xu, Jun-Zhi Cui, Xin-Peng Du, and Hao Jiang. Multiscale computation method for an advection-diffusion equation. Applied Mathematics and Computation, 218(14):7369–7374, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014597.

Su:2011:MSM

[SXCD11] F. Su, Z. Xu, J. Z. Cui, and Q. L. Dong. Multi-scale method for the quasi-periodic structures of composite materials. Applied Mathematics and Computation, 217(12):5847–5852, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310012701.

Su:2011:MCM

[SXDJ11] Fang Su, Zhan Xu, Qiao-Li Dong, and Hao Jiang. Multiscale computation method for parabolic problems of composite materials. Applied Mathematics and Computation, 217(21):8337–8342, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311003821.

Sun:2011:MRA

[SXGR11] Zhixin Sun, Yuhua Xu, Jing Gong, and Zhiguang Ren. A multicast routing algorithm based on searching in directed graph. Applied Mathematics and Computation, 218(7):3723–3732, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311011830.

Sun:2011:BDE

[SXH11] Taixiang Sun, Hongjian Xi, and Qiuli He. On boundedness of the difference equation x_{n+1} = p_{n} +x_{n−3k+1}/x_{n−k+1} with period-k coefficients. Applied Mathematics and Computation, 217(12): 5994–5997, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300310012385.

Shi:2012:EWS

[SXL12] Qihong Shi, Qian Xiao, and Xiaojun Liu. Extended wave solutions for a nonlinear Klein–Gordon– Zakharov system. Applied Mathematics and Computation, 218(19):9922–9929, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003372.

Shao:2013:CWF

[SXL13] Xinhui Shao, Guanyu Xue, and Changjun Li. A conservative weighted finite difference scheme for regularized long wave equation. Applied Mathematics and Computation, 219(17):9202–9209, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313003299.

Shen:2010:PET

[SY10a] Chunxia Shen and Minsheng You. Permanence and extinction of a threespecies ratio-dependent food chain model with delay and prey diffusion. Applied Mathematics and Computation, 217(5):1825–1830, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310002092.

Shoukaku:2010:ONH

[SY10b] Yutaka Shoukaku and Norio Yoshida. Oscillations of nonlinear hyperbolic equations with functional arguments via Riccati method. Applied Mathematics and Computation, 217(1):143–151, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005746.

Song:2010:ETW

[SY10c] Ming Song and Chenxi Yang. Exact traveling wave solutions of the Zakharov–Kuznetsov– Benjamin–Bona–Mahony equation. Applied Mathematics and Computation, 216(11):3234–3243, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004765.

Soderlind:2012:ISG

[SY12] Gustaf Söderlind and Arjun Singh Yadaw. The impact of smooth W -grids in the numerical solution of singular perturbation two-point boundary value problems. Applied Mathematics and Computation, 218(10):6045–6055, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014329.

Shen:2013:SCC

[SY13a] Tao Shen and Zhugang Yuan. Stability criterion for a class of fixedpoint digital filters using two’s complement arithmetic. Applied Mathematics and Computation, 219(9):4880–4883, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312010806.

Shi:2013:SSF

[SY13b] Dongyang Shi and Zhiyun Yu. Superclose and superconvergence of finite element discretizations for the Stokes equations with damping. Applied Mathematics and Computation, 219(14):7693–7698, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313000982.

Shi:2014:PAO

[SY14a] Qiuhong Shi and Xiaoping Yang. Pricing Asian options in a stochastic volatility model with jumps. Applied Mathematics and Computation, 228(??):411–422, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012782.

Sogabe:2014:NFB

[SY14b] Tomohiro Sogabe and Fatih Yilmaz. A note on a fast breakdown-free algorithm for computing the determinants and the permanents of k-tridiagonal matrices. Applied Mathematics and Computation, 249(??):98–102, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314014064.

Solis:2014:CPP

[SY14c] Francisco J. Solis and Carlos Yebra. Cyclic pursuit problems in the two dimensional sphere. Applied Mathematics and Computation, 244(??):91–100, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314009394.

Song:2014:BIG

[SY14d] Guang-Jing Song and Shaowen Yu. The block independence in the generalized inverse A_{T,S}^{(2)} for some ordered matrices and applications. Applied Mathematics and Computation, 232(??):399–410, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314000125.

Song:2013:OPI

[SYB13] Hua-Ming Song, Hui Yang, and Alain Bensoussan. Optimizing production and inventory decisions in a supply chain with lot size, production rate and lead time interactions. Applied Mathematics and Computation, 224(??):150–165, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313009223.

Shah:2010:UCF

[SYK10] Abdullah Shah, Li Yuan, and Aftab Khan. Upwind compact finite difference scheme for time-accurate solution of the incompressible Navier–Stokes equations. Applied Mathematics and Computation, 215(9):3201–3213, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300309008947.

Shen:2014:RDC

[SYl14] Peiping Shen, Linpeng Yang, and Yanchao liang. Range division and contraction algorithm for a class of global optimization problems. Applied Mathematics and Computation, 242(??):116–126, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314006894.

Shi:2013:CSM

[SYW13] Lei Shi, Jing-Ping Yi, and Zhi-Gang Wang. On certain subclass of meromorphic close-to-convex functions. Applied Mathematics and Computation, 219 (20):10143–10149, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313003482.

Song:2010:ESWa

[SYZ10] Ming Song, Chenxi Yang, and Bengong Zhang. Exact solitary wave solutions of the Kadomtsov– Petviashvili–Benjamin–Bona– Mahony equation. Applied Mathematics and Computation, 217(4):1334–1339, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300309004445.

Shi:2011:HAA

[SZ11] Dongyang Shi and Yadong Zhang. High accuracy analysis of a new nonconforming mixed finite element scheme for Sobolev equations. Applied Mathematics and Computation, 218(7):3176–3186, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311010940.

Shen:2014:DHB

[SZ14a] Zuolin Shen and Chunrui Zhang. Double Hopf bifurcation of coupled dissipative Stuart–Landau oscillators with delay. Applied Mathematics and Computation, 227(??):553–566, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313012125.

Shi:2014:ASA

[SZ14b] Yanchao Shi and Peiyong Zhu. Asymptotic stability analysis of stochastic reaction-diffusion Cohen– Grossberg neural networks with mixed time delays. Applied Mathematics and Computation, 242(??):159–167, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007218.

Smaoui:2014:DCD

[SZ14c] Nejib Smaoui and Mohamed Zribi. Dynamics and control of the 2-D Navier–Stokes equations. Applied Mathematics and Computation, 237(??):461–473, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005219.

Sun:2014:GSS

[SZ14d] Gui-Quan Sun and Zi-Ke Zhang. Global stability for a sheep brucellosis model with immigration. Applied Mathematics and Computation, 246(??):336–345, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011229.

Szablowski:2013:ART

[Sza13] Paweł J. Szablowski. On affinity relating two positive measures and the connection coefficients between polynomials orthogonalized by these measures. Applied Mathematics and Computation, 219(12):6768–6776, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313000246.

Szepietowski:2010:FTE

[Sze10] Andrzej Szepietowski. Faulttolerant edge and vertex pancyclicity in alternating group graphs. Applied Mathematics and Computation, 217(6):2827–2832, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300310008726.

Szepietowski:2011:FTV

[Sze11] Andrzej Szepietowski. Fault tolerance of vertex pancyclicity in alternating group graphs. Applied Mathematics and Computation, 217(16):6785–6791, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031100169X.

Szepietowski:2012:FTE

[Sze12] Andrzej Szepietowski. Fault tolerance of edge pancyclicity in alternating group graphs. Applied Mathematics and Computation, 218(19):9875–9881, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312003220.

Santos:2012:NEM

[SZG12] Juan E. Santos, Fabio I.

Zyserman, and Patricia M. Gauzellino. Numerical electroseismic modeling: a finite element approach. Applied Mathematics and Computation, 218(11):6351–6374, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014615.

Sheng:2013:UOC

[SZH13] Linxue Sheng, Yuanguo Zhu, and Timo Hamalainen. An uncertain optimal control model with Hurwicz criterion. Applied Mathematics and Computation, 224(??):412–421, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313009478.

Sun:2012:PFS

[SZS^{+}12] Gui-Quan Sun, Juan Zhang, Li-Peng Song, Zhen Jin, and Bai-Lian Li. Pattern formation of a spatial predator-prey system. Applied Mathematics and Computation, 218(22):11151–11162, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004882.

Shi:2013:NLO

[SZS13] Dongyang Shi, Jiaquan Zhou, and Dongwei Shi. A new low order least squares nonconforming characteristics mixed finite element method for Burgers’ equation. Applied Mathematics and Computation, 219(24):11302– 11310, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300313005717.

Shi:2014:QWN

[SZW14] Dongyang Shi, Yanmin Zhao, and Fenling Wang. Quasi-Wilson nonconforming element approximation for nonlinear dual phase lagging heat conduction equations. Applied Mathematics and Computation, 243(??):454–464, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314007838.

Si:2011:MSL

[SZZ^{+}11] Xinhui Si, Liancun Zheng, Xinxin Zhang, Min Li, Jianhong Yang, and Ying Chao. Multiple solutions for the laminar flow in a porous pipe with suction at slowly expanding or contracting wall. Applied Mathematics and Computation, 218(7):3515–3521, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311011398.

Shi:2013:GES

[SZZH13] Lin Shi, Hong Zhu, Shouming Zhong, and Liyuan Hou. Globally exponential stability for neural networks with timevarying delays. Applied Mathematics and Computation, 219(21):10487– 10498, July 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313004505.

Temimi:2011:NIT

[TA11] H. Temimi and A. R. Ansari. A new iterative technique for solving nonlinear second order multipoint boundary value problems. Applied Mathematics and Computation, 218(4):1457–1466, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0096300311008563.

Temimi:2013:EAD

[TA13] Helmi Temimi and Slimane Adjerid. Error analysis of a discontinuous Galerkin method for systems of higher-order differential equations. Applied Mathematics and Computation, 219(9):4503–4525, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312010752.

Tadi:2010:EDC

[Tad10] M. Tadi. Evaluation of diffusion coefficient based on multiple forward solutions. Applied Mathematics and Computation, 216(12):3707–3717, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310005710.

Tadi:2012:FGL

[Tad12] M. Tadi. A fixed-grid local method for 1-D Stefan problems. Applied Mathematics and Computation, 219(4):2331–2341, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031200817X.

Tagliani:2011:HMP

[Tag11] A. Tagliani. Hausdorff moment problem and fractional moments: a simplified procedure. Applied Mathematics and Computation, 218(8):4423–4432, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311012495.

Tagliani:2014:HMP

[Tag14]  A. Tagliani.  Ham-

burger moment problem and Maximum Entropy: On the existence conditions. Applied Mathematics and Computation, 231(??):111–116, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313014434.

Tang:2014:TRA

[Tan14] Mingyun Tang. Trust region algorithm with two subproblems for bound constrained problems. Applied Mathematics and Computation, 242(??): 778–789, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314009187.

Triki:2014:SSK

[TB14] Houria Triki and Noureddine Boucerredj. Soliton solutions of the Klein– Gordon–Zakharov equations with power law nonlinearity. Applied Mathematics and Computation, 227(??):341–346, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313011600.

Totur:2013:SMI

[TÇ13] Ümit Totur and Ibrahim Çanak. On the (C, 1) summability method of improper integrals. Applied Mathematics and Computation, 219(24): 11065–11070, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300313005468.

Ting:2014:SFO

[TC14a] Pin-Shou Ting and Kun-Jen Chung. Some formulations for optimal solutions with delays in payment and price-discount offers in the supply chain system. Applied Mathematics and Computation, 230(??):180–192, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313013520.

Tsai:2014:GVG

[TC14b] Huai-Yin Tsai and Jein-Shan Chen. Geometric views of the generalized Fischer-Burmeister function and its induced merit function. Applied Mathematics and Computation, 237(??):31–59, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314004512.

Teng:2011:ELS

[TCBL11] Jinn-Tsair Teng, Leopoldo Eduardo Cárdenas-Barrón, and Kuo-Ren Lou. The economic lot size of the integrated vendor– buyer inventory system derived without derivatives: a simple derivation. Applied Mathematics and Computation, 217(12):5972–5977, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031001218X.

Tchuindjo:2011:CFS

[Tch11] Leonard Tchuindjo. Closedform solutions for pricing credit-risky bonds and bond options. Applied Mathematics and Computation, 217(13):6133–6143, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310013093.

Tosic:2012:ILC

[TCI12] Marina Tosíc and Dragana S. Cvetkovíc-Ilíc. Invertibility of a linear combination of two matrices and partial orderings. Applied Mathematics and Computation, 218(9):4651–4657, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012987.

Tripathy:2013:MTD

[TD13] A. K. Tripathy and G. Devi. Mixed type duality for nondifferentiable multiobjective fractional programming under generalized (dρηθ)-type 1 univex function. Applied Mathematics and Computation, 219(17):9196–9201, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031300297X.

Tavelli:2014:HOS

[TD14a] Maurizio Tavelli and Michael Dumbser. A high order semi-implicit discontinuous Galerkin method for the two dimensional shallow water equations on staggered unstructured meshes. Applied Mathematics and Computation, 234(??):623–644, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002719.

Tavelli:2014:SSI

[TD14b] Maurizio Tavelli and Michael Dumbser. A staggered semi-implicit discontinuous Galerkin method for the two dimensional incompressible Navier– Stokes equations. Applied Mathematics and Computation, 248(??):70–92, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314013149.

Tavana:2014:OIA

[TDSA14] Madjid Tavana, Debora Di Caprio, and Francisco J. Santos-Arteaga. An optimal information acquisition model for competitive advantage in complex multiperspective environments. Applied Mathematics and Computation, 240(??):175–199, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314006146.

Toscano:2012:PPC

[TDT12] E. Toscano, G. Di Blasi, and A. Tortorici. The Poisson problem: a comparison between two approaches based on SPH method. Applied Mathematics and Computation, 218(17):8906–8916, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312001890.

Tang:2014:NNI

[TDZF14] Jingyong Tang, Li Dong, Jinchuan Zhou, and Liang Fang. A new non-interior continuation method for solving the second-order cone complementarity problem. Applied Mathematics and Computation, 236(??):287–299, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314004020.

Temimi:2012:DGF

[Tem12] H. Temimi. A discontinuous Galerkin finite element method for solving the Troesch’s problem. Applied Mathematics and Computation, 219(2):521–529, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312006364.

Tabata:2014:EUC

[TES14] Minoru Tabata, Nobuoki Eshima, and Yuusuke Sakai. Existence, uniqueness, and computation of short-run and longrun equilibria of the Dixit–Stiglitz–Krugman model in an urban setting. Applied Mathematics and Computation, 234(??):339–355, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002537.

Tang:2010:FLI

[TF10] Lei-Yu Tang and Jian-Cong Fan. A family of Liouville integrable lattice equations and its conservation laws. Applied Mathematics and Computation, 217(5):1907–1912, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007307.

Tian:2013:GRI

[TFM13] Yazhou Tian, Min Fan, and Fanwei Meng. A generalization of retarded integral inequalities in two independent variables and their applications. Applied Mathematics and Computation, 221(??):239–248, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007017.

Tewa:2012:MAG

[TFMB12] J. J. Tewa, R. Fokouop, B. Mewoli, and S. Bowong. Mathematical analysis of a general class of ordinary differential equations coming from within-hosts models of malaria with immune effectors. Applied Mathematics and Computation, 218(14):7347–7361, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312000744.

Tang:2014:SSA

[TGL14] Qiulin Tang, Jing Ge, and Zhigui Lin. An SEI–SI avian-human influenza model with diffusion and nonlocal delay. Applied Mathematics and Computation, 247(??):753–761, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012673.

Trejo-Guerra:2013:SID

[TGTCCGRG13] R. Trejo-Guerra, E. Tlelo-Cuautle, V. H. Carbajal-Gómez, and G. Rodriguez-Gómez. A survey on the integrated design of chaotic oscillators. Applied Mathematics and Computation, 219(10):5113–5122, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011757.

Thanh:2014:TWV

[TH14a] Mai Duc Thanh and Nguyen Huu Hiep. On traveling waves in viscouscapillary Euler equations with thermal conductivity. Applied Mathematics and Computation, 234(??):127–141, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314002367.

Tian:2014:NUB

[TH14b] Gui-Xian Tian and Ting-Zhu Huang. A note on upper bounds for the spectral radius of weighted graphs. Applied Mathematics and Computation, 243(??):392–397, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314008066.

Triki:2011:SWS

[THAB11] Houria Triki, T. Hayat, Omar M. Aldossary, and Anjan Biswas. Solitary wave and shock wave solutions to a second order wave equation of Korteweg–de Vries type. Applied Mathematics and Computation, 217(21):8852–8855, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004401.

Teng:2011:OOD

[THCW11] Hui-Ming Teng, Ping-Hui Hsu, Yufang Chiu, and Hui Ming Wee. Optimal ordering decisions with returns and excess inventory. Applied Mathematics and Computation, 217(22):9009–9018, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311004759.

Tang:2011:SNM

[THDF11] Jingyong Tang, Guoping He, Li Dong, and Liang Fang. A smoothing Newton method for secondorder cone optimization based on a new smoothing function. Applied Mathematics and Computation, 218(4):1317–1329, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311008289.

Tang:2010:BTW

[THH10] Shengqiang Tang, Xiaoliang Huang, and Wentao Huang. Bifurcations of travelling wave solutions for the generalized KP– BBM equation. Applied Mathematics and Computation, 216(10):2881–2890, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310004054.

Thai-Hoang:2011:AAF

[THNTNXR11] C. Thai-Hoang, N. Nguyen-Thanh, H. Nguyen-Xuan, and T. Rabczuk. An alternative alpha finite element method with discrete shear gap technique for analysis of laminated composite plates. Applied Mathematics and Computation, 217(17):7324–7348, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311002001.

Thuan:2014:ESN

[THP14] M. V. Thuan, L. V. Hien, and V. N. Phat. Exponential stabilization of nonautonomous delayed neural networks via Riccati equations. Applied Mathematics and Computation, 246(??):533–545, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011540.

Thukral:2010:NEO

[Thu10] R. Thukral. A new eighthorder iterative method for solving nonlinear equations. Applied Mathematics and Computation, 217(1):222–229, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310006077.

Tu:2014:ABS

[THX14] Jianjun Tu, Hanlin He, and Ping Xiong. Adaptive backstepping synchronization between chaotic systems with unknown Lipschitz constant. Applied Mathematics and Computation, 236(??):10–18, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003658.

Tseng:2010:IOI

[THYC10] Kuei-Lin Tseng, Shiow-Ru Hwang, Gou-Sheng Yang, and Yi-Ming Chou. Improvements of the Ostrowski integral inequality for mappings of bounded variation I. Applied Mathematics and Computation, 217(6):2348–2355, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310007794.

Tang:2012:TWS

[THZF12] Jiashi Tang, Feng Han, Minghua Zhao, and Wenbin Fu. Travelling wave solutions for the (2 + 1) dimensional Nizhnik– Novikov–Veselov equation. Applied Mathematics and Computation, 218(22):11083–11088, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312004791.

Triki:2010:SWS

[TI10a] H. Triki and M. S. Ismail. Solitary wave solutions for a coupled pair of mKdV equations. Applied Mathematics and Computation, 217(4):1540–1548, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300309006110.

Triki:2010:SSB

[TI10b] H. Triki and M. S. Ismail. Soliton solutions of a BBM(m, n) equation with generalized evolution. Applied Mathematics and Computation, 217(1):48–54, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310006223.

Tien:2012:RGP

[Tie12] Tzu-Li Tien. A research on the grey prediction model GM(1,n). Applied Mathematics and Computation, 218(9):4903–4916, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311013014.

Tatsumi:2013:CPS

[TIT13] Keiji Tatsumi, Takeru Ibuki, and Tetsuzo Tanino. A chaotic particle swarm optimization exploiting a virtual quartic objective function based on the personal and global best solutions. Applied Mathematics and Computation, 219(17):8991–9011, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313002828.

Tojtovska:2012:GDS

[TJ12] Biljana Tojtovska and Svetlana Janković. On a general decay stability of stochastic Cohen– Grossberg neural networks with time-varying delays. Applied Mathematics and Computation, 219(4):2289–2302, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312008697.

See corrigendum [TJ13].

Tojtovska:2013:CSG

[TJ13] Biljana Tojtovska and Svetlana Janković. Corrigendum on “On a general decay stability of stochastic Cohen–Grossberg neural networks with timevarying delays” [Applied Mathematics and Computation 219 (2012) 2289– 2302]. Applied Mathematics and Computation, 219(10):4963, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300312012246.

See [TJ12].

Tojtovska:2014:SSP

[TJ14] Biljana Tojtovska and Svetlana Janković. On some stability problems of impulsive stochastic Cohen–Grossberg neural networks with mixed time delays. Applied Mathematics and Computation, 239(??):211–226, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005712.

Teng:2013:EMR

[TJZ13] Kaimin Teng, Hongen Jia, and Haifeng Zhang. Existence and multiplicity results for fractional differential inclusions with Dirichlet boundary conditions. Applied Mathematics and Computation, 220(??):792–801, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312008430.

Torabi:2014:ECF

[TJZ14] Keivan Torabi, Adel Jafarzadeh Jazi, and Ehsan Zafari. Exact closed form solution for the analysis of the transverse vibration modes of a Timoshenko beam with multiple concentrated masses. Applied Mathematics and Computation, 238(??):342–357, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005438.

Touma:2012:WBU

[TK12] R. Touma and S. Khankan.

Well-balanced unstaggered central schemes for one and two-dimensional shallow water equation systems. Applied Mathematics and Computation, 218(10):5948–5960, January 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311014020.

Toker:2013:PTP

[TK13] Selma Toker and Selahattin Kaçiranlar. On the performance of two parameter ridge estimator under the mean square error criterion. Applied Mathematics and Computation, 219(9):4718–4728, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312011046.

Tarazaga:2014:CMC

[TK14a] Pablo Tarazaga and Hiroshi Kurata. On cell matrices: a class of Euclidean distance matrices. Applied Mathematics and Computation, 238(??):468–474, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314005591.

Temimi:2014:AAA

[TK14b] H. Temimi and H. Kurkcu.

An accurate asymptotic approximation and precise numerical solution of highly sensitive Troesch’s problem. Applied Mathematics and Computation, 235(??):253–260, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003750.

Thanh:2012:RNM

[TKC12] Mai Duc Thanh, Dietmar Kröner, and Christophe Chalons. A robust numerical method for approximating solutions of a model of two-phase flows and its properties. Applied Mathematics and Computation, 219(1):320–344, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300312006212.

Tamizhmani:2014:ARE

[TKL14] K. M. Tamizhmani, K. Krishnakumar, and P. G. L.

Leach. Algebraic resolution of equations of the Black–Scholes type with arbitrary time-dependent parameters. Applied Mathematics and Computation, 247(??):115–124, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314011977.

Tarasenko:2010:MSR

[TKLH10] Anna Tarasenko, Aleksandr Karelin, Gilberto Pérez Lechuga, and Manuel González Hernández. Modelling systems with renewable resources based on functional operators with shift. Applied Mathematics and Computation, 216(7):1938–1944, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002651.

Tuglu:2011:BFL

[TKS11] Naim Tuglu, E. Gokcen Kocer, and Alexey Stakhov. Bivariate Fibonacci like p-polynomials. Applied Mathematics and Computation, 217(24): 10239–10246, August 15, 2011. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300311006771.

Tan:2010:DDP

[TL10] Qianrong Tan and Zongbing Lin. Divisibility of determinants of power GCD matrices and power LCM matrices on finitely many quasi-coprime divisor chains. Applied Mathematics and Computation, 217(8):3910–3915, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310010131.

Teng:2014:TRC

[TL14] Yuan-Hsiang Teng and Cheng-Kuan Lin. A test round controllable local diagnosis algorithm under the PMC diagnosis model. Applied Mathematics and Computation, 244(??):613–623, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S009630031400993X.

Tan:2013:IGD

[TLDZ13] Qimeng Tan, Naiguang Lu, Mingli Dong, and Lianqing Zhu. Influence of geometrical distribution of common points on the accuracy of coordinate transformation. Applied Mathematics and Computation, 221(??):411–423, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313007200.

Tan:2013:DDP

[TLL13] Qianrong Tan, Miao Luo, and Zongbing Lin. Determinants and divisibility of power GCD and power LCM matrices on finitely many coprime divisor chains. Applied Mathematics and Computation, 219(15):8112–8120, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001501.

Li:2010:DSC

[tLpOjCW10] Yao tang Li, Shun ping Ouyang, Shu juan Cao, and Rui-Wu Wang. On diagonal-Schur complements of block diagonally dominant matrices. Applied Mathematics and Computation, 216(5):1383–1392, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300310002109.

Tao:2011:GMQ

[TLW11] Li Tao, Zaiming Liu, and Zhizhong Wang. The GI/M/1 queue with startup period and single working vacation and Bernoulli vacation interruption. Applied Mathematics and Computation, 218(8):4401–4413, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300311012471.

Tian:2012:DDS

[TLZZ12] Junkang Tian, Yongming Li, Jinzhou Zhao, and Shouming Zhong. Delay-dependent stochastic stability criteria for Markovian jumping neural networks with modedependent time-varying delays and partially known transition rates. Applied Mathematics and Computation, 218(9):5769–5781, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300311014330.

Takasi:2011:PAB

[TM11] Kusano Takaŝi and Jelena Manojlović. Precise asymptotic behavior of solutions of the sublinear Emden–Fowler differential equation. Applied Mathematics and Computation, 217(9):4382–4396, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300310010647.

Tagliani:2013:LTF

[TM13] Aldo Tagliani and Mariyan Milev. Laplace transform and finite difference methods for the Black–Scholes equation. Applied Mathematics and Computation, 220(??):649–658, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300313007613.

Thakur:2014:MRC

[TMJ14] Manoj Thakur, Suraj S. Meghwani, and Hemant Jalota. A modified real coded genetic algorithm for constrained optimization. Applied Mathematics and Computation, 235(??):292–317, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300314003415.

Takasi:2013:ISS

[TMM13] Kusano Takaŝi, Jelena V.

Manojlović, and Jelena Milosević. Intermediate solutions of second order quasilinear ordinary differential equations in the framework of regular variation. Applied Mathematics and Computation, 219(15):8178–8191, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0096300313001239.

Takasi:2014:ISF

[TMM14] Kusano Taka^{s}i, Jelena V.Manojlović, andJelenaMilosević.Intermediatesolutionsoffourthorderquasilineardifferentialequationsintheframeworkofregularvariation.AppliedMathematicsandComputation, 248 (??) : 246−−272, December1, 2014.CODENAMHCBQ.ISSN0096−

3003(print), 1873−5649(electronic).URLhttpwwwsciencedirectcomsciencearticlepiiS0096300314013435.

Tang:2011:WDS

[TMXG11] Yaning Tang, Wen-Xiu Ma, Wei Xu, and Liang Gao. Wronskian determinant solutions of the (3 + 1)-dimensional Jimbo–Miwa equation. Applied Mathematics and Computation, 217(21):8722–8730, July 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311004887.

Toutounian:2014:NMB

[TN14] F. Toutounian and H. Nasabzadeh. A new method based on generalized Taylor expansion for computing a series solution of the linear systems. Applied Mathematics and Computation, 248(??):602–609, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031401368X.

Taleizadeh:2011:MBM

[TNB11] Ata Allah Taleizadeh, Seyed Taghi Akhavan Niaki, and Farnaz Barzinpour. Multiplebuyer multiple-vendor multi-product multiconstraint supply chain problem with stochastic demand and variable lead-time: a harmony search algorithm. Applied Mathematics and Computation, 217(22):9234–9253, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031100542X.

Tomasiello:2012:SRN

[Tom12] S. Tomasiello. Some remarks on a new DQ-based method for solving a class of Volterra integro-differential equations. Applied Mathematics and Computation, 219 (1):399–407, September 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312006303. See corrigendum

[Tom15].

Tomaschitz:2013:BIE

[Tom13] Roman Tomaschitz. Bessel integrals in epsilon expansion: Squared spherical Bessel functions averaged with Gaussian power-law distributions. Applied Mathematics and Computation, 225(??):228–241, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010072.

Tomasiello:2014:NDB

[Tom14] S. Tomasiello. A new dq-based method and its application to the stability of columns. Applied Mathematics and Computation, 226(??):145–156, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010709.

Tomasiello:2015:CSR

[Tom15] Stefania Tomasiello. Corrigendum to “Some remarks on a new DQ-based method for solving a class of Volterra integro-differential equations” [Appl. Math. Comput. 219 (2012) 399– 407]. Applied Mathematics and Computation, 253(??):410, February 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300315000089. See [Tom12].

Tosic:2014:SLC

[Tos14] Marina Tosić. On some linear combinations of commuting involutive and idempotent matrices. Applied Mathematics and Computation, 233(??):103–108, May 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314001556.

Tan:2012:MAP

[TP12] Hu Tan and Minfang Peng. Minimization of ambiguity in parametric fault diagnosis of analog circuits: a complex network approach. Applied Mathematics and Computation, 219(1):408–415, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312006327.

Thangaraj:2010:NMS

[TPA10] Radha Thangaraj, Millie Pant, and Ajith Abraham. New mutation schemes for differential evolution algorithm and their application to the optimization of directional overcurrent relay settings. Applied Mathematics and Computation, 216(2):532–544, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310000950.

Thangaraj:2011:PSO

[TPAB11] Radha Thangaraj, Millie Pant, Ajith Abraham, and Pascal Bouvry. Particle swarm optimization: Hybridization perspectives and experimental illustrations. Applied Mathematics and Computation, 217(12):5208–5226, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300310012555.

Tripathi:2010:PFV

[TPD10] D. Tripathi, S. K. Pandey, and S. Das. Peristaltic flow of viscoelastic fluid with fractional Maxwell model through a channel. Applied Mathematics and Computation, 215 (10):3645–3654, January 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630030900976X.

Thuan:2012:DOF

[TPT12] M. V. Thuan, V. N. Phat, and H. M. Trinh. Dynamic output feedback guaranteed cost control for linear systems with interval time-varying delays in states and outputs. Applied Mathematics and Computation, 218(21):10697–10707, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312004389.

Tan:2014:GST

[TPTW14] Hu Tan, Minfang Peng, Chi K. Tse, and Fangrong Wu. Global similarity tests of physical designs of circuits: a complex network approach. Applied Mathematics and Computation, 230(??):96–103, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313013738.

Tassaddiq:2011:FTD

[TQ11] Asifa Tassaddiq and Asghar Qadir. Fourier transform and distributional representation of the generalized gamma function with some applications. Applied Mathematics and Computation, 218(3):1084–1088, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/

S0096300311004437. Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Tong:2010:SMS

[TQWZ10] Xiaojiao Tong, Liqun Qi, Felix Wu, and Hui Zhou. A smoothing method for solving portfolio optimization with CVaR and applications in allocation of generation asset. Applied Mathematics and Computation, 216(6):1723–1740, May 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309010789.

Tripathy:2013:MFI

[TR13] Binod Chandra Tripathy and Gautam Chandra Ray. Mixed fuzzy ideal topological spaces. Applied Mathematics and Computation, 220(??):602–607, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313006528.

Tracina:2012:FSC

[Tra12] Rita Tracinà. Fundamental solution in classical elasticity via Lie group method. Applied Mathematics and Computation, 218 (9):5132–5139, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311013257.

Tavakoli:2013:DSG

[TRA13] M. Tavakoli, F. Rahbarnia, and A. R. Ashrafi. Distribution of some graph invariants over hierarchical product of graphs. Applied Mathematics and Computation, 220 (??):405–413, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313006218.

Tresaco:2013:NAP

[TRE13] E. Tresaco, A. Riaguas, and A. Elipe. Numerical analysis of periodic solutions and bifurcations in the planetary annulus problem. Applied Mathematics and Computation, 225(??):645–655, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010862.

Trif:2012:DOT

[Tri12]  Damian Trif.  Direct operatorial tau

method for pantograph-type equations. Applied Mathematics and Computation, 219 (4):2194–2203, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312008582.

Tripathy:2014:HOD

[Tri14] A. K. Tripathy. Higher order duality in multiobjective fractional programming with square root term under generalized higher order (F, α, β, ρ, σ, d)-∨-type I univex functions. Applied Mathematics and Computation, 247(??):880–897, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314012612.

Tan:2010:OPC

[TS10a] Xueyuan Tan and Yongzhong Song. Optimal Parameters for 2-cyclic AOR. Applied Mathematics and Computation, 216(5):1428–1442, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300310002183.

Tasic:2010:DGI

[TS10b] Milan B. Tasić and Predrag S. Stanimirović. Differentiation of generalized inverses for rational and polynomial matrices. Applied Mathematics and Computation, 216(7):2092–2106, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300310002857.

Tsoulos:2010:EPM

[TS10c] Ioannis G. Tsoulos and Athanassios Stavrakoudis. Enhancing PSO methods for global optimization. Applied Mathematics and Computation, 216(10):2988–3001, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310004315.

Tang:2012:TFE

[TS12a] Wensheng Tang and Yajuan Sun. Time finite element methods: a unified framework for numerical discretizations of ODEs. Applied Mathematics and Computation, 219 (4):2158–2179, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312008557.

Tomovski:2012:FWE

[TS12b] Zivorad Tomovski and Trifce Sandev. Fractional wave equation with a frictional memory kernel of Mittag-Leffler type. Applied Mathematics and Computation, 218 (20):10022–10031, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312002949.

Toutounian:2013:IMC

[TS13] F. Toutounian and F. Soleymani. An iterative method for computing the approximate inverse of a square matrix and the Moore–Penrose inverse of a non-square matrix. Applied Mathematics and Computation, 224(??):671–680, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313009545.

Tang:2014:CRK

[TS14] Wensheng Tang and Yajuan Sun. Construction of Runge–Kutta type methods for solving ordinary differential equations. Applied Mathematics and Computation, 234 (??):179–191, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314002811.

Tsaur:2011:DRA

[Tsa11] Ruey-Chyn Tsaur. Decision risk analysis for an interval TOPSIS method. Applied Mathematics and Computation, 218 (8):4295–4304, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311012318.

Tsai:2013:RDD

[Tsa13a] Zhi-Ren Tsai. Robust digital design of continuous-time nonlinear control systems using adaptive prediction and randomlocal-optimal NARMAX model. Applied Mathematics and Computation, 221 (??):152–163, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313006929.

Tsao:2013:DCN

[Tsa13b] Yu-Chung Tsao. Distribution center network design under trade credits. Applied Mathematics and Computation, 222(??):356–364, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300313007789.

Tsai:2014:NBA

[Tsa14] Hsing-Chih Tsai. Novel Bees Algorithm: Stochastic self-adaptive neighborhood. Applied Mathematics and Computation, 247 (??):1161–1172, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314013046.

Tseng:2012:IOI

[Tse12]  Kuei-Lin Tseng.  Improvements of the

Ostrowski integral inequality for mappings of bounded variation II. Applied Mathematics and Computation, 218 (10):5841–5847, January 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311013907.

Taoudi:2010:ISM

[TSG10] M. A. Taoudi, N. Salhi, and B. Ghribi. Integrable solutions of a mixed type operator equation. Applied Mathematics and Computation, 216(4):1150–1157, April 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310001736.

Tsitouras:2011:GRB

[Tsi11a]  Ch. Tsitouras.  Greatest remainder bi-

proportional rounding and the Greek parliamentary elections of 2007. Applied Mathematics and Computation, 217(22):9254–9260, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311005431.

Tsitouras:2011:HHW

[Tsi11b] Ch. Tsitouras. Hybrid Hamilton–Webster and the Greek apportionment. Applied Mathematics and Computation, 218 (7):3957–3961, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311011817.

Tsitouras:2014:FMR

[Tsi14] Ch. Tsitouras. On fitted modifications of Runge-Kutta–Nyström pairs. Applied Mathematics and Computation, 232(??):416–423, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300314001064.

Tomashchuk:2013:MDW

[TSJ13] I. Tomashchuk, P. Sallamand, and J. M. Jouvard. The modeling of dissimilar welding of immiscible materials by using a phase field method. Applied Mathematics and Computation, 219(13):7103–7114, March 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031200077X.

Tasy:2013:NEO

[TSN13] Ashley Tasy, Gina-Louise Santamaria, and Bogdan G. Nita. Numerical evaluation of a one-dimensional inverse scattering algorithm for simultaneous seismic imaging and inversion. Applied Mathematics and Computation, 219(10):5049–5069, January 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312011666.

Tasgetiren:2010:EDD

[TSP10a] M. Fatih Tasgetiren, P. N. Suganthan, and Quan-Ke Pan. An ensemble of discrete differential evolution algorithms for solving the generalized traveling salesman problem. Applied Mathematics and Computation, 215(9):3356–3368, January 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309009321.

Tasic:2010:AGI

[TSP10b] Milan B. Tasić, Predrag S. Stanimirović, and Selver H. Pepić. About the generalized LMinverse and the weighted Moore–Penrose inverse. Applied Mathematics and Computation, 216(1):114–124, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310000329.

Togun:2014:NSL

[TSS^{+}14] Hussein Togun, M. R. Safaei, Rad Sadri, S. N. Kazi, A. Badarudin, K. Hooman, and E. Sadeghinezhad. Numerical simulation of laminar to turbulent nanofluid flow and heat transfer over a backward-facing step. Applied Mathematics and Computation, 239(??):153–170, July 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314005918.

Toutounian:2013:NBM

[TT13] F. Toutounian and E. Tohidi. A new Bernoulli matrix method for solving second order linear partial differential equations with the convergence analysis. Applied Mathematics and Computation, 223(??):298–310, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313008503.

Tang:2012:TNW

[TTM12] Yaning Tang, Junyi Tu, and Wen-Xiu Ma. Two new Wronskian conditions for the (3 + 1)-dimensional Jimbo–Miwa equation. Applied Mathematics and Computation, 218 (20):10050–10055, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312002950.

Tuan:2010:NCP

[TTQ10] Nguyen Huy Tuan, Dang Duc Trong, and Pham Hoang Quan. A note on a Cauchy problem for the Laplace equation: Regularization and error estimates. Applied Mathematics and Computation, 217(7):2913–2922, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310009665.

Tuan:2011:MIE

[TTQ11] Nguyen Huy Tuan, Dang Duc Trong, and Pham Hoang Quan. A modified integral equation method of the semilinear backward heat problem. Applied Mathematics and Computation, 217(12):5177–5185, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310011872.

Tsai:2013:GPS

[TTWL13] Hsing-Chih Tsai, Yaw-Yauan Tyan, Yun-Wu Wu, and Yong-Huang Lin. Gravitational particle swarm. Applied Mathematics and Computation, 219(17):9106–9117, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313003597.

Tuncer:2012:NRB

[TÜ12] Yilmaz Tunçer and Serpil Ünal. New representations of Bertrand pairs in Euclidean 3-space. Applied Mathematics and Computation, 219(4):1833–1842, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312008168.

Tuan:2013:CSN

[Tua13] Nguyen Huy Tuan. Corrigendum to: “A new regularized method for two dimensional nonhomogeneous backward heat problem” [Appl. Math. Comput. 215 (2009) 873– 880]. Applied Mathematics and Computation, 219(9):4931–4932, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312010594. See [?].

Tuna:2013:SPD

[Tun13] Hüseyin Tuna. On spectral properties of dissipative fourth order boundary-value problem with a spectral parameter in the boundary condition. Applied Mathematics and Computation, 219(17):9377–9387, May 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313002634.

Tuna:2014:CRD

[Tun14] Hüseyin Tuna. Completeness of the rootvectors of a dissipative Sturm–Liouville operators on time scales. Applied Mathematics and Computation, 228(??):108–115, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313012411.

Turkyilmazoglu:2014:EAN

[Tur14a] M. Turkyilmazoglu. An effective approach for numerical solutions of high-order Fredholm integro-differential equations. Applied Mathematics and Computation, 227 (??):384–398, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313011363.

Turkyilmazoglu:2014:HON

[Tur14b]  M. Turkyilmazoglu.  High-order nonlin-

ear Volterra–Fredholm–Hammerstein integrodifferential equations and their effective computation. Applied Mathematics and Computation, 247(??):410–416, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314011849.

Theaker:2013:HRE

[TV13] Kyle A. Theaker and Robert A. Van Gorder. Hyper-Riccati equations and integrable reductions permitting stationary solutions for complex hyperbolic field equations. Applied Mathematics and Computation, 219 (16):8525–8541, April 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313002579.

Tortosa:2010:MST

[TVZ10] L. Tortosa, J. F. Vicent, and A. Zamora. A model to simplify 2D triangle meshes with irregular shapes. Applied Mathematics and Computation, 216(10):2937–2946, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031000425X.

Triki:2010:BSS

[TW10a] Houria Triki and Abdul-Majid Wazwaz. Bright soliton solution to a generalized Burgers–KdV equation with time-dependent coefficients. Applied Mathematics and Computation, 217(2):466–471, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310006375.

Triki:2010:SSD

[TW10b] Houria Triki and Abdul-Majid Wazwaz. Soliton solutions for (2 + 1)-dimensional and (3 + 1)-dimensional K(m, n) equations. Applied Mathematics and Computation, 217 (4):1733–1740, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309010145.

Tian:2011:SAD

[TW11a] Yanling Tian and Peixuan Weng. Stability analysis of diffusive predator-prey model with modified Leslie–Gower and Holling-type III schemes. Applied Mathematics and Computation, 218(7):3733–3745, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311011842.

Triki:2011:DSC

[TW11b] Houria Triki and Abdul-Majid Wazwaz. Dark solitons for a combined potential KdV and Schwarzian KdV equations with t-dependent coefficients and forcing term. Applied Mathematics and Computation, 217(21):8846–8851, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300311004188.

Tian:2013:RBL

[TW13] Yongge Tian and Hongxing Wang. Relations between least-squares and least-rank solutions of the matrix equation AXB = C. Applied Mathematics and Computation, 219(20):10293–10301, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313003998.

Tang:2014:HCM

[TW14a] Ping Tang and Xiaohuan Wang. Hölder continuity for a modified Camassa–Holm system. Applied Mathematics and Computation, 234(??):63–68, May 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314002744.

Tao:2014:PTA

[TW14b] Youtian Tao and Dongyin Wang. Padé-type approximation method for two dimensional Fredholm integral equations of the second kind. Applied Mathematics and Computation, 246(??):71–78, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314010972.

Tian:2014:GSH

[TX14] Xiaohong Tian and Rui Xu. Global stability and Hopf bifurcation of an HIV-1 infection model with saturation incidence and delayed CTL immune response. Applied Mathematics and Computation, 237(??):146–154, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314004536.

Tian:2014:IDP

[TXX14] Junkang Tian, Wenjun Xiong, and Fang Xu. Improved delay-partitioning method to stability analysis for neural networks with discrete and distributed time-varying delays. Applied Mathematics and Computation, 233(??):152–164, May 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031400174X.

Tan:2014:CPF

[TYCZ14] Jieqing Tan, Yangang Yao, Huijuan Cao, and Li Zhang. Convexity preservation of fivepoint binary subdivision scheme with a parameter. Applied Mathematics and Computation, 245(??):279–288, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314010285.

Tollu:2014:FSS

[TYT14] D. T. Tollu, Y. Yazlik, and N. Taskara. On fourteen solvable systems of difference equations. Applied Mathematics and Computation, 233(??):310–319, May 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314002331.

Tian:2011:IDD

[TZ11] Junkang Tian and Shouming Zhong. Improved delay-dependent stability criterion for neural networks with time-varying delay. Applied Mathematics and Computation, 217 (24):10278–10288, August 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311006989.

Tian:2012:ASD

[TZJZ12] Shou-Fu Tian, Sheng-Wu Zhou, Wu-You Jiang, and Hong-Qing Zhang. Analytic solutions, Darboux transformation operators and supersymmetry for a generalized one-dimensional time-dependent Schrödinger equation. Applied Mathematics and Computation, 218(13):7308–7321, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312000288.

Tang:2014:RBO

[TZS14] Xiaomin Tang, Yang Zhang, and Qiong Sun. Rota–Baxter operators on 4-dimensional complex simple associative algebras. Applied Mathematics and Computation, 229 (??):173–186, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313013088.

Tang:2014:EPP

[TZYH14] Zhenjun Tang, Xianquan Zhang, Chunqiang Yu, and Dan He. Efficient point pattern matching algorithm for planar point sets under transform of translation, rotation and scale. Applied Mathematics and Computation, 232(??):624–631, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314001246.

Tang:2011:EAA

[TZZX11] Qian Tang, Zhihua Zhong, Guiyong Zhang, and Xu Xu. An efficient adaptive analysis procedure for node-based smoothed point interpolation method (NS-PIM). Applied Mathematics and Computation, 217(21):8387–8402, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311004024.

Udovicic:2013:SLI

[UB13] Z. Udovicić and D. Barrera. A second look at the interpolatory background of the Euler–Maclaurin quadrature formula. Applied Mathematics and Computation, 220 (??):608–615, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031300787X.

Uddin:2013:RPS

[Udd13] Marjan Uddin. RBF–PS scheme for solving the equal width equation. Applied Mathematics and Computation, 222(??):619–631, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313007819.

Udwadia:2010:PBE

[Udw10] Firdaus E. Udwadia. Photograph, biosketch and editorial for special issue. Applied Mathematics and Computation, 217(3):953–955, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300310005631.

Siraj-ul-islam:2014:NSC

[uiH14] Siraj ul islam and Nadeem Haider. Numerical solution of compartmental models by meshless and finite difference methods. Applied Mathematics and Computation, 238(??):408–435, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300314005384.

Umbarkar:2014:MPD

[UJH14] A. J. Umbarkar, M. S. Joshi, and Wei-Chiang Hong. Multithreaded Parallel Dual Population Genetic Algorithm (MPDPGA) for unconstrained function optimizations on multi-core system. Applied Mathematics and Computation, 243(??):936–949, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314008698.

Ufuktepe:2011:UCT

[UK11] Ünal Ufuktepe and Sinan Kapçak. Unification of calculus on Time Scales with Mathematica. Applied Mathematics and Computation, 218(3):1102–1106, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/

S0096300311003857. Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ursini:2014:SAS

[UMMT14] Edson L. Ursini, Paulo S. Martins, Regina L. Moraes, and Varese S. Timóteo. n-steps ahead software reliability prediction using the Kalman filter. Applied Mathematics and Computation, 245(??):116–134, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314009758.

Unal:2013:WMF

[Üna13] Basri Ünal. Weak magnetic field solution of Schrödinger equation in two dimensions and continuous density of states. Applied Mathematics and Computation, 225 (??):562–572, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313010370.

Uyanik:2010:SPF

[UOK10] Neslihan Uyanik, Shigeyoshi Owa, and Ekrem Kadioğlu. Some properties of functions associated with close-to-convex and starlike of order α. Applied Mathematics and Computation, 216(2):381–387, March 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310000354.

urRehman:2013:EUS

[uRE13] Mujeeb ur Rehman and Paul W. Eloe. Existence and uniqueness of solutions for impulsive fractional differential equations. Applied Mathematics and Computation, 224 (??):422–431, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313009569.

Upadhyay:2011:CLH

[URR11] R. K. Upadhyay, V. Rai, and S. N. Raw. Challenges of living in the harsh environments: a mathematical modeling study. Applied Mathematics and Computation, 217 (24):10105–10117, August 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311006618.

Ullah:2014:EMI

[USAF14] M. Zaka Ullah, F. Soleymani, and A. S. Al-Fhaid. An efficient matrix iteration for computing weighted Moore–Penrose inverse. Applied Mathematics and Computation, 226(??):441–454, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031301103X.

Ulukok:2013:IUB

[UT13] Zübeyde Ulukök and Ramazan Türkmen. Improved upper bounds for the solution of the continuous algebraic Riccati matrix equation. Applied Mathematics and Computation, 225(??):306–317, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010035.

Udwadia:2010:HPF

[UW10] Firdaus E. Udwadia and Thanapat Wanichanon. Hamel’s paradox and the foundations of analytical dynamics. Applied Mathematics and Computation, 217(3):1253–1265, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310002018.

Ucar:2011:IWT

[UY11] Faruk Uçar and Osman Yürekli. Identities for the Widder transform and transforms with Bessel function kernels. Applied Mathematics and Computation, 218(3):1096–1101, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311005133. Special Issue in

Honour of Hari M. Srivastava on his 70th birth anniversary.

Vidal:2012:SSS

[VA12] Antonio M. Vidal and Pedro Alonso. Solving systems of symmetric Toeplitz tridiagonal equations: Rojo’s algorithm revisited. Applied Mathematics and Computation, 219 (4):1874–1889, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312008235.

Valls:2011:DIN

[Val11] Claudia Valls. Darboux integrability of a nonlinear financial system. Applied Mathematics and Computation, 218(7):3297–3302, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S009630031101109X.

VanGorder:2010:HON

[Van10] Robert A. Van Gorder. High-order nonlinear boundary value problems admitting multiple exact solutions with application to the fluid flow over a sheet. Applied Mathematics and Computation, 216(7):2177–2182, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031000295X.

VanGorder:2011:ANB

[Van11a] Robert A. Van Gorder. Analysis of nonlinear BVPs motivated by fluid film flow over a surface. Applied Mathematics and Computation, 217(20):8068–8079, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311003420.

VanGorder:2011:CRC

[Van11b] Robert A. Van Gorder. Computing the region of convergence for power series in many real variables: a ratio-like test. Applied Mathematics and Computation, 218 (5):2310–2317, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311010009.

Vitela:2012:SNM

[VC12] Javier E. Vitela and Octavio Castaños. A sequential niching memetic algorithm for continuous multimodal function optimization. Applied Mathematics and Computation, 218(17):8242–8259, May 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007211.

Velimirovic:2014:GSC

[VC14a] Ljubica S. Velimirović and Milica D. Cvetković. Gaudi surfaces and curvature based functional variations. Applied Mathematics and Computation, 228 (??):377–383, February 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313012733.

Viswanathan:2014:FPM

[VC14b] P. Viswanathan and A. K. B. Chand. A fractal procedure for monotonicity preserving interpolation. Applied Mathematics and Computation, 247(??):190–204, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314009345.

Velimirovic:2012:AGS

[VCĆV12] Ljubica S. Velimirović, Milica D. Cvetković, Marija S. Ćirić, and Nikola Velimirović. Analysis of Gaudi surfaces at small deformations. Applied Mathematics and Computation, 218(13):6999–7004, March 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311014639.

Velimirovic:2013:VSO

[VCNV13] Ljubica S. Velimirović, Milica D. Cvetković, Marija S. Najdanović, and Nikola M. Velimirović. Variation of shape operator under infinitesimal bending of surface. Applied Mathematics and Computation, 225 (??):480–486, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313010059.

Vaz:2014:NRK

[VD14a] F. A. Vaz and A. L. De Bortoli. A new reduced kinetic mechanism for turbulent jet diffusion flames of bioethanol. Applied Mathematics and Computation, 247(??):918–929, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314012880.

Vitanov:2014:SWS

[VD14b] Nikolay K. Vitanov and Zlatinka I. Dimitrova. Solitary wave solutions for nonlinear partial differential equations that contain monomials of odd and even grades with respect to participating derivatives. Applied Mathematics and Computation, 247(??):213–217, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314012119.

Vita:2012:UIM

[VDFV12] Maria C. Vita, Samuele De Bartolo, Carmine Fallico, and Massimo Veltri. Usage of infinitesimals in the Menger’s Sponge model of porosity. Applied Mathematics and Computation, 218(16):8187–8195, April 15, 2012.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311008265.

Vitanov:2010:MMS

[VDK10] Nikolay K. Vitanov, Zlatinka I. Dimitrova, and Holger Kantz. Modified method of simplest equation and its application to nonlinear PDEs. Applied Mathematics and Computation, 216(9):2587–2595, July 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310003565.

Vitanov:2013:AMS

[VDK13] Nikolay K. Vitanov, Zlatinka I. Dimitrova, and Holger Kantz. Application of the method of simplest equation for obtaining exact traveling-wave solutions for the extended Korteweg–de Vries equation and generalized Camassa–Holm equation. Applied Mathematics and Computation, 219 (14):7480–7492, March 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031300074X.

Vargas-De-Leon:2012:LFT

[VDL12a] Cruz Vargas-De-León. Lyapunov functions for two-species cooperative systems. Applied Mathematics and Computation, 219 (5):2493–2497, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312008776.

Vargas-De-Leon:2012:SAM

[VDL12b] Cruz Vargas-De-León. Stability analysis of a model for HBV infection with cure of infected cells and intracellular delay. Applied Mathematics and Computation, 219 (1):389–398, September 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312006285.

Vargas-De-Leon:2014:CCG

[VDL14] Cruz Vargas-De-León. The complete classification for global dynamics of a model for the persistence of HTLV-I infection. Applied Mathematics and Computation, 237 (??):489–493, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314005098.

Vargas-De-Leon:2014:ADH

[VDLEK14] Cruz Vargas-De-León, Lourdes Esteva, and Andrei Korobeinikov. Age-dependency in host-vector models: the global analysis. Applied Mathematics and Computation, 243 (??):969–981, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314008789.

Vishal:2013:SFS

[VDOG13] K. Vishal, S. Das, S. H. Ong, and P. Ghosh. On the solutions of fractional Swift Hohenberg equation with dispersion. Applied Mathematics and Computation, 219 (11):5792–5801, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312013082.

Vejchodsky:2013:CEB

[Vej13] Tomás Vejchodský. Complementary error bounds for elliptic systems and applications. Applied Mathematics and Computation, 219(13):7194–7205, March 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007910.

Verma:2010:NAP

[Ver10] Ram U. Verma. New approach to the η-proximal point algorithm and nonlinear variational inclusion problems. Applied Mathematics and Computation, 217 (7):3155–3165, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310009124.

Verrelli:2011:FSE

[Ver11] Cristiano Maria Verrelli. Fourier series expansion for synchronization of permanent magnet electric motors. Applied Mathematics and Computation, 217(9):4502–4515, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310011136.

Verma:2012:RIO

[Ver12] Ram U. Verma. Role of (ρηA)-invexity to ϵ-optimality conditions for multiple objective fractional programming. Applied Mathematics and Computation, 218(17):8444–8452, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312001178.

Verma:2013:WEC

[Ver13] Ram U. Verma. Weak ϵ-efficiency conditions for multiobjective fractional programming. Applied Mathematics and Computation, 219(12):6819–6827, February 15, 2013.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313000568.

Veress:2013:CPI

[VFGP13] Árpád Veress, Attila Felföldi, Tamás Gausz, and László Palkovics. Coupled problem of the inverse design and constraint optimization. Applied Mathematics and Computation, 219(13):7115–7126, March 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311011519.

Vasin:2014:ALR

[VG14a] Vladmir Vasin and Santhosh George. An analysis of Lavrentiev regularization method and Newton type process for nonlinear ill-posed problems. Applied Mathematics and Computation, 230(??):406–413, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313013945.

Vishwakarma:2014:CCR

[VG14b] Gajendra K. Vishwakarma and Raj K. Gangele. A class of chain ratio-type exponential estimators in double sampling using two auxiliary variates. Applied Mathematics and Computation, 227(??):171–175, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313011946.

Verma:2012:SAP

[VGA12] D. K. Verma, Vijay Gupta, and P. N. Agrawal. Some approximation properties of Baskakov–Durrmeyer–Stancu operators. Applied Mathematics and Computation, 218 (11):6549–6556, February 5, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311015050.

Villadelprat:2010:PFC

[Vil10] J. Villadelprat. On the period function in a class of generalized Lotka–Volterra systems. Applied Mathematics and Computation, 216(7):1956–1964, June 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310002675.

Villadelprat:2012:BLC

[Vil12] J. Villadelprat. Bifurcation of local critical periods in the generalized Loud’s system. Applied Mathematics and Computation, 218(12):6803–6813, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311015347.

Virdi:2013:SEI

[Vir13] Jasvinderpal Singh Virdi. Search of exact invariants for P T and non-P T -symmetric complex Hamiltonian systems. Applied Mathematics and Computation, 219(18):9731–9736, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313003202.

Vishwakarma:2014:TWP

[Vis14] Sumit Kumar Vishwakarma. Torsional wave propagation in a self-reinforced medium sandwiched between a rigid layer and a viscoelastic half space under gravity. Applied Mathematics and Computation, 242(??):1–9, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314006663.

Vitovec:2014:COC

[Vít14] Jirí Vítovec. Critical oscillation constant for Euler-type dynamic equations on time scales. Applied Mathematics and Computation, 243(??):838–848, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314009096.

Vasilova:2011:SGA

[VJ11] Maja Vasilova and Miljana Jovanović. Stochastic Gilpin–Ayala competition model with infinite delay. Applied Mathematics and Computation, 217(10):4944–4959, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310011720.

Verdiere:2013:IIP

[VJBDV13] Nathalie Verdière, Ghislaine Joly-Blanchard, and Lilianne Denis-Vidal. Identifiability and identification of a pollution source in a river by using a semi-discretized model. Applied Mathematics and Computation, 221(??):1–9, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313006358.

Verdejo:2014:SIR

[VKV14] Humberto Verdejo, Wolfgang Kliemann, and Luis Vargas. Stability indices for randomly perturbed power systems. Applied Mathematics and Computation, 231(??):386–394, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314000617.

Vlcek:2012:CDA

[VL12] J. Vlcek and L. Luksan. A conjugate directions approach to improve the limited-memory BFGS method. Applied Mathematics and Computation, 219(3):800–809, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312006418.

Vazquez-Leal:2013:ECF

[VLMHK^{+}13] H. Vazquez-Leal, A. Marin-Hernandez, Y. Khan, A. Yildirim, U. Filobello-Nino, R. Castaneda-Sheissa, and V. M. Jimenez-Fernandez. Exploring collision-free path planning by using homotopy continuation methods. Applied Mathematics and Computation, 219(14):7514–7532, March 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313000775.

Vidal:2014:MPB

[VMRA14] J. Vidal, C. Muriel, J. L. Romero, and J. J. Alonso. A Maple procedure based on λsymmetries for second-order ordinary differential equations. Applied Mathematics and Computation, 249(??):147–163, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314013800.

Vampa:2010:HMU

[VMS10] Victoria Vampa, María T. Martin, and Eduardo Serrano. A hybrid method using wavelets for the numerical solution of boundary value problems on the interval. Applied Mathematics and Computation, 217 (7):3355–3367, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031000932X.

Vynnycky:2012:ANS

[VO12] M. Vynnycky and M. A. O’Brien. An asymptotic and numerical study of slow, steady ascent in a Newtonian fluid with temperaturedependent viscosity. Applied Mathematics and Computation, 219(6):3154–3177, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312009484.

Vynnycky:2014:SSA

[VO14] M. Vynnycky and M. A. O’Brien. The slow, steady ascent of a hot solid sphere in a Newtonian fluid with strongly temperaturedependent viscosity. Applied Mathematics and Computation, 231(??):231–253, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313013611.

Vodicka:2013:APD

[Vod13] Roman Vodicka. An asymptotic property of degenerate scales for multiple holes in plane elasticity. Applied Mathematics and Computation, 220(??):166–175, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031300619X.

Vrscay:2014:SCF

[VOL14] E. R. Vrscay, D. Otero, and D. La Torre. A simple class of fractal transforms for hyperspectral images. Applied Mathematics and Computation, 231(??):435–444, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031400037X.

Vajravelu:2012:AMH

[VPS12] K. Vajravelu, K. V. Prasad, and S. R. Santhi. Axisymmetric magneto-hydrodynamic (MHD) flow and heat transfer at a non-isothermal stretching cylinder. Applied Mathematics and Computation, 219 (8):3993–4005, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312010351.

Vetro:2012:NQC

[VR12] Francesca Vetro and Stojan Radenović. Nonlinear ψ-quasi-contractions of Ćirić-type in partial metric spaces. Applied Mathematics and Computation, 219(4):1594–1600, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300312007916.

Vijayakumar:2014:CRC

[VRMT14] V. Vijayakumar, C. Ravichandran, R. Murugesu, and J. J. Trujillo. Controllability results for a class of fractional semilinear integro-differential inclusions via resolvent operators. Applied Mathematics and Computation, 247(??):152–161, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314011904.

Velimirovic:2011:RFA

[VRZ11] Ljubica S. Velimirović, Svetozar R. Rančić, and Milan Lj. Zlatanović. Rigidity and flexibility analysis of a kind of surfaces of revolution and visualization. Applied Mathematics and Computation, 217(9):4612–4619, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310011240.

Vinh:2010:ESE

[VS10] Pham Chi Vinh and Géza Seriani. Explicit secular equations of Stoneley waves in a non-homogeneous orthotropic elastic medium under the influence of gravity. Applied Mathematics and Computation, 215 (10):3515–3525, January 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309009631.

Vrabel:2013:NPE

[VS13] Vladimír Vrábel’ and Marián Slodicka. Nonlinear parabolic equation with a dynamical boundary condition of diffusive type. Applied Mathematics and Computation, 222 (??):372–380, October 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313008072.

Vijayakumar:2014:CCF

[VSM14] V. Vijayakumar, A. Selvakumar, and R. Murugesu. Controllability for a class of fractional neutral integro-differential equations with unbounded delay. Applied Mathematics and Computation, 232(??):303–312, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314000666.

Vajravelu:2013:CIV

[VSS13] K. Vajravelu, S. Sreenadh, and R. Saravana. Combined influence of velocity slip, temperature and concentration jump conditions on MHD peristaltic transport of a Carreau fluid in a non-uniform channel. Applied Mathematics and Computation, 225(??):656–676, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010692.

VanGorder:2010:ASC

[VSV10] Robert A. Van Gorder, Erik Sweet, and K. Vajravelu. Analytical solutions of a coupled nonlinear system arising in a flow between stretching disks. Applied Mathematics and Computation, 216(5):1513–1523, May 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310002419.

Vulanovic:2013:MSD

[VT13] Relja Vulanović and Ljiljana Teofanov. A modification of the Shishkin discretization mesh for one-dimensional reaction-diffusion problems. Applied Mathematics and Computation, 220(??):104–116, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313006024.

Valian:2014:IGH

[VTM14] Ehsan Valian, Saeed Tavakoli, and Shahram Mohanna. An intelligent global harmony search approach to continuous optimization problems. Applied Mathematics and Computation, 232(??):670–684, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314001234.

VanGorder:2010:GCC

[VV10] Robert A. Van Gorder and K. Vajravelu. A general class of coupled nonlinear differential equations arising in self-similar solutions of convective heat transfer problems. Applied Mathematics and Computation, 217(2):460–465, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310006363.

VanGorder:2011:CHT

[VV11a] Robert A. Van Gorder and K. Vajravelu. Convective heat transfer in a conducting fluid over a permeable stretching surface with suction and internal heat generation/absorption. Applied Mathematics and Computation, 217 (12):5810–5821, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310012658.

Vilcu:2011:SGP

[VV11b] Alina Daniela Vîlcu and Gabriel Eduard Vîlcu. On some geometric properties of the generalized CES production functions. Applied Mathematics and Computation, 218 (1):124–129, September 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311007314.

Verdejo:2012:SLS

[VVK12] Humberto Verdejo, Luis Vargas, and Wolfgang Kliemann. Stability of linear stochastic systems via Lyapunov exponents and applications to power systems. Applied Mathematics and Computation, 218(22):11021–11032, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312004626.

Vinogradova:2011:SGM

[VZ11] Polina Vinogradova and Anatoli Zarubin. A study of Galerkin method for the heat convection equations. Applied Mathematics and Computation, 218(2):520–531, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311007764.

Wagener:2010:CPL

[Wag10] F. O. O. Wagener. On conjugate points and the Leitmann equivalent problem approach. Applied Mathematics and Computation, 217(3):1266–1276, October 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310002043.

Wahba:2013:NSG

[Wah13] E. M. Wahba. Non-systematic grid refinement procedures for computational fluid dynamics. Applied Mathematics and Computation, 225(??):829–842, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010795.

Walczak:2010:ATG

[Wal10] Zbigniew Walczak. Approximation theorems for a general class of truncated operators. Applied Mathematics and Computation, 217(5):2142–2148, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007587.

Wang:2010:CSI

[Wan10a] Chuan-Long Wang. Convergences of splitting iterative methods for symmetric indefinite linear systems. Applied Mathematics and Computation, 216(6):1687–1693, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300309010820.

Wang:2010:ICK

[Wan10b] Deng-Shan Wang. Integrability of a coupled KdV system: Painlevé property, Lax pair and Bäcklund transformation. Applied Mathematics and Computation, 216(4):1349–1354, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310001980.

Wang:2010:SNG

[Wan10c] Jing Wang. Some new and general solutions to the compound KdV–Burgers system with nonlinear terms of any order. Applied Mathematics and Computation, 217(4):1652–1657, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300309008182.

Wang:2010:NDC

[Wan10d] Kun Wang. A new defect correction method for the Navier–Stokes equations at high Reynolds numbers. Applied Mathematics and Computation, 216(11):3252–3264, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310004789.

Wang:2010:ASS

[Wan10e] Qi Wang. Approximate solution for system of differential-difference equations by means of the homotopy analysis method. Applied Mathematics and Computation, 217 (8):4122–4128, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310010623.

Wang:2010:ISO

[Wan10f] Yuan-Ming Wang. The iterative solutions of 2n th-order nonlinear multi-point boundary value problems. Applied Mathematics and Computation, 217(5):2251–2259, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310007708.

Wang:2011:PPSb

[Wan11a] Haiyan Wang. Positive periodic solutions of singular systems of first order ordinary differential equations. Applied Mathematics and Computation, 218(5):1605–1610, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031100871X.

Wang:2011:EDA

[Wan11b] Hongchun Wang. An efficient dimensionadaptive uncertainty propagation approach. Applied Mathematics and Computation, 218 (7):3230–3237, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311011039.

Wang:2011:HSP

[Wan11c] Jianfeng Wang. How to solve the polynomial ordinary differential equations. Applied Mathematics and Computation, 218 (6):2421–2438, November 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311010022.

Wang:2011:EPP

[Wan11d] Zhen Wang. Existence of positive periodic solution of n-dimensional Lotka–Volterra system with delays. Applied Mathematics and Computation, 218(5):1934–1940, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311009258.

Wang:2013:ERE

[Wan13a]  Qi Wang.  Extended rational expan-

sion method for differential-difference equation. Applied Mathematics and Computation, 219(17):8965–8972, May 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313003585.

Wang:2013:PSP

[Wan13b] Quan-Fang Wang. Predication of single particle quantum controlling for Yukawa interaction. Applied Mathematics and Computation, 219(15):8078–8094, April 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313001604.

Wang:2013:CCC

[Wan13c] Shuang Wang. Convergence and certain control conditions for hybrid iterative algorithms. Applied Mathematics and Computation, 219(20):10325–10332, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313004025.

Wang:2013:SSN

[Wan13d] Wansheng Wang. Stability of solutions of nonlinear neutral differential equations with piecewise constant delay and their discretizations. Applied Mathematics and Computation, 219(9):4590–4600, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312010867.

Wang:2013:PAP

[Wan13e] Yijie Wang. Periodic and almost periodic solutions of a nonlinear single species discrete model with feedback control. Applied Mathematics and Computation, 219 (10):5480–5486, January 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312011319.

Wang:2014:EES

[Wan14a] Chao Wang. Existence and exponential stability of piecewise mean-square almost periodic solutions for impulsive stochastic Nicholson’s blowflies model on time scales. Applied Mathematics and Computation, 248 (??):101–112, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314012715.

Wang:2014:NSH

[Wan14b] Dongshu Wang. On a non-selective harvesting prey-predator model with Hassell–Varley type functional response. Applied Mathematics and Computation, 246(??):678–695, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314011916.

Wang:2014:MRR

[Wan14c] Hongxing Wang. The minimal rank of A − BX with respect to Hermitian matrix. Applied Mathematics and Computation, 233(??):55–61, May 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314001611.

Wang:2014:SEC

[Wan14d] Hongxing Wang. Some equivalent conditions for block two-by-two matrices to be nonsingular. Applied Mathematics and Computation, 249(??):142–146, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314014131.

Wang:2014:PIA

[Wan14e] Yingguang Wang. A path integration algorithm for stochastic structural dynamic systems. Applied Mathematics and Computation, 228(??):423–431, February 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031301285X.

Wang:2014:SSH

[Wan14f] Zenggui Wang. Symmetries and solutions of hyperbolic mean curvature flow with a constant forcing term. Applied Mathematics and Computation, 235(??):560–566, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313014240.

Wazwaz:2010:CLT

[Waz10a] Abdul-Majid Wazwaz. The combined Laplace transform–Adomian decomposition method for handling nonlinear Volterra integro– differential equations. Applied Mathematics and Computation, 216(4):1304–1309, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310001918.

Wazwaz:2010:FDI

[Waz10b]  Abdul-Majid Wazwaz.  Four (2 + 1)-

dimensional integrable extensions of the Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 215 (10):3631–3644, January 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309009758.

Wazwaz:2010:MKS

[Waz10c] Abdul-Majid Wazwaz. Multiple kink solutions for M-component Burgers equations in (1 + 1)-dimensions and (2 + 1)-dimensions. Applied Mathematics and Computation, 217 (7):3564–3570, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310008842.

Wazwaz:2010:MSSb

[Waz10d] Abdul-Majid Wazwaz. Multiple soliton solutions for the Bogoyavlenskii’s generalized breaking soliton equations and its extension form. Applied Mathematics and Computation, 217(8):4282–4288, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310010088.

Wazwaz:2010:SSI

[Waz10e] Abdul-Majid Wazwaz. N-soliton solutions for the integrable bidirectional sixth-order Sawada–Kotera equation. Applied Mathematics and Computation, 216(8):2317–2320, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310003267.

Wazwaz:2010:NIV

[Waz10f] Abdul-Majid Wazwaz. Non-integrable variants of Boussinesq equation with two solitons. Applied Mathematics and Computation, 217(2):820–825, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310007071.

Wazwaz:2010:SKG

[Waz10g] Abdul-Majid Wazwaz. A study on KdV and Gardner equations with time-dependent coefficients and forcing terms. Applied Mathematics and Computation, 217(5):2277–2281, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S009630031000723X.

Wazwaz:2011:SSS

[Waz11a] Abdul-Majid Wazwaz. N-soliton solutions for shallow water waves equations in (1 + 1) and (2 + 1) dimensions. Applied Mathematics and Computation, 217(21):8840–8845, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311004164.

Wazwaz:2011:RTS

[Waz11b] Abdul-Majid Wazwaz. A reliable treatment of singular Emden–Fowler initial value problems and boundary value problems. Applied Mathematics and Computation, 217 (24):10387–10395, August 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311006540.

Wazwaz:2012:DME

[Waz12a] Abdul-Majid Wazwaz. (2 + 1)-Dimensional mKdV (N) equations by the mKdV recursion operator: Multiple soliton and multiple singular soliton solutions. Applied Mathematics and Computation, 219(5):2535–2544, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031200882X.

Wazwaz:2012:SMS

[Waz12b] Abdul-Majid Wazwaz. Structures of multiple soliton solutions of the generalized, asymmetric and modified Nizhnik–Novikov–Veselov equations. Applied Mathematics and Computation, 218(22):11344–11349, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312005206.

Wazwaz:2013:DBE

[Waz13] Abdul-Majid Wazwaz. (2 + 1)-dimensional Burgers equations BE(m + n + 1): Using the recursion operator. Applied Mathematics and Computation, 219(17):9057–9068, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313003548.

Waz:2014:NIC

[Waz14] Piotr Waz. New indicators of chaos. Applied Mathematics and Computation, 227 (??):449–455, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031301196X.

Wang:2013:ARO

[WaZW13] Bo Wang, Guang an Zou, and Qiang Wang. Application of the restrained optimal perturbation method to study the backward heat conduction problem. Applied Mathematics and Computation, 221(??):703–709, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313007327.

Wang:2011:FVM

[WaZZW11] Bo Wang, Guang an Zou, Peng Zhao, and Qiang Wang. Finite volume method for solving a one-dimensional parabolic inverse problem. Applied Mathematics and Computation, 217(12):5227–5235, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310009793.

Wu:2011:UBD

[WC11] Hongwu Wu and Sui Sun Cheng. Upper bounds for the distances between adjacent zeros of solutions of delay differential equations. Applied Mathematics and Computation, 218(7):3379–3386, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311011222.

Wang:2013:MTR

[WC13] Long Wang and Jianlong Chen. Mixed-type reverse-order laws of (AB)^{(1,3)}, (AB)^{(1,2,3)} and (AB)^{(1,3,4)}. Applied Mathematics and Computation, 222(??):42–52, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313007637.

Wang:2014:GDN

[WC14a] Yi Wang and Jinde Cao. Global dynamics of a network epidemic model for waterborne diseases spread. Applied Mathematics and Computation, 237(??):474–488, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314005190.

Wang:2014:NGS

[WC14b] Yi Wang and Jinde Cao. A note on global stability of the virose equilibrium for network-based computer viruses epidemics. Applied Mathematics and Computation, 244(??):726–740, October 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031401008X.

Wei:2014:HBP

[WC14c] Chunjin Wei and Lansun Chen. Homoclinic bifurcation of prey-predator model with impulsive state feedback control. Applied Mathematics and Computation, 237(??):282–292, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314004950.

Walraevens:2014:ASS

[WCB14] J. Walraevens, D. Claeys, and H. Bruneel. Analysis of steady-state and transient delay in discrete-time single-arrival and batch-arrival systems. Applied Mathematics and Computation, 240(??):62–71, August 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314006110.

Wang:2011:NMS

[WCL11] Yulan Wang, Xuejun Cao, and Xiaona Li. A new method for solving singular fourthorder boundary value problems with mixed boundary conditions. Applied Mathematics and Computation, 217(18):7385–7390, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311001597.

Wang:2012:CMF

[WCL12] Fuzhang Wang, Wen Chen, and Leevan Ling. Combinations of the method of fundamental solutions for general inverse source identification problems. Applied Mathematics and Computation, 219(3):1173–1182, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312007278.

Wang:2014:BTW

[WCL14] Juan Wang, Longwei Chen, and Changfu Liu. Bifurcations and travelling wave solutions of a (2 + 1)-dimensional nonlinear Schrödinger equation. Applied Mathematics and Computation, 249(??):76–80, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314013915.

Wen:2011:AIF

[WCN11] You-Wei Wen, Wai-Ki Ching, and Michael Ng. Approximate inverse-free preconditioners for Toeplitz matrices. Applied Mathematics and Computation, 217(16):6856–6867, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311000464.

Wu:2013:SMT

[WCW^{+}13] Chin-Chia Wu, Shuenn-Ren Cheng, Wen-Hsiang Wu, Yunqiang Yin, and Wen-Hung Wu. The single-machine total tardiness problem with unequal release times and a linear deterioration. Applied Mathematics and Computation, 219(20):10401–10415, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313004189.

Wang:2011:ENS

[WCZ11] Feng Wang, Yujun Cui, and Fang Zhang. Existence of nonnegative solutions for second order m-point boundary value problems at resonance. Applied Mathematics and Computation, 217(10):4849–4855, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630030800221X.

Wang:2013:SAS

[WCZ13] Lasheng Wang, Tao Cheng, and Qimin Zhang. Successive approximation to solutions of stochastic differential equations with jumps in local non-Lipschitz conditions. Applied Mathematics and Computation, 225(??):142–150, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313009983.

Wu:2012:NBD

[WD12] Fengxia Wu and Zhengde Dai. New bright and dark solitons for quintic nonlinear derivative Schrödinger equation. Applied Mathematics and Computation, 218(18):9305–9309, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031200241X.

Wang:2013:DND

[WD13] Yuanyuan Wang and Xiaohua Ding. Dynamics of numerical discretization in a delayed diffusive Nicholson’s blowflies equation. Applied Mathematics and Computation, 222(??):589–603, October 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313008321.

Wang:2014:BTM

[WD14a] Chuanjian Wang and Zhengde Dai. Breathertype multi-solitary waves to the Kadomtsev– Petviashvili equation with positive dispersion. Applied Mathematics and Computation, 235(??):332–337, May 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314003622.

Wang:2014:DSC

[WD14b] Xinwei Wang and Guohui Duan. Discrete singular convolution element method for static, buckling and free vibration analysis of beam structures. Applied Mathematics and Computation, 234(??):36–51, May 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314002185.

Wu:2014:CSE

[WD14c] Kaining Wu and Xiaohua Ding. Convergence and stability of Euler method for impulsive stochastic delay differential equations. Applied Mathematics and Computation, 229(??):151–158, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313013179.

Wallace:2014:MMR

[WDBM14] Dorothy Wallace, Erin R. Dauson, Benjamin A. Bier, and Clyde Martin. A mathematical model of the repopulation of southern high plains playas by Ambystoma tigrinum following Antevs’ Altithermal. Applied Mathematics and Computation, 233(??):187–202, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300313014495.

Wang:2010:ETW

[WDL10] Chuanjian Wang, Zhengde Dai, and Lin Liang. Exact three-wave solution for higher dimensional KdV-type equation. Applied Mathematics and Computation, 216(2):501–505, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310000706.

Wang:2012:MGB

[WDL12a] Xiang Wang, Lin Dai, and Dan Liao. A modified gradient based algorithm for solving Sylvester equations. Applied Mathematics and Computation, 218(9):5620–5628, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311013981.

Wang:2012:SRL

[WDL12b] Xinzeng Wang, Huanhe Dong, and Yuxia Li. Some reductions from a Lax integrable system and their Hamiltonian structures. Applied Mathematics and Computation, 218(20):10032–10039, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312003293.

Wang:2010:BTS

[WDLM10a] Chuanjian Wang, Zhengde Dai, Songqing Lin, and Gui Mu. Breather-type soliton and twosoliton solutions for modified Korteweg–de Vries equation. Applied Mathematics and Computation, 216(1):341–343, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310000214.

Wang:2010:CNE

[WDLM10b] Chuanjian Wang, Zhengde Dai, Songqing Lin, and Gui Mu. Comments on “New exact kink solutions, solitons and periodic form solutions for a combined KdV and Schwarzian KdV equation”. Applied Mathematics and Computation, 216(2):361–363, March 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310001323. See [?].

Wilkie:2010:TSE

[WDS10] K. P. Wilkie, C. S. Drapaca, and S. Sivaloganathan. A theoretical study of the effect of intraventricular pulsations on the pathogenesis of hydrocephalus. Applied Mathematics and Computation, 215(9):3181–3191, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300309008832.

Wilkie:2011:NVF

[WDS11] K. P. Wilkie, C. S. Drapaca, and S. Sivaloganathan. A nonlinear viscoelastic fractional derivative model of infant hydrocephalus. Applied Mathematics and Computation, 217(21):8693–8704, July 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311004838.

Wang:2013:URK

[WDT^{+}13] Yulan Wang, Mingjing Du, Fugui Tan, Ziyang Li, and Tingfang Nie. Using reproducing kernel for solving a class of fractional partial differential equation with non-classical conditions. Applied Mathematics and Computation, 219(11):5918–5925, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312012763.

Wang:2012:EFD

[WDZ12] Xiaofeng Wang, Jovana Dzunić, and Tie Zhang. On an efficient family of derivative free three-point methods for solving nonlinear equations. Applied Mathematics and Computation, 219(4):1749–1760, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312008077.

Wang:2013:HBA

[WDZ13] Xuedi Wang, Lianwang Deng, and Wenli Zhang. Hopf bifurcation analysis and amplitude control of the modified Lorenz system. Applied Mathematics and Computation, 225(??):333–344, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010369.

Wu:2013:DBC

[WE13] Hong-Wu Wu and Lynn Erbe. On the distance between consecutive zeros of solutions of first order delay differential equations. Applied Mathematics and Computation, 219(16):8622–8631, April 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313002026.

Webb:2010:RNB

[Web10] J. R. L. Webb. Remarks on nonlocal boundary value problems at resonance. Applied Mathematics and Computation, 216 (2):497–500, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/

S009630031000069X. See addendum [Web12].

Webb:2012:ARN

[Web12] J. R. L. Webb. Addendum to “Remarks on nonlocal boundary value problems at resonance”. Appl. Math. Comput. 216(2) (2010) 497–500. Applied Mathematics and Computation, 218(10):6176, January 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031101352X. See [Web10].

Wei:2010:BAF

[Wei10a] Hsiu-Chuan Wei. On the bifurcation analysis of a food web of four species. Applied Mathematics and Computation, 215 (9):3280–3292, January 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309009163.

Wei:2010:NES

[Wei10b] Long Wei. New exact solutions to some variable coefficients problems. Applied Mathematics and Computation, 217(4):1632–1638, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300309007905.

Wei:2011:ABR

[Wei11a] Gengping Wei. Asymptotic behavior results for nonlinear neutral delay difference equations. Applied Mathematics and Computation, 217(17):7184–7190, May 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311001615.

Wei:2011:MPS

[Wei11b] Long Wei. Multiple periodic-soliton solutions to Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 218 (2):368–375, September 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007533.

Weichen:2012:HOC

[Wei12] Shi Weichen. Higher order conservation integrals in elasticity and application. Applied Mathematics and Computation, 219 (4):1522–1535, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312007825.

Wei:2014:MNM

[Wei14] Hsiu-Chuan Wei. A modified numerical method for bifurcations of fixed points of ODE systems with periodically pulsed inputs. Applied Mathematics and Computation, 236(??):373–383, June 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314004160.

Wen:2010:CNE

[Wen10a] Xiaoyong Wen. Construction of new exact rational form non-travelling wave solutions to the (2 + 1)-dimensional generalized Broer– Kaup system. Applied Mathematics and Computation, 217(4):1367–1375, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630030900455X.

Wen:2010:EJE

[Wen10b] Xiaoyong Wen. Extended Jacobi elliptic function expansion solutions of variant Boussinesq equations. Applied Mathematics and Computation, 217(6):2808–2820, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310008696.

Wen:2012:ILH

[Wen12] Xiao-Yong Wen. An integrable lattice hierarchy, associated integrable coupling, Darboux transformation and conservation laws. Applied Mathematics and Computation, 218 (9):5796–5805, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311014408.

Wang:2012:GPD

[WES12] Xia Wang, Ahmed Elaiw, and Xinyu Song. Global properties of a delayed HIV infection model with CTL immune response. Applied Mathematics and Computation, 218 (18):9405–9414, May 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312002561.

Wang:2010:PEP

[WF10] Lin-Lin Wang and Yong-Hong Fan. Permanence and existence of periodic solutions for a generalized system with feedback control. Applied Mathematics and Computation, 216(3):902–910, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031000127X.

Wang:2011:MPS

[WF11] Lin-Lin Wang and Yong-Hong Fan. Multiple periodic solutions for a non-autonomous delayed predator-prey model with harvesting terms. Applied Mathematics and Computation, 217(23):9552–9561, August 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311005510.

Wang:2012:SKC

[WF12] Yanxin Wang and Qibin Fan. The second kind Chebyshev wavelet method for solving fractional differential equations. Applied Mathematics and Computation, 218(17):8592–8601, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300312001440.

Wang:2014:NSS

[WF14] Ruihua Wang and Shumin Fei. New stability and stabilization results for discrete-time switched systems. Applied Mathematics and Computation, 238(??):358–369, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314005402.

Wu:2011:ISK

[WFDL11] Ai-Guo Wu, Gang Feng, Guang-Ren Duan, and Wanquan Liu. Iterative solutions to the Kalman–Yakubovich-conjugate matrix equation. Applied Mathematics and Computation, 217(9):4427–4438, January 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031001101X.

Wang:2014:EFP

[WFL14a] Qi Wang, Yayun Fang, and Dicheng Lu. Existence of four periodic solutions for a generalized delayed ratio-dependent predator-prey system. Applied Mathematics and Computation, 247(??):623–630, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314012685.

Wang:2014:ABG

[WFL14b] Shu Wang, Yue-Hong Feng, and Xin Li. The asymptotic behavior of globally smooth solutions of non-isentropic Euler–Maxwell equations for plasmas. Applied Mathematics and Computation, 231(??):299–306, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314000228.

Wang:2010:RSI

[WFW10] Peiguang Wang, Yongyan Fan, and Yonghong Wu. On relative ϕ_{0}-stability of impulsive hybrid systems via perturbing Lyapunov functions. Applied Mathematics and Computation, 216(10):3050–3055, July 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310004431.

Weng:2012:LRA

[WFwC12] Peter Chang-Yi Weng, Hung-Yuan Fan, and Eric King wah Chu. Low-rank approximation to the solution of a nonsymmetric algebraic Riccati equation from transport theory. Applied Mathematics and Computation, 219(2):729–740, October 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312006741.

Weng:2013:ATG

[WFZ13] Zhifeng Weng, Xinlong Feng, and Shuying Zhai. Analysis of two-grid method for semilinear elliptic equations by new mixed finite element scheme. Applied Mathematics and Computation, 219(9):4826–4835, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312011381.

Wang:2012:ELC

[WFZL12] Shiying Wang, Kai Feng, Shurong Zhang, and Jing Li. Embedding long cycles in faulty kary 2-cubes. Applied Mathematics and Computation, 218(9):5409–5413, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311013567.

Wang:2014:ASS

[WFZL14] Wansheng Wang, Qin Fan, Yuan Zhang, and Shoufu Li. Asymptotic stability of solution to nonlinear neutral and Volterra functional differential equations in Banach spaces. Applied Mathematics and Computation, 237(??):217–226, June 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314004822.

Wang:2011:SCU

[WG11] Guodong Wang and Cishui Ge. Semidiscrete central-upwind scheme for conservation laws with a discontinuous flux function in space. Applied Mathematics and Computation, 217(17):7065–7073, May 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310012956.

Wen:2012:SSE

[WG12] Xiao-Yong Wen and Yi-Tian Gao. N-soliton solutions and elastic interaction of the coupled lattice soliton equations for nonlinear waves. Applied Mathematics and Computation, 219(1):99–107, September 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312005401.

Wang:2010:PSN

[WGS10] Jian Wang, Wenjie Gao, and Menglong Su. Periodic solutions of non-Newtonian polytropic filtration equations with nonlinear sources. Applied Mathematics and Computation, 216(7):1996–2009, June 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310002729.

Wen:2011:DTE

[WGW11] Xiao-Yong Wen, Yi-Tian Gao, and Lei Wang. Darboux transformation and explicit solutions for the integrable sixth-order KdV equation for nonlinear waves. Applied Mathematics and Computation, 218(1):55–60, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311007156.

Wang:2010:NCD

[WGZ10] Tingchun Wang, Boling Guo, and Luming Zhang. New conservative difference schemes for a coupled nonlinear Schrödinger system. Applied Mathematics and Computation, 217(4):1604–1619, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309006833.

Wen:2011:MAP

[WH11] Chengming Wen and Ting-Zhu Huang. A modified algorithm for the Perron root of a nonnegative matrix. Applied Mathematics and Computation, 217(9):4453–4458, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310011082.

Wang:2012:CAH

[WH12]  Kun Wang and Yinnian He.  Conver-

gence analysis for a higher order scheme for the time-dependent Navier–Stokes equations. Applied Mathematics and Computation, 218(17):8269–8278, May 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031200094X.

Wang:2013:ALS

[WH13] Yan Wang and Chuanjiang He. An adaptive level set evolution equation for contour extraction. Applied Mathematics and Computation, 219(24):11420–11429, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313005961.

Wang:2014:PSSb

[WH14] Xinhui Wang and Canyun Huang. Permanence of a stage-structured predator-prey system with impulsive stocking prey and harvesting predator. Applied Mathematics and Computation, 235(??):32–42, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314003403.

Wang:2010:SCN

[WHC10] Shuang Wang, Changsong Hu, and Guoqing Chai. Strong convergence of a new composite iterative method for equilibrium problems and fixed point problems. Applied Mathematics and Computation, 215(11):3891–3898, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S009630030901025X.

Wang:2014:MLC

[WHF14] Qinlong Wang, Wentao Huang, and Jingjing Feng. Multiple limit cycles and centers on center manifolds for Lorenz system. Applied Mathematics and Computation, 238(??):281–288, July 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314005670.

Wu:2013:FPF

[WHG13] Lihua Wu, Guoliang He, and Xianguo Geng. The full positive flows of Manakov hierarchy, Hamiltonian structures and conservation laws. Applied Mathematics and Computation, 220(??):20–37, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313005973.

Walker:2012:APA

[WHI12] Sean G. Bernath Walker, Keith William Hipel, and Takehiro Inohara. Attitudes and preferences: Approaches to representing decision maker desires. Applied Mathematics and Computation, 218(12):6637–6647, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311014482.

Wang:2011:LCS

[WHL11] Qinlong Wang, Wentao Huang, and Bai-Lian Li. Limit cycles and singular point quantities for a 3D Lotka–Volterra system. Applied Mathematics and Computation, 217(21):8856–8859, July 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311004814.

Wu:2014:CPL

[WHL14] Jong-Wuu Wu, Ching-Wen Hong, and Wen-Chuan Lee. Computational procedure of lifetime performance index of products for the Burr XII distribution with upper record values. Applied Mathematics and Computation, 227(??):701–716, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313011934.

Witula:2014:CT

[WHS14a] Roman Witula, Edyta Hetmaniok, and Damian Slota. Corrigendum to: “Generalized Gregory’s series” [Applied Mathematics and Computations 237 (2014) 203–216, https://doi.org/10.1016/j.amc.2014.03.101].

Applied Mathematics and Computation, 241(??):13, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314006584. See [WHS14b].

Witula:2014:GGS

[WHS14b] Roman Witula, Edyta Hetmaniok, and Damian Slota. Generalized Gregory’s series. Applied Mathematics and Computation, 237(??):203–216, June 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031400472X. See corrigendum

[WHS14a].

Wang:2010:RDA

[WHW10] Zengyun Wang, Lihong Huang, and Yaonan Wang. Robust decentralized adaptive control for a class of uncertain neural networks with time-varying delays. Applied Mathematics and Computation, 215(12):4154–4163, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300309010650.

Wang:2013:NPI

[WHW13a] Chao Wang, Ting-Zhu Huang, and Chun Wen. A new preconditioner for indefinite and asymmetric matrices. Applied Mathematics and Computation, 219(23):11036–11043, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313002695.

Wang:2013:LRF

[WHW13b] Qinlong Wang, Wentao Huang, and Haitao Wu. Linear recursion formulas of generalized focus quantities and applications. Applied Mathematics and Computation, 219 (10):5233–5240, January 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312011988.

Wee:2014:EMP

[WHWC14] Hui-Ming Wee, Yen-Deng Huang, Wan-Tsu Wang, and Yung-Lung Cheng. An EPQ model with partial backorders considering two backordering costs. Applied Mathematics and Computation, 232(??):898–907, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314001519.

Wu:2014:CSB

[WHZ14] Qiang Wu, Lin Hu, and Zujin Zhang. Convergence and stability of balanced methods for stochastic delay integro-differential equations. Applied Mathematics and Computation, 237(??):446–460, June 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314004895.

Williams:2010:CMS

[Wil10a] M. M. R. Williams. Corrigendum to “A method for solving stochastic eigenvalue problems” [Appl. Math. Comput. 215 (2010) 3906– 3928]. Applied Mathematics and Computation, 217(6):2895, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310008623. See [Wil10b].

Williams:2010:MSS

[Wil10b] M. M. R. Williams. A method for solving stochastic eigenvalue problems. Applied Mathematics and Computation, 215 (11):3906–3928, February 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300309010285. See corrigendum

[Wil10a].

Williams:2011:QDM

[Wil11] Paul Williams. Quadrature discretization method in tethered satellite control. Applied Mathematics and Computation, 217 (21):8223–8235, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311003146.

Wilson:2012:LDM

[Wil12]  Nicholas E. Wilson.  On the Leray-

deconvolution model for the incompressible magnetohydrodynamics equations. Applied Mathematics and Computation, 218 (21):10557–10571, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031200402X.

Williams:2013:MSS

[Wil13] M. M. R. Williams. A method for solving stochastic eigenvalue problems II. Applied Mathematics and Computation, 219 (9):4729–4744, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312011058.

Wang:2011:QNC

[WiN11] Quan-Fang Wang and Shin ichi Nakagiri. Quantum numerical control for free elementary particle. Applied Mathematics and Computation, 217(12):5695–5701, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031001249X.

Wang:2012:SOC

[WiN12] Quan-Fang Wang and Shin ichi Nakagiri. Sensitivity of optimal control for diffusion Hopfield neural network in the presence of perturbation. Applied Mathematics and Computation, 219(8):3793–3808, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312009952.

Ward:2011:IMC

[WJ11] Alan Ward and Josep Jorba. An iterative method for the creation of structured hexahedral meshes over complex orography. Applied Mathematics and Computation, 218 (7):3847–3855, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311012045.

Wang:2012:CCH

[WJ12] Tao Wang and Nuo Jia. Chaos control and hybrid projective synchronization of several new chaotic systems. Applied Mathematics and Computation, 218(13):7231–7240, March 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312000197.

Wang:2014:PNS

[WJ14] Lijuan Wang and Hongling Jiang. Properties and numerical simulations of positive solutions for a variable-territory model. Applied Mathematics and Computation, 236(??):647–662, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300314004421.

Wang:2012:IRC

[WKLW12] J. Wang, Y. Khan, L. X. Lu, and Z. W. Wang. Inner resonance of a coupled hyperbolic tangent nonlinear oscillator arising in a packaging system. Applied Mathematics and Computation, 218(15):7876–7879, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031200118X.

Wang:2010:IGE

[WL10a] Jiayu Wang and Xiaodi Li. Improved global exponential stability for delay difference equations with impulses. Applied Mathematics and Computation, 217(5):1933–1938, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310007332.

Wu:2010:ARK

[WL10b] Boying Wu and Xiuying Li. Application of reproducing kernel method to third order three-point boundary value problems. Applied Mathematics and Computation, 217 (7):3425–3428, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310009562.

Wu:2010:INA

[WL10c] Chunhong Wu and Linzhang Lu. Implicit numerical approximation scheme for the fractional Fokker–Planck equation. Applied Mathematics and Computation, 216(7):1945–1955, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300310002663.

Wu:2010:TWD

[WL10d] Shi-Liang Wu and San-Yang Liu. Traveling waves for delayed non-local diffusion equations with crossing-monostability. Applied Mathematics and Computation, 217 (4):1435–1444, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309005542.

Wang:2011:OCS

[WL11] Peiguang Wang and Ping Li. k th Order convergence for a semilinear elliptic boundary value problem in the divergence form. Applied Mathematics and Computation, 217(21):8547–8551, July 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311004279.

Wang:2012:NSS

[WL12a] Shuli Wang and Jinsheng Liu. Nontrivial solutions of a second order difference systems with multiple resonance. Applied Mathematics and Computation, 218(18):9342–9352, May 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312002494.

Wang:2012:IPS

[WL12b]  Zewen Wang and Jijun Liu.  Identifi-

cation of the pollution source from onedimensional parabolic equation models. Applied Mathematics and Computation, 219 (8):3403–3413, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300308001550.

Wang:2013:ASL

[WL13a] Di Wang and Fucheng Liao. Absolute stability of Lurie direct control systems with time-varying coefficients and multiple nonlinearities. Applied Mathematics and Computation, 219(9):4465–4473, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312010661.

Wang:2013:PCB

[WL13b] Juan Wang and Junping Li. The properties of collision branching processes with immigration and instantaneous resurrection. Applied Mathematics and Computation, 224(??):96–107, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313009065.

Wu:2013:NPB

[WL13c] Shi-Liang Wu and Cui-Xia Li. A note on parameterized block triangular preconditioners for generalized saddle point problems. Applied Mathematics and Computation, 219(14):7907–7916, March 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313000799.

Wu:2013:SBD

[WL13d] Zikai Wu and An Li. Stability and boundedness of differential equations relative to initial time difference by employing vector Lyapunov functions. Applied Mathematics and Computation, 225(??):358–365, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313010060.

Wang:2014:SMB

[WL14a] Jian-Jun Wang and Ya-Jing Liu. Singlemachine bicriterion group scheduling with deteriorating setup times and job processing times. Applied Mathematics and Computation, 242(??):309–314, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314007139.

Wang:2014:EMD

[WL14b] Xinxin Wang and Shengqiang Liu. An epidemic model with different distributed latencies and nonlinear incidence rate. Applied Mathematics and Computation, 241 (??):259–266, August 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314006973.

Wang:2010:NMS

[WLCW10] Y. L. Wang, Z. Y. Li, Y. Cao, and X. H. Wan. A new method for solving a class of mixed boundary value problems with singular coefficient. Applied Mathematics and Computation, 217(6):2768–2772, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310008659.

Wu:2011:IAS

[WLD11] Ai-Guo Wu, Lingling Lv, and Guang-Ren Duan. Iterative algorithms for solving a class of complex conjugate and transpose matrix equations. Applied Mathematics and Computation, 217(21):8343–8353, July 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311003845.

Wang:2010:EUS

[WLH10] Youyu Wang, Guofeng Liu, and Yinping Hu. Existence and uniqueness of solutions for a second order differential equation with integral boundary conditions. Applied Mathematics and Computation, 216(9):2718–2727, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310003838.

Wu:2011:FIA

[WLH11] Ai-Guo Wu, Lingling Lv, and Ming-Zhe Hou. Finite iterative algorithms for a common solution to a group of complex matrix equations. Applied Mathematics and Computation, 218(4):1191–1202, October 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311007776.

Wang:2012:HSD

[WLH12] Chi-Chang Wang, Wu-Jung Liao, and Yin-Sung Hsu. Hybrid spline difference method for the Burgers’ equation. Applied Mathematics and Computation, 219(3):1031–1039, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312007072.

Wu:2010:SSE

[WLHZ10] Hongxia Wu, Xiaojun Liu, Yehui Huang, and Yunbo Zeng. Solving soliton equations with self-consistent sources by method of variation of parameters. Applied Mathematics and Computation, 217(4):1445–1454, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300309005578.

Wan:2011:FEU

[WLJ11] Minghua Wan, Zhihui Lai, and Zhong Jin. Feature extraction using two-dimensional local graph embedding based on maximum margin criterion. Applied Mathematics and Computation, 217(23):9659–9668, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311006205.

Wei:2014:LCB

[WLL14] Lijun Wei, Feng Liang, and Shiping Lu. Limit cycle bifurcations near a generalized homoclinic loop in piecewise smooth systems with a hyperbolic saddle on a switch line. Applied Mathematics and Computation, 243(??):298–310, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314007061.

Wu:2013:CPQ

[WLLD13] Ai-Guo Wu, Wanquan Liu, Caiheng Li, and Guang-Ren Duan. On j-conjugate product of quaternion polynomial matrices. Applied Mathematics and Computation, 219 (24):11223–11232, August 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313005390.

Wang:2014:FTS

[WLLY14] Yan-Wu Wang, Meng Liu, Zhi-Wei Liu, and Jing-Wen Yi. Formation tracking of the second-order multi-agent systems using position-only information via impulsive control with input delays. Applied Mathematics and Computation, 246(??):572–585, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314011680.

Wang:2010:STD

[WLMJ10] Chun-Ni Wang, Shi-Rong Li, Jun Ma, and Wu-Yin Jin. Synchronization transition in degenerate optical parametric oscillators induced by nonlinear coupling. Applied Mathematics and Computation, 216(2):647–654, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310001256.

Wloch:2013:SIG

[Wło13] Andrzej Włoch. Some identities for the generalized Fibonacci numbers and the generalized Lucas numbers. Applied Mathematics and Computation, 219(10):5564–5568, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312011976.

Wang:2014:DIG

[WLP^{+}14] Xiaoxu Wang, Yan Liang, Quan Pan, Chunhui Zhao, and Feng Yang. Design and implementation of Gaussian filter for nonlinear system with randomly delayed measurements and correlated noises. Applied Mathematics and Computation, 232(??):1011–1024, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314000071.

Wei:2010:PSG

[WLQ^{+}10] Fengying Wei, Yangrui Lin, Lulu Que, Yingying Chen, Yunping Wu, and Yuanfu Xue. Periodic solution and global stability for a nonautonomous competitive Lotka–Volterra diffusion system. Applied Mathematics and Computation, 216(10):3097–3104, July 15, 2010.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310004662.

Wang:2014:PSSa

[WLW14] Ying Wang, Lishan Liu, and Yonghong Wu. Positive solutions for singular semipositone boundary value problems on infinite intervals. Applied Mathematics and Computation, 227(??):256–273, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313011697.

Wang:2011:SCM

[WLX11] Zhi-Gang Wang, Zhi-Hong Liu, and Ri-Guang Xiang. Some criteria for meromorphic multivalent starlike functions. Applied Mathematics and Computation, 218(3):1107–1111, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311004474. Special Issue in

Honour of Hari M. Srivastava on his 70th birth anniversary.

Wu:2010:BCO

[WLXL10] Guochang Wu, Dengfeng Li, Huimin Xiao, and Zhanwei Liu. The M-band cardinal orthogonal scaling function. Applied Mathematics and Computation, 215(9):3271–3279, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300309009151.

Wang:2014:NCM

[WLXX14] Shao-Fu Wang, Xiao-Cong Li, Fei Xia, and Zhan-Shan Xie. The novel control method of three dimensional discrete hyperchaotic Hénon map. Applied Mathematics and Computation, 247(??):487–493, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314012296.

Wang:2011:CBN

[WLY11] Qing-Wen Wang, Xin Liu, and Shao-Wen Yu. The common bisymmetric nonnegative definite solutions with extreme ranks and inertias to a pair of matrix equations. Applied Mathematics and Computation, 218 (6):2761–2771, November 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311010447.

Wu:2012:PAA

[WLY^{+}12] Yongqing Wu, Changpin Li, Aili Yang, Lunji Song, and Yujiang Wu. Pinning adaptive antisynchronization between two general complex dynamical networks with non-delayed and delayed coupling. Applied Mathematics and Computation, 218(14):7445–7452, March 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312000252.

Wee:2013:BLV

[WLYC13] H. M. Wee, M. C. Lee, P. C. Yang, and R. L. Chung. Bi-level vendor-buyer strategies for a time-varying product price. Applied Mathematics and Computation, 219 (18):9670–9680, May 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313002567.

Wang:2011:SMC

[WLZ11] Fei Wang, Wei Li, and HongQing Zhang. A systematic method to construct polynomials possessing linear superposition principle and dispersion relation. Applied Mathematics and Computation, 218(7):3925–3933, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311012124.

Wang:2012:TRC

[WLZ12a] Jizhong Wang, Jin Li, and Yueting Zhou. The trapezoidal rule for computing supersingular integral on interval. Applied Mathematics and Computation, 219(4):1616–1624, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312007953.

Wu:2012:ESD

[WLZ12b] Shu-Lin Wu, Ke-Lin Li, and Jin-Shan Zhang. Exponential stability of discrete-time neural networks with delay and impulses. Applied Mathematics and Computation, 218 (12):6972–6986, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311015694.

Wazwaz:2010:VET

[WM10] Abdul-Majid Wazwaz and Mona S. Mehanna. A variety of exact travelling wave solutions for the (2+1)-dimensional Boiti–Leon–Pempinelli equation. Applied Mathematics and Computation, 217(4):1484–1490, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309005888.

Wang:2013:SPA

[WM13a] Minghui Wang and Wenhao Ma. A structurepreserving algorithm for the quaternion Cholesky decomposition. Applied Mathematics and Computation, 223(??):354–361, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313008941.

Weckner:2013:VMM

[WM13b] Olaf Weckner and Nik Abdullah Nik Mohamed. Viscoelastic material models in peridynamics. Applied Mathematics and Computation, 219(11):6039–6043, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031201257X.

Wang:2012:CAI

[WMH12] Jin Wang, Chairat Modnak, and Gene Hou. Convergence analysis of an iterative algorithm for a class of constrained dynamic problems. Applied Mathematics and Computation, 219(3):1200–1221, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312007308.

Wang:2011:EFI

[WMJW11] Chun-Ni Wang, Jun Ma, Wu-Yin Jin, and Ying Wu. Electric Field-induced dynamical evolution of spiral wave in the regular networks of Hodgkin–Huxley neurons. Applied Mathematics and Computation, 218 (8):4467–4474, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311012574.

Wang:2014:HWM

[WMM14] Lifeng Wang, Yunpeng Ma, and Zhijun Meng. Haar wavelet method for solving fractional partial differential equations numerically. Applied Mathematics and Computation, 227(??):66–76, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313011648.

Mao:2011:FMD

[wMwWsT11] Bing wei Mao, Fu wei Wang, and Nai shuo Tian. Fluid model driven by an M/G/1 queue with multiple exponential vacations. Applied Mathematics and Computation, 218 (8):4041–4048, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031101112X.

Withers:2011:PSS

[WN11] Christopher S. Withers and Saralees Nadarajah. Power series solutions to Volterra integral equations. Applied Mathematics and Computation, 218(5):2353–2363, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311009933.

Withers:2012:CFE

[WN12a] Christopher S. Withers and Saralees Nadarajah. Cornish–Fisher expansions about the F - distribution. Applied Mathematics and Computation, 218(15):7947–7957, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312001348.

Withers:2012:EBF

[WN12b] Christopher S. Withers and Saralees Nadarajah. Expansions for the beta function and its reciprocal when one parameter is large. Applied Mathematics and Computation, 219 (4):2345–2349, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312008387.

Withers:2014:DQR

[WN14a] Christopher S. Withers and Saralees Nadarajah. The distribution and quantiles of the range of a Wiener process. Applied Mathematics and Computation, 232(??):766–770, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314001933.

Withers:2014:SSL

[WN14b] Christopher S. Withers and Saralees Nadarajah. Series solutions to linear integral equations. Applied Mathematics and Computation, 230(??):218–230, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313014501.

Woods:2020:CRS

[Woo20]  R. C. Woods.  Comments on “Rela-

tivistic symmetries in Yukawa-type interactions with Coulomb-like tensor”. Applied Mathematics and Computation, 365(??): Article 124675, January 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300319306678. See [IHF13].

Wozny:2012:SAC

[Woź12] Paweł Woźny. Simple algorithms for computing the Bézier coefficients of the constrained dual Bernstein polynomials. Applied Mathematics and Computation, 219 (5):2521–2525, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312008806.

Wozny:2013:SNJ

[Woź13] Paweł Woźny. A short note on Jacobi– Bernstein connection coefficients. Applied Mathematics and Computation, 222 (??):53–57, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313007790.

Wongsaijai:2014:TLA

[WP14] B. Wongsaijai and K. Poochinapan. A threelevel average implicit finite difference scheme to solve equation obtained by coupling the Rosenau–KdV equation and the Rosenau– RLW equation. Applied Mathematics and Computation, 245(??):289–304, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314010327.

Wang:2014:GTD

[WPKE14] Jinliang Wang, Jingmei Pang, Toshikazu Kuniya, and Yoichi Enatsu. Global threshold dynamics in a five-dimensional virus model with cell-mediated, humoral immune responses and distributed delays. Applied Mathematics and Computation, 241(??):298–316, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314006717.

Wu:2011:DAS

[WPSC11] Zheng-Guang Wu, Ju H. Park, Hongye Su, and Jian Chu. Dissipativity analysis for singular systems with time-varying delays. Applied Mathematics and Computation, 218 (8):4605–4613, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311012896.

Wu:2012:ADA

[WPSC12] Zheng-Guang Wu, Ju H. Park, Hongye Su, and Jian Chu. Admissibility and dissipativity analysis for discrete-time singular systems with mixed time-varying delays. Applied Mathematics and Computation, 218(13):7128–7138, March 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311015645.

Wang:2013:DDR

[WPSW13] Jing Wang, Ju H. Park, Hao Shen, and Jian Wang. Delay-dependent robust dissipativity conditions for delayed neural networks with random uncertainties. Applied Mathematics and Computation, 221(??):710–719, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313007674.

Wang:2011:QAR

[WQ11a] Jing Wang and Xin Qu. Qualitative analysis for a ratio-dependent predator-prey model with disease and diffusion. Applied Mathematics and Computation, 217(24):9933–9947, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S009630031100587X.

Wang:2011:BPC

[WQ11b] Ren-Hong Wang and Jiang Qian. Bivariate polynomial and continued fraction interpolation over ortho-triples. Applied Mathematics and Computation, 217(19):7620–7635, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311002487.

Wei:2013:APS

[WQ13] Xuerui Wei and Zhipeng Qiu. Anti-periodic solutions for BAM neural networks with time delays. Applied Mathematics and Computation, 221(??):221–229, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313007029.

Wang:2010:MMR

[WQW^{+}10] Zhigang Wang, Xinqiang Qin, Guo Wei, Lijun Su, Lihua Wang, and Baoyan Fang. Meshless method with ridge basis functions. Applied Mathematics and Computation, 217 (5):1870–1886, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007277.

Wang:2010:TDI

[WQX^{+}10] Xinhua Wang, Shaobo Qu, Zhuo Xu, Hua Ma, Jiafu Wang, Chao Gu, and Xiang Wu. Three-dimensional invisible cloaks with arbitrary shapes based on partial differential equation. Applied Mathematics and Computation, 216(2):426–430, March 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310000469.

Wang:2011:HBW

[WqXlC11] Hong Wang, Da quan Xian, and Han lin Chen. Homoclinic breather-wave solutions and doubly periodic wave solutions for coupled KdV equations. Applied Mathematics and Computation, 218(2):610–615, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311007958.

Woon:2010:CRD

[WR10] Siew Fang Woon and Volker Rehbock. A critical review of discrete filled function methods in solving nonlinear discrete optimization problems. Applied Mathematics and Computation, 217(1):25–41, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031000528X.

Wazwaz:2013:ADM

[WRD13] Abdul-Majid Wazwaz, Randolph Rach, and Jun-Sheng Duan. Adomian decomposition method for solving the Volterra integral form of the Lane–Emden equations with initial values and boundary conditions. Applied Mathematics and Computation, 219 (10):5004–5019, January 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312011654.

Wright:2011:NER

[Wri11] S. E. Wright. A note on the equality of rank and trace for an idempotent matrix. Applied Mathematics and Computation, 217(16):7048–7049, April 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311001585.

Wang:2012:OOE

[WRZ12] Hongmei Wang, Hongxing Rui, and Shuai Zhang. An optimal-order error estimate for the mass-conservative characteristic finite element scheme. Applied Mathematics and Computation, 218(20):10271–10278, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312003876.

Wauer:2010:DRT

[WS10] J. Wauer and B. Schweizer. Dynamics of rotating thermoelastic disks with stationary heat source. Applied Mathematics and Computation, 215(12):4272–4279, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309011102.

Wang:2011:IDD

[WS11] Cheng Wang and Yi Shen. Improved delaydependent robust stability criteria for uncertain time delay systems. Applied Mathematics and Computation, 218(6):2880–2888, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300311010575.

Wang:2012:LST

[WS12] Zuolei Wang and Xuerong Shi. Lag synchronization of two identical Hindmarsh–Rose neuron systems with mismatched parameters and external disturbance via a single sliding mode controller. Applied Mathematics and Computation, 218(22):10914–10921, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312004535.

Wu:2013:MSP

[WS13] Xuedong Wu and Zhihuan Song. Multi-step prediction of chaotic time-series with intermittent failures based on the generalized nonlinear filtering methods. Applied Mathematics and Computation, 219(16):8584–8594, April 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313002099.

Wang:2014:SMM

[WS14] Liqun Wang and Liwei Shi. A simple method for matrix-valued coefficient elliptic equations with sharp-edged interfaces. Applied Mathematics and Computation, 242 (??):917–930, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031400928X.

Wang:2014:EEI

[WSCH14] J. F. Wang, F. X. Sun, Y. M. Cheng, and A. X. Huang. Error estimates for the interpolating moving least-squares method. Applied Mathematics and Computation, 245 (??):321–342, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314010297.

Wei:2012:ASL

[WSD12] Gao-Feng Wei, Guo-Hua Sun, and Shi-Hai Dong. Application of the Sturm–Liouville theorem and shape invariance formalism to the Dirac equation with hyperbolic like potential. Applied Mathematics and Computation, 218(22):11171–11176, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312004912.

Wang:2013:ELN

[WSD13] Jing Wang, Pengjian Shang, and Keqiang Dong. Effect of linear and nonlinear filters on multifractal analysis. Applied Mathematics and Computation, 224(??):337–345, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313009053.

Witula:2012:SPF

[WSK12] Roman Witula, Damian Slota, and Marian Kampik. Some properties of the full matrices. Applied Mathematics and Computation, 219(3):1222–1231, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031200731X.

Wang:2013:SCG

[WSL^{+}13] Ming Wang, Wen-Rui Shan, Xing Lü, Yu-Shan Xue, Zhi-Qiang Lin, and Bo Tian. Soliton collision in a general coupled nonlinear Schrödinger system via symbolic computation. Applied Mathematics and Computation, 219(24):11258–11264, August 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031300413X.

Wang:2014:AOM

[WSY14] Kuo-Hsiung Wang, Je-Hung Su, and Dong-Yuh Yang. Analysis and optimization of an M/G/1 machine repair problem with multiple imperfect coverage. Applied Mathematics and Computation, 242(??):590–600, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314008327.

Wazwaz:2010:MSSa

[WT10] Abdul-Majid Wazwaz and Houria Triki. Multiple soliton solutions for the sixth-order Ramani equation and a coupled Ramani equation. Applied Mathematics and Computation, 216(1):332–336, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310000809.

Wang:2011:CMT

[WT11a] Kuo-Hsien Wang and Che-Tsung Tung. Construction of a model towards EOQ and pricing strategy for gradually obsolescent products. Applied Mathematics and Computation, 217(16):6926–6933, April 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311001391.

Wang:2011:ART

[WT11b] Yue-Peng Wang and Su-Lin Tao. Application of regularization technique to variational adjoint method: a case for nonlinear convectiondiffusion problem. Applied Mathematics and Computation, 218(8):4475–4482, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311012586.

Wu:2011:DME

[WT11c] Hsin-Hung Wu and Ya-Ning Tsai. A DE-MATEL method to evaluate the causal relations among the criteria in auto spare parts industry. Applied Mathematics and Computation, 218(5):2334–2342, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311010034.

Wu:2011:ACL

[WT11d] Yan Wu and Jonathan S. Tanner. Adaptive control of linear time invariant systems via a wavelet network and applications to control Lorenz chaos. Applied Mathematics and Computation, 218(1):22–31, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007065.

Wenpeng:2012:ISR

[WT12] Zhang Wenpeng and Wang Tingting. The infinite sum of reciprocal Pell numbers. Applied Mathematics and Computation, 218 (10):6164–6167, January 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311014366.

Wei:2013:SPS

[WTFC13] Minzhi Wei, Shengqiang Tang, Hualiang Fu, and Guangxi Chen. Single peak solitary wave solutions for the generalized KP-MEW (2, 2) equation under boundary condition. Applied Mathematics and Computation, 219(17):8979–8990, May 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313002609.

Wang:2014:BAS

[WTH^{+}14] Zhaocai Wang, Jian Tan, Dongmei Huang, Yingchao Ren, and Zuwen Ji. A biological algorithm to solve the assignment problem based on DNA molecules computation. Applied Mathematics and Computation, 244(??):183–190, October 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314009424.

Wang:2011:LPB

[WTL^{+}11] Pan Wang, Bo Tian, Wen-Jun Liu, Xing Lü, and Yan Jiang. Lax pair, Bäcklund transformation and multi-soliton solutions for the Boussinesq–Burgers equations from shallow water waves. Applied Mathematics and Computation, 218(5):1726–1734, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311008861.

Wu:2014:OCO

[WTQU14] Zhiyou Wu, Jing Tian, Jing Quan, and Julien Ugon. Optimality conditions and optimization methods for quartic polynomial optimization. Applied Mathematics and Computation, 232(??):968–982, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314001118.

Wei:2011:NMO

[WTZ11] W. Wei, K. L. Teo, and Z. Zhan. A numerical method for an optimal control problem with minimum sensitivity on coefficient variation. Applied Mathematics and Computation, 218(4):1180–1190, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311007740.

Wu:2010:CAP

[Wu10a] Honglin Wu. On computing of arbitrary positive powers for one type of anti-tridiagonal matrices of even order. Applied Mathematics and Computation, 217(6):2750–2756, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310008647.

Wu:2010:NET

[Wu10b] Jianping Wu. New explicit traveling wave solutions for three nonlinear evolution equations. Applied Mathematics and Computation, 217(4):1764–1770, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309006572.

Wu:2010:ABP

[Wu10c]  Yong-Qi Wu.  Asymptotic behavior of

the periodic wave solution for the (3+1)- dimensional Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 216(11):3154–3161, August 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310004509.

Wu:2012:LSM

[Wu12a] Jinming Wu. Least squares methods for solving partial differential equations by using Bézier control points. Applied Mathematics and Computation, 219(8):3655–3663, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300312009629.

Wu:2012:RAC

[Wu12b] Qingtai Wu. Reliability analysis of a cold standby system attacked by shocks. Applied Mathematics and Computation, 218 (23):11654–11673, August 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312005668.

Wu:2012:DCA

[Wu12c] Xiaoqian Wu. A divide and conquer algorithm on the double dimensional inverse eigenvalue problem for Jacobi matrices. Applied Mathematics and Computation, 219 (8):3840–3846, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312010004.

Wu:2012:QPW

[Wu12d] Yong-Qi Wu. Quasi-periodic wave solution and asymptotic behavior for the (2 + 1)- dimensional Toda lattice equation. Applied Mathematics and Computation, 219 (6):3065–3072, November 25, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312009320.

Wu:2013:ABC

[Wu13] Zhonglin Wu. Asymptotic behavior for a coupled Petrovsky and wave system with localized damping. Applied Mathematics and Computation, 224(??):442–449, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313009338.

Wu:2014:ERI

[Wu14a] Jibo Wu. Efficiency of the restricted r−d class estimator in linear regression. Applied Mathematics and Computation, 236(??):572–579, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300314003725.

Wu:2014:REL

[Wu14b] Jibo Wu. The relative efficiency of Liutype estimator in a partially linear model. Applied Mathematics and Computation, 243 (??):349–357, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314008030.

Wu:2014:DFO

[Wu14c] Yan Wu. Discretization of fractional order differentiator over Paley–Wiener space. Applied Mathematics and Computation, 247 (??):162–168, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314011990.

Wang:2014:SGO

[WV14] Zhibo Wang and Seakweng Vong. On some generalizations of an Ostrowski–Grüss type integral inequality. Applied Mathematics and Computation, 229(??):239–244, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313013349.

Wang:2010:GEC

[WW10] Yu-Zhu Wang and Yin-Xia Wang. Global existence of classical solutions to the minimal surface equation with slow decay initial value. Applied Mathematics and Computation, 216(2):576–583, March 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310001037.

Wang:2011:UPT

[WW11a] Dongshu Wang and Quanyi Wang. Uniform persistence for a two-species ratio-dependent predator–prey system with diffusion and variable time delays. Applied Mathematics and Computation, 217(10):4900–4910, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310011434.

Wang:2011:NAM

[WW11b] Fusheng Wang and Yanping Wang. Nonmonotone algorithm for minimax optimization problems. Applied Mathematics and Computation, 217(13):6296–6308, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311000063.

Wang:2011:DSS

[WW11c] Hongwei Wang and Shubin Wang. Decay and scattering of small solutions for Rosenau equations. Applied Mathematics and Computation, 218(1):115–123, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311007302.

Wang:2011:PMS

[WW11d] Ji-Bo Wang and Cai-Min Wei. Parallel machine scheduling with a deteriorating maintenance activity and total absolute differences penalties. Applied Mathematics and Computation, 217(20):8093–8099, June 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311003493.

Wang:2011:EQS

[WW11e] Zhiguo Wang and Yiqian Wang. Existence of quasiperiodic solutions and Littlewood’s boundedness problem of super-linear impact oscillators. Applied Mathematics and Computation, 217(13):6417–6425, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311000531.

Wu:2011:MRP

[WW11f] Ke Wu and Xian Wu. Multiplicity results of periodic solutions for systems of first order delay differential equation. Applied Mathematics and Computation, 218(5):1765–1773, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300311009052.

Wu:2011:RAT

[WW11g] Qingtai Wu and Shaomin Wu. Reliability analysis of two-unit cold standby repairable systems under Poisson shocks. Applied Mathematics and Computation, 218 (1):171–182, September 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007703.

Wu:2011:OIP

[WW11h] Shaomin Wu and Wenbin Wang. Optimal inspection policy for three-state systems monitored by control charts. Applied Mathematics and Computation, 217(23):9810–9819, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031100645X.

Wang:2012:ARE

[WW12] MeiJiao Wang and Qi Wang. Application of rational expansion method for stochastic differential equations. Applied Mathematics and Computation, 218(9):5259–5264, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311013403.

Wang:2013:ANT

[WW13a] Fu-Sheng Wang and Chuan-Long Wang. An adaptive nonmonotone trust-region method with curvilinear search for minimax problem. Applied Mathematics and Computation, 219(15):8033–8041, April 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031300146X.

Wang:2013:PDR

[WW13b] Jin-Liang Wang and Huai-Ning Wu. Passivity of delayed reaction-diffusion networks with application to a food web model. Applied Mathematics and Computation, 219 (24):11311–11326, August 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313005602.

Wang:2013:LCM

[WW13c] Keyan Wang and Qisheng Wang. Lagrange collocation method for solving Volterra–Fredholm integral equations. Applied Mathematics and Computation, 219 (21):10434–10440, July 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313004177.

Wang:2013:FPC

[WW13d] Mingbin Wang and Gang Wang. Four properties for complex potentials in power series form. Applied Mathematics and Computation, 219(9):4185–4193, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312011629.

Wang:2013:FVA

[WW13e] Xinwei Wang and Yongliang Wang. Free vibration analysis of multiple-stepped beams by the differential quadrature element method. Applied Mathematics and Computation, 219 (11):5802–5810, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312013136.

Wang:2013:DQA

[WW13f]  Xinwei Wang and Zhe Wu.  Differen-

tial quadrature analysis of free vibration of rhombic plates with free edges. Applied Mathematics and Computation, 225 (??):171–183, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313009909.

Wang:2013:EST

[WW13g] Yuxiao Wang and Zhi-Cheng Wang. Entire solutions in a time-delayed and diffusive epidemic model. Applied Mathematics and Computation, 219(10):5033–5041, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312011964.

Wong:2013:VSL

[WW13h] Tat Wing Wong and Hoi Ying Wong. Valuation of stock loans using exponential phasetype Lévy models. Applied Mathematics and Computation, 222(??):275–289, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313007996.

Wu:2013:CBC

[WW13i] Xiaoqin P. Wu and Liancheng Wang. Codimension-2 bifurcations of coupled BVP oscillators with hard characteristics. Applied Mathematics and Computation, 219 (10):5303–5320, January 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031201096X.

Wu:2013:OCO

[WW13j] Ying Wu and Yapeng Wang. On the optimal constants of Ostrowski-like inequalities involving n knots. Applied Mathematics and Computation, 219(14):7789–7794, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313001203.

Wang:2014:TPM

[WW14a] Keyan Wang and Qisheng Wang. Taylor polynomial method and error estimation for a kind of mixed Volterra–Fredholm integral equations. Applied Mathematics and Computation, 229(??):53–59, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313012903.

Wang:2014:HOC

[WW14b] Tao Wang and Yuan-Ming Wang. A higherorder compact LOD method and its extrapolations for nonhomogeneous parabolic differential equations. Applied Mathematics and Computation, 237(??):512–530, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314005104.

Wang:2014:DES

[WW14c] Yu-Zhu Wang and Keyan Wang. Decay estimate of solutions to the sixth order damped Boussinesq equation. Applied Mathematics and Computation, 239(??):171–179, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314005852.

Wang:2014:DMM

[WW14d] Yuanshi Wang and Hong Wu. Dynamics of a mutualism model with saturated response. Applied Mathematics and Computation, 240(??):16–29, August 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314005943.

Wen:2014:OSS

[WW14e] Xiao-Yong Wen and Deng-Shan Wang. Odd-soliton solutions and inelastic interaction for the differential-difference nonlinear Schrödinger equation in nonlinear optics. Applied Mathematics and Computation, 244(??):598–605, October 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314009849.

Wei:2012:NSO

[WWC12] Xin Wei, Yi-Zhong Wu, and Li-Ping Chen. A new sequential optimal sampling method for radial basis functions. Applied Mathematics and Computation, 218(19):9635–9646, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031200207X.

Wang:2013:IMC

[WWG13] Keyan Wang, Qisheng Wang, and Kaizhong Guan. Iterative method and convergence analysis for a kind of mixed nonlinear Volterra–Fredholm integral equation. Applied Mathematics and Computation, 225 (??):631–637, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313010485.

Wang:2013:ESM

[WWH13] Minghui Wang, Musheng Wei, and Shanrui Hu. The extremal solution of the matrix equation X^{s} + A^{∗}X^{−q}A = I. Applied Mathematics and Computation, 220(??):193–199, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313006851.

Wee:2014:EPQ

[WWK^{+}14] Hui-Ming Wee, Wan-Tsu Wang, Tsai-Chi Kuo, Yung-Lung Cheng, and Yen-Deng Huang. An economic production quantity model with non-synchronized screening and rework. Applied Mathematics and Computation, 233(??):127–138, May 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314001969.

Wang:2011:CGM

[WWLL11] Guangbin Wang, Hao Wen, Liangliang Li, and Xue Li. Convergence of GAOR method for doubly diagonally dominant matrices. Applied Mathematics and Computation, 217(18):7509–7514, May 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311002499.

Wang:2014:USA

[WWQ14] Lingzhi Wang, Yonghai Wang, and Yuming Qin. Upper semicontinuity of attractors for nonclassical diffusion equations in H^{1}(R^{3}). Applied Mathematics and Computation, 240(??):51–61, August 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314006328.

Wang:2010:EMS

[WWS10] Yanan Wang, Yuwen Wang, and Junping Shi. Exact multiplicity of solutions to a diffusive logistic equation with harvesting. Applied Mathematics and Computation, 216(5):1531–1537, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300310002444.

Wang:2014:RHS

[WWS^{+}14] Zuolei Wang, Chao Wang, Xuerong Shi, Jun Ma, Keming Tang, and Hongsheng Cheng. Realizing hybrid synchronization of time-delay hyperchaotic 4D systems via partial variables. Applied Mathematics and Computation, 245(??):427–437, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314010686.

Wang:2013:SRP

[WWT13] Guangbin Wang, Ting Wang, and Fuping Tan. Some results on preconditioned GAOR methods. Applied Mathematics and Computation, 219(11):5811–5816, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312012970.

Wei:2012:LIR

[WWW12] Zhiqiang Wei, Yu-Zhu Wang, and Yin-Xia Wang. Logarithmically improved regularity criteria for the Navier–Stokes equations in Lorentz spaces. Applied Mathematics and Computation, 218(19):9848–9852, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312003001.

Wang:2013:OCG

[WWW13a] Chuan-Long Wang, Fu-Sheng Wang, and Yan-Ping Wang. The optimality conditions for generalized minimax programming. Applied Mathematics and Computation, 219(17):8881–8892, May 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313002944.

Wang:2013:PSO

[WWW13b] Hui Wang, Wenjun Wang, and Zhijian Wu. Particle swarm optimization with adaptive mutation for multimodal optimization. Applied Mathematics and Computation, 221 (??):296–305, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313007236.

Wang:2013:NRC

[WWW13c] Xiaoyun Wang, Li Wang, and Yunjin Wu. Novel results for a class of singular perturbed slow-fast system. Applied Mathematics and Computation, 225(??):795–806, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313010503.

Wang:2013:PPM

[WWW13d] Yuanshi Wang, Hong Wu, and Shikun Wang. A predator-prey model characterizing negative effect of prey on its predator. Applied Mathematics and Computation, 219(19):9992–9999, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300313004116.

Wang:2014:OSS

[WWWJ14] Zhiguo Wang, Luping Wang, Deng-Shan Wang, and Yan Jin. Optimal system, symmetry reductions and new closed form solutions for the geometric average Asian options. Applied Mathematics and Computation, 226(??):598–605, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031301076X.

Wen:2010:GNM

[WWY10] Rui-Ping Wen, Chuan-Long Wang, and Xi-Hong Yan. Generalizations of the nonstationary multisplitting iterative method for symmetric positive definite linear systems. Applied Mathematics and Computation, 216(6):1707–1714, May 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309010807.

Wang:2010:SMS

[WWZ10a] Ji-Bo Wang, Dan Wang, and Guo-Dong Zhang. Single-machine scheduling with learning functions. Applied Mathematics and Computation, 216(4):1280–1286, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310001888.

Wang:2010:CDI

[WWZ10b] Yi Wang, Jianzhong Wang, and Li Zhang. Cross diffusion-induced pattern in an SI model. Applied Mathematics and Computation, 217(5):1965–1970, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031000737X.

Wang:2013:NCC

[WWZ13] Shanshan Wang, Tingchun Wang, and Luming Zhang. Numerical computations for N-coupled nonlinear Schrödinger equations by split step spectral methods. Applied Mathematics and Computation, 222 (??):438–452, October 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313008102.

Wang:2014:DAP

[WWZ14a] Xiaoyun Wang, Lijuan Wei, and Juan Zhang. Dynamical analysis and perturbation solution of an SEIR epidemic model. Applied Mathematics and Computation, 232(??):479–486, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300314001271.

Wu:2014:NIB

[WWZ14b] Jinming Wu, Renhong Wang, and Xiaolei Zhang. Numerical integration based on a multilevel quartic quasi-interpolation operator. Applied Mathematics and Computation, 227(??):132–138, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313011612.

Wang:2012:CST

[WX12] Hong Xia Wang and Gen Qi Xu. A cold system with two different components and a single vacation of the repairman. Applied Mathematics and Computation, 219(5):2634–2657, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300312008909.

Wang:2013:CLS

[WX13a] Hui Wang and Tiecheng Xia. Conservation laws and self-consistent sources for a super KN hierarchy. Applied Mathematics and Computation, 219(10):5458–5464, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031201209X.

Xu:2013:GPM

[wX13b] Wei wei Xu. A generalization of preconditioned MHSS iteration method for complex symmetric indefinite linear systems. Applied Mathematics and Computation, 219 (21):10510–10517, July 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313003949.

Xu:2013:EBT

[wX13c] Wei wei Xu. On eigenvalue bounds of two classes of two-by-two block indefinite matrices. Applied Mathematics and Computation, 219(12):6669–6679, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312013240.

Wen:2013:MSS

[WX13d] Xiao-Yong Wen and Xiao-Ge Xu. Multiple soliton solutions and fusion interaction phenomena for the (2 + 1)-dimensional modified dispersive water-wave system. Applied Mathematics and Computation, 219 (14):7730–7740, March 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313001112.

Wang:2010:CSM

[WXA10] Zhi-Gang Wang, Neng Xu, and Mugur Acu. Certain subclasses of multivalent analytic functions defined by multiplier transforms. Applied Mathematics and Computation, 216(1):192–204, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310000445.

Wang:2013:ISL

[WXLW13] Qing Wang, Pengcheng Xu, Edmund Y. Lam, and Ngai Wong. Iterative solution to linear matrix inequality arising from time delay descriptor systems. Applied Mathematics and Computation, 219(9):4176–4184, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312011708.

Wang:2014:APC

[WXWY14] Guan-Sheng Wang, Jiang-Wen Xiao, Yan-Wu Wang, and Jing-Wen Yi. Adaptive pinning cluster synchronization of fractionalorder complex dynamical networks. Applied Mathematics and Computation, 231 (??):347–356, March 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314000605.

Wang:2010:TAH

[WY10a] Cheng-Chi Wang and Her-Terng Yau. Theoretical analysis of high speed spindle air bearings by a hybrid numerical method. Applied Mathematics and Computation, 217 (5):2084–2096, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007538.

Wei:2010:ACH

[WY10b] Zhouchao Wei and Qigui Yang. Anticontrol of Hopf bifurcation in the new chaotic system with two stable node-foci. Applied Mathematics and Computation, 217 (1):422–429, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310005795.

Wu:2010:ETR

[WY10c] Jie Wu and Hong Yan. An effective transformation in ranking using l_{1}-norm in data envelopment analysis. Applied Mathematics and Computation, 217(8):4061–4064, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310010465.

Wang:2011:HPP

[WY11a] I-Lin Wang and Hui-E Yang. Haplotyping populations by pure parsimony based on compatible genotypes and greedy heuristics. Applied Mathematics and Computation, 217(23):9798–9809, August 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311006436.

Wang:2011:CCN

[WY11b] Jufeng Wang and Huizhong Yang. H_{∞} control of a class of networked control systems with time delay and packet dropout. Applied Mathematics and Computation, 217 (18):7469–7477, May 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311002384.

Wang:2011:GDT

[WY11c] Lei Wang and Wei Yang. Global dynamics of a two-patch SIS model with infection during transport. Applied Mathematics and Computation, 217(21):8458–8467, July 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311004139.

Wang:2012:ESC

[WY12a] Jufeng Wang and Huizhong Yang. Exponential stability of a class of networked control systems with time delays and packet dropouts. Applied Mathematics and Computation, 218(17):8887–8894, May 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312001749.

Wang:2012:SOP

[WY12b] Kai-Yong Wang and Guo-Fu Yu. On a set of orthogonal polynomials associated with a quantized physical model. Applied Mathematics and Computation, 218 (11):6275–6279, February 5, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311014494.

Wangkeeree:2012:EAB

[WY12c] Rabian Wangkeeree and Panu Yimmuang. Existence and algorithms for the bilevel new generalized mixed equilibrium problems in Banach spaces. Applied Mathematics and Computation, 219(6):3022–3038, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312009290.

Wang:2013:GBS

[WY13a] Qing-Wen Wang and Juan Yu. On the generalized bi (skew-) symmetric solutions of a linear matrix equation and its procrust problems. Applied Mathematics and Computation, 219(19):9872–9884, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313003147.

Wang:2013:GBH

[WY13b] Xunhua Wang and Xuefeng Yan. Global best harmony search algorithm with control parameters co-evolution based on PSO and its application to constrained optimal problems. Applied Mathematics and Computation, 219(19):10059–10072, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031300372X.

Wu:2014:NEN

[WY14] Lan Wu and Yuehan Yang. Nonnegative elastic net and application in index tracking. Applied Mathematics and Computation, 227(??):541–552, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313012174.

Wu:2012:GOC

[WYBT12] Zhiyou Wu, Yongjian Yang, Fusheng Bai, and Jing Tian. Global optimality conditions and optimization methods for quadratic assignment problems. Applied Mathematics and Computation, 218(11):6214–6231, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311014111.

Watanabe:2011:RZF

[WYK11] Kohtaro Watanabe, Hiroyuki Yamagishi, and Yoshinori Kametaka. Riemann zeta function and Lyapunov-type inequalities for certain higher order differential equations. Applied Mathematics and Computation, 218 (7):3950–3953, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311011465.

Wang:2010:CCE

[WYL10a] Shu Wang, Jianwei Yang, and Dang Luo. Convergence of compressible Euler–Poisson system to incompressible Euler equations. Applied Mathematics and Computation, 216 (11):3408–3418, August 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310004558.

Wu:2010:MCR

[WYL10b] Jie Wu, Feng Yang, and Liang Liang. A modified complete ranking of DMUs using restrictions in DEA models. Applied Mathematics and Computation, 217(2):745–751, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310006867.

Wille:2011:DCA

[WYL11] Emilio Carlos Gomes Wille, Eduardo Yabcznski, and Heitor Silvério Lopes. Discrete capacity assignment in IP networks using particle swarm optimization. Applied Mathematics and Computation, 217(12):5338–5346, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310011896.

Wang:2013:NFN

[WYT13] G. Q. Wang, C. J. Yu, and K. L. Teo. A new full Nesterov–Todd step feasible interiorpoint method for convex quadratic symmetric cone optimization. Applied Mathematics and Computation, 221(??):329–343, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313007030.

Wang:2014:IBS

[WYTL14] Deng-Shan Wang, Shujuan Yin, Ye Tian, and Yifang Liu. Integrability and bright soliton solutions to the coupled nonlinear Schrödinger equation with higher-order effects. Applied Mathematics and Computation, 229(??):296–309, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313013404.

Wen:2014:CNS

[WYW14] Rui-Ping Wen, Xi-Hong Yan, and Chuan-Long Wang. On the convergence of a new splitting iterative method for non-Hermitian positive definite linear systems. Applied Mathematics and Computation, 248(??):118–130, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314013101.

Wang:2011:RIP

[WYY11] Zhongwen Wang, Qingzhi Yang, and Yuning Yang. The relaxed inexact projection methods for the split feasibility problem. Applied Mathematics and Computation, 217 (12):5347–5359, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310011902.

Wang:2014:AOB

[WYZ14] Meilin Wang, Dansheng Yu, and Ping Zhou. On the approximation by operators of Bernstein–Stancu types. Applied Mathematics and Computation, 246(??):79–87, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314011096.

Wang:2010:GGB

[WZ10] Wu-Sheng Wang and Xiaoliang Zhou. A generalized Gronwall–Bellman–Ou–Iang type inequality for discontinuous functions and applications to BVP. Applied Mathematics and Computation, 216(11):3335–3342, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310004893.

Wang:2011:BOR

[WZ11a] Guan Jun Wang and Yuan Lin Zhang. A bivariate optimal replacement policy for a cold standby repairable system with preventive repair. Applied Mathematics and Computation, 218(7):3158–3165, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311010903.

Wang:2011:EMS

[WZ11b] JinRong Wang and Yong Zhou. Existence of mild solutions for fractional delay evolution systems. Applied Mathematics and Computation, 218(2):357–367, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007521.

Wang:2011:EAO

[WZ11c] Jinting Wang and Feng Zhang. Equilibrium analysis of the observable queues with balking and delayed repairs. Applied Mathematics and Computation, 218(6):2716–2729, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311010381.

Wang:2011:PPSa

[WZ11d] Rui Wang and Xiaosheng Zhang. Positive periodic solution for a neutral Logarithmic population model with feedback control. Applied Mathematics and Computation, 217(19):7692–7702, June 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311002955.

Wang:2011:SSO

[WZ11e] Shanshan Wang and Luming Zhang. Split-step orthogonal spline collocation methods for nonlinear Schrödinger equations in one, two, and three dimensions. Applied Mathematics and Computation, 218(5):1903–1916, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311009192.

Wang:2011:RHA

[WZ11f] Zhujun Wang and Detong Zhu. A reduced Hessian algorithm with line search filter method for nonlinear programming. Applied Mathematics and Computation, 217 (19):7679–7691, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311002785.

Wong:2011:ABM

[WZ11g] Hoi Ying Wong and Jing Zhao. An artificial boundary method for the Hull–White model of American interest rate derivatives. Applied Mathematics and Computation, 217(9):4627–4643, January 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310011276.

Wrobel:2011:LBF

[WZ11h] Iwona Wróbel and Krystyna Zieţak. On the Legendre-based filters of Persson and Strang. Applied Mathematics and Computation, 218 (8):4216–4233, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311012252.

Wu:2011:PST

[WZ11i] Yanyan Wu and Zengqin Zhao. Positive solutions for third-order boundary value problems with change of signs. Applied Mathematics and Computation, 218(6):2744–2749, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311010411.

Wu:2012:BLC

[WZ12] Yusen Wu and Cui Zhang. Bifurcation of limit cycles and pseudo-isochronous center at degenerate singular point in a septic system. Applied Mathematics and Computation, 218(17):8513–8525, May 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312001312.

Wan:2013:ABS

[WZ13a] Li Wan and Qinghua Zhou. Asymptotic behaviors of stochastic Cohen–Grossberg neural networks with mixed time-delays. Applied Mathematics and Computation, 225 (??):541–549, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010564.

Wang:2013:HBT

[WZ13b] Ling Wang and Hongyong Zhao. Hopf bifurcation and Turing instability of 2-D Lengyel–Epstein system with reactiondiffusion terms. Applied Mathematics and Computation, 219(17):9229–9244, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313003329.

Wang:2013:CSA

[WZ13c] Zhi-Gang Wang and Zhi-Hua Zhang. Comments on “On analytic functions with generalized bounded Mocanu variation” by K.I. Noor and A. Muhammad [Appl. Math. Comput. 196 (2008) 802–811]. Applied Mathematics and Computation, 219(11):5754–5756, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300312012465.  See

[NM08].

Wu:2013:ISS

[WZ13d] Jinming Wu and Xiaolei Zhang. Integro sextic spline interpolation and its super convergence. Applied Mathematics and Computation, 219(12):6431–6436, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313000039.

Wang:2014:BAT

[WZ14a] Yongfu Wang and Qin Zhang. Blowup analysis for two-dimensional viscous compressible, heat-conductive Navier–Stokes equations. Applied Mathematics and Computation, 232(??):719–726, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314001404.

Wang:2014:PAS

[WZ14b] Zhujun Wang and Detong Zhu. Projected affine-scaling interior-point Newton’s method with line search filter for box constrained optimization. Applied Mathematics and Computation, 230(??):484–495, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313013817.

Wu:2010:ISE

[WZDW10] Ai-Guo Wu, Xianlin Zeng, Guang-Ren Duan, and Wei-Jun Wu. Iterative solutions to the extended Sylvester-conjugate matrix equations. Applied Mathematics and Computation, 217(1):130–142, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310005734.

Wang:2012:SNR

[WZG12] Wu-Sheng Wang, Xiaoliang Zhou, and Zhonghai Guo. Some new retarded nonlinear integral inequalities and their applications in differential-integral equations. Applied Mathematics and Computation, 218 (21):10726–10736, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312004456.

Wang:2010:PAB

[WzH10] Ling Wang and Fu zhuo Huang. Parameter analysis based on stochastic model for differential evolution algorithm. Applied Mathematics and Computation, 217(7):3263–3273, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300310009240.

Wen:2013:OBC

[WZH13] Shiping Wen, Zhigang Zeng, and Tingwen Huang. Observer-based H_{∞} control of discrete time-delay systems with random communication packet losses and multiplicative noises. Applied Mathematics and Computation, 219(12):6484–6493, February 15, 2013.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312013057.

Wang:2013:CEC

[WZJ13] Tingchun Wang, Luming Zhang, and Yong Jiang. Convergence of an efficient and compact finite difference scheme for the Klein–Gordon– Zakharov equation. Applied Mathematics and Computation, 221(??):433–443, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031300698X.

Wu:2012:CFS

[WZL12] Ai-Guo Wu, Enze Zhang, and Fuchun Liu. On closed-form solutions to the generalized Sylvester-conjugate matrix equation. Applied Mathematics and Computation, 218(19):9730–9741, June 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312002524.

Wang:2014:NCI

[WZL14a] JinRong Wang, Yong Zhou, and Zeng Lin. On a new class of impulsive fractional differential equations. Applied Mathematics and Computation, 242(??):649–657, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314008364.

Wu:2014:NZA

[WZL14b] Juanjuan Wu, Yongkang Zhang, and Cuiping Li. On the number of zeros of Abelian integrals for a kind of quartic Hamiltonians. Applied Mathematics and Computation, 228(??):329–335, February 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313012617.

Wang:2014:PUL

[WZSW14] Jinfeng Wang, Xin Zhang, Junping Shi, and Yuwen Wang. Profile of the unique limit cycle in a class of general predator-prey systems. Applied Mathematics and Computation, 242(??):397–406, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314006857.

Wang:2012:PCS

[WZW12] Zhi-Gang Wang, Guo-Wei Zhang, and Feng-Hua Wen. Properties and characteristics of the Srivastava–Khairnar–More integral operator. Applied Mathematics and Computation, 218(15):7747–7758, April 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312000768.

Wu:2014:APS

[WZW14] Ruihua Wu, Xiaoling Zou, and Ke Wang. Asymptotic properties of stochastic hybrid Gilpin–Ayala system with jumps. Applied Mathematics and Computation, 249 (??):53–66, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031401409X.

Wang:2013:LES

[WZWW13] Junyi Wang, Huaguang Zhang, Zhanshan Wang, and Binrui Wang. Local exponential synchronization in complex dynamical networks with time-varying delay and hybrid coupling. Applied Mathematics and Computation, 225(??):16–32, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313009946.

Wang:2014:SBO

[WZY14] Guoqiu Wang, Qingyun Zou, and Mengyun Yang. Sub-band operators and saddle point wavelets. Applied Mathematics and Computation, 227(??):27–42, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313011338.

Wang:2011:ADP

[WZZ11] Zhen Wang, Li Zou, and Zhi Zong. Adomian decomposition and Padé approximate for solving differential-difference equation. Applied Mathematics and Computation, 218 (4):1371–1378, October 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311008460.

Wang:2013:RDS

[WZZ13] Jinhuan Wang, Lizhong Zhao, and Sining Zheng. A reaction-diffusion system with mixed-type coupling. Applied Mathematics and Computation, 219(9):4219–4224, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312011599.

Wang:2014:CPS

[WZZ14] Guan-Wei Wang, Chun-Xia Zhang, and Jian Zhuang. Clustering with Prim’s sequential representation of minimum spanning tree. Applied Mathematics and Computation, 247 (??):521–534, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314012442.

Wang:2011:SFM

[WZZH11] Xiangling Wang, Zhibin Zhu, Shuangyong Zuo, and Qingqun Huang. An SQPfilter method for inequality constrained optimization and its global convergence. Applied Mathematics and Computation, 217 (24):10224–10230, August 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311006746.

Wang:2012:STS

[WZZL12] Zhaocai Wang, Yiming Zhang, Weihua Zhou, and Haifeng Liu. Solving traveling salesman problem in the Adleman–Lipton model. Applied Mathematics and Computation, 219 (4):2267–2270, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312008661.

Xiao:2013:CPO

[XBC^{+}13] Manyu Xiao, Piotr Breitkopf, Rajan Filomeno Coelho, Catherine Knopf-Lenoir, Pierre Villon, and Weihong Zhang. Constrained Proper Orthogonal Decomposition based on QRfactorization for aerodynamical shape optimization. Applied Mathematics and Computation, 223(??):254–263, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313008424.

Xiao:2014:POD

[XBC^{+}14] Manyu Xiao, Piotr Breitkopf, Rajan Filomeno Coelho, Pierre Villon, and Weihong Zhang. Proper orthogonal decomposition with high number of linear constraints for aerodynamical shape optimization. Applied Mathematics and Computation, 247(??):1096–1112, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314012934.

Xiaolu:2014:MSS

[XBYF14] Liu Xiaolu, Bai Baocun, Chen Yingwu, and Yao Feng. Multi satellites scheduling algorithm based on task merging mechanism. Applied Mathematics and Computation, 230(??):687–700, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313013994.

Xian:2010:SRNb

[XC10a] Da-Quan Xian and Han-Lin Chen. Symmetry reduced and new exact non-traveling wave solutions of potential Kadomtsev–Petviashvili equation with p-power. Applied Mathematics and Computation, 217(4):1340–1349, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300309004457.

Xian:2010:SRNa

[XC10b] Da-Quan Xian and Han-Lin Chen. Symmetry reductions and new exact non-traveling wave solutions of potential Kadomtsev–Petviashvili equation with p-power. Applied Mathematics and Computation, 216(1):70–79, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310000044.

Xiang:2010:RRC

[XC10c] Zhengrong Xiang and Qingwei Chen. Robust reliable control for uncertain switched nonlinear systems with time delay under asynchronous switching. Applied Mathematics and Computation, 216(3):800–811, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310001086.

Xu:2013:DDG

[XC13] Aimin Xu and Zhongdi Cen. Divided differences and a general explicit formula for sequences of Mansour–Mulay–Shattuck. Applied Mathematics and Computation, 224 (??):719–723, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313009764.

Xu:2014:MSN

[XC14] Liping Xu and Haibo Chen. Multiplicity of small negative-energy solutions for a class of nonlinear Schrödinger–Poisson systems. Applied Mathematics and Computation, 243(??):817–824, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314008790.

Xing:2014:EMS

[XCH14] Hui Xing, Hongbin Chen, and Xibing He. Exact multiplicity and stability of solutions of second-order Neumann boundary value problem. Applied Mathematics and Computation, 232(??):1104–1111, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314001647.

Xu:2013:SCE

[XCS13] Qing-Hua Xu, Qiao-Min Cai, and H. M. Srivastava. Sharp coefficient estimates for certain subclasses of starlike functions of complex order. Applied Mathematics and Computation, 225(??):43–49, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031300996X.

Xu:2010:GES

[XCT10] Honglei Xu, Yuanqiang Chen, and Kok Lay Teo. Global exponential stability of impulsive discrete-time neural networks with timevarying delays. Applied Mathematics and Computation, 217(2):537–544, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310006569.

Xie:2011:NMG

[XD11a] Manlin Xie and Xuanhao Ding. A new method for a generalized Hirota–Satsuma coupled KdV equation. Applied Mathematics and Computation, 217(17):7117–7125, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300311000646. See com-

ment [Aba12].

Xu:2011:TEA

[XD11b] Shengqiang Xu and Jinqiao Duan. A Taylor expansion approach for solving partial differential equations with random Neumann boundary conditions. Applied Mathematics and Computation, 217(23):9532–9542, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311005170.

Xie:2012:DRF

[XDG12] Hui Xie, Hong-Liang Dai, and Zhi-Yang Guo. Dynamic response of a FGPM hollow cylinder under the coupling of multifields. Applied Mathematics and Computation, 218(21):10492–10499, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312003967.

Xue:2014:NSI

[XDZ14] Yunhua Xue, Huoyuan Duan, and Qiang Zhang. A new and simple implementation of the element-local L ^{2}-projected continuous finite element method. Applied Mathematics and Computation, 228(??):170–183, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313012654.

Xu:2012:BSR

[XFS12] Guang-Hui Xu, Kun-Fu Fang, and Jian Shen. Bounds on the spectral radii of digraphs in terms of walks. Applied Mathematics and Computation, 219(8):3721–3728, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312009861.

Xu:2012:MSA

[XFW^{+}12] Min Xu, Qin Fang, Ren-Hong Wang, Zi-Wu Jiang, and Ming-Zeng Liu. A multinomial spline approximation scheme using spline quasi-interpolants. Applied Mathematics and Computation, 218(9):5081–5089, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311013191.

Xingong:2010:SMG

[XG10] Zhang Xingong and Yan Guangle. Singlemachine group scheduling problems with deteriorated and learning effect. Applied Mathematics and Computation, 216 (4):1259–1266, April 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310001864.

Xu:2013:AFP

[XG13] Ding Xu and Xin Guo. Application of fixed point method to obtain semi-analytical solution to Blasius flow and its variation. Applied Mathematics and Computation, 224 (??):791–802, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031300934X.

Xu:2011:SSB

[XGG^{+}11] Peng-Bo Xu, Yi-Tian Gao, Xiao-Ling Gai, De-Xin Meng, Yu-Jia Shen, and Lei Wang. Soliton solutions, Bäcklund transformation and Wronskian solutions for the extended (2+1)- dimensional Konopelchenko–Dubrovsky equations in fluid mechanics. Applied Mathematics and Computation, 218(6):2489–2496, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300311010113.

Xu:2010:SNF

[XGM10] Youji Xu, Chenghua Gao, and Ruyun Ma. Solvability of a nonlinear fourth-order discrete problem at resonance. Applied Mathematics and Computation, 216(2):662–670, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310001360.

Xia:2013:ANP

[XGZW13] Lijun Xia, Shujing Gao, Qin Zou, and Jianping Wang. Analysis of a nonautonomous plant disease model with latent period. Applied Mathematics and Computation, 223(??):147–159, October 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313008552.

Xu:2010:NLC

[XH10] Weijiao Xu and Maoan Han. On the number of limit cycles of a Z_{4}-equivariant quintic polynomial system. Applied Mathematics and Computation, 216(10):3022–3034, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310004340.

Xu:2011:CCL

[XH11] Qiujin Xu and Wentao Huang. The center conditions and local bifurcation of critical periods for a Liénard system. Applied Mathematics and Computation, 217(15):6637–6643, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311000518.

Xiang:2013:IDG

[XH13a] Shuhuang Xiang and Kaixian He. On the implementation of discontinuous Galerkin methods for Volterra integral equations with highly oscillatory Bessel kernels. Applied Mathematics and Computation, 219(9):4884–4891, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300312010892.

Xu:2013:MAR

[XH13b] Houbao Xu and Weiwei Hu. Modelling and analysis of repairable systems with preventive maintenance. Applied Mathematics and Computation, 224(??):46–53, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313009193.

Xiong:2014:LCB

[XH14] Yanqin Xiong and Maoan Han. Limit cycle bifurcations in a class of perturbed piecewise smooth systems. Applied Mathematics and Computation, 242(??):47–64, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314007000.

Xie:2011:PTW

[XHG11] Shaolong Xie, Xiaochun Hong, and Bin Gao. The periodic traveling-wave solutions of the short-pulse equation. Applied Mathematics and Computation, 218(6):2542–2548, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300311010186.

Xu:2012:MFE

[XHXH12] Youcai Xu, Bing Hu, Xiaoping Xie, and Jinsong Hu. Mixed finite element analysis for dissipative SRLW equations with damping term. Applied Mathematics and Computation, 218(9):4788–4797, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311012501.

Xiang:2010:ASK

[Xia10a] Chunhuan Xiang. Analytical solutions of KdV equation with relaxation effect of inhomogeneous medium. Applied Mathematics and Computation, 216(8):2235–2239, June 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310003310.

Xiang:2010:TWS

[Xia10b] Chunhuan Xiang. Travelling wave solutions of reduced super-KdV equation: a perspective from Lamé equation. Applied Mathematics and Computation, 217(1):42–47, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310005850.

Xiao:2010:ISC

[Xia10c] Xiaoyong Xiao. Improving speed of convergence for the prices of European options in binomial trees with even numbers of steps. Applied Mathematics and Computation, 216(9):2659–2670, July 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310003656.

Xiao:2012:ROL

[Xia12] Ying-Xiong Xiao. Remarks on Ostrowski-like inequalities. Applied Mathematics and Computation, 219(3):1158–1162, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312007254.

Xiao:2013:CAF

[Xia13] Haibin Xiao. Compact attractors of fourth order parabolic equations on R^{n}. Applied Mathematics and Computation, 219 (18):9827–9837, May 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313003822.

Xiang:2014:LTA

[Xia14] Shuhuang Xiang. Laplace transforms for approximation of highly oscillatory Volterra integral equations of the first kind. Applied Mathematics and Computation, 232(??):944–954, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300314000915.

Xie:2011:NMS

[Xie11a]  Dongxiu Xie.  A numerical method of

structure-preserving model updating problem and its perturbation theory. Applied Mathematics and Computation, 217 (13):6364–6371, March 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311000269. See corrigendum

[Xie12a].

Xie:2011:SSF

[Xie11b] Shenglan Xie. Stability of sets of functional differential equations with impulse effect. Applied Mathematics and Computation, 218(2):592–597, September 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007922.

Xie:2012:CNM

[Xie12a] Dongxiu Xie. Corrigendum to “A numerical method of structure-preserving model updating problem and its perturbation theory” [Appl. Math. Comput. 217 (13) (2011)

6364–6371]. Applied Mathematics and Computation, 218(19):10007, June 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312002779. See corrigendum

[Xie11a].

Xie:2012:WES

[Xie12b] Zheng Xie. Wave equation simulation on manifold by unconditional stable schemes. Applied Mathematics and Computation, 218 (15):7967–7971, April 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312001506.

Xin:2010:EFR

[Xin10] Hua Xin. The exponential function rational expansion method and exact solutions to nonlinear lattice equations system. Applied Mathematics and Computation, 217 (4):1561–1565, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309006237.

Xiao:2014:MFH

[XK14] Xuejiao Xiao and Chang Nyung Kim. Magnetohydrodynamic flows in a hairpin duct under a magnetic field applied perpendicular to the plane of flow. Applied Mathematics and Computation, 240(??):1–15, August 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031400589X.

Xu:2010:IAA

[XKH10] Haiyan Xu, D. Marc Kilgour, and Keith W. Hipel. An integrated algebraic approach to conflict resolution with three-level preference. Applied Mathematics and Computation, 216(3):693–707, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310000676.

Xia:2011:ECV

[XL11] Ning-Mao Xia and Ai-Hong Lin. Existence conditions and variational approach for adapted solutions of the two-point boundary value problem of stochastic differential equations. Applied Mathematics and Computation, 217(20):7900–7912, June 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311003249.

Xu:2014:ISC

[XL14] Na Xu and Wenbin Liu. Iterative solutions for a coupled system of fractional differentialintegral equations with two-point boundary conditions. Applied Mathematics and Computation, 244(??):903–911, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314010005.

Xie:2012:HOC

[XLGW12] Shusen Xie, Peng Li, Zhen Gao, and Hui Wang. High order compact finite difference schemes for a system of third order boundary value problem. Applied Mathematics and Computation, 219(5):2564–2573, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312008843.

Li:2013:BMC

[xLhYlZ13] Zhao xiang Li, Zhong hua Yang, and Hai long Zhu. Bifurcation method for computing the multiple positive solutions to p-henon equation. Applied Mathematics and Computation, 220(??):593–601, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313006346.

Xiao:2011:IPP

[XLL11] Cuie Xiao, Jinbo Liu, and Yiliang Liu. An inverse pollution problem in porous media. Applied Mathematics and Computation, 218 (7):3649–3653, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031101157X.

Xie:2014:NSC

[XLLL14] Dongmei Xie, Qingli Liu, Liangfu Lv, and Songying Li. Necessary and sufficient condition for the group consensus of multi-agent systems. Applied Mathematics and Computation, 243(??):870–878, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314009126.

Xiang:2013:DLV

[XLS13a] Zhongyi Xiang, Dan Long, and Xinyu Song. A delayed Lotka–Volterra model with birth pulse and impulsive effect at different moment on the prey. Applied Mathematics and Computation, 219(20):10263–10270, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313003913.

Xu:2013:CEI

[XLS13b] Qing-Hua Xu, Chun-Bo Lv, and H. M. Srivastava. Coefficient estimates for the inverses of a certain general class of spirallike functions. Applied Mathematics and Computation, 219(12):7000–7011, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312013318.

Xu:2011:ESG

[XLW11] Jieqiong Xu, Qunhong Li, and Nan Wang. Existence and stability of the grazing periodic trajectory in a two-degree-of-freedom vibroimpact system. Applied Mathematics and Computation, 217(12):5537–5546, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310012294.

Xie:2010:GBL

[XLY10] Li Xie, Yanjun Liu, and Huizhong Yang. Gradient based and least squares based iterative algorithms for matrix equations AXB + CX^{T} D = F. Applied Mathematics and Computation, 217(5):2191–2199, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310007642.

Xu:2014:ESG

[XLY14] Yongjia Xu, Yongzeng Lai, and Haixiang Yao. Efficient simulation of Greeks of multiasset European and Asian style options by Malliavin calculus and quasi-Monte Carlo methods. Applied Mathematics and Computation, 236(??):493–511, June 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314004196.

Xin:2010:SRE

[XLZ10] Xiang-Peng Xin, Xi-Qiang Liu, and Lin-Lin Zhang. Symmetry reduction, exact solutions and conservation laws of the Sawada–Kotera–Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 216(4):1065–1071, April 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310001517.

Xu:2010:BIP

[XM10] Jia Xu and Ruyun Ma. Bifurcation from interval and positive solutions for second order periodic boundary value problems. Applied Mathematics and Computation, 216(8):2463–2471, June 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310003462.

Xie:2011:SLM

[XM11] Yajun Xie and Changfeng Ma. A smoothing Levenberg–Marquardt algorithm for solving a class of stochastic linear complementarity problem. Applied Mathematics and Computation, 217(9):4459–4472, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310011094.

Xu:2012:ATB

[XM12] Yingxiang Xu and Vital D. Mabonzo. Analysis on Takens–Bogdanov points for delay differential equations. Applied Mathematics and Computation, 218(24):11891–11899, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312005693.

Xiang:2014:HHE

[XMA14] Wanli Xiang, Shoufeng Ma, and Meiqing An. hABCDE: a hybrid evolutionary algorithm based on artificial bee colony algorithm and differential evolution. Applied Mathematics and Computation, 238(??):370–386, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314004172.

Xia:2014:ECT

[XMWD14] Chengyi Xia, Qing Miao, Juan Wang, and Shuai Ding. Evolution of cooperation in the traveler’s dilemma game on two coupled lattices. Applied Mathematics and Computation, 246(??):389–398, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031401100X.

Xiao:2012:GSF

[XP12] Wei Xiao and Yan Ping. Global solutions and finite time blow up for some system of nonlinear wave equations. Applied Mathematics and Computation, 219(8):3754–3768, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312009927.

Xia:2014:ITU

[XPLZ14] Jianwei Xia, Ju H. Park, Tae H. Lee, and Baoyong Zhang. H_{∞} tracking of uncertain stochastic time-delay systems: Memory state–feedback controller design. Applied Mathematics and Computation, 249 (??):356–370, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314013952.

Xiong:2011:CRN

[XQ11a] Zhiping Xiong and Yingying Qin. The common Re-nnd and Re-pd solutions to the matrix equations AX = C and XB = D. Applied Mathematics and Computation, 218 (7):3330–3337, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311011155.

Xiong:2011:CPE

[XQ11b] Zhiping Xiong and Yingying Qin. On the characteristic polynomials and eigenvalues of some specially bordered matrices. Applied Mathematics and Computation, 218 (4):1407–1416, October 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311008502.

Xiong:2012:ISS

[XQ12] ZhiPing Xiong and Yingying Qin. On the inverse of a special Schur complement. Applied Mathematics and Computation, 218 (14):7679–7684, March 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312000835.

Xiong:2013:NRO

[XQ13] Zhiping Xiong and Yingying Qin. A note on the reverse order law for reflexive generalized inverse of multiple matrix products. Applied Mathematics and Computation, 219(9):4255–4265, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312011678.

Xia:2013:NMR

[XRBZ13] Liang Xia, Balaji Raghavan, Piotr Breitkopf, and Weihong Zhang. Numerical material representation using proper orthogonal decomposition and diffuse approximation. Applied Mathematics and Computation, 224(??):450–462, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031300920X.

Xiang:2011:RRSb

[XSC11] Zhengrong Xiang, Ya-Nan Sun, and Qingwei Chen. Robust reliable stabilization of uncertain switched neutral systems with delayed switching. Applied Mathematics and Computation, 217(23):9835–9844, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311006527.

Xie:2010:NRT

[XSL10] Pingli Xie, Dongyang Shi, and Hua Li. A new robust C^{0}-type nonconforming triangular element for singular perturbation problems. Applied Mathematics and Computation, 217(8):3832–3843, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310010027.

Xiao:2012:GGC

[XSX12] Minqing Xiao, Hongye Su, and Weihua Xu. Generalized guaranteed cost control with Dstability and multiple output constraints. Applied Mathematics and Computation, 218 (24):12013–12027, August 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312005930.

Xie:2014:CPB

[XT14] Li Xie and Hongjiong Tian. Continuous parallel block methods and their applications. Applied Mathematics and Computation, 241(??):356–370, August 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314006912.

Xue:2012:DTH

[XTAJ12] Yu-Shan Xue, Bo Tian, Wen-Bao Ai, and Yan Jiang. Darboux transformation and Hamiltonian structure for the Jaulent–Miodek hierarchy. Applied Mathematics and Computation, 218(24):11738–11750, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312004894.

Xu:2010:SBA

[XTL10] Changjin Xu, Xianhua Tang, and Maoxin Liao. Stability and bifurcation analysis of a delayed predator–prey model of prey dispersal in two-patch environments. Applied Mathematics and Computation, 216(10):2920–2936, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310004108.

Xu:2010:SWD

[Xu10a] Jianhong Xu. Simulating weighted, directed small-world networks. Applied Mathematics and Computation, 216(7):2118–2128, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310002870.

Xu:2010:ICH

[Xu10b] Xi-Xiang Xu. An integrable coupling hierarchy of the Mkdv integrable systems, its Hamiltonian structure and corresponding nonisospectral integrable hierarchy. Applied Mathematics and Computation, 216(1):344–353, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310000299.

Xu:2010:FAS

[Xu10c] Xiangjian Xu. A fast algorithm for solving a Toeplitz system. Applied Mathematics and Computation, 217(5):1944–1948, November 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310007356.

Xu:2011:NTE

[Xu11] Gui-Qiong Xu. New types of exact solutions for the fourth-order dispersive cubic– quintic nonlinear Schrödinger equation. Applied Mathematics and Computation, 217 (12):5967–5971, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310012087.

Xu:2012:BOO

[Xu12a] Gongxian Xu. Bi-objective optimization of biochemical systems by linear programming. Applied Mathematics and Computation, 218(14):7562–7572, March 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312000562.

Xu:2012:GDS

[Xu12b] Rui Xu. Global dynamics of an SEIS epidemic model with saturation incidence and latent period. Applied Mathematics and Computation, 218(15):7927–7938, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312001324.

Xu:2013:APT

[Xu13a] Gang Xu. An adaptive parameter tuning of particle swarm optimization algorithm. Applied Mathematics and Computation, 219(9):4560–4569, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312010831.

Xu:2013:ENP

[Xu13b] Jia Xu. Existence and nonexistence of positive solutions for nonlocal problems with inhomogeneous strong Allee effect. Applied Mathematics and Computation, 219 (23):11029–11035, August 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313004529.

Xu:2013:ABD

[Xu13c] Xiaoming Xu. Anticipated backward doubly stochastic differential equations. Applied Mathematics and Computation, 220 (??):53–62, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313006012.

Xu:2014:PDC

[Xu14a] Ling Xu. A proportional differential control method for a time-delay system using the Taylor expansion approximation. Applied Mathematics and Computation, 236(??):391–399, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S009630031400335X.

Xu:2014:GDS

[Xu14b] Rui Xu. Global dynamics of an SEIRI epidemiological model with time delay. Applied Mathematics and Computation, 232 (??):436–444, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314001374.

Xu:2014:CCF

[Xu14c] Wenwen Xu. Cell-centered finite difference method for parabolic equation. Applied Mathematics and Computation, 235 (??):66–79, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314003051.

Xie:2010:CGK

[XW10] Shaolong Xie and Lin Wang. Compacton and generalized kink wave solutions of the CH–DP equation. Applied Mathematics and Computation, 215(11):4028–4039, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309010571.

Xiang:2013:NSV

[XW13] Shuhuang Xiang and Qinghua Wu. Numerical solutions to Volterra integral equations of the second kind with oscillatory trigonometric kernels. Applied Mathematics and Computation, 223(??):34–44, October 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313008254.

Xiang:2011:RRSa

[XWC11] Zhengrong Xiang, Ronghao Wang, and Qingwei Chen. Robust reliable stabilization of stochastic switched nonlinear systems under asynchronous switching. Applied Mathematics and Computation, 217(19):7725–7736, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311002992.

Xu:2011:DTP

[XwK11] Chendong Xu and Tae wan Kim. Determining tension parameters in rational gammaspline interpolation. Applied Mathematics and Computation, 217(12):5243–5258, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310011252.

Xiao:2012:APC

[XWMW12] Jiang-Wen Xiao, Zhi-Wei Wang, Wen-Tuan Miao, and Yan-Wu Wang. Adaptive pinning control for the projective synchronization of drive-response dynamical networks. Applied Mathematics and Computation, 219 (5):2780–2788, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312009046.

Xie:2010:AHT

[XWWC10] Kanghe Xie, Yulin Wang, Kun Wang, and Xin Cai. Application of Hankel transforms to boundary value problems of water flow due to a circular source. Applied Mathematics and Computation, 216(5):1469–1477, May 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310002377.

Xiao:2010:NAS

[XWWF10] Xuan Xiao, Bo Wang, Shunting Wang, and Mengyin Fu. Noise analysis and suppression method in attitude determination using the global positioning system (GPS). Applied Mathematics and Computation, 217 (8):3985–3992, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310010222.

Xu:2011:NNS

[XWZF11] Min Xu, Ren-Hong Wang, Ji-Hong Zhang, and Qin Fang. A novel numerical scheme for solving Burgers’ equation. Applied Mathematics and Computation, 217(9):4473–4482, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310011100.

Xu:2010:GBD

[XX10] Meirong Xu and Sipeng Xu. Global behavior of a discrete haematopoiesis model with several delays. Applied Mathematics and Computation, 217(1):441–449, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310006351.

Xia:2011:CPN

[XX11] Suxia Xia and Guixiang Xu. On the Cauchy problem for the NLS equations in R^{1+1}. Applied Mathematics and Computation, 217(19):7737–7752, June 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311003006.

Xiang:2014:ESD

[XX14a] Mei Xiang and Zhengrong Xiang. Exponential stability of discrete-time switched linear positive systems with time-delay. Applied Mathematics and Computation, 230 (??):193–199, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313014082.

Xiao:2014:NRA

[XX14b] Jian Xiao and Weiming Xiang. New results on asynchronous H_{∞} control for switched discrete-time linear systems under dwell time constraint. Applied Mathematics and Computation, 242(??):601–611, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314007978.

Xu:2014:NEC

[XX14c] Zhenhua Xu and Shuhuang Xiang. Numerical evaluation of a class of highly oscillatory integrals involving Airy functions. Applied Mathematics and Computation, 246(??):54–63, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314011163.

Xu:2012:CGS

[XXS12] Qing-Hua Xu, Hai-Gen Xiao, and H. M. Srivastava. A certain general subclass of analytic and bi-univalent functions and associated coefficient estimate problems. Applied Mathematics and Computation, 218 (23):11461–11465, August 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031200519X.

Xu:2014:SAD

[XXS14] Qing-Hua Xu, Hai-Gen Xiao, and H. M. Srivastava. Some applications of differential subordination and the Dziok–Srivastava convolution operator. Applied Mathematics and Computation, 230(??):496–508, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313013489.

Xu:2010:FSA

[XXZH10] Tao Xu, Tianshuang Xu, Wenjie Zuo, and Liang Hao. Fast sensitivity analysis of defective system. Applied Mathematics and Computation, 217(7):3248–3256, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310009215.

Xiao:2010:TAS

[XY10] Yu Xiao and Bo Yu. A truncated aggregate smoothing Newton method for minimax problems. Applied Mathematics and Computation, 216(6):1868–1879, May 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309010236.

Xu:2011:CBF

[XY11] Dehua Xu and Yunqiang Yin. Comments on “A bicriteria flowshop scheduling problem with setup times”. Applied Mathematics and Computation, 217(17):7361–7364, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311001408. See [EG06].

Xiao:2011:LFN

[XYC11] Liu-Chao Xiao, Yong-Qin Yang, and Shao-Chun Chen. Locking-free nonconforming finite elements for three-dimensional elasticity problem. Applied Mathematics and Computation, 217(12):5790–5797, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310012634.

Xu:2010:CNM

[XYL10] Dehua Xu, Yunqiang Yin, and Hongxing Li. Comments on “A note on minimizing maximum lateness in an m-machine scheduling problem with a learning effect”. Applied Mathematics and Computation, 217 (2):939–943, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310006752. See [?].

Xue:2014:GSM

[XYL14] Yakui Xue, Xinpeng Yuan, and Maoxing Liu. Global stability of a multi-group sei model. Applied Mathematics and Computation, 226(??):51–60, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010291.

Xiao-Yong:2012:GLP

[XYY12] Zhang Xiao-Yong and Li Yan. Generalized Laguerre pseudospectral method based Laguerre interpolation. Applied Mathematics and Computation, 219(5):2545–2563, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312008831.

Xu:2012:PSI

[XZ12] Yumei Xu and Haijun Zhang. Positive solutions of an infinite boundary value problem for n th-order nonlinear impulsive singular integro-differential equations in Banach spaces. Applied Mathematics and Computation, 218(9):5806–5818, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311014469.

Xu:2013:MAD

[XZ13a] Rui Xu and Shihua Zhang. Modelling and analysis of a delayed predator-prey model with disease in the predator. Applied Mathematics and Computation, 224(??):372–386, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313009351.

Xu:2013:IDS

[XZ13b] Ying Xu and Ruguang Zhou. Integrable decompositions of a symmetric matrix Kaup– Newell equation and a symmetric matrix derivative nonlinear Schrödinger equation. Applied Mathematics and Computation, 219 (9):4551–4559, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312010818.

Xue:2013:NFD

[XZ13c] Guan-Yu Xue and Luming Zhang. A new finite difference scheme for generalized Rosenau– Burgers equation. Applied Mathematics and Computation, 222(??):490–496, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313008023.

Xie:2011:CAP

[XZF11] Liping Xie, Jianchao Zeng, and Richard A. Formato. Convergence analysis and performance of the extended artificial physics optimization algorithm. Applied Mathematics and Computation, 218(8):4000–4011, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311002669.

Xie:2013:TTB

[XZH13] Shaolong Xie, Yuzhong Zhang, and Jianghua He. Two types of bounded traveling-wave solutions of a two-component Camassa–Holm equation. Applied Mathematics and Computation, 219(20):10271–10282, June 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313003925.

Xie:2010:BTW

[XZT10] Yongan Xie, Bowen Zhou, and Shengqiang Tang. Bifurcations of travelling wave solutions for the generalized (2 + 1)-dimensional Boussinesq–Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 217 (6):2433–2447, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007897.

Xiao:2011:WHH

[XZW11] Zhen-Gang Xiao, Zhi-Hua Zhang, and Yu-Dong Wu. On weighted Hermite–Hadamard inequalities. Applied Mathematics and Computation, 218(3):1147–1152, October 1, 2011.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/

S0096300311004498. Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Xiao:2014:TCI

[XZY14a] Aiguo Xiao, Gengen Zhang, and Xing Yi. Two classes of implicit-explicit multistep methods for nonlinear stiff initial-value problems. Applied Mathematics and Computation, 247(??):47–60, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031401176X.

Xue:2014:SMS

[XZY14b] Pengfei Xue, Yulin Zhang, and Xianyu Yu. Single-machine scheduling with piecerate maintenance and interval constrained position-dependent processing times. Applied Mathematics and Computation, 226 (??):415–422, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313010916.

Xiao:2012:PFS

[XZZK12] Yi-Yong Xiao, Ren-Qian Zhang, Qiu-Hong Zhao, and Ikou Kaku. Permutation flow shop scheduling with order acceptance and weighted tardiness. Applied Mathematics and Computation, 218(15):7911–7926, April 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312001245.

Xue:2014:EAS

[XZZX14] Yu Xue, Shuiming Zhong, Yi Zhuang, and Bin Xu. An ensemble algorithm with self-adaptive learning techniques for high-dimensional numerical optimization. Applied Mathematics and Computation, 231(??):329–346, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313014203.

Younis:2014:SWS

[YA14] Muhammad Younis and Safdar Ali. Solitary wave and shock wave solitons to the transmission line model for nano-ionic currents along microtubules. Applied Mathematics and Computation, 246(??):460–463, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031401162X.

Yuksel:2010:PWB

[YAAS10] Aydin Yuksel, Levent Akdeniz, and Aslihan Altay-Salih. On the performance of West’s bubble test: a simulation approach. Applied Mathematics and Computation, 217 (7):3236–3247, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310009203.

Yadav:2013:SSF

[Yad13] Manoj K. Yadav. Solutions of a system of forced Burgers equation. Applied Mathematics and Computation, 225(??):151–157, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313009910.

Yadav:2014:AMS

[Yad14] Rani Yadav. Approximation by modified Srivastava–Gupta operators. Applied Mathematics and Computation, 226(??):61–66, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313010965.

Yaman:2014:ATH

[Yam14] Irem Yaman. Affine transformational HDMR and linearised rational least squares approximation. Applied Mathematics and Computation, 246(??):761–770, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314011746.

Yang:2010:ACS

[Yan10a] Chi-Ching Yang. Adaptive control and synchronization of identical new chaotic flows with unknown parameters via single input. Applied Mathematics and Computation, 216(4):1316–1324, April 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310001943.

Yang:2010:TTL

[Yan10b] Min Yang. Two-time level ADI finite volume method for a class of second-order hyperbolic problems. Applied Mathematics and Computation, 215(9):3239–3248, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300309009114.

Yang:2010:CES

[Yan10c] Ming-Chien Yang. Cycle embedding in star graphs with conditional edge faults. Applied Mathematics and Computation, 215 (10):3541–3546, January 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309009679.

Yang:2010:ECV

[Yan10d] Ming-Chien Yang. Embedding cycles of various lengths into star graphs with both edge and vertex faults. Applied Mathematics and Computation, 216(12):3754–3760, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310005825.

Yang:2010:SMS

[Yan10e] Suh-Jenq Yang. Single-machine scheduling problems with both start-time dependent learning and position dependent aging effects under deteriorating maintenance consideration. Applied Mathematics and Computation, 217(7):3321–3329, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310009288.

Yang:2010:SJS

[Yan10f] Weifeng Yang. Symmetries of the Julia sets of Newton’s method for multiple root. Applied Mathematics and Computation, 217 (6):2490–2494, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310008064.

Yan:2011:ESN

[Yan11a] Zhenya Yan. Exact solutions of nonlinear dispersive K(m, n) model with variable coefficients. Applied Mathematics and Computation, 217(22):9474–9479, July 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311006175.

Yang:2011:ESN

[Yan11b] Chi-Ching Yang. Exponential synchronization of a new Lorenz-like attractor with uncertain parameters via single input. Applied Mathematics and Computation, 217 (14):6490–6497, March 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311000403.

Yang:2011:TMI

[Yan11c] Fan Yang. The truncation method for identifying an unknown source in the Poisson equation. Applied Mathematics and Computation, 217(22):9334–9339, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031100573X.

Yang:2011:RPC

[Yan11d] Shyi-Kae Yang. Robust pole-clustering of structured uncertain systems in union of disjointed circular regions. Applied Mathematics and Computation, 217(18):7488–7495, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311002438.

Yang:2011:COS

[Yan11e] Weifeng Yang. Composition operators from F (p, q, s) spaces to the n-th weighted-type spaces on the unit disc. Applied Mathematics and Computation, 218(4):1443–1448, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031100854X.

Yang:2012:NPC

[Yan12a] Liping Yang. A note on a paper “Convergence theorem for the common solution for a finite family of ϕ-strongly accretive operator equations”. Applied Mathematics and Computation, 218(21):10367–10369, July 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312004365.

Yang:2012:EEQ

[Yan12b] Min Yang. L^{2} error estimation of a quadratic finite volume element method for pseudoparabolic equations in three spatial dimensions. Applied Mathematics and Computation, 218(13):7270–7278, March 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312000227.

Yang:2012:SCB

[Yan12c] Ming-Chien Yang. Super connectivity of balanced hypercubes. Applied Mathematics and Computation, 219(3):970–975, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312007011.

Yang:2012:GWC

[Yan12d] Weifeng Yang. Generalized weighted composition operators from the F (pqs) space to the Bloch-type space. Applied Mathematics and Computation, 218(9):4967–4972, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311013087.

Yang:2012:EEP

[Yan12e] Wengui Yang. Existence of an exponential periodic attractor of periodic solutions for general BAM neural networks with time-varying delays and impulses. Applied Mathematics and Computation, 219(2):569–582, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312006431.

Yang:2012:PSS

[Yan12f] Zhilin Yang. Positive solutions for a system of nonlinear Hammerstein integral equations and applications. Applied Mathematics and Computation, 218(22):11138–11150, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312004870.

Yang:2013:OIC

[Yan13a] Chi-Ching Yang. One input control of exponential synchronization for a four-dimensional chaotic system. Applied Mathematics and Computation, 219(10):5152–5161, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312011563.

Yang:2013:HBT

[Yan13b] Gaoxiang Yang. Hopf bifurcation of traveling wave solutions of delayed Fisher–KPP equation. Applied Mathematics and Computation, 220(??):213–220, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313006905.

Yang:2013:DNI

[Yan13c] Wensheng Yang. Diffusion has no influence on the global asymptotical stability of the Lotka–Volterra prey-predator model incorporating a constant number of prey refuges. Applied Mathematics and Computation, 223(??):278–280, October 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313008692.

Yang:2013:DTA

[Yan13d] Xi Yang. Discrete-time accelerated block successive overrelaxation methods for timedependent Stokes equations. Applied Mathematics and Computation, 222(??): 519–538, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313008199.

Yang:2013:HOL

[Yan13e] Xiao Yang. A higher-order Levenberg– Marquardt method for nonlinear equations. Applied Mathematics and Computation, 219 (22):10682–10694, July 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313004487.

Yang:2014:ENM

[Yan14a]  Changqing Yang.  An efficient numeri-

cal method for solving Abel integral equation. Applied Mathematics and Computation, 227(??):656–661, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313012095.

Yang:2014:PSN

[Yan14b] Wengui Yang. Positive solutions for nonlinear semipositone fractional q-difference system with coupled integral boundary conditions. Applied Mathematics and Computation, 244(??):702–725, October 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314009965.

Yao:2010:SIM

[Yao10] Qingliu Yao. Successively iterative method of nonlinear Neumann boundary value problems. Applied Mathematics and Computation, 217(6):2301–2306, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310007824.

Yasar:2010:NTW

[Yaş10] Emrullah Yaşar. New travelling wave solutions to the Ostrovsky equation. Applied Mathematics and Computation, 216(11):3191–3194, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300310004649. See note

[Par10a].

Yang:2010:GPH

[YAW10] Ai-Li Yang, Jing An, and Yu-Jiang Wu. A generalized preconditioned HSS method for non-Hermitian positive definite linear systems. Applied Mathematics and Computation, 216(6):1715–1722, May 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309010790.

Yayenie:2011:NGF

[Yay11] Omer Yayenie. A note on generalized Fibonacci sequences. Applied Mathematics and Computation, 217(12):5603–5611, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310012403.

Yasar:2012:APP

[YB12] Meral Yasar and Durmus Bozkurt. Another proof of Pell identities by using the determinant of tridiagonal matrix. Applied Mathematics and Computation, 218 (10):6067–6071, January 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311014354.

Yang:2012:BMC

[YC12] ZhengHong Yang and BingFei Cui. On the Bezoutian matrix for Chebyshev polynomials. Applied Mathematics and Computation, 219(3):1183–1192, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031200728X.

Yang:2014:HAI

[YC14] Xinsong Yang and Jinde Cao. Hybrid adaptive and impulsive synchronization of uncertain complex networks with delays and general uncertain perturbations. Applied Mathematics and Computation, 227(??):480–493, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313011922.

Yin:2012:TAS

[YCC^{+}12] Yunqiang Yin, Shuenn-Ren Cheng, T. C. E. Cheng, Chin-Chia Wu, and Wen-Hsiang Wu. Two-agent single-machine scheduling with assignable due dates. Applied Mathematics and Computation, 219(4):1674–1685, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300312008004.

You:2013:NEA

[YCF13] Xiong You, Zhaoxia Chen, and Yonglei Fang. New explicit adapted Numerov methods for second-order oscillatory differential equations. Applied Mathematics and Computation, 219 (11):6241–6255, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312013021.

Chen:2012:NMI

[yCfW12] Ching yu Chen and Tsung fang Wu. The Nehari manifold for indefinite semilinear elliptic systems involving critical exponent. Applied Mathematics and Computation, 218 (22):10817–10828, July 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312004250.

Yu:2014:ABO

[YCJ^{+}14] Xin Yu, Rongjun Cheng, Huachen Jiang, Qian Zhang, and Chao Xu. The approximation for the boundary optimal control problem of Burgers–Fisher equation with constraints. Applied Mathematics and Computation, 243(??):889–898, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314008534.

Yi:2013:EMS

[YCK13] Dokkyun Yi, Booyong Choi, and Eun-Youn Kim. An effective method for solving nonlinear equations and its application. Applied Mathematics and Computation, 220 (??):568–579, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313006474.

Yang:2011:MAC

[YCS11] Shu-Ling Yang, Jian-Feng Cai, and Hai-Wei Sun. Multigrid algorithm from cyclic reduction for Markovian queueing networks. Applied Mathematics and Computation, 217 (16):6990–7000, April 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311001500.

Yang:2011:NSN

[YCTD11] Liu Yang, Yanping Chen, Xiaojiao Tong, and Chunlin Deng. A new smoothing Newton method for solving constrained nonlinear equations. Applied Mathematics and Computation, 217(24):9855–9863, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311006151.

Yang:2012:GNN

[YCXL12] Yongqing Yang, Jinde Cao, Xianyun Xu, and Jiao Liu. A generalized neural network for solving a class of minimax optimization problems with linear constraints. Applied Mathematics and Computation, 218 (14):7528–7537, March 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312000537.

Yang:2014:SRO

[YCZ14] Junqi Yang, Yantao Chen, and Fanglai Zhu. Singular reduced-order observer-based synchronization for uncertain chaotic systems subject to channel disturbance and chaos-based secure communication. Applied Mathematics and Computation, 229 (??):227–238, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313013283.

Yadav:2013:SHB

[YD13] Anupam Yadav and Kusum Deep. Shrinking hypersphere based trajectory of particles in PSO. Applied Mathematics and Computation, 220(??):246–267, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313006206.

Yuksel:2014:VTA

[YD14] Ismet Yüksel and Ülkü Dinlemez. Voronovskaja type approximation theorem for q-Szász-beta operators. Applied Mathematics and Computation, 235(??):555–559, May 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314003737.

Dai:2014:CRS

[yDpY14] Dong yan Dai and Yu ping Yuan. The classification and representation of single traveling wave solutions to the generalized Fornberg– Whitham equation. Applied Mathematics and Computation, 242(??):729–735, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314008856.

Yin:2014:EEW

[YDT^{+}14] Jiuli Yin, Shanyu Ding, Lixin Tian, Xinghua Fan, and Xiaoyan Deng. Existence of exotic waves for the nonlinear dispersive mKdV equation. Applied Mathematics and Computation, 229(??):499–504, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313013192.

Yu:2011:IMQ

[YDTW11] Bo Yu, Ning Dong, Qiong Tang, and Feng-Hua Wen. On iterative methods for the quadratic matrix equation with M-matrix. Applied Mathematics and Computation, 218 (7):3303–3310, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311011118.

Yang:2010:HSH

[YDXC10] Hong-Xiang Yang, Jun Du, Xi-Xiang Xu, and Jin-Ping Cui. Hamiltonian and super-Hamiltonian systems of a hierarchy of soliton equations. Applied Mathematics and Computation, 217(4):1497–1508, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309005918.

Yin:2011:ASA

[YDZ11] Hongxia Yin, Fei Ding, and Jianzhong Zhang. Active set algorithm for mathematical programs with linear complementarity constraints. Applied Mathematics and Computation, 217(21):8291–8302, July 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031100350X.

Yuan:2014:NMS

[YDZH14] Xiaofang Yuan, Xiangshan Dai, Jingyi Zhao, and Qian He. On a novel multi-swarm fruit fly optimization algorithm and its application. Applied Mathematics and Computation, 233(??):260–271, May 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314002379.

Yang:2011:EES

[YDZS11] Xingdong Yang, Zhiying Ding, Jiajing Zhang, and Suya Sun. On the eigenvalue estimation for solution to Lyapunov equation. Applied Mathematics and Computation, 217(16):6974–6980, April 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311001482.

Ye:2011:HGT

[YG11] Haiping Ye and Jianming Gao. Henry– Gronwall type retarded integral inequalities and their applications to fractional differential equations with delay. Applied Mathematics and Computation, 218(8):4152–4160, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300311012185.

Yin:2013:OMC

[YG13] Peng-Yeng Yin and Yi-Ming Guo. Optimization of multi-criteria Website structure based on enhanced tabu search and Web usage mining. Applied Mathematics and Computation, 219(24):11082–11095, August 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313005663.

Yang:2011:MCD

[YGL11] Fan Yang, HengZhen Guo, and XiaoXiao Li. The method of central difference for the inverse time-dependent heat source problem. Applied Mathematics and Computation, 218(7):3025–3034, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311011829.

Yuan:2014:ORS

[YGX14] Wenzhi Yuan, Lina Guo, and Genqi Xu. Optimal repair strategies for a two-unit deteriorating standby system. Applied Mathematics and Computation, 227(??):102–111, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313011375.

Yu:2014:DNL

[YGZW14] Guanglong Yu, Shu-Guang Guo, Rong Zhang, and Yarong Wu. The domination number and the least Q-eigenvalue. Applied Mathematics and Computation, 244(??):274–282, October 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314009205.

Yang:2010:RGF

[YH10] Zonghang Yang and Y. C. Hon. A rational generalization of Fan’s method and its application to generalized shallow water wave equations. Applied Mathematics and Computation, 216(7):1984–1995, June 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310002717.

You:2012:HAS

[YH12] Peng-Sheng You and Yi-Chih Hsieh. A heuristic approach to a single stage assembly problem with transportation allocation. Applied Mathematics and Computation, 218(22):11100–11111, July 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031200481X.

Yin:2013:LRI

[YH13] Feng Yin and Guang-Xin Huang. Left and right inverse eigenvalue problem of (RS)- symmetric matrices and its optimal approximation problem. Applied Mathematics and Computation, 219(17):9261–9269, May 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313003123.

Yan:2014:SBB

[YH14a] Hui-Yin Yan and Yu-Mei Huang. Splittingbased block preconditioning methods for block two-by-two matrices of real square blocks. Applied Mathematics and Computation, 243 (??):825–837, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314008765.

Yang:2014:ISC

[YH14b] Chang-Bo Yang and Ting-Zhu Huang. Improved stability criteria for a class of neural networks with variable delays and impulsive perturbations. Applied Mathematics and Computation, 243(??):923–935, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314008819.

Yang:2014:LSR

[YH14c] Shaojie Yang and Cuncai Hua. Lie symmetry reductions and exact solutions of a coupled KdV-Burgers equation. Applied Mathematics and Computation, 234(??):579–583, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314000812.

Yi:2014:WOM

[YH14d] Mingxu Yi and Jun Huang. Wavelet operational matrix method for solving fractional differential equations with variable coefficients. Applied Mathematics and Computation, 230(??):383–394, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313012186.

Yu:2014:IAR

[YH14e] Vincent F. Yu and Kuo-Jen Hu. An integrated approach for resource allocation in manufacturing plants. Applied Mathematics and Computation, 245(??):416–426, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314010819.

Yang:2011:SLH

[YHH11] Jaw-Yen Yang, Li-Hsin Hung, and Sheng-HsinX Hu. Semiclassical lattice hydrodynamics of rarefied channel flows. Applied Mathematics and Computation, 217 (11):5151–5159, February 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310008805.

Yuan:2008:HGA

[YHL08] Quan Yuan, Zhiqing He, and Huinan Leng. A hybrid genetic algorithm for a class of global optimization problems with box constraints. Applied Mathematics and Computation, 197(2):924–929, April 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300307008703. See erratum [?].

Yan:2014:CBT

[YHL14] Lanlan Yan, Xuli Han, and Jiongfeng Liang. Conversion between triangular Bézier patches and rectangular Bézier patches. Applied Mathematics and Computation, 232(??):469–478, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300314000678.

Yuan:2013:ATW

[YHS13] Wenjun Yuan, Yong Huang, and Yadong Shang. All traveling wave exact solutions of two nonlinear physical models. Applied Mathematics and Computation, 219 (11):6212–6223, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312012994.

Yi:2013:BPO

[YHW13] Mingxu Yi, Jun Huang, and Jinxia Wei. Block pulse operational matrix method for solving fractional partial differential equation. Applied Mathematics and Computation, 221(??):121–131, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313006292.

Yao:2011:SNI

[YHWZ11] Yuqin Yao, Yehui Huang, Yuan Wei, and Yunbo Zeng. Some new integrable systems constructed from the bi-Hamiltonian systems with pure differential Hamiltonian operators. Applied Mathematics and Computation, 218(2):583–591, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007867.

Yigit:2011:ODS

[Yig11] Faruk Yigit. One-dimensional solidification of pure materials with a time periodically oscillating temperature boundary condition. Applied Mathematics and Computation, 217(14):6541–6555, March 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031100049X.

Yin:2010:CPB

[Yin10] Jun-Feng Yin. A class of preconditioners based on matrix splitting for nonsymmetric linear systems. Applied Mathematics and Computation, 216(6):1694–1706, May 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309010819.

Yusong:2010:PEL

[YJ10] Cao Yusong and Xu Jin. Proportional and excess-of-loss reinsurance under investment gains. Applied Mathematics and Computation, 217(6):2546–2550, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031000826X.

Yao:2014:AEA

[YJ14a] Changhui Yao and Shanghui Jia. Asymptotic expansion analysis of nonconforming mixed finite element methods for timedependent Maxwell’s equations in Debye Medium. Applied Mathematics and Computation, 229(??):34–40, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313012927.

Yinying:2014:HBE

[YJ14b] Zhang Yinying and Jia Jianwen. Hopf bifurcation of an epidemic model with a nonlinear birth in population and vertical transmission. Applied Mathematics and Computation, 230(??):164–173, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031301374X.

Yuan:2014:NAE

[YJCC14] Xiaohui Yuan, Bin Ji, Zhijun Chen, and Zhihuan Chen. A novel approach for economic dispatch of hydrothermal system via gravitational search algorithm. Applied Mathematics and Computation, 247(??):535–546, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314012351.

Yang:2011:SDJ

[YK11a] Dar-Li Yang and Wen-Hung Kuo. Scheduling with deteriorating jobs and learning effects. Applied Mathematics and Computation, 218(5):2069–2073, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311009568.

Yunus:2011:BNC

[YK11b] Rossita M. Yunus and Shahjahan Khan. The bivariate noncentral chi-square distribution — a compound distribution approach. Applied Mathematics and Computation, 217(13):6237–6247, March 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310013299.

Yildirim:2012:CTC

[YK12] Isa Yildirim and Safeer Hussain Khan. Convergence theorems for common fixed points of asymptotically quasi-nonexpansive mappings in convex metric spaces. Applied Mathematics and Computation, 218(9):4860–4866, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031101294X.

Yadav:2013:ICR

[YK13a] Subhash Kumar Yadav and Cem Kadilar. Improved class of ratio and product estimators. Applied Mathematics and Computation, 219(22):10726–10731, July 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313004839.

Yoon:2013:GGO

[YK13b] Yourim Yoon and Yong-Hyuk Kim. Geometricity of genetic operators for real-coded representation. Applied Mathematics and Computation, 219(23):10915–10927, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313005377.

Yadav:2014:TPV

[YK14a] Subhash Kumar Yadav and Cem Kadilar. A two parameter variance estimator using auxiliary information. Applied Mathematics and Computation, 226(??):117–122, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313011016.

Yao:2014:EOM

[YK14b] Kai Yao and Hua Ke. Entropy operator for membership function of uncertain set. Applied Mathematics and Computation, 242(??):898–906, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314009254.

Yadani:2010:SVD

[YKI10] Kenichi Yadani, Koichi Kondo, and Masashi Iwasaki. A singular value decomposition algorithm based on solving hyperplane constrained nonlinear systems. Applied Mathematics and Computation, 216(3):779–790, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310001001.

Yang:2012:LTIb

[YKL12a] Xiaojing Yang, Yong-In Kim, and Kueiming Lo. Lyapunov-type inequality for a class of linear differential systems. Applied Mathematics and Computation, 219(4):1805–1812, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300312008120.

Yang:2012:LTIa

[YKL12b] Xiaojing Yang, Yong-In Kim, and Kueiming Lo. Lyapunov-type inequality for quasilinear systems. Applied Mathematics and Computation, 219(4):1670–1673, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312007990.

Yang:2014:LTI

[YKL14] Xiaojing Yang, Yong-In Kim, and Kueiming Lo. Lyapunov-type inequality for a class of even-order linear differential equations. Applied Mathematics and Computation, 245(??):145–151, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314010492.

Yang:2010:CAF

[YL10a] Ding-Gong Yang and Jin-Lin Liu. On a class of analytic functions with missing coefficients. Applied Mathematics and Computation, 215(9):3473–3481, January 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309009485.

Yang:2010:LTI

[YL10b] Xiaojing Yang and Kueiming Lo. Lyapunovtype inequality for a class of even-order differential equations. Applied Mathematics and Computation, 215(11):3884–3890, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300309010212.

Yu:2010:NIC

[YL10c] Fajun Yu and Li Li. Noncommutative integrable coupling system of TC equation hierarchy. Applied Mathematics and Computation, 215(11):3796–3801, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630030901008X.

Yan:2011:EBM

[YL11a] Lanlan Yan and Jiongfeng Liang. An extension of the Bézier model. Applied Mathematics and Computation, 218(6):2863–2879, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300311010563.

Yang:2011:CIP

[YL11b] Ding-Gong Yang and Jin-Lin Liu. Certain inequality properties of multivalent analytic functions involving the Dziok–Srivastava operator. Applied Mathematics and Computation, 218(3):924–928, October 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/

S0096300311002359. Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Yang:2011:MTR

[YL11c] Hu Yang and Xifu Liu. Mixed-type reverseorder laws of (AB)^{(1,2,3)} and (AB)^{(1,2,4)}. Applied Mathematics and Computation, 217 (24):10361–10367, August 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311007168.

Yang:2011:QRF

[YL11d] Li-Ying Yang and Guan-Ting Liu. Quasirational function solutions of an elliptic equation and its application to solve some nonlinear evolution equations. Applied Mathematics and Computation, 217(18):7377–7384, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311000130.

Yang:2011:PDN

[YL11e] Wensheng Yang and Xuepeng Li. Permanence of a discrete nonlinear N-species cooperation system with time delays and feedback controls. Applied Mathematics and Computation, 218(7):3581–3586, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311011477.

Yuan:2011:SPT

[YL11f] Shao-Mou Yuan and Zai-Ming Liu. Some properties of two subclasses of k-fold symmetric functions associated with Srivastava– Attiya operator. Applied Mathematics and Computation, 218(3):1136–1141, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311004486. Special Issue in

Honour of Hari M. Srivastava on his 70th birth anniversary.

Yu:2013:CLI

[YL13] Fajun Yu and Li Li. Continuous limits for an integrable coupling of the Kac–Van Moerbeke hierarchy. Applied Mathematics and Computation, 219(11):5772–5778, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312012787.

Yang:2014:GAS

[YL14a] Wensheng Yang and Xuepeng Li. Global asymptotical stability of a n + 1 dimensional Leslie–Gower predator-prey model. Applied Mathematics and Computation, 235 (??):377–382, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314003506.

Yi:2014:LCE

[YL14b] Son-Young Yi and Sunmi Lee. A locally conservative Eulerian–Lagrangian finite difference method for the forced KdV equation. Applied Mathematics and Computation, 230(??):276–289, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313014094.

Lan:2014:CAN

[yL14c] Heng you Lan. Convergence analysis of new over-relaxed proximal point algorithm frameworks with errors and applications to general A-monotone nonlinear inclusion forms. Applied Mathematics and Computation, 230(??):154–163, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313013040.

Yu:2014:NIC

[YL14d] Fajun Yu and Li Li. A nonlinear integrable couplings of C–KdV soliton hierarchy and its infinite conservation laws. Applied Mathematics and Computation, 233(??):413–420, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314002562.

Ye:2014:MPN

[YLAT14] H. Ye, F. Liu, V. Anh, and I. Turner. Maximum principle and numerical method for the multi-term time-space Riesz–Caputo fractional differential equations. Applied Mathematics and Computation, 227 (??):531–540, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031301182X.

Liu:2010:ACF

[yLbLjH10] Qing you Liu, Zhi bin Liu, and Nan jing Huang. Approximating the common fixed points of two sequences of uniformly quasi-Lipschitzian mappings in convex metric spaces. Applied Mathematics and Computation, 216(3):883–889, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/

S0096300310001207. See comments [Liu14d].

Yu:2010:MVR

[YLC10] Vincent F. Yu, Shih-Wei Lin, and Shuo-Yan Chou. The museum visitor routing problem. Applied Mathematics and Computation, 216(3):719–729, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310000792.

Yang:2011:UPD

[YLC11] Xitao Yang, Yaqin Liu, and Jing Chen. Uniform persistence for a discrete predator-prey system with delays. Applied Mathematics and Computation, 218(4):1174–1179, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031100717X.

Yuan:2012:MCP

[YLC12] Hua Yuan, Guannan Liu, and Guoqing Chen. On modeling the crowding and psychological effects in network-virus prevalence with nonlinear epidemic model. Applied Mathematics and Computation, 219(5):2387–2397, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300312007898.

Yu:2014:DPD

[YLC14] Jing-Rung Yu, Wen-Yi Lee, and Wan-Jiun Paul Chiou. Diversified portfolios with different entropy measures. Applied Mathematics and Computation, 241(??):47–63, August 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300314005293.

Yang:2014:PCM

[YLD14] Ximei Yang, Hongwei Liu, and Xiaoliang Dong. Polynomial convergence of Mehrotratype prediction-corrector infeasible-IPM for symmetric optimization based on the commutative class directions. Applied Mathematics and Computation, 230(??):616–628, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313014355.

Yong:2014:VSS

[YLGW14] Xuelin Yong, Xiaotong Liu, Jianwei Gao, and Hui Wang. Variable separation solutions to the coupled integrable dispersionless equations. Applied Mathematics and Computation, 235(??):358–363, May 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314003154.

Yu:2014:SCR

[YLJ14] Yanhua Yu, Huili Liu, and Seoung Dal Jung. Structure and characterization of ruled surfaces in Euclidean 3-space. Applied Mathematics and Computation, 233(??):252–259, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314002380.

Yun:2014:NPF

[YLK14] Ana Yun, Yibao Li, and Junseok Kim. A new phase-field model for a water-oil-surfactant system. Applied Mathematics and Computation, 229(??):422–432, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313013374.

Yang:2013:FMG

[YLLL13] Suhua Yang, Xingjun Luo, Fanchun Li, and Guangqing Long. A fast multiscale Galerkin method for the first kind ill-posed integral equations via iterated regularization. Applied Mathematics and Computation, 219(21):10527–10537, July 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031300444X.

Yao:2013:CEF

[YLMZ13] Haixiang Yao, Yongzeng Lai, Qinghua Ma, and Huabao Zheng. Characterization of efficient frontier for mean-variance model with a drawdown constraint. Applied Mathematics and Computation, 220(??):770–782, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313007595.

Yang:2013:IDB

[YLQ13] Weihua Yang, Huiqiu Lin, and Chengfu Qin. On the t/k-diagnosability of BC networks. Applied Mathematics and Computation, 225(??):366–371, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313010424.

Yao:2013:SAN

[YLS13] Changhui Yao, Jinfa Li, and Dongwei Shi. Superconvergence analysis of nonconforming mixed finite element methods for timedependent Maxwell’s equations in isotropic cold plasma media. Applied Mathematics and Computation, 219(12):6466–6472, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312013161.

Yang:2010:SGK

[YlTA10] Yang Yang, Zhao ling Tao, and Francis R. Austin. Solutions of the generalized KdV equation with time-dependent damping and dispersion. Applied Mathematics and Computation, 216(4):1029–1035, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309011163.

Yu:2012:CEA

[YLTB12] Q. Yu, F. Liu, I. Turner, and K. Burrage. A computationally effective alternating direction method for the space and time fractional Bloch–Torrey equation in 3-D. Applied Mathematics and Computation, 219 (8):4082–4095, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312010727.

Yu:2014:DML

[YLWH14] Jing Yu, Mingqiang Li, Yongli Wang, and Guoping He. A decomposition method for large-scale box constrained optimization. Applied Mathematics and Computation, 231(??):9–15, March 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314000083.

Yan:2013:SDD

[YLWU13] Shuling Yan, Xinzhe Lian, Weiming Wang, and R. K. Upadhyay. Spatiotemporal dynamics in a delayed diffusive predator model. Applied Mathematics and Computation, 224 (??):524–534, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313009132.

Yin:2013:RAS

[YLZ13] Jinyan Yin, Yangrong Li, and Huijun Zhao. Random attractors for stochastic semi-linear degenerate parabolic equations with additive noise in L^{q}. Applied Mathematics and Computation, 225(??):526–540, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010308.

Yang:2010:GCO

[YM10a] Weifeng Yang and Xiaoge Meng. Generalized composition operators from F (p, q, s) spaces to Bloch-type spaces. Applied Mathematics and Computation, 217(6):2513–2519, November 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300310008088.

Yu:2010:ESCa

[YM10b] Yaodong Yu and Hong-Cai Ma. Exact solutions of the combined KdV–Burgers equation with variable coefficients. Applied Mathematics and Computation, 215 (10):3534–3540, January 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309009667.

Yu:2010:ESD

[YM10c] Yaodong Yu and Hong-Cai Ma. Explicit solutions of (2 + 1)-dimensional nonlinear KP-BBM equation by using Exp-function method. Applied Mathematics and Computation, 217(4):1391–1397, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309005293.

Yang:2011:PPS

[YMG11] Bianxia Yang, Ruyun Ma, and Chenghua Gao. Positive periodic solutions of delayed differential equations. Applied Mathematics and Computation, 218(8):4538–4545, December 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311012811.

Yan:2010:REF

[YMZS10] Huaicheng Yan, Max Q.-H. Meng, Hao Zhang, and Hongbo Shi. Robust H_{∞} exponential filtering for uncertain stochastic time-delay systems with Markovian switching and nonlinearities. Applied Mathematics and Computation, 215(12):4358–4369, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310000056.

Yagoub:2011:VSV

[YNBV11] Hemza Yagoub, Truong Nguyen-Ba, and Rémi Vaillancourt. Variable-step variable-order 3-stage Hermite–Birkhoff–Obrechkoff DDE solver of order 4 to 14. Applied Mathematics and Computation, 217(24):10247–10255, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300311006783.

Yao:2010:NHI

[YNL10] Yonghong Yao, Muhammad A. Noor, and Yeong-Cheng Liou. A new hybrid iterative algorithm for variational inequalities. Applied Mathematics and Computation, 216(3):822–829, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310001116.

Yordzhev:2014:FSB

[Yor14] Krasimir Yordzhev. Factor-set of binary matrices and Fibonacci numbers. Applied Mathematics and Computation, 236(??):235–238, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300314004354.

Yoshida:2013:PIQ

[Yos13] Norio Yoshida. Picone identity for quasilinear elliptic equations with p(x)-laplacians and Sturmian comparison theory. Applied Mathematics and Computation, 225 (??):79–91, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313009880.

Ouyang:2014:OSP

[yOZ14] Zheng yong Ouyang and Shan Zheng. Orbital stability of peakons for a generalized CH equation. Applied Mathematics and Computation, 232(??):183–190, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031400099X.

Yan:2014:PSN

[YP14] Dongming Yan and Zhigang Pan. Positive solutions of nonlinear operator equations with sign-changing kernel and its applications. Applied Mathematics and Computation, 230(??):675–679, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313014227.

Yang:2011:SBM

[YPSW11] Yu-Fei Yang, Zhi-Feng Pang, Bao-Li Shi, and Zhi-Guo Wang. Split Bregman method for the modified lot model in image denoising. Applied Mathematics and Computation, 217(12):5392–5403, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310012099.

Yero:2010:NPD

[YRV10] Ismael G. Yero and Juan A. Rodríguez-Velázquez. A note on the partition dimension of Cartesian product graphs. Applied Mathematics and Computation, 217 (7):3571–3574, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310008921.

Yakhno:2010:SIV

[YS10a] V. G. Yakhno and A. Sevimlican. Solving an initial value problem in inhomogeneous electrically and magnetically anisotropic uniaxial media. Applied Mathematics and Computation, 215(11):3839–3850, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309010169.

Yetiskin:2010:OCP

[YS10b] Hakan Yetişkin and Murat Subaşı. On the optimal control problem for Schrödinger equation with complex potential. Applied Mathematics and Computation, 216(7):1896–1902, June 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S009630031000281X.

Yan:2011:NDA

[YS11a] Huiyun Yan and Yimin Shi. Nonlinear dynamical analysis and demonstration on Chinese business cycle. Applied Mathematics and Computation, 218(5):1547–1552, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311008538.

Yang:2011:RUM

[YS11b] XiaoLi Yang and ZhongKui Sun. Recovering unknown model parameters contained in a class of time-variant chaotic dynamical systems. Applied Mathematics and Computation, 217(17):7311–7317, May 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311001986.

Yang:2012:NSN

[YS12] Chao Yang and Jiachang Sun. New sixnode and seven-node hexagonal finite elements. Applied Mathematics and Computation, 218(24):11763–11774, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312005115.

Yuzbasi:2013:NSS

[YS13a] Suayip Yüzbasi and Niyazi Sahin. Numerical solutions of singularly perturbed onedimensional parabolic convection-diffusion problems by the Bessel collocation method. Applied Mathematics and Computation, 220 (??):305–315, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313006401.

Yuzbasi:2013:CMF

[YS13b] Suayip Yüzbasi and Mehmet Sezer. A collocation method to find solutions of linear complex differential equations in circular domains. Applied Mathematics and Computation, 219(22):10610–10626, July 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313003093.

Yuksel:2011:ULP

[YSB11] Saziye Yuksel, Tugba Han Simsekler, and Berrak Bilik. Upper and lower pre-strong na continuous multifunctions. Applied Mathematics and Computation, 218(3):1142–1146, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300311005741. Special

Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Yilmaz:2011:SID

[YSI11] M. M. Yilmaz, S. K. Serenbay, and E. Ibikli. On singular integrals depending on three parameters. Applied Mathematics and Computation, 218(3):1132–1135, October 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/

S0096300311004358. Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Yu:2011:MSA

[YSW11] Zeng Yu, Hu Shao, and Guodong Wang. Modified self-adaptive projection method for solving pseudomonotone variational inequalities. Applied Mathematics and Computation, 217(20):8052–8060, June 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311003407.

Yang:2012:CCN

[YSZ12] Shujie Yang, Bao Shi, and Qiang Zhang. Complete controllability of nonlinear stochastic impulsive functional systems. Applied Mathematics and Computation, 218(9):5543–5551, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300311013737.

Yagi:2011:DSP

[YT11] Atsushi Yagi and Ta Viet Ton. Dynamic of a stochastic predator-prey population. Applied Mathematics and Computation, 218 (7):3100–3109, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311010770.

Yazlik:2013:ICM

[YT13a] Yasin Yazlik and Necati Taskara. On the inverse of circulant matrix via generalized khoradam numbers. Applied Mathematics and Computation, 223(??):191–196, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031300828X.

Ye:2013:PSS

[YT13b] Yiwei Ye and Chun-Lei Tang. Periodic solutions for second-order discrete Hamiltonian system with a change of sign in potential. Applied Mathematics and Computation, 219 (12):6548–6555, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313000544.

Youssef:2013:MSO

[YT13c] I. K. Youssef and A. A. Taha. On the modified successive overrelaxation method. Applied Mathematics and Computation, 219 (9):4601–4613, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312010879.

Yang:2013:GSS

[YTC13] Jin Yang, Sanyi Tang, and Robert A. Cheke. Global stability and sliding bifurcations of a non-smooth Gause predator-prey system. Applied Mathematics and Computation, 224(??):9–20, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313008734.

Yang:2012:IPS

[YTCY12] Cheng-Hong Yang, Sheng-Wei Tsai, Li-Yeh Chuang, and Cheng-Huei Yang. An improved particle swarm optimization with doublebottom chaotic maps for numerical optimization. Applied Mathematics and Computation, 219(1):260–279, September 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031200598X.

Yagubov:2012:OCP

[YTS12a] Gabil Yagubov, Fatma Toyoglu, and Murat Subasi. An optimal control problem for two-dimensional Schrödinger equation. Applied Mathematics and Computation, 218 (11):6177–6187, February 5, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311015025.

Yurdakadim:2012:AFS

[YTS12b] T. Yurdakadim, E. Tas, and I. Sakaoglu. Approximation of functions by the sequence of integral operators. Applied Mathematics and Computation, 219(8):3863–3871, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312010077.

Yang:2014:VCH

[YTY^{+}14] Hao Yang, Keming Tang, Jianjiang Yu, Licai Zhu, Hua Xu, and Yingying Cao. Virtual coordinates in hyperbolic space based on Ricci flow for WLANs. Applied Mathematics and Computation, 243(??):537–545, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314006882.

Yu:2011:GKD

[Yu11a] Fajun Yu. The generalized Kupershmidt deformation for the fifth-order coupled KdV equations hierarchy. Applied Mathematics and Computation, 217(17):7138–7145, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311001470.

Yu:2011:ROL

[Yu11b] Yaoming Yu. The reverse order law for the generalized inverse A_{T,S}^{(2)} over a skew field. Applied Mathematics and Computation, 217 (24):10158–10165, August 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311006655.

Yu:2012:NFD

[Yu12a] Fajun Yu. New form (2 + 1)-dimensional integrable lattice hierarchies with two shift operators Ẽ and Ê. Applied Mathematics and Computation, 219(6):3002–3008, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312009265.

Yu:2012:SSK

[Yu12b] Shuimeng Yu. N-soliton solutions of the KP equation by Exp-function method. Applied Mathematics and Computation, 219 (8):3420–3424, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310013123.

Yu:2013:MRW

[Yu13a] Fajun Yu. Multi-rogue waves for a higherorder nonlinear Schrödinger equation in optical fibers. Applied Mathematics and Computation, 220(??):176–184, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031300564X.

Yu:2013:TDE

[Yu13b] Fajun Yu. Three-dimensional exact solutions of Gross–Pitaevskii equation with variable coefficients. Applied Mathematics and Computation, 219(11):5779–5786, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312012568.

Yu:2014:NAS

[Yu14] Haodong Yu. A new active-set strategy for NCP with degenerate solutions. Applied Mathematics and Computation, 236(??):118–128, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300314003270.

Yuan:2010:LSS

[Yua10] Yongxin Yuan. Least-squares solutions to the matrix equations AX = B and XC = D. Applied Mathematics and Computation, 216(10):3120–3125, July 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031000408X.

Yuan:2011:IMU

[Yua11] Yongxin Yuan. An iterative method for updating gyroscopic systems based on measured modal data. Applied Mathematics and Computation, 218(7):3753–3762, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311011866.

Yuan:2012:RAG

[Yua12] Li Yuan. Reliability analysis for a k-out-of-n:G system with redundant dependency and repairmen having multiple vacations. Applied Mathematics and Computation, 218 (24):11959–11969, August 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312005851.

Yuan:2013:ADM

[Yua13] Quan Yuan. Alternating direction method for structure-persevering finite element model updating problem. Applied Mathematics and Computation, 223(??):461–471, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313008990.

Yue:2014:NSF

[Yue14]  Chaohui Yue.  Neutral stochastic func-

tional differential equations with infinite delay and Poisson jumps in the C_{g} space. Applied Mathematics and Computation, 237 (??):595–604, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031400441X.

Yun:2010:IAA

[Yun10a] Beong In Yun. Intuitive approach to the approximate analytical solution for the Blasius problem. Applied Mathematics and Computation, 215(10):3489–3494, January 15, 2010.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309008753.

Yun:2010:TMF

[Yun10b] Beong In Yun. Transformation methods for finding multiple roots of nonlinear equations. Applied Mathematics and Computation, 217(2):599–606, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310006673.

Yun:2011:SNE

[Yun11a] Beong In Yun. Solving nonlinear equations by a new derivative free iterative method. Applied Mathematics and Computation, 217 (12):5768–5773, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310012579.

Yun:2011:CRP

[Yun11b] Jae Heon Yun. Comparison results of the preconditioned AOR methods for L-matrices. Applied Mathematics and Computation, 218 (7):3399–3413, December 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031101126X.

Younis:2014:TWS

[YuRI14] Muhammad Younis, Hamood ur Rehman, and Muzammal Iftikhar. Travelling wave solutions to some time-space nonlinear evolution equations. Applied Mathematics and Computation, 249(??):81–88, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314013381.

Yusufoglu:2010:EAS

[Yus10] Elçin Yusufoḡlu. An efficient algorithm for solving generalized pantograph equations with linear functional argument. Applied Mathematics and Computation, 217 (7):3591–3595, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310009525.

Yuzbasi:2013:NSF

[Yüz13] Suayip Yüzbasi. Numerical solutions of fractional Riccati type differential equations by means of the Bernstein polynomials. Applied Mathematics and Computation, 219 (11):6328–6343, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312012647.

Yuzbasi:2014:LAS

[Yüz14] Suayip Yüzbasi. Laguerre approach for solving pantograph-type Volterra integro-differential equations. Applied Mathematics and Computation, 232(??):1183–1199, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031400112X.

Yang:2011:EPS

[YW11a] Jingbao Yang and Zhongli Wei. On existence of positive solutions of Sturm–Liouville boundary value problems for a nonlinear singular differential system. Applied Mathematics and Computation, 217(13):6097–6104, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310013044.

Yang:2011:MTW

[YW11b] Shu-Hui Yang and Ji-Bo Wang. Minimizing total weighted completion time in a twomachine flow shop scheduling under simple linear deterioration. Applied Mathematics and Computation, 217(9):4819–4826, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310011495.

Ye:2011:EAL

[YWCH11] Yichao Ye, Lihong Wang, Zhaowei Chang, and Jingsong He. An efficient algorithm of logarithmic transformation to Hirota bilinear form of KdV-type bilinear equation. Applied Mathematics and Computation, 218 (5):2200–2209, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311009696.

Yuan:2014:AMS

[YWCH14] Wenjun Yuan, Yonghong Wu, Qiuhui Chen, and Yong Huang. All meromorphic solutions for two forms of odd order algebraic differential equations and its applications. Applied Mathematics and Computation, 240(??):240–251, August 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314006390.

Yuan:2013:SQM

[YWD13] Shi-Fang Yuan, Qing-Wen Wang, and Xue-Feng Duan. On solutions of the quaternion matrix equation AX = B and their applications in color image restoration. Applied Mathematics and Computation, 221 (??):10–20, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313006462.

Yang:2011:EAF

[YWK11] Dong-Yuh Yang, Kuo-Hsiung Wang, and Yu-Ting Kuo. Economic application in a finite capacity multi-channel queue with second optional channel. Applied Mathematics and Computation, 217(18):7412–7419, May 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311002219.

Yu:2013:FVE

[YWL13] Changhua Yu, Yanhe Wang, and Yonghai Li. The finite volume element methods for Poisson equation based on Adini’s element. Applied Mathematics and Computation, 219(10):5537–5549, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312012027.

Yang:2011:GMM

[YWLX11] Shuiping Yang, Jianxin Wu, Min Li, and Aiguo Xiao. Generations and mechanisms of multi-stripe chaotic attractors of fractional order dynamic system. Applied Mathematics and Computation, 217(13):6219–6229, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031001324X.

Yan:2014:MPS

[YWT14] Sheng-Hua Yan, Xing-Ping Wu, and Chun-Lei Tang. Multiple periodic solutions for second-order discrete Hamiltonian systems. Applied Mathematics and Computation, 234(??):142–149, May 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314002136.

Yang:2013:ACE

[YWW13] Jianwei Yang, Shu Wang, and Fuqiang Wang. Approximation of a compressible Euler– Poisson equations by a non-isentropic Euler– Maxwell equations. Applied Mathematics and Computation, 219(11):6142–6151, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312012556.

Yi:2014:DAH

[YWW^{+}14] Guosheng Yi, Jiang Wang, Xile Wei, Bin Deng, Huiyan Li, and Chunxiao Han. Dynamic analysis of Hodgkin’s three classes of neurons exposed to extremely low-frequency sinusoidal induced electric field. Applied Mathematics and Computation, 231 (??):100–110, March 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314000204.

Yang:2013:ICU

[YWY^{+}13] Wankou Yang, Zhenyu Wang, Jun Yin, Changyin Sun, and Karl Ricanek. Image classification using kernel collaborative representation with regularized least square. Applied Mathematics and Computation, 222(??):13–28, October 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313007741.

Yang:2012:HCO

[YWYY12] Yimin Yang, Yaonan Wang, Xiaofang Yuan, and Feng Yin. Hybrid chaos optimization algorithm with artificial emotion. Applied Mathematics and Computation, 218 (11):6585–6611, February 5, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311011970.

Yan:2010:UAI

[YWZ10] Xingjie Yan, Ying Wu, and Chengkui Zhong. Uniform attractors for impulsive reaction– diffusion equations. Applied Mathematics and Computation, 216(9):2534–2543, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310003498.

Wang:2014:MPS

[yWZhL14] Xing yuan Wang, Hao Zhang, and Xiao hui Lin. Module-phase synchronization in hyperchaotic complex Lorenz system after modified complex projection. Applied Mathematics and Computation, 232(??):91–96, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314000307.

Yuan:2011:DSR

[YX11] Li Yuan and Jian Xu. A deteriorating system with its repairman having multiple vacations. Applied Mathematics and Computation, 217(10):4980–4989, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031001177X.

Yuan:2012:MDS

[YX12a] Wen Zhi Yuan and Gen Qi Xu. Modelling of a deteriorating system with repair satisfying general distribution. Applied Mathematics and Computation, 218(11):6340–6350, February 5, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311014603.

Yusong:2012:OPR

[YX12b] Cao Yusong and Zeng Xianquan. Optimal proportional reinsurance and investment with minimum probability of ruin. Applied Mathematics and Computation, 218 (9):5433–5438, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311013610.

Yuan:2014:NIS

[YXO14] Xiaoqing Yuan, Yong-Hui Xia, and Donal O’Regan. Nonautonomous impulsive systems with unbounded nonlinear terms. Applied Mathematics and Computation, 245 (??):391–403, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314010674.

Yu:2010:PEC

[YXZ10] Guozhu Yu, Xiaoping Xie, and Xu Zhang. Parameter extension for combined hybrid finite element methods and application to plate bending problems. Applied Mathematics and Computation, 216(11):3265–3274, August 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310004807.

Yuan:2010:EAB

[YY10] Junli Yuan and Zuodong Yang. Existence and asymptotic behavior of radially symmetric ground states of quasi-linear singular elliptic equations. Applied Mathematics and Computation, 216(1):213–220, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310000494.

Yang:2012:SCV

[YY12] Lu-Xing Yang and Xiaofan Yang. The spread of computer viruses under the influence of removable storage devices. Applied Mathematics and Computation, 219(8):3914–3922, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300312010284.

Yang:2013:ASD

[YY13a] Qing Yang and Yirang Yuan. An approximation of semiconductor device by mixed finite element method and characteristics-mixed finite element method. Applied Mathematics and Computation, 225(??):407–424, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313010461.

Yuan:2013:PHG

[YY13b] Quan Yuan and Zhixin Yang. On the performance of a hybrid genetic algorithm in dynamic environments. Applied Mathematics and Computation, 219(24):11408–11413, August 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313006139.

Yang:2014:KDE

[YY14a] Fan Yang and Zhufeng Yue. Kernel density estimation of three-parameter Weibull distribution with neural network and genetic algorithm. Applied Mathematics and Computation, 247(??):803–814, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314012909.

Yu:2014:NRW

[YY14b] Fajun Yu and Zhenya Yan. New rogue waves and dark-bright soliton solutions for a coupled nonlinear Schrödinger equation with variable coefficients. Applied Mathematics and Computation, 233(??):351–358, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314002628.

Ye:2014:NOS

[YYJ14] Hailong Ye, Jingxue Yin, and Chunhua Jin. Non-oscillatory solutions for a nonlinear neutral delay differential equation. Applied Mathematics and Computation, 235(??):283–291, May 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314003531.

Yang:2013:ECV

[YYL^{+}13] Lu-Xing Yang, Xiaofan Yang, Jiming Liu, Qingyi Zhu, and Chenquan Gan. Epidemics of computer viruses: a complex-network approach. Applied Mathematics and Computation, 219(16):8705–8717, April 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313001537.

Yang:2014:FCS

[YYL14] Yun Yang, Yanhua Yu, and Huili Liu. Flat centroaffine surfaces with parallel second fundamental form in R^{4}. Applied Mathematics and Computation, 243(??):775–788, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300314009059.

Yuan:2012:IPC

[YYW12] Xiaofang Yuan, Yimin Yang, and Hui Wang. Improved parallel chaos optimization algorithm. Applied Mathematics and Computation, 219(8):3590–3599, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312009526.

Youjun:2010:MSQ

[YYY10] Wang Youjun, Zhang Yimin, and Shen Yaotian. Multiple solutions for quasilinear Schrödinger equations involving critical exponent. Applied Mathematics and Computation, 216(3):849–856, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310001153.

Yin:2010:PDS

[yYyL10] Xiao yan Yin and San yang Liu. On positive definite solutions of nonlinear matrix equation X^{s} − A^{⋆}X^{−t}A = Q. Applied Mathematics and Computation, 216(1):27–34, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300309010972.

Ye:2010:NES

[YZ10] Caier Ye and Weiguo Zhang. New explicit solutions for the Klein–Gordon equation with cubic nonlinearity. Applied Mathematics and Computation, 217(2):716–724, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310006831.

Ye:2011:NES

[YZ11] Caier Ye and Weiguo Zhang. New exact solutions for the nonlinear wave equation with fifth-order nonlinear term. Applied Mathematics and Computation, 218 (5):2155–2166, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311009659.

Yang:2012:NSK

[YZ12a] Yang Yang and Jihui Zhang. Nontrivial solutions on a kind of fourth-order Neumann boundary value problems. Applied Mathematics and Computation, 218 (13):7100–7108, March 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311015591.

Yuan:2012:CTP

[YZ12b] J. Y. Yuan and D. D. Zontini. Comparison theorems of preconditioned Gauss–Seidel methods for M-matrices. Applied Mathematics and Computation, 219(4):1947–1957, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312008302.

Yin:2013:AGA

[YZ13] Jun-Feng Yin and Yu Zhang. Alternative gradient algorithms for computing the nearest correlation matrix. Applied Mathematics and Computation, 219(14):7591–7599, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313000854.

Zhai:2014:GOF

[yZ14a] Jun yong Zhai. Global output feedback stabilization for a class of nonlinear time-varying delay systems. Applied Mathematics and Computation, 228(??):606–614, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313012952.

Yu:2014:ECL

[YZ14b] Xiang Yu and Xueqing Zhang. Enhanced comprehensive learning particle swarm optimization. Applied Mathematics and Computation, 242(??):265–276, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314007097.

Yu:2014:TWC

[YZ14c] Zhi-Xian Yu and Hai-Kun Zhao. Traveling waves for competitive Lotka–Volterra systems with spatial diffusions and spatio-temporal delays. Applied Mathematics and Computation, 242(??):669–678, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314009011.

Yuan:2014:NSU

[YZ14d] Yongxin Yuan and Kezheng Zuo. A no spillover updating method for undamped structural systems. Applied Mathematics and Computation, 238(??):13–20, July 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314005025.

Yang:2013:DDC

[YZC13] Xinbo Yang, Caidi Zhao, and Juan Cao. Dynamics of the discrete coupled nonlinear Schrödinger–Boussinesq equations. Applied Mathematics and Computation, 219 (16):8508–8524, April 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313000933.

You:2014:NFS

[YZC14] Xiong You, Yuanmei Zhou, and Xiaohao Cheng. A novel family of P -stable symmetric extended linear multistep methods for oscillators. Applied Mathematics and Computation, 249(??):597–610, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314014362.

Yang:2010:IMS

[YZD10] Minbo Yang, Fukun Zhao, and Yanheng Ding. Infinitely many stationary solutions of discrete vector nonlinear Schrödinger equation with symmetry. Applied Mathematics and Computation, 215(12):4230–4238, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309010923.

Zhang:2010:NESa

[yZhLjMzC10] Zai yun Zhang, Zhen hai Liu, Xiu jin Miao, and Yue zhong Chen. New exact solutions to the perturbed nonlinear Schrödinger’s equation with Kerr law nonlinearity. Applied Mathematics and Computation, 216(10):3064–3072, July 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310004467.

Yang:2012:HPM

[YZK12] Li Yang, Fuhai Zhao, and Jong Kyu Kim. Hybrid projection method for generalized mixed equilibrium problem and fixed point problem of infinite family of asymptotically quasiϕ-nonexpansive mappings in Banach spaces. Applied Mathematics and Computation, 218 (10):6072–6082, January 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311014378.

Yang:2010:PPM

[YZL10] Haijian Yang, Lijun Zhou, and Qingguo Li. A parallel projection method for a system of nonlinear variational inequalities. Applied Mathematics and Computation, 217 (5):1971–1975, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310007381.

Zhang:2012:SRE

[yZqLwW12] Ying yuan Zhang, Xi qiang Liu, and Gang wei Wang. Symmetry reductions and exact solutions of the (2 + 1)-dimensional Jaulent– Miodek equation. Applied Mathematics and Computation, 219(3):911–916, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312006881.

Yang:2010:ESP

[YZW^{+}10] Fengjian Yang, Chaolong Zhang, Dongqing Wu, Xinming Chen, and Jianfu Yang. Exponential stability and periodicity of impulsive cellular neural networks with time delays. Applied Mathematics and Computation, 217(7):3417–3424, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310009550.

Yu:2013:MTW

[YZX13] Zhi-Xian Yu, Ziheng Zhang, and Jing Xia. Monotone traveling waves of a population model with non-monotone terms. Applied Mathematics and Computation, 219 (15):7973–7981, April 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313000817.

Zuo:2010:IGC

[yZxGyM10] Xian yu Zuo, Tong xiang Gu, and Ze yao Mo. An improved GPBi-CG algorithm suitable for distributed parallel computing. Applied Mathematics and Computation, 215 (12):4101–4109, February 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309010339.

You:2010:SSF

[YZZ10] Fucai You, Jiao Zhang, and Jianbing Zhang. The N-soliton solutions of the fifth-order KdV equation under Bargmann constraint. Applied Mathematics and Computation, 217 (4):1321–1333, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309004433.

Yang:2012:DED

[YZZ12] Lu Yang, Zhenbing Zeng, and Weinian Zhang. Differential elimination with Dixon resultants. Applied Mathematics and Computation, 218(21):10679–10690, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312004353.

Yuan:2014:RIE

[YZZ^{+}14] Jinlong Yuan, Xi Zhu, Xu Zhang, Hongchao Yin, Enmin Feng, and Zhilong Xiu. Robust identification of enzymatic nonlinear dynamical systems for 1,3-propanediol transport mechanisms in microbial batch culture. Applied Mathematics and Computation, 232(??):150–163, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314000642.

Yang:2014:CGD

[YZZP14a] Bo Yang, Wei-Guo Zhang, Hai-Qiang Zhang, and Sheng-Bing Pei. Corrigendum to ‘Generalized Darboux transformation and rational soliton solutions for Chen–Lee–Liu equation’ [Applied Mathematics and Computation 242 (2014) 863–876]. Applied Mathematics and Computation, 245(??):591, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314011564. See [YZZP14b].

Yang:2014:GDT

[YZZP14b] Bo Yang, Wei-Guo Zhang, Hai-Qiang Zhang, and Sheng-Bing Pei. Generalized Darboux transformation and rational soliton solutions for Chen–Lee–Liu equation. Applied Mathematics and Computation, 242 (??):863–876, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314008844. See corrigendum

[YZZP14a].

Zayed:2011:ESN

[ZA11] E. M. E. Zayed and M. A. M. Abdelaziz. Exact solutions for the nonlinear Schrödinger equation with variable coefficients using the generalized extended tanh-function, the sinecosine and the exp-function methods. Applied Mathematics and Computation, 218 (5):2259–2268, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311009908.

Zeng:2014:SMC

[ZA14a] X. Zeng and M. Anitescu. Sequential Monte Carlo sampling in hidden Markov models of nonlinear dynamical systems. Applied Mathematics and Computation, 233(??):507–521, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300314002513.

Zenkour:2014:TSP

[ZA14b] Ashraf M. Zenkour and Ibrahim A. Abbas. Thermal shock problem for a fiber-reinforced anisotropic half-space placed in a magnetic field via G–N model. Applied Mathematics and Computation, 246(??):482–490, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314011618.

Zenkour:2014:RNT

[ZAA^{+}14] A. M. Zenkour, A. E. Abouelregal, K. A. Alnefaie, N. H. Abu-Hamdeh, and E. C. Aifantis. A refined nonlocal thermoelasticity theory for the vibration of nanobeams induced by ramp-type heating. Applied Mathematics and Computation, 248(??):169–183, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314013009.

Zhang:2011:SSM

[ZaFT11] Wenbing Zhang, Jian an Fang, and Yang Tang. Stochastic stability of Markovian jumping genetic regulatory networks with mixed time delays. Applied Mathematics and Computation, 217(17):7210–7225, May 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311001688.

Zahra:2011:SAB

[Zah11] W. K. Zahra. A smooth approximation based on exponential spline solutions for nonlinear fourth order two point boundary value problems. Applied Mathematics and Computation, 217(21):8447–8457, July 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311004097.

Zakeri:2010:ART

[Zak10] Gholam-Ali Zakeri. Analysis of ray tube area in geometrical shock dynamics. Applied Mathematics and Computation, 217 (2):830–836, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007095.

Zoraghi:2012:EFP

[ZANA12] Nima Zoraghi, Babak Abbasi, Seyed Taghi Akhavan Niaki, and Mehrzad Abdi. Estimating the four parameters of the Burr III distribution using a hybrid method of variable neighborhood search and iterated local search algorithms. Applied Mathematics and Computation, 218(19):9664–9675, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312002342.

Zayed:2011:NMS

[Zay11] Elsayed M. E. Zayed. A note on the modified simple equation method applied to Sharma– Tasso–Olver equation. Applied Mathematics and Computation, 218(7):3962–3964, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S009630031101191X.

Zhang:2014:BIA

[ZAY^{+}14] Xiaoge Zhang, Andrew Adamatzky, Hai Yang, Sankaran Mahadaven, Xin-She Yang, Qing Wang, and Yong Deng. A bioinspired algorithm for identification of critical components in the transportation networks. Applied Mathematics and Computation, 248(??):18–27, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314012806.

Zemouche:2011:OSM

[ZB11] Ali Zemouche and Mohamed Boutayeb. Observer synthesis method for Lipschitz nonlinear discrete-time systems with time-delay: an LMI approach. Applied Mathematics and Computation, 218(2):419–429, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311007624.

Zhang:2012:GIM

[ZB12] Kuize Zhang and Changjiang Bu. Group inverses of matrices over right Ore domains. Applied Mathematics and Computation, 218 (12):6942–6953, February 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031101558X.

Zahran:2013:INO

[ZB13] Yousef Hashem Zahran and Mohammed M. Babatin. Improved ninth order WENO scheme for hyperbolic conservation laws. Applied Mathematics and Computation, 219 (15):8198–8212, April 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313001410.

Zhang:2014:ESO

[ZB14] Zhi-Hai Zhang and Danyu Bai. An extended study on an open-shop scheduling problem using the minimisation of the sum of quadratic completion times. Applied Mathematics and Computation, 230(??):238–247, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313014173.

Zhao:2010:VNS

[ZBM10] Qiu Hong Zhao, Jack Brimberg, and Nenad Mladenović. A variable neighborhood search based algorithm for finite-horizon Markov Decision Processes. Applied Mathematics and Computation, 217(7):3480–3492, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310009653.

Zhu:2011:SCM

[ZC11] Song-Ping Zhu and Wen-Ting Chen. A spectral-collocation method for pricing perpetual American puts with stochastic volatility. Applied Mathematics and Computation, 217(22):9033–9040, July 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311004784.

Zhang:2012:NGC

[ZC12] Li Zhang and Xinlong Chen. A note on “A globally convergent BFGS method with nonmonotone line search for non-convex minimization”. Applied Mathematics and Computation, 219(2):764–766, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312006972.

Zhang:2013:SLS

[ZC13a] Xiaojiao Zhang and Baotong Cui. Synchronization of Lurie system based on contraction analysis. Applied Mathematics and Computation, 223(??):180–190, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313008308.

Zhang:2013:HSS

[ZC13b] Zhi-Yong Zhang and Temuer Chaolu. Homotopy series solutions of perturbed PDEs via approximate symmetry method. Applied Mathematics and Computation, 225 (??):92–101, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031301000X.

Zhao:2014:CQR

[ZC14a] Yulin Zhao and Yanyan Chen. The cyclicity of quadratic reversible systems with a center of genus one and non-Morsean point. Applied Mathematics and Computation, 231(??):268–275, March 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314000368.

Zhou:2014:SBC

[ZC14b] Liangqiang Zhou and Fangqi Chen. Subharmonic bifurcations and chaos for the traveling wave solutions of the compound KdV– Burgers equation with external and parametrical excitations. Applied Mathematics and Computation, 243(??):105–113, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314007292.

Zhou:2013:CAG

[ZCC13] Zhaojie Zhou, Fengxin Chen, and Huanzhen Chen. Convergence analysis for H ^{1}-galerkin mixed finite element approximation of one nonlinear integro-differential model. Applied Mathematics and Computation, 220 (??):783–791, September 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313007832.

Zhou:2010:STM

[ZCD10] Jianwei Zhou, Yanping Chen, and Yongquan Dai. Superconvergence of triangular mixed finite elements for optimal control problems with an integral constraint. Applied Mathematics and Computation, 217 (5):2057–2066, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007514.

Zhu:2013:BCC

[ZCF13] Weiyi Zhu, Xiaochun Chen, and Jishan Fan. A blow-up criterion for compressible nematic liquid crystal flows. Applied Mathematics and Computation, 219(14):7365–7368, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312013343.

Zhang:2010:PNC

[ZCG10] Yi Zhang, Li-Jin Chu, and Bo-Ling Guo. Positons, negatons and complexitons of the mKdV equation with non-uniformity terms. Applied Mathematics and Computation, 217 (4):1463–1469, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309005608.

Zhang:2011:SRB

[ZCL11] Dali Zhang, Elena Cherkaev, and Michael P. Lamoureux. Stieltjes representation of the 3D Bruggeman effective medium and Padé approximation. Applied Mathematics and Computation, 217(17):7092–7107, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311000361.

Zhang:2013:MAD

[ZCL13a] L. W. Zhang, Y. M. Cheng, and K. M. Liew. A mathematical analysis of DNA damage induced G2 phase transition. Applied Mathematics and Computation, 225 (??):765–774, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313010825.

Zhou:2013:PDS

[ZCL13b] Bin Zhou, Guang-Bin Cai, and James Lam. Positive definite solutions of the nonlinear matrix equation X + A^{H} X̄^{−1}A = I. Applied Mathematics and Computation, 219 (14):7377–7391, March 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313000593.

Zhang:2014:MMP

[ZCL14a] L. W. Zhang, Y. M. Cheng, and K. M. Liew. Mathematical modeling of p53 pulses in G2 phase with DNA damage. Applied Mathematics and Computation, 232(??):1000–1010, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300314001659.

Zhang:2014:MSR

[ZCL14b] L. W. Zhang, Y. M. Cheng, and K. M. Liew. A mathematical study of the robustness of G2/M regulatory network in response to DNA damage with parameters sensitivity. Applied Mathematics and Computation, 232(??):365–374, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300314001180.

Zecevic:2010:CDL

[ZCŠ10] A. I. Zečević, E. Cheng, and D. D. Šiljak. Control design for large-scale Lur’e systems with arbitrary information structure constraints. Applied Mathematics and Computation, 217(3):1277–1286, October 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310001438.

Zhou:2013:FTF

[ZCS13] Xiaojian Zhou, Xin Chen, and Yongzhong Song. Families of third and fourth order methods for multiple roots of nonlinear equations. Applied Mathematics and Computation, 219(11):6030–6038, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312013173.

Zhang:2011:SES

[ZCT11] Lina Zhang, Aiyong Chen, and Jiade Tang. Special exact soliton solutions for the K(2, 2) equation with non-zero constant pedestal. Applied Mathematics and Computation, 218 (8):4448–4457, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311012550.

Zhanlav:2014:TSA

[ZCU14] T. Zhanlav, O. Chuluunbaatar, and V. Ulziibayar. Two-sided approximation for some Newton’s type methods. Applied Mathematics and Computation, 236(??):239–246, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314004305.

Zhou:2014:ASC

[ZCY14] Liqun Zhou, Xiubo Chen, and Yixian Yang. Asymptotic stability of cellular neural networks with multiple proportional delays. Applied Mathematics and Computation, 229 (??):457–466, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313013441.

Zhong:2013:FTF

[ZCZ^{+}13a] Qishui Zhong, Jun Cheng, Yuqing Zhao, Jianhua Ma, and Bo Huang. Finite-time H_{∞} filtering for a class of discrete-time Markovian jump systems with switching transition probabilities subject to average dwell time switching. Applied Mathematics and Computation, 225(??):278–294, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010230.

Zhou:2013:SIN

[ZCZ13b] Duanmei Zhou, Guoliang Chen, and Xiangyun Zhang. Some inequalities for the nonlinear matrix equation X^{s} + A^{∗}X^{−t}A = Q: Trace, determinant and eigenvalue. Applied Mathematics and Computation, 224(??):21–28, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300313008680.

Zhang:2014:PEE

[ZCZ14] Bei Zhang, Shaochun Chen, and Jikun Zhao. A posteriori error estimation based on conservative flux reconstruction for nonconforming finite element approximations to a singularly perturbed reaction-diffusion problem on anisotropic meshes. Applied Mathematics and Computation, 232(??):1062–1075, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031400191X.

Zhu:2014:MPI

[ZCZP14] Huihui Zhu, Jianlong Chen, Xiaoxiang Zhang, and Pedro Patrício. The Moore–Penrose inverse of 2 × 2 matrices over a certain *-regular ring. Applied Mathematics and Computation, 246(??):263–267, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314011205.

Zhang:2012:HBD

[ZCZS12] Guodong Zhang, Boshan Chen, Lulu Zhu, and Yi Shen. Hopf bifurcation for a differentialalgebraic biological economic system with time delay. Applied Mathematics and Computation, 218(15):7717–7726, April 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312000586.

Zaghdani:2010:CLF

[ZD10] A. Zaghdani and C. Daveau. On the coupling of LDG–FEM and BEM methods for the three dimensional magnetostatic problem. Applied Mathematics and Computation, 217(5):1791–1810, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007460.

Zheng:2011:MLM

[ZD11a] Baojing Zheng and Baodong Dai. A meshless local moving Kriging method for twodimensional solids. Applied Mathematics and Computation, 218(2):563–573, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311007818.

Zhukovsky:2011:ENS

[ZD11b] K. V. Zhukovsky and G. Dattoli. Evolution of non-spreading Airy wavepackets in time dependent linear potentials. Applied Mathematics and Computation, 217 (20):7966–7974, June 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311003110.

Zheng:2010:NHS

[ZDB10] Song Zheng, Gaogao Dong, and Qinsheng Bi. A new hyperchaotic system and its synchronization. Applied Mathematics and Computation, 215(9):3192–3200, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309008893.

Zhou:2012:CDD

[ZDC12] Chuan Zhou, Mingli Du, and Qingwei Chen. Co-design of dynamic scheduling and Hinfinity control for networked control systems. Applied Mathematics and Computation, 218(21):10767–10775, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312004341.

Zhao:2010:ENE

[ZDH10] Zhanhui Zhao, Zhengde Dai, and Song Han. The EHTA for nonlinear evolution equations. Applied Mathematics and Computation, 217 (8):4306–4310, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310010751.

Zhu:2014:MMM

[ZDH14] Bohui Zhu, Yongsheng Ding, and Kuangrong Hao. Multiclass maximum margin clustering via immune evolutionary algorithm for automatic diagnosis of electrocardiogram arrhythmias. Applied Mathematics and Computation, 227(??):428–436, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313011958.

Zhao:2010:ETW

[ZDW10] Zhanhui Zhao, Zhengde Dai, and Chuanjian Wang. Extend three-wave method for the (1+2)-dimensional Ito equation. Applied Mathematics and Computation, 217 (5):2295–2300, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007447.

Zhao:2011:NEL

[ZDW11] Weidong Zhao, Huanhe Dong, and Hui Wang. New extended Lie algebra and the generalized integrable Liouville hierarchy. Applied Mathematics and Computation, 218 (2):643–650, September 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311008009.

Zhang:2010:GGC

[ZDZ10] Jianhua Zhang, Hua Dai, and Jing Zhao. Generalized global conjugate gradient squared algorithm. Applied Mathematics and Computation, 216(12):3694–3706, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310005709.

Zeidan:2011:NRC

[Zei11] D. Zeidan. Numerical resolution for a compressible two-phase flow model based on the theory of thermodynamically compatible systems. Applied Mathematics and Computation, 217(11):5023–5040, February 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310007988.

Zenkour:2012:DBA

[Zen12] Ashraf M. Zenkour. Dynamical bending analysis of functionally graded infinite cylinder with rigid core. Applied Mathematics and Computation, 218(17):8997–9006, May 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312002020.

Zhou:2010:EDE

[ZF10] P. Zhou and L. W. Fan. An extension to data envelopment analysis with preference structure for estimating overall inefficiency. Applied Mathematics and Computation, 216(3):812–818, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310001098.

Zhang:2011:PEA

[ZF11] Tie Zhang and Nan Feng. A posteriori error analysis of the discontinuous finite element methods for first order hyperbolic problems. Applied Mathematics and Computation, 218(5):1752–1764, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311009040.

Zhao:2013:ROS

[ZF13a] Yun-Bin Zhao and Masao Fukushima. Rankone solutions for homogeneous linear matrix equations over the positive semidefinite cone. Applied Mathematics and Computation, 219(10):5569–5583, January 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312011769.

Zhou:2013:SDA

[ZF13b] Guangming Zhou and Chunsheng Feng. The steepest descent algorithm without line search for p-Laplacian. Applied Mathematics and Computation, 224(??):36–45, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313008576.

Zuo:2013:NMR

[ZFCW13] Zhiqiang Zuo, Youhua Fu, Yinping Chen, and Yijing Wang. A new method of reachable set estimation for time delay systems with polytopic uncertainties. Applied Mathematics and Computation, 221(??):639–647, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313007480.

Zhai:2014:NMD

[ZFH14] Shuying Zhai, Xinlong Feng, and Yinnian He. A new method to deduce high-order compact difference schemes for two-dimensional Poisson equation. Applied Mathematics and Computation, 230(??):9–26, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313013866.

Zhang:2012:SPS

[ZG12] Xuemei Zhang and Weigao Ge. Symmetric positive solutions of boundary value problems with integral boundary conditions. Applied Mathematics and Computation, 219 (8):3553–3564, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312009368.

Zhang:2010:DDT

[ZGH10] Ping Zhang, Shangjiang Guo, and Yigang He. Dynamics of a delayed two-coupled oscillator with excitatory-to-excitatory connection. Applied Mathematics and Computation, 216(2):631–646, March 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310001219.

Zhang:2012:IDS

[ZgJhD12] Yi Zhang, Li gang Jin, and Kang hui Dong. Integrable discretizations and soliton solution of KdV and mKdV equations. Applied Mathematics and Computation, 218(18):9555–9562, May 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312002834.

Zhang:2012:GWS

[ZgJlK12] Yi Zhang, Li gang Jin, and Ying li Kang. Generalized Wronskian solutions for the (3 + 1)-dimensional Jimbo–Miwa equation. Applied Mathematics and Computation, 219 (5):2601–2610, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312008879.

Zhang:2014:TDH

[ZGW14a] Yufeng Zhang, Jing Gao, and Guangming Wang. Two (2 + 1)-dimensional hierarchies of evolution equations and their Hamiltonian structures. Applied Mathematics and Computation, 243(??):601–606, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314008480.

Zhou:2014:IOC

[ZGW14b] Xiaoliang Zhou, Zhonghai Guo, and Wu-Sheng Wang. Interval oscillation criteria for nonlinear impulsive delay differential equations with damping term. Applied Mathematics and Computation, 249(??):32–44, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300314014428.

Zhu:2010:PPS

[ZGY^{+}10] Shun-Hui Zhu, Yi-Tian Gao, Xin Yu, Zhi-Yuan Sun, Xiao-Ling Gai, and De-Xin Meng. Painlevé property, soliton-like solutions and complexitons for a coupled variable-coefficient modified Korteweg–de Vries system in a twolayer fluid model. Applied Mathematics and Computation, 217(1):295–307, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031000620X.

Zhou:2014:CPD

[ZGY14] Xiaojun Zhou, David Yang Gao, and Chunhua Yang. Canonical primal-dual algorithm for solving fourth-order polynomial minimization problems. Applied Mathematics and Computation, 227(??):246–255, January 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313011739.

Zhang:2014:RPD

[ZGZ14] Ju-Li Zhang, Chuan-Qing Gu, and Ke Zhang. A relaxed positive-definite and skew-Hermitian splitting preconditioner for saddle point problems. Applied Mathematics and Computation, 249(??):468–479, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314014349.

Zhao:2010:PSN

[ZGZG10] Junfang Zhao, Fengjie Geng, Jianfeng Zhao, and Weigao Ge. Positive solutions to a new kind Sturm–Liouville-like four-point boundary value problem. Applied Mathematics and Computation, 217(2):811–819, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310007058.

Zhang:2010:BGS

[ZH10a] Tong Zhang and Yinnian He. Blow-up and global solutions for a class of nonlinear parabolic equations with different kinds of boundary conditions. Applied Mathematics and Computation, 217(2):801–810, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310007046.

Zhao:2010:FTN

[ZH10b] Na Zhao and Zheng-Hai Huang. Finite termination of a Newton-type algorithm for a class of affine variational inequality problems. Applied Mathematics and Computation, 217(7):3368–3378, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310009331.

Zemanek:2012:FEO

[ZH12] Petr Zemánek and Petr Hasil. Friedrichs extension of operators defined by even order Sturm–Liouville equations on time scales. Applied Mathematics and Computation, 218 (22):10829–10842, July 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312004262.

Zhu:2013:CBB

[ZH13] Yuanpeng Zhu and Xuli Han. A class of αβγ-Bernstein–Bézier basis functions over triangular domain. Applied Mathematics and Computation, 220(??):446–454, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313006826.

Zhong:2014:AST

[ZH14] Xian-Ci Zhong and Qiong-Ao Huang. Approximate solution of three-point boundary value problems for second-order ordinary differential equations with variable coefficients. Applied Mathematics and Computation, 247(??):18–29, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314011862.

Zhang:2010:NEC

[Zha10a] Huiqun Zhang. A note on exact complex travelling wave solutions for (2 + 1)- dimensional B-type Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 216(9):2771–2777, July 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310003607.

Zhang:2010:EFM

[Zha10b] Sheng Zhang. Exp-function method for Riccati equation and new exact solutions with two arbitrary functions of (2 + 1)- dimensional Konopelchenko–Dubrovsky equations. Applied Mathematics and Computation, 216(5):1546–1552, May 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031000247X.

Zhang:2010:MBC

[Zha10c] Xuecang Zhang. Mapped barycentric Chebyshev differentiation matrix method for the solution of regular Sturm–Liouville problems. Applied Mathematics and Computation, 217(5):2266–2276, November 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310007721.

Zhao:2010:ISW

[Zha10d] Hong Zhao. Interactions of solitary waves under the conditions of the (3 + 1)-dimensional Kadomtsev–Petviashvilli equation. Applied Mathematics and Computation, 215 (9):3383–3389, January 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309009369.

Zhao:2010:NAP

[Zha10e] Hong Zhao. A new algebraic procedure to construct exact solutions of nonlinear differential– difference equations. Applied Mathematics and Computation, 216(4):1219–1225, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310001797.

Zhao:2010:IHS

[Zha10f] Junjian Zhao. Interpolatory Hermite splines on rectangular domains. Applied Mathematics and Computation, 216(10):2799–2813, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310003954.

Zhang:2011:CRS

[Zha11a] Guizhou Zhang. Convergence rate of the solution toward boundary layer solution for initial-boundary value problem of the 2-D viscous conservation laws. Applied Mathematics and Computation, 217(19):7799–7805, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311003109.

Zhang:2011:NIS

[Zha11b] Jian-Jun Zhang. A note on the iterative solutions of general coupled matrix equation. Applied Mathematics and Computation, 217(22):9380–9386, July 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311005832.

Zhang:2011:SFV

[Zha11c] Tong Zhang. The semidiscrete finite volume element method for nonlinear convectiondiffusion problem. Applied Mathematics and Computation, 217(19):7546–7556, June 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311001330.

Zhang:2012:GEB

[Zha12a] Jia-Fang Zhang. Global existence of bifurcated periodic solutions in a commensalism model with delays. Applied Mathematics and Computation, 218(23):11688–11699, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312005711.

Zhao:2012:NLE

[Zha12b]  Hong Zhao.  New localized excitations

and their interaction behaviors in (3 + 1)- dimensional system. Applied Mathematics and Computation, 219(1):345–350, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312006224.

Zhao:2012:PSS

[Zha12c] Zengqin Zhao. Positive solutions of semipositone Hammerstein integral equations and applications. Applied Mathematics and Computation, 219(5):2789–2797, November 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312009083.

Zhang:2013:CLD

[Zha13a] Li-Hua Zhang. Conservation laws of the (2 + 1)-dimensional KP equation and Burgers equation with variable coefficients and cross terms. Applied Mathematics and Computation, 219(9):4865–4879, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/

S009630031201079X. See comments [Ibr14].

Zhang:2013:UBC

[Zha13b] Yanhong Zhang. Uniform boundedness and convergence of global solutions to a stronglycoupled parabolic system with three competitive species. Applied Mathematics and Computation, 221(??):720–726, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313007005.

Zhao:2013:BSN

[Zha13c] Dianli Zhao. The boundedness and stability of neutral delay differential equations with damped stochastic perturbations via fixed point theory. Applied Mathematics and Computation, 219(12):6792–6803, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313000453.

Zhao:2013:CFD

[Zha13d] Jichao Zhao. Compact finite difference methods for high order integro-differential equations. Applied Mathematics and Computation, 221(??):66–78, September 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313006450.

Zhao:2013:NIC

[Zha13e] Yun-Bin Zhao. New and improved conditions for uniqueness of sparsest solutions of underdetermined linear systems. Applied Mathematics and Computation, 224(??):58–73, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313008540.

Zhang:2014:PSF

[Zha14a] Hui Zhang. Properties of solution of fractional backward stochastic differential equation. Applied Mathematics and Computation, 228(??):446–453, February 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313012502.

Zhang:2014:RAS

[Zha14b] Maozhu Zhang. Regular approximation of singular Sturm–Liouville problems with transmission conditions. Applied Mathematics and Computation, 247(??):511–520, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314012636.

Zhang:2014:CGSa

[Zha14c] Xiang Zhang. Characterization for the general solution to a system of matrix equations with quadruple variables. Applied Mathematics and Computation, 226(??):274–287, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313010886.

Zhang:2014:MSC

[Zha14d] Xu Zhang. Multibump solutions of a class of second-order discrete Hamiltonian systems. Applied Mathematics and Computation, 236(??):129–149, June 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314003257.

Zhao:2014:RRA

[Zha14e] Wen-Qiang Zhao. Regularity of random attractors for a degenerate parabolic equations driven by additive noises. Applied Mathematics and Computation, 239(??):358–374, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314006535.

Zhou:2011:SKT

[ZHC11] Shaobo Zhou, Shigeng Hu, and Liqun Cen. Stochastic Kolmogorov-type system with infinite delay. Applied Mathematics and Computation, 218(2):207–218, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007879.

Zhao:2014:DDA

[ZHC^{+}14] Chuanli Zhao, Chou-Jung Hsu, Shuenn-Ren Cheng, Yunqiang Yin, and Chin-Chia Wu. Due date assignment and single machine scheduling with deteriorating jobs to minimize the weighted number of tardy jobs. Applied Mathematics and Computation, 248 (??):503–510, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314013290.

Zheng:2011:TWS

[Zhe11] Bin Zheng. Travelling wave solutions of two nonlinear evolution equations by using the \frac{G′}{G} - expansion method. Applied Mathematics and Computation, 217(12):5743–5753, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310012543.

Zheng:2012:AMF

[Zhe12]  Song Zheng.  Adaptive modified func-

tion projective synchronization of unknown chaotic systems with different order. Applied Mathematics and Computation, 218 (10):5891–5899, January 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311013646.

Zheng:2013:IMA

[Zhe13] Liang Zheng. An improved monkey algorithm with dynamic adaptation. Applied Mathematics and Computation, 222(??):645–657, October 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313008175.

Zheng:2014:CGG

[Zhe14] Xiao-Jun Zheng. A connectivity graph generation approach for Manhattan path calculation in detailed facility layout. Applied Mathematics and Computation, 237(??):238–251, June 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314004627.

Zheng:2014:DMP

[ZHGP14] Hong-Chan Zheng, Shu-Pei Huang, Feng Guo, and Guo-Hua Peng. Designing multiparameter curve subdivision schemes with high continuity. Applied Mathematics and Computation, 243(??):197–208, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314008133.

Zheng:2014:NBN

[ZhGQ14] Quan Zheng, Xiu hui Guo, and Feng Qin. A natural BEM with non-uniform grids on the elliptic boundary and its coupling method. Applied Mathematics and Computation, 228(??):552–562, February 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313012861.

Zhang:2013:SKP

[ZHH^{+}13] Xiaoge Zhang, Shiyan Huang, Yong Hu, Yajuan Zhang, Sankaran Mahadevan, and Yong Deng. Solving 0–1 knapsack problems based on amoeboid organism algorithm. Applied Mathematics and Computation, 219(19):9959–9970, June 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313004232.

Zhi:2011:BQB

[Zhi11] Yuanhong Zhi. The boundary quenching behavior of a semilinear parabolic equation. Applied Mathematics and Computation, 218(2):233–238, September 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311006576.

Zhigljavsky:2012:CSC

[Zhi12] Anatoly Zhigljavsky. Computing sums of conditionally convergent and divergent series using the concept of grossone. Applied Mathematics and Computation, 218 (16):8064–8076, April 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311015086.

Zhou:2011:DSS

[Zho11a] Jian Zhou. The differential systems with the same reflecting function. Applied Mathematics and Computation, 218(7):3144–3148, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300311010873.

Zhou:2011:ESL

[Zho11b] Yunkai Zhou. On the eigenvalues of specially low-rank perturbed matrices. Applied Mathematics and Computation, 217 (24):10267–10270, August 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311006953.

Zhou:2011:RBQ

[Zho11c] Zhengxin Zhou. On the relationship between quadratic polynomial differential system and the Bernoulli equation. Applied Mathematics and Computation, 217(21):8716–8721, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311004863.

Zhou:2012:CFD

[Zho12a] Shuming Zhou. The conditional fault diagnosability of (nk)-star graphs. Applied Mathematics and Computation, 218(19):9742–9749, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300312002536.

Zhou:2012:NNI

[Zho12b] Weijun Zhou. A note on “A new iteration method for the matrix equation AX = B”. Applied Mathematics and Computation, 218(21):10639–10641, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312004249.

Zhou:2012:RPS

[Zho12c] Zhengxin Zhou. Research on the properties of some planar polynomial differential equations. Applied Mathematics and Computation, 218(9):5671–5681, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311014056.

Zhong:2013:NNT

[Zho13] Xian-Ci Zhong. A new Nyström-type method for Fredholm integral equations of the second kind. Applied Mathematics and Computation, 219(17):8842–8847, May 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313002890.

Zhang:2013:IHD

[ZHT13] Yufeng Zhang, Zhong Han, and Hon-Wah Tam. An integrable hierarchy and Darboux transformations, bilinear Bäcklund transformations of a reduced equation. Applied Mathematics and Computation, 219 (11):5837–5848, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312012532.

Zhu:2010:CSN

[Zhu10a] Congxu Zhu. Control and synchronize a novel hyperchaotic system. Applied Mathematics and Computation, 216(1):276–284, March 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310000664.

Zhu:2010:CHH

[Zhu10b] Congxu Zhu. Controlling hyperchaos in hyperchaotic Lorenz system using feedback controllers. Applied Mathematics and Computation, 216(10):3126–3132, July 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310004443.

Zhu:2010:GFJ

[Zhu10c] Ling Zhu. A general form of Jordan-type double inequality for the generalized and normalized Bessel functions. Applied Mathematics and Computation, 215(11):3802–3810, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300309010091.

Zhu:2010:CDD

[Zhu10d] Shaohong Zhu. Conservative domain decomposition procedure with unconditional stability and second-order accuracy. Applied Mathematics and Computation, 216 (11):3275–3282, August 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310004820.

Zhu:2010:WCO

[Zhu10e] Xiangling Zhu. Weighted composition operators from area Nevanlinna spaces into Bloch spaces. Applied Mathematics and Computation, 215(12):4340–4346, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310000020.

Zhu:2011:SFD

[Zhu11a] Bo Zhu. The space fractional diffusion equation with Feller’s operator. Applied Mathematics and Computation, 217(24):9971–9977, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300311006357.

Zhu:2011:ERT

[Zhu11b] Ling Zhu. Extension of Redheffer type inequalities to modified Bessel functions. Applied Mathematics and Computation, 217 (21):8504–8506, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311004218.

Zhuang:2011:LSE

[Zhu11c] Qingqu Zhuang. A Legendre spectral-element method for the one-dimensional fourth-order equations. Applied Mathematics and Computation, 218(7):3587–3595, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311011489.

Zhu:2012:SIH

[Zhu12a] Ling Zhu. Sharp inequalities for hyperbolic functions. Applied Mathematics and Computation, 219(1):416–420, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312006340.

Zhu:2012:GLH

[Zhu12b] Mu-Zheng Zhu. A generalization of the local Hermitian and skew-Hermitian splitting iteration methods for the non-Hermitian saddle point problems. Applied Mathematics and Computation, 218(17):8816–8824, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312001671.

Zou:2010:BTU

[ZHW10] Shaofen Zou, Lihong Huang, and Yaonan Wang. Bifurcation of a three-unit neural network. Applied Mathematics and Computation, 217(2):904–917, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007198.

Zheng:2011:PSA

[ZHW11] Fangying Zheng, Congying Han, and Yongli Wang. Parallel SSLE algorithm for large scale constrained optimization. Applied Mathematics and Computation, 217(12): 5377–5384, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310012051.

Zhang:2011:ESI

[ZHX11] Ya-Xuan Zhang, Zhong-Jie Han, and Gen-Qi Xu. Expansion of solution of an inverted pendulum system with time delay. Applied Mathematics and Computation, 217 (14):6476–6489, March 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311000397.

Zhang:2014:DDS

[ZHXM14] Xiao-Bing Zhang, Hai-Feng Huo, Hong Xiang, and Xin-You Meng. Dynamics of the deterministic and stochastic SIQS epidemic model with non-linear incidence. Applied Mathematics and Computation, 243 (??):546–558, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314008406.

Zhang:2011:KWI

[ZHY11] Defei Zhang, Ping He, and Xinsong Yang. Kunita–Watanabe inequalities and Tanaka formula for multi-dimensional G-Brownian motion. Applied Mathematics and Computation, 217(12):5774–5783, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310012610.

Zhang:2013:QBS

[ZHY13] Haixiang Zhang, Xuli Han, and Xuehua Yang. Quintic B-spline collocation method for fourth order partial integro-differential equations with a weakly singular kernel. Applied Mathematics and Computation, 219 (12):6565–6575, February 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313000477.

Zhao:2012:NAA

[ZHZ12] Lei Zhao, Dingjiang Huang, and Shuigeng Zhou. A new algorithm for automatic computation of solitary wave solutions to nonlinear partial differential equations based on the Exp-function method. Applied Mathematics and Computation, 219(4):1890–1896, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300312008247.

Zilinskas:2012:SHT

[Zil12] Antanas Zilinskas. On strong homogeneity of two global optimization algorithms based on statistical models of multimodal objective functions. Applied Mathematics and Computation, 218(16):8131–8136, April 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311009982.

Zimbidis:2014:IPU

[Zim14] Alexandros A. Zimbidis. Insurance pricing using H_{∞}-control. Applied Mathematics and Computation, 232(??):685–697, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314001428.

Zitnan:2011:DWL

[Zit11] P. Zitnan. Discrete weighted least-squares method for the Poisson and biharmonic problems on domains with smooth boundary. Applied Mathematics and Computation, 217(22):8973–8982, July 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311004711.

Zhang:2013:FSW

[ZJ13] Li Zhang and Shuyuan Jian. Further studies on the Wei–Yao–Liu nonlinear conjugate gradient method. Applied Mathematics and Computation, 219(14):7616–7621, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S009630031300088X.

Zhao:2014:SLS

[ZJ14a] Meixiang Zhao and Zhigang Jia. Structured least-squares problems and inverse eigenvalue problems for (P, Q)-reflexive matrices. Applied Mathematics and Computation, 235(??):87–93, May 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314003464.

Zhao:2014:TSS

[ZJ14b] Yanan Zhao and Daqing Jiang. The threshold of a stochastic SIS epidemic model with vaccination. Applied Mathematics and Computation, 243(??):718–727, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314008248.

Zhou:2014:SNG

[ZJ14c] Jianwei Zhou and Zhaolin Jiang. The spectral norms of g-circulant matrices with classical Fibonacci and Lucas numbers entries. Applied Mathematics and Computation, 233(??):582–587, May 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314002598.

Zhang:2012:HLS

[ZJD12] Zhening Zhang, Jie Jia, and Ruifeng Ding. Hierarchical least squares based iterative estimation algorithm for multivariable Box-Jenkins-like systems using the auxiliary model. Applied Mathematics and Computation, 218(9):5580–5587, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311013944.

Zhang:2013:APSb

[ZjHO13] Cong Zhang, Nan jing Huang, and Donal O’Regan. Almost periodic solutions for a Volterra model with mutual interference and Holling type III functional response. Applied Mathematics and Computation, 225 (??):503–511, December 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313010266.

Zhang:2014:LTB

[ZJLO14] Qiumei Zhang, Daqing Jiang, Zhenwen Liu, and Donal O’Regan. The long time behavior of a predator–prey model with disease in the prey by stochastic perturbation. Applied Mathematics and Computation, 245(??):305–320, October 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314010558.

Zhang:2010:BAD

[ZJS10] Fen-Fen Zhang, Zhen Jin, and Gui-Quan Sun. Bifurcation analysis of a delayed epidemic model. Applied Mathematics and Computation, 216(3):753–767, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310000986.

Zhu:2010:NSB

[ZK10a] Chun-Gang Zhu and Wen-Sheng Kang. Numerical solution of Burgers–Fisher equation by cubic B-spline quasi-interpolation. Applied Mathematics and Computation, 216 (9):2679–2686, July 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031000367X.

Zhu:2010:GGA

[ZK10b] Guopu Zhu and Sam Kwong. Gbest-guided artificial bee colony algorithm for numerical function optimization. Applied Mathematics and Computation, 217(7):3166–3173, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300310009136.

Zhou:2011:IFE

[ZK11] Ruguang Zhou and Ying Kui. Integrable fermionic extensions of the Garnier system and the anharmonic oscillator. Applied Mathematics and Computation, 217(22):9261–9266, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311005443.

Zubik-Kowal:2014:NAG

[ZK14] B. Zubik-Kowal. Numerical algorithm for the growth of human tumor cells and their responses to therapy. Applied Mathematics and Computation, 230(??):174–179, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313013003.

Zhang:2010:SDB

[ZL10a] Li Zhang and Shu Tang Liu. Spatial dynamical behavior of narrow composite beams. Applied Mathematics and Computation, 216 (12):3627–3633, August 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310005291.

Zhang:2010:NSC

[ZL10b] Xinguang Zhang and Lishan Liu. A necessary and sufficient condition of positive solutions for nonlinear singular differential systems with four-point boundary conditions. Applied Mathematics and Computation, 215 (10):3501–3508, January 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309009497.

Zheng:2010:SVI

[ZL10c] Xiuyun Zheng and Hongwei Liu. Solving variational inequalities by a modified projection method with an effective stepsize. Applied Mathematics and Computation, 216(12):3778–3785, August 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310006284.

Zhou:2010:HBR

[ZL10d] Ming-Chun Zhou and Zong-Yu Liu. Hopf bifurcations in a Ricardo–Malthus model. Applied Mathematics and Computation, 217 (6):2425–2432, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310007885.

Zhu:2010:NMB

[ZL10e] Hailong Zhu and Zhaoxiang Li. Newton’s method’s basin of attraction for sign-changing solutions of concave and convex nonlinearities. Applied Mathematics and Computation, 217(7):2937–2943, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310008817.

Zhu:2010:FRR

[ZL10f] Jiehua Zhu and Xiezhang Li. A full rowrank system matrix generated by the stripbased projection model in discrete tomography. Applied Mathematics and Computation, 216(12):3536–3540, August 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310005138.

Zi-LiangLi:2010:CNE

[ZL10g] Zi-LiangLi. Constructing of new exact solutions to the GKdV–mKdV equation with any-order nonlinear terms by \frac{G′}{G} -expansion method. Applied Mathematics and Computation, 217(4):1398–1403, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309005311.

Zhang:2011:HTI

[ZL11a] Lai Zhang and Zhigui Lin. A Hölling’s type II prey–predator model with stage structure and nonlocal delay. Applied Mathematics and Computation, 217(10):5000–5010, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310011835.

Zhang:2011:NGT

[ZL11b] Li Zhang and Junli Li. A new globalization technique for nonlinear conjugate gradient methods for nonconvex minimization. Applied Mathematics and Computation, 217 (24):10295–10304, August 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007016.

Zhao:2011:CEA

[ZL11c] Ning Zhao and Winston T. Lin. A copula entropy approach to correlation measurement at the country level. Applied Mathematics and Computation, 218(2):628–642, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311007983.

Zhang:2012:UMP

[ZL12a] Ping-An Zhang and Wan-Tong Li. Uniqueness of monostable pulsating wave fronts for time periodic reaction-diffusion equations. Applied Mathematics and Computation, 219(3):1300–1305, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312007370.

Zhang:2012:WMF

[ZL12b] Qi Zhang and Yongshun Liang. The Weyl– Marchaud fractional derivative of a type of self-affine functions. Applied Mathematics and Computation, 218(17):8695–8701, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312001531.

Zhao:2012:FDF

[ZL12c] Zhengang Zhao and Changpin Li. Fractional difference/finite element approximations for the time-space fractional telegraph equation. Applied Mathematics and Computation, 219 (6):2975–2988, November 25, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312009216.

Zhu:2012:VVS

[ZL12d] Song-Ping Zhu and Guang-Hua Lian. On the valuation of variance swaps with stochastic volatility. Applied Mathematics and Computation, 219(4):1654–1669, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312007989.

Zhang:2013:PMS

[ZL13a] Minjiao Zhang and Chengxin Luo. Parallelmachine scheduling with deteriorating jobs, rejection and a fixed non-availability interval. Applied Mathematics and Computation, 224(??):405–411, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313007510.

Zhang:2013:ZEM

[ZL13b] Xuemin Zhang and Yeping Li. Zero-electron-mass limit and zero-relaxation-time limit in a multi-dimensional stationary bipolar Euler– Poisson system. Applied Mathematics and Computation, 219(10):5174–5184, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312011575.

Zhou:2013:VFP

[ZL13c] Ling Zhou and Zuhan Liu. Vortex-filaments pinning for inhomogeneous Ginzburg–Landau equations. Applied Mathematics and Computation, 219(10):5321–5338, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312011952.

Zhu:2013:GIG

[ZL13d] Jiehua Zhu and Xiezhang Li. A generalized ℓ_{1} greedy algorithm for image reconstruction in CT. Applied Mathematics and Computation, 219(10):5487–5494, January 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312012192.

Zhu:2013:OCS

[ZL13e] Shengkun Zhu and Shengjie Li. Optimality conditions of strict minimality in optimization problems under inclusion constraints. Applied Mathematics and Computation, 219(9):4816–4825, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031201137X.

Zhu:2013:ENS

[ZL13f] Tao Zhu and Gang Li. Existence of nondecreasing solutions of a quadratic integral equation of Volterra type. Applied Mathematics and Computation, 221(??):214–220, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313007054.

Zeng:2014:ZDS

[ZL14a] Cheng Zeng and Shan Liang. Zero dynamics of sampled-data models for nonlinear multivariable systems in fractional-order hold case. Applied Mathematics and Computation, 246(??):88–102, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314010959.

Zhang:2014:EFB

[ZL14b] L. W. Zhang and K. M. Liew. An elementfree based solution for nonlinear Schrödinger equations using the ICVMLS-Ritz method. Applied Mathematics and Computation, 249 (??):333–345, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031401399X.

Zhang:2014:IML

[ZL14c] L. W. Zhang and K. M. Liew. An improved moving least-squares Ritz method for two-dimensional elasticity problems. Applied Mathematics and Computation, 246 (??):268–282, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314009588.

Zhang:2014:DCS

[ZL14d] Wentao Zhang and Yang Liu. Distributed consensus for sampled-data control multiagent systems with missing control inputs. Applied Mathematics and Computation, 240(??):348–357, August 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314006171.

Zhang:2014:CMI

[ZL14e] Zuhua Zhang and Hui Liang. Collocation methods for impulsive differential equations. Applied Mathematics and Computation, 228(??):336–348, February 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031301254X.

Zhao:2014:NMN

[ZL14f] Guihua Zhao and Mingzhu Liu. Numerical methods for nonlinear stochastic delay differential equations with jumps. Applied Mathematics and Computation, 233(??):222–231, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300314000137.

Zhang:2010:GFE

[ZLA10] H. Zhang, F. Liu, and V. Anh. Galerkin finite element approximation of symmetric space-fractional partial differential equations. Applied Mathematics and Computation, 217 (6):2534–2545, November 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310008118.

Zhao:2010:DPI

[ZLC10] Zhong Zhao, Ying Li, and Lansun Chen. Dynamics of product inhibition on lactic acid fermentation. Applied Mathematics and Computation, 217(1):175–184, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310005837.

Zakynthinaki:2013:DCB

[ZLC^{+}13] Maria S. Zakynthinaki, Alfonso López, Carlos A. Cordente, Jonathan A. Ospina Betancurt, Manuel Sillero Quintana, and Javier Sampedro. Detecting changes in the basin of attraction of a dynamical system: Application to the postural restoring system. Applied Mathematics and Computation, 219(17):8910–8922, May 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031300283X.

Zhao:2011:BAC

[ZLD11] Huitao Zhao, Yiping Lin, and Yunxian Dai. Bifurcation analysis and control of chaos for a hybrid ratio-dependent three species food chain. Applied Mathematics and Computation, 218(5):1533–1546, November 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311007089.

Zhang:2011:BTS

[ZlDxX11] Yi Zhang, Xiao lan Dang, and Hong xian Xu. Bäcklund transformations and soliton solutions for the KdV6 equation. Applied Mathematics and Computation, 217 (13):6230–6236, March 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310013251.

Zhang:2014:NBE

[ZLG^{+}14] Y. M. Zhang, Z. Y. Liu, X. W. Gao, V. Sladek, and J. Sladek. A novel boundary element approach for solving the 2D elasticity problems. Applied Mathematics and Computation, 232(??):568–580, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314001088.

Zhang:2011:LSS

[ZLGZ11] Fengxia Zhang, Ying Li, Wenbin Guo, and Jianli Zhao. Least squares solutions with special structure to the linear matrix equation AXB = C. Applied Mathematics and Computation, 217(24):10049–10057, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311006515.

Zheng:2011:OST

[ZLH11] Quan Zheng, Jingya Li, and Fengxi Huang. An optimal Steffensen-type family for solving nonlinear equations. Applied Mathematics and Computation, 217(23):9592–9597, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311005923.

Zhang:2011:CCB

[ZLJ11] Lina Zhang, Yirong Liu, and Xuejiao Jiang. The center conditions and bifurcation of limit cycles at the infinity for a cubic polynomial system. Applied Mathematics and Computation, 218(4):1360–1370, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311008459.

Zhang:2014:GSA

[ZLJZ14] Hui Zhang, Yingqi Li, Bin Jing, and Weizhou Zhao. Global stability of almost periodic solution of multispecies mutualism system with time delays and impulsive effects. Applied Mathematics and Computation, 232(??):1138–1150, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314001763.

Zeng:2010:NAP

[ZLL10] Xiao-Ming Zeng, Diwang Lin, and Lingling Li. A note on approximation properties of q-Durrmeyer operators. Applied Mathematics and Computation, 216(3):819–821, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310001104.

Zhang:2011:CVZ

[ZLL11a] Yunong Zhang, Zhan Li, and Kene Li. Complex-valued Zhang neural network for online complex-valued time-varying matrix inversion. Applied Mathematics and Computation, 217(24):10066–10073, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311006552.

Zhao:2011:SCF

[ZLL11b] Xiaoshan Zhao, Zhenbo Li, and Shuang Li. Synchronization of a chaotic finance system. Applied Mathematics and Computation, 217(13):6031–6039, March 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310007629.

Zhao:2013:TIL

[ZLL13] Qiu-Lan Zhao, Xin-Yue Li, and Fa-Sheng Liu. Two integrable lattice hierarchies and their respective Darboux transformations. Applied Mathematics and Computation, 219 (10):5693–5705, January 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312012209.

Zhang:2014:SCS

[ZLL14] Qimin Zhang, Yating Liu, and Xining Li. Strong convergence of split-step backward Euler method for stochastic age-dependent capital system with Markovian switching. Applied Mathematics and Computation, 235(??):439–453, May 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314000289.

Zhao:2012:IHA

[ZLLL12] Yinlong Zhao, Zhiliang Lin, Zeng Liu, and Shijun Liao. The improved homotopy analysis method for the Thomas–Fermi equation. Applied Mathematics and Computation, 218(17):8363–8369, May 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312001117.

Zhang:2012:RFTa

[ZLM12] Yingqi Zhang, Caixia Liu, and Xiaowu Mu. Robust finite-time H_{∞} control of singular stochastic systems via static output feedback. Applied Mathematics and Computation, 218(9):5629–5640, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311014007.

Zhang:2012:COB

[ZLMG12] Yingqi Zhang, Caixia Liu, Xiaowu Mu, and Xiulan Guo. Comment on “Observer-based finite-time control of time-delayed jump systems”. Applied Mathematics and Computation, 218(17):9027–9032, May 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312001658. See [HL10].

Zhang:2012:CMN

[ZLS12a] Kai Zhang, Jie Li, and Haiming Song. Collocation methods for nonlinear convolution Volterra integral equations with multiple proportional delays. Applied Mathematics and Computation, 218(22):10848–10860, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312004444.

Zhang:2012:RFTb

[ZLS12b] Yingqi Zhang, Caixia Liu, and Huixia Sun. Robust finite-time H_{∞} control for uncertain discrete jump systems with time delay. Applied Mathematics and Computation, 219(5):2465–2477, November 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312008752.

Zhang:2013:GTA

[ZLSW13] Chunmei Zhang, Wenxue Li, Huan Su, and Ke Wang. A graph-theoretic approach to boundedness of stochastic Cohen–Grossberg neural networks with Markovian switching. Applied Mathematics and Computation, 219(17):9165–9173, May 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313003019.

Zhu:2010:PSS

[ZLW10] Fengling Zhu, Lishan Liu, and Yonghong Wu. Positive solutions for systems of a nonlinear fourth-order singular semipositone boundary value problems. Applied Mathematics and Computation, 216(2):448–457, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310000512.

Zhang:2012:EPS

[ZLW12a] Xinguang Zhang, Lishan Liu, and Yonghong Wu. The eigenvalue problem for a singular higher order fractional differential equation involving fractional derivatives. Applied Mathematics and Computation, 218(17):8526–8536, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S009630031200135X.

Zhang:2012:ERM

[ZLW12b] Xinguang Zhang, Lishan Liu, and Yonghong Wu. Existence results for multiple positive solutions of nonlinear higher order perturbed fractional differential equations with derivatives. Applied Mathematics and Computation, 219(4):1420–1433, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031200762X.

Zhang:2013:ISN

[ZLWL13] Xinguang Zhang, Lishan Liu, Yonghong Wu, and Yinan Lu. The iterative solutions of nonlinear fractional differential equations. Applied Mathematics and Computation, 219(9):4680–4691, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312010983.

Zhang:2014:ECS

[ZLWW14] Xinguang Zhang, Lishan Liu, Benchawan Wiwatanapataphee, and Yonghong Wu. The eigenvalue for a class of singular p-Laplacian fractional differential equations involving the Riemann–Stieltjes integral boundary condition. Applied Mathematics and Computation, 235(??):412–422, May 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314003014.

Zhang:2012:SCP

[ZLX12] G. Y. Zhang, G. R. Liu, and X. Xu. A strain-constructed point interpolation method and strain field construction schemes for solid mechanics problems using triangular mesh. Applied Mathematics and Computation, 219 (4):2067–2086, November 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312008454.

Zhang:2011:HBT

[ZLY11a] Jia-Fang Zhang, Wan-Tong Li, and Xiang-Ping Yan. Hopf bifurcation and Turing instability in spatial homogeneous and inhomogeneous predator-prey models. Applied Mathematics and Computation, 218 (5):1883–1893, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311009167.

Zhao:2011:NMP

[ZLY11b] W. Zhao, J. K. Liu, and J. J. Ye. A new method for parameter sensitivity estimation in structural reliability analysis. Applied Mathematics and Computation, 217 (12):5298–5306, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310011823.

Zhu:2011:ESS

[ZLY11c] Quanxin Zhu, Xiaodi Li, and Xinsong Yang. Exponential stability for stochastic reaction–diffusion BAM neural networks with time-varying and distributed delays. Applied Mathematics and Computation, 217 (13):6078–6091, March 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310012944.

Zhou:2010:NSH

[ZLYL10] Yue-Ting Zhou, Jin Li, De-Hao Yu, and Kang Yong Lee. Numerical solution of hypersingular equation using recursive wavelet on invariant set. Applied Mathematics and Computation, 217(2):861–868, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310007149.

Zhang:2012:MSC

[ZLZ12] Qi Zhang, Fuyi Li, and Xiaoli Zhu. Multiple sign-changing solutions to the Sturm– Liouville boundary value problem with resonance. Applied Mathematics and Computation, 219(3):1061–1072, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312007126.

Zhang:2014:NAN

[ZLZ^{+}14a] H. Zhang, F. Liu, P. Zhuang, I. Turner, and V. Anh. Numerical analysis of a new space– time variable fractional order advectiondispersion equation. Applied Mathematics and Computation, 242(??):541–550, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314008376.

Zhang:2014:ACF

[ZLZ14b] Xuxi Zhang, Xianping Liu, and Qidan Zhu. Adaptive chatter free sliding mode control for a class of uncertain chaotic systems. Applied Mathematics and Computation, 232(??):431–435, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314001313.

Zhang:2014:NSE

[ZLZ14c] Zujin Zhang, Peng Li, and Dingxing Zhong. Navier–Stokes equations with regularity in two entries of the velocity gradient tensor. Applied Mathematics and Computation, 228(??):546–551, February 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313012873.

Zhao:2014:EPS

[ZLZ14d] Xinchao Zhao, Wenqiao Lin, and Qingfu Zhang. Enhanced particle swarm optimization based on principal component analysis and line search. Applied Mathematics and Computation, 229(??):440–456, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313013519.

Zhou:2012:KWS

[ZLZY12] Yuqian Zhou, Qian Liu, Weinian Zhang, and Fuchun Yang. Kink wave and saddlespiral shock wave of a partial differential equation model of excitable media. Applied Mathematics and Computation, 219 (8):3496–3503, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312009228.

Zhanlav:2010:LIC

[ZM10a] T. Zhanlav and R. Mijiddorj. The local integro cubic splines and their approximation properties. Applied Mathematics and Computation, 216(7):2215–2219, June 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310002791.

Meng:2010:SNS

[zM10b] Xin zhu Meng. Stability of a novel stochastic epidemic model with double epidemic hypothesis. Applied Mathematics and Computation, 217(2):506–515, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310006429.

Zu:2010:IAE

[ZM10c] Jian Zu and Masayasu Mimura. The impact of Allee effect on a predator–prey system with Holling type II functional response. Applied Mathematics and Computation, 217 (7):3542–3556, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310009768.

Zhang:2012:NMS

[ZM12] Jianjun Zhang and Musa Mammadov. A new method for solving linear ill-posed problems. Applied Mathematics and Computation, 218(20):10180–10187, June 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031200361X.

Zhang:2014:ETR

[ZM14a] Huiqun Zhang and Wen-Xiu Ma. Extended transformed rational function method and applications to complexiton solutions. Applied Mathematics and Computation, 230 (??):509–515, March 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031301446X.

Zhanlav:2014:IQS

[ZM14b] T. Zhanlav and R. Mijiddorj. Integro quintic splines and their approximation properties. Applied Mathematics and Computation, 231(??):536–543, March 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314000800.

Zhao:2014:FGM

[ZM14c] Xinran Zhao and Mehrdad Massoudi. Flow of granular materials with slip boundary condition: a continuum-kinetic theory approach. Applied Mathematics and Computation, 242(??):518–527, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314007930.

Zheng:2014:CAU

[ZM14d] Qingqing Zheng and Changfeng Ma. A class of accelerated Uzawa algorithms for saddle point problems. Applied Mathematics and Computation, 247(??):244–254, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314012144.

Zheng:2014:NSL

[ZM14e] Qingqing Zheng and Changfeng Ma. A new SOR-like method for the saddle point problems. Applied Mathematics and Computation, 233(??):421–429, May 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314002501.

Zhu:2010:DMI

[ZMC10] Guanghu Zhu, Xinzhu Meng, and Lansun Chen. The dynamics of a mutual interference age structured predator–prey model with time delay and impulsive perturbations on predators. Applied Mathematics and Computation, 216(1):308–316, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310000779.

Zdravkovic:2014:MDC

[ZMD^{+}14] S. Zdravković, A. Maluckov, M. Dekić, S. Kuzmanović, and M. V. Satarić. Are microtubules discrete or continuum systems? Applied Mathematics and Computation, 242 (??):353–360, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314007334.

Zheng:2013:HBE

[ZMG13] Hong-Chan Zheng, Wen-Xiu Ma, and Xiang Gu. Hirota bilinear equations with linear subspaces of hyperbolic and trigonometric function solutions. Applied Mathematics and Computation, 220(??):226–234, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313006322.

Zhou:2011:IQH

[ZMZ11] Shouming Zhou, Chunlai Mu, and Rong Zeng. Incomplete quenching of heat equations with coupled singular logarithms boundary fluxes. Applied Mathematics and Computation, 217 (12):5491–5499, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310012208.

Zhang:2012:GDN

[ZMZS12] Tongqian Zhang, Xinzhu Meng, Tonghua Zhang, and Yi Song. Global dynamics for a new high-dimensional SIR model with distributed delay. Applied Mathematics and Computation, 218(24):11806–11819, August 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300312005395.

Zhu:2014:FIT

[ZnXwY14] Tao Zhu, Shou ne Xiao, and Guang wu Yang. Force identification in time domain based on dynamic programming. Applied Mathematics and Computation, 235(??):226–234, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314003610.

Zhou:2014:ESS

[ZnZ14] Tong Zhou and Zuo nong Zhu. Explicit solutions for a semidiscrete Boussinesq system. Applied Mathematics and Computation, 249(??):121–132, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314014076.

Zerarka:2010:FVM

[ZOA10] A. Zerarka, S. Ouamane, and A. Attaf. On the functional variable method for finding exact solutions to a class of wave equations. Applied Mathematics and Computation, 217(7):2897–2904, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310009343.

Zegeye:2010:CTE

[ZOS10] Habtu Zegeye, Eric U. Ofoedu, and Naseer Shahzad. Convergence theorems for equilibrium problem, variational inequality problem and countably infinite relatively quasinonexpansive mappings. Applied Mathematics and Computation, 216(12):3439–3449, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310002420.

Zou:2012:UFC

[Zou12] Changwu Zou. Using Fourier coefficients to determine the unstable points of Hill equation. Applied Mathematics and Computation, 218(18):9519–9527, May 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312002755.

Zhang:2012:EIU

[ZP12] Bo Zhang and Jin Peng. Euler index in uncertain graph. Applied Mathematics and Computation, 218(20):10279–10288, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312003888.

Zhu:2013:RFF

[ZP13] Xiaojing Zhu and Dingguo Pu. A restorationfree filter SQP algorithm for equality constrained optimization. Applied Mathematics and Computation, 219(11):6016–6029, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S009630031201260X.

Zeng:2014:IDD

[ZPXX14] Hong-Bing Zeng, Ju H. Park, Jian-Wei Xia, and Shen-Ping Xiao. Improved delay-dependent stability criteria for T – S fuzzy systems with time-varying delay. Applied Mathematics and Computation, 235(??):492–501, May 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314003580.

Zhao:2014:CGC

[ZPZ14] Huanyu Zhao, Ju H. Park, and Yulin Zhang. Couple-group consensus for second-order multi-agent systems with fixed and stochastic switching topologies. Applied Mathematics and Computation, 232(??):595–605, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314000551.

Zhang:2014:MRK

[ZQ14] Chengjian Zhang and Tingting Qin. The mixed Runge–Kutta methods for a class of nonlinear functional-integro-differential equations. Applied Mathematics and Computation, 237(??):396–404, June 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314005141.

Zhang:2013:CNS

[ZR13] Qimin Zhang and A. Rathinasamy. Convergence of numerical solutions for a class of stochastic age-dependent capital system with random jump magnitudes. Applied Mathematics and Computation, 219 (14):7297–7305, March 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313000155.

Zhuang:2014:NAN

[ZR14] Qingqu Zhuang and Quanwei Ren. Numerical approximation of a nonlinear fourthorder integro-differential equation by spectral method. Applied Mathematics and Computation, 232(??):775–783, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314002033.

Zhang:2010:CUS

[ZS10a] Jianjun Zhang and Jijuan Shang. A class of Uzawa-SOR methods for saddle point problems. Applied Mathematics and Computation, 216(7):2163–2168, June 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310002936.

Zhou:2010:BTW

[ZS10b] Hongxian Zhou and Jianwei Shen. Bifurcations of travelling wave solutions for modified nonlinear dispersive phi-four equation. Applied Mathematics and Computation, 217(4):1584–1597, October 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309006481.

Zhang:2012:SBE

[ZS12] Kaijun Zhang and Yangfeng Su. Structured backward error analysis for sparse polynomial eigenvalue problems. Applied Mathematics and Computation, 219(6):3073–3082, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300312009344.

Zdravkovic:2013:NDD

[ZS13a] Slobodan Zdravković and Miljko V. Satarić. Nonlinear DNA dynamics: Nonlinearity versus dispersion. Applied Mathematics and Computation, 225(??):401–406, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300313010436.

Zhang:2013:APSa

[ZS13b] Hong Zhang and Jianying Shao. Almost periodic solutions for cellular neural networks with time-varying delays in leakage terms. Applied Mathematics and Computation, 219 (24):11471–11482, August 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313005936.

Zegeye:2014:AFC

[ZS14a] H. Zegeye and N. Shahzad. An algorithm for finding a common point of the solution set of a variational inequality and the fixed point set of a Bregman relatively nonexpansive mapping. Applied Mathematics and Computation, 248(??):225–234, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314011813.

Zhang:2014:DAL

[ZS14b] Jia-Fang Zhang and Yu-Juan Sun. Dynamical analysis of a logistic equation with spatiotemporal delay. Applied Mathematics and Computation, 247(??):996–1002, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314012922.

Zheng:2014:SNM

[ZS14c] Xiuyun Zheng and Jiarong Shi. Smoothing Newton method for generalized complementarity problems based on a new smoothing function. Applied Mathematics and Computation, 231(??):160–168, March 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314000095.

Zhou:2014:CRO

[ZS14d] Xiaojian Zhou and Yongzhong Song. Convergence radius of Osada’s method under center–Hölder continuous condition. Applied Mathematics and Computation, 243 (??):809–816, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314009114. See note [HMGC18].

Zhou:2014:DBS

[ZS14e] Xiaojian Zhou and Yongzhong Song. Dynamical behaviors of some iterative methods for multiple roots of nonlinear equations. Applied Mathematics and Computation, 246(??):661–677, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314011898.

Zolfaghari:2014:RUT

[ZS14f] R. Zolfaghari and A. Shidfar. Reconstructing an unknown time-dependent function in the boundary conditions of a parabolic PDE. Applied Mathematics and Computation, 226(??):238–249, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313011314.

Zhang:2015:ANP

[ZS15] Shuyi Zhang and Xiaoguang Song. Another note on a paper “Convergence theorem for the common solution for a finite family of φ-strongly accretive operator equations”. Applied Mathematics and Computation, 258(??):367–371, May 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300315001563.

Zhu:2010:NSP

[ZSD10] Hongqing Zhu, Huazhong Shu, and Meiyu Ding. Numerical solutions of partial differential equations by discrete homotopy analysis method. Applied Mathematics and Computation, 216(12):3592–3605, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310005230.

Zhao:2011:PSB

[ZSHZ11] Yige Zhao, Shurong Sun, Zhenlai Han, and Meng Zhang. Positive solutions for boundary value problems of nonlinear fractional differential equations. Applied Mathematics and Computation, 217(16):6950–6958, April 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311001421.

Zhao:2010:ESE

[ZSL10] Guihua Zhao, M. H. Song, and M. Z. Liu. Exponential stability of Euler–Maruyama solutions for impulsive stochastic differential equations with delay. Applied Mathematics and Computation, 215(9):3425–3432, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300309009424.

Zhang:2011:ASU

[ZSL11] Zhengqiang Zhang, Hao Shen, and Junling Li. Adaptive stabilization of uncertain unified chaotic systems with nonlinear input. Applied Mathematics and Computation, 218 (8):4260–4267, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311012288.

Zdravkovic:2014:NMD

[ZSMB14] S. Zdravković, M. V. Satarić, A. Maluckov, and A. Balaz. A nonlinear model of the dynamics of radial dislocations in microtubules. Applied Mathematics and Computation, 237(??):227–237, June 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314004846.

Zhou:2010:NGA

[ZSQ10] Qunyan Zhou, Wenyu Sun, and Liqun Qi. A nonmonotone globalization algorithm with preconditioned gradient path for unconstrained optimization. Applied Mathematics and Computation, 217(8):4257–4264, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300310011021.

Zhang:2012:ROO

[ZSWS12] Zhengqiang Zhang, Hanyong Shao, Zhen Wang, and Hao Shen. Reduced-order observer design for the synchronization of the generalized Lorenz chaotic systems. Applied Mathematics and Computation, 218 (14):7614–7621, March 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312000641.

Sun:2012:LDS

[zSWZW12] Zhi zhong Sun, Xiaonan Wu, Jiwei Zhang, and Desheng Wang. A linearized difference scheme for semilinear parabolic equations with nonlinear absorbing boundary conditions. Applied Mathematics and Computation, 218(9):5187–5201, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311013324.

Zhang:2012:SIS

[ZSZ12] Shuorui Zhang, Jitao Sun, and Yu Zhang. Stability of impulsive stochastic differential equations in terms of two measures via perturbing Lyapunov functions. Applied Mathematics and Computation, 218(9):5181–5186, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311013312.

Zhao:2011:IHC

[ZT11] Tingting Zhao and Sanyi Tang. Impulsive harvesting and by-catch mortality for the theta logistic model. Applied Mathematics and Computation, 217(22):9412–9423, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311005881.

Zhang:2012:EHO

[ZT12] Qiongfen Zhang and X. H. Tang. Existence of homoclinic orbits for a class of asymptotically p-linear aperiodic p-Laplacian systems. Applied Mathematics and Computation, 218(13):7164–7173, March 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311015748.

Zhang:2013:NMP

[ZT13] Xingyong Zhang and Xianhua Tang. A note on the minimal periodic solutions of nonconvex superlinear Hamiltonian system. Applied Mathematics and Computation, 219 (14):7586–7590, March 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313000842.

Zhang:2014:EMSa

[ZT14] Xianghong Zhang and Sanyi Tang. Existence of multiple sliding segments and bifurcation analysis of Filippov prey-predator model. Applied Mathematics and Computation, 239(??):265–284, July 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314006389.

Zhang:2010:DBB

[ZTD10] Li Zhang, Jieqing Tan, and Zhiyuan Dong. The dual bases for the Bézier–Said–Wang type generalized Ball polynomial bases and their applications. Applied Mathematics and Computation, 217(7):3088–3101, December 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310008957.

Zhou:2010:NET

[ZTF10] Jiangbo Zhou, Lixin Tian, and Xinghua Fan. New exact travelling wave solutions for the K(2, 2) equation with osmosis dispersion. Applied Mathematics and Computation, 217(4):1355–1366, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300309004548.

Zhang:2010:BTW

[ZTW10] Kelei Zhang, Shengqiang Tang, and Zhaojuan Wang. Bifurcations of travelling wave solutions for the nonlinear dispersion Drinfel’d– Sokolov (D(m, n)) system. Applied Mathematics and Computation, 217(4):1620–1631, October 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300309006912.

Zheng:2010:NFF

[ZTX10] Z. C. Zheng, B. K. Tan, and Y. Xu. Nearfield fluctuations and far-field noise of a threeelement airfoil system by a discrete vortex method. Applied Mathematics and Computation, 216(4):1072–1086, April 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310001530.

Zhang:2010:CHP

[ZTY^{+}10] Xianquan Zhang, Zhenjun Tang, Jinhui Yu, Mingming Guo, and Lianyuan Jiang. Convex hull properties and algorithms. Applied Mathematics and Computation, 216 (11):3209–3218, August 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310004728.

Zhang:2014:EMSb

[ZTZ14] Jian Zhang, Xianhua Tang, and Wen Zhang. Existence of multiple solutions of Kirchhoff type equation with sign-changing potential. Applied Mathematics and Computation, 242(??):491–499, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314007358.

Zufiria:2012:MFN

[Zuf12] Pedro J. Zufiria. A mathematical framework for new fault detection schemes in nonlinear stochastic continuous-time dynamical systems. Applied Mathematics and Computation, 218(23):11391–11403, August 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312005097.

Zuo:2010:AEE

[Zuo10]  Jin-Ming Zuo.  Application of the ex-

tended (\frac{G′}{G})-expansion method to solve the Pochhammer–Chree equations. Applied Mathematics and Computation, 217 (1):376–383, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310006314.

Zuo:2013:GSH

[Zuo13] Wenjie Zuo. Global stability and Hopf bifurcations of a Beddington–DeAngelis type predator-prey system with diffusion and delays. Applied Mathematics and Computation, 223(??):423–435, October 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313008977.

Zhang:2010:OSO

[ZW10a] Shao-Yan Zhang and Qi-Ru Wang. Oscillation of second-order nonlinear neutral dynamic equations on time scales. Applied Mathematics and Computation, 216(10):2837–2848, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300310004005.

Zhao:2010:ICS

[ZW10b] Qiu-Lan Zhao and Xin-Zeng Wang. The integrable coupling system of a 3 × 3 discrete matrix spectral problem. Applied Mathematics and Computation, 216(3):730–743, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310000937.

Zhou:2010:SD

[ZW10c] Jieyong Zhou and Hui Wei. A simpler DGM-RES. Applied Mathematics and Computation, 217(1):124–129, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310005680.

Zhou:2010:GRA

[ZW10d] Qinghua Zhou and Li Wan. Global robust asymptotic stability analysis of BAM neural networks with time delay and impulse: an LMI approach. Applied Mathematics and Computation, 216(5):1538–1545, May 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310002456.

Zhang:2011:SSE

[ZW11a] Qin Zhang and Qing-Wen Wang. The (P, Q)-(skew)symmetric extremal rank solutions to a system of quaternion matrix equations. Applied Mathematics and Computation, 217(22):9286–9296, July 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311005522.

Zhang:2011:ERP

[ZW11b] Yuan Lin Zhang and Guan Jun Wang. An extended replacement policy for a deteriorating system with multi-failure modes. Applied Mathematics and Computation, 218 (5):1820–1830, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311009118.

Zheng:2011:SRM

[ZW11c] G. H. Zheng and T. Wei. Spectral regularization method for solving a time-fractional inverse diffusion problem. Applied Mathematics and Computation, 218(2):396–405, September 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300311007570.

Zhang:2012:PCC

[ZW12] Pei-Guang Zhang and Jian-Ping Wang. A predictor-corrector compact finite difference scheme for Burgers’ equation. Applied Mathematics and Computation, 219 (3):892–898, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312006728.

Zhang:2013:GPI

[ZW13a] Guo-Feng Zhang and Shan-Shan Wang. A generalization of parameterized inexact Uzawa method for singular saddle point problems. Applied Mathematics and Computation, 219(9):4225–4231, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312011526.

Zhang:2013:NSC

[ZW13b] Qimin Zhang and Dongmei Wei. Necessary and sufficient conditions for near-optimal harvesting control problem of stochastic agedependent system. Applied Mathematics and Computation, 221(??):394–402, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313006875.

Zhang:2013:FFT

[ZW13c] Sumei Zhang and Lihe Wang. Fast Fourier Transform option pricing with stochastic interest rate, stochastic volatility and double jumps. Applied Mathematics and Computation, 219(23):10928–10933, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313005274.

Zhang:2013:ESS

[ZW13d] Yang Zhang and Rong-Hao Wang. The exact solution of a system of quaternion matrix equations involving η-hermicity. Applied Mathematics and Computation, 222 (??):201–209, October 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313007753.

Zhang:2013:IUS

[ZW13e] Z. Q. Zhang and T. Wei. Identifying an unknown source in time-fractional diffusion equation by a truncation method. Applied Mathematics and Computation, 219 (11):5972–5983, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312013008.

Zhang:2014:SSM

[ZW14a] Xianghua Zhang and Ke Wang. Stochastic SEIR model with jumps. Applied Mathematics and Computation, 239(??):133–143, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300314006018.

Zhao:2014:CCF

[ZW14b] Meichun Zhao and Junwei Wang. H_{∞} control of a chaotic finance system in the presence of external disturbance and input timedelay. Applied Mathematics and Computation, 233(??):320–327, May 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313013751.

Zhao:2014:NAD

[ZW14c] Zhong Zhao and Xianbin Wu. Nonlinear analysis of a delayed stage-structured predatorprey model with impulsive effect and environment pollution. Applied Mathematics and Computation, 232(??):1262–1268, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314000332.

Zong:2014:GSG

[ZW14d] Yidan Zong and Li Wang. Global simpler GM-RES for nonsymmetric systems with multiple right-hand sides. Applied Mathematics and Computation, 248(??):371–379, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314013332.

Zhang:2012:ASS

[ZWD12] Zhengqiang Zhang, Yanxia Wang, and Zhenbin Du. Adaptive synchronization of singledegree-of-freedom oscillators with unknown parameters. Applied Mathematics and Computation, 218(12):6833–6840, February 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311015438.

Zong:2014:CSS

[ZWH14] Xiaofeng Zong, Fuke Wu, and Chengming Huang. Convergence and stability of the semitamed Euler scheme for stochastic differential equations with non-Lipschitz continuous coefficients. Applied Mathematics and Computation, 228(??):240–250, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313012691.

Zhao:2014:CSB

[ZWL14a] Jiakun Zhao, Ying Wu, and Qingfang Liu. Chaos synchronization between the coupled systems on network with unknown parameters. Applied Mathematics and Computation, 229(??):254–259, February 25, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313013490.

Zou:2014:CPZ

[ZWL14b] Xiaoling Zou, Ke Wang, and Meng Liu. Can protection zone potentially strengthen protective effects in random environments? Applied Mathematics and Computation, 231 (??):26–38, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314000113.

Zhu:2010:SPP

[ZWLW10] Li-Mei Zhu, Ai-Ling Wang, Yong-Jiang Liu, and Biao Wang. Stationary patterns of a predator–prey model with spatial effect. Applied Mathematics and Computation, 216 (12):3620–3626, August 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310005266.

Zhao:2013:CCT

[ZWP13] Zhi-Wen Zhao, De-Hui Wang, and Cui-Xin Peng. Coefficient constancy test in generalized random coefficient autoregressive model. Applied Mathematics and Computation, 219 (20):10283–10292, June 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313003974.

Zhang:2014:EPS

[ZWS14] Xingqiu Zhang, Lin Wang, and Qian Sun. Existence of positive solutions for a class of nonlinear fractional differential equations with integral boundary conditions and a parameter. Applied Mathematics and Computation, 226(??):708–718, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313011569.

Zhou:2012:EAC

[ZWT12] Han Zhou, Yu-Jiang Wu, and WenYi Tian. Extrapolation algorithm of compact ADI approximation for two-dimensional parabolic equation. Applied Mathematics and Computation, 219(6):2875–2884, November 25, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312009058.

Zhao:2012:EES

[ZWZ12] Yumei Zhao, Ling Wang, and Hongyong Zhao. Existence and exponential stability of almost periodic solutions for a neutral multi-species Logarithmic population model. Applied Mathematics and Computation, 218(9):5346–5356, January 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311013488.

Zhang:2013:IST

[ZWZ13] YanXia Zhang, ZhongZhi Wang, and Xuefeng Zhang. The intersection of the spectra of two Sturm–Liouville equations. Applied Mathematics and Computation, 219 (9):4232–4238, January 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312008983.

Zong:2013:FTS

[ZWZH13] Guangdeng Zong, Ruihua Wang, Wei Xing Zheng, and Linlin Hou. Finite-time stabilization for a class of switched time-delay systems under asynchronous switching. Applied Mathematics and Computation, 219 (11):5757–5771, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312012453.

Zhang:2010:PSD

[ZX10] Dong Zhang and Jian Xu. Projective synchronization of different chaotic time-delayed neural networks based on integral sliding mode controller. Applied Mathematics and Computation, 217(1):164–174, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310005813.

Zhang:2011:VCJ

[ZX11a] Sheng Zhang and Tiecheng Xia. Variablecoefficient Jacobi elliptic function expansion method for (2 + 1)-dimensional Nizhnik–Novikov–Vesselov equations. Applied Mathematics and Computation, 218 (4):1308–1316, October 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311008277.

Zhu:2011:MSF

[ZX11b] Xing-Hua Zhu and Jian-Zhong Xiao. Minimum selections and fixed points of set-valued operators in Banach spaces with some uniform convexity. Applied Mathematics and Computation, 217(12):6004–6010, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310012993.

Zhang:2012:ODI

[ZX12] Shu Zhang and Jian Xu. Oscillatory dynamics induced by time delay in an Internet congestion control model with ring topology. Applied Mathematics and Computation, 218(22):11033–11041, July 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312004638.

Zhang:2013:TWG

[ZX13] Shihua Zhang and Rui Xu. Travelling waves and global attractivity of an SIRS disease model with spatial diffusion and temporary immunity. Applied Mathematics and Computation, 224(??):635–651, November 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313009776.

Zhou:2014:NSC

[ZX14] Fengying Zhou and Xiaoyong Xu. Numerical solution of the convection diffusion equations by the second kind Chebyshev wavelets. Applied Mathematics and Computation, 247 (??):353–367, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314012016.

Zhang:2010:RGS

[ZXCL10] Yijun Zhang, Shengyuan Xu, Yuming Chu, and Jinjun Lu. Robust global synchronization of complex networks with neutral-type delayed nodes. Applied Mathematics and Computation, 216(3):768–778, April 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310000998.

Zhou:2014:ASE

[ZXF14] Shaobo Zhou, Songfa Xie, and Zheng Fang. Almost sure exponential stability of the backward Euler–Maruyama discretization for highly nonlinear stochastic functional differential equation. Applied Mathematics and Computation, 236(??):150–160, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314003634.

Zhang:2014:SNR

[ZxHW14] Jie Zhang, Su xiang He, and Quan Wang. A SAA nonlinear regularization method for a stochastic extended vertical linear complementarity problem. Applied Mathematics and Computation, 232(??):888–897, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314001660.

Zhu:2011:UCC

[ZXZ11] Peng Zhu, Ziqing Xie, and Shuzi Zhou. A uniformly convergent continuous–discontinuous Galerkin method for singularly perturbed problems of convection–diffusion type. Applied Mathematics and Computation, 217 (9):4781–4790, January 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310011458.

Zhao:2012:EDP

[ZY12] Yongxia Zhao and Chuancun Yin. The expected discounted penalty function under a renewal risk model with stochastic income. Applied Mathematics and Computation, 218 (10):6144–6154, January 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311014470.

Zhang:2013:NMM

[ZY13a] Jun Zhang and Yu-Fei Yang. Nonlinear multigrid method for solving the LLT model. Applied Mathematics and Computation, 219 (10):4964–4976, January 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312012271.

Zhang:2013:ULD

[ZY13b] Tie Zhang and Shun Yu. An upwind-like discontinuous Galerkin method for hyperbolic systems. Applied Mathematics and Computation, 219(16):8610–8621, April 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313002142.

Zheng:2013:CPT

[ZY13c] Ning Zheng and Jun-Feng Yin. On the convergence of projected triangular decomposition methods for pricing American options with stochastic volatility. Applied Mathematics and Computation, 223(??):411–422, October 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313008710.

Zhao:2014:NPE

[ZY14a] Dianli Zhao and Sanling Yuan. A note on persistence and extinction of a randomized food-limited logistic population model. Applied Mathematics and Computation, 246 (??):599–607, November 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314011801.

Zhao:2014:AAD

[ZY14b] Kang Zhao and Guozhu Yao. Application of the alternating direction method for an inverse monic quadratic eigenvalue problem. Applied Mathematics and Computation, 244(??):32–41, October 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314009680.

Zhang:2012:CAS

[ZYC12] Zhi-Yong Zhang, Xue-Lin Yong, and Yu-Fu Chen. Classification and approximate solutions to perturbed diffusion-convection equations. Applied Mathematics and Computation, 219(3):1120–1124, October 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312007199.

Zhi:2014:CLL

[ZYC14] Hongyan Zhi, Zhongyuan Yang, and Hui Chang. Conservation laws, Lie symmetry and Painlevé analysis of the variable coefficients NNV equation. Applied Mathematics and Computation, 249(??):174–181, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314014052.

Zhou:2014:NSI

[ZYG14] Xiaojun Zhou, Chunhua Yang, and Weihua Gui. Nonlinear system identification and control using state transition algorithm. Applied Mathematics and Computation, 226(??):169–179, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313010345.

Zheng:2012:VMM

[ZYLS12] Haibiao Zheng, Jiaping Yu, Kaitai Li, and Feng Shi. A variational multiscale method with bubble stabilization for the Oseen problem based on two local Gauss integrations. Applied Mathematics and Computation, 219 (8):3701–3708, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312009848.

Zhu:2011:QCC

[ZYM11] Qingbao Zhu, Zhijun Yang, and Wei Ma. A quickly convergent continuous ant colony optimization algorithm with Scout Ants. Applied Mathematics and Computation, 218 (5):1805–1819, November 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311009106.

Zhang:2013:BDE

[ZYP13] Qiang Zhang, Dongming Yan, and Zhigang Pan. Bifurcation from double eigenvalue for nonlinear equation with third-order nondegenerate singularity. Applied Mathematics and Computation, 220(??):549–559, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300313006371.

Zheng:2011:DLW

[ZYSQ11] Xiaoyang Zheng, Xiaofan Yang, Hong Su, and Liqiong Qiu. Discontinuous Legendre wavelet element method for elliptic partial differential equations. Applied Mathematics and Computation, 218(7):3002–3018, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031101085X.

Zhu:2012:OCC

[ZYYZ12] Qingyi Zhu, Xiaofan Yang, Lu-Xing Yang, and Chunming Zhang. Optimal control of computer virus under a delayed model. Applied Mathematics and Computation, 218 (23):11613–11619, August 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312005589.

Zhang:2012:LDG

[ZYZ12] Rongpei Zhang, Xijun Yu, and Guozhong Zhao. The local discontinuous Galerkin method for Burger’s–Huxley and Burger’s– Fisher equations. Applied Mathematics and Computation, 218(17):8773–8778, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300312001622.

Zhang:2014:ADT

[ZYZ14a] Gui-Qing Zhang, Qiu-Ying Yang, and Ai-Zhen Zhang. Avalanche duration time in a simple heterogeneous Olami–Feder–Christensen model. Applied Mathematics and Computation, 242(??):346–352, September 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314007383.

Zhao:2014:ISC

[ZYZ14b] Dianli Zhao, Sanling Yuan, and Tiansi Zhang. Improved stability conditions for a class of stochastic Volterra–Levin equations. Applied Mathematics and Computation, 231(??):39–47, March 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314000599.

Zhu:2014:BRM

[ZYZ14c] Shihui Zhu, Han Yang, and Jian Zhang. Blow-up rate, mass concentration and asymptotic profile of blow-up solutions for the nonlinear inhomogeneous Schrödinger equation. Applied Mathematics and Computation, 242(??):889–897, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314009047.

Zhang:2010:CPG

[ZZ10a] Guo-Feng Zhang and Jing-Yu Zhao. On constraint preconditioners for generalized saddle point matrices. Applied Mathematics and Computation, 216(6):1837–1844, May 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300309010686.

Zhao:2010:GSQ

[ZZ10b] Jiakun Zhao and Kecun Zhang. A general scheme for Q–S synchronization of chaotic systems with unknown parameters and scaling functions. Applied Mathematics and Computation, 216(7):2050–2057, June 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310002778.

Zhao:2010:SAE

[ZZ10c] Lei Zhao and Shuigeng Zhou. Symbolic analysis and exact travelling wave solutions to a new modified Novikov equation. Applied Mathematics and Computation, 217 (2):590–598, September 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310006661.

Zhou:2010:BLC

[ZZ10d] Hongxian Zhou and Yanmin Zhao. Bifurcations of limit cycles in a Z_{4}-equivariant planar polynomial vector field of degree

7. Applied Mathematics and Computation, 216(1):35–50, March 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300309011126.

Zhou:2011:ADZ

[ZZ11a] Jian-Rong Zhou and Yu-Qiu Zhao. Asymptotic distributions of the zeros of certain classes of Gauss hypergeometric polynomials. Applied Mathematics and Computation, 218(3):1153–1159, October 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/

S0096300311007892. Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Zhu:2011:NPC

[ZZ11b] Song-Ping Zhu and Jin Zhang. A new predictor–corrector scheme for valuing American puts. Applied Mathematics and Computation, 217(9):4439–4452, January 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310011045.

Zuo:2011:AEM

[ZZ11c] Jin-Ming Zuo and Yao-Ming Zhang. Application of the \frac{G′}{G} -expansion method to solve coupled MKdV equations and coupled Hirota–Satsuma coupled KdV equations. Applied Mathematics and Computation, 217 (12):5936–5941, February 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310013214.

Zhong:2012:MAD

[ZZ12a] Yuan-Guang Zhong and Yong-Wu Zhou. The model and algorithm for determining optimal ordering/trade-credit policy of supply chains. Applied Mathematics and Computation, 219(8):3809–3825, December 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312009964.

Zhou:2012:GRF

[ZZ12b] Jian Zhou and Shiyin Zhao. Generalized reflective function and periodic solution of differential systems with small parameter. Applied Mathematics and Computation, 218(13):7204–7211, March 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311015797.

Zhang:2013:HBT

[ZZ13a] Zhizhou Zhang and Lingling Zhang. Hopf bifurcation of time-delayed feedback control for maglev system with flexible guideway. Applied Mathematics and Computation, 219(11):6106–6112, February 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312013215.

Zhao:2013:FLS

[ZZ13b] Huihong Zhao and Chenghui Zhang. H_{∞} fixed-lag smoothing for linear discrete timevarying systems with uncertain observations. Applied Mathematics and Computation, 224(??):387–397, November 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313009247.

Zhuang:2013:LDG

[ZZ13c] Chijie Zhuang and Rong Zeng. A local discontinuous Galerkin method for 1.5-dimensional streamer discharge simulations. Applied Mathematics and Computation, 219(19):9925–9934, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300313003792.

Zaizai:2014:UMC

[ZZ14a] Yan Zaizai and Hong Zhimin. Using the Monte Carlo method to solve integral equations using a modified control variate. Applied Mathematics and Computation, 242 (??):764–777, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314009230.

Zhang:2014:BSC

[ZZ14b] Fuchen Zhang and Guangyun Zhang. Boundedness solutions of the complex Lorenz chaotic system. Applied Mathematics and Computation, 243(??):12–23, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314008029.

Zhang:2014:BME

[ZZ14c] Songlin Zhang and Kun Zhang. On a basic multivariate EIV model with linear equality constraints. Applied Mathematics and Computation, 236(??):247–252, June 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/

S0096300314004081. See corrigendum [ZZ15].

Zunic:2014:SEH

[ZZ14d] Dragisa Zunić and Jovisa Zunić. Shape ellipticity from Hu moment invariants. Applied Mathematics and Computation, 226 (??):406–414, January 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313011193.

Zhang:2015:CBM

[ZZ15] Songlin Zhang and Kun Zhang. Corrigendum to ‘On a basic multivariate EIV model with linear equality constraints’ [Applied Mathematics and Computation 236 (2014) 247– 252]. Applied Mathematics and Computation, 265(??):1155, August 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300315007778. See [ZZ14c].

Zhang:2010:HBS

[ZZC10] Guodong Zhang, Lulu Zhu, and Boshan Chen. Hopf bifurcation and stability for a differential-algebraic biological economic system. Applied Mathematics and Computation, 217(1):330–338, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300310006247.

Zhou:2013:NTR

[ZZC13] Qunyan Zhou, Fen Zhou, and Fengxue Cao. A nonmonotone trust region method based on simple conic models for unconstrained optimization. Applied Mathematics and Computation, 225(??):295–305, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313010175.

Zhou:2012:NSS

[ZZF12] Anwa Zhou, Zhibin Zhu, and Hao Fan. A new semi-supervised PSVM classifier. Applied Mathematics and Computation, 219 (8):4006–4012, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300312010387.

Zhu:2010:FSM

[ZZG10] Zhibin Zhu, Weidong Zhang, and Zhenjie Geng. A feasible SQP method for nonlinear programming. Applied Mathematics and Computation, 215(11):3956–3969, February 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300309010340.

Zhang:2014:CIR

[ZZGL14a] Li-Tao Zhang, Yong-Wei Zhou, Tong-Xiang Gu, and Xing-Ping Liu. Convergence improvement of relaxed multisplitting USAOR methods for H-matrices linear systems. Applied Mathematics and Computation, 247 (??):225–232, November 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314012168.

Zhang:2014:ICT

[ZZGL14b] Li-Tao Zhang, Xian-Yu Zuo, Tong-Xiang Gu, and Xing-Ping Liu. Improved convergence theorems of multisplitting methods for the linear complementarity problem. Applied Mathematics and Computation, 243(??):982–987, September 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314008741.

Zheng:2011:FFO

[ZZH11] Quan Zheng, Peng Zhao, and Fengxi Huang. A family of fourth-order Steffensen-type methods with the applications on solving nonlinear ODEs. Applied Mathematics and Computation, 217(21):8196–8203, July 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300311001342.

Zhu:2014:LWI

[ZZHZ14] Hong Zhu, Ligang Zhang, Jiecai Han, and Yumin Zhang. Love wave in an isotropic homogeneous elastic half-space with a functionally graded cap layer. Applied Mathematics and Computation, 231(??):93–99, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S009630031400006X.

Zhao:2014:NSC

[ZZJ14] Jiandong Zhao, Zhengchun Zhang, and Jing Ju. Necessary and sufficient conditions for permanence and extinction in a three dimensional competitive Lotka–Volterra system. Applied Mathematics and Computation, 230(??):587–596, March 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313014306.

Zhang:2013:PEE

[ZZL13] Tong Zhang, Xin Zhao, and Gang Lei. A posteriori error estimates of stabilized finite element method for the steady Navier–Stokes problem. Applied Mathematics and Computation, 219(17):9081–9092, May 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313003160.

Zhou:2014:BAS

[ZZL14] Tingting Zhou, Weipeng Zhang, and Qiuying Lu. Bifurcation analysis of an sis epidemic model with saturated incidence rate and saturated treatment function. Applied Mathematics and Computation, 226(??):288–305, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300313010758.

Zhang:2010:NAT

[ZZQ10] Jian Zhang, Kecun Zhang, and Shaojian Qu. A nonmonotone adaptive trust region method for unconstrained optimization based on conic model. Applied Mathematics and Computation, 217(8):4265–4273, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310011033.

Zhou:2013:MSA

[ZZQ13] Hui Zhou, Zongfu Zhou, and Zongmin Qiao. Mean-square almost periodic solution for impulsive stochastic Nicholson’s blowflies model with delays. Applied Mathematics and Computation, 219(11):5943–5948, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300312012635.

Zhao:2011:ENC

[ZZR11] Xuebin Zhao, Zhao Zhang, and Qin Ren. Edge neighbor connectivity of Cartesian product graph G. Applied Mathematics and Computation, 217(12):5508–5511, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310012221.

Zhao:2014:BTS

[ZZR14] Zhonglong Zhao, Yufeng Zhang, and Wenjuan Rui. Bäcklund transformation and solutions of a (3 + 1)-dimensional nonlinear evolution equation. Applied Mathematics and Computation, 248(??):456–462, December 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314013903.

Zhu:2014:RAP

[ZZS14] Song Zhu, Kai Zhong, and Yi Shen. Robustness analysis for parameter matrices of global exponential stability time varying delay systems. Applied Mathematics and Computation, 237(??):386–395, June 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314004925.

Zhang:2011:MHB

[ZZW11a] Chunrui Zhang, Baodong Zheng, and Liancheng Wang. Multiple Hopf bifurcations of three coupled van der Pol oscillators with delay. Applied Mathematics and Computation, 217(17):7155–7166, May 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311001524.

Zhou:2011:PAP

[ZZW11b] Hui Zhou, Zongfu Zhou, and Qi Wang. Positive almost periodic solution for a class of Lasota–Wazewska model with infinite delays. Applied Mathematics and Computation, 218 (8):4501–4506, December 15, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300311012616.

Zhang:2012:MCS

[ZZW^{+}12] Xiaoyuan Zhang, Jianzhong Zhou, Changqin Wang, Chaoshun Li, and Lixiang Song. Multiclass support vector machine optimized by inter-cluster distance and self-adaptive deferential evolution. Applied Mathematics and Computation, 218(9):4973–4987, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300311013099.

Zhang:2013:OPL

[ZZW13] Yanxia Zhang, Xuefeng Zhang, and Zhongzhi Wang. Oscillation problem of Left-Definite Sturm–Liouville Problems with coupled BCs.

Applied Mathematics and Computation, 219 (18):9709–9716, May 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031300307X.

Zhang:2014:GLN

[ZZW14] Yufeng Zhang, Zhonglong Zhao, and Guangming Wang. On generating linear and nonlinear integrable systems with variable coefficients. Applied Mathematics and Computation, 244(??):672–682, October 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314009874.

Zhangchun:2013:MET

[ZZWF13] Tang Zhangchun, Lu Zhenzhou, Pan Wang, and Zhang Feng. A mean extrapolation technique for high reliability analysis. Applied Mathematics and Computation, 222(??):82–93, October 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300313007972.

Zhao:2014:CCT

[ZZWY14] Shouwei Zhao, Zihou Zhang, Tianbo Wang, and Weiqin Yu. Controllability for a class of time-varying controlled switching impulsive systems with time delays. Applied Mathematics and Computation, 228(??):404–410, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0096300313012721.

Zheng:2014:LSI

[ZZX14] Yutao Zheng, Bing Zheng, and Zhiping Xiong. The least squares g-inverse for the sum of multiple matrices. Applied Mathematics and Computation, 231(??):591–600, March 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300313014525.

Zhao:2012:MPM

[ZZY12] Jinling Zhao, Yanjun Zhang, and Qingzhi Yang. Modified projection methods for the split feasibility problem and the multiple-sets split feasibility problem. Applied Mathematics and Computation, 219(4):1644–1653, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0096300312007977.

Zhao:2014:SIS

[ZZY^{+}14a] Feng Zhao, Qingling Zhang, Xing-Gang Yan, Min Cai, and Junchao Ren. Stochastic input-to-state stability and H_{∞} filtering for a class of stochastic nonlinear systems. Applied Mathematics and Computation, 227(??):672–686, January 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S009630031301240X.

Zhou:2014:SSD

[ZZY14b] Yanli Zhou, Weiguo Zhang, and Sanling Yuan. Survival and stationary distribution of a SIR epidemic model with stochastic perturbations. Applied Mathematics and Computation, 244(??):118–131, October 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/
S0096300314009448.

Zhang:2012:CLP

[ZZYX12] Weiguo Zhang, Yong Zhang, Xingyu Yang, and Weijun Xu. A class of on-line portfolio selection algorithms based on linear learning. Applied Mathematics and Computation, 218(24):11832–11841, August 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S009630031200553X.

Zhang:2011:HBA

[ZZZ11] Lan Zhang, Chengjian Zhang, and Dongming Zhao. Hopf bifurcation analysis of integro-differential equation with unbounded delay. Applied Mathematics and Computation, 217(10):4972–4979, January 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300310011756.

Zhang:2014:CGSb

[ZZZ14] Xianmin Zhang, Xianzhen Zhang, and Min Zhang. On the concept of general solution for impulsive differential equations of fractional order q ∈ (0, 1). Applied Mathematics and Computation, 247(??):72–89, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/article/
                                        pii/S0096300314011795.

Zhu:2014:HBS

[ZZZL14] Mu-Zheng Zhu, Guo-Feng Zhang, Zhong Zheng, and Zhao-Zheng Liang. On HSSbased sequential two-stage method for non-Hermitian saddle point problems. Applied Mathematics and Computation, 242 (??):907–916, September 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/article/pii/
                                        S0096300314009278.

Zheng:2010:VSS

[ZZZM10] Quan Zheng, Peng Zhao, Li Zhang, and Wenchao Ma. Variants of Steffensen-secant method and applications. Applied Mathematics and Computation, 216(12):3486–3496, August 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/article/
pii/S0096300310004868.